A Complete Bibliography of Publications in the British Journal for the History of Science: 1962–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.18

Title word cross-reference

1/2 [Hei71b, Hen84c]. $10 [Dol74, Woo73]. $10.00 [Ait86, Rav69]. $10.50 [Kni75]. $10.95 [Wei85a, Mor77a]. $100 [Rud86]. $100.00 [Nic85]. $105.95 [Dor79]. $11 [Lev73, Web72]. $11.00 [Gou84a]. $11.90 [GG75c]. $11.95 [Rus69]. $110.00 [Hil89b]. $12 [Mol74, Rav67]. $12-50 [Har77]. $12.00 [Rav69, Hos67, Jor78]. $12.50 [Dun69a, Dur79a, Hei75, Huf85, Lor77, Sch69a, Web66c, Web69b]. $12.95 [Mor72a]. $13.50 [Cha78a, Kni69b, Wil84d]. $14 [Har75]. $14.25 [Por87c]. $14.50 [Buc68]. $14.95 [Coc75, Fis81b, Har79b]. $15 [Cro89a, Bow74]. $15.00 [Cal77, Gab83a, Hac74, Nor71b, Sin76, Swe76]. $15.95 [Hen81, Rus87a, Hes74]. $16.95 [Fur85].

$17.50 [Hen81, Hop73, Fin68, Kel73, Mac80, Sch79a, Tur78c]. $17.80 [Gra83b]. $18.95 [Cun88, Hal80c, Smi80a]. $18.95/$7.95 [Kni81]. $19.25 [Nev78]. $19.50 [Nut84, Sch78b, Coo84a, Tho76b]. $19.75/£ [Por79b]. $19.95 [Pic84a, Pum87, Dur82, Lew85]. $2 [Fox74]. $2.25 [Lar74]. $2.45 [Rav72]. $2.50 [Kel68, Ste70]. $20 [Ste62c, Hop75, Mor84b]. $20.00 [Hen81, Hop73, LM74, Mor87a, Nay85, Por81, Rat69, Tre88]. $20.00/£ [Goo81b]. $21.00 [Goo82]. $21.95 [Roc88b]. $22.50 [Mol79, Sha86]. $22.75 [Col81b]. $23.50 [Mol84c]. $24.95 [O’H86, Edg87c, O’H86]. $25 [Hei74a]. $25.00 [Bad89, Far77, Grm69, Nak75, Nor85b, Sch74a, Tre87b, War89]. $25.95/514.95 [Wor88]. $26.95 [Cah88]. $27.00 [Nor72]. $27.50 [Bro85a, For87c, Lud80]. $27.50/£ [Joh81]. $27.75 [Nad83]. $28 [Fre86]. $28.00 [Cha80b, Gil88, Pin87b, Whi84]. $28.50 [Wal89]. $29.00 [Fra87c, Gra86a, Fri87]. $29.50 [Ver89b, McG68, Nun88]. $29.95 [Bow88a, Bro86d, Hal86d, Kau84, Pic88a, Ric89]. $3.50 [Hen82c, Lay74]. 3/4 [Cal63]. $30 [Goo85b, Bro85e, Goo85b, Smi85d]. $30.00 [Can80b, Hac77, Kee88, Sei85]. $31.50 [Can84]. $32.50 [Hen88b, War88b]. $32.50/£ [Mor83b]. $32.95 [Bro89f, Hen88a, Kee89]. $34.00 [Kni89d]. $34.50 [Hal85b, Hal85a, Tre87a, Wil87c]. $35 [Wei85a, Tha85]. $35.00 [Bow87a, Col83, Bud84]. $35.00/£ [Jor85b]. $35.95 [Hoc88b]. $36.00 [Bro86a]. $36.10 [Hen84j]. $37 [Ped85]. $37.00 [Hen89e, Kni88b, Mol85]. $37.50 [Rud86, Sec84]. $38.35 [Rud84]. $38.45 [Por87c]. $39 [Bro85d]. $39.00 [Lin86]. $39.50 [Mac85b, Aas87, Cla86a, Dea86a, HS88, Jac89b, Mül87, Nee85, Tei88, Whi89a]. $39.90 [Cla86b]. $39.95/£ [Coo81]. $4.95 [Byn78, Koh78]. $40 [Rat66a]. $40.00/£ [Can80c, Gra81]. $42 [Bod84]. $42.00 [Bow89b, Cas82]. $42.50 [Pic84a, Sti89]. $44.50 [Cas85, Rus87a, Coo89a, Sho86a]. $44.95 [Bro89e]. $45 [Whi85a]. $45.00 [Nye86, RH88, Rup89]. $47.50 [Geo87]. $49.00 [Hen87c]. $49.50 [Fur85, Pum87, Byn88, Cro87a, Fra87b, Gol87b, Por89g]. $49.95 [Mol84c, Fie86]. $5.00 [Kni83a, GG77b, Sch68a]. $5.95 [Koh78]. $51.00 [Wil69, Ser88]. $52 [Hal84d]. $54.50 [Bro89c, Gou85]. $58 [Buc85a, Hal86b]. $58.50 [Eme87]. $59 [Cla86b]. $59.00 [Kni84a, Kni84b]. $59.50 [Sho88, Bro86b, Edg84, Goo85a, Goo88a]. $6 [Und66c]. $6.00 [Lar74]. $6.50 [Ait86, Cha65]. $6.75 [Buc66]. $6.95 [Far71b]. $62.00 [Blo79]. $64.00 [Hen88e, Kni89b, Ner89]. $65.00 [Hen83b, Hen89e]. $7.50 [Col88a, May66, Gin84, May69, Sha83a].

$7.95 [Koh78, GG72, Olb70]. $7.95/$4.95 [Wea79]. $70 [Wuj85]. $70.00 [Mor83c]. $75.00 [Mül89, Sch80b]. $77.00/£ [Tur83c]. $79.00 [Sha89, Whe89]. $79.50 [Jam86, Bro89b, Smi87c]. $8.00 [Gid82, Tur83a, Woo83]. $8.50 [Rav72, Smi78b]. $8.95 [Koh78]. $80.00 [Hil89a]. $9 [Arm66]. $9.50 [Nut84, Mac72]. $9.75 [Duc84, May67]. $9.95 [Hal85b, Ver89b]. $99.50 [Bow87b, Hal88b]. $A1.50 [Dol69]. γ [Stu71b]. π [Tee89]. × [Cal63].

-Rays [Stu71b].

/1235 [Pic84a]. /1235-2 [Pic84a]. /1323 [Pic84a]. /1323-5 [Pic84a]. /539 [Wor88]. /539-X [Wor88]. /Cambridge [Des88]. /£ [Hac82b, Hen82a, Wea83].

0 [Bow89c, Bra84, Bro84a, Hen83a, Hen83b, Hen84d, Hen84f, Hen84e, Ink88, Sec83, Sin84, Wei86]. 0-00-217555-X [Kni88a]. 0-00-686019-2 [Ink87]. 0-09-145180-9 [Sec85]. 0-09-146170-7 [Blo84]. 0-09-146180-4 [Blo84]. 0-09-151450-9 [Blo84]. 0-11-290324-X [Ben84]. 0-11-290423-8 [Wil87b]. 0-11-290435-1 [Hac88a]. 0-11-580088-3 [All84b]. 0-11-630942-3 [Por89d]. 0-12-273150-6 [Kni89e]. 0-19-287587-6 [Out85]. 0-19-287588-4 [Out85]. 0-19-287624-4 [Bow85]. 0-19-287625-2 [Bow85]. 0-19-287635-X [Gab88]. 0-19-287636-8 [Gab88]. 0-19-503964-5 [GG87]. 0-19-822606-3 [Mor84d]. 0-19-822901-1 [Pep86]. 0-19-822907-0 [Rud84]. 0-19-822947-X [Por89c]. 0-19-851997-4 [Hen89a]. 0-19-853741-7 [Rus89a]. 0-19-853904-5 [Sin84]. 0-19-853907-X [Red84]. 0-19-853924-X [Bra89a]. 0-19-854430-8 [Rup85]. 0-19-854431-6 [Rup85]. 0-19-858157-2 [Wil85b]. 0-19-858159-9 [Kel84]. 0-19-858164-5 [Sch85a]. 0-19-858173-4 [Swa86]. 0-19-858174-2 [Wil84c]. 0-226-07882-5 [Roc87]. 0-226-16228-1 [Nay85]. 0-226-24823-2 [Lin89b]. 0-226-27915-4 [Hen89d]. 0-226-36050-4 [Ver89a]. 0-226-46583-7 [Hoc88b]. 0-226-46950-6 [Bro89f]. 0-226-64562-2 [Bro84c]. 0-226-68474-1 [Bow87c]. 0-226-70946-9 [Bud84]. 0-226-72835-8 [Rap86a]. 0-226-73101-4 [Mor87b]. 0-226-81304-5 [Bro89e]. 0-226-84881-7 [Wil87c]. 0-226-89721-4 [Hen88a]. 0-231-05676-1 [Bro85d]. 0-231-05996-5 [Eme87]. 0-231-05997-3 [Eme87]. 0-231-06016-5 [Fri87]. 0-231-06442-X [RH88]. 0-252-01406-5 [War89]. 0-262-01083-6 [Fra87a]. 0-262-02179-X [Out84a]. 0-262-08114-8 [Mea84]. 0-262-11083-0 [Olb85]. 0-262-11117-9 [Wil89a]. 0-262-13194-3 [Bea85b]. 0-262-16094-3 [Hen85a]. 0-262-19223-3 [Con86]. 0-262-19237-3 [Edg87a]. 0-262-19242-X [Sec88]. 0-262-23109-3 [Gra85a]. 0-262-23112-3 [Hal84c]. 0-262-23123-9 [Asp88]. 0-262-62040-5 [Rav84]. 0-299-09170-8 [Rus85]. 0-299-09430-8 [Bro85e]. 0-299-09450-2 [Lew85]. 0-299-09900-8 [Bow86]. 0-299-09980-6 [Col87b]. 0-299-11110-5 [Fel89]. 0-299-11114-8 [Fel89]. 0-300-01465-1 [Bro86c]. 0-300-03392-3 [Hen87b]. 0-312-00485-0 [Jud88]. 0-330-28112-7 [Nor85a]. 0-333-22087-0 [Hen84d]. 0-333-29272-3 [Ins86]. 0-333-36942-4 [Can86]. 0-333-39320-1 [Hen87a]. 0-333-41742-9 [Mor89a]. 0-333-45710-2 [Kni89e]. 0-335-15175-2 [Ear87]. 0-340-23805-4 [Stu85]. 0-340-26325-3 [GG85]. 0-387-13610-X [Gra87b]. 0-387-90626-6 [Gra84c]. 0-387-90942-7 [Gra86b]. 0-387-96257-3 [Mol88]. 0-387-96403-7 [Tur88]. 0-394-50702-9 [Hoc88a]. 0-394-55394-2 [Kee88]. 0-415-00359-8 [Mar89]. 0-416-01341-4 [Rog88a]. 0-416-73480-4 [Kil84b]. 0-420-04836-9 [Fre86]. 0-434-603953-3 [Bru86]. 0-444-86897-6 [Fra87c, Gra86a]. 0-520-04318-9 [Hal84d]. 0-520-04415-0 [Hal84d]. 0-520-04592-2 [Rus87b]. 0-520-04621-8 [Bod84]. 0-520-04871-7 [RH86]. 0-520-04970-5 [Dea86b]. 0-520-05160-2 [McC85b]. 0-520-05213-7 [Smi87b]. 0-520-05538-1 [Rus87b]. 0-520-05694-9 [Har89a]. 0-520-05710-4 [San87]. 0-520-05792-9 [Bow88a]. 0-520-05830-5 [Bow89b]. 0-521-08690-6 [Hal86c]. 0-521-22743-7 [Law87]. 0-521-23164-7 [Hen85c]. 0-521-23212-0 [BH84b]. 0-521-23419-0 [Smi85c]. 0-521-23741-6 [Bra84]. 0-521-23930-3 [Bal84]. 0-521-24005-0 [Cas85]. 0-521-24172-3 [Cla86a]. 0-521-24206-1 [GG84b]. 0-521-24245-2 [Bro89c]. 0-521-24308-4 [Pum84]. 0-521-24349-9 [Jor87]. 0-521-24477-3 [Hun84]. 0-521-24600-8 [Hen84f]. 0-521-24608-3 [Hen83a]. 0-521-24647-4 [Lie85]. 0-521-24689-X [Wea84]. 0-521-24746-2 [Law85]. 0-521-24807-8 [Hen84k]. 0-521-24876-0 [Hen84m]. 0-521-24935-X [Mül87]. 0-521-25076-5 [Sho87a]. 0-521-25098-6 [Mac85b]. 0-521-25104-4 [Hut89]. 0-521-25121-4 [Nee85]. 0-521-25226-1 [Ait85a]. 0-521-25248-2 [Hal86a]. 0-521-25272-5 [Pum87]. 0-521-25314-4 [Fur85]. 0-521-25404-5 [Cro87a]. 0-521-25444-2 [Osb85]. 0-521-25548-1 [Smi87c]. 0-521-25565-1 [Sho88]. 0-521-25587-2 [Rud86]. 0-521-25588-0 [Rud88b]. 0-521-25593-7 [Kau84]. 0-521-257336 [Wil86b]. 0-521-25746-8 [Huf85]. 0-521-25752-2 [Hal85a]. 0-521-25804-9 [Moo85]. 0-521-25858-8 [Hal85b]. 0-521-25879-0 [Dea86a]. 0-521-25953-3 [Bow87b]. 0-521-25978-9 [Hal87b]. 0-521-26305-0 [Sch87]. 0-521-26505-3 [Sho86a]. 0-521-26582-7 [Huf86]. 0-521-26584-3 [Rec86]. 0-521-26590-8 [Sch86a]. 0-521-26707-2 [Hen86e]. 0-521-26720-X [Aas87]. 0-521-26746-3 [Rob87]. 0-521-26806-0 [Gol87b]. 0-521-27243-2 [Pye87]. 0-521-27307-2 [Fur85]. 0-521-27784-1 [Hal87b]. 0-521-28619-0 [Jor87]. 0-521-28812-6 [Hen84f]. 0-521-30358-3 [Hal88b]. 0-521-30408-3 [Gol88b]. 0-521-30529-2 [Edg87c]. 0-521-30532-2 [Coo89a]. 0-521-30657-4 [Geo87]. 0-521-30999-9 [Pic88a]. 0-521-31560-3 [Pum87]. 0-521-31683-9 [Gol88b]. 0-521-31742-8 [Rus87a]. 0-521-32016-X [Pin88]. 0-521-32021-6 [Hil89b]. 0-521-32029-1 [Tei88]. 0-521-32105-0 [War88b]. 0-521-32279-0 [Byn88]. 0-521-32523-4 [Sti89]. 0-521-32997-3 [Pum88]. 0-521-33163-3 [Tho89]. 0-521-33328-8 [War88a]. 0-521-33360-1 [Hen88b]. 0-521-33512-4 [Hil89a]. 0-521-34013-6 [Hen89b]. 0-521-34017-9 [Hen89b]. 0-521-34093-4 [Bow88b]. 0-521-34143-4 [Eva89]. 0-521-34254-6 [Jac89b]. 0-521-34255-4 [Ric89]. 0-521-35055-7 [Des88]. 0-521-36712-3 [Tho89]. 0-56500-850-1 [Tha84b]. 0-56500-850-1/0-7201-1655-4 [Tha84b]. 0-566-05147-8 [Wil89b]. 0-582-49133-9 [Hen84a]. 0-600-38423-3 [Hen84c]. 0-631-13134-5 [Wil84e]. 0-631-14188-X [Liv86]. 0-631-15641-0 [Hen89c]. 0-631-15739-5 [Fal89]. 0-631-16176-7 [Rob89a]. 0-632-01791-0 [Smi89c]. 0-632-01792-9 [Smi89c]. 0-674-04525-4 [Hut84]. 0-674-15420-7 [Por87b]. 0-674-16985-9 [Hac89]. 0-674-40340-1 [Hod89a]. 0-674-48050-3 [Sch84b]. 0-674-50507-7 [Bad89]. 0-674-76777-2 [Wil86a]. 0-674-88330-6 [Pic88b]. 0-691-04694-8 [Coo84a]. 0-691-05137-2 [Cun89]. 0-691-05335-0 [Wei85b]. 0-691-05375-8 [Bro84a]. 0-691-05379-0 [Hen84l]. 0-691-06491-1 [Jor85a]. 0-691-08299-5 [Kil84a].

0-691-08321-5 [Car84]. 0-691-08355-X [Dea86c]. 0-691-08356-8 [Smi89e]. 0-691-08393-2 [Jon87]. 0-691-08407-6 [Hen88d]. 0-691-08414-9 [Nun88]. 0-691-08416-5 [Hod89a]. 0-691-08417-3 [Tha87]. 0-691-08497-1 [Por89g]. 0-7043-3954-4 [Out87]. 0-7099-1206-4 [Mor84c]. 0-7099-1209-9 [HS86a]. 0-7099-1836-4 [Por88]. 0-7099-2104-7 [Ron86]. 0-7099-3383-5 [Por89b]. 0-7099-3962-0 [Rup88]. 0-7100-0009-X [Har84b]. 0-7100-9208-3 [Har84b]. 0-7100-9530-9 [Hen85b]. 0-7100-9602-X [Hea87b]. 0-7100-9809-1 [For87b]. 0-7100-9845-6 [Sha87]. 0-7102-0279-2 [Rog88b]. 0-7102-0576-7 [Ric88a]. 0-7102-0919-3 [Law88]. 0-7108-0451-2 [Hen84f]. 0-7108-1042-3 [Jud88]. 0-7126-0722-6 [Kni86c]. 0-7156-1588-2 [Van86]. 0-7156-1693-5 [Whi85b]. 0-7156-2100-9 [Hut88]. 0-7190-0868-9 [Har84a]. 0-7190-0938-3 [Smi85a]. 0-7190-1077-2 [Cro85]. 0-7190-1741-6 [Smi87a]. 0-7190-1756-4 [Jam87b]. 0-7190-1809-0 [Web87]. 0-7190-2338-6 [Ait89]. 0-7190-2537-0 [Law89]. 0-7190-2621-0 [Bow89a]. 0-7197-1070-5 [McL86]. 0-7201-1642-2 [Hen84g]. 0-7201-1643-0 [Hen84g]. 0-7201-1655-4 [Tha84b]. 0-8014-1409-1 [Sch84c]. 0-8014-1444-X [Hen84j]. 0-8014-1467-9 [Rud84]. 0-8014-1583-7 [Hal86d]. 0-8014-1770-8 [For87c]. 0-8014-1875-5 [Wal89]. 0-8018-2380-3 [Jac85]. 0-8018-2932-1 [Nor85b]. 0-8018-3229-2 [Sha86]. 0-8018-3375-2 [Yeo88]. 0-8018-3558-5 [Mor89a]. 0-8047-1075-9 [Nic85]. 0-8047-1119-4 [Sec84]. 0-8047-1364-2 [Rup89]. 0-8071-1293-3 [Col88b]. 0-8122-8057-1 [Woo89]. 0-8122-8092-X [Bow89c]. 0-8130-0700-3 [All84a]. 0-8130-0760-7 [Smi85d]. 0-8135-1118-6 [Bow87a]. 0-8135-1193-3 [Gil88]. 0-8138-1159-7 [Whi89b].

0-8142-0351-5 [Kei86]. 0-8142-0360-4 [Bro85a]. 0-8173-0295-6 [Cah88]. 0-8173-0305-7 [Kee89]. 0-8176-3317-0 [Gre88]. 0-8176-3318-9 [Ser88]. 0-8191-2607-1 [Nut84]. 0-8214-0692-2 [Hal84e]. 0-8214-0703-1 [Hal84e]. 0-8240-8557-4 [Kni88b]. 0-8240-9224-4 [Rud86]. 0-8240-9267-8 [Tha85]. 0-8240-9278-3 [Bud84]. 0-8240-9283-X [Hen83b]. 0-8248-1120-8 [Kni89d]. 0-8419-0837-0 [Wei86]. 0-8419-0977-6 [Tre87a]. 0-85066-253-2 [Kni86d]. 0-85131-379-5 [Pel85a]. 0-85186-917-3 [Bro86a]. 0-85224-417-7 [Law85]. 0-85224-458-4 [Hen86b]. 0-85264-283-0 [Mah89]. 0-85274-274-6 [Hac88b]. 0-85274-416-1 [Ben84]. 0-85274-423-4 [Whi88a]. 0-85274-456-0 [Hun86]. 0-85274-470-6 [Hal87a]. 0-85274-524-9 [Kei89]. 0-85274-526-5 [Buc89a]. 0-85274-527-3 [Fie89]. 0-85274-545-1 [Whi84]. 0-85274-555-9 [Tre87b]. 0-85274-563-X [The87]. 0-85274-569-9 [Tur87a]. 0-85274-759-4 [Stu85]. 0-85274-761-6 [Stu85]. 0-85274-774-8 [Dra86]. 0-85274-779-9 [Pin87b]. 0-85274-801-9 [Nye86]. 0-85496-191-7 [Por89a]. 0-85634-121-5 [Tha84a]. 0-85667-170-3 [McC85b]. 0-85976-072-3 [Hil86b]. 0-85976-107-X [Aus86]. 0-86027-014-9 [Ham84]. 0-86193-103-3 [Pat86]. 0-86228-026-5 [Smi85b]. 0-86228-056-7 [Lie86]. 0-86228-065-6 [Bro86e]. 0-86341-101-0 [Hun89a]. 0-87049-517-8 [Cun88]. 0-87169-744-0 [Ait86]. 0-87169-983-4 [Mor84b]. 0-87338-282-X [Whi85a]. 0-87474-540-3 [Wor88]. 0-87474-785-6 [Ink88]. 0-88706-910-X [Ver89b]. 0-88844-068-5 [Lin86]. 0-88920-152-8 [GG89]. 0-89835-242-8 [Mor87a]. 0-89859-342-5 [Fie86]. 0-89874-573-X [Ait86]. 0-89874-939-5 [Rud89]. 0-900733-29-2 [Jam87a]. 0-901805-28-9 [Hen89g].

0-904357-53-8 [Cla86b]. 0-904357-64-3 [Cla86b]. 0-905484-15-0 [Lay85]. 0-905555-60-0 [Wil85a]. 0-906450-02-0 [Wil84d]. 0-906450-03-9 [Gou84a]. 0-906716-01-2 [Sec83]. 0-906783-05-4 [O’H86]. 0-906783-06-2 [O’H86]. 0-906795-36-2 [Bro87]. 0-907582-67-2 [Mac89]. 0-913870-64-1 [Jar84b]. 0-918102-05-7 [Hen83b]. 0-918102-15-4 [Tre88]. 0-934235-11-2 [Cro89a]. 0-936508-09-4 [Kni86b]. 0-938228-06-4 [Sei85]. 0-939454-03-3 [Roc88b]. 0-941901-05-X [Col88a]. 0-9506207-1-8 [Kni89e]. 0-9507276-1-X [For85a]. 0-9513586-1-8 [Kni89c]. 000 [Por79a, Sch79b, Sch80c]. 0068 [Ree85]. 0068-0729 [Ree85]. 008 [Kni83a]. 008-022-00165-4 [Kni83a]. 0269-8897 [Hen89e]. 0435-2866 [Duc88]. 0729 [Ree85]. 0734-1512 [Buc85a].

1 [Dur80a, Han85a, Har87b]. 1-2 [Bur74b]. 1-55553-030-3 [Cro89b]. 1-85027-003-1 [Fau89]. 1-85196-005-8 [Twe89]. 1-85401-008-9 [For87a]. 1-90 [Ano82n, Ano82m, Bro73a, Fox79]. 1-95 [Mor77c]. 1.20 [Mor72b]. 1.25 [Bro83a, Ora74]. 1.30 [Olb71b]. 1.50 [Ben79, Eag78]. 1.75 [For71a, Mor74a, Web71b]. 1/0 [Tha84b]. 1/2 [Gua89]. 10 [Bry74, For72, Har82d, Til75a, Whi72b]. 10.00 [Blu79, Goo81b, Hac76, Rav84, Smi76a, Tei83a, Wea78, Yox83a]. 10.00/$23.00 [Kni83c]. 10.00/£ [Goo81a]. 10.25 [Ree80]. 10.50 [Gow78, Hal77b, Har76, Kel85a, Nay80, Nev83]. 10.70 [Hop72]. 10.80 [Irw83b]. 1000 [Mac83]. 105s [Sty68]. 10s [Cow68, Kni65, Nor68a, Nor69a, Nor70a, Tur70b]. 11.00 [Can77, Duf75, Nor77]. 11.50 [Bro80b, Wal77]. 11.50/$20.60 [Rid83]. 11.55 [Gur78]. 11.75 [Out83]. 11.95 [Jor79]. 110/$25.00 [Bal83]. 114s [Kub70].

115.00 [Pic87a]. 116e [Ano93]. 116th [Ano93]. 118.00 [Bra88]. 12.00 [Col76, Lev80, Por82, Sin77, Smi76c, Til76, Tup82, Tyl75, Wig74, Wig78]. 12.00/$21.00 [All82]. 12.00/£ [Dur81, Fie83]. 12.25 [Law80b, Law81, Out84a]. 12.25/$21.85 [Coo83]. 12.50 [Dol83, GG75a, Kel84, Nor83, Rus79b, Sin81, Tur77a, Wil86b]. 12.50/$17.95 [Ora80]. 12.75 [Por85a]. 12.75/522.10 [Por85a]. 12.95 [GG85, Hen84c]. 123 [Hom86b]. 1235-2 [Pic84a]. 125 [Gou84b]. 128 [Sch83b]. 128.00 [Bra87]. 12s [McG65]. 13 [Ait72, Cap81b, Jar84b]. 13.95 [Woo85, Yox83b]. 1323-5 [Pic84a]. 135/$59.00 [Jar84a]. 14 [Bro82a, Jac85]. 14-22 [BG^{+}92]. 14-80 [Bro82a]. 14.00 [Bro84c, Hal80a, Joh79, Law80a, Mea84, Smi78a]. 14.50 [Mor82a, Mor84c, Yox82a]. 14.75 [Por79b]. 14.75/$30.60 [Pin83]. 14.80 [Hut84]. 14.85 [Hai77]. 14.95 [Cun83a, Hac82a, Sha87]. 140 [Por86a]. 140.00 [Hal86e, Out88a]. 140.00/$56.00 [Stu78]. 140s [Cla68]. 142s [Ear70]. 145.00 [Jor88, Sho89]. 145/$72.50 [Web84]. 1450 [Lin92]. 14s [Tur70c, You66]. 15 [Ins86, Red84, Ril85]. 15.00 [Blo84, Bry81, Bus85, Fre74, Ham84, Kei89, Kni86d, Lie86, Liv86, Out82a, Rus85, Sch81a, Tur81, Wil84e, Woo80a, Woo84]. 15.00/$25.00 [Gli83b]. 15.00/$31.25 [Kni83b]. 15.25 [San87]. 15.40/$23.50 [Yox83a]. 15.75 [Clo80]. 15.95 [Fis83a, Hea87b, Hen85b, Lay85, Ron86, Tha84a, Wil87b]. 150 [Ben89, Gil84, Mol81]. 150.00 [Ege89, Mül89, Osb89b]. 158.00 [Tur88]. 15s [Fox68b, Ste62b, Vea64]. 15th [Ano82m, NW82a]. 16.00 [Har83a, Les83]. 16.25 [Sch76]. 16.75 [Hun86]. 16.90 [Ros79]. 16.95 [Kni86c]. 16.95/$19.95 [Kei84]. 160 [Rus74]. 1610 [Wey92]. 1610/11 [Kra91]. 1660s [U^{+}91]. 1670 [Nor69a]. 1673-April [Woo80c]. 1674-September [Woo80c]. 1676 [LM23].

168.00 [Mor81]. 1683 [LHH91]. 1690s [Str87]. 16s [Hew66, Sme69]. 16th [BW90, Rat68, Woo82]. 17.00 [Dur83, Mor84f]. 17.10 [Duf83]. 17.30 [Dea86b]. 17.40 [Moo84]. 17.50 [Bra89a, Dea83b, Gra83a, Kni88a, Law85, Olb85, Por85b, Sec85, Smi78c, Stu85, Swa86]. 17.60 [Nor80]. 170 [Kra86]. 1700 [KAZ92]. 175 [Bad83]. 179 [Dun86]. 1797-October [Hac74]. 17th [Can80c, Hac75, Kle94, Rat68, Woo82, Grm90, Pic87c, Poi04]. 18 [Pac72, Pep72, Sch78b, Sha77b, Smi85b]. 18.00 [Bal84, Rus77c]. 18.40 [Mor80]. 18.50 [Coo81, Por89a, Wil85b]. 18.55 [Her80]. 18.95 [Bar81, Bea85b, HS86a]. 18/$29 [Dra86]. 1817 [Ink79]. 1832-December [Hac77]. 1836-December [Goo85b]. 1837 [OS92b]. 1838-December [Goo85b]. 184.£ [Sin81]. 1855-’56-’57 [Kni83a]. 1864-5 [BM76]. 1868 [BMLC84]. 1872 [vLvHH^{+}88]. 1877 [EJR94, EJR95]. 188.00 [Von88]. 18s [Hei68, Web68a]. 18th [Bry81, Can80c, Hac75, Kle94, Les83, Poi04, Tur73c, Bec91b, BaJANJ84, FHR90, Roc88a, Sch01, SGŚ00]. 18th-Century [Hac75]. 19 [BH84b]. 19.00 [Har84a, Joh81, Pum84]. 19.00/£ [Nay80]. 19.10 [Jor85a]. 19.50 [Ben84, GG84a, Hen84m, Mor83b, Nut82b]. 19.95 [Fie89, Law88, Por87b, Ric88a, Wei83a]. 1913-65 [Ano82n]. 1914 [Dau98]. 1930 [Sok87]. 1930s [Hoc88b, Kuz87]. 1939 [HP85, PvM85]. 1940 [Pes84, Rei86]. 1945 [PCA91]. 1950s [BDH89]. 1960s [Fis98]. 1963 [Cal63, Din63a, Had63]. 1964 [Spa65]. 1965 [Deb66]. 1968 [FGJ98]. 1968/1969 [Fox70]. 1971 [Ora73]. 1971-2 [WE72]. 1973-4 [Sha78]. 1973/1974 [CE74]. 1974 [Mor76a]. 1974-5 [Ano75o]. 1976 [Sme78b]. 1977 [Car78, Nee78]. 1978 [Cha80a, Sme79, Tor80]. 1979 [Ano82l, HE^{+}82, Sim84, Yox79]. 1981-2 [Ano82o]. 1982 [Cas85, Cro84a, KK85]. 1982-3 [Ano83m]. 1983 [Deb84, Hal85c, Hoc86, Sme84b]. 1984 [Bro85b, De 85, GD85]. 1985 [A^{+}87, Sme87]. 1986 [Hen86f, Mad88, N^{+}88, Sme88, Tor87]. 1987 [Cha88a, CM90, WV92]. 1988 [BG^{+}92, EK91, EK92]. 1990s [U^{+}91]. 1991 [Ano93]. 1992 [Ped92b]. 1993 [CT95]. 19s [Phi69, Sch70]. 19th [Ano82m, Bry81, CM90, Kni89a, KI92, NW82a, Dau98, Jun89, Man88, PP95, Sch89]. 19th-Century [Man88]. 1s [FF66]. 1st [PCA91].

2 [Hen84f, Hil86a, Sec83, Web65a]. 2-13-041414-8 [Ege89]. 2-213-02122-8 [Por89e]. 2-222-03094-3 [Sch84c]. 2-222-04052-3 [Jen89]. 2-222-04104-X [Ben89]. 2-7011-0640-0 [Out88a]. 2-7011-0641-9 [Out88b]. 2-7011-1085-8 [Cro88]. 2-7116-0825-5 [Sch86b]. 2-7116-2018-2 [Sch84c]. 2-7132-0875-0 [Fox87b]. 2-7177-1643-2 [Hen84b]. 2-7178-1323-3 [Gua89]. 2-7302-0019-3 [Sme84a]. 2-7342-0146-1 [Gua89]. 2-7342-0146-1/2-7178-1323-3 [Gua89]. 2-7355-0126-4 [Osb89a]. 2-8257-0099-1 [Hom86a]. 2-87027-180-8 [Jor88]. 2-900927-19-6 [Osb89b]. 2-903528-40-3 [Hil86a]. 2-903928-08-8 [Stu87]. 2.00 [Nev84]. 2.25 [Mar73]. 2.50 [Fee77, GG71, Hei74b, OM75, Smi71]. 2.70 [Olb72]. 2.75 [Bro72b, Cro72, Dau75, Gre71, Kel71]. 2.80 [Sha73a]. 2.95 [Fri74a, Ora77]. 2/3 [Sch75]. 20 [Hen86b, Mer78, Nor73a, Smi85a]. 20.00 [All81, Blo82b, Bro89a, Bro78a, Bro86c, Gra81, Hen84k, Sha79, Wil84c]. 20.95 [Yeo88]. 200/$30.00 [Sar81]. 204 [Mol84b]. 20th [CM90, Und75a, HW81, Lin93, Neb95]. 20th-Century [Lin93]. 21 [Hal84c, Hei73]. 21.25 [Col88b, Hac89]. 21.40 [Und75b]. 210F [Sch84c]. 21s [Arm64, Nor68b, Wil65a, Fox69b, Kuh68, Nor68b, Rav68, Smi66, Web66b, Whi68b]. 22 [Kra84a, Nor72]. 22.30 [Hai82]. 22.50 [Ben82b, Har83b, Hen87b, Kei86, Mor84d, Mor89a, Rec86, Rud84, Smi85c, Tei84]. 22.75 [Wil86a]. 22.80 [Wei85b]. 226/226a [Rav65]. 226a [Rav65]. 23-26 [Wil87d]. 23.50 [Che85, Wil89b]. 23.75 [Hen85a]. 23.95 [RH86]. 232 [Hen83c]. 24 [Sme68a]. 24.00 [Can80c, Hun89a, Pel85b]. 24.50 [Hac75]. 24.50/$43.75 [Tha83a]. 24.70 [Kil84a]. 24/$30 [Kei84]. 248 [HS86b]. 24DM [Mai85]. 25 [Cro85, GG84b, Hil86b, McL86, Sch86a, Whi73]. 25.00 [Ait89, All84b, Bro78b, Bry83, GG87, Hei76b, Hun84, Law87, Mor76a, Pep86, Por83a, Por89b, Rob87, Rog83, Rog88b, Sin78, Smi87a, War88a, Wea84]. 25.00/£ [Sha81]. 25.50 [Bow87c]. 250 [Rud83]. 250th [AL87, Hen89g]. 253 [MdLD86]. 258.00 [Mol88]. 25F [Sch84c]. 25s [Chi64, Chi65b, Und64a, Law65, Ste62a, Wil68]. 26.00 [Hen84l]. 26.25 [Jor85b]. 262 [Har86]. 267 [Eas86]. 269 [Fre82]. 27.50 [Bow89a, Che78, Hen86e, Jam87b, Kri83, Law89, Pic88b, Por88, Pum88]. 273 [Bro75]. 28 [Law80c, Pat84, Por84a]. 28.00 [Dor80, Jar83]. 28.00/$56.00 [Fox79]. 28.50 [Gra85a, Jud88]. 283.£ [Sha81]. 29 [Ano93, Rus79a]. 29.50 [Fal89, Smi89c, Whi85b, Whi88a]. 29.75 [Smi87b]. 29.95 [Edg87b, Hut88]. 290 [Mol78a]. 295.00 [Stu87]. 298.00 [Hen86a]. 2nd [Cha78a, Cro89a, EK92, Goo89b, Had63, Hah70, Her80, Nor72, Osb85]. 2O4 [Blu79]. 2s [Whi68c, Ant64, Cla66].

3 [Gra83b, Har87b, Hen83a, Hil86a, Olb71a]. 3-11-008912-2 [Hal85d]. 3-11-010317-6 [Bra87]. 3-11-010825-9 [Von88]. 3-406-31559-3 [Fig88]. 3-437-10939-1 [Har87b]. 3-540-10662-6 [Gra83b]. 3-540-13609-6 [Hen86a]. 3-540-17888-0 [Bro89b]. 3-540-90642-8 [Hen86b]. 3-540-90667-3 [Hen86b]. 3-540-90674-6 [Hen86b]. 3-540-90675-4 [Hen86b]. 3-540-90680-0 [Hen86b]. 3-540-96257-3 [Mol88]. 3-540-96284-0 [Hen88c]. 3-540-96403-7 [Tur88]. 3-540-96541-6 [Bra88]. 3-7643-1840-6 [Mah89]. 3-7643-3317-0 [Gre88]. 3-7643-3318-9 [Ser88]. 3-8047-0984-2 [Jac89a]. 3-8067-0902-5 [Mau87]. 3-88208-012-4 [Bro89a]. 3.00 [Buc73a, Ilt77, Nev80, Ste75]. 3.15 [Buc73b, Cla71, Kni71c]. 3.25 [Buc85b, Cri71, Hil71b, Joh78, Mar74]. 3.30 [Bar76b, Buc75]. 3.50 [Bar76a, Byn72, Dol71b, Dol73a, Dyb78, Hah73, Mos77a]. 3.55 [Rob75]. 3.60 [Bar75, Bur71]. 3.75 [Dol71a, Kni80a, Smi74a]. 3.80 [Bib74, Mos77b]. 3.85 [Bro72a]. 3.90 [Has72]. 3.95 [HS83, Kil84b, Mus74]. 3.95/£ [Fox82]. 30 [Eva89, Mac89]. 30.00 [Ear87, Hac88a, Hal87a, Har79a, Hen89b, Hen83d, Kni89e, Moo85, Pin88, Por89c, Smi83, The87]. 30.00/£ [Rud88b]. 30.20 [Car84]. 30.50 [Byn87]. 30/$11.95 [Byn80]. 3000 [Mol72b]. 30s [Arm62, Arm64, Smi66, Web66b, Far68a, Pag66, She63]. 31.40 [Cun89]. 31.95 [Ver89a]. 31s [Nee69]. 32 [Lin89b, Whi68a]. 32.00 [Joh83]. 32.50 [Ait85a]. 325.£ [Clo80]. 32F [Jor83a]. 33 [Bro73a]. 334.ISBN [Smi85d]. 339.50 [Out84b]. 35 [Hom74, Tur74a]. 35.00 [Buc89a, Can86, Hac88b, Hen87a, Huf86, Kni80b, Lie85, Mar89, Tur83c, Tur87a, Web87]. 35.50 [Bow88b]. 35.95 [Wil89a]. 350FF [Gua89]. 350th [PS82, vLP82]. 35O [Buc77]. 35s [Arm67, Bro69, Far67, Hal68d, Rat66b, Rus68]. 36 [Mol72a]. 36.60 [Col87b]. 36.75 [Mor87b, Rap86a]. 360 [Pug84]. 362 [Por87a]. 36s [Bod65]. 37.00 [Tha87]. 37.50 [Rud88b]. 37.95 [Con86]. 38 [Wil84b]. 38.00 [Cro89b]. 38/$18.50 [Sch84a]. 39.25 [Edg87a, Fra87a]. 397.00 [Dou84, Sch84b]. 3s [Had63].

4 [Bad71, Buc71, Can75, Kni83a, Sch73b, Tur69]. 4-20 [Fig77b]. 4.00 [McC77a, Pye76]. 4.00/$7.45 [Rid83]. 4.20 [Hei71a, Mor71]. 4.25 [Bar73a]. 4.40 [Cow74]. 4.50 [Bro73b, Cha77b, Fis81a, For74a, Han76, McG71a, Mon78, Mor73b, Tur73a, Tur74c]. 4.65 [Dub72, Fox71b]. 4.70 [Chi71]. 4.75 [Hal71, Lor73, Rap71, Tha74]. 4.95 [Har84b, Ink87, Mor74b, Out87, Sch74b, Smi73b, Tur78b]. 40 [Fis74]. 40.00 [All79, Sch87, Sch85a]. 40s [Geo66, Hal66a, Bro67, Ede65, Mor69a]. 42.45 [Pin87a]. 42s [McG68, Arm70, Cri66, For66, Hal68b, Rus66, Tha66a, Tha67, Web66a]. 43.00 [Jon87]. 44.00 [Fis83b]. 44.50 [Mil87]. 45 [Mac78]. 45.00 [Agr83b, Hal86c, McN83]. 45.80 [Dea86c]. 45/US [Mol84c]. 45s [Arm69, Fox69a, Jos67, Sme66]. 460.00 [Cro88]. 461 [Cun83b]. 47s [Whi70a]. 48 [Fau89, Nor71a]. 48.00 [Fig88, Gre88, Mol86]. 48s [Web65b]. 496 [Tee84]. 4s [Chi65a, Hig65, Rav66b].

5

[Boy70, Bru72, Cla70, Far71a, Ged71, Pac70, Rud69, Rud70, Sin74, Sin84, Tur73c, Wig69]. 5-October [Kel68]. 5.00 [For80, Kar76, Nay77, Smi74b, Tha83b, Und80]. 5.25 [Bro71a, Byn74]. 5.50 [Byn73, Dol73b, Kel76, Kni72, Sha73b, Smi75b, Wat73, Web76, Wig77]. 5.65 [Gow73]. 5.75 [Kni74]. 5.80 [Rav65]. 5.90 [Asq80]. 5.95 [Bur80, Rav71, Sey77]. 50 [Cro70b, Tha84b]. 50.00 [Hen85c, Hut89, Nut83, Sho87a, Und79b]. 500 [Out80b, Twe89]. 50s [Eyl68, Gre63, Hod70, McG66, Poy70]. 52 [Fox71a, Scr68a]. 522.10 [Por85a]. 52s [Goo68, Was69]. 53.50 [Jam86, Smi89b]. 531 [Bar83]. 539-X [Wor88]. 55.00 [Rus89a, Van86]. 55s [Kni69a, McG70a]. ’56 [Kni83a]. 56s [Und66b]. ’57 [Kni83a]. 57s [May66, Olb69]. 58 [Gra85b]. 59.50 [Roc87]. 5s [Chi65c, Din64, Got66, Scr68b, Tur68].

6 [Bra84, Dav71, Gin74, Gra83b, Ink88, Kni67, Mor73a, Rap75, Rau75, Sme70, Stu71a]. 6.00 [Cha78b, Jam87a, Sha77a]. 6.10 [Dor75]. 6.25 [Hei77, Lor75, Mor75a, Mor75c, Ran74, Und75a]. 6.30 [McG71b]. 6.50 [Goo75, Kni76, Law83, Loh75, Mor75b, Und79a, Wal74]. 6.55 [Web75]. 6.75 [Pic76]. 6.95 [Dav81, Jor87, Pic79, Woo80e]. 6/5 [Pep89]. 60 [Bry78b]. 60-521-2494-X [Dun87]. 60-521-2494-X/8 [Dun87]. 60.00 [Str73]. 600 [Lin92]. 60p [Rus72a]. 60s [Hea64]. 63 [Bry78c, Kni71b]. 63s [Buc69, Cro70a, Rav66a, Sti70, Woo67]. 64s [Cro68a]. 65 [Hil75]. 65.00 [Des88]. 65/$24.50 [Byn80]. 65/$29.50 [Web84]. 65/$34.00 [Sch84a]. 65s [Ant63, Dol70b]. 66s [Gow69, Hal69a, Men67]. 67 [Mor78]. 68.00 [Hom86a]. 69 [Tur75]. 69.50 [Asp88]. 6d [Hew66, Ear70, Gow69, Hei68, Mol69, Whi70a, Whi68c]. 6EL [Hal88a]. 6s [Fin68, Wel70, Car70, Hah70, McC69, Swi70].

7 [Ait73]. 7.00 [Bar79, Hei76a, Por77]. 7.30 [Mor77b]. 7.35 [Ait71, Sme71, Whi72a]. 7.50 [Bus74, Col81a, Mol71, Pat86]. 7.55 [Bar77]. 7.75 [Dur77]. 7.85 [Rud71]. 7.90 [For85a, Vic88]. 7.95 [Bro82b, Fis83c, Mac79, Rup85]. 7.95/£ [Sch80d]. 70.00 [Edw78]. 70F [Hai80]. 70p [Rus72a, Sol83]. 70s [Wil65b, Ait67, Dun69b, Lau68]. 72.00 [Gra86b]. 72A [Vic88]. 72s [Swi67]. 74 [Por84b]. 74.00 [Mol89]. 74s [Web68b]. 75 [Fis81c, Hal86a]. 75.00 [Mol78b]. 75.00/$175.00 [Rud80]. 75s [Car67a, Cla67, Rat68]. 76s [Web70]. 78 [Fis78]. 7s [Han70, Und66a, Web69a].

8 [Bro84a, Sch71, Yat72]. 8.00 [Fre78]. 8.25 [Men74]. 8.50 [Aus86, Bro86e, Fig78, McC77b, Wal81b, Whi75, Whi77]. 8.95 [Goo83a, Rap79]. 80 [Bro82a]. 800 [C^{+}95]. 80f [Wor79]. 81s [Dol70a, Far69]. 83-222 [Rat66b]. 84-00-06699-5 [Kni88c].

84-335-0029-5 [Goo89b]. 84s [Got67, Sch69b, Sco67]. 85.00 [Fra87c, Gra86a]. 85s [Mol69]. 88 [BG^{+}92]. 88-06-05632-8 [She85]. 8s [Hew66]. 8vo [Ste62a, Ste62c].

9-02772-417-2 [Ner89]. 9-11 [Hom74]. 9.00 [Agr83a, Tei78]. 9.30 [Gab75, Smi82a]. 9.45 [Cro71, Web78]. 9.50 [Cun78, Fre81b, Hol71, Pel85a, Tur78a, Web71a]. 9.75 [Lef78, Pic80, Wil78]. 9.95 [Bru86, For82, Hen89g, McV80, Rog88a, Yox82b]. 90 [Han85a, Hen84e, Pag69]. 90-227-2224-2 [Fra87b]. 90-247-2500-3 [Dou84, Sch84b]. 90-247-2508-9 [Gil84]. 90-247-2823-1 [Han85a]. 90-247-2897-5 [Mil87]. 90-247-3425-8 [Goo88a]. 90-277-1215-8 [Mol84c]. 90-277-1216-6 [Mol84c]. 90-277-1223-9 [Can84]. 90-277-1234-4 [Pic84a]. 90-277-1303-0 [Hen84e]. 90-277-1322-7 [Pic84a]. 90-277-1356-1 [Edg84]. 90-277-1363-4 [Kni84a, Kni84b]. 90-277-1389-8 [Ped85]. 90-277-1602-1 [Kni85]. 90-277-1624-2 [Bro86b]. 90-277-1637-4 [Gou85]. 90-277-1648-X [Che85]. 90-277-1682-X [Hal86e]. 90-277-1686-2 [Por86a]. 90-277-1693-5 [Hal86b]. 90-277-1713-3 [Byn87]. 90-277-1720-6 [Jam86]. 90-277-1945-4 [Hen87c]. 90-277-2152-1 [Pin87a]. 90-277-2265-X [Sha89]. 90-277-2351-6 [Pic87a]. 90-277-2424-5 [Mol89]. 90-277-2428-8 [Hen88e]. 90-277-2498-9 [Whe89]. 90-277-2541-1 [Kni89b]. 90-6191-621-6 [Smi89b]. 90-6191-681-X [Mül89]. 90-6203-824-7 [Hal84a]. 90-6203-917-0 [Wil87d]. 90-6292-070-5 [Wil87d]. 90.00 [Kni85, Out88b]. 90s [Car69, Cro69, Jev67, McG69, Web67a]. 91-22-00385-1 [Bro84d]. 91-22-00716-4 [HS86b]. 91-22-00827-6 [Nic88b]. 92 [Wea80]. 92-3-102294-6 [HS88]. 93s [Wel70]. 95/US [Mol84c]. 95p [Pac71]. 95s [Fox70, Kel70, Rav70]. 96377-4 [Hen88c].

973-1051 [Sar81]. 978-0-8133-0452-6 [Whi89a]. 97s [Pra70]. 98 [Gra84c, Hal85d]. 99F [Sch86b]. 99s [Men70]. 9th [PCA91].

= [AHY90, B^{+}89, GP92, KA81, SLT88, VB0 ].

âge [BBC88, Gol83b].

A-M. [Blo82a]. A. [For85a, Fre82, Hal88a, Hen87c, Jam86, Jam87a, Out88a, Wil87c, Wuj85, PHvMW93]. A.A [Mül89]. A.D [Bod84]. A.D. [C^{+}95, Lin92, Mac83]. a.o [Hen86a]. aan [Hom80, Phi67]. Aaron [FF66]. abacus [Eva77]. Abaris [Jar84b, Mor87a]. Abbé [Hom86a]. Abbey [BaJANJ84, FM00]. Abbot [Has72, Nev83, Rus79b, Sey77, Tur78a]. Aberdeen [Ant64]. Aberration [Wil72]. Abhandlungen [Nor72, Pag69]. Ablex [Fis81b]. Abominations [Blo78, Wor79]. Abraham [Bar75, Hah70, Nut84, Dau95, Hen89a, Red84]. abridgement [NR86]. abroad [Bro87]. Absolute [Bar89a, Ear89]. absolutism [Bia93]. Abstammung [Pat84]. abstract [Gra85a, Wus84]. Abstractions [Wig73]. Abu [Sar81]. Abyss [Rap86a, Ros84b]. Academic [All82, Bar76a, Buc85a, Dor80, Hom83a, Kni89e, Mol89, Mor77b, Mor89a, Rud80, Swa88, Und80, Wat73, Wes74, Yox83a, Eng83, Mor84d, Pau84]. Academics [Tan82]. Academie [Phi67, Cro88, Out80a, Str87, Stu95]. Academy [Baa90, Cha80c, Cow68, Cro70b, Cro92, Dol69, Frä89, Gil84, Mac88a, Mol81, Mor83c, Nut84, Olb82, Out83, Ped92b, RH86, Sch85b, Swe76, Tur74a, Wuj85, Hey83, Kal95a, Vuc84]. Accademia [Hop72, Mid75]. acceleration [Dra70]. accompanied [Bro87]. according [Hen91]. Account [Bar73b, Car67a, Con86, GG74a, Tur72c, Tur81, Ver89a, Ben75, Hul88, Nut82b, Sta84a, Yox83a]. Accounting [Pow96]. Achievement [Hei68, Gou95]. Achinstein [Fra87a, Wil89a]. Achromatic [Bec75]. Acid [Lar74, Ben90]. Acids [Cro69]. Ackerknecht [Und75a]. Ackerman [Bow88b]. Ackrill [Hen84d]. Acot [Ege89]. acoustics [Tur77b]. Acquaintances [RG65, GG65]. across [Pan91]. Act [Ink79]. Acta [Mer78]. actes [Ano93, BG^{+}92, Pic87c]. Action [Cla71, Rus76b, Ste91b]. active [Cal90]. Activists [Hoc88b, Kuz87]. Activity [May66, Rav67]. Acts [Wei80, BG^{+}92, Pic87c]. Actualism [Can76]. actuels [B^{+}95]. acupuncture [Agr83b]. AD [Sar81, Wil84e]. Ada [Sec88, Ste85]. Adam [Ben84, Buc89a, Dra86, Fie89, Hac88b, Hal87a, Hun86, Kei89, Nye86, Pin87b, Sec83, Stu85, The87, Tre87b, Tur83c, Tur87a, Whi84, Whi88a, Wig69, Mol89, Spe82, Tom65]. Adams [ATKL08]. Adaptation [Bow89d, Cri71]. adaptations [Sul82a]. Additional [Bry71]. Additions [Rud89, Sar87]. Address [Bro81, Hal69b, Mar64, Nee78, Wat80]. Addresses [Mol71, Mor72b]. Adelard [Mol85, BEoCF92, ECL92]. Adelmann [Und79b]. Adil [Fre86]. Adler [Fig88, Kni86c]. administration [Don93]. ado [Hen83d]. Adriaan [Hen86d, Mül89]. Adrian [Nee69, Tha84a]. Adrien [Hen89b, KW87]. Adult [For87a]. Advancement [Mor72b, Mor77a, Mor84e, Ora72, Ora76, Pat86, Pye87, Hol86]. Advent [Law80a, Pin87b, Pye85]. Adventurer [Cri66, Hal92]. advisors [Yor89]. advocate [Joy87]. Aeri [Jon87]. aeris [SSH85]. aerodynamics [HK82, Mea84]. aeronautical [Pen89]. Aeronautics [She63]. aesthetics [Ano04, WM85]. Aether [Mar74, Gol83a, Smi85b]. Aether-Drift [Mar74]. Affair [Dun86]. Affiliation [Kal95b]. Affinität [Kle94]. affinities [NPH92]. Affinity [Bro73b, Adl87, Kle94]. Africa [G^{+}09]. African [Fan83]. After [Fre78, Swa86, AS96, Can83, Cun89, Sir87, Smi85a, Wil84f]. Against [Duc88, Ait85a, JK84]. Agassi [Gow73]. Agassiz [Bar83, Win91]. Age [Bod84, Dea86b, Ink87, McG71a, Pay74, Por77, Rob89a, Smi76a, Bea85b, Ber85, BBC88, Bow89b, Bur83, Cau89, Cer89, Cor88a, Hak95, Ili95, Joy87, Kni86a, Kni88d, Kni89a, Mor84a, Tur96, Wai90, Hol78]. agency [Goo90]. agenda [Rai91]. Agent [Ger71, Kni80b]. Ages [Ano82n, Hen83d, Hos67, Kni80b, Lef78, Lor75, McV75, Mol78b, Mol81, Mon78, Nor71b, Nor77, Sch69b, Sch76, Sch80c, Sch80b, Ste75, Wea78, Wea80, Mor83d, Tur83a, Whi85b, Gol91, NW82b, Sor83]. Agnes [Yox83a]. Agnosticism [Yeo88, Lig87]. Agreeable [Bec75]. Agricola [Ber93a]. Agricultural [Far76, Coo89b, Pri83]. Agriculture [Rus77a]. Agriculturist [Hal86e, McC84]. Agrimensores [Has72]. Ahmad [Sar81, HS88]. AI [Cre93]. Aikin [Tor83]. aim [Blo84, PB83]. Aims [Hen89c, Max84]. Ainsworth [Kri83]. Air [Dau74, Jon87, PF67, Pen89, SSH85]. air-pump [SSH85]. Airs [Nut84]. Aiton [Gin74, Hal87a, Jar84b]. Aix [GG77b]. Akademgorodok [Jos97]. Akademie [Wes70]. Akademisk [Bad83]. Aktiengesellschaft [Mor84g]. Al [Ear70, HS88, Kee89, Sar81, Cro72, Gra87b, Hoc86, Kel75, LN95, Osb89b, Sch81b, AW71, AS89, vdBLW85]. Al-Biruni [Sar81]. Al-Hassan [HS88]. Al-Haytham [AW71, AS89]. al-Jazzari [Kel75]. Al-Khaledy [Ear70]. al-Khwarizmi [Gra87b, vdBLW85]. al-Razzaz [Kel75]. Al-Samarqandi [Ear70]. Alabama [Cah88, Kee89]. Alamos [Hen82b]. Alan [Dol71b, Aas87, Bro85a, Hal86a, Woo89]. Albany [Hac74, Ver89b]. Albert [Bro72a, Dug77, GG72, GG75b, Hal72b, Hei75, Hen82c, Hen88d, Kil84a, Rap71, Red84, Smi78c, Tur72a, BH87, HE^{+}82, Pai82, Sei85, SCS^{+}87]. Albertus [Col83, Sch69b, Ste75, Zam92]. Albin [Sch72]. alchemical [Fis83a, Mor91a, Per89a, RCRC98, Wey92]. Alchemie [Wey92]. alchemist [New94]. alchemistische [Wey92]. alchemists [Pat94]. Alchemy [Hal77b, Olb85, Wey92, Win75, Dob91, GGC90, Kri82, Smi94b, Gol83b, Pin83, Sch83b, Sha83a]. alchimia [GGC90]. alchimiques [Gol83b]. Alchimistes [Tur73b]. Alcmaeon [Fre74]. Alden [Fre70]. Alderotti [Cun83b]. Aldershot [Wil89b]. Aleksandr [Bro89b, FS88]. Aleotti [Kel67]. Alex [Sin84]. Alexander [Bar75, Hen86e, Jac89a, Kau84, Por83a, Aus86, Din64, Hei88, Mol88, NP77, RH86, TFC93]. Alexandra [Gur78]. Alexandre [Ano64l, For74a, Hal68c, McG66, Tur73b, Whi73, Her65a]. Alexandria [Mol88, Jon86, Lon88, VoC89]. Alfred [Bro85d, Goo85a, Hoc88a, Kee88, Sme68a, Und66a, Bro84b, Dur79b, Hug89, Smi72, WS91a]. Algarotti [Ara91]. Algebrâ [Viè83]. Algebra [Cow68, Gra87b, Whi85a, Cor96, KI92, Sin91, Viè83, vdBLW85]. algèbre [GG77b, Sin91]. algorism [Eva77]. alibi [Mol81]. Alic [Out87]. Alice [Hen81]. Alier [Fal89]. Alison [Wil85a]. alla [Sch81b]. Allan [Tur73c, Wil85a, Pin88]. Alle [Sch80c]. Alleged [Cro73]. allemand [Tur73b]. Allen [Bro87, Dur77, Gra81, Har77, Jos67, Kel76, Lay74, Nor80, Pel85a, Rob75, Sch80d, Wal81b, Whi77, Wig74, Wil69, Kni85]. allgemeinen [GG77a]. Allibone [Stu85]. Allied [Tur73c, Sto88]. alloys [Mad88]. Almagest [Mer78, Van86, PT84]. Almanac [For67a]. Almanacs [Woo80a]. Almquist [Bro84d]. Almqvist [Cha80c, Cro67, Fis81c, HS86b, Mor78, Nic88b, Rus67]. Alpacas [Kni89e, GL88]. Alpha [Bar76a]. Alphen [Hom80]. also [Hen89d, Mor69a]. Alter [Mac89]. alteration [Poo98]. alternative [Coo88]. Alternativen [Jun89]. alternatives [Jun89]. altertumswissenschaftliche [Scr68a]. altruism [Cro91a]. Altschule [Lud80]. Alvarez [Bro89e, AT87]. Álvaro [Fre90]. Alvey [OO89]. Alvi [Wuj85]. Alwyne [Kni83c]. am [AH93, Cap81b, Sch92b]. Ambiguity [Koh89]. ambiguous [Per92]. Ambix [Rat66b]. Ambroise [Bro84c]. Amedeo [Bro86b, Mor84h]. America [Bud84, Byn78, Cas82, Cor81, Gab83a, Hoc88b, Jam86, Koh78, Mac72, Mea84, Mor89a, Nut84, Pic88b, Por81, BEFH10, Doe96, Edw96, FL88, HK82, Kuz87, Lun94, Mis95, NBR^{+}95, Nye90, RR81, RT95, Seg94, Swa88, T^{+}85, War86]. American [Ait86, Bow89c, Coo83, Coo84a, Dur82, Edg87c, Edg87b, Far71a, For72, Gid82, Gil88, Gin74, GG77b, Gur78, Huf86, Kee88, Kee89, Kei84, Kel68, Kni83a, Koh78, Lev73, Mac80, Mac88a, Mol84b, Mor84b, Mor77a, Mor89a, Mor83d, Nad83, Olb82, Pel85b, Pep89, Por78b, RBM88, Rap75, Sch73b, Sch80b, Sei85, Sok87, Ste70, Tur78c, Tur83a, War89, Wea79, Web78, Whi72b, Woo83, Woo73, Wor88, Bru87a, Cal77, Deg91, Gos87, Gur78, HL86, Kal95b, Koh78, Lie87, Liv87, Mai91, Mar62a, MR89, Mit92, MM10, Mor77a, Mor89b, Mos77a, Mos77b, Moy83, New94, Ost84, Pic76, Pit93, Rai91, Rei85, SB92, Web78, Who82, Wri94]. Americana [Coo83, Duf83, Gli83b, Kni83a, Nad83]. Americans [Far76]. Americas [Hac82b]. Amersham [Mor83a]. Ames [Whi89b]. Ami [Rav67]. Among [Mor87b, Rud85, Van93]. Ampère [Hen84b, Blo82a, Can80a, GG65, Hof87, RG65]. Amrine [Sha89]. AMST [Duf75]. Amsterdam [De 85, Dor79, Eyl68, Fis74, Fra87c, Fre82, Gra86a, Hal84a, Hea64, Nev78, Rat66a, Sme70, Wil87d]. Analecta [Cla70]. Analogy [Got67]. Analyseos [Whi85a, Viè83]. Analysis [Dub72, GG75a, Smi87c, Dau95, Ham90, Hea86, Hea90, Pum89]. analyst [BJB92]. Analytic [Whi85a, Kni83b, Viè83]. Analytical [Kni67, Sme69, Sha98]. Analytics [Sch78a, Wal92]. anatomical [BW90, Cun97, RT92]. anatomized [Tha83a]. Anatomy [Lev73, Lon88, Ste67, Arm83, Law85, OS92b]. ancestors [Des82a, Des82b, Des84, Tha84a]. Ancestry [Bar73b]. Ancient [Bar94, Cla99a, Cla99b, Cla95a, Cla9 , Cla2 , Fre74, Fre94, Fur85, Lor77, Mol78a, Phi69, CK90, HW87, Hen84j, Kre82, Llo83, Pen89]. ancients [Cun97]. Anders [Fis81c]. Anderson [Fis81a, Hen89g, Ped85, Sha86]. André [Fis83c, Tur78b, Fox87b, GG65, Yox83a, RG65]. Andre-Marie [GG65, RG65]. Andreas [Rat65a]. Andrew [Hen86b, Law85, Wal81b, Woo84]. Andrews [Bro83b, Hom86b, Mor69a, Tha83b, Und80, Woo80b]. Anfänge [Gol91]. Anfängen [Rav65]. Angeles [Bod84, Bow88a, Bro83b, Bro86d, Can80c, Grm69, Hal84d, Har89a, Hol71, Hom86b, Law80c, Nor77, Olb69, Out83, Por79b, Rat65a, Rus87b, Sei85, Sin76, Whi77, Woo80b]. Angeles/London [Har89a]. Angeli [Mor89c]. Angelic [Mol80]. anglaise [Pet84]. angles [Cha90c]. Anglican [Bro89c, Cor88b]. Anglo [Dur82]. Anglo-American [Dur82]. angular [Cha94]. Angus [For66, Nor68b, Sme78b]. Anhang [Whi68a]. Animal [Bod65, Cro91b, Geo87, Mau77, Ost89, BW90, Nyh95, Per92, Rus86b]. Animals [Asq80, Mor82b, Nut84, Ste91b]. Anita [Ben84, Nut82b, Tur81]. Ann [Har75, Mol74, Rav72, Roc88b, Kni88a]. Anna [Hen88d]. Annales [Gel76]. Anne [Gua89, Tan82, Tei83a, Whi73, CMNH92]. Anne-Marie [Gua89]. années [Cas85]. Annette [Por78b, Har86]. annexes [CAMR87, CLR87]. anniversary [AL87, Hen89g, PS82, Por85a, vLP82]. annotated [Ant64, Coo89c, DeV82, Hen83b, Ogi86, Por83b, Tha85]. Annotations [Rat66b]. annotée [McC80, McC83]. Announcement [Ano77a]. Annual [Arm62, Arm64, Mor83a, Sin77, Smi66, Mor77b, Wei92]. annum [Wea83]. Annus [Ait73]. Anomaly [Hof87]. Anonymous [Per70]. ant [Cro91a]. Antarctic [Fog92]. ante [Gli84]. Antenore [Mol72b, Sch73a, Sch80c]. anthology [NBR^{+}95, WS91b, WS91a]. Anthony [Tur77a, Wea78, Hut88, Rup85]. Anthropological [Byn87, ME81, Ver85]. Anthropologie [Jor88, MD90, Atr86]. Anthropology [Bow86, MD90, Atr86, Cru90, Kuk91, Kuk93, Sto84, Web84]. Anti [Nor85b, Wil87b, Bow83b, Hac84]. Anti-Darwinian [Nor85b, Bow83b]. Anti-Submarine [Wil87b, Hac84]. Antibiosis [Fri74b]. antibiotics [Wai90]. Anticipations [KK87, Woo89]. antievolution [BN95, RT95, RD95]. Antievolutionism [Num95a]. antiken [Bos72]. Antikythera [Smi78b]. Antique [Fox82, Wil77, Law80c]. antiques [SB92]. Antiquities [Nut83]. Antiquity [Ant63, Got71, Grm69, Mol72a, Out87, Phi69, Scr68a, Was69, Wil69, Ali86, Cro90b, Gre92a, O’M89, Ogi86, Rai91, Sor83, Whi85b, Bro84c, Dyb78, Hen84j, Hen84k, Hut84, Lor77, Mer78, Mol78a, Mol84c, Nut84, Sch78a, Sim84, Whi90, Wil77, Smi78b]. Antoine [Cro71, Sch83b, Don93, Poi93]. Anton [vBDGO93]. Antoni [Hal84a, PS82, vLP82]. Antonia [Smi74a]. Antonino [Sch73a]. Antony [Hig65]. Antwerp [Rat66b]. Anwendungen [Sch89]. Anziehungskraft [Adl87, Fig88]. Ape [CT95, CT95]. apeman [CT95]. Apian [Nor66]. Apologetics [Woo80d]. Apologia [JK84]. Apology [Din77]. Apostle [Got71]. Apothecarie [Hal88a]. Apothecaries [Vea64]. apparantium [Rav69]. Apparatus [Hof87, Tur70b, Agr83b]. apparitione [Rav69]. Appendix [Mor77a, Und75a, vLvHH^{+}88]. Appleton [Chi64, May69]. Appleton-Century-Crofts [May69]. applicability [Lin82]. Application [Hom74, Gra84a, Mol68]. Applications [Fau89, Fox70, Hen88c, Kni81, Sch89, WW86]. Applied [Col76, Hac76]. Appraisal [Cla66, Gow78, Kni85]. Appraisals [Kni81]. Appreciating [Rob86]. Appreciation [Wil62]. Apprentice [Fox74]. Apprenticeship [Ben85]. apprenticeships [WW88a]. Approach [Tre79, Gie88]. Approaches [Bur87b, Goo88a, Kni89b, Ner87]. Approaching [Mic95]. appropriated [Wal92]. Appropriating [Fri89]. approx [Gra83b]. approximate [Tee89]. après [Gay92]. April [Ano64l, Ano93, HWW86, LB73, Sme88, Woo80c, EJR94, EJR95]. Arab [Lor73]. Arabic [Ear70, Fre86, Wuj85, Dun90, RM96, Win75]. Arabo [Hos67]. Arabo-Latin [Hos67]. Aran [Pau76]. Arbeiten [Hom74]. Arbor [Har75, Mol74, Rav72]. Arch [Rog88a, Old86]. Archaeoastronomers [Kno85]. archaeological [Tri89]. Archaeology [Fig77b, MMVZ90, Gut89, PW89a]. archeology [CT93]. Archer [Mor81]. Archetypes [Des82a, Des82b, Des84, Tha84a]. Archimede [Sim84]. Archimedes [Sch80b, Hos67, Sch80b, Sim88a]. Architectural [Far77]. Architecture [Sch78b, Tur78a, For86, MB^{+}08, Tur78a]. Archival [Olb82]. Archive [Ano77m, PS78b, Sch75, WG81]. Archives [Can74, Col88a, Dou84, Gil84, Han85a, Sch84b, Stu87, BKH90, GG86a, GG86b, PLP92, TW87]. Archon [Blu79, Swe76]. Arctic [Ora77, Lev88b, Lev93]. arcu [Rav69]. Arcueil [Sme69, Har67]. Areas [FB75, Old77]. argent [Poi04]. Argument [Hod77, Blo84, Dea91, PB82].

Argumentation [Got67]. Arguments [Sha87, Des89b]. Arising [Lar74]. Aristarchus [Wil87c, Van85]. Aristot [Mol72b]. Aristot-elianism [Mol72b]. Aristotelean [Eva89, Fur88]. Aristotelianism [Hen88e, Bru87b]. aristotelica [Sch73a]. Aristotle [Dyb78, Got71, Hut84, Jud88, Nut84, Sch78a, Dun90, Eva87, Fre95, Liv85, McK78, Mor82b, Sch83c, Wal92, War90]. arithmetic [Eva77]. Arithmetical [LY66, Mah89, Scr68b, Edw87, yL66]. arithmetice [BJ91]. Arlautic [Fie83]. Armament [Swi71]. Armand [Hal85c, McC80, McC83]. Armed [Wil84e, Mac83]. Armitage [For66, Nor68b, Sme78b]. Armstrong [Bib74, Law85]. Army [Kni83a]. Arno [Cas82]. Arnold [Hei71a, Kel76, Kni74, Law80a]. Arrhenius [Cra96]. Arrow [Olb71b]. Art [Cla86a, Con86, Fie86, Fie89, Poi04, Sch70, Whi85a, Bea91, BB84, BW90, Cro96, Edg91, Fie97, Gil83, Hal65c, HA87, PS91, Sta84a, Sta91, Tom84, Viè83, VoC89]. Arte [Gou84b, Goo89b, Lan86]. Artful [Sta94]. Arthur [Car74, Mor84d, Nee69, Sho88, Whi68b, Woo83, BN95, Cro89a, Fra78b, Tor83]. Arti [Sch81b]. Articles [Dyb78]. Artifacts [Jen87]. artificial [Cre93]. Artillery [Hom74]. Artis [Mor83d, BJ91]. Artisan [Kar76]. Arts [Hut88, Kni75, Olb82, Por84b, Swe76, GJ86, Whi90, Sch85b]. Artz [Cro68a]. Arzneimittellehre [Ree85]. Asa [Web65a]. Asberg [Rau75]. Aschendorff [Mol86]. Asclepius [Cop92]. ascribed [Mol85]. Ash [Sti89]. Ashford [Dra86]. Ashmole [Web68a]. Ashworth [Fre78, Had63, Han70]. Asia [Agr83a, Mar10]. Asian [Kni86c, Nak75, Nee69, Ald85]. Asimov [Wil65a]. Aspect [Lay65]. Aspects [Hal68b, Her66, Stu78, Buc85d, Lin93, Whi68c].

Assen [Sme68b]. Assessments [Kni74].

Assessorato [Sch81b]. assistance [Ben79, Cla67, GG75a, Hei73, Hei76b, Nor68a, Nor69a, Nor70a, Nor73a, Pep72, Sme66, Whi73]. assisted [Rav67, Tur70b, Vic88]. associate [Hut89]. Associates [Fie86]. Association [Goo74, Gow71, Mor72b, Mor77a, Mor84e, Ora72, Ora76, Pat86, Sol83, Fis83c, Hac82b, Mus74, Rus77b, Smi89e, Tup82]. Astrarium [Kel68]. Astrologi [Bra87, Zam86]. astrologia [GGC90]. Astrological [Lor73]. Astrology [Jev64a, Jev64b, Bar94, GGC90, Kit89, Smo94, Zam92, Woo80a]. Astronautical [Mac88a]. Astronautics [Mac88a]. astronomer [FFLW97, How89, Tal62]. astronomiae} [Zam92]. Astronomical [DTS^{+}87, For74a, Hun86, Kri88, STT87, Smi87c, Smi89c, Swe76, Tay87, TST87, Ger89, Hea86, Hea90, Mur85]. Astronomie [Ben89, For67b, MdLD86, Pic87c]. astronomisches [Ger89]. Astronomy [Arm70, Bra88, BH84b, Cla62, For71a, Hal69a, Hom74, Mol78a, Mol80, Nee85, Rav70, Rya70, Sti70, Stu71a, Swi71, Whi89b, Wil86b, Ben87, Cha94, Cla95a, Cla95b, DeV82, Doe96, Edg84, EBT96, For73, Hen83b, HK84, Het88, Hir83, Hon89, Nor94, Pic87c, Rob92, Sch80a, Smi82b, Ste87, Ste94, Sul82b, TW95, Kul63, Kul64, Arm69, Arm70, Bar76a, Bry78a, For69a, Rav69, Rav70, Sey77, Sti70, Swe76, Tho76a, Tho76b, Til76, McC69]. Astrophysicist [Huf86, Cha83, Huf85, Ost84]. Astrophysics [Huf86, DeV82, Hen83b, Ost84, TW95]. Asylum [Cla86a, Tom84]. Asylum-Keeping [Cla86a, Tom84]. Athens [Hal84e]. Athlone [Byn73, Chi64, Pic79, Wea78]. Atlantic [Ben82b, Cha77a, Fie81, Eri94]. Atlantis [Jos97]. Atlas [Kni88c]. Atmospheres [Hom72]. Atom [Bad89, Far68b, Tha67, HA84, Wil87e, Stu85].

Atomic [Far67, Far71b, Fis74, Fuj86, Har81, Jon75b, Kel83, Sin77, Sin84, Smy89, Tha66b, War88a, Fis81c, Kra84a]. Atomism [Bro85a, Hen82d, McG68, Pac70, Roc84]. atomist [Joy87]. Atomists [Bro69, Sch68b]. Atoms [Far68a, Hei71a, Osl91]. atomteorin [Fis81c]. atop [Dra81, Nay85]. Atran [Jor88, Osb89b]. Attacks [Per70]. Attempt [Hum67]. attitude [Por84a]. attitudes [DH90, Maz95a, Ros91]. Attractions [Dun69a, Sme71]. attractive [Adl87]. attributed [Per89a]. Auctions [All77]. Audience [Gol88a, Rap86b, Tor83]. Auflage [Wes70]. Auftrag [Wes70]. Auftreten [Rav65]. augmentée [GG70]. August [Ano67i, Bro89a, Frä90, Nee78, Sme79, Str73]. Augustan [Por79b]. Augustus [Scr68b]. Ault [Hei77]. auspices [Smy89]. Aussagen [Gro89]. Austin [Fie89, Hen82c, Mar74, Por82, Tur81]. Australasia [Cor81]. Australia [Kni88a, Fin93, Moy86]. Australian [Dol69, Hom88, Nor89, Old77]. Author [Bry78c, Ede65]. Authority [Whi89a, Cro85, Out84c, Red87, Sto88]. Authors [Ano88o, Ano89p, Roc88b]. Autobiographical [Web68a, MR89]. Autobiographies [Bar76b, Rus88d]. Autobiography [Kau84, Kni69b, GH84, GH96, Hal85a, Tod83, MB^{+}08]. autograph [Til75a]. Automata [Kel67]. Automation [Bur87a]. autour [B^{+}95, BD95]. Auvergne [Rud74]. aux [Cas85, GG77b]. Available [Kni89c, Bry78a, For87a, Hen88d]. avait [Lan87, Out88b]. Avebury [Mor83a, U^{+}91]. Avery [Rus88c]. Aviation [Car84, Gil83]. Avicenna [Cun89, Sir87]. Avogadro [Bro86b, Mor84h]. avoiding [CGBE95]. avril [Ano93]. avril-4 [Ano93]. awareness [Whi88b]. Axial [Tur74b]. axiomatischen [GG81]. Ayala [Smi76c].

B [All81, Bow83a, Bow87c, Bro89b, Buc89a, Cro68a, Fig77a, Fra87c, Fre70, Gow71, Gra86a, Gre71, Her80, Hut84, Hut89, Kra86, Lar74, McC76, Men70, Mol72b, Mos77b, Nor68b, Nut83, Out83, Por89b, Pra70, Roc88b, Rud80, Sch74a, Sme68a, Smi78a, Smi87c, Sne74, Tho76b, Tom65, Und79b, Wea79, Wel70, Wil86b]. B. [Gra87b]. B.C [Bod84, C^{+}95]. B.C. [Lin92, Lon88, RL97, Smi78b]. Babbage [Car69, Twe89, BCKR89, GG79a, Lin90, Rei68]. Babylonians [Cla62]. Bachelard [Out80b]. Back [Ano62b, Ano63a, Ano63c, Ano64a, Ano64c, Ano65a, Ano65c, Ano66a, Ano66c, Ano67a, Ano67c, Ano68a, Ano68c, Ano69a, Ano69c, Ano70a, Ano70c, Ano71a, Ano71c, Ano72a, Ano72c, Ano73a, Ano73c, Ano74a, Ano74e, Ano74c, Ano75a, Ano75c, Ano75e, Ano76a, Ano76c, Ano76e, Ano77b, Ano77d, Ano77f, Ano78a, Ano78c, Ano78e, Ano79a, Ano79c, Ano79e, Ano80a, Ano80c, Ano80e, Ano81a, Ano81c, Ano81e, Ano82a, Ano82c, Ano82e, Ano83a, Ano83c, Ano83e, Ano84a, Ano84c, Ano84e, Ano85a, Ano85c, Ano85e, Ano86a, Ano86c, Ano86e, Ano87a, Ano87c, Ano87e, Ano87g, Ano88a, Ano88c, Ano88e, Ano88g, Ano89a, Ano89c, Ano89e, Ano89g, Bea85a, BH84a, Cap81a, Fre81a, Geo78a, Gli83a, Hea87a, Kni71a, Lin89a, Mol84a, Nut82a, PS78a]. Back [Sme78a, Tur70a, Roc90]. Background [Ait67, Fox68a, Kni87, Mül87, Rav70, Ste81, Can93b, Gow78, McC81, McI85, Mol68, Roc81a, Wal92, Whi68c]. backwardness [Bro82a]. Bacon [Gol88b, Por83a, Sch85a, Vic88, BL83, Lin82, Mar92, PR88, Vic87]. Baconian [Smi94a]. Bad [GG77a, Rus72b]. Badash [Hei71b, Hen82b, Bro86c]. Baden [Bro89c, Cor88b]. Baer [BEFH10, vBDGO93]. Baeriana [Geo78b]. Baglivi [Und75b]. Bailey [Eyl68]. Baker [Ged71, Hai77, Kra86]. Bakken [OS92a]. Balaguer [Mal95]. balance [BD95, HE92]. balances [Bry81]. Baldini [Sch79b]. Baliani [Ait68]. Balkema [Mül89, Smi89b].

Ball [Mos77b]. Ballad [Ste81]. Ballistics [Ros68]. ballooning [Gil83]. Baltimore [Bow74, Bro71a, Buc75, Cal77, Chi68, Chi71, Clo80, Coo83, Gel76, Gur78, Hal80a, Hop72, Jac85, Kni76, Law81, Mor83b, Mor89a, Nor85b, Pic76, Pic80, Pin83, Por85a, Ree80, Ros79, Sch70, Sha86, Web78, Wil89a, Yeo88, KT87]. Baltzer [Olb69]. Band [Bos72, Bro82a, Mol72a, Mol75, Mol86, Rav65]. bang [SN93]. bank [Hag92]. Banks [Nev83]. Bannister [Dur82]. Baptista [Hen84k, Pag82]. Baptiste [Ait68, Fra78a, Stu86]. Barbara [Hen84l, Rus67, Tur70b, Con86]. Barber [Fre81b, Kni86b, Smi87a]. Barber-Surgeon [Fre81b]. Barber-Surgeons [Fre81b]. Barbèra [Hat87]. Barbers [Fre81b]. Barcelona [Goo89b]. Barnard [She95]. Barnes [Bus83, Dyb78, Goo81a, Jar83, Mor77c, Rus79b, Sch78a, Wea84]. Barocas [Stu71a]. Barometer [Spa65]. Baron [Boy70, Smi78a]. Baroque [Eri94]. Barraclough [Cla86b]. Barret [Des88]. Barrow [Hal88a, Ora73, Whi72a, Fei90]. Barry [Goo81a, Mor77c, Hen88e]. Barthélemy [Out84a]. Barthélemy-Madaule [Out84a]. Bartram [All84a, BB82]. Barzun [McC77a]. Basalla [Mor72b]. Based [Joh79, BDH89, EK88, EK92, Gun84, May66, Mol71, Wil85a]. Basel [Gra85b, Gre88, Kra86, Mah89, Ser88, Whi68c]. Basel-Stadt [Gra85b]. Basel/New [Whi68c]. Basic [Fin68]. Basil [Bro69, Fal89, Har83a, Hen89c, Liv86, Rob89a, Wil84e, You83]. Basin [Tee88]. Basingstoke [Can86, Rud83, Woo80e]. Basis [Dol73a, GG81, Sch92b]. Bates [Men67]. Bateson [Nor82, Olb87]. Bath [Ben79, Mol85]. Baton [Col88b]. Batsford [Bry74, Wal74]. Battersby [Eyl68]. Bau [For69a]. bauer [Bos72, Nor89]. Baumeister [Sch92b]. Bäumler [Wei86]. Baustatik [Sch89]. Bayerischen [Wes70]. Bayet [Out80a]. BBC [Yox82b]. BBC-2 [Yox82b]. BBC2 [Yox79]. Bd [GG77a, HP85, Hom74, LM23, Mah89, Scr68a]. Be [Bra89a, Hey95, F^{+}88, Goo89a, Hal69b, Moo77, Mos77b, Roc88b]. Beagle [Bur75, DK88, Gru69]. ‘Beame [Ost89]. Bearman [Olb82]. Beaude [Sch72]. Beaulieu [Hal85c, McC80, McC83]. Beaumarchais [vPvP90]. Beaumont [The78]. Because [Bea91]. Beche [Nev80]. Bechler [Hen84e]. Bechmann [Tee89]. Beck [Fig88, Hen88d]. Beckenham [HS86a, Por88]. Beckinsale [Hig65]. Beck’sche [Wes70]. Beddoes [Por86a, Sta84b, Lev84]. Bedeutung [Fri91, GG81]. Bedini [Kel68, Tur78c]. Bedlam [Sch80a]. bedside [Coo83]. Beecher [Lud80]. Beer [Bar76b, Hod70]. Before [Fei90, FC96, Gos87, Jac85, Por89d, War89, Bar84, Bro84b, Bro85d, Cla86b, Deb66, Edw78, Ger89, Hig65, Hod89a, Mor82a, NC81, Num95a, Nut82b, Rus76b, Rus77a, Ste67, Sti86, Web88]. Beginning [Por85b, Rec86, Mad88]. Beginnings [Bod84, Rus88b, Cra84, Gol91, Hen82a, Lin92]. Begriffe [GG77a]. Begründung [Pag69]. behalf [Hei73]. behavior [Gra87a]. Behaviour [Asq80, Ric82, RH88]. behavioural [Cou96]. Behram [Hen89b]. Being [Dol71a, Whi72a, Win75, Gra81]. Beitrag [For69b, Pag69]. Beiträge [Bro82a, Cap81b, Gra85b, Nor71a]. Beiträgen [Gol91]. belge [Tan82]. Belgian [Phi67]. Belgie [Phi67]. Belgium [Gol83b]. Belief [BM76, Duf75, Liv86, Rus77b, Dur85, Hoo97]. Beliefs [Yat72]. Believe [Por85b]. Belin [Cro88, Out88a, Out88b]. Belknap [Hod89a]. Bell [Edg87c, Lar74, Mor72b, Rei85]. Belle [Mor89c, Sch81b]. Belles [Wil77]. Bellone [Smi82a]. Belonged [Far63]. Beloved [Ste62a]. Below [Bry88, Por89c]. Belvey [Din63a]. Ben [Rav67, Rav73, Swe76]. Ben-Ami [Rav67]. Ben-David [Rav73].

Benett [Hen83a]. Bengal [Bal91]. Benguigui [Hom86a]. Beniger [Hac89]. Benison [Cla68]. Benjamin [Bry78c, Moo85, Bro89d, Cro84b, Mil73, Fox69b]. Bennett [Hal88a, Nee69]. Bennion [Law80c]. Benoît [Rap71]. Benson [Bow89c]. Bentham [dCC09]. Bentley [Gid82]. Berg [Fie81, GG77a, Ink87, Mac89, Por89a]. Bergman [Cro67, Dun69a, Sme71]. Berkeley [All84a, Bod84, Bow88a, Bow89b, Bro86d, Can80c, Dea86b, Fre86, Grm69, Hal84d, Har89a, Hen83b, Kra86, Law80c, McC85b, Nor77, Olb69, Out83, Por79b, Rat65a, Rid83, RH86, Rus87b, San87, Sin76, Smi87b, Tre88, Whi77, Ath90]. Berkeley/Los [Har89a]. Berkley [Hol71]. Berkson [Goo75]. Berlin [Bos72, Bra88, Bra87, Bro89b, For69b, GG75c, GG81, Gra83b, Gra86b, Gra87b, Hal85d, Har81, Hen86b, Hen86a, Hen88c, Kra86, Von88, Mol72a, Mol78a, Mol88, Mor84g, Nor71a, Nor72, Rav65, Sch75, Sch83b, Scr68a, Sme68a, Tur88, HRP90]. Berlin-Heidelberg-New [Gra83b]. Berliner [Bro82a]. Bermannus [AHY90]. Bern [Pac71]. Bernal [Rav66b]. Bernard [Ait72, Bow87a, Byn74, Cha88a, GG77a, Hal77a, Hen86f, Jar84b, Lor75, Mau77, Swe76, Til76, Whi73, Wil86a, Woo80a, Bro85b, GGSGB94, Yeo88]. Berne [Mol75]. Bernet [All77, Law80c, Wal77]. Bernhard [For69a, Goo85a, Rus74]. Bernoulli [PF67]. Bert [Lin86]. Berthollet [Cro79, Sme84a, Fuj86, SG80]. Bertram [Swi67]. Bertrand [Bro72a]. Berzelius [Fis81c, Rus67, Fis81c, Frä76, MF92]. Besoin [Out88b, Lan87]. besorgt [Wes70]. Bessemer [Cla86b, Bar84]. Best [Ste75]. Bethesda [Sch68a]. Betty [Hal77b]. Between [And84, Bro75, Cha77a, Gil84, Hai78, Hey83, Mil83, Mor87a, Sha86, Ano90, BM82, BaJANJ84, Ben82b, Ber95, Bro72a, Clu88, Gre86, Hen91, Hen84l, Lan91, LN86, NPH92, Out84a, Pac70, Pag76, Rap69, Rap71, Ros86, Rus87b, Sch81a, Sch89, Sha83b, STW90, BBC88]. Beverly [Goo81a, Irw82, Irw83a, Sha79]. Beyond [Ald85, Aus86, Cer89, Hoc88b, Kni86c, Kuz87, Rob92, SG90, Sme66, Wil84a, FS84, Tay96]. Beziehung [Sch89]. Bianchini [Ger89, Ger89]. Biarnais [Sch84c]. Bibby [Fri74a]. Bible [Har98b]. Biblical [Cun78]. Bibliofilia [Gou84b]. bibliog [Hei68]. Bibliographia [Wes70]. bibliographical [BW90]. Bibliographies [Ano82n, Bro73a, Bro87, Fox79, NW82b, NW82a, Ano82m]. Bibliography [Ano82n, Ano82m, Ant64, Bro73a, Cha78b, Cha78a, Fox79, Fre82, Har79a, Kni83c, Kra86, LM74, Mol72b, Rud83, Rud89, Sha73b, Tha66a, Web69a, Bud84, DeV82, Fis83b, Fru82, Hen83b, Hen84g, Le 90, Mor84b, NW82b, NW82a, Ogi86, Por83b, Rot93, Sar87, Tha85, Wuj85, Blu79, Che78, For80, Fox79, Har79a, Olb82]. Bibliophily [Blu79, Fox79, Har79a]. Bibliotheca [Mer78]. Bibliotheque [Hen84b, Bur74a]. bicentenaire [BG^{+}92]. Bicentenary [For67a, Kni89e, Wat68, BG^{+}92, GS88, WBP97]. Bicentennial [Por81, Way87]. Bielefeld [JO81]. Bienowska [Tho76b]. Big [GH92, SN93]. Biggs [Tor87]. Bijvoet [Tur70c, Tur73d]. Bildung [Dau98, For69a]. Bills [Sha73b]. bio [Fru82, Mor84b]. bio-bibliography [Fru82, Mor84b]. Biobibliography [Eag78, Fau89, WW86, Coo89c]. biochemical [Fru82, Fru90, Mor84b]. Biochemist [Tei83a]. Biochemistry [Har81, Kau84, Mor84b, Mor84f, Olb71a, Pni82, Sme84a, Tei84, Olb82]. Biographical [DM96, Kni83a, Law89, Wig69, Wil69, Bry83, Büh81, Gra83b, O’C88, Ogi86, Web68a]. Biographie [Mai85]. Biography [Bow87a, Bro89b, GG72, Hal87a, Hei75, Jor83a, Mor72a, Pep86, Ait85b, Car89, FS88, FIF86, Kra90, Lev85, Mor84h, Pen89, Rog83, Shi83, SY96, Spe82, TC90, Wal91, Wal90, Cun83a, For82, Gid82, McN83, Nev83, Out83, Por82, Rog83]. Biologen [Dei92]. Biological [Cla67, Geo66, Gin84, Hai78, Koh89, Mil89, Slo79, Smi66, Ste62c, Und66a, Web66a, Web67a, Whi68c, You66, Ben82b, Grm90, NLH^{+}84, Cla71, Cri71, Olb71a, Sch71]. biologie [CM93, Hai80, MC^{+}93]. biologique [Grm90]. Biologist [Olb69, WV92, Wei91]. biologiste [Cla70]. Biologists [Dei92]. Biology [Bar89c, Bow87b, Bow87c, Bow89c, Bro85e, Cro66a, Dav71, Gel76, Har83b, HWW86, Jor79, Lin74, Men65, NLH^{+}84, Nyh95, Olb71b, Ric82, Rud71, Rus88c, Rus88d, Slo79, Smi73a, Smi76c, Stu88, Tei78, Und66c, Und66b, Und80, Ver89b, Wil65a, Yox78, Bow83a, Kni84a, Kni84b, Len82, Mai91, MC^{+}93, Olb82, Pro86, Reh83, Rus88a, Sap90, Yox83a, Cla70, Fre70, Grm69, Hai80, Hod70, Olb69, Olb70, Poy70, Ree80, RBM88]. Biomedical [Byn87, Irw83a, Tei84, Ver85]. Biometric [Olb89]. Bionomica [Bro89a]. Biot [Fra78a]. Biotechnology [Hen87b, Ken86, Nor85a, Ben90, Bud93, Yox83c, Bau95]. Birch [For85a, Gun84]. Bird [Hea87b]. birefringent [Bar91]. Birkhauser [Ser88, Gra85b, Gre88, Kra86, Mah89]. Birkhoff [GG75a]. Birmingham [Fox68b, Tup82, Wil65b]. Birnie [CHG91]. Birth [BH83, Cla86b, Fig77b, Ait85a, AL87, Cas85, Hen89g, JK84]. birthday [vLP82, PS82]. Biruni [Sar81]. bis [Gol91, Rav65]. Bjerknes [Fri89]. BJH [Ano62a, Ano62b, Ano62c, Ano63a, Ano63b, Ano63c, Ano63d, Ano64a, Ano64b, Ano64c, Ano64d, Ano65a, Ano65b, Ano65c, Ano65d, Ano66a, Ano66b, Ano66c, Ano66d, Ano67a, Ano67b, Ano67c, Ano67d, Ano68a, Ano68b, Ano68c, Ano68d, Ano69a, Ano69b, Ano69c, Ano69d, Ano70a, Ano70b, Ano70c, Ano70d, Ano71a, Ano71b, Ano71c, Ano71d, Ano72a, Ano72b, Ano72c, Ano72d, Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74e, Ano74b, Ano74f, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano80a]. BJH [Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h]. BJHS [Ano88p, Ano88q]. Black [Hal88a, Hei76a, Lar74, Mor71, Wig69, For71a]. black-and-white [For71a]. Blacker [Bal84, Vig82]. Blackwell [Dol70a, Fal89, Fre81b, Har83a, Hen89c, Liv86, Rob89a, Smi89c, Wil84e, You83]. Blagrave [Nor66]. Blaise [Mah89, Loe87]. Blake [Hei77]. Blanchard [Bur74b, Dug77, GG75b, Hal72b, Jor83a, Tur72a]. Blanche [Edw78]. Blandford [Fox82]. Blatt [Fie86]. Blay [Sch86b]. blessings [Seg94]. Bliss [Lie86]. Blister [Cla86b]. Bloch [Sch74a]. Blond [Tha84a]. Blondel [Hen84b]. Blood [Sch73b, Und66a, WKB88]. Bloomington [Buc66, Dol74, Got71, GG84a]. Bloor [Joh78, Wor79]. Blum [Kra86, Olb71b]. Blut [WKB88]. Board [Bur75, Ano88q]. Boas [Buc66, Hop73, Hop75, Rob87, Web66c, Web69b, Woo80c]. Bob [Fie83]. Bochum [Bro82a]. Bodies [Hod71, HM66, Hum67, Nay85, Sin91, Dra69, Sha93, Dra81]. Bodily [Fre86]. Bodleian [Scr64b]. Body [Sch69a, Sta91, Yox79, Arm83, Law85, RT92].

Boehme [PW89b]. Boerhaave [Fre78, Han70, Wil87d, Ano68e]. Boethius [Pag69, Whi68a]. Bohemica [Rav69]. Bohme [Hoc86]. Bohr [Bad83, Che85, Dor79, Hen86b, Hen86a, HP85, PHvMW93, PvM85, Aas90, Hen84i, HP85, MR82f, MR82g, PHvMW93, PvM85]. Bokhara [Kni86c, Ald85]. Boltwood [Hei71b]. Bolzano [GG77a]. bomb [Pou93, Smy89]. Bonaventure [McV75]. Bond [Koh75]. Bone [Wel67]. Bonelli [Tur72b]. bones [Bry92]. Bonnet [Duc88, Ped85, And82]. Book [Aas87, Agr83a, Agr83b, Ait67, Ait68, Ait71, Ait72, Ait73, Ait85a, Ait86, Ait89, All77, All79, All81, All82, All84a, All84b, Ano87i, Ant63, Ant64, Arm62, Arm64, Arm66, Arm67, Arm69, Arm70, Asp88, Asq80, Aus86, Bad71, Bad83, Bad89, Bal83, Bal84, Bar77, Bar79, Bar81, Bar83, Bar73a, Bar76a, Bar75, Bar76b, Bea85b, Ben79, Ben84, Ben89, Ben82b, Bib74, Blo79, Blo82b, Blo84, Blu79, Bod65, Bod84, Bos70, Bos72, Bow74, Bow87a, Bow89a, Bow83a, Bow85, Bow86, Bow87b, Bow87c, Bow88a, Bow88b, Bow89b, Bow89c, Boy70, Bra88, Bra84, Bra87, Bra89a, Bro67, Bro69, Bro71a, Bro73a, Bro82a, Bro84a, Bro85a, Bro86a, Bro86b, Bro89a, Bro89b, Bro80b, Bro72a, Bro76, Bro89c, Bro72b, Bro73b, Bro75, Bro82b, Bro78a, Bro78b, Bro86c, Bro89e, Bro83b, Bro83a, Bro84c]. Book [Bro84d, BH84b, Bro85d, Bro85e, Bro86d, Bro86e, Bro87, Bro89f, Bru72, Bru86, Bry74, Bry76, Bry77, Bry78a, Bry78b, Bry78c, Bry83, Buc77, Buc69, Buc71, Buc73a, Buc75, Buc85a, Buc85b, Buc66, Buc68, Buc89a, Buc73b, Bud84, Bur80, Bur74a, Bur74b, Bur71, Bus83, Bus85, Bus74, Byn72, Byn73, Byn74, Byn78, Byn80, Byn87, Byn88, Cah88, Cal63, Cal77, Can74, Can75, Can77, Can80b, Can80c, Can84, Can86, Cap81b, Car67a, Car69, Car70, Car84, Car74, Cas82, Cas85, Cha65, Cha77a, Cha80b, Cha77b, Cha80c, Che78, Che85, Chi64, Chi65b, Chi65c, Chi65a, Chi68, Chi71, Cla86a, Cla66, Cla67, Cla68, Cla70, Cla71, Cla86b, Clo80, Coc75, Col81a, Col83, Col76, Col87b, Col88a, Col88b, Col81b, Con86, Coo81, Coo83, Coo89a, Coo84a, Cor81, Cow74, Cow68, Cri66]. Book [Cri71, Cro66b, Cro67, Cro68a, Cro68b, Cro69, Cro70a, Cro70b, Cro71, Cro79, Cro85, Cro87a, Cro88, Cro89a, Cro89b, Cro72, Cun78, Cun83a, Cun83b, Cun88, Cun89, Dau75, Dav71, Dav81, Dea83a, Dea83b, Dea86a, Dea86b, Dea86c, Deb66, Des88, Din63a, Din64, Dol69, Dol70a, Dol70b, Dol71a, Dol71b, Dol73a, Dol73b, Dol74, Dol83, Dor75, Dor79, Dor80, Dou84, Dra86, Dub72, Duc84, Duc88, Duf81, Duf83, Duf75, Dug77, Dun69a, Dun69b, Dun86, Dun87, Dur77, Dur79a, Dur80a, Dur81, Dur82, Dur83, Dyb78, Eag78, Ear70, Ear87, Eas86, Ede65, Edg84, Edg87a, Edg87c, Edg87b, Edw78, Ege89, Eme87, Eva89, Eyl68, Fal89, FF66, Far67, Far68a, Far69, Far71a, Far71b, Far76, Far77, Fau89, Fee77, Fel89, Fie81, Fie83, Fie86, Fie89, Fig88, Fig77a, Fig77b]. Book [Fig78, Fin68, Fis74, Fis78, Fis81a, Fis81b, Fis83b, Fis83a, Fis83c, For66, For67b, For68, For69a, For69b, For71b, For71a, For72, For74a, For80, For85a, For87a, For82, For87b, For87c, Fox68b, Fox69a, Fox69b, Fox70, Fox71a, Fox71b, Fox74, Fox79, Fox82, Fox87b, Fra87a, Fra87b, Fra87c, Fre70, Fre74, Fre78, Fre81b, Fre86, Fri74a, Fri87, Fur85, Gab83b, Gab83a, Gab75, Gab83c, Gab88, Gau83, Ged71, Gel76, Geo66, Geo78b, Geo87, Gid82, Gil88, Gil84, Gin74, Gin84, Gli83b, Gli84, Gol83b, Gol87b, Gol88b, Goo75, Goo81a, Goo81b, Goo82, Goo83a, Goo85a, Goo85b, Goo88a, Goo68, Goo89b, Got66, Got67, Got71, Gou82, Gou84a, Gou85, Gou84b, Gow71, Gow73, Gow78, Gow69, Gra81, GG70, GG71, GG72, GG74a, GG75a, GG75b, GG75c, GG77a]. Book [GG77b, GG81, GG84a, GG84b, GG85, GG87, GG89, Gra83b, Gra83a, Gra84c, Gra85a, Gra85b, Gra86a, Gra86b, Gra87b, Gre63, Gre71, Gre88, Grm69, Gua89, Gur78, Hac74, Hac75, Hac76, Hac77, Hac82b, Hac82a, Hac88a, Hac88b, Hac89, Had63, Hah70, Hah73, Hai77, Hai80, Hai82, Hal65a, Hal66a, Hal68c, Hal68d, Hal68b, Hal69a, Hal71, Hal72b, Hal77a, Hal77b, Hal80a, Hal80c, Hal84e, Hal84c, Hal84d, Hal84a, Hal85a, Hal85b, Hal85d, Hal85c, Hal86a, Hal86e, Hal86b, Hal86c, Hal86d, Hal87a, Hal87b, Hal88a, Hal88b, Ham84, Han85a, Han70, Han76, Har77, Har84a, Har84b, Har75, Har89a, Har76, Har83a, Har86, Har79a, Har79b, Har81, Har82d, Har83b, Har87b, Has72, Hat87, Hea87b, Hea64, Hei68, Hei75, Hei71a, Hei71b, Hei73, Hei74a, Hei74b, Hei76a, Hei76b, Hei77]. Book [Hen81, Hen82c, Hen82b, Hen82a, Hen83a, Hen83b, Hen84a, Hen84c, Hen84b, Hen84d, Hen84f, Hen84e, Hen84g, Hen85a, Hen86b, Hen86a, Hen87a, Hen87b, Hen88a, Hen88b, Hen88d, Hen88c, Hen89a, Hen89b, Hen89c, Hen89d, Hen89e, Hen83c, Hen83d, Hen84j, Hen84k, Hen84l, Hen84m, Hen85b, Hen85c, Hen86e, Hen87c, Hen88e, Hen89g, Hes74, Hew66, Hig65, Hil71a, Hil71b, Hil75, Hil86b, Hil86a, Hil89b, Hil89a, Hoc86, Hoc88a, Hoc88b, Hod70, Hod89a, HS83, HS86a, HS86b, Hol71, HS88, Hom74, Hom80, Hom86a, Hom86b, Hop72, Hop73, Hop75, Hos67, Huf85, Huf86, Hun89a, Hun84, Hun86, Hut80, Hut84, Hut88, Hut89, Ilt77, Ink87, Ink88, Ins86, Irw82, Irw83a, Irw83b, Jac89a, Jac89b, Jac85, Jam86, Jam87a, Jam87b, Jar83, Jar84b, Jar84a, Jen89, Jev67, Joh79, Joh81, Joh83, Joh78, Jon86]. Book [Jon87, Jor78, Jor79, Jor83b, Jor85a, Jor85b, Jor87, Jor88, Jos67, Jud88, Kar76, Kau84, Kee88, Kee89, Kei84, Kei86, Kei89, Kel71, Kel68, Kel70, Kel73, Kel75, Kel76, Kel78, Kel84, Kel85a, Kil84b, Kil84a, Kni65, Kni67, Kni69b, Kni69a, Kni71c, Kni71b, Kni72, Kni74, Kni75, Kni76, Kni80a, Kni80b, Kni81, Kni83a, Kni83b, Kni83c, Kni84a, Kni84b, Kni85, Kni86b, Kni86c, Kni86d, Kni88a, Kni88c, Kni88b, Kni89b, Kni89c, Kni89e, Kni89d, Koh78, Kra84a, Kra86, Kri83, Kub70, Kuh68, Lar74, Lau68, Law65, Law80a, Law80b, Law81, Law83, Law85, Law87, Law88, Law89, Lay74, Lay85, Lef78, Les83, Lev73, Lev80, Lew85, Lie85, Lie86, Lin86, Lin89b, Lin74, Liv86, Loh75, Lor73, LM74, Lor75, Lor77, Lud80, Mac78, Mac80, Mac85b, Mac72, Mac79]. Book [Mac88a, Mac89, Mah89, Mai85, Mar74, Mar73, Mar89, Mau77, Mau87, May66, May67, May69, McC77a, McC77b, McC80, McC83, McC85b, McC69, MdLD86, McG65, McG66, McG68, McG69, McG70a, McG71a, McG71b, McL86, McN83, McV75, Mea71, Mea84, Men67, Men70, Men74, Mer78, Von88, Mil87, Mol69, Mol71, Mol72a, Mol72b, Mol74, Mol75, Mol78a, Mol78b, Mol79, Mol81, Mol84c, Mol84b, Mol85, Mol86, Mol88, Mol89, Mon78, Moo84, Moo85, Mor87a, Mor84b, Mor69a, Mor71, Mor72a, Mor72b, Mor73a, Mor73b, Mor74a, Mor74b, Mor75a, Mor75b, Mor75c, Mor76a, Mor77a, Mor77b, Mor77c, Mor78, Mor80, Mor81, Mor82a, Mor83a, Mor83b, Mor83c, Mor84c, Mor84d, Mor87b, Mor84f, Mor84g, Mor89a, Mor83d, Mor89c, Mos77a, Mos77b, Mül87, Mül89, Mus74, Nad83, Nak75, Nay75]. Book [Nay77, Nay80, Nay85, Nee85, Nee69, Ner89, Nev78, Nev80, Nev83, Nev84, Nic85, Nic88b, Nor68a, Nor68b, Nor69a, Nor70a, Nor71a, Nor71b, Nor72, Nor73a, Nor77, Nor80, Nor82, Nor83, Nor85a, Nor85b, Nun88, Nut82b, Nut83, Nut84, Nye86, OM75, O’H86, Olb69, Olb70, Olb71a, Olb71b, Olb72, Olb82, Olb85, Ora73, Ora74, Ora76, Ora77, Ora80, Osb85, Osb89a, Osb89b, Osp74, Out80a, Out80b, Out82a, Out83, Out84a, Out84b, Out85, Out87, Out88a, Out88b, Pac70, Pac71, Pac72, Pag66, Pag69, Pat86, Pat84, Pau76, Ped85, Pel85a, Pel85b, Pep71, Pep72, Pep86, Phi67, Phi69, Pic84a, Pic87a, Pic76, Pic80, Pic88a, Pic88b, Pin83, Pin87a, Pin87b, Pin88, Por89a, Por77, Por78a, Por78b, Por79a, Por79b, Por81, Por82, Por83a, Por84a, Por84b, Por85a, Por85b, Por86a, Por87b]. Book [Por87a, Por87c, Por88, Por89d, Por89e, Por89b, Por89c, Por89g, Poy70, Pra70, Pum84, Pum87, Pum88, Pye76, Pye87, Ran74, Rap71, Rap75, Rap79, Rap86a, Rat65a, Rat66a, Rat66b, Rat68, Rat69, Rau75, Rav65, Rav66a, Rav66b, Rav67, Rav68, Rav69, Rav70, Rav71, Rav72, Rav73, Rav84, Rec86, Red84, Ree80, Ree85, Rei86, Ric88a, Ric89, Rid83, Ril85, Rob75, Rob80a, Rob87, Rob89a, Roc87, Roc88b, Rog83, Rog88a, Rog88b, RH86, RH88, Ron86, Ros79, Rud69, Rud70, Rud71, Rud80, Rud83, Rud84, Rud86, Rud88b, Rud89, Rup85, Rup88, Rup89, Rus87a, Rus89a, Rus66, Rus67, Rus68, Rus69, Rus72a, Rus72b, Rus74, Rus77a, Rus77b, Rus77c, Rus79a, Rus79b, Rus85, Rus87b, Rya70, San87, Sar81, Sch81a, Sch83a, Sch84b, Sch84c, Sch84a, Sch86a]. Book [Sch86b, Sch87, Sch69a, Sch68a, Sch69b, Sch70, Sch71, Sch72, Sch73a, Sch73b, Sch74b, Sch74a, Sch75, Sch76, Sch78a, Sch78b, Sch79a, Sch80c, Sch80b, Sch80d, Sch81b, Sch83b, Sch85a, Sch78c, Sco67, Scr68a, Scr68b, Sec83, Sec84, Sec85, Sec88, Sei85, Ser88, Sey77, Sha78, Sha89, Sha73a, Sha77a, Sha79, Sha81, Sha83a, Sha86, Sha87, Sha77b, Sha73b, She85, She63, Sho87a, Sho86a, Sho88, Sho89, Sim84, Sin74, Sin76, Sin77, Sin78, Sin81, Sin84, Sme66, Sme68a, Sme69, Sme70, Sme71, Sme84a, Sme68b, Smi66, Smi71, Smi73a, Smi73b, Smi74b, Smi74a, Smi75b, Smi75a, Smi76c, Smi76a, Smi78b, Smi78a, Smi78c, Smi80a, Smi80b, Smi82a, Smi83, Smi85a, Smi85b, Smi85c, Smi85d, Smi87b, Smi87c, Smi87a, Smi89e, Smi89b, Smi89c, Sne74, Sne87, Sol83, Spa65, Ste62a, Ste62c, Ste62b]. Book [Ste70, Ste75, Sti89, Sti70, Str73, Str75, Stu71a, Stu78, Stu85, Stu87, Sty68, Swa86, Swe76, Swi70, Swi71, Swi77, Swi67, Tan82, Tee84, Tei78, Tei83b, Tei83a, Tei84, Tei88, Tha66a, Tha67, Tha74, Tha83a, Tha83b, Tha84a, Tha84b, Tha85, Tha87, The87, Tho89, Tho76a, Tho76b, Til75a, Til76, Tre87a, Tre87b, Tre88, Tup82, Tur68, Tur69, Tur70b, Tur70c, Tur72a, Tur72b, Tur72c, Tur73d, Tur73c, Tur73a, Tur73b, Tur74a, Tur74c, Tur75, Tur77a, Tur78a, Tur78b, Tur78c, Tur81, Tur82, Tur83a, Tur83c, Tur87a, Tur88, Twe89, Tyl75, Und64a, Und66c, Und66a, Und66b, Und75a, Und75b, Und79a, Und79b, Und80, Van86, Vea64, Ver89a, Ver89b, Vic88, Wal89, Wal74, Wal77, Wal81b, War88a, War88b, War89, Was69, Wat73, Wea78, Wea80, Wea83, Wea84, Web75, Web76]. Book [Web78, Web84, Web65a, Web65b, Web66a, Web66b, Web66c, Web67a, Web68a, Web68b, Web69b, Web69a, Web70, Web71a, Web71b, Web72, Web87, Wei83a, Wei85a, Wei85b, Wei86, Wel70, Wes74, Wes70, Whe89, Whi70a, Whi68a, Whi72a, Whi72b, Whi73, Whi75, Whi85a, Whi89a, Whi68b, Whi84, Whi85b, Whi88a, Whi89b, Whi68c, Whi77, WM85, Wig69, Wig74, Wig77, Wig78, Wil65a, Wil75, Wil65b, Wil68, Wil69, Wil77, Wil84e, Wil84d, Wil84b, Wil84c, Wil85a, Wil85b, Wil86a, Wil86b, Wil87d, Wil87c, Wil87b, Wil78, Wil89b, Wil89a, Win75, Woo80a, Woo80b, Woo80c, Woo83, Woo84, Woo89, Woo67, Woo73, Woo85, Wor88, Wuj85, Yat72, Yeo88, You66, You83, Yox82b, Yox82a, Yox83a, Yox83b, Cla99a, Cla99b, Cla95a, Cla9 , Cla2 , OW10, Pat94, Ped92a]. Books [AS89, All81, Ano62d, Ano63e, Ano63f, Ano64e, Ano64f, Ano65e, Ano65f, Ano66e, Ano66f, Ano67e, Ano67f, Ano68f, Ano68g, Ano69e, Ano69f, Ano70e, Ano70f, Ano71e, Ano71f, Ano72e, Ano72f, Ano73e, Ano73f, Ano74g, Ano74i, Ano74h, Ano75g, Ano75h, Ano75i, Ano76g, Ano76h, Ano77h, Ano77i, Ano77j, Ano78g, Ano78h, Ano78i, Ano79g, Ano79h, Ano79i, Ano80g, Ano80h, Ano80i, Ano81g, Ano81h, Ano81i, Ano82g, Ano82h, Ano82i, Ano83g, Ano83h, Ano83i, Ano83l, Ano84g, Ano84h, Ano84i, Ano84l, Ano85g, Ano85h, Ano85i, Ano85l, Ano86g, Ano86h, Ano86l, Ano87j, Ano87k, Ano87l, Ano87q, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ant64, Asq80, Bar79, Ben79, Blu79, Bry81, Bus83, Deb66, Fie81]. Books [Fie83, Fin68, Fis81c, Fis83c, Gol79, Gre63, Hen82c, Her80, Jar83, Jar84b, Kni88b, Law80c, Lay85, McV80, Mol85, Mor87a, Nor85a, OS92a, Pic79, Rus72a, Sch68a, Sch79b, Scr68b, Stu71a, Swe76, Tur81, Wal74, Wea79, Woo80e, For73, Got92, Mas98, CL87]. Boole [O’H86, Mac85a]. Boon [Cro70a, Far71a]. borders [YK94]. Born [Bro72a, Bro72a]. Borodin [Bro89b, FS88]. Borrowed [Rap82]. Bos [Sch83a]. Bostock [Hen84d]. Boston [Bar83, Blo79, Bro86b, Byn80, Byn87, Can84, Cha80b, Che85, Cro89b, Dea83a, Din63a, Gau83, Goo88a, Gou85, Gre88, Hal87a, Han85a, Hen82c, Hen82b, Hen84e, Hen88e, Hoc86, Hun86, Kel75, Kni85, Kni89b, Kra86, Law83, Mah89, May67, Mol84c, Mol89, Mor77c, Mül89, Nye86, Olb82, Out84b, Pic84a, Pin87b, Por86a, Sch84a, Ser88, Sha89, Smi89b, Stu78, The87, Tre87b, Web84, Whe89, Woo80a, May67]. Botanic [All82]. Botanical [Bro84d, Bro87, Edw78, Nic85, Sty68, GEMM83, All80, All86]. Botanik [Jun89]. Botanist [Nev83, Web70]. Botanists [All79, Bro87, All86]. Botany [Jun89, Mül89, Ste62d, Sty68, Dea83b, Hen86d, MR96, Sht96, All79]. Boulder [Har82d, Pel85b, Whi89a]. Bound [Hen89a, Ait71, Arm64, Bro89a, Pai86, Smi66]. boundaries [Mal05]. Bourbaki [GG70]. Bourgeoisie [Har83a]. Boussingault [Hal86e, McC84]. Boveri [Olb69]. Bowen [Chi64, Por83a]. Bowler [Bro86d, Nor85b, Por78a]. Box [Roc88b]. Boyd [Ant64]. Boyer [Pra70]. Boyle [Buc66, Car74, Hal80c, Hen86e, Jon87, Web71a, Ale85, BKH90, BHW94, Hoo97, Ost89, SSH85]. Brace [Bar77]. Bradbury [Hal68b, Tur69]. Bradford [Rus77c, Mor85]. Bradley [Col81b]. Bradwardine [BM89]. Bragg [Kni71b, Stu71b]. Brahe [Mor87a, Hel63, Ros86, TC90]. Brain [Cri71, Grm69, Jac85, Smi75b, Har87a, NC81, Smi92, Sta89b]. Branch [Duc84]. Brandeis [Win75]. Braunschweig [Ree85]. Bray [Sho87a]. Breakthrough [Hac76]. Bredon [Tal62]. Breeding [Geo87, Rus86b]. Brepols [Gol83b]. Brewster [Jam87a]. Brian [Lor75, McV80, Pye76, Sha81, Wea80, Wil84d, Dea86a, Gol88b, GG85]. Bridge’ [Hol77]. Bridgewater [Car69]. Bridgman [Wal90]. Briefe [GG75c, LM23, LS76]. Briefen [Bro89a, Mai85, vLvHH^{+}88]. Briefwechsel [Ger89, Hen86a, HP85, LHH91, PHvMW93, PvM85, LHB04, LS76, vL76]. brieven [Phi67]. Briggs [Tha84a]. Bright [Kni88a, Moy86]. Brightman [Ste75]. Brighton [Bar81, Ben82b, Blo82b, Hen84f, Jar83, Sha81]. Bringing [HK82, Mea84]. Bristol [Ben84, Buc89a, Dra86, Fie89, Hac88b, Hal87a, Hun86, Kei89, Mus74, Nye86, Pin87b, Stu85, The87, Tre87b, Tur83c, Tur87a, Whi84, Whi88a, KH86]. Britain [Can74, Fie81, Gre71, Hei73, Hen84d, Ink87, Kei86, Law87, Mac78, Mac89, McL86, Nor83, Por89c, Rap79, Smi85a, Wei83a, Wil84d, Alt87, Ano69h, Arm83, BB92, Ber85, Bry88, Buc85c, Can83, Coo84b, DH90, Fox88, FL88, Gol92b, Ham90, H^{+}82, Hei72, Hun95, Jam88, Lay65, Luc90, Maz92, MJ84, McG84, OO89, Rei68, Ste92, Sut74, TBDA10, Tyl64, Wil94a, Wyn82, Yeo86, Yeo93]. Britannica [Wil85a, FC82]. British [All77, All79, All81, All82, All86, Ano62h, Bow89a, Bow86, Bow87b, Bro85e, Bro87, Bur80, Can77, Cor81, Dav81, Des88, Eag78, Edw78, Fau89, Gou84a, Hei68, HS83, HWW86, Law89, Lie85, McG71a, Mor72b, Mor75b, Nut83, Ora76, Pat86, Pel85a, Rud70, Sch71, Sec85, Smi78c, Tha83a, Tha84b, Vic88, Web84, Wil84d, Wil85b, Wil89b, Wor79, Ano62k, Ano84j, Ano85j, Ano86i, Ano87m, B.69, Ban88, Bow89d, Cal90, CC66, CH67, CE73, Coo89b, Coo89c, FB82, FB75, Gun84, Hen89f, Kuk91, Kuk93, Lie88, Mac88b, Mar62a, Mor84e, O’C88, Ora72, Pat83b, Poc88, Reh83, She86, Sto84, WW75, WW86, WW93, Wil82a, Wil91a, Wil81a, Wri94, FC82]. Broad [Sec83]. Broadcasting [Hei68, Woo92]. Broberg [Bro84d]. Brocchi [McC76]. Brock [Bib74, Far67, Kni86d, Mor81, Nye86]. Brockmeyer [Bro82a]. Brockway [All82]. Brodhead [Duf83]. Broida [Hen82b]. Broke [Sme88]. Bromberg [Kei84]. Bronowski [Dur79a]. Brontë [Shu96, Shu04]. Bronze [Bod84]. Brooke [Duf75, Koh78, Wor88]. Brooks [Bro85d]. Brothers [Car84, Gil83]. Brouw [Dea83a]. Brown [Ait89, Cas85, Fox69b, Fox70, Gur78, Swi71, Wil78, BN95]. Browne [Dur83, Web69a]. Bruce [Cal77, Dau75, Hen83b, Kee88, Mea71, Bra88, Mac85b]. Bruijn [Bry76, Bry78b]. Brundell [Hen88e]. Bruno [For74a, Irw82]. Brunschwig [Wea84]. Brunswick [Bow87a, Gil88, Ree85]. Brush [Lay74, Sch74a, Smi80a]. Brussels [Rat65a, Tan82]. Bruxelles [Jor88]. Bryder [Por89c]. BSHS [Bur80, Wil84d]. Bubonic [Sch71]. Bücher [Whi68a]. Buchgesellschaft [Sim84]. Buchreihe [Mol86]. Buchwald [Roc87]. Bucke [Pic88a, Sho86b]. Buckland [Rud84, Rup83]. Bucks [Gou84a, Vic88, Wil84d]. Bud [Jam86, McL86]. Buffalo [LM74, Lor77]. Buffetaut [Rup88]. Buffon [BG^{+}92, Por84b, BG^{+}92, Slo79]. Bugge [Fox71b]. Bühler [Gra83b, Gra86b]. builder [Sch92b]. building [Gil83, Web84, Wri87]. Buildings [Bry78a]. built [HS83]. Bukharin [Cro72]. Bulletin [Cro68b]. Bulzoni [Sch81b]. Bunge [Sin81]. Burchfield [Smi76a]. Burdens [Pye87, Hol86]. Bureau [Cro91b]. bürgerliche [Dau98]. Burgers [May66, Tur70c, Tur73d]. Burke [Dea86b, Goo68, Cro87b]. Burkhardt [Jor79, Rud86, Rud88b]. Burndy [Pac71]. Burnett [Fis83c, Mor73b]. Burnyeat [Wea84]. Burrow [Pic80]. Burt [Goo82, Hal80c, Smi80a]. Burwick [Kni89b]. Busard [Mol85]. Business [Edg87c, Nor85a, Wil89b, CK89, Lie88, Rei85, Smi94b, Yox83c]. Busvine [Pic79]. Butler [Und71]. Butlerov [Roc81b]. Butterworths [Cha78b, GG71]. Büttner [Hal85d]. Butts [Nay80]. Buxton [Bar75]. Bynum [Dur83, Gol87b]. Byron [NN95].

C

[Aas87, Arm64, Bow87b, Car84, Cas85, Chi64, Chi65c, Cla86b, Coo89a, Coo84a, Dea83a, Din63a, Dor80, Duf75, Dur82, Far71b, Fox69b, Fox70, Fre74, Gra84a, Grm69, Gur78, Hai82, Hal84a, Hal88a, Har84b, Hen82c, Hol71, Hut84, Kni72, Kra86, Kri83, Law65, McV75, MR89, Mor73a, Mus74, Nay80, NN95, Ner89, Nor71b, Ora76, Phi69, Rat65a, Rat66b, Rat68, Rav66b, Rav67, Rid83, Rud80, Rus77b, Rus87b, Sch85a, Sha86, She63, Smi66, Swi71, Tei88, Tur75, Und66b, Vea64, Web66b, Web68a, Wel70, Wes70, Whi68a, Whi68b, Whi77, Woo80e, YK94, Sar81, Tal62, Wig78, WD82, Car86, Ric91]. C-8 [Cas85]. C. [Bro86a, Pum87, Sha98, Whi88a]. C.H [Fig88]. C.U.P [Dun87]. C.W [Hen89b]. ca [Bod84, Got92, Smi78b]. cabala [GGC90]. Cachées [Sho89]. Caedmon [Ora77]. Caesar [RL97]. Cahiers [Sch84c]. Cajal [Cla67]. calcul [Bea91]. Calculating [Hah70, Neb95, MKW92]. calculation [Bea91]. calculations [DR97]. Calculatores [Wal69]. Calculus [Boy70, Gra83a]. Calendar [Cor81, Rud86, Smi78b, BSK^{+}85, BSK^{+}94, Poo98]. Calendars [Cla95b, Cla95a]. Calendrical [DR97]. Calgary [Sha83a]. Calif [McC85b]. California [Arm67, Bod84, Bow88a, Bow89b, Bro83b, Bro86d, Can80c, Dea86b, Fre86, Goo68, GG74a, Grm69, Hal66a, Hal84d, Har89a, Hen83b, Hol71, Hom86b, Kra86, Law80c, McC85b, Nor77, Olb69, Out83, Por79b, Rat65a, Rid83, RH86, Rus87b, San87, Sch78b, Sch79a, Sec84, Sin76, Smi87b, Whi77]. Called [Kni80a, BEoCF92, ECL92, PTS^{+}96, vGA94]. calling [Ole91]. Caloric [Bru70, Bru72, Wat84, Mor72e]. Cambrian [Tha87, Sec86]. Cambrian-Silurian [Tha87]. Cambridge [Aas87, Agr83b, Ait72, Ait73, Ait85a, Arm62, Arm64, Arm67, Asp88, Bad89, Bal84, Bar79, Bea85b, Bec86, Bib74, Bow83a, Bow87b, Bow88b, Bra84, Bro89c, Bro78a, Bro78b, BH84b, Byn74, Byn88, Car67a, Cas85, Che78, Chi65a, Cla86a, Cla67, Cla68, Coc75, Col81a, Con86, Coo81, Coo89a, Cow68, Cro68a, Cro87a, Cun78, Dea83b, Dea86a, Des88, Dol71b, Dub72, Dun87, Dur79a, Dur81, Ede65, Edg87a, Edg87c, Edg87b, Eva89, Fee77, For68, Fox69a, Fox71b, Fra87a, Fur85, Geo87, Gol87b, Gol88b, Goo68, Got67, Gou82, Gow71, Gow78, Gra81, GG75a, GG84b, Gra83a, Gra85a, Hac89, Hal66a, Hal68b, Hal77b, Hal84c, Hal85a, Hal85b, Hal86a, Hal86c, Hal87b, Hal88b, Har85a, Har83b, Hei73, Hei76b, Hen81, Hen83a, Hen84c, Hen84f, Hen85a, Hen88b, Hen89b, Hen89e, Hen83d, Hen84k, Hen84m]. Cambridge [Hen85c, Hen86e, Her80, Hil71b, Hil89b, Hil89a, Hod89a, HS88, Huf85, Huf86, Hun84, Hut84, Hut89, Jac89b, Jam87b, Jev67, Joh81, Jor79, Jor87, Kau84, Kei84, Kel73, Kel85a, Kni69b, Kni71c, Kni74, Kni80b, Kri83, Law80b, Law83, Law85, Law87, Les83, Lie85, Lor73, Mac85b, Mau77, May66, McG70a, McN83, Mea84, Men67, Mol71, Moo85, Mor80, Mor82a, Mül87, Nak75, Nee85, Nor68a, Nor69a, Nor70a, Nor73a, Olb71a, Olb85, Osb85, Out82a, Out84a, Pac71, Pac72, Pep72, Pic88a, Pic88b, Pin88, Por83a, Por85b, Por87b, Pum84, Pum87, Pum88, Pye87, Rap79, Rat65a, Rat68, Rav84, Rec86, Ric89, Rob87, Rog83, Rud86, Rud88b, Rus87a, Rus66, Rus69, Sch81a, Sch84b, Sch86a, Sch87, Sch71, Sch76, Sch80d, Sco67, Sec88, Sha78, Sha73a, Sho87a, Sho86a]. Cambridge [Sho88, Smi66, Smi78a, Smi82a, Smi83, Smi85c, Smi87c, Sti89, Tei84, Tei88, Tho89, Til75a, Tur70b, Tur87a, Tyl75, Und79a, War88a, War88b, Wea84, Web84, Wel70, Whi73, Whi75, Wig77, Wil86a, Wil86b, Wil89a, GG72, CW87, GA98, Hen84h, KBRS90, Ron83, SSK88, Sel10, Stu85, YED97]. Camden [Pat86]. Camera [Fie89, HA87]. Cameron [Vea64]. Campanella [Ern91]. Campbell [Twe89, Woo84]. Campbell-Kelly [Twe89]. Camps [Mil83]. Can [Hal69b]. Canada [Duc84, Gin84, Gow69, GG89, Pep89, Lev88a]. Canadian [Cal77, Mor77a, Mos77a, Mos77b, Lev93]. Canberra [Dol69, Mor84c]. Cancer [Irw83a, Law81, Fuj96, Ric91]. Canguilhem [Bur74a, Law83, CD94]. Cannizzaro [Bro85a, Roc84]. Cannon [Gra84c, Ora80]. Cannstatt [GG77a, Rus72b]. Cano [Cla67]. Canon [Ano64l, Cun89, Sir87]. Cantor [Joh81, Smi83, Smi85a]. Cape [Bry81, For82]. capitalist [Cou96]. Caplan [Sho88]. Capp [Woo80a]. Cappellari [Kra86]. Captain [DJS97, Wri87]. Cardano [Sir97]. Cardiff [Nev80]. Cardwell [Buc73a, Kar76, Kni69a, Ora74]. Care [Dav81, Mac79, Por89d, Bak96, Mor91b, Wea79, Web88, Wie93]. Career [Fra78a, Gab75, Gou68, Mor76b, She91]. Career-making [Fra78a]. careers [Dei92]. Carey [Rup89]. Carezzi [Rap71]. Cargill [Mac88a]. Carl [GG72, Pra70, Rau75, Ste70, Goo85a]. Carlid [Cro67]. Carlisle [Rav67]. Carlo [Arm67, Wig77]. Carmichael [Und80]. Carnot [Hut80, Men74, Smi87a, Ler85, Men81]. Carol [Aas87, Dur81]. Carolina [Nad83, Por78b]. Carolyn [Law83]. Carozzi [Por84a]. Carpenter [Tei88]. Carroll [Jam86]. Carruccio [Rav66a]. Carson [Tur70b]. Carteggio [Hat87]. Carter [Sti70, Wei85a]. Cartesian [Gil84, Cla89, Gro91, Hey83]. Cartography [HW87, HW03]. cas [Jor83a]. Casaubon [Hen83c]. Case [Bec75, Cro73, For70, Gol88a, Lem77, Per73, Pic87a, Rap78, Rav67, Sho88, Stu86, Des89b, EC87, Fra78a, Kus95, Mei91, Pet94, Sch84d, Zan86]. cased [Blo84, Hen84f]. Casimir [Cla70]. Casimir-Joseph [Cla70]. caso [Des89b]. Caspar [Wes70]. Cass [Ait71, Arm69, Car69, Cro69, Cro71, Cro72, Dol73a, Hah70, McC69, McG69, McG71b, Ora73, Sme71, Wel70, Whi72a]. Cassette [Hen84c]. Cassiano [FPLJ89]. Cassidy [Kra86]. Castillejo [Gou84b]. Castle [Wey92]. catalogazione [Olm92]. cataloging [Olm92]. Catalogue [Che78, For80, Kni88b, Rya70, Sch68a, AL87, Ber95, Bry83, CL87, FC82, Hen89g, Nut83, RWD90, Smi78a]. Catalogues [Wal77]. Catcott [NP77]. Cathedrals [SP88]. Cathode [Fal87, Dah97]. Catholic [Maz95a, Nye76]. Catholicism [Hen82d]. Catholics [Las81]. Catoptrica [EiT92, ET92, Tak92]. Cattenius [Kra91, Kra91]. Cattermole [Tur87a]. Cauchy [Gra83a]. Causality [Har75, Mol74, vR95]. Cause [Dra81, Lin68, Nay85, For82]. causis [Lin86, OH85, Rav69]. Cavendish [Chi65b, Mor76a, Cro74a, Cro74b, Far63, McC68, McC69]. Cayley [She63]. Cecil [Por81]. Cecilia [GH84, GH96, Hal85a]. celebrating [AL87, Hen89g]. Celebration [Hen89b, Kil87, Por85a]. Celebrations [Ano66g]. céleste [GP92]. Celestial [Agr83b, Arm69, Cha90c, GP92]. celi [Mol73]. Celia [Bus83]. Cell [Jev67, Tei83a, Wei91]. cellulaire [Duc87]. cellular [Duc87]. celluloid [Fri83, Rus85]. census [CHU87, Tre88]. Centenary [Ano66g, F.76, Hen89b, KK85, Hen82c, Kil87, Sar92, Spr97, Olb65]. centennial [HE^{+}82]. Center [Col88a, Hen82c, KK85, Lin93]. centers [Kri88]. Centigrade [Bry71]. Central [Hen84c, Ste62b, Tre79, Lan87]. Centrale [Bur74a, Out88b, Lan87]. Centrality [Ait86, Gra84b]. Centre [Ben89, Gol79, Hal65a, Hal85c, Hut80, McC80, McC83, Sch84c, Tup82, Ano82l]. Centres [Nay75]. Centro [Sch73a]. Centuries [Ait68, Ait72, Ait73, Arm66, Bro73b, Bro78a, Bro78b, Bru72, Bry74, Bry81, Buc66, Bur74a, Byn72, Can77, Can80c, Car69, Cha77b, Duf81, Dur81, Eag78, Edw78, Fis81b, For68, For72, For74a, Fri74a, Gin74, GG74a, Hac74, Hac75, Hac77, Hah73, Hai77, Hal66a, Hal77b, Har81, Hei68, Hew66, Hom80, Hop72, Hop73, Jor78, Kel71, Kle94, Kni69a, Kni72, Kni74, Kub70, Lar74, Mac72, McG66, McG70a, Men74, Mor69a, Mor72a, Mor72b, Mor73a, Mor74a, Mor81, Nay77, Nor68b, Olb72, Ora77, Pep71, Pep72, Por79b, Rat66b, Rat68, Rau75, Rav72, Rus72b, Rus74, Rus77c, Rya70, Sch81a, Sch68a, Sch70, Sch73b, Sch74b, Sch78c, Sha77b, Sha73b, Sme68b, Smi73a, Smi74a, Sne74, Str73, Tur72a, Tur73b, Tur74a, Web66c, Web68b, Web70, Web71b, Web72, Wes74, Wes70, Whi72b]. Centuries [Yat72, NW82a, Woo82, Ano82m, Asq80, Bar76b, Bib74, Bro76, Bry78b, Buc77, Bur74b, Byn78, Can74, Can75, Cap81b, Cor81, Cow74, Cro79, Dau75, Dol69, Dor75, Dor79, Dug77, Dur79a, Far76, Fis78, Fox74, Geo78b, Goo75, Goo81b, Gow78, Gow69, GG75c, Har77, Har79b, Hei75, Hei74a, Hei77, Hey95, Joh79, Jor79, Kni71b, Mac78, Mac80, Mac79, Mar74, Mau77, Mor74b, Mor75b, Mor77b, Mus74, Nor80, Ora74, Ora76, Pau76, Pic76, Pic80, Poi04, Por78a, Por79a, Pye76, Rob75, Rus79a, Sha78, Sha77a, Sin74, Sin76, Sin77, Sin78, Smi76a, Smi78a, Smi78c, Smi80a, Ste62d, Stu78, Tei78, Und79a, Web75, Web76, MP89]. Century [Ait71, Ait86, Ant64, Bar73a, Bec86, Bro71a, Bro85a, Bro72a, Bro75, Bro85e, Bro86e, Bry77, Buc73b, Byn74, Byn80, Byn88, Car74, Cha80b, Coo81, Cri71, Cro70b, Dav81, Dol70b, Dol71a, Dol74, Dor80, Eme87, Far71b, Fie81, Fin68, Fis81a, For69b, Fox71a, Gab75, Gol87b, Gou84a, Gou84b, Gow73, Grm69, Hac75, Hal68c, Hal68d, Hal69a, Hal77a, Hal80a, Hal84a, Ham90, Han70, Har77, Har84a, Har84b, Har89a, Hei71a, Hei73, Hei76b, Hen81, Hen84e, Hen84m, Hil75, Hom74, Hop75, HM66, Hun84, Hun86, Hut80, Hut88, Ilt77, Ito88, Jam87b, Jar84b, Joh81, Jun89, Kel71, Kni69b, Kni71c, Kni86c, Kni89a, Kuh68, Law80b, Law87, Len82, Lev80, Lon88, Lor75, Mac72, Mac79, Mar74, Mar73, May69, McC77b, McC85b, McG68, McG69, McG71a, McG71b, Men70, Mol75, Mor71, Mor73b, Mor80]. Century [Mor89a, Nak75, Nay75, Nee69, Nev80, Nev83, Ora73, Ora80, Out80a, Out87, Pic76, Pic88a, Por89a, Por78a, Por85b, Por89c, Pum84, Rap71, Rat69, Rav68, Rob80a, Rob87, Rob89a, Rud80, Rud84, Rus89a, Rus68, Rus88d, Sch68a, Sch74a, Sch75, Sei85, Sha78, Sha86, Smi73a, Smi73b, Smi78c, Smi80a, Smi80b, Smi85c, Ste62a, Ste70, Sti89, Str75, Tei78, Til75a, Tur70c, Tur73c, Tur81, Und66b, Und75a, Und75b, Und79b, Und80, Wal81b, Web68a, Web69b, Web69a, Web71a, Web72, Whi72b, Whi75, Whi68b, Whi68c, Whi77, Wil62, Wil68, Wil87d, Agr83a, Ara91, Ben79, BW90, BS79, BPH85, Can84, Dea83b, FM00, Gab83b, G^{+}09, Gli83b, GJ86, Hen83b, Hen84f, Hen84l, Her88, Hui92, JO81, Kel84, Kni84a, Kni84b, KI92, Les83, Lev93, McC85a, Mol84c, Mor84d, NBR^{+}95]. century [Ogi86, Rap82, Roc81a, Rud88a, Sch84a, Sch65, Sch85b, Sha83b, Sti84, Tur83a, Tur83c, Web70, Wei83a, Wil82a, Wil65b, Wil84a, Yat72, Ali86, And84, Ano04, Arm83, AW87, BB92, Bec91b, BaJANJ84, Bos83, Bra89b, Bro81, Bry88, Buc85d, Buc89b, Bur99, BP02, Can91, Cha82, CJ87, CM90, Coo84b, Dau98, DeK73, Dem83, Din66, Dun88, Eng83, Far65, Fin71, Fin93, Fox87a, Fox88, FHR90, Fre92, GA98, Gol87a, Gol88a, Gra81, Gra84b, Gre82, Gre92b, Grm90, Hal84b, Har82b, Har87a, HW81, Her66, Hol89, Hun95, Kni86a, Kni88d, Kro91, Lev90, Lin96, Lin93, Mal05, Man88, Mas98, Maz95a, Men65, Mol68, Mor91b, Moy83, Mur85, Neb95, PP95, Pic87c, Pri83, PW89b, Rap78, RCRC98, Reh83, Ric82, RT95, Roc88a, Roc84, Ros91, Rus88b, Sch88, Sch81b]. Century [Sch89, Sch01, Sha81, SP88, Sho86b, SGŚ00, Sla82, Smi94a, Swa88, TBDA10, Tur83b, Tur87c, Wal81a, WW88a, WR92, Whi90, WD82, Yeo86, De 85, Har85a, Wal69, Sin81]. Cerebellum [Sho87b]. Certain [Hen85c, MT83]. certainty [Hen84l, Mat95, Sha83b, Sta89b]. Châtelet [ZH06]. Chêne [Pau76]. Chadwell [Bad89]. Chadwick [Sha78, Bru88a]. Chain [Rud84, Mor84f, Rup83, War90]. Chair [AH93]. Chalfont [Bar76a, Bur80, Cor81, Gou84a, Vic88, Wil84d]. Challenge [Wil89b, Deb91, Lie88]. Chalmers [All77, Kni80a, Smi79]. Chalmers-Hunt [All77]. Chalubinski [Löw90]. Chamber [Cha80c]. Chambers [Han85a, Hod70, M^{+}89, Mil90, Ogi75]. Chambéry [Ano93]. Chambre [Hil86a, Cha84]. Champ [Hil86a]. Champier [Wea80]. Champion [Cri66]. Chance [Hod92, EM94, G^{+}89, Hac90]. Chances [Hah70]. Chandra [Wal91].

Chandrasekhar [Huf85, Wal91]. Change [Bal83, Chi65c, Edg87b, Fie86, Hal80c, Ins86, Kal95a, Kni80b, Sei85, Web71b, Wor88, AS96, BB84, Fox95, Fox96, HL86, Hun95, LeG88, Mar10, McC93, Mor83a, Moy83, RA91, Rus83b, Sch93, War90, vR95]. changed [G^{+}89, Goo83a]. Changes [Ner89, PP87]. Changing [Can70, Mea87, Rud84, CT95, Dol96, Gre82, RCRC98]. Channel [Hen84c]. chantiers [B^{+}95]. Chaos [Bru86, PS84a, PS84b]. Chapel [Por78b]. chapitres [BN82, Sch84c]. Chaplain [Hil86a]. Chapman [Kni65, McG66, Ste62c, Tho89, Wil85a]. Chaptal [Gra84a]. chapter [Ben90, Men70]. Chard [Pen89]. charismatic [Nol94]. Charlene [Bro89b]. Charles [Ano64l, Bal84, Bro83b, Buc75, Bus74, Car69, Coo81, Des88, GG77b, GG84a, Had63, Har83b, Has72, Hei75, Hen81, Kni89e, KK85, Law65, Lev73, Mah89, Men74, Mol72b, Mor83a, Nev83, Out83, Ped85, Por85b, Rud71, Rud86, Rud88b, Rus79b, Sch70, Sey77, Sti70, Tur78a, Tur78c, Twe89, Web78, Web65b, Wig69, Wig78, Woo89, Abe86, And82, Ano77k, Bur00, BCKR89, BGH^{+}87, Bar76b, Bow90, Bro95b, BSK^{+}85, BS10, BSK^{+}94, But65, Can76, Cla66, Cow74, DK88, DDG90, F.76, GL88, Hen86f, Hut89, Lin90, McC76, Moo81, Olb63, Pag76, Por76a, Por79a, Por89d, Rud76, Rus76a, Tom65, Vig82]. Charleton [McG68]. Charlotte [For67b, Shu96, Shu04]. Charlton [Smi85c]. Charltoniana [McG68]. Charmasson [Gua89]. Charon [For71a]. Chartres [Lem77]. charts [DJS97]. Chatto [Twe89, Wes74]. check [WW75]. check-list [WW75]. checklist [Got92]. ’Chemical [FTG91, Bro85a, Bro73b, Col76, Col88a, Cro71, Cro89a, Dun88, Far65, Hal86c, Hal88b, Hil89b, Jam88, Mor91b, Roc81b, Roc84, Sin77, Sme70, Aft91, B^{+}95, BV86, Bro97, Bro95a, Deb91, Don88, Fis81c, Fru90, Gra84a, GGSGB94, Hal84d, Huf82, KN88, Mor91a, NPYGDL86, NhTWL87, Nye93, Pep89, SG80, TW87, Clo80, Deb66, Fuj86, Tha66b]. Chemie [Adl87, Fig88, Fis78, Fri91, Kle94, Pag69, Sch84d, Sch92b]. Chemiekonzern [Mor84g]. Chemist [Bro89b, Cro83, Ear87, Ede65, Hal86e, Kau84, FS88, Kni85, McC84, Rus86a, Sta84b, Tod83, Wil90b, HE92]. Chemistry [Bro67, Bro86a, Bro71b, Bro72b, Col87b, Col88a, Col88b, Cri66, Cro67, Cro69, Cro70a, Dun69b, FF66, Far67, Far68a, Far68b, Far69, Far71a, Fis81a, Fis81b, Gol88a, Gre63, Hal86c, Hal88b, Har81, Hei71a, Hei76a, Hil89b, Jam86, Jam88, Ken64, Kni67, Kni76, Lev80, Mau77, McL86, Mor69b, NPYGDL86, Olb71a, Pag71, Roc88b, Rus66, Rus67, Rus69, Rus88b, Sha86, Sme66, Sme68a, Tei83b, T^{+}85, Tre88, Adl87, And84, Ber93a, Bro92, Büt83, CHU87, Fri91, Gol92b, Hal85d, Hol85b, Hol89, Kle94, Kni92b, Ler85, MF92, Nye93, RCRC98, RA91, R^{+}85, Sch84d, Sch92b, Sha93, TT86, Tei84, TW87, Bro72b, Col76, Cro70a, Cro71, Dun69a, Far71a, Fis74, Hal72b, Hei76a, Kau84, Kni76, KN88, Mor84b, Mor84f, NhTWL87, Pag69, Rus69, Sme70, Sme71, Sme84a, Tei84, WH91, BVA95]. chemistry [Tra87]. Chemists [Kni76]. cheque [Bro87]. Chester [BEoCF92, ECL92]. chez [BH90]. Chicago [Bow87c, Bro89e, Bro84c, Bro89f, Buc73b, Bud84, Col81b, Duf83, Gow73, Hai77, Hei77, Hen88a, Hen89d, Hoc88b, Kra84a, Lin89b, McC77a, Mor75a, Mor75c, Mor77a, Mor87b, Nay85, Rap86a, Roc87, Ver89a, War89, Web75, Wil69, Wil87c]. Chichester [Law80a]. chief [GG87]. Childbed [Wei85a]. childbirth [Wil94b]. Chilton [Cha80a]. Chimie [Sme69]. Chimique [BV82, Sme84a, B^{+}95, BG80, SG80, SG80]. Chimiques [Hal72b]. chimiste [Cro79]. Ch’in [Nak75]. China [Agr83b, Hal86c, Hal88b, Hil89b, Mad88, Mol71, Nee69, Pac72, Pum87, Sho87a, Tyl75, Ano87s, Car66, Huf93, Huf95, KN88, Mol71, Nak75, Nee69, NLH^{+}84, NR86, NhTWL87, Pac72, yQ85, RA91, Tyl75]. Chinese [Agr83a, Nak75, Rav70, yL66, LY66]. Chiu [Nak75]. Chiu-shao [Nak75]. Chiuchang [Nak75]. choisies [Hal68c]. choisis [Dug77]. choix [Hal68c]. Cholera [Pel85b]. Chomel [Stu86]. Chorley [Hig65]. Chotkowski [Rav84]. Chouet [Gil84, Hey83]. Christiaan [Sch83a]. Christian [Arm69, Sch83b, Hoo97, Smo94]. Christianity [LN86, Rus87b]. Christie [Jam87a]. Christine [Hen84b]. Christopher [Hen83a, Hen89g, Vic88, Ben82a]. chronicled [Dav99]. chronicles [Cau89]. Chronologic [Fis78]. Chronologies [Hen88a, Wil87a]. chuang [Agr83a]. Chubin [Irw83a]. Chur [Buc85a]. Church [Rav72, FC94, Fan96]. Chymia [Bro67, Far69]. chymiatria [Mor91b]. Chymicall [Kni88b, CL87]. Ciampini [Mid75]. Cie [Hom86a]. Ciencia [Gli84, Kni88c, Lan87, OE^{+}90, LN95]. Ciencias [Gli84]. Cientificas [Kni88c]. científico [Gli84]. Ciliary [Tei70]. Cimelia [Rav69]. Cimento [Hop72]. Cinquecento [Sch81b]. Cipolla [Wig77]. circa [For87a, Mor77b]. Circle [Byn80, Ben87, Cha90c, Cha94, Mor91a]. Circular [Mol73]. Circulation [Sch73b, Und66a, Web67b]. circumstance [Pet89]. Cirey [dG01]. Cité [CM90, Hil86a, Cha84]. citing [Kni83a]. citizens [Irw83b]. città [Sch81b]. City [Dun86, Hay74, Cha84, Jos97, CM90]. Civil [Gur78, Pac72, Dau98]. Civilisation [Cal63, Hal86c, Hal88b, Hil89b, Pac72, Pum87, Tyl75, Agr83b, Ano87s, KN88, NLH^{+}84, NR86, NhTWL87]. Civilization [Car66, Mol69, Sho87a, Mei91, Jam89]. Civilizations [Ano82n, Ano82m, NW82b, NW82a, Hen84g]. Civilizing [Pye93]. Clagett [Hos67, Sch80b]. Claim [She70]. Claire [Out80a]. Clarendon [Ait67, Bro69, Bro72b, Bro73b, Bru72, Buc77, Car70, Cri71, Dol70b, Dol71a, Edw78, Fre74, Got66, Had63, Hah73, Har76, Hei74b, Loh75, McG68, Mol78b, Mor73a, Mor73b, Nor68b, Pep86, Por89c, Red84, Rud84, Sch69b, Sch78a, Sch85a, Sha73b, Ste75, Swa86, Tha74, Web67a, Web68a, Web69a, Wig78, Wil65b, Woo67]. Clark [Bro83b, For82, Grm69, Hac76, Hom86b, Kni83a, Woo80b, Woo67, CMN95]. Clarke [Buc85b, Byn73, Har84a, Har89a, Jac85, Smi75b]. Classical [Cro70a, Das88, Far71a, GG75a, HS86a, Kel85a, Mol74, Phi69, Por89g, Was69, BBC88, Har82c, Hen84f, Hil84, Kra84b, Sta90, Got71, Scr68a]. Classics [Edg84, Kni86b, Sul82b, Gje84]. Classification [Bar75, Jor83b]. Classify [Byn88, Gol87a]. classifying [Fis83c]. classique [BBC88]. Claude [Cro79, Sme84a, Mau77, SG80]. Claude-Louis [Cro79, Sme84a, SG80]. Clausius [Pay74]. claw [Mir94]. Clayton [Kei84]. cleanliness [WKB88]. Cleevely [Tha84b]. Clément [Ler85]. clergyman [Eng83, Mor84d]. Clerk [Gol83a, Har02, Har98a, Hen86c, McC69, Smi85b, The87, Chi64]. Clerks [Mol71]. Clerq [Wil87d]. Cliffs [Rav73]. climate [Mar10]. Clinic [Fig77b]. Clinical [AJ^{+}89, Coo83, BP95, Büt83, Hal85d, PB99, Rob89b]. Clio [McC77a, Dij90, Kit89]. Clive [Smi87b]. clock [Sir97]. Clockmaking [Bee89]. clocks [Cla95a, Cla95b]. cloning [Yox83b]. clos [Bla93]. Close [LSS90]. closed [Bla93, Edw96]. Closure [Sho88, EC87]. Cloth [Arm64, Dou84, Jar84a, Sch84b, Smi66, Wei85a]. Cloyd [Mor73b]. Club [Ora73, Jen70]. CNRS [Jen89]. Co [Bry81, For69b, Goo82, Gou85, Joh83, McC69, Rat66a, Rud63, Scr68a, Sme68a, Yox83b]. Co-operative [Rud63]. Co-ordination [Sme68a]. Coadaptation [Rid82]. Coast [Man88]. coastal [DJS97]. Cochrane [Rap86a]. Cockcroft [HA84, Stu85]. Cocks [Gre71]. Codell [Wei85a]. coelesti [Rav69]. Coelestis [Wil85a, FC82]. Cognitive [Ric89, Gie88, MH88]. Cohen [Ait72, Blo79, Byn74, Gou85, Her80, Jar84b, May67, Til76, Whi73, Wil86a]. Cohn [Mon78]. Coimbra [Kni85]. Coincidence [Per73]. coins [Bry81]. Cold [Hui92, Edw96]. Cole [Fre70, Mor75a]. Coleman [Fel89, Geo66, Jor85b, Mor72b, Smi73a]. Cólera [Kni88c]. Coleridge [Wyl89]. Colette [Hen82c]. Colin [Gow78, Hen84c, Nay77, Nor68b, Tee84, Ear87, Ins86, Sec83]. collaboration [Fie86, Hal88b, McC80, Tei83a]. collaborative [Mol71]. Collard [Und79a]. Colleagues [Bro89e, AT87]. Collected [Ano77k, Asp88, Dor79, Fox70, GF^{+}03, GG87, Hen88d, Hen85b, Kra86, Nor77, Swi71, BH87, DD95, Kni89e, SCS^{+}87, Tee84, KvDC^{+}37, LM23, LS76, vPSBD^{+}04]. Collection [Fis81a, Jon86, Mol88, Tur70b, McC81, O’C88, For73, SH73]. Collections [Ano69h, Cha80c, Col83, Col88a, Deb66, For74b, Got86a, Jar83, Kni83b, Rya70, Smi78a, Smi83, Cle83, HAB92, Tha84b, TW87, GG86b, WG81]. Collective [Fig78, Tha74]. Colleen [Col88a]. College [Gra81, Kni89c, N^{+}88, Pat86, PS78b, Rat69, Woo67]. Collegio [Dea86c, Wal84]. Collezionismo [Sch81b]. colliding [Ste90b]. Collier [Und75a]. Collins [Dol83, Kni88a, Pic87b]. Colloque [BG^{+}92, Cas85, A^{+}87, Pic87c, Bur74b]. Colloquium [Ano69i, Dol71b, Hof62, HS63, May67, Scr64a]. Colonial [Gur78, Kni88a, Mül89, All82, Gro95, Hen86d, Moy86, SP88]. Colonialism [Ink88, Osb94, RR87]. colonisation [San91]. Color [Jac89b, Bla83, Sep88]. Colorado [Har82d, Pel85b, Whi89a]. colored [Sha93]. colorum [Rav69]. Colour [Tur83c, For71a, Law80c]. Colours [Ham84, Wel70, Her83]. Columbia [Bro85d, Eme87, Fri87, Men70, RH88]. Columbus [Bro85a, Can80b, Kei86, Mol79]. Combination [Dun88]. comburentibus [Sch85a, BL83]. Comedy [Wil78]. comes [Bea85b, Mor84a]. comet [RL97]. Comité [Osb89a, Cro88]. commemorate [Kar76]. commemorating [PS82, vLP82]. commemoration [GT92, Hen82c]. Commemorative [Sar81, Chi64, Col83, Hen82c]. Commencé [McC80, McC83]. commensurabilitate [Mol73]. commentaries [Bro72a, Dav99, Got71]. Commentary [Bro89e, Got66, Jar84b, Joh79, Kni69b, Lin86, Mol81, Nut84, AT87, BM89, Eas86, EiT92, ET92, Hal69a, Mol78b, Mol88, OH85, Sin77, Tak92, Vic88, Yox83a]. commentated [ZH92]. commercial [Cat88, Pow96]. Commitments [Cro66a]. Committee [AB89]. Common [Bos83, Clo80, Kel71]. commonly [Mol85]. Commonwealth [Whi68b]. communicating [Jen91]. Communication [Bry77, Hal65c, Bec91b, Cha78b]. Communications [Gol79, Mar62a, Pri83]. communis [Nor71b]. communities [Doe96]. Community [Mor77a, Smi74c, War89, Gol95, Gos87, Hal84d, Huf82, Mit92]. Companies [Fre81b]. Companion [BP93b, O^{+}90, Sel10, YED97]. Company [Ait86, Ant63, Bro86b, Byn87, Can84, Car74, Che85, Din63a, Fre81b, Ged71, Hal88a, Hea64, Lie85, Mah89, Mor72a, Mor83a, Out84b, Rud89, Sch69a, Sha89, Tur87a, Und66c, CW87, FB82]. Comparative [Buc89a, Gli87, McC76, Rav73, Whe89, OS92b, Wil87g, Win91]. Compared [Rud62]. Comparison [Ink88, RR87]. Compiled [All77, Che78, Geo78b, Sch68a, Smi78a]. Complete [Arm66, Por87b, Tur74a, FMF85, Kit92, VB0 ]. Completed [B.69].

Completeness [May69]. Complètes [Bos70, VB0 ]. completion [MR00]. Complex [Far68b, For87b, Hen87b, Ken86]. Complexe [Jor88]. Complexity [Far65]. components [Cal90]. composite [FS91]. Composition [Gro89]. Compressed [PF67]. Compton [Bro76]. Computer [CKA96, Ran74, Smi78b, Bea85b, Cer89, Hen89f, Mor84a, NOF96]. computers [Edw96, HS83]. Computing [Sch93, Asp90, OO89]. Comte [Hai78]. Comune [Sch81b]. Concept [Cla71, Goo81b, Hal68d, Mül87, Osb89b, Rus66, Wei85a, A^{+}87, CM93, Goo80, Gra85a, Hon89, McI85, MC^{+}93, Row90a, Wus84]. Conception [Wea78, You66, Gau82, Had94]. Conceptions [Smi83, Hak95]. Concepts [Blo84, Bow89d, Din63a, Dun88, Har89a, Hen84f, Kil84b, Kil84a, Kra84a, Men65, Pac74, Red84, Tha74, Wil84c, Car67b, Cas85, CJ87, Law83]. Conceptual [Bec91a, May66, Sei85, Ver89a, Har82b, Hen84f, Hul88, Moy83]. Conceptualisation [Sch86b, Bla83]. conceptualization [Bla83]. concern [You83]. Concerning [Ano64g, Hen83c, MC81, Sch69a]. concise [M^{+}89, Mil90]. Condenser [Swi70]. Condorcet [Hai77]. conduct [Gol95]. conductors [BC92]. confer [Kit89]. Conference [Ano87n, Che78, Chr93, Dun86, EK91, EK92, KK85, Mac88a, Mad88, Por81, A^{+}87, BG^{+}92, CM90, Hol85a, Pic87c, WV92, EK88, HS89]. Conferences [Stu78]. Confidence [Cla86a, Tom84]. Conflict [Pac70]. Confronting [Fra87b]. Congrès [Ano93]. Congreso [Gli84]. Congress [Ano68j, Ano93, Cro72, G.62, Nee78, Sar81]. conjunction [Bro73a, Fis83b, Fox79, Rus77c]. Conn [Kni83a, Mor72a]. Connecticut [Blu79, Gab83b, Hei74a, Swe76]. connection [Hug89, Kle94]. Connell [Gol79]. Connexion [Bro81, Can80a]. Connor [Hac88a]. connus [vPvP90]. Conquest [Llo63, Que88, Sha77a]. Conrad [Por85b]. Consciousness [HS85, Fox88, Tay96]. Consejo [Kni88c]. Consellaria [Kni88c]. Consent [Law87, Coo84b]. consequences [Ric92a, Wol87]. Consequent [GG74a]. Conservation [Can75, Pac70, Pac74, Can93a, Mac88b]. conservative [Mei91, Coo89b]. Conservés [Bur74a]. Console [Byn88, Gol87a]. Constable [Mor71, Web66b]. Constitution [Cro69]. Constructing [Dan90, Hen86b, Pic84b]. Construction [Tur81, Dol83, Fri89, Irw82, Mac90, Nor83, Nut82b, Ric92b, Rus89b, Tee89, WT82]. Constructive [Cha88b, Tho89]. consultations [AJ^{+}89]. Consum [Kni88c]. consummatio [Mor83d]. Consumption [BP93a]. Contemporaines [Stu87]. contemporaries [Zam92]. Contemporary [Bec82, Goo88a, Hen80, Hen84e, Rav66a, Tur74b, Web68a, Ner87, PW89a, Tor83]. content [Wal92]. contentious [ZH92]. Contents [Ano82j, Ano83j, Sme69]. contested [Liv93]. Context [Byn87, Cro77, For86, Hen89e, Ito88, Kni87, Mor81, BVA95, Goo82, Har82d, Kum91, Lin92, PT92, TR92, Ver85]. contexts [Irw83a]. Continent [Lev88b, Sch84c, Gue81]. Continental [Fra78b, Lev88b]. continents [LeG88, Ste90b]. Contingency [Pic87b]. Continuity [BB84, Fie86, Mor83a, BA91, Hen84j, Kre82]. Continuum [Whi85b, Sor83]. Contra [Jac89b, JK84, Sep88]. Contraction [Bro79, Hun88]. Contrasts [Fru90]. Contribution [Hen88f, Mül89, Cou96, Gra85a, Hen86d, Hen85b, Wei91, Wus84, Yat83]. Contributions [Mar62a, Wil68, Ben79, Gol91]. Contributor [Ano81j, Ano81k]. Contributors [Ano62f, Ano62g, Ano63j, Ano63k, Ano64j, Ano64k, Ano65h, Ano65i, Ano66i, Ano66j, Ano67g, Ano67h, Ano68h, Ano68i, Ano69j, Ano69k, Ano70h, Ano70i, Ano71i, Ano71j, Ano72g, Ano72h, Ano73g, Ano73h, Ano74j, Ano74k, Ano74l, Ano75l, Ano75m, Ano75n, Ano76i, Ano76j, Ano76k, Ano77n, Ano77o, Ano78j, Ano78k, Ano78l, Ano79j, Ano79k, Ano79l, Ano80k, Ano80l, Ano81l, Ano81m, F.77, Ano80j]. Control [Hac89, Ben86, Cro92, Hac82b, Nor85a, Yox83c]. controverses [BD95]. Controversies [Sho88, BD95, EC87, Hal83b, Rup85, Sho96]. Controversy [Cow74, Fuj86, Mac80, Mor87b, Olb85, Olb89, Ree80, Sec86, Tha87, Fri91, Kri82, Mos93, Per92, Rud85, Sch84b, Sul82a, Tur94]. Conversation [Arm66]. Conversations [Rap71]. Conway [Und66c, CMNH92]. Cook [DJS97, Wat68]. Cooke [McC92]. Cooper [Kni88b, CL87]. Cooperative [Mor89a, Swa88]. Cooter [Law87]. Copenhagen [Ant63, Bad83, Str73]. Copenhaver [Wea80]. Copernican [Cro90b, GB95, Mos93]. Copernicana [Eas86, Mol81]. Copernicanism [Ait86, FC94, Fan96, Gra84b]. copernicano [GB95]. Copernicu [Tho76a, Tur74a]. Copernicus [Ait86, Duf75, Ger89, Tho76b, Tur74a, Ger89, Hen91, Ros84a, Rus79b]. Copper [BC92]. Coral [Sto76]. Cord [Grm69, Jac85, NC81]. Cordelia [Ora77]. Cornelis [Hal85c, McC80, McC83]. Cornell [For74a, For87c, Gra86b, Hal86d, Hen84j, Kni83b, Por87c, Rud84, Sch84c, Sch78b, Tha83a, Und75b, Und79b, Wal89, Wil75]. cornered [Nor68b]. Coronado [Rus77a]. corpis [Per89a]. Corporation [Arm66, Buc68, Dun69a, Hei68, McG68, Rat69, Rav69, Sch74a, Sme69, Ste62c, Web69b]. corporis [Wea78]. Corps [Kni83a, Sin91, Fis98, GG92]. Corpus [Cop92]. Corpuscles [Fal87, Hen86e, Ale85]. Corpuscular [Hal71, Stu71b]. Correction [Ano89n]. Correspondance [BaJANJ84, Hal85c, Hom86a, McC83, LF97, McC80]. Correspondence [Ano64g, Ano04, Ano05, Bar73a, BaJANJ84, Bro72a, Bro78a, Bro78b, Cro67, Cro88, Dor79, Duc88, For68, Hei73, Hen86a, Hop73, Hop75, Joh79, Kni69b, McG71b, Mon87, Mor69a, Pat86, Rud86, Rud88b, Und75b, Und79b, Web66c, Web68a, Wes79, Woo80c, vBDGO93, Bur00, BSK^{+}85, BS10, BSK^{+}94, CHG91, CMNH92, FFLW97, GGSGB94, HRP90, HP85, LF97, LS76, Mor84e, PHvMW93, STW90, vL76]. Corresponding [Ora76]. corrigée [GG70]. Corsi [Bow89b, Bro89c]. corte [Sch81b]. cosas [ML89]. cose [Ara91]. Coser [Bar77]. Cosmic [Fur89, Wil87c, Van85]. Cosmographical [Sch79a]. Cosmographicum [Jar84b, KA81]. Cosmologists [War88a, Fur87]. Cosmology [BBBM90, Din66, For71a, For74a, Lin89b, Sch78b, Fie88, Nor94, Pau93]. Cosmos [vH97, Gou84b, Gra94]. Costabel [Bos70, Hal77a, Ilt77]. Cote [Pau76]. Cotes [Bra84, McG71b, Gow83]. Cotes-Natural [Bra84]. Cotter [Sti70]. Cottingham [Dun87]. Coues [Duf83]. Could [Moo77]. Couleur [Sch86b, Bla83]. Coulomb [BD95, Fra75]. Coulson [Chi64]. Coulston [Men74]. Council [Duc84, Gin84, Gow69, Ano62h, Ano64n, Ano75o, Ano76l, Ano77p, Ano78m, Ano80m, Ano81n, Ano82o, Ano83m, Ano84j, Ano85j, Ano86i, Ano87m, Ano88r, Ano89q, AB89, BU63, CM64, CC65, CC66, CH68, CW69, CE74, EW70, Hal79, WE71, WE72]. Count [Fox69b, Fox70, Swi71, ABR93, Wey92]. counting [Ger89]. Countries [Gou82, Hai82, Har82d, Yox82a, Ano90, MP89]. Country [Bry78c]. County [LN95]. Courier [vPvP90]. Cours [Out88b, Lan87]. Course [Dor80, Por84b, Duf75]. courses [Lan87]. court [Mor91a]. courtier [Bia93]. Courting [Sho87b]. Couturat [GG75b]. Cover [Ano62a, Ano62b, Ano62c, Ano63a, Ano63b, Ano63c, Ano63d, Ano64a, Ano64b, Ano64c, Ano64d, Ano65a, Ano65b, Ano65c, Ano65d, Ano66a, Ano66b, Ano66c, Ano66d, Ano67a, Ano67b, Ano67c, Ano67d, Ano68a, Ano68b, Ano68c, Ano68d, Ano69a, Ano69b, Ano69c, Ano69d, Ano70a, Ano70b, Ano70c, Ano70d, Ano71a, Ano71b, Ano71c, Ano71d, Ano72a, Ano72b, Ano72c, Ano72d, Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74e, Ano74b, Ano74f, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano80a]. Cover [Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Gra87b]. Covetous [Sou93]. Cowan [Dol73a]. Cower [Pat86]. Cowing [Bra84]. Coyne [Dun86]. Cr [Cha80c]. crabs [PR95]. Cracow [Dun86, Tur74a]. Cradle [Kel83, Sin84]. Crafting [Fuj96]. Craftsmen [Mol71, FJ93]. Craib [Ada89]. Craig [Rud80, Sch74a]. Craigie [Cla67]. Crane [Hen81]. Cravens [Mac80]. Crawford [For80, Goo85a, Rec86, Tre88, For73]. Crawforth [Bry81]. crayfishes [PR95]. Creation [Che85, Hal68d, Hen88c, Hod70, Kal95b, Lef78, Moo81, Num95b, Whi85b, Hen84b, Buc94, CS94, Dau95, Hen84i, Sor83, Blo82a].

Creation-evolution [Num95b].

Creationism [NBR^{+}95, RT95]. creationist [NN95]. creationists [Num92]. Creativity [Col87b, Hol85b]. Crew [Web66b]. criminal [Bur99]. Crise [Fox87b, Gre86]. crisis [Gre86, Wal90, vR95]. Cristaux [Tur72a]. Critical [Ano82n, Ano82m, Bro73a, Fox79, Kni74, Lin86, McG65, Mol72b, Mol89, PW89a, Sch85a, Sec84, Smi87a, Tre79, BL83, Bud84, CF86, Cha90c, Cha94, EiT92, ET92, Gle82, Gow78, KRS83, Men75, NW82b, Nor71b, OH85, Rot93, SG80, Tak92, WW88b, Yox83a]. Criticism [Cun78, Dol71b, Jun89, Sta91, WM85]. Critics [Cow74, Smi74c, War88a, Maz92]. Critique [Str75, Ait68, CS92, Hal77a, Hey95, Mei91, SG80, Sch78c, Sme84a]. Crofts [May69]. Crombie [Arm62, Arm64, Chi65c, Smi66]. Crook [Moo85]. Crookes [DeK73]. Croom [Bus83, Fig78, HS86a, Mac79, Mor84c, Por88, Por89b, Ron86, Rup88, Sha77a]. Crop [Rus77a]. Crosby [Rus77c]. Crosland [Fox71b, Gre63, Rus72a, Sme69]. Cross [Ink88, Sha79, RR87]. Cross-Cultural [Ink88, RR87]. Cross-Disciplinary [Sha79]. Crosse [Ste81]. Crossroads [Cro72, Din77, Pye76]. Crowds [vG92]. Crowe [Sch87]. Crowther [Mor76a]. Crucial [Per89b]. Crucible [Cla86b, Hen84c]. Crusaders [Coo84a, Who82]. crystal [Bar91]. Crystallography [Goo68, Tur70c, Tur73d, Sch89, Tur73d]. Crystals [ES90, Goo68, Tur70c, Tur73d, Tei78]. CSE [Fie81, Fie83]. cuiuslibet [Mor83d]. Cullen [Hei76a, Gol88a]. Cult [Sha98, Nol94]. Cultivating [Sht96]. Cultivation [Han85a, Pat83a]. Cultural [Byn80, Ink88, Law87, Rob80b, TBDA10, Bee96, Coo84b, Gou95, HE^{+}82, Kil84a, MB^{+}08, MF92, Pan91, RR87, Rud88a, TR92, Wal90].

Culture [Bod84, Bro84a, Can84, Dur81, For87a, Gil84, Goo81a, Jar84a, Kee89, Kni89b, Moo84, Mor84d, Ora80, Pat84, Por89e, Rav84, Rus87a, Sch81b, Sec85, Smi80a, Tur78b, Ash95, Baa90, BB92, Bia93, Bur87b, CK90, Dau98, Gol92b, Har96, Høy94, KC95, Kro91, Liv87, Mei91, OS06, Ols90, Roc88a, Smi94b, Sut95, You85, vHC95]. Culturelles [Sho89]. Cultures [JSS96, ME81, Pic84a]. Culverhouse [Und71]. Cumulative [Ano82n, Ano82m, Bro73a, Fox79, Fis83b, Hen84g, NW82b, NW82a]. cure [Wai90, Wea90]. Curie [Tre87a, Gir86]. curiosities [PR95]. Curiosity [BS91, Bow88a, SL87]. Curious [Mos77a]. Curl [Tur78a]. Current [Ano62i, Ano62j, Ano63m, Ano63l, Ano64m, Ano64o, Ano65k, KH95, Sie91]. Curriculum [Rob75]. Curves [Fra87c, Gra86a, Eng84]. Cusanus [Mol86]. Cusanus-Gesellschaft [Mol86]. Cuthbertson [Bry77]. Cutright [Duf83]. Cuvier [Cor81, Cro85, Geo66, Osb89a, Lau87b, Out84c]. cxliv [Web75]. Cybernetics [WM85, Hei91]. Cycle [Whi70a]. cycles [Cou96]. Cyclic [Ito88]. Cynthia [Byn78]. Cyril [Fri74a]. Cytochrome [Jev67].

D [Ait89, Ant64, Bar75, Ben79, Blo79, Bro86b, Bro82b, Bry77, Buc73a, Bur71, Byn80, Byn87, Can84, Cha80b, Che85, Col88a, Dea83a, Din63a, Duf75, Dun87, Dur80a, Edg84, Far77, Fis74, Fox71a, Fra87b, Gau83, Got66, Gou85, GG77b, Gre63, Grm69, Hac88a, Hal86b, Hal88a, Han76, Hei76b, Hen82c, Hen82b, Hen84e, Hen88e, Hoc86, Hol71, Hom80, Jam86, Jam87a, Jar84a, Jud88, Kar76, Kel73, Kel75, Kni69a, Kni84a, Kni84b, Kni85, Kni89b, Kuh68, Law83, Lud80, Mar89, Mol78b, Mol84c, Mor74a, Mus74, Nay80, Ner89, Nev78, Nor68a, Nor69a, Nor70a, Nor73a, Ora74, Out84b, Pep72, Pic87a, Pic88a, Por86a, Rat65a, Rav66b, Rud80, Sch84a, Sch71, Sha89, Smi74b, Smi76a, Stu78, Til75a, Tur82, Tur83c, Und64a, Wea83, Web76, Web84, Web66a, Whe89, Whi77, Woo80e, You83]. D [Yox83a, Yox83b, Lef78, Sme69]. D. [Moo85, WHS92]. D.C [Hac74, Hac77, Kni83a, Woo85, Wor88]. D.D [But65, For85a, Gun84]. D.M [Bro82a, Nor71a]. D.M. [Wes70]. D.SC [But65]. d’Ailly [Smo94]. daily [Gun84]. Dakota [Kni83a]. d’Alembert [Hah73]. dalesman [Sec83, Spe82]. Dalton [Bro85a, Far71b, GG77b, Kni69a, Kni74, Tha66a, Tha67, Fuj86, Roc84, Tha66b, Tre69]. Dame [Dol70a, Kni81, Lef78, Nay80, Por84b]. Dan [Str73]. Daniel [Buc69, Ste70, Hoc88a, PF67, Por89e]. Daniels [Woo73]. Danish [Ped92b]. Dannenfeldt [Web70]. dans [A^{+}87, Hai80, Osb89b, Tan82, Wil77, dG01]. D’après [Bur74a]. D’Archimède [Sho89]. Darien [Mor72a]. Dario [Sch81b]. Dark [Rap86a, Ros84b]. Darmstadt [Sim84]. Darnton [Mas98]. Darwin [Bar76b, Bro83a, Byn80, Cow74, Des88, Duf75, KK85, McN83, Mor82a, Osb89a, Por79a, Por85b, Rud86, Rud88b, Rus87a, Tur87a, Web65b, Ano77k, Bur00, BGH^{+}87, Bec92, Bow90, Bro95b, BSK^{+}85, BS10, BSK^{+}94, Bur75, Byn78, CW87, Cro91a, DK88, DDG90, DM91, Gay92, Hod77, Koh89, Lau87b, Mon87, Moo81, Olb63, Pan91, Ric92b, Rit90, Smi74c, Sto76, Wil91b, Wil81b, You85]. Darwinian [Bro82b, KK85, Nor85b, Smi89e, Bow83b, Bow88c, Cit90]. Darwiniana [Ben82b, Bro82b, Dur82, Mor82a, Nor82]. Darwiniana. [Web65a]. Darwinism [Ben82b, Dur82, Liv86, Nor85b, Web65a, Bec92, Bow83b, Bow93, Deg91, Dur85, Jun89, Moo77, Out80c, Tor92, Wei84, Wei91]. darwinisme [Bec92, Tor92]. Darwinismus [Jun89]. Darwinist [Moo85, Cro84b]. Darwinizing [Moo89a]. Daryl [Irw83a].

Daston [Por89g]. data [EMMS93a, EMMS93b]. Dauben [Joh81]. daughters [Sht96]. Dauler [Har84b]. Daumas [Bry74, Cal63, Hil71a]. Davaine [Cla70]. Davenport [Buc69]. David [Blu79, Cha65, Cha78a, Cro70a, Des88, Far68a, Far71a, Fis83c, Gou84b, GG75c, Har81, Has72, Hen83b, Jam87a, Jev67, Joh78, Jor78, Kel70, Law65, Law85, Lev80, Lor73, Mac79, McV75, Moo84, Mor69a, Nev83, Nor71b, Nor72, Olb82, Por84b, Rav73, Rob75, Rus79b, Sey77, Sha77a, Smi78a, Swi77, Tur70b, Tur78a, Tur78b, Wal74, Bro87, Buc89a, Coo89a, Eag77, Hen87a, Kee89, Kra86, Rob89a, Rog88a, Rus87b, Sch85a, Smi89e, Stu85, Ver89a, War88a, Wor79]. Davies [Bus83, Ham84, Rud70]. Davis [Hig65, Tre87a]. Davy [Kni71c, Mil83, Rus68, Cro73, Kni92a, SD76]. Dawson [All81, Blu79, Goo81b, Har79a, Har82d, Lev80, Ora80, Sin81, Smi78c, Til76]. Day [Lar74, Mor73a, BW90]. Days [Lay85, Mac82]. DDT [Irw83b]. Deacon [Wat73]. Dead [Lev90]. Dean [Roc88b]. Deane [Wel70]. Dear [Joh79, Wal89, Joh79]. dearth [BR95]. Death [Jor85b, Law88, Ric87c, BG^{+}92, DH90]. Debate [Bro89c, Jen88, Olb89, Sch87, BH84b, Cor88b, Cro86, Dic82, Les83, Mas98, Pum84, Yeo93, Smi82b]. Debates [Far67, Sha81, Ilt73, Num95b, B^{+}95]. Débats [B^{+}95]. Debt [Cla62, Smi79]. Debus [Gra81, Jos67, Kni85, Sch80d, Whi77, Wig74, Wil69]. début [Hal72b]. Débuts [Ben89, Out88b, Tan82, Lan87, Pic87c]. decades [Bow83b, Nor85b]. Decay [Rud70]. December [Ben79, EJR94, EJR95, Goo85b, Hac74, Hac77, Sar81]. Decembre [Cas85]. Decisions [Hen85a, Pug84, Wil84d, Wyn82]. Decker [Smi80b]. Declaration [BM76]. Decline [Wil75, Yat72, Deg91, Law93]. découverte [Jen89, Six87]. découvertes [Cas85]. d’Edition [Wil77]. Deduktion [GG81]. Dee [Rat66b, Clu88, DF98, RWD90]. deep [Ben84, McC82, Rud92]. Defence [Jon75b, Ait85a, JK84]. Defense [Ait86, Gra84b]. Deficiency [Bar89c]. Defied [Bod84, Ols90]. Defining [Yeo93]. definition [Ber93a]. Deified [Bod84, Ols90]. deist [Sch84b, Sul82a]. del [Fre90, GB95, Goo89b, Hop72, Kni88c, Lan86, Olm92, Rap86a, Rey90, Sch79b, Sch81b]. deleted [Bec88, Tee89]. Delft [PS82, Tei83b, vLP82]. Delhi [Hil75, Wuj85]. Delisle [Sme84b]. delivered [Bal83]. della [Olm92, Out80b, Por79a, Sch73a, Tur72b]. dell’aristotelismo [Sch80c]. dell’atomo [Mor89c]. Delusions [Jon75b]. democrat [Sta84b]. Democritus [Pum84, Dic82]. demon [Gol83a, Smi85b]. Demons [Mon78]. demonstrated [Whi72a]. demonstration [Sut95]. Dempster [Bro86e]. Denis [Kni89c]. Dennis [Jac89b]. Dent [Cun83a]. Denyse [Wei83a]. Department [For87a, Kni89c, Nut83, Sch68a]. deque [Rav69]. Derbyshire [Bar75]. Derek [Bry78a, Cha77b, Sha79, Smi78b, Kni86b, Rog88b]. Desargues [Tur88, DFG87]. Descarte [Har84a]. Descartes [Byn74, Dun87, Gab88, Har84b, Jar83, Jar84a, Lau68, Rav65, Ree80, BH90, Buc63, Cla82, Gab78, Gar92, Lov75, She91, Sor87, Web69c, Wil82b, vR95]. Described [Fox71b]. Description [Bar75, vH97]. Design [Pet94, Smi79, Web68b, Bro83b, Hod92, Pet89, Sch89]. Desire [GG74a]. Desmond [All79, All84b, Har84a, McN83, O’H86, Tha84a]. d’Espèce [Osb89b, A^{+}87]. Destitute [Law88, Ric87c]. Determinants [Abe86]. Determinism [Nye76, SM94]. Detroit [Hes74]. Deutsch [Fis83c, Tur78b]. deutsche [Dau98]. Deutschen [CM90]. Deutschland [CM90, Jun89, WKB88]. Deux [Fox87b, Gre86]. Deuxième [GG70]. developing [Lie85, Wri94]. Development [All82, Bod65, Bow89c, Bow89d, Bro71a, Bro82a, Bry77, Cha65, Cha78a, Col81b, Cro68a, Dub72, Fan83, FF66, Fie86, Fre82, Goo75, GG71, Gra86b, Har76, Hei71a, Hei74b, Hen86b, Hen88c, Huf86, Jac85, Jam89, Kni86d, Mil73, Out82a, Pic87a, Rav71, She86, Stu78, Sut74, Tan82, Tup82, Und79a, Ver89a, Web87, BEFH10, BB84, BBC88, BM84, Cal90, Car67b, Cat88, Cha88b, Cha90c, Cha94, CM90, Dav99, Duc84, Har82b, Hen84f, HRP90, Hul88, MMVZ90, MKW92, MR82d, MR82e, NC81, Ost84, RBM88, SO85, SP88, Smy89, Web84, Zan86]. Developmental [Bow87b, HWW86, Tei78]. Developments [Bro86a, R^{+}85]. Developpement [MdLD86, BBC88]. Devices [Hal86b, Kel75, Swi71, Ole84]. devil [Des94]. Devonian [Mor87b, Rud85]. Dewhurst [Smi75b]. Df [Cro70b]. DF1 [Che85, Hen84e]. Df120/$9.95 [Hen82b]. Df150.00/$26.50 [Hen82c]. Df150/$26.50 [Hen82b]. Df155.00/$21.25 [Rob80a]. Df189/$44.50 [Hom80]. DFL [Kni84a, Kni84b, Ped85, Bro86b, Bry78b, Bry78c, Can84, Dou84, Edg84, Fra87c, Gil84, Gou85, Gra86a, Hal86e, Hal86b, Jam86, Jar84a, Kni85, Mac78, Mil87, Mol84c, Mül89, Pic84a, Pic87a, Por86a, Sch84b, Sch84a, Sha89, Stu78, Web84]. DFl.135.00 [Goo88a]. DFl.160.00 [Kni89b]. Dfl.25.00/$12.95 [Dea83a]. Dfl.37.50/$14.95 [Gau83]. Dfl.55.00/$29.00 [Dea83a]. Dfl.65 [Sch83a]. Dfl.80.00/$34.95 [Gau83]. Dfl30./$14.50 [Law83]. Dfl60./$29.00 [Law83]. Dfl84.50 [Tur82]. Dharampal [Hil75]. d’Histoire [Bur74a, Bur74b, Cro68b, GG70, Por84a, Sch84c, Ano93, Han85a]. Diagrams [Sch79a, Sco67, Bec88]. Dialexis [Rav69]. Dialog [GB95]. Diálogo [GB95]. Dialogue [Ano05, Bru86, Che85, AHY90, Dra81, Hen84i, Nay85, PS84a, PS84b]. Dialogues [Hal68c]. Dialogus [Jon87, SSH85]. Diamond [GG74a]. diaries [DF98]. ’Diary [Gun84, DK88, PCA91]. Dibner [Pac71]. Dick [Pum84]. Dickinson [Kni75, Und66a]. Dickreiter [Mol75]. Diction [Lay65]. Dictionaries [Dur83, Fis83b, Fis83a, Lay65, Mor83a, Tur83a, Wea83, Wig69, Wil69]. Dictionary [All79, Dur83, Fis83a, Law89, Wig69, Wil69, DM96, Kni83a, M^{+}89, Mil90, O’C88, Ogi86, PB^{+}89, Gun84]. Didactic [Kni76, Mor91b]. Diderot [Dul88, OS06, dSR95]. Diego [Mac88a]. Diesel [Cah88, Tho87]. Dieter [Nor71a, Hil89b, Ree85, Wil86b]. Dietetics [Und75a]. Difference [Por87c, Lin90]. Different [Gou82, Hai82, Har82d, Yox82a, Gol92a]. Differential [Ser88, Gra85c]. Differentiation [Fra87c, Gra86a, Eng84]. difficult [WHS92]. Difficultés [CMDM82]. Difficulties [Hen86b, MR82f, MR82g, CMDM82]. Diffraction [Tur70c, Tur73d]. Diffusion [Lay65, Cro66b, Sch01]. Dijksterhuis [Ano65j, Pep71]. Dijon [BG^{+}92]. dilemma [Pou93, SM94]. Dilemmas [San87, Hei86]. Dilke [Has72]. Dimensions [Olb89, Sho89, Wil87c, Van85]. Diminution [Rap71]. d’influences [Ano90]. Dingle [Pye76, Sme66, Sme79]. Dione [Wil85a]. dioramas [Won93]. Diputación [Gli84]. Dirac [Hen89b, DD95, Kra90, KW87]. direct [AS89]. Directed [Rus88d, Roc88b]. direction [Cal63]. directory [HAB92]. Dirk [Sch84a]. Disagreeable [Bec75]. Discepoli [Hat87]. disciple [Ait68, Des94]. disciples [Web84]. Disciplinary [Sha79]. discipline [Ole91, Out84b, Tei84]. disciplines [Nye93]. Discontinuity [Bar81]. Discours [dSR95]. Discourse [Ait89, AW71, Buc68, vLMB88, Edw96, dSR95, Old72]. Discover’d [Arm69]. discoverer [SN93]. Discoveries [Hei76a]. Discovering [AT87, Bro89e]. Discovery [Can75, Dol70a, Fal87, For87c, Fox68a, Hen86b, Hum67, Hun89a, Lie86, Agr83b, Bar89a, Bar89b, Bli82, DTF97, Des92, MF86, MR82a, OP87, PB82, PB83, Pum89, Sch84d, Six87, Tha83b, Van93, Wal92, Wea90, Yox83a]. discussioni [Des89b]. disease [DPP88, EE95, Jor85b, Law83, Mic95, Nev84]. Disenchanted [Por89a, Sch88]. Disgraceful [Rus88b]. Diskussionen [Rus74]. disorder [Wei83a]. disorders [BP95, PB99]. Displacement [Tur74b, Sie91]. disputata [Law93]. Dispute [Tha87, Sec86, Tur77b, ZH92]. Disputes [Rap69, Sho88, EC87, ZH92]. Dissection [Law88, Ric87c]. dissent [AL87, Hen89g]. Dissertation [Dun69a, Sch69a, Sme71, Woo83]. Dissertations [B.69, Osp74]. Dissolution [Smi79]. distances [Hen91]. distinguished [Cha83, Huf85]. Distortion [Osp76]. distributed [Swi77]. distributor [Whi85a]. Diuinity’ [Ost89]. Diversity [Rei86]. Divide [Sto88]. divided [Ben87]. Dividing [Cha90c, Cha94]. Divine [Mar89, PW89b, Gol91, Jon88]. Divinity [Liv86, Dur85]. Division [Dav81, GG89, Kni83a, HF87, yL66, LY66]. dix [Sch01]. dix-huitiéme [Sch01]. Djen [Agr83b, Hal88b, Mol71, Pac72]. DM [Bra88, Bra87, Bro89a, Cap81b, Fig88, Fis78, GG75c, GG77a, Gra84c, Gra86b, Gra87b, Hal85d, Hen86b, Hen86a, Hen88c, Hom74, Jac89a, Von88, Mol72a, Mol78a, Mol86, Mol88, Mor84g, Nor72, Pag69, Pat84, Rav65, Rus74, Rus79a, Sch83b, Scr68a, Sme68a, Tur88, Wea80, Whi68a, Wil84b]. DM39.00 [Gra83b]. DM60/$33.00 [Har81]. DM78 [GG81]. DNA [Olb85, Yox83a, Yox83b, Kri82]. Dobbs [Hal77b]. Dobrzycki [Tho76a]. Dobson [Fre81b, Poy70]. Dobzhansky [Gid82, Smi76c]. Doctor [Twy88]. Doctors [McC77a, Pes84, Wea79]. Doctrine [Bar76c, Hah70, Bos83]. Doctrines [Hal72b, Hei76a]. Document [Per89b, Rei68]. Documentary [Bur87a, Ano87n, Bud84, RR81, Yox83b]. Documentation [Bro84d, Hun89a, OP87]. Documents [Fie81, Jen87, NBR^{+}95, vPvP90]. Dodgson [Abe86]. Doerner [Har83a]. Does [SM94]. Dogma [Rup89, Car88]. Dohrn [vBDGO93]. Doing [Rei87]. Dollars [Mac72]. Dols [Fre86]. Domb [Chi64]. Dominique [Stu87]. Don [Hal84e, Eng83, Mor84d, Ada89]. Donald [Aus86, Hei77, Hil86b, HS88, Kel75, Nor83, Olb70, Woo84, You66, Cah88, Hen88a, HS86a, Huf86]. Donau [Rus79a]. Donau-Verlag [Rus79a]. Dondi [Kel68]. Donley [Aas87]. Donna [Tei78]. Donne [Slo85]. Donovan [Cro89a, Hei76a, Woo83, Cha80a]. Dora [Men70]. Dordrecht [Bar83, Blo79, Bro86b, Byn80, Byn87, Can84, Che85, Dea83a, Edg84, Fra87b, Gau83, Goo88a, Gou85, Hal86e, Hal86b, Han85a, Hen82c, Hen82b, Hen84e, Hen87c, Hen88e, Hoc86, Jam86, Jar84a, Kel75, Kni84a, Kni84b, Kni85, Kni89b, Law83, Mil87, Mol84c, Mol89, Nay80, Ner89, Out84b, Ped85, Pic84a, Pic87a, Pin87a, Por86a, Sch84a, Sha89, Stu78, Tho76a, Tho76b, Web84, Whe89]. Dordrecht/Boston [Whe89]. Dordrecht/Boston/Lancaster [Byn87, Che85]. Dordrecht/Boston/Lancaster/Tokyo [Sha89]. Dordrechtand [Cha80b]. d’organisation [Hai80]. Dorinda [Cor81, Cro85]. Dorothea [Ano64l, Olb69, Und65]. Dorothy [All84a, Hei75, Sch69b, Und75b, Por86a, Sec88, Und74]. Dostrovsky [Gra84c]. Dottrina [Sch73a]. Double [Ros79, Har87a, Wil84a, Yox83a]. Doubleday [Far71b, Mor72b]. Douglas [Ano67i, Mor71, Tur74c]. Dover [Web87]. Dowd [Lev73]. Dr [Bro82a, CHG91, Por79b, Rus74, Yox79]. Dr. [Bau75, Lev84]. Drachmann [Ant63]. draft [McG70b]. Drafts [Byn80]. Drake [Gab83c, Nay75, Nay85, Rav72, LS90]. Drakenstein [Mül89, Hen86d]. drapery [Cha84]. Drapière [Hil86a, Cha84]. drawing [Bec88]. Drawings [Rud80, Cri66, Nor89]. Dream [Hal69a, Hal66a, Kit92, SC88]. Dressenhaus [Fee77]. Driffield [Lay85]. Drift [Fra78b, Mar74]. Drifting [LeG88, Ste90b]. dritte [LS76]. dritten [Cap81b]. drive [SM94]. Droz [Har86, Kel78]. Dualism [Mac85b, Whe83]. Dualität [Sch89]. duality [Sch89]. Duane [Smi74b]. Dubbey [GG71]. Dublin [Ham84, Kel71, Kni89c, N^{+}88, O’H86]. ducal [But96]. Duchenne [EE95]. Duchesneau [Dou84, Sch84b]. Duckworth [Dyb78, Hut88, Joh79, Joh83, Van86, Whi85b]. Due [Mol84b, SB92, Sch81b]. Duhem [Mil87, Out80b, Jak84]. Duke [Nad83]. Dulong [Fox68a]. d’un [Dug77, SG80]. Duncan [Jar84b, Rus79b, Sme71]. d’une [BV93, Sch72]. Dunlap [Irw83b]. Dunn [Hig65]. Dunsink [Way87]. Dupree [Web65a]. Durant [Liv86]. durch [Wes70]. Durham [Nad83, BK75]. During [Hom86b, GGSGB94, Nev78, Sch84a, SP88, Str87, Swa83]. Durling [Sch68a]. Dürr [Kra86]. Dutch [Cro70b, Mül89, Sch84a, Sme68b, Bur99, Dij90, Hen86d, Und63]. Dutrochet [Pic78]. Dyeing [MC63]. Dyes [MC63]. dyestuffs [Tra93]. dynamical [Bar89a, Men75, Smi76b, Smi89a]. Dynamics [Ilt77, Whi72b, Bar89b, CD81, Gou95, Gra84c, Nye93]. dystrophy [EE95].

E. [Hal87a, MdLD86]. e.V [Lil85]. e.V. [Har87b]. Each [Ant64, Dun87, Fox70, Swi71]. Eagle [Byn78]. Eales [Fre70]. Earle [Ano64l, But65]. Earliest [Coc75, McV80, War88a]. Early [BH87, Bow89a, Bro85e, Bry78a, Bur86, Can80c, Cha77b, Ear70, Ear87, Fel89, For71b, Fre70, Geo87, Got86b, Got67, Gou68, Gow69, Gur78, Hac88b, Hal85b, Hen88d, Hen88c, HS83, Hop76a, Hop76b, Huf86, Koh78, Law80b, Mar64, Mol74, Nay80, Ora80, Ost89, Pac74, Pat86, Pay74, Per70, Rap69, Ric82, She86, Sho87b, TP66, Tei70, Til75b, Tur68, Tur70c, Tur73d, Tur78c, Wal69, Was69, Bad83, BA91, Buc89b, Can84, Car67b, CMN95, Col87a, Cro90a, Deb91, EJR94, EJR95, Eis83, Eis93, FJ93, Fox88, Gin84, Gou84a, Gou95, Gra84a, Had94, Hen89f, Hen84m, Hey95, Huf93, Huf95, Hun82, Hun89b, Jac83, JO81, Jor85b, Kal95b, Kro91, MR89, Mil89, Mor91b, Mor83a, Mor84e, Ost84, Phi87, Poc88, Poo98, PJP91, Reh83, Rud92, Rus86a, Rus86b, Sir90, Sor83]. early [SCS^{+}87, VoC89, Whi64, Whi85b, Wil84a, Wil95, Yeo93, Got86a, Sin78, Stu86]. Earth [Ait86, Bow83a, Bro83a, Bur80, Dea83b, Kni83c, Mor72a, Nev78, Rap79, Rap86a, Rud70, Rud80, Rud83, Rup89, Sec83, Sec84, Smi76a, Tei83a, Tha83a, Tha83b, Tha85, Yox83a, Yox83b, Car88, Duf81, Gle82, NPH92, Por83b, Ros84b, Car88, Gra84b, Ito88, Mac90, Rap73, She70, Tur74b]. Earth1 [Hen91]. Earthquakes [Rap86b, Old72]. Easlea [Sha81]. East [Ear70, Nak75, Nee69, Sec83, Tup82, MM10]. Eastlake [Wel70]. Eastwood [Hac76]. easy [Sha93]. EC4V [Hal88a]. Eccentric [Tei83a]. Échanges [Ano90]. Eckhard [Hut89]. eclettico [Sch81b]. eclipse [Bow83b, Nor85b, Sta94]. Ecole [Out88a, Out88b, Fox87b, Sme84a, GG86a, GG86b, Lan87]. Ecological [Fal89, MS87]. écologie [Ege89]. ecologist [Duf81]. Ecologists [Nic88b, Cro91b, Söd86]. ecologization [Nic88b, Söd86]. Ecology [Bra89b, Mül87, She86, Cit90, Hag92, McI85, Mit92]. econometric [Mor90a]. economia [Por79a]. Economic [BEFH10, Hac89, Mar73, Rat68, Ben86, EM94, Gou95, II90, Mir94]. Economica [Gua89]. Economics [Fal89, Tre79, MS87, Mir89, Sch90]. economies [Cou96]. Economy [Rat68, Rud74, Jor85b, Wei92]. ecosystem [Hag92]. Ed [Ait73, Bad71, Bal83, Bar73a, Bar76b, Bib74, Bos70, Bro73a, Bry76, Bry78b, Bur74a, Bur74b, Bur71, Byn73, Byn74, Can74, Cap81b, Cha77a, Col81b, Cor81, Cro66b, Cro70a, Cro70b, Deb66, Dol70b, Dol71b, Dol73b, Dol74, Dur77, Dur80a, Dur81, Far67, Far69, Far71a, Fig77a, Fox70, Fox71a, Fox71b, Gel76, GG75a, GG75c, GG77a, Gur78, Hac74, Hac76, Hac77, Hei68, Hei71b, Hei73, Hei76b, Hil71a, Hol71, Hom74, Hop73, Hop75, Kar76, Kel73, Kni69a, Kni71b, Kni81, Koh78, Lay74, Loh75, Lor73, Mar73, McC69, McG70a, McV75, Mor71, Mor72b, Mor75c, Mos77a, Mos77b, Mus74, Nor69a, Nor70a, Nor72, Nor73a, OM75, Olb71a, Osp74, Phi69, Por81, Rob80a, Rud69, Rud71, Rus69, Rus72a, Rus74, Rus77b, Sar81, Sch72, Sch74a, Sch76, Sha73a, Sha77b, Sin77]. Ed [Smi74b, Smi76c, Ste75, Str75, Swi71, Tho76a, Tho76b, Til76, Tur70c, Tur73d, Tur73c, Und75b, Web75, Web69b, Whi73, Whi77, Wig69, Wig74, Wil69, Win75, Woo80c, Ait89, Asp88, Ben89, Bow83a, Bow86, Bow88a, Bra87, Bro86a, Bro89a, Bro89e, Bro84c, Bro84d, Bry74, Bus83, Car67a, Col83, Col88b, Cro89a, Cro89b, Cun88, Dea86a, Dea86b, Edg84, Edg87b, Fis83b, For87a, Fox87b, Fre82, Gid82, Gil88, Gli84, Gol88b, Goo85b, Goo88a, GG87, Hah70, Hal85d, Hal86a, Har82d, Har83b, Hen82c, Hen82b, Hen82a, Hen84e, Hen84g, Hen88d, Hen89b, Hen84j, Hen87c, Hom86a, Jam87b, Jar83, Kei89, Kel75, Kil84a, Kni83a, Kni83b, Kni85, Kni88b, Kni89b, Lew85, Liv86, Mac88a, McN83, Mol85, Mol88, Mor83a, Mor83d, Nev84, Olb82, Out82a, Pat84, Por79a, Por85a, Por87b, Rap71, Rav84]. ed [Ric89, Rus89a, Sha87, Smi89e, Smi89c, Stu85, Tee84, Tha83a, Twe89, Wea83, Whe89, Whi84, Wil87d, Wil89b, You83, Yox83a]. ed-in-chief [GG87]. Eddington [Huf85, Whi68b, Cha83, Kil94]. Edelstein [Phi69]. Edens [Gro95]. Edgar [Hal88a]. Edge [BS77]. Edgeworth [EM94]. Ediciones [Gou84b]. Edidit [Mer78]. Edinburgh [Ant64, Aus86, Bro86e, Bus85, Fis81a, For80, Fre78, Hei76a, Hen86b, Hil86b, Jam87a, Lar74, Law85, Lie86, Nee78, Nor83, Por85b, Poy70, Rud80, Sha78, Smi85b, Tha83b, Und80, Wes74, Wil65a, Woo84, Eme79, Eme81, Eme85, Eme88, For73, Mor72d, Sha74]. Edison [Bow87a, EJ^{+}07, ER93, EJR95, FIF86]. Edited [Arm67, Bar77, Blo79, Bro67, Bro78a, Bro78b, Buc69, Bur80, Can80b, Chi64, Chi65c, Coo81, Cow68, Cro67, Dor79, Dor80, Dyb78, Fin68, For68, For80, Fox79, Goo81a, Got66, Gow78, GG77b, Hal68b, Hal85a, Hal85c, Hea64, Hei74a, Hen81, Her80, Hoc86, Kni69b, Kuh68, Law80a, Mac79, May67, McG68, Mol79, Mor77b, Nay80, Nor68a, Nor71b, Pep72, Rat68, Rav69, Sch70, Sha79, Sin81, Sme66, Smi66, Ste62c, Und66a, Und79b, Wal77, Wea79, Web65a, Web68a, Wei85a, Whi85a, Wig78, Wil85a, You66, Fie81, Geo78b, Jac85, Nic85, Nor68b, Smi87a]. Éditer [Dul88]. Editing [Bur87a, Mon87, Ano87n, Sta87]. Editio [Rav69]. Edition [Bar73a, Buc68, Cro71, Dol71a, Had63, Hal77a, Hew66, Lin86, Men67, Mol81, Mol89, Olb71b, Rav66b, Sch85a, Sme69, Smi87a, Web69a, Whi73, ZH92, BL83, Car69, CF86, Cha78a, Cla67, Coc75, Dol73a, Eas86, EiT92, ET92, Fee77, GG77a, Hen89a, McC69, McG68, Mol78b, Nor72, Ora74, OH85, Osb85, Rat66a, Rat69, Rob89a, SG80, Sme71, Tak92, Tur72a, VoC89, Web65b, WW88b, You83, Yox83a, GG70, SG80, Sme84a]. Editions [Cro88, Hut80, Jen89, Jor88, McC80, McC83, Osb89a, Out82a, Pau76, Stu87, Fox87b, Sch72]. Editor [Wal77, Gun84, Hut89, Lay85, Mai85, Rud80, Ano88p, Hal80b, Pir83]. Editore [Dur80a, Out80b]. Editorial [Ano82k, Ano83k, Ano84k, Ano85k, Ano86j, Ano86k, Ano87o, Ano88n, Ano89o, Goo89b, Ano88q]. Editors [Arm62, Arm64, Cas85, Hut89]. Editrice [Sch73a, Sch80c]. Edizioni [Sch81b]. Edleston [McG71b]. Edmond [For66]. Edmund [All84a]. edn [Bos72, Bro83b, Goo89b, Her80]. Eds [Kra86, Aus86, Bow87b, Bow89c, Bra89a, Bro89a, Bud84, Can84, Dea83a, Des88, Dun86, Dur83, Fie83, Fra87a, Fre86, Gol87b, Goo85a, Hal84a, Hal87b, Hal88a, Hat87, Hen87a, Hen89b, Hen89g, Ink88, Jam87a, Kni83c, Kni89c, Kni89d, Law85, Mor83c, Ner89, Pat86, Pic84a, Pin87a, Por87a, Rud86, Rud88b, Rus87b, Sch83a, Sch83b, Sec85, Sha89, Sha83a, Sho88, Smi83, Sti89, Wea84, Wil89a, Woo84, Woo89]. Edsall [Olb82]. Eduard [Tei83b]. Education [Cro68a, For87a, FG93, Hol71, Hut88, Lay85, Sch68a, Dau98, GG84b, Hal84c, How82, Mor82a, Sta94, Sta90, Wei82, Cro68a]. Educational [Asq80, Gre63, Hen82c, Stu71a]. educator [Gol92a]. Edvard [Lin90]. Edward [Ait86, Arm66, Byn80, Chi64, Ear87, Eyl68, Hac75, Hal69a, Hen83d, Kni89e, Rus86a, Sch76, Tur83a, Wal88, Ait86, Fis91b, Kra84a, Mor87a, Nic85, Nor85a, She95, WBP97]. Edwardian [Tro89, Ubb63]. Edwards [Mah89, Wei86]. Edwin [Byn73, Grm69, Jac85, Smi75b, Bru88a, Chr95, Har89a, SN93]. eeuw [Bur99]. Effect [Bro76, Cra96]. Effective [Wei80, NPH92]. effects [Buc94, MJ89, Smi94c, Woo85]. Egerton [Nic85]. egg [Cor97]. Egypt [Fre86]. Egyptian [Cla99a, Cla99b, Cla95a, Cla9 , Cla2 ]. Ehrlich [Wei86, Bäu84]. Eichholz [Got66]. Eighteenth [Bro78a, Bro78b, Bry74, Bry77, Can80c, Cha77b, Eag78, Edw78, Eme87, Fis81b, Gol87b, Hac75, Hai77, Hal77b, Hil75, Hom80, Jor78, Law80b, Mar73, Nay77, Nev83, Por79b, Rau75, Rus77c, Sch81a, Sch78c, Sha86, Sha77b, Ste62d, Und66b, Wil65b, Ben79, BPH85, Gab83b, Gli83b, McC85a, Rap82, Sch85b, Wil82a, And84, Ano04, BB92, Bar73a, Bro73b, Bru72, Bur74a, BP02, Can77, Car69, Cro70b, Duf81, Dun88, Dur81, Fin71, Fis81a, For69b, For72, Fox71a, Fox87a, Fox88, Fri74a, Gol88a, Hac74, Hac77, Hah73, Hal80a, Han70, Har81, Hei71a, Hey95, Hol89, Hom74, Kni69b, Kni69a, Kni72, Kni74, Kro91, Lar74, Law80b, Lev90, Lin96, Mac72, Mal05, Mas98, McG69, McG71a, Men74, Mor69a, Mor71, Mor72a, Mor72b, Mor73a, Mor73b, Mor74a, Mor81, Olb72, Ora77, Out80a]. Eighteenth [Rap71, Rap78, Rud80, Smi73a, Sne74, Ste70, TBDA10, Tur72a, Tur73c, Tur81, WW88a]. Eighteenth-Century [Gol87b, Sha86, Und66b, Wil65b, Ben79, BPH85, Rap82, Wil82a, And84, BB92, BP02, Fin71, Fox88, Gol88a, Hol89, Lin96, Rap78]. Eighth [Bow87b, Chi65a, HWW86, HS63]. Eilhard [Sch92b]. Einaudi [She85]. Einleitung [GG77a]. Einstein [Aas87, Bro72a, Goo75, Goo82, Hei68, Hen82c, Hen86b, Hen86a, Hen88d, HP85, HE^{+}82, Kil84a, Von88, PHvMW93, PvM85, Pin87b, Red84, Sho89, Wil84c, Wol87, BH87, FD85, Gol70, HP85, MR82f, MR82g, Pai82, PHvMW93, PvM85, Pye85, Ros82, SCS^{+}87, Sta87, Wol87]. Eire [HAB92]. Eisenstein [Hal85b, Kni80b]. Elborg [Gel76]. Eleanor [Har76, Ste62b]. Elective [Dun69a, Sme71]. Electra [Tur72b]. Electric [Bow87a, Fri87, Buc94, FIF86, Wis85]. Electrical [Bry77, Hom72, Hom80, Mar62a, McC69, Wil68, BC92, Tun92, BaJANJ84]. Electricity [Can80c, Hei71c, Hom80, McE79, Bea97, H^{+}82, Hen84d, Hom83a, Luc90, Per92]. electrification [Coo92]. Electrifying [Nye90]. Électriques [Hom86a, BaJANJ84]. électrodynamique [Blo82a, Hen84b].

Electromagnetic [Hun89a, The87, Buc85d, Hen86c, OP87, Sie91]. Electron [Fal87, Spr97, Dah97, DTF97]. Electrons [Fal87, ES90]. Electrostatic [Fin71]. Elek [Web76]. Élement [BV82]. elementis [BJ91]. Elements [Ano92, Bro73b, DeK73, Far65, Far68a, Jam88, McG69, Mol85, Sme70, GG70, BV86, Ano92, Ord76, BEoCF92, ECL92]. Elgin [CHG91]. elianism [Mol72b]. Elias [Pic84a, Web68a]. Eliot [Por85b]. Elisabeth [Goo85a, Kra86, Law80c, Tre88]. Elite [Har79b]. Elizabeth [Far71b, Kni80b, Hal85b, Han85a, Rec86]. Elizabethan [Pum88, Par87]. Elkana [Can75, Cha77a, Kil84a, Law80a, Pic84a]. Elliott [Duf83, Kni83a, Scr68b]. Elliston [Bro87]. Elmer [Bow74]. Éloges [Out83]. eloquence [NS02, NS02]. Elsevier [Eyl68, Far71a, Fis74, For72, Fre82, Gin74, Hea64, Kni71b, Lev73, Nev78, Pac70, Rap75, Sch73b, Sin74, Sme70, Whi72b]. Elton [Cro91b]. Elusive [Sey77, Pin83]. embryo [Duf81, Dun90]. Embryology [Bow87b, HWW86, Les83]. Emergence [Ant64, Bro84a, Far76, Nee85, Ric82, Tha66b, Hir83, Out84b, Wei83b]. Emergent [Bli92]. emerging [NPH92]. Emerson [She95, Hen85a]. Emerton [Hal86d]. emigre [AS96]. Emil [Bro89a, Bro89a]. Emilie [ZH06]. Emilio [Buc73b]. Eminent [Wal77]. Emmy [Gra87b, vdBLW85]. emphasis [Law93]. Empire [Cro87a, HJF89, Kum91, RH86, Vuc84, G^{+}89, Mac88b, MR96, Sta89a, Cit90, HJF89, Smi94b, GS94]. Empirical [Mac85b, Mor75c, Whe83]. Empiricism [GG74a, Por83a, Wil91b]. empiricist [Goo89a]. Empirisme [Dou84, Sch84b, Duc82]. enclosed [Sch69a]. Encounter [Rus87b, Bee96, LN86]. encounters [LSS90]. Encyclopedia [Dol71a, RM96, Ste62c, BP93b, dSR95]. Encyclopédie [dSR95]. End [Bra87, Bro81, Eme88, Hen89d, Gal87, Hai82, PP95, Slo85, Zam86]. Endeavour [Ste70]. endless [Her90]. Ends [Whi89a, Red87]. Enduring [Kni86b, Gje84]. Energy [Can75, Dau75, Goo81b, Har82b, Hen84f, Kei84, Tre79, Can93a, Car67b, Her90, KH95, MS87, Row90a, Smy89]. Eng [Huf86]. engagement [BC98]. Engel [Mor84d]. Engelhardt [Sho88]. Engelsman [Fra87c, Gra86a]. Engend’ring [Geo87, Rus86b]. Engine [Buc85b, Law65, Ler85, Pac74, Por82]. Engineer [Bow74, Bal84, Gol92a, Vig82]. Engineering [Bro81, Buc85c, Bur71, Car66, Goo83b, HS86a, Pac72, Hal84c, Hil84, Pep89, Pet94, Wei82, Wri94, Buc69, Fox69a]. Engineers [Car67a, Gre86, H^{+}82, Hen84d, Kni83a]. Engines [Dau74, HL86, Wor88, Lin90]. England [Bro75, Cun88, For85a, For87b, Gab83a, Geo87, Gou82, GG84b, Hac88a, Hen84l, Hod89a, Kau84, KAZ92, Lay74, Lay85, Lud80, Mac79, McG68, Mor84c, Ora74, Por77, Por79b, Por81, Pum88, Rob75, Sco67, Sho86a, Smi74a, Wal81b, Web72, Whi77, Wil65b, Win75, Woo89, Yat72, Bru88a, Con87, Gou95, Hey81, How82, Ito88, KK87, Mac82, Num87, Olb87, Opp85, Par87, Poo98, Por89f, Ric88b, Rus86b, Sha83b, Sht96, Ste91a, Thé66, Tho74b, Wil94b]. Englewood [Rav73]. English [Arm70, Ben82b, Bry81, Dol73a, Fis81c, Gol88b, Gra87b, Hal80c, Har76, Hen84c, Hen88d, Jos67, Mol78b, Mol81, Nay85, Nor71b, Rud84, Rup83, Sch69a, Sch85a, Sin76, Tur78a, Wal74, Woo80a, BL83, BM89, Cop92, Dra81, EiT92, ET92, Gab78, GG65, Mac88c, RG65, Tak92, Vic87, Woo82]. Engraving [Lan87]. Engstrand [Gli83b]. Enigma [Lan82, Whi84, Zam92]. enlightened [Wil84b]. Enlightenment [Ano07, Byn72, Gil84, Hah73, Hei76a, Hen84c, Jor87, Mor73a, Por77, Por84b, Por89g, Tur73a, Woo89, Ber93a, Gol92b, Hen84m, Sta94, Sut95, Woo84, Das88, Dou84, Eme88, Gib84, Han85b, Hey83, Jac83, KK87, LL09, LP04, MF92, Out95, Por84a, Sch84b, Sta91, ZH06]. enough [Ear89]. Enquiries [PL72, BGH^{+}87, CGBE95]. Enquiry [Mon78, Ben91, BDH96, For86]. Enrico [Buc73b, Dor80, Smi82a]. enseignement [Gua89, HJF89, Cro66b]. entangled [Hag92]. Entdeckung [Mau87, Sch84d]. Enterprise [Edg87b, Mor80, Liv93, Hol89]. enters [Cer89, Mor91b]. entertainment [Sta94]. Enthroned [Wes79]. enthusiasm [Hey95]. Entre [BBC88, Ano90, BaJANJ84, Fox87b, Gli84, Gre86, Hom86a]. Entrepreneur [Nev83]. Entstehung [GG81, Mol86]. Entwicklung [CM90, HRP90]. Environment [Mac80, Luc90, MS87]. Environmental [Bur80, Bow92, Sho93b]. environmentalism [Gro95]. Epic [Hal88b, NPYGDL86]. Epicurean [Osl91]. Epicuro [McG68]. Epicuro-Gassendo-Charltoniana [McG68]. Epicurus [LL09]. Epidemiográfico [Kni88c]. Epidemiology [Fel89, Col87a]. Episodes [Rap75, Whi89b, Het88, Liv93]. Epistemologia [Out80b]. Epistemological [Can84, Des89b, JO81]. epistemologiche [Des89b]. épistémologie [PP95]. epistemology [PP95]. Epistolario [VG98]. epistolieres [SGŚ00, SGŚ00]. Epochen [Sch83b]. Époque [Mor89c]. Equality [Tan62]. Equations [Hen86b, Ser88, Gra85c, MR82c, MR82e, MR87b]. Equilibrium [Hof87, II90]. Equivalence [Ber93b]. Era [Bod84, Fox71b, Pic80, Wei85b, Fri81, MF92]. Erasmus [McN83]. Eretico [She85, Red83]. Eric [Bar73a, Bry78a, Bry83, Hac76, Hom74, Mor71, Sha77b, Smi71, GD85, Rup88]. Erich [Sch83b]. Ericsson [Dau74]. Erik [Rus67]. Erlbaum [Fie86]. Erlenmeyer [Bro89a]. Ernest [But64a, But64b, Dol73b, Nor77]. Ernst [Mai85, Web65b, Bow83a, GG74b, Wei86]. ero [Kni88c]. error [Hon89, Pet94, Por85d]. erste [GG77a]. Erster [Bos72]. Erträge [Sim84]. eruption [Woo85]. Erwin [Hen88c, Und75a, MR87a]. Eschatology [Hen82d]. Espa [Kni88c]. Espaäna [Rey90, Lan86]. España [Gli84, Goo89b, Lan89]. español [Gli84]. Española [Gli84]. Esquisse [Sch72]. Essai [BV82, BG80, Sme84a, Sin91]. Essais [Sho89]. Essay [Ano66h, Blu79, Buc66, Car74, Col72, Cro69, Gue68, Hal72a, Hei71a, Her73, Mor72c, Pag67, Rat65b, Und64b, Und64c, Und64d, Whi89a, Whi65, BEFH10, MJ89, May66, Red87, Sin91, Wil90a, Col83, Jar83, Kni83b, Smi83]. Essays [Blo79, Bow86, BM76, Bur80, Col88b, Cro71, Cro89a, Dol73b, Dur79a, For87a, Hac82a, Hal80a, Hen87a, Kar76, LS90, Lew85, Liv86, McG66, Ner89, Nor70b, PP87, Ric89, Roc88b, Ros79, Rus87b, TT86, Und64a, Wal81a, Web65a, Wes89, Whi70b, Wig74, Ait85a, ABR93, AB89, BA91, Col83, Don88, Dur85, Fur89, Gau83, GJ85, GR88, Hen85b, Hol86, Jam89, JK84, Kni83b, Kum91, LN86, MH88, Mol84c, Nor68b, PW89a, Rid83, SG90, Sho93b, SY96, Sto83, Sto84, VoC89, WHS92, WBP97, vHC95, Tra87, Wig78]. essor [Hut80]. Establishing [Hun89b]. Establishments [Mor77b, Pic84a, EMW82]. Estes [Gab83a]. Estrange [ABR93]. Estudios [Rey90, Lan89]. etc [Aas87, Buc85a, For73, Rud86]. Ethel [Fie86]. Ether [Can70, Wil72, Smi83]. Ethereal [Mar74]. etherians [Can80a]. Ethics [WM85, SG96]. Ethnographic [Lew85, Sto83]. Ethnography [Fan83]. Ethnology [MD90]. Ethology [Asq80]. ethos [Pau84]. Etiology [Wei85a]. etrenne [Sch78c]. Ettore [Rav66a]. Etudes [B^{+}89, Fox87b, Hal68c]. Euclid [Mol85, BEoCF92, EiT92, EMMS93a, EBT96, ECL92, ET92, EMMS93b, Tak92, The78].

Eudiometer [Far63]. Eudoxos [Scr68a]. Eugene [Hen89b]. Eugenic [Cro89b, Raf88]. Eugenics [Bar89c, Har89c, Hoc88a, Maz92, Olb70, Olb77, RH89, Wei89, WKB88, Mac78, Maz92]. Eugenik [WKB88]. Euler [Bar73a, Dub72, Gra85b, Gau82, Hak95]. Euler-Mayer [Bar73a]. Eurich [Web68b]. Europa [Sch81b]. Europe [Cor81, Fre82, GGR96, Hal85b, Hut88, Mon78, Rat68, BS91, Bro89f, C^{+}92, Eis83, Eis93, FJ93, FG93, GR91, GJ86, Had94, HW87, KRS94, Mar10, PB94, Tra93, vPvP90]. European [BVA95, Col81a, Jor85a, Nut83, Sha78, Tup82, Whi85a, Ano90, Chr93, Des92, Dun90, Lin92, MD88, MM10, OS06, Osl91, RT92, San88, Tyt82]. européens [Ano90]. Eurospan [Whi85a]. Eva [Ano67j, Sme66]. evaluative [CT95]. Evans [Jud88]. Evanston [FF66]. eve [Edg91, Cor97]. Evelyn [Duf81, Nor82]. Events [Ano64h, Hah70]. Everett [Bod65, Pic84a]. everyday [BC89, CB^{+}90, G^{+}89]. everyone [CP93]. Evidence [Gou68, She70]. evils [MJ89]. Evolution [Bar73b, Bro85d, Bro86d, Fis74, Geo66, Har89c, Hul67, Liv86, Mac80, Olb71b, Por78a, Tha74, Tur69, Und66c, Whi88a, Bas88, Bli92, Bow83b, CD81, Cha88b, Des89a, Dur85, Gra84c, Kal95b, Lau87b, Maz95a, Nor85b, Num95b, Ray87, Ric92b, Tho89, Whi88b, Bro84b, Smi74c, Lau87b, Osb89a]. Evolutionary [Bar73b, Bow87c, Bow89d, Jor79, Ric82, Ver89a, Bow83a, Bow89b, CS94, Cor88a, Hul88, Pit93, Pro86, Rid83, Sar92]. Evolutionism [Ric87a, Bec92]. évolutionnisme [Bec92]. Evolving [Sal85]. Exact [Nor70b, Whi70b, Pye93, Sch75, WHS92]. Exakte [Nor71a]. Exakten [Mol86, Pag69]. Examination [Gou68, Men75]. Examinations [Lay85, Mac82]. Examined [Rus77c]. examining [WT82]. Example [HM66]. exchange [Ger89, Had94].

Exchanges [Ano90]. excluded [Can89]. excluding [Hac82b]. exemplified [Old72]. Exhibiting [Bea97]. exhibition [AL87, Ben79, Ber95, Hen89g]. Exile [Kra86]. exotic [Osb94]. expand [TFC93]. Expanding [Rat68, BH84b, Gou84b, Smi82b]. Expansion [Bro79, Hil71a, All82, Gro95, Pye93]. expeditions [Gli83b]. Experience [Edg87b, Pel85b, Por78b, Ait68, Gra98, Hun89b, San91, NS02, Out80a, May66]. Experiment [Fra87a, Goo90, LS90, Nay85, Pin88, AH85, Buc89b, Dra81, Fra86, Gol70, Hal90, Hal91, Jam89, NS02]. Experimental [Hof87, Jac85, Jam88, Jon87, Til75b, Web69b, Hal91, Hen83c, Hon89, Kre90, NC81, Pin83, SSH85, Tur90]. Experimente [Pag69]. Experimenter [Tre69]. Experimenters [Hop72]. Experiments [Bar91, Fri91, Hen89d, Mar74, Rob86, Gal87, Goo83a, vGA94]. explained [Whi72a]. Explaining [Gie88]. Explanation [Har75, Hes74, Men65, Mol74, Kra84a]. explanatory [Jar84a, Sch69a]. Exploration [Col87b, Hol85b, Lev93, Mol68, Sta89a]. explorations [Kni83a]. Explorer [Kni83a, Kni86c, Nev83, Ald85]. Exploring [Dur80b, Huf91]. explosive [Mei91]. exposures [Cau89]. expulsion [Dei92]. Extent [Kni89d, MR88]. external [Ale85]. extra [Tro89]. extra-ordinary [Tro89]. extraits [Dug77]. Extraordinary [Fri74a, Dol83, Wil90b]. Extraterrestrial [Sch87, Cro86, Dic82, Pum84]. Extreme [GG89, HF87]. eye [Tur94]. Eyles [Mor72a, Tor80, Tor87].

F [Ait71, Arm69, Bow89a, Bow88a, Bro85e, Buc85b, Car69, Car74, Cri66, Cro69, Cro71, Dav71, Dur83, For68, For80, Gol87b, Gou85, Hah70, Hal68c, Hal88a, Hen84b, Hew66, Kni80a, Law83, Mah89, Mar74, McG71b, McL86, Nee85, Nor68b, Nut83, Olb71b, Rud69, Sch71, Sme66, Sme71, Tur87a, Und64a, Web66b, Wel70, Bal84, Fox71a, Ait68, Cro66b, Cro68b, Hal68c]. F. [Can86, Hal86e, Kni89b, Por89b, War89]. F.B.A [But65]. F.R.S [Chi65b, For85a, Sha73b, Sin77, Wal88, Web71a]. F.R.S. [Gun84]. Faber [Nor68b, Rav66a, Woo80a]. Fabric [Edg87a, RT92, Sta85]. fabrication [Cha90a, Cha90b]. Fabrics [Tei78]. Face [Ros79]. Faces [Jor85a, Sho89, Dob91, Tyt82]. facs [Wes70]. Facsimile [Bar75, Bro83b, Buc68, Can74, Col76, Fee77, Rat66a, Rat69, Sch69a, Sme69, Sin77, Til75a, Tur74a, Web65b]. facsimiledrawings [Rud80]. Fact [Col81b, Smi94a]. Factor [Bry71]. factors [Car67b]. facts [Irw82]. failings [Maz92]. failure [Hen89f, Lin90]. Fairfield [Rai91]. Faith [Dun86]. Faivre [Sch83b]. Fall [Cun78, Mor74b, Nay74]. Fallgesetz [Mol72a]. Falling [HM66, Hum67, Dra69]. Fallstudie [Mei91, Sch84d]. Fälschung [Wol87]. Fame [She70]. Families [Eng84, Fra87c, Gra86a]. Family [Bow88a, Cro89b, Raf88, SL87]. fang [Agr83a]. Fanlac [Jor83a]. Faraday [Col76, Goo75, Gow73, Hei73, Hen87a, Kni65, Wil68, BS91, Can89, CGJ91, Can91, FTG91, Goo80, GJ85, Hei71c, Lev68]. Farben [Sto88]. Farbenindustrie [Mor84g]. Farber [Out84b]. Farina [Sch79b]. Farleigh [Kni75]. Farley [Ree80]. Farmer [Bro86e, Dem83]. Farrar [Car78]. Fascicule [Cro88, Gol83b]. fast [Adl87, Fig88]. Fate [Sch80b]. Father [CMDM82]. Faulhaber [Sch93, Sch93]. Faure [Jor83a]. Faust [Hay94]. Fauvel [Bra89a]. Fawcett [Law83]. Fay [Hal88a]. Fayard [Por89e, Sho89]. Faye [Ora80]. fear [Wea88]. February [Cha80a, Chr93, Sme78b, Yox79]. Fechners [Hei93, Hei93]. Feeble [Nol95]. Feeble-minded [Nol95]. feeding [Wie93]. feeling [Adl87, Smi94a]. Feferman [GG87]. Feisenberger [Wal77]. Felix [Ste62a, Dug77]. Fellows [Gou84a, Hun82]. Feltham [Hen84c]. Feltrinelli [Out80b]. Femmes [SGŚ00]. Fenton [Aus86]. Ferdinand [For67b, Nor89]. Ferenc [Kni67]. Ferguson [Tur73c]. Fermat [Gab75, Gra86b, Rav65, SO85]. Fermi [Buc73b, Dor80]. Fermilab [BDH89]. Ferran [Mal95]. Ferrier [Lie85]. Fertig [Web72]. Fever [Fel89, Wei85a, Col87a]. Feyer [Und74]. Feyerabend [Blo79]. FF [Ben89, Cro88, Ege89, Jor88, Osb89b, Out88a, Out88b, Por89e, Sho89, Stu87, Fox87b, Osb89a]. FF110.00 [Jen89]. fiasco [Hui92]. Fiction [Por85b, LSS90]. fictions [Sha98]. Field [Lin89b, Mos77a, Tur88, Goo80, Hen86c, The87]. Field-Book [Mos77a]. Fields [Tei78, Bee96, Goo75]. Fieldwork [Lew85, Sto83]. fifteen [H^{+}82, Hen84d]. Fifteenth [Hut88, Rya70, GJ86]. fifteenth-and [GJ86]. Fifth [Mor77b]. fiftieth [Por85a]. Fifty [Bro87, Smi87c, All86, Hea86, Hea90]. fig [Gre63]. Figs [Fox69a, Fox70, Hal88a, Sti70, Ait68, Ait71, Arm69, Arm70, Bro71a, Dol70b, Fig88, For69b, Han70, Kel68, Sty68, Web70, Whi70a]. figures [Cal63, Chi64, Chi65b, FF66, For66, Sme66, Sme70, Ste62a, Ste62c]. Figurovsku [Bro89b]. Fin [TP90, Hal72b, PP95, Har89b]. Finalization [Hoc86]. Financial [Ano63g, Ano64n, Ano69g, Ano70g, Ano71g, CM64, CC65, CH68]. Findlay [Dol71a]. Finn [Bow87a]. Finocchiaro [Hes74]. Fire [Smi87a, CF86, New94, She95]. Firenze [Dur80a, Por79a, Sch81b, Tur72b]. Firmament [Arm70]. First [Ano71h, Arm66, Bar73a, Byn80, Gou85, Lud80, Mol85, Ord76, Sme70, Til75a, Wil62, Coh84, FFLW97, Grm90, H^{+}82, Hen84d, Her65b, Mac88a, Olm92, Poi04, Wal88, Web65b, WH91]. Fischer [Har87b]. Fischler [GG89]. fisheries [Smi94c]. Fishes [PR95, Tha83b]. fishing [Smi94c]. Fishman [Und66a]. Fisicomatematica [Mid75].

Fiske [Woo85]. fitness [Coo84a, Who82]. Fits [Sha93]. Fitzgerald [Hei73, Hun88]. FitzRoy [Bur86]. Flammarion [Out80a]. Flamsteed [Sha77b, Wil85a, FC82, FFLW97, GW74]. Flash [Dah97]. Fleck [Löw90]. Flek [Col81b]. Fleming [You66]. Fliess [Por87b, FMF85]. Flight [Smi87b, Cer89, Har85b]. Flood [Bra89a, Rap78]. flora [Agr83a, Des92, Sht96]. Florence [Hat87, KK85, Sch79b, Sch81b, But96, Out80c]. Florey [Swa86, Wil84f]. Florida [Ait86, All84a, Din63a, Fox74, Jor78, Rud89, Smi85d]. Flourens [Jac85, NC81]. Floyd [Wri87]. Fludd [Gra81]. folding [Bry81, Sme71]. Folgen [Von88, Wol87]. Folia [Geo78b]. Folio [Roc88b]. Folkestone [Blu79, Goo81b, Har79a, Lev80, Ora80, Sin81, Smi78c, Til76]. Folklore [Fur85, Swi67, Llo83]. Folkstone [Har82d]. Follette [Rav84]. fondamentaux [BN82, Sch84c]. Fondements [Atr86, Jor88]. font [Lil85]. Fontana [Bro83a, Ink87, Bow92, Bro92, Car94, Nor94, Smi97]. food [BR95]. Foods [Lar74]. Forbes [Bar73a, Bry78a, Bry83, Cro70b, Fox71a, GD85, Hom74, LB73, Mor69a, Sha77b, Wil84b]. Forbidden [CT93]. Force [Gan95, Goo75, Nay75, Sch86a, Whi72b, Wil84e, Adl87, Gau82, Gou84b, Har82b, Hen84f, Mac83, Nad83]. Forced [AS96]. Forces [Hen89a, Pai86]. Ford [Bar75]. Foreground [Kni87]. Foreign [Cro67, Wuj85]. Foreward [Sch84c]. Foreword [Ano87p, Bro72a, Byn74, Col76, Cro72, Dol71a, Han70, Lev73, Mac85b, Sch84c, Sch69a, Ste62a, Und62]. Forgotten [Bru70, For70]. Forlag [Bad83]. Form [Eva89, Hal86d, Kni89b, Smi73a, BBC88, Bur87b, Eme84, Fre95, Fur88, Nyh95]. Forman [Mor77b]. Formation [KH86, Mor77a, Mor89a, War88a, BBC88, Hal84d, Huf82, Lie87]. formatione [Wea78]. forme [BBC88]. Formed [NW82b, NW82a, Ano82n, Ano82m, Bro73a, Fox79]. Forms [Pic87b]. formulae [Ste91b]. Formulary [Ear70]. Formulation [Hen86b, MR82b]. Forrest [Web76]. Forrester [Ril85]. Forschung [Bro82a, CM90, Sim84, Tur75, Dei92]. Forster [Gel76]. Forthcoming [Ano64h]. Fortschritt [Sch83b]. fortune [dCC09]. Fortunes [Col81a, Gab78, Jor85a, Tyt82]. Forum [Hol71]. fossil [Ste90a, Tha83b]. Fossils [Rap75, Por78a]. Foster [Mor80]. Foucault [Fig77b, AO77, Gut89, Ste91b]. Foulis [Tur74c]. found [Ear70]. Foundation [Duc88, Goo85a, Hea64, Kar76, Rud63, Sar81, Kle94, MR82f, MR82g]. Foundations [Bro89f, Can77, Dol74, Dub72, GG74a, Hal77b, Hen86b, Web66b, Wes62, Lau87a]. Founder [Eyl68, Sme68a]. Founders [Law89, O’C88, Sar92]. Founding [Sha74, Ver88, Sha98]. Four [PTS^{+}96, Cra92, Hen82c, Nev84]. Fourcroy. [Ede65]. Fourcy [Out88a]. Fourier [Buc77]. Fourteenth [HM66, Fre92, Mol68]. Fourteenth-Century [HM66, Fre92]. Fourth [De 85, Pat86, Wil87d, Agr83a]. Fox [Bru72, Out82a, Smi87a]. foxglove [Gab83a]. Fr [Hom86a]. fra [Sch81b]. Fragment [Car69, Dra69, McG70b, Twy88]. Fragmente [Scr68a]. Frames [Dol83]. Francais [Pet84]. Francaise [Cro68b]. France [Bow89b, Bro84a, Cal63, Cro66b, Cro68a, Cro85, Cro87a, Cun88, Ege89, EK91, EK92, Fox71b, Hai82, Hal65a, Hal72b, Hil71a, Jor78, Jor85b, Lau87b, McC77b, Num87, Osb89a, Out82a, Pes84, Sch84c, Sha86, Stu87, Wea80, Woo89, And84, Ano82l, CM90, Cor88a, Deb91, Fox71b, Fra78a, KK87, Lat88, Lau87b, Moo77, Out84c, Poi04, Pri83, Ram88, Sch01, TBDA10, Wei83b, Wil94a, dCC09, dG01]. Frances [Hen84d, Tur73a, Hen85b]. Francesca [Sho87a]. Francesco [Sch79b, Ara91, GGC90]. Franchini [Sch81b]. Francis [All79, Bry78a, Cha77b, Chi65a, Dol73a, Kel68, Kni86d, Por83a, Sha89, Sin77, Sin78, Tur73c, Vic88, Wea83, Web71a, Wig78, BM84, Fan83, Mar92, PR88, Und71, Web76]. Franciscan [McV75]. Francisco [All82, Byn78, Smi76c, Yox83a, Yox83b, Ara91, Fre90, GGC90]. Francke [Mol75]. Franco [Mor89c]. Francois [Dou84, Olb72, Sch84b, Scr68a, Whi85a]. Francoise [Sch84c, Tre87a]. francs [GG77b, Bur74a, Bur74b, Cro79, Dug77, GG70, GG75b, Hal65a, Hal72b, Hal77a, Hut80, McC80, Mol75, Pau76, Phi67, Sch72, Tur72a, Wil77]. Frank [Bro75, Can74, Cro71, Cro72, Dav81, Dol73a, Hen87a, Kel73, Kni83a, Kub70, May67, McC69, McG69, Mor84f, Nic85, Ora73, Ste75, Tha67, Tur74c, Whi72a, CMN95]. Frankfurt [Cap81b]. Frankfurt-am-Main [Cap81b]. Frankland [Ear87, Rus86a]. Franklin [Cal77, Goo82, Hal80c, Pin88, Smi80a, Hom72]. Frankreich [CM90]. Franz [Pag69, Rus72b, Whi68a, Sap90]. Französische [For69b]. Fraternity [Jen70]. Frazer [Bow88b, Ack87]. Fred [Col81b, Fis83a]. Frederic [Mau77, Col87b]. Frederick [Buc68, Cha80b, Cro68a, Sin77, Ste62a, Tro89, Fre79, Rud86, Rud88b, Sha89]. Free [Har79b, Nay74, BDH96, Bro87]. Freedom [Nye76, SG96]. Freeman [All81, Byn78, Sme88, Yox83b]. Freiburg [For69a]. Freies [Pat84]. French [Bla83, Byn88, Fox74, Hal84c, Hen82c, Rap71, Tre87a, AHY90, Ano87r, Ano90, Ano92, Ano93, B^{+}95, Atr86, A^{+}87, Bea91, BG^{+}92, Bec91b, Bec92, BaJANJ84, BV93, BH90, Bla93, BBC88, BD95, CMDM82, Cha84, Cor78, Cro92, Deb91, Duc87, FPLJ89, FM00, GP92, G^{+}09, Gay92, Gol87a, GGR96, Gre86, Grm90, Har67, Her66, HJF89, Lan87, Las81, Lau87b, LF97, MC^{+}93, NS02, Osb94, Out80d, PP95, Pes84, Pic87c, Poi04, Pye93, Roc88a, SG80, Sch01, SGŚ00, Sin91, Six87, Tor92, Wei82, dCC09, dGB^{+}89, dG01, vPvP90]. fresh [Whi64]. Freud [For82, For87c, Por87b, Smi80b, War89, Bru95, Cra93, FMF85, Gos87, Kit92, MF86, Rit90]. friar [FC96]. Friars [Hal88a]. Frictional [Hom80]. Friday [Ora76]. Friedel [Bow87a, Rus85]. Frieden [Wei85b]. Friedman [Aas87]. Friedmann [TFC93]. Friedrich [Dor75, GG72, GG77a, Mei91, Pat84, Rus72b]. Friends [Nev80, CMNH92]. Fritz [Olb69, Rus74, Mol86]. Froderberg [GG72]. frog [Per92]. Frommann [GG77a, Rus72b]. Front [Ano62a, Ano62c, Ano63b, Ano63d, Ano64b, Ano64d, Ano65b, Ano65d, Ano66b, Ano66d, Ano67b, Ano67d, Ano68b, Ano68d, Ano69b, Ano69d, Ano70b, Ano70d, Ano71b, Ano71d, Ano72b, Ano72d, Ano73b, Ano73d, Ano74b, Ano74f, Ano74d, Ano75b, Ano75d, Ano75f, Ano76b, Ano76d, Ano76f, Ano77c, Ano77e, Ano77g, Ano78b, Ano78d, Ano78f, Ano79b, Ano79d, Ano79f, Ano80b, Ano80d, Ano80f, Ano81b, Ano81d, Ano81f, Ano82b, Ano82d, Ano82f, Ano83b, Ano83d, Ano83f, Ano84b, Ano84d, Ano84f, Ano85b, Ano85d, Ano85f, Ano86b, Ano86d, Ano86f, Ano87b, Ano87d, Ano87f, Ano87h, Ano88b, Ano88d, Ano88f, Ano88h, Ano89b, Ano89d, Ano89f, Ano89h]. frontier [Duf83]. frontiers [Pan91]. frontis [Chi64]. Frontispiece [Hea64, Sme66]. FRS [Mor81, Und80]. Frueh [Hal66a]. fruits [Duf81]. Fruton [Mor84b]. Fry [Ash85, Dra86]. Fryer [Nee69]. Fu [Whi70a]. Fuchs [Bad89, Wil87e]. Führer [Wes70]. Fulke [Bau75]. Fullmer [Kni71c]. Function [Buc69, Smi73a, Smi75b, For86]. Functionalism [Bow86, Sto84]. Fundament [Sch92b]. Fundamental [Hen86b, Kil94, MR82c, MR82e]. fundamentals [MR87b]. Funding [Mor77b, Str87]. funeral [RL97]. Furley [War88a]. Furth [Eva89]. Fusion [Her90, Kei84, Hui92]. Fussell [Rus77a]. Future [Dur79a, Seg94, MJ89].

G [Ack87, Ano67j, Ant63, Arm62, Bar73a, Bow88b, Bry76, Bry78a, Bry78b, Bry83, Bus74, Can74, Can80b, Chi64, Col88b, Cro72, Cun83b, Cun89, Dea86b, Dol73a, Duf81, Fis81a, Goo68, Got67, Got71, Gra81, GG75a, Gra86b, Gre88, Hac75, Hac82a, Hal68b, Hal88a, Han70, Har87b, Hei68, Hen82c, Hen83c, Hen89g, Hom74, Hun89a, Jos67, Jud88, Kni71b, Kni85, Kri83, Lay74, Mac78, McL86, Mol79, Mor76a, Mor84g, Nor82, Pic76, Pic80, Por85a, Rap86a, Rob80a, Rud71, Rud80, Rus66, Rus77a, Sch78b, Sch80d, Sco67, Sha98, Sha83a, Sha77b, Sin76, Sme66, Smi80a, Smi83, Sti89, Sto88, Tur70c, Tur73d, Tur74c, Tur82, Tur87a, War89, Whi77, Wig74, Wil69, Wil84b, Yox83a]. G. [Ait89, BT96, Dun86, Fur85, Koh75, McC76, Osb85, Pum88, Smi85a, Van86]. Gabriel [Smi78a, Wil72]. Gabriele [Kni89e]. Gage [Kni89e]. Gainesville [Fox74, Jor78]. Gainsville [Smi85d]. galaxy [Pau93]. Galen [Fre74, Nut84]. Galenism [Wil75]. Galilean [Nay85, Dra81]. Galilée [Ait68, Hal68c, Hal77a]. galiléenne [dGB^{+}89]. Galiléennes [Hal68c]. Galilei [Hat87, Nay75, Rav68]. Galileo [Arm66, Dea86c, Dun86, Gab83c, Hat87, Kel78, Mol84c, Nay75, Nay77, Nay80, Nay85, Rav68, Rav72, Sch74b, Sch75, She85, Ano64g, Arm67, Bia93, Dra69, Dra70, Dra81, FC94, Fan96, Hal65b, Hal72a, Huf91, Hum67, LS90, MW94, Nay74, Red83, Ros68, Seg91, She70, SB09, Wal81a, Wal84, Wal92, Wes89]. Galison [Hen89d]. Galley [yL66, LY66]. gallic [dGB^{+}89]. Gallimard [Olb72, Tur73b]. Galluzzi [Hat87]. Galton [Dol73a, Web76, Fan83]. Galvani [Per92]. Gamal [Fre86]. game [Pou93, Wei92]. games [RL97]. Gaposchkin [Hal85a, GH84, GH96]. Garbrecht [Smi89b]. garden [Wea83]. Gardens [All82]. Garland [Bud84, Har77, Hen83b, Mor82a, Nor80, Rud86, Tha85]. Garma [Gli84]. Garret [Sch69a]. Garry [Kni86c]. Gary [Dur77, Kni81]. Gas [Wil85b, Wil81a]. Gases [Bru72, TP66, Men75]. Gassendi [Hen88e, Sch74a, Bru87b, Joy87, SG96]. Gassendo [McG68]. Gaston [Str73]. gatekeeper [Bro97]. Gaukroger [Jar83]. Gauss [GG72, Gra83b, Büh81]. Gautrey [Des88]. Gavin [Bar76b, Hod70]. Gaylord [Bow88a]. GE [Edg87c, Rei85]. Gears [Smi78b]. gebruik [Tur82]. Gedanken [Pag69]. Gedenkband [Gra85b, Nor72]. Gedichte [Hüb90]. gedruckte [Wes70]. Gehennical [New94]. Geiger [Swi67]. Geison [Mor80]. Geistesgeschichte [Pag69]. Geistesleben [Pat84]. Gelehrten [ZH92]. Gelfand [Gab83b]. Gems [Car74]. Gender [Dur81, Ben91, Lun94, Sut95]. Gene [Nor85a, WKB88, Yox83b, Yox83c]. General [Dav81, EK91, Fin68, Fri87, Tha74, Wal77, EK92, Hut89, Whi88b, Bro73a, Buc73a, Can80b, Cha77a, Cro72, Dol73a, Dol73b, Duf75, Dur77, Fee77, Goo81a, Har75, Hes74, Kni75, Kni80a, Lay74, Mol74, Mol79, Mor75a, Mor78, Por77, Rus72a, Rus77b, Sch72, Sha79, Smi74b, Swi77, Tha74, Tur78b, Wal74, Wal77, Wig74, Wis85, Gun84]. Générale [Cal63, Hil71a]. generated [Tha83a]. Generation [Hod89b, Ree80, Les83, Mor82b, Nut84]. Generatione [Web67b]. generations [Cun83b]. Generous [Cla86a, Tom84]. genes [WKB88]. Genese [Mei91, Duc87, Sch84c, Tur72a, BN82]. Genesis [Col81b, Goo82, Law81, Lef78, Gra85a, Mei91, Wus84, Duc87]. Genetic [Kri82, Olb85, EE95, Wri94]. Geneticists [RH89]. Genetics [Bar89c, Coc75, Har89c, Hog74, Olb77, Web66b, Fuj96, Maz92, Nor82, Sar92, Coc75]. Geneva [Gil84, Har86, Hom86a, Kel78, Por84a, Hey83]. Genevan [BaJANJ84]. Genevois [Hom86a, BaJANJ84]. Genius [For70, Mil87, Mur89, Web76, Dob91, Jak84, Sim88b, Wil83, Stu85]. Gens [Por89e, Roc88a]. Gentleman [Bro86e, Dem83]. Gentleman-Farmer [Bro86e, Dem83]. Gentlemanly [Mor87b, Rud85]. Gentlemen [Mor84e, Pat86]. Geochemie [Fri91]. geochemistry [Fri91]. Geoffrey [Aus86, GG84b, Sha73b, Web67a, Web69a]. geographical [Liv93, Por83a]. Geographies [OW10]. Geography [Hom74, You83, CL96, Dor91, GS94]. geoheliocentric [Sch65]. Geological [Frä76, Ger71, Rud62, Sto76, Hal83b, Rup85, Tor83, Her95, Rud63]. Geologist [Osp74, SD76, MS95, Sec83, Spe82]. Geologists [Rud83, Rud89, Sar87]. Geology [AO77, Bro89f, Bus74, Dav71, Eyl68, Gre82, NP77, Por76a, Por81, Rap78, Rap79, Rud69, Rud83, Rud84, Rud89, Sch69a, Tha87, BGH^{+}87, Lau87a, LeG88, LR91, LS97, Por84a, Rap82, Rup83, Sar87, Sec86, Bar75, Bur80, Bus74, Dav71, Nev78, Olb71b, Osp74, Por81, Rap75, Rap79, Rud70, Rud71, Smi89a, Whi70a, NPH92, Cor78, Ord76, Osp76]. Geomagnetic [Nut82b]. Geometer [Kno85]. Geometria [BM89]. Geometrical [Lin89b, Tur88, DFG87, Fie88, Mol68]. Geometrie [Mol84b, Mor83d, dGB^{+}89]. Geometry [Whi85a, Edg91, Gan95, Mor83d, Ric88b, Viè83, dGB^{+}89]. Geomorphology [Hig65, Rud70]. Georg [Hom86a, Joh81, Har87b, Lin90, Mai85]. George [Bow88a, Bro84a, Bro89b, Bro83a, Chi65b, Col76, Dur77, Joh83, Mor69a, Mor72b, O’H86, Out82a, Phi67, Por85b, Ric89, Rob75, Sch69a, She63, Sme68a, Smi78a, Swi67, Tre87b, Web75, Woo67, Woo73, Bow86, Can93b, Col88a, Fri87, Lev85, Lew85, Mac85a, MH88, New94, PN95, Rud62, SL87, Wil72]. Georges [Bur74a, Cor81, Cro85, Geo66, Law83, CD94, Out84c]. geoscience [Ste90b].

Geoteleology [Whi70a]. Gerald

[Joh83, Mor80, Kil84a, Pye87]. Gerard [Fox82, ABR93, McC85b]. Gereon [GG81, Von88]. Gerlof [Byn87]. German [Fox74, Gra86b, Hal84d, Har81, Kni84a, Kni84b, Mor74b, Olb69, San87, Adl87, AH93, Ash95, AS96, Cit90, CM90, Dau98, Dei92, Fri91, Ger89, Gol91, Gre92b, Gro89, Har96, Hei93, Hei86, Hei88, HRP90, HP85, Hüb90, Huf82, Jun89, KvDC^{+}37, Kle94, Kra91, LM23, LHB04, LS76, LHH91, Len82, Lil85, Mai88, MD90, Mei91, Mor91a, Nyh95, Pau84, PHvMW93, PvM85, Sch93, Sch89, Sch84d, Sch92b, Slo79, Ver85, Wei89, WKB88, Wey92, Wol87, ZH92, vL76, vLvHH^{+}88, vPSBD^{+}04]. German-Jewish [Har81]. German-speaking [AS96]. Germanic [Sne74]. Germany [Cah88, Cha80b, Smi80b, Wil84b, Woo89, CM90, Geu92, Gow71, Jun89, KK87, Lin96, Tho87, Wei91, WKB88]. Gernot [Hoc86]. Gerson [Swe76]. Gersonides [Fre92]. Gerstenberg [Mau87, Rus74]. Gesammelte [Kra86, Nor72, KvDC^{+}37, vPSBD^{+}04]. Gesamtausgabe [GG77a]. Geschichte [AH93, Bos72, Pag69, Sch84d, WKB88, Rav65]. Geschidenis [Bry77]. Gesellschaft [Mol86]. Gestalt [Ash95]. Getting [Kni87]. Gettings [Fis83a]. Gewichten [Tur82]. Ghori [Wuj85]. Giacconi [Smi82a]. Giacomo [Sch73a]. Giambattista [Kel67]. Giancarlo [Sch79b]. giants [Mer93]. Gibbon [WBP97]. Gibbs [Cri66, She63]. Giere [Dol74]. Gifford [BC98, Sha78]. Gift [Clo80]. gifts [Agr83b]. Gilbert [Por82, Rat66a]. Giles [Bar76a, Bur80, Cor81, Gou84a, Vic88, Wea78, Wil84d]. Gillispie [Car84, Hai82, Men74]. Gilman [Por87c]. Gindikin [Gre88]. Giordano [For74a]. Giorgio [Hal68c, Mol69]. Giotto [Edg91]. Giovanni [Kel68, Mid75]. Girard [Tur88, DFG87]. Girolamo [Sir97]. Giroud [Tre87a]. Girvetz [Swi67]. Giuliano [Por79a]. Giunti [Hat87]. Giunti-Barbèra [Hat87]. Giustino [Mid75, Sha77a]. given [Gra87b, Jor83b, Kni85, Kni88c, Kni89e, Bow86, Duc88, Goo89b, GG89, Har89a, Hat87, Hen89c, Hoc88a, Hod89a, Kni88c, Mah89, Mau87, Nic88b, Por89d, Rud89, Whi89b, Wil85a, Wil87d, Woo89]. Gjertsen [Kni86b, Rog88b]. Glanvill’ [Ste81]. Glas [Tei83b]. Glasgow [BC98, Bro83a, Moo84, Bro81]. Glass [Gid82, Hom80, Sch79a]. Glassmaking [Har76, Tur74c]. Glaxo [DHS92]. gleanings [Dra69, Dra70]. Glen [Sec84]. Glick [Whe89]. Glimpsing [Hir83, Nee85]. Glisson [Bos83]. Global [Mar10, Nun88, Men86]. Glory [Lin90, NP77]. Glossary [Tur78a, Got71]. Glyn [Bar76a]. Gnomonics [Rya70]. gns [For68]. Gobineau [Har86]. God [Rus87b, Whi70a, LN86, McG78, vR95]. Gödel [GG87]. Godfrey [Har76]. Gods [Web78]. Goethe [Fig88, Jac89b, Sha89, Sne74, Wel70, Adl87, AZW87, Sep88, Wel67]. Goethes [Adl87]. gold [Bry81]. golden [Sch84a, Wai90]. Goldin [Far77]. Goldman [Smi85b]. Goldstein [Byn88, Swe76]. Goldstine [Ran74]. golem [CP93]. Golino [Arm67]. Gollancz [Und64a]. good [Goo89a, GR88]. Goodchild [Yox82b]. Gooding [Hen87a]. Goodman [Coo89a, Duf75, Mus74, Por84b]. goods [BP93a]. Gorcum [Sme68b]. Gordon [Dor75, Ham84, Rud70, CHG91]. Gore [Dol73a]. Göschen [Rav65]. Gosling [War89]. Göte [Cro67]. Gotland [Rau75]. Göttingen [Hom74, Wil84b]. göttliche [Gol91]. Gottlob [Bar75]. Goulven [Osb89a]. Goupil [Cro79, Sme84a]. Gouverneur [Kni83a]. Government [Coo89a, HE92, Kni83a, Bru88a, Goo88b, Hen89f, Smy89]. Gower [Wil89b]. Gowing [GR88]. Gowles [Web75]. Grabado [Kni88c, Lan87]. Grabiner [Gra83a]. Grace [Far77]. Graduate [Kar76]. Graeme [Jor85a]. Graf [Wey92]. Grafton [Hut88]. Graham [RH88, Hea87b, Hil89a, Vic88]. Gramiccia [Kni89e]. Grand [Pau76, Cla70]. Granit [Cla66]. Granite [Mor72a]. Grant [Ait86, Hen83d, Sch76, Wei86]. graphes [GG77b]. Graphs [Til75b]. Grassé [Bur74a]. Grassland [She86]. Grattan [Dub72, Joh79]. Grattan-Guinness [Dub72, Joh79]. Graunt [Sha73b]. Graves [Tay89]. Gravesend [Kni86b]. Gravitation [Wes67]. Gravity [Nay75]. Gray [Ser88, Ste62c, Tur88, Web65a]. Great [Can74, Goo83a, Gre71, Hal83b, Hei73, Hen89b, Hew66, Kei86, Mon78, Mor87b, Olb69, Rei87, Ron86, Rud84, Rup85, BH84b, KW87, Por82, yQ85, Rud85, Rup83, Ano69h, Lay65, MJ84, McG84, Rei68, Smi82b]. Greatest [GG74a, Kni89d, MR88]. Greece [Fur85, Han76, Llo83]. Greek [Ant63, Fre74, Got67, Hal86b, Han76, War88a, Cla62, Cop92, Fre74, Fur87, Fur89, Hon89, Llo91, Ole84]. Greeks [Smi78b]. Green [Gro95, Can93b]. Greenaway [Can74, Cro71, Tha67]. Greene [Hal77b, Nic85, Rid83, Rud84]. greenhouse [Cra96]. Greenough [Rud62]. Greenwich [Bry78a, Cha77b, Wil85a]. Greenwood [Gab83b, Kni83a]. Gregor [Ora76]. Gregory [Cha80b, Mor77a, Eag77]. Grelon [Fox87b]. Gresham [Rat69, Sha77b]. Greta [Ben82b, Gol79]. Grew [Le 90]. Gridgeman [Gin84]. Griffin [Mah89]. Grmek [Dur80a]. Groom [Wei83a]. Gross [Har82d]. Grössenlehre [GG77a]. Grounds [Cav96]. Group [Ser88, Ste62b, Gra85a, Gra85c, Hei91, Sch89, Wus84]. groups [Fru90]. Grove [Por82]. grown [HS89]. Growth [Cla86b, Dol71b, Mar73, Bru88a, GH92, Mei91, Whi75]. Gründerzeit [Mei91]. Grundlegung [Kle94]. Gruppe [Sch89]. Gruyter [Bra87, For69b, GG81, Hal85d, Von88, Mol72a, Nor71a, Rav65, Scr68a]. Guathier [Osb89b]. Guathier-Villars [Osb89b].

Guenter [Fig77a, Wea79]. Guericke [vGA94]. Guerlac [Hal80a, Kni83b, Sch84c]. Guerres [Fox87b, Gre86]. Guicciardini [GGC90]. Guide [C^{+}92, Col88a, Rav67, BW90, Bru88b, Olb82, TW87]. guilders [Mor73a, Pep71, Sme68b]. Guildford [Jon87]. Guinness [Dub72, Joh79]. Gummett [Yox82a]. Gunnar [Bro84d, Cha80c]. Gunpowder [Hal88b, NPYGDL86]. Gunther [For85a, GG77a, Smi78c, Tur68, Und80, Rus72b, Smi89b]. Gustaf [Goo85a]. Gustav [Hei93, Osp74, Hei93]. Gustina [Kel73]. Gutting [Kni81]. Guy [Din63a, Stu85, Ben89]. Guyton [GGSGB94]. Gwei [Agr83b, Hal88b, Mol71, Pac72]. Gwei-Djen [Agr83b, Hal88b, Mol71, Pac72]. Gweneth [Sch73b]. Gyula [Kni67].

H [All82, Asq80, Bal84, Ben79, Bib74, Bro76, Bro73b, Bry76, Bry83, Buc69, Bud84, Byn78, Can84, CHG91, Cow68, Des88, Dol83, Duf75, Far67, Fee77, Fie89, Fis81b, Fox71a, Got71, Gow71, GG75c, GG81, Hac82b, Hal68d, Hal84e, Hal84a, Hal88a, Kel70, Kni83b, Kni83c, Kni86d, Kra86, Lef78, Mor72b, Mor81, Mos77a, Nev78, Nye86, Out84a, Pin83, Ran74, Rat66b, Rat68, Rus74, Sch83a, Sch69a, She63, Smi73b, Smi74b, Sne74, Ste75, Stu71b, Und75a, Vea64, Wal77, Web84, Web68a, Web70, Wes70, Wil84e, Woo84, Woo73, Yox83b]. H. [Bib74, Gou85, Mol85, Nun88, Sho88, Sin76, Tom65]. H.-P [Kra86]. H.M.S.O [Por89d, Swi70]. Haar [Kuh68]. Haarlem [Cro70b, Fox71a]. Habitat [Won93]. Hackmann [Bry77, Hom80, Wil87b]. Haeckel [Mai85]. Hafner [Ant63, Bar75, Car74, Mor72a, Sch69a, Und75a]. Hagecius [Rav69]. Hagen [Jac89a]. Hägg [Tur70c, Tur73d]. Hagner [Car74]. Hague [Dou84, Gil84, Hac75, Han85a, Hen83c, Mor73a, Pep71, Sch84b, Sch74a, Tur75, Wea80]. Hahn [Cro89b, Mol84b]. Hakim [Sar81]. Halberstam [Cow68]. Haldane [Stu88]. Halesworth [For85a]. Half [Rus89a, Bry77, Chi65b, Cri66, Her88]. Half-Century [Rus89a, Her88]. half-tone [Bry77, Chi65b, Cri66]. Hall [Bro78a, Bro78b, Buc66, Byn74, Gab83a, GG72, Hen84a, Hop73, Hop75, Kni65, Mac88a, McG66, Nev78, Rav73, Rob87, Sch81a, Sch84c, Scr68b, Ste62c, Web66c, Web69b, Woo80c]. Hallam [Rup85]. Hallenx [Sch78c]. Haller [Les83]. Haller-Wolff [Les83]. Halleux [Gol83b, Wil77]. Halley [For66, Wil87c, Van85]. hallucinati [Bra87, Zam86]. Hamburg [Por89a]. Hamdard [Sar81]. Hamden [Blu79, Swe76]. Hamilton [Cow68, Mac80, Mor83b, O’H86, O’D83]. Hammer [Rus72b]. Hammersmith [Ste62b]. Hammond [Fie89]. Hampshire [Gab83a, Lay74, Lud80, Por81, Web87, Win75]. Hampson [Dav89]. Hamshaw [But63]. hand [II90]. Handbook [Ano69h, Mai88, Rog88b, Gje86]. Handbuch [Mai88]. Handlist [All81]. Handschriften [Whi68a]. Hankins [Hah73, Jor87, Mor83b]. Hanle [Mea84]. Hanna [McG69]. Hannah [Hen84d, Smi80b]. Hannaway [Fra87a, Kni76]. Hannen [Sha83a]. Hanover [Gab83a, Lay74, Lud80, Por81, Win75]. Hanoverian [Sco67]. Hans [Coc75, Gra86b, Pag69, Sin78, Tei83a, Gra85a, Mah89, Mau87, Tei83a]. Hansen [Lin86]. Hanson [Lar74]. Hantz [Swi67]. happiness [MJ89]. Haramundanis [Hal85a]. Haraway [Tei78]. Harcourt [Bar77, Wil78]. hard [Gra87b]. hardback [Hen89a, Hen89c, Hen89d, Rob89a]. hardcover [Blo84, Bow85, Dun87, Eme87, Fel89, Gab88, Gol88b, Hal84e, Hal84d, Hal87b, Hen88c, Jor87, Kni86b, Mol84c, Out85, Pic84a, Por87c, Pum87, Rup85, Rup88, Rus87a, Rus87b, Sec88, Sho88, Smi89e, Smi89c, Tho89, Wor88]. Hardie [Nor68b]. hardware [Bea85b, Mor84a]. Hardwick [Vic88]. Hariot [McG68]. Harley [Pic88b]. Harman [Hen84f, Jam87b]. harmonic [Ste94]. Harmony [Sch78b]. Harold [Bow73, Bro72b, Din63a, Olb71b, Rus68, Rus77c, Str75, Swi67, CMN95]. Harper [FF66, Swi67]. Harrap [Dor75, Gre71]. Harré [Dol70b, Goo83a]. Harriet [Har79b, Law80a, Nut82b, Tur77a]. Harriot [Loh75, Pep86, Pep67, Roc81a, Shi83]. Harris [Bro86e, Fre74, Lie86, Smi85b]. Harrison [Pum88, Par87]. Harrower [Smi85d]. Harry [Fox74, Swi67, Cro87a, Pic87b, Whi89a, YK94]. Hart [Smi87b]. Hartcup [Stu85]. Hartley [Bro72b, Rus68, Bow73]. Hartmann [Mor91b]. Harvard [Ait72, Bad89, Bod65, Bow83a, Byn74, Chi65a, Cla71, Dav71, Fox70, Geo66, GG75a, Hac89, Hal71, Hei71a, Hen81, Her80, Hod89a, Hut84, Joh81, Jor79, Kel70, Kel85a, Kni71c, Kni74, Kub70, Law80b, Lor73, Lud80, Mau77, Mol69, Nee69, Pic88b, Por87b, Rav70, Rud69, Rus69, Sch84b, Sch76, Swi71, Tur70b, Web65a, Web65b, Web68b, Web70, Whi73, Wil86a, You66]. Harvester [Bar81, Ben82b, Blo82b, Bro80b, Hen84f, Jar83, Jud88, Sha81]. Harvey [Coo83, Sch73b, Sha89, Web66a, Web67a, Whi68c, Web67b]. Harwood [Buc85a]. Hassan [HS88]. Hassocks [Bro80b]. Hauniensis [Mer78]. Hauser [Kni88c]. Hautes [Fox87b]. Haven [Bro75, Bro86c, Bus74, Duf81, Far76, Far77, Hei71b, Hen87b, Kei84, Nor82, Rud71, Smi73b, Tei78, Und66b]. Hawaii [Kni89d]. Hawkins [Rav71]. Hayck [Rav69]. Haytham [AS89, AW71]. hazard [EM94]. healing [Cra93, Lin96]. Health [Lin96, Nev84, Por89d, Por89f, Sch68a, Web98, Wig77, Coo84a, Jor85b, Lan89, Wea79, Web88, Who82, Coo81, Mac79]. Hearnshaw [Ric88a, Smi87c]. Heart [Fre74]. Heat [Bod65, Bru70, Fox70, Ger71, Jam82, Pac74, CGBE95, Fre95, Gil83, Mir89, Bod65, Gol77, Smi76b, Fox69b]. Heath [Din63a]. Heathcote [Sme87]. Heathfield [Sec83]. Heavenly [Rus79b]. heavens [Mos93, Ste94, FC82]. Hebraica [Lev90]. Hedwig [Bro72a]. Heffer [Arm62, Arm64, Ede65, Hal68b, Smi66]. Heft [Nor71a]. Hegel [Dol71a]. Heidelberg [Bos72, Gra83b, Gra87b, Har81, Hen86b, Hen86a, Hen88c, Kra86, Mol78a, Mol88, Sch75, Sme68a, Tur88]. Heidelberger [Mol75]. Heike [GG91]. Heilbron [Can80c, Hen83b, Hom86b, Mor77b, San87, Sin76, Tre88]. Hein [Jac89a]. Heinemann [Asq80, Bru86, Buc73a, Chi65c, Gre63, Hen82c, Mon78, OM75, Ora74, Smi74a, Stu71a, Tha67, Wig74]. Heinrich [Buc94]. heirs [Sto88]. Heisenberg [Bro72a, Hen86a, HP85, Kra86, PHvMW93, PvM85, HP85, PHvMW93, PvM85]. held [AL87, Ber95, Chi65c, Frä90, Hen89g, Hol85a, JO81, N^{+}88, Pep89, Sar81, WV92, Whi77]. Helden [Wil87c]. Hele [McN83]. Helen [War88b]. Hélène [Tur72a]. Helge [Hen88b]. Heliocentric [Tho76a]. helix [Yox83a]. Hellenistic [B^{+}82, EBT96, Ano82p]. Heller [Dun86]. Helm [Bus83, Fig78, HS86a, Mac79, Mor84c, Por88, Por89b, Ron86, Rup88, Sha77a, Wei83a]. Helmholtz [Hei74a, Tur83c, Hat90, Tur77b, Tur94]. helminthology [Gro90]. Helmont [Hen84k, Bos83, Pag82]. Helmut [Hen86b, Hen88c, Mor84g]. Helmuth [Mor84g]. Helped [Byn80]. Hendrik [Mül89, Hen86d]. Hendry [Che85, Stu85, The87]. Heniger [Mül89]. Heninger [Sch78b, Sch79a]. Henk [Dea83a, Pic87a]. Henley [Tur74c]. Henley-on-Thames [Tur74c]. Henrey [Edw78]. Henri [Dug77, For74a, Hal68c]. Henrietta [Whi85a]. Henry [Bar76b, Bro67, Cro71, Far69, Fri74a, Goo81b, Goo85b, Hac74, Hac77, Hal71, Hal80a, Hop73, Hop75, Jen91, Kni83b, Lef78, Lud80, McC69, Mor73a, Nev80, Rus69, Sch84c, Web66c, Web69b, Woo80c, But64a, But64b, CMNH92, Far63, FB97, Hal90, Hen88f, HC90, Jac83, McC68, Rai91, Web69c]. Her [All84b, Ben84, Hac88a, She63, Wil87b]. Herapath [Men75, TP66]. Herausgegeben [Scr68a, Wes70]. Herbert [Cap81b, Des88, Hen82b, Pye76, Sme66, Und71, Sme79]. Hercules [Sin84, Kel83]. Herder [For69a, Gow71]. Hérédité [Olb72]. Heredity [Geo87, Mac80, Rus86b]. Heredity-Environment [Mac80]. Hering [Tur94]. Heritage [Dea86c, Out87, Smi89e, Ali86, Edg91, KK85, Wal84]. Herivel [Ait67, Buc77]. Herken [Rus89a]. Herman [Han70, Ran74]. Hermann [Ait68, Cro66b, For74a, GG70, Hal68c, Hei74a, GG75c, Mid65, Tur77b]. hermetic [Fis83a]. Hermetica [Cop92]. Hermeticism [Woo80b]. Hermeticum [Cop92]. Hermetische [Sch83b]. Herminio [Pic84a]. Herophilus [VoC89]. Herries [Ham84]. Herrmann [Wil86b]. Herschel [Buc68, Wil86b, HK84, Sch92a, Sch80a, Sut74]. Hersey [Sch78b]. Herstmonceux [Bry78a]. Hertwig [Wei91]. Hertz [Hun89a, Buc94, OP87]. Hertzsprung [Wil86b, HK84]. Herz [GG89]. Herz-Fischler [GG89]. Herzberger [Rat66a]. Hesse [Blo82b]. Hessen [Mor91a]. Hétène [Hal72b]. Hetherington [Whi89b]. Heuser [Bro89a]. Hevesy [Tre87b, Lev85]. Hewson [Wea78]. Hexenprozess [Hüb90]. Heyck [Mor84c]. Heyd [Gil84]. Heyningen [Pel85b]. Hfl [Byn80, Fis74, Hac75, Sch74a, Tur75, Bry76]. Hidden [Rei86, CT93, GGC90]. hierarchical [Sal85]. hierarchies [EMW82, Pic84a]. Hieroglyphica [Rat66b].

High [Arm70, McC93, Mor83d]. Higher [Sta90]. Highlands [Ben82b, Cha77a, Fie81, Fie83]. Highlights [Har81]. Hilbert [Nor72]. Hilde [Tre87b]. Hildesheim [Mau87, Rus74]. Hilger [Ben84, Buc89a, Dra86, Fie89, Hac88b, Hal87a, Hun86, Kei89, Nye86, Pin87b, Stu85, The87, Tre87b, Tur83c, Tur87a, Whi84, Whi88a]. Hilken [Fox69a]. Hill [HS86a, HS88, Kel75, Por78b]. Hills [Buc71, Goo81a, Irw82, Irw83a, Sha79]. Hillsdale [Fie86]. him [Tha66a]. Hindle [Koh78, Wor88]. hints [FTG91]. Hippocratic [Nut84]. Hirschfelder [Hen82b]. Hirsh [Nee85]. Hirst [Mor81]. Hirt [For67b]. Hispánica [Kni88c, Lan87]. Hist [Wor79]. Histoire [A^{+}87, CM93, Ege89, FD87, GG70, Har86, Hil71a, Jor88, MC^{+}93, Olb72, Osb89b, Out88a, Sch72, Atr86, BBC88, B^{+}89, Cas85, Gay92, Hai80, Lau87b, Osb89a, Sin91, Cal63]. histoires [GGR96]. Historia [Gli84, Wil85a, LN95, ML89, FC82]. Historian [Ken64, Duf83]. Historians [Wat80]. historic [Dav99]. Historica [Cla70, Mer78]. Historical [AO77, AB89, Bad71, Ben84, Bod84, CM90, Cro66a, Eag78, Fau89, Fox68b, Goo81a, Gra86b, Grm69, Hal68b, Hen86b, Hen88c, Jac85, Jam86, Jar84a, Kni83a, McC81, McG65, Mol89, Nor82, Pat86, Ric88a, Rus72a, Rus77b, Rus87b, Rus88d, Sch84b, Slo79, Smi89b, Sol83, Tur70b, Und64a, Vea64, Web68a, Wil89a, ABR93, Bal91, Ben75, Bow88c, Coo89c, Dan90, Dea91, Gar87, Gau83, GH93, HE^{+}82, Jor83b, KH95, KT87, Kil84a, Kuh68, LN86, Lin93, MR82d, MR82e, ME81, NC81, Ols90, OE^{+}90, SO85, SY86, Tei83b, T^{+}85, Whi68c, ZH92, AH93, Gre63, Mor77b]. históricas [OE^{+}90]. Historicized [Bow86, Sto84]. histories [Fre94, Pet94]. Historiographical [Hod89b, Coh94]. Historiography [Hen88b, Hen87c, McC76, Roc81b, Chr93, Hol85a, Kra87, McC77a, Out80b]. Historiques [Osb89a]. Historischen [AH93, CM90]. historisches [ZH92]. History [All77, All81, All84a, All84b, AL87, Ano62h, Ano62k, Ano63m, Ano68j, Ano77m, Ano84j, Ano85j, Ano86i, Ano87m, Ano05, A^{+}87, B.69, Ban88, Bar76a, Ber79, Blu79, Bow87b, Bro69, Bro73a, Bro86a, Bro76, Bro72b, Bro84d, Bro85e, Bro86d, Bro87, Bru88c, Bry71, Bry78a, Buc73a, Buc85b, Bur80, BK75, Bur86, Büt83, Byn73, Cal77, Cam82, Can80b, Can86, Car70, Cha65, CC66, CH67, CE73, Cha78a, Cha77b, Cha80c, Chi68, Cla68, Clo80, Coc75, Col88a, Col88b, Coo81, Cor81, Cro77, Cro89a, Cro72, Dav81, Day85, Des88, Dor75, Dor80, Dou84, Dou88, DTS^{+}87, Dur80a, Edg87a, Far77, Fig77a, Fin68, FB75, Fis83b, Fox69b, Fox79, Fre81b, G.62, Ged71, Gel76, Geo66, Gil84, Goo81b, Gou84a, Gow71, Gow69, Gra81, GG89, Gra87b, Gre71, Hac82b, Hac88b]. History [Had63, Hal69b, Hal80a, Hal85d, Hal86b, Hal87b, HSHSJ93, Ham84, Har69, Hen80, Hen84c, Hen86b, Hen87c, Hen89g, Hes74, Hig65, Hod70, Hod89a, Hof62, HS63, HS86a, Hol71, HS88, Hom80, Hut89, Jen87, Jev67, Kei84, Kip91, Kit89, Kni67, Kni72, Kni83c, Kni84a, Kni84b, Kni85, Kni89e, Koh78, KK85, Kra86, Kul63, Kul64, Law81, Lay65, Lay74, Lev88a, LS90, Lie85, Lor73, LM74, Mac88a, Mar74, MJ84, McC77a, McV80, Mol71, Mol78a, Mol79, Mol88, Moo89a, Mor78, Mor83a, Mos77a, Mos77b, Mül89, Nee78, Nev78, Nic85, Nor69a, Nor69b, Nor70b, Nun88, Nye76, Olb71a, Olb85, Old77, Ora80, Out84a, Out84b, Out87, Pag71, Ped85, PS78b, Pic79, Por76a, Por78a, Por78b, Por79b, Por81, Por84b, Por89c, Pra70, Rap75, Rap86a, Rat68, Rav66a, Rav66b, Rav70, Rei86]. History [Ric87b, Rob80a, Roc88b, Rog88a, Ros79, Rud70, Rud83, Rud84, Rud89, Rup88, Rup89, Rus66, Rus72a, Rus88b, Rus88d, STT87, Sch84b, Sch70, Sch71, Sch79c, Scr64a, Sec84, Sei85, Sho87c, Sin78, Sin81, Sme70, Smi66, Smi74b, Smi75b, Smi78b, Smi80a, Smi85c, Smi89c, Smo94, Sne87, Spa65, Ste62b, Ste67, Sti70, Swi77, Tan62, Tay87, Tei88, Tha83a, Tha84a, Tha84b, Tor88, Tre88, TST87, Tur68, Tur74c, Tur78b, Tur87a, Und63, Und75a, Vea64, Vic88, Web75, Web87, WKB88, Whi70b, Whi89b, Wil65a, Wil65b, Wil84d, Wil85b, Wil86b, Woo80d, WG81, Woo92, Woo67, Woo73, Yox78, Aft91, Agr83b, Ait85a, AH93, Ali86, All86, Ano82n, Ano82m, Ano93, Ara91, Atr86, AB89, BA91, Ben84, Ben87, Ben90, BP95, BBC88, B^{+}89, Bow92, Bra89b, Bro92, Bro95a]. history [Bru88b, Bry92, Bry88, Bud84, Bud93, Buf87, Bus83, BP93b, CK89, CKA96, Can88, C^{+}92, Car94, Car88, Car86, CW87, CS94, Cha82, Chi65c, Coo84a, Coo88, CT93, Cre93, Cro91b, Dah97, DHS92, DM96, DeV82, Dea83b, Dur83, EK91, EK92, EE95, FB97, FB82, Fin93, Fis83c, Fog92, For73, Fox95, Fox96, Fre94, Fru82, GS88, GA98, Gau92, Gay92, GGR96, GG84b, GG85, Gra85a, GEMM83, Gro90, Hac82a, HE92, HW03, Har83a, HH86, HH92, aHH86b, HW81, Hen83b, Hen86d, Hen84l, HK84, HF87, Het88, Hil84, Hil94, Hol85a, Hom83b, HWW86, How82, Hun85, Ing84, II90, JK84, JSS96, Joh84, Joy87, Kel84, KRS83, Kni83b, Kni92b, Kri83, Kri82, Kuk91, Kuk93, Lau87b, Lea90, Lea94, Liv93, MB^{+}08, MKW92, Maz95b, McC82]. history [McC92, Men86, Mer89, Mil89, ML89, Mor84b, Mor90a, Mor90b, Moy83, Mur89, NW82b, NW82a, Nor89, Nor94, Olb82, O^{+}90, Old86, Ole84, Pat94, Ped92b, Pet89, Phi87, Pic84b, PR95, PW89a, Por85a, PT86, PB99, RM96, RR81, Rid83, Roc90, Ron83, Ros84b, Rot93, Rup83, R^{+}85, Sar87, Sch80a, Sch75, SSK88, SG90, Sch84d, SH73, Sha83b, SP88, SW89, Sho93b, Sin91, Smi83, Smi92, SM94, Smi97, Sol83, Ste91b, Sti86, TT86, Tra87, Tri89, TW87, Twy88, Way87, Wea79, Wea83, Wea88, Wea90, Web98, Wei83a, Wei92, WR92, WHS92, Whi88b, Who82, Wil81a, WD82, Wil94a, Won93, Wus84, Yox83b, aHH86a, aHH88, aHH92, vdBLW85, All84a, All84b, Arm62, Arm64, Bad71, Bal84, Ben84, Bra84, Bro67, Bro84d, Buc85a, Buc68, Car67a]. History [Car84, Cla67, Cla68, Cow68, Cri66, Cro67, Cro68b, Cro69, Dun69b, Edg84, Eyl68, FF66, Far67, Far68a, Far69, For67b, Geo66, Goo68, Got67, GG84a, Gra84c, Hal84e, Hal84c, Ham84, Hen84a, Hen84c, Hen84b, Hen84l, Kei84, Kel68, Kel84, Kni67, Kni84a, Kni84b, Lau68, LN95, Mea84, MC^{+}93, Nor68a, Nor71a, Out84a, Out84b, Phi67, Por84b, Rai91, Rav67, Rei86, Rud69, Rud84, Rus66, Rus67, Sch84c, Sch69a, Scr68b, Sec84, Sme66, Sme68a, Smi66, Sne87, Sty68, Swi67, Tee84, Tha84a, Tha84b, Und66a, Web66a, Web67a, Wei92, Whi68a, Wil84e, Wil84d, You66]. Hitherto [Jam88]. Hitler [Dei92, Dei92, Har96]. Hobbes [Jon87, Rus77c, Sha87, Str75, Jon75a, SSH85, Sor86]. Hobbs [Sch69a, Tha83a]. Hodder [GG85, McV80, Stu85, Swi77]. Hoddeson [Cas85]. Hodge [Smi83]. Hoeveler [Moo84]. Hofmann [Bro89a, Rav65, Whi75]. Hohenheim [PTS^{+}96]. Hohenlohe [Wey92, Wey92]. Holbrook [Bry74]. Holburne [Ben79]. holism [Ash95, Har96]. Holistic [Tre79, Woo80e]. Holland [Dor79, Fra87c, Gra86a, Hen87c, Kel75, Tho76a, Tho76b]. Hollis [Bar79, Sti70]. Holmes [Col87b, Mau77, Tre87a, Wei86, Fra78b]. Holton [Kil84a, Pye87]. Holy [Smi94b]. Holzboog [GG77a, Rus72b]. Home [Kni85, Mor73a, Wea79]. Homine [Jon75a]. Homme [Out82a]. Homoeopathy [Nic88a, Por88]. Hon [But65]. Honderich [Sha87]. Honest [Mos77b, Sch86a, For85b]. Hong [Agr83a]. Honigsbaum [Dav81]. Honnecourt [Bec88, Bec91b, Tee89]. Honolulu [Kni89d]. Honor [Wig74, BEFH10, Kni83b, MH88, Ric89, SG90, Whi77]. Honour [Bar77, Dol73b, LS90, May67, GR88, WHS92]. Honourable [McC69, Web71a]. Hook [Her95]. Hooke [HS89, Rap86a, Ben75, Ito88, Old72, Rap86b, Ros84b, Tur74b, Wes67]. Hoover [PLP92]. Hooykaas [Wes74]. Hopkin [Cal77]. Hopkins [Bow74, Bro71a, Buc75, Chi68, Chi71, Clo80, Coo83, Gel76, Gur78, Hal80a, Hop72, Jac85, Kni76, Law81, McG65, Mor83b, Mor89a, Nor85b, Phi69, Pic76, Pic80, Pin83, Por85a, Ree80, Ros79, Sch70, Sha86, Spa65, Web78, Yeo88, Smi89a]. Hoppen [Kel71]. Horace [Tur87a, CW87]. Horder [Bow87b]. Horn [Bry81]. Horne [Cla67]. horoscope [GGC90]. Horsky [Rav69, Rya70]. Horticultural [Edw78]. Horticulturists [All79]. Hortus [Hen86d, Mül89]. Hoskin [Arm62, Arm64, For72, Hei76b, Nor68a, Nor69a, Nor70a, Nor73a, Pep72, Smi66]. Hospital [Dav81, Far77, Web87]. Hostile [Mil83]. Hot [Dau74]. Houben [Tur82]. house [Frä90]. Howard [Bro83a, Hen82c, Mac72, McC77b, Swa86, Und79b, Web72, Wil84f]. Howlett [Bea85b]. Howse [Bry78a, Cha77b]. Howson [Gow78, GG84b]. Hsuin [Hal86c]. Hubble [Hal84c, Chr95, SN93]. Hubert [GG75c, Nor72]. Hudson [Fin68, Han76, Por77]. Hufbauer [Hal84d]. Hugh [McG68, Scr68b, Sme87, Web71b, But63, MS95, Sho96]. Hughes [Bow74, Car67a]. Hugo [Dea83a]. Hui [Agr83a]. Hui-Lin [Agr83a]. huitiéme [Sch01]. Hull [Ver89a]. Hulst [Dea83a]. humaines [CM93, MC^{+}93]. Human [Buc69, Edg87a, Grm69, Olb77, RH88, Tha74, Wil78, CT93, Deg91, Dun90, Goo90, Gra87a, Gro90, MJ89, Maz92, Ros94, Smi97, Sta85, Van93]. humani [Wea78]. Humanism [Bro80b, Hut88, Smi74a, GJ86]. humanist [Slo85]. Humanistic [Sch69a]. Humanists [Kel78, FJ93]. Humanities [Ben82b, Cha77a, Fie81, Fie83, Gow71, Hen82c, Hut88, May67, Nut84, Sha83a, GJ86, MMVZ90]. Humboldt [Din64, Hei88, Jac89a, Por83a, Hei88, Thé66]. Hume [Hei74b]. Humor [Lar74]. Humphry [Kni71c, Mil83, Rus68, Cro73, Kni92a, SD76]. Hund [Dor75]. Hundred [Bro87, Din63b, Por81, Sme70, Smi87c, Tur81, All86, Hea86, Hea90, Pep89, vLP82]. Hunt [All77, Bro84d, Cha88a, Mon78, Nor80, OM75, Wea83]. Hunter [Gol87b, Gou82, Gou84a, Lev73, Poy70, Web65a, Ste90a, BP02]. Hunterian [OS92b]. Hunting [Hal77b, Sha81, Mac88b]. Huntington [Fig77a, Sch78b, Sch79a]. Hurst [Ric89]. Hussein [Mol69]. Hutchinson [Blo84, Can75, Duf81, Far68a, Lin74, Nor82, Sec85]. Hutton [Eyl68, Mor72a, Rud80, Woo83, Ger71, Rud62]. Hutton-the [Eyl68]. Huxley [Bar76b, Fri74a, WV92, CHG91, Des94, Fri74b, Jen88, Jen91]. Huygens [Arm69, Sch83a, Sch84c, Sch84a, Ano82l]. Hydraulic [Smi89b, Gar87]. Hydraulics [Smi89b, Gar87]. Hydrographic [Man88]. Hydrologic [Whi70a]. Hygiene [Pic79, Pro88]. Hypatia [Ali86, Out87]. hypnosis [CS92]. hypnotism [Gau92]. hypothèse [Gay92]. hypotheses [Cla89]. Hypothesis [Fra87a, AH85, Gau83, Gay92, Sch65, Ogi75]. Hysteria [For87c, MF86, Mic95].

I-VIII [Mol85]. I. [Sto88, Wil86a]. Iain [Pel85a]. Ian [Bro82a, Col81a, Gol79, Mor73a, Web84, For87a, Sec85]. IBM [Hen85a, Pug84, PJP91]. Ibn [Fre86, Kel75, Sar81, Win75, AS89, AW71]. IBSN [Hoc88a]. Icelandic [Bar91]. ICL [CK89]. Ida [Bud84]. Idea [Bar77, Bro86d, Por78a, Sch87, Ste67, Cro86, Mur89, PR88]. ideal [Mor82a]. Ideas [Ale85, Ber79, Cor78, Dou84, Frä76, Gil84, Hen86e, Kni92b, Law81, Sch84b, Und66a, Whi68c, Ano05, Fre90, Hen82c, Lau87b, Mor90a, Ric92a, Rid83, Wie93]. Idée [Sch72]. idees [Han85a, Lau87b, Osb89a]. identités [GGR96]. Identity [Pic88b, Fox88, GGR96, Tur96, War86]. ideological [Ric92b]. Ideologie [Cap81b]. Ideology [Can88, Fur85, Wer75, BM82, Llo83, Out84a, Rid83, You83]. If [Bur75, Bro87, Kni89e]. Ignace [Str73]. Ignaz [Wei85a]. Ihde [FF66]. ihre [Gro89, Von88, Wol87]. ihren [Bro89a, vLvHH^{+}88]. ihrer [For69b]. II [Ait71, Ant64, Arm64, Can74, Coo89a, Fau89, Fox70, Fox71a, GG77a, Har75, Hat87, Hen85b, Mac88a, May67, Von88, Nic85, Nor69a, Sme69, Tur72b, Tur73d, Wal77, Wey92, Wol87, Bro82a, Bry78a, McC80, Mor89c, Rap79, BEoCF92, Cla95a, ECL92, Goo83b, Goo88b, Har96, Hop76b, Jev64b, Swa83, WW86, Wol87]. III [Cow68, Edg84, Eme87, Hil71a, Nor70a, Out88b, Rav69, Sme69, Swi71, Bry78a, Dor79, Fox79, Kni86b, Mar89, Sch80b, Smi78a, AS89, Lan87]. ijkwezen [Tur82]. Ikonologie [For69b]. Ilkley [Can74]. Illingworth [Lev73]. Illinois [Duf83, Mor77a, Rap71, War89]. illness [Law83]. Ills [Fre86, Hal86b]. Illus [Chi71, Fin68, Pac71, Swi71, Web71b, Bro89b, For67b, For69a, Rus67]. illusions [Won93]. Illustr [Far71b, Hil71a, Men70, Pac70, Web69a, Goo68, Gre71]. Illustrated [Con86, Grm69, Hen84c, HS88, Smi75b, Tur78a, HH86, HH92, aHH86b, Nut82b, Ron83, Sta84a, aHH86a, aHH88, aHH92, Chi68]. Illustration [Blu79, BW90]. illustrations [Ant63, CD81, Can86, Dav99, For71a, Hew66, Law65, Nun88, RT92, Sch79a, Tur88, Web66b, Yox82b]. Ilomets [Geo78b]. Ilya [Hen82c, Bru86]. Im [Wes70, CM90, Dau98, Jun89, Kra91, Sch83b, Whi68a]. Image [Cro87b, Fox74, Dol96, For86]. Imagery [Joh78]. Images [Ano04, Bur80, Jen87, FL88, Mir94, PLP92, Wea88]. imagination [HS95, Smi94a, Wil90a]. imaginative [CL96]. imaging [Sta91]. immagine [Sch81b]. Immobility [Ait86, Gra84b]. immortal [She95]. immortality [Out83]. immunology [Maz95b, TC91]. Impact [Hac76, Rav70, Hen82c]. impacts [Bro82b]. impenetrability [Gau82]. imperfect [Seg94]. Imperial [Mor87a, Ros86, Wei91]. Imperialism [Bal91, Gro95, Mac88b, Sta89a]. Impex [Hil75]. implications [Ver85]. Impolite [Gol95]. Importance [Cor78, Gil83]. impulses [Ros94]. impurity [Ham90]. Imre [Dol71b, Blo79]. Ina [Sha79]. Inadequacy [Rid82]. inches [Cal63]. Incident [Fri74b]. incl [Kni89e, Yox82b]. Inclined [Hum67]. Including [Jon87, Ber93b, Bry77, Dol73a, KK85, Nay75, SSH85, Yox83a]. incognitae [Rav69]. incommensurabilitate [Mol73]. incorporated [Ste91a]. incorporating [Dra81, Nay85]. incubator [Bak96]. Independence [Rud63]. indeterminism [Blo84, PB82]. Index [Ano83l, Ano84l, Ano85l, Ano86l, Ano87q, Ano88o, Ano89p, Eag78, Kni71b, Mor69a, WW93, Ait71, Blo79, Coo89c, Cro69, McG71b, WW75, WR92, Yox82b]. India [All84b, Ear70, Hil75, Wuj85, Des82c, Hil75]. Indian [Hil75, Rap71, Wuj85, Des92, Kum91, San88, San91]. Indiana [Buc66, Dol74, Got71, GG84a]. Indians [ML89]. Indias [ML89]. Indicators [Jam86, Law80a, T^{+}85]. Indies [PR95]. Indigenous [Ear70]. individual [Can74]. individuals [Mor83a, Wea83, Hen89e]. Indivisibilibus [Mol89, WW88b]. Industrial [Buc71, Cah88, Edg87c, Fri87, Hen87b, Nev78, Pay74, Smi71, Web87, Wor88, CM90, CK90, ES90, FG93, HL86, Hun85, Jam89, Jor85b, Ken86, Ler85, Rei85, Tho87, Tup82, Wis85, Buc71, Hil71a, Mac88c, Smi71]. Industrialisation [Por89a, Sch88, Tup82]. industrialization [Ste91a]. industrie [CM90, Tan82]. industrielle [CM90]. industries [Buc85b]. Industry [Bow89a, Cla86b, Clo80, CM90, For87a, Gre71, Hil89a, Ink87, Mor75b, Mor89a, Wil65b, Wil85b, Aft91, Ber85, Fox84, Hen85a, Hen89f, Lie87, Mor90b, Poc88, Pug84, Swa88, Tra93, Wes88, Wil81a, Wil94a]. Inedita [Bar73a]. Inédits [Bur74a, vPvP90]. Inequality [Tan62]. Inertia [Ron86, yQ85]. Infancy [Kel83, Sin84]. Inference [Lor77]. infini [Bla93, GG75b]. Infinite [Joh81]. Infinitesimal [Boy70]. Infinity [Hen84j, Kre82, Bla93, Fie97]. Influence [GG74b, Por85b, Ste62d, Ano90, Bow90, Rit90, Tun92]. Influences [Byn80]. Influential [Hel63]. Information [Hac89, Ora76, Bau95, Ben86, CKA96, NOF96]. Information/Publishing [Ora76]. informed [Nic88b, Söd86]. infortunes [dCC09]. Ingeneis [Kel73]. Ingénieurs [Fox87b, Gre86]. Ingenious [Kel75]. Ingenuity [Kel76]. Inglis [GG85, McV80]. Ingram [Cow68]. inhalation [SE91]. Inhibition [Smi92]. initiative [OO89]. Inkster [Bro82a, For87a, Sec85]. inland [Wuj85]. Inner [Gow69, Kra86, Mon78]. innere [Hei93]. Innovating [Hen89f]. Innovation [Hil89a, Ink87, Sie91, Wes88, Wil89b, Ber85, Lie88, McG95]. Innovations [McC77b]. Inquiries [Ano77l]. Inquiry [Ano75j, Ano75k, Wil84d, Coh94, MJ89, Pow96, Wyn82]. Insanity [Wal81b, Bus85, Har83a]. Insects [Pic79]. Inspired [Mon78]. Institut [AH93, CM90, Gua89, AH93]. Institute [AL87, Bad83, Bro84d, Cal77, CM90, Col83, Hen89g, Kar76, Lin86, McV75, Mol85, Por85a, Sco67, Sha83a, Sne74]. Institution [Can74, Col76, Duc88, Goo85b, Hac74, Hac76, Hac77, Hei73, Ink88, Kar76, Kni71b, Out80a, Rec86, Woo85, Wor88, Cra84, Gou84a, Hun82, Hay74, Mar62b, Mar64, MJ84, PLP92, Rob80b]. institutional [Lin92]. institutionalization [Frä90]. institutionalizing [GJ86]. Institutions [GG77b, Mor76a, Mor76b, Cas85, Fis83b, Gab83b, Hen89e, Mor83a, Nol95, Kar76]. instruction [Bry92]. Instrument [Bry78c, CW87, De 85, Har87b, Lil85, McC92, Tur87a, Wil87d]. Instrument-Maker [Bry78c]. Instruments [Ano69h, Bal83, Bry74, Bry76, Bry77, Bry78a, Bry78c, Cha80c, Chi68, Chi71, Hal68b, HS95, HS83, Law80c, McC85b, Mor83d, Nut83, Rya70, Sol83, Tur68, Tur69, Tur70b, Tur72b, Tur77a, Tur78c, Wil87d, ABR93, Ben84, Ben87, B^{+}89, De 85, Fox82, Hac82b, HAB92, McC81, McC82, Nut82b, Tur83b, Tur90, Wil94a, Bry76, Fox82, Hac82b, Hac82a, Nut82b, Tur82]. Insulin [Lie86, Bli82]. Integration [Cla71, Rav71]. Integrative [Cla71]. Intellect [For87a]. Intellectual [Byn78, Col81a, GG74a, Hal80c, Hew66, Sch74b, Hey81, Hun95, Moo84, Mor84c, Wal90, dG01]. intellectuelle [dG01]. intelligence [Cre93]. intended [McG70b]. intensive [Bak96]. Inter [Bar89c, Hai78, Luc90]. Inter-Relations [Hai78]. Inter-War [Bar89c, Luc90]. Interaction [Cha77a, Ben82b]. Interactions [Ubb63]. interdisciplinary [Doe96, Kit92]. Interest [Ano62e, Ano63h, Ano63i, Ano64i, Ano65g]. interference [Kip91]. Intermaxillary [Wel67]. Intermedia [Rav67]. internal [Hei93]. International [Ano68j, Bry78b, Bry78c, Buc85a, Bur74b, Cro72, Dol71b, Dor80, Dou84, EK91, EK92, G.62, Gil84, Hal65a, Mac78, Mac88a, Mad88, Nee78, Rob80a, Rud83, Rud89, Sar81, Sch84b, Smi80b, Tur70c, Tur73d, Aft91, A^{+}87, BG^{+}92, Cas85, CM90, N^{+}88, Sar87, Woo92]. Internationales [Rus74, CM90, Han85a]. Internationalism [Wei89, Cra92]. Internet [Tur96]. Interpretation [Slo79, Jam84, Maz95b]. interpretations [Mic95]. Interregnum [Hen82d].

Interrelations [Can80b, Mol79].

Interscience [Sme66]. Intertional [Bry76]. intr [Kni88c]. intro [Bro83b]. Introd [Dur77]. Introduced [Moo77, Fre86, Ste62a, Wil85a]. Introduction [Ait72, Ant64, Arm66, Bro72a, Buc85b, Buc68, Car74, Cro71, Cro72, Dol73a, Dug77, Fee77, For72, For85a, Gil84, Got86b, Got66, Gra81, Hal71, Hal77a, Has72, Hei74a, Hen88b, Hod70, HS86b, Jar84b, Kri83, Lar74, Lev73, Mol81, Mol89, Mor72a, Nee69, Olb87, Rat66b, Ric88a, Sch85a, Scr68b, Sha81, Sme69, Sme71, Tha74, Til75a, Wei85a, Wel70, BL83, Bar75, BM89, Bro82b, Bro84c, Car67a, CAMR87, CLR87, Cop92, CK90, Dun69a, Eas86, EiT92, ET92, Got86a, Gun84, Hal68c, Hal68d, Hal93, Hey83, Kra87, Kuh68, Lin84, McV75, Nor71b, Nor82, Rus79b, Sch69a, Sir90, Tak92, Wal77, Web65b, Web68a]. introductions [Mol78b]. Introductory [Rog88a, Old86]. Inventing [Mac88c]. Invention [Bow87a, Bry77, Car84, Chi71, Hac76, Hof87, Kei84, Scr68b, Sho89, Wie93, Agr83b, Bow89e, Fie97, FC96, FIF86, Gil83, Sch92a, Wil95]. Inventor [Bow74]. Inventors [Lev65]. Inventory [WR92, Olm92]. Investigaciones [Kni88c]. Investigation [Tei70, WHS92]. Investigations [Mor75c]. investigative [Hol89]. Investment [Cou96]. Invisible [Nee85, Hir83, II90, Wil95]. Inward [Hen89a, Pai86]. ionic [Cra96]. Iowa [Whi89b]. IR [Ham84, O’H86]. Iraq [MM10]. Ireland [Ham84, Kni89c, N^{+}88, Smi85a, Can83, Her83, Her95]. Irene [Bro72a]. Iris [Gli83b]. Irish [All79, Cow68]. IR£ [Kni89c]. Isaac [Ait73, Bro78a, Bro78b, For68, Hal86a, Hal88a, Hei76b, Kub70, McG69, McG71b, Nor68a, Nor69a, Nor70a, Nor73a, Pep72, Rog83, Til75a, Til76, Whi72a, Whi73, Wil65a, Fei90, Hal92, Her80, Hom86a, NS84, New89]. Isabel [Rav66a]. Isabelle [Bru86]. ISBN [Aas87, Ait85a, Ait86, Ait89, All84a, All84b, Asp88, Aus86, Bad89, Bal84, Bea85b, Ben84, Ben89, Blo84, Bod84, Bow87a, Bow89a, Bow85, Bow86, Bow87b, Bow87c, Bow88a, Bow88b, Bow89b, Bow89c, Bra88, Bra84, Bra87, Bra89a, Bro84a, Bro85a, Bro86a, Bro86b, Bro89a, Bro89b, Bro89c, Bro86c, Bro89e, Bro84c, Bro84d, BH84b, Bro85d, Bro85e, Bro86e, Bro87, Bro89f, Bru86, Buc89a, Bud84, Byn87, Byn88, Cah88, Can84, Can86, Car84, Cas85, Che85, Cla86a, Cla86b, Col87b, Col88a, Col88b, Con86, Coo89a, Coo84a, Cro85, Cro87a, Cro88, Cro89a, Cro89b, Cun88, Cun89, Dea86a, Dea86b, Dea86c, Des88, Dou84, Dra86, Duc88, Dun87, Ear87, Edg84, Edg87a, Edg87c, Ege89, Eme87, Eva89, Fal89, Fau89, Fel89, Fie86, Fie89, Fig88, For85a, For87a, For87b, For87c, Fox87b, Fra87a]. ISBN [Fra87b, Fra87c, Fre86, Fri87, Fur85, Gab88, Geo87, Gil88, Gil84, Gol87b, Gol88b, Goo88a, Goo89b, Gou84a, Gou85, GG84b, GG85, GG87, GG89, Gra83b, Gra84c, Gra85a, Gra86a, Gra86b, Gra87b, Gre88, Gua89, Hac88a, Hac88b, Hac89, Hal84e, Hal84c, Hal84d, Hal84a, Hal85a, Hal85b, Hal85d, Hal86a, Hal86e, Hal86b, Hal86c, Hal86d, Hal87a, Hal87b, Hal88b, Ham84, Han85a, Har84a, Har84b, Har89a, Har87b, Hea87b, Hen83a, Hen83b, Hen84a, Hen84c, Hen84b, Hen84d, Hen84f, Hen84e, Hen84g, Hen85a, Hen86b, Hen86a, Hen87a, Hen87b, Hen88a, Hen88b, Hen88d, Hen88c, Hen89a, Hen89b, Hen89c, Hen89d, Hen84j, Hen84k, Hen84l, Hen84m, Hen85b, Hen85c, Hen86e, Hen87c, Hen88e, Hen89g, Hil86b, Hil86a, Hil89b, Hil89a, Hoc88b, Hod89a, HS86a, HS86b, HS88, Hom86a, Huf85, Huf86, Hun89a, Hun84]. ISBN [Hun86, Hut84, Hut88, Hut89, Ink87, Ink88, Ins86, Jac89a, Jac89b, Jac85, Jam86, Jam87a, Jam87b, Jar84b, Jen89, Jon87, Jor85a, Jor87, Jor88, Jud88, Kau84, Kee88, Kee89, Kei86, Kei89, Kel84, Kil84b, Kil84a, Kni84a, Kni84b, Kni85, Kni86b, Kni86c, Kni86d, Kni88a, Kni88c, Kni88b, Kni89b, Kni89c, Kni89e, Kni89d, Law85, Law87, Law88, Law89, Lay85, Lew85, Lie85, Lie86, Lin86, Lin89b, Liv86, Mac85b, Mac89, Mah89, Mar89, Mau87, McC85b, McL86, Mea84, Von88, Mil87, Mol84c, Mol88, Mol89, Moo85, Mor87a, Mor84b, Mor84c, Mor84d, Mor87b, Mor89a, Mül87, Mül89, Nay85, Nee85, Ner89, Nic85, Nic88b, Nor85a, Nor85b, Nun88, Nut84, Nye86, O’H86, Olb85, Osb85, Osb89a, Osb89b, Out84a, Out85, Out87, Out88a, Out88b, Pat86, Ped85, Pel85a]. ISBN [Pep86, Pic84a, Pic87a, Pic88a, Pic88b, Pin87a, Pin87b, Pin88, Por89a, Por86a, Por87b, Por88, Por89d, Por89e, Por89b, Por89c, Por89g, Pum84, Pum87, Pum88, Pye87, Rap86a, Rav84, Rec86, Red84, Ric88a, Ric89, Rob87, Rob89a, Roc87, Roc88b, Rog88a, Rog88b, RH86, RH88, Ron86, Rud84, Rud86, Rud88b, Rud89, Rup85, Rup88, Rup89, Rus87a, Rus89a, Rus85, Rus87b, San87, Sch84b, Sch84c, Sch86a, Sch86b, Sch87, Sch85a, Sec83, Sec84, Sec85, Sec88, Sei85, Ser88, Sha89, Sha86, Sha87, She85, Sho87a, Sho86a, Sho88, Sin84, Sme84a, Smi85a, Smi85b, Smi85c, Smi87b, Smi87c, Smi87a, Smi89e, Smi89b, Smi89c, Sti89, Stu85, Stu87, Swa86, Tei88, Tha84a, Tha84b, Tha85, Tha87, The87, Tho89, Tre87a, Tre87b, Tre88, Tur87a, Tur88, Twe89, Van86, Ver89a, Ver89b]. ISBN [Wal89, War88a, War88b, War89, Wea84, Web87, Wei85b, Wei86, Whe89, Whi85a, Whi89a, Whi84, Whi85b, Whi88a, Whi89b, Wil84e, Wil84d, Wil84c, Wil85a, Wil85b, Wil86a, Wil86b, Wil87d, Wil87c, Wil87b, Wil89b, Wil89a, Woo89, Wor88, Yeo88]. Isis [Bro73a, Fis83b, Fox79, Hen84g, Ano82n, Ano82m, NW82b, NW82a]. Islam [Mol69, Huf93, Huf95, Mol69, Sta90]. Islamic [Fre86, HS88, Smi75a, HH86, HH92, aHH86b, Sar81, aHH86a, aHH88, aHH92]. island [Gro95]. Isles [All77, Bro87, Sch71, All86, Coo89c]. isomorphicism [Sch84d]. Isomorphismus [Mau87, Sch84d]. Israel [Bow87a, Nut84]. ISSN [Buc85a, Hen89e, Ree85]. issue [Ano62a, Ano62b, Ano62c, Ano63a, Ano63b, Ano63c, Ano63d, Ano64a, Ano64b, Ano64c, Ano64d, Ano65a, Ano65b, Ano65c, Ano65d, Ano66a, Ano66b, Ano66c, Ano66d, Ano67a, Ano67b, Ano67c, Ano67d, Ano68a, Ano68b, Ano68c, Ano68d, Ano69a, Ano69b, Ano69c, Ano69d, Ano70a, Ano70b, Ano70c, Ano70d, Ano71a, Ano71b, Ano71c, Ano71d, Ano72a, Ano72b, Ano72c, Ano72d, Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74e, Ano74b, Ano74f, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano80a]. issue [Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Dol69]. Issues [Smi80b, Hen89e, KH95]. it. [Nay85]. Italia [Por79a, Sch79b]. Italian [Cun83b, Cun89, Hen85b, Kel78, Sch78b, Tur83a, Ara91, Des89b, Ern91, GGC90, Lan91, MP89, Maz95a, Olm92, Out82b, Sir87, Yat83]. Italy [Cun89, Ber95, Pan91, Sir87, Tor88]. Items [Ano62e, Ano63h, Ano63i, Ano64i, Ano65g]. Ithaca [For74a, For87c, Hal86d, Hen84j, Kni83b, Por87c, Rud84, Sch84c, Sch78b, Tha83a, Und75b, Und79b, Wal89, Wil75]. IV [Asp88, Bry76, Cro88, For68, GG70, Pep72, Rat68, Swi71, Ben84, Cor81, Cro89a, Fox68b, Gol79, Hal85b, Hen82a, HS83, Rap75, Rud83, Sme69, Woo80b]. Ivo [Sim84]. Ivor [Dub72]. IX [Hil89b, Hop75, Kra86, Mol86, Ait85a, Bar77, Bar76a, Bos70, Bro89c, Bro89e, Buc71, Buc66, Bus85, Cro87a, Dea83b, Dor80, Dun87, Ear87, Fal89, Far67, Fox69b, Fre82, Geo87, Hal84e, Hea87b, Hen84f, Hen88a, Hen83c, Hod89a, Joh81, Jon87, Kni67, Lay85, Liv86, Mor69a, Mor73a, Mor82a, Mor89a, Out87, Pic80, Pum87, Pum88, Rec86, Sch84c, Smi85a, Smi87a, Smi89b, Sti89, Tei84, Van86, Wil68].

J [Aas87, All77, Ano65j, Arm70, Bea85b, Ben79, Bos70, Bow87b, Bow88b, Bro85a, Bro83b, Bro86d, Bro86e, Bry76, Bry78a, Bry78b, Bur80, Bur74b, Bur71, Can80c, Cha78b, Chi65b, Cla67, Col81b, Cow74, Cro70b, Cro79, Cun83a, Dah97, DTF97, Des88, Dol70a, Dol71a, Dol83, Dor79, Duf83, Duf75, Dun69a, Dun86, Dun87, Dur83, Fal87, FF66, Fau89, Fee77, Fie86, For68, For71b, For80, For87c, Fox69a, Fox71a, Frä76, Fre82, Gab83a, Ged71, Gin74, Goo88a, GG71, GG74a, GG77b, GG81, Gra83a, Gre71, Hac88b, Hal77a, Hal80c, Hal86e, Hal87a, Hei68, Hen82c, Hen83a, Hen83b, Hen87a, Hen88a, Hen84l, Hen87c, Hew66, Hig65, Hil71b, Hoc88a, Hoc88b, Hut81, Jam86, Jam87a, Jar84b, Kni86d, Kni88c, Kni88b, Law81, Law89, LB73, Loh75, May66, McC67, McG71b, Men74]. J [Mol78b, Moo84, Mor84b, Mor76a, Mor84d, Mos77b, Nev80, Nev84, Nor68b, Ora76, Osb85, Out85, Pel85a, Pep71, Pic79, Por78a, Por81, Por82, Pum84, Pum88, Rap75, Rav65, Rav66b, Rav68, Rav69, Rav72, Rus67, Sch83a, Sch84a, Sch87, Sch68a, Sch71, Sha89, Sha83a, Sin76, Sin77, Sme70, Sme68b, Smi83, Smi89c, Str73, Str75, Swi70, Tei88, Tha84b, Tho76a, Tre88, Tur70c, Tur72c, Tur73d, Tur87a, Tur88, Und64a, Und66c, Was69, Wea83, Whi68a, Woo80b, Woo89, Wor79, Yox82b, Yox83a, Yox83b]. J. [Ack87, Buc85b, Chi65b, Dah97, DTF97, Fal87, Frä76, Gra84a, Hac89, Hen85c, Hom86b, Hun89a, Jud88, Lin89b, McC67, MC81, Mül89, Pin87a, Roc87, San87, Smi87c, Tha87, Tur72c, Tur88, Wil89b, Yox83b]. J.S [Mor87b]. Jaar [Tur82, Bur99]. Jaargang [Phi67]. Jac [Rus67]. Jack [Gol79, Sec85, Ste62a, Pat86]. Jackson [Bur71, Gab83a]. Jacob [Dur79a, Hal80c, Hen84m, Kni75, Olb72, Mid65, PW89b]. Jacques [McC77a, Wea84, You66, Yox83a]. Jacquot [Str75]. Jacyna [Har89a]. Jagdish [Stu78, Hen86b, Hen88c]. Jahnke [Can84]. Jahre [AH93, For67b]. Jahrhundert [CM90, Dau98, Kle94]. Jahrhunderts [For69b, Pag69, Jun89, Sch89]. Jaki [Mil87, Sey77]. Jallabert [Hom86a]. James [Col83, Coo84a, Eyl68, Hea87b, Hei74b, Hen87a, Hen84m, Huf86, Kni83c, Kni89e, McC69, Mor69a, Mor71, Mor72a, Mor73b, Mor83c, Pic80, Pye87, Rob89b, Rud80, Smi85b, Swi70, Tur78a, Tur81, Wal88, War88b, Web75, Woo83, Bir86, But66, Car89, Col88b, Cro84a, Eme87, Ger71, Gol83a, Goo81b, Har02, Har98a, Hen86c, Ost84, Por87a, Sch86a, Ste62b, Tay96, The87, Wri87]. Jan [GG77a, Byn88]. Janet [Sho86a]. Janis [Bro84c, Out88b]. Jannis [Tur75]. January [Cha80a, Goo85b, LB73, Sme84b, Sme87]. Janus [Ros79, Dob91]. Japan [Bro82a, Gou95, MR89, Nak91, YK94]. Japan-Forschung [Bro82a]. Japanese [Rav70, Bro81]. Jaramillo [Sec84, Gle82]. Jardine [Ait85a, Hut88]. Jarrell [Mos77a]. Jauregg [Whi93]. Jazzari [Kel75]. Jean [Ait68, Ben89, Bow89c, Cla70, For71a, Gil84, Hah73, Hen88b, Hom86a, Pau76, Sin74, Str75, Stu86, Fra78a, Hey83, Hil86a, Jor83a, Pic87c]. Jean-Baptiste [Ait68, Fra78a]. Jean-Michael [Hil86a]. Jean-Pierre [Jor83a]. Jean-Robert [Gil84, Hey83]. Jeanne [Tei78]. Jefferson [Hol86]. Jeffrey [Por87b]. Jenner [Fis91b]. Jennett [Ste62a]. Jeremy [Fig88, Ser88]. Jerome [Rav72]. Jerrold [Gol92a]. Jersey [Cha78a, Cun89, Hai82, Hen84l, Hod89a, Jar83, Jon87, McC77b, Moo84, Olb71b, Por87a, Wei85b]. Jerusalem [Cun78, Gil84, HE^{+}82, Nut84]. Jessie [Fre81b, Poy70]. Jesuit [Kra91]. Jesuiten [Kra91]. Jet [Kei84]. Jevons [Sch90]. Jewish [Har81, Lor77, Fre92, Pat94, Rud88a]. Jill [Fie89, Hut89, Tur81]. Jo [Hal77b, Sin74]. Joachim [Pag69]. Joan [Hen84k, Kei84, Ste62b, Pag82, Tor87]. Joe [Smi76a]. Johan [Cro67]. Johann [Sch78c, Lin90]. Johannes [Jar84b, Mol72a, Mol75, Nor68b, Nut84, Rus72b, Wes70, Hal85d, Mor91b, Sch93]. Johannus [Rav69]. John [Ait67, Ait71, All84a, Aus86, Bar81, Bow74, Bro85b, Bry77, Bry78c, Buc77, Buc68, Chi64, Deb66, Eag78, Far71b, Far77, Fis83b, Goo81b, Goo68, GG77b, Gra84c, Hal66a, Hal84c, Har77, Hew66, Hil86b, Kni69a, Kni74, Mac79, McG69, McG70a, McV75, Mor77b, Mor89a, Mus74, Nee69, Nor68b, Nor71b, Olb82, Ora73, Pel85b, Por77, Por84b, Por85a, Poy70, Pra70, Rat66b, Rat69, Ree80, Rid83, Sch84b, Sch69a, Sha77b, Sha73b, Sme66, Smi73a, Stu85, Sut74, Whi68a, Whi72a, Woo84, ATKL08, Asp90, BB82, Bra89a, Bro85d, Che85, Clu88, Dea86b, DF98, Edg87a, Fie89, FC82, FFLW97, Hal86b, Hen88d, Hil86b, Kni89c, Liv86, McC68, Pen89, Pep86, Pic88b, Pou93, Ril85, RWD90, Smi79, Stu88, Sul82a, Tha66a, Tha67, The87, The78, Web87, Wil85a]. Johns [Bro71a, Buc75, Cal77, Chi68, Chi71, Clo80, Coo83, Gel76, Gur78, Hal80a, Hop72, Jac85, Kni76, Law81, McG65, Mor83b, Mor89a, Nor85b, Phi69, Pic76, Pic80, Pin83, Por85a, Ree80, Ros79, Sch70, Sha86, Spa65, Web78, Yeo88]. Johnson [Arm66, Buc68, Can86, Dun69a, McG68, Rat69, Rav69, Sch74a, Sme69, Web69b, Wil85a]. Jonas [Wil93]. Jonathan [Bow89b, Bro83a, Bry77, Dyb78, For82, Mor75a, Sch78a, Wea84, Yox79, Kil84b, Mor89a]. Jones [Bar76a, Ben82b, Gol79, Mar89, Mol88, Rus77c, Str75]. Jordanova [Bur80, Out85]. Jordanus [BJ91, Tho74a]. Jorpes [Rus67]. Jose [Smi76c, Fig78, Goo89b, Kni88c]. Joseph [Agr83b, Bar83, Cla70, Cri66, Cro72, Dur77, Goo85b, Gow73, Hac74, Hac77, Hai80, Hei76a, Hen82b, Hen89g, Joh81, Jor83b, Kni69b, Lar74, Mol71, Mor71, Nay80, Nev83, Olb71a, Pac72, Por79b, Por85b, Rav73, Sch72, Ste81, Tyl75, Whi75, Woo83, AL87, Buc77, Cro83, Gol88a, Hal86c, Hal88b, NR86, Ner89, Sho87a]. Joshua [Law80a]. Josiah [Cha75]. Josten [Rat66b, Web68a]. Joule [Goo81b, Car89, MC81]. Jourdain [Joh79]. Journal [Buc85a, Cas85, Fox68b, Ste62a, Wea83, Ano71h]. Journals [Mor81, Rud71, MD88, Dur83, Fis83b, Fis83a, Mor83a, Tur83a, Wea83]. Journey [Rau75]. jours [Ano87r, CAMR87, CLR87, Gua89]. Jovanovich [Bar77]. JPL [Kei84]. Jr [Cah88, Din63a, Goo82, Hen82c, Lew85, Mar74, Moo84, Sch78b, Sch79a, Sho88, Web75]. Juan [Fal89]. Judd [Wel70]. Judgement [Lay85, Mac82]. judgment [Pet94]. Judith [Cla71, Wea79]. Judy [Hen84d]. juillet [Bur74b]. juin [BG^{+}92]. Jules [Jen89]. Julian [WV92]. Julius [Dau75, Kra84b, Whi93]. July [Ano64l, Ano67j, Chi65c, CT95, Sme78b]. June [Ano64l, BG^{+}92, CT95, Kni71c, KK85, LHH91, OS92b, Rat66b, Sme84b, Sme87, Woo80c]. Jung [Nol94, Sha98, YED97]. Jungiu [Pag69]. Juni [LHH91]. Just [Bro84b, Bro85d]. Justus [Bro89a, Far76, ZH92, Bro97, ZH92, vLvHH^{+}88].

K. [Cla86b]. Kaavere [Geo78b]. Kabbalah [Cou95, Rud88a]. Kahl [Hei74a]. Kaleckian [Cou96]. Kamerlingh [GG91]. Kames [Mor73a]. Kangro [Pag69, Sin78]. Kansas [Rus77a]. Kant [Har84b, Pum84, Sch87, Cro86, Dic82, Hat90, Wal78]. Kantons [Gra85b]. Kapitza [Bad85, Bro86c, KBRS90]. Karachi [Sar81]. Karger [Whi68c]. Kargon [McG68, Mor72b, Wil89a]. Karl [Hal84d, Pat84, Por82, Rus74, Web70, Blo84, Mei91, vBDGO93]. Karrieren [Dei92]. Katherine [Cas82, Hal85a]. Kauffman [Bro89b, Sme68a]. Kearney [Web71b]. Keele [Web66a, SH73]. Keeler [Huf86, Ost84]. Keepers [Smi78c]. Keeping [Cla86a, Tom84]. Kegan [Dol83, Fis83a, For87b, Goo75, Hal68d, Har84b, Hea87b, Hen85b, Joh78, Kel71, Law88, Mor75b, Mor77c, Ric88a, Rog88b, Sha87, Tur73a]. Keilin [Jev67]. Keith [Bow89c, Hai77, Yat72]. Kekulé [Roc81b]. Keller [Sin84]. Kellner [Din64]. Kelly [Rat66a, Twe89, YK94]. Kelvin [Buc89a, Smi76a, Smi78a, Wil89a, Tun92, Wil87g]. kemiska [Fis81c]. Kemp [Cun83a]. Kennedy [Hal84e, Lor73]. Kenneth [Bar76a, Irw83a, LM74, Mar74, Smi75b, Web66a, Tei88]. Kenney [Hen87b]. kennissen [Phi67]. Kent [Whi85a]. Kepler [Ait85a, Arm66, Bra88, Hal66a, Hal69a, Jar84b, Lin89b, Mol75, Mor87a, Nor68b, Rus72b, Rus74, Sch78c, Wes70, Fie88, Her65b, JK84, Ros86, Rus64, Ste87, Ste94, Tho74b]. Kepler-Symposium [Rus74]. Kepleriana [Wes70]. Keratoscope [Lev65]. Kessler [Hut89]. Kevin [Wil86b]. Kevles [Hoc88a]. Key [Gra81, Rob86, Kil94]. Keynes [Bro82b, Duf75, Ear87, Kni80a, Por84b, Sha73b, Web67a, Web69a]. Keynes/Philadelphia [Ear87]. Khaledy [Ear70]. Khalid [Win75]. Khan [Cun78, Wuj85]. Kholodinin [Hen82c]. Khwarizmi [Gra87b, vdBLW85]. Kidd [Moo85, Cro84b]. Kiel [For67b]. Kieler [For67b]. Kilmister [Hen89b, Whi68b].

Kimball [Hen81]. Kinckhuysen [Scr64b]. Kinckhuysen-translation [Scr64b]. kindly [Duf81]. Kinematics [Mol73]. Kinetic [Mid65, TP66]. King [Byn72, Law80b, Lay74, McN83, Rus72a]. King-Hele [McN83]. Kingdom [AB89, HAB92]. Kinnon [Hen82c]. kinship [Web84]. Kipp [Pat84]. Kirk [Osb85]. Kirkbride [Cla86a, Tom84]. Kirkby [Whi72a]. Kirkwood [Hal66a]. Kirtley [Dav71, Rud69]. Kirwan [Cro69, GGSGB94]. Klasse [Phi67]. klassieke [Bur99]. Klaus [Bad89, For69b, Har83a, Wil87e]. Kluwer [Mol89]. Knidos [Scr68a]. Knight [Blu79, Cro70a, Far68a, Far71a, Fis83c, Lev80, Rob89a, Swi77, Tur78b, Wal74]. Knopf [Hoc88a, Kee88, Woo73]. know [CP93]. Knowledge [Baa90, Col81b, Cro87a, Dol71b, Gow74, Gur78, Hen89c, Joh78, Kel75, Lay65, Lay85, McE79, Mor77c, Mor81, Mor87b, Pic88b, Rog88a, RH86, Yeo88, Arm83, Bea91, Dol96, G^{+}09, GG91, Gre92a, Hac82b, Hue88, Law85, Lig87, Lun94, Mac82, Max84, Nor83, Old86, Olm92, PR88, Rud85, Sir90, Vuc84, War86, WT82, Yeo93]. Knowles [Chi68, Chi71, Din63a, Duc84, Hop72, Spa65]. Known [Ped85, And82, SLT88, vPvP90]. Knoxville [Cun88]. København [Bad83]. Koffka [Smi85d, HK83]. Kogan [Dav81]. Kohler [Tei84]. Kohlstedt [Mor77a]. Kohn [Des88, Smi89e]. Kolloquium [CM90]. Kolodny [Por78b]. Kombinationslehre [Whi68a]. Kommentare [Scr68a]. Kommentare-eine [Scr68a]. kommentiert [Scr68a]. kommentierte [ZH92]. Komposition [Gro89]. Kong [Agr83a]. Königsberg [Ole91]. Koning [Tur82]. Koninklijke [Phi67]. konservativer [Mei91]. Kontroverse [Fri91]. Koppes [Kei84]. Korrespondenz [HRP90]. Kors [Woo89]. Korshin [Woo89]. Kosmologie [Mol72a]. Kovalevsky [Hal84e, Ken83].

Koyré [Ano64l, For74a, Hal68c, McG66, Tur73b, Whi73, Her65a]. Kr [Bal83, Fis81c, Mor78, Str73, Cro67]. Krafft [Rus74]. Kragh [Hen88b]. Krakatau [Woo85]. Kräupl [Law83]. Kraus [Kel70]. Kraye [Hut89]. Krebs [Tei83a]. Kremlin [Bad85, Bro86c]. Kretzmann [Hen84j]. Krieger [Ait86, Fig77a, Rud89]. Krige [Bar81]. Krimsky [Olb85]. Krisciunas [Wil86b]. Kristallographie [Sch89]. Kritik [Jun89]. Kronick [Cha65, Cha78a]. Kt [Web69a]. Kubla [Cun78]. Kuhn [Hil89b, Kni81, Ber79, Lug79, Mac85b, Mul79, Sch79c]. Kultur [Dau98]. Kunsten [Phi67]. Kursunoglu [Hen89b]. Kurt [GG87, Smi85d, HK83]. kurz [Ger89]. Kuznick [Hoc88b].

L [Abe86, Arm67, Bro84c, Buc71, Buc73a, Bur80, Can80c, Cun88, Dav71, Din63a, Din64, GG75c, Gra87b, Hah73, Hal84a, Hal85b, Ham84, Har89a, Har82d, Hei73, Hei76a, Hen83b, Hen87a, Hom86b, Jac89b, Jam86, Jor87, Kar76, Kni65, Kni69a, Kni71b, Law65, Law81, Lay74, Lor77, Mil87, Mol72b, Mol84b, Mol85, Mor73b, Mor77b, Mor80, Mor83b, Nor68b, Nor73b, Ora74, Out80b, Pac70, Por79a, Por87c, Rud70, Rus87b, San87, Sch75, Sch78b, Sch79b, Sch80c, Sey77, Sho88, Sin76, Tha66a, Tre88, Van86, Ver89a, Wal77, Wat84, Wea79, Wel70]. L. [Nor73b, Out85, Ric88a]. L22 [Sch81b]. L25 [Sch81b]. L’abbé [FM00]. Labor [Goo89b]. Laboratory [Chi65b, Hoc88b, Irw82, Kei84, Mor76a, Sha86, And84, Cit90, Jam89, Kuz87, Cro74a, Cro74b, HE92]. labour [Fie83]. Lacan [CS92]. Lacer [Cro68b]. lactic [Ben90]. L’Afrique [G^{+}09]. Laissus [Bur74a]. Lakatos [Blo79, Dol71b, GG79b, Sch79c]. Lake [All84a, BB82]. Lamarck [Bow89b, Bur74a, Bur74b, Jor79, Jor83a, Osb89a, Out84a, Out85, BM82, Cor88a, Hod71, Jor84, Lau87b]. Lambert [GG81].

Lamp [Kni86d, BM84]. Lancashire [Nev84]. Lancaster [Bro86b, Byn87, Che85, Goo88a, Gou85, Han85a, Hen88e, Hoc86, Hol85a, Kni85, Kni89b, Por86a, Sha89]. Lancastrian [Ear87, Rus86a]. lancets [Agr83b]. Land [Has72, Kni88a, Por78b, FS84, Moy86, Sch84a]. Landforms [Hig65]. Landmarks [GEMM83, Nic85]. Landsat [Mac90]. Landsberg [Whi84]. Lane [Hal88a, Kel76, Wal81b]. Langdon [Bro85d]. Langenstein [Lef78]. Langford [Rav72]. Langham [Web84]. Langins [Out88b]. Language [Byn80, Gre63, Ril85, Sha86, And84, BVA95, NPH92]. languages [Hun84, Sla82]. Lapidibus [Got66]. Laplace [Des89b, Des89b]. Laporte [Bow88a]. large [GH92]. large-scale [GH92]. largely [Mol71]. Largs [Smi78a, STW90]. Larry [Gau83]. Lasserre [Scr68a]. last [Buc85d]. Late [Bro75, DeK73, Lay85, Mor80, Out87, Pic88a, Sei85, Ali86, Ern91, Gol91, Hun95, Mac82, Moy83, O’M89, SP88, Sho86b, Sme66, Mol72a]. Later [Nut83]. Latin [Eas86, Hos67, Mol81, Mol85, Rus79b, Sty68, BEFH10, BM89, Cop92, EiT92, ET92, Tak92]. Latour [Irw82, Ste91b]. Laudan [Bro89f, Gau83]. Launching [Kee88, Bru87a]. Laura [Bro78a, Bro78b]. Laureates [Har79b]. Laurent [Osb89a, Poi93]. Laurie [Cas85]. Laurier [Duc84, Gin84, GG89, Sha83a]. Lavington [HS83]. Lavoisier [Bru72, Col87b, Cro71, Cro88, Rus79a, B^{+}95, BV93, BVA95, Ber93a, CS92, Don93, Gou68, Hol85b, LF97, Mor72e, Per73, Per89b, Poi93, Rap69, Rap73]. Law [Fox68a, Sin78, Swi70, Bur99, Har89b, Hen84l, Sha83b, Hom68, Hum67, Tho74b, Wes67]. Lawlor [Kni89c]. Lawrence [Bro86c, Col87b, Fie86, Hei71b, Hen82b, Hen84d, Hen89g, Kel78, Mau77, Out84b, Rus77a]. Laws [Coc75, Whi71, Rus64]. Lay [Por78b]. Layton [Rob75]. Lazare [Men74]. Lead [Tyl64]. leading [Gun84, Hen85a, Pug84]. Lear [Hal66a]. learn [Bro83b]. Learned [Gur78, Gra98]. Learning [Kni86d, Wal89, Whi72a, BM84, Cun83b, Dea88, Gol95, Hal91, Ped92b]. Leavitt [Wea79]. Leben [Gra85b]. Lebensborn [Har87b, Lil85]. Lebesgue [Dug77, Rav71]. Lebrun [Bro83b]. l’École [FD87]. Lecture [Rud76, Cro83, Hol86]. Lectures [Gra86b, Hal86a, Hal88a, Lev73, Mol71, Olb71a, Rap86b, Sha78, Sha77b, Ste67, Til75a, Wei80, Whi72a, Wil89a, BC98, Chi64, Hen82c, KT87, Kra91, Ler85, NS84, OS92b, SO85, Sch01]. Lederberg [Law80a]. Ledger [Kni89e, GL88]. Lee [Nic85, Ste62b, Win75]. Leeds [WG81]. Leeuwenhoeck [vLP82]. Leeuwenhoek [Hal84a, PS82]. Lefebvre [Bry76, Bry78b]. Legacy [Sec88, BT96, Ste85, TP90]. Lehmann [Mor73a]. Lehrstuhl [AH93]. Leibniz [Ait89, Hal87a, Ilt77, Sch81a, Whi75, Ait85b, Ber93b, Cou95]. Leibnizian [Ilt73]. Leicester [Arm70, Bro67, Far67, Far69, Gol79, Hal71, Hod70, Kni72, Rus69, Goo74]. Leiden [CT95, Hac75, Wil87d]. Leipzig [Mai85]. Lelorrain [Gua89]. Lenoble [Sch72]. Lenoir [Kni84a, Kni84b]. L’Enseignement [CAMR87, CLR87, Ano87r]. Leo [Dur80a, Bow88a]. Leona [Hen81]. Leonard [Bus74, Rud71, Web70, Edg87c, Gra85b]. Leonardo [Cun83a, Kel70, Sim88a]. Lerner [Hal77a]. Lesley [Hun86]. Leslie [Hen84d, Cha88a]. L’esprit [Pet84]. Less [Bec75]. L’essai [SG80, SG80]. Lester [Byn72, Law80b, Rus72a]. Let [Bra89a, F^{+}88, Mos77b]. L’Éther [BV82]. Letter [Hal80b, Pep67, Pir83]. lettere [GGC90]. Letteren [Phi67]. Letters [Ano77k, Bow88a, BKH90, Bro72a, Cor81, Hei71b, Hen81, Mor71, Ped92b, Por78b, Por87b, Sin76, Tom65, vLvHH^{+}88, AJ^{+}89, CMNH92, FMF85, Ger89, Gid82, GGC90, Har02, Her80, KBRS90, LM23, LHB04, LS76, LHH91, McN83, PvM85, Roc88a, SGŚ00, SL87, Tee84, WR92, FPLJ89, Gol95].

Lettre [Wil77]. Lettres [FPLJ89, Hal68c, Por89e, Roc88a, SGŚ00]. L’Europe [GGR96, vPvP90]. Level [Por84b]. levels [Bli92]. Levere [Bro73b, Bry76, Mos77a]. Levey [Ear70]. Levi [Swe76, Tre87b]. Leviathan [Jon87, SSH85]. Levidow [Fie83]. Levine [Por79b]. Leviticus [Wor79, Blo78]. Lewis [Bar77, Dra86, Pep89, Smi73b, Tup82, Ash85, Koh75, Pin87b]. Leyden [Bry76, Bry77, Bry78b, Bry78c, Fre78, Mac78, Rob80a, Ste62d]. Li [Agr83a, Bro78a]. Libbrecht [Nak75]. Libby [Hen81]. Liber [Mol81, Rav69, Eas86]. Liberal [Hut88, GJ86, Mor82a]. liberalism [Fre90]. liberalismo [Fre90]. Libraire [Har86]. Librairie [Bos70, Cro79, Dug77, Kel78]. Libraries [Ban88, BK75, Che78, Deb66, Wal77, Mor83a, Wea83]. Library [Ano77m, Ant64, Bro83b, Cam82, Che78, Col76, Day85, For80, Fre70, Hac76, Hew66, Hom86b, Jar84a, Kni71b, Mol89, Mor69a, Pac71, PS78b, Por85a, Sch68a, Sch78b, Sch79a, Sha86, Smi78a, Ste62b, Til75a, Und75b, Whi68b, Woo80b, WG81, And84, RWD90, You83, Goo74, Got86a, Got86b, Got92, Scr64b]. licences [WW88a]. Licensing [Wei80]. Lichtenberg [Bar73a]. Lick [War88b, Wri87]. Liebenau [Mor89a, Wil89b]. Liebig [Bro89a, Far76, Rus77a, Rus79a, ZH92, Bro97, ZH92]. lies [Sap90]. Lieut [Kni83a]. Life [Bod65, Bow83a, Bow88b, Bro83a, Bry76, Bry78b, Cla67, Cla68, Col81a, Col87b, Cro70b, Dea83b, Dra86, For85a, Fox71a, Fri74a, Fur85, GG74b, Hen87a, Jon87, Kni74, Kni83c, Kni86c, Kni89e, Kra86, Men81, Mil87, O’H86, Olb69, Olb71a, Pic87b, Por78b, Rud83, Rus67, Sch87, Sec83, Sec88, Sin76, Tei83a, Tha83a, Tha83b, Tre87a, Tre87b, Tur96, Tur81, Und80, Web76, Web67a, Web68a, Web71a, Wei86, Wig78, You66, Yox83a, Yox83b, Ack87, Ald85, All84a, Ash85, A^{+}87, BT96, Bäu84, BB82, BC89, BH90, Bro89d, Bud93, Can88, Can93b, Cha91, CB^{+}90, Cro86, Dic82, Dur80a, FB97, Fei90, G^{+}89, Gir86, Gol92a, GJ85, GL88, Gun84, Hey81, Hol85b, Irw82, Jak84, KBRS90, Kni84a, Kni84b, Kre90, Len82, Les83, Lev85, Lil85, Llo83, Mac85a, Moo89b, Mor84c]. life [Pai82, PS82, Por95, Pum84, Rav68, Red84, Ros90, SSH85, She95, Sou93, Ste85, Ste90a, dG01, vLP82, Agr83a, Bro83b, Gel76, Har77, Jor83b, Kri83, Les83, Rid83, Smi76c, Tei83b]. Lifting [Smi75a, Hal86b, Ole84]. Light [AO77, AW71, Bow87a, Hei75, McC67, Por89a, Stu71a, Swi71, Bec88, Buc89b, Can83, FIF86, G^{+}09, Hak95, Hal93, Jam84, Kip91, Mir89, Roc88a, Sch88, Sie91, Smi85a, Ste91b]. Lightman [Yeo88]. Lights [Lau68]. liii [Wig78]. Like [Geo87, Rus86b, Fis98]. Lilian [Kni86b, Phi69, Smi87a]. Lilienthal [Har87b]. Lillian [Cas85]. Limited [Gre63, Swi77, Tee84]. Limits [Yeo88, Cer89, Lig87]. Lin [Agr83a]. Linacre [Wig78]. Linda [Har82d, Por89c]. Lindberg [McV75, Nor71b, Rus87b, Sch85a]. Lindeboom [Han70]. Linden [Kni88b]. Lindqvist [HS86b]. Lindroth [Mor78]. Lindsay [Dau75, Mea71, Ste62a]. line [Chi65b, Cri66]. Linear [Gra85c, Ser88, Abe86]. Ling [Hal88b, Mol71, Pac72]. linguistics [NPH92]. Linnaeus [Bro84d, Rau75, Bro80a, FL^{+}94]. Linnean [Bro80a, Bro84d, Kni89e, Rau75, Wal88, Dou88, GS88]. L’inventario [Olm92]. Lipetz [Rav67]. Lire [Mor89c, Sch73a, She85]. Lisa [Hut88, Kei84]. Lischi [Sch81b]. Lisse [Sch83a]. List [FB75, Old77, Wes70, WW75]. Lister [Tur72c]. Lists [RW72]. Literary [Por87a, Dea91, Mor83c, Smi94a, WM85]. literate [Kro91]. Literature [Arm62, Arm64, Arm70, Cun78, Edw78, HW81, Hen83b, Lor77, Pet90, Por78b, Smi66, Wil78, Fox87a, Hay94, Hen84l, Maz95a, Nad83, Sha83b]. Lithium [Can86, Joh84]. Little [Hal84e, Ken83, vPvP90]. little-known [vPvP90]. liv [Kni88b]. Liverpool [Row90b]. Lives [Car67a, Rat69, New94, SY96, Smi78c, Bry83, Dea83a, Gab83c, Hen83c, McC83, Sch83a]. Living [Hod71]. Livingston [Hei75]. Livingstone [Kee89, Poy70]. Livorno [Sch81b]. Lloyd [Fur85, Got67, Mar74, Por85a]. Loads [Aus86, FS84]. Local [Ste62b]. locali [Sch75]. Locating [Jef64, Pic78]. Lochem [Tur82]. Lochem-Poperinge [Tur82]. Locke [Hen86e, McG70a, Ale85, Fox88, Sch01]. Lockwood [Rus77c]. Lodge [Row90b]. Loeb [You66]. Loeffel [Mah89]. L’Oeuvre [Hal65a, dGB^{+}89]. Logic [Byn80, Joh79, Nor77, Rav66a, KI92, PB82, PB83, Wal92]. logical [Des89b]. logiche [Des89b]. Logique [Olb72, GG77b]. Lomonosov [Hal71]. London [Ait67, Ait71, Ait72, Ait73, All77, All79, All81, All82, All84b, Ant63, Arm66, Arm67, Arm69, Asp88, Asq80, Bad89, Bar73a, Bar76b, Bea85b, Ben84, Blo84, Bod65, Bod84, Bow74, Bow83a, Bow87c, Bro69, Bro71a, Bro73a, Bro86a, Bro89b, Bro72a, Bro75, Bro78b, Bro84c, Bro86d, Bro89f, Bru86, Bry74, Bry78a, Bry81, Bry83, Buc73a, Buc75, Buc85a, Buc68, Buc73b, Bud84, Bus83, Bus74, Byn72, Byn73, Byn74, Cal77, Can75, Can80c, Can84, Can86, Car67a, Car69, Car74, Cha78b, Che78, Chi64, Chi65b, Chi65c, Chi68, Cla66, Cla67, Cla68, Cla86b, Clo80, Coc75, Col76, Col88b, Col81b, Con86, Cow74, Cow68, Cri66, Cro68a, Cro69, Cro70a, Cro71, Cro72, Cun83a, Dau75, Dav81, Dea83a, Dea86b, Des88, Dol70a, Dol70b, Dol71b, Dol73a, Dol73b, Dol83, Dor75, Dor80]. London [Dub72, Duf81, Duf83, Dun69a, Dun69b, Dur77, Dur79a, Dur83, Dyb78, FF66, Far68a, Far71a, Far76, Far77, Fee77, Fie81, Fie83, Fie86, Fig77b, Fig78, Fin68, Fis83b, Fis83a, Fis83c, For66, For71a, For72, For74a, For82, For87b, Fox69a, Fox71b, Fox79, Fre81b, Fre86, Gab83b, Gau83, Ged71, Gel76, Geo66, Gin74, Gli83b, Gol88b, Goo75, Goo81a, Got67, Got71, Gow73, GG71, GG72, GG85, Gra85a, Gre63, Gre71, Gur78, Hac76, Hac82b, Hac88a, Hac89, Hah70, Hal66a, Hal68d, Hal69a, Hal80a, Hal84e, Hal84c, Hal84d, Hal86d, Hal88a, Hal88b, Han70, Han76, Har77, Har84b, Har89a, Hea87b, Hea64, Hei68, Hei71b, Hei77, Hen81, Hen82c, Hen82b, Hen82a, Hen84a, Hen84d, Hen84g, Hen85a, Hen87a, Hen87b, Hen88a, Hen84j, Hen85b, Hen89g, Her80, Hew66, Hig65, Hil71b]. London [Hoc88b, Hod89a, Hol71, Hop72, Hop73, Hop75, Hun89a, Hut84, Hut88, Ilt77, Ink87, Ins86, Irw82, Irw83a, Jac85, Jev67, Joh79, Joh81, Joh83, Joh78, Jor79, Jos67, Kel71, Kel73, Kel76, Kil84b, Kni65, Kni69b, Kni71b, Kni76, Kni81, Kni83a, Kni83b, Kni83c, Kni86c, Kni86d, Kni88b, Kni89e, Lau68, Law65, Law80b, Law80c, Law81, Law83, Law88, Lef78, Lev73, Lin74, Lor73, Mac72, Mac79, Mah89, Mar73, Mar89, May66, McC77a, McC85b, McC69, McG65, McG66, McG68, McG69, McG70a, McG71b, McV80, Mea84, Men70, Mol71, Mol84c, Mol89, Mon78, Mor71, Mor72b, Mor74b, Mor75a, Mor75b, Mor75c, Mor76a, Mor77a, Mor77c, Mor81, Mor83b, Mor84c, Mor87b, Mor84f, Mor89a, Nak75, Nay77, Nay80, Nor68a, Nor68b, Nor69a, Nor70a, Nor73a, Nor77, Nor82, Nor85a, Nut82b]. London [Nut83, OM75, Olb71a, Olb85, Ora73, Ora74, Ora76, Out83, Out84a, Out84b, Out87, Pac70, Pac71, Pac72, Pag66, Pat86, Pel85a, Pep72, Pic84a, Pic76, Pic79, Pin83, Por77, Por79b, Por82, Por84b, Por85a, Por87b, Por89d, Por89b, Poy70, Pra70, Pye76, Ran74, Rap71, Rap75, Rap86a, Rat68, Rat69, Rau75, Rav66a, Rav66b, Rav73, Rav84, Ree80, Ric88a, Ril85, Rob75, Rog88a, Rog88b, Ron86, Ros79, Rud70, Rud71, Rud83, Rud84, Rud86, Rup88, Rus66, Rus68, Rus72a, Rus77b, Rus77c, Rus87b, Sch84a, Sch69a, Sch69b, Sch71, Sch73b, Sch74b, Sch74a, Sch78b, Sch80d, Scr68b, Sec85, Sha73a, Sha77a, Sha79, Sha86, Sha87, Sha77b, Sin74, Sin76, Sin77, Sin78, Sme66, Sme69, Sme70, Sme71, Smi73b, Smi74a, Smi76c, Smi76a, Smi82a, Sne74, Sol83, Spa65]. London [Ste62a, Ste62c, Ste62b, Sti70, Stu71a, Stu85, Sty68, Swi70, Swi77, Swi67, Tei78, Tha67, Tha83a, Tha84a, Tha84b, Tha85, Tup82, Tur70b, Tur73c, Tur73a, Tur74a, Tur77a, Tur78b, Tur81, Tur88, Twe89, Und64a, Und66b, Und75a, Und75b, Und79b, Van86, Vea64, Ver89a, Wal89, Wal74, Wal77, Wal81b, Wat73, Wea78, Wea83, Web75, Web76, Web78, Web84, Web65a, Web65b, Web66a, Web66b, Web67a, Web68b, Web69b, Web69a, Web71a, Web71b, Wei83a, Wei85a, Wei86, Wel70, Wes74, Whi72a, Whi72b, Whi85a, Whi85b, Whi77, Wig69, Wig74, Wig77, Wil65a, Wil75, Wil87c, Wil87b, Wil89a, Woo80a, Woo80c, Woo67, Woo80e, Wor88, Yat72, Yeo88, You66, Yox82b, Yox83a, All80, Bro81, Des82b, Des84, Des89a, Dou88, GS88, Gle91, Hay74, Ink79, Mil83, Mor76b, PS78b, Rud63]. London [Wal88]. Long [Hod77, HS85, TBDA10]. Longman [Hen84a]. look [Gue84, Roc90, Whi64]. López [Goo89b, Kni88c]. Lord [Mea71, Mor73a, Mor73b, Pai82, Red84, Smi76a, TC90, Tun92]. L’ordine [Des89b]. Loren [RH88]. L’organisation [HJF89]. Lorin [Ped85]. Lorraine [Por89g]. Lost [Rud80, Gre92b]. Loughborough [Hal68a]. Louis [Cro79, GG75b, Sme84a, Cla89, Gei95, GGSGB94, PR95, SG80]. Louis-Bernard [GGSGB94]. Louisiana [Col88b]. Louviers [Hil86a, Cha84]. Lovers [Ped92b]. Low [Jam87a, Rus79a, MP89]. Lowell [Sch87, Cro86]. Loyd [Goo82, Hen82c]. Ltd [Ant63, Aus86, Ben84, Cro70a, Ede65, Hac82b, Hun89a, Joh83, Lay85, Mah89, Nye86, Ora76, Por89a, Scr68b, Tur83c, Twe89, Wea83, Web66a, Whi85a, Woo84]. Lu [Agr83b, Hal88b, Mol71, Pac72]. Lubar [Wor88]. Lubbock [GG77b]. Lubrano [Har82d]. Lucida [Fie89, HA87]. Lucienne [Dug77]. Lucile [All82]. Ludwig [Phi69, Wes70]. Ludwik [Col81b, Löw90]. Luigi [GGC90]. Luis [Bro89e, AT87]. Luisa [Tur72b]. Lull [Per89a]. Lumières [Por89e, Dou84, Duc82, G^{+}09, Roc88a, Sch84b]. Luminiferous [Wil72]. Luminy [EK91, EK92]. Lunacy [Pic88a, Sho86b]. Lunar [Fox68b, GW74, Hal69a, Wil65b]. Lundgren [Fis81c]. Lundskjær [Hen88b]. Lundskjær-Nielsen [Hen88b]. luoghi [Olm92]. Luther [Bra87, Zam86]. lvi [Cro67, Kel68]. lvii [Pac72, Ste70]. Lwoff [Yox83a]. lxi [Rud71]. lxii [Wel70]. lxvi [Hen88d, McC69]. lxviii [Bro73a]. lxxii [Hil75]. lxxxi [Sch85a]. lxxxiv [Tur83a]. lxxxv [Mol84b]. Lydia [Rap86a, Tre87a]. Lyell [Bus74, F.76, Rud71, Bar73b, Bar76c, Can76, Cor78, McC76, Mor76b, Ord76, Osp76, Pag76, Por76a, Rud76, Rus76a, Sto76, Tom65]. Lyellian [Rud74]. Lynn [Pag66]. Lyon [Hal77b, Por84b]. Lyte [Nev83].

M [All77, Arm62, Arm64, Bar75, Blo79, Bro67, Bur71, Can84, Cas85, Col76, Cri71, Cro70a, Cun83a, Dea83b, DD95, Dol83, Dra86, Dun86, Dur80a, Edg87a, Far68a, Far69, Far71a, Fig77b, Fre86, Gel76, Geo78b, Gil88, GG71, Gre63, Hal71, Hal84a, Har84a, Hei76b, Hen82c, Hen84f, Hen83c, Hod89a, Hun84, Hut81, Jam87b, Jar84b, Jor78, Kni88c, Kra84a, Lev80, May66, McC77b, McG65, Mor81, Mor84f, Nor68a, Nor69a, Nor70a, Nor73a, Ora76, Osb85, Out84a, Pep89, Pep72, Pic88a, Por79b, Rus87a, Rus69, Rus72a, Rus79b, Sch83a, Sho86b, Sme71, Smi66, Smi83, Smi85c, Tha83b, Tur70c, Tur73d, Tur82, Und66c, Und75b, Wal74, Wea78, Wuj85]. M. [Blo82a, Tur87a].

M.D [Deb84, Wal88, Web69a]. M.D. [Sta84b]. M.I.T [Ait73, Car67a, Cla67, Cla68, Coc75, Cro68a, Dub72, Fee77, Fox71b, GG72, Gra83a, Kel73, Kni69b, May66, Men67, Nak75, Out84a, Pac71, Pep89, Wel70]. MA [Asp88, Fra87a, Pic88b]. MacCormack [Dur81]. MacDonald [Pac70, Byn72, For72, Gin74, Rap75, Rud70, Sch73b, Sin74, Whi72b]. Mach [Von88, Wol87, GG74b, Wol87]. MacHale [O’H86]. Machamer [Can80b, Mol79]. Machian [Bar89a]. Machine [Bry77, Car70, Hom80, dLT96, Rus89a, Bak96, CKA96, Cha91, Her88, Tan82]. machinery [Ric92a]. Machines [Fin71, Koh78, MKW92]. Machinisme [Hil71a]. MacKenzie [Nor83]. MacKinnon [Kra84a]. MacLaurin [Tee84]. Maclean [Col81a]. MacLeod [Lay85, Mor81, Ora76, Kni89d]. Maclure [MD88]. MacMillan [Hen84d, Ril85, Bar73a, Bro72a, Can86, Dol73b, Dur83, Hen87a, Ins86, Kni89e, Mor74b, Mor76a, Mor84f, Mor89a, Rud83, Sch74b, Smi76c, Smi76a, Tup82, Tur74a, Und75a, Woo80e]. Madaule [Out84a]. Maddison [For74a, Ilt77, Kel68, Rus85, Web71a, Wig78]. made [TFC93]. Madeleine [Out84a]. Madison [Bow86, Bro85e, Col87b, Deb66, Fel89, Hal69a, Hop73, Hop75, Hos67, Jor85b, Kni75, Lew85, Nay75, Nor71b, Rav71, Web66c, Wei85a]. Madmen [Har83a]. Madness [Por87c, Wal81b, Har89b]. Madrid [Gli84, Gou84b, Kni88c]. Magazine [Dav99]. Magda [Bro73a, Fox79, Hen84g]. Magdeburg [vGA94]. Magi [Gol91]. Magic [Por89c, Sch78b, Sha81, Yat72, Adl87, Bry88, Gol91, Hal90, Rud88a, She91]. Magie [Chi65a, Gol91]. Magie-Verständnis [Gol91]. Magier [Gol91]. Magierin [Gol91]. magische [Adl87, Fig88]. Magnes [Gil84, Pum89]. Magnesia [Lar74]. Magnet [Hom74]. Magnetic [Pic87a, Cra93, Goo80, Pum89, Zan86]. Magnetism [Hom74]. Magnus [Col83, Sch69b, Ste75, Zam92]. Mahoney [Gab75]. mai [Ano93, A^{+}87]. Maienschein [Bow89c]. Maillet [Rap71]. Main [Cap81b, Hen88d]. Mainz [Kra91, Kra91]. Maiocchi [Mor89c]. Maison [Out82a]. Majesty [All84b, Ben84, Hac88a, She63, Wil87b]. Make [Por85b]. Make-Believe [Por85b]. Maker [Bry78c, PR88]. Makers [Bry74, Hal88a, Wil87d, De 85, McC92, Tra93, Tur82, Wil94a]. Making [ABR93, Edg87c, FB97, Fis98, Kei89, Rap79, Ric89, Ben90, Dav99, Fra78a, Fri83, Goo90, Hal84c, MH88, Mis95, Rei85, Rus85, Tei84, Wei82, Wil87f, Wil94b, Hom88]. Malabar [Ait86, Rud89]. Malabaricus [Mül89, Hen86d]. Malcolm [Dyb78, Wea84]. Malebranche [Bos70, CMDM82]. Malheureux [BV82]. Maloine [Hai80, Jor83b]. Malpighi [Und79b]. mammoths [Van93]. Man [Bar79, Bar73b, Bru86, Buc77, Byn74, Gel76, Han70, Hea64, Hei68, MD90, Mea71, Nor80, Nye76, San87, Smi72, Web75, App93, Bow90, CT95, Cun83a, Fis83c, For82, FL^{+}94, Hal84c, Hei86, dLT96, PS84a, PS84b, Sta87, TFC93, Wei82, Wil94b, Jev64b, Sch80d]. man-midwifery [Wil94b]. Management [She86]. Manager [Can74]. managers [H^{+}82, Hen84d]. Manchester [Ait89, Bow89a, Buc71, Bur71, Cro85, Har84a, HS83, Jam87b, Kar76, Kni69a, Law89, McL86, Nev84, Smi71, Smi85a, Smi87a, Tha66a, Web87, Yox82a, Ano77m]. Mandarin [Pau84]. Mandelbaum [Bow89b, McG65]. Mandelbrojt [Dug77]. Mandelker [Wei85b]. Mandrou [Bro80b]. Manier [Byn80]. Manierisme [Sch81b]. Manipulation [Col76]. Mannheim [Bro89a]. Mansell [Bro73a, Che78, Fis83b, Fox79, Hen84g, Mor81, Ora76, Sha77b, Tha84b, Wal77, Woo80c]. Mantova [Sch81b]. Manual [LM74, Bry92]. Manuel [Kub70]. Manufacture [Hil71b]. Manufactures [Ink87, Ber85]. Manuscript [Col88a, Cor81, For74b, Jef64, Lev68, Olb63, Smi78a, Tur74a, Bec88, Gra81, Tee89, TW87, Vic88]. Manuscripts [Smi87a, Ber93b, CF86, For73, New89, NJ91, OS92a]. Manuscrits [Bur74a]. Many [Ham84, Her83]. map [Ste62b]. Mapping [Ham84, Her83]. Marangon [Sch80c]. Marc [Bud84, Ink88]. Marcel [Rav84]. Marcello [Ner89, Und79b]. March [Deb84, HE^{+}82, N^{+}88, Tor80, Woo80b]. Marci [Rav69]. Marconi [Ged71]. Marcus [Rav69]. Marek [Rav69]. Margaret [Boy70, Far76, Har84b, Hen84d, Sha83a, Wat73, Wig78, GR88, Out87]. Margarita [Por83a]. Margery [Hal68d]. margin [Tay96]. marginados [Rey90]. marginalia [DDG90]. Margot [Kni89e]. Marguerite [Por84a]. Maria [Con86, Tur72b, Goo89b]. Mariano [Kel73]. Marie [Buc66, Gua89, Hop73, Hop75, Rau75, Sch84c, Tre87a, Web66c, Web69b, Woo80c, GG65, Gir86, RG65, Rob87]. Marie-Francoise [Sch84c]. Marilyn [Dur81]. Marin [Hal77a, McC80, McC83, Hal85c]. Marine [Buc85b, Und80]. mariner [Chr95]. Marino [Sch78b, Sch79a]. Mario [Sin81, Bro86b]. Maritime [Wil85a]. Mark [Eas86, Lud80]. market [Pri83]. Markets [Mir94]. Marquis [Wil69]. Marquis-Who [Wil69]. mars [Ano82l]. Marseille [EK91]. Marsh [CMN95]. Marsha [Sha83a]. Marshall [Hen81, Hos67, Sch80b]. Marsilio [Sch81b]. Martha [Kra86, Mor82a, Wes70]. Martin [Ait89, Bar79, Bry78c, Cun83a, Ear70, Hen85c, Jud88, Nad83, Pye76, Rap75, Sme66, Tei83a, Gal72, Hen87b, Mil73, Mor87b, Smi85b, Twe89]. Martinez [Fal89]. Martinez-Alier [Fal89]. Martins [Pic84a]. Martinus [Bry76, Cro70b, Dou84, Fox71a, Gil84, Goo88a, Hac75, Han85a, Hen83c, Mil87, Mor73a, Pep71, Sch84b, Sch74a, Bry78b]. Martyr [BMLC84, Jam87a]. Marum [Bry76, Bry78b, Cro70b, Fox71a]. marvellous [Cun83a]. Marvels [Bro84c, Lin86, OH85, PP82]. Marx [May67]. Marxist [Fie83]. Mary [Blo82b, Bry74, For80, Han85a, Sin74, Sol83, Tor87, Pat69, Pat83a]. Maryland [Sch68a]. Masani [Asp88]. Masha’allah [Lor73]. Maskelyne [How89]. Mason [Dav71, Nor68b]. Mass [Ait72, Ait73, Bea85b, Bow83a, Byn74, Car67a, Cla67, Cla68, Coc75, Con86, Cro68a, Dub72, Dur79a, Fee77, Fox71b, GG72, GG75a, Gra83a, Gra85a, Hac89, Hal84c, Hen81, Hen85a, Her80, Hut84, Joh81, Jor79, Kel73, Kel75, Kel85a, Kni69b, Kni71c, Kni74, Law80b, Lor73, Mau77, May66, Mea84, Nak75, Olb85, Out84a, Pac71, Por87b, Rav67, Rav84, Rus69, Sch76, Sec88, Smi82a, Tur70b, Wel70, Whi73, Wil86a, Row90a]. mass-energy [Row90a]. Massachusetts [Bad89, Chi65a, Edg87a, Edg87b, Hod89a, Kei84, Men67, Sch84b, Wil89a]. Massebegriff [Mol72a]. Masson [Bur74a, Por87b]. Master [Gab88, Hei75, Sch93, Sch92b]. Masters [Bow85, Out85]. Matematika [KI92]. material [CK90, Fre95, Kro91, Sti84]. Materialism [Cha80b, McG71a]. Matérialisme [Sch74a]. Materials [Cor81, Wuj85]. Mathematica [Bos70, Mah89, Whi73]. mathematicae [Viè83]. Mathematical [Ait71, Bra84, Cow68, Dub72, Gab75, GG84a, GG89, Gra84c, Hal84e, Hei76b, Hen84a, Hen84c, Hen84b, Hen84l, Jam87b, Mol78a, Mol80, Mol81, Nor68a, Nor69a, Nor70a, Nor73a, Pep72, Ric88b, Sch84c, Sco67, Tee84, Whi72a, Ben82a, Bla93, Cor96, Dau95, Gue84, Had94, Hen83a, HF87, KI92, LS76, Sch90, Sch65, Whi75, Bad83, GGR96, Goo74, Gra83b, Gra83a, Hen83a, Hen83b, Hen83d, Joh83, LHB04, LHH91, Mor83b, Tur83c, vL76]. Mathematician [Ben85, Can93b, Tal62]. Mathematicians [Eag78, Gre88, Kel78, Mor87a, WW93, Gin88, Ros86, WW75]. mathématicien [Dug77]. Mathematics [Cow68, Fau89, GG71, GG74a, GG77b, GG79a, GG79b, Hai77, Joh81, Kel78, Kno85, LM74, Mol78a, Mol88, Nak75, Nor68a, Nor69a, Pra70, Scr68b, Tan62, Tee88, Whi68a, Whi85a, Fie97, For73, GG84b, How82, Høy94, Jar83, Joh83, Lin82, Mol68, O’M89, Por85d, Sch89, SH73, WW86, Wil93, Bos70, Bos72, Boy70, Dub72, EK88, GG70, GG71, GG72, GG74a, GG75a, GG75b, GG77a, GG77b, Hah70, Hof62, HS63, KI92, Kra91, LM74, Nor70a, Nor72, Pra70, Rav66a, Rav71, Scr64a, Whi72a]. Mathematik [Bos72, Rav65, Sch89, Whi68a, Kra91]. mathématique [Bla93, GGR96, GG75b]. Mathematiques [GG70]. mathematischen [Bos72]. mathematischer [LS76, LHB04, LHH91, vL76]. Mather [Dav71, Rud69]. Mathias [Sha73a]. Matrix [Hen86b, MR82b]. Matter [Bea85a, Bec75, Bro73b, BH84a, Cap81a, Fre79, Fre81a, Geo78a, Gli83a, Hea87a, Hei71a, Hei71c, Hen89a, Kni71a, Les83, Lev68, Lin89a, Lov75, Mol79, Mol84a, Nut82a, Nye86, PS78a, Rus72a, Sme78a, Tur70a, Wil82a, Ano62a, Ano62b, Ano62c, Ano63a, Ano63b, Ano63c, Ano63d, Ano64a, Ano64b, Ano64c, Ano64d, Ano65a, Ano65b, Ano65c, Ano65d, Ano66a, Ano66b, Ano66c, Ano66d, Ano67a, Ano67b, Ano67c, Ano67d, Ano68a, Ano68b, Ano68c, Ano68d, Ano69a, Ano69b, Ano69c, Ano69d, Ano70a, Ano70b, Ano70c, Ano70d, Ano71a, Ano71b, Ano71c, Ano71d, Ano72a, Ano72b, Ano72c, Ano72d, Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74e, Ano74b, Ano74f, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano76a, Ano76b, Ano76c]. matter [Ano76d, Ano76e, Ano76f, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e]. matter [Ano89f, Ano89g, Ano89h, Ber93a, Bro85c, Bro95a, Har82b, Hen84f, Nye93, Pai86, Sti84]. Matter-Theory [Hei71a]. Matthew [Bro86e, Dem83]. maturity [Whi75]. Maurice [Bry74, Cal63, For69b, Fox71b, Hen89b, Hes74, Hil71a, KW87, Sme69]. Maurizio [Hat87]. Mauskopf [Pin83]. Max [Bro72a, Bur74a, San87, Wes70, Hei86, Jac85]. máximos [GB95]. maximum [GB95]. Maxine [Fie81, Ink87]. Maxwell [Chi64, Hen89c, McC69, Roc87, The87, Tur83c, Wil68, Buc85d, Gol83a, Har02, Har98a, Hen86c, Sie91, Smi85b]. Maxwellians [Hun89a, Hun91, OP87]. May [Ano65j, Ano93, Dun86, LM74, PCA91, Woo80c, A^{+}87, OS92b]. Mayer [Bar73a, Dau75, Hom74, Wil84b, Can93a, For70]. Mayr [Bow83a, Koh78, Web65b]. Maze [Kel76]. Mazzolini [Duc88]. McCann [Fis81b]. McCartney [Nev80]. McClellan [Eme87]. McConnell [Ben84, Nut82b, Tur81]. McCormmach [Bad71, Kel85a, Mor77b]. McCosh [Hal86e, Moo84]. McDowell [Bur71]. McGehee [Coo83]. McGrath [For87c]. McGucken [Bro71a, Kei86]. McGuire [Hen85c, Woo80b]. McIntosh [Mül87]. McIntyre [Rud80]. McKenna [Kni89c]. McKenzie [Bro83b]. McKie [Ano67i, Mor71]. McKinney [Smi73b]. McLaughlin [Rav67]. McLean [Smi74a]. McMackin [Mor82a]. McMillan [Kni89c, Mor84f]. McNeill [Wil84e]. McVaugh [Pin83]. MD [Por86a, Und80]. Meadows [Arm70, Bry78a, Cha78b, For71b, Fre82, Gol79, Kni86d]. Mean [GG89, HF87]. Meaning [Koh89, Law87, Rap75, Coo84b, Dol83, Fri91, Goo90, Ric92b, Smi92]. meanings [Nye90]. Meant [Jev64a]. Measure [Hen88a, Høy94, Wil87a]. Measurement [Hod89a, Cha90c, Cha94, GG91, Sti86]. measurements [Tun92]. Measures [Hac88a, Con87, Tur83a]. Measuring [Hac82b, Van85, Wil87c, Smi94c]. Mechanic [Kar76]. Mechanical [Ant63, Car66, Hal86b, Kel68, Kel75, Ole84, Whi90]. Mechanics [Cal77, Che85, CK90, Hal66b, Hen86b, Hen88c, Hom68, Hom74, CGBE95, Hen84i, Kni84a, Kni84b, Len82, MR82a, MR82c, MR82b, MR82e, MR87a, MR87b, MR00]. Mechanism [McG71a, Smi78b]. Mechanistic [You66, SG96]. médecin [Cla70]. Médecine [Cro68b]. media [Gol79]. Mediaeval [Col83, Lin86, Mol85, Wuj85]. Medical [Byn72, Col81a, Cun89, Deb66, Ear70, Fig77b, Fig78, Gol87b, Hai78, Hol71, Jor78, Lar74, Law80c, Lie87, Mor89a, OM75, Pic88b, Por88, Ste62a, Und64a, Vea64, War89, Wig77, Wil75, Arm83, BPH85, BP02, C^{+}95, Cun83b, Deb91, DPP88, FB97, Gab83b, Gab83a, Gos87, Hue88, Law85, Mat95, Nic88a, Pes84, Por85a, Ram88, Sir87, Tei84, War86, Woo83, WT82, Agr83b, AB89, Bus83, Cun83b, Gab83b, Gab83a, Irw83a, Law83, Wei83a, Woo83]. Medical-Historical [Und64a]. Medicee [Sch81b]. Medici [Sch81b]. medicinal [ML89]. medicinalium [Mer78]. Medicine [Bro85e, Byn73, Cla67, Cla68, Cla70, Cla71, Coo81, Cri71, Cro68b, Cun88, Dav81, Ear70, Fre70, Fre74, Fre86, Geo66, Grm69, Had63, Han76, Har87a, Hod70, Law80b, Löw90, Mac79, Olb69, Olb70, Olb71a, Phi69, Pic80, Por85a, Poy70, Ros79, Rus72a, Sch68a, Sch71, Smi66, Und66a, Was69, Web66a, Web67a, Web87, Whi77, Wig74, WG81, You66, AL87, Bal91, Bus85, BP93b, Coo83, Coo88, Des89a, FL88, Gab83b, Got92, Grm90, Har89b, Har83b, Hen84k, Hen89g, HH97, Law93, Lev84, Mor91a, Nev84, Pag82, Pro88, Ram88, Ric91, Sir90, Sir97, Sta91, VoC89, WT82, Grm90, AL87, Byn73, Far77, Fig77a, Fig77b, Fig78, Fre78, Hol71, Law81, Lev73, Lud80, Num87, OM75, Ros79, Smi75b, Und75a, Wea79, Wig77, Wil75]. Medico [Cla70, Hue88, Rey90]. Medico-Historica [Cla70]. médico-sociales [Rey90]. medics [WW88a]. Medieval [Fre86, HS86a, Kre69, Lin68, Lor77, Mol74, Mol84b, Nor77, Nut83, Sch76, Sch80b, Sir90, Wea78, Bea91, BJ91, Cro90a, Cro96, EiT92, ET92, Eva77, Gra94, HW87, Hen84j, Hil84, Kre82, Mol84c, Tak92, Wal81a]. médiévaux [Bea91]. Mediterranean [HW87]. Medvedev [Har82d]. Meeting [Dun86]. Meetings [Can74, Ink79, Wei80]. Mehra [Hen86b, Hen88c, Stu78]. Mehrtens [Cap81b]. Meier [Tre87a, Wei86]. Meiji [MR89]. Mel [Gow69]. Melanchthon [Kus95]. Melbourne [Bro89c, Gol88b, Jac89b, Sch80d]. Melders [Sch83a]. Mellor [Fis74]. Members [All80, Cor81, Sol83, Stu95]. memoir [MS95]. Mémoires [Sme69, BV93]. Memoirs [Sch80b, BV93, Hom83a]. Memorial [Bro83b, Cro83, Hom86b, Woo80b]. Memories [Hen85a, Pug84, Sho87c]. Memory [Blo79]. Men [Din63a, Dol73a, Fre78, GG77b, Tur78c, Und66a, Van93, Ros91]. Menard [Nun88]. Mendel [Bow85, Coc75, Ano69i, Cam85, Olb65, Ore84, Wil62]. Mendéléev [BV82, BV86]. Mendelian [Olb89]. Mendelism [Web66b, Olb87]. Mendelsohn [Bod65, Pic84a]. Mendelssohn [Mor74b]. Mendenhall [MR89]. Méndez [Fre90]. Menlo [EJR94, EJR95]. Menno [Rat66a]. mens [Phi67]. Menschen [MD90, Pat84]. Menstrie [Ben79]. Mensuration [Mol84b]. Mental [Bar89c, Ric92a]. Mentalities [Dea86a, Vic84]. mentally [Nol95]. Mercator [Scr64b]. merchants [Had94]. Mercury [Wil84c, Ros82]. Meric [Hen83c]. merken [Tur82]. Merritt [Edg87b]. Merry [Nic88b, Söd86]. Mersenne [Hal77a, Hal85c, McC80, McC83, Wal89, Dea88]. Merton [Bar77, Col81b, Law80a, Mol68, Mor75c, Web72, CDS90]. Meryon [EE95]. Merzbach [GG75a]. Mesmer [Cra93]. message [SB09, Dug77]. messager [GP92]. Messenger [Arm66, Ber95, GP92, GV89]. metabasis [Liv85]. Metals [Cla86b, Mad88]. Metaphor [Por78b, Rus87a, Sti84, TC91, You85]. Metaphors [Lea90, Lea94, Tei78]. metaphysical [BGH^{+}87, Gar92]. Metaphysics [Ait89, Eva89, Har82c, Har82a, Hen84f, Hul67, Fur88, Gau82, McG95, vLMB88]. Métaphysique [Sch74a]. métaux [Wil77]. Metcalfe [Wat84]. Metchnikoff [TC91]. Meteorological [Chi71, Bur86]. Meteorology [Chi68, Fri89, Neb95]. Meteoroscope [Nor66]. Method [Dol73b, Dol74, Hah70, LY66, Pin87a, Bro83b, Gro91, yL66, SY86, Sha93, Gow78]. Methode [GG81, Out80a]. Methodology [Old72, Rog88a, Woo80d, Gau83, Old86]. Methods [Byn73, Fel89, Hei74b, Hen89c, Llo91, Col87a, Fox95, Fox96, Max84, Nev84]. Methuen [For74a, Ged71, Han70, Hig65, Ilt77, Kil84b, Mar73, Rog88a]. Metres [Cal90]. Metric [Law80a]. Metropolis [Sec85, Baa90]. Metuchen [Cha78a]. Metzger [Hal72b, Tur72a]. Meyer [Rus74].

Michael [Bra89a, Bry81, Buc68, Duf83, For72, For80, Gab75, Gil84, Gou82, Gou84a, Hai77, Hei73, Hen87a, Lin74, Mol72a, Mol75, Mor75b, Mor80, Pin83, Ste75, Can91, FTG91, Fre86, Gil88, GJ85, Hil86a, Kni65, Lie86, Pag66, Sch87, Tho89, Ver89b]. Michel [Fig77b, For74a, Hal68c, Hal77a, Sch72, AO77, Gut89, Sch86b]. Michel-Pierre [Hal77a]. Michelet [Dol71a, Jor80]. Michell [McC68]. Michelle [Cro79, Sme84a]. Michelson [Hei75, Mar74]. Michelson-Morley-Miller [Mar74]. Michigan [Har75, Mol74, Rav72, Sch69a]. Microbiology [Und79a]. microchip [Que88]. microfiche [Hen88d, WR92]. microfiches [Mor81]. Microphysics [Roc87, Buc85d]. Microscope [Tur69, Hac82a, Wil95]. Microscopic [Fri74b]. Microscopy [Hal68b]. microwaves [Cal90]. Micshigan [Rav72]. mid [RCRC98]. mid-Twentieth [RCRC98]. Middle [Ano82n, Hen83d, Hos67, Lef78, Sch80b, Mor83d, Tur83a, Whi85b, Gol91, Kni80b, Lef78, Lor75, McV75, MM10, Mol78b, Mol81, Mon78, NW82b, Nor71b, Nor77, Sch69b, Sch76, Sch80c, Sch80b, Sor83, Ste75, Wea78, Wea80]. Middleton [Chi68, Chi71, Duc84, Hop72, Spa65]. Middletown [Hei74a]. Middx [Hen84c]. midst [Nol95]. midwifery [Wil94b]. Mighty [Web68b]. migration [AS96]. Mikhail [Hal71]. Mikulas [Hal87b]. Milan [Out80b, Tur72b]. Milano [Mor89c]. mild [Buc85b]. Mildred [Fee77]. Milestones [Nor82]. Military [Edg87b, Hal88b, Smy89, NPYGDL86]. Milky [Sey77, Pau93]. Millburn [Bry78c]. Millenary [Sar81]. Miller [Bro75, Dol71a, GG77b, Jar84a, Mac72, Mar74, Ric89, Yox79, MH88, MS95, Sho96]. Mills [Cro70a, Far71a, Tee84]. Milnes [Fre81b]. Milton [Bro82b, Duf75, Ear87, Kni80a, Por84b, Slo85]. Milwaukee [Deb66, Hal69a, Hop73]. Mind [Cri71, Jac81, Sha78, Sha77a, Har87a, Kit92, Smi92, Sta89b, Tur94]. minded [Nol95]. miner [AHY90]. Mineral [AO77]. Mineralogical [Tor83]. Mineralogie [Sch84d]. Mineralogy [Bro89f, Lau87a, Sch84d]. Minerals [Sch69b]. Minerva [Ili95]. mines [AHY90]. mineur [AHY90]. Minime [McC80, McC83]. minister [Bru88a]. Minkowski [GG75c, Gra86b, SO85]. Minneapolis [Lar74]. Minnesota [Lar74]. M’Intosh [Und80]. Minutes [Can74, Hen84c]. Mirabilis [Ait73]. mirabilium [Lin86, OH85]. Miracle [Wai90]. Mirella [Smi82a]. mirror [Sir97]. misfortune [dCC09]. Mission [Irw83a, Pye93]. Missionary [Rap71]. Misunderstanding [IW96]. Mitarbeit [Wes70]. Mitchell [Sti89]. mitigation [MJ89]. Mitscherlich [Sch92b, Sch92b]. mixed [Seg94]. Mobility [Fig78]. model [Bro82a]. modèles [Dou84, Duc82, Sch84b, Sin91]. Models [Bar79, Men65, Pic87a, Zan86, Sin91]. moderado [Fre90]. moderate [Fre90]. Modern [BBBM90, Bod84, Bro84a, Buc63, Byn73, Can80c, Chi64, Cor96, Eyl68, FF66, Fra87a, Geo87, Gow71, GG71, GG74a, Hac76, Hal80a, Hal85b, Hen88f, Hoo87, Kee88, Ped85, Pni82, Ric88a, Stu86, Wes74, Wil89a, AH85, App93, Asp90, BA91, BR95, Bru87a, Bru88b, CDS90, Cro90a, Cro96, DeV82, Deb91, Eis83, Eis93, FJ93, Fri89, Gab83b, Gou95, Gow78, Had94, Hea87c, Hen82c, Hen83b, Huf93, Huf95, KT87, Kle94, LeG88, Lun94, Mis95, Mor83a, O^{+}90, Olm92, Poi04, Poo98, Rus86b, Wei83b, Wil95]. moderna [Olm92]. modernes [Poi04]. Modernist [Ros94]. modernization [Pri83]. Modes [Fie86, BB84]. Modest [Mos77b]. Modifications [Hen86b, MR82b]. Moewus [Sap90]. Mohammed [Sar81]. Moivre [Hah70]. Molecular [Rus88c, Wri94, Yox78, Olb82, Sap90, Sie91, Yox83a, Sin74]. molecules [CGBE95]. Moll [Rei68]. Molly [Smi85d]. Moments [Ros90]. Monas [Rat66b]. Monboddo [Mor73b]. monde [Bla93, Lov75]. mondo [Olm92]. money [Poi04]. Monge [Per73]. Monnet [Rap69]. Monod [Yox83a]. Monograph [Smi80b, Swi70]. Monographs [Bur80, Coo81, Nee69, Wil84d, Wil85a]. Monopoly [Pic80]. Monsters [Bro84c, PP82]. montagne [Ano93]. Montbard [BG^{+}92]. Montesquieu [Ano07]. Montgolfier [Car84, Gil83]. Montgomery [Eva89]. Montpellier [Ste62a]. Monument [War88b, Wri87]. Monumental [Fri74b]. Moody [Nor77]. moon [Hen91, Til76]. Moorcroft [Kni86c, Ald85]. Moore [For71a, Wil93]. Moorland [Bar75]. Moral [Jac81, Nye76]. Morality [Dol73a]. Morals [Jor80]. morbus [Law83]. Moreau [Bea85b]. Morgan [Nor80, Scr68b]. Morgenbesser [Dol73b]. Morienus [Win75]. Moritz [Mor91a]. Morley [Mar74]. Morphological [Tei70, Ric92b]. Morphology [Bow89d, Des89a, Gou84a, Hun82, Nyh95]. Morrell [Pat86, Sec85]. Morris [Kra86, Swi67]. Morrison [Jam87a]. Morrison-Low [Jam87a]. Morselli [Bro86b]. Morsink [Nut84]. mort [BG^{+}92]. Mortality [Coo81, Sha73b]. Morton [Dol73b]. Moscovici [Ait68]. Moscow [KBRS90]. Moseley [Sin76]. most [Cha83, Huf85]. mothers [HH97]. Motion [Hal65b, Mol73, Mol79, Sch75, Tei70, Whi71, Bar89a, Her65b, She91, Til76, Whi64, Rus64]. Motions [Gin74]. Motive [Smi87a, CF86]. Motor [Cla71]. Mott [Hen82a, Rud84]. motu [Sch75, BJB92]. motuum [Mol73]. Moulds [Dur77]. Mountain [Por89c, Ano93, Bry88]. Moussaieff [Por87b]. Mouton [Tur75, Wea80]. Mouvement [Ait68]. Move [Nay85, Pic80, Dra81]. Movement [Kei86, RH89, McG84, Nol94]. Moves [She70]. Movies [Sho87c]. Moving [Pac71]. moxa [Agr83b]. Moyal [Kni88a]. moyen [Gol83b]. Moyer [Sei85]. Mozans [Fee77]. MR [Wor79]. MRC [AB89]. Mrs [Ano64l]. MS [Rus67]. mu [Agr83a]. Much [Hen83d]. Muhammad [Sar81]. Mulhouse [Fox84]. Muller [Ste62a, Lin90]. Multhauf [Clo80, Dun69b]. Multiple [Cau89]. Multiplicatione [Sch85a, BL83]. Munby [Wal77]. München [CM90, Fig88, Rus79a]. Mundo [Rat66a, Rus77c, Str75, GB95]. Mundy [Din63a]. Munich [CM90, Mol75, Wes70]. Münster [Mol86]. Münsterische [Nor71a]. MUP [Law89]. Murchison [Pag76, Sta89a]. Murders [Har89b]. Murdin [Hun86]. Murdoch [Dun87]. Murray [Hew66]. Murthy [Wuj85]. muscular [EE95]. Muse [Aas87, FD85]. Museet [Bal83]. Museo [Tur72b]. Museum [All84b, Ben79, Bry76, CM90, Day85, Des88, Fis81a, Hac82b, Hen89g, Jam87a, Lar74, Nev80, Nut83, Rai91, Rya70, She63, Smi78c, Swi70, Tha83a, Tha83b, Tha84b, Vea64, Wil85a, Wil87d, Bur74a, Bur74b, Ban88, Des82c, Gun84, McC81, Win91]. Museums [Chi68, Hal68b, Tur68, Wal81b, But92, SP88, Won93]. Musgrave [Dol71b]. Music [Gou85, Ben79, Coh84, Cro90a, Ste94]. Musiktheoretiker [Mol75]. Musikwissenschaft [Mol75]. Musson [Mar73, Smi71]. my [Cla67, Sta87]. Myles [Wea84]. mysterious [Twy88]. Mysterium [Jar84b, KA81]. Mystiques [Tur73b]. Myth [Aas87, Lor75, Per89b, Bow88c, Dur82, FD85, GGR96]. mythes [GGR96]. mythical [BM82, Out84a]. Mythological [Cun78].

N [Ait89, Bro82a, Can84, Cro72, Dea83a, Dol71a, Dol74, Fox69a, Gli83b, GG70, Hen82c, Hen89b, Kee89, Kni83a, Koh75, Lev73, May66, Men74, Mos77b, Nic85, Smi83, Smi85a, Wal77]. N.

[Ait85a, Bro89b, Whi89b, Wil84c]. n.d [Cla67]. N.J [Ben82b, Bus83, Can77, Cha77a, Fie86, Gab75, Lor75, Rav73]. N.Y [For74a, McV75, Und75b]. Nachlass [GG77a]. Nachmansohn [Har81]. Nachträge [Bro89a, vLvHH^{+}88]. Nachum [Lor77]. Nafferton [Lay85]. Nagel [Dol73b, Mol86]. Nakayama [Rav70]. Nambor [GG74a]. Name [Hoc88a]. Nan [Agr83a, Mol84b]. Nan-fang [Agr83a]. Nancy [Cun83b, Cla86a, Cun89, Goo88a, Wei85b]. Nantics [Pac72]. Nantwich [Tee84]. Napier [Bry92]. Napoleon [Har67]. narcotism [SE91]. narrative [Nic88b, Söd86]. NASA [Mac88a, McC93, Smi89d]. Nashville [Olb70, Roc88b]. Nasr [Mol69]. Nathan [Agr83b, Bud84, Goo85b, Hac74, Hac77, Koh78, Goo85b, Ink88]. Nathaniel [Kee89, Liv87, Pep67]. Nation [Nic88b, Mis95, Söd86]. National [Ben89, Bur74a, Bur74b, Cro77, Dol73a, Duc84, Gin84, Gow69, Gua89, Hal85c, Hut80, McC80, McC83, Nev80, Por89d, Rya70, Sar81, Sch68a, Wil85a, Wuj85, Ano82l, Ano93, PT92, YK94, Bud84, Coo84a, Duc84, Gin84, Gli84, GG84b, Hal84d, Hen84d, Lil85, Mor84c, Mor84g, Nev84, Pic84a, Sch84a, Web84, Web88, Web98, Wil84b, Woo84]. Nationale [Hen84b, Sch84c]. nationalised [H^{+}82, Hen84d]. Nationalism [Cra92, Wei89]. Nationalsozialistischer [Har87b, Lil85]. Nations [Rap86a, Ros84b]. Natur [Gol91, Hei93, Mai88, MD90]. Natur- [Mai88]. Natura [Jon87, Ern91, Olm92, Por79a, Rav69, Sch81b, SSH85]. Natural [AH93, All77, All81, All84a, All84b, Asq80, Ban88, Bra84, Bro84d, Bro86e, Buc66, Buc68, Byn80, Des88, Dur77, Dur79a, Goo81a, Gow73, Gre92a, Hai77, Hal72a, Ham84, Hen82d, Hen86e, Hen88e, Hod70, Ilt73, Kni83c, Kni84a, Kni84b, Lev68, Mai88, McG71a, Mor81, Out84a, Out84b, Por84b, Rai91, Rid82, Rud84, Sec84, Smi73b, Smi76b, Tha83a, Tha84a, Tha84b, Tur73c, Vic88, Wal74, Wil82a, Yeo86, Bru87b, Cav96, CS94, Cla89, Clu88, Dem83, FJ93, Fin93, Fre94, FC96, Gay92, Gol91, Gow83, Gre92b, Gro89, Har82c, Har98a, Har82a, Har98b, Hat90, Hen84f, Hen83c, Hen84l, Her80, JSS96, Kal95a, Kus95, Mar92, Mer89, Mol68, Nor89, PR95, Pow96, Sch80a, Sha83b, SP88, Ste91b, Ste92, Won93, Yeo93, AH93, All77, All81, McV80, Mir94]. natural-philosophical [Gro89]. Naturales [Gro89]. Naturalis [Whi73]. Naturalism [Bro75, Bro79, Dur79b]. Naturalist [Bro86e, Bur75, Gru69, Mar89, Ste70, Dem83, Duf83, Gid82, Jon88]. Naturalists [Bro85e, Nic88b, Und66b, Reh83, Söd86]. naturalium [Mer78]. Nature [Bru86, Cit90, Con86, Dol71a, Dol73a, Dur81, Fox69b, Fox70, Fra87b, Gre92b, Hei72, Hop76a, Hop76b, Jac81, Jor80, Kni89d, LS90, Lin86, Mai88, MD90, McE79, Nye76, Nye86, Osb94, PW89b, Rus87a, Rus87b, Sch69a, Sch72, Sch80d, Sch85a, Smi72, Ste91a, Stu71a, Tur78b, Wes62, BL83, BDH96, Bro85c, BC98, Cro96, Cun83a, Deg91, Ern91, Fre94, Fur89, Gol91, Had94, Hak95, Hei93, Lin82, LN86, Mac88b, McG95, MR96, Mir89, Mit92, Olm92, Ped92a, Por84b, PS84a, PS84b, Sho93a, Sho93b, Sta84a, Win91, Woo84, Wyl89, You85, vR95, MR88, OH85]. Naturelle [Bur74a, Bur74b, Har86, Jor88, Por84a, Atr86, Gay92]. Naturhistoriska [Ber95]. Naturmagie [Gol91]. Naturmystik [Sch83b]. naturnal [Atr86]. Naturphilosophie [Ree85]. naturphilosophische [Gro89]. Naturwissenschaft [AH93, Gol91, AH93, Cap81b]. naturwissenschaftliche [Dau98]. naturwissenschaftlicher [LHB04, LS76, LHH91, vL76]. Natuurwetenschappen [Bry77]. Nautical [Sti70, For67a]. Naval [Man88]. navegar [Goo89b, Lan86]. Navigation [Arm70, Rav70, Sco67, Sti70, Ben87, Roc81a]. Navy [Wil87b, Hac84]. Nazi [Geu92]. Nazis [Pro88]. Near [Ear70, CHG91]. nearly [Adl87]. Nebraska [Kni83a]. Nebulae [Bar76a, Chr95]. Nebular [Bro79, Ogi75]. Neck [Tan82]. Nederlanden [Tur82]. née [Tor87]. needed [Lan87]. Needham [Agr83b, Cro72, Duc88, Dur77, Hal86c, Hal88b, Mol71, Olb71a, Pac72, Sho87a, Tyl75, NR86]. Neglect [Pin88, Fra86]. negotiating [BVA95]. Nehemiah [Le 90]. neige [Sch78c]. Neil [Can86]. nel [Sch73a]. Nell [Web68b, Sch81b]. nella [Olm92, Sch81b]. Nellie [Fre70]. Nelson [Chi65b, Cla66, Cri66, For66, Hac76, Pag66, Rus68, Sty68, Web66a, Wil65a, NN95]. Nemore [BJ91, Tho74a]. Neptune [Clo80, Fri91]. Neptunismus [Fri91]. Nernst [Mor74b]. Nersessian [Goo88a]. net [Arm64, McG68, Arm62, Arm64, Chi64, Chi65b, Chi65c, For68, Geo66, Hal66a, Hea64, Rat65a, Und64a, Wil65a, Wil65b, Wil69]. Netherlands [CT95, Pep71, Sme68b, Sne87]. networks [Pri83]. Neu [Deb66, Fis83b]. Neuburger [Jac85]. Neue [Mol75]. Neufeldt [Fis78]. Neuge [Bos72]. Neuge-bauer [Bos72]. Neugebauer [Mol78a]. Neumann [Ran74, Asp90, Pou93]. Neurasthenia [War89, Gos87]. Neuroscientific [Har89a, CJ87]. neutron [Jen89, Six87]. neuzeitlichen [Kle94]. Never [Rog83]. Nevil [How89]. Nevill [Hen82a]. Newberry [Got86a, Got86b, Got92]. newborn [Bak96]. Newcastle [Buc85b, Eag78, Fau89, Cam82]. Newcomen [Law65]. Newnes [Hen84c]. Newton [Ait67, Ait72, Ait73, Bra89a, Bro78a, Bro78b, Dea86b, For68, Gol88b, Gou84b, Gue84, Hal77b, Hal86a, Har79a, Has72, Hei76b, Hei77, Hen84c, Hen85c, Her80, Hom86b, Hun86, Jac89b, Kub70, Lau68, McG68, McG69, McG71b, Nev83, Nor68a, Nor69a, Nor70a, Nor73a, Pep72, Rog83, Rog88b, Rus79b, Sch81a, Sch84c, Sey77, Smi85a, Til75a, Til76, Tur78a, Whi72b, Whi73, Ano92, Bec75, Bec91a, Ber93b, Bur83, Can83, Dob91, F^{+}88, Fei90, Gan95, Gje86, Gue81, Hal92, Hal93, Her65b, Hol78, Hom68, McG70b, McG78, MT83, McG95, Mur85, NS84, New89, NJ91, Row90a, Rus76b, Sep88, Sha93, SC88, Vic87, Wes62, Wes79, Whi64, WHS92, dG01]. Newtonian [Hei71a, Hen84e, Hen88a, McG66, Sch86a, Bec82, Bla83, Eag77, For85b, Ilt73, Mer79, Nor73a, Ste92, Whi73, Wil87a]. Newtoniana [Har79a]. newtonieene [Sch86b]. newtonienne [Bla83]. Nicholas [Lef78, Geo87, Hen89c]. Nicholls [Por88]. Nicholson [Nay77]. Nicolaas [Rud84]. Nicolaidis [MdLD86]. Nicolas [Bos70, Tur74a, Ler85]. Nicolaus [Mol86, Sch69a]. Nicole [Lin86, Cro89b, Mol73, OH85]. Nicolin [Dol71a]. Nicolson [For71a, For82, Web71b, Yat72, Yox83a]. Niedersächsischen [Hom74]. Niels [Bad83, Aas90, Dor79, Sme87]. Nielsen [Dor79, Hen88b]. Night [Kel85a, Por89a, Sch88]. nihil [SLT88]. Nijhoff [Dou84, Gil84, Goo88a, Hac75, Han85a, Hen83c, Mil87, Mor73a, Pep71, Sch84b, Sch74a]. Nikolas [For87b]. Nimbus [Tur81]. Nine [Whi85a, Maf91, Viè83]. Nineteenth [Bro71a, Bro85a, Bro73b, Bro85e, Bro86e, Bru72, Bur74a, Byn88, Can77, Car69, Cha80b, Cri71, Dol74, Fie81, For72, Fri74a, GG74a, Hac74, Hac77, Hah73, Har89a, Jam87b, Kni69a, Kni72, Kni74, Lar74, Law87, Mac72, Mac79, Mar74, McC85b, Men74, Mor69a, Mor72a, Mor72b, Mor73a, Mor74a, Nee69, Olb72, Ora77, Out87, Pic76, Pic88a, Por89a, Por78a, Por85b, Rob80a, Rob87, Rob89a, Rud84, Rya70, Sei85, Sha78, Smi73a, Smi80a, Smi85c, Sne74, Tur72a, Wal81b, Wil87d, Can84, Har85a, Hen84f, JO81, Kni84a, Kni84b, Mor84d, Ogi86, Tur83c, Wei83a, Wil84a, Ali86, Asq80, Bar76b, Bec86, Bib74, Bro81, Bro76, Bro75, Bry78b, Buc77, Buc85d, Buc89b, Bur74b, Byn78, Byn80, Can74, Can75, Can91, Cap81b, Cha80b, Cha82, CJ87, Coo84b, Cor81, Cow74, Cro79]. Nineteenth [Dau75, De 85, DeK73, Dem83, Dol69, Dol71a, Dor75, Dor79, Dug77, Dur79a, Eng83, Far65, Far71b, Far76, Fie81, Fin93, Fis78, Fox74, Geo78b, Gol87a, Goo75, Goo81b, Gow73, Gow78, Gow69, GG75c, Gre82, Gre92b, Hal84b, Ham90, Har77, Har82b, Har87a, Har79b, Hei75, Hei73, Hei74a, Hei77, Her66, Hut80, Joh79, Jor79, Kni71c, Kni71b, Kni86a, Kni88d, Len82, Lev80, Mac78, Mac80, Mac79, Mar74, Mau77, Maz95a, Men65, Men70, Mor74b, Mor75b, Mor77b, Mor80, Moy83, Mus74, Nev80, Nor80, Ora73, Ora74, Ora76, Ora80, Pau76, Pic76, Pic80, Por78a, Por79a, Pri83, Pye76, Reh83, Ric82, Rob75, Rob80a, Roc84, Ros91, Rus68, Rus79a, Rus88b, Sch88, Sha78, Sha77a, SP88, Sho86b, Sin74, Sin76, Sin77, Sin78, Smi73b, Smi76a, Smi78a, Smi78c]. Nineteenth [Smi80a, Smi94a, Stu78, Tei78, Tur83b, Und79a, Und80, Wal81b, Web75, Web76, Wil68, Yeo86]. Nineteenth-Century [Bro71a, Bro85e, Cha80b, Fie81, Har89a, Mac72, Mac79, McC85b, Pic88a, Wal81b, Hen84f, Mor84d, Wil84a, CJ87, Coo84b, Eng83, Far65, Fin93, Har82b, Har87a, Maz95a, Men65, Pri83, Reh83, Ros91, Sho86b, Smi94a, Tur83b, Yeo86, De 85]. ninetenth [Tur83a]. Ninth [Car69, Scr64a]. Nisbet [Gow71, Sne74]. Nistri [Sch81b]. Nistri-Lischi [Sch81b]. NJ [Fie81, Fis81b, Gil88]. No [All77, All84a, Bal83, Ben84, Bos72, Bow86, Bro76, Bro83b, Bro84d, Bry77, Buc68, Can74, Car74, Cha77a, Dol69, Duc88, Far76, Fie81, Fig77a, Gab83b, Gli84, Gol79, Gol83b, Goo89b, Gow71, GG70, GG74a, GG89, Hac82b, Hal84a, Har89a, Hat87, Hen83b, Hen89c, Hoc88a, Hod89a, Hol71, Irw83a, Jor83b, Kel75, Kel78, Kni85, Kni88c, Kni89e, Lay74, Mah89, Mau77, Mau87, McC83, McC82, McV75, Mor83d, Nay75, Nic88b, Nut84, Olb82, Ora76, Osp74, Out80a, Phi67, Pic84a, Por78b, Por89d, Ree85, Rud89, Rus72b, Rus77a, Sch84b, Sch84c, Sch75, Sch81b, Sch78c, Sme69, Smi75a, Smi80b, Sne74, Str75, Tan82, Tei83b, Tho76a, Tur72b, Tur73d, Tur73b, Web78, Whi89b, Wil75, Wil85a, Wil87d, Win75, Woo80b, Woo80c, Woo89, Bow87a, Bow83a, Buc85b, Cas85, Duf81]. no [Fox68b, Nee69, Pac71, Vea64, Wor79, YK94]. Noah [Rap78]. Nobel [Goo85a, Har79b, Hea64, Rec86, Tre88, Cra84, CHU87, Cra92, Hea64]. Noble [Bus83, Jar83, Rus79b]. Noel [Fig78]. Noether [Gra87b, vdBLW85]. Nollet [Hom86a]. nombres [Bea91]. Nominalisme [Sch74a]. Nominators [Tre88, CHU87]. Nominees [Tre88, CHU87]. Non [Bar76c, Bow88c]. non-Darwinian [Bow88c]. Non-Progress [Bar76c]. Non-Progression [Bar76c]. Nonmathernatical [GG74a]. nonstandard [Dau95]. Noordhoff [Bry76, Bry78b, Bry78c, Hom80, Mac78, Rob80a]. Norbert [Asp88, Pic84a]. Nordic [KH95]. Nordström [Cro67]. Norma [Hal86d]. normal [Law83]. Norman [Bry83, Gin84, Hen84j, Kni89c, Mon78, Mor75c, Smi74b]. normative [Hat90]. North [Dor79, Fel89, Fra87c, Gra86a, Her95, Mol78b, Nad83, Por78b, Col87a]. North-Holland [Dor79, Fra87c, Gra86a]. Northeastern [Cro89b]. Northern [Hil71b, Kni83a]. Norton [Bro85b, Pep89]. Norwood [Fis81b]. Nos [Rat66b, Ano87r, CAMR87, CLR87]. nostro [Rat66a]. Notable [Wil69]. Note [Dav89, Far63, GW74, Rei68, Tee89, The78]. Notebaart [Tur75]. Notebook [Hen85c, MT83]. Notebooks [Des88, Nay80, BGH^{+}87]. Notes [Ano62f, Ano62g, Ano63j, Ano63k, Ano64j, Ano64k, Ano65h, Ano65i, Ano66i, Ano66j, Ano67g, Ano67h, Ano68h, Ano68i, Ano69j, Ano69k, Ano70h, Ano70i, Ano71i, Ano71j, Ano72g, Ano72h, Ano73g, Ano73h, Ano74j, Ano74k, Ano74l, Ano75l, Ano75m, Ano75n, Ano76i, Ano76j, Ano76k, Ano77n, Ano77o, Ano78j, Ano78k, Ano78l, Ano79j, Ano79k, Ano79l, Ano80j, Ano80k, Ano80l, Ano81l, Ano81m, Ano81j, Ano81k, Arm66, F.77, Mol89, Sch78a, Sch85a, Web68a, Wil90a, BL83, Bro84c, Bry83, CAMR87, CLR87, Cop92, FTG91, Hug89, Jar84a, McV75, MR89, Nor71b, Rus79b, Sch69a]. nothing [Hen83d, SLT88]. Notice [Fre81b]. Notices [Bar79, Ben79, Bry81, Fis81c, Gol79, Her80, Law80c, McV80, Pic79, Sch79b, Wea79, Woo80e]. Notion [Col81a, BBC88, BDH96, Hai80]. Notre [Dol70a, Kni81, Lef78, Nay80, Por84b]. Nottingham [For87a, HWW86]. Noury [Ear70]. nous [Gua89]. nouvelle [BN82, Sch84c, dGB^{+}89]. Nouvelles [Hal77a]. Nov [JO81]. novâ [Viè83]. Nova [Rat66a, Smi89e, Whi85a]. novae [Rav69]. Novel [Jor85a, Tyt82]. Novelties [Mos93]. novelty [Bli92]. November [Deb84, Hac77, Sar81, Yox79]. Noxon [Hei74b]. NS [Cap81b]. NS-Ideologie [Cap81b]. Nuclear [Pic87a, Tre79, Wea88, Aas90, Bau95, Wil84d, Wyn82, Zan86]. Nudos [Kni89c]. nuestras [ML89]. number [Buc85a, BJ91, Høy94, Kni83a, KI92, Sch90, Sch75]. numbered [Nun88]. Numbers [Cun88, Gra86b, Rus87b, Wea79, Bea91, Cru90, Por95, SO85, She91, Buc85a]. Numerical [Ver88]. numerous [Cal63, Chi64, Ste62c]. nuncius [GP92, SB09, GV89]. Nuova [Por79a, Sch79b]. Nursery [Ste62b, Bak96]. nursing [Bus83]. Nurture [Dol73a]. Nutrition [Rus77a, KC95]. NY [Und79b]. Nye [Sin74].

O [Bos72, Has72, Hen82b, LM74, Mol78a, Mos77b, Pum89]. O.S.B [Rat66a]. Oak [Sec83]. Obelisks [Pac71]. Obervatory [Maf91]. Obituaries [BC72, WM85]. Obituary [Ano64l, Ano65j, Ano67i, Ano67j, Bro85b, But64b, But65, Car78, Cha80a, Cha88a, Cro84a, Deb84, GD85, Hen86f, Her65a, LB73, Sme78b, Sme79, Sme84b, Sme87, Sme88, Tor80, Tor87, Und65, Und71]. Objection [Kre69]. Objectivity [BS77, Whi89b, Ash95, Het88, Meg94, Por95]. objects [FTG91]. Observation [AH85, Fra87a, Goo90]. Observations [Mor72a, Sha73b, Tor88, Nev80]. Observatoire [MdLD86]. Observatory [Bry78a, Cha77b, Dun86, For73, For80, Way87, Wri87]. Observed [Aus67, Lew85, Sto83]. observer [Nev80]. Observers [Lew85, Sto83]. Occasion [Sar81]. Occasional [Buc85b]. occasions [MJ89]. occidental [Gol83b]. Occidentales [ML89]. Occult [Cla89, Dea86a, Vic84, Fis83a, Mor91a]. occulte [GGC90]. Occultist [Wea80]. Ocean [Nun88, Men86]. oceanographic [Ben84, McC82]. oceanography [Ben84, McC81, Mil89, Wat73]. Ockham [Liv85]. O’Connor [Law89]. October [Ano65j, CM90, De 85, Hac74, Kel68, Mad88, Rat66b, Tor80, Wil87d]. Odense [Mer78, Osp74, Smi75a, Str73]. O’Donnell [O’H86]. odyssey [Dau95]. Oersted [Can80a]. Oeuvre [Cro79]. Oeuvres [Bos70, VB0 , LF97]. Öffentlichkeit [Dau98]. Office [All84b, Ben84, Hac88a, Hen83b, Kni83a, Kra86, She63, Tre88, Wil87b, Bur86, Man88]. Officers [Ano62h, Mar63]. Offices [Pat86]. official [Smy89]. officiels [CAMR87, CLR87]. O’Hanlon [Por85b]. O’Hara [Hun89a]. Ohio [Bro85a, Can80b, Hal84e, Kei86, Mol79, Whi85a]. Ohm [Tur77b]. oil [EK88]. Oilman [Fis81b]. O’Keefe [Pye76]. Oklahoma [Smi74b].

Oktober [CM90]. Olaf [Mer78]. Öland [Rau75]. Olaus [Eri94]. Olby [Web66b]. Old [Kuh68, Hai82]. Oldbourne [Dun69b, Jos67, Lau68]. Oldenburg [Hop73, Hop75, Ree85, Web66c, Woo80c, Hal65c, Hen88f]. Oldroyd [Bro82b, Rog88a]. Oleson [Gur78, Hal86b]. Oliver [Ant64, Hei73, Dra86, Row90b]. Olivier [Sch74a]. Olschki [Dur80a]. Olson [Bod84, Can77]. O’Malley [Grm69, Hol71, Rat65a, Whi77]. Omer [Din63a]. One [Mol74, Smi87c, Ait71, Bro89a, Hea86, Hea90, Pep89]. only [Bry78a, Hen88d, Mor83a]. Onnes [GG91]. Ont [Duc84]. Ontario [All84a, Gin84, Gou85, Sha83a, BB82]. ontogenèse [CM93, MC^{+}93]. ontogenesis [MC^{+}93]. Ontogeny [Slo80]. Oort [Dea83a]. Oosthoek [Tur70c, Tur73d]. OP [Col83]. Oparin [Ree80]. Open [Bee96, Bro82b, Duf75, Ear87, Kni80a, Mus74, Por84b, Rus77b, Blo84, PB82]. opened [HJF89]. opening [Nee78]. Opera [Bar73a]. operative [Rud63]. Opere [Hat87]. Opinion [Tur74b]. Opolka [Gra86b]. Oppenheim [Sho86a]. Oppenheimer [Hen81, Yox82b, Yor89]. Opper [Kni75]. Opposition [Per70]. Optical [Hal86a, Hal88a, Buc89b, NS84]. Opticks [McG70b, Hal93]. Optics [AS89, Bad62, Can83, Hak95, Lau68, Lin68, Nor71b, Smi85a, Til75a, Cro90a, Wel70]. Opus [Whi85a, Viè83]. Oral [Cla68]. Orange [Mor74a]. orbits [KH92]. orbs [Gra94]. Order [AO77, Bru86, Goo81a, Jac81, Ped85, PS84a, PS84b, And82, Des89b, Fox84, Ste91b]. Ordering [Fis83c]. orders [GG77a, Hal88a, Roc88b]. ordinary [Tro89]. ordination [Sme68a]. Orel [Bow85]. Oresme [Lin86, Mol73, OH85]. Orest [Gel76]. Orfila [Hue88]. Organic [Bro71b, Col88b, Kni89b, Roc88b, Bur87b, Cha91, TT86, Tra87]. Organicism [Tei78].

Organisation [Law87, Ora74]. Organization [Lay85, Wal81b, Coo84b, Frä90, HJF89, Mac82, Out82a]. organizational [McC93]. organized [CM90, Pic80]. Organological [Sho87b]. Orientation [Hoc86]. origenes [Lan89]. Origin [Bro85d, Car74, Cow74, Hod89b, LY66, PW89b, Rus88c, Web65b, Wei84, BP95, Bro84b, Gra85a, yL66, PB99, Wus84]. Original [For72, Hen89a, Hen89c, NR86, Rob89a, Yox83a]. Origines [Cal63]. origini [Sch80c]. Origins [Asq80, Boy70, Bry78a, Cha65, Cha78a, Duh91, Dun69b, Fra87c, Fri87, Goo68, Gra86a, Hac89, Har89a, Hen88f, Hog74, Hun88, Kni76, Kno85, Mar62b, Ora72, Rav71, Rob75, Rud83, Rud89, Web67b, Yeo88, Yox83a, Asp90, Bak96, Ben86, Cat88, Cha88b, CJ87, Dan90, Dic82, Eng84, Gra83a, Gro95, Hag92, Hal84c, Lig87, Lan89, Nol94, Pic87c, Pin83, Pum84, Ric92a, Ril85, Sap90, Sar87, TC91, Tra93, Tur77b, Wei82, Wis85, Woo84, Atr86, Web66b]. ornithology [Out84b]. oroscopo [GGC90]. Orrery [Mil73]. Orthodoxy [Gil84, Rap78, Hey83, Hun95]. ortu [Rav69]. Osborn [Rai91]. Osborne [Bur71]. Oscar [Wei91]. Oscillography [Hac88b, Phi87]. Osiris [Cro89a]. Osler [Sha83a, Und75b]. Ospovat [Bar75]. Ossolineum [Eas86, Mol81]. Osterbrock [Huf86]. Ostwald [Hol77]. Oswald [Rus88c]. Other [Ano64i, Ano65g, Sho86a, Ste67, Web78, Web68a, Ano62e, Ano63h, Ano63i, CS94, GH84, GH96, Gun84, Hal85a, Hol86, dLT96, Opp85, Ros79, Und64a]. Otilie [Rya70]. Otte [Can84]. Otto [Koh78, Kra91, Tei83a, Kra91, vGA94]. our [Ano87r, Goo83a, MJ89, ML89, Nol95, Whi88b]. Output [Fin71]. Outram [Cor81, Cro85]. ouverte [HJF89]. Ouvres [Cro88]. ovary [Cor97]. overseas [Hac82b, Hen82a, Pye93]. Overtures [Und66b]. Owen [Fra87a, Kni76, Sha78, Bro77, Des85, GT92, Ric87a]. Owsei [Phi69, Ros79, Wil75]. Oxford [Ait67, Bad71, Bar76b, Bod65, Bow85, Boy70, Bra89a, Bro67, Bro69, Bro72b, Bro73b, Bro83a, Bru72, Buc77, Byn74, Car70, Chi65c, Chi68, Cla70, Cla71, Cri71, Dav71, Din64, Dol70b, Dol71a, Duc88, Ear70, Edw78, Fal89, For71b, Fox69b, Fox70, Fre74, Fre81b, Fre82, Fri74a, Gab83c, Gab88, Geo66, Goo83a, Goo85a, Got66, Gow69, GG87, Hac82a, Had63, Hah73, Hal71, Har76, Har83a, Har82d, Hei71a, Hei74b, Hen89a, Kel70, Kel84, Kni67, Kni71c, Kub70, Kuh68, Liv86, Loh75, Lor73, Mac89, Mah89, Mar74, McG65, McG68, McG71a, Mea71, Mol69, Mol78b, Mor73a, Mor73b, Mor84d, Mos77a, Mos77b, Nee69, Nev78, Nor68b, Out85, Pep86, Phi69, Por89a, Por89c, Ran74, Rav68, Rav70, Red84, Rob89a, Rud69, Rud84, Rup85, Rus89a, Sch69b, Sch78a, Sch85a, Sha73b]. Oxford [Sin84, Smi75b, Smi89c, Spa65, Ste75, Swa86, Swi71, Tei83a, Tha74, Tur68, Tur69, Tur70b, Und66a, Vea64, Was69, Web65a, Web65b, Web66b, Web67a, Web68a, Web68b, Web69a, Web70, Whi70a, Whi68b, Wig78, Wil65b, Wil68, Wil84e, Wil84c, Wil85b, Woo67, You66, You83, Yox83b, Cru79, Eng83, PB^{+}89, Scr64b]. Oxfordshire [Bee89]. Oxygen [Fis81b]. Oye [Phi67].

P [Ant64, Blo79, Bur74a, Cla71, Clo80, Cro72, Dun69b, Dur81, Fau89, Fis74, Fox71b, Hac82b, Hen82c, Hen82b, Hen89b, Hig65, Hil71b, Jam86, Jar84a, Kel70, Kra86, McC80, McC83, MC81, Moo85, Mor89a, Mül87, Rav67, Rus72a, Sch83a, Sha83a, Sme69, Tur70b, Und66a, Wea80, Whi84, For87a, Kni89e]. P. [Asp88, DD95, Jam87b, Wil87d]. P.L.S [Wal88]. Pacey [Kel76]. Pacific [Kni89d, MR88, Tee88]. Pacifica [Smi89e]. Pacini [Sch81b]. Padova [Sch73a]. padovano [Sch80c]. Padua [Mol72b, Sch80c]. Page [Dav81]. Pagel [Deb84, Hen84k, Whi68c, Wig74]. pages [Cri66, Mac88a, Tur74a, Mac88a]. Paget [Chi65b, Rob89b]. Pairing [Cam85]. Pais [Hen89a, Red84, LN95]. Pakistan [Sar81]. Palaces [Sch78b]. palaeontological [Cle83, Tha84b]. Palaeontology [Des82b, Rap75, Rup88, Buf87, Des84, Tha84a]. Palais [Tan82]. Paléontologie [Lau87b, Osb89a]. Paleontology [Por78a, Rai91, Lau87b]. Pallister [Bro84c]. Palm [Hal84a]. Palmer [Rus66]. Palter [Ait73]. pamphlets [RT95, RD95]. Pan [Nor85a]. Pancaldi [Por79a]. Paneth [Sme66]. Pangenesis [Olb63]. Paola [Bra87]. Paolo [Sch79b, Sch80c, Hat87, Rap86a]. Paper [Arm64, Hal86c, NhTWL87, Rob89a, Smi66, Hen84f, Hen89c, Hen89d, Smi82a, Wei85a]. Paperback [Fur85, Hal68c, Hal85b, Hen89a, Ink87, Osb85, Sha78, Ait86, Bow85, Bus83, Byn78, Byn80, Col83, Col88a, Cro89a, Dea83a, Dun87, Dur81, Eme87, Fel89, Fig77b, For71b, For87b, Fox82, Gab83c, Gab88, Gau83, Gel76, Gol88b, Goo81a, Got71, Gou82, Hal84e, Hal84d, Hal87b, Hen82b, Hen84a, Hen84b, Hen88d, Hen88c, Irw82, Jac89a, Jen89, Jor83b, Jor87, Kni81, Kni86b, Koh78, Lar74, Law83, Law85, Lin74, Mea71, Mol84c, Mor77c, Mor89c, Nay80, Nor82, Nor85a, Nut84, Out85, Pac71, Pic84a, Por87c, Por89e, Pum87, Pye87, Rav72, Rav73, Rav84, Rid83, Rup85, Rus87a, Rus77b, Rus87b, Sch84a, Sch86b, Sch78a, Sch80d, Sec83, Sec88, Sha81, Sho88, Smi73a, Smi89e, Smi89c, Swi77, Tha83a, Tho89, Wea79, Web84, Wes74, Wor88, You83, Fie83]. Paperbacks [Bro83a]. Papers [ATKL08, Ano62i, Ano62j, Ano63m, Ano63l, Ano64m, Ano64o, Ano65k, Bal83, Bar77, Buc85b, Cow68, Cro70a, Cro72, Far71a, Goo85b, Hac74, Hac77, Hal80a, Hal86a, Hei76b, Hen88d, Hen89g, Kra86, Mor69a, Nor68a, Nor69a, Nor70a, Nor73a, Nor77, Pep72, Phi69, Rei87, Tur70c, Tur73d, Wil87d, Woo80b, AL87, BH87, BKH90, Dav99, De 85, EJ^{+}07, ER93, EJR95, GG91, Gou84b, Har02, Her80, JO81, Mad88, Nev84, NS84, NJ91, Pep89, SCS^{+}87, Sta87, Yox83a]. Pappus [Mol88, Jon86]. Paracelsians [Jos67, Deb91]. Paracelsus [Gol91, Jev64a, Jev64b, PTS^{+}96]. Paradigmatic [Fis81b]. Paradigms [Kni81, Pet94, Ste90b]. Paradis [Mor83c, Por87a]. Paradise [Whi90, Fin93]. Paranormal [McV80, GG85, Ing84]. Parapsychology [Gol83b, Pin83, Sch83b, Sha83a]. parascience [GG85, Ing84]. parchments [Lev90]. Pardies [Str73]. Paré [Bro84c]. Paris [Ait68, Ano82l, BG^{+}92, Ben89, Bos70, Bur74a, Bur74b, Cal63, CM90, Cro66b, Cro68b, Cro79, Cro88, Dug77, Ege89, For74a, Fox87b, Gab83b, GG70, GG75b, Gua89, Hai80, Hal68c, Hal72b, Hal77a, Hal85c, Har86, Hen84b, Hil71a, HS88, Hut80, Jen89, Jor83a, Jor83b, McC80, McC83, MdLD86, Olb72, Osb89a, Osb89b, Out80a, Out82a, Out83, Out88a, Out88b, Por89e, Sch84c, Sch86b, Sch72, Sch78c, Sho89, Sme69, Sme84a, Str75, Stu87, Tur72a, Tur73b, Tur75, Tur88, Wea80, Whi75, Wil77, BG^{+}92, CM90]. Parisian [Str87]. Park [EJR94, EJR95]. Parke [All77, Bow74, Car67a, Law80c, Wal77]. Parkes [Rob92]. paroxysms [Sha93]. Parry [Fig78, Pum88]. Part [Car66, Dol71a, Eag78, Fau89, Hal86c, Hal88b, Hen88c, Hil89b, Lev80, MR87a, MR87b, Mol84b, Nic85, Pac72, Sho87a, Tha74, Tyl75, WW93, Hei74a, Kel68, Ric92a, Ste70, Wor79, Yox82b, Hop76a, Hop76b, NPYGDL86, WW86]. Partes [Mol84b]. Partial [Fra87c, Gra86a, Eng84]. Particle [BH83, Hen86b, BDH89, Cas85, Mac85b, Pic84b, Whe83]. Particles [Dun88]. particules [Cas85]. Partington [Ano65j, But66]. Partners [Mor71].

Partridge [Buc68]. parts [Hen86b, Mac88a]. Pasadena [GG74a]. Pascal [Hal65a, Mah89, Ran74, Edw87, Ege89, Loe87]. Passions [Sho89, Bro83b, Sha93]. Past [Bow85, Gab88, Hen88a, Out85, Bow89e, MJ89, Wil87a]. Pasteur [Gei95]. pasteurization [Lat88]. patent [Mac88c]. Paterson [PCA91]. Pathobiology [Lar74]. pathological [Law83]. Pathology [Por87c, Kri83]. paths [GG92]. Patricia [Gel76, Hal66a]. Patrick [Bro86e, Dol73b, For71a, Und79a, Dem83]. Patron [Mac89, Alt87]. Patronage [Goo83b, Rob80a, Sha74, Doe96]. Patrons [Mac72, Wei77]. Pattern [Dur77]. Patterson [Far71b, Han85a]. Pattison [Pel85a]. Paul [Bow87a, Bro86e, Bro87, Cro87a, Dol83, Duf83, Fis83a, For74a, For87b, Fox74, Goo75, Hal68c, Hal68d, Har84b, Hea87b, Hen89b, Hen85b, Joh78, Kel71, Kel78, Law88, Lie86, McC80, McC83, Mea84, Mor75b, Mor77b, Mor77c, Nev80, Out83, Out84b, Phi67, Ric88a, Rog88b, Sha87, Smi85b, Tur73a, Tur83c, Woo89, Bäu84, Des88, Ink81, KW87]. Paul-Henri [For74a, Hal68c]. Pauli [Che85, Hen86a, Hen84i, HP85, PHvMW93]. Payne [Hal85a, GH84, GH96]. Payne-Gaposchkin [Hal85a, GH84, GH96]. pays [Ano90]. Peace [GG92, Hod92, Fis98]. peacock [Cro91a]. Pearce [Hei73, Kni65]. Pecham [McV75, Nor71b]. Pédagogique [Gua89]. Pedersen [Mer78]. Peierls [Chi64]. Peiling [Wig78]. Peirce [Des89b, GG77b, GG84a, Des89b]. pencil [Pet89]. Pendule [For69b]. Penguin [Rus72a]. Penicillin [Swa83, Swa86, Wai90, Wil84f]. Pennsylvania [Bad71, Bow89c, Bro67, Cro89a, Far69, Mac80, Woo89]. Pensée [GG70, Bec91b]. Pensées [Hal77a]. Pentagon [NOF96]. Penury [Coo89a, Goo88b]. People [Hen81, Rob75, Bea85b, HS83, Mor84a, Roc88a]. Pera [Ner89]. Perception [Can80b, Fig77b, McG65, Hat90]. Percussion [Nay75]. Percy [Wal90]. père [CMDM82]. Peregrinus [Hac82b, Hun89a]. perfectly [BS91]. performance [FG93]. Pergamon [Boy70, Buc69, Cla70, Dau75, For71b, Fox69b, Fri74a, Goo85a, Kni67, Kuh68, Mar74, Mea71, Rav68, Tur69, Web66b, Whi68b, Wil68]. Périgueux [Jor83a]. perihelion [Ros82, Wil84c]. Period [Ora80, Sta90, Stu86]. Periodic [Sme70, BV86]. Periodical [Dur80a, LP04]. Periodicals [Ano62i, Ano62j, Ano63m, Ano63l, Ano64m, Ano64o, Ano65k, Cha65, Cha78a, Che78]. periods [Ano82n, Ano82m, Hen84g, NW82b, NW82a]. periphery [Lin93]. Permutations [Kni89b, Bur87b]. Perrin [Sin74]. Persehouse [Sme84b]. Persian [Wuj85]. person [Dea83a]. Personal [Nun88, CGBE95, Dau95, Men86, Yox83a]. personalities [Fis83b]. Personality [Rav72]. Personnel [Mor77b]. Persons [Wal77]. Perspectiva [Eas86, McV75, Mol81, Nor71b]. Perspectivae [Eas86, Mol81]. perspectivas [OE^{+}90]. Perspective [Cro83, Mon87, Pic88b, Sha79, Sin74, Bal91, Kil84a, War86, Whi88b]. Perspectives [Edg87b, McG70a, Nay80, Por87a, Smi74b, Wil89a, AB89, Bow83a, HE^{+}82, KH95, KT87, Mal05, Mor83c, OE^{+}90, Ste90b, GH93]. Persuading [PS91]. persuasion [Lun94]. Persuasive [PLP92]. Pertaining [Web65a]. peste [Sch81b]. Pestre [Stu87]. Peter [Bad83, Bro89e, Can80b, Cow74, Des88, Eag78, Hac82b, Hal86b, Har79a, Hen84f, Hun89a, Mol79, Por78a, Sha73a, Ste62c, Wil89a, Yox82b, Bro86d, Fra87a, Hen84f, Hen89d, Hen86e, Hoc88b, Law85, Mac89, Nor85b, Wal89]. Petersburg [Sch85b]. Petersen [Mos77b]. Petit [Fox68a].

Petrarch [Kel78]. Petroleum [FB82, Lie85]. Petrus [Sme68b]. Petty [Sha73b]. peu [vPvP90]. Pflanzenchemie [Rus79a]. Ph.D. [FB75, Old77]. Phaenomena [EBT96]. Phaidon [Goo83a]. Pharmaceutical [Deb66, Mor89a, Lie87, Swa88, Wea90]. Pharmacology [Ear70]. Pharmacy [Hei88, Mor91b]. Pharmazie [Jac89a, Hei88]. Pharmaziegeschichtlichen [Ree85]. Phases [Smi80a]. Phénomènes [Sch86b, Bla83]. phenomenon [Bla83]. Philadelphia [Ait86, Bow89c, Cal77, Col88a, Cro89a, Dur82, Ear70, Ear87, Far69, Gid82, Mac80, Mol84b, Mor84b, Mor83d, Olb82, Sch80b, Ste70, Tur83a, Whi88a, Woo83, Woo89]. philanthropy [Aas90]. Philip [Coo89a, Joh79, Kni89d, Und66b, Was69, Yox82a, Bro85e, Goo83b, Goo88b, Kus95]. Philipp [Kni88c, May67]. Philistines [Und64a]. Phillip [Por84b, Por88]. Phillips [Hac88b, Han76, Woo80e]. Philoponus [Mol72a]. Philosopher [Bra84, Cal77, Gow73, Rap71, Fis91a, Fre92, Gow83]. philosopher-scientist [Fre92]. Philosophers [Koh78, Mor74a, Osb85, Rus76a, Sch81a, Sha87, FJ93, KRS83, Wyl89]. Philosophia [Rat66a]. Philosophiae [Whi73]. Philosophic [Cro87b]. Philosophical [Ait71, Ait86, Bro85e, Dol71a, Dun87, Gid82, Goo88a, Hei74b, Hei76a, Hen85c, Kel71, Kel68, Mol84b, Mor84b, Mor74a, Mor83d, Olb82, Sch80b, SB92, Ste67, Ste70, Tur83a, Wil89a, Woo83, Gro89, Hei93, KT87, Lin92, May66, MT83, Mor82b, Ner87, Nut84, Reh83, Tei83b, Ver85, vHC95, Dav99, Eme79, Eme81, Eme85, Eme88]. Philosophicall [Gra81]. Philosophie [HRP90, Mol72a, Sch84c, Sch74a, Ano92, Hen83c]. Philosophique [Cro79]. philosophische [Hei93]. Philosophy [Bad71, Blo82b, Bro73b, Buc63, Buc66, Buc68, Can77, Can80b, Cha77a, Cla79, Dol71a, Dol71b, Dol73b, Dur77, Dur79a, For67b, Got67, Gou85, Gow71, GG79b, Hai77, Hal72a, Har69, Hei72, Hen82d, Hen87c, Hen88e, Hut89, Jar84a, Joh81, Kil84b, Kni81, Kni85, KK85, Law80b, Lin74, Löw90, Lug79, Mai88, May67, McG69, McG71a, Mol79, Mol84c, Nor69b, Nor71a, Nor77, Phi67, Pow82, Rav67, Rog88a, RH88, Rus88a, Sch85a, Sha89, Sha81, Smi72, Smi76c, Swi67, Tur73c, Ver89b, Vic88, Web69b, Wes62, Wil75, Ait85a, Ano92, BL83, BA91, Bar83, Bro95a, Bru87b, Cav96, Cla82, Cla89, Clu88, Dur80a, FC96, Fur89, GA98, Gra87a, Har82c, Har84a, Har98a, Har82a, Hen84f, HRP90, Her80, Hom83b, Ilt73, Jar83, JK84, Kal95a, Kus95, Mar92, Mol68, Mor91a, Nye93, O’M89]. philosophy [Old86, PW89a, SSK88, Smi76b, Ste92, Tur90, Wil82a, Wil95, ZH06, McG70b, Bar81, Blo79, Dol70a, Dol71b, Dol74, GG81, Hei74b, Hol85a, Kni81, KAZ92, Lin74, May69, McG65]. Philosphique [Bos70]. Phlogiston [Cro69, Fis81b, Per70]. Photo [Lev65]. Photo-Keratoscope [Lev65]. Photographing [FL88]. photography [Sch92a]. Phrenological [Sho87b]. Phrenology [Coo89c, Law87, Sha77a, Coo84b]. phylogenèse [CM93, MC^{+}93]. Phylogeny [Slo80, MC^{+}93]. Physical [Bad71, Bad62, Bra88, Cro71, Dol70a, Duh91, Fra87a, Hei71b, Hen89a, Kni71b, MD90, Men65, Mol78a, Mol88, Mor77b, Nay80, Row90b, Web75, AH85, Gow78, Jam84, Pai86, Ste87, vH97, Bro71a, Din63a, For71b, Mea71]. Physician [Mar89, Web70, BaJANJ84, Jon88, Rud88a, Sta84b, Tal62]. Physicians [Pic76, Wei85b, Woo67, Fri81]. Physicien [Hom86a, Str73, BaJANJ84]. Physiciens [Stu87, Pes84]. Physicist [Buc77, Buc73b, Hen89b, Kel85a, Sin76, Can93b, KBRS90, KW87]. Physicists [Gre88, Jam87b, Kei89, Roc90, Gin88, Har85a, Wil87f]. Physics [Arm69, Bec91a, Blo84, Bro69, Bro76, BH83, Buc89a, Can77, Can80c, Cas82, Cha80c, Chi65a, Dor80, For69a, For71b, Fox69b, Goo82, Got71, Hac75, Har81, Hei72, Hei77, Hen82c, Hen82b, Hen84f, Hen86b, Her66, Hom86b, HM66, Jam87b, Kel83, Kil84b, Kil84a, Kni89c, Kra84a, Lay74, McC69, Mor77b, Ole91, Rav69, Red84, Sei85, Sin81, Sin84, Smi82a, Stu78, Stu85, Tre88, Wal69, Whi72b, Whi68b, Wil84c, Wil89a, Yox82b, Aas90, BDH89, Bru88c, BS79, Cas85, CHU87, Gar92, Gol70, Har82b, Har82c, Har85a, HW81, Hen82a, Hen83b, Hen84h, Hom83a, Jar83, KT87, KH86, Lin93, Mir89, Moy83, Pic84b, Pop82, Pow82, Roc90, Sha93, WR92, Wil87g, For71a, Fox70, Pac70, Pes84, Stu71a, Swi71, War90]. physicus [SSH85]. Physik [For67b]. Physiognomy [Jor85a, Tyt82]. Physiologia [McG68]. physiologic [Dou84, Sch84b]. Physiological [Men65, Stu88, Tur77b]. physiologie [Duc82, Grm90]. Physiologist [Tei83a]. Physiology [Bos83, Fig77a, Jac81, Jac85, Jor83b, Law89, Mor80, Tei70, Grm90, Kre90, NC81, O’C88, Tei83b]. Physique [Por84a, Sme69, Stu87, Cas85, Pes84]. physischen [MD90]. Pi [Kni88c]. Piaget [Cha88b]. Picard [Ben89, Pic87c]. Pickens [Olb70]. Pickering [Hen86b, Twe89]. Pickstone [Nev84, Web87]. Picolet [Ben89]. pictorial [Rud92]. picture [PB^{+}89]. Pictures [Blu79]. Pierre [Bos70, Bur74a, Gab75, GG77b, Hal77a, Hen88e, Ilt77, Mil87, Sch74a, Sho89, Bru87b, Jak84, Jor83a, Sho89, Smo94, Stu86]. Pierre-Jean-Baptiste [Stu86]. Pierre-P [Bur74a]. Pietro [Out80b, Bow89b, Bro89c, She85]. Pinch [Dol83, Fra87b]. Piñero [Goo89b, Kni88c]. Ping [Agr83b, Hal88b, Mol71]. Ping-Yü [Agr83b, Hal88b, Mol71]. Pingree [Lor73]. Pioneer [Fri83, Huf86, Kni86c, Rus85, Tur68, Und80, Wil84b, Ald85, Ost84, Pen89]. Pioneers [Har81]. Pions [BDH89]. Pipping [Cha80c]. Pisa [Sch81b]. pitfalls [DPP88]. Pitt [Nay80, Ner89]. Pittsburgh [Bro84d]. Pl [Nut83, Ait68, Gre63]. Place [Cha77b, Rus87a, Hom83b, Jam89, Kni85, McG78, You85]. Places [Nut84, Olm92]. Plague [Sch71, Und64a]. plains [Kni83a]. Planck [Hen86b, San87, Sin78, Hei86, MR82f, MR82g]. Planes [Hum67]. Planetary [Gin74, Rus64, TW95, Sch65, Whi64]. Planets [Gra94, Lev90, Wil84a]. Plant [BW90, Cit90, Dea83b, Kri83]. Plants [Ste67, Ste91b]. Plastic [Rus85, Fri83]. Plate [Bry77, Fre70]. Plates [Cla70, Cro67, Dun69a, Fox68b, Ged71, Hew66, Kni69a, Pag69, Phi67, Poy70, Rat69, Sti70, Web66a, Web67a, Bro78b, Bry77, Bry78a, Bry78b, Cal63, Car70, Cha80c, Chi64, Chi65b, Cri66, Cro66b, Cro68b, Cro70b, Cro79, Dol70b, Dun69b, Ede65, Edw78, FF66, Fig88, For66, For69b, Hal68d, Han70, Law65, Law80c, Mol71, Mol78b, Mol81, Pag66, Pep71, Rat65a, Rud70, Rus68, Rya70, Sch71, Sch78b, Sme71, Ste62a, Swi70, Tur74a, Web66b, Web68a, Whi68a, Whi68c, Wil68, Woo67, Por78a]. Plato [Sti84]. Platter [Ste62a]. Play [Hah70]. Playfair [Fis81a, Mor72a]. pleasure [CL96]. Plenitude [Yeo86]. plenty [BR95]. Plurality [Dic82, Pum84, Sch87, Cro86]. plus [Bro87, For87a, Ste81]. pneuma [Fre95, Osl91]. Pneumatic [Cro83]. Pneumatics [Kel67]. Pocock [Bow89a]. poder [Hue88]. Poems [Hüb90]. Poetics [Sch78b]. Pöggeler [Dol71a]. Poincaré [Ser88, Gol70, Gra85c]. Point [Bro76, Bar89a]. points [Ben75]. Poisson [Hom83a]. polar [Ber95]. Polarity [Got67]. Polemic [Ano04]. Polemics [Jac89b, Sep88]. policies [Mar10]. policy [Coo89b, Hen89f, Irw83b, Lil85, Wri94]. Polish [Löw90, Mol81, Tur74a]. polite [Sut95]. politica [Sch81b]. Political [Arm83, Hoc88b, Rud74, Jor85b, Kuz87, Law85, Web98, Wei92]. politicians [H^{+}82, Hen84d]. Politics [For87b, For87c, Jon75b, Kei84, Mac78, Out80d, Pin87a, Sha74, Yox81, Bru88a, Bru95, Cat88, Coo89b, Des89a, Edw96, GR88, Hom83a, Lev84, Man88, MF86, PW89a, Ric91, SY86, Smi89d, Wil91a, Wri94, vG92]. polity [Hai82]. Polyhedra [Blo78, Wor79]. Polymath [Hen89b, Kil87]. Polytechnic [Buc85b]. Polytechnical [Lan87]. Polytechnique [Out88a, Out88b, Sme84a, Lan87, GG86a, GG86b, FD87]. Pontifical [Col83, Lin86, Mol85]. Poole [Fox82]. Poperinge [Tur82]. popes [Maf91]. Popper [Blo84, GG79b]. Poppi [Sch73a]. Popular [Law87, Mac79, Woo80a, Yat72, Coo84b, Ram88, vHC95]. Popularization [Lay65, Dau98]. Population [Hog74, Tre88, CHU87, Cra92, MJ89, Cro91b]. Porfir’evich [Bro89b, FS88]. porphyry [But96]. Porter [Bur80, Col76, Dur83, Gol87b, Hal87b, Hod89a, Kni71b, Rap79, Tha83a, Tha85]. portfolio [Rud80]. Portrait [Hal71, Hal84e, Kub70, Moo85, Mor72b, O’H86, Ada89, Cro84b, FS91, HK83, Ken83, O’D83, Smi85d, Ste62b]. portraits [Bry83, Sme70]. positivism [Lan91]. positivismo [Lan91]. possession [SB92]. post [Bro87, Cro85, Fra78a, Gay92, Out84c]. post-Darwin [Gay92]. post-revolutionary [Cro85, Fra78a, Out84c]. postage [Bro87]. Posterior [Sch78a, Wal92]. posters [PLP92]. Postface [Bur74a]. Posthumous [Hal69a]. Postlewait [Mor83c, Por87a]. Postscript [PB82, Mer93, PB83]. postwar [Nak91, YK94]. potions [Lev90]. Poulett [Rud74]. Poulshot [Hen84c]. Pour [Jor88, Atr86, BG^{+}92]. Powell [Bro89c, Cor88b]. Power [Buc71, Coo89a, Cro87a, Goo88b, Hen81, Hil86b, Hil94, HS86b, Smi87a, Web69b, Wil84e, Bau95, Car67b, CF86, FL88, Hue88, Hun85, Kni92a, Lin84, Luc90, Lun94, Mac83, Sha84]. Powers [Hei71a, Kil84b, Cla89]. Pozzo [FPLJ89]. PP [Fig77b, Har75, Hen81, Lin74, Aas87, Agr83a, Agr83b, Ait67, Ait68, Ait71, Ait72, Ait73, Ait85a, Ait86, Ait89, All77, All79, All81, All82, All84a, All84b, Ant63, Ant64, Arm62, Arm64, Arm66, Arm67, Arm69, Arm70, Asp88, Asq80, Aus86, Bad71, Bad83, Bad89, Bal83, Bal84, Bar77, Bar79, Bar81, Bar83, Bar73a, Bar76a, Bar76b, Bea85b, Ben79, Ben84, Ben89, Ben82b, Bib74, Blo79, Blo82b, Blu79, Bod65, Bod84, Bos70, Bos72, Bow74, Bow87a, Bow89a, Bow83a, Bow85, Bow86, Bow87b, Bow87c, Bow88a, Bow89b, Bow89c, Boy70, Bra88, Bra84, Bra87, Bro67, Bro69, Bro71a, Bro73a, Bro82a, Bro84a, Bro85a, Bro86a, Bro86b, Bro89a, Bro89b, Bro80b, Bro72a, Bro76, Bro89c, Bro72b, Bro73b, Bro75, Bro82b, Bro78a, Bro78b, Bro86c, Bro89e, Bro83b, Bro83a, Bro84c, Bro84d]. Pp [BH84b, Bro85d, Bro85e, Bro86d, Bro86e, Bro87, Bro89f, Bru72, Bru86, Bry74, Bry76, Bry77, Bry78a, Bry78b, Bry78c, Bry81, Bry83, Buc77, Buc69, Buc71, Buc73a, Buc75, Buc85a, Buc85b, Buc66, Buc89a, Buc73b, Bud84, Bur80, Bur74a, Bur74b, Bur71, Bus85, Bus74, Byn72, Byn73, Byn74, Byn78, Byn80, Byn87, Byn88, Cah88, Cal63, Cal77, Can74, Can75, Can77, Can80b, Can80c, Can84, Can86, Cap81b, Car70, Car84, Car74, Cas82, Cas85, Cha65, Cha78b, Cha78a, Cha77a, Cha80b, Cha77b, Cha80c, Che78, Che85, Chi64, Chi65b, Chi65c, Chi65a, Chi68, Chi71, Cla86a, Cla66, Cla67, Cla68, Cla70, Cla71, Cla86b, Clo80, Coc75, Col81a, Col83, Col76, Col87b, Col88a, Col81b, Con86, Coo81, Coo83, Coo89a, Coo84a, Cor81, Cow74, Cow68, Cri66, Cri71, Cro66b]. Pp [Cro67, Cro68a, Cro68b, Cro69, Cro70a, Cro70b, Cro71, Cro79, Cro87a, Cro88, Cro89a, Cro89b, Cro72, Cun78, Cun83a, Cun83b, Cun88, Cun89, Dau75, Dav71, Dea83a, Dea83b, Dea86a, Dea86b, Dea86c, Deb66, Des88, Din63a, Din64, Dol69, Dol70a, Dol70b, Dol71a, Dol71b, Dol73a, Dol73b, Dol74, Dol83, Dor75, Dor79, Dor80, Dou84, Dra86, Dub72, Duc84, Duc88, Duf81, Duf83, Duf75, Dug77, Dun69a, Dun69b, Dun86, Dur77, Dur79a, Dur81, Dur82, Dur83, Dyb78, Eag78, Ear70, Ear87, Eas86, Ede65, Edg84, Edg87a, Edg87c, Edg87b, Edw78, Eme87, Eva89, Eyl68, Fal89, FF66, Far67, Far68a, Far69, Far71a, Far71b, Far76, Far77, Fau89, Fee77, Fel89, Fie81, Fie86, Fie89, Fig88, Fig77a, Fig78, Fin68, Fis74, Fis78, Fis81a, Fis81b, Fis81c, Fis83b, Fis83a, Fis83c]. Pp [For66, For67b, For68, For69a, For69b, For71b, For71a, For72, For74a, For80, For85a, For87a, For82, For87b, For87c, Fox68b, Fox69a, Fox69b, Fox70, Fox71a, Fox71b, Fox74, Fox79, Fox82, Fox87b, Fra87a, Fra87b, Fra87c, Fre70, Fre74, Fre78, Fre82, Fre86, Fri74a, Fri87, Fur85, Gab83b, Gab83a, Gab75, Gab83c, Gab88, Gau83, Ged71, Gel76, Geo66, Geo78b, Geo87, Gid82, Gil88, Gil84, Gin74, Gin84, Gli83b, Gli84, Gol79, Gol83b, Gol87b, Gol88b, Goo75, Goo81b, Goo83a, Goo85a, Goo85b, Goo88a, Goo68, Goo89b, Got66, Got67, Got71, Gou82, Gou84a, Gou85, Gou84b, Gow71, Gow73, Gow78, Gow69, Gra81, GG70, GG71, GG72, GG74a, GG75a, GG75b, GG75c, GG77a, GG77b, GG84a, GG84b, GG85, GG89, Gra83b, Gra83a]. Pp [Gra84c, Gra85a, Gra85b, Gra86a, Gra86b, Gra87b, Gre63, Gre71, Gre88, Grm69, Gua89, Gur78, Hac74, Hac75, Hac76, Hac77, Hac82b, Hac82a, Hac88a, Hac88b, Hac89, Had63, Hah70, Hah73, Hai77, Hai80, Hai82, Hal65a, Hal66a, Hal68c, Hal68d, Hal68b, Hal69a, Hal71, Hal72b, Hal77a, Hal77b, Hal80a, Hal80c, Hal84c, Hal84d, Hal84a, Hal85a, Hal85b, Hal85d, Hal86a, Hal86e, Hal86b, Hal86d, Hal87a, Hal87b, Hal88a, Hal88b, Ham84, Han85a, Han70, Han76, Har77, Har84a, Har84b, Har89a, Har76, Har83a, Har86, Har79a, Har79b, Har81, Har82d, Har83b, Har87b, Has72, Hat87, Hea87b, Hea64, Hei68, Hei75, Hei71a, Hei71b, Hei73, Hei74a, Hei74b, Hei76a, Hei76b, Hei77, Hen82c, Hen82b, Hen82a, Hen83a, Hen83b, Hen84a, Hen84c, Hen84b, Hen84d, Hen84f, Hen84e, Hen85a, Hen86a, Hen87a, Hen87b, Hen88a]. Pp [Hen88b, Hen88d, Hen88c, Hen89a, Hen89b, Hen89c, Hen89d, Hen83c, Hen83d, Hen84j, Hen84k, Hen84l, Hen84m, Hen85b, Hen85c, Hen86e, Hen87c, Hen88e, Hen89g, Her80, Hes74, Hew66, Hig65, Hil71a, Hil71b, Hil75, Hil86b, Hil86a, Hil89b, Hil89a, Hoc88a, Hoc88b, Hod70, Hod89a, HS83, HS86a, HS86b, Hol71, HS88, Hom74, Hom80, Hom86a, Hom86b, Hop72, Hop73, Hop75, Hos67, Huf85, Huf86, Hun89a, Hun84, Hun86, Hut80, Hut84, Hut88, Hut89, Ilt77, Ink87, Ink88, Ins86, Irw82, Irw83a, Irw83b, Jac89a, Jac89b, Jac85, Jam86, Jam87a, Jam87b, Jar83, Jar84b, Jar84a, Jen89, Jev67, Joh79, Joh81, Joh83, Joh78, Jon87, Jor78, Jor79, Jor83a, Jor83b, Jor85a, Jor85b, Jor87, Jor88, Jos67, Jud88, Kar76, Kau84, Kee88, Kee89, Kei84, Kei86, Kei89, Kel71, Kel68, Kel70]. Pp [Kel73, Kel75, Kel76, Kel78, Kel84, Kel85a, Kil84b, Kil84a, Kni65, Kni67, Kni69b, Kni69a, Kni71c, Kni71b, Kni72, Kni74, Kni75, Kni76, Kni83a, Kni83b, Kni83c, Kni84a, Kni84b, Kni85, Kni86b, Kni86c, Kni86d, Kni88a, Kni88c, Kni88b, Kni89b, Kni89c, Kni89e, Kni89d, Koh78, Kra84a, Kri83, Kub70, Kuh68, Lar74, Lau68, Law65, Law80a, Law80b, Law80c, Law83, Law85, Law87, Law88, Law89, Lay74, Lay85, Lef78, Les83, Lev73, Lev80, Lew85, Lie85, Lie86, Lin86, Lin89b, Liv86, Loh75, LM74, Lor75, Lor77, Lud80, Mac78, Mac80, Mac85b, Mac72, Mac79, Mah89, Mai85, Mar74, Mar73, Mar89, Mau77, Mau87, May66, May67, May69, McC77a, McC77b, McC80, McC69, MdLD86, McG65, McG66, McG68, McG69, McG70a, McG71a, McG71b, McL86, McN83, McV75].

Pp [McV80, Mea71, Mea84, Men67, Men70, Men74, Mer78, Mil87, Mol69, Mol71, Mol72a, Mol72b, Mol74, Mol75, Mol78a, Mol78b, Mol79, Mol81, Mol84c, Mol84b, Mol85, Mol86, Mol88, Mol89, Mon78, Moo84, Moo85, Mor87a, Mor84b, Mor69a, Mor71, Mor72a, Mor72b, Mor73a, Mor73b, Mor74a, Mor74b, Mor75a, Mor75b, Mor75c, Mor76a, Mor77a, Mor77b, Mor77c, Mor78, Mor80, Mor82a, Mor83a, Mor83b, Mor83c, Mor84c, Mor84d, Mor87b, Mor84f, Mor84g, Mor89a, Mor83d, Mor89c, Mos77a, Mos77b, Mül87, Mül89, Mus74, Nad83, Nak75, Nay75, Nay77, Nay80, Nay85, Nee85, Nee69, Ner89, Nev78, Nev80, Nev83, Nev84, Nic85, Nic88b, Nor68b, Nor69a, Nor70a, Nor71a, Nor71b, Nor72, Nor73a, Nor77, Nor80, Nor82, Nor83, Nor85a, Nor85b, Nun88, Nut82b, Nut83, Nut84, Nye86, OM75, O’H86]. Pp [Olb69, Olb70, Olb71a, Olb71b, Olb72, Olb82, Olb85, Ora74, Ora76, Ora77, Ora80, Osb85, Osb89a, Osb89b, Osp74, Out80a, Out80b, Out82a, Out83, Out84a, Out84b, Out85, Out87, Out88a, Out88b, Pac70, Pac71, Pac72, Pag66, Pag69, Pat86, Pat84, Pau76, Ped85, Pel85a, Pel85b, Pep71, Pep72, Pep86, Phi67, Phi69, Pic84a, Pic87a, Pic76, Pic79, Pic80, Pic88a, Pic88b, Pin83, Pin87a, Pin87b, Pin88, Por89a, Por77, Por78a, Por78b, Por79a, Por79b, Por81, Por82, Por83a, Por84a, Por84b, Por85a, Por85b, Por86a, Por87b, Por87a, Por87c, Por88, Por89d, Por89e, Por89b, Por89c, Por89g, Poy70, Pra70, Pum84, Pum87, Pum88, Pye76, Pye87, Ran74, Rap71, Rap75, Rap79, Rap86a, Rat65a, Rat66a, Rat66b, Rat68, Rat69, Rau75, Rav65, Rav66b, Rav68, Rav69, Rav70]. Pp [Rav71, Rav72, Rav73, Rav84, Rec86, Red84, Ree80, Ree85, Ric88a, Ric89, Rid83, Ril85, Rob75, Rob80a, Rob87, Rob89a, Roc87, Roc88b, Rog83, Rog88a, Rog88b, RH86, RH88, Ron86, Ros79, Rud70, Rud71, Rud83, Rud84, Rud86, Rud89, Rup85, Rup88, Rup89, Rus87a, Rus89a, Rus66, Rus67, Rus68, Rus69, Rus72a, Rus72b, Rus74, Rus77a, Rus77b, Rus77c, Rus79a, Rus79b, Rus85, Rus87b, Rya70, San87, Sar81, Sch81a, Sch83a, Sch84b, Sch84c, Sch84a, Sch86a, Sch86b, Sch87, Sch69a, Sch68a, Sch69b, Sch70, Sch71, Sch72, Sch73a, Sch73b, Sch74b, Sch74a, Sch75, Sch76, Sch78a, Sch78b, Sch79a, Sch79b, Sch80c, Sch80b, Sch80d, Sch83b, Sch85a, Sch78c, Sco67, Scr68a, Scr68b, Sec83, Sec84, Sec85, Sec88, Sei85, Ser88, Sey77, Sha78, Sha89, Sha73a, Sha77a]. Pp [Sha79, Sha81, Sha83a, Sha86, Sha87, Sha77b, Sha73b, She85, She63, Sho86a, Sho88, Sho89, Sin74, Sin76, Sin77, Sin78, Sin81, Sin84, Sme66, Sme68a, Sme69, Sme70, Sme71, Sme84a, Sme68b, Smi66, Smi71, Smi73a, Smi73b, Smi74b, Smi74a, Smi75b, Smi75a, Smi76c, Smi76a, Smi78b, Smi78a, Smi78c, Smi80a, Smi80b, Smi82a, Smi83, Smi85a, Smi85b, Smi85c, Smi85d, Smi87b, Smi87c, Smi87a, Smi89e, Smi89b, Sne74, Sol83, Spa65, Ste62a, Ste62c, Ste62b, Ste70, Ste75, Sti89, Sti70, Str73, Str75, Stu71a, Stu78, Stu85, Stu87, Sty68, Swa86, Swe76, Swi70, Swi71, Swi77, Swi67, Tan82, Tee84, Tei78, Tei83b, Tei83a, Tei84, Tei88, Tha66a, Tha67, Tha74, Tha83a, Tha83b, Tha84a, Tha84b, Tha85, Tha87, The87, Tho76a, Tho76b, Til75a, Til76, Tre87a, Tre87b, Tre88, Tup82]. Pp [Tur68, Tur69, Tur70b, Tur70c, Tur72a, Tur72b, Tur73d, Tur73c, Tur73a, Tur73b, Tur74a, Tur74c, Tur75, Tur77a, Tur78a, Tur78b, Tur78c, Tur81, Tur82, Tur83a, Tur83c, Tur87a, Tur88, Tyl75, Und64a, Und66c, Und66a, Und66b, Und75a, Und75b, Und79a, Und79b, Und80, Van86, Vea64, Ver89a, Ver89b, Wal89, Wal74, Wal77, Wal81b, War88a, War88b, War89, Was69, Wat73, Wea78, Wea79, Wea80, Wea84, Web75, Web76, Web78, Web84, Web65a, Web65b, Web66a, Web66b, Web66c, Web67a, Web68a, Web68b, Web69b, Web69a, Web70, Web71a, Web71b, Web72, Web87, Wei83a, Wei85a, Wei85b, Wei86, Wel70, Wes74, Wes70, Whe89, Whi70a, Whi68a, Whi72a, Whi72b, Whi73, Whi75, Whi85a, Whi89a, Whi68b, Whi85b, Whi88a, Whi89b, Whi68c, Whi77, Wig69, Wig74, Wig77, Wig78, Wil65a]. Pp [Wil75, Wil65b, Wil68, Wil69, Wil77, Wil84d, Wil84b, Wil84c, Wil85a, Wil85b, Wil86a, Wil86b, Wil87d, Wil87c, Wil87b, Wil78, Wil89b, Wil89a, Win75, Woo80a, Woo80b, Woo80c, Woo83, Woo84, Woo89, Woo67, Woo73, Woo80e, Woo85, Wor88, Wuj85, Yat72, Yeo88, You66, You83, Yox82b, Yox82a, Yox83a, Yox83b, Cro85, Dav81, Dun87, Fie83, Goo81a, Goo82, GG81, GG87, Hal84e, Hal86c, Hen84g, Hen86b, Kni80a, Kni80b, Kni81, Kra86, Law81, McC83, McC85b, Rav67, Rud80, Sch81b, Whi84]. Pp.ix [Col88b]. Pp.x [Bow88b]. Pp.xix [Fre81b]. Pp.xvi [Lor73, Nay80]. Practical [Cro83, Fox70, Fre79, Jam88, Mor83d, Wil93]. Practice [Dav81, Goo88a, McL86, Pic88a, Pic88b, Rus77a, Ano82p, Bia93, Eva77, Gab83a, Ner87, Ole91, Pet90, Ram88, Sho86b, Sir90, War86, B^{+}82]. Practices [Hun86, Mur85]. Practitioners [Sco67]. Prag [Hei76b, Nor70a, Nor73a, Pep72]. Prager [Kel73]. Pragopress [Rav69]. Prague [Rav69, Rya70]. Pratike [Mor83d]. Pre [Hen88a, CK90, Wil87a]. pre-industrial [CK90]. Pre-Newtonian [Hen88a, Wil87a]. Précision [Ben89, Pic87c, Wil94a, Pic87c, Wis95]. preclassical [Gre92a]. precursor [BM82, Out84a]. predecessor [AB89]. Predecessors [TP66, Lin82]. Preface [Bur74a, Jar84b, Wil85a, Dug77, FC82, McG70b, Mol69, Hal68c]. Prefaced [Tur81]. preformation [Cor97]. Prehistoric [Kno85, HW87, Rud92]. Prehistory [Law65, NW82b, Smi87b, Ano82n, Har85b, Van93]. préliminaire [dSR95]. Preliminary [Buc68, dSR95, For86, Kni83a, New89].

Prelude [Mol84c, Wal81a]. première [Grm90]. premiers [Poi04]. Prentice [Rav73]. Prentice-Hall [Rav73]. Prentiss [Woo83]. Prescott [Goo81b]. Present [Scr68b, Wil62, Wil69, App93, BW90, Dea83b, MJ89, Pep89]. presentation [Gol79]. Presented [Wil87d, ABR93, De 85, Pep89, Yox79]. Presenting [Sin77]. Preserved [Cor81, HAB92]. Presidencey [Hom86b]. Presidency [Mil83]. president [Wal88]. Presidential [Mor72b, Bro81, Hal69b, Mar64, Wat80]. Presocratic [Osb85, KRS83]. Press [Aas87, Agr83a, Agr83b, Ait67, Ait72, Ait73, Ait85a, Ait89, All82, Arm67, Arm70, Asp88, Bad71, Bad89, Bal84, Bar79, Bar81, Bar76b, Bea85b, Ben82b, Bib74, Blo82b, Bod65, Bod84, Bow74, Bow87a, Bow89a, Bow83a, Bow85, Bow86, Bow87b, Bow87c, Bow88a, Bow88b, Bow89b, Bow89c, Boy70, Bra84, Bra89a, Bro67, Bro69, Bro71a, Bro84a, Bro85a, Bro80b, Bro89c, Bro72b, Bro73b, Bro75, Bro82b, Bro78a, Bro78b, Bro86c, Bro89e, Bro83a, Bro84c, BH84b, Bro85d, Bro85e, Bro86d, Bro89f, Bru72, Buc77, Buc69, Buc71, Buc75, Buc66, Buc73b, Bud84, Bur71, Bus85, Bus74, Byn73, Byn74, Byn88, Cah88, Cal77, Can74, Can77, Can80b, Can80c, Car67a, Car70, Car84, Cas82, Cas85, Cha65, Cha78a, Cha77a, Chi64, Chi65a, Chi68, Chi71, Cla86a, Cla67, Cla68, Cla70, Cla71, Clo80]. Press [Coc75, Col81a, Col87b, Col88b, Col81b, Con86, Coo81, Coo83, Coo89a, Coo84a, Cow74, Cow68, Cri71, Cro68a, Cro85, Cro87a, Cro89b, Cun78, Cun83b, Cun88, Cun89, Dau75, Dav71, Dea83b, Dea86a, Dea86b, Dea86c, Deb66, Des88, Din64, Dol70a, Dol70b, Dol71a, Dol71b, Dol74, Dor80, Dub72, Duc84, Duf81, Duf83, Duf75, Dur79a, Dur81, Dur82, Dur83, Ear70, Ear87, Edg87a, Edg87c, Edg87b, Edw78, Eme87, Eva89, Far67, Far69, Far76, Far77, Fee77, Fel89, Fie81, Fie83, For68, For71b, For74a, For85a, For87c, Fox69a, Fox69b, Fox70, Fox71b, Fox74, Fox82, Fra87a, Fre70, Fre74, Fre78, Fre86, Fri74a, Fri87, Fur85, Gab83b, Gab83a, Gab75, Gab83c, Gab88, Gel76, Geo66, Geo87, Gil88, Gil84, Gin84, Gli83b, Gol87b, Gol88b, Goo83a, Goo85a, Goo85b, Goo68, Got66, Got67]. Press [Got71, Gou82, Gow73, Gow78, Gow69, GG72, GG74a, GG75a, GG84a, GG84b, GG87, GG89, Gra85a, Grm69, Gur78, Hac74, Hac77, Hac89, Had63, Hah73, Hai77, Hai82, Hal66a, Hal69a, Hal71, Hal77b, Hal80a, Hal84e, Hal84c, Hal84d, Hal85a, Hal85b, Hal86a, Hal86c, Hal86d, Hal87b, Hal88b, Har84a, Har75, Har89a, Har76, Har79b, Har82d, Har83b, Hei71a, Hei71b, Hei73, Hei74a, Hei74b, Hei76a, Hei76b, Hei77, Hen81, Hen83a, Hen84f, Hen85a, Hen86b, Hen87b, Hen88a, Hen88b, Hen88d, Hen89a, Hen89b, Hen89d, Hen89e, Hen83d, Hen84j, Hen84k, Hen84l, Hen84m, Hen85c, Hen86e, Her80, Hes74, Hil71b, Hil89b, Hil89a, Hoc88b, Hod70, Hod89a, HS83, Hol71, HS88, Hop72, Hop73, Hop75, Hos67, Huf85, Huf86, Hun84, Hut84, Hut89, Ink88, Ins86, Irw83b, Jac89b, Jac85]. Press [Jam87b, Jar83, Jev67, Joh81, Jon87, Jor78, Jor79, Jor85a, Jor85b, Jor87, Jos67, Jud88, Kar76, Kau84, Kee89, Kei84, Kei86, Kel70, Kel73, Kel84, Kel85a, Kil84a, Kni67, Kni69b, Kni69a, Kni71c, Kni72, Kni74, Kni75, Kni76, Kni80a, Kni80b, Kni81, Kni83a, Kni83b, Kni85, Kni86b, Kni89e, Kni89d, Kra84a, Kri83, Kub70, Kuh68, Lau68, Law80b, Law80c, Law81, Law83, Law85, Law87, Lay74, Lef78, Les83, Lew85, Lie85, Lin89b, Loh75, Lor73, LM74, Lor75, Lor77, Lud80, Mac80, Mac85b, Mac72, Mah89, Mar74, Mau77, May66, May67, McC77a, McC77b, McC85b, McG65, McG68, McG70a, McG71a, McL86, McN83, Mea71, Mea84, Men67, Men70, Men74, Mer78, Mol69, Mol71, Mol74, Mol78b, Mol79, Mol81, Mon78, Moo84, Moo85, Mor73a, Mor73b, Mor75a]. Press [Mor75c, Mor77a, Mor77b, Mor80, Mor82a, Mor83b, Mor84d, Mor87b, Mor89a, Mos77a, Mos77b, Mül87, Mus74, Nad83, Nak75, Nay75, Nay85, Nee85, Nee69, Nic85, Nor68a, Nor68b, Nor69a, Nor70a, Nor71b, Nor73a, Nor77, Nor80, Nor82, Nor83, Nor85b, Nun88, Nut84, O’H86, Olb69, Olb70, Olb71a, Olb71b, Olb85, Ora77, Osb85, Osp74, Out82a, Out83, Out84a, Out85, Out87, Pac71, Pac72, Pel85b, Pep72, Pep86, Phi69, Pic76, Pic79, Pic80, Pic88a, Pic88b, Pin83, Pin88, Por78b, Por79b, Por81, Por82, Por83a, Por84b, Por85a, Por85b, Por87b, Por87a, Por87c, Por89c, Por89g, Pum84, Pum87, Pum88, Pye87, Ran74, Rap71, Rap79, Rap86a, Rat65a, Rat68, Rav68, Rav70, Rav71, Rav72, Rav84, Rec86, Red84, Ree80, Ric89, Rid83, Rob87, Roc87, Rog83, RH86, RH88]. Press [Ros79, Rud69, Rud71, Rud80, Rud84, Rud86, Rud88b, Rup85, Rup89, Rus87a, Rus89a, Rus66, Rus69, Rus77a, Rus77b, Rus77c, Rus85, Rus87b, San87, Sch81a, Sch84b, Sch84c, Sch86a, Sch87, Sch69b, Sch70, Sch71, Sch76, Sch78a, Sch78b, Sch80d, Sch85a, Sco67, Sec83, Sec84, Sec88, Sha78, Sha73a, Sha83a, Sha86, Sha73b, Sho87a, Sho86a, Sho88, Sin76, Sin84, Smi71, Smi73b, Smi74b, Smi75a, Smi80b, Smi82a, Smi83, Smi85a, Smi85c, Smi87b, Smi87c, Smi87a, Smi89e, Spa65, Ste75, Sti89, Str73, Swa86, Swi71, Tei78, Tei83b, Tei83a, Tei84, Tei88, Tha66a, Tha74, Tha83a, Tha87, Tho89, Tur69, Tur70b, Tyl75, Und66c, Und66a, Und66b, Und75a, Und75b, Und79a, Und79b, Und80, Vea64, Ver89a, Ver89b, Wal89, War88a, War88b, War89, Was69, Wat73, Wea78, Wea84, Web75]. Press [Web78, Web65a, Web65b, Web66b, Web66c, Web67a, Web68a, Web68b, Web69a, Web70, Web87, Wei85a, Wei85b, Wel70, Wes74, Whi70a, Whi73, Whi75, Whi85a, Whi89a, Whi68b, Whi89b, Whi77, Wig77, Wig78, Wil75, Wil65b, Wil68, Wil84c, Wil85b, Wil86a, Wil86b, Wil87c, Wil78, Wil89a, Win75, Woo80a, Woo89, Woo67, Woo85, Wor88, Yeo88, You66, Yox82a, Yox83a, Gra83a, LP04, Nay80]. Presses [All84a, Cal63, Ege89, Hal65a, Hil71a, Smi85d]. presuppositions [Ver85]. preussischen [ZH92]. Prevention [Fre86]. Prévost [FM00, FM00]. Price [Ait86, Arm62, Arm64, Bos70, Chi64, Edg87c, For67b, For69a, For69b, Fre70, Hil71a, Kni83c, Mor84g, Pic88a, Rus67, Rya70, Sha79, Sme84a, Smi78b, Ste62a, Ste62c, Ste62b, All77, All84a, Bos72, Bow87a, Bow83a, Bow86, Bro76, Bro83b, Bro84d, Bry77, Can74, Car74, Cha77a, Duc88, Duf81, Duf75, Far76, Fie81, Fig77a, Gab83b, Gli84, Gol79, Gol83b, Goo89b, Gow71, GG74a, GG89, Gra87b, Hal84a, Har89a, Hat87, Hen83b, Hen89c, Hoc88a, Hod89a, Irw83a, Jor83b, Kel75, Kel78, Kni85, Kni88c, Kni89e, Lay74, Mah89, Mau77, Mau87, McC83, McV75, Mor83d, Nay75, Nic88b, Nut84, Olb82, Ora76, Osp74, Out80a, Pic84a, Por78b, Por89d, Rud89, Rus72b, Rus77a, Sch75, Sch81b, Sch78c, Smi75a, Smi80b, Sne74, Str75, Tei83b, Tho76a, Tur72b, Tur73d, Tur73b, Whi89b, Wil75]. price [Wil85a, Wil87d, Win75, Woo80b, Woo80c, Woo89, PN95, Tan82]. priced [Ree85]. Pricha [Hun89a]. Prichard [Web75]. Priestley [Cri66, Hen89g, Kni69b, AL87, Cro83, Cro87b, Gol88a, McE79]. Prigogine [Bru86, Hen82c]. prima [Olm92]. Primary [Gol79]. Primitives [Und75a]. Primus [Mol81]. prince [But96, Dew63]. Princeton [Bro84a, Can77, Car84, Coo84a, Cun83b, Cun89, Dea86c, Gab75, Goo85b, Hac77, Hai82, Hen88d, Hen84l, Hod89a, Irw83b, Jon87, Jor85a, Kil84a, Lor75, McC77b, McG71a, Men74, Moo84, Mor77b, Mor80, Nor80, Nun88, Olb71b, Por89g, Ran74, Smi89e, Tha87, Wei85b]. Principal [Car67a]. Principales [Mol84b]. principe [BH90]. Principia [Ait72, Sch84c, Whi73, Ait67, Ber93b, BN82, Gan95, New89]. Principle [Dau74, Dor79, Yeo86, BH90, Kip91, MJ89]. Principles [Jar84a, LR91, LS97, Por76a, McG70b, Cor78, Osp76]. Printed [Blu79, Cor81, Ede65, Sch68a, Tur68, Deb66, Got92, Hal88a]. Printing [Hal85b, Hal86c, Kni80b, Kni83a, Scr68b, Chi65a, Eis83, Eis93, NhTWL87]. priority [Ber93b]. Prisoner [Pou93]. prius [Rav69]. Private [For74b, Gei95]. Prize [Ano88s]. Prizes [Hea64, Rec86, Tre88, Cra84, CHU87]. pro [JK84]. probabilità [Des89b]. Probabilities [Hah70]. Probability [Hen84l, Lor77, Por89g, Sha83b, Das88, Des89b, G^{+}89, KI92]. Problem [Moo81, Nay74, CS92, Gro91, Har82c, Hen84f, Her65b, Law85, WT82]. Probleme [MD90, Whi68a, Wil77]. Problems [Gow74, Jam88, MD90, McG70a, Nor82, Por89d, Smi73a, Smi76c, Stu86, Web88, Can84, Fur89, JO81, Llo91, Nev84, Rap82]. Proceedings [Ano62k, Che78, Dol71b, Dor80, Dun86, Mac88a, May67, Nut84, Por81, Sar81, Chr93, CT95, EK88, EK91, EK92, KK85, N^{+}88, WV92]. Process [Goo88a, Sch69a, Ver89a, Fie83, Hul88, Ner87]. processing [NOF96]. Prodigy [O’H86, O’D83]. Prodromus [Sch69a]. production [BR95, EK88]. Productivity [Mor77b]. Profession [Byn88, Fig78, Fox87b, Pel85a, Por88, Wig77, Eng83, Gol87a, Gre86, Mor84d, Nic88a, Pat83b]. Professional [Ram88]. professionalization [Geu92]. Professionalizing [Gab83b]. Professor [Ano64l, Ano65j, Dra86, McG71b, Mor69b, Sme66, Ash85, Ano65j, Ano67i, Ano67j, LB73]. Professors [Rat69]. program [McC93, Mor89b]. Programme [Kei84]. Progress [B.69, Bar76c, Dol73a, Hoc86, Kni69a, McE79, Por78a, Bow89e, Bro80a, Bro84d, Cat88, DPP88]. Progression [Bar76c]. Progressive [Pic80, Por78a]. Progressives [Olb70]. Project [Eag78, Fau89, Jac89b, SG80, Sep88]. projects [Cun97]. projet [SG80]. Promoting [Hal91]. Proof [She70, Wal92]. Propaganda [McC77b]. Properly [Ric87a]. Property [Sha74]. prophecy [Smo94]. Prophet [Ash85, Dau75, Dra86]. Prophets [Wei77]. Prophylaxis [Wei85a]. proposed [CMDM82]. proposées [CMDM82]. Propulsion [Kei84]. Prose [Gol88b]. Prospect [Bar73b]. Prospects [GG79b, Bro80a, Bro84d, MJ89, Mar10]. prosper [Sto88]. Protestant [Pum88, Par87]. protestantism [Hen84m, Har98b, Jac83]. Protestants [Wil81b]. Proton [Nye86, Bro85c]. Protyle [Nye86, Bro85c]. Proust [Fuj86]. Prout [Nye86, Bro85c]. provenance [Ait85a, JK84]. Provence [GG77b]. Province [Sec85]. Provinces [Hil71b]. Provincial [Gli84, Wil65b]. Provincials [Mor74a]. Provine [Bow83a, Bow87c]. PRs [Sar81]. Prussian [Bru88a, ZH92]. pseudo [Sha83a]. pseudo-science [Sha83a]. pseudosciences [Got92]. Psyche [Eva89, Fur88]. Psychiatric [Byn88, BP95, Gol87a, Lun94, PB99]. Psychiatry [Byn87, Pic88a, Sho87c, Ver85, BP95, Har83a, PB99, Sho86b]. Psychical [Sho86a, Opp85, Pin83]. Psycho [McC77a]. Psycho-History [McC77a]. Psychoanalysis [For87c, Sho87c, TR92, Wei84, Bru95, MF86, Ril85]. psychoanalytic [CS92]. Psychological [For87b, Gil88, Smi80b, Cra93, Dan90, Ric92a, Sok87]. Psychology [AW87, Fan83, For87b, Fox87a, Ric87b, Ric88a, Sti89, Ash95, Geu92, Hea87c, Ilt73, Lea90, Lea94, Sha98, Shu96, Shu04, Sim88b, vG92]. pt [Agr83b, Sch69a, Woo83, KN88]. Ptolemaic [GB95]. ptolemaico [GB95]. Ptolemy [Kre69, Van86, Hen91, PT84]. Pubilée [McC80]. Public [Jor85b, McC77b, Mor85, Sch68a, Wig77, Dau98, Gol92b, GR88, Irw83b, IW96, Lan87, Lan89, OE^{+}90, Por95, Ste92, Yeo93]. publica [Lan89]. Publication [Cro73, Mac88a, Pac71, Hei76b, Nor68a, Nor69a, Nor70a, Nor73a, Pep72]. Publications [Ano62e, Ano63h, Ano63i, Ano64i, Ano65g, Bar76a, Bro76, Bro85e, Cha77b, Fre81b, Goo81b, Gra81, GG87, Hen84c, Irw82, Koh78, McC85b, Nut83, Ora80, Por78a, Sin81, Smi75b, Smi78b, Smi89c, Vea64, Wal77, Wea79, Yox82b, Tor83]. Publick [Whi72a]. público [OE^{+}90]. Publics [Out88b, Lan87]. Publiée [McC83]. Published [Cal63, For68, Kni71c, Sch75, Smi89c, Mac88a, Tur70c]. publisher [For87a]. Publishers [Aus86, Buc85a, Col76, Fre82, Goo88a, Hac76, Han85a, Mol89, Por89a, Roc88b, Tur74a, Twe89, Woo84]. Publishing [Ait86, Ant63, Bro86b, Bry78b, Bry78c, Bud84, Byn87, Can84, Car74, Che85, Fis81b, Gou85, Hac82a, Hea64, Hen83b, Kni86c, Kni86d, Mac78, Mor72a, Mor83a, Ora76, Out84b, Rob80a, Rud86, Rud89, Sch69a, Sha89, Smi85b, Ste62c, Tee84, Tha85, BM84, Fre82]. Pugh [Hen85a]. Pump [Jon87, SSH85]. punishment [DH90]. pupils [Cun83b]. purchasers [Hen88d]. Pure [Per73]. Puritan [Rab78]. Puritanism [CDS90]. Purkyne [Tei70]. purple [Gab83a]. purposes [Smy89]. Pursuit [Gur78, Wil84e, Mac83, Por95, Ric88b, FTG91]. Purver [Hal68d]. puzzle [Pou93]. Pyenson [Pin87b]. Pyne [Por82]. Pythagoras [Was69, O’M89]. Pythagorean [Sch78b]. Pythagoreanism [Was69].

Qian [Ron86]. quackery [Por89f].

Quadrane [Mol84b]. Quaestio [Law93]. quaestiones [Gro89]. qualitative [Bli92]. Qualities [Hen86e, Ale85]. Quality [LaF82, Rav84]. Quantification [Mat95].

Quantifying [Coh84, Gou85, FHR90]. Quantitative [Mon87]. Quanto [McC77a]. Quanto-History [McC77a]. Quantum [Blo84, Cas82, Che85, Dor75, Hen86b, Hen88c, Kuh68, Pop82, Hen84i, MR82a, MR82c, MR82d, MR82e, MR82f, MR82g, MR87b, MR00]. Quarks [Hen86b, BDH89, Pic84b]. quarrel [Sch81a]. quarter [Buc85d]. que [ML89]. Quellen [Mol72a, Gro89]. Quelques [Cas85]. Quentin [Hut89]. quest [Ash95, Fuj96, Mat95, Sta89b, vdVV91]. Question [Ric87a, Rud71, Yox79]. Questions [Hen85c, Luc90, Nay80, Nye76, MT83]. Quetelet [Por85d]. Quigly [Rav66a]. Quinine [Kni89e, GL88]. quintius [Eas86]. Quod [SLT88].

R [Ait89, All81, Ano65j, Ano67j, Blo79, Blo84, Bow89c, Bro82b, Bry78c, Chi64, Coc75, Cow68, Cro68b, Cro69, Cro70b, Dol70b, Dun87, Fis81a, For74a, Fox71a, Fre74, Fre81b, Fri87, Fur85, Gol92a, Got67, GG74a, Hac89, Hal68d, Hal80c, Hen81, Hen83b, Hen83c, Hen84m, HS88, Ilt77, Irw83b, Jam86, Jam87a, Jar84a, Kei84, Kel68, Kel75, Kni85, Law83, LB73, Mac78, Mac85b, Mar89, Mor75a, Mos77b, Nad83, Ora76, Pel85b, Pic79, Pin83, Pin87a, Por84b, Pum88, RH88, Sch83a, Sch74b, Sco67, Sin81, Sme66, Ste70, Ste75, Swe76, Swi70, Tur74c, Web84, Web71a, Wes74, Wil68, Wil87d, Wis85, Woo84, You83]. R. [Bea85b, Dun87, Fau89, Hac88a, Hen89g, Mac88a, McL86, Smi89c]. Rabdology [Nap90]. Rabinovitch [Lor77]. Race [Olb77, Por87c, CT93, WKB88]. Rachel [Hen84d, Bro89f]. Rachootin [Nor82]. Racial [Pro88, Lil85, WKB88]. Radar [Duc84, Cal90]. Radiation [Hun89a, Sin78, Cau89, OP87]. Radical [Can76, Des89a, Hen84m, Jac83, Sch93, Gib84]. Radio [Bow89a, Duc84, Edg84, Poc88, Rob92, Sul82b]. Radioactive [Fre79]. Radioactivity [Sin77, Sin77, Hei71b]. Radiochemistry [Bad79]. Rafter [Cro89b]. ragione [Ern91]. Ragnar [Cla66]. Rahman [Wuj85]. Raihan [Sar81]. railroads [Gil83]. rainbow [Tra93]. Rainger [Bow89c]. raison [Bea91]. raisons [Bla93]. Raj [Kum95]. Rajkumari [Kei89]. Ralph [Har84b, Men67]. Ramón [Cla67]. Ramus [Sme68b]. Randall [Chi64]. Rankine [Hut81]. Ranum [Gel76]. rarest [PR95]. Raspail [Men70]. Rasse [WKB88]. Rassenhygiene [WKB88]. Rassenkunde [MD90]. Rassenpolitik [Har87b, Lil85]. Rather [Law81]. Ratio [GG89, HF87]. Rational [Ner89, PP87]. rationale [Agr83b]. rationalist [CD94]. Rationality [Wil84d, Wyn82, Can88]. Rauschenberg [Ste70]. Rauwolf [Web70]. Raven [Ano64l, But65, Osb85]. Ray [All79, All84b, Dea83b, Nee85, Whi88a, Hir83]. Rayleigh [Mea71]. Rayleigh-the [Mea71]. Raymond [Bra89a, Rav68, Per89a]. Rays [Fal87, Stu71b, Dah97, Tur70c, Tur73d]. Razzaz [Kel75]. Reaction [Ait86, Bro75, Des85, Ger71, Rus76b, Smi80b, Gra84b]. read [Mir94, Ste67, Whi72a, Woo80b]. Reading [Fre70, Win91]. Readings [Rus72a, Whi73, Whi68b, Sch01]. real [Sin91]. Realism [Blo84, PB83, Nol88]. Realities [Mea87]. Reality [Sin74, Bli92]. Reappraisal [AZW87, Bro71b, Sha89, Wes67, Sar92]. Reappraisals [LW90]. Reappraised [Arm67]. Reason [Hol78, McG71a, Por77, CS92, Ern91, Gut89, Pow96]. reasonable [Hey95]. Reasoning [Ner89, PP87]. Reasons [Wil62, Bla93]. Reassessed [Ric87a]. Reassessment [Gol77, Olb87]. Rebecca [Tre88]. recapitulation [MC^{+}93, CM93, MC^{+}93]. Received [Ano62d, Ano63e, Ano63f, Ano64e, Ano64f, Ano65e, Ano65f, Ano66e, Ano66f, Ano67e, Ano67f, Ano68f, Ano68g, Ano69e, Ano69f, Ano70e, Ano70f, Ano71e, Ano71f, Ano72e, Ano72f, Ano73e, Ano73f, Ano74g, Ano74h, Ano75g, Ano75h, Ano75i, Ano76g, Ano76h, Ano77h, Ano77i, Ano77j, Ano78g, Ano78h, Ano78i, Ano79g, Ano79h, Ano79i, Ano80g, Ano80h, Ano80i, Ano81g, Ano81h, Ano81i, Ano82g, Ano82h, Ano82i, Ano83g, Ano83h, Ano83i, Ano84g, Ano84h, Ano84i, Ano85g, Ano85h, Ano85i, Ano86g, Ano86h, Ano87j, Ano87i, Ano87k, Ano87l, Ano88i, Ano88j, Ano88k, Ano88l, Ano89i, Ano89j, Ano89k, Ano89l, BDH96]. Recently [B.69]. Reception [Gli87, Hen86b, Jun89, Smi74c, Tho76a, Wea80, Whe89, MR82c, MR82e, MR87b, dG01, dG01]. Rechenberg [Hen86b, Hen88c, Kra86]. Rechenmaschinen [MKW92]. Rechenmeister [Sch93]. Recherche [Ben89, Gua89, Hal85c, Hut80, McC80, McC83, Sch84c, Ano82l]. Rechnen [Ger89]. Recht [Nor71a]. Rechts [Nor71a]. Recieved [Ano74i]. Recollections [Cla67, Hen81, Kra86, Duf81, GH84, GH96, Hal85a]. Recombinant [Olb85, Kri82]. reconsidered [U^{+}91]. Reconstructing [BC98, Hum67]. reconstruction [IW96, Ric92b, YK94]. Reconstructions [Blo82b]. Records [Mea87, Dol69]. redaction [BEoCF92, ECL92]. Redirecting [Aas90]. Rediscovered [Hen87a, GJ85]. Rediscovery [Coc75, Wil62]. Redmond [Por85b]. Redner [Whi89a]. Redondi [Out80b, She85]. Reduction [Smi76c, Gro91]. Redwood [Por77]. Reece [Dor75]. Reede [Hen86d, Mül89]. Reefs [Sto76]. Reeling [Hil89b, KN88]. réelle [Sin91]. Reenchanted [Har96]. Rees [Vic88]. Referate [Rus74]. reference [Olb82]. reflection [Sha93]. Reflections [Cla68, Lev68, Grm90]. Reflexes [Cla71]. Reflexions [BM76, CF86, Smi87a, Grm90]. Reform [Dur79b, Kni89c, Mor72d, Wei85b, Des89a, Fri81, Gle91, Hen85b, Mar92, N^{+}88, Poo98, Yat83]. Reformation [Pum88, Par87]. Reformer [Men70, Hen84k, Pag82, Mor69b]. reformers [Coo84a, Who82]. Refraction [Bec75, Lin68]. Rega [Mol89]. Regenerator [Dau74]. regime [Hai82]. Regiment [Roc81a]. Regine [Bro89a]. Regiomontanus [Ger89, Ger89]. Regiomontanus-Bianchini [Ger89]. Region [Web87]. Regions [Sta89b]. Register [All77]. Regnault [Bru72]. Regulation [Stu88]. regulatory [Wri94]. Rehbock [Bro85e, Kni89d]. Reich [Edg87c]. Reichenbach [HRP90]. Reichs [Cap81b]. Reidel [Bar83, Blo79, Bro86b, Byn80, Byn87, Can84, Cha80b, Che85, Dea83a, Edg84, Fra87b, Gau83, Gou85, Hal86e, Hal86b, Hen82c, Hen82b, Hen84e, Hen87c, Hen88e, Hoc86, Jam86, Jar84a, Kel75, Kni84a, Kni84b, Kni85, Kni89b, Law83, Mol84c, Nay80, Ner89, Out84b, Ped85, Pic84a, Pic87a, Pin87a, Por86a, Sch84a, Sha89, Stu78, Tho76a, Tho76b, Web84, Whe89]. Reidemeister [Nor72]. Reihe [GG77a, LS76, Scr68a]. Reingold [Bud84, Goo85b, Hac74, Hac77, Ink88, Koh78]. Reinhard [Rus79a]. Reinhold [Ste62c, Hen91]. Reinterpretation [Cro89a, Hal86d, Don88, Eme84]. reinterpreting [Bow88c]. Related [Ait89, FB75, Old77, Smi76c, McG70b, vLMB88]. Relating [Stu86, Web68a, Ant64]. Relation [Jev64b]. relational [Ear89]. Relations [Ear70, Hai78, Kei86, BM82, Ber95, Jor83b, McG84, MM10, Out84a, Tei83b]. Relationship [PW89b, Sch89]. Relationships [Kni75, Hen84l, Sha83b]. Relative [Hen88a, Bar89a, Bro82a, Wil87a]. Relativism [Nol88]. Relativitätstheorie [Wol87]. Relativity [EK91, Pin87b, Whe89, Whi88a, EK92, Gol70, Goo82, Pye85, Gli87, Ray87]. Relevance [Buc63, Web67b]. Religieux [McC80, McC83]. Religion [Bau75, Bro89c, Bro75, Ern91, Gol91, Hol78, Por77, Wes74, Yat72, BC98, Can91, CMDM82, Clu88, Cor88b, Hal90, Hen84l, KAZ92, Lan91, Sha83b]. Religione [Ern91, Lan91]. Religious [Liv86, Rus77b, Dur85, Lin92]. Reluctant [Mac89, Alt87]. Remember [Kau84, Tod83]. Reminiscence [Hen80]. Reminiscences [Hen82b, Hen89b, KW87, Hod92]. removal [MJ89]. Renacimiento [Goo89b, Lan86]. Renaissance [Ant64, Bro84c, Col81a, Cun89, Dea86a, Gol91, Hen84j, Hen84k, Hen85b, Hut84, Hut89, Kel78, Loh75, Mol72b, Mol84c, Nut84, Sch70, Sch78b, Sch79a, Sch80d, Sim84, Wea80, Wig74, Wig77, Cun97, AO77, Bro80b, Col81a, Coo81, DH90, Ern91, FJ93, Fie97, Gol91, Gra81, Kel70, Kel73, Kel78, Kni80b, Loh75, Lan86, Mol72b, Rus79b, Sch73a, Sch78b, Sch79a, Sch80b, Sch80d, Sch81b, Sch83c, SSK88, Sha81, Sir87, Sir90, Sir97, TW95, Tur73a, Vic84, Wea80, Wig78, Yat83]. Renard [PR95]. Renato [Duc88]. Renaudot [McC77b]. René [Byn74, Cro66b, Fin68, Jar84a, Sch84c, Sch74a, She91]. Reorganized [Eme87, McC85a]. replications [BD95, BD95]. Reply [Ink81, Wor79]. Report [Ano64n, Ano75o, Ano76l, Ano77p, Ano78m, Ano80m, Ano81n, Ano82o, Ano83m, Ano84j, Ano85j, Ano86i, Ano87m, Ano88r, Ano89q, BU63, CM64, CC65, CC66, CH68, CW69, CE74, EW70, Hal79, Hof62, HS63, Scr64a, WE71, WE72, Kni83a, Olb82, Smy89, WR92]. Reports [Sin77, Wil85a]. Representation [Fie86, BB84, Sal85]. representations [G^{+}09, Hay94, MR96, Rud92, G^{+}09]. Reprint [Arm66, Buc68, Car74, Cla67, Dol73a, Dun69a, For72, GG75b, Lar74, McC69, McG68, Ora73, Pac71, Rat69, Rav69, Sch74a, Sme69, Web69b, Whi72a, Bos72, Col76, Kni88c]. Reprinted [McG68, Scr68b]. Reprints [Bar75, Bru88c]. reproduction [Sin77]. Republic [Gur78, Sch84a, Bro89d, Cat88, Gol95, Lan87]. Republicains [Por89e, Roc88a]. Republicans [Roc88a]. Republik [Mor84g]. République [Out88b, Lan87]. Research [Arm62, Arm64, AB89, CM90, Duc84, Edg87c, Fri87, Gin84, Gol79, Gow71, Gow69, Hen82c, Hen84d, LM74, Mac72, Mor89a, Rud63, Sho86a, Smi66, Smi89b, Bec82, Bro84d, Coo83, Dan90, Dea83a, Dei92, ES90, Ern91, Fru90, Gab83a, GH92, Gar87, Hen84e, Irw83a, KH86, Mor89b, Opp85, Pin83, Por82, Rei85, Sta89b, Swa88, Tup82, Wis85, GH93, Bro80a]. Researches [McC69, Web75]. Reservoir [Bos83]. Resistance [Bau95]. Resolution [Sho88, EC87]. Resonance [Pic87a, Zan86]. Resources [Ano77m, Ban88, Cam82, Day85, Dou88, Lay74, MJ84, PS78b, WG81]. respecting [MJ89]. Respiration [Jev67]. Response [Bar73b, Byn78, Hei77, Hen88c, Hof87, San88]. Responses [Bar76c]. Responsibility [Tre79, Bus85]. rest [Rog83]. Restaging [BD95, BD95]. Restitutae [Whi85a, Viè83]. restoration [Gou82, Wil93, Kro91]. restored [Whi90]. Resurrected [Smi82c]. resurrection [Cun97]. retarded [Nol95]. Rethinking [Meg94]. retirees [GG92]. Retirement [Mar63]. Retreat [Por89b, Smi88]. Retrospect [BBBM90]. Return [Jen88]. Rev [Ait71, Ano64l, For85a]. Rev. [CHG91, Gun84]. revealed [Fin93]. Revelations [Win75]. reverence [SB92]. Review [Aas87, Agr83a, Agr83b, Ait67, Ait68, Ait71, Ait72, Ait73, Ait85a, Ait89, All77, All79, All81, All82, All84a, All84b, Ant63, Ant64, Arm62, Arm64, Arm66, Arm67, Arm69, Arm70, Asp88, Asq80, Aus86, Bad68, Bad71, Bad83, Bad89, Bal83, Bal84, Bar77, Bar79, Bar81, Bar83, Bar73a, Bar76a, Bar75, Bar76b, Bea85b, Ben79, Ben89, Ben82b, Bib74, Blo79, Blo82b, Blu79, Bod65, Bod84, Bos70, Bos72, Bow74, Bow87a, Bow89a, Bow83a, Bow85, Bow86, Bow87b, Bow87c, Bow88a, Bow88b, Bow89b, Bow89c, Boy70, Bra88, Bra84, Bra87, Bra89a, Bro67, Bro69, Bro71a, Bro73a, Bro82a, Bro84a, Bro85a, Bro86a, Bro86b, Bro89a, Bro89b, Bro80b, Bro72a, Bro76, Bro89c, Bro72b, Bro73b, Bro75, Bro82b, Bro78a, Bro78b, Bro86c, Bro89e, Bro83b, Bro83a, Bro84c, Bro84d, BH84b, Bro85d]. Review [Bro85e, Bro86d, Bro86e, Bro87, Bro89f, Bru72, Bru86, Bry74, Bry76, Bry77, Bry78a, Bry78b, Bry78c, Bry83, Buc77, Buc69, Buc71, Buc73a, Buc75, Buc85a, Buc85b, Buc66, Buc68, Buc89a, Buc73b, Bur80, Bur74a, Bur74b, Bur71, Bus83, Bus85, Bus74, Byn72, Byn73, Byn74, Byn78, Byn80, Byn87, Byn88, Cah88, Cal63, Cal77, Can74, Can75, Can77, Can80b, Can80c, Can84, Can86, Cap81b, Car67a, Car69, Car70, Car84, Car74, Cas82, Cas85, Cha65, Cha77a, Cha80b, Cha77b, Cha80c, Che78, Che85, Chi64, Chi65b, Chi65c, Chi65a, Chi68, Chi71, Cla86a, Cla66, Cla67, Cla68, Cla70, Cla71, Cla86b, Clo80, Coc75, Col81a, Col83, Col76, Col87b, Col88a, Col88b, Col81b, Con86, Coo81, Coo83, Coo89a, Coo84a, Cor81, Cow74, Cow68, Cri66, Cri71, Cro66b, Cro67]. Review [Cro68a, Cro68b, Cro69, Cro70a, Cro70b, Cro71, Cro79, Cro85, Cro87a, Cro88, Cro89a, Cro89b, Cro72, Cun78, Cun83a, Cun83b, Cun88, Cun89, Dau75, Dav71, Dav81, Dea83a, Dea83b, Dea86a, Dea86b, Dea86c, Deb66, Des88, Din63a, Din64, Din77, Dol69, Dol70a, Dol70b, Dol71a, Dol71b, Dol73a, Dol73b, Dol74, Dol83, Dor75, Dor79, Dor80, Dou84, Dra86, Dub72, Duc84, Duc88, Duf81, Duf83, Duf75, Dug77, Dun69a, Dun69b, Dun86, Dun87, Dur77, Dur79a, Dur80a, Dur81, Dur82, Dur83, Dyb78, Eag78, Ear70, Ear87, Eas86, Ede65, Edg84, Edg87a, Edg87c, Edg87b, Edw78, Ege89, Eme87, Eva89, Eyl68, Fal89, FF66, Far67, Far68a, Far69, Far71a, Far71b, Far76, Far77, Fau89, Fee77, Fie81, Fie83, Fie86, Fie89, Fig88, Fig77a, Fig77b, Fig78, Fin68, Fis74, Fis78]. Review [Fis81a, Fis81b, Fis83b, Fis83a, Fis83c, For66, For67b, For68, For69a, For69b, For71b, For71a, For72, For74a, For80, For85a, For87a, For82, For87b, For87c, Fox68b, Fox69a, Fox69b, Fox70, Fox71a, Fox71b, Fox74, Fox79, Fox82, Fox87b, Fra87a, Fra87b, Fra87c, Fre70, Fre74, Fre78, Fre81b, Fre86, Fri74a, Fri87, Gab83b, Gab83a, Gab75, Gab83c, Gau83, Ged71, Gel76, Geo66, Geo78b, Geo87, Gid82, Gil88, Gil84, Gin74, Gin84, Gli83b, Gli84, Gol83b, Gol87b, Gol88b, Goo75, Goo81a, Goo81b, Goo82, Goo83a, Goo85a, Goo85b, Goo88a, Goo68, Goo89b, Got66, Got67, Got71, Gou82, Gou84a, Gou85, Gou84b, Gow71, Gow73, Gow78, Gow69, Gra81, GG70, GG71, GG72, GG74a, GG75a, GG75b, GG75c, GG77a, GG77b, GG81, GG84a, GG84b]. Review [GG85, GG87, GG89, Gra83b, Gra83a, Gra84c, Gra85a, Gra85b, Gra86a, Gra86b, Gra87b, Gre63, Gre71, Gre88, Grm69, Gua89, Gue68, Gur78, Hac74, Hac75, Hac76, Hac77, Hac82b, Hac82a, Hac88a, Hac88b, Hac89, Had63, Hah70, Hah73, Hai77, Hai80, Hai82, Hal65a, Hal66a, Hal68c, Hal68d, Hal68b, Hal69a, Hal71, Hal72b, Hal72a, Hal77a, Hal77b, Hal80a, Hal80c, Hal84c, Hal84d, Hal84a, Hal85a, Hal85b, Hal85d, Hal85c, Hal86a, Hal86e, Hal86b, Hal86c, Hal86d, Hal87a, Hal87b, Hal88a, Hal88b, Ham84, Han85a, Han70, Han76, Har77, Har84a, Har75, Har89a, Har76, Har83a, Har86, Har79a, Har79b, Har81, Har82d, Har83b, Har87b, Has72, Hat87, Hea87b, Hea64, Hei68, Hei75, Hei71a, Hei71b, Hei73, Hei74a, Hei74b, Hei76a, Hei76b, Hei77, Hen81, Hen82c, Hen82b, Hen82a, Hen83a, Hen84a]. Review [Hen84c, Hen84b, Hen84d, Hen84e, Hen85a, Hen86a, Hen87a, Hen87b, Hen88a, Hen88b, Hen88d, Hen88c, Hen89a, Hen89b, Hen89c, Hen89d, Hen89e, Hen83c, Hen83d, Hen84j, Hen84k, Hen84l, Hen84m, Hen85b, Hen85c, Hen86e, Hen87c, Hen88e, Hen89g, Her73, Hes74, Hew66, Hig65, Hil71a, Hil71b, Hil75, Hil86b, Hil86a, Hil89b, Hil89a, Hoc86, Hoc88a, Hoc88b, Hod70, HS83, HS86a, HS86b, Hol71, HS88, Hom74, Hom80, Hom86a, Hom86b, Hop72, Hop73, Hop75, Hos67, Huf85, Huf86, Hun89a, Hun84, Hun86, Hut80, Hut84, Hut88, Hut89, Ilt77, Ink87, Ink88, Ins86, Irw82, Irw83a, Irw83b, Jac89a, Jac89b, Jac85, Jam86, Jam87a, Jam87b, Jar83, Jar84b, Jar84a, Jen89, Jev67, Joh79, Joh81, Joh83, Joh78, Jon87, Jor78, Jor79, Jor83b, Jor85a, Jor85b, Jor88, Jos67, Jud88, Kar76, Kau84]. Review [Kee88, Kee89, Kei84, Kei86, Kei89, Kel71, Kel68, Kel70, Kel73, Kel75, Kel76, Kel78, Kel84, Kel85a, Kel85b, Kil84b, Kil84a, Kni65, Kni67, Kni69b, Kni69a, Kni71c, Kni71b, Kni72, Kni74, Kni75, Kni76, Kni80a, Kni80b, Kni81, Kni83a, Kni83b, Kni83c, Kni84a, Kni84b, Kni85, Kni86b, Kni86c, Kni86d, Kni88a, Kni88b, Kni89b, Kni89c, Kni89d, Koh78, Kra84a, Kra86, Kri83, Kub70, Kuh68, Lar74, Lau68, Law65, Law80a, Law80b, Law81, Law83, Law87, Law88, Law89, Lay74, Lay85, Lef78, Les83, Lev73, Lev80, Lew85, Lie85, Lie86, Lin86, Lin89b, Lin74, Liv86, Loh75, Lor73, LM74, Lor75, Lor77, Lud80, Mac78, Mac80, Mac85b, Mac72, Mac79, Mac88a, Mac89, Mah89, Mai85, Mar74, Mar73, Mar89, Mau77, Mau87, May66, May67, May69, McC77a, McC77b]. Review [McC80, McC83, McC69, MdLD86, McG65, McG66, McG68, McG69, McG70a, McG71a, McG71b, McL86, McN83, McV75, Mea71, Mea84, Men67, Men70, Men74, Mer78, Von88, Mil87, Mol69, Mol71, Mol72a, Mol72b, Mol74, Mol75, Mol78a, Mol78b, Mol79, Mol81, Mol84b, Mol85, Mol86, Mol88, Mol89, Mon78, Moo84, Moo85, Mor87a, Mor84b, Mor69a, Mor71, Mor72a, Mor72b, Mor73a, Mor73b, Mor74a, Mor74b, Mor75a, Mor75b, Mor75c, Mor76a, Mor77a, Mor77b, Mor77c, Mor78, Mor80, Mor81, Mor82a, Mor83a, Mor83b, Mor83c, Mor84c, Mor84d, Mor87b, Mor84f, Mor84g, Mor83d, Mor89c, Mos77a, Mos77b, Mül87, Mül89, Mus74, Nad83, Nak75, Nay75, Nay77, Nay80, Nay85, Nee85, Nee69, Ner89, Nev78, Nev80, Nev83, Nev84, Nic85, Nic88b, Nor68a, Nor68b, Nor69a, Nor70a, Nor71a]. Review [Nor71b, Nor72, Nor73a, Nor77, Nor80, Nor82, Nor83, Nor85a, Nor85b, Nun88, Nut82b, Nut83, Nut84, Nye86, OM75, Olb69, Olb70, Olb71a, Olb71b, Olb72, Olb82, Olb85, Ora73, Ora74, Ora76, Ora77, Ora80, Osb85, Osb89a, Osb89b, Osp74, Out80a, Out80b, Out82a, Out83, Out84a, Out84b, Out85, Out87, Out88a, Out88b, Pac70, Pac71, Pac72, Pag66, Pag67, Pag69, Pat86, Pat84, Pau76, Ped85, Pel85a, Pel85b, Pep71, Pep72, Pep86, Phi67, Phi69, Pic87a, Pic76, Pic80, Pic88a, Pic88b, Pin83, Pin87a, Pin87b, Pin88, Por89a, Por77, Por78a, Por78b, Por79a, Por79b, Por81, Por82, Por83a, Por84a, Por84b, Por85a, Por85b, Por85c, Por86a, Por87b, Por87a, Por87c, Por88, Por89d, Por89e, Por89b, Por89c, Por89g, Poy70, Pra70, Pum84, Pum87, Pum88, Pye76, Pye87, Ran74, Rap71]. Review [Rap75, Rap79, Rap86a, Rat65a, Rat66a, Rat66b, Rat68, Rat69, Rau75, Rav65, Rav66a, Rav66b, Rav67, Rav68, Rav69, Rav70, Rav71, Rav72, Rav73, Rav84, Rec86, Red84, Ree80, Ree85, Rei86, Ric88a, Ric89, Rid83, Ril85, Rob75, Rob80a, Rob87, Rob89a, Roc87, Roc88b, Rog83, Rog88a, Rog88b, RH86, RH88, Ron86, Ros79, Rud69, Rud70, Rud71, Rud80, Rud83, Rud88b, Rud89, Rup88, Rup89, Rus87a, Rus89a, Rus66, Rus67, Rus68, Rus69, Rus72a, Rus72b, Rus74, Rus77a, Rus77b, Rus77c, Rus79a, Rus79b, Rus85, Rus87b, Rya70, SG80, San87, Sar81, Sch81a, Sch83a, Sch84a, Sch86a, Sch86b, Sch87, Sch69a, Sch68a, Sch69b, Sch70, Sch71, Sch72, Sch73a, Sch73b, Sch74b, Sch74a, Sch75, Sch76, Sch78a, Sch78b, Sch79a, Sch80c, Sch80b, Sch80d].

Review [Sch81b, Sch83b, Sch85a, Sch78c, Sco67, Scr68a, Scr68b, Sec83, Sec84, Sec85, Sec88, Sei85, Ser88, Sey77, Sha78, Sha89, Sha73a, Sha77a, Sha79, Sha81, Sha83a, Sha86, Sha87, Sha77b, Sha73b, She85, She63, Sho87a, Sho86a, Sho88, Sho89, Sim84, Sin74, Sin76, Sin77, Sin78, Sin81, Sin84, Sme66, Sme68a, Sme69, Sme70, Sme71, Sme84a, Sme68b, Smi66, Smi71, Smi73a, Smi73b, Smi74b, Smi74a, Smi75b, Smi75a, Smi76c, Smi76a, Smi78b, Smi78a, Smi78c, Smi80a, Smi80b, Smi82a, Smi83, Smi85a, Smi85b, Smi85c, Smi85d, Smi87b, Smi87c, Smi87a, Smi89e, Smi89b, Sne74, Sne87, Sol83, Spa65, Ste62a, Ste62c, Ste62b, Ste70, Ste75, Sti89, Sti70, Str73, Str75, Stu71a, Stu78, Stu87, Sty68, Swa86, Swe76, Swi70, Swi71, Swi77, Swi67, Tan82, Tee84, Tei78, Tei83b, Tei83a, Tei84]. Review [Tei88, Tha66a, Tha67, Tha74, Tha83a, Tha83b, Tha84a, Tha84b, Tha85, Tha87, The87, Tho89, Tho76a, Tho76b, Til75a, Til76, Tre87a, Tre87b, Tre88, Tup82, Tur68, Tur69, Tur70b, Tur70c, Tur70d, Tur72a, Tur72b, Tur73d, Tur73c, Tur73a, Tur73b, Tur74a, Tur74c, Tur75, Tur77a, Tur78a, Tur78b, Tur78c, Tur81, Tur82, Tur83a, Tur83c, Tur87a, Tur88, Twe89, Tyl75, Und64a, Und64b, Und66c, Und66a, Und66b, Und75a, Und75b, Und79a, Und79b, Und80, Van86, Vea64, Ver89a, Ver89b, Vic88, Wal89, Wal74, Wal77, Wal81b, War88a, War88b, War89, Was69, Wat73, Wea78, Wea80, Wea83, Wea84, Web75, Web76, Web78, Web84, Web65a, Web65b, Web66a, Web66b, Web66c, Web67a, Web68a, Web68b, Web69b, Web69a, Web70, Web71a, Web71b, Web72, Web87, Wei83a, Wei85a, Wei85b, Wei86]. Review [Wel70, Wes74, Wes70, Whe89, Whi70a, Whi68a, Whi72a, Whi72b, Whi73, Whi75, Whi85a, Whi89a, Whi68b, Whi84, Whi85b, Whi88a, Whi89b, Whi68c, Whi77, Wig69, Wig74, Wig77, Wig78, Wil65a, Wil75, Wil65b, Wil68, Wil69, Wil77, Wil84e, Wil84d, Wil84b, Wil84c, Wil85a, Wil85b, Wil86a, Wil86b, Wil87c, Wil87b, Wil78, Wil89b, Wil89a, Win75, Woo80a, Woo80b, Woo80c, Woo83, Woo84, Woo89, Woo67, Woo73, Woo85, Wor88, Wuj85, Yat72, Yeo88, You66, You83, Yox79, Yox82b, Yox82a, Yox83a, Yox83b, Gar87]. Reviewed [Ano83l, Ano84l, Ano85l, Ano86l, Ano87q, Ano88m, Ano89m]. Reviews [Ait86, Ano66h, Ben84, Blo84, BM76, Bud84, Col72, Fel89, Fur85, Gab88, Hal84e, Har84b, Hen83b, Hen84f, Hen84g, Hen86b, Hod89a, Jor87, Kni88c, Kni89e, Law85, McC85b, Mol84c, Mor72c, Mor89a, O’H86, Pic84a, Rat65b, Rud84, Rud86, Rup85, Sch84b, Sch84c, Smi89c, Stu85, Und64c, Und64d, Web65a, Whi65, Wil87d, WM85, Yox83a]. Revised [Ora74]. Revisions [Lov75]. Revisited [Per89b, Sch79c, Frä90, Jos97]. revival [Deg91]. revived [O’M89]. Revolution [Ait86, BA91, Bar81, Bec91a, Bos83, Buc71, Bus74, Coh85, Cro87b, Cro89a, For74a, Gou85, Hac89, Hal80c, Hal85b, Hal87b, Hen84a, Hen89c, Hen83d, Hew66, Hil71a, Jor78, McC80, Nay80, PT86, Rav72, Sch74b, Sec84, Sha81, Smi71, Smi80b, Wei85b, Wil86a, Woo80b, Woo80c, Wor88, Gua89, Ano87r, B^{+}95, Ath90, Ben86, BR95, Bow88c, Bro82b, Bru88b, Coh84, Coh94, Cro90b, Don88, Don93, Edg91, Eis83, Eis93, FJ93, Fri81, Gle82, Grm90, GGSGB94, Hal83a, HL86, LeG88, LW90, LP04, Mas98, Max84, New94, PLP92, PT92, Ros84a, Smi82a, Ste90b, B^{+}95, BV93, Grm90, Mac88c, NS02, Ano87r, CAMR87, CLR87, NS02]. Revolutionary [Ede65, Fox71b, Cro85, Fra78a, Gab83a, Lan87, Out84c]. Revolutionnaires [Out88b, Lan87]. Revolutions [Blo82b, Kni81, Nev78, Rus79b, BV93, Tur74a]. Revue [BG80, SG80, Sme84a, GG70, Nye76]. Reweaving [Edg87a, Sta85]. Rewriting [Bus83, ZH06]. Reynolds [Bur71]. Rezeption [Jun89]. Rhetoric [Hen88a, Pin87a, Mos93, PS91, SY86, Slo85, Ste92, Wil87a]. rhetorical [NS02]. rhétorique [NS02]. Riccardo [Smi82a]. Rice [WV92]. ricerche [Ern91, Sch79b]. Rich [Rat68]. Richard [Bod84, Buc71, Can77, Dol70a, Dol74, Dyb78, Hig65, Jor79, May69, Mol78b, Mos77a, Pic84a, Pic88a, Rog83, Sch68a, Whi72b, Whi85a, Woo85, Wri87, Bro77, Des85, GT92, GGSGB94, Nee85, Ric87a, Sho86b, Sme88, Whi85b]. Richards [Mac79, Und66a]. Richardson [Dra86, Hen82c, Law88, Ash85]. Richter [Cap81b]. Riddick [But66]. Ridicule [Por77]. Ridwan [Fre86]. Riemann [Dub72, Ser88, Gra85c]. Righini [Tur72b]. Right [Ait71]. Rights [Ric87a]. rigorous [Gra83a]. Rijksmuseum [Bry77]. Rijn [Hom80]. riksmuseet [Ber95]. Riley [RT95]. Rimmer [RD95]. Rinascimento [Sch81b, Ern91]. Ring [Gow69]. Ripa [Gua89]. Rise [Asq80, Buc85c, Cro87a, Fig78, Hen86b, Hen88c, Hil89a, Hod89a, Hoo87, Hut81, Mor74b, Mus74, Pay74, Smi74a, Wes74, Wil75, Aas90, Buc89b, CDS90, Cor96, Eng83, GR91, Har98b, Huf93, Huf95, Law93, MR82f, MR82g, MR87a, Mor84d, Por86b, Rob89b, Sch90, Ste92, TW95, Wes88]. Risse [Fig77a, Wea79]. Ritterbush [Und66b]. ritual [Wil84d, Wyn82]. Rivalry [Pag76, Man88]. River [All84a, BB82]. Rivers [Cla68, Web84]. Road [Aus86, Byn72, Sec84, Sey77, Gle82]. Roads [Aus86, FS84]. Robert [Ait73, Ait86, Bar77, Bro80b, Bru72, Buc66, Can80b, Car74, Clo80, Col81b, Cri71, Dau75, Dun69b, Dur82, Gel76, Gil84, Gol83b, Hen81, Hig65, Hod70, Kni69b, Law80a, May67, McG68, McG71a, Mea71, Mol79, Mor72b, Mor75c, Nay80, Out82a, Rud89, Sch84b, Sch72, Sch78c, Sha83a, Smi87a, Ste81, Tei84, Tup82, Tur68, Web66b, Web71a, Web72, Wil65b, Wil77, Woo80b, Yox82b, Bad89, Bow87a, Bow88b, BKH90, BHW94, BH84b, Bur86, BEoCF92, Can93a, ECL92, Gra81, Hal80c, Hey83, Hoo97, HS89, Kee88, Mül87, Ogi75, Old72, Ost89, Rus87a, Rus85, SG90, Tay89, Wil90b, Wil89a]. Roberto [Mor89c]. Roberts [McL86]. Robertson [Bad83]. Robin [Bra89a]. Robinson [Bry83, Mor71, Smi71, Dau95, Wil90b]. Robison [Smi79]. Roche [Por89e]. Rochelle [Bro89c, Gol88b, Jac89b]. Rochot [Hal77a]. Rocke [Bro85a]. Rocketry [Mac88a]. Rockey [Wei83a]. Rocks [Bar75, Ham84, Her83]. Roderick [Sta89a, Pag76]. Rodopi [Hal84a, Wil87d]. Roe [Duc88, Edg87b, Les83]. Roger [Bra84, Bro76, Sch85a, BL83, Bus85, Gow83, GG89, Law87, Lin82]. Role [Buc69, Can70, Fan83, Gol70, Rav73, Tan62, Tre79, All82, AB89, Dob91]. roles [RCRC98]. Rolf [Sch83b, Rus89a]. Roller [Smi74b]. Rolt [Law65]. Rom [Goo83a]. Roman [Ant63, Hal86b, Has72, Hil71b, Smi75a, Smi94b, Fur89, Ole84, Tyl64]. romance [Mer89]. Romano [Dea86c, Wal84]. Romantic [Jor80, Bal84, MF92, Vig82]. Romanticism [PB94, CJ90]. Romantische [Ree85]. Rome [Sch81b, Wea78, Has72, Mid75]. Ron [Tha84b]. Ronald [Bra84, Dol74, For82, Hac76, Rus87b, Tur83a, Und66c, Wea79, Bow89c, Cun88, Lie85]. Ronan [Hen84c, Nay77, Pum87]. Ronchi [Stu71a]. ronde [Ano82l]. room [Cha84]. Roots [Dur80b, Mac85b, Cra93, Whe83]. Roper [Hal68d]. Rose [For87b, Kel78]. Rosemary [Hei73]. Rosen [Ait86, Arm66, Hal69a, Mor87a]. Rosenberg [Web78]. Rosenthal [Buc69]. Roseveare [Wil84c]. Rosicrucian [Tur73a]. Rösing [Sha79]. Roskilde [Chr93]. Ross [Mor73a]. Rossi [Rap86a]. Rossiter [Far76]. Roswitha [Tei83a]. Rothenberg [Bud84, Ink88]. Rothenfelder [Wes70].

Rothshuh [Fig77a]. Rothstein [Pic76]. Rotterdam [Mül89, Smi89b]. roubles [Geo78b]. Rouge [Col88b]. Rousseau [Bur74a, Sch01]. Routledge [Dol83, Fis83a, For87b, Goo75, Hal68d, Har84b, Hea87b, Hen85b, Joh78, Kel71, Law88, Mar89, Mor75b, Mor77c, Ric88a, Rog88b, Sha87, Tur73a]. roving [Gid82]. Row [FF66, Swi67]. Rowan [Cow68, Mor83b, O’D83, O’H86]. Rowland [Bow89a]. Roy [Bur80, Dur83, Gol87b, Mor81, Rap79, Hal87b, Kni89d, Lay85, Tha83a, Tha85]. Royal [All82, Bro86a, Bro78a, Bro78b, Bry78a, Bry83, Can74, Cha80c, Col76, Cow68, Fis81a, For68, For80, Gou84a, Hac76, Hal68d, Ham84, Hei73, Hen82a, Hen83c, Hen89g, Hom86b, Jam87a, Kni71b, Mil83, Ora73, Pat86, Rob87, Smi89c, Tha83b, Wil87b, Woo67, FFLW97, Ped92b, AL87, Bad62, BKH90, DTS^{+}87, For73, Frä89, Gle91, Gun84, Hac84, Hal66b, Hal84b, Hal91, Hop76a, Hop76b, Hun82, Hun89b, HH97, Mar62b, Mar64, MJ84, Rob80b, STT87, Sha74, Ste67, Str87, Tay87, TST87, WH91, Woo80d]. royale [Out80a, Str87]. Rs [Hil75]. Rs.200 [Wuj85]. rubber [Mor89b]. Rud [Dor79]. Rudbeck [Eri94]. Rude [Rus88b]. Rüdenberg [GG75c]. Rudnick [Olb69]. Rudwick [Mor87b, Rap75, Sch83a]. Ruestow [Hac75]. ruled [Sch90]. Rumford [Fox69b, Fox70, Swi71, Gol77]. Rupert [Bro78a, Bro78b, Hen84a, Hop73, Hop75, Sch81a, Sch84c, Scr68b, Web66c, Woo80c, Dew63]. Rupke [Rud84]. Ruprecht [Hom74, Wil84b]. rural [Pri83]. Ruse [Lin74, Ver89b]. Russak [Hen81]. Russel [Bro84b, Bro85d, Cas82, Dur79b, Hug89, Smi72]. Russell [Bad71, Bro86a, Bro72a, Duf83, Duf75, Ear87, Geo87, Hei74a, Ins86, Joh79, Kni72, Mor77b, Mus74, Kel85a]. Russell-Dear [Joh79]. Russett [Byn78]. Russia [Coo92, Hal71]. Russian [Bro89b, Kel70, Tup82, Wei85b, Fri81, Gra98, KBRS90]. Russification [Sch85b]. Rutgers [Bow87a, Gil88, Por87a]. Ruth [Dol73a, Har79a, Law88]. Rutherford [Bro86c, Kni75, Bad85, BS79, Stu85, Wil83, Hei71b, Sin81].

S [Ben79, Blo79, Bow87a, Buc73a, Byn72, Cla67, Cun78, Dol74, Dug77, Dur80a, Edg87c, Fie86, Fre74, Fre86, Goo82, GG84a, Hal68b, Har84b, Har89a, Har76, Hen82c, Kar76, Kni69a, Kra86, Law80b, Lor73, Mac85b, Mac72, Mar74, May67, Men67, Mor84b, Ora74, Osb85, Poy70, Rap75, Ric88a, Rog83, Rud83, Rud89, Rus72a, Sch83a, Sch70, Sch78b, Sch79a, Smi80b, Smi83, Str73, Tur69, Whi72b, Whi89b, Whi68c, Woo80b, Woo84, Woo85, Wuj85, Yox83a]. S. [Dea83b, GG87, Gre88, Huf85, Pic88a, Rup89, Tha83b, Wal91, Wat84]. S.J [Str73]. Sa [Cro79, Bec91b]. saber [Hue88]. Sabetai [Mol81]. Sabra [Lau68]. Sadi [Hut80, Ler85, Men81, Smi87a]. Sadoun [Cro79, Sme84a]. Sadoun-Goupil [Cro79, Sme84a]. saga [Wri87]. Sage [Goo81a, Irw82, Irw83a, Sha79]. Saggi [Sch73a, Mid69]. Said [Sar81]. Saint [Hai78]. Saint-Simon [Hai78]. Salamander [Por85b]. Sale [Wal77, Por89f]. Sales [All77]. Salitter [PW89b]. Sally [Mor77a, Por85b]. Salomon [Out80a]. Salomon-Bayet [Out80a]. Salt [Clo80]. salud [Lan89]. Samarqandi [Ear70]. Samba [Wuj85]. Sammlung [Rav65]. Sämtliche [LM23, LS76]. Samuel [Car67a]. San [All82, Byn78, Mac88a, Sch78b, Sch79a, Yox83a, Yox83b]. Sanborn [Fox69b, Fox70, Gur78, Swi71]. Sandemanian [Can91]. Sandemanians [Can89]. Sander [Por87c]. Sanders [GG77b]. Sanderson [Mor75b]. Sandford [Smi75b]. Sandra [Des88]. Sandwich [Twy88]. sanitarias [Fre90]. sanitary [Fre90]. Sanitat [Kni88c]. Sanskrit [Wuj85]. Santiago [Gli84]. Santillana [Hal68c, Mol69]. sapere [Olm92]. Sarjeant [Rud83, Rud89]. Sarton [Phi67]. satellite [Mac90]. Satire [Por79b]. satisfyed [BS91]. Saul [Cla68]. Savage [Kni88a, Kuk91, Kuk93, Moy86]. Savant [Men74]. savantes [Ano93]. Savants [Out88b, Lan87, PP95]. Saviours [Nic88b, Söd86]. Savoir [Sho89]. savoirs [Bea91, G^{+}09]. Scaglia [Kel73]. scale [GH92]. Scaling [Smi94c]. Scán [Ste62a]. Scandal [GG74a]. Scandinavia [RH89]. Scarecrow [Cha65, Cha78a]. Scenes [Rud92]. Schafer [Hoc86]. Schaffer [Hen84c, Jon87, SSH85]. Schaffner [Mar74]. Scharlau [Gra86b]. Schegel [May69]. Scheme [Rud63]. Scherz [Osp74]. Scheutz [Lin90]. Schiller [Bur74b, Hai80, Jor83b]. Schiøler [Smi75a]. schism [Blo84, Pop82]. Schloss [Wey92]. Schmid [Tei83a]. Schmidt [For67b, Sch83b]. Schmidt-Schönbeck [For67b]. Schmitt [Hen86f, Hut84, Hut89, Mol72b, Mor83a]. Schneer [Por81]. Schneider [Sim84]. Schofield [Dyb78, Kni69b, McG71a, Osb85, Wea84, Wil65b, SG90]. scholar [ES90]. scholars [AS96, FJ93, ZH92]. Scholastic [Ait86, Gra84b, Law93]. Scholasticism [Col81a]. Schönbeck [For67b]. Schone [Phi67]. Schonland [Bro69]. School [Cun78, Dor80, Lar74, Löw90, Mor80, Rob75, Rud84, KH86, Lan87, Mol68, Rup83, Lan87]. Schools [Wal89, Whi72a, GH93, Dea88]. Schöpfung [Pat84]. Schriften [LM23, LS76]. Schrifttum [Wes70]. Schrödinger [Hen88c, Hen89b, Kil87, MR87a, Moo89b]. Schubert [Kni83a]. Schufle [Dun69a]. Schullian [Und75b]. Schuster [Pin87a]. Schütt [Mau87]. Schwartz [Dol73a]. Sci [Wor79]. Science [Agr83b, All82, AL87, Ano63m, Ano77m, Ano82p, Ano85j, Ano86i, Ano87m, Ano87s, Arm62, Arm64, Asq80, B.69, Bad71, Ban88, Bar77, Bar81, Bar83, Bar76a, Bau75, Bec86, Ben91, Ben79, BC89, Bib74, Blo79, Blo82b, Bod84, Bro73a, BM76, Bro80b, Bro76, Bro89c, Bro75, Bro85e, Bro89f, Buc73a, Buc85b, Buc63, Buc68, Bud84, Bur80, BK75, Bur71, But92, Cal77, Cam82, Can80b, Car66, Cha78b, Cha77a, Cha90a, Cha90b, CB^{+}90, Cha77b, Cha80c, Chi64, Cla79, Cla68, CM90, Col81a, Col76, Con86, Coo83, Coo89a, Coo84a, Cor81, Cor88b, Cri66, Cro90a, Cro96, Cro68a, Cro77, Cro87a, Cro87b, Cro89a, Cro92, Cro72, Dau98, Dav99, Day85, Dea86b, Dea86c, Din63a, Din77, Dol69, Dol70a, Dol71b, Dol73a, Dol73b, Dou88, Duc84, Duc88, Duf75, Dun86, Dyb78, Eag77]. Science [Edg87c, Eme87, Far76, Fee77, Fie83, Fie89, Fin68, FB75, Fis83b, For67b, Fox71b, Fox74, Fox79, Fox84, Frä89, Fra87a, Fre74, Fur85, Gau83, Gil84, Gin84, Gli84, Gol88b, Gol92b, Goo81b, Goo85a, Goo88a, Goo68, Goo83b, Got86b, Got67, Gou68, Gou82, Gou84a, Gou85, Gow71, GR88, Gra81, Gra87a, GG79b, GG81, GG84b, GG85, Gur78, Hac76, Hac82b, Hah73, Hai82, Hal65b, Hal69b, Hal80a, Hal84d, Hal86c, Hal88b, Han85a, Han85b, Har77, Har69, Har67, Har81, Hay74, Hen80, Hen83b, Hen84a, Hen84c, Hen84d, Hen88b, Hen89c, Hen88f, Hen89g, Hes74, Het88, Hil75, Hil89b, Hoc86, Hod71, HAB92, Hol78, Hom83b, Hoo87, Hul88, Hun95, Ing84, Ink79, Ink81, Ins86, Jac89b, Jam87a, Jam88, Jef64, Joh78, Jor80, Jor87, Jor88, Kar76, Kee88, Kee89, Kei84]. Science [Kei86, Kni69a, Kni71b, Kni74, Kni75, Kni80a, Kni81, Kni85, Kni86b, Kni86d, Kni89b, Kni89c, Kni89d, Koh78, KK85, Kra86, KN88, Kum91, Kum95, Las81, Law80a, Law87, Lay65, Lay85, Lef78, Lem77, Lev84, Lev88a, LS90, Lev93, Llo83, Lor73, Lor75, Lug79, Mac72, Mac89, Mar73, May67, May69, MJ84, McC77b, McC85a, McG65, McL86, Mid75, Mol69, Mol71, Mol74, Mol79, Mol84c, Mor71, Mor72b, Mor72d, Mor74b, Mor75a, Mor75c, Mor77a, Mor77c, Mor84c, Mor85, Mor84g, Mor89a, Mos77a, Mos77b, Mul79, Nak75, Nak91, Nay85, Nee69, Nee78, NLH^{+}84, NhTWL87, Ner89, Nev84, Nor69b, Nor71a, Nor71b, Nor77, Nor80, Nor82, Nye76, Old72, Old77, Ols90, Ora74, Ora76, Ora80, OE^{+}90, Out80b, Out80d, Out82b, Out83, Out87, Pac72, Pat86, Pau76, Ped85, Pep71, PS78b]. Science [Phi67, Pic84a, Pic87b, Pic76, PP87, Por76b, Por78a, Por79b, Por85b, Por87a, Pum87, Pye87, Rab78, Rap79, Rav67, Rav72, Rec86, Rei68, RR81, Rei86, Ric89, Rid83, Rob75, Rob80a, Rob80b, Rob89a, Rog88a, RH88, Rus76a, Rus72a, Rus77a, Rus77b, Rus83a, Rus83b, Rus87b, San88, San91, Sar81, Sch84a, Sch70, Sch75, Sch76, Sch79c, Sec84, Sec85, Sey77, Sha89, Sha73a, Sha74, Sha79, Sha83a, She63, Sho87a, Sho88, Sho89, Sho93a, Sho93b, Sin81, Sme69, Smi66, Smi71, Smi74b, Smi74a, Smi78b, Smi80a, Smi80b, Sne87, Stu95, Sut95, Swi70, Swi77, Swi67, Tor88, Tre79, Tre88, Tup82, Tur68, Tur70b, Tur72a, Tur74a, Tur78b, Tur78c, Tyl75, Ubb63, Ver89a, Vic88, Wal74, Wal90, Wea79, Wea84, Web78, Web84, Web68b, Web71b, Web72, Wei80, Wer75]. Science [Wes74, Whi72b, Whi89a, Whi89b, Wig74, Wil65b, Wil69, Wil84d, Wil84b, Wil86a, Wil78, Win75, WG81, Woo84, Woo73, Woo80e, Wuj85, YK94, vHC95, Aas90, AH85, Ait85a, AH93, Ali86, Alt87, Ano82n, Ano82m, Ano93, App93, Ara91, Atr86, A^{+}87, Baa90, BA91, BM82, Bee96, Ben82a, Bia93, Blo84, BMLC84, BW90, BM84, BC98, Bru87a, Bru88b, Bur83, Bur87b, Can91, Chi65c, Cla99a, Cla99b, Cla95a, Cla9 , Cla2 , Cla82, Clu88, CDS90, Coh84, Coh85, CP93, Coo84b, Cra84, Cra92, Cro85, Doe96, Dol83, Dol96, Don93, Dor91, Dra81, Dur82, Dur83, Edg91, EC87, Eri94, Fis91a, Fog92, Frä90, Fre82, FC96, Fuj96, Gab83b, GH92, Gei95, Gie88, G^{+}89, Gil83, Gje84, Gle82, Gol91, Gol79, Goo88b, Got86a, Got92, Gow78, Gra98]. science [Gre92b, GGC90, Hal83a, Ham90, HA87, Har84a, Har96, Har98b, Har82d, Hei86, Hen83a, Hen84k, Hen84l, Hey83, Hod92, Hol86, Hom88, Hoo97, Huf91, Huf93, Huf95, HJF89, Hun89b, HH97, II90, IW96, JK84, Jen91, Jos97, Joy87, KC95, Kit92, Kni83a, Kni83b, Kni86a, Kni88d, Kni89a, Kni92a, Kni92b, Kra87, KAZ92, LaF82, LSS90, Lan91, Lau87a, Law93, Lie87, Lin92, Liv87, Llo91, Lan87, LN95, Mac82, MR88, MW94, Man88, Max84, McG84, Mei91, MF92, MH88, Mor83c, Mor84e, Mos93, NR86, Ner87, NW82b, NW82a, Nol88, Ogi86, O^{+}90, Old86, Osb94, Out80a, Out82a, Out84a, Out84c, Pag82, Pai82, Pan91, Pat83a, Pau84, PS91, PR88, Pet90, Pet84, Pin83, PB94, PB83, Por95, Pow96, RM96, Rav84, Red84, Red87, Rei85, Ron83, Ros91]. science [Rot93, Rud88a, Rus89b, SG90, Sep88, Sha83b, SP88, SW89, Sho96, SY96, Sht96, Sim88b, Smi89d, Smi94b, Smi94a, Smi94c, Sta84a, Sta94, Ste92, Van93, Vic87, Wal84, WV92, WM85, Wil91a, Wil77, Wil93, Wil91b, Yeo93, Zam92, ZH06, dGB^{+}89, Ano62h, Ano62k, Ano68j, Ano84j, Ant64, B^{+}82, Bar83, Blo82b, Bod84, Bro84a, Can84, CC66, CH67, CE73, Cun78, Dol83, Fie83, Fis83c, G.62, Gau83, Gil84, Goo83a, Gou82, Hai82, Har83a, Har82d, Har83b, Hen82a, Hen84g, Hen89e, Irw82, Irw83b, Jar84a, Mar73, McC81, Mol84b, Moo84, Mor83c, Mor84d, Mus74, Nor83, N^{+}88, Ora72, Pat84, Phi69, Por78b, Por83a, Pye76, Rav66b, Rav73, Rav84, Rob75, San87, Sha73a, Sin84, Und63, Was69, Wea84, Whi84, Wil68, Wil78, Woo73, Wuj85, You83]. Science [Yox82a, LN86, Ste91b]. Science-Physical [Kni71b]. Sciences [Agr83a, Agr83b, AZW87, Ant64, Bad71, Bad83, Bow83a, Bra84, Bro71a, Bro83b, Bro83a, Bur80, Bus83, Cha80c, Cla67, Cla71, Cri71, Cro70b, Cro88, Cun83b, Dea83b, Dol70a, Dol71a, Fin68, For71b, Fox87b, Fur85, Gab83b, Gab83a, Gel76, Geo66, GG84a, Gra83b, Gra83a, Gra84c, Hal84e, Hei71b, Hen83a, Hen83b, Hen84a, Hen84c, Hen84b, Hen83d, Hen84l, Hol71, Irw83a, Joh83, Jor83b, Kni71b, Kni83c, Kri83, Law83, Les83, LS90, McK78, Mea71, ME81, Mol78a, Mol81, Mol88, Mor77b, Mor83b, Mor83c, Nay75, Nev78, Nor70b, Nut84, Olb71a, Olb82, Osb89b, Out82a, Out83, Ped92b, Pic84a, Rid83, RH86, Rud83, Rup89, Sch84c, Sch71, Sch85b, Sec83, Sha89, Smi66, Smi76c, Ste62c, Swe76, Tee84, Tei83b, Tei83a, Tha83a, Tha83b, Tur83c, Und66a, Vuc84, Web66a, Web67a]. Sciences [Wei83a, Whi70b, Woo83, You66, Yox83a, Yox83b, Ano93, A^{+}87, Bow92, Can88, Can84, Car88, CM93, CM90, Col83, Cro66b, CJ90, Dur80a, Fru82, Fru90, Gin84, Gow78, HJF89, JO81, Lan87, Liv85, Llo83, MMVZ90, MC^{+}93, Mor84b, NPH92, Out80a, Por83b, Pye93, Rit90, Ros94, Sch75, Sho93b, Smi92, Smi97, Tha85, WHS92, Baa90, Cro92, Frä89, Str87, Stu95]. scientiarum [Mer78]. Scientific [Ait86, Ano69h, Ano93, Bec91a, Bro75, Bry74, Bry76, Buc85c, Can77, CW87, Cha65, Cha78a, Cha80b, Che78, Col87b, Col81b, Cro70a, Dea86a, Dol70b, Dol73a, Dol74, Dur79b, EMW82, Eme87, EC87, Far71a, Fre81b, Fri74a, Goo81a, Goo88a, Gou85, Gow74, Gur78, Hac76, Hal65c, Hal80c, Hal86d, Har75, Hen86a, Hen83d, Hoc86, Ink88, Kal95b, Kni69b, Kni86b, Kra84a, Lay65, Mai88, McC80, McC85b, Mol74, Mor77a, Mor77c, Mor80, Mor87b, Nay80, Ner89, Nev78, Olb89, Pel85b, Pic84a, Pic87a, Pin87a, PP87, Por85b, PW89b, Rav67, Rob89a, Rud71, Sha81, Sho88, Sim88b, Sin74, Smi74c, Smi89c, Sne74, Tho76b, Tur70b, Tur74a, Tur77a, Tur87a, Tur87b, Tur90, Wei89, Whi89a, Wil87d, Woo80b, Woo80c, ABR93, Ano90, AS96, Bea91, Ben91, Ber95, B^{+}89, Bru88b, Buc94]. scientific [CF86, CS92, Coh84, Coh94, CHG91, Dau98, Dav99, De 85, Dea91, Deb91, Edg91, Eme84, For86, Fox82, Fru90, Gau83, Gje84, Goo83a, Goo90, GR91, Gou84a, Gut89, GGSGB94, Har02, Hei93, Hen82c, HRP90, HP85, HAB92, Hol85b, Hui92, Hun82, Hun84, Irw82, Kni86a, Kni88d, Kni89a, Kra90, LeG88, LHB04, LS76, LHH91, LW90, Lin92, McC85a, Mor84h, Ner87, New94, Nor83, Nut83, Out84b, PHvMW93, PS91, PB82, PB83, PT92, yQ85, Red87, Ros84a, Rud85, SY86, She91, SY96, Sla82, Smi82a, Sta89a, Sta89b, Tur83b, Vic84, Woo82, Zan86, vL76, Bry74, Bry77, Bry78c, Cha80c, Chi65c, Chi68, Chi71, Fox68b, Hal68b, Har79b, Men67, PvM85, RR87, Rya70, Sme69, Tur68, Tur69, Tur70b, Tur72b, Tur77a, Tur78c]. scientific-philosophical [Hei93]. scientifica [Lev90]. Scientifique [Ben89, Hal65a, Hal85c, Hut80, McC80, McC83, Sch84c, Sho89, Ano82l, Dug77]. Scientifiques [Osb89a, Ano90, Ano93, Bea91, B^{+}89, Nye76]. Scientist [BM76, Can91, Cha75, Dew63, Fox74, Fri74a, Kel71, Loh75, Men70, Ora77, Rav73, Sta89a, Wei86, Yox81, Bäu84, Fre92, Gol92a, Hay94, MR89, PS82, PCA91, Ros90, vLP82]. Scientists [Hen81, Hoc88b, Irw83b, Jen70, Kei86, Kni88a, Mac80, McG84, Mor89a, Rei87, Rob87, RW72, Stu86, Wal77, Wat73, Wat80, Wig69, Wil69, Yox82a, Ano93, AS96, Gli83b, Hal84b, Kuz87, MP89, M^{+}89, Mil90, Moy86, Swa88, PP95]. Scienza [Sch73a, Tur72b, Ara91, Lan91, Out80b, Sch81b]. scienze [GGC90]. scitur [SLT88]. Scolar [Can74]. Scoresby [Ora77]. Scot [Pag66]. Scotland [Aus86, Hil86b, Mor73a, Tha83b, FS84, For74b, Rud62, Sha84]. Scott [Cla66, For68, Pac70, Stu88, Jor88, Osb89b]. Scottich [Rud80]. Scottish [Can77, Fis81a, Hei76a, Jam87a, Moo84, Mor73a, Tha83b, Und80, Wes74, Woo84, Eme88, Mor72d]. Screen [Sho87c, Tur96]. Scriba [Whi68a]. Scriblerians [Fox88]. Scribner [Hei75, Hen81, Tur78c]. Scrolls [Lev90].

Scrope [Rud74]. Scrutiny [Kni85, Hom83b]. Scull [Wal81b]. sculptor [But96]. Scurvy [Llo63, Tei88, Car86]. se [ML89]. Sea [Lev90, Rap71, Wat73, Ben84, McC82, O’H86]. Seacraft [Lan86]. Seal [Pel85b]. seaman [How89]. Seamen [Wat80]. Seamless [Bud89]. Sean [Gab75]. Search [Bar76a, Jam88, Swa83, Ada89, Cre93, Deg91, Her90, Kil94, Twy88, Wea90]. Searle [Mac78]. seas [PR95]. Seattle [Gli83b, Mac72]. sec [Sch80c]. secolo [Ara91]. Second [Car69, Coc75, Cor78, Cro71, Cro72, Dol73a, EK91, Far71a, HJF89, Mad88, Por84b, Sme71, Web69a, Bru88b, SG80, Smi82a, HJF89, Tho74b]. seconde [SG80]. Secord [Tha87]. secret [Jar84b, KA81]. Secretary [Gun84]. Secrets [Ste75]. Section [Ano93]. Sects [Pic76]. Secular [Cun78, Pum89]. Secularization [Koh89, Sha78]. Sedgwick [Sec83, Spe82, Tom65]. Sedition [Wei80]. Seditious [Ink81, Ink79]. Seebeck [Tur77b]. Seeger [Rav68]. Seek [Hac84, Wil87b]. Segni [Rap86a]. Segrè [Buc73b]. seicento [Sch79b]. seiner [Adl87, Fig88]. Seite [Hei93]. Seiten [Von88]. SEK [HS86b]. Selbstzeugnisse [Rus72b]. Selected [Bro89e, Hei73, Kni69b, Phi69, PN95, Rus72a, Rus72b, Sch74a, Whi68b, Whi68c, AT87, Bru88c, CD94, DeV82, GG91, Hen83b, Hei74a]. Selection [Bro86e, Byn80, Rid82, Rus77b, Smi73b, Cro91a, Dem83, Gay92, KRS83, Sme66, Gay92]. Selections [Buc66, Car67a, Gel76]. Selective [Dol70b, Bud84, Rot93]. Self [Mor72b, HK83, Smi85d]. Self-Portrait [Mor72b, HK83, Smi85d]. selling [Fri83, Rus85]. semiconductor [Mor90b]. Seminar [Ree85, Woo80b, Ole91]. Semmelweis [Wei85a]. Senecas [Gro89]. Senecio [Hac82a]. Sense [Dur79a, McG65]. sensibility [BB92, Ben91]. Separate [Swi70]. separately [Bry78a]. Separation [Dav81]. Sepper [Jac89b]. September [Ben79, EK91, Hol85a, PCA91, Sme79, Sme88, WV92, Woo80c]. Serge [Ait68]. serial [Yox82b]. série [Sch84c]. Series [Far71a, Gou85, Hac82b, Hen84c, Kel68, Kra86, Mac88a, Nak75, Pat86, Ste70, Cro89a, Gab88, LS76, Sha87, Vea64]. Service [Por89d, Sch68a, Kni83c, Maf91, Web88, Web98]. Services [Por89d]. session [Nee78]. set [Hen84g, Rat66a]. setting [MF92]. seu [Viè83, Whi85a]. Seventeenth [Ait68, Ait72, Ait73, Ait86, Arm66, Bry74, Buc66, Byn72, For68, For74a, Gin74, Gra81, Hal66a, Hei68, Hew66, Hop72, Hop73, Hun86, Kel71, Kub70, McC77b, McG66, McG70a, Nor68b, Pep71, Pep72, Rat66b, Rat68, Rav72, Rus72b, Rus74, Sch68a, Sch70, Sch73b, Sch74b, Sch81b, Sha81, Sha73b, Sme68b, Smi74a, Str73, Tur73b, Tur74a, Web66c, Web68b, Web70, Web71b, Web72, Wes74, Wes70, Whi72b, Whi77, Yat72, Hen84l, Hun84, Sha83b, Ait71, Bos83, Bro78a, Bro78b, Byn74, Can80c, Car74, Cha77b, Eag78, Edw78, Fis81b, Gab75, GA98, Gou84a, Gou84b, Gra84b, Hac75, Hai77, Hal68c, Hal68d, Hal69a, Hal77a, Hal77b, Hal84a, Har84a, Har84b, Hei76b, Hen84e, Hen84m, Hey95, Hom80, Hop75, Hun84, Hun95, Ilt77, Ito88, Jar84b, Jor78, McC77b, McG68, McG71b, Mol75, Mor91b]. Seventeenth [Mur85, Nay75, Nay77, Por79b, PW89b, Pum84, Rat69, Rau75, Rav68, Rus77c, Sch81a, Sch74a, Sch75, Sch78c, Sha77b, Sla82, Ste62d, Str75, Til75a, Und75b, Und79b, Web68a, Web69b, Web69a, Web71a, Whi75, Whi68c, Whi77]. Seventeenth-Century [Web72, Hen84l, Sha83b, Bos83, GA98, Hun95]. Seventh [Hof62]. Several [Ste67]. Sewall [Bow87c, Pro86]. sewn [Sec83]. Sex [DH90, BB92]. sexangulaire [Sch78c]. Sexual [Rus89b, Cro91a]. Sexuality [Por87c, Sho87b, DH90]. Seymour [Pin83]. Seyyed [Mol69]. SFr [Gra85b, Gre88, Mah89]. SFRp [Por84a].

Shadow [Hun86, Mur85]. shadows [Sch92a]. Shaffer [Cun78]. Shaler [Kee89, Liv87]. Shandean [Mer93]. Shank [Out84a]. shao [Nak75]. Shape [Byn80, Hun95, Win91]. shaped [Hen85a, Pug84]. Shapin [Goo81a, Jon87]. Shaping [Mor87b, Ric88a, BC92, Dea83b, Hea87c, Rud85, Smi89a]. Shapiro [Hal86a, Hen84l]. Shark [Mac85b, Whe83]. Shatterer [Yox82b]. Shaw [Hil86b]. Shea [Sch74b, Sin81]. Sheets [Ham84, Her83]. Sheldon [Olb85]. Shelf [Lev88b]. Sheridan [Fig77b]. Sheridan-Smith [Fig77b]. Sherman [Tur83c]. Sherrington [Cla66, Cla71]. Shield [Por79b]. Shift [Fis81b]. shifting [LeG88]. Shigeru [Rav70]. Shipbuilding [Buc85b]. Shirely [Duc88]. Shirley [Dav71, Les83, Loh75, Pep86]. Shiva [Tee84]. Shivelbusch [Por89a]. Shop [Tur87a, CW87]. Short [AW71, Bar75, Bry81, Fis81c, Gol79, Had63, Her80, Law80c, McV80, Pic79, Rup88, Sch79b, Wea79, Wil65a, Woo80e, Buf87, Dea83b, Kel84, WD82, Fre81b]. Shorter [Pum87, NR86, WS91b, WS91a, Bar79, Ben79]. Shortland [Bra89a]. shortly [Ger89]. Shortt [Pic88a]. should [CP93, Nor85a, Yox83c]. shoulders [Had94, Mer93]. Showman [Bry78c]. shown [Gou84b, Yox82b]. Shrewsbury [Sch71, Und64a]. Shri [Hil75]. Shropshire [Tor83]. Shu [Nak75]. Shu-shu [Nak75]. Shuttleworth [Por85b]. Siberian [Jos97]. side [Hei93]. Sidereal [Arm66, GV89, SB09]. Sidereus [GP92, GP92, GV89, SB09]. Sidney [Dol73b, Jac89b, Fie86]. siebzehnten [Pag69]. Siècle [Ben89, Hal72b, Hom86a, Por89e, Bec91b, BaJANJ84, Cro66b, FM00, G^{+}09, Grm90, PP95, Pet84, Pic87c, Roc88a, Sch01, SGŚ00, Tur73b, TP90, Har89b]. siècles [Poi04]. Sieghard [Fis78]. Sigalia [Gra84c]. Sigerist [FB97]. sigils [Fis83a]. siglo [Rey90]. Sigmund [Por87b, FMF85].

Significance [Bod84, Fre79, Rob80b, Sto76, Ait85a, Ols90, JK84]. Sijthoff [Hom80]. silence [Gol70]. Silliman [Nor82, Bro89d]. Silurian [Tha87, Sec86]. Silvester [Lor75]. Silvio [Kel68, Tur78c]. Simkin [Woo85]. Simms [McC92]. Simon [Hen84c, HS83, Jon87, Pep71, Hai78, SSH85, Tal62]. Simone [Men70]. Simple [Bow88a, SL87, Stu85, Wil83]. Simples [Ear70]. Simpson [Bow88a, Mor73a, SL87]. Since [Por89d, CT95, DH90, FL88, Fru82, Huf91, Koh78, Lie88, Mac83, Mor84b, Mus74, Stu78, Tur78b, Wil91a, Wil84e, Wil89b]. Sinclair [Cal77, Mos77b]. Singer [Ano64l, Had63, Und65, Ano88s]. Singleton [Sch70]. Sir [Ait73, Bro69, Bro72b, Chi64, Chi65b, Col76, Cow68, Ear87, Hac76, Jam87a, Kni71c, Kni89e, Lev73, McG69, McG71b, Mor83b, Nev83, Ora73, Rud71, Rus68, Sha73b, Smi78a, Und75b, Web67a, Web69a, Whi68b, Woo67, Bow73, DeK73, Mil83, NJ91, PCA91, Rob89b, Rus86a, She63, Sta89a, STW90, Sut74, Tro89, Wal88, Wil93]. Siraisi [Cun83b, Cun89]. Sirks [Und66c]. sistemas [GB95]. Sister [Rat66a]. sisters [HH97]. Sivin [Agr83b]. Six [Chi64, Jen89, Nor68b, Tur81, Ede65]. Six-cornered [Nor68b]. sixteen [vLP82]. Sixteenth [Ant64, Coo81, Hut88, Sch68a, Ste62a, Web70, GJ86, Mol84c, Roc81a, Rud88a, Sch65, Yat72, Ait68, Ait72, Ait73, Arm66, Buc66, Byn72, For68, For74a, Gin74, Hal66a, Hei68, Hew66, Hop72, Hop73, Kel71, Kub70, McG66, McG70a, Nor68b, Pep71, Pep72, Rat66b, Rat68, Rav72, Rus72b, Rus74, Sch68a, Sch70, Sch73b, Sch74b, Sha73b, Sme68b, Smi74a, Str73, Tur73b, Tur74a, Wal69, Wal81a, Web66c, Web68b, Web70, Web71b, Web72, Wes74, Wes70, Whi72b, Yat72]. Sixteenth-Century [Ant64, Hut88, Sch68a, Web70, GJ86, Mol84c, Roc81a, Rud88a, Wal81a, Wal69].

Sixth [Mac88a]. sketch [Dea83a, Sch84a, Tee89, vH97]. Sketches [Ora73, Bec88, O’C88]. skies [Rob92]. Skinner [Hut89, Woo84]. Skopová [Rya70]. Slaughter [Hun84]. sleep [Cra93]. Slinn [Hen84d]. Sloan [Por84b]. Smeaton [Ede65]. Smelting [Gre71]. Smiles [Car67a]. Smith [All84a, BH84b, Bus85, Chi64, Des88, Eas86, Edg87b, Fig77b, Har86, Hen81, Kni89e, Por89b, Rud86, Rud88b, She63, Wal88]. Smithsonian [Goo85b, Hac74, Hac77, Ink88, Woo85, Wor88]. Smyth [For80, Tha66a]. Snelders [Hal84a]. Snell [Pat84]. Snow [Hen82c]. Snowflake [Nor68b]. so-called [BEoCF92, ECL92, vGA94]. sober [Hey95]. sobre [GB95, Lan89, Rey90]. Social [Bal83, Bar77, Ben82b, Bow86, Dur79b, Dur82, Far77, Fig78, Gow74, Hai77, Hai78, Hal80c, Hoc86, HS85, Ins86, Joh78, Kei86, Moo85, Mor75a, Mor81, Olb85, Por89c, Rob80b, Sha77a, Stu86, Stu88, Tre79, Ver89a, Wal81b, Web78, Wil65b, Woo73, Woo80e, Bro82a, Bry88, Can84, Cro84b, Deg91, Dol83, Fox84, Hal84c, Har83a, Har82d, Hul88, Ilt73, Irw82, Irw83a, JO81, Jor85b, Kri82, Kuk91, Kuk93, Mac90, McG84, MC^{+}93, Mir89, Mit92, Nor83, Nye90, Ram88, Rus83a, Rus83b, Sto84, Stu95, Web84, Wei91, Wei82, WT82, You83]. social- [Bro82a]. Sociales [Fox87b, Rey90]. Socialism [Yox81]. Socialist [Lil85]. socialists [Pit93]. Sociedad [Gli84]. Société [Cro68b, Por84a, Sme69, Wil77, Tor92]. sociétés [Ano93]. Societies [Eme87, For87a, Fox68b, Gur78, Men67, Ora76, Sme69, For86, McC85a, Ano93]. Society [Ait86, All80, All86, Ano62h, Ano62k, Ano84j, Ano85j, Ano86i, Ano87m, Bar83, Bow87b, Bro73a, Bro86a, Bro78a, Bro78b, Bro87, Bry83, Buc69, Bur80, Cah88, CC66, CH67, CE73, Cla86b, Cor81, Cro89a, Dol83, Dou88, Edg87a, Eme79, Eme81, Eme85, Eme88, Fie83, Fis83b, Fis83c, For68, For87b, Fox68b, Fox79, GS88, Gau83, Gid82, Gil88, Gle91, Goo83a, Goo85a, Gou84a, Hac89, Hal68d, Ham84, Har83a, Har83b, Hen82a, Hen83c, Hen89g, Hom86b, HWW86, Irw83b, Kei86, Kel71, Kel68, Kni83c, Kni89e, Mac88a, Mar73, Mil83, Mol84b, Mor84b, Mor74a, Mor80, Mor83c, Mor83d, Mos77a, Mos77b, Nor83, OM75, Olb82, Ora73, Pat86, Pep89, Por83a, Rau75, Rav73, Rob87, Rud63, Sch80b, Sha73a, Sha74, Sha79, Sin77, Smi89c, Ste70, Sti89, Tur83a, Und63, Vea64, Vic88]. Society [Wal88, Web72, Web87, WH91, Wig74, Wil65b, Wil84e, Wil84d, Woo83, Woo73, You83, AW87, BB92, Ben86, Gou82, Har89b, Hen82c, Lev84, Mac83, MS87, McG84, Nak91, Por85d, Sha83a, Sok87, Sut95, Tho87, Tor92, WM85, Wil90a, AL87, Bad62, BKH90, DTS^{+}87, Gun84, Hal66b, Hal68a, Hal84b, Hal91, Har67, Hop76a, Hop76b, Hun82, Hun89b, HH97, Maz92, Row90b, STT87, SB92, Ste67, Tay87, TST87, Woo80d]. socio [Bal91, PW89a]. socio-historical [Bal91]. socio-politics [PW89a]. Sociobiology [Dur80b, Moo89a]. sociohistory [Fuj96]. Sociological [Hen86b, Mor77c, Pic84b, Pum89]. sociologically [Nic88b, Söd86]. Sociology [Moo89a, Mor75c, Mul79, Bar83, Bar77, Col81b, Joh78, Law80a, Mor75c, Mor77c]. Soddy [Sin77, Cru79, Fre79, Tre79]. Söderqvist [Nic88b]. software [Bea85b, Mor84a]. Sokal [Gil88]. Solander [Ste70]. Solar [Doe96, For71b, Jam82, Huf91]. Solid [Sch69a, Hen82a, KH86]. Solla [Sha79, Smi78b]. Solomon [Har82d, McC77b, Frä90]. Solovev [Bro89b]. solution [Her65b]. Solvay [Stu78]. Some [Car67b, Gou68, Mor69a, Pac74, TP66, Tur74b, Web69c, Wil62, Bec88, Ben75, Hug89]. Somerville [Han85a, Pat69, Pat83a]. Sommerfeld [Hen86b, MR82f, MR82g]. Somnium [Hal69a]. Son [Mor69a, Ste62a, Cro79]. Sonar [Wil87b, Hac84]. Sonderweg [Wei89]. Sons [Cri66, Cro70b, Ede65, Har77, Hei75, Hen81, Mus74, Pra70, Sty68, Tur78c, Web66a]. Sophia [Hal84e, Ken83]. Sopka [Cas82]. Sorabji [Dyb78, Whi85b]. Sörbom [Bal83, Goo85a]. Sorcerer [Fox74]. Sorell [Gab88]. Sorrell [Sha87]. sort [Gol92a]. Sotheby [Law80c, McC85b, Wal77]. Sotherby [All77]. soul [App93, Fre95]. Sound [Jen87]. Source [Chi65a, Dav71, GG75a, Kei84, Rud69, Rus69, Sch76, Vic88, Cla99a, Cla99b, Cla95a, Cla9 , Cla2 , McG78, Pat94, Wuj85]. Sources [Buc68, Dea86c, Duh91, Jam82, Jef64, Mol88, Rus88d, Sme69, Swi77, Wal81a, Web68a, Web69c, Gol83b, Gro89, Gun84, Maz92, Mol84c, Nev84, Olb82, Wal84, WR92]. South [Nol95]. Southeast [Agr83a]. southern [Rob92]. Southgate [Kee89, Liv87]. Soviet [Har82d, RH88, Gra87a, Kul63, Wer75]. Space [Far68b, Kei84, Mol79, Dra70, Ear89, Hen83d, KRS94, McC93, Smi89d]. space-time [Ear89]. Spagyrical [Agr83b]. Spain [Coo89a, Cun88, Num87, Goo88b, Lan86, Lan89]. Spanish [Gli83b, Fre90, GB95, Hue88, Lan86, Lan87, Lan89, LN95, Mal95, ML89, Ord76, OE^{+}90]. Sparrow [Hal84e, Ken83]. spatial [Hat90]. Spätmittelalter [Gol91]. speaking [AS96]. Speakman [Sec83]. Speaks [Rud76]. special [Law93]. Specialists [Mor87b, Rud85]. speciarum [Sch85a]. Specie [Cow74]. specierum [BL83]. Species [Hod89b, Maz95b, Rud71, Slo79, Web65b, A^{+}87, BGH^{+}87]. specificity [Maz95b]. Spectacle [Hal88a]. Spectra [Bro71a]. Spectro [Jam88]. Spectro-Chemistry [Jam88]. Spectroscopy [Bro71a, DeK73, Din63b, Smi87c, Sut74, Hea86, Hea90]. Speculation [Wea84, Ano82p, B^{+}82]. Speculations [Far65, Und66b]. speculativa [BM89]. speculis [BL83, Sch85a, EiT92, ET92, Tak92]. Speculum [Zam92]. Speech [Wei83a]. Spencer [Hen81, Mor77b]. sperm [Cor97]. Sperry [Bow74]. Spheres [Rus79b]. spherical [EBT96]. Spiegel [Sha79]. Spiegel-Rösing [Sha79]. Spiller [Hen83c]. Spinal [Grm69, Jac85, NC81]. Spinning [Hil89b, KN88]. Spirit [Jor79, Wil82a, FHR90, Kni83b]. Spiritualism [Sho86a, Opp85]. Spirituels [Tur73b]. Spokesman [San87, Hei86]. Spontaneous [Ree80]. Sprat [Woo80d]. Springer [Bos72, Bra88, Bro89b, GG75c, Gra83b, Gra84c, Gra86b, Gra87b, Har81, Hen86b, Hen86a, Hen88c, Kra86, Mol78a, Mol88, Nor72, Sch75, Sme68a, Tur88]. Springer-Verlag [Bos72, Bra88, Bro89b, GG75c, Gra83b, Gra87b, Har81, Hen86b, Hen88c, Kra86, Mol78a, Mol88, Nor72, Sch75, Sme68a, Tur88]. Spronsen [Sme70]. Spy [Bad89, Wil87e]. Squares [Hum67]. St [All84a, Bar76a, BB82, Bro87, Bur80, Cor81, Gou84a, Jud88, McV75, Mor69a, Und80, Vic88, Wil84d]. St. [Sch85b]. Staats [Hom74]. Staats- [Hom74]. Staatswissenschaft [Nor71a]. stability [Kal95a]. Stacey [Lar74]. Stachel [Hen88d]. Stadt [Gra85b, Rus74]. Stafford [Con86]. Stage [Bru70, Gou85, Coh84]. stagnation [yQ85]. Stallybrass [Hei73]. Stamp [Ora77]. Stan [Nor82]. Stand [Tur75]. Stanford [Nic85, Rup89, Sec84]. Stanley [Roc88b, Sch90, Sey77, Mil87]. Stansfield [Por86a]. Stanton [Kni88b]. Staples [Rus77c]. Stapleton [But64a, But64b]. Star [Jev64a, Ber95]. Starkey [New94]. Starlight [Smi87c, Hea86, Hea90]. starry [Ber95]. Stars [AW71, Bra87, Jev64b, McC68, Gra94, Smo94, Wil84a, Zam86]. start [Lan87]. State [Bro85a, Can80b, Col88b, Hes74, Kei86, Mac89, Mol79, Rei68, Ver89b, Whi85a, Whi89b, Alt87, Hen82a, KH86, Mar92, McG84, Mit92]. stated [All84a, Bow83a, Bro83b, Bro84d, Can74, Duf81, Gab83b, Gli84, Hal84a, Hen83b, Lay74, McC83, Mor83d, Nut84, Olb82, Pic84a, Ree85, Sch81b, Sme84a, Tei83b, Woo80c, All77, Bos70, Bos72, Bro76, Bry77, Car74, Cha77a, Far76, Fie81, Fig77a, For67b, For69a, For69b, Fre70, Gol79, Gol83b, Gow71, GG74a, Hil71a, Irw83a, Kel75, Kel78, Mau77, McV75, Nay75, Ora76, Osp74, Out80a, Por78b, Rus67, Rus72b, Rus77a, Rya70, Sch75, Sch81b, Sch78c, Smi75a, Smi80b, Sne74, Str75, Tho76a, Tur72b, Tur73d, Tur73b, Wil75, Win75, Woo80b, Woo80c]. Statement [Ano63g, Ano64n, Ano69g, Ano70g, Ano71g, CM64, CC65, CH68]. statements [Gro89]. States [Bud84, Cor81, Har79b, Men67, Rei86, Roc88b, Smy89, Hun85, Rot93, TT86]. statesman [WV92]. Statics [Duh91]. Station [Roc88b]. Stationery [All84b, Ben84, Hac88a, She63, Wil87b]. Statique [Sme84a, BG80, SG80]. Statistical [Hod89a, Lor77, CGBE95, Pau93, Por86b]. Statistics [Hod89a, Nor83, EM94, Por85d, Sti86, SG80]. Status [Stu86, Stu95]. Staudenmaier [Edg87a]. Stavenhagen [Win75]. Stay [Nay85, Dra81]. Steam [For87a, HS86b, Law65, Ler85, Lin84]. Stearn [Kni89e, Rau75, Sty68]. Steeds [Car70]. Steel [Cla86b, Buc85b, Mis95]. Steel-The [Cla86b]. Steele [Byn74, Rus67]. Steelmaking [Bar84, Cla86b]. Stefansson [Lev88b]. Steffen [Cap81b]. Steffens [Goo81b]. Stegmüller [Cla79]. Stein [Hen82c, Sec88]. Steinberg [Bro89b]. Steiner [Pag69, Whi68a]. Stell [Aus86]. Stella [Tee84]. stellae [Rav69]. Stellar [Wil72, Sch80a]. stelsels [Tur82]. Sten [Mor78]. Stendhal [Pau76]. Steneck [Lef78]. Stengers [Bru86]. Steno [Osp74, Sch69a]. Stent [Yox83a]. stepchild [Dij90]. Stephanie [Hen84d]. Stephanitz [Nor71a]. Stephen [Hod89a, Jar83, Lay74, Mor75a, Mor83d, Por82, Smi80a, Tha74]. Stephenson [Bra88]. Stereotypes [Por87c]. Sterling [Cro84a, Bow87a]. Steve [Irw82]. Steven [Goo81a, Pum84, Wor88, Fra87c, Gra86a, Jon87]. Stevens [Tur78a]. Stevenson [Por85a]. Stevin [Pep71, Sch84a]. Sticker [For69a, Rus74]. stiefkind [Dij90]. Stigler [Hod89a]. Stillman [Gab83c, Nay75, Rav72, LS90, Nay85]. Stirling [Dau74]. Stock [Lor75, Kni83a]. Stockholm [Bal83, Bro84d, Cha80c, Cro67, Frä90, Mor78, Nic88b, Rus67, Ber95]. Stocking [Bow86, Lew85, Web75]. Stoddard [You83]. Stoic [Osl91]. Stokes [Buc89a, Smi78a, STW90, Wil72, Wil87g]. Stollard [Can74]. stonemason [MS95]. Stoothoff [Dun87]. Storer [Mor75c]. Storia [Tur72b, Ara91, Out80b, Por79a, Sch73a]. Storr [Buc85b]. Story [Cla86a, Tom84, Yox83b, Gol83a, HS83, Smi85b, Wai90]. Storytellers [Edg87a, Sta85]. Stoughton [GG85, McV80, Stu85, Swi77]. strafrecht [Bur99]. straighteners [Mor84f]. Strangelove [Hay94]. Strata [Bar75]. strategic [OO89]. Strategy [Hod77, Rap86b, Wat80, Kni84a, Kni84b, Len82]. Strathern [Dur81]. Stratification [Mor75a]. Straubing [Rus79a]. Street [Pat86, Whi85a]. streitbaren [ZH92]. Strike [Wil87b, Hac84]. Stringfellow [Pen89]. structural [Sch89]. Structure [Bar77, Bec91a, Hod77, McC67, Roc81b, Bar89a, BN82, Dea91, Sal85, Sch84c, Yox83a]. Structures [Smi85c, Cha82, Cor96, Pri83]. Struggle [Fri74b, Rud63]. Struik [Sch84a]. Stuart [Ait89]. Stubbe [Coc75, Hen84m, Jac83]. Student [Hew66, Ste62a]. Students [Bro89e, AT87]. Studer [Irw83a]. Studia [Eas86, Mol81].

Studien [GG81, Mol72a, Mol75, Wey92, Whi68a]. Studienplan [Kra91]. Studienverlag [Bro82a]. Studies [AO77, Bad71, Blo82b, Bro72b, Can80b, Col83, Coo88, Cro89b, Fre92, Fur85, Gol88a, Goo81a, Gre63, Hen82a, Hen84g, Irw82, Jam87b, Kel78, Kra91, Kul63, Lay85, Lin86, May67, McG65, Mol72b, Mol78a, Mol84c, Mol84b, Mol85, Mor77b, Nor73a, Nor77, Ped85, Rav67, Rav72, Rus77b, Sch83a, Sch84b, Sch69a, Sha89, Sho88, Sin84, Smi76c, Sne74, Tup82, Wea84, Whi73, Whi85a, Whi84, Yat72, Ano82p, Ano07, B^{+}82, Bar83, Bro95a, Cra92, Dea91, EC87, Fie83, GT92, Har85a, Høy94, HS89, HC90, Llo83, Lan89, ME81, PS82, Raf88, Roc90, SY86, Smi82a, Smi83, Viè83, Web84, Wey92, Wil84a, vLP82, B^{+}89, EK91, EK92, Kul64]. Studio [Tur77a]. Study [Ant64, Buc68, Buc89a, Byn80, Can77, Can80c, Col81a, Fig78, Geo66, Grm69, Hal80c, Hei74b, Hig65, Hop72, Hun89a, Kel71, Kni75, Kni86b, Law81, Lin86, Lor75, McC76, Rav73, Rog88a, Sch75, Stu86, Wea78, Bar89a, BM82, Büh81, Can91, Cha91, Coo89b, ES90, EBT96, Gje84, Gra83b, Har87a, Hen84l, Hoo97, Le 90, Mei91, Mor82b, Nut84, OP87, Old86, OH85, Out84a, RA91, Sch84d, Sha83b, Smi89d, Sul82a, War90, Wil87g]. Stuewer [Bro76]. Stupidity [Bar89c]. Sturchio [Jam86]. stuttering [Wei83a]. Stuttgart [GG77a, Har87b, Jac89a, Kra86, Mah89, Pat84, Rus72b, AH93]. Stuttgart-Bad [GG77a, Rus72b]. Stuttgart/New [Har87b]. Style [Ano04, Can77, Fru90, Hom83a]. subalternate [Liv85]. subject [Dan90]. Subjects [Fox79]. Sublunari [Rat66a]. Submarine [Wil87b, Hac84]. Subordinate [McK78]. Subscribers [Tur68]. Subscription [Hen89e, RW72, Mor83a]. subscriptions [WW88a]. Substance [Con86, Eva89, Fur88, PW89b, Fre95, Har82c, Hen84f, Sta84a]. Subtle [Pai82, Red84]. Success [Bad79]. Suez [MM10]. suffering [Ost89]. Suffolk [For85a]. suggested [Tee89]. Suggestion [Hod89b, MC81]. suggestions [FTG91]. Suhrkamp [Cap81b]. Suicidal [Bad79]. sulla [Des89b]. sull’atomismo [Sch79b]. Sullivan [Edg84, Sch84b]. Summary [Cor81, Fis81c]. sun [Hen91, Huf91, Roc81a]. Sunt [Mol84b]. Sunyer [Mal95]. superbomb [Yor89]. Superior [Kni88c]. Supernatural [McV80]. Suppes [Dol73b]. Supplement [For80, Rud89, BSK^{+}94, Cas85, Wei92, Sar87]. supply [H^{+}82, Hen84d]. Supposed [Far63]. Suppressed [Cro73]. Supracelestial [Lem77]. Surgeon [Fre81b, Kni86c, Ald85]. Surgeons [Fre81b, Gab83b]. surgery [Rob89b]. Surrey [Jon87]. Survey [Arm70, Dol70b, Her95, Mer78, Mol72b, Rus89a, Rus72a, Tor83, Her88, Nic88b, Söd86, Man88]. surveying [Ben87]. Surveyors [Has72]. Surviving [Smi87a, CF86]. Susan [Har82d, Kni89c, Ora80, Tur74c]. Sussex [Bro80b, Hen84c, Jar83, Jud88, Mon78, Sec83]. Susskind [Buc75]. Sutt [Geo78b]. Suzanne [Rat66a]. Svante [HS86b]. Svehla [Kni67]. svolte [Out80b]. Sw [Bal83, Fis81c, Mor78]. Swann [Mor89a]. Swazey [Cla71]. Sweden [Frä90, HS86b, Nic88b, Ber95, Frä89, Lin84, Söd86]. Swedish [Cha80c, Cro67, Frä89, Rus67, Lin93]. Sweet [Sch78b]. Swenson [Goo82, Hen82c, Mar74]. Swets [Sch83a]. Swineshead [HM66]. Swiss [Mol75, Pau76]. Switzerland [Pau76]. Sydenham [Hac82b]. Sydney [All82, Bro89c, Des88, Gol88b, Har77, Kni88a, Rud86, Rud88b, Sme66]. Symbols [Wil82a]. Symmetrie [Sch89]. Symmetrien [Mai88]. Symmetry [Mai88, Sch89]. Symphorien [Wea80]. Symposia [Mac88a]. Symposium [Bal83, Bow87b, Chi65c, De 85, HWW86, Rus74, Whi77, Wil87d, CT95, HE^{+}82, N^{+}88, Pep89, F.76, Hal68a, BDH89]. Synthese [Jar84a, Mol89, Hal65a]. Synthesis [Bro71b, Nye76, Per73, Bow83a, vdVV91]. Synthetic [Swa83, Mor89b, Tra93]. System [Fre74, Jor79, Mor72a, Sme70, Ben75, BV86, Doe96, Hen85a, Mac88c, Mac90, Pug84]. System/360 [Pug84]. systematics [Kni83c]. systems [Abe86, Cal90, GB95, KH95, PJP91, Sal85]. Szabadváry [Kni67].

T [Bro83b, Bry76, Coc75, Dol83, Fox69a, Geo78b, Gin84, GG77b, Gra84c, Hei76b, Jam86, Kni89e, Law65, Mor84c, Nor68a, Nor69a, Nor70a, Nor73a, Olb82, Pep72, Rau75, Rud84, Stu85, Sty68, Til75a, Tur68, Tur74c, Wal81b, Whi85a, WHS92, Whi84, Wil84c]. T. [Bow87b, CHG91, Smi85c, Whe89]. Table [Sme69, Ano82l, Web70]. Tables [GW74, Swe76, Cha65, Sch78b, Sme70, Sme71]. Taccola [Kel73]. Tacitus [Hüb90, Hüb90]. Tacitus-Übersetzung [Hüb90]. Taddeo [Cun83b]. takes [Nyh95]. Talbot [Sch92a]. tale [GG92, Rit90]. Tales [Gin88, Gre88]. Tallahassee [All84a]. Tallinn [Geo78b]. taming [Hac90]. Tammen [Mor84g]. Tamny [Hen85c]. Tannery [McC80, McC83]. tapestry [KH92]. Tarbell [Roc88b]. tardo [Ern91]. Tarwater [GG77b]. Tasker [Sol83]. Taton [Cro66b, Fin68, Sch84c]. Tavistock [Fig77b]. Taxonomy [Ver88, Hun84, Sla82]. Tayler [Smi89c]. Taylor [All79, Ano67j, Bry78a, Duc88, Kni86d, Law83, Sco67, Scr68b, Sin77, Sin78, Tur73c, Wea83, Web71a, BT96, BM84]. Teaching [Arm62, Arm64, Bib74, Cun89, Fis81a, Lay74, SW89, Smi66, Sol83, Ano87r, HJF89, Law93, Sir87]. Teaneck [Kni75]. Tech [GG77b]. Technical [Cha65, Cha78a, Cro68a, Rya70, Ano90, Bec88, Bec91b, CK89, LHB04, LS76, LHH91, vL76, Ano87r]. Technik [AH93, Cap81b]. Technique [Cal63, Ano87r, Bec91b, CAMR87, CLR87, Dug77, Gua89]. Techniques [Cal63, Hil71a, Swi71, Ano90, Ano93]. Technischen [Ree85]. technischer [LHB04, LS76, LHH91, vL76]. Technological [Cha65, Edg87b, Fox95, Hac89, Ben86, Bro82a, Hal84c, SM94, Wei82, Fox96]. Technologie [CM90]. Technology [Ant63, Bal83, Bal84, Ben84, Buc73a, Buc75, Buc85a, Buc85b, Cah88, Car67a, Car84, Cla86b, Col88a, Coo89a, Cro68a, Cro72, Dou88, Edg84, Edg87a, Fie81, Goo85a, Hac82b, Hal84c, Hal86b, Hal86c, Hal88b, Hen83b, Hil75, Hil89b, HS86b, HS88, Jen87, Kar76, Kei84, Kel68, Kel84, Kra86, Lin84, Mar73, MJ84, Mea84, Mol71, Mos77b, Mus74, Nee69, San88, Sha79, Sho88, Smi71, Smi74b, Sta85, Tre88, Ubb63, Web72, Wil84e, Wil84d, Wil89b, AH93, Ano93, Bak96, Bas88, Bau95, BC89, BC92, But92, C^{+}92, Car94, Cha91, CB^{+}90, Chr93, CM90, CK90, DM96, EC87, Fie83, FG93, Fox95, Goo88b, Got92, Gra98, HSHSJ93, HH86, HH92, aHH86b, Hil94, IW96, KN88, Lie88, Mac83, McC93, Nak91, NLH^{+}84, NPYGDL86, NhTWL87, NOF96, Nye90, Ole84]. technology [Out82a, Rot93, San91, Sch84a, Seg94, Smi89d, SM94, Sol83, Ste92, Tho87, TW87, WD82, Wil90a, Wuj85, YK94, aHH86a, aHH88, aHH92, AH93, All82, Bal83, Bow74, Bro82a, Buc75, Bur71, Car70, Ged71, Gre71, Hac76, Har76, Hil71b, HS83, Kel75, Kel76, Mor83d, Nut83, Out82a, Pac71, Ran74, Smi75a, Sol83, Swi70, Tan82, Tup82, Tur74c, Tur75, Fox96]. Tectonics [Nun88, Men86]. Ted [Sha87]. Teeter [Hal77b]. Teich [Hal87b]. Tekniska [Bal83]. teksten [Bur99]. Telegraphy [Hac76]. teleology [Kni84a, Kni84b, Len82]. telescope [Rob92, Smi89d]. Television [Yox79, Yox82b]. Teller [Yor89]. Telliamed [Rap71]. Telling [SY96]. Temkin [Phi69, Ros79, Wil75]. tempera [Pum89]. Temperature [Smi80a]. Temple [Dur82, Joh83]. Tempo [Rap86a]. temporal [Whi88b]. temps [Poi04]. Ten [Ano71h, NBR^{+}95]. ten-volumes [NBR^{+}95]. Tennessee [Cun88, Olb70, Roc88b]. Tenu [Bur74b]. teoriche [Out80b]. ter [Kuh68]. tercentenary [HC90, Ano68e]. Terrestrial [Hom74]. territorio [Sch81b]. Test [SG80]. Testament [Win75]. testi [Sch73a]. Testing [Gil88, Sok87]. Texas [GG77b, Hen82c, Mar74, Por82]. Text [Dol71a, Per89b, BM89, Cal63, Chi64, Fre86, NR86, Ste62a, Ste62c, Yox83a]. text-figures [Ste62c]. Texte [Pag69, Scr68a]. textes [CAMR87, CLR87, Gol83b]. Textile [Hil71b, Hil89b, KN88]. Texts [Gol88b, Ilt77, Mol85, Bur99, KRS83]. Textual [Mol89]. Teyler [Bry76]. Thackray [Hei71a, Jam86, Kni74, Law80a, Pat86]. Thaddeus [Col81b, Rav69, Sin77]. Thames [Fin68, Han76, Por77, Tur74c]. Thaumantias [Rav69]. Theatre [Rud76]. Their [Car67a, Dol73a, Ubb63, Bry74, CMNH92, De 85, Gro89, MKW92, Nut82b, RCRC98, Sal85, Tei83b, Wil87d, Wol87, vLvHH^{+}88]. them [Fis98, HS83]. Theme [Whi70a]. Themes [Bro85e, Fox95, Fox96, Osl91, Reh83, Spe82]. Theodor [Hei93, Olb69]. Theodore [Kel71, Hod89a]. Théodoridès [Cla70, Pau76]. Theodosius [Gid82, Smi76c]. Theologica [Hüb90]. theological [Gre92b]. Theology [Lem77, Lev68, Yeo86, Zam92]. Theology-Reflections [Lev68]. Théophraste [McC77b]. Theophrastus [Got66, PTS^{+}96]. Theoretical [Her66, Hog74, KH86, Mor75c, Wil89a, Jun89, KT87, Nye93, Sch89]. theoretische [Jun89]. theoretischer [Sch89]. Theorie [GG81, Duc87]. Theories [Car88, Cro90b, For71a, Hei71c, Lau68, Lug79, Mar74, PW89b, Rup89, Smi85a, TP66, BaJANJ84, Hom86a, Agr83b, Bar89a, BaJANJ84, Bow83b, Bow89b, Can83, Cor88a, Ear89, Hat90, Hen83d, LeG88, Nor85b, Sha93, Smi83, Sor83, Whi85b, Dou84, Duc82, Sch84b]. Theory [Bod65, Bro85d, Bru70, Bru72, Byn80, Cam85, Can76, Dor75, Duh91, Eva77, Far71b, Fis74, For72, Geo66, Gin74, Gra84a, Gra86b, Hal71, Hei71a, Hen86b, Hen88c, Hom74, Ito88, Koh75, Kuh68, Lov75, Mid65, Mor72a, Mor77c, Mül87, Pac70, Per70, Rav71, Roc81b, Ros68, Rud80, Ser88, Sin77, Smi85c, Stu71b, Stu88, The87, Tho76a, Til76, Tur74b, War88a, Wat84, Wea78, Web67b, Wel70, Ano82p, B^{+}82, Ben82b, Blo84, Buc85d, Buc89b, BJ91, CD81, Cha82, Cra96, Duc87, Fis81c, Fre95, Gol70, Gra84c, Gra85a, Gra85c, Hen86c, Jam84, Kil94, KI92, Liv85, McI85, MR82d, MR82e, MR82f, MR82g, Men75, Pet90, Pop82, Pou93, Ric92b, SO85, Sch65, Sie91, Smi76b, TC91, Tur83c, Wei92, Wus84, Bro84b, Fuj86, Gol77, Mor72e, Rap73, Smi74c, Tha66b]. Therapeutic [Pic88b, War86]. Therapeutics [Und75a]. Therapy [Can86, Joh84]. there [Hon89]. Therésè [Gua89]. thermochemistry [Kra84b]. Thermodynamics [Bru70, Hut81, Jam82, Pay74, Kre90, Smi76b]. thermodynamique [Hut80]. Thermometer [Bry71, Chi68, Tur81]. Theses [B.69, FB75, Old77]. thesis [CDS90]. Thibaud [GG77b]. Thing [Kni80a]. Things [Kni87, Ara91, WHS92, AO77, Ste91b]. Thinkers [Tur78c, Tur87c]. Thinking [Bro95a, Hod89a, Por86b]. Third [Mac88a, Men67, Olb71b, Rav66b, Whi73, LS76, Hom68, Lon88]. Thirteenth [Nak75, Whi90]. thirties [Hen84h, Stu85]. thirty [vLP82]. thirty-two [vLP82].

Thistle [Gow69]. Thomas [Bar76b, Bow74, Byn74, Cah88, Car67a, Cla86a, Cla66, Cri66, Cro71, For66, For72, For85a, Fox71b, Fri74a, Gun84, Hah73, Irw83b, Jon87, Kni81, Lar74, Lev84, Mac85b, Mor81, Mor83b, Mor83c, OM75, Pag66, Pep86, Por86a, Por87a, Rav71, Rus77c, Str75, Sty68, Web66a, Web69a, Wig78, Yat72, BM89, But63, EJ^{+}07, ER93, EJR95, Gal72, Jen91, Jon75a, Jor87, Ken64, Law65, Loh75, MR89, Mor69b, Nic88b, Nor80, Pep67, SSH85, Shi83, Smi79, Sou93, Sta84b, Tom84, Woo80d]. Thompson [Far77, Fox69b]. Thompson-Count [Fox69b]. Thomsen [Kra84b]. Thomson [Chi65b, Mor69b, Smi78a, STW90, Dah97, DTF97, Fal87, Goo80, Ken64, McC67, Smi76b]. Thorkild [Smi75a]. Thorndike [Pag66]. Thorpe [Asq80]. those [Ear70]. Thought [Dol70b, Dur79b, Got67, Hai78, Hel63, Jac81, McE79, Ost89, Por85b, Rap78, Rav66a, Sha77a, Sti89, Wal81a, Web78, Woo80e, AW87, Bec91b, Ben82b, Cha88b, Cro90a, Cro96, Deg91, Dob91, Dur82, Hal92, Har87a, Hen84j, Kre82, Mir94, Mit92, Mol84c, Moo89b, Osl91, Pit93, Por83a, Tri89]. Thoughts [Kel85a, Whi64]. Threat [Cro87b]. Three [Mor87a, Ros86, GT92]. thru [Sar81]. Thuillier [Sho89]. Tides [She70]. Tiger [Mac85b, Whe83]. Tijdstroom [Tur82]. Tilling [Bro78a, Bro78b]. Timaeus [Sti84]. Time [Bra87, Goo85a, GG79a, Hen88a, Kau84, Lan82, Mer79, Mol79, Olb71b, Poo98, Rap86a, Rud74, Scr68b, Sor83, Whi84, Whi85b, Whi88b, Adl87, Ano87r, Cha83, Dea83b, Dra70, Ear89, Gol92a, Har67, Huf85, McG78, Poi04, Ros84b, Rud92, Tod83, Wil87a, Zam86]. Times [Coc75, Gur78, Hen88a, HS86a, McV80, FB97, Fei90, Hil84, Olm92, RCRC98, Wil87a]. Timothy [Kni84a, Kni84b]. Tinkers [Tur78c]. TISC [Chr93]. TISC-Conference [Chr93]. Tisseurs [Hil86a, Cha84]. Title [Ano63m, Ano63l, Mid69, Gre86]. Titles [Ano64o]. Titre [Fox87b, Gre86]. Tjeenk [Cro70b, Fox71a]. TN [Roc88b]. Tobias [Bar73a, Hom74, Wil84b, For70]. Toby [Gab83b]. Today [Bur71, Rei86, Sne87, Ver89b, Cro91a, MMVZ90, Rus88a]. Todd [Kau84]. together [AL87, Hen89g, Mor72a]. Toil [Fie81]. Tokyo [Gra87b, Hen88e, Kni89b, Kra86, Mol88, Sha89, Tur88, Bro81]. Toland [Sch84b, Sul82a]. Toleration [GG74a]. Tom [Cla68, Ora77, Gab88, Sha87, Woo85]. Tomash [Sei85]. Tome [Bos70, Cal63, Cas85, CLR87, Cro68b, Gua89, Hil71a]. Tomes [Cla86a, Sch80b]. Tommaso [Ern91]. tone [Bry77, Chi65b, Cri66]. too [Ben84, McC82]. Tools [Car70, But96, Hac82b]. Toomer [Van86]. tooth [Mir94]. Tooze [Yox83b]. Topics [Tur73c]. Torbern [Cro67, Dun69a, Sme71]. Tord [GG72]. Toronto [All82, Col83, Gow69, Har77, Lin86, LM74, Lor77, Mol85, Mos77a, Mos77b, Pep89, Was69, Whi70a, Wil78]. Torporley [Pep67]. Torrens [Ben79]. Torricelli [dGB^{+}89]. Torrini [Hat87]. torsion [BD95]. Toscana [Sch81b]. tot [Hen86d, Mül89]. Total [Duf75]. Totowa [Bus83, Jar83]. Touches [Sch78b]. Toulmin [Tha74]. Toulouse [Tho74a]. Tour [Rud62]. tournant [FM00]. toutes [SGŚ00]. Toys [Tur87b]. tr [Ait89, Kni67, Mol88, Por87b, Van86]. tra [Lan91]. Tractatus [McV75, Mol89, WW88b, Mol73]. Tracy [Roc88b]. Trading [Bry81]. Tradition [Hos67, Kni89c, McG95, Rav72, Sch83b, Sne74, Wea80, Ano07, C^{+}95, Deb91, Lin92, Liv93, Moo84, N^{+}88, PR88]. traditional [Bur99, CK90, yQ85]. Traditions [Bud84, Coo84a, Duc84, Gin84, Gli84, GG84b, Hal84d, Hen84d, Mor84c, Mor84g, Nev84, Pic84a, Sch84a, Web84, Wil84b, Woo84, Dun90, EiT92, ET92, Gre92b, Mai91, PW89a, RT92, Tak92]. tradizione [Sch73a]. Traduction [Sch78c, BN82, Hal68c, Sch84c, Sch01]. traen [ML89]. tranquillity [Osl91]. Trans [Bro72a, Bry74, Cla67, Coc75, Cro71, Dor75, Fig77b, For71a, For74a, Gel76, Ilt77, Kel75, Lar74, Law83, McG69, Nay75, Rap71, Rau75, Stu71a, Bar73a, Bro83b, Byn74, Fig77a, Rus67, Sch74a, Smi82a, Win75, Woo80c]. Transactions [Kel68, Mol84b, Ste70, Swe76, Woo83]. Transcend [GG74a]. Transcendental [Lev80]. Transcription [Gra81, Vic88]. transfer [Bro82a]. transfers [TBDA10]. Transformation [Smi73a, GG86a, GG86b, Hey81, Kus95, Mor84c]. Transformed [Fis81b, FGJ98]. Transforming [Mai91, NOF96]. Transformist [Cor78]. transistors [ES90]. Transition [Bru70, Sei85, Moy83]. Transl [Tei83a]. Translated [Bea85b, Bow89b, Bro89b, Col81b, Dol71a, Dun87, Fre86, Got71, Gra86b, Hal66a, Hal69a, Hal71, Hal88a, Hen88b, Jac85, Jar84a, Kel70, Kra86, Out84a, Rus77c, Sch69b, Sch78a, Smi87a, Ste62a, Tre87a, Wei85a, Wei86, Wel70, Whi72a, Wil86b, Bro84c, Dun69a, Hei74a, Nor68b, Olb69, Rav66a]. Translation [Arm66, Bar73a, Eas86, Got66, Jar84b, Jon87, Lin86, Mol81, Mol85, Mol89, Ord76, Rat66b, Sch85a, BL83, Bar75, BM89, Cop92, Dra81, EiT92, EBT96, ET92, Hen88d, Hüb90, Mol78b, Nay85, Nor71b, OH85, Rap86a, Rus79b, Sch01, Scr64b, SSH85, Tak92, Vic88, VoC89, Wil85a]. Transmutation [Des85, BGH^{+}87]. Transport [Aus86, Bal83, FS84]. Transports [CL96]. Trapped [Mor87a, Ros86]. Trash [Cro89b, Raf88]. Travaux [Osb89a, Out88b, Lan87]. Travel [Con86, CL96, Gid82, Sta84a]. Traveler [Web70]. Traveling [GG92]. travels [All84a, BB82, BS91]. Traynham [Col88b]. Treacher [Law85]. Treatise [Byn74, Car69, Fre86, Nut84, BJ91, EBT96]. treatises [PTS^{+}96, Wal92]. Trenn [Col81b, Sin77]. Treves [Tro89]. Trevisani [Sch79b]. Trevor [Bar75, Bro73b, Hal68d, Kel84, Mos77a, Wig69, Fra87b, Swa86, Wil85b]. Trevor-Roper [Hal68d]. Trial [Bus85, HS86b, Lin84, Wes89]. trials [Bus85, Hüb90]. Triangle [Mah89, Edw87]. tribute [Yox83a]. tricentenaire [Pic87c]. tricentennial [Pic87c]. Tricker [Wil68]. Trigonometric [Ger89]. Trigonometric-astronomical [Ger89]. Trigonometrisch [Ger89]. Trigonometrisch-astronomisches [Ger89]. Trigonometry [Whi85a, Viè83]. Trinity [Gra81, Hen85c, Kni89c, N^{+}88, MT83]. Tristram [Sho88, Tom65]. Triumph [Mac80, But96]. tropical [Gro95]. Troughton [McC92]. Trower [Bro89e]. True [Lev65]. Trust [Hen89g, Por95]. Trust/Science [Hen89g]. Trustee [Gun84]. Truth [Cri66, Nun88, Ada89, Men86, Sap90, Sou93]. ts’ao [Agr83a]. ts’ao-mu [Agr83a]. Tselos [Col88a]. Tsien [Hal86c]. Tsuen [Hal86c]. Tsuen-Hsuin [Hal86c]. Tuan [Whi70a]. Tuberculosis [Por89b, Por89c, Bry88, Smi88]. Tudor [Mar89, Smi74a, Jon88]. tumultuous [Cre93]. Tunnelling [Hil89a, Wes88]. Turin [She85]. Turing [Rus89a, Her88]. Turku [Kal95a]. Turn [Sin81, BS79, FM00, Kle94]. Turnbull [Can80b, Mol79]. Turner [Ben79, Bro75, Bry76, Edg84, Fox82, Hac82a, Hal68b, Mar89, McC85b, Rob80a, ABR93, Jon88, SH73]. Turnhout [Gol83b]. Turning [Bro76]. Tuscaloosa [Cah88, Kee89]. TV [Hen84c]. Twelfth [Lor75, Sti84]. Twelfth-century [Sti84]. Twelve [Kni86b, Gje84, Spe82]. Twentieth [Asq80, Bar76b, Bib74, Bro76, Bry78b, Buc77, Bur74b, Byn78, Can74, Can75, Cap81b, Cor81, Cow74, Cro79, Dau75, Dol69, Dor75, Dor79, Dor80, Duf81, Dug77, Dur79a, Dur81, Far76, Fin68, Fis78, Fox74, Geo78b, Goo75, Goo81b, Gow78, Gow69, GG74a, GG75c, Grm69, Har77, Har79b, Har81, Hei75, Hei74a, Hei77, Joh79, Jor79, Kni71b, Mac78, Mac80, Mac79, Mar74, Mau77, Mor74b, Mor75b, Mor77b, Mor81, Mor89a, Mus74, Nor80, Ora74, Ora76, Pau76, Pic76, Pic80, Por78a, Por79a, Por89c, Pye76, Rob75, Rus79a, Rus88d, Sha78, Sha77a, Sin74, Sin76, Sin77, Sin78, Smi76a, Smi78a, Smi78c, Smi80a, Sti89, Stu78, Tei78, Und79a, Web75, Web76, Hen83b, Kel84, NBR^{+}95, Arm83, AW87, Bra89b, Bro72a, Bry88, Buc73b, Bur99, Dav81, Din66, Dol70b]. Twentieth [Dor80, Fin68, Hen81, Joh81, Kuh68, RCRC98, RT95, Smi80b, Swa88, Tur70c, Tur87c, WR92, Whi68b, WD82]. Twentieth-Century [Dor80, Por89c, Sti89, Tei78, Kel84, NBR^{+}95, AW87, Bry88, Din66, RT95, Swa88, WR92, WD82]. twintigste [Bur99]. Two [Ait71, Bar76c, Dol71a, Got67, Hen89e, Jev64a, Jev64b, Mac88a, Nay75, Per70, Por81, Tur81, Cun83b, GB95, GG92, Gre86, Hen86b, Rit90, Smi78c, vLP82]. Two-Way [Jev64a, Jev64b]. Tycho [Ait85a, JK84, Mor87a, Hel63, Ros86, TC90]. Tychone [JK84]. Tyne [Buc85b, Eag78, Fau89, Cam82]. type [Sta87]. Types [Got67]. typescript [Hal88a]. Typologie [Gol83b]. Tyranny [Wig73]. Tytler [Jor85a]. Tytus [Löw90].

U [GG75a, PHvMW93, PvM85]. u.a [Hen86a, HP85]. U.C.L.A [Hol71]. U.K [Bro87, Hac82b, Hen82a, Jac89b]. U.M.I.S.T [Nev84]. U.R.S.S [MdLD86]. U.S [Bro85d, Fri87, Gra87b, Hen87c, Kni83a, Sch68a]. U.S. [Man88, McC93, Wis85]. U.S.A. [Roc88b]. übersetzt [Scr68a]. Übersetzung [Hüb90]. UCLA [Whi77]. Ugo [Sch79b]. uit [Phi67]. Uitgeversmaatschappij [Tur70c, Tur73d, Tur82]. UK [McC85b, Yox79, Yox82b]. Ullman [Yox83a].

Ullrich [Tre88]. Ulrich [Nak75]. Ultra [Ste81]. Umbruchs [Sch93]. un-numbered [Nun88]. Unbelief [Yeo88, Lig87]. Unbelievers [Duc88]. Uncertain [Dol96]. Uncertainty [Hod89a, Sti86]. uncharted [GG92]. underground [Wil90a]. Understanding [App93, Bro71a, Buc75, Dur77, Goo88a, Tha74, vdVV91, Gol91, Ner87]. Underwood [Fre78, Had63, Han70, Vea64]. Uneasy [Jak84, Mil87]. UNESCO [HS88]. Unguru [Mol81]. Unification [Bro71b, Bow83a]. Uniform [Dra70]. Uniformity [Rei86]. Union [Che78, RH88, Tur70c, Tur73d, Gra87a, Wer75]. United [Bud84, Cor81, Har79b, Men67, Rei86, Roc88b, AB89, HAB92, Hun85, Rot93, Smy89, TT86]. units [Tun92]. Unity [Jac81]. Univ [JO81, McC85b]. univers [Bla93]. Universal [Hun84, Rus89a, Sla82, Wes67, Her88]. Universe [For71a, For72, Hei72, Kel68, Nee85, Rup89, Sch79a, Wil87c, Bla93, Blo84, BH84b, Car88, Dea83a, Hir83, Jar84b, KA81, Kil94, Nad83, PB82, Rud88a, SG96, SN93, Smi82b, TFC93, Van85, vH97]. Universitaires [Cal63, Ege89, Hal65a, Hil71a]. Universität [AH93, For67b, Kra91, Ree85]. Universitäten [ZH92]. Universitätsbibliothek [Hom74]. Université [GG77b]. Universitetsforlaget [Bad83]. Universities [B.69, Bro84a, Cun89, Mor75b, Old77, Por76b, Smi80b, Mor83a, Nyh95, Sir87, Wei83b, ZH92, FB75]. University [Aas87, Agr83a, Agr83b, Ait67, Ait72, Ait85a, Ait89, All84a, Ant64, Arm67, Arm70, Bad71, Bad89, Bal84, Bar79, Bar76b, Ben79, Bib74, Bod65, Bod84, Bow87a, Bow89a, Bow83a, Bow85, Bow86, Bow87b, Bow87c, Bow88a, Bow88b, Bow89b, Bow89c, Bra84, Bra89a, Bro67, Bro69, Bro84a, Bro85a, Bro89c, Bro72b, Bro73b, Bro75, Bro82b, Bro78a, Bro78b, Bro86c, Bro89e, Bro83b, Bro83a, Bro84c, BH84b, Bro85d, Bro85e, Bro86d, Bro89f, Bru72, Buc77, Buc71, Buc75, Buc66, Buc73b, Bud84, Bur71, Bus85, Bus74, Byn74, Byn88, Cah88, Cal77, Can77, Can80b, Can80c, Car70, Car84, Cas85, Che78, Chi64, Chi65a, Chi68, Cla86a, Cla71, Clo80, Col81a, Col87b, Col88b, Col81b, Coo81, Coo83, Coo89a, Coo84a, Cow74, Cow68, Cri71, Cro85, Cro87a, Cro89a, Cro89b, Cun78, Cun83b, Cun88]. University [Cun89, Dav71, Dea83b, Dea86a, Dea86b, Dea86c, Deb66, Des88, Din63a, Din64, Dol70a, Dol70b, Dol71a, Dol71b, Dol74, Duc84, Duf81, Duf83, Duf75, Dur81, Dur82, Eag78, Ear70, Ear87, Edg87c, Edw78, Eme87, Eva89, Far67, Far69, Far76, Far77, Fel89, Fis81a, Fis81c, For68, For74a, For87a, For87c, Fox68b, Fox69a, Fox70, Fox74, Fre70, Fre74, Fre78, Fre86, Fri87, Fur85, Gab83a, Gab75, Gab83c, Gab88, Gel76, Geo66, Geo87, Gil88, Gin84, Gli83b, Gol87b, Gol88b, Goo68, Got67, Got71, Gou82, Gou85, Gow73, Gow78, Gow69, GG75a, GG77b, GG84a, GG84b, GG87, GG89, Grm69, Gur78, Hac89, Had63, Hah73, Hai77, Hai82, Hal66a, Hal69a, Hal71, Hal77b, Hal80a, Hal84e, Hal84d, Hal85a, Hal85b, Hal86a, Hal86c, Hal86d, Hal87b, Hal88b, Har84a, Har75, Har89a]. University [Har76, Har82d, Har83b, Hei71a, Hei71b, Hei73, Hei74a, Hei74b, Hei76a, Hei76b, Hei77, Hen81, Hen82c, Hen83a, Hen83b, Hen84f, Hen86b, Hen87b, Hen88a, Hen88b, Hen88d, Hen89a, Hen89b, Hen89d, Hen89e, Hen83d, Hen84j, Hen84k, Hen84l, Hen84m, Hen85c, Hen86e, Her80, Hes74, Hil71b, Hil89b, Hil89a, Hoc88b, Hod70, Hod89a, Hol85a, HS83, Hol71, HS88, Hom86b, Hop73, Hop75, Hos67, Huf85, Huf86, Hun84, Hut84, Hut89, Irw83b, Jac89b, Jac85, Jam87b, Jev67, Joh81, Jon87, Jor78, Jor79, Jor85a, Jor85b, Jor87, Kar76, Kau84, Kee89, Kei84, Kei86, Kel70, Kel84, Kel85a, Kil84a, Kni69a, Kni71c, Kni72, Kni74, Kni75, Kni76, Kni80a, Kni80b, Kni81, Kni83b, Kni85, Kni89d, Kra84a, Kra86, Kri83, Kub70, Lar74, Law80b, Law80c, Law81, Law83, Law85, Law87, Lay74]. University [Lef78, Les83, Lew85, Lie85, Lin89b, Loh75, Lor73, LM74, Lor75, Lor77, Lud80, Mac80, Mac85b, Mac72, Mah89, Mar74, Mau77, McC77a, McC77b, McG65, McG68, McG70a, McG71a, McL86, McN83, Men70, Men74, Mer78, Mol69, Mol71, Mol74, Mol78b, Mol79, Mon78, Moo84, Moo85, Mor69a, Mor69b, Mor72d, Mor73a, Mor73b, Mor75a, Mor75c, Mor77a, Mor77b, Mor78, Mor80, Mor82a, Mor83b, Mor84d, Mor87b, Mor89a, Mos77a, Mos77b, Mül87, Mus74, Nad83, Nay75, Nay80, Nay85, Nee85, Nee69, Nic85, Nor68a, Nor68b, Nor69a, Nor70a, Nor71b, Nor73a, Nor77, Nor80, Nor82, Nor83, Nor85b, Nun88, Nut84, Olb69, Olb70, Olb71a, Olb71b, Osb85, Osp74, Out82a, Out83, Out85, Pac72, Pat86, Pep72, Phi69, Pic76, Pic80, Pic88a, Pic88b, Pin83, Pin88, Por78b, Por79b, Por81]. University [Por82, Por83a, Por84b, Por85a, Por85b, Por87b, Por87a, Por87c, Por89g, Pum84, Pum87, Pum88, Pye87, Ran74, Rap71, Rap79, Rap86a, Rat65a, Rat68, Rav70, Rav71, Rav72, Rec86, Ree80, Ric89, Rid83, Rob87, Roc87, Roc88b, Rog83, RH86, RH88, Ros79, Rud69, Rud71, Rud84, Rud86, Rud88b, Rup85, Rup89, Rus87a, Rus89a, Rus66, Rus69, Rus77b, Rus77c, Rus85, Rus87b, San87, Sch81a, Sch84b, Sch84c, Sch86a, Sch87, Sch69a, Sch69b, Sch71, Sch76, Sch78a, Sch78b, Sch80d, Sco67, Sec84, Sha78, Sha73a, Sha83a, Sha86, Sha73b, Sho87a, Sho86a, Sho88, Sin76, Sin84, Smi71, Smi73b, Smi74b, Smi75a, Smi78a, Smi83, Smi85a, Smi85c, Smi85d, Smi87b, Smi87c, Smi87a, Smi89e, Sne74, Spa65, Ste75, Sti89, Str73, Swi71, Tei78, Tei83b, Tei83a, Tei84, Tei88, Tha66a]. University [Tha74, Tha83a, Tha87, Tho89, Til75a, Tup82, Tur70b, Tyl75, Und66a, Und66b, Und75b, Und79a, Und79b, Und80, Vea64, Ver89a, Ver89b, Wal89, War88a, War88b, War89, Was69, WV92, Wea84, Web75, Web78, Web65a, Web65b, Web66c, Web67a, Web68a, Web68b, Web69a, Web70, Web87, Wei85a, Wei85b, Whi70a, Whi73, Whi75, Whi85a, Whi89b, Whi77, Wig77, Wig78, Wil75, Wil84c, Wil85b, Wil86a, Wil86b, Wil87c, Wil78, Win75, Woo89, Woo67, Yeo88, You66, Yox82a, Whi72a, Ken86, Mor91b, AH93, Bec86, Goo74, Kra91, PS78b, SH73, Tho74a, WG81]. University-Industrial [Hen87b, Ken86]. Unknown [Jam88, Bec88, vPvP90]. unnumbered [Tur74a]. Unpublished [Hom74, Lev68, Til75a, Ber93b, Gou84b, McG78]. Unseen [Hei72, Sta91]. Untersuchungen [Mol72a]. untitled [Bad68, Kel85b, Por85c, Tur70d]. Unwin [Dur77, Rob75]. unwitting [HK83, Smi85d]. Upon [Cam82, Buc85b, Eag78, Fau89, May66]. Uppsala [Fis81c, HS86b, Mor78]. Upright [San87, Hei86]. Upton [Vic88]. Urania [Mai85, Kit89]. Urania-Verlag [Mai85]. Uraniborg [TC90]. Uranium [Hen81]. Uranus [Aus67]. Urbana [Duf83, Mor77a, Rap71, War89]. Ursprung [Tur75]. Ursum [JK84]. Ursus [Ait85a, JK84, Mor87a, Ros86]. USA [McC85b]. usages [BD95]. Uschmann [Mai85]. Use [Tha74, Tur81, Buc85b, Chi68, Ger71, Law93, Mad88, Nut82b, Wal92]. Usefulness [Whi72a]. Uses [BD95, Dea86b, Bud93, Bur83]. USSR [RH86, Tup82, Vuc84]. utero [Wea78]. utilitarianism [dCC09]. utilitarisme [dCC09]. Utility [Gol88a]. Utopia [Web68b]. Utrecht [Tur70c, Tur73d].

V [Agr83b, Chi64, Dol71a, Dun86, Eas86, Fau89, Geo78b, Gra83a, Hal86c, Jar84a, Kee88, Kel70, Lin89b, Mac79, Mor72a, Nev84, Nor73a, Rap71, Stu71a, Tur88, Tyl75, Web87, Wuj85, Bra88, Bro78a, Bro78b, Fig78, Gou84a, GG74a, Har83a, Kni83c, Mol78b, Nor71a, Nor72, Sha78, Tre87a, Woo80e]. V. [Hac88b]. Vachon [Bur74a]. Vacuum [Kel67, Hen83d]. Valencia [Kni88c, LN95, LN95]. Valency [Kni72, Rus66]. Valentin [Tei70]. Valgus [Geo78b]. Vallon [Hil86a]. Valt [Geo78b]. Values [Por87a, Mor83c, Wis95]. Vandenhoeck [Hom74, Wil84b]. Vanderbilt [Olb70, Roc88b]. vapeur [Tan82]. vapours [SE91]. Variant [Whi73]. Variation [Por85d, Pum89]. Various [Bar75]. Vasco [Stu71a]. Vascular [Fre74]. Vasil’evich [Hal71]. Vatican [Dun86, Maf91]. Vaudrey [Sme87]. veka [KI92]. vel [Mol73]. Veneto [Sch73a]. Venezia [Sch81b, Sch81b]. Venice [Sch81b]. veranstaltet [CM90]. Verbindung [CM90, Kle94]. Verdenk [Sme68b]. Verhandelingen [Phi67]. Verlag [Bos72, Bra88, Bro89a, Bro89b, Cap81b, Fig88, Fis78, For67b, For69a, GG75c, GG77a, Gra83b, Gra85b, Gra87b, Har81, Hen86b, Hen88c, Kra86, Mah89, Mai85, Mau87, Mol75, Mol78a, Mol88, Mor84g, Nor72, Pat84, Rus72b, Rus74, Rus79a, Sch75, Sch83b, Sme68a, Tur75, Tur88, Whi68a]. Verlagsbuchhandlung [Wes70]. Verlagsgessellschaft [Jac89a]. Vernon [Car78]. Veröffentlichungen [Ree85]. Verrier [Wil84c, Ros82]. vers [PP95]. Version [Sch69a, Til75a, BEoCF92, ECL92, The78]. Verständnis [Gol91]. versus [Ber93b, Cro87b, Ear89, McL86, Moo89a]. Vertebrate [Rup88, Buf87, Rai91]. Vertreibung [Dei92]. Verwey [Byn87]. Vesalius [Rat65a]. Vess [Jor78]. Vestiges [CS94, Hod70, Bro77]. Veterinary [Kni86c, Pel85a, Ald85, Pat83b]. Vetus [Mol84b]. VI [Hei76b, Arm64, Arm69, Bow86, Bro78b, Bry74, Che78, Dug77, Got67, Gra84c, Gre71, Hai80, Hen84e, Hen86b, Hen89c, Hod70, Hom74, Kni85, McG71a, Mol85, Por85a, Rob80a, Sar81, Sch83a, Und66c, Whi68b, Wil85a, You83]. VI.3 [Nut84]. vibration [CD81, Gra84c]. Vickers [Dea86a, Gol88b]. Vico [Rap86a, Ros84b]. Victor [Mor83d, Und64a, Tor80]. Victorian [Bro75, Buc89a, Bus85, Can77, Lay85, Mac82, McL86, Mor72b, Mor80, Mor83c, Mor84c, Ora80, Pic88a, Por87a, Rus87a, Sha77a, Sta89a, Tha84a, Tha87, Web76, Yeo88, Des82b, Des84, Hei72, Hey81, Jac81, Jen70, Lig87, Mer89, Pen89, Ric88b, Rus89b, Sec86, Sho96, Sho86b, Shu96, Shu04, Van93, Wil87g, Yeo93, You85]. Victorians [Bow89e]. Video [Hen84c]. Vie [Cro79, Osb89b, A^{+}87, BH90, dG01]. Vienna [Hen88c, MR87a]. Viète [Whi85a]. View [Goo75, Rob89a, Rus88b, Abe86, Bar89a, Goo83a, Har67, Hei93, Kni86a, Kni88d, Kni89a, MJ89, Rid83, Sti84]. Viewing [Mac90]. Views [Koh75, Rud84, Sho87b, CT95, DJS97, Gre82]. Vignoles [Bal84, Vig82]. VII [Hop73, Pag69, Aus86, Bar79, Bow85, Chi68, Dol70b, Fis78, Fre78, Gid82, Gow78, GG71, Har75, Hen86b, LM74, Mac78, May66, McG70a, Mol86, Mol89, Moo85, Nic88b, Nor72, Rav67, Ric88a, Rup85, Rus79a, Sch69a, Sch70, Sha86, Sme84a, Swi71, Tre88, Whe89, War90]. VIII [Cro79, Hop73, Mol85, Ait73, Arm62, Ben79, Bow89a, Boy70, Bro69, Bro72b, Can74, Can77, Col81a, Col76, Col88a, Cro85, Dau75, Dea83a, Deb66, Des88, Dol71b, Far71b, Fis74, Fis81a, For71b, For87b, Fox70, Fox74, Gab88, Got66, Hac82a, Hac88b, Hal71, Hal84d, Har89a, Har87b, Hen84a, Hen86b, Hen88b, Hen88c, Hen84m, Hew66, Hom74, Hun86, Hut84, Jam87b, Jor87, Kar76, Kau84, Kei84, Kei89, Kni71c, Kni81, Kni89e, Law80b, Law85, Lev80, McG66, McG68, Mea71, Mol78b, Mor78, Mor83a, Nee85, Nor72, Nor83, Nut82b, Por78a, Rau75, Rus66, Rus68, Rus77b, Sch84b, Scr68a, Sha73a, Sha77a, Sin81, Smi71, Smi85c, Tei83a, Tei88, Til76, Und79a, War88a, Wea78, Wig77, Wil84c, Wil87c, Woo84, Woo89, WG81]. Vilhelm [Fri89]. Vilhjalmur [Lev88b]. Villard [Bec88, Bec91b, Tee89]. Villars [Osb89b]. Vinci [Cun83a, Kel70]. Vineyard [Ste62b]. vintage [HS83]. Virginia [Sel10]. Virtues [Car74]. virtuosi [Woo82]. Vis [PF67]. Viscountess [CMNH92]. Vision [Pum88, AS89, Ath90, Eri94, Par87, Tur83c, Tur94]. Visionary [Hei77]. Visions [MR96, Ric88b]. Visser [Sch83a]. Vista [Tur77a]. visu [The78]. visual [Sta94]. Vita [Mah89]. vital [CD94, Cha91]. Vitamin [Ric91, Tei88, Car86]. Vitezslav [Bow85]. Viva [PF67]. Vivant [Olb72, Out80a]. Vlaamse [Phi67]. vocabulary [Rap82]. Vocation [Cro85, Out80d, Out84c]. Voetius [vR95]. voie [HJF89]. Vol [Agr83b, Ait71, Ant64, Bad71, Bro78a, Cla86b, Cro67, Dol69, Fie83, Hal86c, Hal88b, Hig65, Hil89b, Jar84a, Mor83a, Rat66b, Rat68, Rav69, Sha89, Sho87a, Sme69, Ste70, Tur68, Vea64, Wal77, Whi68a, Woo83, Ano82n, Ano82m, Bro89a, Cro89a, Far69, Gou85, GG84a, Hal86a, Kel68, Kni71b, Pag69, Ree85, Sch69a, Sch73a, Sch80b, Smi80b, Ste62b, War88a, BH87, EJR95, KN88, MR82e, NLH^{+}84, STT87, SCS^{+}87]. volcanic [Woo85]. Volcanism [Fri91]. Volcanoes [Rud74]. Vols [Wig74, Che78, Dun87, Edw78, Hal77a, Kni71b, Kni80b, Mol78a, Mol78b, Mol88, Nor72, Rud83, Rud89, Smi78a, Twe89, Und79b, Web68a]. Volta [Per92]. Voltaire [Duc88, McG69, Por84a, VB0 ]. Volume [Ano82j, Ano83j, Arm62, Arm64, Asp88, Bro67, Bro78b, Bry76, Bry78a, Buc85a, Car66, Cor78, Cow68, Cro70b, Dol71b, Dor79, Dyb78, Fis83b, For68, Fox68b, Fox70, Fox71a, Fox79, GG87, Hac74, Hac77, Hal86a, Har75, Hat87, Hei76b, Hen84g, Hen86b, Hen88d, Hen88c, Hen89e, Hom74, Hop73, Hop75, Hos67, LM23, Lie85, Mac88a, May67, McC80, McC83, Mol74, Mol84c, Mol84b, Mol89, Mor83a, Nak75, Nor68a, Nor69a, Nor70a, Nor73a, Pac72, Pat86, Pep72, Por89d, Pum87, Rud86, Rud88b, Sar81, Sch80b, Smi66, Smi89c, Swi71, Tha74, Tur73d, Tur74a, Tyl75, Web66c, Woo80c, Woo67, Ano62a, Ano62b, Ano62c, Ano63a, Ano63b, Ano63c, Ano63d, Ano64a, Ano64b, Ano64c, Ano64d, Ano65a, Ano65b, Ano65c, Ano65d, Ano66a, Ano66b, Ano66c, Ano66d, Ano67a, Ano67b]. volume [Ano67c, Ano67d, Ano68a, Ano68b, Ano68c, Ano68d, Ano69a, Ano69b, Ano69c, Ano69d, Ano70a, Ano70b, Ano70c, Ano70d, Ano71a, Ano71b, Ano71c, Ano71d, Ano72a, Ano72b, Ano72c, Ano72d, Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74e, Ano74b, Ano74f, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d]. volume [Ano83e, Ano83f, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Bar83, Goo85b, Hen82c, Hen85b, HP85, Mor77b, Spr97, Tay87, Wei92, Cla95a, NS84, Hen88d]. Volumes [Can74, Ait71, Bro73a, Duf75, Hei73, Hen84g, Mac88a, NBR^{+}95, Whi73, Can74]. Voluntary [Bec86]. volunteers [Fis98]. vom [CM90, Gol91, Kle94]. voor [Bry77, Phi67]. vor [Ger89].

Vorgriechische [Bos72]. Vorkin [Hen83b]. Vorkommen [Tur75]. Vorlesung [Kra91]. Vorlesungen [Bos72, Pat84]. vormen [Tur82]. Vortex [Gin74]. vortices [Sie91]. Vorzimmer [Cow74]. Voyage [Con86, Sta84a]. voyages [DJS97, MR96]. Voyaging [Bro95c]. vrienden [Phi67]. Vrin [Bos70, Cro79, Hal77a, Sch84c, Sch86b, Sch78c, Str75]. vs [Man88]. Vucinich [RH86]. Vulcan [But96]. vulgarly [BDH96]. Vulkanismus [Fri91]. Vulkanismus-Neptunismus-Kontroverse [Fri91]. Vulkanismusstreit [Fri91]. Vygotsky [vdVV91].

W [Ait89, AT87, Ant64, Bib74, Blo79, Bow86, Bro89e, Bry77, Byn78, Car70, Chi68, Chi71, CMN95, Coc75, Cri66, Cro87a, Din63a, Duc84, Dur83, Ede65, Far67, Far76, Fis81a, For74a, For82, Fox74, Fre86, Ged71, Gli83b, Gra83b, Gra86b, Hac76, Hal86e, Has72, Hen85a, Hen89g, Hom80, Hop72, Ilt77, Jor79, Kni71b, Kni85, Kni89e, Kra86, Lew85, Loh75, Mah89, May67, McG70a, Mor75c, Mor84c, Nun88, Pep86, Rus66, Sch83a, Sch69a, Sme70, Smi66, Spa65, Sti89, Tur70c, Tur73d, Tur74c, Und66a, Web75, Web76, Web84, Web71a, Whi68b, Yox83b]. W. [Bro89e, Bro86e, Gol87b, Hut81, Kni86d, Law89, Nye86, Pel85b]. W.6 [Ste62b]. Waard [Hal85c, McC80, McC83]. Waerden [Gra87b]. Wagner [Whi93]. Wagner-Jauregg [Whi93]. Wahlverwandtschaften [Adl87, Fig88]. Wake [BM76, Seg91]. Wales [Nev80]. Waley [Ano64l, Und65]. Walker [Fre81b, Har84b, Kni89e]. Wall [Vea64]. Wallace [Bro85d, Dea86c, Har75, Mol74, Mol84c, Nay80, Smi73b, Bro84b, Dur79b, Hug89, Smi72, WS91a]. Wallingford [Mol78b]. Wallis [Eag78, Fau89, Har79a, Whi68a]. Walter [Bra87, GG81, Hal85d, Hen84k, McG68, Mol72a, Nor71a, Scr68a, Whi68c, Wig74, Deb84]. Walters [Dea83b, Gre71]. Walther [Mor74b]. Walton [Scr68b]. Waltzer [Wea79]. Wang [Hal88b, Mol71, Pac72]. War [Bar89c, Edw96, Gur78, Num95a, Por89d, Sim88a, Swa83, Hod92, Luc90, PLP92, PCA91, Sch81a]. Warburg [Tei83a]. Ward [Nut83, Rat69]. Warfare [Wil87b, Hac84]. warming [Mar10]. Warner [Pic88b]. Warren [Joh81, Kni83a, Rup89]. wars [Gre86]. Warsaw [Tur74a]. Wartofsky [Blo79, May67]. Was [Hel63, Bur75, Can89, Gru69, Hal93]. Washington [Gli83b, Goo85b, Hac74, Hac77, Ink88, Kni83a, Mac72, Nut84, Woo85, Wor88]. wasn’t [Bur75]. Wasserstein [Nut84]. Water [Hal86b, Hil86b, Nay85, Per73, Sha84, Smi75a, Dra81, Ham90, Ole84]. Water-Lifting [Smi75a, Hal86b, Ole84]. Waterloo [Duc84, Gin84, GG89, Sha83a]. Waters [Coc75, Lem77, Nut84, MW94, Ste91a]. Waterson [Rud80]. Watson [Bar75, Yox83a, Yox83b]. Watt [Mor71, Swi70, Pay74]. Watts [Rav66b]. Wave [Bru70, Hen88c, Hun89a, Buc89b, Jam84, Mac85b, MR87a, OP87, Whe83]. wave-particle [Mac85b, Whe83]. Waveforms [Hac88b, Phi87]. Waves [HS85, Buc94]. Way [Jev64a, Jev64b, Sey77, Gue84, HJF89, Pau93]. Wayne [Hes74]. WC [Pat86]. WC2E [Whi85a]. Weapons [Sim88a]. Weare [Kni89c]. Weart [Hen81, Mor77b]. weather [Fri89, Neb95]. weavers [Cha84]. Web [Bud89]. Webster [Coo81, Har83b, Por89d, Wig78]. Wedgwood [Cha75]. Weekes [Twy88]. Weidenfeld [For71a, For82, Nay77, Web71b, Yat72, Yox83a]. Weighed [HE92]. Weighing [McC68, Bry81]. weight [Høy94]. Weights [Hac88a, Con87, Tur83a]. Weikersheim [Wey92, Wey92]. Weil [Rus74]. Weimarer [Mor84g]. Weindling [Ink81]. Weiner [Dor80, Hen81, Men70]. Weinheim [Fis78]. Weisheipl [Col83]. Weiss [Hal84c]. Weisz [Bro84a, Out82a]. Welch [Por85a]. Welfare [McC77b, Sch68a]. Wellcome [Hen89g, Vea64, AL87]. Wellsprings [Gou95]. Welsh [Hug89]. Welt [Sch93]. Weltalls [For69a]. Weltauffassung [Hei93]. Wen [Ron86]. Wen-Yuan [Ron86]. Wende [Kle94]. Were [Wei80]. Werk [Gra85b, Phi67]. Werke [Kra86, KvDC^{+}37, vPSBD^{+}04]. Werke/Collected [Kra86]. Werner [Bar75, Bro72a, Kra86, Mau87, Sme68a, BEFH10, Nor66, Osp76, Rud62]. Werskey [Cro72, Dur77]. Wesleyan [Hei74a]. West [Hew66, Hil89a, Mol71, Huf93, Huf95]. Western [Bod84, Gou85, Kni89d, MR88, Nee69, Rav70, Tur78b, Hay94, Tra93, C^{+}95, Lin92, ML89, Ols90]. Westfall [Dol74, Rog83, Whi72b]. Westman [Woo80b]. Westport [Gab83b, Kni83a]. Westview [Har82d, Pel85b, Whi89a]. Wetenschappen [Phi67]. Weyant [Sha83a]. Wheatland [Tur70b]. Wheaton [Hen83b, Mac85b]. Wheeler [Kni83c, Sha89]. Wheels [Smi75a]. Where [BS77, Sap90]. Whewell [Bro77, FS91, Fis91a, Har69, Nor69b, Yeo93]. which [Byn80, MJ89]. Whiston [Sch86a, For85b]. Whitby [Ora77]. White [Bar75, Cro89b, Dol73b, GG77b, Lar74, Raf88, Rus77c, Sch69a, Str75, For71a, Sou93]. Whitehall [Yox82a]. Whitehead [May66]. Whiteside [Hei76b, Nor68a, Nor69a, Nor70a, Nor73a, Pep72, Til75a, WHS92]. Whitley [Pic84a]. Whitman [Whi73]. Whitmore [Rus77c, Str75]. Whitney [Fri87, Mar89, Wis85]. Whitrow [Bro73a, Fox79, Hei68, Hen84g]. Whitteridge [Sch73b]. Who [Gru69, Nor85a, Wil69, HS83, TFC93, Yox83c]. whole [Dea83b, GG77a]. Whorton [Coo84a]. Whyte [Nor68b, Nor73b]. Wickey [Col88a]. Wiener [Asp88].

Wiesbaden [Pag69, Whi68a]. Wightman [Ant64, Sme84b]. Wigner [Hen89b]. Wiksell [Bro84d, Cha80c, Cro67, Fis81c, HS86b, Mor78, Nic88b, Rus67]. Wilberforce [Jen88]. Wilcox [Hen88a]. Wild [Hil71b]. Wilda [Sha86]. wilderness [Won93]. Wiley [Har77, Law80a, Pra70, Sme66, Smi73a]. Wilfred [GG89, Car78]. Wilfrid [Duc84, Gin84, Sha83a]. Wilhelm [Bro89a, Por87b, FMF85, Har96, Hol77, Smi82c]. Wilkie [Cro84a]. Wilkins [Ait71]. Willem [Dea83a, Wil87b]. William [Asq80, Bau75, Bow83a, Bro71a, Bro83b, Buc69, Buc68, Chi65a, Cow68, DeK73, Geo66, Gol87b, Goo75, Har75, Hea87b, Hei76a, Hom86b, Kni86c, Kni88b, Mar89, Mol74, Mol84c, Mor72b, Mor73a, Mor81, Mor83b, Nay80, Nor82, Nye86, O’H86, Ora77, Pic76, Pum88, Rat66a, Rau75, Rud84, Rup83, Sch86a, Sch69a, Sch73b, Sch74b, Sec84, Sha73b, Sin81, Smi73a, Smi78a, Sti89, Sty68, Tha83a, Und75b, Und80, Web66a, Web67a, Whi68c, Wil84e, Woo80b, Ald85, Bir86, Bow87c, Bro85c, Bro77, BPH85, BP02, CL87, Dav89, Dea86c, Fel89, FS91, Fis91a, For85b, For87c, Gol88a, Har69, Jon88, Jor85b, Kei86, Liv85, MD88, Nor69b, O’D83, Olb87, Par87, Ric89, RT95, Rud83, Rud89, Sch90, Sme84b, Smi76b, Smi89a, Stu71b, Tay96, Yeo93]. Williams [Bad89, Bro83b, Hei73, Kel84, Kni65, Swa86, Wal90, Wig69]. Williamson [Kei89]. Willibald [Hun89a]. Willink [Cro70b, Fox71a]. Willis [Fri87, Wis85]. Willson [Ste62b]. Wilma [Bro83a]. Wilson [Bra89a, Buc89a, Bus74, Har84b, Pac70, Rat68, Rud71, Smi78a, Stu85]. Wiltshire [Hen84c]. Winchester [Bro87]. wind [Hil94]. windmill [Hil94]. Windmühlen [Tur75]. Windscale [Wil84d, Wyn82]. Windus [Wes74]. Winfried [Gra86b]. Winifred [Sch75]. Winkelteilungen [Whi68a]. Winter [Sch69a]. wire [BC92]. Wireless [Hac76]. wirtschaftswissenschaftlichen [Bro82a]. Wisan [Sch75]. Wisconsin [Bow86, Bro85e, Col87b, Deb66, Fel89, Hal69a, Hop73, Hop75, Hos67, Jor85b, Lew85, Nay75, Nor71b, Rav71, Rus85, Web66c, Wei85a]. Wisdom [Hen89c, Whi70a, Max84]. Wise [Fri87]. wiskunde [Sme68b]. Wissenschaft [Nor71a, Pag69, Mor85]. Wissenschaften [Bos72, Mol86, Pag69, Wes70]. wissenschaftlich [Hei93]. wissenschaftlich-philosophische [Hei93]. Wissenschaftliche [Jac89a, Sim84]. wissenschaftlichen [HRP90, Sch83b]. Wissenschaftlicher [Hen86a, HP85, PHvMW93, PvM85]. Wissenschaftsgeschichte [Cap81b]. wissenschaftshistorische [Mol72a]. Wissenschaftskultur [Mei91]. Wissenschaftsphilosophie [Mai88]. Wissenschaftspopularisierung [Dau98]. Witch [Mon78, Hüb90, Sha81]. Witch-Hunt [Mon78]. Witelo [Eas86, Mol81]. Witelo. [Eas86]. Witelonis [Mol81]. within [Kuk91, Kuk93, Sch69a, She95]. without [Wea79]. Witkowski [Bow87b]. Witmer [Whi85a]. Wittop [Tur82]. Wodeham [Mol89]. Woerden [Dea83a]. Wolf [Hoc86]. Wolfcang [Jac89a]. Wolfcang-Hagen [Jac89a]. Wolfe [Tur87a]. Wolff [Les83, Mol72a]. Wolfgang [Wey92, Hen86a, HP85, PHvMW93, Por89a]. Wolters [GG81, Von88]. Woman [Col81a]. womanhood [Rus89b]. Women [All80, Fee77, HH97, Ogi86, Out87, RCRC98, SGŚ00, Ali86, Ano07, Sht96]. Wood [Mol89]. Woodfield [Ben79]. Woodruff [Edg84]. Woodward [Mac79, Por79b, Sti89]. Woolf [Kni83b, Sel10]. Woolgar [Irw82]. Word [Kni76, Gil83, Kro91]. Words [Jen87, Rap82]. Work [Bow88b, Bry76, Bry78b, Cro70b, DeK73, Dra86, Fox71a, Fri74a, GG74b, Hal69a, Han70, Hen87a, Ink87, Kel67, Mea71, Men81, MC81, Mil87, O’H86, Olb69, Rus67, Sin74, Tre87b, Tur88, Web76, Web68a, Wil62, Ack87, Ash85, Ber85, Can93b, Car67b, CHG91, DFG87, FL^{+}94, GJ85, Gra84a, Jak84, Lev85, Mac85a, Mol71, PS82, She95, Sou93, dGB^{+}89, vLP82]. Working [Tyl64]. Works [Ait71, Asp88, Bro89e, Bry83, Car67a, Dea83a, Dor79, Fox70, Gab83c, GG87, Hen83c, Kni71c, Kni86b, Kra86, McC83, Sch83a, Sch74a, Swi71, Tha66a, Tur74a, Tur81, Twe89, Wig78, AT87, B^{+}95, BCKR89, BN95, Cun83a, DD95, Gje84, GF^{+}03, KvDC^{+}37, Lan87, LM23, PN95, Rav68, VB0 , vPSBD^{+}04, LF97]. workshop [JO81]. World [Bra87, Cle83, Cun88, Ear89, Gli83b, Gol87b, Goo75, Hen84c, Hen89a, Hen86e, Mor74b, Num87, Rob89a, Rud84, Sho86a, Tha84b, Tho76b, Wil69, Ale85, Ben75, Bla93, BP93a, BP02, Cro90b, Dol96, Edw96, FGJ98, Fis83c, GB95, Goo83a, Gre82, Hei93, Kni86a, Kni88d, Kni89a, Kri88, McC92, Mor91a, Mor90b, Olm92, Opp85, Pai86, Ram88, Rid83, Ron83, Rud92, Sch90, Sch93, Smi82a, Wil95, Zam86, Num95a, Swa83, BPH85]. World-New [Cun88]. world-view [Kni86a, Kni88d, Kni89a]. Worlds [Arm69, Sch87, Yox82b, Cro86, Dic82, Pum84]. Worse [DPP88]. Worshipful [Fre81b, Hal88a, Vea64]. Worstedopolis [Mor85]. Worth [Gab83a]. Wranglers [Jam87b, Har85a]. Wren [Hen83a, Ben75, Ben82a]. Wright [Bow87c, For72, Law85, Mus74, War88b, Pro86]. Writer [Bro86e, Dem83]. Writings [Ait89, Bau75, Buc66, Dun87, GG84a, Grm69, Hei74a, Hom74, Kra86, CD94, CS94, CMN95, dLT96, Le 90, LS76, May66, Mol78b, NN95, Sme66, WS91b, WS91a, vLMB88]. Written [Yox79]. Wroclaw [Eas86]. Wrolcaw [Mol81]. Wundt [Smi82c]. Wussing [Gra85a]. Wyckoff [Sch69b]. Wylie [Bow87b]. Wynne [Wil84d]. Wynter [Nut82b, Tur77a].

X [Aas87, Ait86, Aus86, Ben84, Ben89, Bra89a, Che85, Col88a, Dea86c, Gab88, Gra87b, Hal86e, Har84b, Hea87b, Hen84j, Kni88a, Liv86, Mül87, Mül89, Nee85, Out84a, Pin88, Por89c, Red84, RH88, Sec88, Sha89, The87, Tur70c, Tur73d, Wea84, Whi85a, Wor88, Agr83a, Arm67, Arm70, Bad71, Bad89, Bar73a, Bea85b, Bib74, Bod65, Bra84, Bra87, Bro86a, Bro75, Bro83a, Bry81, Buc75, Bud84, Byn72, Can75, Can80b, Chi64, Coc75, Cro68a, Cro89b, Cun78, Din63a, Dol73b, Dol74, Dol83, Dur79a, Dur81, Dur82, Fis81b, For69b, For85a, Fra87a, Fra87c, Fre74, Gab83c, Gau83, Gel76, Gin74, Goo81b, GG84b, Gra83a, Gra86a, Hac89, Hal87b, Hei76a, Hen81, Hen85a, Hen89b, Hen84l, Hen87c, Hen88e, Hoc88a, Hoc88b, HS86a, Hun84, Kee88, Kuh68, Law83, Les83, Loh75, Mar74]. x [Mar73, Mol72a, Mol88, Mor72b, Mor75b, Mor77c, Mos77a, Nad83, Out82a, Out85, Pel85a, Pep71, Por79b, Por89d, Pum84, Ric89, Rid83, Roc88b, RH86, Ros79, Sha83a, She85, Sho88, Sme68b, Smi73a, Smi74b, Smi83, Ste62b, Swi67, Tha67, Til75a, Tur69, Tur88, Und66b, Und75a, Ver89b, Wal74, Was69, Web76, Web68b, Whi88a, Wil65b, Wil84d, Wil86b, Wil89b, Woo80c, Yeo88, Yox83a, Bow89c, Hen83b, Hir83, Jen70, Stu71b]. X-Ray [Nee85, Hir83]. X-rays [Tur70c, Tur73d, Stu71b]. X.36 [Mol85]. X.36-XV.2 [Mol85]. X/8 [Dun87]. xciv [Fox79]. xcviii [McG71b]. XI [Fox68b, Ait89, All84b, Asq80, Blo79, Bro86b, Bro72a, Bro86c, Bry78a, Bry78b, Buc73a, Car84, Cla66, Cla67, Cro70b, Dol73a, Fox69a, Fox70, Fra87b, Fur85, Gil88, Gol87b, Got71, Gra86b, Gra87b, Gre88, Hac76, Hei75, Hen81, Hen83b, Hen84g, Hen86b, Hen84k, Hil86b, Jar83, Lin86, Mac72, May69, McC77a, McG65, Mor84d, Mor89a, Nev78, Nor85b, Pau76, Pic84a, Pye87, Rav68, Rav84, Rus77c, Rus87b, Sch69a, Sch71, Sey77, Sha79, Smi80b, Stu85, Web66a, Web70, Whi75, Wig69, Woo80c, Yox82a]. XII [Sch80c, Ait71, All77, Bal84, Blu79, Bos72, Bow89c, Bro71a, Bro84a, Bro76, Bro89f, Bry78c, Buc77, Buc69, Buc73b, Bur71, Byn80, Cah88, Chi65b, Chi65c, Coo89a, Dol70a, Dor79, Duc84, Dun87, Dyb78, Eyl68, FF66, Fie89, For66, For80, For82, Fri74a, Gol88b, Gou82, Gow71, Gra81, GG72, GG75a, Hah73, Hai82, Hal86e, Har84b, Har76, Har82d, Hea64, Hei73, Hen89d, Hen85b, Hen85c, Hen89g, Her80, Hod89a, Hol71, Huf86, Hut89, Joh78, Jud88, Kel71, Kel73, Kni72, Kni74, Kni84a, Kni84b, Kri83, Law80c, Law85, Lay74, Mac88a, Mah89, McC83, Men74, Mil87, Mol74, Mol79, Mor73b, Mor83d, Mus74, Nee69, Nor71a, Nye86, Olb69, Olb70, Olb82, Ora74, Ora77, Ora80, Out83, Pep86, Pin88, Ran74, Rat66b, Rav73, Rob87, Ron86, Rud84, Rus74, Sch68a]. xii [Sch74b, Sha81, Sha87, Sho86a, Sho89, Sin74, Smi76a, Smi80a, Smi89e, Sne74, Sti70, Stu71a, Stu85, Tha74, Tho76b, Tur70b, Tur73c, Tur74a, Tur74c, Ver89a, Web87, Wei85a, Whi68a, Whi72b, Whi84, Whi89b, Wil89a, Woo73, GG86b]. xii-f [Bal84]. XII-XIII [Sch80c]. XIII [Sch80c, Ben84, Bow89b, Bro85d, Bud84, Byn88, Dub72, Far71a, Goo75, GG81, Grm69, Hen81, Hen84d, Hen83d, Ink88, Jor83b, Kei86, Kel78, Kni86c, Kni89d, Lor77, McC80, Mor84b, Mor77a, Mor83c, Mül87, Nay80, Ner89, O’H86, Olb85, Osb85, Osb89b, Ped85, Pic87a, Por87a, Rav70, San87, Sch73b, Sch75, Smi82a, Tur68, Tur83c, Wal89, Web78, Web66b, Whi70a, Whi77, Got86a]. XIIIe [Bec91b]. xiiii [Bro83b]. XIIth [Ano68j]. XIV [McC83, Por78a, RH88, Aas87, Ben82b, Bow83a, Bro89b, Bro82b, BH84b, Bro86d, Bus74, Byn73, Cal77, Can80c, Chi65a, Chi71, Col83, Coo81, Cri71, Dea86a, Dea86c, Dra86, Duf81, Eva89, Far76, Fox71b, Fre70, Gil84, Gli83b, Goo88a, Hah70, Hai77, Hal86b, Hal87a, Ham84, Han85a, Har84a, Hat87, Hei74b, Hen82c, Hen87a, Hen89a, HS88, Hom80, Hom86b, Hun89a, Kee89, Kni69b, Koh78, Lar74, Law80a, Law81, Law87, Mea84, Men70, Moo84, Mor75a, Nev80, Nun88, Osb89a, Pac70, Phi69, Pin87b, Por78b, Por82, Por84b, Por85b, Rap71, Ree80, Rog88b, Rus89a, Sch81a, Sch74a, Sin76, Sty68, Tup82, Tur81, Und66a, Vea64, Wes74, Whi89a, Cla89]. XIX [KI92, Bry78a, Byn87, Cla86b, Coo83, Duc88, Fie86, Fig77b, Hal80a, Hal86a, Hen88c, Jev67, Jor85a, Kel84, Kni89b, Mol71, Mor80, Por86a, Rus85, Sch87, Sch78a, Sec88, Smi76c, Tee84, Tha87, The87, Tur78c, Woo67, Dra70, Rey90]. XIXe [PP95]. XIXème [Pet84]. xl [Hac77, Sme71, Web66c]. xlii [Can84, Sch69b, Whi73, You66]. xlv [Swe76]. xlvi [Hei74a, Sme69]. xlvii [Hen86b]. xlviii [Agr83b, Byn74, Ste75, Wel70]. Xth [G.62]. XV [Hal77b, Hal85c, Von88, Mol81, Bod84, Bro85e, Bro87, Bru72, Bry78a, Cla86a, Cro79, Cro88, Duf83, Fre86, Har79b, Har82d, Hei68, Hei77, Hil89a, Kni76, Lud80, McC77b, Out84a, Pic76, Por87b, Por89c, Pra70, Rav71, Ril85, Roc87, Sco67, Sha89, Sin77, Sme68a, Sme70, Sme71, Smi87c, Stu87, Tur68, Web69a, Wes70, Nee78]. XV.2 [Mol85]. xvi [Ait67, Ait71, All82, All84a, Ant64, Bow87a, Bow87c, Buc89a, Cal63, Cha78a, Dav81, Edg87c, Edw78, Fel89, Goo82, Gow73, GG87, GG89, Hac88a, Had63, Hen84f, Hen87b, Hig65, Hop72, Hut88, Joh83, Kni65, Kni86d, Kub70, Lor75, McG69, Mol84c, Mon78, Mor71, Mor76a, Mor84f, Mos77b, Mül89, Nor68b, Nor72, O’H86, Pic88a, Pin83, Rap86a, Rat65a, Red84, Rud70, Smi85d, Swa86, Tei83b, Tha66a, Tur70c, Tur73a, Tur87a, War88b, Wat73, Wei86, Wil69, Wil77]. XVIe [Tur73b]. XVII [Bos70, Bow74, Bro73b, Can86, Cha77a, Cha80b, Edw78, Fee77, Gab83a, Goo85a, Gou85, Gre63, Ins86, Kil84b, Kni80a, Kni83a, Kni85, Nor71b, Nor80, Pic84a, Por81, Poy70, Rus87a, Rus69, Sch78b, Sha77b, Smi87b, Wei85b, Wil75, Wil85b, Wil87b]. XVII-2 [Bos70]. XVIIe [Ben89, Hal72b, Grm90, Pic87c, Poi04]. XVIII [Por89e, Bro85a, Cla71, Clo80, Gab83b, Gab75, Hac82b, Hal68d, Hal84c, Har84b, Hen84g, Hen89b, Law88, OM75, Pel85b, Por89g, Rav66b, Rog83, Rup89, Sch76, Sha73b, Sin78, Tha85, War89, Web67a, Whi85b, Dra69]. XVIIIe [Cro66b, Hal72b, Hom86a, Poi04, Roc88a, SGŚ00]. XVIIIème [BaJANJ84]. XVIIIth [MP89]. XVIIth [MP89]. xx [Ant64, Arm66, Asp88, Bur80, Car70, Cas82, Cla86b, Cow74, Han70, Har81, Hen82c, Kel70, Lin89b, Mau77, Nor77, Sch84a, Sch79a, Sec84, Sei85, Smi73b, Spa65, Tur73d, Web68a, Wil86a, Wil78]. xxi [Ait72, Bar83, Cla68, Dou84, Ede65, Gin84, Hen82b, Jor85b, Kni69a, Kni80b, Mol78a, Mor83b, Out84b, Sch84b, Sme66]. xxii [Agr83b, Cas85, Dav71, Dea86b, Fig77a, Hei71b, Hil71b, Jac85, Mar74, Mol78b, Nor69a, Nor82, Ora76, Ste62c, Und75b, Und79b, Web71a]. xxiii [Con86, Cro69, Cun83b, Fau89, Hal69a, Hei71a, Jam86, Mor72a, Sch86a, She63]. XXIV [Mol78b, Bow87b, Col87b, Cow68, Cro70a, Edg84, Fin68, GG75b, Har79a, Hil71a, Mac80, Mac85b, Nor73a, Rat69, Tur81, Wea84, Web65a, Woo80c, Yox83b]. xxix [Hen86a]. xxv [Gur78, Hal86c, Har77, McG68, Pag69, Ser88]. xxvi [All79, Bar76b, Blo82b, Edw78, Kei84, Kel75, Law83, Mor73a]. XXVII [Phi67, Dun69a, Edg87a, Hen83b, Hop73, Sho87a, Sho89, Web65b, Web69b, Woo80c]. xxviii [Car74, Col81b, Eag78, Far77, Hop73, Jar84a, Nay85, Scr68b]. xxx [Cro72, Fis83b, Hac74, Hop75, Lie85, Mer78, Olb71a, Sch74a]. xxxi [Bru86, Dol71a, Mor75c, Nak75, Wuj85]. xxxii [Bro84c, Cro71, Eme87, For68, Hos67, Kil84a, McN83, Rat68, Rud86, Rud89, Stu78, Web84, Web72, Whi72a]. xxxiii [Cro71, Gow69, Hal88b, Mor87b]. xxxiv [Dur83, Goo85b, Hil89b, Kni83a, May67, Pep72]. xxxix [Nay75, Pin87a]. xxxv [Hal88a, Hei76b]. xxxvii [GG84a, Nor70a]. xxxviii [Bro78b, For72].

Yale [Bro75, Bro86c, Bus74, Duf81, Far76, Far77, Hei71b, Hen87b, Kei84, Nor82, Rud71, Smi73b, Tei78, Und66b]. Yannick [Gua89]. Yates [Hen85b, Tur73a]. Yazid [Win75]. Year [Ano75o, Ano76l, Ano77p, Ano78m, Ano80m, Ano81n, Ano82o, Ano83m, Ano84j, Ano85j, Ano86i, Ano87m, Ano88r, Ano89q, CC66, CE74, Hal79, Per89b, WE72, Lan87]. Years [AH93, Ano71h, Aus86, Bar89c, Bro87, Bus74, Cow74, Din63b, Ear87, Goo85b, Hac74, Hac77, Hen88d, Lud80, Mar64, Por81, Sec84, Sme70, Smi87c, Tur81, Wil62, All86, Bad83, BH87, Bur99, EJR94, EJR95, FS84, Gin84, Gle91, Gle82, H^{+}82, Hea86, Hea90, Hen84d, Her95, Joh83, Lie85, Maf91, Pep89, Rus86a, SCS^{+}87, WH91]. Yehuda [Can75, Kil84a, Law80a, Pic84a]. Yellow [Col87a, Fel89]. Yeo [Pin87a]. yesterday [MMVZ90]. Yi [Whi70a]. Yi-Fu [Whi70a]. Yoho [Din63a]. Yolton [McG70a]. York [Ait89, All82, Arm66, Baa90, Bar77, Bar75, Bos72, Bow89a, Bra87, Bro89b, Bro76, Bro89c, Bro85d, Buc69, Buc68, Bud84, Byn88, Car74, Cas82, Cha65, Cha77b, Dor80, Dun69a, Eme87, FF66, Far71a, Far71b, Fig77a, Fin68, Fis78, For72, For87c, Fre82, Fri87, Gab88, Gin74, Gol88b, Goo81b, Goo82, Goo85a, Gra81, GG81, GG87, Gra83b, Gra84c, Gra86b, Gra87b, Hal80c, Hal85d, Har77, Har79b, Har81, Har87b, Hea64, Hei75, Hen81, Hen83b, Hen86b, Hen86a, Hen88c, Hoc88a, Jac89b, Jar84b, Jud88, Kee88, Kni86b, Kni88b, Koh78, Kra86, Law80a, Law88, Lev73, Mah89, Mar89, May67, May69, McG68, Men70, Von88, Mol78a, Mol88, Mor87a, Mor72b, Mor74a, Mor83c, Nev78, Ora80, Pac70, Por89a, Por78a, Pra70, Rap75, Rat69, Rog88a, Rog88b, RH88, Rud86, Rus79b]. York [Sch69a, Sch73b, Sch74a, Sch75, Sch80d, Sha87, Sin74, Sin81, Sme66, Sme68a, Sme69, Sme70, Smi73a, Smi78b, Smi80a, Smi80b, Smi87a, Ste62c, Swi67, Tha85, Tre87a, Tur78c, Tur88, Und66c, Und66a, Und75a, Ver89b, Wat73, Wea79, Wea80, Web69b, Web72, Wei86, Whi72b, Whi68c, Woo73, Yox83a]. Yorkshire [Mor74a]. Young [Byn80, Cri71, Fie83, Pin87b, Rus87a, Tur83c, Wyl89, Bro89d, Pye85, Can70]. Yoxen [Nor85a]. Yu [Bro89b, Agr83b, Hal88b, Mol71]. Yuan [Ron86]. Yyves [Bur74a].

Z  [Kni71c,  Rav69,  Roc87].

Zabarella [Sch73a]. Zacharias [Gol92a]. zahlentheoretische [Whi68a]. Zambelli [Bra87]. Zandvoort [Pic87a]. Zanier [Sch79b]. Zarach [Hen84d]. Zassenhaus [GG75c]. Zdenek [Rya70]. Zeit [Fig88, Adl87]. Zeitlinger [Sch83a]. Zev [Hen84e]. Zhengzhou [Mad88]. Zhores [Har82d]. Ziggelaar [Str73]. zijn [Phi67]. Zimmermann [Sch83b]. Zinc [Gre71]. Zirkle [Und66c]. Zivilisationskritik [Mei91]. Zl [Mol81]. Zöllner [Mei91, Mei91]. Zoological [Blu79]. Zoologist [Geo66]. Zoology [Fre70, Smi78c, Win91]. Zott [Bro89a]. Zubov [Kel70]. Zucker [Sha89]. Zuckerman [Har79b, Law80a]. zum [Gol91, Kle94, Rav65]. Zupko [Tur83a]. zur [Bro82a, Cap81b, Gol91, GG77a, GG81, HRP90, Mai88, Mei91, Mol72a, Mol75, Nor71a, Pag69, Sch89, Sch84d, Whi68a]. Zurich [Hen88c, MR87a]. Zusätze [Dol71a]. Zwei [Mol72a]. Zweite [Wes70]. zwischen [Rus79a, Sch89]. Źyciński [Dun86].

References

Atran:1987:HDC

[A^{+}87] Scott Atran et al., editors. Histoire du concept d’espèce dans les sciences de la vie: colloque international, mai 1985. (French) [History of the concept of species in science and life: international conference, May 1985]. Editions de la Fondation Singer-Polignac, Paris, France, 1987. ISBN 2-900927-19-6. xiii + 324 pp. LCCN QH83 .H57 1987.

Aaserud:1987:BRA

[Aas87]  Finn Aaserud.  Book re-

view: Alan J. Friedman & Carol C. Donley. Einstein as Myth and Muse. Cambridge, etc.: Cambridge University Press, 1985. Pp. xiv + 224. ISBN 0-521-26720-X. £25.00, $39.50. British Journal for the History of Science, 20(1): 105, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                                    URL http://www.jstor.org/
                                                    stable/4026459.

Aaserud:1990:RSN

[Aas90]  Finn Aaserud.  Redirecting

science: Niels Bohr, philanthropy, and the rise of nuclear physics. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-35366-1. xiii + 356 pp. LCCN QC789.2.D4 A27

                                                    1990. URL http://www.loc.
                                                    gov/catdir/description/
                                                    cam024/89048317.html;

http://www.loc.gov/
catdir/toc/cam028/
89048317.html.

Austoker:1989:HPR

[AB89] Joan Austoker and Linda Bryder, editors. Historical perspectives on the role of the MRC: essays in the history of the Medical Research Council of the United Kingdom and its predecessor, the Medical Research Committee, 1913–1953. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1989. ISBN 0-19-261651-X. xi + 259 pp. LCCN R772.A6 H57 1989. US$45.00. URL

http://www.loc.gov/
catdir/enhancements/
fy0637/89003122-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0637/89003122-t.html.

Abeles:1986:DLS

[Abe86] Francine Abeles. Determinants and linear systems: Charles L. Dodgson’s view. British Journal for the History of Science, 19(3):331–335, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026509.

Anderson:1993:MIC

[ABR93]  R.  G.  W.

(Robert Geoffrey William) Anderson, J. A. (James A.) Bennett, and W. F. (William Francis) Ryan, editors. Making instruments count: essays on historical scientific instruments presented to Gerard L’Estrange Turner. Variorum, Aldershot, Hampshire, UK, 1993. ISBN 0-86078-394-4. xix + 492 pp. LCCN Q185 .M255 1993.

Ackerman:1987:JGF

[Ack87]  Robert Ackerman.  J. G.

Frazer: his life and work. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34093-4. x + 348 + 10 pp. LCCN GN21.F65 A27

1987. URL http://www.loc.
gov/catdir/description/
cam023/87009343.html;
http://www.loc.gov/
catdir/toc/cam023/
87009343.html.

Adams:1989:STP

[Ada89] E. B. (Edward Barry) Adams. In search of truth: a portrait of Don Craib. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1989. ISBN 1-85315-119-X, (paperback), 1-85315-118-1 (paperback). xi + 123 pp. LCCN R654.C73 A35

1989. URL http://catalog.
hathitrust.org/api/
volumes/oclc/20798932.
html.

Adler:1987:FMA

[Adl87] Jeremy Adler. “Eine fast magische Anziehungskraft”: Goethes “Wahlverwandtschaften” und die Chemie seiner Zeit. (German) [A nearly magic attractive force: Goethe’s “Wahlverwandtschaften” [feeling of affinity] and the chemistry of its time]. Beck, München, Germany, 1987. ISBN 3-406-31559-3. 260 pp. LCCN ???? URL

http://www.gbv.de/dms/bs/
toc/024186546.pdf.

Aftalion:1991:HIC

[Aft91] Fred Aftalion. A history of the international chemical industry. The Chemical sciences in society series. University of Pennsylvania Press, Philadelphia, PA, USA, 1991. ISBN 0-8122-8207-8 (hardcover), 0-8122-1297-5 (paperback). xix + 411 pp. LCCN HD9650.5 .A3813 1991.

Agren:1983:BRL

[Agr83a] Hans Agren. Book review: Life Sciences: Li Hui-Lin, Nan-fang ts’ao-mu chuang: a fourth century flora of Southeast Asia. Hong Kong: The Chinese University Press, 1979. Pp. x + 168. £9.00. British Journal for the History of Science, 16 (2):212–213, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026253.

Agren:1983:BRM

[Agr83b] Hans Agren. Book review:

Medical Sciences: Joseph Needham, Ho Ping-Yü, Lu Gwei-Djen & Nathan Sivin, Science and civilisation in China. Vol. V, pt. 4. Spagyrical discovery and invention: apparatus, theories and gifts. Cambridge: Cambridge University Press, 1980. Pp. xlviii + 772. £48.00. Lu Gwei-Djen & Joseph Needham, Celestial lancets: a history and rationale of acupuncture and moxa. Cambridge: Cambridge University Press, 1980. Pp. xxii + 427. £45.00. British Journal for the History of Science, 16 (1):81–84, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026096.

Achinstein:1985:OEH

[AH85] Peter Achinstein and Owen Hannaway, editors. Observation, experiment, and hypothesis in modern physical science. Studies from the Johns Hopkins Center for the History and Philosophy of Science. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-01083-6. x + 379 pp. LCCN QC6 .O26 1985.

Albrecht:1993:NTG

[AH93] Helmuth Albrecht and Armin Hermann, editors. Naturwissenschaft und Technik in der Geschichte: 25 Jahre Lehrstuhl für Geschichte der Naturwissenschaft und Technik am Historischen Institut der Universität Stuttgart. (German) [Natural science and technology in history: 25 Years of the Chair of History in Natural Science and Technology at the Historical Institut of the University of Stuttgart ]. Verlag für Geschichte der Naturwissenschaften und der Technik, Stuttgart, Germany, 1993. ISBN 3-928186-15-9. 401 pp. LCCN Q183.4.G293 U556 1993.

al-Hassan:1986:ITI

[aHH86a] Ahmad Y. al Hassan and Donald R. Hill. Islamic technology: an illustrated history. Cambridge University Press, Cambridge, UK, 1986. ISBN 92-3-102294-6 (UNESCO), 0-521-26333-6 (Cambridge). xiv + 304 pp. LCCN ????

Hassan:1986:ITI

[aHH86b] Ahmad Y. al Hassan and Donald R. Hill. Islamic technology an illustrated history. Cambridge University Press, Cambridge, UK, 1986. ISBN 92-3-102294-6, 0-521-26333-6. xiv + 304 pp. LCCN ????

al-Hassan:1988:ITI

[aHH88] Ahmad Y. al Hassan and Donald R. (Donald Routledge) Hill, editors. Islamic technology: an illustrated history. Cambridge University Press, Cambridge, UK, 1988. ISBN 92-3-102294-6 (UNESCO), 0-521-26333-6 (Cambridge). xv + 304 pp. LCCN ????

al-Hassan:1992:ITI

[aHH92] Ahmad Yusuf al Hassan and Donald Routledge Hill. Islamic technology: an illustrated history. Cambridge University Press, Cambridge, UK, 1992. ISBN 92-3-102294-6 (UNESCO, hardback), 92-3-102733-6 (UNESCO, paperback), 0-521-26333-6 (Cambridge, hardback), 0-521-42239-6 (Cambridge, paperback). xv + 304 pp. LCCN ????

Agricola:1990:BMD

[AHY90] Georg Agricola, Robert Halleux, and Albert Yans. Bermannus = Le mineur: un dialogue sur les mines. (French) [Bermannus: the miner: a dialogue on mines]. Science et humanisme, 0761-2885. Belles Lettres, Paris, France, 1990. ISBN 2-251-34504-3. xxx + 185 pp. LCCN TN144 .A3414 1990. 250F.

Aiton:1967:BRB

[Ait67] E. J. Aiton. Book review: The Background to Newton’s Principia. By John Herivel. London: Clarendon Press: Oxford University Press. Pp. xvi + 337. 1965. 70s. British Journal for the History of Science, 3(3):298–299, June 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025059.

Aiton:1968:BRS

[Ait68] E. J. Aiton. Book review: Sixteenth and Seventeenth Centuries: L’Expérience du Mouvement: Jean-Baptiste Baliani, disciple et critique de Galilée.

By Serge Moscovici. Paris: Hermann. 1967. Pp. 263. 9 pl., 26 figs. 15 F. British Journal for the History of Science, 4(2): 191–192, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025164.

Aiton:1971:BRS

[Ait71] E. J. Aiton. Book review:

Seventeenth Century: The Mathematical and Philosophical Works of the Right Rev. John Wilkins. London: F. Cass. 1970. Two volumes bound as one: Vol. I, Pp. xii + 261; Vol. II, Pp. xvi + 265, figs., index. £7.35. British Journal for the History of Science, 5 (3):308–309, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025338.

Aiton:1972:BRS

[Ait72] E. J. Aiton. Book review: Sixteenth and Seventeenth Centuries: Introduction to Newton’s ‘Principia’. By I. Bernard Cohen. London: Cambridge University Press, and Cambridge, Mass.: Harvard University Press, 1971. Pp. xxi + 380. £13. British Journal for the History of Science, 6 (1):94–95, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025276.

Aiton:1973:BRS

[Ait73] E. J. Aiton. Book review: Sixteenth and Seventeenth Centuries: The Annus Mirabilis of Sir Isaac Newton, 1666– 1966. Ed. by Robert Palter. Cambridge, Mass, and London: M.I.T. Press, 1971. Pp. viii + 351. £7. British Journal for the History of Science, 6(3):322–323, June 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025457.

Aiton:1985:BRJ

[Ait85a] E. J. Aiton. Book review: N. Jardine, The birth of history and philosophy of science. Kepler’s A defence of Tycho against Ursus with essays on its provenance and significance. Cambridge: Cambridge University Press, 1984. Pp. ix + 301. ISBN 0-521-25226-1. £32.50. British Journal for the History of Science, 18(2):252, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026354.

Aiton:1985:LB

[Ait85b] E. J. Aiton. Leibniz: a biography. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-470-6. xiv + 370 pp. LCCN B2597 .A67

                                        1985. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0745/85238184-d.html.

Aiton:1986:BRE

[Ait86] E. J. Aiton. Book reviews:

Edward Grant. In Defense of the Earth’s Centrality and Immobility: Scholastic Reaction to Copernicanism in the Seventeenth Century. Philadelphia: American Philosophical Society 1984. Pp. 69. ISBN 0-87169-744-0. Price $10.00 (paperback). Edward Rosen. Copernicus and the Scientific Revolution. Malabar, Florida: Robert E. Krieger Publishing Company, 1984. Pp. 220. ISBN 0-89874-573-X. Price $6.50 (paperback). British Journal for the History of Science, 19(1): 126–127, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026498.

Aiton:1989:BRG

[Ait89] E. J. Aiton. Book review: G.

W. Leibniz, Discourse on Metaphysics and Related Writings, (ed. and tr. by R. N. D. Martin & Stuart Brown). Manchester and New York: Manchester University Press, 1988. Pp. xi + 170. ISBN 0-7190-2338-6. £25.00. British Journal for the History of Science, 22(1):100, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026700.

Albertini-:1989:CCL

[AJ^{+}89] Ippolito Francesco Albertini, Saul Jarcho, et al., editors. Clinical consultations and letters. Francis A. Countway Library of Medicine, Boston, MA, USA, 1989. ISBN 0-88135-089-3. lxix + 356 pp. LCCN RC66 .C54 1989.

Anderson:1987:SMD

[AL87] R. G. W. (Robert Geoffrey William) Anderson and Christopher Lawrence, editors. Science, medicine, and dissent: Joseph Priestley (1733–1804): papers celebrating the 250th anniversary of the birth of Joseph Priestley together with a catalogue of an exhibition held at the Royal Society and the Wellcome Institute for the History of Medicine. Wellcome Trust/Science Museum, London, UK, 1987. ISBN 0-901805-28-9. xi + 105 pp. LCCN QD22.P8 S35x 1987.

Alder:1985:BBL

[Ald85] Garry Alder. Beyond Bokhara: the life of William Moorcroft, Asian explorer and pioneer veterinary surgeon, 1767– 1825. Century Pub., London, UK, 1985. ISBN 0-7126-0722-6. xiii + 417 + 8 pp. LCCN DS327.7.M67 A43 1985. US$16.95.

Alexander:1985:IQC

[Ale85] Peter Alexander. Ideas, qualities, and corpuscles: Locke and Boyle on the external world. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-26707-2.

336 pp. LCCN B1297 .A28

1985. URL http://www.loc.
gov/catdir/description/
cam031/84017457.html;
http://www.loc.gov/
catdir/enhancements/
fy0903/84017457-t.html.

Alic:1986:HHH

[Ali86] Margaret Alic. Hypatia’s heritage: a history of women in science from antiquity to the late Nineteenth Century. Women’s Press, London, UK, 1986. ISBN 0-7043-3954-4 (paperback). ix + 230 + 8 pp. LCCN Q130 .A48 1986. US$4.95.

Allen:1977:BRN

[All77] D. E. Allen. Book review: Natural History Natural History Auctions 1700–1972. A Register of Sales in the British Isles. Compiled by J. M. Chalmers-Hunt. London: Sotherby Parke Bernet, 1976. Pp. xii + 189. No price stated. British Journal for the History of Science, 10(3):257–258, November 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025848.

Allen:1979:BRB

[All79]  D. E. Allen.  Book re-

view: Botany Dictionary of British and Irish Botanists and Horticulturists. By Ray Desmond. London: Taylor & Francis, 1977. Pp. xxvi + 747. £40.00. British Journal for the History of Science, 12(1): 92–94, March 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025669.

Allen:1980:WMB

[All80] D. E. Allen. The women members of the Botanical Society of London, 1836–1856. British Journal for the History of Science, 13(3):240– 254, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026198.

Allen:1981:BRN

[All81] D. E. Allen. Book review: Natural, History British Natural History Books, 1495–1900: a Handlist. By R. B. Freeman. London: Dawson, 1980. Pp.

437. £20.00. British Journal for the History of Science, 14 (1):86–87, March 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026076.

Allen:1982:BRT

[All82]  D. E. Allen.  Book re-

view: Technology and Development Lucile H. Brockway, Science and colonial expansion: the role of the British Royal Botanic Gardens. New York, London, Toronto, Sydney & San Francisco: Academic Press, 1979. Pp. xvi + 215.

£12.00/$21.00. British Journal for the History of Science, 15 (1):91–93, March 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025938.

Allen:1984:BRHa

[All84a] D. E. Allen. Book review: History of Natural History: Edmund Berkeley and Dorothy Smith Berkeley, The life and travels of John Bartram. From Lake Ontario to the River St. John, Tallahassee: University Presses of Florida, 1982. Pp. xvi + 376. ISBN 0-8130-0700-3. (No price stated.). British Journal for the History of Science, 17(3):311–312, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Allen:1984:BRHb

[All84b] D. E. Allen. Book review:

History of Natural History: Ray Desmond, The India Museum 1801–1879. London: Her Majesty’s Stationery Office, 1982. Pp. xi+ 215. ISBN 0-11-580088-3. £25.00. British Journal for the History of Science, 17(3):323–324, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026638.

Allen:1986:BHB

[All86] David Elliston Allen. The botanists: a history of the Botanical Society of the British Isles through a hundred and fifty years. St Paul’s Bibliographies, Winchester, UK, 1986. ISBN 0-906795-36-2. xv + 232 + 8 pp. LCCN QK1 .A3647 1986. US$15.00.

Alter:1987:RPS

[Alt87] Peter Alter. The reluctant patron: science and the state in Britain, 1850–1920. Berg, Oxford, UK, 1987. ISBN 0-907582-67-2. 292 pp. LCCN Q127.G4 A6413 1987. US$45.00.

Anderson:1982:CBO

[And82] Lorin Anderson. Charles Bonnet and the order of the known, volume 11 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1982. ISBN 90-277-1389-8. xv + 159 pp. LCCN QH31.B72 A8 1982.

Anderson:1984:BLL

[And84] Wilda C. Anderson. Between the library and the laboratory: the language of chemistry in Eighteenth-Century France. The Johns Hopkins University Press, Baltimore, MD, USA, 1984. ISBN 0-8018-3229-2. 190 + 1 pp. LCCN PQ618 .A5 1984. US$22.50.

Anonymous:1962:BVIa

[Ano62a] Anonymous. BJH volume 1 issue 1: Cover and front matter. British Journal for the History of Science, 1(1):f1–f2, June 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1962:BVIb

[Ano62b] Anonymous. BJH volume 1 issue 2 cover and back matter. British Journal for the History of Science, 1(2):b1, December 1962. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1962:BVIc

[Ano62c] Anonymous. BJH volume 1 issue 2 cover and front matter. British Journal for the History of Science, 1(2):f1–f3, December 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1962:BR

[Ano62d] Anonymous. Books received.

British Journal for the History of Science, 1(2):192, December 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1962:IIO

[Ano62e] Anonymous. Items of interest from other publications. British Journal for the History of Science, 1(2):195–196, December 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1962:NCa

[Ano62f] Anonymous. Notes on contributors. British Journal for the History of Science, 1(1):88, June 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1962:NCb

[Ano62g] Anonymous. Notes on contributors. British Journal for the History of Science, 1(2):183, December 1962. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1962:OCB

[Ano62h]  Anonymous.  Officers and

Council of the British Society for the History of Science 1961– 1962. British Journal for the History of Science, 1(1):2, June 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1962:PCPa

[Ano62i] Anonymous. Papers in current periodicals. British Journal for the History of Science, 1(1):91–95, June 1962. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025079.

Anonymous:1962:PCPb

[Ano62j] Anonymous. Papers in current periodicals. British Journal for the History of Science, 1(2):193–195, December 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025139.

Anonymous:1962:RPB

[Ano62k] Anonymous. Recent proceedings of the British Society for the History of Science. British Journal for the History of Science, 1(1):79–87, June 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:BVIa

[Ano63a] Anonymous. BJH volume 1 issue 3 cover and back matter. British Journal for the History of Science, 1(3):b1, June 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:BVIb

[Ano63b] Anonymous. BJH volume 1 issue 3 cover and front matter. British Journal for the History of Science, 1(3):f1–f4, June 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:BVIc

[Ano63c] Anonymous. BJH volume 1 issue 4 cover and back matter. British Journal for the History of Science, 1(4):b1–b2, December 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:BVId

[Ano63d] Anonymous. BJH volume 1 issue 4 cover and front matter. British Journal for the History of Science, 1(4):f1–f3, December 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:BRa

[Ano63e] Anonymous. Books received. British Journal for the History of Science, 1(3):289, June 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:BRb

[Ano63f] Anonymous. Books received. British Journal for the History of Science, 1(4):387, December 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:FS

[Ano63g] Anonymous. Financial statement. British Journal for the History of Science, 1(4):394– 395, December 1963. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:IIOa

[Ano63h] Anonymous. Items of interest from other publications. British Journal for the History of Science, 1(3):290–292, June 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024932.

Anonymous:1963:IIOb

[Ano63i] Anonymous. Items of interest from other publications. British Journal for the History of Science, 1(4):390, December 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:NCa

[Ano63j] Anonymous. Notes on contributors. British Journal for the History of Science, 1(3): 284, June 1963. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:NCb

[Ano63k] Anonymous. Notes on contributors. British Journal for the History of Science, 1(4):381, December 1963. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:TPCb

[Ano63l] Anonymous. Title of papers in current periodicals. British Journal for the History of Science, 1(4):388–390, December 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1963:TPCa

[Ano63m] Anonymous. Title of papers in current periodicals on the history of science. British Journal for the History of Science, 1(3): 290, June 1963. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:BVIa

[Ano64a] Anonymous. BJH volume 2 issue 1 cover and back matter. British Journal for the History of Science, 2(1):b1–b3, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:BVIb

[Ano64b] Anonymous. BJH volume 2 issue 1 cover and front matter. British Journal for the History of Science, 2(1):f1–f3, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:BVIc

[Ano64c] Anonymous. BJH volume 2 issue 2 cover and back matter. British Journal for the History of Science, 2(2):b1–b4, December 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:BVId

[Ano64d] Anonymous. BJH volume 2 issue 2 cover and front matter. British Journal for the History of Science, 2(2):f1–f3, December 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:BRa

[Ano64e] Anonymous. Books received. British Journal for the History of Science, 2(1):82, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:BRb

[Ano64f] Anonymous. Books received.

British Journal for the History of Science, 2(2):172–173, December 1964. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:CCG

[Ano64g] Anonymous. Correspondence concerning Galileo. British Journal for the History of Science, 2(2):173, December 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:FE

[Ano64h]  Anonymous.  Forthcoming

events. British Journal for the History of Science, 2(1):95–96, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:IIO

[Ano64i] Anonymous. Items of interest from other publications. British Journal for the History of Science, 2(1):83, June 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:NCa

[Ano64j] Anonymous. Notes on contributors. British Journal for the History of Science, 2 (1):94, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025097.

Anonymous:1964:NCb

[Ano64k] Anonymous. Notes on contributors. British Journal for the History of Science, 2(2):176, December 1964. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:OPA

[Ano64l] Anonymous. Obituary: Professor Alexandre Koyré (28 April 1964), Mrs. Dorothea Waley Singer (24 June 1964), Rev. Canon Charles Earle Raven (8 July 1964). British Journal for the History of Science, 2 (2):173, December 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:PCP

[Ano64m] Anonymous. Papers in current periodicals. British Journal for the History of Science, 2(2):174–175, December 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:RCF

[Ano64n] Anonymous. Report of Council and financial statement. British Journal for the History of Science, 2(1):90–93, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1964:TPC

[Ano64o] Anonymous. Titles of papers in current periodicals. British Journal for the History of Science, 2(1):83, June 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1965:BVIa

[Ano65a] Anonymous. BJH volume 2 issue 3 cover and back matter. British Journal for the History of Science, 2(3):b1–b4, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1965:BVIb

[Ano65b] Anonymous. BJH volume 2 issue 3 cover and front matter. British Journal for the History of Science, 2(3):f1–f3, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1965:BVIc

[Ano65c] Anonymous. BJH volume 2 issue 4 cover and back matter. British Journal for the History of Science, 2(4):b1–b4, December 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1965:BVId

[Ano65d] Anonymous. BJH volume 2 issue 4 cover and front matter. British Journal for the History of Science, 2(4):f1–f3, December 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1965:BRa

[Ano65e] Anonymous. Books received. British Journal for the History of Science, 2(3):272, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1965:BRb

[Ano65f] Anonymous. Books received. British Journal for the History of Science, 2(4):365, December 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1965:IIO

[Ano65g] Anonymous. Items of interest from other publications. British Journal for the History of Science, 2(3):274, June 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1965:NCa

[Ano65h] Anonymous. Notes on contributors. British Journal for the History of Science, 2(3): 275, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024956.

Anonymous:1965:NCb

[Ano65i] Anonymous. Notes on contributors. British Journal for the History of Science, 2(4):366, December 1965. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1965:OPJ

[Ano65j] Anonymous. Obituary: Professor E. J. Dijksterhuis (18 May 1965) and Professor J. R. Partington (9 October 1965). British Journal for the History of Science, 2(4):364, December 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024900.

Anonymous:1965:PCP

[Ano65k] Anonymous. Papers in current periodicals. British Journal for the History of Science, 2(3): 273, June 1965. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1966:BVIa

[Ano66a] Anonymous. BJH volume 3 issue 1 cover and back matter. British Journal for the History of Science, 3(1):b1–b7, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1966:BVIb

[Ano66b] Anonymous. BJH volume 3 issue 1 cover and front matter. British Journal for the History of Science, 3(1):f1–f3, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1966:BVIc

[Ano66c] Anonymous. BJH volume 3 issue 2 cover and back matter.

British Journal for the History of Science, 3(2):b1–b7, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1966:BVId

[Ano66d] Anonymous. BJH volume 3 issue 2 cover and front matter. British Journal for the History of Science, 3(2):f1–f3, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1966:BRa

[Ano66e] Anonymous. Books received. British Journal for the History of Science, 3(1):95–96, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025125.

Anonymous:1966:BRb

[Ano66f] Anonymous. Books received. British Journal for the History of Science, 3(2):200–201, December 1966. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1966:CC

[Ano66g] Anonymous. Centenary celebrations. British Journal for the History of Science, 3(1):72, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1966:ER

[Ano66h] Anonymous. Essay reviews.

British Journal for the History of Science, 3(1):73–78, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025105;   http://
www.jstor.org/stable/
4025106;   http://www.
jstor.org/stable/4025107.

Anonymous:1966:NCa

[Ano66i] Anonymous. Notes on contributors. British Journal for the History of Science, 3(1):96, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1966:NCb

[Ano66j] Anonymous. Notes on contributors. British Journal for the History of Science, 3(2):202, December 1966. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1967:BVIa

[Ano67a] Anonymous. BJH volume 3 issue 3 cover and back matter. British Journal for the History of Science, 3(3):b1–b10, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1967:BVIb

[Ano67b] Anonymous. BJH volume 3 issue 3 cover and front matter. British Journal for the History of Science, 3(3):f1–f3, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1967:BVIc

[Ano67c] Anonymous. BJH volume 3 issue 4 cover and back matter. British Journal for the History of Science, 3(4):b1–b8, December 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1967:BVId

[Ano67d] Anonymous. BJH volume 3 issue 4 cover and front matter. British Journal for the History of Science, 3(4):f1–f3, December 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1967:BRa

[Ano67e] Anonymous. Books received. British Journal for the History of Science, 3(3):305–306, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1967:BRb

[Ano67f] Anonymous. Books received. British Journal for the History of Science, 3(4):409–410, December 1967. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1967:NCa

[Ano67g] Anonymous. Notes on contributors. British Journal for the History of Science, 3(3):

307, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025068.

Anonymous:1967:NCb

[Ano67h] Anonymous. Notes on contributors. British Journal for the History of Science, 3(4):410, December 1967. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1967:OPD

[Ano67i] Anonymous. Obituary: Professor Douglas McKie (28 August 1967). British Journal for the History of Science, 3 (4):408, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1967:OPE

[Ano67j] Anonymous. Obituary: Professor Eva G. R. Taylor (5 July 1966). British Journal for the History of Science, 3(3):306, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1968:BVIa

[Ano68a] Anonymous. BJH volume 4 issue 1 cover and back matter. British Journal for the History of Science, 4(1):b1–b8, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1968:BVIb

[Ano68b] Anonymous. BJH volume 4 issue 1 cover and front matter. British Journal for the History of Science, 4(1):f1–f4, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1968:BVIc

[Ano68c] Anonymous. BJH volume 4 issue 2 cover and back matter. British Journal for the History of Science, 4(2):b1–b6, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1968:BVId

[Ano68d] Anonymous. BJH volume 4 issue 2 cover and front matter. British Journal for the History of Science, 4(2):f1–f4, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1968:BT

[Ano68e] Anonymous. Boerhaave Tercentenary. British Journal for the History of Science, 4(2): 200, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1968:BRa

[Ano68f] Anonymous. Books received. British Journal for the History of Science, 4(1):91–93, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1968:BRb

[Ano68g] Anonymous. Books received.

British Journal for the History of Science, 4(2):197–198, December 1968. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1968:NCa

[Ano68h] Anonymous. Notes on contributors. British Journal for the History of Science, 4 (1):94, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025008.

Anonymous:1968:NCb

[Ano68i] Anonymous. Notes on contributors. British Journal for the History of Science, 4(2): 206, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025172.

Anonymous:1968:XIC

[Ano68j] Anonymous. XIIth International Congress of the History of Science. British Journal for the History of Science, 4 (2):200, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1969:BVIa

[Ano69a] Anonymous. BJH volume 4 issue 3 cover and back matter. British Journal for the History of Science, 4(3):b1–b4, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1969:BVIb

[Ano69b] Anonymous. BJH volume 4 issue 3 cover and front matter. British Journal for the History of Science, 4(3):f1–f4, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1969:BVIc

[Ano69c] Anonymous. BJH volume 4 issue 4 cover and back matter. British Journal for the History of Science, 4(4):b1–b4, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1969:BVId

[Ano69d] Anonymous. BJH volume 4 issue 4 cover and front matter. British Journal for the History of Science, 4(4):f1–f4, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1969:BRa

[Ano69e] Anonymous. Books received. British Journal for the History of Science, 4(3):304–305, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1969:BRb

[Ano69f] Anonymous. Books received. British Journal for the History of Science, 4(4):424–425, December 1969. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1969:FS

[Ano69g] Anonymous. Financial statement. British Journal for the History of Science, 4(4):426– 427, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1969:HCS

[Ano69h] Anonymous. Handbook to the collections of scientific instruments in Great Britain. British Journal for the History of Science, 4(3):306–307, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1969:MC

[Ano69i] Anonymous. Mendel colloquium. British Journal for the History of Science, 4(4):423, December 1969. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1969:NCa

[Ano69j] Anonymous. Notes on contributors. British Journal for the History of Science, 4(3): 308, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025499.

Anonymous:1969:NCb

[Ano69k] Anonymous. Notes on contributors. British Journal for the History of Science, 4(4):423, December 1969. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1970:BVIa

[Ano70a] Anonymous. BJH volume 5 issue 1 cover and back matter. British Journal for the History of Science, 5(1):b1–b7, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1970:BVIb

[Ano70b] Anonymous. BJH volume 5 issue 1 cover and front matter. British Journal for the History of Science, 5(1):f1–f4, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1970:BVIc

[Ano70c] Anonymous. BJH volume 5 issue 2 cover and back matter. British Journal for the History of Science, 5(2):b1–b10, December 1970. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1970:BVId

[Ano70d] Anonymous. BJH volume 5 issue 2 cover and front matter. British Journal for the History of Science, 5(2):f1–f4, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1970:BRa

[Ano70e] Anonymous. Books received.

British Journal for the History of Science, 5(1):107–109, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1970:BRb

[Ano70f] Anonymous. Books received.

British Journal for the History of Science, 5(2):209–210, December 1970. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1970:FS

[Ano70g] Anonymous. Financial statement. British Journal for the History of Science, 5(2):214– 215, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1970:NCa

[Ano70h] Anonymous. Notes on contributors. British Journal for the History of Science, 5(1): 110, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025376.

Anonymous:1970:NCb

[Ano70i] Anonymous. Notes on contributors. British Journal for the History of Science, 5(2): 216, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025441.

Anonymous:1971:BVIa

[Ano71a] Anonymous. BJH volume 5 issue 3 cover and back matter. British Journal for the History of Science, 5(3):b1–b7, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1971:BVIb

[Ano71b] Anonymous. BJH volume 5 issue 3 cover and front matter. British Journal for the History of Science, 5(3):f1–f4, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1971:BVIc

[Ano71c] Anonymous. BJH volume 5 issue 4 cover and back matter. British Journal for the History of Science, 5(4):b1–b5, December 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1971:BVId

[Ano71d] Anonymous. BJH volume 5 issue 4 cover and front matter. British Journal for the History of Science, 5(4):f1–f4, December 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1971:BRa

[Ano71e] Anonymous. Books received. British Journal for the History of Science, 5(3):316–318, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1971:BRb

[Ano71f] Anonymous. Books received.

British Journal for the History of Science, 5(4):423–425, December 1971. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1971:FS

[Ano71g] Anonymous. Financial statement. British Journal for the History of Science, 5(4):430– 431, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1971:FTY

[Ano71h] Anonymous. The first ten years of the Journal. British Journal for the History of Science, 5(4):425, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1971:NCa

[Ano71i] Anonymous. Notes on contributors. British Journal for the History of Science, 5(3): 321, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025348.

Anonymous:1971:NCb

[Ano71j] Anonymous. Notes on contributors. British Journal for the History of Science, 5(4):426, December 1971. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1972:BVIa

[Ano72a] Anonymous. BJH volume 6 issue 1 cover and back matter. British Journal for the History of Science, 6(1):b1–b5, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1972:BVIb

[Ano72b] Anonymous. BJH volume 6 issue 1 cover and front matter. British Journal for the History of Science, 6(1):f1–f4, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1972:BVIc

[Ano72c] Anonymous. BJH volume 6 issue 2 cover and back matter. British Journal for the History of Science, 6(2):b1–b9, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1972:BVId

[Ano72d] Anonymous. BJH volume 6 issue 2 cover and front matter. British Journal for the History of Science, 6(2):f1–f4, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1972:BRa

[Ano72e] Anonymous. Books received.

British Journal for the History of Science, 6(1):101–103, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1972:BRb

[Ano72f] Anonymous. Books received.

British Journal for the History of Science, 6(2):224–226, December 1972. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1972:NCa

[Ano72g] Anonymous. Notes on contributors. British Journal for the History of Science, 6(1): 107, June 1972. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1972:NCb

[Ano72h] Anonymous. Notes on contributors. British Journal for the History of Science, 6(2):229, December 1972. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1973:BVIa

[Ano73a] Anonymous. BJH volume 6 issue 3 cover and back matter. British Journal for the History of Science, 6(3):b1– b8, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025226.

Anonymous:1973:BVIb

[Ano73b] Anonymous. BJH volume 6 issue 3 cover and front matter. British Journal for the History of Science, 6(3):f1–f4, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1973:BVIc

[Ano73c] Anonymous. BJH volume 6 issue 4 cover and back matter. British Journal for the History of Science, 6(4):b1–b4, December 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1973:BVId

[Ano73d] Anonymous. BJH volume 6 issue 4 cover and front matter. British Journal for the History of Science, 6(4):f1–f4, December 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1973:BRa

[Ano73e] Anonymous. Books received. British Journal for the History of Science, 6(3):333–335, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1973:BRb

[Ano73f] Anonymous. Books received. British Journal for the History of Science, 6(4):456–458, December 1973. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1973:NCa

[Ano73g] Anonymous. Notes on contributors. British Journal for the History of Science, 6(3): 341, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025469.

Anonymous:1973:NCb

[Ano73h] Anonymous. Notes on contributors. British Journal for the History of Science, 6(4):463, December 1973. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1974:BVIa

[Ano74a] Anonymous. BJH volume 7 issue 1 cover and back matter. British Journal for the History of Science, 7(1):b1–b4, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1974:BVIc

[Ano74b] Anonymous. BJH volume 7 issue 1 cover and front matter. British Journal for the History of Science, 7(1):f1–f4, March 1974. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1974:BVIe

[Ano74c] Anonymous. BJH volume 7 issue 2 cover and back matter. British Journal for the History of Science, 7(2):b1–b4, July 1974. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1974:BVIf

[Ano74d] Anonymous. BJH volume 7 issue 2 cover and front matter. British Journal for the History of Science, 7(2):f1–f4, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1974:BVIb

[Ano74e] Anonymous. BJH volume 7 issue 3 cover and back matter. British Journal for the History of Science, 7(3):b1–b2, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1974:BVId

[Ano74f] Anonymous. BJH volume 7 issue 3 cover and front matter. British Journal for the History of Science, 7(3):f1–f4, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1974:BRa

[Ano74g] Anonymous. Books received. British Journal for the History of Science, 7(1):98–99, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1974:BRc

[Ano74h] Anonymous. Books received. British Journal for the History of Science, 7(3):303, November 1974. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1974:BRb

[Ano74i] Anonymous. Books recieved.

British Journal for the History of Science, 7(2):199–200, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1974:NCa

[Ano74j] Anonymous. Notes on contributors. British Journal for the History of Science, 7(1): 104, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025197.

Anonymous:1974:NCb

[Ano74k] Anonymous. Notes on contributors. British Journal for the History of Science, 7(2): 203, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025226.

Anonymous:1974:NCc

[Ano74l] Anonymous. Notes on contributors. British Journal for the History of Science, 7(3):305, November 1974. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:BVIa

[Ano75a] Anonymous. BJH volume 8 issue 1 cover and back matter. British Journal for the History of Science, 8(1):b1–b4, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:BVIb

[Ano75b] Anonymous. BJH volume 8 issue 1 cover and front matter. British Journal for the History of Science, 8(1):f1–f4, March 1975. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:BVIc

[Ano75c] Anonymous. BJH volume 8 issue 2 cover and back matter. British Journal for the History of Science, 8(2):b1–b4, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:BVId

[Ano75d] Anonymous. BJH volume 8 issue 2 cover and front matter. British Journal for the History of Science, 8(2):f1–f4, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:BVIe

[Ano75e] Anonymous. BJH volume 8 issue 3 cover and back matter. British Journal for the History of Science, 8(3):b1–b5, November 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:BVIf

[Ano75f] Anonymous. BJH volume 8 issue 3 cover and front matter. British Journal for the History of Science, 8(3):f1–f4, November 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:BRa

[Ano75g] Anonymous. Books received.

British Journal for the History of Science, 8(1):92–93, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:BRb

[Ano75h] Anonymous. Books received.

British Journal for the History of Science, 8(2):190–191, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:BRc

[Ano75i] Anonymous. Books received.

British Journal for the History of Science, 8(3):265–266, November 1975. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:Ia

[Ano75j] Anonymous. Inquiry. British Journal for the History of Science, 8(2):192, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025659.

Anonymous:1975:Ib

[Ano75k] Anonymous. Inquiry. British Journal for the History of Science, 8(3):283, November 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025576.

Anonymous:1975:NCa

[Ano75l] Anonymous. Notes on contributors. British Journal for the History of Science, 8(1): 100, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025838.

Anonymous:1975:NCb

[Ano75m] Anonymous. Notes on contributors. British Journal for the History of Science, 8(2): 192, July 1975. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:NCc

[Ano75n] Anonymous. Notes on contributors. British Journal for the History of Science, 8(3):279, November 1975. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1975:RCY

[Ano75o] Anonymous. Report of Council for the year 1974-5. British Journal for the History of Science, 8(3):280–282, November 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1976:BVIa

[Ano76a] Anonymous. BJH volume 9 issue 1 cover and back matter. British Journal for the History of Science, 9(1):b1–b7, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1976:BVIb

[Ano76b] Anonymous. BJH volume 9 issue 1 cover and front matter. British Journal for the History of Science, 9(1):f1–f4, March 1976. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1976:BVIc

[Ano76c] Anonymous. BJH volume 9 issue 2 cover and back matter. British Journal for the History of Science, 9(2):b1–b8, July 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1976:BVId

[Ano76d] Anonymous. BJH volume 9 issue 2 cover and front matter. British Journal for the History of Science, 9(2):f1–f4, July 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1976:BVIe

[Ano76e] Anonymous. BJH volume 9 issue 3 cover and back matter. British Journal for the History of Science, 9(3):b1–b5, November 1976. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1976:BVIf

[Ano76f] Anonymous. BJH volume 9 issue 3 cover and front matter. British Journal for the History of Science, 9(3):f1–f4, November 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1976:BRa

[Ano76g] Anonymous. Books received. British Journal for the History of Science, 9(1):87–88, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1976:BRb

[Ano76h] Anonymous. Books received. British Journal for the History of Science, 9(3):338–341, November 1976. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1976:NCa

[Ano76i] Anonymous. Notes on contributors. British Journal for the History of Science, 9(1):89, March 1976. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1976:NCb

[Ano76j] Anonymous. Notes on contributors. British Journal for the History of Science, 9(2): 241–242, July 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025811.

Anonymous:1976:NCc

[Ano76k] Anonymous. Notes on contributors. British Journal for the History of Science, 9(3): 345, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025701.

Anonymous:1976:RCY

[Ano76l] Anonymous. Report of the Council for the year 1975–1976. British Journal for the History of Science, 9(3):342–344, November 1976. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:A

[Ano77a] Anonymous. Announcement.

British Journal for the History of Science, 10(2):185, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025888.

Anonymous:1977:BVIa

[Ano77b] Anonymous. BJH volume 10 issue 1 cover and back matter. British Journal for the History of Science, 10(1):b1– b2, March 1977. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:BVIb

[Ano77c] Anonymous. BJH volume 10 issue 1 cover and front matter. British Journal for the History of Science, 10(1):f1–f4, March 1977. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:BVIc

[Ano77d] Anonymous. BJH volume 10 issue 2 cover and back matter. British Journal for the History of Science, 10(2):b1– b4, July 1977. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:BVId

[Ano77e] Anonymous. BJH volume 10 issue 2 cover and front matter. British Journal for the History of Science, 10(2):f1–f4, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:BVIe

[Ano77f] Anonymous. BJH volume 10 issue 3 cover and back matter. British Journal for the History of Science, 10(3):b1–b6, November 1977. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:BVIf

[Ano77g] Anonymous. BJH volume 10 issue 3 cover and front matter. British Journal for the History of Science, 10(3):f1–f4, November 1977. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:BRa

[Ano77h] Anonymous. Books received.

British Journal for the History of Science, 10(1):87–88, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:BRb

[Ano77i] Anonymous. Books received.

British Journal for the History of Science, 10(2):183–184, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:BRc

[Ano77j] Anonymous. Books received.

British Journal for the History of Science, 10(3):274–276, November 1977. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:CLC

[Ano77k] Anonymous. The collected letters of Charles Darwin. British Journal for the History of Science, 10(1):93, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:I

[Ano77l] Anonymous. Inquiries. British Journal for the History of Science, 10(2):185, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:LAR

[Ano77m]  Anonymous.  Library and

archive resources in the history of science: Manchester. British Journal for the History of Science, 10(1):89–92, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:NCa

[Ano77n] Anonymous. Notes on contributors. British Journal for the History of Science, 10(2):186, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:NCb

[Ano77o] Anonymous. Notes on contributors. British Journal for the History of Science, 10(3):276, November 1977. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1977:RCY

[Ano77p] Anonymous. Report of Council for the year 1976–1977. British Journal for the History of Science, 10(3):277– 280, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025862.

Anonymous:1978:BVIa

[Ano78a] Anonymous. BJH volume 11 issue 1 cover and back matter. British Journal for the History of Science, 11(1):b1– b3, March 1978. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:BVIb

[Ano78b] Anonymous. BJH volume 11 issue 1 cover and front matter. British Journal for the History of Science, 11(1):f1–f4, March 1978. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:BVIc

[Ano78c] Anonymous. BJH volume 11 issue 2 cover and back matter. British Journal for the History of Science, 11(2):b1– b3, July 1978. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:BVId

[Ano78d] Anonymous. BJH volume 11 issue 2 cover and front matter. British Journal for the History of Science, 11(2):f1–f4, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:BVIe

[Ano78e] Anonymous. BJH volume 11 issue 3 cover and back matter. British Journal for the History of Science, 11(3):b1–b5, November 1978. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:BVIf

[Ano78f] Anonymous. BJH volume 11 issue 3 cover and front matter. British Journal for the History of Science, 11(3):f1–f4, November 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:BRa

[Ano78g] Anonymous. Books received. British Journal for the History of Science, 11(1):95–96, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:BRb

[Ano78h] Anonymous. Books received. British Journal for the History of Science, 11(2):188–190, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:BRc

[Ano78i] Anonymous. Books received. British Journal for the History of Science, 11(3):297–299, November 1978. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:NCa

[Ano78j] Anonymous. Notes on contributors. British Journal for the History of Science, 11(1): 100–101, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:NCb

[Ano78k] Anonymous. Notes on contributors. British Journal for the History of Science, 11(2):190, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:NCc

[Ano78l] Anonymous. Notes on contributors. British Journal for the History of Science, 11(3):299, November 1978. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1978:RCY

[Ano78m] Anonymous. Report of Council for the year 1977–1978. British Journal for the History of Science, 11(3):300–303, November 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1979:BVIa

[Ano79a] Anonymous. BJH volume 12 issue 1 cover and back matter. British Journal for the History of Science, 12(1):b1– b5, March 1979. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1979:BVIb

[Ano79b] Anonymous. BJH volume 12 issue 1 cover and front matter. British Journal for the History of Science, 12(1):f1–f4, March 1979. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1979:BVIc

[Ano79c] Anonymous. BJH volume 12 issue 2 cover and back matter. British Journal for the History of Science, 12(2):b1– b8, July 1979. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1979:BVId

[Ano79d] Anonymous. BJH volume 12 issue 2 cover and front matter. British Journal for the History of Science, 12(2):f1–f4, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1979:BVIe

[Ano79e] Anonymous. BJH volume 12 issue 3 cover and back matter. British Journal for the History of Science, 12(3):b1–b4, November 1979. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1979:BVIf

[Ano79f] Anonymous. BJH volume 12 issue 3 cover and front matter. British Journal for the History of Science, 12(3):f1–f4, November 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1979:BRa

[Ano79g] Anonymous. Books received. British Journal for the History of Science, 12(1):105–106, March 1979. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1979:BRb

[Ano79h] Anonymous. Books received.

British Journal for the History of Science, 12(2):239–241, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1979:BRc

[Ano79i] Anonymous. Books received.

British Journal for the History of Science, 12(3):342–345, November 1979. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1979:NCa

[Ano79j] Anonymous. Notes on contributors. British Journal for the History of Science, 12(1): 107, March 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025680.

Anonymous:1979:NCb

[Ano79k] Anonymous. Notes on contributors. British Journal for the History of Science, 12(2): 244, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025775.

Anonymous:1979:NCc

[Ano79l] Anonymous. Notes on contributors. British Journal for the History of Science, 12(3): 350–351, November 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026011.

Anonymous:1980:BVIa

[Ano80a] Anonymous. BJH volume 13 issue 1 cover and back matter. British Journal for the History of Science, 13(1):b1– b6, March 1980. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1980:BVIb

[Ano80b] Anonymous. BJH volume 13 issue 1 cover and front matter. British Journal for the History of Science, 13(1):f1–f4, March 1980. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1980:BVIc

[Ano80c] Anonymous. BJH volume 13 issue 2 cover and back matter. British Journal for the History of Science, 13(2):b1– b2, July 1980. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1980:BVId

[Ano80d] Anonymous. BJH volume 13 issue 2 cover and front matter. British Journal for the History of Science, 13(2):f1–f4, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1980:BVIe

[Ano80e] Anonymous. BJH volume 13 issue 3 cover and back matter. British Journal for the History of Science, 13(3):b1–b2, November 1980. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1980:BVIf

[Ano80f] Anonymous. BJH volume 13 issue 3 cover and front matter. British Journal for the History of Science, 13(3):f1–f4, November 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1980:BRa

[Ano80g] Anonymous. Books received.

British Journal for the History of Science, 13(1):86–87, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1980:BRb

[Ano80h] Anonymous. Books received.

British Journal for the History of Science, 13(2):184–187, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1980:BRc

[Ano80i] Anonymous. Books received.

British Journal for the History of Science, 13(3):279–281, November 1980. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1980:NCa

[Ano80j] Anonymous. Notes on contributors. British Journal for the History of Science, 13(1):

90, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026141.

Anonymous:1980:NCb

[Ano80k] Anonymous. Notes on contributors. British Journal for the History of Science, 13(2): 188, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025918.

Anonymous:1980:NCc

[Ano80l] Anonymous. Notes on contributors. British Journal for the History of Science, 13(3): 288, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026216.

Anonymous:1980:RCY

[Ano80m] Anonymous. Report of Council for the year 1979–1980. British Journal for the History of Science, 13(3):284–287, November 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:BVIa

[Ano81a] Anonymous. BJH volume 14 issue 1 cover and back matter. British Journal for the History of Science, 14(1):b1– b4, March 1981. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:BVIb

[Ano81b] Anonymous. BJH volume 14 issue 1 cover and front matter. British Journal for the History of Science, 14(1):f1–f4, March 1981. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:BVIc

[Ano81c] Anonymous. BJH volume 14 issue 2 cover and back matter. British Journal for the History of Science, 14(2):b1– b3, July 1981. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:BVId

[Ano81d] Anonymous. BJH volume 14 issue 2 cover and front matter. British Journal for the History of Science, 14(2):f1–f4, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:BVIe

[Ano81e] Anonymous. BJH volume 14 issue 3 cover and back matter. British Journal for the History of Science, 14(3):b1–b4, November 1981. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:BVIf

[Ano81f] Anonymous. BJH volume 14 issue 3 cover and front matter. British Journal for the History of Science, 14(3):f1–f4, November 1981. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:BRa

[Ano81g] Anonymous. Books received. British Journal for the History of Science, 14(1):105–107, March 1981. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:BRb

[Ano81h] Anonymous. Books received. British Journal for the History of Science, 14(2):222–224, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:BRc

[Ano81i] Anonymous. Books received. British Journal for the History of Science, 14(3):297–299, November 1981. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:NCc

[Ano81j] Anonymous. Notes on contributor. British Journal for the History of Science, 14(3):299, November 1981. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:NCd

[Ano81k] Anonymous. Notes on contributor. British Journal for the History of Science, 14(3):303, November 1981. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:NCa

[Ano81l] Anonymous. Notes on contributors. British Journal for the History of Science, 14(1): 108, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026090.

Anonymous:1981:NCb

[Ano81m] Anonymous. Notes on contributors. British Journal for the History of Science, 14(2):224– 225, July 1981. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1981:RCY

[Ano81n] Anonymous. Report of Council for the year 1980–1981. British Journal for the History of Science, 14(3):300–303, November 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1982:BVIa

[Ano82a] Anonymous. BJH volume 15 issue 1 cover and back matter. British Journal for the History of Science, 15(1):b1– b2, March 1982. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1982:BVIb

[Ano82b] Anonymous. BJH volume 15 issue 1 cover and front matter. British Journal for the History of Science, 15(1):f1–f4, March 1982. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1982:BVIc

[Ano82c] Anonymous. BJH volume 15 issue 2 cover and back matter. British Journal for the History of Science, 15(2):b1– b2, July 1982. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1982:BVId

[Ano82d] Anonymous. BJH volume 15 issue 2 cover and front matter. British Journal for the History of Science, 15(2):f1–f2, July 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1982:BVIe

[Ano82e] Anonymous. BJH volume 15 issue 3 cover and back matter. British Journal for the History of Science, 15(3):b1–b4, November 1982. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1982:BVIf

[Ano82f] Anonymous. BJH volume 15 issue 3 cover and front matter. British Journal for the History of Science, 15(3):f1–f2, November 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1982:BRa

[Ano82g] Anonymous. Books received. British Journal for the History of Science, 15(1):95–97, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025941.

Anonymous:1982:BRb

[Ano82h] Anonymous. Books received.

British Journal for the History of Science, 15(2):210, July 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025979.

Anonymous:1982:BRc

[Ano82i] Anonymous. Books received.

British Journal for the History of Science, 15(3):315–316, November 1982. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1982:CV

[Ano82j] Anonymous. Contents of volume 15. British Journal for the History of Science, 15(3):321, November 1982. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1982:E

[Ano82k] Anonymous. Editorial. British Journal for the History of Science, 15(1):1–2, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1982:HFT

[Ano82l] Anonymous. Huygens et la France: table ronde du Centre national de la recherche scientifique, Paris, 27–29 mars 1979. L’Histoire des sciences. Textes et études. J. Vrin, Paris, France, 1982. ISBN 2-7116-2018-2. ix + 268 pp. LCCN Q143.H96 H88 1982. 210F.

Anonymous:1982:ICBb

[Ano82m] Anonymous. ISIS Cumulative Bibliography: a bibliography of the history of science formed from ISIS Critical Bibliographies 1-90, 1913– 1965: vol. 5. Civilizations and periods 15th to 19th Centuries. Mansell Pub. with the History of Science Society, London, UK, 1982. ISBN 0-7201-1643-0. xi + 593 pp. LCCN Z7405.H2 I2 1982 vol. 5.

Anonymous:1982:ICBa

[Ano82n] Anonymous. ISIS Cumulative Bibliography: a bibliography of the history of science formed from ISIS Critical Bibliographies 1-90, 1913-65: vol. 4. Civilizations and periods: prehistory to Middle Ages. Mansell Pub. with the History of Science Society, London, UK, 1982. ISBN 0-7201-1642-2, 0-7201-0549-8 (set). xviii + 457 pp. LCCN Z7405.H6 I2 1982 vol. 4.

Anonymous:1982:RCY

[Ano82o] Anonymous. Report of Council for the year 1981-2. British Journal for the History of Science, 15(3):317–319, November 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1982:SSS

[Ano82p] Anonymous. Science and speculation: studies in Hellenistic theory and practice. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24689-X, 2-901725-47-3. 351 pp. LCCN ????

Anonymous:1983:BVIa

[Ano83a] Anonymous. BJH volume 16 issue 1 cover and back matter. British Journal for the History of Science, 16(1):b1– b2, March 1983. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1983:BVIb

[Ano83b] Anonymous. BJH volume 16 issue 1 cover and front matter. British Journal for the History of Science, 16(1):f1–f3, March 1983. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1983:BVIc

[Ano83c] Anonymous. BJH volume 16 issue 2 cover and back matter. British Journal for the History of Science, 16(2):b1– b2, July 1983. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1983:BVId

[Ano83d] Anonymous. BJH volume 16 issue 2 cover and front matter. British Journal for the History of Science, 16(2):f1–f2, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1983:BVIe

[Ano83e] Anonymous. BJH volume 16 issue 3 cover and back matter. British Journal for the History of Science, 16(3):b1–b2, November 1983. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1983:BVIf

[Ano83f] Anonymous. BJH volume 16 issue 3 cover and front matter. British Journal for the History of Science, 16(3):f1–f2, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1983:BRa

[Ano83g] Anonymous. Books received. British Journal for the History of Science, 16(1):80–81, March 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1983:BRb

[Ano83h] Anonymous. Books received. British Journal for the History of Science, 16(2):220– 222, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026261.

Anonymous:1983:BRc

[Ano83i] Anonymous. Books received.

British Journal for the History of Science, 16(3):323–325, November 1983. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1983:CV

[Ano83j] Anonymous. Contents of volume 16, 1983. British Journal for the History of Science, 16 (3):331, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1983:E

[Ano83k] Anonymous. Editorial. British Journal for the History of Science, 16(1):ii, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1983:IBR

[Ano83l] Anonymous. Index of books reviewed. British Journal for the History of Science, 16(3): 333–334, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1983:RCY

[Ano83m] Anonymous. Report of Council for the year 1982-3. British Journal for the History of Science, 16(3):327–330, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1984:BVIa

[Ano84a] Anonymous. BJH volume 17 issue 1 cover and back matter. British Journal for the History of Science, 17(1):b1– b3, March 1984. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1984:BVIb

[Ano84b] Anonymous. BJH volume 17 issue 1 cover and front matter. British Journal for the History of Science, 17(1):f1–f3, March 1984. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1984:BVIc

[Ano84c] Anonymous. BJH volume 17 issue 2 cover and back matter. British Journal for the History of Science, 17(2):b1– b2, July 1984. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1984:BVId

[Ano84d] Anonymous. BJH volume 17 issue 2 cover and front matter. British Journal for the History of Science, 17(2):f1–f2, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1984:BVIe

[Ano84e] Anonymous. BJH volume 17 issue 3 cover and back matter. British Journal for the History of Science, 17(3):b1–b2, November 1984. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1984:BVIf

[Ano84f] Anonymous. BJH volume 17 issue 3 cover and front matter. British Journal for the History of Science, 17(3):f1–f2, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1984:BRa

[Ano84g] Anonymous. Books received.

British Journal for the History of Science, 17(1):123–125, March 1984. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1984:BRb

[Ano84h] Anonymous. Books received.

British Journal for the History of Science, 17(2):253– 254, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026587.

Anonymous:1984:BRc

[Ano84i] Anonymous. Books received.

British Journal for the History of Science, 17(3):338–340, November 1984. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1984:BSH

[Ano84j] Anonymous. The British Society for the History of Science report of Council for the year 1983–1984. British Journal for the History of Science, 17(3): 341–345, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1984:E

[Ano84k] Anonymous. Editorial. British Journal for the History of Science, 17(1):ii, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1984:IBR

[Ano84l] Anonymous. Index of books reviewed. British Journal for the History of Science, 17(3): 348–350, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1985:BVIa

[Ano85a] Anonymous. BJH volume 18 issue 1 cover and back matter. British Journal for the History of Science, 18(1):b1– b2, March 1985. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1985:BVIb

[Ano85b] Anonymous. BJH volume 18 issue 1 cover and front matter. British Journal for the History of Science, 18(1):f1–f3, March 1985. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1985:BVIc

[Ano85c] Anonymous. BJH volume 18 issue 2 cover and back matter. British Journal for the History of Science, 18(2):b1– b2, July 1985. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1985:BVId

[Ano85d] Anonymous. BJH volume 18 issue 2 cover and front matter. British Journal for the History of Science, 18(2):f1–f2, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1985:BVIe

[Ano85e] Anonymous. BJH volume 18 issue 3 cover and back matter. British Journal for the History of Science, 18(3):b1–b3, November 1985. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1985:BVIf

[Ano85f] Anonymous. BJH volume 18 issue 3 cover and front matter. British Journal for the History of Science, 18(3):f1–f2, November 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1985:BRa

[Ano85g] Anonymous. Books received.

British Journal for the History of Science, 18(1):122– 126, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026303.

Anonymous:1985:BRb

[Ano85h] Anonymous. Books received. British Journal for the History of Science, 18(2):253– 254, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026356.

Anonymous:1985:BRc

[Ano85i] Anonymous. Books received. British Journal for the History of Science, 18(3):376–379, November 1985. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1985:BSH

[Ano85j] Anonymous. British Society for the History of Science Report of Council for the year 1984– 5. British Journal for the History of Science, 18(3):373–375, November 1985. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1985:E

[Ano85k] Anonymous. Editorial. British Journal for the History of Science, 18(1):ii, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1985:IBR

[Ano85l] Anonymous. Index of books reviewed. British Journal for the History of Science, 18(3): 380–381, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1986:BVIa

[Ano86a] Anonymous. BJH volume 19 issue 1 cover and back matter. British Journal for the History of Science, 19(1):b1– b2, March 1986. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1986:BVIb

[Ano86b] Anonymous. BJH volume 19 issue 1 cover and front matter. British Journal for the History of Science, 19(1):f1–f2, March 1986. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1986:BVIc

[Ano86c] Anonymous. BJH volume 19 issue 2 cover and back matter. British Journal for the History of Science, 19(2):b1– b2, July 1986. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1986:BVId

[Ano86d] Anonymous. BJH volume 19 issue 2 cover and front matter. British Journal for the History of Science, 19(2):f1–f6, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1986:BVIe

[Ano86e] Anonymous. BJH volume 19 issue 3 cover and back matter. British Journal for the History of Science, 19(3):b1–b6, November 1986. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1986:BVIf

[Ano86f] Anonymous. BJH volume 19 issue 3 cover and front matter. British Journal for the History of Science, 19(3):f1–f2, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1986:BRa

[Ano86g] Anonymous. Books received. British Journal for the History of Science, 19(2):236– 240, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026619.

Anonymous:1986:BRb

[Ano86h] Anonymous. Books received. British Journal for the History of Science, 19(3):373–374, November 1986. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1986:BSH

[Ano86i] Anonymous. British Society for the History of Science Report of Council for the year 1985–1986. British Journal for the History of Science, 19(3):

379–384, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026547.

Anonymous:1986:Ea

[Ano86j] Anonymous. Editorial. British Journal for the History of Science, 19(1):1, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1986:Eb

[Ano86k] Anonymous. Editorial. British Journal for the History of Science, 19(3):241, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1986:IBR

[Ano86l] Anonymous. Index of books reviewed. British Journal for the History of Science, 19(3): 375–377, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:BVIa

[Ano87a] Anonymous. BJH volume 20 issue 1 cover and back matter. British Journal for the History of Science, 20(1):b1–b2, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:BVIb

[Ano87b] Anonymous. BJH volume 20 issue 1 cover and front matter. British Journal for the History of Science, 20(1):f1–f3, January 1987. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:BVIc

[Ano87c] Anonymous. BJH volume 20 issue 2 cover and back matter. British Journal for the History of Science, 20(2):b1– b2, April 1987. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:BVId

[Ano87d] Anonymous. BJH volume 20 issue 2 cover and front matter. British Journal for the History of Science, 20(2):f1–f2, April 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:BVIe

[Ano87e] Anonymous. BJH volume 20 issue 3 cover and back matter. British Journal for the History of Science, 20(3):b1– b5, July 1987. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:BVIf

[Ano87f] Anonymous. BJH volume 20 issue 3 cover and front matter. British Journal for the History of Science, 20(3):f1–f2, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:BVIg

[Ano87g] Anonymous. BJH volume 20 issue 4 cover and back matter. British Journal for the History of Science, 20(4):b1–b10, October 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:BVIh

[Ano87h] Anonymous. BJH volume 20 issue 4 cover and front matter. British Journal for the History of Science, 20(4):f1–f2, October 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:BRb

[Ano87i] Anonymous. Book received.

British Journal for the History of Science, 20(2):239– 240, April 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026326.

Anonymous:1987:BRa

[Ano87j] Anonymous. Books received.

British Journal for the History of Science, 20(1):126– 128, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026478.

Anonymous:1987:BRc

[Ano87k] Anonymous. Books received.

British Journal for the History of Science, 20(3):371–373, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:BRd

[Ano87l] Anonymous. Books received. British Journal for the History of Science, 20(4):488–489, October 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:BSH

[Ano87m] Anonymous. British Society for the History of Science Report of Council for the year 1986–1987. British Journal for the History of Science, 20(4): 493–498, October 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026429.

Anonymous:1987:CDE

[Ano87n] Anonymous, editor. Conference on documentary editing. Cambridge University Press, Cambridge, UK, 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). 1–128 pp. Papers from the international conference held in Tarrytown, N.Y., April 11–14, 1985, British J. Hist. Sci. 20 (1987), no. 1(64).

Anonymous:1987:E

[Ano87o] Anonymous. Editorial. British Journal for the History of Science, 20(4):375, October 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:F

[Ano87p] Anonymous. Foreword. British Journal for the History of Science, 20(1):1, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:IBR

[Ano87q] Anonymous. Index of books reviewed. British Journal for the History of Science, 20(4): 490–492, October 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1987:ETR

[Ano87r] Anonymous. L’enseignement technique de la Révolution à nos jours. (French) [Technical teaching in the revolution of our time]. Economica, Paris, France, 1987. ISBN 2-7342-0146-1, 2-7178-1323-3. 783 pp. LCCN ????

Anonymous:1987:SCC

[Ano87s] Anonymous. Science and civilisation in China. Cambridge University Press, Cambridge, UK, third edition, 1987. ISBN 0-521-08690-6. xxv + 485 pp. LCCN ????

Anonymous:1988:BVIa

[Ano88a] Anonymous. BJH volume 21 issue 1 cover and back matter. British Journal for the History of Science, 21(1):b1– b5, March 1988. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:BVIb

[Ano88b] Anonymous. BJH volume 21 issue 1 cover and front matter. British Journal for the History of Science, 21(1):f1–f2, March 1988. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:BVIc

[Ano88c] Anonymous. BJH volume 21 issue 2 cover and back matter. British Journal for the History of Science, 21(2):b1– b2, June 1988. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:BVId

[Ano88d] Anonymous. BJH volume 21 issue 2 cover and front matter. British Journal for the History of Science, 21(2):f1–f2, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:BVIe

[Ano88e] Anonymous. BJH volume 21 issue 3 cover and back matter. British Journal for the History of Science, 21(3):b1– b4, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:BVIf

[Ano88f] Anonymous. BJH volume 21 issue 3 cover and front matter. British Journal for the History of Science, 21(3):f1–f2, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:BVIg

[Ano88g] Anonymous. BJH volume 21 issue 4 cover and back matter. British Journal for the History of Science, 21(4):b1–b8, December 1988. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:BVIh

[Ano88h] Anonymous. BJH volume 21 issue 4 cover and front matter. British Journal for the History of Science, 21(4):f1–f2, December 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:BRa

[Ano88i] Anonymous. Books received.

British Journal for the History of Science, 21(1):137–140, March 1988. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:BRb

[Ano88j] Anonymous. Books received.

British Journal for the History of Science, 21(2):270– 271, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026999.

Anonymous:1988:BRc

[Ano88k] Anonymous. Books received. British Journal for the History of Science, 21(3):387–390, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026858.

Anonymous:1988:BRd

[Ano88l] Anonymous. Books received. British Journal for the History of Science, 21(4):500–501, December 1988. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:BRe

[Ano88m] Anonymous. Books reviewed. British Journal for the History of Science, 21(4):502–504, December 1988. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:E

[Ano88n] Anonymous. Editorial. British Journal for the History of Science, 21(4):391, December 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:IA

[Ano88o] Anonymous. Index of authors. British Journal for the History of Science, 21(4):505, December 1988. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:NEBa

[Ano88p] Anonymous. New Editor for BJHS. British Journal for the History of Science, 21(2):273, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:NEBb

[Ano88q] Anonymous. New Editorial Board for BJHS. British Journal for the History of Science, 21(4):392, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1988:RCY

[Ano88r] Anonymous. Report of Council for the year 1987–1988. British Journal for the History of Science, 21(4):507– 512, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026974.

Anonymous:1988:SP

[Ano88s] Anonymous. The Singer Prize.

British Journal for the History of Science, 21(1):141, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026893.

Anonymous:1989:BVIa

[Ano89a] Anonymous. BJH volume 22 issue 1 cover and back matter. British Journal for the History of Science, 22(1):b1– b3, March 1989. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:BVIb

[Ano89b] Anonymous. BJH volume 22 issue 1 cover and front matter. British Journal for the History of Science, 22(1):f1–f2, March 1989. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:BVIc

[Ano89c] Anonymous. BJH volume 22 issue 2 cover and back matter. British Journal for the History of Science, 22(2):b1– b2, July 1989. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:BVId

[Ano89d] Anonymous. BJH volume 22 issue 2 cover and front matter. British Journal for the History of Science, 22(2):f1–f2, July 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:BVIe

[Ano89e] Anonymous. BJH volume 22 issue 3 cover and back matter. British Journal for the History of Science, 22(3):b1– b4, September 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:BVIf

[Ano89f] Anonymous. BJH volume 22 issue 3 cover and front matter. British Journal for the History of Science, 22(3):f1–f4, September 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:BVIg

[Ano89g] Anonymous. BJH volume 22 issue 4 cover and back matter. British Journal for the History of Science, 22(4):b1–b13, December 1989. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:BVIh

[Ano89h] Anonymous. BJH volume 22 issue 4 cover and front matter. British Journal for the History of Science, 22(4):f1–f2, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:BRa

[Ano89i] Anonymous. Books received.

British Journal for the History of Science, 22(1):125–127, March 1989. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:BRb

[Ano89j] Anonymous. Books received.

British Journal for the History of Science, 22(2):254– 256, July 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026673.

Anonymous:1989:BRc

[Ano89k] Anonymous. Books received. British Journal for the History of Science, 22(3):397–399, September 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026914.

Anonymous:1989:BRd

[Ano89l] Anonymous. Books received. British Journal for the History of Science, 22(4):487–489, December 1989. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:BRe

[Ano89m] Anonymous. Books reviewed. British Journal for the History of Science, 22(4):491–493, December 1989. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:C

[Ano89n]  Anonymous.  Correction.

British Journal for the History of Science, 22(4):439, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:E

[Ano89o] Anonymous. Editorial. British Journal for the History of Science, 22(1):1, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:IA

[Ano89p] Anonymous. Index of authors.

British Journal for the History of Science, 22(4):494, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1989:RCY

[Ano89q] Anonymous. Report of Council for the year 1988–1989. British Journal for the History of Science, 22(4):495– 512, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026955.

Anonymous:1990:EIS

[Ano90] Anonymous.

Échanges d’influences scientifiques et techniques entre pays européens de 1780 à 1830. (French) [Exchanges of scientific and technical influence between the European countries from 1780 to 1830], volume 5 of Actes du 114e Congrès national des sociétés savantes, Paris, 3-9 avril 1989. Section histoire des sciences et des techniques; Colloques du CTHS, 0993-8192. Editions du CTHS, Paris, France, 1990. ISBN 2-7355-0208-2. 170 pp. LCCN Q127.E8 C663 1989. 200F.

Anonymous:1992:EPN

[Ano92] Anonymous.

Elements de la philosophie de Newton. (French) [Elements of the philosophy of Newton], volume 15 of The complete works of Voltaire. The Voltaire Foundation Taylor Institution, Oxford, UK, 1992. ISBN 0-7294-0374-2. xxii + 805 pp. LCCN ????

Anonymous:1993:SMA

[Ano93] Anonymous. Les scientifiques et la montagne: actes du 116e Congrès national des sociétés savantes, Chambéry, 29 avril-4 mai 1991, Section d’histoire des sciences et des techniques. (French) [The scientists and the mountain: actes of the 116th National Congress of Scientific Societies, 29 April–4 May, 1991, section for the history of science and technology]. Actes du 116e Congrès national des sociétés savantes. Section d’histoire des sciences et des techniques, 0998-5654. Editions du CTHS, Paris, France, 1993. ISBN 2-7355-0272-4. 293 pp. LCCN Q127.E8 C663 1991.

Anonymous:2004:CIE

[Ano04] Anonymous. Correspondence: Images of the Eighteenth Century: Polemic: Style and aesthetics, volume 2004:07 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2004. ISBN 0-7294-0844-2. v + 293 pp. LCCN DC33.4 .C84 2004.

Anonymous:2005:CDH

[Ano05] Anonymous. Correspondence:

Dialogue: History of ideas, volume 2005:07 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2005. ISBN 0-7294-0861-2. v + 342 pp. LCCN DC33.4 .C82 2005.

Anonymous:2007:ETW

[Ano07] Anonymous. Enlightenment and tradition: women’s studies: Montesquieu, volume 2007:06 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2007. ISBN 0-7294-0908-2 (paperback). xi + 316 pp. LCCN PQ261 .E55 2007.

Anthony:1963:BRM

[Ant63] H. D. Anthony. Book review:

The Mechanical Technology of Greek and Roman Antiquity. By A. G. Drachmann. Pp. 220. 72 illustrations. Copenhagen, 1963. Hafner Publishing Company Ltd., London. 65s. British Journal for the History of Science, 1(3):285–286, June 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024927.

Anthony:1964:BRS

[Ant64] H. D. Anthony. Book review:

Science and the Renaissance. By W. P. D. Wightman. Vol. I: An Introduction to the Study of the Emergence of the Sciences in the Sixteenth Century. Pp. xvi + 327. Vol. II: An annotated Bibliography of the Sixteenth-Century Books relating to the Sciences in the Library of the University of Aberdeen. Pp. xx + 293. Oliver and Boyd, Edinburgh. 1963. Each vol. £2 2s. British Journal for the History of Science, 2(1):76–77, June 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025089.

Albury:1977:RMS

[AO77] W. R. Albury and D. R. Oldroyd. From Renaissance mineral studies to historical geology, in the light of Michel Foucault’s the order of things. British Journal for the History of Science, 10(3):187– 215, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025840.

Appleyard:1993:UPS

[App93] Bryan Appleyard. Understanding the present: science and the soul of modern man. Picador. Pan Books, London, UK, 1993. ISBN 0-330-32013-0. xix + 283 pp. LCCN ????

Arato:1991:SDC

[Ara91] Franco Arato. Il secolo delle cose: scienza e storia in Francesco Algarotti. (Italian) [The century of things: science and history in Francisco Algarotti], volume 25 of Ricerche, studi, strumenti; Letteratura.

Marietti, Genova, Italy, 1. edition, 1991. ISBN 88-211-8566-X, 88-211-9566-X. 173 pp. LCCN Q127.I8 A72 1991; PQ4683.A2 Z565 1991.

Armitage:1962:BRH

[Arm62] A. Armitage. Book review:

History of Science: An Annual Review of Literature, Research and Teaching. Volume 1: 1962. Editors: A. G. Crombie and M. A. Hoskin. Pp. viii + 133. Cambridge: Heffer, 1962. Price 30s. net. British Journal for the History of Science, 1(2):187–188, December 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025134.

Armitage:1964:BRH

[Arm64] A. Armitage. Book review:

History of Science: An Annual Review of Literature, Research and Teaching. Volume II: 1963. Editors: A. C. Crombie and M. A. Hoskin. Pp. vi + 177. Cambridge: Heffer, 1963. Price: Paper bound, 21s. net; Cloth, 30s. net. British Journal for the History of Science, 2(1):80–81, June 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025092.

Armitage:1966:BRS

[Arm66] A. Armitage. Book review: Sixteenth and Seventeenth Centuries: Kepler’s Conversation with Galileo’s Sidereal Messenger. First Complete Translation, with an Introduction and Notes. By Edward Rosen. Pp. xx + 164. New York and London: Johnson Reprint Corporation, 1965. $9. British Journal for the History of Science, 3(1):80–81, June 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025109.

Armitage:1967:BRG

[Arm67] A. Armitage. Book review: Galileo Reappraised. Edited by Carlo L. Golino. University of California Press; London: Cambridge University Press. Pp. x + 110. 1966. 35s. British Journal for the History of Science, 3(3):297–298, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025058.

Armitage:1969:BRA

[Arm69] A. Armitage. Book review: Astronomy and Physics The Celestial Worlds Discover’d. By Christian Huygens. London: F. Cass. 1968. Pp. (vi) + vi + 160, with 5 figs. 45s. British Journal for the History of Science, 4(4):406–407, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025238.

Armitage:1970:BRA

[Arm70] A. Armitage. Book review: Astronomy and Navigation The High Firmament: A Survey of Astronomy in English Literature. By A. J. Meadows. Leicester: Leicester University Press. 1969. Pp. x + 207. 4 figs. 42s. British Journal for the History of Science, 5(1):92, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025360.

Armstrong:1983:PAB

[Arm83] David Armstrong. Political anatomy of the body: medical knowledge in Britain in the Twentieth Century. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24746-2. xii + 146 pp. LCCN R486 .A75 1983.

Alhazen:1989:OIA

[AS89] Alhazen and A. I. Sabra.

The Optics of Ibn al-Haytham. Books I–III, On direct vision, volume 40 of Studies of the Warburg Institute. Warburg Institute, University of London, London, UK, 1989. ISBN 0-85481-072-2. ???? pp. LCCN QC353 .A3313 1989.

Ash:1996:FMS

[AS96] Mitchell G. Ash and Alfons Söllner, editors. Forced migration and scientific change: emigre German-speaking scientists and scholars after 1933. Publications of the German Historical Institute. German Historical Institute, Washington, DC, USA, 1996. ISBN 0-521-49741-8 (hardcover). xviii + 301 pp. LCCN E184.G3 F73

1996. URL http://www.loc.
gov/catdir/description/
cam027/95024894.html;
http://www.loc.gov/
catdir/samples/cam031/
95024894.html;   http://
www.loc.gov/catdir/toc/
cam026/95024894.html.

Ashford:1985:PPL

[Ash85] Oliver M. Ashford. Prophet — or professor?: the life and work of Lewis Fry Richardson. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-774-8. xiv + 304 pp. LCCN QC858.R53 A9 1985. US$17.00. URL http:

//www.loc.gov/catdir/
enhancements/fy0745/
84235055-d.html.

Ash:1995:GPG

[Ash95] Mitchell G. Ash. Gestalt psychology in German culture, 1890–1967: holism and the quest for objectivity. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-47540-6 (hardcover). xii + 513 pp. LCCN BF203 .A84

1995. URL http://www.loc.
gov/catdir/description/
cam026/94036273.html;
http://www.loc.gov/
catdir/toc/cam026/
94036273.html.

Aspray:1988:BRP

[Asp88] William Aspray. Book review: P. Masani (ed.). Norbert Wiener: Collected Works. Volume IV. Cambridge, MA, and London: MIT Press, 1985. Pp. xx + 1083. ISBN 0-262-23123-9. £69.50. British Journal for the History of Science, 21(1):116, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026871.

Aspray:1990:JNO

[Asp90] William Aspray. John von Neumann and the origins of modern computing. History of computing. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-01121-2. xvii + 376 pp. LCCN QA76.5 .A77446 1990.

Asquith:1980:BRN

[Asq80] Pamela J. Asquith. Book review: Nineteenth and Twentieth Centuries: The Origins and Rise of Ethology: The Science of the Natural Behaviour of Animals. By William H. Thorpe. London: Heinemann Educational Books, 1979. Pp. xi + 174. £5.90. British Journal for the History of Science, 13(3): 273–274, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026209.

Alvarez:1987:DAS

[AT87] Luis W. Alvarez and W. Peter Trower. Discovering Alvarez: selected works of Luis W. Alvarez, with commentary by his students and colleagues. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-81304-5. ix + 272 + 16 pp. LCCN QC71 .A462

1987. URL http://www.loc.
gov/catdir/enhancements/
fy0608/87012766-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/87012766-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/87012766-b.html.

Atherton:1990:BRV

[Ath90] Margaret Atherton. Berkeley’s revolution in vision. Cornell University Press, Ithaca, NY, USA, 1990. ISBN 0-8014-2358-9. xii + 249 pp. LCCN B1339 .A72 1990.

Adams:1977:PJA

[ATKL08] John Adams, Robert Joseph Taylor, Mary-Jo Kline, and Gregg L. Lint. Papers of John Adams. The Adams papers: Series III, General correspondence and other papers of the Adams statesmen. Belknap Press of Harvard University Press, Cambridge, MA, USA and London, UK, 1977– 2008. ISBN 0-674-65441-2 (vol. 1–2), 0-674-65442-0 (vol. 3– 4), 0-674-65443-9 (vol. 5–6), 0-674-65444-7 (vol. 7–8), 0-674-65445-5 (vol. 9–10), 0-674-01136-8 (vol. 11), 0-674-01281-X (vol. 12), 0-674-01812-5 (vol. 13), (v., 14,, ), 0-674-02607-1 (vol. 14). ???? pp. LCCN E302 .A26 1977. URL http://www.

loc.gov/catdir/toc/fy052/
77004707.html.

Atran:1986:FHN

[Atr86] Scott Atran. Fondements de l’histoire naturelle: pour une anthropologie de la science. (French) [Origins of naturnal history: towards an anthropology of science]. Collection “Le Genre humain”. Editions Complexe, Bruxelles, Belgium, 1986. ISBN 2-87027-180-8. 244 pp. LCCN QH83 .A87 1986.

Austin:1967:UO

[Aus67] R. H. Austin. Uranus observed. British Journal for the History of Science, 3(3): 275–284, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025052.

Austen:1986:BRA

[Aus86]  Brian Austen.  Book re-

view: Alexander Fenton & Geoffrey Stell, eds. Loads and Roads in Scotland and Beyond: Road Transport over 6000 Years. Edinburgh: John Donald Publishers Ltd, 1984 Pp. vii + 144. ISBN 0-85976-107-X £8.50. British Journal for the History of Science, 19(2):210, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026598.

Arafat:1971:LSS

[AW71] W. ‘Arafat and H. J. J. Winter. The light of the stars — A short discourse by Ibn Al-Haytham. British Journal for the History of Science, 5 (3):282–288, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025317.

Ash:1987:PTC

[AW87]  Mitchell  G.

Ash and William Ray Woodward, editors. Psychology in Twentieth-Century thought and society. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-32523-4. ix + 320 pp. LCCN BF105 .P79 1987.

Amrine:1987:GSR

[AZW87] Frederick Amrine, Francis J. Zucker, and Harvey Wheeler, editors. Goethe and the Sciences: a Reappraisal, volume 97 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1987. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346. xvii + 442 pp. LCCN Q174 .B67 vol. 97. US$79.00. URL https://

                                        link.springer.com/book/
                                        10.1007/978-94-009-3761-1.

B:1969:TDH

[B.69] W. H. B. Theses and dissertations in history of science in British universities in progress, or recently completed. British Journal for the History of Science, 4(4):430, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Barnes:1982:SSS

[B^{+}82] Jonathan Barnes et al., editors. Science and speculation: studies in Hellenistic theory and practice. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24689-X (Cambridge), 2-901725-47-3 (Éd. de la Maison des Sciences de l’Homme). xxvi + 351 pp. LCCN Q175 .S4175

1982. URL http://www.loc.
gov/catdir/enhancements/
fy0632/82004221-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0632/82004221-t.html.

Blondel:1989:SHS

[B^{+}89] Christine Blondel et al., editors. Studies in the history of scientific instruments = Etudes sur l’histoire des instruments scientifiques. Rogers Turner Books, London, UK, 1989. ISBN 0-9502557-8-5. LCCN Q184 .S44 1987. URL http:

//catalog.hathitrust.org/
api/volumes/oclc/
22878352.html. Papers pre-

sented at the 7th Symposium of the Scientific Instruments Commission of the Union internationale d’histoire et de philosophie des sciences.

Anonymous:1995:DCA

[B^{+}95] Patrice Bret et al., editors. Débats et chantiers actuels autour de Lavoisier et de la révolution chimique. (French) [Debates and works about Lavoisier and the chemical revolution], volume 68, 1–2 of Revue d’histoire des sciences. 48e année, Section d’histoire des sciences. Presses universitaires de France, Paris, France, 1995. ISBN 2-13-046972-8. 230 pp. LCCN ????

Barker:1991:RCE

[BA91] Peter Barker and Roger Ariew, editors. Revolution and continuity: essays in the history and philosophy of early modern science, volume 24 of Studies in philosophy and the history of philosophy. Catholic University of America Press, Washington, DC, USA, 1991. ISBN 0-8132-0738-X. v + 222 pp. LCCN B21 .S78 vol. 24.

Baatz:1990:KCS

[Baa90] Simon Baatz. Knowledge, culture, and science in the metropolis: the New York Academy of Sciences, 1817–1970, volume 584 of Annals of the New York Academy of Sciences, 0077-8923. New York Academy of Sciences, New York, NY, USA, 1990. ISBN 0-89766-545-7, 0-89766-546-5 (paperback). ix + 269 pp. LCCN Q11 .N5 vol. 584.

Badcock:1962:POR

[Bad62] A. W. Badcock. Physical optics at the Royal Society 1660– 1800. British Journal for the History of Science, 1(2):99– 116, December 1962. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025127.

Badash:1968:RU

[Bad68] Lawrence Badash. Review:

[untitled]. British Journal for the History of Science, 4(2): 194–195, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025167.

Badash:1971:BRH

[Bad71] Lawrence Badash. Book review: History and Philosophy of Science Historical Studies in the Physical Sciences. Vol. I. Ed. by Russell McCormmach. University of Pennsylvania Press & Oxford University Press. 1969 [1970]. Pp. x + 314. £4. British Journal for the History of Science, 5 (3):292–293, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025319.

Badash:1979:SSR

[Bad79] Lawrence Badash. The suicidal success of radiochemistry. British Journal for the History of Science, 12(3):245– 256, November 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025999.

Badash:1983:BRM

[Bad83] Lawrence Badash. Book review: Mathematical Sciences: Peter Robertson, The early years: the Niels Bohr Institute, 1921–1930. Copenhagen: Akademisk Forlag, Universitetsforlaget i København, 1979. Pp. 175. British Journal for the History of Science, 16(3):296–297, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026166.

Badash:1985:KRK

[Bad85] Lawrence Badash. Kapitza, Rutherford, and the Kremlin. Yale University Press, New Haven, CT, USA, 1985. ISBN 0-300-01465-1. xi + 129 pp. LCCN QC16.K25 B3 1985.

Badash:1989:BRR

[Bad89] Lawrence Badash. Book review: Robert Chadwell Williams, Klaus Fuchs, Atom Spy. Cambridge, Massachusetts, and London: Harvard University Press, 1987. Pp. x + 267. ISBN 0-674-50507-7. $25.00. British Journal for the History of Science, 22(2):247–248, July 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026666.

Benguigui:1984:TED

[BaJANJ84] Isaac Benguigui, abbé Jean Antoine Nollet, and Jean Jallabert. Théories électriques du XVIIIème siècle. Correspondance entre l’abbé Nollet (1700–1770) et le physicien genevois Jean Jallabert (1712–1768). (French) [Electrical theories of the 18th Century. Correspondence between the Abbey Nollet (1700–1770) and the Genevan physician Jean Jallabert (1712–1768)]. Ed. Georg, Genève, Switzerland, 1984. ISBN 2-8257-0099-1. 234 + 1 pp. LCCN QC516 .B46 1984. URL http:

//catalog.hathitrust.org/
api/volumes/oclc/
11454840.html.

Baker:1996:MNI

[Bak96] Jeffrey P. Baker. The machine in the nursery: incubator technology and the origins of newborn intensive care. The Johns Hopkins University Press, Baltimore, MD, USA, 1996. ISBN 0-8018-5173-4. x + 247 pp. LCCN RJ253.5 .B34

1996. URL http://www.loc.
gov/catdir/bios/jhu052/
95040320.html;
http://www.loc.gov/
catdir/description/
jhu051/95040320.html.

Baldwin:1983:BRT

[Bal83] Mark Baldwin. Book review: Technology and Instruments Transport Technology and Social Change. Papers delivered at Tekniska Museet Symposium No 2, Stockholm, 1979. Ed. by Per Sörbom. Stockholm: Tekniska Museet, 1980. Pp 296. Sw Kr 110/$25.00. British Journal for the History of Science, 16(2):210–211, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026251.

Baldwin:1984:BRH

[Bal84] Mark Baldwin. Book review: History of Technology: K. H. Vignoles, Charles Blacker Vignoles: romantic engineer. Cambridge: Cambridge University Press, 1982. Pp. xiif 187. ISBN 0-521-23930-3. £18.00. British Journal for the History of Science, 17(3): 329, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026644.

Bala:1991:IMB

[Bal91] Poonam Bala. Imperialism and medicine in Bengal: a sociohistorical perspective. Sage Publications, New Delhi, India, 1991. ISBN 81-7036-245-8 (India), 0-8039-9100-2 (US). 174 pp. LCCN R607.B46 B35

                                        1991. URL http://www.loc.
                                        gov/catdir/enhancements/

fy0657/91022537-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0657/91022537-t.html.

Banks:1988:RHS

[Ban88] R. E. R. Banks. Resources for the history of science in the libraries of the British Museum (natural history). British Journal for the History of Science, 21(1):91–97, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026864.

Barocas:1973:BRE

[Bar73a] V. Barocas. Book review:

Eighteenth Century: Tobias Mayer’s Opera Inedita. The First Translation of the Lichtenberg Edition of 1775. Ed. by Eric G. Forbes. Macmillan: London, 1971. Pp. x + 166. £6.90. The Euler-Mayer Correspondence (1751–1755) Ed. and trans, by Eric G. Forbes. London: Macmillan, 1971. Pp. x + 118. £4.25. British Journal for the History of Science, 6(4):448–449, December 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025519.

Bartholomew:1973:LEA

[Bar73b] Michael Bartholomew. Lyell and evolution: An account of Lyell’s response to the prospect of an evolutionary ancestry for man. British Journal for the History of Science, 6(3):261– 303, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025445.

Bartholomew:1975:BRG

[Bar75] Michael J. Bartholomew. Book review: Geology Short Classification and Description of the Various Rocks (1786). By Abraham Gottlob Werner. Facsimile, with translation and introduction by Alexander M. Ospovat. New York: Hafner, 1971. £6.95. The Strata of Derbyshire (1811). By White Watson. Facsimile, with introduction by Trevor D. Ford. Buxton: Moorland Reprints, 1973. £3.60. British Journal for the History of Science, 8(3): 252, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025564.

Barocas:1976:BRA

[Bar76a] V. Barocas. Book review: Astronomy The Search for the Nebulae. By Kenneth Glyn Jones. Chalfont St Giles: Alpha Academic, Science History Publications, 1975. Pp. ix + 84. £3.50. British Journal for the History of Science, 9(3): 327–328, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025689.

Bartholomew:1976:BRN

[Bar76b] Michael Bartholomew. Book review: Nineteenth and Twentieth Centuries: Charles Darwin: Thomas Henry Huxley. Autobiographies. Ed. by Gavin de Beer. London: Oxford University Press, 1974. Pp. xxvi + 123. £3.30. British Journal for the History of Science, 9(1):85, March 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025719.

Bartholomew:1976:NPN

[Bar76c] Michael Bartholomew. The non-progress of nonprogression: Two responses to Lyell’s doctrine. British Journal for the History of Science, 9(2):166–174, July 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025804.

Barnes:1977:BRS

[Bar77] Barry Barnes. Book review:

Sociology of Science The Idea of Social Structure. Papers in Honour of Robert K. Merton. Edited by Lewis A. Coser, New York: Harcourt Brace Jovanovich, 1975. Pp. ix + 547. £7.55. British Journal for the History of Science, 10(3):254– 255, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025846.

Barnes:1979:BRS

[Bar79] Barry Barnes. Book review: Shorter Notices of Books Models of Man. By Martin Hollis. Cambridge: Cambridge University Press, 1977. Pp. vii + 198. £7.00. British Journal for the History of Science, 12 (1):104, March 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025678.

Barnes:1981:BRP

[Bar81] Barry Barnes. Book review: Philosophy of Science: Science, Revolution and Discontinuity. By John Krige. Brighton: Harvester Press, 1980. Pp. 231. £18.95. British Journal for the History of Science, 14(2): 206–208, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025950.

Barnes:1983:BRS

[Bar83]  Barry Barnes.  Book re-

view: Science and Society Joseph Agassiz, Science and society: studies in the sociology of science. Boston studies in the philosophy of science, volume 65. Dordrecht and Boston: Reidel, 1981. Pp. xxi + 531. British Journal for the History of Science, 16(3): 299, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026168.

Barraclough:1984:SBB

[Bar84] Kenneth Charles Barraclough, editor. Steelmaking before Bessemer, volume 279 of Book / Metals Society. The Metals Society, London, UK, 1984. ISBN 0-904357-53-8, 0-904357-64-3. ???? pp. LCCN ????

Barbour:1989:ARM

[Bar89a] Julian B. Barbour, editor. Absolute or relative motion?: a study from a Machian point of view of the discovery and the structure of dynamical theories. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-32467-X. xv + 746 pp. LCCN QC133 .B37 1989.

Barbourger:1989:DD

[Bar89b] Julian B. Barbour. The discovery of dynamics, volume 1 of Absolute or relative motion? Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-32467-X. xv + 746 pp. LCCN ????

Barker:1989:BSG

[Bar89c] David Barker. The biology of stupidity: Genetics, eugenics and mental deficiency in the inter-war years. British Journal for the History of Science, 22 (3):347–375, September 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026901.

Bartholin:1991:EBI

[Bar91] Erasmus Bartholin. Experiments on birefringent Icelandic crystal, volume 40 of Acta historica scientiarum naturalium et medicinalium, 0065-1311. Danish National Library of Science and Medicine, Copenhagen, Denmark, 1991. ISBN 87-7709-010-1. 62, 60 pp. LCCN IN PROCESS.

Barton:1994:AA

[Bar94] Tamsyn Barton. Ancient astrology. Sciences of antiquity. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1994. ISBN 0-415-08066-5, 0-415-11029-7 (paperback). xxv + 245 pp. LCCN BF1671 .B378

                                        1994. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0648/93032441-d.html.

Basalla:1988:ET

[Bas88] George Basalla. The evolution of technology. Cambridge history of science. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-22855-7, 0-521-29681-1 (paperback). viii + 248 pp. LCCN T15 BAS 1988; 509; T; Hist; 609; 125; T15.BAS.

Bauckham:1975:SRW

[Bau75] Richard Bauckham. Science and religion in the writings of Dr. William Fulke. British Journal for the History of Science, 8(1):17–31, March 1975.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025814.

Baumler:1984:PES

[Bäu84] Ernst Bäumler. Paul Ehrlich, scientist for life. Holmes and Meier, New York, NY, USA, 1984. ISBN 0-8419-0837-0. xvi + 288 + 32 pp. LCCN R512.E4 B3313 1984.

Bauer:1995:RNT

[Bau95] Martin W. Bauer, editor. Resistance to new technology: nuclear power, information technology, and biotechnology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-45518-9 (hardcover). LCCN T174.5 .R48 1995. URL http:

//www.loc.gov/catdir/
description/cam026/
94026745.html;   http://
www.loc.gov/catdir/toc/
cam021/94026745.html.

Berkeley:1982:LTJ

[BB82]  Edmund  Berke-

ley and Dorothy Smith Berkeley. The life and travels of John Bartram from Lake Ontario to the River St. John. University Presses of Florida, Tallahassee, FL, USA, 1982. ISBN 0-8130-0700-3. xv + 376 pp. LCCN QK31.B3 B47.

Blatt:1984:CCA

[BB84] Sidney J. (Sidney Jules) Blatt and Ethel S. Blatt. Continuity and change in art: the development of modes of representation. L. Erlbaum, Hillsdale, NJ, USA, 1984. ISBN 0-89859-342-5. xix + 411 pp. LCCN N7430.5 .B55

1984. URL http://www.loc.
gov/catdir/enhancements/
fy0742/84004201-d.html.

Barker-Benfield:1992:CSS

[BB92] G. J. Barker-Benfield. The culture of sensibility: sex and society in Eighteenth-Century Britain. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-03713-4. xxxiv + 520 pp. LCCN HQ1593 .B37

1992. URL http://www.loc.
gov/catdir/description/
uchi052/91047945.html;
http://www.loc.gov/
catdir/toc/uchi051/
91047945.html.

Bertotti:1990:MCR

[BBBM90] B. Bertotti, R. Balbinot, S. Bergia, and A. Messina, editors. Modern Cosmology in Retrospect. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37213-5. xx + 426 pp. LCCN QB981 .M774

1990. URL http://www.loc.
gov/catdir/description/
cam024/90041803.html;
http://www.loc.gov/
catdir/toc/cam028/
90041803.html.

Bloch:1988:EFH

[BBC88] Olivier Bloch, Bernard Balan, and Paulette Carrive, editors. Entre forme et histoire: la formation de la notion de développement à l’âge classique. (French) [Between form and history: the formation of the notion of development in the classical age]. Philosophie. Meridiens Klincksieck, Paris, France, 1988. ISBN 2-86563-212-1. 310 pp. LCCN B818 .E58 1988.

Bernal:1972:O

[BC72] J. D. Bernal and J. G.

Crowther. Obituaries. British Journal for the History of Science, 6(1):104–105, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025283.

Bessel:1989:STE

[BC89] Richard Bessel and Colin Chant, editors. Science, technology and everyday life, 1870– 1950. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1989. ISBN 0-415-00037-8, 0-415-03557-0. 391 pp. LCCN ????

Blake-Coleman:1992:CWE

[BC92] B. C. (Barrie Charles) Blake-Coleman. Copper wire and electrical conductors: the shaping of a technology. Harwood Academic Publishers, Chur, Switzerland and New York, NY, USA, 1992. ISBN 3-7186-5200-5. xx + 284 pp. LCCN TS270 .B58

1991. URL http://www.loc.
gov/catdir/enhancements/
fy0653/91040910-d.html.

Brooke:1998:RNE

[BC98] John Hedley Brooke and Geoffrey N. Cantor. Reconstructing nature: the engagement of science and religion: Glasgow Gifford lectures. Glasgow Gifford Lectures. T and T Clark, Edinburgh, Scotland, 1998. ISBN 0-567-08600-3. xii-367 pp. LCCN ????

Babbage:1989:WCB

[BCKR89] Charles Babbage, Martin Campbell-Kelly, and Peter M. Roget, editors. The works of Charles Babbage. The Pickering masters. W. Pickering, London, UK, 1989. ISBN 1-85196-005-8 (set), 1-85196-501-7 (vol. 1), 1-85196-502-5 (vol. 2), 1-85196-503-3 (vol. 3), 1-85196-504-1 (vol. 4), 1-85196-505-X (vol. 5), 1-85196-506-8 (vol. 6), 1-85196-507-6 (vol. 7), 1-85196-508-4 (vol. 8), 1-85196-509-2 (vol. 9), 1-85196-510-6 (vol. 10), 1-85196-511-4 (vol. 11). LCCN QA37.2. 11 volumes.

Blondel:1995:RCU

[BD95] Christine Blondel and Matthias Dörries, editors. Restaging Coulomb: usages, controverses et réplications autour de la balance de torsion. (French) [Restaging Coulomb: Uses, controversies, and replications of the torsion balance], volume 15 of Biblioteca di Nuncius. Studi e testi. Leo S. Olschki, Firenze, Italy, 1995.

ISBN 88-222-4196-7. 167 pp. LCCN ????

Brown:1989:PQP

[BDH89] Laurie M. Brown, Max Dresden, and Lillian Hoddeson, editors. Pions to quarks: particle physics in the 1950s: based on a Fermilab symposium. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-30984-0. LCCN QC793.16 .P56

1989. URL http://www.loc.
gov/catdir/description/
cam023/88025644.html;
http://www.loc.gov/
catdir/toc/cam028/
88025644.html.

Boyle:1996:FEV

[BDH96] Robert Boyle, Edward Bradford Davis, and Michael Cyril William Hunter. A free enquiry into the vulgarly received notion of nature. Cambridge texts in the history of philosophy. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56100-0 (hardcover), 0-521-56796-3 (paperback). xxxvi + 171 pp. LCCN Q155 .B68

1996. URL http://www.loc.
gov/catdir/description/
cam027/96010997.html;
http://www.loc.gov/
catdir/toc/cam024/
96010997.html.

Beauchamp:1985:BM

[Bea85a]  K. G. Beauchamp.  Back

matter. British Journal for the History of Science, 18(3): 382, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026411.

Beauchamp:1985:BRR

[Bea85b] K. G. Beauchamp. Book review: R. Moreau, The computer comes of age. The people, the hardware, and the software. Translated by J. Howlett. Cambridge, Mass, and London: MIT Press, 1984. Pp. x + 227. ISBN 0-262-13194-3. £18.95. British Journal for the History of Science, 18(3): 372, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026408.

Beaujouan:1991:PRN

[Bea91] Guy Beaujouan. Par raison de nombres: l’art du calcul et les savoirs scientifiques médiévaux. (French) [Because of numbers: the art of calculation and medieval scientific knowledge], volume CS344 of Variorum collected studies series. Variorum, Aldershot, Hampshire, UK, 1991. ISBN 0-86078-281-6. ???? pp. LCCN Q124.97 .B43 1991.

Beauchamp:1997:EE

[Bea97] K. G. Beauchamp. Exhibiting electricity, volume 21 of IEE history of technology series. Institution of Electrical Engineers, London, UK, 1997.

ISBN 0-85296-895-7. xii + 337 pp. LCCN TK6.A1 B43 1997.

Bechler:1975:LAM

[Bec75] Zev Bechler. ‘A less agreeable matter’: The disagreeable case of Newton and achromatic refraction. British Journal for the History of Science, 8(2):101–126, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025636.

Bechler:1982:CNR

[Bec82] Z. Bechler, editor. Contemporary Newtonian research, volume 9 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1982. ISBN 90-277-1303-0. 241 pp. LCCN QC16.N7 C66 1982.

Becher:1986:VSN

[Bec86] Harvey W. Becher. Voluntary science in Nineteenth Century Cambridge University to the 1850’s. British Journal for the History of Science, 19 (1):57–87, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026485.

Bechmann:1988:AST

[Bec88] Roland Bechmann. About some technical sketches of Villard de Honnecourt’s manuscript.

New light on deleted diagrams: an unknown drawing. British Journal for the History of Science, 21(3):341– 361, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026833.

Bechler:1991:NPC

[Bec91a] Zev Bechler. Newton’s Physics and the Conceptual Structure of the Scientific Revolution, volume 127 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346. xviii + 588 pp. LCCN Q174 .B67

vol. 127. URL https://link.
springer.com/book/10.
1007/978-94-011-3276-3.

Bechmann:1991:VHP

[Bec91b] Roland Bechmann. Villard de Honnecourt: la pensée technique au XIIIe siècle et sa communication. (French) [Villard de Honnecourt: technical thought in the 18th Century and its communication]. Picard, Paris, France, 1991. ISBN 2-7084-0367-2. 384 pp. LCCN NA1053.V6 B4 1991. 420F.

Becquemont:1992:DDE

[Bec92]  Daniel  Bec-

quemont. Darwin, darwinisme, évolutionnisme. (French) [Darwin, Darwinism, and evolutionism]. Histoire des idées, théorie politique et recherches en sciences sociales. Kimé, Paris, France, 1992. ISBN 2-908212-13-7. 341 pp. LCCN QH365.O8 B435 1992. 200F.

Beeson:1989:CO

[Bee89] C. F. C. Beeson. Clockmaking in Oxfordshire, 1400–1850. Museum of the History of Science, Oxford, UK, third edition, 1989. ISBN 0-903364-06-9. 212 pp. LCCN ???? With a new introduction and index by A. V. Simcock.

Beer:1996:OFS

[Bee96] Gillian Beer. Open fields:

science in cultural encounter. Clarendon Press, New York, NY, USA, 1996. ISBN 0-19-818369-0. viii + 341 pp. LCCN PR468.S34 B44

1996. URL http://www.loc.
gov/catdir/enhancements/
fy0639/95043234-d.html.

Baer:2010:EDL

[BEFH10] Werner Baer, Hadi Salehi Esfahani, Giovanni Facchini, and Geoffrey Hewings, editors. Economic development in Latin America: essay in honor of Werner Baer. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2010. ISBN 0-230-22381-8. xxii + 309 pp. LCCN HC125 .E369 2010.

Bennett:1975:HWS

[Ben75] J. A. Bennett. Hooke and Wren and the system of the world: some points towards an historical account. British Journal for the History of Science, 8 (1):32–61, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025815.

Bennett:1979:BRS

[Ben79] J. A. Bennett. Book review: Shorter Notices of Books Science and music in eighteenthcentury Bath. An exhibition in the Holburne of Menstrie Museum, Bath, 22 September 1977–29 December 1977. By A. J. Turner, with the assistance of I. D. Woodfield and contributions by H. S. Torrens. Bath: University of Bath, 1977. Pp. viii + 131. £1.50. British Journal for the History of Science, 12(1):104, March 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bennett:1982:MSC

[Ben82a] J. A. (James A.) Bennett. The mathematical science of Christopher Wren. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24608-3. 148 pp. LCCN Q143.W68 B46

1982. URL http://www.loc.
gov/catdir/description/
cam022/81018045.html;
http://www.loc.gov/
catdir/samples/cam031/
81018045.html;   http://
www.loc.gov/catdir/toc/
cam031/81018045.html.

Benton:1982:BRD

[Ben82b] Ted Benton. Book review:

Darwiniana Greta Jones, Social Darwinism and English thought: the interaction between biological and social theory. Brighton: Harvester Press. Atlantic Highlands N.J.: Humanities Press, 1980. Pp. xiv + 234. £22.50. British Journal for the History of Science, 15(1):74–76, March 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025926.

Bennett:1984:BRH

[Ben84]  J. A. Bennett.  Book re-

views: History of Technology: Anita McConnell, Historical instruments in oceanography. London: Her Majesty’s Stationery Office, 1981. Pp. iv + 51. ISBN 0-11-290324-X. £3.95. Anita McConnell, No sea too deep: the history of oceanographic instruments. Bristol: Adam Hilger Ltd, 1982. Pp. xiii + 162. ISBN 0-85274-416-1. £19.50. British Journal for the History of Science, 17(3): 332, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026648.

Bennett:1985:MA

[Ben85] J. A. Bennett. The mathematicians’ apprenticeship. British Journal for the History of Science, 18(2):212–218, July 1985.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026331.

Beniger:1986:CRT

[Ben86] James R. (James Ralph) Beniger. The control revolution: technological and economic origins of the information society. Harvard University Press, Cambridge, MA, USA, 1986. ISBN 0-674-16985-9. x + 493 pp. LCCN HM258 .B459 1986.

Bennett:1987:DCH

[Ben87] J. A. (James A.) Bennett. The divided circle: a history of instruments for astronomy, navigation, and surveying. Phaidon, Christie’s, Oxford, UK, 1987. ISBN 0-7148-8038-8. 224 pp. LCCN QB85.8 .B46 1987.

Bennett:1989:BRG

[Ben89] J. A. Bennett. Book review: Guy Picolet (ed.). Jean Picard et les Débuts de l’Astronomie de Précision au XVIIe Siècle. Paris: Centre National de la Recherche Scientifique, 1987. Pp. 382. ISBN 2-222-04104-X. FF 150. British Journal for the History of Science, 22(4): 467–468, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026942.

Benninga:1990:HLA

[Ben90] H. (Harm) Benninga. A history of lactic acid making: a chapter in the history of biotechnology. Chemists and chemistry. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0625-2 (US). xxi + 478 pp. LCCN TP248.L3 B46 1990.

Benjamin:1991:SSG

[Ben91] Marina Benjamin, editor. Science and sensibility: gender and scientific enquiry, 1780– 1945. Blackwell, Oxford, UK, 1991. ISBN 0-631-16649-1. x + 295 pp. LCCN Q130 .S327 1991.

Busard:1992:RCR

[BEoCF92] H.  L.  L.

(Hubertus Lambertus Ludovicus) Busard, Euklides, Robert of Chester, and Menso Folkerts. Robert of Chester’s (?) redaction of Euclid’s Elements, the so-called Adelard II version, volume 9 of Science networks: historical studies, 99-0922970-6. Birkhäuser Verlag, Basel, Switzerland, 1992. ISBN 3-7643-2727-8, 3-7643-2728-6 (set). 442–959. pp. LCCN ????

Berry:1979:KHI

[Ber79] Christopher J. Berry. Kuhn and the history of ideas. British Journal for the History of Science, 12(3):295–298, November 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026004.

Berg:1985:AMH

[Ber85] Maxine Berg. The age of manufactures: industry, innovation and work in Britain, 1700– 1820. Fontana history of England. Fontana Press, London, UK, 1985. ISBN 0-00-686019-2. 378 pp. LCCN HD9720.5 .B47 1985.

Beretta:1993:EMD

[Ber93a] Marco Beretta. The enlightenment of matter: the definition of chemistry from Agricola to Lavoisier, volume 15 of Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1993. ISBN 0-88135-152-0. xvi + 396 pp. LCCN QD7 .B52 1993.

BertoloniMeli:1993:EPN

[Ber93b] Domenico Bertoloni Meli. Equivalence and priority: Newton versus Leibniz; including Leibniz’s unpublished manuscripts on the Principia. Oxford science publications. Clarendon Press, New York, NY, USA, 1993. ISBN 0-19-853945-2 (hardcover). ix + 318 pp. LCCN QA803 .B47 1993. US$35.00 (US$52.50

                                        U.S.). URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0639/92018967-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0639/92018967-t.html.

Beretta:1995:SMP

[Ber95] Marco Beretta, editor. The starry messenger and the polar star: scientific relations between Italy and Sweden from 1500 to 1800: catalogue of an exhibition held at the Naturhistoriska riksmuseet — Stockholm, volume 22 of Uppsala studies in history of science. Giunti, Prato, Italy, 1995. ISBN 88-09-20793-9. ISSN 0282-1036. 188 pp. LCCN Q127.I8 B466 1995.

Berthollet:1980:RES

[BG80] Claude-Louis Berthollet and Michelle Goupil. Revue de l’“Essai de statique chimique”. École polytechnique, Palaiseau, France 91128 cedex, 1980. ISBN 2-7302-0019-3. 204 pp. LCCN ????

Beaune:1992:BAD

[BG^{+}92] Jean-Claude Beaune, Jean Gayon, et al., editors. Buffon 88: actes du Colloque international pour le bicentenaire de la mort de Buffon: (Paris, Montbard, Dijon, 14–22 juin 1988). (French) [Buffon 88: acts of the international conference for the bicentenary of the death of Buffon: (Paris, Montbard, Dijon, 14-22 June 1988)]. Science, histoire, philosophie. J. Vrin, Paris, France, 1992. ISBN 2-7116-9755-X. ix + 771 pp. LCCN QH31.B88 B84 1988.

Barrett:1987:CDN

[BGH^{+}87] Paul  H.  Bar-

rett, Peter Jack Gautrey, Sandra Herbert, David Kohn, and Sydney Smith, editors. Charles Darwin’s notebooks, 1836–1844: geology, transmutation of species, metaphysical enquiries. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-35055-7 (hardcover), 0-521-09975-7 (paperback). viii + 747 pp. LCCN QH31.D2; QH365.Z9 B37 1987. US$65.00.

Brown:1983:BPP

[BH83] Laurie M. Brown and Lillian Hoddeson, editors. The Birth of Particle Physics. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Browne:1984:BM

[BH84a] Janet Browne and John Hendry. Back matter. British Journal for the History of Science, 17(1):126, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026068.

Browne:1984:BRR

[BH84b] Janet Browne and John Hendry. Book review: Robert Smith, The expanding universe. Astronomy’s ‘great debate’, 1900– 1931. Cambridge: Cambridge University Press, 1982. Pp. xiv + 220. ISBN 0-521-23212-0. £19. British Journal for the History of Science, 17(1):120– 123, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026058;   http://
www.jstor.org/stable/
4026059; http://www.
jstor.org/stable/4026061;
 http://www.jstor.org/
stable/4026062; http://
www.jstor.org/stable/
4026065;   http://www.
jstor.org/stable/4026066.

Beck:1987:CPAa

[BH87] Anna Beck and Peter Havas.

The collected papers of Albert Einstein. Vol. 1. The Early years, 1879–1902. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-08475-0 (paperback), 0-691-08463-7 (microfiche). xxii + 196 pp. LCCN QC16.E5 A2 1987a. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/
catdir/description/
prin031/87160800.html;
http://www.loc.gov/
catdir/toc/prin031/
87160800.html.

Bitbol-Hesperies:1990:PVC

[BH90] Annie Bitbol-Hespériès. Le principe de vie chez Descartes. (French) [Descartes principle of life]. Bibliothèque d’histoire de la philosophie. Nouvelle série. J. Vrin, Paris, France, 1990. ISBN 2-7116-1034-9. 236 pp. LCCN B1875 .B48 1990. 198F.

Boyle:1994:RB

[BHW94]  Robert  Boyle,

Michael Cyril William Hunter, and William Wotton. Robert Boyle. W. Pickering, London, UK, 1994. ISBN 1-85196-085-6. cvii + 188 pp. LCCN Q143.B77 A3 1994.

Biagioli:1993:GCP

[Bia93] Mario Biagioli. Galileo, courtier: the practice of science in the culture of absolutism. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1993. ISBN 0-226-04559-5 (hardcover). xii + 402 pp. LCCN QB36.G2 B54

                                        1993. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi051/92033736.html;
                                        http://www.loc.gov/
                                        catdir/toc/uchi051/
                                        92033736.html.

Bibby:1974:BRN

[Bib74] Cyril Bibby. Book review: Nineteenth and Twentieth Centuries: H. E. Armstrong and the Teaching of Science. Ed. by W. H. Brock. Cambridge: Cambridge University Press, 1973. Pp. x + 161. £3.80. British Journal for the History of Science, 7(3):299–300, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025547.

Bird:1986:WJ

[Bir86] Graham Bird. William James.

The Arguments of the philosophers. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1986. ISBN 0-7100-9602-X. vi + 221 pp. LCCN B945.J24 B39 1986. US$18.95 (US$45.00 U.S. : est.).

Busard:1991:JNE

[BJ91] H. L. L. (Hubertus Lambertus Ludovicus) Busard and Nemorarius Jordanus. Jordanus de Nemore, De elementis arithmetice artis: a medieval treatise on number theory, volume 22, 1–2 of Boethius. F. Steiner Verlag, Stuttgart, Germany, 1991. ISBN 3-515-05214-3. ???? pp. LCCN QA241 .B835 1991.

Berkeley:1992:MA

[BJB92] George Berkeley, Douglas M.

Jesseph, and George Berkeley. De motu and, The analyst, volume 41 of The New synthese historical library. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1520-0. x + 230 pp. LCCN Q175 .B4674 1992.

Burnett:1975:HSD

[BK75] A. D. Burnett and D. M.

Knight. History of science in Durham libraries. British Journal for the History of Science, 8(1):94–99, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025837.

Boyle:1990:LPR

[BKH90] Robert Boyle, Paul Kesaris, and Michael Cyril William Hunter. Letters and papers of Robert Boyle: from the archives of the Royal Society. Collections from the Royal Society. University Publications of America, Frederick, MD, USA, 1990. ISBN 1-55655-170-3 (microfilm), 1-55655-217-3 (guide). LCCN Q143.B77; Microfilm 95/4348 (Q). US$1,395.00, US$60.00. 16 microfilm reels.

Bacon:1983:RBP

[BL83] Roger Bacon and David C. Lindberg. Roger Bacon’s philosophy of nature: a critical edition, with English translation, introduction, and notes, of De multiplicatione specierum and De speculis comburentibus. Clarendon Press, New York, NY, USA, 1983. ISBN 0-19-858164-5. lxxxi + 420 pp. LCCN QC353 .B313 1983. US$74.00.

Blay:1983:CNP

[Bla83] Michel Blay. La conceptualisation newtonienne des phénomènes de la couleur. (French) [The Newtonian conceptualization of the phenomenon of color]. L’Histoire des sciences. Textes et études. J. Vrin, Paris, France, 1983. ISBN 2-7116-0825-5. 304 pp. LCCN QC494.7 .B57 1983. 99.00F.

Blay:1993:RID

[Bla93]  Michel Blay.  Les

raisons de l’infini: du monde clos à l’univers mathématique. (French) [The reasons for infinity: from a closed world to a mathematical universe]. nrf essais. Gallimard, Paris, France, 1993. ISBN 2-07-072992-3. 258 pp. LCCN BD411 .B55

1993. URL http://catalog.
hathitrust.org/api/
volumes/oclc/29673358.
html.

Bliss:1982:DI

[Bli82] Michael Bliss. The discovery of insulin. Paul Harris publishing, Edinburgh, Scotland, 1982. ISBN 0-86228-056-7. 304 + 16 pp. LCCN ????

Blitz:1992:EEQ

[Bli92] David Blitz. Emergent evolution: qualitative novelty and the levels of reality, volume 19 of Episteme. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1658-4. ix + 239 pp. LCCN B818 .B65

1992. URL http://www.loc.
gov/catdir/enhancements/
fy0823/92004950-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0823/92004950-t.html.

Bloor:1978:PAL

[Blo78] David Bloor. Polyhedra and the abominations of Leviticus. British Journal for the History of Science, 11(3):245– 272, November 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025779. See reply

[Wor79].

Bloor:1979:BRP

[Blo79] David Bloor. Book review: Philosophy of Science Essays in Memory of Imré Lakatos. Edited by R. S. Cohen, P. K. Feyerabend, and M. W. Wartofsky. Dordrecht & Boston: D. Reidel, 1976. Pp. xi + 754 + index. $62.00. British Journal for the History of Science, 12(1):91–92, March 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025668.

Blondel:1982:MAC

[Blo82a]  Christine  Blondel.

A-M. Ampère et la création de l’électrodynamique, 1820–1827, volume 10 of Mémoires de la Section des sciences, 0553-0734. Bibliothèque nationale, Paris, France, 1982. ISBN 2-7177-1643-2. 202 pp. LCCN Q46 .F73216 vol. 10.

Bloor:1982:BRS

[Blo82b] David Bloor. Book review: Science Studies Mary Hesse, Revolutions and Reconstructions in the Philosophy of Science. Brighton: Harvester Press, 1980. Pp. xxvi + 271. £20.00. British Journal for the History of Science, 15(3): 306–309, November 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026225.

Bloor:1984:BRP

[Blo84]  David Bloor.  Book

reviews: Physics and its concepts: Karl R. Popper, Realism and the aim of science. London: Hutchinson, 1983. ISBN 0-09-151450-9 cased. The open universe. An argument for indeterminism. London: Hutchinson, 1982. ISBN 0-09-146180-4 (hardcover). £15.00. Quantum theory and the schism in physics. London, Hutchinson, 1982. ISBN 0-09-146170-7 (hardcover). £15.00. British Journal for the History of Science, 17(2):217–221, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026555.

Blunt:1979:BRB

[Blu79]  Wilfrid Blunt.  Book re-

view: Bibliography and Bibliophily Zoological Illustration: An Essay Towards a History of Printed Zoological Pictures. By David Knight. Folkestone: Dawson; Hamden, Connecticut: Archon Books, 1977. Pp. xii + 2O4. £10.00. British Journal for the History of Science, 12(2):220–221, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025758.

Brock:1976:SDR

[BM76] W. H. Brock and R. M. Macleod. The scientists’ declaration: Reflexions on science and belief in the wake of essays and reviews, 1864-5. British Journal for the History of Science, 9(1):39–66, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025705.

Barthelemy-Madaule:1982:LMP

[BM82]  Madeleine  Barthélemy-

Madaule. Lamarck, the mythical precursor: a study of the relations between science and ideology. MIT Press, Cambridge, MA, USA, 1982. ISBN 0-262-02179-X. xv + 174 pp. LCCN QH31.L2 B3613 1982.

Brock:1984:LLT

[BM84] W. H. (William Hodson) Brock and A. J. (Arthur Jack) Meadows. The lamp of learning: Taylor and Francis and the development of science publishing. Taylor and Francis, Boca Raton, FL, USA, 1984. ISBN 0-85066-253-2. xv + 240 pp. LCCN Z286.S4 B76 1984.

Bradwardine:1989:TBG

[BM89]  Thomas  Bradwardine

and George Molland. Thomas Bradwardine, Geometria speculativa: Latin text and English translation with an introduction and a commentary, volume 18 of Boethius. F. Steiner Verlag, Stuttgart, Germany, 1989. ISBN 3-515-05020-5. 176 pp. LCCN QA444 .B7313 1989.

Brewster:1984:MSS

[BMLC84] Sir David Brewster, A. D.

Morrison-Low, and J. R. R. Christie, editors. “Martyr of science”: Sir David Brewster 1781–1868, Royal Scottish Museum studies. Royal Scottish Museum, Edinburgh, Scotland, 1984. ISBN 0-900733-29-2 (paperback). LCCN QC16.B7; Q143.B.

Biarnais:1982:PN

[BN82] Marie Françoise Biarnais and Sir Isaac Newton. Les “Principia” de Newton: genèse et structure des chapitres fondamentaux avec traduction nouvelle, volume N.S. 2 of Cahiers d’histoire et de philosophie des sciences. Centre de documentation sciences humaines, Paris, France, 1982. ISBN 2-222-03094-3. 287 pp. LCCN ????

Brown:1995:AWA

[BN95] Arthur I. Brown and Ronald L.

Numbers. The antievolution works of Arthur I. Brown, volume 3 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1804-9. xiii + 209 pp. LCCN BS651 .B7874 1995.

Bodemer:1965:BRH

[Bod65] Charles W. Bodemer. Book review: Heat and Life. The Development of the Theory of Animal Heat. By Everett Mendelsohn. Pp. x + 208. Harvard University Press; London: Oxford University Press, 1964. 36s. British Journal for the History of Science, 2(3):268– 269, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024950.

Bodson:1984:BRS

[Bod84] Liliane Bodson. Book review: Science in Culture: Richard Olson, Science Deified and Science Defied. The Historical Significance of Science in Western Culture from the Bronze Age to the Beginnings of the Modern Era ca. 3500 B.C. to ca. A.D. 1640. Berkeley– Los Angeles–London: University of California Press, 1983. Pp. xv + 329 ISBN 0-520-04621-8. £27.50. $42. British Journal for the History of Science, 17(2):244–245, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026577.

Bos:1970:BRM

[Bos70] H. J. M. Bos. Book review:

Mathematics: Nicolas Malebranche, Oeuvres Complètes, Tome XVII-2, Mathematica. Ed. by Pierre Costabel. Paris: Librairie Philosphique J. Vrin. 1968. Pp. ix + 375. Price not stated. British Journal for the History of Science, 5 (1):91, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025359.

Bos:1972:BRM

[Bos72] H. J. M. Bos. Book review: Mathematics: Vorlesungen uber Geschichte der antiken mathematischen Wissenschaften. Erster Band, Vorgriechische Mathematik. By O. Neuge-bauer. Berlin, Heidelberg, and New York: Springer-Verlag, reprint (1969) of 1934 edn. Pp. xii + 212. No price stated. British Journal for the History of Science, 6(2): 206–207, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025295.

Boss:1983:HGD

[Bos83] Jeffrey Boss. Helmont, Glisson, and the Doctrine of the Common Reservoir in the Seventeenth-Century revolution in physiology. British Journal for the History of Science, 16(3):261– 272, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026146.

Bowen:1973:SHH

[Bow73] E. J. Bowen. Sir Harold Hartley. British Journal for the History of Science, 6(3):338– 339, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025467.

Bowers:1974:BRT

[Bow74] Brian Bowers. Book review: Technology Elmer Sperry, Inventor and Engineer. By Thomas Parke Hughes. Baltimore and London: John Hopkins Press, 1971. Pp. xvii + 348. $15. British Journal for the History of Science, 7(3):291, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025540.

Bowler:1983:BRL

[Bow83a] Peter J. Bowler. Book review: Life and earth sciences: Ernst Mayr and William B. Provine, ed., The evolutionary synthesis: perspectives in the unification of biology. Cambridge, Mass, and London: Harvard University Press, 1980. Pp. xiv + 487. (no price stated.). British Journal for the History of Science, 16(3):283–284, November 1983.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026153.

Bowler:1983:EDA

[Bow83b] Peter J. Bowler. The eclipse of Darwinism: anti-Darwinian evolution theories in the decades around 1900. The Johns Hopkins University Press, Baltimore, MD, USA, 1983. ISBN 0-8018-2932-1. xi + 291 pp. LCCN QH361 .B68 1983.

Bowler:1985:BRV

[Bow85] Peter J. Bowler. Book review: Vitezslav Orel, Mendel. Oxford: Oxford University Press (Past Masters), 1984. Pp. vii + 111. ISBN 0-19-287625-2, £7.95 (hardcover). 0-19-287624-4, £1.95 (paperback). British Journal for the History of Science, 18(3):348– 349, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026390.

Bowler:1986:BRG

[Bow86] Peter J. Bowler. Book review:

George W. Stocking, ed. Functionalism Historicized. Essays on British Social Anthropology. Madison: University of Wisconsin Press, 1984. Pp. vi + 244. ISBN 0-299-09900-8. No price given. British Journal for the History of Science, 19 (1):120–121, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026493.

Bowers:1987:BRR

[Bow87a] Brian Bowers. Book review: Robert Friedel & Paul Israel, With Bernard S. Finn. Edison’s Electric Light, Biography of an Invention. New Brunswick: Rutgers University Press, 1986. Pp. xvi + 263. ISBN 0-8135-1118-6. $35.00 (no sterling price). British Journal for the History of Science, 20 (2):227–228, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026312.

Bowler:1987:BRJ

[Bow87b] Peter J. Bowler. Book review: T. J. Horder, J. A. Witkowski & C. C. Wylie (eds). A History of Embryology (The Eighth Symposium of the British Society for Developmental Biology). Cambridge: Cambridge University Press, 1986. Pp. xxiv + 475. ISBN 0-521-25953-3. £60.00, $99.50. British Journal for the History of Science, 20(1):125, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026476.

Bowler:1987:BRW

[Bow87c] Peter J. Bowler. Book review: William B. Provine. Sewall Wright and Evolutionary Biology. Chicago and London: University of Chicago Press, 1986. Pp. xvi + 545. ISBN 0-226-68474-1. £25.50. British Journal for the History of Science, 20(3):359–360, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026370.

Bowler:1988:BRL

[Bow88a] Peter J. Bowler. Book review: Leo F. Laporte, (ed.). Simple Curiosity: Letters from George Gaylord Simpson to his Family, 1921–1970. Berkeley and Los Angeles: University of California Press, 1987. Pp. + 340. ISBN 0-520-05792-9. $29.95. British Journal for the History of Science, 21(3): 366, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026837.

Bowler:1988:BRR

[Bow88b] Peter J. Bowler. Book review: Robert Ackerman, J. G. Frazer: his Life and Work. Cambridge: Cambridge University Press, 1987. Pp.x + 348. ISBN 0-521-34093-4. £35.50. British Journal for the History of Science, 21(3): 367, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026839.

Bowler:1988:NDR

[Bow88c] Peter J. Bowler. The non-Darwinian revolution: reinterpreting a historical myth. The Johns Hopkins University Press, Baltimore, MD, USA, 1988. ISBN 0-8018-3678-6. x + 238 pp. LCCN QH361 .B693

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        jhu051/88009738.html.

Bowers:1989:BRR

[Bow89a] Brian Bowers. Book review: Rowland F. Pocock, The Early British Radio Industry. Manchester and New York: Manchester University Press, 1988. Pp: viii + 184. ISBN 0-7190-2621-0. £27.50. British Journal for the History of Science, 22(2):252, July 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026670.

Bowler:1989:BRP

[Bow89b] Peter J. Bowler. Book review: Pietro Corsi. The age of Lamarck: evolutionary theories in France, 1790–1830. Translated by Jonathan Mandelbaum. Berkeley: University of California Press, 1988. Pp. xiii + 360. ISBN 0-520-05830-5. $42.00. British Journal for the History of Science, 22(4): 462–463, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026938.

Bowler:1989:BRR

[Bow89c] Peter J. Bowler. Book review:

Ronald Rainger, Keith R. Benson and Jean Maienschein, (eds). The American Development of Biology. Philadelphia: University of Pennsylvania Press, 1988. Pp. xii + 380. ISBN 0-8122-8092-X. British Journal for the History of Science, 22(3):391–392, September 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026908.

Bowler:1989:DAE

[Bow89d] Peter J. Bowler. Development and adaptation: Evolutionary concepts in British morphology, 1870–1914. British Journal for the History of Science, 22(3): 283–297, September 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026897.

Bowler:1989:IPV

[Bow89e] Peter J. Bowler. The invention of progress: the Victorians and the past. Blackwell, Oxford, UK, 1989. ISBN 0-631-16107-4. xi + 231 pp. LCCN GN360 .B67 1989. US$39.00 (U.S. : est.).

Bowler:1990:CDM

[Bow90] Peter J. Bowler. Charles Darwin: the man and his influence. Blackwell scientific biographies. Blackwell, Oxford, UK, 1990. ISBN 0-631-16818-4. xii + 250 pp. LCCN QH31.D2 B743 1991. US$34.00.

Bowler:1992:FHE

[Bow92] Peter J. (Peter John) Bowler. The Fontana history of the environmental sciences. Fontana history of science. Fontana Press, London, UK, 1992. ISBN 0-00-686184-9. xv + 634 pp. LCCN GE50 .B68 1992.

Bowler:1993:D

[Bow93] Peter J. Bowler. Darwinism, volume 6 of Twayne’s studies in intellectual and cultural history. Twayne Publishers, New York, NY, USA, 1993. ISBN 0-8057-8613-9, 0-8057-8638-4 (paperback). xii + 118 pp. LCCN QH366.2 .B675 1993. US$24.95, US$14.95.

Boyer:1970:BRM

[Boy70] Carl B. Boyer. Book review: Mathematics: The Origins of the Infinitesimal Calculus. By Margaret E. Baron. Oxford: Pergamon Press. 1969. Pp. viii + 304. £5. British Journal for the History of Science, 5 (1):89–91, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025358.

Brewer:1993:CWG

[BP93a] John Brewer and Roy Porter, editors. Consumption and the world of goods. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1993. ISBN 0-415-03712-3. xix + 564 pp. LCCN HC79.C6 C673 1993.

Bynum:1993:CEH

[BP93b] W. F. (William F.) Bynum and Roy Porter, editors. Companion encyclopedia of the history of medicine. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1993. ISBN 0-415-04771-4 (set), 0-415-09242-6 (vol. 1), 0-415-09243-4 (vol. 2). ???? pp. LCCN R133 .E5

1993. URL http://www.loc.
gov/catdir/enhancements/
fy0648/92026013-d.html.

Berrios:1995:HCP

[BP95] German E. Berrios and Roy Porter, editors. A history of clinical psychiatry: the origin and history of psychiatric disorders. Athlone Press, London, UK, 1995. ISBN 0-485-24011-4 (hardcover), 0-485-24211-7 (paperback). 684 pp. LCCN RC455.2.C4 H57 1999.

Bynum:2002:WHE

[BP02] W. F. (William F.) Bynum and Roy Porter, editors. William Hunter and the Eighteenth-Century medical world. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-52517-9, 0-521-26806-0. xi + 424 pp. LCCN R484 .W55 2002x.

Bynum:1985:WHE

[BPH85] W. F. (William Frederick) Bynum, Roy Porter, and William Hunter, editors. William Hunter and the eighteenth-century medical world. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-26806-0. LCCN R484 .W55

1985. URL http://www.loc.
gov/catdir/description/
cam022/84021378.html;
http://www.loc.gov/
catdir/samples/cam031/
84021378.html;   http://
www.loc.gov/catdir/toc/
cam031/84021378.html.

Blaxter:1995:DPM

[BR95] Sir K. L. (Kenneth Lyon) Blaxter and Noel F. Robertson. From dearth to plenty: the modern revolution in food production. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-40322-7 (hardcover). xvii + 296 pp. LCCN S494.5.I5 B55

1995. URL http://www.loc.
gov/catdir/description/
cam026/94025259.html;
http://www.loc.gov/
catdir/toc/cam026/
94025259.html.

Brackenridge:1984:BRH

[Bra84] J. Bruce Brackenridge. Book review: History of Mathematical Sciences: Ronald Cowing, Roger Cotes-Natural Philosopher, Cambridge: Cambridge University Press, 1983. Pp. x + 210. £22.50. ISBN 0-521-23741-6. British Journal for the History of Science, 17(2): 231–232, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026563.

Brackenridge:1987:BRP

[Bra87] J. Bruce Brackenridge. Book review: Paola Zambelli (ed.), ‘Astrologi hallucinati’ Stars and the End of the World in Luther’s Time. Berlin & New York: Walter de Gruyter, 1986. Pp. x + 293. ISBN 3-11-010317-6. DM 128.00. British Journal for the History of Science, 20(4):479–481, October 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026421.

Brackenridce:1988:BRB

[Bra88] J. Bruce Brackenridce. Book review: Bruce Stephenson. Kepler’s Physical Astronomy. Berlin: Springer-Verlag, 1987. Pp. v + 216. ISBN 3-540-96541-6. DM 118.00. British Journal for the History of Science, 21 (3):372–374, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026845.

Brackenridge:1989:BRJ

[Bra89a] J. Bruce Brackenridge. Book review: John Fauvel, Raymond Flood, Michael Shortland, & Robin Wilson (eds). Let Newton Be!. Oxford: Oxford University Press, 1988. ISBN 0-19-853924-X. £17.50. British Journal for the History of Science, 22(4):475– 477, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026946.

Bramwell:1989:ETC

[Bra89b] Anna Bramwell. Ecology in the Twentieth Century: a history. Yale University Press, New Haven, CT, USA, 1989. ISBN 0-300-04343-0. xiii + 292 pp. LCCN QH541 .B7 1989.

Brock:1967:BRH

[Bro67] W. H. Brock. Book review: History of Chemistry Chymia. Volume 11. Edited by Henry M. Leicester. Pp. 208. University of Pennsylvania Press and Oxford University Press, 1966. 40s. British Journal for the History of Science, 3(4):404– 405, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024974.

Brock:1969:BRH

[Bro69]  W. H. Brock.  Book re-

view: History of physics. The Atomists (1805–1933). By Sir Basil Schonland. Pp. viii + 198. London: Oxford University Press: Clarendon Press. 1968. 35s. British Journal for the History of Science, 4 (3):291–292, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025482.

Brock:1971:BRP

[Bro71a] W. H. Brock. Book review:

Physical Sciences Nineteenth-Century Spectroscopy. Development of the Understanding of Spectra, 1803–1897. By William McGucken. Baltimore and London: Johns Hopkins Press. 1969. Pp. xii + 233. 9 figs. £5.25. British Journal for the History of Science, 5 (3):299–300, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025327.

Brooke:1971:OSU

[Bro71b] John Hedley Brooke. Organic synthesis and the unification of chemistry — A reappraisal. British Journal for the History of Science, 5(4): 363–392, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025380.

Bromberg:1972:BRT

[Bro72a] Joan Bromberg. Book review: Twentieth Century The Born–Einstein Letters. Correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955 with commentaries by Max Born. Trans. by Irene Born. Foreword by Bertrand Russell. Introduction by Werner Heisenberg. London: Macmillan, 1971. Pp. xi + 240. £3.85. British Journal for the History of Science, 6(2):222–223, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025309.

Brooke:1972:BRC

[Bro72b] John Hedley Brooke. Book review: Chemistry Studies in the History of Chemistry. By Sir Harold Hartley. Oxford: Clarendon Press: Oxford University Press, 1971. Pp. viii + 243. £2.75. British Journal for the History of Science, 6(1):90–91, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025271.

Brock:1973:BRG

[Bro73a] W. H. Brock. Book review: General Isis Cumulative Bibliography. A Bibliography of the History of Science formed from ISIS Critical Bibliographies 1-90, 1913–1965. Ed. by Magda Whitrow. 2 volumes. London: Mansell, in conjunction with the History of Science Society, 1971. Pp. lxviii + 664 and 789. £33. British Journal for the History of Science, 6(4): 433–434, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025506.

Brooke:1973:BRE

[Bro73b] John Hedley Brooke. Book review: Eighteenth and Nineteenth Centuries: Affinity and Matter. Elements of Chemical Philosophy 1800–1865. By Trevor H. Levere. Oxford: Clarendon Press: Oxford University Press, 1971. Pp. xvii + 230. £4.50. British Journal for the History of Science, 6(3):329–330, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025462.

Brooke:1975:BRN

[Bro75] J. H. Brooke. Book review:

Nineteenth Century: Between Science and Religion. The Reaction to Scientific Naturalism in Late Victorian England. By Frank Miller Turner. New Haven and London: Yale University Press, 1974. Pp. x + 273. British Journal for the History of Science, 8(3):262– 264, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025572.

Bromberg:1976:BRN

[Bro76] Joan Bromberg. Book review:

Nineteenth and Twentieth Centuries: The Compton Effect. Turning Point in Physics. By Roger H. Stuewer. New York: Science History Publications, 1975. Pp. xii + 367. No price stated. British Journal for the History of Science, 9(3): 335–336, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025697.

Brooke:1977:ROW

[Bro77] John Hedley Brooke. Richard Owen, William Whewell, and the Vestiges. British Journal for the History of Science, 10 (2):132–145, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025866.

Brown:1978:BRSa

[Bro78a] G. Burniston Brown. Book review: Seventeenth and Eighteenth Centuries: The Correspondence of Isaac Newton. Vol. v: 1709–1713. Pp. li + 439. Edited by A. Rupert Hall and Laura Tilling. Cambridge: Cambridge University Press for the Royal Society, 1975. £20.00. British Journal for the History of Science, 11 (2):182–183, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025741.

Brown:1978:BRSb

[Bro78b] G. Burniston Brown. Book review: Seventeenth and Eighteenth Centuries: The Correspondence of Isaac Newton. Volume vi: 1713–1718. Edited by A. Rupert Hall and Laura Tilling. Cambridge: Cambridge University Press for the Royal Society of London, 1976. Pp. xxxviii + 499 + v plates. £25.00. British Journal for the History of Science, 11(3): 292, November 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025790.

Brooke:1979:NCE

[Bro79] J. H. Brooke. Nebular contraction and the expansion of naturalism. British Journal for the History of Science, 12 (2):200–211, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025754.

Broberg:1980:LPP

[Bro80a] Gunnar Broberg, editor. Linnaeus: progress and prospects in Linnean research, volume 1978 of Svenska Linnésällskapets årsskrift. Almqvist and Wiksell International, Stockholm, Sweden, 1980. ISBN 0-913196-31-2, 91-22-00385-1. LCCN QH44 .L53

1980. URL http://books.
google.com/books?id=
h2gLAQAAIAAJ.

Brockliss:1980:BRR

[Bro80b] Laurence Brockliss. Book review: Renaissance: From Humanism to Science, 1480–1700. By Robert Mandrou. Hassocks, Sussex: Harvester Press, 1979. Pp. 329. £11.50. British Journal for the History of Science, 13(3):267–268, November 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026205.

Brock:1981:JCE

[Bro81] W. H. Brock. The Japanese connexion: Engineering in Tokyo, London, and Glasgow at the end of the Nineteenth Century (Presidential address, 1980). British Journal for the History of Science, 14(3):227– 244, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025981.

Brock:1982:BRT

[Bro82a] W. H. Brock. Book review: Technology and Development Ian Inkster, Japan as a development model? relative backwardness and technological transfer. Berliner Beiträge zur social- und wirtschaftswissenschaftlichen Japan-Forschung: Band 7 Bochum: Studienverlag Dr N. Brockmeyer, 1980. Pp. ii + 101. D.M. 14-80. British Journal for the History of Science, 15 (1):88–90, March 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025936.

Brooke:1982:BRD

[Bro82b] J. H. Brooke. Book review:

Darwiniana D. R. Oldroyd, Darwinian impacts: an introduction to the Darwinian revolution. Milton Keynes: The Open University Press, 1980. Pp. xiv + 398. £7.95. British Journal for the History of Science, 15(1):72–74, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025925.

Browne:1983:BRLb

[Bro83a]  Janet Browne.  Book re-

view: Life and earth sciences: Wilma George, Darwin. Glasgow: Fontana Paperbacks, 1982. Pp. 160. £1.75. Jonathan Howard, Darwin. Oxford: Oxford University Press Paperbacks, 1982. Pp. x + 102. £1.25. British Journal for the History of Science, 16(3):284– 285, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026154.

Browne:1983:BRLa

[Bro83b]  Janet Browne.  Book re-

view: Life Sciences: Charles Lebrun, A method to learn to design the passions (1734). trans, by J. Williams. Facsimile edn. with intro. by A. T. McKenzie, Los Angeles: University of California, William Andrews Clark Memorial Library, 1980. Pp. xiiii + 98. (No price stated.). British Journal for the History of Science, 16 (2):219–220, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026259.

Brock:1984:BRS

[Bro84a] W. H. Brock. Book review: Science in Culture: George Weisz, The Emergence of Modern Universities in France, 1863–1914. Princeton: Princeton University Press, 1983. Pp. xii + 397. £30.20. ISBN 0-691-05375-8. British Journal for the History of Science, 17(2):249– 250, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026583.

Brooks:1984:JBO

[Bro84b] John Langdon Brooks. Just before the origin: Alfred Russel Wallace’s Theory of Evolution. Columbia University Press, New York, NY, USA, 1984. ISBN 0-231-05676-1. xiii + 284 + 1 pp. LCCN QH31.W2 B76 1984. US$25.00.

Browne:1984:BRA

[Bro84c] Janet Browne. Book review: Antiquity to the Renaissance Janis L. Pallister (ed), Ambroise Paré On Monsters and Marvels: translated with an introduction and notes. Chicago and London: University of Chicago Press, 1982. Pp. xxxii + 224. ISBN 0-226-64562-2. £14.00. British Journal for the History of Science, 17(1): 107, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026043.

Browne:1984:BRH

[Bro84d] Janet Browne. Book review:

History of Natural History: Gunnar Broberg (ed), Linnaeus: progress and prospects in Linnean research. Stockholm and Pittsburgh: Almquist & Wiksell with the Hunt Institute for Botanical Documentation, 1978. Pp 318. ISBN 91-22-00385-1. (No price stated.). British Journal for the History of Science, 17(3):322, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026636.

Brock:1985:BRA

[Bro85a] W. H. Brock. Book review:

Alan J. Rocke, Chemical Atomism in the Nineteenth Century: From Dalton to Cannizzaro. Columbus: Ohio State University Press, 1984. Pp. xviii + 386. ISBN 0-8142-0360-4. $27.50. British Journal for the History of Science, 18(3): 345–347, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026387.

Brock:1985:OBJ

[Bro85b]  W. H. Brock.  Obituary:

Bernard John Norton: 1945– 1984. British Journal for the History of Science, 18(3):342– 344, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026386.

Brock:1985:PPW

[Bro85c] W. H. (William Hodson) Brock. From protyle to proton: William Prout and the nature of matter, 1785–1985. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-801-9. xii + 252 pp. LCCN QD22.P85 B76 1985.

Browne:1985:BRJ

[Bro85d] Janet Browne. Book review: John Langdon Brooks, Just before the Origin: Alfred Russel Wallace’s Theory of Evolution. New York: Columbia University Press, 1984. Pp. xiii + 284. ISBN 0-231-05676-1. U.S. $39. British Journal for the History of Science, 18(3):364– 365, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026402.

Browne:1985:BRP

[Bro85e]  Janet Browne.  Book re-

view: Philip F. Rehbock, The Philosophical Naturalists: Themes in Early Nineteenth-Century British Biology. (Wisconsin Publications in the History of Science and Medicine, 3). Madison, Wisconsin: University of Wisconsin Press, 1983. Pp xv + 281. ISBN 0-299-09430-8. $30. British Journal for the History of Science, 18(2):251, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026353.

Brock:1986:BRC

[Bro86a] W. H. Brock. Book review: C.

A. Russell, ed. Recent Developments in the History of Chemistry. London: The Royal Society of Chemistry, 1985. Pp. x + 333. ISBN 0-85186-917-3. £27.50, $36.00. British Journal for the History of Science, 19(3):358–359, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026528.

Brock:1986:BRM

[Bro86b] W. H. Brock. Book review:

Mario Morselli. Amedeo Avogadro. Dordrecht, Boston, Lancaster: D. Reidel Publishing Company, 1984. ISBN 90-277-1624-2. Pp. xi + 375. Dfl, 160; $59.50. British Journal for the History of Science, 19 (2):205–206, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026595.

Brown:1986:BRB

[Bro86c]  Laurie M. Brown.  Book

review: Badash Lawrence. Kapitza, Rutherford, and the Kremlin. New Haven: Yale University Press, 1985. Pp. xi + 129. ISBN 0-300-01465-1. £20.00. British Journal for the History of Science, 19(3): 372, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026544.

Browne:1986:BRP

[Bro86d] Janet Browne. Book review: Peter J. Bowler. Evolution: the History of an Idea. Berkeley, Los Angeles and London: University of California Press, 1984. Pp. xiv + 412. $29.95. British Journal for the History of Science, 19(3): 345–346, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026516.

Browne:1986:BRW

[Bro86e]  Janet  Browne.

Book review: W. J. Dempster. Patrick Matthew and Natural Selection: Nineteenth Century Gentleman-Farmer, Naturalist and Writer. Edinburgh: Paul Harris, 1983. Pp 156. ISBN 0-86228-065-6. £8.50. British Journal for the History of Science, 19 (1):118–119, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026491.

Browne:1987:BRD

[Bro87]  Janet Browne.  Book re-

view: David Elliston Allen. The Botanists: A History of the Botanical Society of the British Isles Through a Hundred and Fifty Years. Winchester: St Paul’s Bibliographies, 1986. Pp. xv + 232. ISBN 0-906795-36-2. £15.00 (post free in U.K. if accompanied by cheque; or plus £1.50 postage if from abroad). British Journal for the History of Science, 20(2):230–231, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026316.

Brock:1989:BRE

[Bro89a] W. H. Brock. Book review:

Emil Heuser & Regine Zott (eds). Justus von Liebig und August Wilhelm Hofmann in ihren Briefen. Nachträge 1845– 1869. Emil Heuser (ed.). Justus von Liebig und Emil Erlenmeyer in ihren Briefen von 1861–1872. Mannheim: Bionomica Verlag, 1988. Pp. 54 + 32 bound in one vol. ISBN 3-88208-012-4. DM 20.00. British Journal for the History of Science, 22(1):99, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026698.

Brock:1989:BRF

[Bro89b] W. H. Brock. Book review: N. A. Figurovsku & Yu. I. Solovev. Aleksandr Porfir’evich Borodin. A Chemist’s Biography. (Translated from the Russian by Charlene Steinberg & George B. Kauffman.) Berlin, New York and London: Springer-Verlag, 1988. Pp. xiv + 171, illus. ISBN. 3-540-17888-0, $79.50. British Journal for the History of Science, 22(4):449, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026927.

Brook:1989:BRP

[Bro89c] J. H. Brook. Book review: Pietro Corsi, Science and Religion: Baden Powell and the Anglican Debate, 1800–1860. Cambridge, New York, New Rochelle, Melbourne, Sydney: Cambridge University Press, 1988. Pp: ix + 346. ISBN 0-521-24245-2, £32.50, $54.50. British Journal for the History of Science, 22(4):439– 442, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026922.

Brown:1989:BSL

[Bro89d] Chandos Michael Brown. Benjamin Silliman: a life in the young republic. Princeton University Press, Princeton, NJ, USA, 1989. ISBN 0-691-08533-1. xvi + 377 pp. LCCN Q143.S56 B76 1989. US$29.50.

URL http://www.loc.gov/
catdir/description/
prin031/88032086.html.

Brown:1989:BRW

[Bro89e] Laurie M. Brown. Book review:

W. Peter Trower (ed.). Discovering Alvarez: Selected Works of Luis W. Alvarez, with Commentary by His Students and Colleagues. Chicago: University of Chicago Press, 1987. Pp. ix. + 272. ISBN 0-226-81304-5. £29.95, $44.95. British Journal for the History of Science, 22 (3):383–384, September 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026904.

Browne:1989:BRR

[Bro89f] Janet Browne. Book review:

Rachel Laudan. From Mineralogy to Geology: the Foundations of a Science, 1650–1830. Chicago and London: University of Chicago Press, 1987. Pp. xii + 278. ISBN 0-226-46950-6. £21.95; $32.95 in Europe. British Journal for the History of Science, 22(1):105– 106, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026704.

Brock:1992:FHC

[Bro92] William H. (William Hodson) Brock, editor. The Fontana history of chemistry. Fontana history of science. Fontana Press, London, UK, 1992. ISBN 0-00-686173-3. xxiii + 744 pp. LCCN QD11 .B75 1992.

Brooke:1995:TAM

[Bro95a] John Hedley Brooke. Thinking about matter: studies in the history of chemical philosophy, volume CS502 of Collected studies series. Variorum, Aldershot, Hampshire, UK, 1995. ISBN 0-86078-464-9. ???? pp. LCCN QD6 .B76 1995.

Browne:1995:CD

[Bro95b] E. J. (E. Janet) Browne. Charles Darwin. Alfred A. Knopf, New York, NY, USA, 1995. ISBN 0-224-04202-5 (vol. 1), 0-224-04212-2 (vol. 2). xiii + 605 + 32 pp. LCCN ????

Browne:1995:VCJ

[Bro95c] Janet Browne. Voyaging, volume 1 of Charles Darwin. Cape, London, UK, 1995. ISBN 0-224-04202-5. 605 pp. LCCN ????

Brock:1997:JLC

[Bro97] W. H. (William Hodson) Brock. Justus von Liebig: the chemical gatekeeper. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56224-4 (hardcover). xiv + 374 pp. LCCN QD22.L7 B76

1997. URL http://www.loc.
gov/catdir/description/
cam027/96025468.html;
http://www.loc.gov/
catdir/samples/cam031/
96025468.html;   http://
www.loc.gov/catdir/toc/
cam031/96025468.html.

Brush:1970:WTH

[Bru70] Stephen G. Brush. The wave theory of heat: A forgotten stage in the transition from the caloric theory to thermodynamics. British Journal for the History of Science, 5(2): 145–167, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Brush:1972:BRE

[Bru72] Stephen G. Brush. Book review: Eighteenth and Nineteenth Centuries: The Caloric Theory of Gases from Lavoisier to Regnault. By Robert Fox. Oxford: Clarendon Press: Oxford University Press, 1971. Pp. xv + 378. £5. British Journal for the History of Science, 6(2):218–220, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025306.

Brush:1986:BRI

[Bru86] Stephen G. Brush. Book review: Ilya Prigogine & Isabelle Stengers. Order Out of Chaos: Man’s New Dialogue with Nature. London: Heinemann, 1984. Pp. xxxi + 349. ISBN 0-434-603953-3. £9.95. British Journal for the History of Science, 19(3):371– 372, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026543.

Bruce:1987:LMA

[Bru87a] Robert V. Bruce. The launching of modern American science, 1846–1876. The Impact of the Civil War. Knopf, New York, NY, USA, 1987. ISBN 0-394-55394-2. x + 446 + 4 pp. LCCN Q127.U6 B78 1987.

Brundell:1987:PGA

[Bru87b]  Barry  Brun-

dell. Pierre Gassendi: from Aristotelianism to a new natural philosophy, volume 30 of Synthese historical library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1987. ISBN 90-277-2428-8. x + 251 pp. LCCN B1887 .B84 1987.

Brundage:1988:EPM

[Bru88a] Anthony Brundage. England’s “Prussian minister”: Edwin Chadwick and the politics of government growth, 1832–1854. Pennsylvania State University Press, University Park, PA, USA, 1988. ISBN 0-271-00629-3. 208 pp. LCCN DA565.C38 B78 1988. US$24.95.

Brush:1988:HMS

[Bru88b] Stephen G. Brush. The history of modern science: a guide to the second scientific revolution, 1800–1950. The Iowa State University Press series in the history of technology and science. Iowa State University Press, Ames, IA, USA, 1988. ISBN 0-8138-0883-9. xv + 544 pp. LCCN Q125 .B88 1988.

Brush:1988:HPH

[Bru88c] Stephen G. Brush, editor.

History of physics: selected reprints. American Association of Physics Teachers, College Park, MD, USA, 1988. ISBN 0-917853-29-6. 235 pp. LCCN QC7.5 .H57 1988.

Brunner:1995:FPP

[Bru95] José Brunner. Freud: the politics of psychoanalysis. Blackwell, Oxford, UK, 1. publ edition, 1995. ISBN 0-631-16404-9. xiv + 238 pp. LCCN BF109.F74 B78 1995.

Bryden:1971:AFH

[Bry71]  D. J. Bryden.  An addi-

tional factor in the history of the centigrade thermometer. British Journal for the History of Science, 5(4):393– 396, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025381.

Bryden:1974:BRS

[Bry74] D. J. Bryden. Book review: Scientific Instruments Scientific Instruments of the Seventeenth and Eighteenth Centuries and their Makers. By Maurice Daumas. Trans, and ed. by Mary Holbrook. London: Batsford, 1972. Pp. vi + 361. £10. British Journal for the History of Science, 7(1):87–88, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025184.

Bryden:1976:BRI

[Bry76] D. J. Bryden. Book review: Instruments: Van Marum’s Scientific Instruments in Teyler’s Museum. By G. L’E. Turner and T. H. Levere. Volume IV of Martinus Van Marum: Life and Work, Ed. by E. Lefebvre and J. G. De Bruijn. Leyden: Noordhoff Intertional, 1973. Pp.

401. 65 Hfl. British Journal for the History of Science, 9 (1):69–70, March 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025708.

Bryden:1977:BRS

[Bry77] D. J. Bryden. Book review: Scientific Instruments John and Jonathan Cuthbertson. The Invention and Development of the Eighteenth Century Plate Electrical Machine. By W. D. Hackmann. Communication No. 142 from the Rijksmuseum voor de Geschidenis der Natuurwetenschappen. Leyden, 1973. Pp. 72, including 16 half-tone plates. No price stated. British Journal for the History of Science, 10(1):77, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025590.

Bryden:1978:BRA

[Bry78a] D. J. Bryden. Book review: Astronomy Greenwich Observatory. The Royal Observatory at Greenwich and Herstmonceux, 1675–1975. London: Taylor & Francis, 1975. £25.00. (Volume iii only is available separately at £12.0.) Volume i: Origins and Early History (1675–1835). By Eric G. Forbes. Pp. xv + 204 + 8 plates. London: Taylor & Francis, 1975. £25.00. (Volume iii only is available separately at £12.0.) Volume ii: Recent History (1836–1975). By A. J. Meadows. Pp. xi + 135 + 14 plates. London: Taylor & Francis, 1975. £25.00. (Volume iii only is available separately at £12.0.) Volume iii: The Buildings and Instruments. By Derek Howse. Pp. xix + 178 + 130 plates. London: Taylor & Francis, 1975. £25.00. (Volume iii only is available separately at £12.0.). British Journal for the History of Science, 11(2):173–174, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.

                                        jstor.org/stable/4025732.

Bryden:1978:BRN

[Bry78b] D. J. Bryden. Book review: Nineteenth and Twentieth Centuries: Martinus Van Marum, Life and Work. Ed. by E. Lefebvre and J. G. De Bruijn. Leyden: Noordhoff International Publishing, 1976. Pp. xi + 435 + 23 plates. Dfl. 60. British Journal for the History of Science, 11(1):82–83, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025623.

Bryden:1978:BRS

[Bry78c] D. J. Bryden. Book review: Scientific Instruments Benjamin Martin: Author, Instrument-Maker, and ‘Country Showman’. By John R. Millburn. Leyden: Noordhoff International Publishing, 1976. Pp. xii + 244. Dfl. 63. British Journal for the History of Science, 11(3):284–285, November 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025785.

Bryden:1981:SNB

[Bry81] D. J. Bryden. Short notices of books: Weighing coins: English folding gold balances of the 18th and 19th Centuries. By Michael A. Crawforth, London: Cape Horn Trading Co., 1979. Pp x + 194. £15.00. British Journal for the History of Science, 14(1):104, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026088.

Bryden:1983:BRL

[Bry83] D. J. Bryden. Book review:

Lives and Works Norman H. Robinson, The Royal Society catalogue of portraits. With biographical notes by Eric G. Forbes. London: Royal Society, 1980. Pp. 343. £25.00. British Journal for the History of Science, 16(2):198–199, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026242.

Bryder:1988:BMM

[Bry88] Linda Bryder. Below the magic mountain: a social history of tuberculosis in Twentieth-Century Britain. Oxford historical monographs. Clarendon Press, New York, NY, USA, 1988. ISBN 0-19-822947-X. xiv + 298 pp. LCCN RC316.G7 B78 1988. US$27.00 (US$59.00

U.S.). URL http://www.loc.
gov/catdir/enhancements/
fy0606/87031495-d.html.

Bryden:1992:NBH

[Bry92] D. J. Bryden, editor. Napier’s bones: a history and instruction manual. Harriet Wynter, London, UK, 1992. ISBN 0-9507258-2-X. 23 pp. LCCN QA71 .B79 1992.

Barnes:1977:WEO

[BS77]  Barry  Barnes

and Steven Shapin. Where is the edge of objectivity? British Journal for the History of Science, 10(1):61–66, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025581.

Bunge:1979:RPT

[BS79] Mario Bunge and William R. Shea, editors. Rutherford and physics at the turn of the century. Dawson, Kent, UK, 1979. ISBN 0-88202-184-2 (New York), 0-7129-0918-4 (Cannon House Folkestone). 184 pp. LCCN QC7.5 .R87 1979.

Bowers:1991:CPS

[BS91] Brian Bowers and Lenore Symons, editors. Curiosity perfectly satisfyed: Faraday’s travels in Europe, 1813–1815, volume 16 of IEE history of technology series. P. Peregrinus in association with the Science Museum, London, UK, 1991. ISBN 0-86341-234-3. xvi + 168 pp. LCCN QC16.F2 C87 1991.

Burkhardt:1985:CCD

[BS10] Frederick Burkhardt and Sydney Smith, editors. The correspondence of Charles Darwin. Cambridge University Press, Cambridge, UK, 1985– 2010. ISBN 0-521-25587-2 (vol.

1), 0-521-25588-0 (vol. 2), 0-521-25589-9 (vol. 3), 0-521-25590-2 (vol. 4), 0-521-25591-0 (vol. 5), 0-521-25586-4 (vol. 6), 0-521-84459-2 (vol. 14), 0-521-85931-X (vol. 15 : hardback), 0-521-88195-1 (vol. 16, pt. 1 : hardback), 0-521-88196-X (vol. 16, pt. 2 : hardback), 0-521-51836-9 (vol. 16, 2 vol. hardback set), 0-521-19030-4 (vol. 17), 0-521-76889-6 (vol. 18), 0-521-44241-9. ???? pp. LCCN QH31.D2 A4 1985.

Burkhardt:1985:CCC

[BSK^{+}85] Frederick Burkhardt, Sydney Smith, David Kohn, William Montgomery, and Charles Darwin, editors. A calendar of the correspondence of Charles Darwin, 1821–1882, volume 369 of Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1985. ISBN 0-8240-9224-4. 690 + 1 pp. LCCN QH31.D2 C15 1985.

Burkhardt:1994:CCC

[BSK^{+}94] Frederick Burkhardt, Sydney Smith, David Kohn, William Montgomery, Charles Darwin, et al., editors. A calendar of the correspondence of Charles Darwin, 1821–1882, with supplement. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-43423-8 (hardcover). viii 690, 46 pp. LCCN QH31.D2 C15

1994. URL http://www.loc.
gov/catdir/description/
cam026/94191109.html;

                                        http://www.loc.gov/
                                        catdir/toc/cam029/
                                        94191109.html.

Batchelor:1996:LLG

[BT96] G. K. (George Keith) Batchelor and Sir Geoffrey Ingram Taylor. The life and legacy of G. I. Taylor. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-46121-9. xv + 285 pp. LCCN Q143.T29

                                        B38 1996. URL http://www.
                                        loc.gov/catdir/
                                        description/cam027/
                                        95047579.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam027/95047579.html.

Butler:1963:RC

[BU63] F. H. C. Butler and E. Ashworth Underwood. Report of Council. British Journal for the History of Science, 1(4): 392–393, December 1963. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024918.

Buchdahl:1963:RDP

[Buc63] Gerd Buchdahl. The relevance of Descartes’s philosophy for modern philosophy of science. British Journal for the History of Science, 1(3): 227–249, June 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024922.

Buchdahl:1966:BRS

[Buc66] Gerd Buchdahl. Book review: Sixteenth and Seventeenth Centuries: Robert Boyle on Natural Philosophy. An Essay with Selections from His Writings. By Marie Boas Hall. Pp. ix + 406. Bloomington: Indiana University Press, 1965. $6.75. British Journal for the History of Science, 3(1): 82–84, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025112.

Buchdahl:1968:BRH

[Buc68] Gerd Buchdahl. Book review: History and Philosophy of Science A Preliminary Discourse on the Study of Natural Philosophy. By John Frederick William Herschel. A Facsimile of the 1830 Edition. With a new Introduction by Michael Partridge. The Sources of Science, No. 17. New York and London, Johnson Reprint Corporation. 1966. $14.50. British Journal for the History of Science, 4(2): 173–175, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025147.

Buchanan:1969:BRE

[Buc69] R. A. Buchanan. Book review: Engineering: its Role and Function in Human Society. Edited by William H. Davenport and Daniel Rosenthal.

New York: Pergamon Press. 1967. Pp. xii + 284. 63s. British Journal for the History of Science, 4(4):415, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025247.

Buchanan:1971:BRI

[Buc71] R. A. Buchanan. Book review: Industrial Revolution Power in the Industrial Revolution. By Richard L. Hills. Manchester: Manchester University Press. 1970. Pp. ix + 274. £4. British Journal for the History of Science, 5(3):297–298, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025325.

Buchanan:1973:BRG

[Buc73a] R. A. Buchanan. Book review: General Technology, Science and History. By D. S. L. Cardwell. London: Heinemann, 1972. Pp. xi + 244. £3.00. British Journal for the History of Science, 6(3): 314–315, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025448.

Buck:1973:BRT

[Buc73b] Barbara Reeves Buck. Book review: Twentieth Century Enrico Fermi Physicist. By Emilio Segrè. Chicago and London: University of Chicago Press, 1970. Pp. xii + 276. £3.15. British Journal for the History of Science, 6(3): 331–332, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025464.

Buchanan:1975:BRT

[Buc75] R. A. Buchanan. Book review: Technology Understanding Technology. By Charles Susskind. Baltimore and London: The Johns Hopkins University Press. 1973. Pp. x + 164. £3.30. British Journal for the History of Science, 8 (2):179–180, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025648.

Bucciarelli:1977:BRN

[Buc77] L. Bucciarelli. Book review:

Nineteenth and Twentieth Centuries: Joseph Fourier: The Man and the Physicist. By John Herivel. Oxford: Clarendon Press: Oxford University Press, 1975. Pp. xii + 35O. British Journal for the History of Science, 10(3):270– 271, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025857.

Buchanan:1985:BRH

[Buc85a] R. A. Buchanan. Book review: History and Technology:

An International Journal. Volume 1, number 1. Chur, London etc: Harwood Academic Publishers, 1983. Pp. 112. ISSN 0734-1512. $58 per volume of 4 numbers. British Journal for the History of Science, 18 (2):247–248, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026349.

Buchanan:1985:BRJ

[Buc85b] R. A. Buchanan. Book review: J. F. Clarke and F. Storr, The Introduction of the use of mild steel into the Shipbuilding and Marine Engine industries. Newcastle upon Tyne Polytechnic, ‘Occasional Papers in the History of Science and Technology no. 1’, 1983, Pp. 97. £3.25. British Journal for the History of Science, 18(2):248, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026350.

Buchanan:1985:RSE

[Buc85c] R. A. Buchanan. The rise of scientific engineering in Britain. British Journal for the History of Science, 18(2):218– 233, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026332.

Buchwald:1985:MMA

[Buc85d]  Jed Z. Buchwald.  From

Maxwell to microphysics: aspects of electromagnetic theory in the last quarter of the Nineteenth Century. University of Chicago Press, Chicago, IL, USA, 1985. ISBN 0-226-07882-5 (hardcover), 0-226-07883-3 (paperback). xiii + 339 pp. LCCN QC670 .B83 1985. US$55.00.

Buchwald:1989:BRD

[Buc89a] Jed Z. Buchwald. Book review:

David B. Wilson. Kelvin and Stokes. A Comparative Study in Victorian Physics. Bristol: Adam Hilger, 1987. Pp. xvi + 253. ISBN 0-85274-526-5. £35.00. British Journal for the History of Science, 22(3): 384–388, September 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026905.

Buchwald:1989:RWT

[Buc89b] Jed Z. Buchwald. The rise of the wave theory of light: optical theory and experiment in the early Nineteenth Century. University of Chicago Press, Chicago, IL, USA, 1989. ISBN 0-226-07886-8 (paperback), 0-226-07884-1 (hardcover). xxiv + 474 pp. LCCN QC403 .B83

1989. URL http://www.loc.
gov/catdir/description/
uchi052/88018647.html;
http://www.loc.gov/
catdir/enhancements/

fy0608/88018647-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/88018647-b.html.

Buchwald:1994:CSE

[Buc94] Jed Z. Buchwald. The creation of scientific effects: Heinrich Hertz and electric waves. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-07887-6, 0-226-07888-4 (paperback). xiv + 482 pp. LCCN QC661 .B85 1994. URL http://www.

loc.gov/catdir/
description/uchi052/
93041783.html;   http://
www.loc.gov/catdir/toc/
uchi051/93041783.html.

Bud:1984:BRN

[Bud84] Robert Bud. Book reviews: National Traditions in Science Nathan Reingold and Ida H. Reingold (eds) Science in America. A documentary history, 1900–1939. Chicago and London University of Chicago Press, 1981. Pp. x + 490. ISBN 0-226-70946-9. £26.25. Marc Rothenberg, The history of science in the United States: a critical and selective bibliography. New York and London: Garland Publishing, Inc., 1982. Pp xiii + 242. ISBN 0-8240-9278-3. $35.00. British Journal for the History of Science, 17(1):91–92, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026027.

Bud:1989:SW

[Bud89] Robert Bud. The seamless web. British Journal for the History of Science, 22(1):75– 78, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026680.

Bud:1993:ULH

[Bud93] Robert Bud. The uses of life: a history of biotechnology. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-38240-8. xvii + 299 pp. LCCN TP248.18 .B83

1993. URL http://www.loc.
gov/catdir/description/
cam032/92019513.html;
http://www.loc.gov/
catdir/toc/cam032/
92019513.html.

Buffetaut:1987:SHV

[Buf87] Eric Buffetaut. A short history of vertebrate palaeontology. Croom Helm, London, UK, 1987. ISBN 0-7099-3962-0. 233 + 16 pp. LCCN QE841 .B84 1987. US$52.50.

Buhler:1981:GBS

[Büh81] W. K. (Walter Kaufmann)

Bühler. Gauss: a biographical study. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 3-540-10662-6, 0-387-10662-6. viii + 208 pp. LCCN QA29.G3

B83. URL http://catalog.

hathitrust.org/api/
volumes/oclc/7197827.html.

Burstall:1971:BRT

[Bur71] Aubrey Burstall. Book review: Technology Osborne Reynolds and Engineering Science Today. Ed. by D. M. McDowell & J. D. Jackson. Manchester: Manchester University Press. 1970. Pp. xii + 263. £3.60. British Journal for the History of Science, 5(4):411– 412, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025397.

Burkhardt:1974:BRE

[Bur74a] Richard W. Burkhardt. Book review: Eighteenth and Nineteenth Centuries: Inédits de Lamarck. D’après les Manuscrits Conservés à la Bibliothèque Centrale du Muséum National d’Histoire Naturelle de Paris. Ed. by Max Vachon, Georges Rousseau, and Yyves Laissus. Preface by Georges Canguilhem. Postface by Pierre-P. Grassé. Paris: Masson 1972. Pp. 312. 80 francs. British Journal for the History of Science, 7(2):192–194, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025218.

Burkhardt:1974:BRN

[Bur74b] Richard W. Burkhardt. Book review: Nineteenth and Twentieth Centuries: Colloque International ‘Lamarck’. Tenu au Muséum National d’Histoire Naturelle, Paris, les 1-2 et 3 juillet 1971. Ed. by J. Schiller. Paris: Blanchard, 1971. Pp.

263. 28 francs. British Journal for the History of Science, 7(1):93–94, March 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025190.

Burstyn:1975:IDW

[Bur75] Harold L. Burstyn. If Darwin wasn’t the Beagle’s naturalist, why was he on board? British Journal for the History of Science, 8(1):62–69, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025816.

Burchfield:1980:BRG

[Bur80] Joe D. Burchfield. Book review: Geology Images of the Earth: Essays in the History of the Environmental Sciences. Edited by L. J. Jordanova and Roy Porter. Chalfont St Giles: British Society for the History of Science, 1979. BSHS Monographs, 1. Pp. xx + 282. £5.95. British Journal for the History of Science, 13 (2):162–163, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025897.

Burke:1983:USA

[Bur83] John G. Burke, editor. The uses of science in the age of Newton, volume 8 of Publications from the Clark Library professorship, UCLA. University of California Press, Berkeley, CA, USA, 1983. ISBN 0-520-04970-5. 204 pp. LCCN Q127.G4 U83 1983.

Burton:1986:RFE

[Bur86] Jim Burton. Robert FitzRoy and the early history of the Meteorological Office. British Journal for the History of Science, 19(2):147–176, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026590.

Burke:1987:ADE

[Bur87a] Frank G. Burke. Automation and documentary editing. British Journal for the History of Science, 20(1):73– 80, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026438.

Burwick:1987:AOF

[Bur87b] Frederick Burwick, editor. Approaches to organic form: permutations in science and culture, volume 105 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1987. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346. xix + 286 pp. LCCN Q174 .B67 vol. 105. US$64.00(U.S.). URL https:

//link.springer.com/book/
10.1007/978-94-009-3917-2.

Buruma:1999:JSK

[Bur99] Y. Buruma, editor. 100 jaar strafrecht: klassieke teksten van de twintigste eeuw. (Dutch) [100 years of criminal law: traditional texts of the Twentieth Century ]. Amsterdam University Press, Amsterdam, The Netherlands, 1999. ISBN 90-5356-362-8, 90-5356-361-X (paperback). 440 pp. LCCN KKM3800 .A17 1999.

Anonymous:2000:CCD

[Bur00] Frederick Burkhardt. The correspondence of Charles Darwin, 1837–1843. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-25588-0. xxxiii + 603 pp. LCCN ????

Bush:1974:BRG

[Bus74] Rachel Bush. Book review: Geology Charles Lyell. The Years to 1841: The Revolution in Geology. By Leonard G. Wilson. New Haven and London: Yale University Press, 1972. Pp. xiv + 553. £7.50. British Journal for the History of Science, 7(1):87, March 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025183.

Busfield:1983:BRM

[Bus83] Joan Busfield. Book review: Medical Sciences: Celia Davies (ed.), Rewriting nursing history. London: Croom Helm. Totowa N.J.: Barnes & Noble Books, 1980. £11.95; £5.95 (paperback). British Journal for the History of Science, 16 (1):92–93, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026102.

Busfield:1985:BRR

[Bus85]  Joan Busfield.  Book re-

view: Roger Smith, Trial by medicine: insanity and responsibility in Victorian trials. Edinburgh: Edinburgh University Press, 1981. Pp. ix + 238. £15.00. British Journal for the History of Science, 18 (1):89–90, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026273.

Butler:1963:HHT

[But63]  F. H. C. Butler.  Hugh

Hamshaw Thomas. British Journal for the History of Science, 1(3):280–283, June 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024925.

Butler:1964:HES

[But64a] F. H. C. Butler. Henry Ernest Stapleton. British Journal for the History of Science, 2:64– 67, 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Butler:1964:OHE

[But64b]  F. H. C. Butler.  Obitu-

ary: Henry Ernest Stapleton 1878–1962. British Journal for the History of Science, 2 (1):64–67, June 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025085.

Butler:1965:OCE

[But65] F. H. C. Butler. Obituary:

Charles Earle Raven: D.D., Hon. D.SC., F.B.A. (1885– 1964). British Journal for the History of Science, 2(3): 254–256, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024940.

Butler:1966:JRP

[But66] F. H. C. Butler. James Riddick Partington. British Journal for the History of Science, 3(1):70–72, June 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025104.

Buttner:1983:HCC

[Büt83] J. (Johannes) Büttner, editor.

History of clinical chemistry. W. de Gruyter, Berlin, Germany, 1983. ISBN 3-11-008912-2. 91 pp. LCCN RB40 .H56 1983. DM98.00 (est.) (US$49.00 U.S. : est.).

Butler:1992:STM

[But92] Stella V. F. Butler. Science and technology museums. Leicester museum studies series. Leicester University Press, Leicester, UK, 1992. ISBN 0-7185-1357-6. xiii + 149 pp. LCCN T179 .B87 1992.

Butters:1996:TVS

[But96] Suzanne B. Butters. The triumph of Vulcan: sculptors’ tools, porphyry, and the prince in ducal Florence, volume 14 of Villa i Tatti. L. S. Olschki, Firenze, Italy, 1996. ISBN 88-222-4411-7. 724 pp. LCCN NB615 .B84 1996.

Bensaude-Vincent:1982:LEC

[BV82] Bernadette Bensaude-Vincent. L’Éther, Élement chimique: Un essai malheureux de mendéléev? British Journal for the History of Science, 15(2):183–188, July 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025966.

Bensaude-Vincent:1986:MPS

[BV86] Bernadette Bensaude-Vincent. Mendeleev’s periodic system of chemical elements. British Journal for the History of Science, 19(1):3–17, March 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026481.

Bensaude-Vincent:1993:LMR

[BV93] Bernadette Bensaude-Vincent.

Lavoisier: mémoires d’une révolution. (French) [Lavoisier: memoirs of a revolutions]. Figures de la science. Flammarion, Paris, France, 1993. ISBN 2-08-211208-X. 469 pp. LCCN QD22.L4 B28 1993. 160F.

Bensaude-Vincent:1995:LEC

[BVA95] Bernadette Bensaude-Vincent and Ferdinando Abbri, editors. Lavoisier in European context: negotiating a new language for chemistry, volume 1 of European studies in science history and the arts. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1995. ISBN 0-88135-189-X. LCCN QD7 .L38 1995.

Bridson:1990:PAA

[BW90] Gavin D. R. Bridson and James J. (James Joseph) White. Plant, animal and anatomical illustration in art and science: a bibliographical guide from the 16th century to the present day. St Paul’s Bibliographies in association with Hunt Institute for Botanical Documentation; Detroit : Omnigraphics, Winchester, UK, 1990. ISBN 0-906795-81-8. xxxvii + 450 pp. LCCN QH46.5 .P43

1990. URL http://catalog.
hathitrust.org/api/

                                        volumes/oclc/24505682.
                                        html.

Bynum:1972:BRS

[Byn72] William F. Bynum. Book review: Sixteenth and Seventeenth Centuries: The Road to Medical Enlightenment, 1650– 1695. By Lester S. King. London: Macdonald, 1970. Pp. x + 209. £3.50. British Journal for the History of Science, 6 (2):214, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025302.

Bynum:1973:BRM

[Byn73] William F. Bynum. Book review: Medicine Modern Methods in the History of Medicine. Ed. by Edwin Clarke. London: Athlone Press, 1971. Pp. xiv + 389. £5.50. British Journal for the History of Science, 6(3):316–317, June 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025451.

Bynum:1974:BRS

[Byn74] William F. Bynum. Book review: Seventeenth Century: Treatise of Man. René Descartes. Ed. and trans. by Thomas Steele Hall. Foreword by I. Bernard Cohen. Cambridge, Mass.: Harvard University Press, and London: Oxford University Press, 1972. Pp. xlviii + 232. £5.25. British Journal for the History of Science, 7(2):189–190, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025214.

Bynum:1978:BRN

[Byn78] W. F. Bynum. Book review:

Nineteenth and Twentieth Centuries: Darwin in America. The Intellectual Response, 1865– 1912. By Cynthia Eagle Russett. San Francisco: W. H. Freeman, 1976. Pp. 228. $4.95 (paperback). British Journal for the History of Science, 11 (2):184, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025743.

Bynum:1980:BRN

[Byn80] W. F. Bynum. Book review: Nineteenth Century: The Young Darwin and His Cultural Circle: A Study of Influences which Helped Shape the Language and Logic of the First Drafts of the Theory of Natural Selection. By Edward Manier. Dordrecht and Boston: D. Reidel, 1978. Pp. xii + 242. Hfl. 65/$24.50; Hfl 30/$11.95 (paperback). British Journal for the History of Science, 13 (2):169–170, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025903.

Bynum:1987:BRG

[Byn87] W. F. Bynum. Book review: Gerlof Verwey. Psychiatry in an Anthropological and Biomedical Context. Dordrecht/Boston/Lancaster: D. Reidel Publishing Company, 1985. Pp. xix + 316. ISBN 90-277-1713-3. £30.50. British Journal for the History of Science, 20(3):363–364, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026373.

Bynum:1988:BRJ

[Byn88] W. F. Bynum. Book review: Jan Goldstein. Console and Classify. The French Psychiatric Profession in the Nineteenth Century. Cambridge and New York: Cambridge University Press, 1987. Pp. xiii + 414. ISBN 0-521-32279-0. £30.00, $49.50. British Journal for the History of Science, 21 (3):369–370, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026842.

Carding:1992:GHT

[C^{+}92] Janet Carding et al., editors. Guide to the history of technology in Europe. Science Museum, London, UK, 1992. ISBN 0-901805-51-3. 142 pp. LCCN Q180.E9 G85 1992.

Conrad:1995:WMT

[C^{+}95] Lawrence I. Conrad et al., editors. The Western medical tradition: 800 B.C.–1800 A.D. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-38135-5. xiv + 556 pp. LCCN R131 .W47

1995. URL http://www.loc.
gov/catdir/description/
cam026/94034823.html;
http://www.loc.gov/
catdir/toc/cam021/
94034823.html.

Cahan:1988:BRD

[Cah88] David Cahan. Book review:

Donald E. Thomas Jr. Diesel: Technology and Society in Industrial Germany. Tuscaloosa, Alabama: The University of Alabama Press, 1987, Pp. xii, 279. ISBN 0-8173-0295-6 $26.95. British Journal for the History of Science, 21(2): 262–263, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026993.

Calvert:1963:BRH

[Cal63] H. R. Calvert. Book review:

Histoire Générale des Techniques. Tome 1. Les Origines de la Civilisation Technique. Published under the direction of Maurice Daumas. 83/4 × 7 inches. Pp. xvi + 652, numerous text figures, 48 plates. Paris: Presses Universitaires de France, 1963. British Journal for the History of Science, 1(3):285, June 1963. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024926.

Calvert:1977:BRA

[Cal77] Monte A. Calvert. Book review: American and Canadian Science Philadelphia’s Philosopher Mechanics. A History of the Franklin Institute 1824– 1865. By Bruce Sinclair. Baltimore and London: Johns Hopkin University Press, 1974. Pp. xiv + 354. $15.00. British Journal for the History of Science, 10(1):78, March 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025592.

Callick:1990:MMB

[Cal90] E. B. Callick. Metres to microwaves: British development of active components for radar systems, 1937–1944, volume 11 of IEE history of technology series. P. Peregrinus Ltd. on behalf of the IEE, London, UK, 1990. ISBN 0-86341-212-2. xi + 240 pp. LCCN TK6574.4.G7 C35 1990.

Campbell:1982:LRH

[Cam82] W. A. Campbell. Library resources for the history of science in Newcastle Upon Tyne. British Journal for the History of Science, 15(3):281– 284, November 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026221.

Campbell:1985:PMT

[Cam85] Margaret Campbell. Pairing in Mendel’s theory. British Journal for the History of Science, 18(3):337–340, November 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026384.

Charmasson:1987:ETR

[CAMR87] Thérèse Charmasson, Leloirain Anne-Marie, and Yannick Ripa. L’Enseignement technique de la Révolution à nos jours: textes officiels avec introduction, notes et annexes. Institut national de recherche pédagogique, Paris, France, 1987. ISBN 2-7342-0146-1 (INRP: vol. 1), 2-7178-1323-3 (Economica: vol. 1). 783 pp. LCCN LB1028.3. ff 350ff.350.

Cantor:1970:CRY

[Can70] Geoffrey Cantor. The changing role of Young’s ether. British Journal for the History of Science, 5(1):44–62, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025352.

Cantor:1974:BRN

[Can74]  G. N. Cantor.  Book re-

view: Nineteenth and Twentieth Centuries: The Archives of the Royal Institution of Great Britain in Facsimile. Minutes of Managers’ Meetings 1799– 1900. Volumes I and II. Ed. by Frank Greenaway and G. A. Stollard. Ilkley: The Scolar Press, 1971. Pp. viii + 280. No price stated for individual volumes. British Journal for the History of Science, 7 (1):92–93, March 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025189.

Cantor:1975:BRN

[Can75]  G. N. Cantor.  Book re-

view: Nineteenth and Twentieth Centuries: The Discovery of the Conservation of Energy. By Yehuda Elkana. London: Hutchinson, 1974. Pp. x + 213. £4. British Journal for the History of Science, 8 (1):87–88, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025832.

Cannon:1976:CLR

[Can76] W. Faye Cannon. Charles Lyell, radical actualism, and theory. British Journal for the History of Science, 9(2): 104–120, July 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025799.

Cantor:1977:BRE

[Can77]  G. N. Cantor.  Book re-

view: Eighteenth and Nineteenth Centuries: Scottish Philosophy and British Physics 1750–1880. A Study in the Foundations of the Victorian Scientific Style. By Richard Olson. Princeton, N.J.: Princeton University Press, 1975. Pp. viii + 350. £11.00. British Journal for the History of Science, 10(1):81–84, March 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025595.

Caneva:1980:AEO

[Can80a] Kenneth L. Caneva. Ampère, the etherians, and the Oersted connexion. British Journal for the History of Science, 13(2):121–138, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025891.

Cantor:1980:BRG

[Can80b] G. N. Cantor. Book review:

General Studies in Perception. Interrelations in the History of Philosophy and Science. Edited by Peter K. Machamer and Robert G. Turnbull. Columbus, Ohio: Ohio State University Press, 1978. Pp. x + 568. $30.00. British Journal for the History of Science, 13 (1):62–63, March 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026123.

Cantor:1980:BRS

[Can80c]  G. N. Cantor.  Book re-

view: Seventeenth and Eighteenth Centuries: Electricity in the 17th and 18th Centuries: A Study of Early Modern Physics. By J. L. Heilbron. Berkeley, Los Angeles and London: University of California Press, 1979. Pp. xiv + 606. $40.00/£24.00. British Journal for the History of Science, 13(3):270–272, November 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026207.

Cantor:1983:OAN

[Can83] G. N. Cantor. Optics after Newton: theories of light in Britain and Ireland, 1704– 1840. Manchester University Press, Manchester, UK, 1983. ISBN 0-7190-0938-3. ix + 257 pp. LCCN QC352 .C36 1983.

Cantor:1984:BRS

[Can84]  G. N. Cantor.  Book re-

view: Science in Culture: H. N. Jahnke and M. Otte (eds), Epistemological and social problems of the sciences in the early nineteenth century. Dordrecht, Boston and London: D. Reidel Publishing Company, 1981. Pp. xlii + 430. ISBN 90-277-1223-9. Dfl. 60.00, $31.50. British Journal for the History of Science, 17 (2):246–247, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026580.

Cantor:1986:BRF

[Can86] David Cantor. Book review:

F. Neil Johnson. The History of Lithium Therapy. London and Basingstoke: Macmillan, 1984. Pp. xvii + 198 illustrations. ISBN 0-333-36942-4. £35.00. British Journal for the History of Science, 19(3): 368–369, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026540.

Canguilhem:1988:IRH

[Can88] Georges Canguilhem. Ideology and rationality in the history of the life sciences. MIT Press, Cambridge, MA, USA, 1988. ISBN 0-262-03137-X. xi + 160 pp. LCCN QH305 .C2613 1988.

Cantor:1989:WWF

[Can89] Geoffrey Cantor. Why was Faraday excluded from the Sandemanians in 1844? British Journal for the History of Science, 22(4): 433–437, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026920.

Cantor:1991:MFS

[Can91] Geoffrey N. Cantor, editor.

Michael Faraday: Sandemanian and Scientist: a study of science and religion in the Nineteenth Century. MacMillan Publishing Company, New York, NY, USA, 1991. ISBN 0-333-55077-3 (hardcover), 0-333-58802-9 (paperback). xi + 359 pp. LCCN QC16.F2.

Caneva:1993:RMC

[Can93a] Kenneth L. Caneva. Robert Mayer and the conservation of energy. Princeton University Press, Princeton, NJ, USA, 1993. ISBN 0-691-08758-X. xxiii + 439 pp. LCCN QC73.8.C6 C36 1993. US$49.50. URL http://www.

loc.gov/catdir/
description/prin031/
92025400.html;   http://
www.loc.gov/catdir/toc/
prin031/92025400.html.

Cannell:1993:GGM

[Can93b] D. M. (Doris Mary) Cannell. George Green: mathematician and physicist, 1793– 1841: the background to his life and work. Athlone Press, London, UK, 1993. ISBN 0-485-11433-X. xxvi + 265 + 8 pp. LCCN QC16.G64 C36 1993. US$70.00.

Caplan:1981:BM

[Cap81a]  Jane Caplan.  Back mat-

ter. British Journal for the History of Science, 14(2):225– 226, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025961.

Caplan:1981:BRN

[Cap81b]  Jane  Caplan.

Book review: Nineteenth and Twentieth Centuries: Naturwissenschaft, Technik und NS-Ideologie. Beiträge zur Wissenschaftsgeschichte des dritten Reichs. Ed. by Herbert Mehrtens and Steffen Richter. Frankfurt-am-Main: Suhrkamp Verlag, 1980 Pp. 289. DM 13. British Journal for the History of Science, 14(2):219– 221, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025959.

Cardwell:1966:SCC

[Car66] D. S. L. Cardwell. Science and civilization in China: Volume 4, part 2: Mechanical engineering. British Journal for the History of Science, 3(2): 183–184, December 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025028.

Cardwell:1967:BRH

[Car67a] D. S. L. Cardwell. Book review:

History of Technology: Selections from Lives of Engineers, with an Account of Their Principal Works. By Samuel Smiles; ed. and with an introduction by Thomas Parke Hughes. Cambridge, Mass., & London; M.I.T. Press. 1966. 75s. British Journal for the History of Science, 3(4):401, December 1967.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024970.

Cardwell:1967:SFE

[Car67b] D. S. L. Cardwell. Some factors in the early development of the concepts of power, work and energy. British Journal for the History of Science, 3 (3):209–224, June 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025048.

Cardwell:1969:BRE

[Car69] D. S. L. Cardwell. Book review: Eighteenth and Nineteenth Centuries: The Ninth Bridgewater Treatise: A Fragment. By Charles Babbage. Second edition. 1837. London: F. Cass. 1967. 90s. British Journal for the History of Science, 4(4): 421–422, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025254.

Cardwell:1970:BRT

[Car70] D. S. L. Cardwell. Book review: Technology A History of Machine Tools, 1700–1910. By W. Steeds. Oxford: Clarendon Press: Oxford University Press. 1969. Pp. xx + 181. 153 plates. £6 6s. British Journal for the History of Science, 5(2):200–202, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025432.

Carr:1974:BRS

[Car74] J. Carr. Book review: Seventeenth Century: An Essay about the Origin and Virtues of Gems (London, 1672). By Robert Boyle. Reprint with Introduction by Arthur F. Hagner. New York: Hafner Publishing Company, 1972. Pp. xxviii + 185. No price stated. British Journal for the History of Science, 7(2):190, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025215.

Cardwell:1978:OWV

[Car78] Donald Cardwell. Obituary:

Wilfred Vernon Farrar, 1920– 1977. British Journal for the History of Science, 11(1):97– 99, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025632.

Cardwell:1984:BRH

[Car84] Donald Cardwell. Book review: History of Technology: C. C. Gillispie, The Montgolfier Brothers and the Invention of Aviation, 1783–1784, Princeton: Princeton University Press, 1983. Pp. xi+ 210. ISBN 0-691-08321-5. £30.20. British Journal for the History of Science, 17(3):326– 327, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026641.

Carpenter:1986:HSV

[Car86] Kenneth J. (Kenneth John) Carpenter. The history of scurvy and vitamin C. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-32029-1. viii + 288 pp. LCCN RC627.S36 C36

                                        1986. URL http://www.loc.
                                        gov/catdir/description/
                                        cam032/85025464.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam031/
                                        85025464.html.

Carey:1988:TEU

[Car88]  S.  Warren

(Samuel Warren) Carey. Theories of the Earth and universe: a history of dogma in the earth sciences. Stanford University Press, Stanford, CA, USA, 1988. ISBN 0-8047-1364-2. xviii + 413 pp. LCCN QE6 .C37 1988. US$45.00.

Cardwell:1989:JJB

[Car89] D. S. L. (Donald Stephen Lowell) Cardwell. James Joule: a biography. Manchester University Press, Manchester, UK, 1989. ISBN 0-7190-3025-0, 0-7190-3479-5 (paperback). ix + 333 pp. LCCN QC16.J68 C368 1989.

Cardwell:1994:FHT

[Car94] D. S. L. (Donald Stephen Lowell) Cardwell. The Fontana history of technology. Fontana history of science. Fontana Press, London, UK, 1994. ISBN 0-00-686176-8 (paperback). xviii + 565 + 8 pp. LCCN T15 .C37 1994.

Cassidy:1982:BRN

[Cas82] David C. Cassidy. Book review: The New Physics Katherine Russel Sopka, Quantum physics in America, 1920–1935. New York: Arno Press, 1980. Pp. xx + 547. $42.00. British Journal for the History of Science, 15(2):202–204, July 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025975.

Cassidy:1985:BRL

[Cas85] David C. Cassidy. Book review: Laurie M. Brown and Lillian Hoddeson, editors, The birth of particle physics. Cambridge: Cambridge University Press, 1983. Pp. xxii + 412. ISBN 0-521-24005-0 £27.50, $44.50. ‘Colloque international sur l’histoire de la physique des particules. Quelques découvertes, concepts, institutions des années 30 aux années 50’. Journal de Physique. Tome 43, Colloque C-8, supplement au no. 12, Decembre 1982. British Journal for the History of Science, 18(2):235– 237, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026335.

Caton:1988:PPO

[Cat88] Hiram Caton. The politics of progress: the origins and development of the commercial republic, 1600–1835. University of Florida Press, Gainesville, FL, USA, 1988. ISBN 0-8130-0847-6. xii + 627 pp. LCCN D217 .C33 1988. URL http:

                                        //www.loc.gov/catdir/toc/
                                        fy0714/86030886.html.

Caufield:1989:MEC

[Cau89] Catherine Caufield. Multiple exposures: chronicles of the radiation age. Secker and Warburg, London, UK, 1989. ISBN 0-436-09478-9. xii + 304 + 8 pp. LCCN RA569 .C38 1989b. US$12.95.

Cavendish:1996:GNP

[Cav96] Margaret Cavendish, Duchess of Newcastle. Grounds of natural philosophy, volume 2 of Women in the sciences. Locust Hill Press, West Cornwall, CT, 1996. ISBN 0-933951-66-3. xx + 311 pp. LCCN QC19 .N57 1996.

Chant:1990:STE

[CB^{+}90] Colin Chant, Richard Bessel, et al., editors. Science, technology and everyday life, 1870– 1950. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1990. ISBN 0-415-00037-8 (hardcover), 0-415-03557-0 (paperback). 391 pp. LCCN ????

Chaldecott:1965:RCF

[CC65] J. A. Chaldecott and A. C.

Crombie. Report of the Council and financial statement. British Journal for the History of Science, 2(4):367– 370, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024903.

Chaldecott:1966:BSH

[CC66] J. A. Chaldecott and A. C.

Crombie. The British Society for the History of Science: Report of Council for the year 1965–1966. British Journal for the History of Science, 3(2): 203–207, December 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cannon:1981:EDV

[CD81] John T. Cannon and Sigalia Dostrovsky. The evolution of dynamics: vibration theory from 1687 to 1742: with 10 illustrations, volume 6 of Studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 0-387-90626-6. ix + 184 pp. LCCN QA865 .C36.

Canguilhem:1994:VRS

[CD94]  Georges  Canguilhem

and Franøcois Delaporte. A vital rationalist: selected writings from Georges Canguilhem. Zone Books, New York, NY, USA, 1994. ISBN 0-942299-72-8. 481 pp. LCCN Q125 .C34 1994. US$26.95.

Cohen-:1990:PRM

[CDS90] I. Bernard Cohen, K. E. Duffin, and Stuart Strickland, editors. Puritanism and the rise of modern science: the Merton thesis. Rutgers University Press, New Brunswick, NJ, USA, 1990. ISBN 0-8135-1529-7, 0-8135-1530-0 (paperback). xiii + 402 pp. LCCN Q127.G4 M4636 1990. US$38.00, US$18.00.

Chaldecott:1973:BSH

[CE73] J. A. Chaldecott and M. P. Earles. The British Society for the History of Science. British Journal for the History of Science, 6(4):464– 466, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025529.

Chaldecott:1974:RCY

[CE74] J. A. Chaldecott and M. P. Earles. Report of Council for the year 1973/1974. British Journal for the History of Science, 7(3):306–309, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025554.

Ceruzzi:1989:BLF

[Cer89]  Paul E. Ceruzzi.  Beyond

the limits: flight enters the computer age. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-03143-4, 0-262-53082-1 (paperback). xi + 270 pp. LCCN TL553 .C38 1989.

Carnot:1986:RMP

[CF86] Sadi Carnot and Robert Fox.

Reflexions on the motive power of fire: a critical edition with the surviving scientific manuscripts. Manchester University Press, Manchester, UK, 1986. ISBN 0-7190-1741-6, 0-936508-16-7. viii + 230 pp. LCCN ???? URL http://

catalog.hathitrust.org/
api/volumes/oclc/
12976087.html.

ClerkMaxwell:1995:MHS

[CGBE95] James Clerk Maxwell, Elizabeth Garber, Stephen G. Brush, and C. W. F. (C. W. Francis) Everitt. Maxwell on heat and statistical mechanics: on “avoiding all personal enquiries” of molecules. Lehigh University Press, Bethlehem, PA, USA, 1995. ISBN 0-934223-34-3. 550 pp. LCCN QC310.2 .M39 1995.

Cantor:1991:F

[CGJ91] G. N. Cantor, David Gooding, and Frank A. J. L. James. Faraday. MacMillan Publishing Company, New York, NY, USA, 1991. ISBN 0-333-54291-6 (paperback). xiii + 109 pp. LCCN QC16.F2 C36 1991. US$8.99.

Chaldecott:1967:BSH

[CH67] J. A. Chaldecott and A. R. Hall. The British Society for the History of Science. British Journal for the History of Science, 3(4):411–414, December 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024981.

Chaldecott:1968:RCF

[CH68] J. A. Chaldecott and A. R. Hall. Report of Council and financial statement. British Journal for the History of Science, 4(2):201–205, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Chaldecott:1965:BRH

[Cha65] J. A. Chaldecott. Book review: A History of Scientific and Technical Periodicals: The Origins and Development of the Scientific and Technological Press 1665–1790. By David A. Kronick. Pp. 274; tables. New York: The Scarecrow Press, 1962. $6.50. British Journal for the History of Science, 2(4): 360–361, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024896.

Chaldecott:1975:JWS

[Cha75] John A. Chaldecott. Josiah Wedgwood (1730–95) — scientist. British Journal for the History of Science, 8(1): 1–16, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025813.

Chalmers:1977:BRG

[Cha77a] A. F. Chalmers. Book review: General The Interaction Between Science and Philosophy. Ed. by Y. Elkana. Atlantic Highlands, N.J.: Humanities Press, 1974. Pp. xvii + 481. No price stated. British Journal for the History of Science, 10(1):68–69, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025583.

Chapman:1977:BRS

[Cha77b] Allan Chapman. Book review:

Seventeenth and Eighteenth Centuries: Francis Place and the Early History of the Greenwich Observatory. By Derek Howse. New York: Science History Publications, 1975. Pp.

64. £4.50. British Journal for the History of Science, 10(3): 266–267, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025854.

Chaldecott:1978:BHS

[Cha78a] J. A. Chaldecott. Bibliography: A History of Scientific & Technical Periodicals: The Origins and Development of the Scientific and Technical Press, 1665– 1790. By David A. Kronick. 2nd edition. Metuchen, New Jersey: The Scarecrow Press, 1976. Pp. xvi + 336. $13.50. British Journal for the History of Science, 11(1):67, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025610.

Chaldecott:1978:BCS

[Cha78b] J. A. Chaldecott. Bibliography: Communication in Science. By A. J. Meadows. London: Butterworths, 1974. Pp.

248. £6.00. British Journal for the History of Science, 11 (1):67–68, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025611.

Chaldecott:1980:ODC

[Cha80a] J. A. Chaldecott. Obituary: Donovan Chilton: 24 February 1909–17 January 1978. British Journal for the History of Science, 13(1):88–89, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026140.

Chant:1980:BRN

[Cha80b] Colin Chant. Book review:

Nineteenth Century: Scientific Materialism in Nineteenth-Century Germany. By Frederick Gregory. Dordrechtand Boston: D. Reidel, 1977. Pp. xvii + 279. $28.00. British Journal for the History of Science, 13(2):173–175, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025907.

Chapman:1980:BRS

[Cha80c] Allan Chapman. Book review: Scientific Instruments The Chamber of Physics: Instruments in the History of Science Collections of the Royal Swedish Academy of Sciences, Stockholm. By Gunnar Pipping. Stockholm. Almqvist and Wiksell, 1976. Pp. 250 + 43 plates. 70 Cr. British Journal for the History of Science, 13 (1):65–67, March 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026126.

Charlton:1982:HTS

[Cha82] T. M. (Thomas Malcolm)

Charlton. A history of theory of structures in the Nineteenth Century. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-23419-0. viii + 194 pp. LCCN TA646 .C45

1982. URL http://www.loc.
gov/catdir/description/

                                        cam022/81015515.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        81015515.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam031/81015515.html.

Chandrasekhar:1983:EMD

[Cha83]  S.  (Subrahmanyan)

Chandrasekhar. Eddington, the most distinguished astrophysicist of his time. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25746-8. 64 pp. LCCN QB36.E33 C48

                                        1983. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0914/84123568-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0914/84123568-t.html.

Chaplain:1984:CTL

[Cha84]  Jean-Michel Chaplain.  La

chambre des tisseurs: Louviers, cité drapière, 1680–1840. (French) [The weavers room: Louviers, drapery city, 1680–1840]. Collection Milieux, 0291-7157. Champ Vallon, Seyssel, France, 1984. ISBN 2-903528-40-3. 302 pp. LCCN HD9862.8.L68 C46 1984. 178F.

Chaldecott:1988:OLB

[Cha88a] J. A. Chaldecott. Obituary: Leslie Bernard Hunt: 1906– 1987. British Journal for the History of Science, 21(1):99– 100, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026865.

Chapman:1988:CEO

[Cha88b] Michael Chapman. Constructive evolution: origins and development of Piaget’s thought. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-33163-3, 0-521-36712-3 (paperback). xii + 459 pp. LCCN BF723.C5 C484

1988. URL http://www.loc.
gov/catdir/description/
cam023/87021816.html;
http://www.loc.gov/
catdir/toc/cam029/
87021816.html.

Chalmers:1990:SF

[Cha90a] A. F. (Alan Francis) Chalmers.

Science and its fabrication. Open University, Milton Keynes, UK, 1990. ISBN 0-335-09318-3 (hardcover), 0-335-09317-5 (paperback). xii + 142 pp. LCCN Q175 .C4457

1990. URL http://catalog.
hathitrust.org/api/
volumes/oclc/20295911.
html.

Chalmers:1990:SFC

[Cha90b] Alan Chalmers. Science and its fabrication. Open Univ. Press, Buckingham, UK, 1990. ISBN 0-335-09318-3, 0-335-09317-5. xii + 142 pp. LCCN ????

Chapman:1990:DCD

[Cha90c] Allan Chapman. Dividing the circle: development and critical measurement of celestial angles (1500–1850). Ellis Horwood library of space science and space technology. Series in astronomy. Ellis Horwood Ltd., New York, NY, USA, 1990. ISBN 0-13-217555-X. 209 pp. LCCN QB147 .C48 1990. US$15.95.

Channell:1991:VMS

[Cha91] David F. Channell. The vital machine: a study of technology and organic life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-506040-7. xi + 192 pp. LCCN T14 .C464 1991. US$22.95. URL http:

//www.loc.gov/catdir/
enhancements/fy0639/
90020273-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0725/90020273-b.html.

Chapman:1994:DCD

[Cha94] Allan Chapman. Dividing the circle: the development of critical angular measurement in astronomy 1500–1850. Wiley in association with Praxis, Chichester, UK, 1994. ISBN 0-13-217555-X. 209 pp. LCCN ????

Cheesman:1978:BRB

[Che78] Brian Cheesman. Book review: Bibliography Union Catalogue of Scientific Libraries in the University of Cambridge: Scientific Conference Proceedings, 1644–1972. Compiled at the Scientific Periodicals Library, University of Cambridge. London: Mansell, 1975. 2 vols. Pp. vi + [610]; [612]. £27.50. British Journal for the History of Science, 11 (1):66–67, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025609.

Chevalley:1985:BRJ

[Che85] Catherine Chevalley. Book review: John Hendry, The Creation of Quantum Mechanics and the Bohr–Pauli Dialogue. Dordrecht/ Boston/Lancaster: D. Reidel Publishing Company, 1984. Pp.

177. ISBN 90-277-1648-X. DF1 95. £23.50. British Journal for the History of Science, 18(3): 362–364, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026401.

Collie:1991:HWS

[CHG91] Michael Collie, Thomas Henry Huxley, and George Gordon. Huxley at work: with the scientific correspondence of T. H. Huxley and the Rev. Dr. George Gordon of Birnie, near Elgin. MacMillan Publishing Company, New York, NY, USA, 1991. ISBN 0-333-51059-3. xii + 158 + 6 pp. LCCN QE22.H94 A4 1991.

Chilton:1964:BRC

[Chi64] D. Chilton. Book review: Clerk Maxwell and Modern Science. Six commemorative lectures by Sir Edward V. Appleton, E. G. Bowen, C. A. Coulson, R. E. Peierls, Sir John Randall, and R. A. Smith. Edited by C. Domb. Pp. x + 118, frontis., 5 plates and numerous text figures. University of London, The Athlone Press. 1963. Price 25s. net. British Journal for the History of Science, 2(1):76, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025088.

Chilton:1965:BRSb

[Chi65a] D. Chilton. Book review: A Source Book in Physics. By William Francis Magie. Eighth printing. Pp. xiv + 620. Harvard University Press, Cambridge, Massachusetts, 1963. £4 4s. British Journal for the History of Science, 2(4): 362–363, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024898.

Chilton:1965:BRJ

[Chi65b] D. Chilton. Book review: J. J. Thomson and the Cavendish Laboratory. By Sir George Paget Thomson, F.R.S. Pp. xii + 186, 20 half-tone plates and 25 line figures. London: Nelson, 1964. 25s. net. British Journal for the History of Science, 2(3):266, June 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024947.

Chilton:1965:BRSa

[Chi65c] D. Chilton. Book review: Scientific Change: Symposium on the history of science held at Oxford 9–15 July 1961. Edited by A. C. Crombie. Pp. xii + 896. London: Heinemann, 1963. £5 5s. net. British Journal for the History of Science, 2(3):263, June 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024943.

Chilton:1968:BRSc

[Chi68] D. Chilton. Book review: Scientific Instruments and Museums: A History of the Thermometer and its use in Meteorology. By W. E. Knowles Middleton. Baltimore, Johns Hopkins Press. London: Oxford University Press. 1966. Pp. vii + 249. Illustrated. British Journal for the History of Science, 4(2):179, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025152.

Chilton:1971:BRSd

[Chi71] D. Chilton. Book review: Scientific Instruments: Invention of the Meteorological Instruments. By W. E. Knowles Middleton. Baltimore, Johns Hopkins Press. 1969. Pp. xiv + 362. Illus. £4.70. British Journal for the History of Science, 5 (3):293–294, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025321.

Christensen:1993:EHT

[Chr93] Dan Charly Christensen, editor. European historiography of technology: proceedings from the TISC-Conference in Roskilde, [February 1992], volume 156 [i.e. 155] of Odense University studies in history and social sciences, 0078-3307. Odense University Press, Odense, Denmark, 1993. ISBN 87-7492-910-0. LCCN T26.A1 T57 1992.

                                        URL http://books.google.
                                        com/books?id=OD3bAAAAMAAJ.

Christianson:1995:EHM

[Chr95] Gale E. Christianson. Edwin Hubble: mariner of the nebulae. Farrar, Straus, Giroux, New York, NY, USA, 1995. ISBN 0-374-14660-8. x + 420 pp. LCCN QB36.H83 C48 1995.

Crawford:1987:NPC

[CHU87] Elisabeth T. Crawford, J. L. Heilbron, and Rebecca Ullrich. The Nobel population 1901– 1937: a census of the nominators and nominees for the prizes in physics and chemistry, volume 11; 4 of Berkeley papers in history of science; Uppsala studies in history of science. Office for History of Science and Technology, University of California at Berkeley, Berkeley, CA, USA, 1987. ISBN 0-918102-15-4 (paperback). 337 pp. LCCN QC28 .C74 1987.

Cittadino:1990:NLD

[Cit90]  Eugene Cittadino.  Nature

as the laboratory: Darwinian plant ecology in the German Empire, 1880–1900. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-34045-4. xi + 199 pp. LCCN QK900.73.G3 C58

1990. URL http://www.loc.
gov/catdir/description/
cam024/90001422.html;
http://www.loc.gov/
catdir/samples/cam031/
90001422.html;   http://
www.loc.gov/catdir/toc/
cam029/90001422.html.

Clarke:1987:NCO

[CJ87] Edwin Clarke and L. S. Jacyna. Nineteenth-Century origins of neuroscientific concepts. University of California Press, Berkeley, CA, USA, 1987. ISBN 0-520-05694-9. 593 pp. LCCN QP353 .C57 1987.

Cunningham:1990:RS

[CJ90] Andrew Cunningham and Nicholas Jardine, editors. Romanticism and the sciences. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-35602-4, 0-521-35685-7 (paperback). xxii + 345 pp. LCCN Q125 .R735

1990. URL http://www.loc.
gov/catdir/description/
cam023/89017310.html;

                                        http://www.loc.gov/
                                        catdir/toc/cam023/
                                        89017310.html.

Campbell-Kelly:1989:IBT

[CK89] Martin Campbell-Kelly. ICL: a business and technical history. Clarendon Press, New York, NY, USA, 1989. ISBN 0-19-853918-5. xii + 409 + 36 pp. LCCN HD9696.C64 I463 1989. US$37.25. URL http://www.

                                        loc.gov/catdir/
                                        enhancements/fy0639/
                                        89016318-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0639/89016318-t.html.

Cotterell:1990:MPI

[CK90] Brian Cotterell and Johan Kamminga. Mechanics of preindustrial technology: an introduction to the mechanics of ancient and traditional material culture. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-34194-9. xv + 325 pp. LCCN QC125.2 .C68

                                        1989. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/88029238.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        88029238.html.

Campbell-Kelly:1996:CHI

[CKA96] Martin Campbell-Kelly and William Aspray. Computer: a history of the information machine. The Sloan technology series. Basic Books, New York, NY, USA, 1996. ISBN 0-465-02989-2. ix + 342 pp. LCCN QA76.17 .C36

1996. URL http://www.loc.
gov/catdir/enhancements/
fy0831/96002098-d.html.

Cooper:1987:WCC

[CL87] William Cooper and Stanton J.

Linden. William Cooper’s A catalogue of chymicall books, 1673–88, volume 670 of Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, verified. edition, 1987. ISBN 0-8240-8557-4. liv + 159 pp. LCCN Z5522 .C77 1987.

Chard:1996:TTP

[CL96] Chloe Chard and Helen Langdon, editors. Transports: travel, pleasure, and imaginative geography, 1600–1830, volume 3 of Studies in British art. Published for the Paul Mellon Centre for Studies in British Art [and] the Yale Center for British Art [by] Yale University Press, New Haven, 1996. ISBN 0-300-06382-2 (hardcover). viii + 341 pp. LCCN PN56.T7 T67 1996.

Clarke:1962:GAD

[Cla62] Leonard W. Clarke. Greek astronomy and its debt to the Babylonians. British Journal for the History of Science, 1 (1):65–77, June 1962. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025075.

Clarke:1966:BRC

[Cla66]  E.  Clarke.

Book review: Charles Scott Sherrington. An Appraisal. By Ragnar Granit. Pp. xi + 188. London: Thomas Nelson, 1966. £2 2s. British Journal for the History of Science, 3(2): 198–199, December 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025043.

Clarke:1967:BRH

[Cla67] Edwin Clarke. Book review: History of Biological Sciences and Medicine S. Ramón y Cajal, Recollections of my Life. Trans, by E. Horne Craigie with the assistance of J. Cano. Pp. xi + 638. Reprint of 1937 edition. Cambridge, Mass, and London: M.I.T. Press, n.d. [1966], 75s. British Journal for the History of Science, 3(4): 407–408, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024977.

Clarke:1968:BRH

[Cla68]  Edwin Clarke.  Book re-

view: History of Medicine Tom Rivers. Reflections on a Life in Medicine and Science. An Oral History. By Saul Benison. Pp. xxi + 682. Cambridge, Mass. and London: M.I.T. Press. 1967. 140s. British Journal for the History of Science, 4(2):

185–186, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025158.

Clarke:1970:BRB

[Cla70]  Edwin Clarke.  Book re-

view: Biology and Medicine Un grand médecin et biologiste: Casimir-Joseph Davaine (1812– 1882). By Jean Théodoridès. (Analecta Medico-Historica 4.) Oxford: Pergamon Press. 1968. Pp. 238. Plates. £5. British Journal for the History of Science, 5(1):97, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025367.

Clarke:1971:BRB

[Cla71]  Edwin Clarke.  Book re-

view: Biological Sciences and Medicine Reflexes and Motor Integration: Sherrington’s Concept of Integrative Action. By Judith P. Swazey. Harvard University Press and Oxford University Press. 1969. Pp. xviii + 273. £3.15. British Journal for the History of Science, 5(3):305–306, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025334.

Clark:1979:SPS

[Cla79] Peter Clark. Stegmüller and the philosophy of science. British Journal for the History of Science, 12(3):337–341, November 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026008.

Clarke:1982:DPS

[Cla82] Desmond M. Clarke. Descartes’ philosophy of science. Studies in intellectual history. Manchester University Press, Manchester, UK, 1982. ISBN 0-7190-0868-9. xii + 249 pp. LCCN Q175 .C58 1982. US$19.00.

Clark:1986:BRN

[Cla86a]  Michael J. Clark.  Book

review: Nancy Tomes. A Generous Confidence. Thomas Story Kirkbride and the Art of Asylum-Keeping, 1840–1883. Cambridge: Cambridge University Press, 1984. Pp. xv + 387. ISBN 0-521-24172-3. £27.50 $39.50. British Journal for the History of Science, 19 (1):127–128, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026500.

Clarke:1986:BRK

[Cla86b] Joe Clarke. Book review: K. C. Barraclough. Steelmaking before Bessemer. Vol. 1 Blister Steel-The Birth of an Industry. Vol. 2 Crucible Steel-the Growth of a Technology. London: The Metals Society, 1984. Pp. xx + 273, xix + 387. ISBN 0-904357-53-8 and 0-904357-64-3. Vol. 1 £19.95, $39.90. Vol. 2 £29.50, $59. British Journal for the History of Science, 19 (1):116–117, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026489.

Clarke:1989:OPH

[Cla89] Desmond M. Clarke. Occult powers and hypotheses: Cartesian natural philosophy under Louis XIV. Clarendon Press, New York, NY, USA, 1989. ISBN 0-19-824812-1. 265 pp. LCCN Q174.8 .C57 1989. US$30.00. URL http://www.

loc.gov/catdir/
enhancements/fy0637/
88031260-d.html.

Clagett:1995:AES

[Cla95a]  Marshall Clagett.  Ancient

Egyptian science. Volume II. Calendars, clocks, and astronomy: a source book, volume 214 of Memoirs of the American Philosophical Society, 0065-9738. American Philosophical Society, Philadelphia, PA, USA, 1995. ISBN 0-87169-214-7. xiv + 575 + 144 pp. LCCN Q11 .P612 vol. 184; QB20 .C62 1995.

Clagett:1995:CCA

[Cla95b] Marshall Clagett, editor. Calendars, clocks, and astronomy, volume Vol. 2; 214 of Ancient Egyptian science: a source book / by Marshall Clagett; Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1995. ISBN 0-87169-214-7. xiv + 575 + 70 pp. LCCN ????

Clagett:1986:AES

[Cla99a]  Marshall Clagett.  Ancient

Egyptian science: a source book, volume 184, 214, 232 of Memoirs of the American philosophical society, 0065-9738. American Philosophical Society, Philadelphia, PA, USA, 1989–1999. ISBN 0-87169-184-1 (vol. 1),0-87169-214-7 (vol. 2), 0-87169-232-5 (vol. 3). 4 t. en 3 vol. pp. LCCN ????

Clagett:1989:AES

[Cla99b]  Marshall Clagett.  Ancient

Egyptian science: a source book, volume 184, 214, 232 of Memoirs of the American Philosophical Society, 0065-9738. American Philosophical Society, Philadelphia, PA, USA, 1989–1999. ISBN 0-87169-184-1 (vol. 1), 0-87169-232-5 (vol. 3). ???? pp. LCCN Q11 .P612 vol. 184, etc.

Clagett:20xx:AESa

[Cla9 ]  Marshall Clagett.  Ancient

Egyptian science: a source book, volume 184 + 214 + 232 of Memoirs series; 0065-9738. American Philosophical Society, Philadelphia, PA, USA, 1989-. ISBN 0-87169-214-7 (vol. 2), 0-87169-232-5 (vol. 3). bd. pp. LCCN ????

Clagett:20xx:AESb

[Cla2 ] Marshall Clagett, editor. Ancient Egyptian science: a source book, volume 184 of Memoirs of the American Philosophical Society, 0065-9738. American Philosophical Society, Philadelphia, PA, USA, 1992-. ISBN 0-87169-184-1 (1), 0-87169-214-7 (2), 0-87169-232-5 (3). bd. pp. LCCN ????

Cleevely:1983:WPC

[Cle83] R. J. Cleevely. World palaeontological collections. British Museum (Natural History), London, UK, 1983. ISBN 0-565-00850-1, 0-7201-1655-4 (Mansell). 365 pp. LCCN QE716.G72 L633 1983.

Clow:1980:BRC

[Clo80] Archie Clow. Book review: Chemical Industry Neptune’s Gift: A History of Common Salt. By Robert P. Multhauf. Baltimore & London: Johns Hopkins University Press, 1978. Pp xviii + 325.£15.75. British Journal for the History of Science, 13(3): 266–267, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026204.

Charmasson:1987:LTR

[CLR87] Thérèse Charmasson, Anne-Marie Lelorrain, and Yannick Ripa. L’Enseignement technique de la Révolution à nos jours: textes officiels avec introduction, notes et annexes. Tome I, De la Révolution à 1926. Economica, Paris, France, 1987. ISBN 2-7342-0146-1 (INRP) 2-7178-1323-3 (Economica). 783 pp. LCCN ????

Clulee:1988:JDN

[Clu88] Nicholas H. Clulee. John Dee’s natural philosophy: between science and religion. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1988. ISBN 0-415-00625-2, 0-415-03122-2 (paperback). xiv + 347 + 24 pp. LCCN BF1598.D5 C58 1988.

Chaldecott:1964:RCF

[CM64] J. A. Chaldecott and Thomas Martin. Report of Council and financial statement. British Journal for the History of Science, 2(2):177–179, December 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025022.

Cohen:1990:FDF

[CM90] Yves Cohen and Klaus Manfrass, editors. Frankreich und Deutschland: Forschung, Technologie und industrielle Entwicklung im 19. und 20. Jahrhundert: internationales Kolloquium, veranstaltet vom Deutschen Historischen Institut Paris in Verbindung mit dem Deutschen Museum München und der Cité des sciences et de l’industrie, Paris: München, 12.–15. Oktober 1987. (German) [ France and Germany: Research, technology and industrial development in the 19th and 20th Century: international conference organized by the German Historical Institute, Paris and the City of Science and Industry, Paris, Munich, 12–15 October 1987]. In Kommission bei C.H. Beck, München, Germany, 1990. ISBN 3-406-35092-5. xl + 491 pp. LCCN HC280.T4 F73 1990.

Ceard:1993:HDC

[CM93] Jean Céard and Paul Mengal, editors. Histoire du concept de récapitulation: ontogenèse et phylogenèse en biologie et sciences humaines. Histoire des sciences. Masson Editeur, Masson, France, 1993. ISBN 2-225-84057-1. 132 pp. LCCN ????

Challesca:1982:DRP

[CMDM82] Robert Challes, Nicolas Malebranche, Frédéric Deloffre, and Melâhat Menemencioglu, editors. Difficultés sur la religion proposées au père Malebranche. (French) [Difficulties in the religion proposed to Father Malebranche], volume 209 of Studies on Voltaire and the eighteenth century, 0435-2866. The Voltaire Foundation at the Taylor Institution, Oxford, UK, 1982. ISBN 0-7294-0289-4. 576 pp. LCCN PQ2105.A2 S8 vol.

Clark:1995:EWH

[CMN95] Harold W. (Harold Willard) Clark, Frank Lewis Marsh, and Ronald L. Numbers. The early writings of Harold W. Clark and Frank Lewis Marsh, volume 8 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1809-X. xxiii + 531 pp. LCCN BS651 .C54 1995.

Conway:1992:CLC

[CMNH92] Anne Conway, Henry More, Marjorie Hope Nicolson, and Sarah Hutton. The Conway letters: the correspondence of Anne, Viscountess Conway, Henry More, and their friends, 1642–1684. Clarendon Press, New York, NY, USA, revised edition, 1992. ISBN 0-19-824876-8. xxix + 592 pp. LCCN DA378.C6 A4 1992. US$50.00. URL http://www.

                                        loc.gov/catdir/
                                        enhancements/fy0605/
                                        91038838-d.html.

Cock:1975:BRG

[Coc75] A. G. Cock. Book review: Genetics History of Genetics from the Earliest Times to the Rediscovery of Mendel’s Laws. By Hans Stubbe. Second edition. Trans, by T. R. W. Waters. Cambridge, Mass., and London: M.I.T. Press, 1972. Pp. x + 356. $14.95. British Journal for the History of Science, 8(2):177–178, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025645.

Cohen:1984:QMS

[Coh84] H. F. Cohen. Quantifying music: the science of music at the first stage of the scientific revolution, 1580–1650, volume 23 of The University of Western Ontario series in philosophy of science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1637-4. xvii + 308 pp. LCCN ML3807 .C63 1984.

Cohen:1985:RS

[Coh85] I. Bernard Cohen. Revolution in science. Belknap Press of Harvard University Press, Cambridge, MA, USA and London, UK, 1985. ISBN 0-674-76777-2. xx + 711 + 1 pp. LCCN Q125 .C542 1985.

Cohen:1994:SRH

[Coh94] H. F. Cohen. The scientific revolution: a historiographical inquiry. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-11279-9, 0-226-11280-2 (paperback). xviii + 662 pp. LCCN Q125 .C538 1994. URL http://www.

loc.gov/catdir/
description/uchi052/
93041784.html;   http://
www.loc.gov/catdir/toc/
uchi051/93041784.html.

Coleman:1972:ER

[Col72] William Coleman. Essay reviews. British Journal for the History of Science, 6(2):202– 204, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025292.

Coley:1976:BRC

[Col76] N. G. Coley. Book review: Chemistry The Royal Institution Library of Science. Chemical Manipulation, M. Faraday, 1827. Foreword by Sir George Porter. Facsimile reprint. London: Applied Science Publishers, 1974. Pp. viii + 656. £12.00. British Journal for the History of Science, 9(3): 329–330, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025692.

Coleman:1981:BRR

[Col81a] Janet Coleman. Book review: Renaissance: The Renaissance Notion of Woman: A Study in the Fortunes of Scholasticism and Medical Science in European Intellectual Life. By Ian Maclean. Cambridge: Cambridge University Press, 1980. Pp. viii + 119. £7.50. British Journal for the History of Science, 14(2):211–213, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025954.

Collins:1981:BRS

[Col81b] H. M. Collins. Book review:

Sociology of Knowledge Genesis and Development of a Scientific Fact. By Ludwik Flek. Ed. by Thaddeus J. Trenn and Robert K. Merton. Translated by Fred Bradley and Thaddeus J. Trenn. Chicago & London: Chicago University Press, 1979. Pp. xxviii + 203. £10.50 $22.75. British Journal for the History of Science, 14(2): 208–209, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025951.

Coleman:1983:BRE

[Col83] Janet Coleman. Book review: Essay Collections James A. Weisheipl, OP (ed), Albertus Magnus and the sciences: commemorative essays 1980. Toronto: Pontifical Institute of Mediaeval Studies, 1980. Pp. xiv + 658. $35.00 (paperback). British Journal for the History of Science, 16 (2):190–192, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Coleman:1987:YFN

[Col87a]  William Coleman.  Yellow

fever in the North: the methods of early epidemiology, volume 6 of Wisconsin publications in the history of science and medicine. University of Wisconsin Press, Madison, WI, USA, 1987. ISBN 0-299-11110-5, 0-299-11114-8 (paperback). xv + 202 pp. LCCN RA649 .C64 1987. US$45.00, US$22.95.

Coley:1987:BRF

[Col87b] N. G. Coley. Book review: Frederic Lawrence Holmes. Lavoisier and the Chemistry of Life. An Exploration of Scientific Creativity. Madison: University of Wisconsin Press, 1985. Pp. xxiv + 565. ISBN 0-299-09980-6. £36.60. British Journal for the History of Science, 20(1): 85–86, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026443.

Coley:1988:BRG

[Col88a] Noei G. Coley. Book review: George D. Tselos & Colleen Wickey. A Guide to Archives and Manuscript Collections in the History of Chemistry and Chemical Technology. Philadelphia: Center for History of Chemistry, 1987. Pp. viii + 198. ISBN 0-941901-05-X. $7.50 (paperback). British Journal for the History of Science, 21(1):126, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026881.

Coley:1988:BRJ

[Col88b] Noel G. Coley. Book review: James G. Traynham (ed.). Essays on the History of Organic Chemistry, Baton Rouge and London, Louisiana State University Press, 1987, Pp.ix + 145. ISBN 0-8071-1293-3. £21.25. British Journal for the History of Science, 21(3): 368, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026840.

Connor:1986:BRB

[Con86] Patrick Connor. Book review:

Barbara Maria Stafford. Voyage into Substance. Art, Science, Nature, and the Illustrated Travel Account, 1760– 1840. Cambridge, Mass., and London: MIT Press, 1984. Pp. xxiii + 645. ISBN 0-262-19223-3. £37.95. British Journal for the History of Science, 19 (2):228–229, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026611.

Connor:1987:WME

[Con87] R. D. Connor. The weights and measures of England. Her Majesty’s Stationary Office, London, England, 1987. ISBN 0-11-290435-1. xxvi + 422 + 41 pp. LCCN QC89.G82 E543 1987. US$30.00.

Cook:1981:BRR

[Coo81] Harold J. Cook. Book review: Renaissance: Health, Medicine and Mortality in the Sixteenth Century. Edited by Charles Webster. Cambridge Monographs on the History of Medicine, 1. Cambridge: Cambridge University Press, 1979. Pp. xiv + 394. $39.95/ £18.50. British Journal for the History of Science, 14(1): 88–90, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026078.

Cook:1983:BRA

[Coo83] H. J. Cook. Book review: Americana A. McGehee Harvey, Science at the bedside. Clinical research in American medicine 1905–1945. Baltimore: Johns Hopkins University Press, 1981. Pp. xix + 554. £12.25/$21.85. British Journal for the History of Science, 16 (3):319, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026188.

Cooter:1984:BRN

[Coo84a] Roger Cooter. Book review: National Traditions in Science James C. Whorton, Crusaders for fitness: the history of American health reformers. Princeton: Princeton University Press, 1982. Pp. 359. ISBN 0-691-04694-8. $19.50. British Journal for the History of Science, 17(1):92–93, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026028.

Cooter:1984:CMP

[Coo84b] Roger Cooter. The cultural meaning of popular science: phrenology and the organization of consent in Nineteenth-Century Britain. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-22743-7. xiii + 418 pp. LCCN Q127.G4

C66 1984. URL http://www.
loc.gov/catdir/
description/cam031/
84012078.html;   http://
www.loc.gov/catdir/toc/
cam031/84012078.html.

Cooter:1988:SHA

[Coo88] Roger Cooter, editor. Studies in the history of alternative medicine. St. Antony’s/Macmillan series. MacMillan Publishing Company, New York, NY, USA, 1988. ISBN 0-333-46213-0. xx + 180 pp. LCCN R733 .S87 1988b.

Cook:1989:BRD

[Coo89a] Harold J. Cook. Book review:

David C. Goodman. Power and Penury: Government, Technology and Science in Philip II’s Spain. Cambridge: Cambridge University Press, 1988. Pp. xii + 275. ISBN 0-521-30532-2. £30.00, $44.50. British Journal for the History of Science, 22(2):245–246, July 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026664.

Cooper:1989:BAP

[Coo89b] Andrew Fenton Cooper. British agricultural policy, 1912–36: a study in Conservative politics. Manchester University Press, Manchester, UK, 1989. ISBN 0-7190-2886-8. viii + 240 pp. LCCN HD1927 .C66 1989.

Cooter:1989:PBI

[Coo89c] Roger Cooter. Phrenology in the British Isles: an annotated historical biobibliography and index. Scarecrow Press, Metuchen, NJ, USA, 1989. ISBN 0-8108-2165-6. xviii + 431 pp. LCCN Z7204.P47 C66 1989.

Coopersmith:1992:ER

[Coo92] Jonathan Coopersmith. The electrification of Russia, 1880– 1926. Cornell University Press, Ithaca, NY, USA, 1992. ISBN 0-8014-2723-1. xii + 274 pp. LCCN HD9685.S762 C66 1992.

Copenhaver:1992:HGC

[Cop92] Brian P. Copenhaver, editor. Hermetica: the Greek Corpus Hermeticum and the Latin Asclepius in a new English translation, with notes and introduction. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-36144-3 (hardcover). lxxxiii + 320 pp. LCCN BF1600 .H475

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/91025703.html;

http://www.loc.gov/
catdir/toc/cam028/
91025703.html.

Corsi:1978:IFT

[Cor78] Pietro Corsi. The importance of French Transformist ideas for the second volume of Lyell’s principles of geology. British Journal for the History of Science, 11(3):221– 244, November 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025778.

Corsi:1981:BRN

[Cor81] Pietro Corsi. Book review:

Nineteenth and Twentieth Centuries: The Letters of Georges Cuvier. A Summary Calendar of Manuscript and Printed Materials Preserved in Europe, the United States of America, and Australasia. Ed. by Dorinda Outram. Chalfont St Giles: The British Society for the History of Science, 1979. Pp. iv + 102. £3.90 (£2.90 for Society members). British Journal for the History of Science, 14(2):216–217, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025957.

Corsi:1988:ALE

[Cor88a] Pietro Corsi. The age of Lamarck: evolutionary theories in France, 1790–1830. University of California Press, Berkeley, CA, USA, rev. and updat edition, 1988. ISBN 0-520-05830-5. xiii + 360 pp. LCCN QH15 .C6713 1988.

Corsi:1988:SRB

[Cor88b] Pietro Corsi. Science and religion: Baden Powell and the Anglican debate, 1800– 1860. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-24245-2. ix + 346 pp. LCCN BL245 .C67

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/87013845.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        87013845.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam021/87013845.html.

Corry:1996:MAR

[Cor96]  Leo Corry.  Modern alge-

bra and the rise of mathematical structures, volume 17 of Science networks. Historical studies. Birkhäuser Verlag, Basel, Switzerland, 1996. ISBN 3-7643-5311-2 (Basel), 0-8176-5311-2 (Boston). 460 pp. LCCN QA151 .C67 1996.

Correia:1997:OEE

[Cor97] Clara Pinto Correia. The ovary of Eve: egg and sperm and preformation. University of Chicago Press, Chicago, IL, USA, 1997. ISBN 0-226-66952-1. xxiii + 396 pp. LCCN Q127.E8 C67 1997. URL http:

                                        //www.loc.gov/catdir/
                                        description/uchi052/

97014011.html;   http://
www.loc.gov/catdir/toc/
uchi051/97014011.html.

Coudert:1995:LK

[Cou95] Allison Coudert. Leibniz and the Kabbalah, volume 142 of Archives internationales d’histoire des idées = International archives of the history of ideas. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3114-1. xvii + 218 pp. LCCN B2599.C33 C68

1995. URL http://www.loc.
gov/catdir/enhancements/
fy0821/94032967-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0821/94032967-t.html.

Courvisanos:1996:ICC

[Cou96] Jerry Courvisanos. Investment cycles in capitalist economies: a Kaleckian behavioural contribution. New directions in modern economics. Edward Elgar Pub., Cheltenham, UK, 1996. ISBN 1-85898-410-6. xi + 272 pp. LCCN HB3720 .C68 1996.

Cowling:1968:BRH

[Cow68] T. G. Cowling. Book review: History of Mathematics The Mathematical Papers of Sir William Rowan Hamilton. Volume III. Algebra. Edited for the Royal Irish Academy by H. Halberstam and R. E. Ingram. Pp. xxiv + 672. London: Cambridge University Press. 1967.

£10 10s. British Journal for the History of Science, 4(1): 86–88, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025003.

Cowan:1974:BRN

[Cow74] Ruth Schwartz Cowan. Book review: Nineteenth and Twentieth Centuries: Charles Darwin: The Years of Controversy. The ‘Origin of Species’ and its Critics 1859–1882. By Peter J. Vorzimmer. London: University of London Press, 1972. Pp. xx + 300. £4.40. British Journal for the History of Science, 7(1):94–95, March 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025191.

Collins:1993:GWE

[CP93] H. M. (Harry M.) Collins and T. J. (Trevor J.) Pinch. The golem: what everyone should know about science. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-35601-6 (hardcover). xii + 164 pp. LCCN Q125 .C552 1993.

Crawford:1984:BNI

[Cra84] Elisabeth T. Crawford. The beginnings of the Nobel institution: the science prizes, 1901– 1915. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-26584-3. ix + 281 + 6 pp. LCCN QC28 .C73

1984. URL http://www.loc.
gov/catdir/description/
cam022/84005844.html;
http://www.loc.gov/
catdir/toc/cam026/
84005844.html.

Crawford:1992:NIS

[Cra92] Elisabeth Crawford. Nationalism and internationalism in science, 1880–1939: four studies of the Nobel population. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40386-3 (hardcover). xii + 157 pp. LCCN Q126.9 .C73 1992.

Crabtree:1993:MFM

[Cra93] Adam Crabtree. From Mesmer to Freud: magnetic sleep and the roots of psychological healing. Yale University Press, New Haven, CT, USA, 1993. ISBN 0-300-05588-9. x + 413 pp. LCCN RC497 .C73 1993.

Crawford:1996:AIT

[Cra96] Elisabeth T. Crawford. Arrhenius: from ionic theory to the greenhouse effect, volume 23 of Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1996. ISBN 0-88135-166-0. xiii + 320 + 9 pp. LCCN Q143.A77 C68 1996.

Crevier:1993:ATH

[Cre93] Daniel Crevier. AI: the tumultuous history of the search for artificial intelligence. Basic Books, New York, NY, USA, 1993. ISBN 0-465-02997-3. xiv + 386 pp. LCCN Q335 .C66 1993. US$27.50, CAN$37.95.

Cridland:1966:BRH

[Cri66] B. H. Cridland. Book review: History of Chemistry Joseph Priestley, Adventurer in Science and Champion of Truth. By F. W. Gibbs. Pp.

258. With 20 pages of half-tone plates and 12 line drawings. London: Thomas Nelson and Sons, 1965. 42s. British Journal for the History of Science, 3(1):87–88, June 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025116.

Critchley:1971:BRB

[Cri71] Macdonald Critchley. Book review: Biological Sciences and Medicine Mind, Brain, and Adaptation in the Nineteenth Century. By Robert M. Young. Oxford: Clarendon Press: Oxford University Press. 1970. Pp. xiv + 278. £3.25. British Journal for the History of Science, 5(3):304–305, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025333.

Crombie:1966:HCB

[Cro66a]  A. C. Crombie.  Histori-

cal commitments of biology. British Journal for the History of Science, 3(2):97–108, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025024.

Crosland:1966:BRE

[Cro66b] Maurice Crosland. Book review: Enseignement et diffusion des sciences en France au XVIIIe siècle. Ed. René Taton. Pp. 780. 20 plates. Paris: Hermann, 1964. 36 F. British Journal for the History of Science, 3 (2):193, December 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025036.

Crosland:1967:BRH

[Cro67] Maurice Crosland. Book review: History of Chemistry Torbern Bergman’s Foreign Correspondence. Vol. I. Edited by Göte Carlid and Johan Nordström. Pp. lvi + 466. Plates. Stockholm: Almqvist & Wiksell. 1965. 70 Swedish kr. British Journal for the History of Science, 3(4):403, December 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024972.

Crosland:1968:BRE

[Cro68a] Maurice Crosland. Book review: Education, Science and Technology The Development of Technical Education in France, 1500– 1850. By Frederick B. Artz.

Cambridge, Mass. and London, M.I.T. Press. 1966. Pp. x + 274. 64s. British Journal for the History of Science, 4(2): 175–176, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025149.

Crosland:1968:BRH

[Cro68b] Maurice Crosland. Book review: History of Medicine Bulletin de la Société Française d’Histoire de la Médecine, Tome i. Pp. 505; plates. Paris: R. Lacer [1902], 1967. 48 F. British Journal for the History of Science, 4(1):89, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025005.

Crosland:1969:BRH

[Cro69] Maurice Crosland. Book review: History of Chemistry An Essay on Phlogiston and the Constitution of Acids. By R. Kirwan. Pp. xxiii + 317 + index. London: F. Cass. [1789]. 1968. 90s. British Journal for the History of Science, 4 (3):293, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025485.

Crosland:1970:BRC

[Cro70a] Maurice Crosland. Book review: Chemistry Classical Scientific Papers: Chemistry. Ed. by David M. Knight. London: Mills & Boon Ltd. 1968. Pp. xxiv + 391. 63s. British Journal for the History of Science, 5(1):95, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025364.

Crosland:1970:BRE

[Cro70b] Maurice Crosland. Book review: Eighteenth Century: Martinus van Marum. Life and Work. Volume 1. Ed. by R. J. Forbes. Haarlem: Tjeenk Willink & Sons for the Dutch Academy of Sciences. Pp. xi + 415 plates. 1969. Df. 50. British Journal for the History of Science, 5(2):207, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025437.

Crosland:1971:BRC

[Cro71] Maurice Crosland. Book review: Chemistry Essays Physical and Chemical. By Antoine Lavoisier. Trans. Thomas Henry. Second Edition. Introduction by Frank Greenaway. London: F. Cass. 1970. Pp. xxxiii + xxxii + 475. £9.45. British Journal for the History of Science, 5(4):405–406, December 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025389.

Crowther:1972:BRG

[Cro72] J. G. Crowther. Book review: General Science at the Crossroads. Papers from the Second International Congress of the History of Science and Technology, 1931. By N. I. Bukharin et al. Foreword by Joseph Needham. Introduction by P. G. Werskey. London: Frank Cass, 1971. Pp. xxx + 235. £2.75. British Journal for the History of Science, 6 (1):85–86, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025267.

Crosland:1973:HDA

[Cro73] Maurice Crosland. Humphry Davy — an alleged case of suppressed publication. British Journal for the History of Science, 6(3):304–310, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025446.

Crowther:1974:CLa

[Cro74a] J. G. (James Gerald) Crowther. The Cavendish Laboratory, 1874–1974. MacMillan Publishing Company, New York, NY, USA, 1974. ISBN 0-333-14166-0. xvi + 464 pp. LCCN QC51.G72 C353 1974b.

Crowther:1974:CLb

[Cro74b] J. G. (James Gerald) Crowther. The Cavendish Laboratory, 1874–1974. Science History Publications, New York, NY, USA, 1974. ISBN 0-88202-029-3 (hardcover), 1-349-01670-5 (e-book). xvi + 464 pp. LCCN QC51.G72 C353 1974.

Crosland:1977:HSN

[Cro77] Maurice Crosland. History of science in a national context. British Journal for the History of Science, 10(2):95– 113, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025864.

Crosland:1979:BRN

[Cro79] Maurice Crosland. Book review: Nineteenth and Twentieth Centuries: Le chimiste Claude-Louis Berthollet, 1748– 1822: Sa Vie, son Oeuvre. By Michelle Sadoun-Goupil. Paris: Librairie Philosophique J. Vrin, 1977. Pp. xv + 392 + VIII plates. 126 francs. British Journal for the History of Science, 12(2):230, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025764.

Crosland:1983:PML

[Cro83] Maurice Crosland. Priestley Memorial Lecture: A practical perspective on Joseph Priestley as a pneumatic chemist. British Journal for the History of Science, 16(3):223– 238, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026144.

Crombie:1984:OJS

[Cro84a] A. C. Crombie. Obituary: James Sterling Wilkie: 1906– 1982. British Journal for the History of Science, 17(3):310– 311, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026625.

Crook:1984:BKP

[Cro84b] D. P. (David Paul) Crook. Benjamin Kidd: portrait of a social Darwinist. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25804-9. vii + 460 pp. LCCN HM106.K52 C76

                                        1984. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0903/83019009-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0903/83019009-t.html.

Crosland:1985:BRD

[Cro85] Maurice Crosland. Book review: Dorinda Outram, Georges Cuvier. Vocation, science and authority in post-revolutionary France. Manchester: Manchester University Press, 1984. pp viii + 299. ISBN 0-7190-1077-2. £25. British Journal for the History of Science, 18 (2):249–250, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026352.

Crowe:1986:ELD

[Cro86] Michael J. Crowe. The extraterrestrial life debate, 1750– 1900: the idea of a plurality of worlds from Kant to Lowell. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-26305-0. xix + 680 pp. LCCN QB54 .C76

1986. URL http://www.loc.
gov/catdir/description/
cam032/85007842.html.

Crosland:1987:BRH

[Cro87a] Maurice Crosland. Book review: Harry W. Paul. From Knowledge to Power. The Rise of the Science Empire in France, 1860– 1939. Cambridge: Cambridge University Press, 1985. Pp. ix + 415. ISBN 0-521-25404-5. £32.50, $49.50. British Journal for the History of Science, 20(1):105–107, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026460.

Crosland:1987:IST

[Cro87b] Maurice Crosland. The image of science as a threat: Burke versus Priestley and the ‘philosophic revolution’. British Journal for the History of Science, 20(3):277–307, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026359.

Crosland:1988:BRC

[Cro88] Maurice Crosland. Book review: Comité Lavoisier De L’Académie Des Sciences. Ouvres de Lavoisier: Correspondence, Fascicule IV, 1784–1786. Paris: Editions Belin, 1986. Pp. xv + 351. ISBN 2-7011-1085-8. FF 460.00. British Journal for the History of Science, 21(3): 365–366, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026836.

Crosland:1989:BRA

[Cro89a] Maurice Crosland. Book review: Arthur Donovan (ed.). The Chemical Revolution. Essays in Reinterpretation. Osiris, 2nd series, vol. iv (1988). Pp. 236. Philadelphia: History of Science Society, University of Pennsylvania. ISBN 0-934235-11-2. $15. (paperback). British Journal for the History of Science, 22(4):458– 459, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026934.

Cross:1989:BRN

[Cro89b] Stephen J. Cross. Book review: Nicole Hahn Rafter (ed.). White Trash: The Eugenic Family Studies, 1877– 1919. Boston: Northeastern University Press, 1988. Pp. x + 382. ISBN 1-55553-030-3. £38.00. British Journal for the History of Science, 22(4): 456–457, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026932.

Crombie:1990:SOM

[Cro90a] A. C. (Alistair Cameron)

Crombie. Science, optics, and music in medieval and early modern thought. Hambledon Press, London [England], 1990. ISBN 0-907628-79-6. xxii + 474 pp. LCCN Q125 .C686 1990. US$55.00. URL http://www.

loc.gov/catdir/
description/hol032/
90030933.html.

Crowe:1990:TWA

[Cro90b] Michael J. Crowe. Theories of the world from antiquity to the Copernican revolution. Dover Publications, Inc., New York, NY, USA, 1990. ISBN 0-486-26173-5. ix + 229 pp. LCCN QB15 .C77 1990. US$5.95.

Cronin:1991:APA

[Cro91a] Helena Cronin. The ant and the peacock: altruism and sexual selection from Darwin to today. Press Syndicate of the University of Cambridge, Cambridge, UK, 1991. ISBN 0-521-32937-X (hardcover). xiv + 490 pp. LCCN QL761 .C76

1991. URL http://www.loc.

gov/catdir/description/
cam032/91007887.html;
http://www.loc.gov/
catdir/toc/cam032/
91007887.html.

Crowcroft:1991:EEH

[Cro91b] Peter Crowcroft. Elton’s ecologists: a history of the Bureau of Animal Population. University of Chicago Press, Chicago, IL, USA, 1991. ISBN 0-226-12146-1, 0-226-12148-8 (paperback). xx + 177 pp. LCCN QL69.G72 U553

1991. URL http://www.loc.
gov/catdir/description/
uchi052/90039469.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/90039469-t.html.

Crosland:1992:SUC

[Cro92] Maurice P. Crosland. Science under control: the French Academy of Sciences, 1795– 1914. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41373-7. xvii + 454 pp. LCCN Q46 .C77

1992. URL http://www.loc.
gov/catdir/description/
cam024/91028748.html;
http://www.loc.gov/
catdir/samples/cam031/
91028748.html;   http://
www.loc.gov/catdir/toc/
cam028/91028748.html.

Crombie:1996:SAN

[Cro96] A. C. (Alistair Cameron) Crombie. Science, art, and nature in medieval and modern thought. Hambledon Press, London, UK, 1996. ISBN 1-85285-067-1 (hardcover). xv + 516 pp. LCCN Q127.E8 C74

1996. URL http://www.loc.
gov/catdir/description/
hol032/95049964.html.

Cruickshank:1979:SO

[Cru79] A. D. Cruickshank. Soddy at Oxford. British Journal for the History of Science, 12(3): 277–288, November 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026002.

Crump:1990:AN

[Cru90] Thomas Crump. The anthropology of numbers, volume 70 of Cambridge studies in social anthropology. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-38045-6. x + 201 pp. LCCN GR933 .C78

1990. URL http://www.loc.
gov/catdir/description/
cam032/89034123.html;
http://www.loc.gov/
catdir/toc/cam031/
89034123.html.

Chertok:1992:CPR

[CS92] Léon Chertok and Isabelle Stengers. A critique of psychoanalytic reason: hypnosis as a scientific problem from Lavoisier to Lacan. Stanford University Press, Stanford, CA, USA, 1992. ISBN 0-8047-1950-0. xxvi + 319 pp. LCCN RC497 .C5313 1992. US$35.00. URL http://www.

loc.gov/catdir/

                                        description/cam024/
                                        91022388.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam027/91022388.html.

Chambers:1994:VNH

[CS94]  Robert  Cham-

bers and James A. Secord. Vestiges of the natural history of creation and other evolutionary writings. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-10072-3, 0-226-10073-1 (paperback). xlviii, vi, 390, 254 pp. LCCN QH363 .C4

                                        1994. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi051/93048415.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0608/93048415-t.html.

Cremo:1993:FAH

[CT93]  Michael  A.

Cremo and Richard L. Thompson. Forbidden archeology: the hidden history of the human race. Bhaktivedanta Institute, San Diego, 1993. ISBN 0-9635309-8-4. xxxviii + 914 pp. LCCN GN741 .C74 1993. US$39.95 (US$49.95 Can.).

Corbey:1995:AMA

[CT95] Raymond Corbey and Bert Theunissen, editors. Ape, man, apeman, changing views since 1600: evaluative proceedings of the symposium Ape, man, apeman: changing views since 1600, Leiden, the Netherlands, 28 June–1 July, 1993. Dept. of Prehistory, Leiden University, Leiden, The Netherlands, 1995. ISBN 90-73368-05-7. LCCN GN280.7 .A64 1995; GN281.4 .A33 1995.

Cunningham:1978:BRS

[Cun78] Adrian Cunningham. Book review: Science and Literature ‘Kubla Khan’ and The Fall of Jerusalem. The Mythological School in Biblical Criticism and Secular Literature, 1770–1880. By E. S. Shaffer. Cambridge: Cambridge University Press, 1975. Pp. x + 361. £9.50. British Journal for the History of Science, 11 (2):177–178, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025736.

Cunningham:1983:BRB

[Cun83a] Andrew Cunningham. Book review: Biography Martin Kemp, Leonardo da Vinci: the marvellous works of nature and man, London: J. M. Dent, 1981. Pp.

384. £14.95. British Journal for the History of Science, 16 (1):109–110, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026113.

Cunningham:1983:BRM

[Cun83b] Andrew Cunningham. Book review: Medical Sciences: Nancy G. Siraisi, Taddeo Alderotti and his pupils: two generations of Italian medical learning. Princeton: Princeton University Press, 1981. Pp. xxiii + 461. British Journal for the History of Science, 16(1): 84–86, March 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026097.

Cunningham:1988:BRR

[Cun88] Andrew Cunningham. Book review: Ronald L. Numbers (ed.). Medicine in the New World-New Spain, New France and New England. Knoxville: The University of Tennessee Press, 1987. Pp. 175. ISBN 0-87049-517-8. $18.95. British Journal for the History of Science, 21 (3):377–378, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026849.

Cunningham:1989:BRN

[Cun89] Andrew Cunningham. Book review: Nancy G. Siraisi. Avicenna in Renaissance Italy. The ‘Canon’ and Medical Teaching in Italian Universities after 1500. Princeton, New Jersey: Princeton University Press, 1987. Pp. 410. ISBN 0-691-05137-2. £31.40. British Journal for the History of Science, 22(1):85–86, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026685.

Cunningham:1997:ARR

[Cun97] Andrew Cunningham. The anatomical renaissance: the resurrection of the anatomical projects of the ancients. Scolar Press, Aldershot, England, 1997. ISBN 1-85928-338-1. xiv + 283 pp. LCCN QM11 .C86 1997.

Chaldecott:1969:RC

[CW69] J. A. Chaldecott and G. J.

Whitrow. Report of Council. British Journal for the History of Science, 4(4):428– 430, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025259.

Cattermole:1987:HDS

[CW87] M. J. G. (Michael J. G.) Cattermole and A. F. (Arthur F.) Wolfe. Horace Darwin’s shop: a history of the Cambridge Scientific Instrument Company, 1878 to 1968. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-569-9. xvi + 285 pp. LCCN Q185 .C365 1987. US$29.50. URL http://www.

loc.gov/catdir/
enhancements/fy0745/
87121867-d.html.

Dahl:1997:FCR

[Dah97] Per F. Dahl. Flash of the cathode rays: a history of J. J. Thomson’s electron. Institute of Physics Publishers, Bristol, UK, 1997. ISBN 0-7503-0453-7 (hardcover), 0-585-20853-0 (ebook). xvii + 526 pp. LCCN 99.E01389.

Danziger:1990:CSH

[Dan90] Kurt Danziger. Constructing the subject: historical origins of psychological research. Cambridge Studies in the History of Psychology. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-36358-6 (hardcover), 0-521-46785-3 (paperback). vii + 254 pp. LCCN BF181 .D27 1990.

Daston:1988:CPE

[Das88] Lorraine Daston. Classical probability in the Enlightenment. Princeton University Press, Princeton, NJ, USA, 1988. ISBN 0-691-08497-1. xviii + 423 pp. LCCN QA273.A4 D37 1988. US$55.00. URL

http://www.loc.gov/
catdir/description/
prin031/88009869.html.

Daub:1974:RPS

[Dau74] Edward E. Daub. The regenerator principle in the Stirling and Ericsson hot air engines. British Journal for the History of Science, 7(3):259– 277, November 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025533.

Daub:1975:BRN

[Dau75] Edward E. Daub. Book review: Nineteenth and Twentieth Centuries: Julius Robert Mayer. Prophet of Energy. By Robert Bruce Lindsay. London: Pergamon Press, 1973. Pp. viii + 238. £2.75. British Journal for the History of Science, 8(1):88–89, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025833.

Dauben:1995:ARC

[Dau95] Joseph Warren Dauben. Abraham Robinson: the creation of nonstandard analysis: a personal and mathematical odyssey. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03745-0. xix + 559 pp. LCCN QA29.R57

D38 1995. URL http://www.
loc.gov/catdir/
description/prin031/
94032715.html;   http://
www.loc.gov/catdir/toc/
prin031/94032715.html.

Daum:1998:WIJ

[Dau98]  Andreas  W.

Daum. Wissenschaftspopularisierung im 19. Jahrhundert: bürgerliche Kultur, naturwissenschaftliche Bildung und die deutsche Öffentlichkeit, 1848– 1914. (German) [Science popularization in the 19th Century: civil culture, scientific education and the German public]. R. Oldenbourg, München, Germany, 1998. ISBN 3-486-56337-8. xii + 617 pp. LCCN QH149 .D39 1998.

                                        URL http://www.h-net.org/
                                        review/hrev-a0b0s0-aa.

Davies:1971:BRG

[Dav71] Gordon L. Davies. Book review: Geology and Biology A Source Book in Geology, 1400– 1900. By Kirtley F. Mather and Shirley L. Mason. Harvard University Press & Oxford University Press. 1970. Pp. xxii + 702. £6. British Journal for the History of Science, 5(4): 407–408, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025392.

Davies:1981:BRT

[Dav81] Celia Davies. Book review: Twentieth Century The Division in British Medicine: A History of the Separation of General Practice from Hospital Care, 1911–1968. By Frank Honigsbaum. London: Kogan Page, 1979, pp. xvi + 445. £6.95. British Journal for the History of Science, 14(1):99– 101, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026085.

Davies:1989:WHN

[Dav89] Mansel Davies. William Hampson (1854–1926): A note. British Journal for the History of Science, 22(1):63–73, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026679.

Davis:1995:SMS

[Dav99] E. A. (Edward Arthur) Davis, editor. Science in the making: scientific development as chronicled by historic papers in the Philosophical Magazine, with commentaries and illustrations. Taylor and Francis, Boca Raton, FL, USA, 1995–1999. ISBN 0-7484-0219-5 (vol. 1), 0-7484-0642-5 (vol. 2), 0-7484-0766-9 (vol. 3), 0-7484-0767-7 (vol. 4). ???? pp. LCCN Q125 .S39

1995. URL http://www.loc.
gov/catdir/enhancements/
fy0745/95201153-d.html.

Day:1985:RHS

[Day85] L. R. Day. Resources for the history of science in the science museum library. British Journal for the History of Science, 18(1):71–76, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026266.

deChamps:2009:BFF

[dCC09] Emmanuelle de Champs and Jean-Pierre Cléro, editors. Bentham et la France: fortune et infortunes de l’utilitarisme. (French) [Bentham and France: fortune and misfortune of utilitarianism], volume 2009:09 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2009.

ISBN 0-7294-0980-5. xv + 311 pp. LCCN JC223.B5 B46 2009.

Dirac:1995:CWP

[DD95] P. A. M. (Paul Adrien Maurice) Dirac and R. H. (Richard Henry) Dalitz. The collected works of P. A. M. Dirac, 1924– 1948. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-36231-8. xxiii + 1310 pp. LCCN QC21.2 .D57

                                        1995. URL http://www.loc.
                                        gov/catdir/description/
                                        cam026/95001010.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy1006/95001010-t.html.

Darwin:1990:CDM

[DDG90] Charles Darwin, Mario A. Di Gregorio, and N. W. (Nick W.) Gill, editors. Charles Darwin’s marginalia, volume 783 of Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1990. ISBN 0-8240-6639-1 (vol. 1). ???? pp. LCCN QH365 .A1 1990.

DeClercq:1985:NCS

[De 85] P. R. (Peter R.) De Clercq, editor. Nineteenth-century scientific instruments and their makers: papers presented at the Fourth Scientific Instrument Symposium, Amsterdam, 23–26 October 1984, volume 17 of Nieuwe Nederlandse bijdragen tot de geschiedenis der geneeskunde en der natuurwetenschappen; Communication 221 of the National Museum for the History of Science and Medicine “Museum Boerhaave”. Rodopi, Amsterdam and Leiden, The Netherlands, 1985. ISBN 90-6292-070-5 (paperback), 90-6203-917-0. ISSN 0168-9827. LCCN Q184 .S44 1984. fl 70.00.

Dean:1983:BRLa

[Dea83a] Colin Dean. Book review:

Lives and Works Hugo van Woerden, Willem N. Brouw, and Henk C. van de Hulst (eds.), Oort and the universe: a sketch of Oort’s research and person. Dordrecht, Boston & London: D. Reidel, 1980. Pp. viii + 210. Dfl.55.00/ $29.00; Dfl.25.00/$12.95 (paperback). British Journal for the History of Science, 16(2): 204–205, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026246.

Dean:1983:BRLb

[Dea83b] John Dean. Book review: Life and earth sciences: S. M. Walters, The shaping of Cambridge botany: a short history of whole plant botany in Cambridge from the time of Ray into the present century. Cambridge: Cambridge University Press, 1981. Pp. ix + 121. £17.50. British Journal for the History of Science, 16(3): 285–286, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026155.

Dear:1986:BRB

[Dea86a] Peter Dear. Book review: Brian Vickers, ed. Occult and Scientific Mentalities in the Renaissance. Cambridge: Cambridge University Press, 1984. Pp. xiv + 408. ISBN 0-521-25879-0. £27.50, $39.50. British Journal for the History of Science, 19(2):201–203, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026592.

Dear:1986:BRJ

[Dea86b] Peter Dear. Book review: John G. Burke, ed. The Uses of Science in the Age of Newton. Berkeley and London: University of California Press, 1984. Pp. xxii + 204. ISBN 0-520-04970-5. £17.30. British Journal for the History of Science, 19(2):203–204, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026593.

Dear:1986:BRW

[Dea86c] Peter Dear. Book review: William A. Wallace. Galileo and his Sources. The Heritage of the Collegio Romano in Galileo’s Science. Princeton: Princeton University Press, 1984. Pp. xiv + 371. ISBN 0-691-08355-X.

£45.80. British Journal for the History of Science, 19(3): 363–364, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026535.

Dear:1988:MLS

[Dea88]  Peter  Robert

Dear, editor. Mersenne and the learning of the schools, Cornell history of science series, 99-0467310-1. Cornell University Press, Ithaca, NY, USA, 1988. ISBN 0-8014-1875-5. LCCN BX4705.M53 D43 1988.

Dear:1991:LSS

[Dea91] Peter Dear, editor. The literary structure of scientific argument: historical studies. University of Pennsylvania Press, Philadelphia, PA, USA, 1991. ISBN 0-8122-8185-3. vi + 211 pp. LCCN Q126.8 .L58 1991.

Debus:1966:BRC

[Deb66] Allen G. Debus. Book review:

Chemical, Medical and Pharmaceutical Books printed before 1800 in the Collections of the University of Wisconsin Libraries. Ed. John Neu. Pp. viii + 280. Madison and Milwaukee: The University of Wisconsin Press, 1965. British Journal for the History of Science, 3(2):191–192, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025034.

Debus:1984:OWP

[Deb84]  Allen G. Debus.  Obitu-

ary: Walter Pagel, M.D. — 12 November 1898–25 March 1983. British Journal for the History of Science, 17(1):83– 84, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026020.

Debus:1991:FPC

[Deb91] Allen G. Debus. The French Paracelsians: the chemical challenge to medical and scientific tradition in early modern France. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-40049-X. xvi + 247 pp. LCCN QD18.F8 D43

                                        1991. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/90026163.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        90026163.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam028/90026163.html.

Degler:1991:SHN

[Deg91] Carl N. Degler. In search of human nature: the decline and revival of Darwinism in American social thought. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-506380-5, 0-19-507707-5 (paperback). x + 400 pp. LCCN HM; HM106 DEG 1991; 300; 301; HM106.DEG.

Deichmann:1992:BHV

[Dei92]  Ute Deichmann.  Biologen

unter Hitler: Vertreibung, Karrieren, Forschung. (German) [Biologists under Hitler: expulsion, careers, research]. Campus, Frankfurt, Germany, 1992. ISBN 3-593-34763-6. 370 pp. LCCN QH305.2.G3 D45 1992. DM38.00.

DeKosky:1973:SEL

[DeK73] Robert K. DeKosky. Spectroscopy and the elements in the late Nineteenth Century: The work of Sir William Crookes. British Journal for the History of Science, 6(4): 400–423, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025503.

Dempster:1983:PMN

[Dem83]  W. J. Dempster.  Patrick

Matthew and natural selection: Nineteenth Century gentlemanfarmer, naturalist, and writer. P. Harris, Edinburgh, Scotland, 1983. ISBN 0-86228-065-6. 156 + 4 pp. LCCN QH31.M34 D45 1983. US$8.95.

Desmond:1982:AA

[Des82a] Adrian Desmond. Archetypes and ancestors. Blond and Briggs, London, UK, 1982. ISBN 0-85634-121-5. 287 pp. LCCN ????

Desmond:1982:AAP

[Des82b] Adrian J. Desmond. Archetypes and ancestors: Palaeontology in Victorian London, 1850– 1875. Blond and Briggs, London, UK, 1982. ISBN 0-85634-121-5. 287 pp. LCCN QE705.G7

                                        D4. URL http://catalog.
                                        hathitrust.org/api/
                                        volumes/oclc/9283761.html.

Desmond:1982:IM

[Des82c] Ray Desmond. The India Museum, 1801–1879. Her Majesty’s Stationary Office, London, England, 1982. ISBN 0-11-580088-3. xv + 215 pp. LCCN AM101.L715 D47 1982. US$25.00.

Desmond:1984:AAP

[Des84] Adrian J. Desmond. Archetypes and ancestors: palaeontology in Victorian London 1850–1875. Blond and Briggs, London, UK, 1984. ISBN 0-85634-121-5. 287 pp. LCCN ????

Desmond:1985:ROR

[Des85] Adrian Desmond. Richard Owen’s reaction to transmutation in the 1830’s. British Journal for the History of Science, 18(1):25–50, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026264.

Desmond:1988:BRP

[Des88] Adrian Desmond. Book review: Paul H. Barret, Peter J. Gautrey, Sandra Herbert, David Kohn & Sydney Smith (eds). Charles Darwin’s Notebooks, 1836–1844. London, Cambridge: British Museum (Natural History)/Cambridge University Press, 1987. Pp. viii + 747. ISBN 0-521-35055-7. £65.00. British Journal for the History of Science, 21(4): 495–496, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026967.

Desmond:1989:PEM

[Des89a]  Adrian J. Desmond.  The

politics of evolution: morphology, medicine, and reform in radical London. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1989. ISBN 0-226-14346-5. x + 503 pp. LCCN QL810 .D47

1989. URL http://www.loc.
gov/catdir/description/
uchi052/89005137.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/89005137-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/89005137-b.html.

Dessi:1989:LCD

[Des89b] Paola Dessì. L’ordine e il caso: discussioni epistemologiche e logiche sulla probabilità da Laplace a Peirce. (Italian) [The order and the case: epistemological and logical arguments on probability from Laplace to Peirce]. Ricerca. Mulino, Bologna, Italy, 1989. ISBN 88-15-02088-8. 229 pp. LCCN BC141 .D44 1991.

Desmond:1992:EDI

[Des92]  Ray Desmond.  The Eu-

ropean discovery of the Indian flora. Clarendon Press, New York, NY, USA, 1992. ISBN 0-19-854684-X. xii + 355 pp. LCCN QK358 .D44 1992. US$60.00 (US$120.00

                                        U.S.). URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0638/91044110-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0638/91044110-t.html.

Desmond:1994:HDA

[Des94] Adrian J. Desmond. Huxley: the devil’s disciple. M. Joseph, London, UK, 1994. ISBN 0-7181-3641-1. xvii + 475 pp. LCCN Q143.H956 D47 1994.

DeVorkin:1982:HMA

[DeV82] David H. DeVorkin. The history of modern astronomy and astrophysics: a selected, annotated bibliography, volume 1; vol. 304 of Bibliographies of the history of science and technology; Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1982. ISBN 0-8240-9283-X. xxvii + 434 pp. LCCN Z5154.H57 D48 1982. US$65.00.

Dewhurst:1963:PRS

[Dew63] Kenneth Dewhurst. Prince Rupert as a scientist. British Journal for the History of Science, 1(4):365–373, December 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024908.

Dee:1998:DJD

[DF98] John Dee and Edward Fenton.

The diaries of John Dee. Day Books, Charlbury, Oxfordshire, UK, 1998. ISBN 0-9532213-0-X. xi + 369 pp. LCCN MLCM 2004/02973 (B).

Desargues:1987:GWG

[DFG87]  Gérard  De-

sargues, Judith Veronica Field, and Jeremy Gray. The geometrical work of Girard Desargues. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96403-7. x + 237 pp. LCCN QA485 .D47 1987. US$64.00.

deGandt:2001:CVI

[dG01] François de Gandt, editor.

Cirey dans la vie intellectuelle: la réception de Newton en France. (French) [Cirey in intellectual life: the reception of Newton in France], volume 2001:11 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2001. ISBN 0-7294-0771-3. v + 253 pp. LCCN B2177 .C57 2001.

deGandt:1989:LTS

[dGB^{+}89]  François  de  Gandt,

Armand Beaulieu, et al., editors. L’œuvre de Torricelli: science galiléenne et nouvelle géométrie. (French) [The work of Torricelli: gallic science and new geometry], volume 1re sér., no 32 (1987) of Publications de la Faculté des lettres et sciences humaines de Nice. Diffusion, Les Belles Lettres, Paris, France, 1989. ISBN 2-252-62032-X. 247 pp. LCCN QA444 .O37 1989.

Davenport-Hines:1990:SDP

[DH90] R. P. T. (Richard Peter Treadwell) Davenport-Hines. Sex, death, and punishment: attitudes to sex and sexuality in Britain since the Renaissance. Collins, London, UK, 1990. ISBN 0-00-217956-3. xv + 439 pp. LCCN HQ18.G78 D38 1990. US$20.00.

Davenport-Hines:1992:GH

[DHS92] R. P. T. (Richard Peter Treadwell) Davenport-Hines and Judy Slinn. Glaxo: a history to 1962. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41539-X. xiv + 406 + 4 pp. LCCN HD9667.9.G53

D38 1992. URL http://www.
loc.gov/catdir/
description/cam024/
91042204.html;   http://
www.loc.gov/catdir/toc/
cam028/91042204.html.

Dick:1982:PWO

[Dic82] Steven J. Dick. Plurality of worlds: the origins of the extraterrestrial life debate from Democritus to Kant. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24308-4. x + 246 pp. LCCN QB54

.D5. URL http://www.loc.
gov/catdir/enhancements/
fy0803/81010165-d.html.

Dijksterhuis:1990:CS

[Dij90] E. J. (Eduard Jan) Dijksterhuis. Clio’s stiefkind. (Dutch) [Clio’s stepchild]. B. Bakker, Amsterdam, The Netherlands, 1990. ISBN 90-351-0786-1. 321 pp. LCCN Q125 .D47 1990. fl 45.00.

Dingle:1963:BRP

[Din63a] Herbert Dingle. Book review:

Physical Science: Men and Concepts. By Guy C. Omer Jr., Harold L. Knowles, Belvey W. Mundy and W. Herbert Yoho, University of Florida. Boston: D. C. Heath and Company. Pp. x + 601. 1963. British Journal for the History of Science, 1(4):386–387, December 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024913.

Dingle:1963:HYS

[Din63b] Herbert Dingle. A hundred years of spectroscopy. British Journal for the History of Science, 1(3):199–216, June 1963.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024920.

Dingle:1964:BRA

[Din64] Herbert Dingle. Book review: Alexander von Humboldt. By L. Kellner. Pp. 247. Oxford University Press, 1963. £1 5s. British Journal for the History of Science, 2(2): 171, December 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025018.

Dingle:1966:TCC

[Din66] Herbert Dingle. Twentieth-Century cosmology. British Journal for the History of Science, 3(2):184–187, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025029.

Dingle:1977:RAS

[Din77] Herbert Dingle. Review: An apology: Science at the crossroads. British Journal for the History of Science, 10(1): 94, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025601.

David:1988:CCV

[DJS97] Andrew David, Rüdiger Joppien, and Bernard Smith, editors. The charts and coastal

views  of  Captain

Cook’s voyages, volume extra ser., no. 43–46 of Hakluyt Society, 0951-6158. Hakluyt Society in association with the Australian Academy of the Humanities, London, UK, 1988–1997. ISBN 0-904180-23-9 (vol. 1), 0-904180-31-X (vol. 2), 0-904180-55-7 (vol. 3). ???? pp. LCCN G1036 .C57 1988.

Darwin:1988:CDB

[DK88] Charles Darwin and R. D.

Keynes. Charles Darwin’s Beagle diary. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-23503-0. xxix + 464 pp. LCCN QH365 .A1

1988. URL http://www.loc.
gov/catdir/description/
cam023/87031169.html;
http://www.loc.gov/
catdir/enhancements/
fy0732/87031169-b.html;
http://www.loc.gov/
catdir/toc/cam025/
87031169.html.

LaMettrie:1996:MMO

[dLT96] Julien Offray de La Mettrie and Ann Thomson. Machine man and other writings. Cambridge texts in the history of philosophy. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-47258-X (hardcover), 0-521-47849-9 (paperback). xxx + 179 pp. LCCN B2062 .E5

1996. URL http://www.loc.
gov/catdir/description/
cam027/95014609.html;
http://www.loc.gov/

                                        catdir/samples/cam034/
                                        95014609.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam024/95014609.html.

Desmond:1991:D

[DM91] Adrian J. Desmond and James R. (James Richard) Moore. Darwin. Michael Joseph, London, UK, 1991. ISBN 0-7181-3430-3. xxi + 807 pp. LCCN QH31.D2 D47 1991.

Day:1996:BDH

[DM96] Lance Day and Ian McNeil, editors. Biographical dictionary of the history of technology. Routledge reference. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1996. ISBN 0-415-06042-7. xiii + 844 pp. LCCN T39 .B49

                                        1996. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0648/95026250-d.html.

Dobbs:1991:JFG

[Dob91] Betty Jo Teeter Dobbs. The Janus faces of genius: the role of alchemy in Newton’s thought. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-38084-7. xii + 359 pp. LCCN QC16.N7 D64

                                        1991. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/91008695.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        91008695.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam026/91008695.html.

Doel:1996:SSA

[Doe96] Ronald Edmund Doel. Solar system astronomy in America: communities, patronage, and interdisciplinary science, 1920–1960. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-41573-X (hardcover). xiv + 280 pp. LCCN QB501 .D64

1996. URL http://www.loc.
gov/catdir/description/
cam027/95023112.html;
http://www.loc.gov/
catdir/toc/cam023/
95023112.html.

Dolby:1969:BRN

[Dol69] R. G. A. Dolby. Book review: Nineteenth and Twentieth Centuries: Records of the Australian Academy of Science, Vol. 1. Canberra. No. 1. 1966. Pp. 69. No. 2. 1967. Pp. 154. $A1.50 per issue. British Journal for the History of Science, 4(3):302–303, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025495.

Dolby:1970:BRP

[Dol70a] R. G. A. Dolby. Book review:

Philosophy of Science Discovery in the Physical Sciences. By Richard J. Blackwell. Notre Dame and London: University of Notre Dame Press. Pp. xii + 240. 1969. 81s. British Journal for the History of Science, 5(2):187, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025418.

Dolby:1970:BRT

[Dol70b] R. G. A. Dolby. Book review: Twentieth Century Scientific Thought, 1900–1960: A Selective Survey. Ed. by R. Harré. London: Clarendon Press: Oxford University Press. 1969. Pp. vii + 277. 44 figs. and 8 plates. 65s. British Journal for the History of Science, 5 (1):105–106, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025374.

Dolby:1971:BRN

[Dol71a] R. G. A. Dolby. Book review: Nineteenth Century: Hegel’s Philosophy of Nature: Being Part Two of the Encyclopedia of the Philosophical Sciences (1830). Translated from Nicolin and Pöggeler’s Edition (1959) and from the Zusätze in Michelet’s Text (1847). By A. V. Miller. Foreword by J. N. Findlay. Oxford: Clarendon Press: Oxford University Press. 1970. Pp. xxxi + 450. £3.75. British Journal for the History of Science, 5 (3):314–315, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025345.

Dolby:1971:BRP

[Dol71b] R. G. A. Dolby. Book review:

Philosophy of Science Criticism and the Growth of Knowledge. Proceedings of the International Colloquium in the Philosophy of Science, London 1965, Volume 4. Ed. by Imre Lakatos and Alan Musgrave. London: Cambridge University Press. Pp. viii + 282. 1970. £3.50. British Journal for the History of Science, 5 (4):400, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025383.

Dolby:1973:BRGa

[Dol73a] R. G. A. Dolby. Book review:

General: English Men of Science: Their Nature and Nurture (London, 1874). By Francis Galton. Second edition with Introduction by Ruth Schwartz Cowan. London: Frank Cass, 1970. Pp xi + 270. £3.75. The Scientific Basis of National Progress including that of Morality (London, 1899). By G. Gore. Reprint. London: Frank Cass, 1970. Pp.

221. £3.50. British Journal for the History of Science, 6 (3):315, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025449.

Dolby:1973:BRGb

[Dol73b] R. G. A. Dolby. Book review: General Philosophy, Science and Method. Essays in Honour of Ernest Nagel. Ed. by Sidney Morgenbesser, Patrick Suppes, and Morton White. London: Macmillan, 1969. Pp. x + 613. £5.50. British Journal for the History of Science, 6(4):434–435, December 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025507.

Dolby:1974:BRP

[Dol74] R. G. A. Dolby. Book review: Philosophy Foundations of Scientific Method: The Nineteenth Century. Ed. by Ronald N. Giere and Richard S. Westfall. Bloomington: Indiana University Press, 1973. Pp. x + 306. $10. British Journal for the History of Science, 7(3):287–288, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025537.

Dolby:1983:BRS

[Dol83] R. G. A. Dolby. Book review: Science and Society H. M. Collins and T. J. Pinch, Frames of meaning: the social construction of extraordinary science. London: Routledge and Kegan Paul, 1982. Pp. x + 210. £12.50. British Journal for the History of Science, 16(3):308–309, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026176.

Dolby:1996:UKI

[Dol96] R. G. A. (Riki G. A.) Dolby.

Uncertain knowledge: an image of science for a changing world. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56004-7. xi + 365 pp. LCCN Q175 .D688

1996. URL http://www.loc.
gov/catdir/description/
cam027/96022588.html;
http://www.loc.gov/
catdir/samples/cam031/
96022588.html;   http://
www.loc.gov/catdir/toc/
cam022/96022588.html.

Donovan:1988:CRE

[Don88] Arthur L. Donovan, editor. The chemical revolution: essays in reinterpretation, volume 2nd ser., vol. 4 (1988) of Osiris, 0369-7827. History of Science Society, Philadelphia, PA, 1988. ISBN 0-934235-10-4, 0-934235-11-2 (paperback). 236 pp. LCCN Q125 .O76 2nd ser., vol. 4, 1988.

Donovan:1993:ALS

[Don93]  Arthur Donovan.  Antoine

Lavoisier: science, administration, and revolution. Blackwell science biographies. Blackwell, Oxford, UK, 1993. ISBN 0-631-17887-2. xv + 351 pp. LCCN QD22.L4 D58 1993.

Dorling:1975:BRN

[Dor75]  Jon  Dorling.

Book review: Nineteenth and Twentieth Centuries: The History of Quantum Theory. By Friedrich Hund. Trans. by Gordon Reece. London: Harrap, 1974. Pp. 260. £6.10. British Journal for the History of Science, 8(2):188–189, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025657.

Dorling:1979:BRN

[Dor79]  Jon  Dor-

ling. Book review: Nineteenth and Twentieth Centuries: Niels Bohr Collected Works. Volume iii: The Correspondence Principle (1918–1923). Edited by J. Rud Nielsen. Amsterdam: North-Holland, 1976. Pp. xii + 702. $105.95. British Journal for the History of Science, 12 (1):100–102, March 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025674.

Dorling:1980:BRT

[Dor80]  Jon  Dor-

ling. Book review: Twentieth Century History of Twentieth-Century Physics. Proceedings of the International School of Physics ‘Enrico Fermi’, Course 57. Edited by C. Weiner. New York and London: Academic Press, 1977. Pp. ix + 457. £28.00. British Journal for the History of Science, 13(2): 179–180, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025911.

Dorn:1991:GS

[Dor91] Harold Dorn. The geography of science. The Johns Hopkins University Press, Baltimore, MD, USA, 1991. ISBN 0-8018-4151-8. xx + 219 pp. LCCN Q125 .D65 1991.

Dougherty:1984:BRE

[Dou84] F. W. P. Dougherty. Book review: Enlightenment François Duchesneau. La physiologic des lumières. Empirisme, modèles et théories (International Archives of the History of Ideas; 95). The Hague: Martinus Nijhoff, 1982. Pp. xxi + 611. ISBN 90-247-2500-3. Cloth, Dfl. 185.00, US 397.00. British Journal for the History of Science, 17(1): 119–120, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026056.

Douglas:1988:HST

[Dou88] Gina L. Douglas. History of science and technology resources at the Linnean Society of London. British Journal for the History of Science, 21(4):489– 493, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026965.

Dutton:1988:WTD

[DPP88]  Diana  Barbara  Dut-

ton, Thomas A. Preston, and Nancy E. Pfund. Worse than the disease: pitfalls of medical progress. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-34023-3. xvi + 528 pp. LCCN RA418.5.M4 D88

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/87034203.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam023/
                                        87034203.html.

Dershowitz:1997:CC

[DR97] Nachum Dershowitz and Edward M. Reingold. Calendrical calculations. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56413-1 (hardcover), 0-521-56474-3 (paperback). xxi + 307 pp. LCCN CE12 .D47 1997. URL http:

                                        //www.loc.gov/catdir/
                                        description/cam028/
                                        96045964.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam023/96045964.html.

Drake:1969:GFF

[Dra69]  Stillman Drake.  Galileo’s

1604 fragment on falling bodies (Galileo gleanings XVIII). British Journal for the History of Science, 4(4):340–358, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025229.

Drake:1970:UAS

[Dra70] Stillman Drake. Uniform acceleration, space, and time (Galileo gleanings XIX). British Journal for the History of Science, 5(1):21–43, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025351.

Drake:1981:CES

[Dra81] Stillman Drake. Cause, experiment, and science: a Galilean dialogue, incorporating a new English translation of Galileo’s Bodies that stay atop water, or move in it. University of Chicago Press, Chicago, IL, USA, 1981. ISBN 0-226-16228-1. xxix + 237 pp. LCCN QC147 .D7. US$20.00.

Drazin:1986:BRO

[Dra86] Philip Drazin. Book review:

Oliver M. Ashford. Prophet or Professor? The Life and Work of Lewis Fry Richardson. Bristol: Adam Hilger, 1985. Pp. xiv + 364. ISBN 0-85274-774-8. £18/$29. British Journal for the History of Science, 19 (2):217, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026603.

dAlembert:1995:EDP

[dSR95] Jean Le Rond d’Alembert, Richard N. Schwab, and Walter E. Rex, editors. Encyclopédie. Discours préliminaire. Preliminary discourse to the Encyclopedia of Diderot. University of Chicago Press, Chicago, IL, USA, 1995. ISBN 0-226-13476-8 (paperback). liv + 170 pp. LCCN B1933.E5 S38

1995. URL http://www.loc.
gov/catdir/description/
uchi052/95011562.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/95011562-t.html.

Davis:1997:JJT

[DTF97] E. A. (Edward Arthur) Davis, J. J. (Joseph John) Thomson, and I. J. (Isobel J.) Falconer. J. J. Thomson and the discovery of the electron. Taylor and Francis, Boca Raton, FL, USA, 1997. ISBN 0-7484-0696-4 (hardcover). xxvii + 243 pp. LCCN QC16.T45 D38 1997.

Dreyer:1923:HRA

[DTS^{+}87] J. L. E. (John Louis Emil) Dreyer, H. H. (Herbert Hall) Turner, R. A. (Ralph Allen) Sampson, R. J. (Roger John) Tayler, et al., editors. History of the Royal Astronomical Society. Royal Astronomical Society, London, UK, 1923–1987. ISBN 0-632-02173-X (vol. 1), 0-632-02175-6 (vol 1), 0-632-01791-0 (vol. 2.), 0-632-01792-9 (vol. , 2.). ???? pp. LCCN QB1.R75 D73 1923.

Dubbey:1972:BRM

[Dub72] J. M. Dubbey. Book review: Mathematics: The Development of the Foundations of Mathematical Analysis from Euler to Riemann. By Ivor Grattan-Guinness. Cambridge, Mass., and London: M.I.T. Press, 1970. Pp. xiii + 186. £4.65. British Journal for the History of Science, 6(1): 88–89, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025269.

Duchesneau:1982:PLE

[Duc82]  François Duchesneau.  La

physiologie des lumières: empirisme, modèles et théories, volume 95 of Archives internationales d’histoire des idées = International archives of the history of ideas. Martinus Nijhoff, Dordrecht and The Hague, The Netherlands, 1982. ISBN 90-247-2500-3. xxi + 611 pp. LCCN QH331 .D79 1982. fl 180.00 (US$90.00 U.S.).

Duchesne:1984:BRN

[Duc84] Raymond Duchesne. Book review: National Traditions in Science W. E. Knowles Middleton, Radar development in Canada: the Radio Branch of the National Research Council of Canada, 1939–1946. Waterloo Ont: Wilfrid Laurier University Press, 1981. Pp. xii + 147. $9.75. British Journal for the History of Science, 17(1):90, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026025.

Duchesneau:1987:GTC

[Duc87] François Duchesneau. Genèse de la théorie cellulaire. (French) [Genesis of cellular theory], volume 1 of Collection Analytiques. Bellarmin, Montréal, PQ, Canada, 1987. ISBN 2-89007-646-6. 388 pp. LCCN QH577 .D83 1987.

Duchesneau:1988:BRR

[Duc88] François Duchesneau. Book review: Renato Mazzolini and Shirely A. Roe. Science Against the Unbelievers: the Correspondence of Bonnet and Needham, 1760–1780. Oxford: The Voltaire Foundation at the Taylor Institution, 1986. Pp. xix + 409. ISBN 0435-2866. No price given. British Journal for the History of Science, 21 (2):268–269, June 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026997.

Duffin:1975:BRG

[Duf75] Lorna E. Duffin. Book review: General Science and Belief: from Copernicus to Darwin. Open University course AMST 283. 6 volumes. By C. A. Russell, D. C. Goodman, and J. H. Brooke. Milton Keynes: Open University Press, 1974. Pp. 123; 95; 110; 87; 126; 69. Total price £11.00. British Journal for the History of Science, 8(3):

250–251, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025562.

Duff:1981:BRE

[Duf81] Andrew Duff. Book review:

Eighteenth to Twentieth Centuries The kindly fruits of the earth: recollections of an embryo ecologist. By G. Evelyn Hutchinson. New Haven & London: Yale University Press, 1979. Pp. xiv + 264. (no price stated). British Journal for the History of Science, 14(3): 295–296, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025993.

Duff:1983:BRA

[Duf83]  Andrew Duff.  Book re-

view: Americana Paul Russell Cutright and Michael J. Brodhead, Elliott Coues: naturalist and frontier historian. Urbana, Chicago & London: University of Illinois Press, 1981. Pp. xv + 509. £17.10. British Journal for the History of Science, 16 (3):320, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026189.

Dugac:1977:BRN

[Dug77] Pierre Dugac. Book review:

Nineteenth and Twentieth Centuries:

Message d’un mathématicien: Henri Lebesgue. Introduction et extraits choisis par Lucienne Félix. Préface de S. Mandelbrojt. Paris: Librairie scientifique et technique Albert Blanchard, 1974. Pp. vi + 259. 65 francs. British Journal for the History of Science, 10 (2):181–182, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025885.

Duhem:1991:OSS

[Duh91] Pierre Maurice Marie Duhem. The Origins of Statics: The Sources of Physical Theory, volume 123 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-5658-7, 94-011-3730-7 (e-book). ISSN 0068-0346. xxxv + 593 pp. LCCN Q174 .B67 vol.

123. URL http://d-nb.info/
101431545X/34;   https://
link.springer.com/book/
10.1007/978-94-011-3730-0.

Dulac:1988:ED

[Dul88] Georges Dulac, editor. Éditer Diderot, volume 254 of Studies on Voltaire and the eighteenth century, 0435-2866. Voltaire Foundation at the Taylor Institution, Oxford, UK, 1988. ISBN 0-7294-0365-3. 555 + 30 pp. LCCN PQ2105.A2 S8 vol. 254.

Duncan:1969:BRC

[Dun69a] A. M. Duncan. Book review: Chemistry Dissertation on Elective Attractions. By Torbern Bergman, translated with an introduction by J. A. Schufle. New York and London: Johnson Reprint Corporation. 1968. Pp. xxvii + 112. Plates. $12.50. British Journal for the History of Science, 4(4):411–412, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025242.

Duncan:1969:BRH

[Dun69b] A. M. Duncan. Book review: History of Chemistry The Origins of Chemistry. By Robert P. Multhauf. Pp. 412. 9 plates. London: Oldbourne. 1967. 70s. British Journal for the History of Science, 4 (3):292–293, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025483.

Duncan:1986:BRG

[Dun86] Alistair Duncan. Book review: G. V. Coyne, M. Heller & J. Źyciński (eds). The Galileo Affair: A Meeting of Faith and Science. Proceedings of the Cracow Conference, May 1984. Vatican City: Vatican Observatory, 1985. Pp.

179. British Journal for the History of Science, 19(3):343, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026513.

Duncan:1987:BRR

[Dun87] Alistair Duncan. Book review: R. Descartes. The Philosophical Writings. Translated by J. Cottingham, R. Stoothoff and D. Murdoch. Cambridge: C.U.P., 1985. 2 vols, pp. xii + 418, ix + 428. ISBN 60-521-2494-X/8. £27.50 each (hardcover); £8.95 each (paperback). British Journal for the History of Science, 20(2): 232–233, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026318.

Duncan:1988:PEC

[Dun88] Alistair M. Duncan. Particles and Eighteenth Century concepts of chemical combination. British Journal for the History of Science, 21(4):447– 453, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026963.

Dunstan:1990:HEA

[Dun90] Gordon R. (Gordon Reginald) Dunstan, editor. The human embryo: Aristotle and the Arabic and European traditions. University of Exeter Press, Exeter, Devon, UK, 1990. ISBN 0-85989-340-5. xi + 235 pp. LCCN QM601 .H8 1990.

Durant:1977:BRG

[Dur77] John R. Durant. Book review: General Moulds of Understanding. A Pattern of Natural Philosophy. By Joseph Needham. Ed. & Introd. by Gary Werskey. London: George Allen & Unwin, 1976. Pp. 320. £7.75. British Journal for the History of Science, 10(3):252– 254, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025845.

Durant:1979:BRN

[Dur79a] John R. Durant. Book review: Nineteenth and Twentieth Centuries: A Sense of the Future: Essays in Natural Philosophy. By Jacob Bronowski. Cambridge, Mass. & London: The MIT Press, 1977. Pp. x + 286. $12.50. British Journal for the History of Science, 12(2):234–235, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025768.

Durant:1979:SNS

[Dur79b] John R. Durant. Scientific naturalism and social reform in the thought of Alfred Russel Wallace. British Journal for the History of Science, 12 (1):31–58, March 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025662.

Durant:1980:BRN

[Dur80a] John R. Durant. Book review: New Periodical History and philosophy of the life sciences. Ed. by M. D. Grmek. Firenze: Leo S. Olschki Editore, 1979, 1. British Journal for the History of Science, 13(2):183, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025916.

Durant:1980:ERS

[Dur80b] John R. Durant. Exploring the roots of sociobiology. British Journal for the History of Science, 13(1):55–60, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026121.

Durant:1981:BRE

[Dur81] John Durant. Book review: Eighteenth to Twentieth Centuries Nature, Culture and Gender. Ed. by Carol P. Mac-Cormack and Marilyn Strathern. Cambridge: Cambridge University Press, 1980. Pp. x + 227. £12.00/£3.95 (paperback). British Journal for the History of Science, 14(3): 293, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025990.

Durant:1982:BRD

[Dur82] John R. Durant. Book review:

Darwiniana Robert C. Bannister, Social Darwinism: science and myth in Anglo-American social thought. Philadelphia: Temple University Press, 1979. Pp. x + 292. $19.95. British Journal for the History of Science, 15(1):76–77, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025927.

Durant:1983:BRJ

[Dur83] John R. Durant. Book review: Journals and Dictionaries W. F. Bynum, E. J. Browne & Roy Porter (eds.), Dictionary of the history of science. London: Macmillan Press, 1981. Pp. xxxiv + 494. £17.00. British Journal for the History of Science, 16(3): 316–317, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026184.

Durant:1985:DDE

[Dur85] John Durant, editor. Darwinism and divinity: essays on evolution and religious belief. Blackwell, Oxford, UK, 1985. ISBN 0-631-14188-X. ix + 210 pp. LCCN BL263 .D27 1985. US$29.95.

Dybikowski:1978:BRA

[Dyb78] J. Dybikowski. Book review: Antiquity Articles on Aristotle, Volume i. Science. Edited by Jonathan Barnes, Malcolm Schofield, and Richard Sorabji. London: Duckworth, 1975. Pp. xii + 224. £3.50. British Journal for the History of Science, 11(1):72–73, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025615.

Eagles:1977:DGN

[Eag77] Christina M. Eagles. David Gregory and Newtonian science. British Journal for the History of Science, 10(3):216– 225, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025841.

Eagles:1978:BRS

[Eag78] Christina Eagles. Book review: Seventeenth and Eighteenth Centuries: An Index of British Mathematicians. Part 2: 1701– 1760. By Peter John Wallis. Newcastle: Project for Historical Biobibliography, University of Newcastle upon Tyne, 1976. Pp. xxviii + 136. £1.50. British Journal for the History of Science, 11(2):184, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Earles:1970:BRP

[Ear70]  M.  P.

Earles. Book review: Pharmacology The Medical Formulary of Al-Samarqandi and the Relations of Early Arabic Simples to those found in the Indigenous Medicine of the Near East and India. By Martin Levey & Noury Al-Khaledy. University of Philadelphia Press and Oxford University Press. 1967. Pp.

382. 142s. 6d. British Journal for the History of Science, 5(1):95–96, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025365.

Earles:1987:BRC

[Ear87] M. P. Earles. Book review:

Colin A. Russell. Lancastrian Chemist: The Early Years of Sir Edward Frankland. Milton Keynes/Philadelphia: Open University Press, 1986. Pp. ix + 187. ISBN 0-335-15175-2. £30.00. British Journal for the History of Science, 20 (3):358, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026368.

Earman:1989:WES

[Ear89] John Earman. World enough and space-time: absolute versus relational theories of space and time. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-05040-4. xiv + 233 pp. LCCN QC173.59.S65 E17 1989.

Eastwood:1986:BRW

[Eas86] Bruce Eastwood. Book review: Witelo. Perspectivae liber quintius. Book V of Witelo’s Perspectiva. Translation, introduction, commentary, and Latin edition by A. Mark Smith. Studia Copernicana, 23 Wroclaw: Ossolineum, 1983. Pp.

267. British Journal for the History of Science, 19(2):208– 209, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026597.

Euclid:1996:EPT

[EBT96] Euclid, J. L. Berggren, and Robert S. D. Thomas. Euclid’s Phaenomena: a translation and study of a hellenistic treatise in spherical astronomy, volume 4 of Sources and studies in the history and philosophy of classical science. Garland Publishing, Inc., New York, NY, USA, 1996. ISBN 0-8153-0493-5. x + 132 pp. LCCN QB144 .E9313 1996.

Engelhardt:1987:SCC

[EC87] H. Tristram (Hugo Tristram) Engelhardt and Arthur L. Caplan, editors. Scientific controversies: case studies in the resolution and closure of disputes in science and technology. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-25565-1, 0-521-27560-1 (paperback). x + 639 pp. LCCN Q181 .S3826

1987. URL http://www.loc.
gov/catdir/description/
cam022/84028480.html;
http://www.loc.gov/
catdir/toc/cam021/
84028480.html.

Euclide:1992:RCR

[ECL92] Euclide, Robertus Castrensis, and Hubertus Lambertus Ludovicus. Robert of Chester’s (?) redaction of Euclid’s Elements, the so-called Adelard II version, volume 8–9 of Science networks historical studies. Birkhäuser Verlag, Basel, Switzerland, 1992. ISBN 0-8176-2728-6 (Boston), 3-7643-2728-6 (Basel), 0-8176-2658-1 (vol. 1, Boston), 3-7643-2658-1 (vol. 1, Basel), 0-8176-2727-8 (vol. 2, Boston), 3-7643-2727-8 (vol. 2, Basel). 959 pp. LCCN ????

Edelstein:1965:BRF

[Ede65] Sidney M. Edelstein. Book review: Fourcroy. Chemist and Revolutionary (1755–1809). By W. A. Smeaton. Pp. xxi + 288, with six plates. Printed for the author by W. Heffer and Sons Ltd., Cambridge. 1962. 40s. British Journal for the History of Science, 2 (3):266–267, June 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024948.

Edge:1984:BRH

[Edg84] David Edge. Book review: History of Technology: Woodruff Turner Sullivan III, (ed.), Classics in radio astronomy. Dordrecht: D. Reidel, 1982. Pp. xxiv + 348. ISBN 90-277-1356-1. Dfl. 140.00, $59.50. British Journal for the History of Science, 17(3):327– 328, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026642.

Edgerton:1987:BRJ

[Edg87a] D. E. H. Edgerton. Book review: John M. Staudenmaier. Technology’s Storytellers: Reweaving the Human Fabric. Cambridge, Massachusetts: The Society for the History of Technology and the MIT Press, 1985. Pp. xxvii + 282. ISBN 0-262-19237-3. £39.25. British Journal for the History of Science, 20(1):102–103, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026456.

Edgerton:1987:BRM

[Edg87b] D. E. H. Edgerton. Book review: Merritt Roe Smith (ed.) Military Enterprise and Technological Change: Perspectives on the American Experience. Cambridge, Massachusetts: The MIT Press, 1985. Pp. 391. £29.95. British Journal for the History of Science, 20(1): 99–100, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026453.

Edgerton:1987:BRL

[Edg87c] David Edgerton. Book review:

Leonard S. Reich. The Making of American Industrial Research: Science and Business at GE and Bell, 1876–1926. Cambridge: Cambridge University Press, 1985. Pp. xvi + 309. ISBN 0-521-30529-2. Price £25.00, $24.95. British Journal for the History of Science, 20(2):228, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026313.

Edgerton:1991:HGG

[Edg91] Samuel Y. Edgerton. The heritage of Giotto’s geometry: art and science on the eve of the scientific revolution. Cornell University Press, Ithaca, NY, USA, 1991. ISBN 0-8014-2573-5 (hardcover). x + 319 pp. LCCN N7430.5 .E34 1991.

Edwards:1978:BRS

[Edw78] P. I. Edwards. Book review:

Seventeenth and Eighteenth Centuries: British Botanical and Horticultural Literature before 1800. By Blanche Henrey. 3 vols. Pp. xxvi + 290; xvi + 178; xvii + 112. 194 plates. Oxford: Clarendon Press: Oxford University Press, 1975.

£70.00. British Journal for the History of Science, 11(1): 80–81, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025621.

Edwards:1987:PAT

[Edw87]  A.  W.

F. (Anthony William Fairbank) Edwards. Pascal’s arithmetical triangle. C. Griffin, London, UK, 1987. ISBN 0-85264-283-0, 0-19-520546-4 (Oxford). xii + 174 pp. LCCN QA241 .E37 1987. US$35.00. URL http:

                                        //www.loc.gov/catdir/
                                        enhancements/fy0635/
                                        86019264-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0635/86019264-t.html.

Edwards:1996:CWC

[Edw96] Paul N. Edwards. The closed world: computers and the politics of discourse in Cold War America. Inside technology. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-05051-X. xx + 440 pp. LCCN QA76.17 .E34 1996.

                                        URL http://www.h-net.org/
                                        review/hrev-a0a5x0-aa.

Emery:1995:HGD

[EE95] Alan E. H. Emery and Marcia L. H. Emery. The history of a genetic disease: Duchenne muscular dystrophy or Meryon’s disease. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1995. ISBN 1-85315-249-8, 1-85315-249-3. xvi + 248 pp. LCCN RC935.M7 E46 1995.

Egerton:1989:BRP

[Ege89] Frank N. Egerton. Book review: Pascal Acot. Histoire de l’écologie. Paris: Presses Universitaires de France, 1988. ISBN 2-13-041414-8. FF 150.00. British Journal for the History of Science, 22(1):91–92, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026691.

Eisenstein:1983:PRE

[Eis83] Elizabeth L. Eisenstein. The printing revolution in early modern Europe. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25858-8, 0-521-27735-3 (paperback). xiv + 297 pp. LCCN Z124 .E374 1983.

Eisenstein:1993:PRE

[Eis93] Elizabeth L. Eisenstein. The printing revolution in early modern Europe. Cambridge University Press, Cambridge, UK, canto edition, 1993. ISBN 0-521-25858-8 (hardcover), 0-521-27735-3 (paperback), 0-521-44770-4 (Canto paperback). xiv + 300 pp. LCCN Z124 .E374 1993.

US$7.95. URL http://www.
loc.gov/catdir/
description/cam026/
94153513.html;   http://

www.loc.gov/catdir/toc/
cam029/94153513.html.

Euclid:1992:MLT

[EiT92] Euclid and Ken ichi Takahashi. The medieval Latin traditions of Euclid’s Catoptrica: a critical edition of De speculis with an introduction, English translation, and commentary. Kyushu University Press, Higashi-ku, Fukuoka-shi, Japan, 1992. ISBN 4-87378-299-6. viii + 373 pp. LCCN QC353 .E813 1992. URL http:

//catalog.hathitrust.org/
api/volumes/oclc/
26940678.html.

Edison:1989:PTEa

[EJ^{+}07] Thomas A. (Thomas Alva) Edison, Reese Jenkins, et al., editors. The papers of Thomas A. Edison. The Johns Hopkins University Press, Baltimore, MD, USA, 1989–2007. ISBN 0-8018-3100-8 (vol. 1), 0-8018-3101-6 (vol. 2), 0-8018-3102-4 (vol. 3), 0-8018-5819-4 (vol. 4), 0-8018-3104-0 (vol. 5), 0-8018-8640-6 (vol. 6). ???? pp. LCCN TK140.E3 A2

1989. URL http://www.loc.
gov/catdir/bios/jhu051/
88009017.html;
http://www.loc.gov/
catdir/description/
jhu051/88009017.html.

Edison:1994:MPE

[EJR94] Thomas Alva Edison, Reese V. Jenkins, and Robert A. Rosenberg. Menlo Park: the early years: April 1876–December 1877, volume Vol. 3 of The papers of Thomas A. Edison / Reese V. Jenkins, dir. and ed. The Johns Hopkins University Press, Baltimore, MD, USA, 1994. ISBN 0-8018-3102-4. xlvii + 727 pp. LCCN ????

Edison:1995:PTE

[EJR95] Thomas Alva Edison, Reese V.

Jenkins, and Robert A. Rosenberg. The papers of Thomas A. Edison. Vol. 3, Menlo Park: the early years, April 1876– December 1877. The Johns Hopkins University Press, Baltimore, MD, USA, 1995. ISBN 0-8018-3102-4. xlvii + 727 pp. LCCN ????

Edwards:1988:MOP

[EK88] Sam F. (Sam Frederick) Edwards and P. R. (Peter R.) King, Dr, editors. Mathematics in oil production: based on the proceedings of a conference, volume 18 (new series) of The Institute of Mathematics and its Applications conference series. Clarendon Press, New York, NY, USA, 1988. ISBN 0-19-853624-0, 0-19-853924-X. LCCN TN871 .M384 1988. US$60.00. URL http://www.

loc.gov/catdir/
enhancements/fy0638/
88029741-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0638/88029741-t.html.

Eisenstaedt:1991:SHG

[EK91] Jean Eisenstaedt and Anne J.

Kox, editors. Studies in the history of General Relativity: proceedings of the Second International Conference on the History of General Relativity, Luminy, Marseille, France, 6–9 September 1988, volume 3 of Einstein studies. Birkhäuser Verlag, Basel, Switzerland, 1991. ISBN 0-8176-3479-7 (Boston), 3-7643-3479-7 (Berlin). LCCN QC173.6 I572 1988.

Eisenstaedt:1992:SHG

[EK92] Jean Eisenstaedt and Anne J. Kox, editors. Studies in the history of general relativity: based on the proceedings of the 2nd International Conference on the History of General Relativity, Luminy, France, 1988, volume 3 of Einstein studies. Birkhäuser Verlag, Basel, Switzerland, 1992. ISBN 0-8176-3479-7, 3-7643-3479-7. LCCN QC173.6 .I572 1988. US$125.00.

Edgeworth:1994:ECE

[EM94] Francis Ysidro Edgeworth and Philip Mirowski. Edgeworth on chance, economic hazard, and statistics. The Worldly philosophy. Rowman and Littlefield, Lanham, MD, USA, 1994. ISBN 0-8476-7751-6 (hardcover). vii + 462 pp. LCCN HB615 .E43 1994.

Emerson:1979:PSE

[Eme79] Roger L. Emerson. The Philosophical Society of Edinburgh 1737–1747. British Journal for the History of Science, 12 (2):154–191, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025751.

Emerson:1981:PSE

[Eme81] Roger L. Emerson. The Philosophical Society of Edinburgh: 1748–1768. British Journal for the History of Science, 14 (2):133–176, July 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025944.

Emerton:1984:SRF

[Eme84] Norma E. Emerton. The scientific reinterpretation of form. Cornell history of science series. Cornell University Press, Ithaca, NY, USA, 1984. ISBN 0-8014-1583-7. 318 pp. LCCN QC173.3 .E47 1984.

Emerson:1985:PSE

[Eme85] Roger L. Emerson. The Philosophical Society of Edinburgh 1768–1783. British Journal for the History of Science, 18(3): 255–303, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026382.

Emerson:1987:BRJ

[Eme87] Roger L. Emerson. Book review: James E. McClellan III. Science Reorganized: Scientific Societies in the Eighteenth Century. New York: Columbia University Press, 1985. Pp. xxxii + 413. ISBN 0-231-05996-5 (hardcover), 0-231-05997-3 (paperback). US $58.50. British Journal for the History of Science, 20(1): 86–88, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026444.

Emerson:1988:SEE

[Eme88] Roger L. Emerson. The Scottish Enlightenment and the end of the Philosophical Society of Edinburgh. British Journal for the History of Science, 21 (1):33–66, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026861.

Euclid:1993:DE

[EMMS93a] Euclid, Heinrich Menge, George L. McDowell, and Merle A. Sokolik. The data of Euclid. Union Square Press, Baltimore, MD, USA, 1993. ISBN 0-9635924-1-6, 0-9635924-0-8. xxi + 207 pp. LCCN QA444 .E9313 1993.

Euclides:1993:DE

[EMMS93b] Euclides, Heinrich Menge, George L. McDowell, and Merle A. Sokolik, editors. The data of Euclid. Union Square Press, Baltimore, MD, USA, 1993. ISBN 0-9635924-1-6, 0-9635924-0-8. xxi + 207 pp. LCCN ????

Elias-:1982:SEH

[EMW82] Norbert Elias, Herminio Martins, and Richard Whitley, editors. Scientific establishments and hierarchies, volume 6 of Sociology of the sciences: a yearbook. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1982. ISBN 90-277-1322-7, 90-277-1323-5 (paperback). xi + 368 pp. LCCN Q180.A1 S345 1982.

Engel:1983:CDR

[Eng83] A. J. (Arthur Jason) Engel.

From clergyman to don: the rise of the academic profession in Nineteenth-Century Oxford. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1983. ISBN 0-19-822606-3. xi + 302 pp. LCCN LF519 .E53 1983. US$45.00.

Engelsman:1984:FCO

[Eng84] Steven B. (Steven Boudewijn)

Engelsman. Families of curves and the origins of partial differentiation, volume 93 of North-Holland mathematics studies. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984. ISBN 0-444-86897-6 (paperback). ix + 238 pp. LCCN QA565 .E54 1984.

Edison:1993:PTE

[ER93]  Thomas  A.  Edi-

son and Robert A. Rosenberg. The papers of Thomas A. Edison. The Johns Hopkins University Press, Baltimore, MD, USA, 1993. ISBN 0-8018-3102-4. xlvii + 727 pp. LCCN ????

Eriksson:1994:AVO

[Eri94] Gunnar Eriksson. The Atlantic vision: Olaus Rudbeck and Baroque science, volume 19 of Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1994. ISBN 0-88135-158-X. viii + 196 pp. LCCN Q125.2 .E75 1994. US$27.95.

Ernst:1991:RRN

[Ern91] Germana Ernst. Religione, ragione e natura: ricerche su Tommaso Campanella e il tardo Rinascimento. (Italian) [Religion, reason and nature: research on Tommaso Campanella and the late Renaissance], volume 41 of Collana di filosofia. F. Angeli, Milano, Italy, 1991. ISBN 88-204-6663-5. 288 pp. LCCN B785.C24 E76 1991. L36000.

Eckert:1990:CET

[ES90] Michael Eckert and Helmut Schubert. Crystals, electrons, transistors: from scholar’s study to industrial research. American Institute of Physics, Woodbury, NY, USA, 1990. ISBN 0-88318-622-5. xxii + 241 pp. LCCN Q125 .E36313 1990.

Euclides:1992:MLT

[ET92] Euclides and Ken’ichi Takahashi. The medieval Latin traditions of Euclid’s Catoptrica: a critical edition of De speculis with an introduction, English translation and commentary. Kyushu University Press, Fukuoka-shi, Japan, 1992. ISBN 4-87378-299-6. viii + 373 pp. LCCN ????

Evans:1977:AAT

[Eva77] Gillian R. Evans. From abacus to algorism: Theory and practice in medieval arithmetic. British Journal for the History of Science, 10(2):114– 131, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025865.

Evans:1987:A

[Eva87] J. D. G. (John David Gemmill) Evans. Aristotle. Philosophers in context. Harvester Press, Brighton, Sussex, UK, 1987. ISBN 0-312-00485-0 (US), 0-7108-1042-3. xii + 208 pp. LCCN B485 .E93

1987. URL http://catalog.
hathitrust.org/api/
volumes/oclc/14691677.
html.

Evans:1989:BRM

[Eva89] J. D. G. Evans. Book review:

Montgomery Furth. Substance, Form and Psyche: an Aristotelean Metaphysics. Cambridge: Cambridge University Press, 1988. Pp. xiv + 300. ISBN 0-521-34143-4. £30. British Journal for the History of Science, 22(4):485–486, December 1989.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026952.

Earles:1970:RC

[EW70]  M.  P.

Earles and G. J. Whitrow. Report of Council. British Journal for the History of Science, 5(2):211–213, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Eyles:1968:BRH

[Eyl68] V. A. Eyles. Book review: History of Geology James Huttonthe Founder of Modern Geology. By Edward Battersby Bailey. Pp. xii + 161. Elsevier, Amsterdam. 1967. 50s. British Journal for the History of Science, 4(2):184–185, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025157.

F:1976:CLC

[F.76] R. F. The Charles Lyell Centenary Symposium. British Journal for the History of Science, 9(2):??, July 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025797.

F:1977:NC

[F.77] R. F. Notes on contributors. British Journal for the History of Science, 10(1):93–94, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fauvel:1988:LN

[F^{+}88] John Fauvel et al., editors. Let Newton be! Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. ISBN 0-19-853924-X. 272 pp. LCCN QC16.N7 L47 1988. US$29.95.

URL http://www.loc.gov/
catdir/enhancements/
fy0639/88009911-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0725/88009911-b.html.

Falconer:1987:CEC

[Fal87] Isobel Falconer. Corpuscles, electrons and cathode rays: J. J. Thomson and the ‘discovery of the electron’. British Journal for the History of Science, 20(3):241–276, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026358.

Falk:1989:BRJ

[Fal89] Jim Falk. Book review: Juan Martinez-Alier. Ecological Economics. Basil Blackwell, Oxford, 1987. Pp. ix + 286. ISBN 0-631-15739-5. £29.50. British Journal for the History of Science, 22(1):119– 120, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026715.

Fancher:1983:FGA

[Fan83] Raymond E. Fancher. Francis Galton’s African ethnography and its role in the development of his psychology. British Journal for the History of Science, 16(1):67–79, March 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026094.

Fantoli:1996:GCC

[Fan96] Annibale Fantoli. Galileo, for Copernicanism and for the Church, volume 3 of Studi Galileiani. Vatican Observatory Publications, Vatican City, second edition, 1996. ISBN 0-268-01032-3 (paperback). xx + 567 pp. LCCN QB41.G178 F3613 1996.

Farrar:1963:NES

[Far63] Kathleen R. Farrar. A note on a eudiometer supposed to have belonged to Henry Cavendish. British Journal for the History of Science, 1(4):375–380, December 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024909.

Farrar:1965:NCS

[Far65] W. V. Farrar. Nineteenth-Century speculations on the complexity of the chemical elements. British Journal for the History of Science, 2(4): 297–323, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024888.

Farrar:1967:BRH

[Far67]  W. V. Farrar.  Book re-

view: History of Chemistry The Atomic Debates. Ed. by W. H. Brock. Pp. ix + 186. Leicester University Press, 1967. 35s. British Journal for the History of Science, 3(4):405– 406, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024975.

Farrar:1968:BRH

[Far68a]  W. V. Farrar.  Book re-

view: History of Chemistry Atoms and Elements. By David M. Knight. Pp. 167. London: Hutchinson. 1967. 30s. British Journal for the History of Science, 4(2):182–183, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025155.

Farrar:1968:CSC

[Far68b] W. V. Farrar. “chemistry in space” and the complex atom. British Journal for the History of Science, 4(1): 65–67, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024988.

Farrar:1969:BRH

[Far69] W. V. Farrar. Book review: History of Chemistry Chymia. vol. 12. Ed. by Henry M. Leicester. Pp. 236. Philadelphia: University of Pennsylvania Press. 1967. 81s. British Journal for the History of Science, 4(3): 293, June 1969. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Farrar:1971:BRC

[Far71a]  W. V. Farrar.  Book re-

view: Chemistry Classical Scientific Papers: Chemistry. Second Series. Ed. by David M. Knight. London: Mills and Boon. New York: American Elsevier. 1970. Pp. xiii + 441. £5. British Journal for the History of Science, 5(4):406– 407, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025391.

Farrar:1971:BRN

[Far71b] W. V. Farrar. Book review: Nineteenth Century: John Dalton and the Atomic Theory. By Elizabeth C. Patterson. New York: Doubleday. 1970. Pp. viii + 348. Illustr. $6.95. British Journal for the History of Science, 5(3):313–314, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025343.

Farrar:1976:BRN

[Far76] W. V. Farrar. Book review:

Nineteenth and Twentieth Centuries: The Emergence of Agricultural Science. Justus Liebig and the Americans, 1840–1880. By Margaret W. Rossiter. New Haven and London: Yale University Press, 1975. Pp. xiv + 275. No price stated. British Journal for the History of Science, 9(1):80–81, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025715.

Farrar:1977:BRM

[Far77] Kathleen R. Farrar. Book review: Medicine The Hospital. A Social and Architectural History. By John D. Thompson and Grace Goldin. New Haven and London: Yale University Press, 1975. Pp. xxviii + 349. $25.00. British Journal for the History of Science, 10(1):76, March 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025589.

Fauvel:1989:BRR

[Fau89] John Fauvel. Book review:

R. V. & P. J. Wallis. Biobibliography of British Mathematics and its Applications, Part II, 1701–1760. Newcastle upon Tyne: Project for Historical Biobibliography, 1986, Pp. xxiii + 502. ISBN 1-85027-003-1. £48. British Journal for the History of Science, 22(1): 118–119, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026714.

Fisher:1975:LPD

[FB75] N. W. Fisher and W. H. Brock. A list of Ph.D. theses in the history of science and related areas in British Universities, 1945–1974. British Journal for the History of Science, 8(3): 267–278, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025574.

Ferrier:1982:HBP

[FB82] R. W. (Ronald W.) Ferrier and J. H. Bamberg. The history of the British Petroleum Company. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24647-4 (vol. 1), 0-521-25950-9 (vol. 2), 0-521-25951-7 (vol. 3: hardback), 0-521-78515-4 (vol. 3: paperback). ???? pp. LCCN HD9571.9.B73 F47 1982; ECON.

Fee:1997:MMH

[FB97] Elizabeth Fee and Theodore M. Brown, editors. Making medical history: the life and times of Henry E. Sigerist. The Johns Hopkins University Press, Baltimore, MD, USA, 1997. ISBN 0-8018-5355-9. xiii + 387 pp. LCCN R133 .M26

1997. URL http://www.loc.
gov/catdir/bios/jhu051/
96008862.html;
http://www.loc.gov/
catdir/description/
jhu051/96008862.html.

Flamsteed:1982:PJF

[FC82] John Flamsteed and Allan Chapman. The preface to John Flamsteed’s Historia coelestis Britannica, or, British catalogue of the heavens (1725), volume 52–1982 of Maritime monographs and reports, 0307-8590. Trustees of the National Maritime Museum, London, UK, 1982. ISBN 0-905555-60-0 (paperback). vi + 222 pp. LCCN VK553 .F5313 1982.

Fantoli:1994:GCC

[FC94] Annibale Fantoli and George V.

Coyne, editors. Galileo: for Copernicanism and for the Church, volume 3 of Studi Galileiani. Vatican Observatory Publications, Vatican City, 1994. ISBN 0-268-01029-3. xix + 540 pp. LCCN QB36.G2 F36713

1994. URL http://catalog.
hathitrust.org/api/
volumes/oclc/31182040.
html.

French:1996:BSI

[FC96] R. K. (Roger Kenneth) French and Andrew Cunningham, Dr. Before science: the invention of the friars’ natural philosophy.

Scolar Press, Aldershot, Hants, 1996. ISBN 1-85928-287-3. 298 pp. LCCN B738.N3 F74 1996.

Friedman:1985:EMM

[FD85] Alan J. Friedman and Carol C. Donley. Einstein as myth and muse. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-26720-X. xiv + 224 pp. LCCN QC16.E5 F76

                                        1985. URL http://www.loc.
                                        gov/catdir/description/
                                        cam031/84023839.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0906/84023839-t.html.

Fourcy:1987:HEP

[FD87] A. (Ambroise) Fourcy and Jean G. Dhombres. Histoire de l’École polytechnique. Librairie du bicentenaire de la Révolution française. Belin, Paris, France, 1987. ISBN 2-7011-0640-0. viii + 516, 198 pp. LCCN T173.P39 F68 1987.

Fee:1977:BRG

[Fee77] Elizabeth Fee. Book review: General Women in Science. By H. J. Mozans. Facsimile of 1913 edition. Introduction by Mildred Dressenhaus. Cambridge, Mass., and London: M.I.T. Press, 1974. Pp. xvii + 452. £2.50. British Journal for the History of Science, 10(1): 69–70, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025584.

Feingold:1990:BNL

[Fei90] Mordechai Feingold, editor.

Before Newton: the life and times of Isaac Barrow. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-30694-9. xi + 380 pp. LCCN QA29.B37 B44

1990. URL http://www.loc.
gov/catdir/description/
cam024/89038508.html;
http://www.loc.gov/
catdir/enhancements/
fy0632/89038508-t.html.

Felling:1989:BRW

[Fel89] Margaret Felling. Book reviews: William Coleman. Yellow Fever in the North: The Methods of Early Epidemiology. Madison, Wisconsin: University of Wisconsin Press, 1987. Pp. xvi + 202. ISBN 0-299-11114-8, £16.65 (paperback). ISBN 0-299-11110-5, £47.05 (hardcover). British Journal for the History of Science, 22(3):382– 383, September 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026903.

Farrar:1966:BRH

[FF66] W. V. Farrar and Kathleen R.

Farrar. Book review: History of Chemistry The Development of Modern Chemistry. By Aaron J. Ihde. Pp. xii + 851, with 217 plates and figures. New York, Evanston, and London: Harper and Row, 1964. £5 1s. British Journal for the History of Science, 3(1):85–87, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025115.

Flamsteed:1995:CJF

[FFLW97]  John  Flam-

steed, Eric G. Forbes, Murdin Lesley, and Frances Willmoth. The correspondence of John Flamsteed, the first astronomer royal. Institute of Physics Publishing, Bristol, UK, 1995– 1997. ISBN 0-7503-0147-3. ???? pp. LCCN QB36.F6 A4

                                        1995. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0668/95047433-d.html.

Fox:1993:ETI

[FG93]  Robert  Fox

and Anna Guagnini, editors. Education, technology, and industrial performance in Europe, 1850–1939. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-38153-3. xiii + 302 pp. LCCN T105 .E33

                                        1993. URL http://www.loc.
                                        gov/catdir/description/
                                        cam025/92023171.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        92023171.html.

Fink:1998:WT

[FGJ98] Carole Fink, Philipp Gassert, and Detlef Junker, editors. 1968, the world transformed, Publications of the German Historical Institute. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-64141-1 (hardcover), 0-521-64637-5 (paperback). LCCN D839.2 .A17

1998. URL http://www.loc.
gov/catdir/description/
cam029/98023253.html;
http://www.loc.gov/
catdir/toc/cam026/
98023253.html.

Frangsmyr:1990:QSC

[FHR90] Tore Frängsmyr, J. L. Heilbron, and Robin E. Rider, editors. The quantifying spirit in the 18th Century, volume 7 of Uppsala studies in history of science. University of California Press, Berkeley, CA, USA, 1990. ISBN 0-520-07022-4. viii + 411 + 21 pp. LCCN Q125 .Q34 1990.

URL http://ark.cdlib.org/
ark:/13030/ft6d5nb455.

Field:1981:BRN

[Fie81] John Field. Book review: Nineteenth Century: Technology and Toil in Nineteenth-Century Britain. Documents edited by Maxine Berg. London: CSE Books; Atlantic Highlands, NJ: Humanities Press, 1979. Pp.

246. No price stated. British Journal for the History of Science, 14(1):93–94, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026081.

Field:1983:BRS

[Fie83] John Field. Book review: Science and Society Les Levidow and Bob Young (eds.), Science, technology and the labour process: Marxist studies, Vol. 1, London: CSE Books. Arlautic Highlands: Humanities Press, 1981. pp. 207, £12.00/£4.95 paperback. British Journal for the History of Science, 16(3): 310–311, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026178.

Field:1986:BRS

[Fie86] J. V. Field. Book review: Sidney J. Blatt. Continuity and Change in Art: the Development of Modes of Representation. In collaboration with Ethel S. Blatt. Hillsdale, N.J., and London: Lawrence Erlbaum Associates, 1984. Pp. xix + 411. ISBN 0-89859-342-5. £35.50, $49.95. British Journal for the History of Science, 19(2):211, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026599.

Field:1988:KGC

[Fie88] Judith Veronica Field. Kepler’s geometrical cosmology. University of Chicago Press, Chicago, IL, USA, 1988. ISBN 0-226-24823-2. xx + 243 pp. LCCN QB981 .F47

1988. URL http://www.loc.
gov/catdir/enhancements/
fy0608/87010837-d.html.

Field:1989:BRJ

[Fie89] J. V. Field. Book review: John H. Hammond & Jill Austin. The Camera Lucida in Art and Science. Bristol: Adam Hilger, 1987. Pp. xii + 201. ISBN 0-85274-527-3. £19.95. British Journal for the History of Science, 22(1):116, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026712.

Field:1997:IIM

[Fie97] Judith Veronica Field. The invention of infinity: mathematics and art in the Renaissance. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1997. ISBN 0-19-852394-7 (hardcover). xii + 250 pp. LCCN N7430.5 .F52

1997. URL http://www.loc.
gov/catdir/enhancements/
fy0636/97006805-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0636/97006805-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0724/97006805-b.html.

Friedel:1986:EEL

[FIF86] Robert D. (Robert Douglas)

Friedel, Paul Israel, and Bernard S. Finn. Edison’s electric light: biography of an invention. Rutgers University Press, New Brunswick, NJ, USA, 1986. ISBN 0-8135-1118-6. xvi + 263 pp. LCCN TK4351 .F75 1986. US$35.00.

Figlio:1977:BRMa

[Fig77a] Karl Figlio. Book review: Medicine History of Physiology. By Karl E. Rothshuh. Ed. and trans. by Guenter B. Risse. Huntington, New York: Krieger, 1973. Pp. xxii + 379. No price stated. British Journal for the History of Science, 10(2):163–164, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025873.

Figlio:1977:BRMb

[Fig77b]  Karl  Figlio.

Book review: Medicine The Birth of the Clinic. An Archaeology of Medical Perception. By Michel Foucault. Trans. by A. M. Sheridan-Smith. London: Tavistock, 1973. PP. xix + 215. £4-20; £3.20 (paperback; 1976). British Journal for the History of Science, 10 (2):164–167, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025874.

Figlio:1978:BRM

[Fig78] Karl Figlio. Book review: Medicine The Rise of the Medical Profession: a Study of Collective Social Mobility. By Noel and José Parry. London: Croom Helm, 1976. Pp. v + 282. £8.50. British Journal for the History of Science, 11(3): 281–283, November 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025783.

Figala:1988:BRJ

[Fig88] Karin Figala. Book review:

Jeremy Adler. ‘Eine fast magische Anziehungskraft’. Goethe’s ‘Wahlverwandtschaften’ und die Chemie seiner Zeit. München: C.H. Beck Verlag, 1987. Pp. 260, 10 figs, and 4 plates. ISBN 3-406-31559-3. DM 48.00. British Journal for the History of Science, 21(1): 123–125, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026879.

Finn:1968:BRT

[Fin68] Bernard S. Finn. Book review: Twentieth Century A General History of the Sciences: Science in the Twentieth Century. Edited by René Taton. Pp. xxiv + 638. Illus. London: Thames and Hudson. 1966. £6 6s. New York: Basic Books. 1966. $17.50. British Journal for the History of Science, 4(1):79–80, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024997.

Finn:1971:OEC

[Fin71] Bernard S. Finn. Output of Eighteenth-Century electrostatic machines. British Journal for the History of Science, 5(3):289–291, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025318.

Finney:1993:PRN

[Fin93] C. M. (Colin Michael) Finney. Paradise revealed: natural history in Nineteenth-Century Australia. Museum of Victoria, Melbourne, Victoria, Australia, 1993. ISBN 0-7306-2494-3. xv + 186 pp. LCCN QH21.A8 F56 1993.

Fisher:1974:BRC

[Fis74] Nicholas Fisher. Book review: Chemistry The Evolution of the Atomic Theory. By D. P. Mellor. Amsterdam: Elsevier, 1971. Pp. viii + 171. Hfl. 40. British Journal for the History of Science, 7(3): 288–289, November 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025538.

Fisher:1978:BRN

[Fis78] Nicholas Fisher. Book review: Nineteenth and Twentieth Centuries: Chronologic Chemie, 1800– 1970. By Sieghard Neufeldt. Weinheim & New York: Verlag Chemie, 1977. Pp. vii + 359. DM 78. British Journal for the History of Science, 11 (2):186–187, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025745.

Fisher:1981:BRE

[Fis81a] Nicholas Fisher. Book review:

Eighteenth Century: The Playfair Collection and the Teaching of Chemistry at the University of Edinburgh, 1713– 1858. By R. G. W. Anderson. Edinburgh: The Royal Scottish Museum, 1978. Pp. viii + 175. £4.50. British Journal for the History of Science, 14 (1):91–92, March 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026080.

Fisher:1981:BRS

[Fis81b] Nicholas Fisher. Book review:

Seventeenth and Eighteenth Centuries: Chemistry Transformed: the Paradigmatic Shift from Phlogiston to Oxygen. By H. Oilman McCann. Norwood, NJ: Ablex Publishing, 1978. Pp. x + 179. $14.95. British Journal for the History of Science, 14(2):214–216, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025956.

Fisher:1981:SNB

[Fis81c] Nicholas Fisher. Short notices of books: Berzelius och den kemiska atomteorin. (With a summary in English: ‘Berzelius and the chemical atomic theory’.) By Anders Lundgren.

Uppsala: Almqvist & Wiksell for the University of Uppsala, 1979. Pp. 210. Sw Kr 75. British Journal for the History of Science, 14(1):104, March 1981. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fisher:1983:BRJb

[Fis83a] Nicholas Fisher. Book review: Journals and Dictionaries Fred Gettings, Dictionary of occult, hermetic and alchemical sigils. London: Routledge & Kegan Paul, 1981. Pp. 410. £15.95. British Journal for the History of Science, 16(3): 317–318, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026185.

Fisher:1983:BRJa

[Fis83b] Nicholas Fisher. Book review: Journals and Dictionaries John Neu (ed.), Isis cumulative bibliography, 1966–1975. Volume i: personalities and institutions. London: Mansell, in conjunction with the History of Science Society, 1980. Pp. xxx + 483. £44.00. British Journal for the History of Science, 16(3): 315–316, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026183.

Fisher:1983:BRS

[Fis83c] Nicholas Fisher. Book review:

Science and Society David Knight, Ordering the world: a history of classifying man. London: Burnett Books in association with André Deutsch, 1981. Pp. 215. £7.95. British Journal for the History of Science, 16(3):306–307, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026174.

Fisch:1991:WWP

[Fis91a] Menachem Fisch. William Whewell, philosopher of science. Clarendon Press, New York, NY, USA, 1991. ISBN 0-19-824240-9. xiv + 220 pp. LCCN Q175 .F5453 1991. US$59.00. URL http://www.

loc.gov/catdir/
enhancements/fy0638/
90046590-d.html.

Fisher:1991:EJ

[Fis91b] Richard B. Fisher. Edward Jenner: 1749–1823. André Deutsch Ltd., London, UK, 1991. ISBN 0-233-98681-2. 361 + 8 pp. LCCN R489.J5 F57 1991x.

Fischer:1998:MTL

[Fis98] Fritz Fischer. Making them like us: Peace Corps volunteers in the 1960s. Smithsonian Institution Press, Washington, DC, USA, 1998. ISBN 1-56098-889-4. viii + 237 pp. LCCN HC60.5 .F57 1998.

Field:1993:RRH

[FJ93] Judith Veronica Field and Frank A. J. L. James, editors. Renaissance and revolution: humanists, scholars, craftsmen, and natural philosophers in early modern Europe. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-43427-0 (hardcover). xv + 291 pp. LCCN Q125.2 .R45

1993. URL http://www.loc.
gov/catdir/description/
cam025/93006624.html;
http://www.loc.gov/
catdir/toc/cam021/
93006624.html.

Fox:1988:PMI

[FL88] Daniel M. Fox and Christopher Lawrence. Photographing medicine: images and power in Britain and America since 1840, volume 21 of Contributions in medical studies, 0886-8220. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1988. ISBN 0-313-23719-0. 357 pp. LCCN R486 .F69 1988.

Frangsmyr:1994:LMH

[FL^{+}94] Tore Frängsmyr, Sten Lindroth, et al., editors. Linnaeus, the man and his work, volume 18 of Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, revised edition, 1994. ISBN 0-88135-155-5 (hardcover), 0-88135-154-7 (paperback). xiv + 206 pp. LCCN QH44 .L56 1994. US$19.95, US$19.95.

Francis:2000:LPA

[FM00] R. A. Francis and Jean Mainil, editors. L’abbé Prévost au tournant du siècle. (French) [The Abbey Prévost at the turn of the century], volume 2000:11 of SVEC, 0435-2866. Voltaire foundation, Oxford, UK, 2000. ISBN 0-7294-0733-0. ix + 390 pp. LCCN PQ2021.Z5 A53 2000.

Freud:1985:CLS

[FMF85] Sigmund Freud, J. Moussaieff (Jeffrey Moussaieff) Masson, and Wilhelm Fliess. The complete letters of Sigmund Freud to Wilhelm Fliess, 1887–1904. Belknap Press of Harvard University Press, Cambridge, MA, USA and London, UK, 1985. ISBN 0-674-15420-7. xv + 505 + 20 pp. LCCN BF173.F85 A4 1985.

Fogg:1992:HAS

[Fog92] G. E. (Gordon Elliott) Fogg.

A history of Antarctic science. Studies in polar research. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-36113-3. xxi + 483 pp. LCCN G860 .F64 1992. URL http:

//www.loc.gov/catdir/
description/cam024/
91048063.html;   http://
www.loc.gov/catdir/toc/
cam029/91048063.html.

Forbes:1966:BRE

[For66] E. G. Forbes. Book review: Edmond Halley. By Angus Armitage. Pp. xii + 220. 27 figures and 16 plates. London: Thomas Nelson, 1966. 42s. British Journal for the History of Science, 3(2):192–193, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025035.

Forbes:1967:BNA

[For67a] Eric G. Forbes. The bicentenary of the Nautical Almanac (1767). British Journal for the History of Science, 3(4): 393–394, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024963.

Forbes:1967:BRH

[For67b] Eric G. Forbes. Book review: History and Philosophy of Science 300 Jahre Physik und Astronomie an der Kieler Universität. By Charlotte Schmidt-Schönbeck. Kiel: Verlag Ferdinand Hirt. Pp. 261. 23 illus. 1965. Price not stated. British Journal for the History of Science, 3(4):400–401, December 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024969.

Forbes:1968:BRS

[For68] Eric G. Forbes. Book review: Sixteenth and Seventeenth Centuries: The Correspondence of Isaac Newton, Volume IV, 1694–1709. Edited by J. F. Scott. Published for the Royal Society at the Cambridge University Press. 1967. Pp. xxxii + 578. 11 gns. net. British Journal for the History of Science, 4(2):193– 194, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025166.

Forbes:1969:BRA

[For69a] Eric G. Forbes. Book review: Astronomy and Physics Bau und Bildung des Weltalls. By Bernhard Sticker. Freiburg: Verlag Herder. Pp. 272. 10 illus. 1967. Price not stated. British Journal for the History of Science, 4(4):407–408, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025239.

Forbes:1969:BRE

[For69b] Eric G. Forbes. Book review: Eighteenth Century: Die Französische Pendule des 18. Jahrhunderts: ein Beitrag zu ihrer Ikonologie. By Klaus Maurice. Berlin: de Gruyter & Co. 1967. Pp. x + 124. 6 figs., 59 plates. Price not stated. British Journal for the History of Science, 4(3): 300–301, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025493.

Forbes:1970:TMC

[For70] Eric Gray Forbes. Tobias Mayer (1723–1762): A case of forgotten genius. British Journal for the History of Science, 5(1):1–20, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025350.

Forbes:1971:BRPb

[For71a]  E. G. Forbes.  Book re-

view: Physics and Astronomy Cosmology: Theories of the Universe. By Jean Charon. Trans. by Patrick Moore. London: Weidenfeld and Nicolson. 1970. Pp. 256. 28 black-and-white and 27 colour illustrations. £1.75. British Journal for the History of Science, 5(4): 404–405, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025388.

Forbes:1971:BRPa

[For71b] Eric G. Forbes. Book review: Physical Sciences Early Solar Physics. By A. J. Meadows. Oxford: Pergamon Press. 1970. Pp. viii + 312. £1.75 (paperback). British Journal for the History of Science, 5 (3):302, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025330.

Forbes:1972:BRE

[For72] Eric G. Forbes. Book review: Eighteenth and Nineteenth Centuries: An Original Theory of the Universe. By Thomas Wright. Reprint with Introduction by Michael A. Hoskin. London: Macdonald, and New York: American Elsevier, 1971. Pp. xxxviii + 178. £10. British Journal for the History of Science, 6 (1):96, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025277.

Forbes:1973:CCB

[For73] Eric G. Forbes. The Crawford Collection of books and manuscripts on the history of astronomy, mathematics, etc., at the Royal Observatory, Edinburgh. British Journal for the History of Science, 6(4): 459–461, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025526.

Forbes:1974:BRS

[For74a] Eric G. Forbes. Book review: Sixteenth and Seventeenth Centuries: The Astronomical Revolution.

By Alexandre Koyré. Trans, by R. E. W. Maddison. Paris: Hermann, London: Methuen, and Ithaca, N.Y.: Cornell University Press, 1973. Pp. 531. £6.50. The Cosmology of Giordano Bruno. By Paul-Henri Michel. Trans, by R. E. W. Maddison. Paris: Hermann, London: Methuen, and Ithaca, N.Y.: Cornell University Press, 1973. Pp. 306. £4.50. British Journal for the History of Science, 7(3):293–294, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025543.

Forbes:1974:PMC

[For74b]  Eric G. Forbes.  Private

manuscript collections in Scotland. British Journal for the History of Science, 7(2):201– 202, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025225.

Forbes:1980:BRB

[For80] Eric G. Forbes. Book review: Bibliography Supplement to the Catalogue of the Crawford Library of the Royal Observatory, Edinburgh. Pp. xii + 112. Edited by Mary F. I. Smyth and Michael J. Smyth. Edinburgh: Royal Observatory, 1977. £5.00. British Journal for the History of Science, 13 (1):63–64, March 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026124.

Forrester:1982:BRB

[For82] John Forrester. Book review:

Biography Ronald W. Clark, Freud: the man and the cause. London: Jonathan Cape & Weidenfeld & Nicolson, 1980. Pp. xii + 652. £9.95. British Journal for the History of Science, 15(1):81–82, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025931.

Forbes:1985:BRG

[For85a] Eric G. Forbes. Book review: A. E. Gunther, An Introduction to the Life of the Rev. Thomas Birch D.D., F.R.S 1705–1766. Halesworth: The Halesworth Press, Suffolk, England 1984. Pp. x + 118. ISBN 0-9507276-1-X. £7.90. British Journal for the History of Science, 18(3):351– 352, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026392.

Force:1985:WWH

[For85b] James E. Force. William Whiston, honest Newtonian. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-26590-8. xxiii + 208 pp. LCCN BX5199.W52 F67

1985. URL http://www.loc.
gov/catdir/description/

cam022/84004316.html;
http://www.loc.gov/
catdir/samples/cam031/
84004316.html;   http://
www.loc.gov/catdir/toc/
cam029/84004316.html.

Forgan:1986:CIF

[For86] Sophie Forgan. Context, image and function: a preliminary enquiry into the architecture of scientific societies. British Journal for the History of Science, 19(1):89– 113, March 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026486.

Forgan:1987:BRI

[For87a] Sophie Forgan. Book review: Ian Inkster (ed.) The Steam Intellect Societies: Essays on Culture, Education and Industry, circa 1820–1914. University of Nottingham: Department of Adult Education, 1985. Pp. 203. ISBN 1-85401-008-9. £15.00 (plus £1.50, p. & p., available from publisher). British Journal for the History of Science, 20(1): 121–122, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026473.

Forrester:1987:BRN

[For87b]  John Forrester.  Book re-

view: Nikolas Rose. The Psychological Complex. Psychology, Politics and Society in England 1869–1939. London: Routledge and Kegan Paul, 1985. Pp. viii + 293. ISBN 0-7100-9809-1. £9.95 (paperback). British Journal for the History of Science, 20(1): 91–93, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026447.

Forrester:1987:BRW

[For87c]  John Forrester.  Book re-

view: William J. McGrath. Freud’s Discovery of Psychoanalysis. The Politics of Hysteria. Ithaca, New York: Cornell University Press, 1986. Pp. 336. ISBN 0-8014-1770-8. $27.50. British Journal for the History of Science, 20(2): 225–227, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026311.

Fox:1968:BDD

[Fox68a] Robert Fox. The background to the discovery of Dulong and Petit’s law. British Journal for the History of Science, 4 (1):1–22, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024983.

Fox:1968:BRS

[Fox68b] Robert Fox. Book review:

Scientific Societies The Lunar Society of Birmingham. University of Birmingham Historical Journal. Volume XI, no. 1 (1967). Pp. iv + 111. Plates. 15s. British Journal for the History of Science, 4 (2):175, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025148.

Fox:1969:BRE

[Fox69a] Robert Fox. Book review: Engineering at Cambridge University 1783–1965. By T. J. N. Hilken. London: Cambridge University Press. 1967. Pp. xi + 276. Figs. 45s. British Journal for the History of Science, 4(4):414, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025246.

Fox:1969:BRH

[Fox69b] Robert Fox. Book review: History of physics. Benjamin Thompson-Count Rumford. Count Rumford on the Nature of Heat. By Sanborn C. Brown. Pp. ix + 207. Oxford: Pergamon Press. 1967. 21s. British Journal for the History of Science, 4(3):290–291, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025481.

Fox:1970:BRP

[Fox70] Robert Fox. Book review: Physics The Collected Works of Count Rumford. Ed. by Sanborn C. Brown. Volume I, The Nature of Heat; Volume II, Practical Applications of Heat. Harvard University Press and Oxford University Press. 1968/ 1969. Pp. xi + 507; viii + 523. Figs. 95s. each. British Journal for the History of Science, 5(2):192, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025424.

Fox:1971:BRE

[Fox71a] Robert Fox. Book review:

Eighteenth Century: Martinus van Marum, Life and Work. Volume II. Ed. by R. J. Forbes. Haarlem: H. D. Tjeenk Willink. 1970. Pp. 401. f. 52. British Journal for the History of Science, 5(4):420– 421, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025407.

Fox:1971:BRF

[Fox71b] Robert Fox. Book review:

France Science in France in the Revolutionary Era. Described by Thomas Bugge. Ed. by Maurice P. Crosland. Cambridge, Mass., and London: M.I.T. Press. 1969. Pp. xiv + 239. £4.65. British Journal for the History of Science, 5(3): 306–307, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025335.

Fox:1974:BRN

[Fox74] Robert Fox. Book review: Nineteenth and Twentieth Centuries: The Sorcerer’s Apprentice. The French Scientist’s Image of German Science 1840–1919. By Harry W. Paul. Gainesville: University of Florida Press, 1972. Pp. viii + 86. $2. British Journal for the History of Science, 7(3): 301–302, November 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025549.

Fox:1979:BRB

[Fox79] Robert Fox. Book review: Bibliography and Bibliophily Isis Cumulative Bibliography. A Bibliography of the History of Science formed from Isis Critical Bibliographies 1-90, 1913– 1965. Volume iii: Subjects. Edited by Magda Whitrow. London: Mansell, in conjunction with the History of Science Society, 1976. Pp. xciv + 678. £28.00/$56.00. British Journal for the History of Science, 12(2):218, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025756.

Fox:1982:BRI

[Fox82] Robert Fox. Book review: Instruments Gerard L’E. Turner, Antique scientific instruments. Poole: Blandford Press, 1980. Pp. 168. £3.95/£2.95 (paperback). British Journal for the History of Science, 15(3): 310, November 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026227.

Fox:1984:SIS

[Fox84] Robert Fox. Science, industry, and the social order in Mulhouse, 1798–1871. British Journal for the History of Science, 17(2):127–168, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026549.

Fox:1987:PLE

[Fox87a] Christopher Fox, editor. Psychology and literature in the Eighteenth Century, volume 8 of AMS studies in the eighteenth century, 0196-6561. AMS Press, New York, NY, USA, 1987. ISBN 0-404-61474-4. xii + 372 + 7 pp. LCCN PR448.P75 P79 1987.

Fox:1987:BRA

[Fox87b] Robert Fox. Book review: Andre Grelon (ed.). Les Ingénieurs de la Crise. Titre et Profession entre les Deux Guerres. Paris: Éditions de l’École des Hautes Études en Sciences Sociales, 1986. Pp. 461. ISBN 2-7132-0875-0. 250 FF. British Journal for the History of Science, 20(3):364–365, July 1987.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026374.

Fox:1988:LSI

[Fox88] Christopher Fox. Locke and the Scriblerians: identity and consciousness in early Eighteenth-Century Britain. University of California Press, Berkeley, CA, USA, 1988. ISBN 0-520-05859-3. x + 174 pp. LCCN BF697 .F68

                                        1988. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0641/88014201-d.html.

Fox:1995:TCM

[Fox95] Robert Fox, editor. Technological change: methods and themes in the history of technology, volume 1 of Studies in the history of science, technology and medicin. Harwood Academic Publishers, Chur, Switzerland and New York, NY, USA, 1995. ISBN 3-7186-5792-9. vi + 271 pp. LCCN ????

Fox:1996:TCM

[Fox96] Robert Fox, editor. Technological change: methods and themes in the history of technology, volume 1 of Studies in the history of science, technology, and medicine. Harwood Academic Publishers, Chur, Switzerland and New York, NY, USA, 1996. ISBN 3-7186-5792-9. LCCN T15 .T417

                                        1996. URL http://www.loc.

gov/catdir/enhancements/
fy0653/99521773-d.html.

FabridePeiresc:1989:LCD

[FPLJ89]  Nicolas  Claude

Fabri de Peiresc, Cassiano Dal Pozzo, Jean-François Lhote, and Danielle Joyal. Lettres à Cassiano dal Pozzo, 1626– 1637. (French) [Letters to Cassiano dal Pozzo, 1626–1637], volume 3 of Amphion. Adosa, Clermont-Ferrand, France, 1989. ISBN 2-86639-079-2. 280 pp. LCCN DC121.8.P4 A4 1989.

Frankel:1975:C

[Fra75] Eugene Frankel. Coulomb. British Journal for the History of Science, 8(2):170–173, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025641.

Frangsmyr:1976:GIJ

[Frä76] Tore Frängsmyr. The geological ideas of J. J. Berzelius. British Journal for the History of Science, 9(2):228–236, July 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025809.

Frankel:1978:CMP

[Fra78a]  Eugene Frankel.  Career-

making in post-revolutionary France: the case of Jean-Baptiste Biot. British Journal for the History of Science, 11 (1):36–48, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025606.

Frankel:1978:AHC

[Fra78b] Henry Frankel. Arthur Holmes and continental drift. British Journal for the History of Science, 11(2):130–150, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025726.

Franklin:1986:NE

[Fra86] Allan Franklin. The neglect of experiment. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-32016-X. xii + 290 pp. LCCN Q182.3 .F73

1986. URL http://www.loc.
gov/catdir/description/
cam023/86002604.html;
http://www.loc.gov/
catdir/samples/cam031/
86002604.html;   http://
www.loc.gov/catdir/toc/
cam024/86002604.html.

Franklin:1987:BRP

[Fra87a] Allan Franklin. Book review: Peter Achinstein & Owen Hannaway (eds). Observation, Experiment, and Hypothesis in Modern Physical Science. Cambridge, MA: MIT Press, 1985. Pp. x + 379. ISBN 0-262-01083-6. £39.25. British Journal for the History of Science, 20(1): 117–118, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026469.

Franklin:1987:BRT

[Fra87b] Allan Franklin. Book review:

Trevor Pinch. Confronting Nature. Dordrecht: D. Reidel, 1986. Pp. xi + 268. ISBN 90-227-2224-2. £34.50, $49.50. British Journal for the History of Science, 20(2):234– 235, April 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026320.

Fraser:1987:BRS

[Fra87c] Craig Fraser. Book review:

Steven B. Engelsman. Families of Curves and the Origins of Partial Differentiation. Amsterdam: North-Holland, 1984. Pp. x + 238. ISBN 0-444-86897-6. US $29.00, Dfl. 85.00. British Journal for the History of Science, 20(2):238, April 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026324.

Frangsmyr:1989:SSR

[Frä89]  Tore  Frängsmyr,

editor. Science in Sweden: the Royal Swedish Academy of Sciences, 1739–1989. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1989. ISBN 0-88135-092-3. viii + 291 + 34 pp. LCCN Q64.S937 S25 1989.

Frangsmyr:1990:SHR

[Frä90]  Tore  Frängsmyr,

editor. Solomon’s house revisited: the organization and institutionalization of science; [held in Stockholm, Sweden, August 14–18, 1989]. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1990. ISBN 0-88135-066-4. LCCN Q10 .N63 1989.

French:1970:BRB

[Fre70] R. K. French. Book review: Biology and Medicine The Cole Library of Early Medicine and Zoology. By Nellie B. Eales. Alden Press for the Library, University of Reading. 1969. Pp. xiv + 425. 1 plate. Price not stated. British Journal for the History of Science, 5(2): 198–199, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025429.

French:1974:BRG

[Fre74] R. K. French. Book review: Greek Science The Heart and the Vascular System in Ancient Greek Medicine from Alcmaeon to Galen. By C. R. S. Harris. Oxford: Clarendon Press: Oxford University Press, 1973. Pp. x + 474. £15.00. British Journal for the History of Science, 7(2):187, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025212.

French:1978:BRM

[Fre78]  Roger French.  Book re-

view: Medicine Boerhaave’s Men at Leyden and After. By E. Ashworth Underwood. Edinburgh: Edinburgh University Press, 1977. Pp. vii + 227. £8.00. British Journal for the History of Science, 11 (3):281, November 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025782.

Freedman:1979:FSP

[Fre79] Michael I. Freedman. Frederick Soddy and the practical significance of radioactive matter. British Journal for the History of Science, 12(3): 257–260, November 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026000.

French:1981:BM

[Fre81a] R. K. French. Back matter. British Journal for the History of Science, 14(3):304, November 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025997.

French:1981:BRS

[Fre81b]  R. K. French.  Book re-

view: Short Notice Barbers and Barber-Surgeons of London: A History of the Barbers’ and Barber-Surgeons’ Companies. By Jessie Dobson and R. Milnes Walker. Oxford: Blackwell Scientific Publications for the Worshipful Company of Barbers, 1979. Pp.xix + 171. £9.50. British Journal for the History of Science, 14(3): 296, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025994.

French:1982:BJM

[Fre82] Roger French. Bibliography: A. J. Meadows (ed), Development of science publishing in Europe. Amsterdam, New York, Oxford: Elsevier Science Publishers, 1980. Pp. ix + 269. British Journal for the History of Science, 15(1):84–85, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025933.

French:1986:BRM

[Fre86] Roger French. Book review: Michael W. Dols & Adil S. Gamal, eds. Medieval Islamic Medicine. Ibn Ridwan’s Treatise ‘On the Prevention of Bodily Ills in Egypt’. Translated and introduced by M. W. Dols, with Arabic text by A. S. Gamal. Berkeley and London: University of California Press, 1984. Pp. xv + 186 + 63. ISBN 0-420-04836-9. $28. British Journal for the History of Science, 19(2):211–212, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026600.

FresquetFebrer:1990:FMA

[Fre90] José Luis Fresquet Febrer.

Francisco Méndez Álvaro (1806–1883) y las ideas sanitarias del liberalismo moderado. (Spanish) [Francisco Méndez Álvaro (1806–1883) and the sanitary ideas of moderate liberalism ], volume 14 of Textos Clásicos espaänoles de la salud pública. Ministerio de sanidad y consumo, Secretaría general técnica Publicaciones, documentación y biblioteca, Madrid, Spain, 1990. ISBN 84-7670-217-5. 212 pp. LCCN ????

Freudenthal:1992:SGF

[Fre92] Gad Freudenthal, editor. Studies on Gersonides: a Fourteenth-Century Jewish philosopher-scientist, volume 36 of Collection de travaux de l’Acadèmie internationale d’histoire des sciences, 0169-7897. E. J. Brill, Leiden, The Netherlands, 1992. ISBN 90-04-09641-8 (hardcover). xxi + 422 pp. LCCN Q143.L57 S78 1992.

French:1994:ANH

[Fre94] R. K. (Roger Kenneth) French.

Ancient natural history: histories of nature. Sciences of antiquity. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1994. ISBN 0-415-08880-1 (hardcover), 0-415-11545-0 (paperback). xxii + 357 pp. LCCN 509; 606; QH15 .F74 1994; M96.F00079; 508.093.

Freudenthal:1995:ATM

[Fre95]  Gad Freudenthal.  Aristo-

tle’s theory of material substance: heat and pneuma, form and soul. Clarendon Press, New York, NY, USA, 1995. ISBN 0-19-824093-7. xi + 235 pp. LCCN B491.S8 F74

                                        1995. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0605/94048637-d.html.

Friday:1974:BRE

[Fri74a] J. R. Friday. Book review: Eighteenth and Nineteenth Centuries: Scientist Extraordinary. The Life and Scientific Work of Thomas Henry Huxley, 1825–1895. By Cyril Bibby. Oxford: Pergamon Press, 1972. Pp. xii + 208. £2.95. British Journal for the History of Science, 7(2):198, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025223.

Friday:1974:MIM

[Fri74b] James Friday. A microscopic incident in a monumental struggle: Huxley and antibiosis in 1875. British Journal for the History of Science, 7(1):61– 71, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025176.

Frieden:1981:RPE

[Fri81] Nancy Mandelker Frieden.

Russian physicians in an era of reform and revolution, 1856– 1905. Princeton University Press, Princeton, NJ, USA, 1981. ISBN 0-691-05335-9, 0-691-05335-0. xvii + 378 pp. LCCN R532 .F74. US$32.50, US$32.50.

Friedel:1983:PPM

[Fri83] Robert D. (Robert Douglas)

Friedel. Pioneer plastic: the making and selling of celluloid. University of Wisconsin Press, Madison, WI, USA, 1983. ISBN 0-299-09170-8. xix + 153 pp. LCCN TP1180.C5 F74 1983. US$19.95.

Friedel:1987:BRG

[Fri87] Robert Friedel. Book review:

George Wise. Willis R. Whitney, General Electric, and the Origins of U.S. Industrial Research. New York: Columbia University Press, 1985. Pp. 375. ISBN 0-231-06016-5. $29.00. British Journal for the History of Science, 20(1):114– 115, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026467.

Friedman:1989:AWV

[Fri89] Robert Marc Friedman. Appropriating the weather: Vilhelm Bjerknes and the construction of a modern meteorology. Cornell University Press, Ithaca, NY, USA, 1989. ISBN 0-8014-2062-8. xx + 251 pp. LCCN QC858.B53 F75 1989.

Fritscher:1991:VGB

[Fri91] Bernhard Fritscher. Vulkanismusstreit und Geochemie: die Bedeutung der Chemie und des Experiments in der Vulkanismus-Neptunismus-Kontroverse. (German) [Volcanism controversy and geochemistry: the meaning of chemistry and experiments in the Volcanism–Neptune controversy], volume 25 of Boethius. F. Steiner Verlag, Stuttgart, Germany, 1991. ISBN 3-515-05865-6. viii + 346 pp. LCCN QE522 .F75 1991.

Fruton:1982:BBH

[Fru82] Joseph S. (Joseph Stewart) Fruton. A bio-bibliography for the history of the biochemical sciences since 1800. American Philosophical Society, Philadelphia, PA, USA, 1982. ISBN 0-87169-983-4. xiii + 885 pp. LCCN Z5524.B54 F78 1982.

Fruton:1990:CSS

[Fru90] Joseph S. (Joseph Stewart) Fruton. Contrasts in scientific style: research groups in the chemical and biochemical sciences, volume 191 of Memoirs series / American Philosophical Society, 0065-9738. American Philosophical Society, Philadelphia, PA, USA, 1990. ISBN 0-87169-191-4. ix + 473 pp. LCCN Q11 .P612 vol. 191.

Fenton:1984:LRS

[FS84] Alexander Fenton and Geoffrey Stell, editors. Loads and roads in Scotland and beyond: land transport over 6000 years. A Scottish history and culture paperback. John Donald, Edinburgh, Scotland, 1984. ISBN 0-85976-107-X (paperback). vii + 144 pp. LCCN HE363.S3 L6. US$8.50. URL http://

catalog.hathitrust.org/
api/volumes/oclc/
11225466.html.

Figurovskii:1988:APE

[FS88] Nikolai Aleksandrovich Figurovskii and I. Urii Ivanovich Solovev. Aleksandr Porfirevich Borodin: a chemist’s biography. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 3-540-17888-0, 0-387-17888-0 (US). LCCN ???? Translated from the Russian by Charlene Steinberg and George B. Kauffman; foreword by Martin D. Kamen.

Fisch:1991:WWC

[FS91] Menachem Fisch and Simon Schaffer, editors. William Whewell, a composite portrait. Clarendon Press, New York, NY, USA, 1991. ISBN 0-19-824900-4. xiv + 403 pp. LCCN B1612.W54 W55

                                        1991. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0605/90041928-d.html.

Faraday:1991:MFC

[FTG91] Michael Faraday, Ryan D. Tweney, and David Gooding. Michael Faraday’s ’Chemical notes, hints, suggestions, and objects of pursuit’ of 1822, volume 17 of IEE history of technology series. P. Peregrinus in association with the Science Museum, London, UK, 1991. ISBN 0-86341-255-6. xvii + 152 pp. LCCN QD3 .F22 1991.

Fujii:1986:BPC

[Fuj86] Kiyohisa Fujii. The Berthollet– Proust controversy and Dalton’s Chemical Atomic Theory 1800–1820. British Journal for the History of Science, 19(2):177–200, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026591.

Fujimura:1996:CSS

[Fuj96] Joan H. Fujimura. Crafting science: a sociohistory of the quest for the genetics of cancer. Harvard University Press, Cambridge, MA, USA, 1996. ISBN 0-674-17553-0 (hardcover). x + 322 pp. LCCN RC268.415 .F84 1996.

Furley:1985:BRG

[Fur85] David Furley. Book reviews:

G. E. R. Lloyd, Science, Folklore and Ideology: Studies in the Life Sciences in Ancient Greece. Cambridge: Cambridge University Press, 1983. Pp. xi + 260. ISBN 0-521-25314-4. £25.00, $49.50. Paperback ISBN 0-521-27307-2. £8.95, $16.95. British Journal for the History of Science, 18 (1):91–93, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026275.

Furley:1987:GC

[Fur87] David J. Furley. The Greek cosmologists. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-33328-8 (vol. 1). ???? pp. LCCN BD495 .F87

1987. URL http://www.loc.
gov/catdir/description/
cam023/86026384.html;
http://www.loc.gov/
catdir/enhancements/
fy0632/86026384-t.html.

Furth:1988:SFP

[Fur88] Montgomery Furth. Substance, form, and psyche: an Aristotelean metaphysics. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-34143-4. xiv + 300 pp. LCCN B491.M4 F87

1988. URL http://www.loc.
gov/catdir/description/
cam023/87023872.html;

                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0707/87023872-t.html.

Furley:1989:CPE

[Fur89]  David J. Furley.  Cosmic

problems: essays on Greek and Roman philosophy of nature. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-33330-X. xiv + 258 pp. LCCN BD581 .F87

                                        1989. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/88011120.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam023/
                                        88011120.html.

G:1962:XIC

[G.62]  F. G.  Xth International

Congress of the History of Science. British Journal for the History of Science, 1(2): 186, December 1962. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025133.

Gigerenzer:1989:ECH

[G^{+}89] Gerd Gigerenzer et al., editors. The empire of chance: how probability changed science and everyday life. Ideas in context. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-33115-3. xvii + 340 pp. LCCN QA273 .E57

                                        1989. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/88016928.html;
                                        http://www.loc.gov/

catdir/toc/cam022/
88016928.html.

Gallouet:2009:LDS

[G^{+}09] Catherine Gallouët et al., editors. L’Afrique du siècle des lumières: savoirs et représentations. (French) [Africa of the century of light: knowledge and representations], volume 2009:05 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2009. ISBN 0-7294-0959-7. xxix + 307 pp. LCCN PQ265 .A37 2009.

Garber:1998:CHS

[GA98] Daniel Garber and Michael Ayers, editors. The Cambridge history of Seventeenth-Century philosophy. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-58864-2 (set). xvii + 1616 pp. LCCN B801 .C35 1998. URL http:

//www.loc.gov/catdir/
description/cam027/
96025475.html;   http://
www.loc.gov/catdir/toc/
cam025/96025475.html.

Gabbey:1975:BRS

[Gab75] Alan Gabbey. Book review:

Seventeenth Century: The Mathematical Career of Pierre de Fermat (1601–1665). By Michael Sean Mahoney. Princeton, N.J.: Princeton University Press, 1973. Pp. xviii + 419. £9.30. British Journal for the History of Science, 8 (1):81–84, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025827.

Gabbey:1978:EFD

[Gab78] Alan Gabbey. The English fortunes of Descartes. British Journal for the History of Science, 11(2):159–164, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025729.

Gabbay:1983:BRMb

[Gab83a] John Gabbay. Book review: Medical Sciences: J. Worth Estes, Hall Jackson and the purple foxglove: medical practice and research in revolutionary America 1760–1820, Hanover, New Hampshire: University Press of New England, 1979. Pp. xvii + 291. $15.00. British Journal for the History of Science, 16(1): 89–91, March 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026100.

Gabbay:1983:BRMa

[Gab83b]  John  Gabbay.

Book review: Medical Sciences: Toby Gelfand, Professionalizing modern medicine: Paris surgeons and medical science and institutions in the eighteenth century. Westport, Connecticut and London: Greenwood Press, 1980. Pp. xviii + 271. (No price stated). British Journal for the History of Science, 16(1):86–88, March 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026098.

Gabbey:1983:BRL

[Gab83c]  Alan Gabbey.  Book re-

view: Lives and Works Stillman Drake, Galileo, Oxford: Oxford University Press, 1980. Pp. x + 100. £0.95 (paperback). British Journal for the History of Science, 16(2):202–203, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026245.

Gabbey:1988:BRT

[Gab88]  Alan Gabbey.  Book re-

views: Tom Sorell. Descartes. Oxford, New York: Oxford University Press, 1987. Pp.

[viii] + 120. ISBN 0-19-287636-8. £9.95 (hardcover). ISBN 0-19-287635-X. £2.95 (paperback). (‘Past Masters’ series.). British Journal for the History of Science, 21(1):125– 126, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026880.

Gale:1972:TM

[Gal72] A. J. V. Gale. Thomas Martin.

British Journal for the History of Science, 6(1):105, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025284.

Galison:1987:HEE

[Gal87] Peter Galison. How experiments end. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-27915-4 (paperback), 0-226-27914-6. xii + 330 pp. LCCN QC6 .G22 1987. US$12.00 (est.), US$36.00 (est.). URL http:/

                                        /www.loc.gov/catdir/
                                        description/uchi052/
                                        87005928.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0642/87005928-b.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0642/87005928-t.html.

Gandt:1995:FGN

[Gan95] François De Gandt. Force and geometry in Newton’s Principia. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03367-6. xiv + 296 pp. LCCN QA803 .G3613 1995. URL http://

                                        www.loc.gov/catdir/
                                        description/prin031/
                                        94039315.html;   http://
                                        www.loc.gov/catdir/toc/
                                        prin031/94039315.html.

Garbrecht:1987:HHR

[Gar87] Günther Garbrecht, editor. Hydraulics and hydraulic research: a historical review. A. A. Balkema, Rotterdam, The Netherlands, 1987. ISBN 90-6191-621-6. LCCN TC15 .H94 1987.

Garber:1992:DMP

[Gar92]  Daniel Garber.  Descartes’

metaphysical physics. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-28217-1 (hardcover), 0-226-28219-8 (paperback). xiv + 389 pp. LCCN QC6 .G25 1992. US$60.00, US$23.95.

URL http://www.loc.gov/
catdir/description/
uchi051/91030190.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/91030190-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/91030190-b.html.

Gaukroger:1982:MIE

[Gau82] Stephen Gaukroger. The metaphysics of impenetrability: Euler’s conception of force. British Journal for the History of Science, 15(2):132–154, July 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025964.

Gaukroger:1983:BRS

[Gau83] S. Gaukroger. Book review:

Science and Society Larry Laudan, Science and hypothesis: historical essays on scientific methodology. Dordrecht, Boston and London: D. Reidel, 1981. Pp. x + 258. Dfl.80.00/ $34.95; Dfl.37.50/$14.95 (paperback). British Journal for the History of Science, 16(3): 311–312, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026179.

Gauld:1992:HH

[Gau92] Alan Gauld. A history of hypnotism. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-30675-2 (hardcover). xvii + 738 pp. LCCN BF1125 .G38

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        cam025/92005106.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam027/
                                        92005106.html.

Gayon:1992:DAD

[Gay92]  Jean Gayon.  Darwin et

l’après Darwin: une histoire de l’hypothèse de sélection naturelle. (French) [Darwin and post-Darwin: a history of the hypothesis of natural selection]. Histoire des idées, théorie politique et recherches en sciences sociales. Editions Kimé, Paris, France, 1992. ISBN 2-908212-14-5. 453 pp. LCCN QH375 .G38 1992. 250F.

Galilei:1995:DSM

[GB95]  Galileo  Galilei  and

Antonio Beltrán. Diálogo sobre los dos máximos sistemas del mundo ptolemaico y copernicano. (Spanish) [Dialog on the two maximum systems of the world: Ptolemaic and Copernican]. Alianza, Madrid, Spain, 1995. ISBN 84-206-9412-6. lxxxiii + 417 pp. LCCN QB41 .G129418 1995.

Greenaway:1985:OEF

[GD85] Frank Greenaway and H. T. Dickinson. Obituary: Eric Forbes: 1933–1984. British Journal for the History of Science, 18(3): 341, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026385.

Geddes:1971:BRT

[Ged71] W. K. E. Geddes. Book review: Technology A History of the Marconi Company. By W. J. Baker. London: Methuen. 1970. Pp. 414. Plates. £5. British Journal for the History of Science, 5(4): 412–413, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025398.

Geison:1995:PSL

[Gei95] Gerald L. Geison. The private science of Louis Pasteur. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03442-7. xiv + 378 + 16 pp. LCCN Q143.P2

G35 1995. URL http://www.
loc.gov/catdir/
description/prin021/
94035338.html;   http://
www.loc.gov/catdir/toc/
prin031/94035338.html.

Gelfand:1976:BRL

[Gel76]  Toby Gelfand.  Book re-

view: Life Sciences: Biology of Man in History. Selections from the ‘Annales’. Ed. by Robert Forster and Orest Ranum. Trans, by Elborg Forster and Patricia M. Ranum. Baltimore and London: Johns Hopkins University Press, 1975. Pp. x + 205. £6.60; £1.65 (paperback). British Journal for the History of Science, 9(3):331– 333, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025694.

Greene:1983:LBH

[GEMM83] Edward Lee Greene, Frank N. Egerton, Robert P. (Robert Patrick) McIntosh, and Rogers Mc-Vaugh. Landmarks of botanical history. Stanford University Press, Stanford, CA, USA, 1983. ISBN 0-8047-1075-9. x + 1139 pp. LCCN QK21.E85 G74 1983.

George:1966:BRH

[Geo66] Wilma George. Book review: History of Biological Sciences and Medicine Georges Cuvier, Zoologist: A Study in the History of Evolution Theory. By William Coleman. Pp. 212. Harvard University Press; London: Oxford University Press, 1964. 40s. net. British Journal for the History of Science, 3(1):90, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025119.

George:1978:BM

[Geo78a] Wilma George. Back matter. British Journal for the History of Science, 11(3):304, November 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025795.

George:1978:BRN

[Geo78b]  Wilma George.  Book re-

view: Nineteenth and Twentieth Centuries: Folia Baeriana I. Compiled by M. Valt and T. Ilomets, and edited by V. Kaavere and T. Sutt. Tallinn: Valgus, 1975. Pp. 181. 1.16 roubles. British Journal for the History of Science, 11(3):294– 296, November 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025792.

George:1987:BRN

[Geo87]  Wilma George.  Book re-

view: Nicholas Russell, Like Engend’ring Like: Heredity and Animal Breeding in Early Modern England. Cambridge: Cambridge University Press, 1986. Pp. ix + 271. ISBN 0-521-30657-4. £27.50, $47.50. British Journal for the History of Science, 20(4):477– 478, October 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026420.

Gerstner:1971:RJH

[Ger71] Patsy A. Gerstner. The reaction to James Hutton’s use of heat as a geological agent. British Journal for the History of Science, 5(4):353–362, December 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025379.

Gerl:1989:TAR

[Ger89] Armin Gerl. Trigonometrischastronomisches Rechnen kurz vor Copernicus: der Briefwechsel Regiomontanus-Bianchini. (German) [Trigonometricastronomical counting shortly before Copernicus: exchange of letters with Regiomontanus Bianchini], volume 21 of Boethius. F. Steiner Verlag, Stuttgart, Germany, 1989. ISBN 3-515-05373-5. 357 pp. LCCN QB145 .G47 1989.

Geuter:1992:PPN

[Geu92] Ulfried Geuter. The professionalization of psychology in Nazi Germany. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-33297-4 (hardcover). xxi + 335 pp. LCCN BF108.G3 G4813 1992. URL http://

www.loc.gov/catdir/

description/cam024/
91026576.html;   http://
www.loc.gov/catdir/toc/
cam029/91026576.html.

Godel:1986:CW

[GF^{+}03] Kurt Gödel, Solomon Feferman, et al. Collected works. Clarendon Press, New York, NY, USA, 1986–2003. ISBN 0-19-503964-5 (vol. 1), 0-19-503972-6 (vol. 2), 0-19-507255-3 (vol. 3), 0-19-850073-4 (vol. 4), 0-19-850075-0 (vol. 5). ???? pp. LCCN QA9 .G5313

1986. URL http://www.loc.
gov/catdir/enhancements/
fy0603/85015501-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0603/85015501-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0725/85015501-b.html.

Gardiner:1965:AMA

[GG65] K. R. Gardiner and D. L. Gardiner. Andre-Marie Ampère and his English acquaintances. British Journal for the History of Science, 2:235–245, 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Grattan-Guinness:1970:BRM

[GG70] Ivor Grattan-Guinness. Book review: Mathematics: Eléments d’Histoire des Mathematiques. By N. Bourbaki. Deuxième édition revue, corrigée, augmentée. (Histoire de la Pensée, No. IV.) Paris:

Hermann. 1969. Pp. 323. 36 francs. British Journal for the History of Science, 5(2): 190–191, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025422.

Grattan-Guinness:1971:BRM

[GG71] I. Grattan-Guinness. Book review: Mathematics: Development of Modern Mathematics. By J. M. Dubbey. London: Butterworths, 1970. Pp. vii + 145. £2.50. British Journal for the History of Science, 5(4):401, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025384.

Grattan-Guinness:1972:BRM

[GG72] I. Grattan-Guinness. Book review: Mathematics: Carl Friedrich Gauss. A Biography. By Tord Hall Albert Froderberg. cambridge, Mass., and London: M.I.T. Press, 1970. Pp. xii + 176. $7.95. British Journal for the History of Science, 6(2):207–208, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025296.

Grattan-Guinness:1974:BRM

[GG74a] I. Grattan-Guinness. Book review: Mathematics: The Nonmathernatical Foundations of Mathematics. An Account of the Desire of Modern Mathematics to Transcend Empiricism and its Consequent Toleration of the Greatest Intellectual Scandal of the Nineteenth and Twentieth Centuries. By R. J. Diamond. Pasadena, California: Nambor Press, 1971. Pp. v + 108. No price stated. British Journal for the History of Science, 7 (2):185–186, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025210.

Grattan-Guinness:1974:EMH

[GG74b] I. Grattan-Guinness. Ernst Mach: His work, life and influence. British Journal for the History of Science, 7(1): 75–79, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025178.

Grattan-Guinness:1975:BRMa

[GG75a] I. Grattan-Guinness. Book review: Mathematics: A Source Book in Classical Analysis. Ed. by G. Birkhoff, with the assistance of U. Merzbach. Cambridge, Mass: Harvard University Press, 1973. Pp. xii + 470. £12.50. British Journal for the History of Science, 8 (2):176–177, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025644.

Grattan-Guinness:1975:BRMb

[GG75b] I. Grattan-Guinness. Book review: Mathematics: De l’infini mathématique (Paris, 1896.) By Louis Couturat. Reprint. Paris: Albert Blanchard, 1973. Pp. xxiv + 668. 56 francs. British Journal for the History of Science, 8(3):251–252, November 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025563.

Grattan-Guinness:1975:BRN

[GG75c] I. Grattan-Guinness. Book review: Nineteenth and Twentieth Centuries: Hermann Minkowski. Briefe an David Hubert. Ed. by L. Rüdenberg and H. Zassenhaus. Berlin: Springer-Verlag, 1973. Pp. 165. DM 32; $11.90. British Journal for the History of Science, 8(1):91, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025835.

Grattan-Guinness:1977:BRMa

[GG77a] I. Grattan-Guinness. Book review: Mathematics: Einleitung zur Grössenlehre und erste Begriffe der allgemeinen Grössenlehre. By Bernard Bolzano. Ed. by Jan Berg. Bernard Bolzano Gesamtausgabe, Reihe II (Nachlass), Bd. 7. Stuttgart-Bad Cannstatt: Friedrich Frommann Verlag (Gunther Holzboog), 1975. Pp. 297, DM 190 (or, with orders for the whole edition, DM. 170). British Journal for the History of Science, 10(1):73–74, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025587.

Grattan-Guinness:1977:BRMb

[GG77b] I. Grattan-Guinness. Book review: Mathematics: La logique de Charles Sanders Peirce. De l’algèbre aux graphes. By Pierre Thibaud. Aix en Provence: Université de Provence, 1975. Pp. 184. 50 francs. Men and Institutions in American Mathematics. Edited by J. Dalton Tarwater, John T. White and John D. Miller. Lubbock, Texas: Texas Tech. University, 1976. Pp. 136. $5.00. British Journal for the History of Science, 10(3):255– 256, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025847.

Grattan-Guinness:1979:BMT

[GG79a]  I.  Grattan-

Guinness. Babbage’s mathematics in its time. British Journal for the History of Science, 12(1):82–88, March 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025665.

Grattan-Guinness:1979:LPM

[GG79b] I. Grattan-Guinness. Lakatos and the philosophy of mathematics and science: On Popper’s philosophy and its prospects. British Journal for the History of Science, 12(3):317–337, November 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026007.

Grattan-Guinness:1981:BRP

[GG81] I. Grattan-Guinness. Book review: Philosophy of Science Basis und Deduktion: Studien zur Entstehung und Bedeutung der Theorie der axiomatischen Methode bei J. H. Lambert (1728–1777). By Gereon Wolters. Berlin and New York: Walter de Gruyter, 1980. pp. xiii + 194. DM78. British Journal for the History of Science, 14(3):284, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025985.

Grattan-Guinness:1984:BRH

[GG84a] I. Grattan-Guinness. Book review: History of Mathematical Sciences: Charles S. Peirce, Writings, vol. 1, 1857–1866. Bloomington, Indiana: Indiana University Press, 1982. Pp. xxxvii + 698. £19.50. British Journal for the History of Science, 17(2):235–236, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026568.

Grattan-Guinness:1984:BRN

[GG84b] I. Grattan-Guinness. Book review: National Traditions in Science Geoffrey Howson, A history of mathematics education in England. Cambridge: University Press, 1982. Pp. x + 294. ISBN 0-521-24206-1. £25. British Journal for the History of Science, 17(1):97– 98, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026033.

Grattan-Guinness:1985:BRB

[GG85] I. Grattan-Guinness. Book review: Brian Inglis, Science and parascience. A history of the paranormal, 1914–1939. London: Hodder and Stoughton, 1984. Pp. 382. ISBN 0-340-26325-3. £12.95. British Journal for the History of Science, 18(2):248–249, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026351.

Grattan-Guinnes:1986:TEP

[GG86a] I. Grattan-Guinnes. On the transformation of the École Polytechnique archives. British Journal for the History of Science, 19(1):45–50, March 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026483.

Grattan-Guinness:1986:CXT

[GG86b] I. Grattan-Guinness. Collections XII: On the transformation of the École Polytechnique archives. British Journal for the History of Science, 19(1(61)):45–50, 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Grattan-Guinness:1987:BRF

[GG87] I. Grattan-Guinness. Book review: S. Feferman (ed-in-chief). Kurt Gödel. Collected Works. Volume 1, Publications 1929–1936. Oxford and New York: Oxford University Press, 1986. xvi + 474 pp. ISBN 0-19-503964-5. £25.00. British Journal for the History of Science, 20(2):223, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026309.

Grattan-Guinness:1989:BRR

[GG89] I. Grattan-Guinness. Book review: Roger Herz-Fischler. A Mathematical History of Division in Extreme and Mean Ratio. Waterloo, Canada: Wilfred Laurier University Press, 1987. Pp. xvi + 191. ISBN 0-88920-152-8. No price given. British Journal for the History of Science, 22(1):84–85, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026684.

Gavroglu:1991:TMK

[GG91] Kostas Gavroglu and Yorgos Goudaroulis, editors. Through measurement to knowledge: the selected papers of Heike Kamerlingh Onnes 1853–1926, volume 124 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346. cxv + 570 pp. LCCN Q174 .B67 vol.

124.  URL https://link.
springer.com/book/10.
1007/978-94-009-2079-8.

Gottfried:1992:TUP

[GG92] Lew Gottfried and Fran Gottfried. Traveling uncharted paths: the tale of two retirees and The Peace Corps. Fairway Press, Lima, OH, USA, 1992. ISBN 1-55673-455-7. 280 pp. LCCN HC60.5 .G68 1992.

Guicciardini:1990:GSO

[GGC90] Francesco Guicciardini, Lodovico Guicciardini, and Raffaella Castagnola. I Guicciardini e le scienze occulte: l’oroscopo di Francesco Guicciardini: lettere di alchimia, astrologia e cabala a Luigi Guicciardini. (Italian) [The hidden Guicciardini and science: the horoscope of Francisco Guicciardini: letters on alchemy, astrology and cabala to Luigi Guicciardini ], volume 19 of Studi e testi / Istituto Nazionale di Studi sul Rinascimento, 0394-4409. L. S. Olschki, Firenze, Italy, 1990. ISBN 88-222-3761-7. vii + 397 + 8 pp. LCCN BF1728.G84 G8425 1990.

Goldstein:1996:EMH

[GGR96] Catherine Goldstein, Jeremy Gray, and Jim Ritter, editors. L’Europe mathématique: histoires, mythes, identités. (French) [Mathematical Europe: history, myth, identity], Librairie européenne des idées. Editions de la Maison des sciences de l’homme, Paris, France, 1996. ISBN 2-7351-0685-3. LCCN ????

GuytondeMorveau:1994:SCD

[GGSGB94] Louis-Bernard Guyton de Morveau, Emmanuel Grison, Michelle Sadoun-Goupil, and Patrice Bret. A scientific correspondence during the chemical revolution: Louis-Bernard Guyton de Morveau and Richard Kirwan, 1782– 1802, volume 17 of Berkeley papers in history of science. Office for History of Science and Technology, University of California at Berkeley, Berkeley, CA, USA, 1994. ISBN 0-918102-21-9. vi + 257 pp. LCCN QD22.G94 A4 1994.

Gaposchkin:1984:CPG

[GH84] Cecilia Helena Payne Gaposchkin and Katherine Haramundanis. Cecilia Payne-Gaposchkin: an autobiography and other recollections. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25752-2. vi + 269 pp. LCCN QB36.G37 A33 1984.

Galison:1992:BSG

[GH92]  Peter  Louis  Gal-

ison and Bruce William Hevly, editors. Big science: the growth of large-scale research. Stanford University Press, Stanford, CA, USA, 1992. ISBN 0-8047-2335-4, 0-8047-1879-2. xi + 392 pp. LCCN ????

Geison:1993:RSH

[GH93] Gerald L. Geison and Frederic Lawrence Holmes, editors. Research schools: historical perspectives, volume 2nd ser., vol. 8 (1993) of Osiris. Dept. of History and Sociology of Science, University of Pennsylvania, Philadelphia, PA, USA, 1993. ISBN 0-226-28545-6 (hardcover), 0-226-28546-4 (paperback). ISSN 0369-7827 (print), 1933-8287 (electronic). LCCN LB2372.3 .R47 1993.

Gaposchkin:1996:CPG

[GH96] Cecilia Helena Payne Gaposchkin and Katherine Haramundanis. Cecilia Payne-Gaposchkin: an autobiography and other recollections. Cambridge University Press, Cambridge, UK, second edition, 1996. ISBN 0-521-48251-8 (hardcover), 0-521-48390-5 (paperback). xxii + 277 pp. LCCN QB36.G37

                                        A33 1996. URL http://www.
                                        loc.gov/catdir/
                                        description/cam027/
                                        95044001.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam021/95044001.html.

Gibbs:1984:RE

[Gib84] G. C. Gibbs. The Radical Enlightenment. British Journal for the History of Science, 17 (1):67–81, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026019.

Giddings:1982:BRB

[Gid82] Luther Val Giddings. Book review: Biography Bentley Glass (ed), The roving naturalist: travel letters of Theodosius Dobzhansky. Philadelphia: The American Philosophical Society, 1980. Pp. vii + 327. $8.00. British Journal for the History of Science, 15 (1):82–84, March 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025932.

Giere:1988:ESC

[Gie88] Ronald N. Giere. Explaining science: a cognitive approach. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1988. ISBN 0-226-29205-3, 0-226-29206-1 (paperback). xxi + 321 pp. LCCN M88.E00693; 500; 501; Q175 .G4889 1988; J.c.81; 0187.

Gillispie:1983:MBI

[Gil83] Charles Coulston Gillispie. The Montgolfier brothers and the invention of aviation, 1783–1784: with a word on the importance of ballooning for the science of heat and the art of building railroads. Princeton University Press, Princeton, NJ, USA, 1983. ISBN 0-691-08321-5. xi + 210 pp. LCCN TL617.M66 G48 1983. US$35.00.

Gilley:1984:BRS

[Gil84] Sheridan Gilley. Book review: Science in Culture: Michael Heyd, Between Orthodoxy and the Enlightenment. Jean-Robert Chouet and the Introduction of Cartesian Science in the Academy of Geneva. (International Archives of the History of Ideas 96). The Hague and Jerusalem: Martinus Nijhoff and the Magnes Press, 1982. Pp. xiv + 308. ISBN 90-247-2508-9. Dfl. 150. British Journal for the History of Science, 17(2):246, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026579.

Gillespie:1988:BRM

[Gil88] Richard Gillespie. Book review:

Michael M. Sokal (ed.). Psychological Testing and American Society, 1890–1930. New Brunswick, NJ: Rutgers University Press, 1987. Pp. xi + 205. ISBN 0-8135-1193-3. $28.00. British Journal for the History of Science, 21(1): 118–119, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026874.

Gingerich:1974:BRS

[Gin74] Owen Gingerich. Book review: Sixteenth and Seventeenth Centuries: The Vortex Theory of Planetary Motions. By E. J. Aiton. London: Macdonald, and New York: American Elsevier, 1972. Pp. x + 282. £6. British Journal for the History of Science, 7(1): 91–92, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025188.

Gingras:1984:BRN

[Gin84] Yves Gingras. Book review: National Traditions in Science Norman T. Gridgeman, Biological sciences at the National Research Council of Canada: the early years to 1952. Waterloo, Ontario: Wilfrid Laurier University Press, 1979. Pp. xxi + 153. $7.50. British Journal for the History of Science, 17(1):91, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026026.

Gindikin:1988:TPM

[Gin88]  S.  G.

(Semen Grigorevich) Gindikin. Tales of physicists and mathematicians. Birkhäuser Verlag, Basel, Switzerland, 1988. ISBN 0-8176-3317-0. x + 157 pp. LCCN QC7 .G5613 1988. US$30.00.

Giroud:1986:MCL

[Gir86] Françoise Giroud. Marie Curie, a life. Holmes and Meier, New York, NY, USA, 1986. ISBN 0-8419-0977-6, 0-8419-0978-4 (paperback). 291 + 16 pp. LCCN QD22.C8 G4813 1986.

Gooding:1985:FRE

[GJ85] David Gooding and Frank A.

J. L. James, editors. Faraday rediscovered: essays on the life and work of Michael Faraday, 1791–1867. MacMillan Publishing Company, New York, NY, USA, 1985. ISBN 0-333-39320-1, 0-943818-91-5. 258 pp. LCCN ????

Grafton:1986:HHI

[GJ86] Anthony Grafton and Lisa Jardine. From humanism to the humanities: the institutionalizing of the liberal arts in fifteenth-and sixteenth-century Europe. Harvard U.P, Cambridge, Mass, 1986. ISBN 0-674-32460-9, 0-7156-2100-9. xvi + 224 pp. LCCN LA106 .G73

Gjertsen:1984:CSS

[Gje84] Derek Gjertsen. The classics of science: a study of twelve enduring scientific works. Lilian Barber Press, New York, NY, USA, 1984. ISBN 0-936508-09-4. iii + 374 pp. LCCN Q125 .G54 1984.

Gjertsen:1986:NH

[Gje86] Derek Gjertsen. The Newton handbook. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1986. ISBN 0-7102-0279-2. xiv + 665 pp. LCCN QC16.N7 G54 1986. US$25.00.

Gramiccia:1988:LCL

[GL88] Gabriele Gramiccia and Charles Kaufmann Ledger. The life of Charles Ledger, (1818–1905): alpacas and quinine. MacMillan Publishing Company, New York, NY, USA, 1988. ISBN 0-333-45710-2. xiv + 222 pp. LCCN F2236.L43 G73

1988. URL http://catalog.
hathitrust.org/api/
volumes/oclc/59101195.
html.

Glen:1982:RJC

[Gle82] William Glen. The road to Jaramillo: critical years of the revolution in earth science. Stanford University Press, Stanford, CA, USA, 1982. ISBN 0-8047-1119-4. xvi + 459 pp. LCCN QE511.4 .G53 1982.

Gleason:1991:RSL

[Gle91] Mary Louise Gleason. The Royal Society of London: years of reform, 1827–1847. Harvard dissertations in the history of science. Garland Publishing, Inc., New York, NY, USA, 1991. ISBN 0-8240-7446-7. ix + 532 pp. LCCN Q41 .G64 1991.

Glick:1983:BM

[Gli83a] Thomas F. Glick. Back matter. British Journal for the History of Science, 16(3):332, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026194.

Glick:1983:BRA

[Gli83b]  Thomas F. Glick.  Book

review: Americana Iris N. W. Engstrand, Spanish scientists in the New World: the eighteenth century expeditions. Seattle and London: University of Washington Press, 1981. Pp. xiv + 220. £15.00/ $25.00. British Journal for the History of Science, 16(3): 322, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026191.

Glick:1984:BRN

[Gli84] Thomas F. Glick. Book review: National Traditions in Science Santiago Garma (ed.), El científico español ante su Historia. La Ciencia en España entre 1750–1850 (I Congreso de la Sociedad Española de Historia de las Ciencias). Madrid: Diputación Provincial, 1980. Pp. 583. (No price stated.). British Journal for the History of Science, 17 (1):93–94, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026029.

Glick:1987:CRR

[Gli87] Thomas F. Glick, editor. The Comparative Reception of Relativity, volume 103 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1987. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346. vii + 412 pp. LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

(US). URL https://link.
springer.com/book/10.
1007/978-94-009-3875-5.

Goldberg:1970:PSE

[Gol70] Stanley Goldberg. Poincaré’s silence and Einstein’s relativity: The role of theory and experiment in Poincaré’s physics. British Journal for the History of Science, 5(1): 73–84, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025354.

Goldfarb:1977:RTH

[Gol77] Stephen J. Goldfarb. Rumford’s Theory of Heat: A reassessment. British Journal for the History of Science, 10 (1):25–36, March 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025579.

Goldman:1979:SNB

[Gol79] M. R. Goldman. Short notices of books: The presentation of science by the media. By Greta Jones, Ian Connell, and Jack Meadows. Leicester: Primary Communications Research Centre, 1978. Pp. iv + 76. No price stated. British Journal for the History of Science, 12(2):236–237, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025771.

Goldman:1983:DAS

[Gol83a] Martin Goldman. The demon in the aether: the story of James Clerk Maxwell. P. Harris, Edinburgh, Scotland, 1983. ISBN 0-86228-026-5. 224 + 12 pp. LCCN QC16.M4 G65 1983. US$15.00.

Golinski:1983:BRA

[Gol83b] J. V. Golinski. Book review: Alchemy and Parapsychology Robert Halleux, Les textes alchimiques. (Typologie des sources du moyen âge occidental, fascicule 32). Turnhout, Belgium: Brepols, 1979. Pp.

153. No price stated. British Journal for the History of Science, 16(1):97, March 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026105.

Goldstein:1987:CCF

[Gol87a] Jan Ellen Goldstein. Console and classify: the French psychiatric profession in the Nineteenth Century. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-32279-0. xiii + 414 pp. LCCN RC450.F7 G65 1987.

Golinski:1987:BRW

[Gol87b] J. V. Golinski. Book review: W. F. Bynum & Roy Porter, (eds.) William Hunter and the Eighteenth-Century Medical World. Cambridge: Cambridge University Press, 1985. Pp. xi + 424. ISBN 0-521-26806-0. £35.00. $49.50. British Journal for the History of Science, 20(1):103, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026457.

Golinski:1988:UAE

[Gol88a] J. V. Golinski. Utility and audience in Eighteenth-Century chemistry: Case studies of William Cullen and Joseph Priestley. British Journal for the History of Science, 21(1): 1–31, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026860.

Golinski:1988:BRB

[Gol88b]  Jan Golinski.  Book re-

view: Brian Vickers (ed.). English Science, Bacon to Newton (Cambridge English Prose Texts). Cambridge, London, New York, New Rochelle, Melbourne, Sydney: Cambridge University Press, 1987. Pp. xii + 244. ISBN 0-521-30408-3, £27.50 (hardcover); 0-521-31683-9, £9.95 (paperback). British Journal for the History of Science, 21(3):362–363, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026834.

Goldammer:1991:GMM

[Gol91] Kurt Goldammer. Der göttliche Magier und die Magierin Natur: Religion, Naturmagie und die Anfänge der Naturwissenschaft vom Spätmittelalter bis zur Renaissance; mit Beiträgen zum Magie-Verständnis des Paracelsus. (German) [ The divine Magi and the Magi nature: Religion, nature magic and the beginnings of the natural science from the late Middle Ages to the Renaissance; with contributions for the magic understanding of Paracelsus], volume 5 of Kosmosophie. F. Steiner Verlag, Stuttgart, Germany, 1991.

ISBN 3-515-05584-3. 136 pp. LCCN BF1598.P2 G65 1991.

Goldstein:1992:DST

[Gol92a] Jack S. Goldstein. A different sort of time: the life of Jerrold R. Zacharias, scientist, engineer, educator. MIT Press, Cambridge, MA, USA, 1992. ISBN 0-262-07138-X. xviii + 373 pp. LCCN QC16.Z23 G65 1992.

Golinski:1992:SPC

[Gol92b] Jan Golinski. Science as public culture: chemistry and enlightenment in Britain, 1760–1820. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-39414-7 (hardcover). xi + 342 pp. LCCN QD18.G7 G65

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/91039024.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam029/
                                        91039024.html.

Goldgar:1995:ILC

[Gol95] Anne Goldgar. Impolite learning: conduct and community in the Republic of Letters, 1680– 1750. Yale University Press, New Haven, CT, USA, 1995. ISBN 0-300-05359-2. xiii + 395 pp. LCCN CB203 .G64 1995.

Goodman:1968:BRH

[Goo68] D. C. Goodman. Book review: History of Crystallography Origins of the Science of Crystals. By John G. Burke. Pp. 198, illustr. University of California Press and Cambridge University Press. 1966. 52s. British Journal for the History of Science, 4(1):88–89, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025004.

Goodstein:1974:MAL

[Goo74] R. L. Goodstein. The Mathematical Association Library at the University of Leicester. British Journal for the History of Science, 7(1):100– 103, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025196.

Gooding:1975:BRN

[Goo75] David Gooding. Book review: Nineteenth and Twentieth Centuries: Fields of Force. The Development of a World View from Faraday to Einstein. By William Berkson. London: Routledge & Kegan Paul, 1974. Pp. xiii + 370. £6.50. British Journal for the History of Science, 8(1):89–91, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025834.

Gooding:1980:FTC

[Goo80]  David Gooding.  Faraday,

Thomson, and the concept of the magnetic field. British Journal for the History of Science, 13(2):91–120, July 1980.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025890.

Gooding:1981:BRG

[Goo81a] David Gooding. Book review: General Natural Order: Historical Studies of Scientific Culture. Edited by Barry Barnes and Steven Shapin. London and Beverly Hills: Sage, 1979. pp. 225. £10.00/£4.25 (paperback). British Journal for the History of Science, 14(1): 84–86, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026075.

Gooding:1981:BRN

[Goo81b] David Gooding. Book review: Nineteenth and Twentieth Centuries: James Prescott Joule and the Concept of Energy. By Henry John Steffens. New York: Science History Publications; Folkestone: Dawson, 1979. Pp. x + 172. $20.00/ £10.00. British Journal for the History of Science, 14(2): 217–219, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025958.

Gooding:1982:BRN

[Goo82] David Gooding. Book review: The New Physics Loyd S. Swenson Jr, Genesis of relativity: Einstein in context.

New York: Burt Franklin & Co., 1979. pp. xvi + 266. $21.00. British Journal for the History of Science, 15(2): 199–200, July 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025972.

Gooding:1983:BRS

[Goo83a] David Gooding. Book review: Science and Society Rom Harré. Great scientific experiments: 20 experiments that changed our view of the world. Oxford: Phaidon Press, 1981. Pp. 224. £8.95. British Journal for the History of Science, 16(3):313–314, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026180.

Goodman:1983:PIP

[Goo83b] David Goodman. Philip II’s patronage of science and engineering. British Journal for the History of Science, 16(1): 49–66, March 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026093.

Gooding:1985:BRC

[Goo85a] David Gooding. Book review:

Carl Gustaf Bernhard, Elisabeth Crawford and Per Sörbom (eds), Science, Technology and Society in the Time of Alfred Nobel. Oxford and New York:

Pergamon Press (for the Nobel Foundation), 1982. Pp. xvii + 426. £29.75, $59.50. British Journal for the History of Science, 18(2):239–240, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026338.

Gooding:1985:BRN

[Goo85b] David Gooding. Book review: Nathan Reingold (ed), The Papers of Joseph Henry, volume 3: January 1836-December 1837, The Princeton Years. Washington: Smithsonian Institution Press, 1979. Pp. xxxiv + 585. $30. Nathan Reingold (ed), The Papers of Joseph Henry, volume 4: January 1838-December 1840, The Princeton Years. Washington: Smithsonian Institution Press, 1981. Pp. xxxiv + 475. $30. British Journal for the History of Science, 18(1): 113–114, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026294.

Gooding:1988:BRN

[Goo88a] David Gooding. Book review: Nancy J. Nersessian, (ed.). The Process of Science. Contemporary Philosophical Approaches to Understanding Scientific Practice. Dordrecht, Boston, Lancaster: Martinus Nijhoff Publishers, 1987. Pp. xiv + 221. ISBN 90-247-3425-8. DFl.135.00, £47.50, $59.50. British Journal for the History of Science, 21(2):254– 255, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026984.

Goodman:1988:PPG

[Goo88b] David C. Goodman. Power and penury: government, technology and science in Philip II’s Spain. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-30532-2, 0-521-52477-6. xii + 275 pp. LCCN Q127.S7; M89.E02119.

Gooding:1989:HGE

[Goo89a] David Gooding. How to be a good empiricist. British Journal for the History of Science, 22(4):419–427, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026918.

Goodman:1989:BRJ

[Goo89b] David Goodman. Book review:

José Mariá López Piñero. El arte de navegar en la España del Renacimiento. Barcelona: Editorial Labor, 2nd edn, 1986. Pp. 285. ISBN 84-335-0029-5. No price given. British Journal for the History of Science, 22 (1):99–100, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026699.

Gooding:1990:EMM

[Goo90] David Gooding. Experiment and the making of meaning: human agency in scientific observation and experiment, volume 5 of Science and philosophy. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0719-4. xviii + 310 pp. LCCN Q182.3 .G69 1990.

Gosling:1987:BFN

[Gos87] Francis G. (Francis George) Gosling. Before Freud: neurasthenia and the American medical community, 1870–1910. University of Illinois Press, Urbana, IL, USA, 1987. ISBN 0-252-01406-5. xviii + 192 pp. LCCN RC552.N5 G67 1987.

Gottschalk:1966:BRT

[Got66] H. B. Gottschalk. Book review: Theophrastus, De Lapidibus. Edited with Introduction, Translation and Commentary by D. E. Eichholz. Pp. viii + 141. Oxford: Clarendon Press, 1965. £2 5s. British Journal for the History of Science, 3(2):188–189, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025031.

Gottschalk:1967:BRH

[Got67] H. B. Gottschalk. Book review: History and Philosophy of Science Polarity and Analogy: Two Types of Argumentation in Early Greek Thought. By G. E. R. Lloyd. Pp. vi + 503. London: Cambridge University Press. 1966. 84s. British Journal for the History of Science, 3(4):398– 399, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024967.

Gottschalk:1971:BRC

[Got71] H. B. Gottschalk. Book review:

Classical Antiquity Aristotle’s Physics. Translated with commentaries and glossary by H. G. Apostle. Bloomington, Indiana and London: Indiana University Press. 1969. Pp. xi + 386. £1.80 (paperback). British Journal for the History of Science, 5(3):308, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025337.

Gottlieb:1986:CXE

[Got86a]  Jean S. Gottlieb.  Collec-

tions XIII. Early science at the Newberry Library: an introduction. British Journal for the History of Science, 19 (3):323–330, 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026508.

Gottlieb:1986:ESN

[Got86b] Jean S. Gottlieb. Early science at the Newberry Library: An introduction. British Journal for the History of Science, 19(3):323–330, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gottlieb:1992:CNL

[Got92] Jean S. Gottlieb. A checklist of the Newberry Library’s printed books in science, medicine, technology, and the pseudosciences ca. 1460–1750, volume 1195 of Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1992. ISBN 0-8240-5171-8. xxiii + 312 pp. LCCN Z7405.H6 G67 1992.

Gough:1968:LEC

[Gou68] J. B. Gough. Lavoisier’s early career in science: An examination of some new evidence. British Journal for the History of Science, 4(1): 52–57, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024986.

Gouk:1982:BRS

[Gou82] Penelope M. Gouk. Book review: Science in Different Countries Michael Hunter, Science and society in restoration England. Cambridge: Cambridge University Press, 1981. Pp. xii + 233. £18.50; £5.95 (paperback). British Journal for the History of Science, 15 (2):193–194, July 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025969.

Gouk:1984:BRS

[Gou84a] Penelope M. Gouk. Book review: Seventeenth Century: Michael Hunter, The Royal Society and its Fellows 1660– 1700: the morphology of an early scientific institution. Chalfont St Giles, Bucks.: The British Society for the History of Science, 1982. Pp. v + 270. ISBN 0-906450-03-9. £5.90, $11.00. British Journal for the History of Science, 17(1):110, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026047.

Gouka:1984:BRS

[Gou84b] Penelope M. Gouka. Book review: Seventeenth Century: David Castillejo, The expanding force in Newton’s cosmos, as shown in his unpublished papers. Madrid: Ediciones de Arte y Bibliofilia, 1981. Pp.

125. British Journal for the History of Science, 17(1):112– 113, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026049.

Gouk:1985:BRH

[Gou85] Penelope M. Gouk. Book review: H. F. Cohen, Quantifying Music: The Science of Music at the First Stage of the Scientific Revolution, 1580–1650 (the University of Western Ontario Series in the Philosophy of Science vol. 23) Dordrecht, Boston & Lancaster: D. Reidel Publishing Co., 1984. Pp. xvii + 308. ISBN 90-277-1637-4. Dfl 145, $54.50. British Journal for the History of Science, 18(3):369–371, November 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026406.

Gouk:1995:WAC

[Gou95] Penelope Gouk, editor. Wellsprings of achievement: cultural and economic dynamics in early modern England and Japan. Variorum, Aldershot, Hampshire, UK, 1995. ISBN 0-86078-465-7. viii + 271 pp. LCCN HD8399.E52 W45 1995.

Gowing:1969:BRN

[Gow69] Margaret Gowing. Book review: Nineteenth and Twentieth Centuries: The Inner Ring: The Early History of the National Research Council of Canada. By Mel Thistle. University of Toronto Press and Oxford University Press. 1968. Pp. xxxiii + 435. 66s. 6d. British Journal for the History of Science, 4(3):303, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025496.

Gower:1971:BRG

[Gow71] B. Gower. Book review: Germany Herder and the Philosophy and History of Science. By H. B. Nisbet. Cambridge: Modern Humanities Research Association. 1970. Pp. xii + 358. No price stated. British Journal for the History of Science, 5(4):414, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025400.

Gower:1973:BRN

[Gow73] B. S. Gower. Book review:

Nineteenth Century: Faraday as a Natural Philosopher. By Joseph Agassi. Chicago and London: University of Chicago Press, 1972. Pp. xvi + 359. £5.65. British Journal for the History of Science, 6(4):452– 453, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025523.

Gowing:1974:SKS

[Gow74] Margaret Gowing. Scientific knowledge and its social problems. British Journal for the History of Science, 7(1):72– 75, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025177.

Gower:1978:BRN

[Gow78]  Barry Gower.  Book re-

view: Nineteenth and Twentieth Centuries: Method and appraisal in the physical sciences. The critical background to modern science, 1800–1905. Edited by Colin Howson. Cambridge: Cambridge University Press, 1976. Pp. vii + 344. £10.50. British Journal for the History of Science, 11(2): 185–186, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025744.

Gowing:1983:RCN

[Gow83] Ronald Gowing. Roger Cotes, natural philosopher. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-23741-6. x + 210 pp. LCCN QA804 .G68 1983. US$30.00

(est.). URL http://www.loc.
gov/catdir/description/
cam031/82001154.html;
http://www.loc.gov/
catdir/samples/cam031/
82001154.html;   http://
www.loc.gov/catdir/toc/
cam031/82001154.html.

Galilei:1992:SNM

[GP92] Galileo Galilei and Isabelle Pantin. Sidereus nuncius, le messager céleste. (French) [Sidereus nuncius = The celestial messenger]. Les Belles lettres, Paris, France, 1992. ISBN 2-251-34505-1. civ + 115 pp. LCCN 95.E11231; QB41 .G1714 1992.

Gowing:1988:SPP

[GR88] Margaret Gowing and Nicolaas A. Rupke, editors. Science, politics and the public good: essays in honour of Margaret Gowing. MacMillan Publishing Company, New York, NY, USA, 1988. ISBN 0-333-44159-1. xv + 260 pp. LCCN Q127.G4 S433 1988.

Goodman:1991:RSE

[GR91] David C. (David Charles) Goodman and Colin A. (Colin Archibald) Russell, editors. The rise of scientific Europe, 1500–1800. Hodder and Stoughton, London, UK, 1991. ISBN 0-340-55861-X. ix + 437 pp. LCCN Q127.E8 R574 1991.

Grafton:1981:BRR

[Gra81] A. T. Grafton. Book review: Renaissance and Seventeenth Century: Robert Fludd and his Philosophicall Key: being a Transcription of the manuscript at Trinity College, Cambridge. With an Introduction by Allen G. Debus. New York: Science History Publications, 1979. Pp. xii +156. $40.00/£20.00. British Journal for the History of Science, 14(3):290–291, November 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025988.

Gray:1983:BRMb

[Gra83a]  Jeremy Gray.  Book re-

view: Mathematical Sciences: J. V. Grabiner, The origins of Cauchy’s rigorous calculus. Cambridge, Mass.: M.I.T. press, 1981. Pp. x + 252. £17.50. British Journal for the History of Science, 16(3): 290–291, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026161.

Gray:1983:BRMa

[Gra83b] Jeremy Gray. Book review: Mathematical Sciences: W. K. Bühler, Gauss. A biographical study. Berlin-Heidelberg-New York: Springer-Verlag, 1981. Pp. 208 DM39.00; approx. US $17.80. ISBN 3-540-10662-6. British Journal for the History of Science, 16(3): 289–290, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026160.

Grand:1984:TAE

[Gra84a] H. E. Le Grand. Theory and application: the early chemical work of J. A. C. Chaptal. British Journal for the History of Science, 17(1):31– 46, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026015.

Grant:1984:DEC

[Gra84b] Edward Grant. In defense of the Earth’s centrality and immobility: scholastic reaction to Copernicanism in the Seventeenth Century, volume 74, part 4 (1984) of Transactions of the American Philosophical Society, 0065-9746. American Philosophical Society, Philadelphia, PA, USA, 1984. ISBN 0-87169-744-0 (paperback). 69 pp. LCCN QB981 .G66 1984.

Gray:1984:BRH

[Gra84c] Jeremy Gray. Book review:

History of Mathematical Sciences: John T. Cannon and Sigalia Dostrovsky, The evolution of dynamics: vibration theory from 1687 to 1742. New York: Springer, 1981. Pp vi + 184. ISBN 0-387-90626-6. DM 98. British Journal for the History of Science, 17(2): 234–235, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026567.

Gray:1985:BRH

[Gra85a] J. J. Gray. Book review: Hans Wussing, The genesis of the abstract group concept. A contribution to the history of the origin of abstract group theory. Cambridge, Mass., and London: MIT Press, 1984. Pp. 331. ISBN 0-262-23109-3. £28.50. British Journal for the History of Science, 18(3): 360, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026398.

Gray:1985:BRL

[Gra85b]  Jeremy  Gray.

Book review: Leonard Euler 1707–1783. Beiträge zu Leben und Werk. Gedenkband des Kantons Basel-Stadt. Basel: Birkhäuser Verlag, 1983. Pp.

555. SFr. 58. British Journal for the History of Science, 18 (1):104–105, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026287.

Gray:1985:LDE

[Gra85c]  Jeremy John Gray.  Lin-

ear differential equations and group theory from Riemann to Poincaré. Birkhäuser Verlag, Basel, Switzerland, 1985. ISBN 0-8176-3318-9, 3-7643-3318-9. xxv + 460 pp. LCCN QA372 .G68 1985.

Gray:1986:BRS

[Gra86a] Jeremy Gray. Book review: Steven B. Engelsman. Families of Curves and the Origins of Partial Differentiation. Amsterdam: North-Holland, 1984. Pp. x + 238. ISBN 0-444-86897-6. US $29.00, Dfl. 85.00. British Journal for the History of Science, 19(3): 365, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026536.

Gray:1986:BRW

[Gra86b] Jeremy Gray. Book review:

Winfried Scharlau & Hans Opolka. From Fermat to Minkowski. Lectures on the Theory of Numbers and Its Historical Development. Translated from the German by W. K. Bühler and G. Cornell. Berlin and New York: Springer, 1985. Pp xi + 184. ISBN 0-387-90942-7. DM 72.00. British Journal for the History of Science, 19 (3):362, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026533.

Graham:1987:SPH

[Gra87a] Loren R. Graham. Science, philosophy, and human behavior in the Soviet Union. Columbia University Press, New York, NY, USA, 1987. ISBN 0-231-06442-X. xiii + 565 pp. LCCN Q127.S696 G73 1987.

Gray:1987:BRB

[Gra87b] J. J. Gray. Book review: B. L.

van der Waerden. A History of Algebra: From al-Khwarizmi to Emmy Noether. Berlin, Heidelberg, New York, Tokyo: Springer-Verlag, 1985. Pp. xi + 271. ISBN 0-387-13610-X (U.S.) DM 98.00 (hard cover). (English price not given). British Journal for the History of Science, 20(1): 96–97, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026450.

Grant:1994:PSO

[Gra94] Edward Grant. Planets, stars, and orbs: the medieval cosmos, 1200–1687. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-43344-4 (hardcover). xxiii + 816 pp. LCCN QB981 .G664

1994. URL http://www.loc.
gov/catdir/description/
cam025/93025899.html;
http://www.loc.gov/
catdir/toc/cam025/
93025899.html.

Graham:1998:WWL

[Gra98] Loren R. Graham. What have we learned about science and technology from the Russian experience? Stanford University Press, Stanford, CA, USA, 1998. ISBN 0-8047-2985-9 (hardcover), 0-8047-3276-0 (paperback). xiii + 177 pp. LCCN Q175.5 .G73 1998. URL http:

//www.loc.gov/catdir/
description/cam028/
97041635.html;   http://
www.loc.gov/catdir/
enhancements/fy0710/
97041635-b.html;   http:/
/www.loc.gov/catdir/toc/
cam026/97041635.html.

Greenaway:1963:BRH

[Gre63] Frank Greenaway. Book review: Historical Studies in the Language of Chemistry. By M. D. Crosland. Heinemann, Educational Books Limited, London, 1963. Pp. xvii + 406, 5 pl., 7 fig. 50s. British Journal for the History of Science, 1(3):287–288, June 1963. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024929.

Greenaway:1971:BRT

[Gre71]  F. Greenaway.  Book re-

view: Technology A History of the Zinc Smelting Industry in Great Britain. By E. J. Cocks and B. Walters. London: Harrap. 1968. Pp. vi + 224, 42 illustr. £2.75. British Journal for the History of Science, 5 (3):294–296, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025323.

Greene:1982:GNC

[Gre82] Mott T. Greene. Geology in the Nineteenth Century: changing views of a changing world. Cornell history of science series. Cornell University Press, Ithaca, NY, USA, 1982. ISBN 0-8014-1467-9. 324 pp. LCCN QE11 .G73 1982.

Grelon:1986:ICT

[Gre86] André Grelon, editor. Les Ingénieurs de la crise: titre et profession entre les deux guerres. (French) [The Engineers of the crisis: title and profession between the two wars], volume 21 of Recherches d’histoire et de sciences sociales = Studies in history and the social sciences, 0249-5619. Editions de l’Ecole des hautes études en sciences sociales, Paris, France, 1986. ISBN 2-7132-0875-0. 461 pp. LCCN TA71 .I55 1986. 250F.

Grey:1988:BRG

[Gre88] Jeremy Grey. Book review: S. G. Gindikin. Tales of Physicists and Mathematicians. Boston and Basel: Birkhäuser, 1988. Pp. xi + 157. ISBN 0-8176-3317-0 and 3-7643-3317-0. SFr 48.00. British Journal for the History of Science, 21(4): 495, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026966.

Greene:1992:NKP

[Gre92a]  Mott T. Greene.  Natural

knowledge in preclassical antiquity. The Johns Hopkins University Press, Baltimore, MD, USA, 1992. ISBN 0-8018-4292-1. xix + 182 pp. LCCN BL435 .G74

1992. URL http://www.loc.
gov/catdir/bios/jhu051/
91026952.html;   http://
www.loc.gov/catdir/
description/jhu051/
91026952.html;   http://
www.loc.gov/catdir/toc/
jhu051/91026952.html.

Gregory:1992:NLN

[Gre92b]  Frederick Gregory.  Nature

lost?: natural science and the German theological traditions of the Nineteenth Century. Harvard University Press, Cambridge, MA, USA, 1992. ISBN 0-674-60483-0. viii + 341 pp. LCCN BL245 .G74 1992.

Grmek:1969:BRB

[Grm69]  M. D. Grmek.  Book re-

view: Biology and Medicine The Human Brain and Spinal Cord. A Historical Study Illustrated by Writings from Antiquity to the Twentieth Century. By Edwin Clark and C. D. O’Malley. Berkeley and Los Angeles: University of California Press. 1968. Pp. xiii + 926. $25.00. British Journal for the History of Science, 4(4): 413–414, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025245.

Grmek:1990:PRB

[Grm90] Mirko D. (Mirko Dražen)

Grmek. La première révolution biologique: réflexions sur la physiologie et la médicine du XVIIe siècle. (French) [The first biological revolution: reflections on the physiology of medicine in the 17th Century]. Bibliothèque scientifique Payot. Editions Payot, Paris, France, 1990. ISBN 2-228-88277-1. 358 + 8 pp. LCCN QP21 .G76 1990. 220F.

Gross:1989:SNQ

[Gro89] Nikolaus Gross. Senecas Naturales quaestiones: Komposition, naturphilosophische Aussagen und ihre Quellen. (German) [Senecas Naturales quaestiones: Composition, naturalphilosophical statements and their sources], volume 27 of Palingenesia. F. Steiner Verlag, Stuttgart, Germany, 1989. ISBN 3-515-05385-9. 335 pp. LCCN PA6661.N3 G76 1989.

Grove:1990:HHH

[Gro90] David I. Grove. A history of human helminthology. C.A.B. International, Wallingford, UK, 1990. ISBN 0-85198-689-7. viii + 848 + 12 pp. LCCN RC119.7 .G76 1990.

Grosholz:1991:CMP

[Gro91] Emily Grosholz. Cartesian method and the problem of reduction. Clarendon Press, New York, NY, USA, 1991. ISBN 0-19-824250-6. viii + 161 pp. LCCN B1875 .G76

                                        1991. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0605/90035814-d.html.

Grove:1995:GIC

[Gro95] Richard (Richard H.) Grove. Green imperialism: colonial expansion, tropical island Edens, and the origins of environmentalism, 1600–1860. Studies in environment and history. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-40385-5 (hardcover). xiv + 540 pp. LCCN GE195 .G76

1995. URL http://www.loc.
gov/catdir/description/
cam025/93028654.html;
http://www.loc.gov/
catdir/toc/cam026/
93028654.html.

Gruber:1969:WWB

[Gru69] Jacob W. Gruber. Who was the Beagle’s naturalist? British Journal for the History of Science, 4(3):266–282, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025475.

Gage:1988:BHL

[GS88] Andrew Thomas Gage and William T. (William Thomas) Stearn. A bicentenary history of the Linnean Society of London. Academic Press, New York, USA, 1988. ISBN 0-12-273150-6. ix + 242 pp. LCCN QH1 .G24 1988.

Godlewska:1994:GE

[GS94] Anne Godlewska and Neil Smith, editors. Geography and empire, volume 30 of The Institute of British Geographers special publications series. Blackwell, Oxford, UK, 1994. ISBN 0-631-19384-7, 0-631-19385-5 (paperback). LCCN G70 .G443 1994.

Gruber:1992:ROC

[GT92] Jacob W. Gruber and John C.

Thackray. Richard Owen commemoration: three studies, volume 1 of Historical studies in the life and earth sciences. Natural History Museum Publications, London, UK, 1992. ISBN 0-565-01109-X. ix + 181 pp. LCCN QL31.O9 G78 1992.

Guagnini:1989:BRT

[Gua89] Anna Guagnini. Book review: Therésè Charmasson, Anne-Marie Lelorrain, & Yannick Ripa. L’enseignement technique de la Révolution à nous jours Tome 1: 1789–1926. Paris: Institut National de la Recherche Pédagogique and Economica, 1987. Pp. 783. ISBN 2-7342-0146-1/2-7178-1323-3. 350FF. British Journal for the History of Science, 22(4):472– 474, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026944.

Guerlac:1968:ER

[Gue68] Henry Guerlac. Essay review. British Journal for the History of Science, 4(2):166– 172, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025146.

Guerlac:1981:NC

[Gue81] Henry Guerlac. Newton on the Continent. Cornell University Press, Ithaca, NY, USA, 1981. ISBN 0-8014-1409-1. 169 pp. LCCN QC16.N7 G82.

Guerlac:1984:NMW

[Gue84] Henry Guerlac. “Newton’s mathematical way”: another look. British Journal for the History of Science, 17(1):61– 64, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026017.

Gunther:1984:ILR

[Gun84] Albert Edward (Albert Everard) Gunther. An introduction to the life of the Rev. Thomas Birch D.D., F.R.S. 1705–1766: leading editor of the General Dictionary’ — 1741: Secretary of the Royal Society and Trustee of the British Museum: based on a daily ’Diary’, 1735–1764 and other sources. Halesworth Press, Suffolk, England, 1984. ISBN 0-9507276-1-X. x + 118 pp. LCCN Q143.B49 G85 1984.

Guralnick:1978:BRA

[Gur78]  Stanley  M.  Gural-

nick. Book review: American Science The Pursuit of Knowledge in the Early American Republic: American Scientific and Learned Societies from Colonial Times to the Civil War. Ed. by Alexandra Oleson and Sanborn C. Brown. Baltimore & London: Johns Hopkins University Press, 1976. Pp. xxv + 372. £11.55. British Journal for the History of Science, 11 (1):69–71, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025613.

Gutting:1989:MFA

[Gut89] Gary Gutting. Michel Foucault’s archaeology of scientific reason. Modern European philosophy. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-36619-4, 0-521-36698-4 (paperback). xiii + 306 pp. LCCN B2430.F724 G87

                                        1989. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/88031881.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam025/
                                        88031881.html.

Galilei:1989:SNS

[GV89] Galileo Galilei and Albert Van Helden. Sidereus nuncius, or, The Sidereal messenger. University of Chicago Press, Chicago, IL, USA, 1989. ISBN 0-226-27902-2 (hardcover), 0-226-27903-0 (paperback). xii + 127 pp. LCCN QB41 .G173

                                        1989. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi052/88025179.html;
                                        http://www.loc.gov/
                                        catdir/toc/uchi051/
                                        88025179.html.

Gingerich:1974:NFL

[GW74] Owen Gingerich and Barbara Welther. Note on Flamsteed’s lunar tables. British Journal for the History of Science, 7(3): 257–258, November 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hannah:1982:EMP

[H^{+}82] Leslie Hannah et al. Engineers, managers, and politicians: the first fifteen years of nationalised electricity supply in Britain. MacMillan Publishing Company, New York, NY, USA, 1982. ISBN 0-333-22087-0. xiii + 336 pp. LCCN HD9685.G7 H36

1982b. URL http://catalog.
hathitrust.org/api/
volumes/oclc/8715072.html.

Hartcup:1984:CA

[HA84] Guy Hartcup and T. E. Allibone. Cockcroft and the atom. Adam Hilger Ltd., Bristol, UK, 1984. ISBN 0-85274-759-4. xii + 320 pp. LCCN QC774.C6 H37 1984. US$18.95.

Hammond:1987:CLA

[HA87] John H. Hammond and Jill Austin. The camera lucida in art and science. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-527-3. xii + 200 pp. LCCN NC730 .H26 1987.

Holbrook:1992:SPD

[HAB92] Mary Holbrook, R. G. W.

(Robert Geoffrey William) Anderson, and D. J. Bryden. Science preserved: a directory of scientific instruments in collections in the United Kingdom and Eire. Her Majesty’s Stationary Office, London, England, 1992. ISBN 0-11-290060-7. 271 pp. LCCN Q185.7 .H65 1992. US$35.

Hackmann:1974:BRE

[Hac74] W. D. Hackmann. Book review: Eighteenth and Nineteenth Centuries: The Papers of Joseph Henry. Volume I. The Albany Years: December 1797-October 1832. Ed. by Nathan Reingold. Washington, D.C.: Smithsonian Institution Press, 1972. Pp. xxx + 496. $15.00. British Journal for the History of Science, 7(2): 195–196, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025220.

Hackmann:1975:BRS

[Hac75] W. D. Hackmann. Book review: Seventeenth and Eighteenth Centuries: Physics at 17th and 18th-Century Leiden. By Edward G. Ruestow. The Hague: Martinus Nijhoff, 1973. Pp. 174. Hfl. 24.50. British Journal for the History of Science, 8(2):183–185, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025653.

Hackmann:1976:BRT

[Hac76] W. D. Hackmann. Book review: Technology The Scientific Breakthrough. The Impact of Modern Invention. By Ronald W. Clark. London: Nelson, 1974. Pp. 208. £4.50.

Wireless Telegraphy. Royal Institution Library of Science. Ed. by Sir Eric Eastwood. London: Applied Science Publishers, 1974. Pp. xi + 391. £10.00. British Journal for the History of Science, 9(1): 68–69, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025707.

Hackmann:1977:BRE

[Hac77] W. D. Hackmann. Book review: Eighteenth and Nineteenth Centuries: The Papers of Joseph Henry. Volume 2. November 1832-December 1835. The Princeton Years. Ed. by Nathan Reingold. Washington, D.C.: Smithsonian Institution Press, 1975. Pp. xl + 524. $30.00. British Journal for the History of Science, 10(1):85, March 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025597.

Hackmann:1982:BRIb

[Hac82a]  Willem  D.

Hackmann. Book review: Instruments G. L’E. Turner, Essays on the history of the microscope. Oxford: Senecio Publishing, 1980. Pp. viii + 245. £14.95. British Journal for the History of Science, 15(3): 312–313, November 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026229.

Hackmann:1982:BRIa

[Hac82b] Willem D. Hackmann. Book review: Instruments P. H. Sydenham, Measuring instruments: tools of knowledge and control. London: Peter Peregrinus Ltd in association with the Science Museum, 1979. History of Technology Series No.

1. Pp. xviii + 512. £19 (U.K.)/ £22 (overseas excluding Americas). British Journal for the History of Science, 15(3):310– 312, November 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026228.

Hackmann:1984:SSS

[Hac84] Willem Dirk Hackmann. Seek and strike: sonar, anti-submarine warfare, and the Royal Navy, 1914–54. Her Majesty’s Stationary Office, London, England, 1984. ISBN 0-11-290423-8. xxxv + 487 + 32 pp. LCCN V214 .H33 1984. US$15.95.

Hackmann:1988:BRR

[Hac88a] W. D. Hackmann. Book review: R. D. Connor. The Weights and Measures of England. London: Her Majesty’s Stationery Office, 1987. Pp. xvi + 422. ISBN 0-11-290435-1. £30.00. British Journal for the History of Science, 21(4): 499, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026971.

Hackmann:1988:BRV

[Hac88b] W. D. Hackmann. Book review:

V. J. Phillips. Waveforms. A History of Early Oscillography. Bristol: Adam Hilger, 1987. Pp. viii + 259. ISBN 0-85274-274-6. £35.00. British Journal for the History of Science, 21 (3):380–381, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026851.

Hackmann:1989:BRJ

[Hac89] W. D. Hackmann. Book review: J. R. Beniger. The Control Revolution. Technological and Economic Origins of the Information Society. Cambridge, Mass., and London: Harvard University Press, 1986. Pp. x + 493. ISBN 0-674-16985-9. £21.25. British Journal for the History of Science, 22(1): 86–87, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026686.

Hacking:1990:TC

[Hac90] Ian Hacking. The taming of chance. Ideas in context. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-38014-6, 0-521-38884-8 (paperback). xiii + 264 pp. LCCN BD595 .H33 1990. URL http:

//www.loc.gov/catdir/
description/cam024/
89036411.html;   http://
www.loc.gov/catdir/toc/
cam022/89036411.html.

Haddow:1963:BRS

[Had63]  Alex Haddow.  Book re-

view: A Short History of Medicine. By Charles Singer and E. Ashworth Underwood. 2nd Edition. Pp. xvi + 854. £3 3s. Clarendon Press; Oxford University Press, 1963. British Journal for the History of Science, 1(4):385–386, December 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024912.

Hadden:1994:SME

[Had94] Richard W. Hadden. On the shoulders of merchants: exchange and the mathematical conception of nature in early modern Europe. SUNY series in science, technology, and society. State University of New York Press, Albany, NY, USA, 1994. ISBN 0-7914-2011-6, 0-7914-2012-4 (paperback). xviii + 191 pp. LCCN Q175.52.E85 H33 1994.

Hagen:1992:EBO

[Hag92] Joel Bartholemew Hagen. An entangled bank: the origins of ecosystem ecology. Rutgers University Press, New Brunswick, NJ, USA, 1992. ISBN 0-8135-1823-7 (hardcover), 0-8135-1824-5 (paperback). x + 245 pp. LCCN QH540.8 .H34 1992. US$38.00, US$16.00.

Hahn:1970:BRM

[Hah70] Roger Hahn. Book review:

Mathematics: The Doctrine of Chances or, A Method of Calculating the Probabilities of Events in Play. By Abraham de Moivre. 2nd ed. [1738]. London, F. Cass, 1967. Pp. xiv + 258. £6 6s. British Journal for the History of Science, 5(2):189–190, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025421.

Hahn:1973:BRE

[Hah73]  Roger  Hahn.

Book review: Eighteenth and Nineteenth Centuries: Jean d’Alembert. Science and the Enlightenment. By Thomas L. Hankins. Clarendon Press: Oxford University Press, 1970. Pp. xii + 260. £3.50. British Journal for the History of Science, 6(3):327–329, June 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025461.

Haines:1977:BRS

[Hai77] Barbara Haines. Book review: Seventeenth and Eighteenth Centuries: Condorcet. From Natural Philosophy to Social Mathematics. By Keith Michael Baker. Chicago: University of Chicago Press, 1975. Pp. xiv + 523. £14.85. British Journal for the History of Science, 10(3):267–269, November 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025855.

Haines:1978:IRB

[Hai78] Barbara Haines. The interrelations between social, biological, and medical thought, 1750–1850: Saint-Simon and Comte. British Journal for the History of Science, 11(1): 19–35, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025605.

Haines:1980:BNO

[Hai80] Barbara Haines. Book review: Biology La notion d’organisation dans l’histoire de la biologie. By Joseph Schiller. Paris: Maloine, 1978. Pp. vi + 133. 70F. British Journal for the History of Science, 13(2):160–161, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025895.

Haines:1982:BRS

[Hai82] Barbara Haines. Book review: Science in Different Countries C. C. Gillispie, Science and polity in France at the end of the old regime. Princeton, New Jersey: Princeton University Press, 1981. Pp. xii + 601. £22.30. British Journal for the History of Science, 15 (2):191–193, July 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025968.

Hakfoort:1995:OAE

[Hak95] Casper Hakfoort. Optics in the age of Euler: conceptions of the nature of light, 1700– 1795. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-40471-1. vi + 243 pp. LCCN QC403 .H3513

1995. URL http://www.loc.
gov/catdir/description/
cam026/94031656.html;
http://www.loc.gov/
catdir/toc/cam023/
94031656.html.

Hall:1965:BRL

[Hal65a] A. Rupert Hall. Book review: L’Oeuvre Scientifique de Pascal. Centre International de Synthèse, Presses Universitaires de France. 1964. Pp. 311. 20 francs. British Journal for the History of Science, 2(4): 359–360, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024894.

Hall:1965:GSM

[Hal65b] A. Rupert Hall. Galileo and the science of motion. British Journal for the History of Science, 2(3):185–199, June 1965.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024934.

Hall:1965:OAS

[Hal65c] Marie Boas Hall. Oldenburg and the art of scientific communication. British Journal for the History of Science, 2(4): 277–290, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024886.

Hall:1966:BRS

[Hal66a] A. Rupert Hall. Book review: Sixteenth and Seventeenth Centuries: Kepler’s Dream… By John Lear. Translated by Patricia Frueh Kirkwood. Pp. 182. University of California Press; London: Cambridge University Press, 1965. 40s. net. British Journal for the History of Science, 3(1):81–82, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025110.

Hall:1966:MRS

[Hal66b] A. Rupert Hall. Mechanics and the Royal Society, 1668– 1670. British Journal for the History of Science, 3(1): 24–38, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025100.

Hall:1968:SSL

[Hal68a]  A. R. Hall.  The Soci-

ety’s Symposium at Loughborough. British Journal for the History of Science, 4(2): 199, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025170.

Hall:1968:BRSc

[Hal68b] A. Rupert Hall. Book review:

Scientific Instruments and Museums: Historical Aspects of Microscopy. Edited by S. Bradbury and G. L’E. Turner. Pp.

227. Cambridge: Heffer. 1967. 42s. British Journal for the History of Science, 4(2):178– 179, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025151.

Hall:1968:BRSa

[Hal68c] A. Rupert Hall. Book review: Seventeenth Century: Galilée, Dialogues: choix, traduction, préface de Paul-Henri Michel; Lettres choisies, introduction de Giorgio de Santillana, traduction de Paul-Henri Michel. Pp. 430. Paris: Hermann, Paperback. 24 F. Études Galiléennes. By Alexandre Koyré. Pp. 342. Paris: Hermann. 1966. Paperback. 18 F. British Journal for the History of Science, 4(1):69– 71, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024990.

Hall:1968:BRSb

[Hal68d] Marie Boas Hall. Book review: Seventeenth Century: The Royal Society: Concept and Creation. By Margery Purver. With an introduction by H. R. Trevor-Roper. Pp. xviii + 246. 12 plates. London: Routledge and Kegan Paul. 1967. 35s. British Journal for the History of Science, 4(1):76–77, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024994.

Hall:1969:BRS

[Hal69a] A. Rupert Hall. Book review: Seventeenth Century: Kepler’s Somnium: The Dream, or Posthumous Work on Lunar Astronomy. Translated with a commentary by Edward Rosen. Madison, Milwaukee and London: University of Wisconsin Press. Pp. xxiii + 255. 1967. 66s. British Journal for the History of Science, 4 (3):298–299, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025490.

Hall:1969:PAC

[Hal69b] A. Rupert Hall. Presidential address: Can the history of science be history? British Journal for the History of Science, 4(3):207–220, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025471.

Hall:1971:BRR

[Hal71] Marie Boas Hall. Book review:

Russia Mikhail Vasil’evich Lomonosov on the Corpuscular Theory. Translated, with an Introduction, by Henry M. Leicester. Harvard University Press and Oxford University Press. 1970. Pp. viii + 289. Portrait. £4.75. British Journal for the History of Science, 5(3):307, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025336.

Hall:1972:ERN

[Hal72a] A. Rupert Hall. Essay review: The natural philosophy of Galileo. British Journal for the History of Science, 6 (1):80–84, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025265.

Hall:1972:BRC

[Hal72b] Marie Boas Hall. Book review: Chemistry Les Doctrines Chimiques en France du début du XVIIe à la fin du XVI-IIe Siècle. By Hétène Metzger. Paris: Albert Blanchard. 1969. Pp. 496.25 francs. British Journal for the History of Science, 6(1):89–90, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025270.

Hall:1977:BRSa

[Hal77a] A. Rupert Hall. Book review: Seventeenth Century: Les Nouvelles Pensées de Galilée. Par Marin Mersenne. Edition critique par Pierre Costabel et Michel-Pierre Lerner. Introduction par Bernard Rochot. 2 vols. Paris: J. Vrin. 1973. Pp.

319. 75 francs. British Journal for the History of Science, 10(2):171, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025878.

Hall:1977:BRSb

[Hal77b] Marie Boas Hall. Book review: Seventeenth and Eighteenth Centuries: The Foundations of Newton’s Alchemy or, ‘The Hunting of the Greene Lyon’. By Betty Jo Teeter Dobbs. Cambridge: Cambridge University Press, 1975. Pp. XV + 300. £10.50. British Journal for the History of Science, 10(3):262–264, November 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025852.

Hall:1979:RCY

[Hal79] A. R. Hall. Report of Council for the year 1978–1979. British Journal for the History of Science, 12(3):346–349, November 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hall:1980:BRE

[Hal80a] A. Rupert Hall. Book review: Eighteenth Century: Essays and Papers in the History of Modern Science. By Henry Guerlac. Baltimore & London: Johns Hopkins University Press, 1977. Pp. xix + 540. £14.00. British Journal for the History of Science, 13 (1):81–82, March 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026136.

Hall:1980:LE

[Hal80b] A. Rupert Hall. Letter to the Editor. British Journal for the History of Science, 13(1):84– 85, March 1980. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hall:1980:BRS

[Hal80c] Marie Boas Hall. Book review:

Scientific revolution: Robert Boyle and the English Revolution: A Study in Social and Intellectual Change. By J. R. Jacob. New York: Burt Franklin, 1977. Pp. 240. $18.95. British Journal for the History of Science, 13(1):76–77, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026133.

Hall:1983:RS

[Hal83a] A. Rupert (Alfred Rupert) Hall. The revolution in science, 1500–1750. Longman, London, UK, third edition, 1983. ISBN 0-582-49133-9 (paperback). viii + 373 pp. LCCN Q125 .H28 1983.

                                        URL http://www.h-net.org/
                                        review/hrev-a0b9f8-aa.

Hallam:1983:GGC

[Hal83b] A. (Anthony) Hallam. Great geological controversies. Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1983. ISBN 0-19-854431-6, 0-19-854430-8 (paperback). vii + 182 pp. LCCN QE11 .H35 1983. US$34.50, US$18.50.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0640/83002357-d.html.

Hall:1984:BRS

[Hal84a] A. Rupert Hall. Book review: Seventeenth Century: L. C. Palm and H. A. M. Snelders (eds.), Antoni van Leeuwenhoek 1632–1723. Amsterdam: Rodopi, 1983. Pp.

212. ISBN 90-6203-824-7. (No price stated). British Journal for the History of Science, 17 (1):115–116, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026053.

Hall:1984:ASN

[Hal84b] Marie Boas Hall. All scientists now: the Royal Society in the Nineteenth Century. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-26746-3. xii + 261 pp. LCCN Q41.L85 H35

1984. URL http://www.loc.
gov/catdir/description/
cam022/84007705.html;
http://www.loc.gov/
catdir/samples/cam031/
84007705.html;   http://
www.loc.gov/catdir/toc/
cam028/84007705.html.

Hall:1984:BRHb

[Hal84c] Marie Boas Hall. Book review: History of Technology: John Hubble Weiss, The making of technological man: the social origins of French engineering education. Cambridge, Mass, and London: MIT Press, 1982. Pp. xviii + 377. ISBN 0-262-23112-3. £21. British Journal for the History of Science, 17(3):329–330, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026645.

Hall:1984:BRN

[Hal84d] Marie Boas Hall. Book review: National Traditions in Science Karl Hufbauer, The formation of the German chemical community (1720–1795). Berkeley, Los Angeles, London: University of California Press, 1982. Pp. viii + 312.

£30, $52. ISBN 0-520-04318-9 (hardcover) and 0-520-04415-0 (paperback). British Journal for the History of Science, 17(1):89, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026023.

Hall:1984:BRHa

[Hal84e] Marie Boas Hall. Book reviews: History of Mathematical Sciences: Don H. Kennedy, Little Sparrow: A Portrait of Sophia Kovalevsky. Athens, Ohio and London: Ohio University Press, 1983. pp. ix + 341. £20.80, ISBN 0-8214-0692-2 (hardcover); £10.40, ISBN 0-8214-0703-1 (paperback). British Journal for the History of Science, 17(2):238, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026571.

Hall:1985:BRC

[Hal85a] Marie Boas Hall. Book review: Cecilia Payne-Gaposchkin, An autobiography and other recollections. Edited by Katherine Haramundanis. Cambridge: Cambridge University Press, 1984. Pp. 269. ISBN 0-521-25752-2. £19.50, $34.50. British Journal for the History of Science, 18(2):238–239, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026337.

Hall:1985:BRE

[Hal85b] Marie Boas Hall. Book review:

Elizabeth L. Eisenstein, The Printing Revolution in Early Modern Europe. Cambridge: Cambridge University Press, 1983. Pp. iv + 297. ISBN 0-521-25858-8. £25, $34.50, (Paperback £7.95, $9.95.). British Journal for the History of Science, 18(1):97–98, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026280.

Hall:1985:BRM

[Hal85c] Marie Boas Hall. Book review: Marin Mersenne, Correspondance, XV (1647). Edited by Cornelis de Waard and Armand Beaulieu. Paris: Centre National de la Recherche Scientifique, 1983. British Journal for the History of Science, 18(1):100, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026282.

Hall:1985:BRJ

[Hal85d] Vance M. D. Hall. Book review: Johannes Büttner (ed.) History of clinical chemistry. Berlin & New York: Walter de Gruyter, 1983. Pp. 91. ISBN 3-11-008912-2. DM 98. British Journal for the History of Science, 18(2):244, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026343.

Hall:1986:BRA

[Hal86a] A. Rupert Hall. Book review: Alan E. Shapiro (ed.). The Optical Papers of Isaac Newton. Volume 1. The Optical Lectures, 1670–1672. Cambridge: Cambridge University Press, 1984. Pp. xix + 627. ISBN 0-521-25248-2 (vol. 1). £75. British Journal for the History of Science, 19(2):218– 219, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026604.

Hall:1986:BRJa

[Hal86b] A. Rupert Hall. Book review: John Peter Oleson. Greek and Roman Mechanical Waterlifting Devices: the History of a Technology. Dordrecht: D. Reidel, 1984. Pp. xiv + 458 + 173 ills. ISBN 90-277-1693-5. Dfl. 345, £71.75, $58. British Journal for the History of Science, 19(2):221–222, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026606.

Hall:1986:BRJb

[Hal86c] Marie Boas Hall. Book review: Joseph Needham. Science and Civilisation in China. Vol. V (Chemistry and Chemical Technology), Part I by Tsien Tsuen-Hsuin, Paper and Printing. Cambridge: Cambridge University Press, 1985. xxv + 485 pp. ISBN 0-521-08690-6. £45.00. British Journal for the History of Science, 19(3): 348–349, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026520.

Hall:1986:BRN

[Hal86d] Marie Boas Hall. Book review: Norma E. Emerton. The Scientific Reinterpretation of Form. Ithaca and London: Cornell University Press, 1984. Pp. 318. ISBN 0-8014-1583-7. $29.95. British Journal for the History of Science, 19(3): 359–360, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026530.

Hall:1986:BRF

[Hal86e] Vance M. D. Hall. Book review:

F. W. J. McCosh. Boussingault: Chemist and Agriculturist. Dordrecht: Reidel, 1984. Pp. xii + 280. ISBN 90-277-1682-X. Dfl. 140.00. British Journal for the History of Science, 19(3):346–347, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026517.

Hall:1987:BRJ

[Hal87a] A. Rupert Hall. Book review:

E. J. Aiton. Leibniz. A Biography. Bristol and Boston: Adam Hilger, 1985. Pp. xiv + 370. ISBN 0-85274-470-6. £30.00. British Journal for the History of Science, 20(1): 97–98, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026451.

Hall:1987:BRR

[Hal87b] A. Rupert Hall. Book review: Roy Porter & Mikulas Teich (eds). Revolution in History. Cambridge: Cambridge University Press, 1986. Pp. x + 341. ISBN 0-521-25978-9, £25.00 (hardcover); 0-521-27784-1, £8.95 (paperback). British Journal for the History of Science, 20(3): 362–363, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026372.

Hall:1988:BRG

[Hal88a] A. Rupert Hall. Book review: A. G. Bennett and D. F. Edgar (eds). Isaac Barrow’s Optical Lectures 1667. Translated by H. C. Fay. London: The Worshipful Company of Spectacle Makers, Pp. 243 + xxxv. Figs. £25.00. (A printed typescript; orders to the Company, Apothecaries’ Hall, Black Friars Lane, London EC4V 6EL). British Journal for the History of Science, 21(2): 255–256, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026986.

Hall:1988:BRJ

[Hal88b] A. Rupert Hall. Book review: Joseph Needham with the collaboration of Ho Ping-Yü, Lu Gwei-Djen and Wang Ling. Science and Civilisation in China. Vol. 5, Chemistry and Chemical Technology. Part 7, Military Technology. The Gunpowder Epic. Cambridge and London: Cambridge University Press, 1986. Pp. xxxiii + 703. ISBN 0-521-30358-3. £50.00, $99.50. British Journal for the History of Science, 21(2):249–252, June 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026982.

Hall:1990:HMM

[Hal90] A. Rupert (Alfred Rupert)

Hall. Henry More — magic, religion, and experiment. Blackwell science biographies. Blackwell, Oxford, UK, 1990. ISBN 0-631-17295-5. xii + 304 pp. LCCN Q143.M77 H35 1990. US$60.00.

Hall:1991:PEL

[Hal91] Marie Boas Hall. Promoting experimental learning: experiment and the Royal Society 1660–1727. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-40503-3. xiii + 207 pp. LCCN Q41 .H33

1991. URL http://www.loc.
gov/catdir/description/
cam024/91027181.html;
http://www.loc.gov/
catdir/samples/cam031/
91027181.html;   http://
www.loc.gov/catdir/toc/
cam028/91027181.html.

Hall:1992:INA

[Hal92] A. Rupert (Alfred Rupert) Hall. Isaac Newton, adventurer in thought. Blackwell science biographies. Blackwell, Oxford, UK, 1992. ISBN 0-631-17906-2. xv + 468 pp. LCCN QC16.N7 H35 1992.

Hall:1993:AWL

[Hal93] A. Rupert (Alfred Rupert) Hall. All was light: an introduction to Newton’s Opticks. Clarendon Press, New York, NY, USA, 1993. ISBN 0-19-853985-1. vi + 252 pp. LCCN QC353 .H35 1993. US$52.50.

URL http://www.loc.gov/
catdir/enhancements/
fy0636/92043731-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0636/92043731-t.html.

Hamilton:1984:BRH

[Ham84] Beryl M. Hamilton. Book review: History of Natural History: Gordon L. Herries Davies, Sheets of Many Colours. The Mapping of Ireland’s Rocks. 1750–1890. Dublin: The Royal Dublin Society, 1983. Pp. xiv + 242 ISBN 0-86027-014-9 IR £15.00. British Journal for the History of Science, 17(3): 313, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026628.

Hamlin:1990:SIW

[Ham90] Christopher Hamlin. A science of impurity: water analysis in Nineteenth Century Britain. Adam Hilger Ltd., Bristol, UK, 1990. ISBN 0-7503-0042-6. xiii + 342 pp. LCCN TD257 .H35 1990.

Hannaway:1970:BRE

[Han70] Owen Hannaway. Book review:

Eighteenth Century: Herman Boerhaave: The Man and His Work. By G. A. Lindeboom. With a Foreword by E. Ashworth Underwood. London: Methuen. 1968. Pp. xx + 452. 35 plates + figs. £7 7s. British Journal for the History of Science, 5(1):102–103, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025372.

Hanson:1976:BRG

[Han76] A. E. Hanson. Book review:

Greece Greek Medicine. By E. D. Phillips. London: Thames and Hudson, 1973. Pp. 240. £4.50. British Journal for the History of Science, 9(1): 72–74, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025711.

Hankins:1985:BRE

[Han85a] Thomas Hankins. Book review: Elizabeth Chambers Patterson, Mary Somerville and the Cultivation of Science, 1815–1840. (Archives internationales d’histoire des idees,

102) Boston, The Hague, Dordrecht, Lancaster: Martinus Nijhoff Publishers, 1983. Pp. xiv + 264. ISBN 90-247-2823-1. British Journal for the History of Science, 18(1):110– 111, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026291.

Hankins:1985:SE

[Han85b] Thomas L. Hankins. Science and the Enlightenment. Cambridge history of science. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-24349-1, 0-521-28619-0 (paperback). viii + 216 pp. LCCN Q125 .H355 1985. URL http:

                                        //www.loc.gov/catdir/
                                        description/cam022/
                                        84016988.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam029/84016988.html.

Hartley:1967:SAV

[Har67] Harold Hartley. The Society of Arcueil. A view of French science at the time of Napoleon I. British Journal for the History of Science, 3(3):293– 295, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025055.

Harre:1969:WWH

[Har69] R. Harré. William Whewell and the history and philosophy of science. British Journal for the History of Science, 4(4):399–400, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025233.

Harre:1975:BRG

[Har75] Rom Harré. Book review:

General Causality and Scientific Explanation. Volume II. By William A. Wallace. Ann Arbor: University of Michigan Press, 1974. PP. vii + 422. $14. British Journal for the History of Science, 8(2): 174–175, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025642.

Harris:1976:BRT

[Har76] J. R. Harris. Book review:

Technology The Development of English Glassmaking 1540– 1640. By Eleanor S. Godfrey. Oxford: Oxford University Press: Clarendon Press, 1975. Pp. xii + 288. £10.50. British Journal for the History of Science, 9(3):330–331, November 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025693.

Haraway:1977:BRN

[Har77] Donna Haraway. Book review: Nineteenth and Twentieth Centuries: Life Science in the Twentieth Century. By Garland E. Allen. New York, London, Sydney and Toronto: John Wiley and Sons, 1975. Pp. xxv + 258. $12-50. British Journal for the History of Science, 10(3): 271–272, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025858.

Harrison:1979:BRB

[Har79a]  John Harrison.  Book re-

view: Bibliography and Bibliophily Newton and Newtoniana, 1672–1975; a Bibliography. By Peter and Ruth Wallis. Folkestone: Dawson, 1977. Pp. xxiv + 362. £30.00. British Journal for the History of Science, 12(2):219–220, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025757.

Harwood:1979:BRN

[Har79b] Jonathan Harwood. Book review: Nineteenth and Twentieth Centuries: Scientific Elite: Nobel Laureates in the United States. By Harriet Zuckerman. New York: The Free Press, 1977. Pp. xv + 335. $14.95. British Journal for the History of Science, 12(2):233– 234, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025767.

Harwood:1981:BRE

[Har81] Jonathan Harwood. Book review: Eighteenth to Twentieth Centuries German-Jewish Pioneers in Science, 1900– 1933: Highlights in Atomic Physics, Chemistry and Biochemistry. By David Nachmansohn. Berlin, Heidelberg, & New York: Springer-Verlag, 1979. Pp. xx + 388. DM60/ $33.00. British Journal for the History of Science, 14(3): 294–295, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025992.

Harman:20xx:MNP

[Har82a] P. M. Harman. Metaphysics and natural philosophy. Harvester Press, Brighton, Sussex, UK, 1982. ISBN 0-7108-0451-2. 192 pp. LCCN ????

Harman:1982:EFM

[Har82b] P. M. (Peter Michael) Harman. Energy, force, and matter: the conceptual development of Nineteenth-Century physics. Cambridge history of science. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24600-8, 0-521-28812-6 (paperback). ix + 182 pp. LCCN QC7

.H257. URL http://www.loc.

                                        gov/catdir/description/
                                        cam032/81017029.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0632/81017029-t.html.

Harman:1982:MNP

[Har82c] Peter M. Harman. Metaphysics and natural philosophy: the problem of substance in classical physics. Harvester Press, Brighton, Sussex, UK, 1982. ISBN 0-389-20305-X, 0-7108-0451-2. xvi + 168 pp. LCCN ????

Harwood:1982:BRS

[Har82d]  Jonathan Harwood.  Book

review: Science in Different Countries Zhores A. Medvedev, Soviet science. Oxford: Oxford University Press, 1979. Pp. xii + 262. £5.95. Linda L. Lubrano & Susan Gross Solomon (ed), The social context of Soviet science. Boulder & Colorado: Westview Press. Folkstone: Dawson, 1980. Pp. xv + 240. £10. British Journal for the History of Science, 15 (2):189–191, July 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025967.

Harris:1983:BRS

[Har83a] Ruth Harris. Book review: Science and Society Klaus Doerner, Madmen and the Bourgeoisie, a social history of insanity and psychiatry. Oxford: Basil Blackwell, 1981. Pp. v + 361. £16.00. British Journal for the History of Science, 16(3): 309–310, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026177.

Harwood:1983:BRS

[Har83b]  Jonathan Harwood.  Book

review: Science and Society Charles Webster (ed.), Biology, medicine and society, 1840– 1940. Cambridge: Cambridge University Press, 1981. Pp. 344. £22.50. British Journal for the History of Science, 16(3): 303–304, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026172.

Harman:1984:BRSa

[Har84a] P. M. Harman. Book review:

Seventeenth Century: Desmond M. Clarke, Descartes’ philosophy of science. Manchester: Manchester University Press, 1982. Pp. xiv + 249. ISBN 0-7190-0868-9. £19.00. British Journal for the History of Science, 17(1):114, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026051.

Harman:1984:BRSb

[Har84b] P. M. Harman. Book reviews:

Seventeenth Century: Margaret Dauler Wilson, Descartes. London: Routledge and Kegan Paul, 1982. Pp. xviii + 255. ISBN 0-7100-9208-3. £5.95. Ralph C. S. Walker, Kant. London: Routledge and Kegan Paul, 1982. Pp. xii + 201. ISBN 0-7100-0009-X. £4.95. British Journal for the History of Science, 17(1):115, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026052.

Harman:1985:WPS

[Har85a] Peter Michael Harman, editor. Wranglers and physicists: studies on Cambridge physics in the nineteenth century, Studies on Cambridge mathematical physics in the nineteenth century. Manchester University Press, Manchester, UK, 1985. ISBN 0-7190-1756-4. LCCN QC9.G7 W73 1985. US$27.00 (U.S.).

Hart:1985:PF

[Har85b] Clive Hart. The prehistory of flight. University of California Press, Berkeley, CA, USA, 1985. ISBN 0-520-05213-7. xvii + 279 pp. LCCN TL516 .H27 1985.

Harris:1986:BRA

[Har86] Ruth Harris. Book review: Annette Smith. Gobineau et l’Histoire Naturelle. Geneva and Paris: Libraire Droz. 1984. Pp. 262. British Journal for the History of Science, 19(3):

356–357, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026526.

Harrington:1987:MMD

[Har87a] Anne Harrington. Medicine, mind, and the double brain: a study in Nineteenth-Century thought. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-08465-3. xiii + 336 pp. LCCN QP385.5 .H37 1987. US$39.50.

Harwood:1987:BRG

[Har87b] Jonathan Harwood. Book review: Georg Lilienthal. Der ‘Lebensborn e.V.’: Ein Instrument Nationalsozialistischer Rassenpolitik. Stuttgart/New York: G. Fischer, 1985. Pp. viii + 264. ISBN 3-437-10939-1. British Journal for the History of Science, 20 (1):108, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026462.

Harrington:1989:BRE

[Har89a] Anne Harrington. Book review: Edwin Clarke & L. S. Jacyna. Nineteenth-Century Origins of Neuroscientific Concepts. Berkeley/Los Angeles/ London: University of California Press, 1987. Pp. viii + 593. ISBN 0-520-05694-9. No price given. British Journal for the History of Science, 22 (1):97–98, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026696.

Harris:1989:MMM

[Har89b] Ruth Harris. Murders and madness: medicine, law and society in the Fin de Siècle. Oxford historical monographs. Clarendon Press, New York, NY, USA, 1989. ISBN 0-19-822991-7, 0-19-820259-8 (paperback). viii + 366 + 4 pp. LCCN HV6133 HAR 1989.

Harwood:1989:GEE

[Har89c] Jonathan Harwood. Genetics, eugenics and evolution. British Journal for the History of Science, 22(3):257–265, September 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026895.

Harrington:1996:RSH

[Har96] Anne Harrington. Reenchanted science: holism in German culture from Wilhelm II to Hitler. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-02142-2 (hardcover). xxiv + 309 pp. LCCN QH501 .H37

                                        1996. URL http://www.loc.
                                        gov/catdir/description/
                                        prin021/95048463.html;
                                        http://www.loc.gov/
                                        catdir/toc/prin031/
                                        95048463.html.

Harman:1998:NPJ

[Har98a] P. M. (Peter Michael) Harman. The natural philosophy of James Clerk Maxwell. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-56102-7, 0-521-00585-X (paperback). xiv + 232 pp. LCCN QC16.M4 H37

1998. URL http://www.loc.
gov/catdir/description/
cam028/97035703.html;
http://www.loc.gov/
catdir/toc/cam021/
97035703.html;   http://
www.zentralblattmath.org/
zmath/en/search/?an=1005.
01006.

Harrison:1998:BPR

[Har98b] Peter Harrison. The Bible, Protestantism, and the rise of natural science. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-59196-1 (hardcover). xi + 313 pp. LCCN BS650 .H37

1998. URL http://www.loc.
gov/catdir/description/
cam028/97020967.html;
http://www.loc.gov/
catdir/enhancements/
fy0732/97020967-b.html;
http://www.loc.gov/
catdir/toc/cam023/
97020967.html.

Harman:1990:SLP

[Har02] P. M. (Peter Michael) Harman, editor. The scientific letters and papers of James Clerk Maxwell. Cambridge University Press, Cambridge, UK, 1990–2002. ISBN 0-521-25625-9 (vol. 1), 0-521-25626-7 (vol. 2), 0-521-25627-5 (vol. 3). xxviii + 748 (vol. 1), xxx + 999 + 15 (vol. 2), xxvii + 932 pp. LCCN QC670 .M385

1990. URL http://www.loc.
gov/catdir/description/
cam023/89000452.html;
http://www.loc.gov/
catdir/enhancements/
fy0732/89000452-b.html;
http://www.loc.gov/
catdir/toc/cam021/
89000452.html.

Hassall:1972:BRR

[Has72] Mark Hassall. Book review: Rome The Roman Land Surveyors. An Introduction to the Agrimensores. By O. A. W. Dilke. Newton Abbot: David & Charles, 1971. Pp. 260. £3.90. British Journal for the History of Science, 6(2): 210, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025298.

Hatch:1987:BRP

[Hat87] Robert A. Hatch. Book review: Paolo Galluzzi & Maurizio Torrini (eds). Le Opere Dei Discepoli di Galileo Galilei. Volume II: Carteggio 1649–1656. Florence: Giunti-Barbèra, 1984. Pp. xiv + 494. No price given. British Journal for the History of Science, 20(3): 360–362, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026371.

Hatfield:1990:NNT

[Hat90] Gary C. (Gary Carl) Hatfield.

The natural and the normative: theories of spatial perception from Kant to Helmholtz. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-08086-9. xii + 366 pp. LCCN BF469 .H34 1990.

Hays:1974:SCL

[Hay74] J. N. Hays. Science in the city: The London Institution, 1819–1840. British Journal for the History of Science, 7 (2):146–162, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025201.

Haynes:1994:FSR

[Hay94]  Roslynn  D.

(Roslynn Doris) Haynes. From Faust to Strangelove: representations of the scientist in western literature. The Johns Hopkins University Press, Baltimore, MD, USA, 1994. ISBN 0-8018-4801-6 (hardcover), 0-8018-4983-7 (paperback). 417 pp. LCCN PN56.5.S35 H39

1994. URL http://www.loc.
gov/catdir/bios/jhu052/
94007023.html;
http://www.loc.gov/
catdir/description/
jhu051/94007023.html.

Hutton:1990:HMT

[HC90] Sarah Hutton and Robert Crocker, editors. Henry More (1614– 1687) tercentenary studies, volume 127 = of Archives internationales d’histoire des idées; International archives of the history of ideas; 127. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0095-5. xi + 254 pp. LCCN B1299.M64 H46 1990.

Holton:1982:AEH

[HE^{+}82] Gerald James Holton, Yehuda Elkana, et al., editors. Albert Einstein: historical and cultural perspectives; the centennial symposium in Jerusalem; Jerusalem Einstein Centennial Symposium 14–23 March 1979. Princeton University Press, Princeton, NJ, USA, 1982. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Hammond:1992:WBH

[HE92] P. W. Hammond and Harold Egan. Weighed in the balance: a history of the Laboratory of the Government Chemist. Her Majesty’s Stationary Office, London, England, 1992. ISBN 0-11-515302-0. 372 pp. LCCN QD51.5.G72 L334 1992. US$9.95.

Heathcote:1964:BRN

[Hea64] N. H. De V. Heathcote. Book review: Nobel: The Man and his Prizes. Edited by the Nobel Foundation. Pp. xii + 690. Frontispiece. Amsterdam, London and New York: Elsevier Publishing Company, 1962. 60s. net. British Journal for the History of Science, 2(1):79–80, June 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025091.

Hearnshaw:1986:ASO

[Hea86] J. B. Hearnshaw. The analysis of starlight: one hundred and fifty years of astronomical spectroscopy. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25548-1. xv + 531 pp. LCCN QB465 .H43 1986.

Hearnshaw:1987:BM

[Hea87a] L. S. Hearnshaw. Back matter. British Journal for the History of Science, 20(3): 374, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026379.

Hearnshaw:1987:BRG

[Hea87b] L. S. Hearnshaw. Book review:

Graham Bird. William James. London: Routledge & Kegan Paul, 1986. Pp. ix + 221. ISBN 0-7100-9602-X. £15.95. British Journal for the History of Science, 20(3):369, July 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026377.

Hearnshaw:1987:SMP

[Hea87c] L. S. (Leslie Spencer) Hearnshaw. The shaping of modern psychology. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1987. ISBN 0-7102-0576-7. viii + 423 pp. LCCN BF81 .H33 1987. US$16.95 (U.K.), (US$21.95 U.S.).

Hearnshaw:1990:ASO

[Hea90] J. B. Hearnshaw. The analysis of starlight: one hundred and fifty years of astronomical spectroscopy. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-39916-5, 0-521-25548-1. xv + 531 pp. LCCN QB465 .H43

                                        1990. URL http://www.loc.
                                        gov/catdir/description/
                                        cam031/91105528.html.

Heilbron:1968:BRS

[Hei68] J. L. Heilbron. Book review: Sixteenth and Seventeenth Centuries: Einstein. The Man and his Achievement. (Ed.) G. J. Whitrow. London: British Broadcasting Corporation. 1967. Pp. xv + 94, bibliog. 18s. 6d. British Journal for the History of Science, 4(2): 195–196, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025168.

Heimann:1971:BRE

[Hei71a] P. M. Heimann. Book review:

Eighteenth Century: Atoms and Powers. An Essay on Newtonian Matter-Theory and the Development of Chemistry. By Arnold Thackray. Harvard University Press & Oxford University Press. 1970 [1971]. Pp. xxiii + 326. £4.20. British Journal for the History of Science, 5(4):419–420, December 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025406.

Heimann:1971:BRP

[Hei71b] P. M. Heimann. Book review:

Physical sciences: Rutherford and Boltwood: Letters on Radioactivity. Ed. by Lawrence Badash. New Haven & London: Yale University Press. Pp. xxii + 378. 1969. £5.621/2. British Journal for the History of Science, 5(3):301–302, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025329.

Heimann:1971:FTM

[Hei71c] P. M. Heimann. Faraday’s theories of matter and electricity. British Journal for the History of Science, 5(3):235–257, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Heimann:1972:UUP

[Hei72] P. M. Heimann. The unseen universe: Physics and the philosophy of nature in Victorian Britain. British Journal for the History of Science, 6(1):73–79, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025264.

Heimann:1973:BRN

[Hei73] P. M. Heimann. Book review: Nineteenth Century: The Selected Correspondence of Michael Faraday. Ed., on behalf of the Royal Institution of Great Britain, by L. Pearce Williams with the assistance of Rosemary Fitzgerald and Oliver Stallybrass. 2 volumes. Cambridge: Cambridge University Press. 1971. Pp. xii + 1079. £21. British Journal for the History of Science, 6(4): 451–452, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025522.

Heimann:1974:BRN

[Hei74a] P. M. Heimann. Book review: Nineteenth and Twentieth Centuries: Selected Writings of Hermann von Helmholtz. Edited, with an Introduction, and translated in part by Russell Kahl. Middletown, Connecticut: Wesleyan University Press, 1971. Pp. xlvi + 542. $25. British Journal for the History of Science, 7(1):95, March 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025192.

Heimann:1974:BRP

[Hei74b] P. M. Heimann. Book review:

Philosophy Hume’s Philosophical Development. A Study of his Methods. By James Noxon. Oxford: Clarendon Press: Oxford University Press, 1973. Pp. xiv + 197. £2.50. British Journal for the History of Science, 7(3):287, November 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025536.

Heilbron:1975:BRN

[Hei75] J. L. Heilbron. Book review:

Nineteenth and Twentieth Centuries: The Master of Light. A Biography of Albert A. Michelson. By Dorothy Michelson Livingston. New York: Charles Scribner’s Sons, 1973. Pp. xi + 376. $12.50. British Journal for the History of Science, 8(2):187–188, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025656.

Heimann:1976:BRC

[Hei76a] P. M. Heimann. Book review:

Chemistry Philosophical Chemistry in the Scottish Enlightenment.

The Doctrines and Discoveries of William Cullen and Joseph Black. By A. L. Donovan. Edinburgh: Edinburgh University Press, 1975. Pp. x + 343. £7.00. British Journal for the History of Science, 9(3):328– 329, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025691.

Heimann:1976:BRS

[Hei76b] P. M. Heimann. Book review: Seventeenth Century: The Mathematical Papers of Isaac Newton. Volume VI. 1684–1691. Ed. by D. T. Whiteside. With the assistance in publication of M. A. Hoskin and A. Prag. Cambridge: Cambridge University Press, 1974. Pp. xxxv + 614. £25.00. British Journal for the History of Science, 9(1):75–77, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025713.

Heimann:1977:BRN

[Hei77] P. M. Heimann. Book review: Nineteenth and Twentieth Centuries: Visionary Physics. Blake’s Response to Newton. By Donald Ault. Chicago and London: University of Chicago Press, 1974. Pp. xv + 229. £6.25. British Journal for the History of Science, 10(2):177, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025882.

Heilbron:1986:DUM

[Hei86] J. L. Heilbron. The dilemmas of an upright man: Max Planck as spokesman for German science. University of California Press, Berkeley, CA, USA, 1986. ISBN 0-520-05710-4. xiii + 238 + 16 pp. LCCN QC16.P6 H45 1986.

Hein:1988:AHP

[Hei88]  Wolfgang  Ha-

gen Hein. Alexander von Humboldt und die Pharmazie. (German) [Alexander von Humboldt and Pharmacy], volume 56 of Veröffentlichungen der Internationalen Gesellschaft für Geschichte der Pharmazie e. V., 0074-9729. Wissenschaftliche Verlagsgesellschaft, Stuttgart, Germany, 1988. ISBN 3-8047-0984-2. 130 pp. LCCN Q143.H9 H38 1988.

Heims:1991:CG

[Hei91] Steve J. Heims. The cybernetics group. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-08200-4. xii + 334 pp. LCCN H62.5.U5 H45 1991.

Heidelberger:1993:ISN

[Hei93] Michael Heidelberger. Die innere Seite der Natur: Gustav Theodor Fechners wissenschaftlichphilosophische Weltauffassung.

(German) [The internal side of nature: Gustav Theodor Fechners scientific-philosophical view of the world], volume 60 of Philosophische Abhandlungen. V. Klostermann, Frankfurt am Main, Germany, 1993. ISBN 3-465-02590-3. 457 + 1 pp. LCCN B3237 .H45 1993.

Hellman:1963:WTB

[Hel63] C. Doris Hellman. Was Tycho Brahe as influential as he thought? British Journal for the History of Science, 1(4): 295–324, December 1963. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024905.

Hendry:1980:RCH

[Hen80] John Hendry. Reminiscence and the contemporary history of science. British Journal for the History of Science, 13(3): 258–262, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026200.

Hendry:1981:BRT

[Hen81] John Hendry. Book review: Twentieth Century The Uranium People. By Leona Marshall Libby. New York: Crane Russak; Charles Scribner’s Sons, 1979. PP. x + 341. $15.95. Scientists in Power. By Spencer R. Weart. Cambridge, Mass. and London: Harvard University Press, 1979. PP. xiii + 343. $17.50. Robert Oppenheimer: Letters and Recollections. Edited by Alice Kimball Smith and Charles Weiner. Cambridge, Mass. and London: Harvard University Press, 1980. Pp. xi + 376. $20.00. British Journal for the History of Science, 14(1):97–99, March 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026084.

Hendry:1982:BRS

[Hen82a] John Hendry. Book review:

Science Studies Nevill Mott (ed.), The beginnings of solid state physics. London: The Royal Society, 1980. Pp. iv + 177. £9.75 (U.K.)/£10.25 (overseas). British Journal for the History of Science, 15(3): 309–310, November 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026226.

Hendry:1982:BRNb

[Hen82b] John Hendry. Book review:

The New Physics Lawrence Badash, Joseph O. Hirschfelder and Herbert P. Broida (ed.), Reminiscences of Los Alamos, 1943–1945. Dordrecht, Boston & London: D. Reidel, 1980. Pp. xxi + 188. Df150/$26.50; Df120/$9.95 (paperback). British Journal for the History of Science, 15 (2):201–202, July 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025974.

Hendry:1982:BRNa

[Hen82c] John Hendry. Book review: The New Physics Loyd S. Swenson Jr, C. P. Snow, Howard Stein and Ilya Prigogine, Albert Einstein: four commemorative lectures. Austin: The Humanities Research Center, The University of Texas at Austin, 1979. Pp. 64. $3.50. A. P. French (ed.), Einstein. A centenary volume. London: Heinemann Educational Books, 1979. Pp. xx + 332. £10.50. Colette M. Kinnon with A. N. Kholodinin and J. G. Richardson, The impact of modern scientific ideas on society: in commemoration of Einstein. Dordrecht, Boston & London: D. Reidel, 1981. Pp. xiv + 203. Df150.00/$26.50. British Journal for the History of Science, 15(2):200–201, July 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025973.

Henry:1982:AEC

[Hen82d] John Henry. Atomism and eschatology: Catholicism and natural philosophy in the Interregnum. British Journal for the History of Science, 15(3): 211–239, November 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026219.

Hendry:1983:BRMa

[Hen83a]  John Hendry.  Book re-

view: Mathematical Sciences: J. A. Benett, The mathematical science of Christopher Wren. Cambridge: Cambridge University Press, 1982. Pp.

148. £15.00. ISBN 0-521-24608-3. British Journal for the History of Science, 16(3):291– 292, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026162.

Hendry:1983:BRMb

[Hen83b] John Hendry. Book reviews:

Mathematical Sciences: J. L. Heilbron & Bruce R. Wheaton, Literature on the history of physics in the twentieth century. Berkeley: University of California Office for History of Science and Technology, 1981. Pp. xi + 485. No price stated. ISBN 0-918102-05-7. David De Vorkin, The history of modern astronomy and astrophysics. A selected, annotated, bibliography. New York: Garland Publishing, 1982. Pp. xxvii + 434. $65.00. ISBN 0-8240-9283-X. British Journal for the History of Science, 16(3):292– 293, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026163.

Henry:1983:BRL

[Hen83c] John Henry. Book review: Lives and Works M. R. G. Spiller, ‘Concerning natural experimental philosophie’: Meric Casaubon and the Royal Society. The Hague: Martinus Nijhoff, 1980. Pp. ix + 232. British Journal for the History of Science, 16(2):205– 207, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026247.

Henry:1983:BRM

[Hen83d] John Henry. Book review: Mathematical Sciences: Edward Grant, Much ado about nothing: theories of space and vacuum from the Middle Ages to the Scientific Revolution. Cambridge: Cambridge University Press, 1981. Pp. xiii + 456. £30.00. British Journal for the History of Science, 16(3): 294–295, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026165.

Hendry:1984:BRHa

[Hen84a] John Hendry. Book review: History of Mathematical Sciences: A. Rupert Hall, The Revolution in Science, 1500–1750. London: Longman, 1983, Pp. viii + 373. ISBN 0-582-49133-9. £8.95 (paperback). British Journal for the History of Science, 17(2):233, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026565.

Hendry:1984:BRHc

[Hen84b]  John Hendry.  Book re-

view: History of Mathematical Sciences: Christine Blondel, Ampère et la Création de l’électrodynámique (1820– 1827). Paris: Bibliotheque Nationale, 1982. Pp. 202. ISBN 2-7177-1643-2. 175 F (paperback). British Journal for the History of Science, 17(2):238, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hendry:1984:BRHb

[Hen84c] John Hendry. Book review:

History of Mathematical Sciences: Simon Schaffer, Science and the English Enlightenment (Video Cassette). Poulshot, Wiltshire: Sussex Publications, 1982. 281/2 minutes. £45.42 Simon Schaffer, Newton, Central TV for Channel 4: Crucible Series, 1982. Colin A. Ronan. The Cambridge Illustrated History of World Science, Feltham, Middx: Newnes, 1983: Pp. 543. ISBN 0-600-38423-3. £12.95. British Journal for the History of Science, 17(2):236–237, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026569.

Hendry:1984:BRN

[Hen84d]  John Hendry.  Book re-

view: National Traditions in Science Leslie Hannah, Engineers, managers and politicians. The first fifteen years of nationalised electricity supply in Britain. Research by Margaret Ackrill, Frances Bostock, Rachel Lawrence, Judy Slinn and Stephanie Zarach. London: MacMillan, 1982. Pp. xiii + 336. £15.95. ISBN 0-333-22087-0. British Journal for the History of Science, 17 (1):99–100, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026035.

Hendry:1984:BRSa

[Hen84e] John Hendry. Book review: Seventeenth Century: Zev Bechler (ed.), Contemporary Newtonian research. Dordrecht and Boston: D. Reidel, 1982. Pp. vi + 241. DF1. 95.00. ISBN 90-277-1303-0. British Journal for the History of Science, 17 (1):113–114, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026050.

Hendry:1984:BRP

[Hen84f] John Hendry. Book reviews: Physics and its concepts: Peter M. Harman, Energy, force and matter. The conceptual development of nineteenth-century physics. Cambridge: Cambridge University Press, 1982. Pp. ix + 182. £13.50 (cased, ISBN 0-521-24600-8), £5.50 (paper, ISBN 0-521-28812-6). Peter M. Harman, Metaphysics and natural philosophy. The problem of substance in classical physics. Brighton: Harvester Press, 1982. Pp. xvi + 168. £18.95. ISBN 0-7108-0451-2. British Journal for the History of Science, 17(2): 221–223, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026556.

Hendry:1984:BRSb

[Hen84g]  John Hendry.  Book re-

views: Science Studies Magda Whitrow (ed), Isis cumulative bibliography 1913–1965, volumes 4 and 5: civilizations and periods. London: Mansell, 1982. Volume 4: pp. xviii + 457, ISBN 0-7201-1642-2. Volume 5: pp. xi + 593, ISBN 0-7201-1643-0. £100 the set. British Journal for the History of Science, 17(3):335– 336, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026652.

Hendry:1984:CPT

[Hen84h] John Hendry, editor. Cambridge physics in the thirties.

Adam Hilger Ltd., Bristol, UK, 1984. ISBN 0-85274-761-6. 209 pp. LCCN QC51.G72 C352 1984.

Hendry:1984:CQM

[Hen84i] John Hendry. The creation of quantum mechanics and the Bohr–Pauli dialogue, volume 14 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1648-X. xi + 177 pp. LCCN QC173.98 .H46 1984.

Henry:1984:BRAa

[Hen84j] John Henry. Book review: Antiquity to the Renaissance Norman Kretzmann (ed.), Infinity and continuity in ancient medieval thought. Ithaca and London: Cornell University Press, 1982. Pp. 367. ISBN 0-8014-1444-X. £20.75, $36.10. British Journal for the History of Science, 17(1): 103–104, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026039.

Henry:1984:BRAb

[Hen84k] John Henry. Book review: Antiquity to the Renaissance Walter Pagel, Joan Baptista Van Helmont: reformer of science and medicine. Cambridge: Cambridge University Press, 1982. Pp. xi + 219. ISBN 0-521-24807-8. £20.00. British Journal for the History of Science, 17(1):106, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026042.

Henry:1984:BRH

[Hen84l] John Henry. Book review: History of Mathematical Sciences: Barbara J. Shapiro, Probability and certainty in seventeenthcentury England: a study of the relationships between natural science, religion, history, law, and literature. Princeton, New Jersey: Princeton University Press, 1983. Pp. x + 347. ISBN 0-691-05379-0. £26.00. British Journal for the History of Science, 17(2):232, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026564.

Henry:1984:BRS

[Hen84m] John Henry. Book review: Seventeenth Century: James R. Jacob, Henry Stubbe, radical protestantism and the early enlightenment. Cambridge: University Press, 1983. Pp. viii + 222. ISBN 0-521-24876-0. £19.50. British Journal for the History of Science, 17(1): 111–112, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026048.

Hendry:1985:BRE

[Hen85a] John Hendry. Book review: Emerson W. Pugh, Memories that shaped an industry. Decisions leading to IBM system 360. Cambridge, Mass., and London: MIT Press, 1984. Pp. x + 323. ISBN 0-262-16094-3. £23.75. British Journal for the History of Science, 18(3): 348, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026389.

Henry:1985:BRF

[Hen85b] John Henry. Book review: Frances A. Yates, Renaissance and reform: the Italian contribution. Collected essays volume II. London: Routledge and Kegan Paul, 1983. Pp. xii + 273. ISBN 0-7100-9530-9. £15.95. British Journal for the History of Science, 18(3): 362, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026400.

Henry:1985:BRJ

[Hen85c] John Henry. Book review: J. E. McGuire and Martin Tamny. Certain Philosophical Questions: Newton’s Trinity Notebook. Cambridge: Cambridge University Press, 1983. Pp. xii + 519. ISBN 0-521-23164-7. £50.00. British Journal for the History of Science, 18(3):354–356, November 1985.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026395.

Hendry:1986:BRW

[Hen86a] John Hendry. Book review: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Scientific Correspondence with Bohr, Einstein, Heisenberg a.o. Berlin, Heidelberg and New York: Springer, 1985. Pp. xxix + 783. ISBN 3-540-13609-6. DM 298.00. British Journal for the History of Science, 19 (3):348, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026519.

Hendry:1986:BRJ

[Hen86b]  John Hendry.  Book re-

views: Jagdish Mehra & Helmut Rechenberg. The Historical Development of Quantum Theory. Volume 1, The Quantum Theory of Planck, Einstein, Bohr and Sommerfeld: Its Foundations and the Rise of its Difficulties 1900–1925. Volume 2, The Discovery of the Quantum Mechanics 1925. Volume 3, The Formulation of Matrix Mechanics and its Modifications 1925–1926. Volume 4, The Fundamental Equations of Quantum Mechanics 1925–1926 and The Reception of the New Quantum Mechanics 1925–1926. New York, Heidelberg and Berlin: Springer-Verlag, 1982. Volume 1 in two parts, pp. xlvii + 372, vi + 506. ISBN 3-540-90642-8, 3-540-90667-3. DM 75, DM 85. Volume 2, pp. vii + 355. ISBN 3-540-90674-6. DM 65. Volume 3, pp. vii + 334. ISBN 3-540-90675-4. DM 75. Volume 4, pp. viii + 322. ISBN 3-540-90680-0. DM 75. Andrew Pickering. Constructing Quarks. A Sociological History of Particle Physics. Edinburgh: Edinburgh University Press, 1984. Pp. xi + 468. ISBN 0-85224-458-4 £20. British Journal for the History of Science, 19 (2):206–208, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026596.

Hendry:1986:JCM

[Hen86c] John Hendry. James Clerk Maxwell and the theory of the electromagnetic field. Adam Hilger Ltd., Bristol, UK, 1986. ISBN 0-85274-563-X. xix + 305 pp. LCCN QC665.E4 H46

                                        1986. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0745/86196851-d.html.

Heniger:1986:HAV

[Hen86d] J. Heniger. Hendrik Adriaan van Reede tot Drakenstein (1636–1691) and Hortus Malabaricus: a contribution to the history of Dutch colonial botany. A. A. Balkema, Rotterdam, The Netherlands, 1986. ISBN 90-6191-681-X. xvi + 295 pp. LCCN QK358.R43 H46

1986. URL http://www.loc.
gov/catdir/enhancements/
fy0647/87170134-d.html.

Henry:1986:BRP

[Hen86e] John Henry. Book review: Peter Alexander. Ideas, Qualities and Corpuscles: Locke and Boyle on the Natural World. Cambridge: Cambridge University Press, 1985. Pp. 331. ISBN 0-521-26707-2. £27.50. British Journal for the History of Science, 19(3):357– 358, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026527.

Henry:1986:OCB

[Hen86f] John Henry. Obituary: Charles Bernard Schmitt: 1933–1986. British Journal for the History of Science, 19(3):337, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026510.

Hendry:1987:BRD

[Hen87a] John Hendry. Book review:

David Gooding & Frank A. J. L. James (eds). Faraday Rediscovered. Essays on the Life and Work of Michael Faraday, 1791–1867. London: Macmillan, 1986. Pp. xiv + 258. ISBN 0-333-39320-1. £35.00.

British Journal for the History of Science, 20(2):237– 238, April 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026323.

Hendry:1987:BRM

[Hen87b] John Hendry. Book review: Martin Kenney. Biotechnology: The University-Industrial Complex. New Haven and London: Yale University Press, 1986. Pp. xvi + 306. ISBN 0-300-03392-3. £22.50. British Journal for the History of Science, 20(3):353–354, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026364.

Henry:1987:BRJ

[Hen87c] John Henry. Book review: A. J. Holland (ed.). Philosophy, Its History and Historiography. Dordrecht: Reidel, 1985. Pp. x + 335. ISBN 90-277-1945-4. £37.50, U.S. $49.00. British Journal for the History of Science, 20(1):120– 121, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026472.

Hendry:1988:BRD

[Hen88a] John Hendry. Book review: Donald J. Wilcox. The Measure of Times Past. Pre-Newtonian Chronologies and the Rhetoric of Relative Time. Chicago and London: University of Chicago Press, 1987. Pp. ix + 292. ISBN 0-226-89721-4. £21.50, $32.95. British Journal for the History of Science, 21(3): 375, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026846.

Hendry:1988:BRH

[Hen88b]  John Hendry.  Book re-

view: Helge Kragh. An Introduction to the Historiography of Science. Translated by Jean Lundskjær-Nielsen. Cambridge: Cambridge University Press, 1987. Pp. viii + 235. ISBN 0-521-33360-1. £19.50, $32.50. British Journal for the History of Science, 21(2): 267–268, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026996.

Hendry:1988:BRJb

[Hen88c]  John Hendry.  Book re-

view: Jagdish Mehra & Helmut Rechenberg. The Historical Development of Quantum Theory. Volume 5. Erwin Schrödinger and the Rise of Wave Mechanics. Part 1. Schrödinger in Vienna and Zurich, 1887–1925 and Part 2. The Creation of Wave Mechanics; Early Response and Applications, 1925– 1926. Berlin, Heidelberg, New York: Springer-Verlag, 1987. Pp. xix + 366 and viii + 615. ISBN 3-540-96284-0 and 96377-4. DM 148.00 (hardcover) and 98.00 (paperback). British Journal for the History of Science, 21(3):371–372, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026844.

Hendry:1988:BRJa

[Hen88d] John Hendry. Book review: John Stachel (ed.). The Collected Papers of Albert Einstein. Volume 1. The Early Years, 1879–1902. Princeton: Princeton University Press, 1987. Pp. lxvi + 433. ISBN 0-691-08407-6. £35. English translation by Anna Beck, £14.10 (paperback), £6.25 (microfiche), available only to purchasers of main volume. British Journal for the History of Science, 21(2):262, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026992.

Henry:1988:BRB

[Hen88e]  John Henry.  Book

review: Barry Brundell. Pierre Gassendi: From Aristotelianism to a New Natural Philosophy. Dordrecht, Boston, Lancaster, Tokyo: D. Reidel, 1987. Pp. x + 251. ISBN 90-277-2428-8. £49.00, $64.00. British Journal for the History of Science, 21(2):260–261, June 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026991.

Henry:1988:OMS

[Hen88f] John Henry. The origins of modern science: Henry Oldenburg’s contribution. British Journal for the History of Science, 21 (1):103–109, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026867.

Hendry:1989:BRA

[Hen89a]  John Hendry.  Book re-

view: Abraham Pais. Inward Bound: Of Matter and Forces in the Physical World. Oxford: Oxford University Press, 1988. Pp. xiv + 665. ISBN 0-19-851997-4, £9.95. (Paperback: original hardback edition 1986.). British Journal for the History of Science, 22 (4):450, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026929.

Hendry:1989:BRB

[Hen89b] John Hendry. Book review:

Behram N. Kursunoglu & Eugene P. Wigner (eds). Reminiscences about a Great Physicist: Paul Adrien Maurice Dirac. Cambridge: Cambridge University Press, 1987. Pp. xviii + 297. ISBN 0-521-34013-6. £30.00. C.W. Kilmister (ed.). Schrödinger: Centenary Celebration of a Polymath. Cambridge: Cambridge University Press, 1987. Pp. x + 253. ISBN 0-521-34017-9. £30.00. British Journal for the History of Science, 22(1):87–88, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026687.

Hendry:1989:BRN

[Hen89c]  John Hendry.  Book re-

view: Nicholas Maxwell, From Knowledge to Wisdom. A Revolution in the Aims and Methods of Science (Basil Blackwell, 1987). Pp. vi + 298. ISBN 0-631-15641-0. (paper: original hardback 1984). No price given. British Journal for the History of Science, 22(2): 246–247, July 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026665.

Hendry:1989:BRP

[Hen89d] John Hendry. Book review: Peter Galison. How Experiments End. Chicago: University of Chicago Press, 1987. Pp. xii + 330. ISBN 0-226-27915-4. £11.95 (paper; also in hardback). British Journal for the History of Science, 22 (1):79–81, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026681.

Hendry:1989:BRS

[Hen89e] John Hendry. Book review:

Science in Context, Volume 1. Cambridge: Cambridge University Press, 1987. Two issues per volume. ISSN 0269-8897. Subscription for volume 1 £31.00; $65.00 for institutions, £19.00; $37.00 for individuals. British Journal for the History of Science, 22(1): 102–105, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026703.

Hendry:1989:IFG

[Hen89f] John Hendry. Innovating for failure: government policy and the early British computer industry. History of computing. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-08187-3. xvi + 240 pp. LCCN HD9696.C63 G763 1989.

Henry:1989:BRR

[Hen89g] John Henry. Book review: R.

G. W. Anderson and Christopher Lawrence, (eds). Science, medicine and dissent: Joseph Priestley (1733–1804). Papers celebrating the 250th anniversary of the birth of Joseph Priestley, together with a catalogue of an exhibition held at the Royal Society and the Wellcome Institute for the History of Medicine. London: Wellcome Trust/Science Museum, 1987. Pp. xii + 105. ISBN 0-901805-28-9. £9.95. British Journal for the History of Science, 22(3): 388–390, September 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026906.

Henderson:1991:DBS

[Hen91] Janice Adrienne Henderson. On the distances between sun, moon, and Earth1 according to Ptolemy, Copernicus, and Reinhold, volume 30 of Studia Copernicana; Brill’s series, 0925-6806; vol. 1. E. J. Brill, Leiden, The Netherlands, 1991. ISBN 90-04-09378-8. xiv + 220 pp. LCCN QB991.C66 H46 1991.

Herivel:1965:OAK

[Her65a] John Herivel. Obituary: Alexandre Koyré (1892–1964). British Journal for the History of Science, 2(3):257–259, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024941.

Herivel:1965:NFS

[Her65b] John W. Herivel. Newton’s first solution to the problem of Kepler motion. British Journal for the History of Science, 2(4): 350–354, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024891.

Herivel:1966:AFT

[Her66]  J.  W.

Herivel. Aspects of French theoretical physics in the Nineteenth Century. British Journal for the History of Science, 3(2):109–132, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Herivel:1973:ER

[Her73] John Herivel. Essay review.

British Journal for the History of Science, 6(4):429–432, December 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025505.

Herivel:1980:SNB

[Her80] J. W. Herivel. Short notices of books: Isaac Newton’s papers and letters on natural philosophy. Edited by I. B. Cohen. 2nd edn. Cambridge, Mass. & London: Harvard University Press, 1978. Pp. 450 + xii. £18.55. British Journal for the History of Science, 13(2):181, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025913.

HerriesDavies:1983:SMC

[Her83] G. L. Herries Davies. Sheets of many colours: the mapping of Ireland’s rocks, 1750–1890, volume 4 of Historical studies in Irish science and technology.

Royal Dublin Society, Dublin, Ireland, 1983. ISBN 0-86027-014-9. xiv + 242 + 4 pp. LCCN QE265 .H55 1983. US$12.00.

Herken:1988:UTM

[Her88] Rolf Herken, editor. The Universal Turing machine: a halfcentury survey. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. ISBN 0-19-853741-7. xiv + 661 pp. LCCN QA267 .U55

1988. URL http://www.loc.
gov/catdir/enhancements/
fy0636/88213652-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0636/88213652-t.html.

Herman:1990:FSE

[Her90] Robin Herman. Fusion: the search for endless energy. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-38373-0. xi + 267 pp. LCCN QC791.73 .H47

1990. URL http://www.loc.
gov/catdir/description/
cam024/90001445.html;
http://www.loc.gov/
catdir/enhancements/
fy0827/90001445-b.html;
http://www.loc.gov/
catdir/toc/cam021/
90001445.html.

HerriesDavies:1995:NHY

[Her95] G. L. Herries Davies. North from the Hook: 150 years of the Geological Survey of Ireland. Geological Survey of Ireland, Dublin, Ireland, 1995. ISBN 1-899702-00-8. xi + 342 pp. LCCN QE265 .H54 1995.

Hesse:1974:BRG

[Hes74] Mary Hesse. Book review:

General History of Science as Explanation. By Maurice A. Finocchiaro. Detroit: Wayne State University Press, 1973. Pp. 286. $15.95. British Journal for the History of Science, 7(2):180–182, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025206.

Hetherington:1988:SOE

[Het88] Norriss S. Hetherington. Science and objectivity: episodes in the history of astronomy. Iowa State University Press, Ames, IA, USA, 1988. ISBN 0-8138-1159-7. xii + 168 pp. LCCN QB32 .H48 1988.

Hewson:1966:BRS

[Hew66] M. A. Hewson. Book review: Sixteenth and Seventeenth Centuries: The Great Intellectual Revolution. By J. F. West. Pp. viii + 132. Plates and illustrations. London: John Murray, 1965. Student’s Edition 8s. 6d. Library Edition 16s. British Journal for the History of Science, 3 (1):85, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025114.

Heyck:1981:TIL

[Hey81] Thomas William Heyck. The transformation of intellectual life in Victorian England. Croom Helm studies in society and history. Croom Helm, London, UK, 1981. ISBN 0-7099-1206-4. 256 pp. LCCN DA533.

Heyd:1983:BOE

[Hey83] Michael Heyd. Between orthodoxy and the Enlightenment: Jean-Robert Chouet and the introduction of Cartesian science in the Academy of Geneva, volume 96 of Archives internationales d’histoire des idées; International archives of the history of ideas; 96. Martinus Nijhoff, Dordrecht and The Hague, The Netherlands, 1983. ISBN 90-247-2508-9. xii + 308 pp. LCCN B4651.C474 H49 1983.

Heyd:1995:SRC

[Hey95] Michael Heyd. Be sober and reasonable: the critique of enthusiasm in the Seventeenth and early Eighteenth Centuries, volume 63 of Brill’s studies in intellectual history, 0920-8607. E. J. Brill, Leiden, The Netherlands, 1995. ISBN 90-04-10118-7 (hardcover). x + 312 pp. LCCN BR112 .H45 1995.

Herz-Fischler:1987:MHD

[HF87] Roger Herz-Fischler. A mathematical history of division in extreme and mean ratio. Wilfrid Laurier University Press, Waterloo, Ont., Canada, 1987.

ISBN 0-88920-152-8. xvi + 191 pp. LCCN QA481 .H47 1987. US$65.00.

Hasan:1986:ITI

[HH86] Ahmad Yusuf Hasan and Donald Routledge Hill. Islamic technology: an illustrated history. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-26333-6 (Cambridge), 92-3-102294-6 (UNESCO). xiv + 304 pp. LCCN ????

Hassam:1992:ITI

[HH92] Ahmad Yusuf Hassam and Donald Routledge Hill. Islamic technology: an illustrated history. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-42239-6 92-3-102294-6 (UNESCO). xv + 304 pp. LCCN ????

Hunter:1997:WSM

[HH97] Lynette Hunter and Sarah Hutton, editors. Women, science and medicine 1500–1700: mothers and sisters of the Royal Society. Sutton Pub., Thrupp, Stroud, Gloucestershire, 1997. ISBN 0-7509-1334-7 (hardcover), 0-7509-1343-6 (paperback). xx + 292 pp. LCCN Q41.L85 W66 1997.

Higham:1965:BRH

[Hig65] N. Higham. Book review: The History of the Study of Landforms. Vol. I. Geomorphology before Davis. By Richard J. Chorley, Antony J. Dunn and Robert P. Beckinsale. Pp. xvi + 678. London: Methuen, 1964. £4 4s. British Journal for the History of Science, 2(3): 269, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024951.

Hills:1971:BRI

[Hil71a] Richard L. Hills. Book review: Industrial Revolution Histoire Générale des Techniques, Tome III. L’Expansion du Machinisme. Ed. by Maurice Daumas. Paris: Presses Universitaires de France. 1968. Pp. xxiv + 884. Illustr. Price not stated. British Journal for the History of Science, 5(3):298–299, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025326.

Hills:1971:BRT

[Hil71b] Richard L. Hills. Book review: Technology Textile Manufacture in the Northern Roman Provinces. By J. P. Wild. London: Cambridge University Press. 1970. Pp. xxii + 190. £3.25. British Journal for the History of Science, 5 (4):411, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025396.

Hills:1975:BRI

[Hil75] Richard Hills. Book review:

India: Indian Science and Technology in the Eighteenth Century. By Shri Dharampal. Delhi: Impex India, 1971. Pp. lxxii + 281. Rs 65. British Journal for the History of Science, 8(1):78–79, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025824.

Hill:1984:HEC

[Hil84] Donald Routledge Hill. A history of engineering in classical and medieval times. Croom Helm, London, UK, 1984. ISBN 0-7099-1209-9. 263 pp. LCCN TA16 .H55x 1984b.

Hills:1986:BRJb

[Hil86a] Richard L. Hills. Book review: Jean-Michael Chaplain. La Chambre des Tisseurs, Louviers: Cité Drapière 1680–1840. Champ Vallon, 1984. Pp. 302. ISBN 2-903528-40-3. British Journal for the History of Science, 19(2):234–235, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026616.

Hills:1986:BRJa

[Hil86b] Richard L. Hills. Book review: John Shaw. Water Power in Scotland, 1550–1870. Edinburgh: John Donald, 1984. Pp. xi + 606 ISBN 0-85976-072-3. £25. British Journal for the History of Science, 19 (2):204–205, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026594.

Hills:1989:BRG

[Hil89a] Richard Hills. Book review: Graham West. Innovation and the Rise of the Tunnelling Industry. Cambridge: University Press, 1988. Pp. xv + 355. ISBN 0-521-33512-4. £40.00; $80.00. British Journal for the History of Science, 22(1): 95–96, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026694.

Hills:1989:BRD

[Hil89b] Richard L. Hills. Book review: Dieter Kuhn. Science and Civilisation in China, Vol 5. Chemistry and Chemical Technology, Part IX. Textile Technology: Spinning and Reeling. Cambridge: Cambridge University Press, 1988. Pp. xxxiv + 520, ISBN 0-521-32021-6, £60.00, $110.00. British Journal for the History of Science, 22(4):446–448, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026925.

Hills:1994:PWH

[Hil94] Richard Leslie Hills. Power from wind: a history of windmill technology. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-41398-2 (hardcover). 324 pp. LCCN TJ823 .H55 1994. URL http:/

/www.loc.gov/catdir/
description/cam025/
93008858.html;   http://
www.loc.gov/catdir/toc/
cam021/93008858.html.

Hirsh:1983:GIU

[Hir83] Richard F. Hirsh. Glimpsing an invisible universe: the emergence of X-ray astronomy. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25121-4. viii + 186 pp. LCCN QB472 .H57

1983. URL http://www.loc.
gov/catdir/description/
cam031/82023633.html;
http://www.loc.gov/
catdir/enhancements/
fy0906/82023633-t.html.

Hulin-Jung:1989:LES

[HJF89] Nicole Hulin-Jung and Robert Fox. L’organisation de l’enseignement des sciences: la voie ouverte par le Second Empire. (French) [The organization of teaching of science: the way opened by the Second Empire], volume 6 of Mémoire de la section d’histoire des sciences et techniques. Ed. du Comité des travaux historiques et scientifiques, Paris, France, 1989. ISBN 2-7355-0171-X. ISSN 0769-2218. 335 pp. LCCN Q183.4.F8 H85

1989. URL http://catalog.
hathitrust.org/api/
volumes/oclc/21276320.
html.

Hanle:1982:BAA

[HK82]  Paul  A.  Hanle  and

Theodore Von Kármán. Bringing aerodynamics to America. MIT Press, Cambridge, MA, USA, 1982. ISBN 0-262-08114-8. xiv + 184 pp. LCCN TL565 .H32 1982.

Harrower:1983:KKU

[HK83] Molly Harrower and Kurt Koffka. Kurt Koffka, an unwitting self-portrait. University Presses of Florida, Gainesville, FL, USA, 1983. ISBN 0-8130-0760-7. xvi + 334 pp. LCCN BF109.K59 H37 1983.

Herrmann:1984:HAH

[HK84] Dieter B. Herrmann and Kevin Krisciunas. The history of astronomy from Herschel to Hertzsprung. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25733-6. x + 220 pp. LCCN QB28 .H4713 1984.

Hindle:1986:ECA

[HL86] Brooke Hindle and Steven D. Lubar. Engines of change: the American industrial revolution, 1790–1860. Smithsonian Institution Press, Washington, DC, USA, 1986. ISBN 0-87474-540-3, 0-87474-539-X (paperback). 309 pp. LCCN HC105 .H66 1986.

Hoskin:1966:SFB

[HM66] M. A. Hoskin and A. G. Molland. Swineshead on falling bodies: An example of fourteenthcentury physics. British Journal for the History of Science, 3(2):150–182, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025027.

Hoch:1986:BRG

[Hoc86] Paul K. Hoch. Book review:

Gernot Bohme et al. Finalization in Science: the Social Orientation of Scientific Progress. Edited by Wolf Schafer. Dordrecht, Boston and Lancaster: D. Reidel, 1983. British Journal for the History of Science, 19(1):116, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026488.

Hoch:1988:BRD

[Hoc88a]  Paul Hoch.  Book review:

Daniel J. Kevles. In the Name of Eugenics. New York: Alfred Knopf, 1985, Pp. x + 430. IBSN 0-394-50702-9. No price given. British Journal for the History of Science, 21 (2):252–254, June 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026983.

Hoch:1988:BRP

[Hoc88b] Paul K. Hoch. Book review:

Peter J. Kuznick. Beyond the Laboratory: Scientists as Political Activists in 1930s America. Chicago and London: University of Chicago Press, 1987. Pp. x + 363. ISBN 0-226-46583-7. £23.95, $35.95. British Journal for the History of Science, 21 (3):378–380, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026850.

Hodge:1970:BRB

[Hod70] M. J. S. Hodge. Book review: Biology and Medicine Vestiges of the Natural History of Creation. By Robert Chambers; with an Introduction by Gavin de Beer. Leicester: Leicester University Press. 1969. Pp. 38 + vi + 390. 50s. British Journal for the History of Science, 5(1):96–97, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025366.

Hodge:1971:LSL

[Hod71] M. J. S. Hodge. Lamarck’s science of living bodies. British Journal for the History of Science, 5(4):323–352, December 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025378.

Hodge:1977:SSD

[Hod77] M. J. S. Hodge. The structure and strategy of Darwin’s ‘long argument’. British Journal for the History of Science, 10(3):237–246, November 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025843.

Hodge:1989:BRT

[Hod89a] M. J. S. Hodge. Book reviews: Theodore M. Porter, The Rise of Statistical Thinking: 1820–1900. Princeton, New Jersey: Princeton University Press, 1986. Pp. xii + 333. ISBN 0-691-08416-5. £23.40. Stephen M. Stigler, The History of Statistics: the Measurement of Uncertainty before 1900. Cambridge, Massachusetts, and London, England: The Belknap Press of Harvard University Press, 1986. Pp. ix + 410. ISBN 0-674-40340-1. No price given. British Journal for the History of Science, 22(1):111– 114, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026709.

Hodge:1989:GOS

[Hod89b] M. J. S. Hodge. Generation and the origin of species (1837– 1937): A historiographical suggestion. British Journal for the History of Science, 22(3): 267–281, September 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026896.

Hodgkin:1992:CDR

[Hod92] A. L. (Alan Lloyd) Hodgkin. Chance and design: reminiscences of science in peace and war. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40099-6. xi + 412 pp. LCCN QP355.2 .H63

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/91027737.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        91027737.html.

Hofmann:1962:RSC

[Hof62] J. E. Hofmann. Report on the Seventh Colloquium on the History of Mathematics. British Journal for the History of Science, 1(2):184–185, December 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025132.

Hofmann:1987:AIE

[Hof87] James R. Hofmann. Ampère’s invention of equilibrium apparatus: A response to experimental anomaly. British Journal for the History of Science, 20(3):309–341, July 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026360.

Hogben:1974:OTP

[Hog74] Lancelot Hogben. The origins of theoretical population genetics. British Journal for the History of Science, 7(2): 176–179, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025205.

Holloway:1971:BRM

[Hol71] S. W. F. Holloway. Book review: Medicine The History of Medical Education. U.C.L.A. Forum in Medical Sciences No.

12. Ed. by C. D. O’Malley. Berkley, Los Angeles & London: University of California Press. 1970. Pp. xii + 548. £9.50. British Journal for the History of Science, 5(4): 410, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025395.

Holt:1977:WOB

[Hol77] Niles R. Holt. Wilhelm Ostwald’s ‘The Bridge’. British Journal for the History of Science, 10(2):146–150, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025867.

Holmes:1978:SRR

[Hol78] Geoffrey Holmes. Science, reason, and religion in the Age of Newton. British Journal for the History of Science, 11 (2):164–171, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025730.

Holland:1985:PHH

[Hol85a] Alan John Holland, editor. Philosophy, its history and historiography: [conference held at the University of Lancaster in September 1983], volume 1983 of Royal Institute of Philosophy conferences. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1985. ISBN 90-277-1945-4. LCCN B791 .P44 1985.

Holmes:1985:LCL

[Hol85b]  Frederic  Lawrence

Holmes. Lavoisier and the chemistry of life: an exploration of scientific creativity, volume 4 of Wisconsin publications in the history of science and medicine. University of Wisconsin Press, Madison, WI, USA, 1985. ISBN 0-299-09980-6, 0-299-09984-9 (paperback). xxiv + 565 pp. LCCN QP511.8.L38 HOL 1985; 920; QP.

Holton:1986:ASB

[Hol86] Gerald James Holton. The advancement of science, and its burdens: the Jefferson Lecture and other essays. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25244-X, 0-521-27243-2 (paperback). xi + 351 pp. LCCN Q173 .H734

1986. URL http://www.loc.
gov/catdir/description/
cam032/86009722.html.

Holmes:1989:ECC

[Hol89] Frederic Lawrence Holmes, editor. Eighteenth-Century chemistry as an investigative enterprise, volume 12; 6; 1 of Berkeley papers in history of science; Uppsala studies in history of science; Bologna studies in history of science. Office for History of Science and Technology, University of California at Berkeley, Berkeley, CA, USA, 1989. ISBN 0-918102-16-2. LCCN QD18.E85 H65 1989. US$16.00.

Home:1968:TLN

[Hom68]  Roderick W. Home.  The

Third Law in Newton’s mechanics. British Journal for the History of Science, 4(1): 39–51, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024985.

Home:1972:FEA

[Hom72] Roderick W. Home. Franklin’s electrical atmospheres. British Journal for the History of Science, 6(2):131–151, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025288.

Home:1974:BRE

[Hom74] Roderick W. Home. Book review: Eighteenth Century: The Unpublished Writings of Tobias Mayer. Volume I: Astronomy and Geography; Volume 2: Artillery and Mechanics; Volume 3: The Theory of the Magnet and its Application to Terrestrial Magnetism. (Arbeiten aus der Niedersächsischen Staatsund Universitätsbibliothek Göttingen, Bd. 9-11.) Ed. by Eric G. Forbes. Göttingen: Vandenhoeck and Ruprecht, 1972. Pp. viii + 227; viii + 136; vi + 104. DM 18, DM 27, and DM 35. British Journal for the History of Science, 7(3):296– 298, November 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025545.

Home:1980:BRS

[Hom80]  R. W. Home.  Book re-

view: Seventeenth and Eighteenth Centuries: Electricity from Glass: The History of the Frictional Electrical Machine, 1600–1850. By W. D. Hackmann. Alphen aan den Rijn: Sijthoff & Noordhoff, 1978. Pp. xiv & 310. Df189/ $44.50. British Journal for the History of Science, 13(3): 268–270, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026206.

Home:1983:PME

[Hom83a] R. W. Home. Poisson’s memoirs on electricity: Academic politics and a new style in physics. British Journal for the History of Science, 16(3): 239–259, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026145.

Home:1983:SUS

[Hom83b] Roderick Weir Home, editor.

Science under scrutiny: the place of history and philosophy of science, volume 3 of Australasian studies in history and philosophy of science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1983. ISBN 90-277-1602-1. xvii + 182 pp. LCCN Q175.3 .S323 1983.

Home:1986:BRI

[Hom86a] R. W. Home. Book review:

Isaac Benguigui (ed.). Théories Électriques du XVIIIe Siècle: Correspondance entre l’Abbé Nollet (1700–1770) et le Physicien Genevois Jean Jallabert (1712– 1768). Geneva: Georg et Cie, 1984. Pp. 234. ISBN 2-8257-0099-1. Fr. 68.00. British Journal for the History of Science, 19(3):344, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026514.

Home:1986:BRJ

[Hom86b] R. W. Home. Book review:

J. L. Heilbron. Physics at the Royal Society During Newton’s Presidencey. Los Angeles: William Andrews Clark Memorial Library, University of California, 1983. Pp. xiv + 123. British Journal for the History of Science, 19(1): 124–126, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026497.

Home:1988:ASM

[Hom88] Roderick Weir Home, editor. Australian science in the making. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-35556-7. LCCN Q127.A8 A86 1988. URL http:

                                        //www.loc.gov/catdir/
                                        description/cam023/
                                        88011866.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam028/88011866.html.

Hon:1989:TCE

[Hon89] Giora Hon. Is there a concept of experimental error in Greek astronomy? British Journal for the History of Science, 22 (2):129–150, July 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026658.

Hooykaas:1987:RMS

[Hoo87] R. Hooykaas. The rise of modern science: When and why? British Journal for the History of Science, 20(4): 453–473, October 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026418.

Hooykaas:1997:RBS

[Hoo97] R. (Reijer) Hooykaas. Robert Boyle: a study in science and Christian belief. Pascal Centre for Advanced Studies in Faith and Science, Redeemer College, Ancaster, Ont., 1997. ISBN 0-7618-0708-X (hardcover). xxiv + 131 pp. LCCN BT1117 .H66 1997.

Hoppen:1972:BRS

[Hop72] K. Theodore Hoppen. Book review: Sixteenth and Seventeenth Centuries: The Experimenters. A Study of the Accademia del Cimento. By W. E. Knowles Middleton. Baltimore and London: The Johns Hopkins Press, 1971. Pp. xvi + 415. £10.70. British Journal for the History of Science, 6(2):214–216, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025303.

Hoppen:1973:BRS

[Hop73] K. Theodore Hoppen. Book review: Sixteenth and Seventeenth Centuries: The Correspondence of Henry Oldenburg. Volume VII, 1670–1671; Volume VIII, 1671–1672. Ed. by A. Rupert and Marie Boas Hall. Madison, Milwaukee, and London: University of Wisconsin Press, 1970 and 1971. Pp. xxviii + 600 and xxvii + 663. $17.50 and $20.00. British Journal for the History of Science, 6 (3):325–326, June 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025459.

Hoppen:1975:BRS

[Hop75] K. Theodore Hoppen. Book review: Seventeenth Century: The Correspondence of Henry Oldenburg. Volume IX, 1672– 1673. Ed. by A. Rupert Hall and Marie Boas Hall. Madison and London: University of Wisconsin Press, 1973. Pp. xxx + 706. $20. British Journal for the History of Science, 8 (1):84–85, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025829.

Hoppen:1976:NERa

[Hop76a] K. Theodore Hoppen. The nature of the early Royal Society: Part I. British Journal for the History of Science, 9 (1):1–24, March 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025703.

Hoppen:1976:NERb

[Hop76b] K. Theodore Hoppen. The nature of the early Royal Society Part II. British Journal for the History of Science, 9(3):

243–273, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025682.

Hoskin:1967:BRA

[Hos67] M. A. Hoskin. Book review:

Archimedes in the Middle Ages. Volume 1: The Arabo-Latin Tradition. By Marshall Clagett. Madison: University of Wisconsin Press. Pp. xxxii + 720. 1964. $12.00. British Journal for the History of Science, 3 (3):296, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025056.

Howson:1982:HME

[How82] A. G. (Albert Geoffrey) Howson. A history of mathematics education in England. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24206-1. x + 294 pp. LCCN QA14.G7 H68. URL http://

www.loc.gov/catdir/
enhancements/fy0838/
82004175-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0838/82004175-t.html.

Howse:1989:NMS

[How89] Derek Howse. Nevil Maskelyne, the seaman’s astronomer. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-36261-X. xiv + 280 pp. LCCN QB36.M378 H69 1989.

Hoyrup:1994:MNW

[Høy94] Jens Høyrup. In measure, number, and weight: studies in mathematics and culture. SUNY series in science, technology, and society. State University of New York Press, Albany, NY, USA, 1994. ISBN 0-7914-1821-9, 0-7914-1822-7 (paperback : acid-free). xviii + 430 pp. LCCN QA10.7 .H69 1994.

Hermann:1985:WPW

[HP85] Armin Hermann and Wolfgang Pauli, editors. Wolfgang Pauli, wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Bd. 2, 1930–1939. (German) [Wolfgang Pauli, scientific correspondence with Bohr, Einstein, Heisenberg and others, volume 2, 1930–1939], volume Vol. 6 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-13609-6 (Berlin), 0-387-13609-6 (New York). xxxix + 783 pp. LCCN ????

Hentschel:1990:KPR

[HRP90] Klaus Hentschel, Hans Reichenbach, and Joseph Petzoldt. Die Korrespondenz Petzoldt–Reichenbach: zur Entwicklung der “wissenschaftlichen Philosophie” in Berlin. (German) [The Petzoldt–Reichenbach correspondence: development “of scientific philosophy” in Berlin], volume 12 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. SIGMA, Berlin, Germany, 1990. ISBN 3-928068-03-2. 91 pp. LCCN Q175.3 .H46 1990.

Hofmann:1963:REC

[HS63] J. E. Hofmann and G. J. Scriba.

Report on the Eighth Colloquium on the History of Mathematics. British Journal for the History of Science, 1(4): 382–384, December 1963. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024910.

Hollister-Short:1983:BRT

[HS83] Graham Hollister-Short. Book review: Technology and Instruments Simon Lavington, Early British computers: the story of vintage computers and the people who built them. Manchester: Manchester University Press, 1980. Pp. iv + 139. £3.95. British Journal for the History of Science, 16(2): 208–209, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026249.

Hollister-Short:1985:LWS

[HS85] Graham Hollister-Short. The long waves of social consciousness. British Journal for the History of Science, 18(1):77– 82, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026267.

Hollister-Short:1986:BRD

[HS86a] Graham Hollister-Short. Book review: Donald Hill. A History of Engineering in Classical and Medieval Times. Beckenham: Croom Helm, 1984. Pp. x + 263. ISBN 0-7099-1209-9. £18.95. British Journal for the History of Science, 19(3): 366–367, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026538.

Hollister-Short:1986:BRS

[HS86b] Graham Hollister-Short. Book review: Svante Lindqvist. Technology on Trial. The Introduction of Steam Power Technology into Sweden, 1715–1736. Uppsala: Almqvist & Wiksell, 1984. Pp. 392. ISBN 91-22-00716-4. SEK 248. British Journal for the History of Science, 19(2):229–231, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026613.

Holuster-Short:1988:BRA

[HS88] Graham Holuster-Short. Book review: Ahmad Y. Al-Hassan & Donald R. Hill. Islamic Technology, An Illustrated History. Cambridge and Paris: Cambridge University Press and UNESCO, 1986. Pp. xiv + 304. ISBN 92-3-102294-6. £25.00, $39.50. British Journal for the History of Science, 21(1): 119–120, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026875.

Hunter:1989:RHN

[HS89] Michael Cyril William Hunter and Simon Schaffer, editors. Robert Hooke: new studies; grown out of a conference. Boydell Press, Woodbridge, Suffolk, UK, 1989. ISBN 0-85115-523-5. LCCN Q143.H7 H86 1989.

Hankins:1995:II

[HS95]  Thomas  L.

Hankins and Robert J. Silverman. Instruments and the imagination. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-02997-0. xiv + 337 pp. LCCN Q185 .H25

1995. URL http://www.loc.
gov/catdir/description/
prin021/95004301.html;
http://www.loc.gov/
catdir/toc/prin031/
95004301.html.

Hall:1993:HT

[HSHSJ93] Alfred Rupert Hall, Norman A. F. Smith, Graham Hollister-Short, and Frank A. J. L. James, editors. History of technology. Mansell, London, UK, 1976–1993. ISBN 0-7201-1770-4 (vol. 10), 0-7201-1861-1 (vol.

11), 0-7201-2075-6 (vol. 12), 0-7201-2114-0 (vol. 13), 0-7201-2133-7 (vol. 14), 0-7201-2160-4 (vol. 15). ???? pp. LCCN ????

Hubner:1990:THT

[Hüb90] Jürgen Hübner, editor. Theologica, Hexenprozess, Tacitus-Übersetzung, Gedichte. (German) [Theologica, witch trials, Tacitus translation, Poems], volume 12 of Gesammelte Werke. Beck, München, Germany, 1990. ISBN 3-406-01660-X. 443 pp. LCCN ????

HuertasGarcia-Alejo:1988:OSP

[Hue88] Rafael Huertas García-Alejo. Orfila, saber y poder medico. (Spanish) [Orfila, knowledge and medical power], volume 2 of Estudios sobre la ciencia. Consejo Superior de Investigaciones Cientificas, Madrid, Spain, 1988. ISBN 84-00-06745-2. 456 pp. LCCN MLCS 93/04632 (R).

Hufbauer:1982:FGC

[Huf82] Karl Hufbauer. The formation of the German chemical community (1720–1795). University of California Press, Berkeley, CA, USA, 1982. ISBN 0-520-04318-9, 0-520-04415-0. viii + 312 pp. LCCN QD18.G3 H83 1982.

Hufbauer:1985:BRC

[Huf85]  Karl Hufbauer.  Book re-

view: S. Chandrasekhar, Eddington: The most distinguished astrophysicist of his time. Cambridge: Cambridge University Press, 1983. Pp.

64. ISBN 0-521-25746-8. £7.50, $12.50. British Journal for the History of Science, 18(3): 356, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026396.

Hufbauer:1986:BRD

[Huf86] Karl Hufbauer. Book review:

Donald E. Osterbrock. James E. Keeler: Pioneer American Astrophysicist and the Early Development of American Astrophysics. Cambridge, Eng.: Cambridge University Press, 1985. Pp. xii + 411. ISBN 0-521-26582-7. £35.00. British Journal for the History of Science, 19(3):350– 351, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026522.

Hufbauer:1991:ESS

[Huf91]  Karl Hufbauer.  Exploring

the sun: solar science since Galileo. New series in NASA history. The Johns Hopkins University Press, Baltimore, MD, USA, 1991. ISBN 0-8018-4098-8. xv + 370 pp. LCCN QB521 .H93 1991.

Huff:1993:REM

[Huf93] Toby E. Huff. The rise of early modern science: Islam, China, and the West. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-43496-3 (hardcover). xiv + 409 pp. LCCN Q127.A5 H84 1993.

Huff:1995:REM

[Huf95] Toby E. Huff. The rise of early modern science: Islam, China, and the West. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-49833-3 (paperback), 0-521-43496-3 (hardcover). xiv + 409 pp. LCCN Q127.A5 H84 1995. URL http:

                                        //www.loc.gov/catdir/
                                        description/cam027/
                                        95237814.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam025/95237814.html.

Hughes:1989:ARW

[Hug89] R. Elwyn Hughes. Alfred Russel Wallace; some notes on the Welsh connection. British Journal for the History of Science, 22(4):401–418, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026917.

Huizenga:1992:CFS

[Hui92] John R. (John Robert) Huizenga. Cold fusion: the scientific fiasco of the century. University of Rochester Press, Rochester, NY, USA, 1992. ISBN 1-878822-07-1. xv + 259 pp. LCCN QC791.775.C64 H85 1992. US$45.00.

Hull:1967:ME

[Hul67] David L. Hull. The metaphysics of evolution. British Journal for the History of Science, 3(4): 309–337, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024958.

Hull:1988:SPE

[Hul88] David L. Hull. Science as a process: an evolutionary account of the social and conceptual development of science. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1988. ISBN 0-226-36050-4. xiii + 586 pp. LCCN Q158.5 .H84 1988. US$39.95.

Humphreys:1967:GFB

[Hum67] W. C. Humphreys. Galileo, falling bodies and inclined planes: An attempt at reconstructing Galileo’s discovery of the Law of Squares. British Journal for the History of Science, 3(3):225–244, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025049.

Hunter:1982:RSF

[Hun82] Michael Cyril William Hunter.

The Royal Society and its Fellows, 1660–1700: the morphology of an early scientific institution, volume 4 of BSHS monographs. British Society for the History of Science, Chalfont St. Giles, Bucks, England, 1982. ISBN 0-906450-03-9. 270 pp. LCCN Q41.L85 H95 1982.

Hunter:1984:BRS

[Hun84] Michael Hunter. Book review: Seventeenth Century: M. M. Slaughter, Universal languages and scientific taxonomy in the seventeenth century. Cambridge: Cambridge University Press, 1982. Pp. x + 277. ISBN 0-521-24477-3. £25.00. British Journal for the History of Science, 17(1): 109–110, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026046.

Hunter:1979:HIP

[Hun85]  Louis C. Hunter.  A his-

tory of industrial power in the United States, 1780–1930. Published for the Eleutherian Mills-Hagley Foundation by the University Press of Virginia, Charlottesville, VA, USA, 1979– 1985. ISBN 0-8139-0782-9 (vol. 1), 0-262-08198-9 (vol. 3). ???? pp. LCCN HD1694.A5 H76.

Hunter:1986:BRL

[Hun86] Michael Hunter. Book review: Lesley Murdin. Under Newton’s Shadow: Astronomical Practices in the Seventeenth Century. Bristol and Boston: Adam Hilger, 1985. Pp. viii + 152. ISBN 0-85274-456-0. £16.75. British Journal for the History of Science, 19(3): 349–350, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026521.

Hunt:1988:OFC

[Hun88] Bruce J. Hunt. The origins of the FitzGerald Contraction. British Journal for the History of Science, 21(1): 67–76, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026862.

Hunt:1989:BRJ

[Hun89a] Bruce J. Hunt. Book review: J.

G. O’Hara & Willibald Pricha. Hertz and the Maxwellians: A Study and Documentation of the Discovery of Electromagnetic Wave Radiation, 1873– 1894. London: Peter Peregrinus Ltd., 1987. Pp. xiv + 154. ISBN 0-86341-101-0. £24.00. British Journal for the History of Science, 22(3): 392–393, September 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026909.

Hunter:1989:ENS

[Hun89b] Michael Cyril William Hunter.

Establishing the new science: the experience of the early Royal Society. Boydell Press, Woodbridge, Suffolk, UK, 1989. ISBN 0-85115-506-5. xiv + 382 pp. LCCN Q41.L85 H93 1989.

Hunt:1991:M

[Hun91]  Bruce J. Hunt.  The

Maxwellians. Cornell history of science series. Cornell University Press, Ithaca, NY, USA, 1991. ISBN 0-8014-2641-3. x + 266 pp. LCCN QC670 .H84 1991.

Hunter:1995:SSO

[Hun95] Michael Cyril William Hunter. Science and the shape of orthodoxy: intellectual change in late Seventeenth-Century Britain. Boydell Press, Woodbridge, Suffolk, UK, 1995. ISBN 0-85115-594-4. xii + 345 pp. LCCN Q127.G4 H853 1995.

Hutchison:1980:BRN

[Hut80] Keith Hutchison. Book review: Nineteenth Century: Sadi Carnot et l’essor de la thermodynamique. Paris: Editions du Centre National de la Recherche Scientifique, 1976. Pp. 435. 128 francs. British Journal for the History of Science, 13(2):168, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025901.

Hutchison:1981:WJM

[Hut81] Keith Hutchison. W. J. M. Rankine and the rise of thermodynamics. British Journal for the History of Science, 14 (1):1–26, March 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026070.

Hutton:1984:BRA

[Hut84] Sarah Hutton. Book review:

Antiquity to the Renaissance C. B. Schmitt, Aristotle and the Renaissance. Cambridge (Mass.) and London: Harvard University Press, 1983. Pp. viii + 187. ISBN 0-674-04525-4. £14.80. British Journal for the History of Science, 17(1): 104–105, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026040.

Hutton:1988:BRA

[Hut88] Sarah Hutton. Book review:

Anthony Grafton & Lisa Jardine. From Humanism to the Humanities: Education and the Liberal Arts in Fifteenth and Sixteenth-Century Europe. London: Duckworth, 1986. Pp. xvi + 224. ISBN 0-7156-2100-9. £29.95. British Journal for the History of Science, 21 (1):117, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026872.

Hutton:1989:BRC

[Hut89]  Sarah Hutton.  Book re-

view: Charles B. Schmitt (general editor), Quentin Skinner, Eckhard Kessler (editors) and Jill Kraye (associate editor). The Cambridge History of Renaissance Philosophy. Cambridge: Cambridge University Press, 1988. Pp. xii + 968. ISBN 0-521-25104-4. £50.00. British Journal for the History of Science, 22(3):377–382, September 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026902.

Heilbron:1981:LHP

[HW81] J. L. Heilbron and Bruce R. Wheaton. Literature on the history of physics in the 20th Century, volume 5 of Berkeley papers in history of science. Office for History of Science and Technology, University of California at Berkeley, Berkeley, CA, USA, 1981. ISBN 0-918102-05-7 (paperback). ISSN 0145-0379. ix + 485 pp. LCCN Z7144.H55 H44; QC7 .H4 1981.

Harley:1987:CPA

[HW87] J. B. (John Brian) Harley and David Woodward, editors. Cartography in prehistoric, ancient, and medieval Europe and the Mediterranean, volume 1 of The history of cartography. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-31633-5. xxi + 599 pp. LCCN GA201 .H53 1987 vol.

1. URL http://www.loc.gov/
catdir/description/
uchi052/98137538.html;
http://www.loc.gov/

catdir/toc/uchi052/
98137538.html.

Harley:1987:HC

[HW03] J. B. (John Brian) Harley and David Woodward, editors. The history of cartography. University of Chicago Press, Chicago, IL, USA, 1987–2003. ISBN 0-226-31633-5 (vol. 1), 0-226-31635-1 (vol. 2, bk. 1), 0-226-31637-8 (vol. 2, bk. 2). ???? pp. LCCN GA201 .H53 1987.

Horder:1986:HEE

[HWW86] Timothy John Horder, Jan Anthony Witkowski, and Christopher Craig Wylie, editors. A history of embryology: the eighth Symposium of the British Society for Developmental Biology, [April 1983 in Nottingham]. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25953-3. LCCN QL953 .B75

1983. URL http://www.loc.
gov/catdir/description/
cam032/85016583.html.

Ingrao:1990:IHE

[II90] Bruna Ingrao and Giorgio Israel. The invisible hand: economic equilibrium in the history of science. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-09028-7. xiii + 491 pp. LCCN HB145 .I5413 1990.

Ilie:1995:AM

[Ili95] Paul Ilie. The age of Minerva. University of Pennsylvania Press, Philadelphia, PA, USA, 1995. ISBN 0-8122-3307-7 (vol. 1), 0-8122-3308-5 (vol. 2). ???? pp. LCCN B802 .I36 1995.

Iltis:1973:LND

[Ilt73] Carolyn Iltis. The Leibnizian– Newtonian debates: Natural philosophy and social psychology. British Journal for the History of Science, 6(4):343– 377, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025501.

Iltis:1977:BRS

[Ilt77] Carolyn Iltis. Book review: Seventeenth Century: Leibniz and Dynamics: The Texts of 1692. By Pierre Costabel. Trans. by R. E. W. Maddison. London: Methuen, 1973. Pp. 141. £3.00. British Journal for the History of Science, 10(2):176–177, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025881.

Inglis:1984:SPH

[Ing84] Brian Inglis. Science and parascience: a history of the paranormal, 1914–1939. Hodder and Stoughton, London, UK, 1984. ISBN 0-340-26325-3. 382 pp. LCCN BF1028 .I54 1984. US$12.95.

Inkster:1979:LSS

[Ink79] Ian Inkster. London science and the Seditious Meetings Act of 1817. British Journal for the History of Science, 12(2): 192–196, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025752.

Inkster:1981:SSR

[Ink81] Ian Inkster. Seditious science: A reply to Paul Weindling. British Journal for the History of Science, 14(2): 181–187, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025946. See [Wei80].

Inkster:1987:BRM

[Ink87] Ian Inkster. Book review: Maxine Berg. The Age of Manufactures. Industry, Innovation and Work in Britain, 1700– 1820. London: Fontana Paperback, 1985. Pp. 378. ISBN 0-00-686019-2. £4.95. British Journal for the History of Science, 20(1):98–99, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026452.

Inkster:1988:BRN

[Ink88] Ian Inkster. Book review: Nathan Reingold & Marc Rothenberg (eds). Scientific Colonialism: A Cross-Cultural Comparison. Washington: Smithsonian Institution Press, 1987. Pp. xiii + 398. ISBN 0-87474-785-6.

British Journal for the History of Science, 21(3):382–384, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026853.

Inster:1986:BRC

[Ins86] Ian Inster. Book review: Colin Russell. Science and Social Change 1700–1900. London: The Macmillan Press, 1983. Pp. xvii + 307. ISBN 0-333-29272-3. £15. British Journal for the History of Science, 19 (1):122–123, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026495.

Irwin:1982:BRS

[Irw82] Alan Irwin. Book review: Science Studies Bruno Latour and Steve Woolgar, Laboratory life: the social construction of scientific facts. Beverly Hills and London: Sage Publications, 1979. Pp. 271. £11.25; £5.50 (paperback). British Journal for the History of Science, 15 (2):208–209, July 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025977.

Irwin:1983:BRM

[Irw83a] Alan Irwin. Book review: Medical Sciences: Kenneth E. Studer and Daryl E. Chubin, The Cancer Mission: social contexts of biomedical research. Beverly Hills & London: Sage, 1980. Pp. 320. No price stated. British Journal for the History of Science, 16 (1):93–95, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026103.

Irwin:1983:BRS

[Irw83b] Alan Irwin. Book review: Science and Society Thomas R. Dunlap, DDT. Scientists, citizens and public policy. Princeton: Princeton University Press, 1981. Pp.

318. £10.80. British Journal for the History of Science, 16 (3):300, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026169.

Ito:1988:HCT

[Ito88] Yushi Ito. Hooke’s cyclic theory of the Earth in the context of Seventeenth Century England. British Journal for the History of Science, 21(3):295– 314, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026831.

Irwin:1996:MSP

[IW96] Alan Irwin and Brian Wynne, editors. Misunderstanding science?: the public reconstruction of science and technology.

Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-43268-5 (hardcover). vi + 232 pp. LCCN Q175.5 .M58

1996. URL http://www.loc.
gov/catdir/description/
cam027/95032980.html;
http://www.loc.gov/
catdir/samples/cam034/
95032980.html;   http://
www.loc.gov/catdir/toc/
cam022/95032980.html.

Jacyna:1981:PMU

[Jac81] L. S. Jacyna. The physiology of mind, the unity of nature, and the moral order in Victorian thought. British Journal for the History of Science, 14 (2):109–132, July 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025943.

Jacob:1983:HSR

[Jac83] James R. (James Randall) Jacob. Henry Stubbe, radical Protestantism and the early Enlightenment. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24876-0. viii + 222 pp. LCCN CT788.S92 J3

1983. URL http://www.loc.
gov/catdir/description/
cam022/82012919.html;
http://www.loc.gov/
catdir/samples/cam031/
82012919.html;   http://
www.loc.gov/catdir/toc/
cam021/82012919.html.

Jacyna:1985:BRM

[Jac85] L. S. Jacyna. Book review:

Max Neuburger, The Historical Development of Experimental Brain and Spinal Cord Physiology Before Flourens (Translated and edited by Edwin Clarke). Baltimore and London: Johns Hopkins University Press, 1982. Pp. xxii + 391. ISBN 0-8018-2380-3. £14. British Journal for the History of Science, 18(1):119– 120, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026298.

Jack:1989:BRW

[Jac89a] D. B. Jack. Book review:

Wolfcang-Hagen Hein. Alexander von Humboldt und die Pharmazie. Stuttgart: Wissenschaftliche Verlagsgessellschaft, 1988. Pp.

130. ISBN 3-8047-0984-2. DM 29 (paperback). British Journal for the History of Science, 22(4):474–475, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026945.

Jackson:1989:BRD

[Jac89b] Myles W. Jackson. Book review: Dennis L. Sepper. Goethe Contra Newton: Polemics and the Project for a New Science of Color. Cambridge (U.K.), New York, New Rochelle, Melbourne, Sidney: Cambridge University Press, 1988. Pp. 222. ISBN 0-521-34254-6. £27.50, $39.50. British Journal for the History of Science, 22(1): 121–123, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026717.

Jaki:1984:UGL

[Jak84] Stanley L. Jaki. Uneasy genius: the life and work of Pierre Duhem, volume 100 of Archives internationales d’histoire des idées = International archives of the history of ideas. Martinus Nijhoff, Dordrecht and The Hague, The Netherlands, 1984. ISBN 90-247-2897-5. xii + 472 pp. LCCN QC16.D925 J34

                                        1984. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1005/83019458-b.html.

James:1982:TSS

[Jam82] Frank A. J. L. James. Thermodynamics and sources of solar heat, 1846–1862. British Journal for the History of Science, 15(2):155–181, July 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025965.

James:1984:PIW

[Jam84] Frank A. J. L. James. The physical interpretation of the wave theory of light. British Journal for the History of Science, 17(1):47–60, March 1984.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026016.

James:1986:BRT

[Jam86] Frank A. J. L. James. Book review: A. Thackray, J. L. Sturchio, P. T. Carroll & R. Bud. Chemistry in America, 1876–1976: Historical Indicators. Dordrecht: D. Reidel, 1985. Pp. xxiii + 564. ISBN 90-277-1720-6. Dfl 210.00, $79.50, £53.50. British Journal for the History of Science, 19(2): 235, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026617.

James:1987:BRD

[Jam87a] Frank A. J. L. James. Book review: A. D. Morrison-Low & J. R. R. Christie (eds), ‘Martyr of Science’: Sir David Brewster, 1781–1868. Edinburgh: Royal Scottish Museum, 1984. Pp. 138. ISBN 0-900733-29-2. £6.00. British Journal for the History of Science, 20(1): 93–94, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026448.

James:1987:BRP

[Jam87b] Frank A. J. L. James. Book review: P. M. Harman (ed). Wranglers and Physicists. Studies on Cambridge [Mathematical] Physics in the Nineteenth Century. Manchester: Manchester University Press, 1985. Pp. viii + 261. ISBN 0-7190-1756-4. £27.50. British Journal for the History of Science, 20(2): 233–234, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026319.

James:1988:PPN

[Jam88] Frank A. J. L. James. The practical problems of ‘new’ experimental science: Spectrochemistry and the search for hitherto unknown chemical elements in Britain 1860–1869. British Journal for the History of Science, 21(2):181– 194, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026978.

James:1989:DLE

[Jam89] Frank A. J. L. James, editor. The Development of the laboratory: essays on the place of experiment in industrial civilization. Macmillan Press Scientific and Medical, Basingstoke, Hampshire, UK, 1989. ISBN 0-333-48331-6. LCCN Q183.A1 D48 1989. URL http://

catalog.hathitrust.org/
api/volumes/oclc/
20993561.html.

Jardine:1983:BRE

[Jar83] N. Jardine. Book review: Essay Collections Stephen Gaukroger (ed), Descartes: philosophy, mathematics and physics. Brighton, Sussex: The Harvester Press. Totowa, New Jersey: Barnes and Noble Books, 1980. Pp. xi + 329. £28.00. British Journal for the History of Science, 16(2):192–195, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026239.

Jardine:1984:BRSb

[Jar84a] N. Jardine. Book review:

Science in Culture: René Descartes, Principles of Philosophy. Translated, with explanatory notes, by V. R. Miller and R. P. Miller. Synthese Historical Library, Vol. 24. Dordrecht: D. Reidel, 1983. Pp xxviii + 325. Cloth Dfl. 135/ $59.00. British Journal for the History of Science, 17(2): 245, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026578.

Jardine:1984:BRSa

[Jar84b] N. Jardine. Book review: Seventeenth Century: Johannes Kepler, Mysterium Cosmographicum: The secret of the universe. Translation by A. M. Duncan, Introduction and Commentary by E. J. Aiton, with a Preface by I. Bernard Cohen. New York: Abaris Books, 1981.

Pp. 267. ISBN 0-913870-64-1. £13. British Journal for the History of Science, 17(1): 108–109, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026045.

Jeffreys:1964:LMS

[Jef64] Alan Jeffreys. Locating the manuscript sources of science. British Journal for the History of Science, 2(2):157–161, December 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025014.

Jensen:1970:XCF

[Jen70] J. Vernon Jensen. The X Club: Fraternity of Victorian scientists. British Journal for the History of Science, 5(1): 63–72, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025353.

Jenkins:1987:WIA

[Jen87] Reese V. Jenkins. Words, images, artifacts and sound: Documents for the history of technology. British Journal for the History of Science, 20(1): 39–56, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026435.

Jensen:1988:RWH

[Jen88] J. Vernon Jensen. Return to the Wilberforce–Huxley debate. British Journal for the History of Science, 21(2):161– 179, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026977.

Jensen:1989:BRJ

[Jen89] Garsten Jensen. Book review:

Jules Six. La découverte du neutron (1920–1936). Paris: Editions du CNRS, 1988. Pp. 175. ISBN 2-222-04052-3. FF110.00 (paperback). British Journal for the History of Science, 22 (3):390–391, September 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026907.

Jensen:1991:THH

[Jen91] J. Vernon (John Vernon)

Jensen. Thomas Henry Huxley: communicating for science. University of Delaware Press, Newark, DE, USA, 1991. ISBN 0-87413-379-3. 253 pp. LCCN QH31.H9 J46 1991.

Jevons:1964:PTWa

[Jev64a] F. R. Jevons. Paracelsus’s two-way astrology I. What Paracelsus meant by ‘stars’. British Journal for the History of Science, 2(2):139–147, December 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025012.

Jevons:1964:PTWb

[Jev64b] F. R. Jevons. Paracelsus’s twoway astrology II. Man’s relation to the stars. British Journal for the History of Science, 2(2):148–155, December 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025013.

Jevons:1967:BRH

[Jev67] F. R. Jevons. Book review: The History of Cell Respiration and Cytochrome. By David Keilin. London: Cambridge University Press. Pp. xix + 416. 1966. 90s. British Journal for the History of Science, 3 (3):302–303, June 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025064.

Jardine:1984:BHP

[JK84] Nicholas Jardine and Johannes Kepler. The birth of history and philosophy of science: Kepler’s A defence of Tycho against Ursus (Apologia pro Tychone contra Ursum), with essays on its provenance and significance. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25226-1. ix + 301 pp. LCCN Q125.2 .J37

1984. URL http://www.loc.
gov/catdir/enhancements/

fy0907/83014186-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0907/83014186-t.html.

Jahnke:1981:ESP

[JO81] H. N. (Hans Niels) Jahnke and Michael Otte, editors. Epistemological and social problems of the sciences in the early nineteenth century: [papers from a workshop held at the Univ. of Bielefeld, Nov. 27–30, 1979]. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1981. ISBN 90-277-1223-9. LCCN Q127.E82 E64.

Johnston:1978:BRS

[Joh78] Ron Johnston. Book review:

Sociology of Science Knowledge and Social Imagery. By David Bloor. London: Routledge and Kegan Paul, 1976. Pp. xii + 156 £3.25. British Journal for the History of Science, 11 (1):65–66, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025608.

Johnson:1979:BRN

[Joh79] Dale M. Johnson. Book review:

Nineteenth and Twentieth Centuries: Dear Russell-Dear Jourdain: a Commentary on Russell’s Logic, Based on his Correspondence with Philip Jourdain. By I. Grattan-Guinness. London: Duckworth, 1977. Pp.

234. £14.00. British Journal for the History of Science, 12 (2):232–233, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025766.

Johnson:1981:BRT

[Joh81] Dale M. Johnson. Book review: Twentieth Century Georg Cantor: His Mathematics and Philosophy of the Infinite. By Joseph Warren Dauben. Cambridge, Mass. & London: Harvard University Press, 1979. Pp. ix + 404. $27.50/£19.00. British Journal for the History of Science, 14(1):101– 103, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026086.

Johnson:1983:BRM

[Joh83] Dale M. Johnson. Book review: Mathematical Sciences: George Temple, 100 years of mathematics. London: Gerald Duckworth & Co. Ltd., 1981. Pp. xvi + 316. £32.00. British Journal for the History of Science, 16(3):293–294, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026164.

Johnson:1984:HLT

[Joh84] F. Neil (Frederick Neil) Johnson. The history of lithium therapy. MacMillan Publishing Company, New York, NY, USA, 1984. ISBN 0-333-36942-4. xvii + 198 + 15 pp. LCCN RC483.5.L5 J635 1984.

Jones:1975:THH

[Jon75a] H. W. Jones. Thomas Hobbes, de homine. British Journal for the History of Science, 8 (2):166–170, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025640.

Jones:1975:DDA

[Jon75b] R. V. Jones. Delusions in defence: Atomic politics 1945– 1952. British Journal for the History of Science, 8(3):246– 249, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025560.

Jones:1986:BCA

[Jon86] Alexander Jones. Book 7 of the Collection of Alexandria Pappus, volume 8 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-96257-3. x + 748 pp. LCCN QA22 .P3713 1986. US$70.00 (set : est.). Two volumes.

Jones:1987:BRS

[Jon87] Harold W. Jones. Book review:

Steven Shapin & Simon Schaffer. ‘Leviathan’ and the Air Pump: Hobbes, Boyle and the Experimental Life. Including a Translation of Thomas Hobbes, ‘Dialogus de Natura Aeris’ by Simon Schaffer. Princeton University Press, New Jersey and Guildford, Surrey, 1985. Pp. ix + 441. ISBN 0-691-08393-2. £43.00. British Journal for the History of Science, 20(1): 122–123, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026474.

Jones:1988:WTT

[Jon88] Whitney R. D. (Whitney Richard David) Jones. William Turner: Tudor naturalist, physician, and divine. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1988. ISBN 0-415-00359-8. 223 pp. LCCN QH31.T88 J66 1988.

Jordanova:1978:BRS

[Jor78] L. J. Jordanova. Book review: Seventeenth and Eighteenth Centuries: Medical Revolution in France, 1789–1796. By David M. Vess. Gainesville, Florida: University of Florida Press, 1975. Pp. 216. $12.00. British Journal for the History of Science, 11(1):81–82, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025622.

Jordanova:1979:BRN

[Jor79] L. J. Jordanova. Book review: Nineteenth and Twentieth Centuries: The Spirit of System: Lamarck and Evolutionary Biology. By Richard W. Burkhardt. Cambridge, Mass. & London: Harvard University Press, 1977. Pp. 285. £11.95. British Journal for the History of Science, 12(2): 228–229, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025763.

Jordanova:1980:RSM

[Jor80] L. J. Jordanova. Romantic science? Michelet, morals, and nature. British Journal for the History of Science, 13(1): 44–50, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026119.

Jordanova:1983:BJP

[Jor83a]  L. J. Jordanova.  Biogra-

phy: Jean-Pierre Faure, Le cas Lamarck, Paris and Périgueux: Blanchard and Fanlac, 1978. Pp. 133. 32F. British Journal for the History of Science, 16(1):110, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026115.

Jordanova:1983:BRL

[Jor83b] L. J. Jordanova. Book review: Life Sciences: Joseph Schiller, Physiology and classification: historical relations. Paris: Maloine, 1980. Pp. xiii + 194. No price given, (paperback). British Journal for the History of Science, 16(2): 215, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026256.

Jordanova:1984:L

[Jor84] L. J. Jordanova. Lamarck. Past masters. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1984. ISBN 0-19-287588-4, 0-19-287587-6 (paperback). viii + 118 pp. LCCN QH31.L2 J67 1984. US$15.90, US$3.90.

Jordanova:1985:BRG

[Jor85a] L. J. Jordanova. Book review: Graeme Tytler, Physiognomy in the European Novel. Faces and Fortunes. Princeton: Princeton University Press, 1982. Pp. xix + 436 ISBN 0-691-06491-1 £19.10. British Journal for the History of Science, 18(1): 109–110, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026290.

Jordanova:1985:BRW

[Jor85b] L. J. Jordanova. Book review: William Coleman, Death is a social disease. Public health and political economy in early industrial France. Madison, Wisconsin: University of Madison Press, 1982. Pp. xxi + 322. $35.00/£26.25. British Journal for the History of Science, 18(1):86–87, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026271.

Jordanova:1987:BRT

[Jor87] L. J. Jordanova. Book reviews: Thomas L. Hankins. Science and the Enlightenment. Cambridge: Cambridge University Press, 1985. Pp. viii + 216. ISBN 0-521-24349-9 (hardcover). £20.00. ISBN 0-521-28619-0 (paperback). £6.95. British Journal for the History of Science, 20(1):108– 110, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026463.

Jordanova:1988:BRS

[Jor88] L. J. Jordanova. Book review: Scott Atran. Fondements de l’Histoire Naturelle. Pour une Anthropologie de la Science. Bruxelles: Editions Complexe, 1986. Pp.

244. ISBN 2-87027-180-8. FF 145.00. British Journal for the History of Science, 21(3): 371, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026843.

Josten:1967:BRE

[Jos67] C. H. Josten. Book review: The English Paracelsians. By Allen G. Debus. London: Oldbourne Press. Pp. 222. 1965. 45s. British Journal for the History of Science, 3(3):296– 297, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025057.

Josephson:1997:NAR

[Jos97]  Paul R. Josephson.  New

Atlantis revisited: Akademgorodok, the Siberian city of science. Princeton University Press, Princeton, NJ, USA, 1997. ISBN 0-691-04454-6. xxii + 351 pp. LCCN Q127.R9 J67

                                        1997. URL http://www.loc.
                                        gov/catdir/description/
                                        prin021/96045577.html;
                                        http://www.loc.gov/
                                        catdir/toc/prin031/
                                        96045577.html.

Joy:1987:GAA

[Joy87] Lynn Sumida Joy. Gassendi, the atomist: advocate of history in an age of science. Ideas in context. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-30142-4. xiii + 311 pp. LCCN B1887 .J69

                                        1987. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/87010899.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam034/
                                        87010899.html;   http://

www.loc.gov/catdir/toc/
cam031/87010899.html.

Jardine:1996:CNH

[JSS96] Nicholas Jardine, James A.

Secord, and E. C. (Emma C.) Spary, editors. Cultures of natural history. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-45394-1 (hardcover), 0-521-55894-8 (paperback). xxi + 501 pp. LCCN QH15 .C85 1996. URL http:

//www.loc.gov/catdir/
description/cam026/
94039445.html;   http://
www.loc.gov/catdir/toc/
cam021/94039445.html.

Judson:1988:BRJ

[Jud88] Lindsay Judson. Book review:

J. D. G. Evans. Aristotle. Sussex and New York: Harvester Press and St. Martin’s Press, 1987. Pp. xii + 208. ISBN 0-7108-1042-3 and 0-312-00485-0. £28.50. British Journal for the History of Science, 21(3): 377, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026848.

Junker:1989:DBR

[Jun89] Thomas Junker. Darwinismus und Botanik: Rezeption, Kritik und theoretische Alternativen im Deutschland des 19. Jahrhunderts. (German) [Darwinism and Botany: Reception, criticism and theoretical alternatives in 19th Century Germany], volume 54 of Quellen und Studien zur Geschichte der Pharmazie. In Kommission, Deutscher Apotheker Verlag, Stuttgart, Germany, 1989. ISBN 3-7692-1225-8. x + 367 pp. LCCN RS61 .Q4 Bd.54.

Kepler:1981:SUM

[KA81] Johannes Kepler and E. J. Aiton. The secret of the universe = Mysterium cosmographicum, volume 9 of Janus series. Abaris Books, New York, NY, USA, 1981. ISBN 0-913870-64-1. 267 pp. LCCN QB41 .K413 1981.

Kallinen:1995:CSN

[Kal95a] Maija Kallinen. Change and stability: natural philosophy at the Academy of Turku, 1640– 1713, volume 51 of Studia historica, 0081-6493. Finnish Historical Society, Helsinki, Finland, 1995. ISBN 951-710-001-6. 439 pp. LCCN QC9.F5 K35 1995.

Kalthoff:1995:CEE

[Kal95b] Mark A. Kalthoff, editor. Creation and evolution in the early American Scientific Affiliation, volume 10 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1811-1. xxxix + 468 pp. LCCN BS659 .C73 1995.

Kargon:1976:BRI

[Kar76] Robert H. Kargon. Book review: Institutions Artisan to Graduate. Essays to commemorate the Foundation in 1824 of the Manchester Mechanics’ Institution, now in 1974 the University of Manchester Institute of Science and Technology. Ed. by D. S. L. Cardwell. Manchester: Manchester University Press, 1974. Pp. viii + 284. £5.00. British Journal for the History of Science, 9(1):70–71, March 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025709.

Kauffman:1984:BRC

[Kau84] George B. Kauffman. Book review: Chemistry and Biochemistry Alexander Todd, A Time to Remember: The Autobiography of a Chemist. Cambridge, England: Cambridge University Press, 1983, Pp. viii + 257. ISBN 0-521-25593-7. £15, $29.95. British Journal for the History of Science, 17 (2):242–243, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026574.

Kroll:1992:PSR

[KAZ92] Richard W. F. Kroll, Richard Ashcraft, and Perez Zagorin, editors. Philosophy, science, and religion in England, 1640–1700. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41095-9 (hardcover). LCCN BX5126 .P48 1991. URL http:

                                        //www.loc.gov/catdir/
                                        description/cam024/
                                        91002387.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam028/91002387.html.

Kapitsa:1990:KCM

[KBRS90] P. L. (Petr Leonidovich) Kapitsa, J. W. Boag, P. E. Rubinin, and David Shoenberg. Kapitza in Cambridge and Moscow: life and letters of a Russian physicist. North-Holland Publishing Co., Amsterdam, The Netherlands, 1990. ISBN 0-444-98749-5 (paperback), 0-444-98753-3 (hardbound). xvi + 429 pp. LCCN QC16.K25 A4 1990.

Kamminga:1995:SCN

[KC95] Harmke Kamminga and Andrew Cunningham, Dr, editors. The science and culture of nutrition, 1840–1940, volume 32 of Clio medica; The Wellcome Institute series in the history of medicine. Rodopi, Amsterdam and Leiden, The Netherlands, 1995. ISBN 90-5183-818-2, 90-5183-819-0 (paperback). vii + 344 pp. LCCN TX353 .S385 1995.

Keeney:1988:BRR

[Kee88]  Elizabeth  Barn-

aby Keeney. Book review: Robert V. Bruce. The Launching of Modern American Science, 1846–1876. New York: Alfred A. Knopf, 1987. Pp. x + 446. ISBN 0-394-55394-2. $30.00. British Journal for the History of Science, 21(1): 134–135, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026889.

Keeney:1989:BRD

[Kee89]  Elizabeth  Barnaby

Keeney. Book review: David N. Livingstone. Nathaniel Southgate Shaler and the Culture of American Science. Tuscaloosa, Al.: The University of Alabama Press, 1987. Pp. xiv + 395. ISBN 0-8173-0305-7. $32.95. British Journal for the History of Science, 22(1): 98–99, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026697.

Keith:1984:BRH

[Kei84] S. T. Keith. Book review: History of Technology: Clayton R. Koppes, JPL and the American Space Programme: A History of the Jet Propulsion Laboratory. New Haven: Yale University Press, 1982. Pp. viii+ 320. £16.95/$19.95. Joan Lisa Bromberg, Fusion: Science, Politics and the Invention of a New Energy Source. Cambridge, Massachusetts: The MIT Press, 1982. Pp. xxvi + 343. £24/$30. British Journal for the History of Science, 17(3):325–326, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026640.

Keith:1986:BRW

[Kei86] S. T. Keith. Book review: William McGucken. Scientists, Society and State. The Social Relations of Science Movement in Great Britain, 1931–1947. Columbus: Ohio State University Press, 1984. Pp. xiii + 381. ISBN 0-8142-0351-5. £22.50. British Journal for the History of Science, 19(3): 351–352, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026523.

Keith:1989:BRR

[Kei89] S. T. Keith. Book review: Rajkumari Williamson (ed.). The Making of Physicists. Bristol: Adam Hilger, 1987. Pp. viii + 200. ISBN 0-85274-524-9. £15.00. British Journal for the History of Science, 22(1): 115–116, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026711.

Keller:1967:PAV

[Kel67] A. G. Keller. Pneumatics, automata and the vacuum in the work of Giambattista Aleotti. British Journal for the History of Science, 3(4):338– 347, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024959.

Keller:1968:BRH

[Kel68] A. G. Keller. Book review: History of Technology: Mechanical Universe. The Astrarium of Giovanni de’ Dondi. By Silvio A. Bedini and Francis R. Maddison. Transactions of the American Philosophical Society, New Series, vol. lvi, part 5-October 1966; Pp. 69, 53 figs. $2.50. British Journal for the History of Science, 4(2): 176–178, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025150.

Keller:1970:BRR

[Kel70] A. G. Keller. Book review: Renaissance: Leonardo da Vinci. By V. P. Zubov. Translated from the Russian by David H. Kraus. Harvard University Press, and Oxford University Press. 1968. Pp. xx + 335. 95s. British Journal for the History of Science, 5(2):203– 204, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025434.

Kelham:1971:BRS

[Kel71] B. B. Kelham. Book review: Sixteenth and Seventeenth Centuries: The Common Scientist in the Seventeenth Century. A Study of the Dublin Philosophical Society. By K. Theodore Hoppen. London: Routledge & Kegan Paul. 1970. Pp. xii + 297. £2.75. British Journal for the History of Science, 5(4): 417–418, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025404.

Keller:1973:BRR

[Kel73] A. G. Keller. Book review: Renaissance: Mariano Taccola and his Book De Ingeneis. Ed. by Frank D. Prager and Gustina Scaglia. Cambridge, Mass., and London: M.I.T. Press, 1972. Pp. xii + 230. $17.50. British Journal for the History of Science, 6(4): 439–440, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025513.

Keller:1975:BRT

[Kel75]  A.  G.

Keller. Book review: Technology The Book of Knowledge of Ingenious Mechanical Devices. By Ibn al-Razzaz al-Jazzari. Trans, and ed. by Donald R. Hill. Dordrecht, Holland and Boston, Mass.: D. Reidel, 1974. Pp. xxvi + 286. No price stated. British Journal for the History of Science, 8 (1):75–77, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025822.

Keller:1976:BRT

[Kel76] A. G. Keller. Book review:

Technology The Maze of Ingenuity. By Arnold Pacey. London: Allen Lane, 1974. Pp.

350. £5.50. British Journal for the History of Science, 9 (1):67–68, March 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025706.

Keller:1978:BRR

[Kel78]  A. G. Keller.  Book re-

view: Renaissance: The Italian Renaissance of Mathematics: Studies on Humanists and Mathematicians from Petrarch to Galileo. By Paul Lawrence Rose. Geneva: Librairie Droz, 1975. Pp. xiii + 316. No price stated. British Journal for the History of Science, 11(3): 289–290, November 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025788.

Keller:1983:IAP

[Kel83]  Alex Keller.  The Infancy

of Atomic Physics: Hercules in His Cradle. Clarendon Press, New York, NY, USA, 1983. ISBN 0-19-853904-5. 230 pp. LCCN QC773 .K4 1983. US$19.50.

Keller:1984:BRH

[Kel84] Alex Keller. Book review: History of Technology: Trevor I. Williams, A short history of twentieth-century technology. Oxford: Oxford University Press, 1982. Pp. xix + 411. ISBN 0-19-858159-9. £12.50. British Journal for the History of Science, 17(3): 328, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026643.

Keller:1985:BRR

[Kel85a] Alex Keller. Book review: Russell McCormmach, Night Thoughts of a Classical Physicist. Cambridge, Mass.: Harvard University Press, 1982. Pp. 217, £10.50. British Journal for the History of Science, 18(2):233–234, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Keller:1985:RU

[Kel85b] Alex Keller. Review: [untitled]. British Journal for the History of Science, 18(2): 234, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026333.

Kent:1964:TTH

[Ken64] Andrew Kent. Thomas Thomson (1773–1852) historian of chemistry. British Journal for the History of Science, 2(1): 59–63, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025084.

Kennedy:1983:LSP

[Ken83] Don H. Kennedy. Little sparrow: a portrait of Sophia Kovalevsky. Ohio University Press, Athens, OH, USA, 1983. ISBN 0-8214-0692-2, 0-8214-0703-1 (paperback). ix + 341 pp. LCCN QA29.K67 K46 1983.

Kenney:1986:BUI

[Ken86] Martin Kenney. Biotechnology: the university-industrial complex. Yale University Press, New Haven, CT, USA, 1986. ISBN 0-300-03392-3. xvi + 306 pp. LCCN TP248.2 .K46 1986.

Keith:1986:FRS

[KH86] S. T. Keith and Paul K. Hoch.

Formation of a research school: Theoretical solid state physics at Bristol 1930–1954. British Journal for the History of Science, 19(1):19–44, March 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026482.

King-Hele:1992:TO

[KH92] Desmond King-Hele. A tapestry of orbits. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-39323-X. x + 244 pp. LCCN TL1080 .K55

1992. URL http://www.loc.
gov/catdir/description/
cam025/92011674.html;
http://www.loc.gov/
catdir/enhancements/
fy0632/92011674-t.html.

Kaijser:1995:NES

[KH95]  Arne  Kai-

jser and Marika Hedin, editors. Nordic energy systems: historical perspectives and current issues. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1995. ISBN 0-88135-164-4. LCCN TJ163.25.S34 N68 1995.

Kolmogorov-:1992:MXV

[KI92] A. N. (Andrei Nikolaevich) Kolmogorov and A. P. (Adolf Pavlovich) Iíushkevich, editors. Matematika XIX veka. Mathematics of the 19th century: mathematical logic, algebra, number theory, probability theory. Birkhäuser Verlag, Basel, Switzerland, 1992. ISBN 3-7643-2552-6 (Basel), 0-8176-2552-6 (Boston). 308 pp. LCCN QA26 .M27813 1992.

Kilmister:1984:BRPb

[Kil84a] C. W. Kilmister. Book review: Physics and its concepts: Gerald Holton and Yehuda Elkana (ed), Albert Einstein: historical and cultural perspective. Princeton: University Press, 1982. Pp. xxxii + 439. ISBN 0-691-08299-5. £24.70. British Journal for the History of Science, 17(2):227, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026561.

Kilmister:1984:BRPa

[Kil84b] C. W. Kilmister. Book review: Physics and its concepts: Jonathan Powers, Philosophy and the new physics. London: Methuen, 1982. Pp xvii + 203. ISBN 0-416-73480-4. £3.95. British Journal for the History of Science, 17(2): 223–224, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026557.

Kilmister:1987:SCC

[Kil87] Clive William Kilmisterr, editor. Schrödinger: centenary celebration of a polymath. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34017-9. LCCN QC16.S265 S36 1987; Q143.S3. URL http://

www.loc.gov/catdir/
enhancements/fy0901/
86028338-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0901/86028338-t.html.

Kilmister:1994:ESF

[Kil94] C. W. (Clive William) Kilmister. Eddington’s search for a fundamental theory: a key to the universe. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-37165-1 (hardcover). xii + 256 pp. LCCN QC6 .K439

                                        1994. URL http://www.loc.
                                        gov/catdir/description/
                                        cam026/94009035.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam021/
                                        94009035.html.

Kipnis:1991:HPI

[Kip91]  Naum S. Kipnis.  History

of the principle of interference of light, volume 5 of Science networks historical studies. Birkhäuser Verlag, Basel, Switzerland, 1991. ISBN 3-7643-2316-7 (Basel), 0-8176-2316-7 (US). 271 pp. LCCN QC411 .K53 1990.

Kitson:1989:HAC

[Kit89] Annabella Kitson, editor. History and astrology: Clio and Urania confer. Unwin Paperbacks, London, UK, 1989. ISBN 0-04-440522-7. ix + 272 pp. LCCN BF1671 .H57 1989. US$8.99.

Kitcher:1992:FDC

[Kit92]  Patricia Kitcher.  Freud’s

dream: a complete interdisciplinary science of mind. MIT Press, Cambridge, MA, USA, 1992. ISBN 0-262-11172-1. 245 pp. LCCN BF109.F74 K58 1992.

Kohn:1985:DHI

[KK85] David Kohn and Malcolm J. Kottler, editors. The Darwinian heritage: including proceedings of the Charles Darwin Centenary Conference, Florence Center for the History and Philosophy of Science, June 1982. Princeton University Press in association with Nova Pacifica, Princeton, NJ, USA, 1985. ISBN 0-691-08356-8. LCCN QH31.D2 D385 1985. US$95.00.

Kors:1987:AEE

[KK87] Alan Charles Kors and Paul J.

Korshin, editors. Anticipations of the Enlightenment in England, France, and Germany. University of Pennsylvania Press, Philadelphia, PA, USA, 1987. ISBN 0-8122-8057-1. viii + 290 pp. LCCN B1302.E65 A57 1987.

Klein:1994:VAG

[Kle94]  Ursula Klein.  Verbindung

und Affinität: die Grundlegung der neuzeitlichen Chemie an der Wende vom 17 zum 18 Jahrhundert. (German) [Connection and affinity: the foundation of modern chemistry at the turn of the 17th to 18th Centuries], volume 14 of Science networks, historical studies. Birkhäuser Verlag, Basel, Switzerland, 1994. ISBN 3-7643-5003-2. viii + 270 pp. LCCN QD11 .K526 1994.

Kuhn:1988:SCC

[KN88] Dieter Kuhn and Joseph Needham, editors. Science and civilisation in China. Vol. 5. Chemistry and chemical technology. Pt. 9, Textile technology: Spinning and reeling. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-32021-6. xxxiv + 520 pp. LCCN ????

Knight:1965:BRM

[Kni65] D. M. Knight. Book review: Michael Faraday. By L. Pearce Williams. Pp. xvi + 531. London: Chapman and Hall, 1965. £3 10s. British Journal for the History of Science, 2(4): 363–364, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024899.

Knight:1967:BRH

[Kni67]  D. M. Knight.  Book re-

view: History of Chemistry History of Analytical Chemistry. By Ferenc Szabadváry, tr. Gyula Svehla. Pp. ix + 418. Oxford: Pergamon Press. 1966. £6. British Journal for the History of Science, 3(4): 402, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024971.

Knight:1969:BREb

[Kni69a]  D. M. Knight.  Book re-

view: Eighteenth and Nineteenth Centuries: John Dalton and the Progress of Science. Ed. by D. S. L. Cardwell. Manchester: Manchester University Press. 1968. Pp. xxi + 352. Plates. 55s. British Journal for the History of Science, 4(4):420–421, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025253.

Knight:1969:BREa

[Kni69b]  D. M. Knight.  Book re-

view: Eighteenth Century: A Scientific Autobiography of Joseph Priestley (1733–1804): Selected Scientific Correspondence. Edited with Commentary by Robert E. Schofield. Cambridge, Mass. and London: M.I.T. Press. 1966. Pp. xiv + 415. $13.50. British Journal for the History of Science, 4 (3):301–302, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025494.

Knight:1971:BM

[Kni71a] David Knight. Back matter. British Journal for the History of Science, 5(4):432, December 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025411.

Knight:1971:BRNb

[Kni71b] David Knight. Book review:

Nineteenth and Twentieth Centuries: The Royal Institution Library of Science-Physical Sciences. Ed. by W. L. Bragg and G. Porter, 10 vols. + Index vol. London: Elsevier. 1970. Pp. 577, 549, 569, 548, 569, 450, 511, 536, 550, 477, 226. £63. British Journal for the History of Science, 5(4):421– 422, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025408.

Knight:1971:BRNa

[Kni71c]  David M. Knight.  Book

review: Nineteenth Century: Sir Humphry Davy’s Published Works. By June Z. Fullmer. Cambridge, Mass.: Harvard University Press and Oxford University Press. Pp. viii + 112. £3.15. British Journal for the History of Science, 5 (3):314, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025344.

Knight:1972:BRE

[Kni72]  D. M. Knight.  Book re-

view: Eighteenth and Nineteenth Centuries: The History of Valency. By C. A. Russell. Leicester: Leicester University Press, 1971. Pp. xii + 373. £5.50. British Journal for the History of Science, 6(2): 220–221, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025307.

Knight:1974:BRE

[Kni74] David M. Knight. Book review: Eighteenth and Nineteenth Centuries: John Dalton: Critical Assessments of his Life and Science. By Arnold Thackray. Cambridge, Mass.: Harvard University Press, 1972. Pp. xii + 190. £5.75. British Journal for the History of Science, 7(2):197–198, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025222.

Knight:1975:BRG

[Kni75]  David Knight.  Book re-

view: General Science and the Arts. A Study in Relationships from 1600 to 1900. By Jacob Opper. Rutherford, Madison, and Teaneck: Farleigh Dickinson University Press. Pp.

226. $10.50. British Journal for the History of Science, 8 (2):175–176, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025643.

Knight:1976:BRC

[Kni76] David Knight. Book review:

Chemistry The Chemists and the Word. The Didactic Origins of Chemistry. By Owen Hannaway. Baltimore and London: The Johns Hopkins University Press, 1975. Pp. xv + 165. £6.50. British Journal for the History of Science, 9 (3):328, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025690.

Knight:1980:BRG

[Kni80a] David Knight. Book review: General What is this Thing Called Science? By A. F. Chalmers. Milton Keynes: The Open University Press, 1978. pp. xvii + 157. £3.75. British Journal for the History of Science, 13(1):61–62, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026122.

Knight:1980:BRM

[Kni80b] David M. Knight. Book review: Middle Ages and Renaissance: The Printing Press as an Agent of Change. By Elizabeth Eisenstein. Cambridge: Cambridge University Press, 1979. pp. xxi + 794 in 2 vols. £35.00. British Journal for the History of Science, 13(2):164–166, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025899.

Knight:1981:BRP

[Kni81] David Knight. Book review: Philosophy of Science Paradigms and Revolutions: Appraisals and Applications of Thomas Kuhn’s Philosophy of Science. Ed. by Gary Gutting. Notre Dame & London: University of Notre Dame Press, 1980. pp. viii + 339. $18.95/$7.95 (paperback). British Journal for the History of Science, 14(3):

283–284, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025984.

Knight:1983:BRA

[Kni83a] David Knight. Book review:

Americana Clark A. Elliott, Biographical dictionary of American science. Westport, Conn. & London: Greenwood Press, 1979. Pp. xvii + 360. Frank N. Schubert (ed.), Explorer on the northern plains: Lieut. Gouverneur K. Warren’s preliminary report of explorations in Nebraska and Dakota, 1855- ’56-’57. Washington, D.C.: Historical Division, U.S. Army Corps of Engineers, 1981. Pp. xxxiv + 125. $5.00; from Government Printing Office, citing stock number 008-022-00165-4. British Journal for the History of Science, 16(3):318– 319, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026187.

Knight:1983:BRE

[Kni83b] David Knight. Book review:

Essay Collections H. Woolf (ed), The analytic spirit: essays in the history of science in honor of Henry Guerlac. Ithaca & London: Cornell University Press, 1981. Pp. 363. £15.00/$31.25. British Journal for the History of Science, 16 (2):195–196, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026240.

Knight:1983:BRL

[Kni83c] David Knight. Book review: Life and earth sciences: Alwyne Wheeler and James H. Price (eds.), History in the service of systematics. London: Society for the Bibliography of Natural History, 1981. Pp. v + 164. £10.00/$23.00. British Journal for the History of Science, 16(3):277–278, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026149.

Knight:1984:BRHa

[Kni84a] David Knight. Book review: History of Natural History: Timothy Lenoir, The strategy of life: teleology and mechanics in nineteenth century German biology. Dordrecht: D. Reidel, 1982. Pp. xii + 314. ISBN 90-277-1363-4. DFL 135.00, $59.00. British Journal for the History of Science, 17 (3):318, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026631.

Knight:1984:BRHb

[Kni84b] David Knight. Book review: History of Natural History: Timothy Lenoir, The strategy of life: teleology and mechanics in nineteenth century German biology. Dordrecht: D. Reidel, 1982. Pp. xii + 314. ISBN 90-277-1363-4. DFL 135.00, $59.00. British Journal for the History of Science, 17(3):318–319, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026632.

Knight:1985:BRA

[Kni85]  David Knight.  Book re-

view: Allen G. Debus, Science and History: a chemist’s appraisal. Coimbra: University Press, 1984. Pp. vi + 72. No price given. R. W. Home (ed), Science under Scrutiny. The place of History and Philosophy of Science. Dordrecht, Boston, Lancaster: D. Reidel, 1983. Pp. xvii + 182. ISBN 90-277-1602-1. Dfl. 90.00. British Journal for the History of Science, 18 (3):361, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026399.

Knight:1986:ASS

[Kni86a] David Knight. The age of science: the scientific worldview in the Nineteenth Century. Blackwell, Oxford, UK, 1986. ISBN 0-631-16176-7. 251 pp. LCCN ????

Knight:1986:BRD

[Kni86b] David Knight. Book review:

Derek Gjertsen. The Classics of Science. A Study of Twelve Enduring Scientific Works. New York and Gravesend: Lilian Barber Press, 1984. Pp. iii + 374. ISBN 0-936508-09-4. £20.00 (hardcover), £12.95 (paperback). British Journal for the History of Science, 19 (3):367–368, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026539.

Knight:1986:BRG

[Kni86c] David Knight. Book review: Garry Adler. Beyond Bokhara: the Life of William Moorcroft, Asian Explorer and Pioneer Veterinary Surgeon, 1767–1825. London: Century Publishing, 1985. Pp. xiii + 417. ISBN 0-7126-0722-6. £16.95. British Journal for the History of Science, 19 (3):359, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026529.

Knight:1986:BRW

[Kni86d] David Knight. Book review: W. H. Brock & A. J. Meadows. The Lamp of Learning. Taylor & Francis and the Development of Science Publishing. London: Taylor & Francis, 1984. Pp. xvi + 240. ISBN 0-85066-253-2. £15.00. British Journal for the History of Science, 19(3): 365–366, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026537.

Knight:1987:BFG

[Kni87] David Knight. Background and foreground: Getting things in context. British Journal for the History of Science, 20(1): 3–12, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026432. Conference

on documentary editing (Tarrytown, N.Y., 1985).

Knight:1988:BRA

[Kni88a] David Knight. Book review:

Ann Moyal. A Bright and Savage Land: Scientists in Colonial Australia. Sydney: Collins. Pp. 192. ISBN 0-00-217555-X. £17.50. British Journal for the History of Science, 21 (2):259, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026989.

Knight:1988:BRS

[Kni88b]  David Knight.  Book re-

view: Stanton J. Linden (ed.). William Cooper. A Catalogue of Chymicall Books, 1673–88. New York and London: Pp. liv + 159. ISBN 0-8240-8557-4. $37.00. British Journal for the History of Science, 21 (2):259–260, June 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026990.

Knight:1988:BRJ

[Kni88c]  David Knight.  Book re-

views: José M. López Pi ero. El Grabado en la Ciencia Hispánica. Madrid: Consejo Superior de Investigaciones Cientificas, 1987. Pp. 140. ISBN 84-00-06699-5. No price given. Philipp Hauser. Atlas Epidemiográfico del Cólera de 1885 en Espa a (1887), reprint intr. J. M. López Piñero. Valencia: Consellaria de Sanitat y Consum, 1987. No ISBN or price given. British Journal for the History of Science, 21(4):496– 497, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026968.

Knight:1988:ASH

[Kni88d] David M. Knight, editor. The age of science: the scientific world-view in the Nineteenth Century. Blackwell, Oxford, UK, 1988. ISBN 0-631-16176-7 (paperback). x + 251 pp. LCCN ????

Knight:1989:ASS

[Kni89a] David Knight. The age of science: the scientific world-view in the 19th Century. Blackwell, Oxford, UK, 1989. ISBN 0-631-16176-7, 0-631-15064-1. x + 251 pp. LCCN ????

Knight:1989:BRF

[Kni89b] David Knight. Book review:

F. Burwick (ed.). Approaches to Organic Form: Permutations in Science and Culture. Dordrecht, Boston, Lancaster & Tokyo: D. Reidel, 1987. Pp. xix + 286. ISBN 90-277-2541-1. £43.95; DFl.160.00; $64.00. British Journal for the History of Science, 22(1): 110–111, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026707.

Knight:1989:BRJ

[Kni89c] David Knight. Book review:

John Nudos, Norman McMillan, Denis Weare & Susan McKenna Lawlor (eds). Science in Ireland 1800–1930: Tradition and Reform. Dublin: Trinity College, 1988. Pp. 208. ISBN 0-9513586-1-8. IR£10.00. (Available from Physics Department, Trinity College.). British Journal for the History of Science, 22(4):457– 458, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026933.

Knight:1989:BRR

[Kni89d] David Knight. Book review:

Roy Macleod & Philip Rehbock, (eds). Nature in its Greatest Extent: Western Science in the Pacific. Honolulu: University of Hawaii Press, 1988. Pp. xiii + 288. ISBN 0-8248-1120-8. $34.00. British Journal for the History of Science, 22(3): 394–395, September 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026911.

Knight:1989:BRM

[Kni89e] David Knight. Book reviews: Margot Walker. Sir James Edward Smith 1759–1828. London: Linnean Society, 1988. Pp. viii + 60. ISBN 0-9506207-1-8. £4.00 incl p. & p.; £3.00 if collected from the Linnean Society. A. T. Gage & W. T. Stearn. A Bicentenary History of the Linnean Society of London. London: Academic Press, 1988. ISBN 0-12-273150-6. No price given. Gabriele Gramiccia. The Life of Charles Ledger (1818–1905): Alpacas and Quinine. London: Macmillan, 1988. ISBN 0-333-45710-2. £30.00. British Journal for the History of Science, 22(4): 477–478, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026947.

Knight:1992:HDS

[Kni92a] David M. Knight. Humphry Davy: science and power. Blackwell science biographies. Blackwell, Oxford, UK, 1992.

ISBN 0-631-16816-8. xiii + 218 pp. LCCN QD22.D3 K65 1992.

Knight:1992:ICH

[Kni92b] David M. Knight, editor. Ideas in chemistry: a history of the science. Athlone Press, London, UK, 1992. ISBN 0-485-11390-2. 213 pp. LCCN QD11 .K65 1992. US$35.00.

Knorr:1985:GAP

[Kno85] W. R. Knorr. The geometer and the archaeoastronomers: On the prehistoric origins of mathematics. British Journal for the History of Science, 18 (2):197–212, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/

4026330. Review of: Geometry and algebra in ancient civilizations [Springer, Berlin, 1983; MR0706258 (85b:01001)] by B. L. van der Waerden.

Kohler:1975:GLV

[Koh75] Robert E. Kohler. G. N.

Lewis’s views on bond theory 1900–1916. British Journal for the History of Science, 8(3): 233–239, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025558.

Kohlstedt:1978:BRA

[Koh78]  Sally  Gre-

gory Kohlstedt. Book review: American Science Early American Science. Ed. by Brooke Hindle. New York: Science History Publications, 1976. Pp. xiv + 213. $7.95; $4.95 (paperback). Science in America since 1820. Ed. by Nathan Reingold. New York: Science History Publications, 1976. Pp.

334. $8.95; $5.95 (paperback). Philosophers and Machines. Ed. by Otto Mayr. New York: Science History Publications, 1976. Pp. 193. $8.95; $4.95 (paperback). British Journal for the History of Science, 11 (1):71–72, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025614.

Kohn:1989:DAS

[Koh89] David Kohn. Darwin’s ambiguity: The secularization of biological meaning. British Journal for the History of Science, 22(2):215–239, July 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026661.

Kragh:1984:BRP

[Kra84a] Helge Kragh. Book review: Physics and its concepts: Edward M. MacKinnon, Scientific explanation and atomic physics. Chicago: The University of Chicago Press, 1982. Pp. 450. £22. British Journal for the History of Science, 17 (2):224–225, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026558.

Kragh:1984:JTC

[Kra84b] Helge Kragh. Julius Thomsen and classical thermochemistry. British Journal for the History of Science, 17(3):255– 272, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026622.

Kragh:1986:BRD

[Kra86]  Helge  Kragh.

Book review: David Cassidy & Martha Baker. Werner Heisenberg. A Bibliography of his Writings (Berkeley Papers in History of Science, IX), Berkeley: Office for History of Science and Technology, University of California, 1984. pp.

153. W. Heisenberg. Gesammelte Werke/Collected Works, Series B. Eds, W. Blum, H.- P. Dürr and H. Rechenberg. Berlin, Heidelberg, New York, Tokyo: Springer-Verlag, 1984. Pp. 947. Elisabeth Heisenberg. Inner Exile. Recollections of a Life with Werner Heisenberg. Translated by S. Cappellari and C. Morris, Boston, Basel, Stuttgart: Birkhäuser, 1984. Pp. 170. British Journal for the History of Science, 19(2):232, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026614.

Kragh:1987:IHS

[Kra87] Helge Kragh. An introduction to the historiography of science. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-33360-1. viii + 235 pp. LCCN Q125 .K686

                                        1987. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0642/86028334-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0642/86028334-t.html.

Kragh:1990:DSB

[Kra90] Helge Kragh. Dirac: a scientific biography. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-38089-8. x + 389 pp. LCCN QC16.D57 K73

                                        1990. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/89017257.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        89017257.html.

Krayer:1991:MIS

[Kra91] Albert Krayer. Mathematik im Studienplan der Jesuiten: die Vorlesung von Otto Cattenius an der Universität Mainz (1610/11). (German) [Mathematics in Jesuit Studies: the lectures of Otto Cattenius at the University of Mainz (1610/ 11)], volume 15 of Beiträge zur Geschichte der Universität Mainz. Steiner, Stuttgart, Germany, 1991. ISBN 3-515-05874-5. ix + 434 pp. LCCN

                                        ????  URL 01; 04; 2;
                                        http://www.gbv.de/dms/

hbz/toc/ht004191413.;
http://www.zentralblatt-
math.org/zmath/en/search/
?an=0790.01006.

Kren:1969:MOP

[Kre69] Claudia Kren. A medieval objection to “Ptolemy”. British Journal for the History of Science, 4(4):378–393, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025231.

Kretzmann:1982:ICA

[Kre82] Norman Kretzmann, editor. Infinity and continuity in ancient and medieval thought. Cornell University Press, Ithaca, NY, USA, 1982. ISBN 0-8014-1444-X. 367 pp. LCCN BD411 .I53.

Kremer:1990:TLE

[Kre90] Richard Lynn Kremer. The thermodynamics of life and experimental physiology, 1770– 1880. Harvard dissertations in the history of science. Garland Publishing, Inc., New York, NY, USA, 1990. ISBN 0-8240-0039-0. xii + 512 pp. LCCN QP517.T48 K74 1990.

Krimsky:1982:GAS

[Kri82]  Sheldon Krimsky.  Genetic

alchemy: the social history of the recombinant DNA controversy. MIT Press, Cambridge, MA, USA, 1982. ISBN 0-262-11083-0. xiii + 445 pp. LCCN QH442 .K74 1982.

Krikorian:1983:BRL

[Kri83] A. D. Krikorian. Book review: Life Sciences: G. C. Ainsworth, Introduction to the history of plant pathology. Cambridge, Cambridge University Press, 1981. Pp. xii + 315. £27.50. British Journal for the History of Science, 16(2):213, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026254.

Krisciunas:1988:ACW

[Kri88] Kevin Krisciunas. Astronomical centers of the world. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-30278-1. x + 320 pp. LCCN QB81 .K75 1988. URL http:

//www.loc.gov/catdir/
description/cam023/
86021542.html;   http://
www.loc.gov/catdir/toc/
cam028/86021542.html.

Kroll:1991:MWL

[Kro91] Richard W. F. Kroll. The material word: literate culture in the Restoration and early Eighteenth Century. The Johns Hopkins University Press, Baltimore, MD, USA, 1991. ISBN 0-8018-4002-3. xxii + 420 pp. LCCN PR437 .K7

1990. URL http://www.loc.
gov/catdir/description/
jhu051/90030586.html.

Kirk:1983:PPC

[KRS83] G. S. (Geoffrey Stephen) Kirk, J. E. (John Earle) Raven, and Malcolm Schofield. The presocratic philosophers: a critical history with a selection of texts. Cambridge University Press, Cambridge, UK, second edition, 1983. ISBN 0-521-25444-2, 0-521-27455-9 (paperback). xiii + 501 pp. LCCN B188 .K5 1983. URL http:/

/www.loc.gov/catdir/
description/cam022/
82023505.html;   http://
www.loc.gov/catdir/toc/
cam028/82023505.html.

Krige:1994:ES

[KRS94] John Krige, Arturo Russo, and Lorenza Sebesta. Europe in space, 1960–1973, volume 1172 of ESA SP. European Space Agency, Noordwijk, The Netherlands, 1994. ISBN 92-9092-125-0. viii + 142 pp. LCCN TL789.8.E9 K7497 1994. 70 Dfl.

Kargon:1987:KBL

[KT87] Robert Kargon and Peter Achinstein William Thomson, Baron Kelvin, editors. Kelvin’s Baltimore lectures and modern theoretical physics: historical and philosophical perspectives. Studies from the Johns Hopkins Center for the History and Philosophy of Science. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-11117-9. x + 547 pp. LCCN QC19.6 .K45 1987.

Kubrin:1970:BRS

[Kub70]  David Kubrin.  Book re-

view: Sixteenth and Seventeenth Centuries: A Portrait of Isaac Newton. By Frank E. Manuel. Harvard University Press and Oxford University Press. Pp. xvi + 478. 1968. 114s. British Journal for the History of Science, 5 (1):100–101, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025370.

Kuhn:1968:BRT

[Kuh68] Thomas S. Kuhn. Book review: Twentieth Century The Old Quantum Theory. By D. ter Haar. Edited with a historical introduction. Pp. x + 206. Oxford: Pergamon Press, 1967. 21s. British Journal for the History of Science, 4 (1):80–81, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024998.

Kuklick:1991:SWS

[Kuk91] Henrika Kuklick. The savage within: the social history of British anthropology, 1885– 1945. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-41109-2. 320 pp. LCCN ????

Kuklick:1993:SWS

[Kuk93] Henrika Kuklick. The savage within: the social history of British anthropology, 1885–1945. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-45739-4, 0-521-41109-2. 325 pp. LCCN ????

Kulikovsky:1963:SSH

[Kul63] P. G. Kulikovsky. Soviet ‘Studies on the History of Astronomy’. British Journal for the History of Science, 1(4):391, December 1963. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kulikovsky:1964:SHA

[Kul64] P. G. Kulikovsky. ‘Studies on the History of Astronomy’. British Journal for the History of Science, 2(1):84–89, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kumar:1991:SEE

[Kum91] Deepak Kumar, editor. Science and Empire: essays in Indian context, 1700–1947. Anamika Prakashan, Delhi, India, 1991. ISBN 81-85150-19-2. xvi + 205 pp. LCCN Q127.I4 S322 1991. Rs200.00 (US$30.00 U.S.).

Kumar:1995:SR

[Kum95] Deepak Kumar. Science and the Raj, 1857–1905. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1995. ISBN 0-19-563562-0. xiv + 273 pp. LCCN Q127.I4 K86 1995.

Kusukawa:1995:TNP

[Kus95] Sachiko Kusukawa. The transformation of natural philosophy: the case of Philip Melanchthon. Ideas in context. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-47347-0 (hardcover). xv + 246 pp. LCCN BR339 .K87

1995. URL http://www.loc.
gov/catdir/description/
cam026/94018291.html;
http://www.loc.gov/
catdir/toc/cam022/
94018291.html.

Kuznick:1987:BLS

[Kuz87] Peter J. Kuznick. Beyond the laboratory: scientists as political activists in 1930s America. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-46583-7, 0-226-46584-5 (paperback). x + 363 pp. LCCN Q175.52.U5 K89 1987. US$37.00, US$15.00.

URL http://www.loc.gov/
catdir/enhancements/
fy0608/87005098-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/87005098-t.html.

Kepler:1937:GWG

[KvDC^{+}37] Johannes Kepler, Walther von Dyck, Max Caspar, Franz Hammer, Volker Bialas, Jürgen Hübner, and Helmuth Grössing, editors. Gesammelte Werke. (German) [Collected works]. Beck’sche Verlagsbuchhandlung, München, Germany, 1937. ISBN 3-406-37510-3 (vol. 11,2), 3-406-37511-1 (vol. 11,2), 3-406-01661-8 (vol. 12), 3-406-01660-X (vol. 12), 3-406-31501-1 (vol. 20,1), 3-406-31502-X (vol. 20,1), 3-406-40592-4 (vol.

20,2), 3-406-40593-2 (vol. 20,2). ???? pp. LCCN ????

Kursunoglu:1987:RAG

[KW87] Behram Kurşunoğlu and Eugene Paul Wigner, editors. Reminiscences about a great physicist: Paul Adrien Maurice Dirac. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34013-6. xviii + 297 pp. LCCN QC16.D57 R46

1987. URL http://www.loc.
gov/catdir/description/
cam023/86033409.html;
http://www.loc.gov/
catdir/toc/cam028/
86033409.html.

LopezPinero:1986:ANE

[Lan86] José Mariá López Piñero. El arte de navegar en la Espaãna del Renacimiento. (Spanish) [Seacraft in Renaissance Spain]. Editorial Labor, Barcelona, Spain, 2a edition, 1986. ISBN 84-335-0029-5. 285 pp. LCCN VK87 .L67 1986.

LopezPinero:1987:GCH

[Lan87] José María López Piñero. El grabado en la ciencia hispánica. (Spanish) [Engraving in Spanish science]. Consejo Superior de Investigaciones Científicas, Madrid, Spain, 1987. ISBN 84-00-06699-5. 80 + 60 pp. LCCN Q127.S7 L655 1987.

LopezPinero:1989:OEE

[Lan89] José María López Piñero. Los origenes en España de los estudios sobre la salud publica.

(Spanish) [The origins in Spain of studies on public health], volume 1 of Textos clásicos de la salud pública. Ministerio de Sanidad y Consumo, Madrid, Spain, 1989. ISBN 84-7670-164-0, 84-7670-165-9. 245 pp. LCCN RA287 .L66 1989.

LaFollette:1982:QS

[LaF82] Marcel C. (Marcel Chotkowski) LaFollette, editor. Quality in science. MIT Press, Cambridge, MA, USA, 1982. ISBN 0-262-12099-2, 0-262-62040-5 (paperback). xi + 169 pp. LCCN Q172.5.S34 Q34 1982.

Landsberg:1982:ET

[Lan82] Peter Theodore Landsberg, editor. The Enigma of Time. Adam Hilger Ltd., Bristol, UK, 1982. ISBN 0-85274-545-1. xii + 248 pp. LCCN QB209 .E54 1982. US$13.95.

Langins:1987:RAB

[Lan87] Janis Langins. La République avait besoin de savants: les débuts de l’École polytechnique — l’École centrale des travaux publics et les cours révolutionnaires de l’an III. (French) [The Republic needed sciences: the start of the Polytechnical School — the central school for public works and the revolutionary courses of the year III]. Librairie du Bicentenaire de la Révolution française. Belin, Paris, France, 1987. ISBN 2-7011-0641-9. 287 pp. LCCN T173.P39 L36 1987.

Lanaro:1991:PTS

[Lan91] Giorgio Lanaro. Il positivismo tra scienza e religione. (Italian) [The positivism between science and religion], volume 39 of Collana di filosofia. Angeli, Milano, Italy, 1991. ISBN 88-204-6373-3. 173 pp. LCCN IN PROCESS. L24000.

Larder:1974:BRE

[Lar74] David F. Larder. Book review: Eighteenth and Nineteenth Centuries: On Acid Humor Arising From Foods and On White Magnesia (Edinburgh, 1754). By Joseph Black. Trans. by Thomas Hanson, with an Introduction by Stacey B. Day. Reprint. Minneapolis: Bell Museum of Pathobiology, University of Minnesota Medical School, 1973. Pp. xiv + 40. $6.00; $2.25 (paperback). British Journal for the History of Science, 7(2): 190–191, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025216.

Lash:1981:SFC

[Las81] Nicholas Lash. Science and the French Catholics. British Journal for the History of Science, 14(2):202–205, July 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025949.

Latour:1988:PF

[Lat88] Bruno Latour. The pasteurization of France. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-65760-8. 273 pp. LCCN QR22.F8 L3813 1988.

Laudan:1968:BRH

[Lau68] Laurens Laudan. Book review: History of Optics Theories of Lights from Descartes to Newton. By A. I. Sabra. London: Oldbourne Press. 1967. Pp. 363. 70s. British Journal for the History of Science, 4(2): 181–182, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025154.

Laudan:1987:MGF

[Lau87a] Rachel Laudan. From mineralogy to geology: the foundations of a science, 1650–1830. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-46950-6. xii + 278 pp. LCCN QE11 .L35

                                        1987. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi052/86030783.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0608/86030783-t.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0609/86030783-b.html.

Laurent:1987:PEF

[Lau87b] Goulven Laurent. Paléontologie et évolution en France de 1800 à 1860: une histoire des idées de Cuvier et Lamarck à Darwin. (French) [Paleontology and evolution in France from 1800 to 1860: a history of the ideas of Cuvier and Lamarck to Darwin], volume 4 of Mémoires de la Section d’histoire des sciences et des techniques, 0769-2218. Editions du Comité des travaux historiques et scientifiques, Paris, France, 1987. ISBN 2-7355-0126-4. xiv + 553 pp. LCCN QE705.F8 L38 1987. 350F.

Law:1965:BRT

[Law65] R. J. Law. Book review:

Thomas Newcomen. The Prehistory of the Steam Engine. By L. T. C. Rolt. Pp. 158, plates and illustrations. London: David and Charles, 1963. 25s. British Journal for the History of Science, 2(3):264– 265, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024945.

Law:1980:BRS

[Law80a] John Law. Book review: Sociology of Science Toward a Metric of Science: The Advent of Science Indicators. Edited by Yehuda Elkana, Joshua Lederberg, Robert K. Merton, Arnold Thackray, and Harriet Zuckerman. New York and Chichester: Wiley, 1978. Pp. xiv + 354. £14.00. British Journal for the History of Science, 13 (3):264, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026202.

Lawrence:1980:BRE

[Law80b] C. J. Lawrence. Book review: Eighteenth Century: The Philosophy of Medicine: The Early Eighteenth Century. By Lester S. King. Cambridge, Mass. & London: Harvard University Press, 1978. Pp. viii + 291. £12.25. British Journal for the History of Science, 13(1): 80–81, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026135.

Lawrence:1980:SNB

[Law80c] C. J. Lawrence. Short notices of books: Antique Medical Instruments. By Elisabeth Bennion. London: Sotheby Parke Bernet; Berkeley & Los Angeles: University of California Press, 1979. Pp. xii + 355. 16 colour plates. £28. British Journal for the History of Science, 13 (2):181–182, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025914.

Lawrence:1981:BRM

[Law81] Christopher Lawrence. Book review: Medicine The Genesis of Cancer: a Study in the History of Ideas. By L. J. Rather. Baltimore & London: Johns Hopkins University Press, 1978, pp. xiv + 262. £12.25. British Journal for the History of Science, 14 (2):210–211, July 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025953.

Lawrence:1983:BRM

[Law83] Christopher Lawrence. Book review: Medical Sciences: Georges Canguilhem, On the normal and the pathological. Trans, by Carolyn R. Fawcett. Dordrecht, Boston, & London: D. Reidel, 1978. Pp. xxvi + 208. Dfl60./$29.00; Dfl30./$14.50 (paperback). F. Kräupl Taylor, The concepts of illness, disease and morbus. Cambridge: Cambridge University Press, 1979. Pp. x + 131. £6.50. British Journal for the History of Science, 16 (1):95–96, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026104.

Lawrence:1985:BRP

[Law85] Christopher Lawrence. Book reviews: Peter Wright and Andrew Treacher (eds), The problem of medical knowledge. Edinburgh University Press, 1982. Pp. viii + 232. ISBN 0-85224-417-7. £12.00 (paperback). David Armstrong, The political anatomy of the body. Cambridge: Cambridge University Press, 1983. Pp. xii + 146. ISBN 0-521-24746-2. £17.50. British Journal for the History of Science, 18(1): 90–91, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026274.

Lawrence:1987:BRR

[Law87] Christopher Lawrence. Book review: Roger Cooter. The Cultural Meaning of Popular Science: Phrenology and the Organisation of Consent in Nineteenth Century Britain. Cambridge: Cambridge University Press, 1985. Pp. xiv + 418. ISBN 0-521-22743-7. £25.00. British Journal for the History of Science, 20(1): 94–96, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026449.

Lawrence:1988:BRR

[Law88] Christopher Lawrence. Book review: Ruth Richardson. Death, Dissection and the Destitute. London and New York: Routledge and Kegan Paul, 1987. Pp. xviii + 426. ISBN 0-7102-0919-3. £19.95. British Journal for the History of Science, 21 (3):385, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026855.

Lawrence:1989:BRW

[Law89] Christopher Lawrence. Book review: W. J. O’Connor. Founders of British Physiology: A Biographical Dictionary, 1820–1885. Manchester: MUP, 1988. Pp. 264. ISBN 0-7190-2537-0. £27.50. British Journal for the History of Science, 22(1):111, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026708.

Lawn:1993:RDS

[Law93] Brian Lawn. The rise and decline of the scholastic Quaestio disputata: with special emphasis on its use in the teaching of medicine and science, volume 2 of Education and society in the Middle Ages and Renaissance, 0926-6070. New York, Leiden, The Netherlands, 1993. ISBN 90-04-09740-6 (hardcover). 176 pp. LCCN AZ321 .L38 1993.

Layton:1965:DDD

[Lay65] David Layton. Diction and dictionaries in the diffusion of scientific knowledge: an aspect of the history of the popularization of science in Great Britain. British Journal for the History of Science, 2(3): 221–234, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024936.

Layton:1974:BRG

[Lay74] David Layton. Book review: General History in the Teaching of Physics. Ed. by Stephen G. Brush and Allen L. King. Hanover, New Hampshire: University Press of New England, 1972. Pp. xii + 116. No price stated. Resources for the History of Physics. Ed. by Stephen G. Brush. Hanover, New Hampshire: University Press of New England, 1972. Pp. 176. $3.50. British Journal for the History of Science, 7(2):184–185, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025209.

Layton:1985:BRR

[Lay85] David Layton. Book review: Roy MacLeod (editor), Days of Judgement. Science, Examinations and the Organization of Knowledge in Late Victorian England. Driffield: Studies in Education Ltd., Nafferton Books, 1982. Pp. ix + 250. ISBN 0-905484-15-0. £15.95. British Journal for the History of Science, 18(1): 113, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026293.

Lefebvre:1973:ORJ

[LB73] E. Lefebvre and J. G. De Bruijn. Obituary: Professor R. J. Forbes (12 April 1900–13 January 1973). British Journal for the History of Science, 6(4):462, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025527.

LeFanu:1990:NGS

[Le 90] W. R. (William Richard) Le Fanu. Nehemiah Grew: a study and bibliography of his writings. St Paul’s Bibliographies, Winchester, UK, 1990. ISBN 0-906795-43-5. xvii + 182 pp. LCCN Z8369.262 .L4 1990.

Leary:1990:MHP

[Lea90]  David E. Leary,  ed-

itor. Metaphors in the history of psychology. Cambridge studies in the history of psychology, 99-1209266-X. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37166-X, 0-521-42152-7. xiii + 383 pp. LCCN ????

Leary:1994:MHP

[Lea94]  David  E.

Leary, editor. Metaphors in the history of psychology. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-42152-7 (paperback), 0-521-37166-X. xiii + 383 pp. LCCN BF38 .M482

Leff:1978:BRM

[Lef78] Gordon Leff. Book review: Middle Ages Science and Creation in the Middle Ages: Henry of Langenstein (d 1397) on Genesis. By Nicholas H. Steneck. Notre Dame & London: University of Notre Dame Press, 1976. Pp. 231. £9.75. British Journal for the History of Science, 11(3):285– 287, November 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025786.

LeGrand:1988:DCS

[LeG88] H. E. (Homer Eugene) LeGrand. Drifting continents and shifting theories: the modern revolution in geology and scientific change. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-32210-3, 0-521-31105-5 (paperback). vi + 313 pp. LCCN QE511.5 .L44

1988. URL http://www.loc.
gov/catdir/description/
cam023/88030254.html;
http://www.loc.gov/
catdir/toc/cam028/
88030254.html.

Lemay:1977:STC

[Lem77] Helen Rodnite Lemay. Science and theology at Chartres: The case of the supracelestial waters. British Journal for the History of Science, 10(3): 226–236, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025842.

Lenoir:1982:SLT

[Len82] Timothy Lenoir. The strategy of life: teleology and mechanics in Nineteenth Century German biology, volume 13 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1982. ISBN 90-277-1363-4. xii + 314 pp. LCCN QH305.2.G3 L46 1982.

Lervig:1985:SCS

[Ler85] Philip Lervig. Sadi Carnot and the steam engine: Nicolas Clément’s lectures on industrial chemistry 1823–1828. British Journal for the History of Science, 18(2):147–196, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026329.

Lesch:1983:BRL

[Les83] John E. Lesch. Book review: Life Sciences: Shirley A. Roe, Matter, life, and generation: 18th century embryology and the Haller-Wolff debate. Cambridge: Cambridge University Press, 1981. Pp. x + 214. £16.00. British Journal for the History of Science, 16(2):216–218, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026257.

Levene:1965:TIK

[Lev65] John R. Levene. The true inventors of the keratoscope and photokeratoscope. British Journal for the History of Science, 2(4): 324–342, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024889.

Levere:1968:FMN

[Lev68] T. H. Levere. Faraday, matter, and natural theologyreflections on an unpublished manuscript. British Journal for the History of Science, 4(2): 95–107, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025141.

Levene:1973:BRM

[Lev73]  Arnold Levene.  Book re-

view: Medicine Hunter’s Lectures of Anatomy. Foreword by Sir Charles Illingworth. Introduction by N. Dowd. London and New York: American Elsevier, 1972. Pp. 320. $11. British Journal for the History of Science, 6(3):315–316, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025450.

Levere:1980:BRN

[Lev80] Trevor H. Levere. Book review: Nineteenth Century: The Transcendental Part of Chemistry.

By David M. Knight. Folkestone: Dawson, 1978. Pp. viii + 289. £12.00. British Journal for the History of Science, 13(2):170–171, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025904.

Levere:1984:DTB

[Lev84] Trevor H. Levere. Dr. Thomas Beddoes (1750–1808): Science and medicine in politics and society. British Journal for the History of Science, 17(2): 187–204, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026551.

Levi:1985:GHL

[Lev85] Hilde Levi. George de Hevesy:

life and work: a biography. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-555-9. 147 pp. LCCN QD22.H45 L48 1985.

Levere:1988:HSC

[Lev88a] Trevor H. Levere. The history of science of Canada. British Journal for the History of Science, 21(4):419– 425, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026961.

Levere:1988:VSC

[Lev88b] Trevor H. Levere. Vilhjalmur Stefansson, the Continental Shelf, and a new Arctic continent. British Journal for the History of Science, 21(2): 233–247, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026981.

Levy:1990:PPP

[Lev90] B. Barry Levy. Planets, potions, and parchments: scientifica Hebraica from the Dead Sea Scrolls to the Eighteenth Century. Published for the Jewish Public Library by McGill-Queen’s University Press, Montréal, Québec, Canada, 1990. ISBN 0-7735-0791-4 (paperback), 0-7735-0793-0 (hardcover). xii + 139 pp. LCCN Q128 .L48 1990.

Levere:1993:SCA

[Lev93] Trevor Harvey Levere. Science and the Canadian Arctic: a century of exploration, 1818–1918. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-41933-6. xiv + 438 pp. LCCN Q115 .L554

1993. URL http://www.loc.
gov/catdir/description/
cam025/92012570.html;
http://www.loc.gov/
catdir/samples/cam031/
92012570.html;   http://
www.loc.gov/catdir/toc/
cam029/92012570.html.

Lewis:1985:BRG

[Lew85] Gilbert Lewis. Book review: George W. Stocking Jr., (ed.), Observers Observed: Essays on Ethnographic Fieldwork. Madison: The University of Wisconsin Press, 1984. Pp. 242. ISBN 0-299-09450-2. $19.95. British Journal for the History of Science, 18(2):245–246, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026345.

Lavoisier:1955:ULC

[LF97] Antoine Laurent Lavoisier and René Fric. Œuvres de Lavoisier: correspondance. (French) [Works of Lavoisier: correspondence]. A. Michel, Paris, France, 1955– 1997. ISBN 2-7011-1085-8 (vol. 4), 2-85367-184-4 (vol. 5), 2-85367-205-0 (vol. 6). ???? pp. LCCN QD22.L4 A4 1997.

Leibniz:1976:MNT

[LHB04] Gottfried Wilhelm Leibniz, Heinz-Jürgen Hess, and Herbert Breger. Mathematischer naturwissenschaftlicher und technischer Briefwechsel. (German) [Mathematical, scientific and technical letters], volume III of Sämtliche Schriften und Briefe. Akademie-Verlag, Berlin, Germany, 1976–2004. ISBN 3-05-000075-9 (set), 3-05-000235-2 (vol. 2), 3-05-000766-4 (vol. 3), 3-05-002602-2 (vol. 4), 3-05-003471-8 (vol. 5), 3-05-004116-1 (vol. 6). ???? pp. LCCN ????

Leibniz:1991:MNT

[LHH91] Gottfried Wilhelm Leibniz, Heinz-Jürgen Hess, and J. E. Hofmann. Mathematischer, naturwissenschaftlicher und technischer Briefwechsel. 3, 1680–Juni 1683. (German) [Mathematical, scientific and technical letters. 3, 1680–June 1683], volume 3 of Sämtliche Schriften und Briefe / Gottfried Wilhelm Leibniz; Reihe 3. Akademie-Verlag, Berlin, Germany, 1991. ISBN 3-05-000766-4. lix + 895 pp. LCCN ????

Liebenau:1985:BRR

[Lie85] Jonathan Liebenau. Book review: Ronald Ferrier, The History of British Petroleum Company, Volume 1: The developing years, 1901–1932. Cambridge: Cambridge University Press, 1982. Pp. xxx + 801. ISBN 0-521-24647-4. £35.00. British Journal for the History of Science, 18(3): 371–372, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026407.

Liebenau:1986:BRM

[Lie86] Jonathan Liebenau. Book review: Michael Bliss. The Discovery of Insulin. Edinburgh: Paul Harris, 1983. Pp. 304. ISBN 0-86228-056-7. £15.00. British Journal for the History of Science, 19(1):117– 118, March 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026490.

Liebenau:1987:MSM

[Lie87] Jonathan Liebenau. Medical science and medical industry: the formation of the American pharmaceutical industry. Studies in business history. MacMillan Press, Basingstoke, Hampshire, 1987. ISBN 0-333-41742-9 (hardcover), 0-8018-3356-6. ix + 207 pp. LCCN HD9666.5 .L54.

Liebenau:1988:CNT

[Lie88] Jonathan Liebenau, editor. The challenge of new technology: innovation in British business since 1850. Business history series. Gower, Aldershot, Hants, England, 1988. ISBN 0-566-05147-8. x + 155 pp. LCCN HC260.T4 C46 1988. US$60.00 (U.S. : est.).

Lightman:1987:OAV

[Lig87] Bernard V. Lightman. The origins of agnosticism: Victorian unbelief and the limits of knowledge. The Johns Hopkins University Press, Baltimore, MD, USA, 1987. ISBN 0-8018-3375-2. x + 249 pp. LCCN BL2747.2 .L54 1987.

Lilienthal:1985:LVI

[Lil85] Georg Lilienthal. Der “Lebensborn e.V.”: ein Instrument nationalsozialistischer Rassenpolitik. (German) [The life font: an instrument of National Socialist racial policy], volume 1 of Forschungen zur neueren Medizin- und Biologiegeschichte. G. Fischer, Stuttgart, Germany, 1985. ISBN 3-437-10939-1. ISSN 0176-330X. vii + 264 pp. LCCN D810.C4 L48 1985.

Lindberg:1968:CRM

[Lin68]  David C. Lindberg.  The

cause of refraction in medieval optics. British Journal for the History of Science, 4(1): 23–38, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024984.

Lindenbaum:1974:BRP

[Lin74] R. H. Lindenbaum. Book review: Philosophy The Philosophy of Biology. By Michael Ruse. London: Hutchinson, 1973 PP 231 £3; £1.85 (paperback). British Journal for the History of Science, 7(3): 286, November 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025535.

Lindberg:1982:AMN

[Lin82] David C. Lindberg. On the applicability of mathematics to nature: Roger Bacon and his predecessors. British Journal for the History of Science, 15 (1):3–25, March 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025921.

Lindqvist:1984:TTI

[Lin84] Svante Lindqvist. Technology on trial: the introduction of steam power technology into Sweden, 1715–1736, volume 1 of Uppsala studies in history of science, 0282-1036: 1. Almqvist och Wiksell, Stockholm, Sweden, 1984. ISBN 91-86836-00-5 (paperback), 91-22-00716-4 (hardcover). 392 pp. LCCN ????

Lindberg:1986:BRB

[Lin86] David C. Lindberg. Book review: Bert Hansen. Nicole Oresme and the Marvels of Nature: A Study of his De causis mirabilium with Critical Edition, Translation, and Commentary. Toronto: Pontifical Institute of Mediaeval Studies, 1985. Pp. xi + 478. ISBN 0-88844-068-5. $39.00. British Journal for the History of Science, 19(3):342– 343, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026512.

Lindberg:1989:BM

[Lin89a] David C. Lindberg. Back matter. British Journal for the History of Science, 22(1): 128, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026720.

Lindberg:1989:BRJ

[Lin89b] David C. Lindberg. Book review: J. V. Field. Kepler’s Geometrical Cosmology. Chicago: University of Chicago Press, 1988. Pp. xx + 243. ISBN 0-226-24823-2. £32. British Journal for the History of Science, 22(1):124, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026718.

Lindgren:1990:GFD

[Lin90] Michael Lindgren. Glory and failure: the difference engines of Johann Müller, Charles Babbage and Georg and Edvard Scheutz. The MIT Press series in the history of computing. MIT Press, Cambridge, MA, USA, [2nd] edition, 1990. ISBN 0-262-12146-8. 414 pp. LCCN QA75 .L493 1990.

Lindberg:1992:BWS

[Lin92] David C. Lindberg. The beginnings of Western science: the European scientific tradition in philosophical, religious, and institutional context, 600 B.C. to A.D. 1450. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-48230-8 (hardcover), 0-226-48231-6 (paperback). xviii + 455 pp. LCCN Q124.95 .L55

1992. URL http://www.loc.
gov/catdir/description/
uchi051/91037741.html;
http://www.loc.gov/

catdir/enhancements/
fy0608/91037741-t.html.

Lindqvist:1993:CPH

[Lin93] Svante Lindqvist, editor. Center on the periphery: historical aspects of 20th-Century Swedish physics, volume 17 of Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1993. ISBN 0-88135-157-1. lv + 516 pp. LCCN QC9.S8 C46 1993. US$50.00.

Lindemann:1996:HHE

[Lin96] Mary Lindemann. Health and healing in Eighteenth-Century Germany, volume 114th series, 4 of The Johns Hopkins University studies in historical and political science. The Johns Hopkins University Press, Baltimore, MD, USA, 1996. ISBN 0-8018-5281-1 (hardcover). xiii + 505 pp. LCCN RA418.3.G3

L54 1996. URL http://www.
loc.gov/catdir/bios/
jhu051/95053137.html;
http://www.loc.gov/
catdir/description/
jhu051/95053137.html.

Livesey:1985:WOS

[Liv85] Steven J. Livesey. William of Ockham, the subalternate sciences, and Aristotle’s theory of metabasis. British Journal for the History of Science, 18 (2):127–145, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026328.

Livingstone:1986:BRJ

[Liv86] David N. Livingstone. Book review: John Durant (ed.). Darwinism and Divinity. Essays on Evolution and Religious Belief. Oxford: Basil Blackwell, 1985. Pp. ix + 210. ISBN 0-631-14188-X. £15.00. British Journal for the History of Science, 19(3):352–353, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026524.

Livingstone:1987:NSS

[Liv87]  David  N.

Livingstone. Nathaniel Southgate Shaler and the culture of American science. History of American science and technology series. University of Alabama Press, Tuscaloosa, AL, USA, 1987. ISBN 0-8173-0305-7. xiv + 395 pp. LCCN G69.S49 L58 1987.

Livingstone:1993:GTE

[Liv93] David N. Livingstone. The geographical tradition: episodes in the history of a contested enterprise. Blackwell, Oxford, UK, 1993. ISBN 0-631-18535-6, 0-631-18586-0 (paperback). viii + 434 pp. LCCN G80 .L54 1993.

Leddy:2009:EE

[LL09] Neven Leddy and Avi Lifschitz, editors. Epicurus in the Enlightenment, volume 2009, 12 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2009. ISBN 0-7294-0987-2 (paperback). vii + 258 pp. LCCN B512 .E68 2009.

Lloyd:1963:CS

[Llo63] C. C. Lloyd. The conquest of scurvy. British Journal for the History of Science, 1(4): 357–363, December 1963. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024907.

Lloyd:1983:SFI

[Llo83]  G.  E.  R.  (Ge-

offrey Ernest Richard) Lloyd. Science, folklore, and ideology: studies in the life sciences in ancient Greece. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25314-4, 0-521-27307-2 (paperback). xi + 260 pp. LCCN R723 .L58

1983. URL http://www.loc.
gov/catdir/description/
cam022/82019808.html.

Lloyd:1991:MPG

[Llo91]  G.  E.  R.  (Ge-

offrey Ernest Richard) Lloyd. Methods and problems in Greek science. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-37419-7. xiv + 457 pp. LCCN Q127.G7 L593

1991. URL http://www.loc.
gov/catdir/description/
cam024/90001539.html;
http://www.loc.gov/
catdir/toc/cam023/
90001539.html.

Leibniz:1923:SSB

[LM23] Gottfried Wilhelm Leibniz and Uwe Bearb Mayer. Sämtliche Schriften und Briefe. Bd. 5. 1674–1676. (German) [Collected works and letters. Volume 5, 1674–1676]. Akad.-Verl., Berlin, Germany, 1923. ISBN 3-05-004578-7, 3-05-000075-9. xxxviii + 664 pp. LCCN ????

Lorch:1974:BRM

[LM74] R. P. Lorch and J. O. Marsh. Book review: Mathematics: Bibliography and Research Manual of the History of Mathematics. By Kenneth O. May. Toronto and Buffalo: University of Toronto Press, 1973. Pp. vii + 818. $20.00. British Journal for the History of Science, 7(2):186–187, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025211.

Lindberg:1986:GNH

[LN86]  David  C.

Lindberg and Ronald L. Numbers, editors. God and nature: historical essays on the encounter between Christianity and science. University of California Press, Berkeley, CA, USA, 1986. ISBN 0-520-05538-1, 0-520-05692-2 (paperback). LCCN BL245 .G63 1986.

LopezPinero:1995:HCA

[LN95] José María López Piñero and Víctor Navarro Brotóns.

Història de la ciència al País Valencià. (Spanish) [History of science in Valencia County]. Edicions Alfons el Magnànim, València, Spain, 1995. ISBN 84-7822-154-9. 661 + 38 pp. LCCN Q127.S7 L67 1995.

Loeffel:1987:BP

[Loe87] Hans Loeffel. Blaise Pascal, 1623–1662, volume 2 of Vita mathematica. Birkhäuser Verlag, Basel, Switzerland, 1987. ISBN 3-7643-1840-6. 175 pp. LCCN QA29.P2 L64 1987.

Lohne:1975:BRR

[Loh75] J. A. Lohne. Book review:

Renaissance: Thomas Harriot, Renaissance Scientist. Ed. by J. W. Shirley. Oxford: Clarendon Press: Oxford University Press, 1974. Pp. x + 181. £6.50. British Journal for the History of Science, 8(2):183, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025652.

Longrigg:1988:AAT

[Lon88] James Longrigg. Anatomy in Alexandria in the Third Century B.C. British Journal for the History of Science, 21(4): 455–488, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026964.

Lorch:1973:BRA

[Lor73] R. P. Lorch. Book review: Arab Science The Astrological History of Masha’allah. Ed. by E. S. Kennedy and David Pingree. Cambridge, Mass.: Harvard University Press, and London: Oxford University Press, 1971. Pp.xvi + 206. £4.75. British Journal for the History of Science, 6(4):438–439, December 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025512.

Lorch:1975:BRM

[Lor75]  Richard Lorch.  Book re-

view: Middle Ages Myth and Science in the Twelfth Century. A Study of Bernard Silvester. By Brian Stock. Princeton, N.J.: Princeton University Press, 1972. Pp. xvi + 331. £6.25. British Journal for the History of Science, 8 (2):180–181, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025649.

Lorch:1977:BRA

[Lor77]  Richard Lorch.  Book re-

view: Antiquity Probability and Statistical Inference in Ancient and Medieval Jewish Literature. By Nachum L. Rabinovitch. Toronto and Buffalo: University of Toronto Press, 1973. Pp. xiii + 205. $12.50. British Journal for the History of Science, 10(2): 170, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025877.

Love:1975:RDM

[Lov75] Rosaleen Love. Revisions of Descartes’s matter theory in Le Monde. British Journal for the History of Science, 8 (2):127–137, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025637.

Lowy:1990:PSP

[Löw90] Ilana Löwy, editor. The Polish School of Philosophy of Medicine: from Tytus Chalubinski (1820–1889) to Ludwik Fleck (1896–1961), volume 37 of Philosophy and medicine. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0958-8. LCCN R536 .L69 1990. URL http:

//catalog.hathitrust.org/
api/volumes/oclc/
22508739.html.

Lusebrink:2004:ERP

[LP04] Hans-Jürgen Lüsebrink and Jeremy D. Popkin, editors. Enlightenment, revolution, and the periodical press, volume 2004:06 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2004. ISBN 0-7294-0841-8. v + 226 pp. LCCN PN4808 .E55 2004.

Lyell:1990:PG

[LR91] Sir Charles Lyell and M. J. S. Rudwick. Principles of geology. University of Chicago Press, Chicago, IL, USA, 1990– 1991. ISBN 0-226-49793-3 (hardcover : vol. 1), 0-226-49794-1 (paperback : vol. 1), 0-226-49796-8 (hardcover : vol. 2), 0-226-49797-6 (paperback : vol. 2), 0-226-49798-4 (hardcover : vol. 3), 0-226-49799-2 (paperback : vol. 3). ???? pp. LCCN QE26 .L956

                                        1990. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0608/90011008-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0634/90011008-t.html.

Leibniz:1976:SSB

[LS76] Gottfried Wilhelm Leibniz and Walter Salvator. Sämtliche Schriften und Briefe: dritte Reihe: mathematischer, naturwissenschaftlicher und technischer Briefwechsel. (German) [Collected writings and correspondence: third series: mathematical, scientific, and technical letters]. Akademie-Verlag, Berlin, Germany, 1976. ISBN 3-05-000075-9 (set), 3-05-000235-2 (Vol. 2), 3-05-000766-4 (Vol. 3), 3-05-002602-2 (Vol. 4), 3-05-003471-8 (Vol. 5), 3-05-004116-1 (Vol. 6). ???? pp. LCCN ????

Levere:1990:NES

[LS90] Trevor Harvey Levere and William R. Shea, editors. Nature, Experiment, and the Sciences: Essays on Galileo and the History of Science in Honour of Stillman Drake, volume 120 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346. xiii + 349 pp. LCCN Q174 .B67

vol. 120. URL https://link.
springer.com/book/10.
1007/978-94-009-1878-8.

Lyell:1997:PG

[LS97] Sir Charles Lyell and James A.

Secord. Principles of geology. Penguin classics. Penguin Books, Harmondsworth, UK, 1997. ISBN 0-14-043528-X. xlvii + 471 pp. LCCN QE26.2 .L94 1997.

Lambourne:1990:CES

[LSS90] Robert J. Lambourne, Michael Shallis, and Michael Shortland. Close encounters?: science and science fiction. Adam Hilger Ltd., Bristol, UK, 1990. ISBN 0-85274-141-3. xiii + 184 pp. LCCN PN3433.6 .L36

1990. URL http://www.loc.
gov/catdir/enhancements/
fy0707/90033934-d.html.

Luckin:1990:QPE

[Luc90]  Bill Luckin.  Questions of

power: electricity and environment in inter-war Britain. Manchester University Press, Manchester, UK, 1990. ISBN 0-7190-3302-0 (hardcover). viii + 200 pp. LCCN HD9685.G7 L83 1990.

Ludmerer:1980:BRM

[Lud80] Kenneth M. Ludmerer. Book review: Medicine at Harvard: the First 300 Years. By Henry K. Beecher and Mark D. Altschule. Hanover, New Hampshire: University Press of New England. 1977. Pp. xv + 587. $27.50. British Journal for the History of Science, 13 (1):64–65, March 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026125.

Lugg:1979:KPS

[Lug79] Andrew Lugg. Kuhn and the philosophy of science: Theories of science. British Journal for the History of Science, 12(3):289–295, November 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026003.

Lunbeck:1994:PPK

[Lun94] Elizabeth Lunbeck. The psychiatric persuasion: knowledge, gender, and power in modern America. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-04804-5. xiv + 431 pp. LCCN RC437.5 .L89 1994. US$29.95. URL http:

                                        //www.loc.gov/catdir/
                                        description/prin031/
                                        93043818.html;   http://

www.loc.gov/catdir/toc/
prin031/93043818.html.

Lindberg:1990:RSR

[LW90]  David  C.

Lindberg and Robert S. Westman, editors. Reappraisals of the scientific revolution. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-34262-7, 0-521-34804-8 (paperback). xxvii + 551 pp. LCCN Q125 .R34 1990. URL http:

//www.loc.gov/catdir/
description/cam041/
89070859.html;   http://
www.loc.gov/catdir/toc/
cam041/89070859.html.

Lay-Yong:1966:COG

[LY66] Lam Lay-Yong. On the Chinese origin of the Galley Method of Arithmetical Division. British Journal for the History of Science, 3(1):66–69, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025103.

Millar:1989:CCD

[M^{+}89] David Millar et al., editors.

Chambers concise dictionary of scientists. Chambers, Cambridge, UK, 1989. ISBN 1-85296-354-9, 0-550-16032-9. 461 pp. LCCN Q141 .M54

1989. URL http://catalog.
hathitrust.org/api/
volumes/oclc/20820593.
html.

MacLeod:1972:BRE

[Mac72] R. M. MacLeod. Book review: Eighteenth and Nineteenth Centuries: Dollars for Research: Science and its Patrons in Nineteenth-Century America. By Howard S. Miller. Seattle and London: University of Washington Press, 1970. Pp. xi + 258. $9.50. British Journal for the History of Science, 6(1):97–99, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025280.

MacKenzie:1978:BRN

[Mac78] Donald MacKenzie. Book review: Nineteenth and Twentieth Centuries: Eugenics and Politics in Britain, 1900–1914. By G. R. Searle. Leyden: Noordhoff International Publishing, 1976. Pp. vii + 147. Dfl.

45. British Journal for the History of Science, 11(1):89– 91, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025628.

Macintyre:1979:BRN

[Mac79] Sally Macintyre. Book review: Nineteenth and Twentieth Centuries: Health Care and Popular Medicine in Nineteenth-Century England. Edited by John Woodward and David Richards. London: Croom Helm, 1977. Pp. V + 195. £7.95. British Journal for the History of Science, 12(1): 98–99, March 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025672.

MacKenzie:1980:BRN

[Mac80] Donald MacKenzie. Book review: Nineteenth and Twentieth Centuries: The Triumph of Evolution. American Scientists and the Heredity-Environment Controversy, 1900–1941. By Hamilton Cravens. Philadelphia: University of Pennsylvania Press, 1978. Pp. xxiv + 351. $17.50. British Journal for the History of Science, 13(3): 274–275, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026210.

MacLeod:1982:DJS

[Mac82] Roy M. MacLeod, editor. Days of judgement: science, examinations, and the organization of knowledge in late Victorian England. Studies in Education, Driffield, N. Humberside, UK, 1982. ISBN 0-905484-15-0. LCCN LB3056.G7 D34 1982. US$15.95.

MacNeill:1983:PPH

[Mac83] William Hardy MacNeill. The pursuit of power: technology, armed force, and society since A.D. 1000. Blackwell, Oxford, UK, 1983. ISBN 0-631-13134-5. x + 405 pp. LCCN ????

MacHale:1985:GBH

[Mac85a] Des MacHale. George Boole: his life and work, volume 2 of Profiles of genius series, 0332-3218. Boole Press, Dublin, Ireland, 1985. ISBN 0-906783-05-4, 0-906783-14-3 (paperback). xiii + 304 pp. LCCN QA29.B685 M33 1985. US$16.00.

MacKinnon:1985:BRB

[Mac85b] Edward MacKinnon. Book review: Bruce R. Wheaton, The Tiger and the Shark: Empirical Roots of wave-particle Dualism. Cambridge: Cambridge University Press, 1983. Pp. xxiv + 355. ISBN 0-521-25098-6. £22.50, $39.50. With a Foreword by Thomas S. Kuhn. British Journal for the History of Science, 18(3):347– 348, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026388.

Mack:1988:BRR

[Mac88a] Pamela E. Mack. Book review: R. Cargill Hall (ed.). History of Rocketry and Astronautics: Proceedings of the Third through the Sixth History Symposia of the International Academy of Astronautics. American Astronautical Society History Series, Volume 7, parts I and II; International Academy of Astronautics History Symposia, Volume 2, parts I and II. San Diego: American Astronautical Society, 1986 (first published as NASA Conference Publication 2014 in 1977). Two volumes: xii + 238 pages and xii + 489 pages. British Journal for the History of Science, 21 (1):127–128, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026883.

MacKenzie:1988:ENH

[Mac88b] John M. MacKenzie. The empire of nature: hunting, conservation, and British imperialism. Studies in imperialism. Manchester University Press, Manchester, UK, 1988. ISBN 0-7190-2227-4. x + 340 pp. LCCN SK21 .M33 1988. US$35.00 (U.S. : est.).

MacLeod:1988:IIR

[Mac88c] Christine MacLeod. Inventing the Industrial Revolution: the English patent system, 1660– 1800. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-30104-1. xii + 302 pp. LCCN T257.P2 M33

1988. URL http://www.loc.
gov/catdir/description/
cam023/87037406.html;
http://www.loc.gov/
catdir/samples/cam031/
87037406.html;   http://
www.loc.gov/catdir/toc/
cam021/87037406.html.

Macleod:1989:BRP

[Mac89] Roy M. Macleod. Book review:

Peter Alter. The Reluctant Patron: Science and the State in Britain, 1850–1920. Oxford: Berg, 1987. ISBN 0-907582-67-2. £30. British Journal for the History of Science, 22(4): 482–485, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026951.

Mack:1990:VES

[Mac90] Pamela Etter Mack. Viewing the Earth: the social construction of the Landsat satellite system. Inside technology. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-13259-1. xii + 270 + 1 pp. LCCN TL796.5.U6 L375 1990.

Maddin:1988:BUM

[Mad88] Robert Maddin, editor. The beginning of the use of metals and alloys: papers from the Second International Conference on the Beginning of the Use of Metals and Alloys, Zhengzhou, China, 21–26 October 1986. MIT Press, Cambridge, MA, USA, 1988. ISBN 0-262-13232-X. LCCN TN16 .I58 1986.

Maffeo:1991:SNP

[Maf91] Sabino Maffeo. In the service of nine popes: 100 years of the Vatican Obervatory. Pontificia Academia Scientiarum, Vatican City, 1991. ISBN 88-7761-046-8. 241 pp. LCCN QB82.V32 V38513 1991.

Mahoney:1989:BRH

[Mah89] Michael S. Mahoney. Book review: Hans Loeffel. Blaise Pascal 1623–1662. (Vita Mathematica, Bd. 2) Basel, Boston, Stuttgart: Birkhäuser Verlag, 1987. Pp. 173. ISBN 3-7643-1840-6. SFr. 40. A. W. F. Edwards. Pascal’s Arithmetical Triangle. London: Charles Griffin & Company Ltd, New York: Oxford University Press, 1987. Pp. xii + 174. ISBN 0-85264-283-0. No price given. British Journal for the History of Science, 22(4):479– 481, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026949.

Maienschen:1985:BRG

[Mai85] Jane Maienschen. Book review:

Georg Uschmann, editor, Ernst Haeckel. Biographie in Briefen. Leipzig: Urania-Verlag, 1983. Pp. 328. 24DM. British Journal for the History of Science, 18 (1):115–116, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026295.

Mainzer:1988:SNH

[Mai88] Klaus Mainzer. Symmetrien der Natur: ein Handbuch zur Natur- und Wissenschaftsphilosophie. (German) [Symmetry of Nature: a Handbook for Natural and Scientific Philosophy]. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 1988. ISBN 3-11-011507-7. xi + 739 pp. LCCN Q172.5.S95 M3

                                        1988. URL http://catalog.
                                        hathitrust.org/api/
                                        volumes/oclc/22415725.
                                        html.

Maienschein:1991:TTA

[Mai91] Jane Maienschein. Transforming traditions in American biology, 1880–1915. The Johns Hopkins University Press, Baltimore, MD, USA, 1991. ISBN 0-8018-4126-7. xi + 366 pp. LCCN QH305.2.U6 M35 1991.

Malet:1995:FSB

[Mal95] Antoni Malet. Ferran Sunyer i Balaguer, 1912–1967. (Spanish) [Ferran Sunyer in Balaguer, 1912–1967]. Societat Catalana de Matemátiques, Barcelona, Spain, 1995. ISBN 84-7283-286-4. 306 pp. LCCN QA29.S793 M35 1995.

Mallinson:2005:ECN

[Mal05]  G. J. Mallinson,  ed-

itor. The Eighteenth Century now: boundaries and perspectives, volume 2005:10 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2005. ISBN 0-7294-0866-3 (paperback). ix + 348 pp. LCCN PQ261 .E44 2005.

Manning:1988:UCS

[Man88] Thomas G. Manning. U.S. Coast Survey vs. Naval Hydrographic Office: a 19th-Century rivalry in science and politics. History of American science and technology series. University of Alabama Press, Tuscaloosa, AL, USA, 1988. ISBN 0-8173-0390-1. xii + 202 pp. LCCN VK597.U6 M36 1988.

Marland:1962:BAC

[Mar62a] E. A. Marland. British and American contributions to electrical communications. British Journal for the History of Science, 1(1):31–48, June 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025073.

Martin:1962:ORI

[Mar62b] Thomas Martin. Origins of the Royal Institution. British Journal for the History of Science, 1(1):49–63, June 1962. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025074.

Martin:1963:RO

[Mar63] Thomas Martin. Retirement of officers. British Journal for the History of Science, 1(4):396, December 1963. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Martin:1964:PAE

[Mar64] Thomas Martin. Presidential address: Early years at the Royal Institution. British Journal for the History of Science, 2(2):99–115, December 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025010.

Marshall:1973:BRS

[Mar73] J. D. Marshall. Book review: Science and Society Science, Technology, and Economic Growth in the Eighteenth Century. Ed. A. E. Musson. London: Methuen, 1972. Pp. x + 211. £2.25. British Journal for the History of Science, 6(4):437–438, December 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025510.

Marsh:1974:BRN

[Mar74] J. O. Marsh. Book review: Nineteenth and Twentieth Centuries: The Ethereal Aether. A History of Michelson-Morley-Miller Aether-Drift Experiments, 1880–1930. By Lloyd S. Swenson Jr Austin: University of Texas Press, 1972. Pp. xxii +361. £4.75. Nineteenth Century Aether Theories. By Kenneth F. Schaffner. Oxford: Pergamon Press, 1972. Pp. x + 278. £3.25. British Journal for the History of Science, 7(1):96–97, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025193.

Martin:1989:BRW

[Mar89]  Julian Martin.  Book re-

view: Whitney R. D. Jones, William Turner: Tudor Naturalist, Physician and Divine. London and New York: Routledge, 1988. Pp. iii + 223. ISBN 0-415-00359-8. £35.00. British Journal for the History of Science, 22(2):249–250, July 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026668.

Martin:1992:FBS

[Mar92] Julian Martin. Francis Bacon, the state and the reform of natural philosophy. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-38249-1 (hardcover). xiii + 236 pp. LCCN B1198 .M36

1992. URL http://www.loc.
gov/catdir/description/
cam024/90024242.html;
http://www.loc.gov/
catdir/toc/cam028/
90024242.html.

MarquinaBarrio:2010:GWC

[Mar10] Antonio Marquina Barrio, editor. Global warming and climate change: prospects and policies in Asia and Europe. Energy, climate and the environment series. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2010. ISBN 0-230-23771-1 (hardcover). xvii + 510 pp. LCCN QC903.2.E85 G56 2010.

Mason:1998:DDB

[Mas98] Haydn Trevor Mason, editor. The Darnton debate: books and revolution in the Eighteenth Century, volume 359 of Studies on Voltaire and the eighteenth century, 0435-2866. Voltaire Foundation, Oxford, UK, 1998. ISBN 0-7294-0566-4. ix + 305 pp. LCCN PQ2105.A2 S8 vol. 359.

Matthews:1995:QQM

[Mat95] J. Rosser Matthews. Quantification and the quest for medical certainty. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03794-9 (hardcover). x + 195 pp. LCCN R853.C55 M38

1995. URL http://www.loc.
gov/catdir/description/
prin031/94024091.html;
http://www.loc.gov/
catdir/toc/prin031/
94024091.html.

Mauskopf:1977:BRN

[Mau77] Seymour H. Mauskopf. Book review: Nineteenth and Twentieth Centuries: Claude Bernard and Animal Chemistry. By Frederic Lawrence Holmes. Cambridge, Mass.: Harvard University Press, 1974. Pp. xx + 541. No price stated. British Journal for the History of Science, 10(2):179–180, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025884.

Mauskopf:1987:BRH

[Mau87] Seymour H. Mauskopf. Book review: Hans Werner Schütt. Die Entdeckung des Isomorphismus. Hildesheim: Gerstenberg Verlag, 1984. Pp. 324. ISBN 3-8067-0902-5. No price given. British Journal for the History of Science, 20(3): 367–369, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026376.

Maxwell:1984:KWR

[Max84] Nicholas Maxwell. From knowledge to wisdom: a revolution in the aims and methods of science. Blackwell, Oxford, UK, 1984. ISBN 0-631-15641-0 (paperback), 0-631-13602-9. 300 pp. LCCN BD220. US$7.95.

Mays:1966:BRE

[May66] Wolfe Mays. Book review: Experience and Conceptual Activity. A philosophical essay based upon the writings of A. N. Whitehead. By J. M. Burgers. Pp. vii + 277. Cambridge: M.I.T. Press, Cambridge, Mass. and London. $7.50 (57s.). British Journal for the History of Science, 3(2): 195–196, December 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025039.

Mays:1967:BRB

[May67] Wolfe Mays. Book review: Boston Studies in the Philosophy of Science. Volume II. In Honour of Philipp Frank. Edited by Robert S. Cohen and Marx W. Wartofsky. Proceedings of the Boston Colloquium for the Philosophy of Science, 1962–1964. New York: Humanities Press. Pp. xxxiv + 475. 1965. $9.75. British Journal for the History of Science, 3(3):304, June 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025065.

Mays:1969:BRP

[May69] Wolfe Mays. Book review: Philosophy of Science Completeness in Science. By Richard Schegel. New York: Appleton-Century-Crofts. 1967. Pp. xi + 280. $7.50. British Journal for the History of Science, 4(3):287, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025477.

Mazumdar:1992:EHG

[Maz92] Pauline M. H. (Pauline Margaret Hodgson) Mazumdar. Eugenics, human genetics, and human failings: the Eugenics Society, its sources and its critics in Britain. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1992. ISBN 0-415-04424-3. x + 373 pp. LCCN HQ755.5.G7 M39 1991. US$35.00. URL http:

//www.loc.gov/catdir/
enhancements/fy0648/
90028804-d.html.

Mazhar:1995:CAE

[Maz95a] Noor Giovanni Mazhar.

Catholic attitudes to evolution in Nineteenth-Century Italian literature, volume 60 of Memorie, 0393-845X. Istituto veneto di scienze, lettere ed arti, Venezia, Italiy, 1995. ISBN 88-86166-24-9. 284 pp. LCCN BX1795.E85 M39 1995.

Mazumdar:1995:SSI

[Maz95b]  Pauline  M.

H. (Pauline Margaret Hodgson) Mazumdar. Species and specificity: an interpretation of the history of immunology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-43172-7 (hardcover). xiii + 457 pp. LCCN QR182 .M39

1995. URL http://www.loc.
gov/catdir/description/
cam025/93031219.html;
http://www.loc.gov/
catdir/samples/cam034/
93031219.html;   http://
www.loc.gov/catdir/toc/
cam023/93031219.html.

Mallinson:2008:ACH

[MB^{+}08] G. J. Mallinson, Hubert Bost, et al., editors. Architecture, cultural history, autobiography, volume 2008:06 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2008. ISBN 0-7294-0934-1 (paperback). LCCN DC33.4 .A73 2008.

Mellor:1963:DD

[MC63] C. M. Mellor and D. S. L. Cardwell. Dyes and dyeing 1775–1860. British Journal for the History of Science, 1 (3):265–279, June 1963. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024924.

Mendoza:1981:SCW

[MC81] E. Mendoza and D. S. L. Cardwell. On a suggestion concerning the work of J. P. Joule. British Journal for the History of Science, 14(2): 177–180, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025945.

Mengaldir:1993:HDC

[MC^{+}93] Paul Mengal, Jean Céard, et al., editors. Histoire du concept de récapitulation: ontogenèse et phylogenèse en biologie et sciences humaines. (French) [History of the concept of recapitulation: ontogenesis and phylogeny in biology and social sciences]. Histoire des sciences, 1243-5155. Masson Editeur, Masson, France, 1993. ISBN 2-225-84057-1. 132 pp. LCCN QH491 .H578 1993.

McCormmach:1967:JJT

[McC67] Russell McCormmach. J. J. Thomson and the structure of light. British Journal for the History of Science, 3(4):362– 387, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024961.

McCormmach:1968:JMH

[McC68] Russell McCormmach. John Michell and Henry Cavendish: Weighing the stars. British Journal for the History of Science, 4(2):126–155, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

McCormmach:1969:BRA

[McC69]  Russell  McCorm-

mach. Book review: Astronomy and Physics: The Electrical Researches of the Honourable Henry Cavendish. Ed. by James Clerk Maxwell. Reprint of 1879 edition. London: Frank Cass & Co. Pp. lxvi + 454. 1967. £6 6s. British Journal for the History of Science, 4(4): 408–409, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025240.

McCartney:1976:CLG

[McC76] Paul J. McCartney. Charles Lyell and G. B. Brocchi: A study in comparative historiography. British Journal for the History of Science, 9(2): 175–189, July 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025805.

McClaughlin:1977:BRH

[McC77a] Trevor McClaughlin. Book review: Historiography Clio and the Doctors. Psycho-History, Quanto-History, and History. By Jacques Barzun, Chicago and London: University of Chicago Press, 1974. Pp. xi + 173. £4.00. British Journal for the History of Science, 10(2):160, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025870.

McClaughlin:1977:BRS

[McC77b] Trevor McClaughlin. Book review: Seventeenth Century: Public Welfare, Science, and Propaganda in Seventeenth Century France. The Innovations of Théophraste Renaudot. By Howard M. Solomon. New Jersey: Princeton University Press, 1972. Pp. xv + 290. £8.50. British Journal for the History of Science, 10 (2):173–175, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025880.

McClaughlin:1980:BRS

[McC80] Trevor McClaughlin. Book review: Scientific revolution: Correspondance du P. Marin Mersenne, Religieux Minime.

Commencé par Paul Tannery. Pubilée et annotée par Cornélis de Waard avec la collaboration de Armand Beaulieu. Volume xiii: 1644–1645. Paris: Editions du Centre National de la Recherche Scientifique, 1977. Pp. ii + 608. 124 francs. British Journal for the History of Science, 13(1):72–73, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026130.

McConnell:1981:HIO

[McC81] Anita McConnell. Historical instruments in oceanography: background to the oceanography collection at the Science Museum. Her Majesty’s Stationary Office, London, England, 1981. ISBN 0-11-290324-X (paperback). iv + 51 pp. LCCN GC41 .M4 1981. US$3.95.

McConnell:1982:NST

[McC82] Anita McConnell. No sea too deep: the history of oceanographic instruments. Adam Hilger Ltd., Bristol, UK, 1982. ISBN 0-85274-416-1. 162 pp. LCCN GC41 .M425 1982. US$19.50.

McClaughlin:1983:BRL

[McC83] Trevor McClaughlin. Book review: Lives and Works Correspondance du P. Marin Mersenne; religieux minime. Commencé par Paul Tannery. Publiée et annotée par Cornélis de Waard et Armand Beaulieu. Volume XIV: 1646. Paris: Editions du Centre National de la Recherche Scientifique, 1980, pp. xii + 769. (No price stated.). British Journal for the History of Science, 16(2):202, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026244.

McCosh:1984:BCA

[McC84]  F.  W.  J.  (Fred-

erick William James) McCosh. Boussingault, chemist and agriculturist. Chemists and chemistry. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1682-X. xv + 280 + 2 pp. LCCN S584.52.B68 M33 1984.

McClellan:1985:SRS

[McC85a] James E. (James Edward) Mc-Clellan. Science reorganized: scientific societies in the eighteenth century. Columbia University Press, New York, NY, USA, 1985. ISBN 0-231-05996-5, 0-231-05997-3 (paperback). xxix + 413 pp. LCCN Q10 .M38 1985.

McConnell:1985:BRG

[McC85b] Anita McConnell. Book reviews: Gerard L’E Turner: Nineteenth-Century Scientific Instruments. London: Sotheby Publications, 1983, and Berkeley, Calif: Univ. of California Press, 1983. 320 pp. ISBN 0-85667-170-3 (UK), £37.50. ISBN 0-520-05160-2 (USA). British Journal for the History of Science, 18(1):121, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026301.

McConnell:1992:IMW

[McC92] Anita McConnell. Instrument makers to the world: a history of Cooke, Troughton and Simms. University of York, York, UK, 1992. ISBN 1-85072-096-7. xii + 116 pp. LCCN HD9706.6 .M12 1992.

URL http://books.google.
com/books?id=w8i1AAAAIAAJ.

McCurdy:1993:INH

[McC93] Howard E. McCurdy. Inside NASA: high technology and organizational change in the U.S. space program. New series in NASA history. The Johns Hopkins University Press, Baltimore, MD, USA, 1993. ISBN 0-8018-4452-5. xiv + 215 pp. LCCN TL521.312 .M33

1993. URL http://www.loc.
gov/catdir/bios/jhu051/
92018753.html;
http://www.loc.gov/
catdir/description/
jhu051/92018753.html.

McEvoy:1979:EKN

[McE79] John G. McEvoy. Electricity, knowledge, and the nature of progress in Priestley’s thought. British Journal for the History of Science, 12(1):

1–30, March 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025661.

McGuire:1965:BRP

[McG65] J. E. McGuire. Book review: Philosophy, Science and Sense Perception: Historical and Critical Studies. By M. Mandelbaum. Pp. xi + 262. Johns Hopkins Press; London: Oxford University Press, 1964. £2 12s. British Journal for the History of Science, 2(3): 263–264, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024944.

McGuire:1966:BRS

[McG66] J. E. McGuire. Book review: Sixteenth and Seventeenth Centuries: Newtonian Essays. By Alexandre Koyré. Pp. viii + 288. London: Chapman and Hall, 1965. 50s. British Journal for the History of Science, 3(1):84–85, June 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025113.

McGuire:1968:BRS

[McG68] J. E. McGuire. Book review: Seventeenth Century: Atomism in England from Hariot to Newton. By Robert Hugh Kargon. London: Clarendon Press: Oxford University Press.

Pp. viii + 168. 1966. 42s. net. Physiologia Epicuro-Gassendo-Charltoniana. By Walter Charleton. Edited by Robert Hugh Kargon. Reprinted from the 1654 edition. New York and London: Johnson Reprint Corporation. Pp. xxv + 491. 1966. $29.50. British Journal for the History of Science, 4(1):73–76, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024993.

McGuire:1969:BRE

[McG69] J. E. McGuire. Book review:

Eighteenth Century: The Elements of Sir Isaac Newton’s Philosophy. By Voltaire. Trans. John Hanna. London: Frank Cass. 1967. Pp. xvi + 363. 90s. British Journal for the History of Science, 4(3):300, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025492.

McGuire:1970:BRS

[McG70a] J. E. McGuire. Book review: Sixteenth and Seventeenth Centuries: John Locke: Problems and Perspectives. Ed. by John W. Yolton. London: Cambridge University Press. 1969. Pp. vii + 278. 55s. British Journal for the History of Science, 5(1):101–102, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025371.

McGuire:1970:NPP

[McG70b] J. E. McGuire. Newton’s ‘Principles of Philosophy’: An intended preface for the 1704 Opticks and a related draft fragment. British Journal for the History of Science, 5(2): 178–186, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025417.

McGuire:1971:BRE

[McG71a] J. E. McGuire. Book review: Eighteenth Century: Mechanism and Materialism. British Natural Philosophy in an Age of Reason. By Robert E. Schofield. Princeton University Press & Oxford University Press. 1970. Pp. vi + 336. £4.50. British Journal for the History of Science, 5(4):418– 419, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025405.

McGuire:1971:BRS

[McG71b] J. E. McGuire. Book review: Seventeenth Century: Correspondence of Sir Isaac Newton and Professor Cotes. By J. Edleston. London: F. Cass. 1969. Pp. xcviii + 316 + index. £6.30. British Journal for the History of Science, 5 (3):309–310, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025339.

McGuire:1978:NPT

[McG78] J. E. McGuire. Newton on place, time, and God: An unpublished source. British Journal for the History of Science, 11(2):114–129, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025725.

McGucken:1984:SSS

[McG84] William McGucken. Scientists, society, and state: the social relations of science movement in Great Britain, 1931– 1947. Ohio State University Press, Columbus, OH, USA, 1984. ISBN 0-8142-0351-5. 381 pp. LCCN Q175.52.G7 M38 1984. US$22.50.

McGuire:1995:TIN

[McG95] J. E. McGuire. Tradition and innovation: Newton’s metaphysics of nature, volume 56 of The University of Western Ontario series in philosophy of science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3617-8. xvi + 290 pp. LCCN B1299.N34 M37

1995. URL http://www.loc.
gov/catdir/enhancements/
fy0821/95030104-d.html;
http://www.loc.gov/

                                        catdir/enhancements/
                                        fy0821/95030104-t.html.

McIntosh:1985:BEC

[McI85] Robert P. (Robert Patrick) McIntosh. The background of ecology: concept and theory. Cambridge studies in ecology. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-24935-X. xiii + 383 pp. LCCN QH541 .M386

                                        1985. URL http://www.loc.
                                        gov/catdir/description/
                                        cam031/84027490.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam031/
                                        84027490.html.

McKirahan:1978:ASS

[McK78]  Richard  D.

McKirahan, Jr. Aristotle’s subordinate sciences. British Journal for the History of Science, 11(3):197–220, November 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025777.

McLeod:1986:BRR

[McL86]  Roy McLeod.  Book re-

view: R. F. Bud & G. K. Roberts. Science versus Practice: Chemistry in Victorian Britain. Manchester: Manchester University Press, 1984, Pp. 236. ISBN 0-7197-1070-5. £25. British Journal for the History of Science, 19(1): 119–120, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026492.

McNeil:1983:BRB

[McN83] Maureen McNeil. Book review: Biography Desmond King-Hele (ed.), The letters of Erasmus Darwin, Cambridge: Cambridge University Press, 1981. Pp. xxxii + 363. £45.00. British Journal for the History of Science, 16(1): 108–109, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026112.

McVaugh:1975:BRM

[McV75] Michael McVaugh. Book review: Middle Ages Tractatus de Perspectiva. By John Pecham. Ed. with introduction and notes by David C. Lindberg. St Bonaventure, N.Y.: The Franciscan Institute, 1972. Pp. 110. No price stated. British Journal for the History of Science, 8(2):181, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025650.

McVaugh:1980:SNB

[McV80] Michael McVaugh. Short notices of books: Natural and Supernatural: A History of the Paranormal from Earliest Times to 1914. By Brian Inglis. London: Hodder and Stoughton, 1977. Pp. 490. £9.95. British Journal for the History of Science, 13(2): 180–181, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025912.

Maclure:1988:EJW

[MD88] William Maclure and John S. Doskey. The European journals of William Maclure, volume 171 of Memoirs of the American Philosophical Society, 0065-9762. American Philosophical Society, Philadelphia, PA, USA, 1988. ISBN 0-87169-171-X. xlix + 815 + 1 pp. LCCN Q11 .P612 vol. 171.

Mann:1990:NMP

[MD90] Gunter Mann and Franz Dumont, editors. Die Natur des Menschen: Probleme der physischen Anthropologie und Rassenkunde (1750– 1850). (German) [The Nature of Man: Problems of Physical Anthropology and Ethnology], volume 6 of Soemmerring-Forschungen, 0176-876X. G. Fischer, Stuttgart, Germany, 1990. ISBN 3-437-11271-6. LCCN GN50.4 .N38 1990.

McCutcheon:1986:BRN

[MdLD86] Robert A. McCutcheon and Elizabelth de Lima Dantas. Book review: E. Nicolaidis. Le Developpement de l’Astronomie en U.R.S.S., 1917–1935. Paris: Observatoire, 1984. Pp. 253. British Journal for the History of Science, 19(2):233–234, July 1986.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026615.

Mendelsohn:1981:SCA

[ME81] Everett Mendelsohn and Yehuda Elkana, editors. Sciences and cultures: anthropological and historical studies of the sciences, volume 5 of Sociology of the sciences. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1981. ISBN 90-277-1234-4, 90-277-1235-2 (paperback). xvii + 270 pp. LCCN Q175.55 .S3.

Meadows:1971:BRP

[Mea71] A. J. Meadows. Book review:

Physical Sciences Lord Rayleigh-the Man and his Work. By Robert Bruce Lindsay. Oxford: Pergamon Press. 1970. Pp. viii + 251. £1.50 (paperback). British Journal for the History of Science, 5(3):300–301, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025328.

Meadows:1984:BRH

[Mea84] Jack Meadows. Book review:

History of Technology: Paul A. Hanle, Bringing aerodynamics to America. Cambridge, Mass, and London: MIT Press, 1982. Pp. xiv + 184. ISBN 0-262-08114-8. £14.00. British Journal for the History of Science, 17(3):330–331, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026646.

Meadows:1987:CRC

[Mea87] A. J. Meadows. Changing records and changing realities. British Journal for the History of Science, 20(1):67–72, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Megill:1994:RO

[Meg94] Allan Megill, editor. Rethinking objectivity. Post-contemporary interventions. Duke University Press, Durham, NC, USA, 1994. ISBN 0-8223-1479-7, 0-8223-1494-0 (paperback). ix + 342 pp. LCCN BD220 .R48 1994.

Meinel:1991:KFZ

[Mei91]  Christoph  Meinel.

Karl Friedrich Zöllner und die Wissenschaftskultur der Gründerzeit: eine Fallstudie zur Genese konservativer Zivilisationskritik. (German) [Karl Friedrich Zöllner and the explosive growth of the culture of science: a case study on the genesis of a more conservative critique of civilization], volume 13 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. Sigma, Berlin, Germany, 1991. ISBN 3-928068-04-0. 59 pp. LCCN Q127.G35 M45 1991.

Mendelsohn:1965:PMP

[Men65] Everett Mendelsohn. Physical models and physiological concepts: Explanation in Nineteenth-Century biology. British Journal for the History of Science, 2(3): 201–219, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024935.

Mendelsohn:1967:BRS

[Men67] Everett Mendelsohn. Book review: Scientific Societies in the United States. By Ralph S. Bates. Third Edition. Cambridge, Massachusetts: The M.I.T, Press. Pp. 326. 1965. 66s. British Journal for the History of Science, 3(3):301– 302, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025063.

Mendelsohn:1970:BRN

[Men70] Everett Mendelsohn. Book review: Nineteenth Century: Raspail: Scientist and Reformer. By Dora B. Weiner (with a chapter by Simone Raspail). New York, London: Columbia University Press. Pp. xiv + 336. Illustr. 1968. 99s. British Journal for the History of Science, 5(1):103–105, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025373.

Mendoza:1974:BRE

[Men74] Eric Mendoza. Book review: Eighteenth and Nineteenth Centuries: Lazare Carnot, Savant. By Charles Coulston Gillispie. Princeton, N. J. Princeton University Press, 1971. Pp. xii + 359. £8.25. British Journal for the History of Science, 7 (2):191–192, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025217.

Mendoza:1975:CEH

[Men75] Eric Mendoza. A critical examination of Herapath’s dynamical theory of gases. British Journal for the History of Science, 8(2):155–165, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025639.

Mendoza:1981:LWS

[Men81] Eric Mendoza. The life and work of Sadi Carnot. British Journal for the History of Science, 14(1):75–78, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026073.

Menard:1986:OTP

[Men86] Henry W. (Henry William) Menard. The ocean of truth: a personal history of global tectonics. Princeton series in geology and paleontology. Princeton University Press, Princeton, NJ, USA, 1986. ISBN 0-691-08414-9. xiv + 353 + 24 pp. LCCN QE501.3 .M45 1986. US$29.50.

Mercier:1978:BRA

[Mer78] Raymond Mercier. Book review: Antiquity A Survey of the Almagest. By Olaf Pedersen. Acta historica scientiarum naturalium et medicinalium, Edidit Bibliotheca Hauniensis, xxx. Odense: Odense University Press, 1974. Pp. 454. £20. British Journal for the History of Science, 11(1):75– 77, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025618.

Mercier:1979:NT

[Mer79] Raymond Mercier. Newtonian time. British Journal for the History of Science, 12 (2):211–217, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025755.

Merrill:1989:RVN

[Mer89] Lynn L. Merrill. The romance of Victorian natural history. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1989. ISBN 0-19-505203-X. x + 288 + 12 pp. LCCN PR756.N38 M47 1989; QH138.A1. URL http://www.

loc.gov/catdir/
enhancements/fy0638/
88009956-d.html.

Merton:1993:SGS

[Mer93] Robert King Merton. On the shoulders of giants: a Shandean postscript. University of Chicago Press, Chicago, IL, USA, Post-Italianate, University of Chicago Press edition, 1993. ISBN 0-226-52086-2 (paperback). xxv + 319 pp. LCCN QC16.N7 M38

1993. URL http://www.loc.
gov/catdir/description/
uchi052/92043871.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/92043871-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/92043871-b.html.

McGrath:1986:FDP

[MF86] William J. McGrath and Sigmund Freud. Freud’s discovery of psychoanalysis: the politics of hysteria. Cornell University Press, Ithaca, NY, USA, 1986. ISBN 0-8014-1770-8. 336 pp. LCCN BF173.F85 M226 1986.

Melhado:1992:ESR

[MF92] Evan Marc Melhado and Tore Frängsmyr, editors. Enlightenment science in the romantic era: the chemistry of Berzelius and its cultural setting, volume 10 of Uppsala studies in history of science. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41775-9. xiv + 246 pp. LCCN QD22.B5

E55 1992. URL http://www.
loc.gov/catdir/
description/cam024/
91045939.html;   http://
www.loc.gov/catdir/toc/
cam029/91045939.html.

Miller:1988:MCS

[MH88] George A. (George Armitage)

Miller and William Hirst, editors. The making of cognitive science: essays in honor of George A. Miller. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-34255-4. x + 284 pp. LCCN BF311 .M196

1988. URL http://www.loc.
gov/catdir/description/
cam023/87012492.html.

Micale:1995:AHD

[Mic95] Mark S. Micale. Approaching hysteria: disease and its interpretations. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03717-5. xii + 327 pp. LCCN RC532 .M53 1995. US$29.95. URL http:

//www.loc.gov/catdir/
description/prin021/
94016596.html;   http://
www.loc.gov/catdir/toc/
prin031/94016596.html.

Middleton:1965:JHK

[Mid65] W. E. Knowles Middleton. Jacob Hermann and the kinetic theory. British Journal for the History of Science, 2(3): 247–250, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024938.

Middleton:1969:TS

[Mid69] W. E. Knowles Middleton. The title of the Saggi. British Journal for the History of Science, 4(3):283–286, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025476.

Middleton:1975:SRA

[Mid75] W. E. Knowles Middleton. Science in Rome, 1675–1700, and the Accademia Fisicomatematica of Giovanni Giustino Ciampini. British Journal for the History of Science, 8(2): 138–154, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025638.

Millburn:1973:BMD

[Mil73] John R. Millburn. Benjamin Martin and the development of the orrery. British Journal for the History of Science, 6(4): 378–399, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025502.

Miller:1983:BHC

[Mil83] David Philip Miller. Between hostile camps: Sir Humphry Davy’s Presidency of The Royal Society of London, 1820–1827. British Journal for the History of Science, 16(1):1–47, March 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026092.

Miller:1987:BRS

[Mil87] Donald G. Miller. Book review: Stanley L. Jaki. Uneasy Genius: The Life and Work of Pierre Duhem. Dordrecht: Martinus Nijhoff, 1985. Pp. xii + 472. ISBN 90-247-2897-5. Dfl. 175.00; £44.50. British Journal for the History of Science, 20(1):110– 111, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026464.

Mills:1989:BOE

[Mil89]  Eric L. Mills.  Biological

oceanography: an early history, 1870–1960. Cornell University Press, Ithaca, NY, USA, 1989. ISBN 0-8014-2340-6. xvii + 378 pp. LCCN QH91.25 .M55 1989.

Millar:1990:CCD

[Mil90] David Mitarb Millar, editor.

Chambers concise dictionary of scientists. Chambers, Cambridge, UK, 1990. ISBN 1-85296-354-9, 0-550-16032-9. 461 pp. LCCN ????

Mirowski:1989:MHT

[Mir89] Philip Mirowski. More heat than light: economics as social physics, physics as nature’s economics. Historical perspectives on modern economics. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-35042-5. xii + 450 pp. LCCN HB74.P49 M57

1989. URL http://www.loc.
gov/catdir/description/
cam032/89009728.html;
http://www.loc.gov/
catdir/toc/cam031/
89009728.html.

Mirowski:1994:NIE

[Mir94] Philip Mirowski, editor. Natural images in economic thought: “Markets read in tooth and claw”, Historical perspectives on modern economics. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-44321-0, 0-521-47884-7. LCCN ????

Misa:1995:NSM

[Mis95] Thomas J. Misa. A nation of steel: the making of modern America, 1865–1925, volume new ser., no. 17 of Studies in the history of technology. The Johns Hopkins University Press, Baltimore, MD, USA, 1995. ISBN 0-8018-4967-5. xxvi + 367 pp. LCCN HD9515 .M576 1995. URL http://www.

loc.gov/catdir/bios/
jhu052/94038681.html;
http://www.loc.gov/
catdir/description/
jhu051/94038681.html.

Mitman:1992:SNE

[Mit92] Gregg Mitman. The state of nature: ecology, community, and American social thought, 1900–1950. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-53236-4 (hardcover), 0-226-53237-2 (paperback). xiv + 290 pp. LCCN QH541.26 .M58

1992. URL http://www.loc.
gov/catdir/description/
uchi051/91045638.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/91045638-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/91045638-b.html.

McCabe:1984:HST

[MJ84] Irena M. McCabe and Frank A.

J. L. James. History of science and technology resources at the Royal Institution of Great Britain. British Journal for the History of Science, 17 (2):205–209, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026552.

Malthus:1989:EPP

[MJ89] T. R. (Thomas Robert)

Malthus and Patricia James. An essay on the principle of population: or a view of its past and present effects on human happiness, with an inquiry into our prospects respecting the future removal or mitigation of the evils which it occasions. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-32361-4 (set), 0-521-32360-6 (vol. 1), 0-521-32363-0 (vol. 2), 0-521-24775-6. 2 v. pp. LCCN HB861 .E7

                                        1989. URL http://www.loc.
                                        gov/catdir/description/
                                        cam041/87009372.html.

Martin:1992:CMR

[MKW92] Ernst Martin, Peggy Aldrich Kidwell, and Michael R. Williams. The calculating machines (Die Rechenmaschinen): their history and development, volume 16 of The Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1992. ISBN 0-262-13278-8. xvii + 367 pp. LCCN QA75 .M3513 1992.

Monardes:1989:HMC

[ML89]  Nicolás  Monardes  and

José María López Piänero. La historia medicinal de las cosas que se traen de nuestras Indias Occidentales (1565–1574). (Spanish) [The medicinal history of our Western Indians]. Ministerio de Sanidad y Consumo, Madrid, Spain, 1989. ISBN 84-7670-191-8. 291 pp. LCCN RS169 .M65 1989.

Mockli:2010:EAR

[MM10] Daniel Möckli and Victor Mauer, editors. European–American relations and the Middle East: from Suez to Iraq. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2010. ISBN 0-415-47664-X (hardcover), 0-203-85144-7 (ebook). ???? pp. LCCN DS63.2.U5 E88 2010.

Malina:1990:AYT

[MMVZ90] Jaroslav Malina, Jaroslav Malina, Zdeněk Vašíček, and Marek Zvelebil. Archaeology yesterday and today: the development of archaeology in the sciences and humanities. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-26621-1 (hardcover), 0-521-31977-3 (paperback). xiv + 320 pp. LCCN CC100 .V3313 1990.

Molland:1968:GBM

[Mol68]  A. G. Molland.  The ge-

ometrical background to the “Merton school”: An exploration into the application of mathematics to natural philosophy in the Fourteenth Century. British Journal for the History of Science, 4(2):108– 125, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025142.

Molland:1969:BRI

[Mol69] A. G. Molland. Book review:

Islam Science and Civilization in Islam. By Seyyed Hussein Nasr, with a preface by Giorgio de Santillana. Harvard University Press and Oxford University Press. 1968. Pp. 384. 85s. 6d. British Journal for the History of Science, 4(4):

416, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025248.

Molland:1971:BRC

[Mol71] A. G. Molland. Book review: China Clerks and Craftsmen in China and the West: Lectures and Addresses on the History of Science and Technology. By Joseph Needham, based largely on collaborative work with Wang Ling, Lu Gwei-Djen and Ho Ping-Yü. London: Cambridge University Press. 1970. Pp. xix + 470. 40 plates. £7.50. British Journal for the History of Science, 5(4): 413–414, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025399.

Molland:1972:BRL

[Mol72a] A. G. Molland. Book review: Late Antiquity Fallgesetz und Massebegriff: Zwei wissenschaftshistorische Untersuchungen zur Kosmologie des Johannes Philoponus. Quellen und Studien zur Philosophie, Band 2. By Michael Wolff. Berlin: Walter de Gruyter, 1971. Pp. x + 159. DM.

36. British Journal for the History of Science, 6(1):91– 92, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025272.

Molland:1972:BRR

[Mol72b] A. G. Molland. Book review: Renaissance: A Critical Survey and Bibliography of Studies on Renaissance Aristotelianism 1958–1969. By Charles B. Schmitt. Padua: Antenore. 1971. Pp. 196. L. 3000. British Journal for the History of Science, 6(2):212–213, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025300.

Molland:1973:NOK

[Mol73] A. G. Molland. Nicole Oresme and the kinematics of circular motion. tractatus de commensurabilitate vel incommensurabilitate motuum celi. British Journal for the History of Science, 6(3):311–313, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025447.

Molland:1974:BRG

[Mol74] A. G. Molland. Book review:

General Causality and Scientific Explanation. Volume One: Medieval and Early Classical Science. By William A. Wallace. Ann Arbor: University of Michigan Press, 1972. Pp. xii + 288. $12. British Journal for the History of Science, 7 (1):83–84, March 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025181.

Molland:1975:BRS

[Mol75] A. G. Molland. Book review: Seventeenth Century: Der Musiktheoretiker Johannes Kepler. By Michael Dickreiter. Neue Heidelberger Studien zur Musikwissenschaft, Band 5. Berne and Munich: Francke Verlag, 1973. Pp. 252. 48 Swiss francs. British Journal for the History of Science, 8(1): 79–80, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025825.

Molland:1978:BRA

[Mol78a] A. G. Molland. Book review: Antiquity A History of Ancient Mathematical Astronomy. By O. Neugebauer. Studies in the History of Mathematics and Physical Sciences, I. Berlin, Heidelberg, & New York: Springer-Verlag, 1975. 3 vols. Pp. xxi + 1456. DM.

290. British Journal for the History of Science, 11(1):74– 75, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025617.

Molland:1978:BRM

[Mol78b] A. G. Molland. Book review: Middle Ages Richard of Wallingford. An edition of his writings with introductions, English translation, and commentary by J. D. North. Oxford: Clarendon Press: Oxford University Press, 1976. 3 vols. Pp. xxii + 567; v + 387; viii + 318 + XXIV plates. £75.00. British Journal for the History of Science, 11(3): 287–288, November 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025787.

Molland:1979:BRG

[Mol79] A. G. Molland. Book review: General Motion and Time, Space and Matter. Interrelations in the History of Philosophy and Science. Edited by Peter K. Machamer and Robert G. Turnbull. Columbus, Ohio: Ohio State University Press, 1976. Pp. xii + 559. $22.50. British Journal for the History of Science, 12 (1):89–90, March 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025666.

Molland:1980:MAA

[Mol80] A. G. Molland. Mathematical and angelic astronomy. British Journal for the History of Science, 13(3):255– 258, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026199.

Molland:1981:BRM

[Mol81] A. G. Molland. Book review: Middle Ages Witelonis Perspectivae Liber Primus. Book 1 of Witelo’s Perspectiva. An English Translation with Introduction and Commentary and Latin Edition of the Mathematical Book of Witelo’s Perspectiva by Sabetai Unguru. Studia Copernicana, XV. Wrolcaw et alibi: Ossolineum, The Polish Academy of Sciences Press, 1977. Pp. 332; 14 plates. Zl.

150. British Journal for the History of Science, 14(1):87– 88, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026077.

Molland:1984:BM

[Mol84a] A. G. Molland. Back matter. British Journal for the History of Science, 17(3):346, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026656.

Molland:1984:BRS

[Mol84b] A. G. Molland. Book review: Science Studies Nan L. Hahn, Medieval Mensuration. Quadrane Vetus and Geometrie Due Sunt Partes Principales … (Transactions of the American Philosophical Society, Volume 72, Part 8.) Philadelphia: American Philosophical Society, 1982. Pp. lxxxv + 204. British Journal for the History of Science, 17(3):336– 337, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026653.

Molland:1984:BRA

[Mol84c] A. George Molland. Book reviews: Antiquity to the Renaissance William A. Wallace, Prelude to Galileo: essays on medieval and sixteenth-century sources of Galileo’s thought. (Boston Studies in the Philosophy of Science, Volume 62.) Dordrecht, Boston & London: D. Reidel, 1981. Pp. xvi + 369. ISBN 90-277-1215-8, Dfl. 95/US $49.95 (hardcover); ISBN 90-277-1216-6, Dfl. 45/US $23.50 (paperback). British Journal for the History of Science, 17(1):105– 106, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026041.

Molland:1985:BRH

[Mol85] A. George Molland. Book review: H. L. L. Busard, ed., The First Latin Translation of Euclid’s Elements commonly ascribed to Adelard of Bath. Books I-VIII and Books X.36-XV.2. (Studies and Texts 64.) Toronto: Pontifical Institute of Mediaeval Studies, 1983. Pp. vi + 425. $37.00. British Journal for the History of Science, 18(1):95, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026277.

Molland:1986:BRF

[Mol86] A. G. Molland. Book review: Fritz Nagel. Nicolaus Cusanus und die Entstehung der Exakten Wissenschaften. Buchreihe der Cusanus-Gesellschaft, Band IX. Münster: Aschendorff, 1984. Pp. vii + 192. DM 48.00. British Journal for the History of Science, 19 (3):345, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026515.

Molland:1988:BRA

[Mol88] A. G. Molland. Book review: Alexander Jones (ed., tr. and commentary). Pappus of Alexandria. Book 7 of the Collection. Sources in the History of Mathematics and Physical Sciences 8. New York, Berlin, Heidelberg, Tokyo: Springer-Verlag, 1986. Pp. x + 748 in 2 vols. ISBN 0-387-96257-3, 3-540-96257-

3. DM 258.00. British Journal for the History of Science, 21(2):255, June 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026985.

Molland:1989:BRA

[Mol89] A. G. Molland. Book review:

Adam De Wodeham. Tractatus de Indivisibilibus: A Critical Edition. Introduction, Translation, and Textual Notes by Rega Wood. Synthese Historical Library, Volume 31. Dordrecht, Boston, London: Kluwer Academic Publishers, 1988. Pp. vii + 333. ISBN 90-277-2424-5. £74.00. British Journal for the History of Science, 22(4): 464–465, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026940.

Monter:1978:BRM

[Mon78] E. William Monter. Book review: Middle Ages Europe’s Inner Demons. An Enquiry Inspired by the Great Witch-Hunt. By Norman Cohn. London: Heinemann for Sussex University Press, 1975. Pp. xvi + 302. £4.50. British Journal for the History of Science, 11(2):180, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025739.

Montgomery:1987:EDC

[Mon87] William Montgomery. Editing the Darwin correspondence: A quantitative perspective. British Journal for the History of Science, 20(1):13– 28, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026433.

Moore:1977:CDI

[Moo77] J. R. Moore. Could Darwinism be introduced in France? British Journal for the History of Science, 10(3):246– 251, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025844.

Moore:1981:CPC

[Moo81] James R. Moore. Creation and the problem of Charles Darwin. British Journal for the History of Science, 14(2): 189–200, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025947.

Moore:1984:BRS

[Moo84] James R. Moore. Book review: Science in Culture: J. David Hoeveler Jr., James Mc-Cosh and the Scottish intellectual tradition: from Glasgow to Princeton. Princeton, New Jersey: Princeton University Press, 1981. Pp. xiv + 374. £17.40. British Journal for the History of Science, 17 (2):248–249, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026582.

Moore:1985:BRD

[Moo85] James R. Moore. Book review: D. P. Crook, Benjamin Kidd, Portrait of a Social Darwinist. Cambridge: Cambridge University Press, 1984. Pp. vii + 460. ISBN 0-521-25804-9. £30.00. British Journal for the History of Science, 18(3): 365–368, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026403.

Moore:1989:DHS

[Moo89a] James Moore. Darwinizing history: Sociobiology versus sociology. British Journal for the History of Science, 22(4): 429–432, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026919.

Moore:1989:SLT

[Moo89b]  Walter  John

Moore. Schrödinger, life and thought. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-35434-X. xi + 513 pp. LCCN QC16.S265 M66

1989. URL http://www.loc.
gov/catdir/description/
cam023/88025807.html;
http://www.loc.gov/
catdir/toc/cam021/
88025807.html.

Morrell:1969:BRE

[Mor69a]  J. B. Morrell.  Book re-

view: Eighteenth and Nineteenth Centuries: An Index to the Correspondence and Papers of James David Forbes (1809–1868) and also to Some Papers of his Son, George Forbes. St. Andrews University Library. 1968. Pp. ix + 123. 40s. British Journal for the History of Science, 4(4): 422–423, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025255.

Morrell:1969:TTP

[Mor69b] J. B. Morrell. Thomas Thomson: Professor of Chemistry and University Reformer. British Journal for the History of Science, 4(3):245– 265, June 1969. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Morrell:1971:BRE

[Mor71] J. B. Morrell. Book review: Eighteenth Century: Partners in Science: Letters of James Watt and Joseph Black. Ed. by Eric Robinson and Douglas McKie. London: Constable. 1970. Pp. xvi + 502. £4.20. British Journal for the History of Science, 5(3): 311–312, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025341.

Morrell:1972:BREa

[Mor72a]  J. B. Morrell.  Book re-

view: Eighteenth and Nineteenth Centuries: James Hutton’s System of the Earth, 1785; Theory of the Earth, 1788; Observations on Granite, 1794; together with Playfair’s Biography of Hutton. Introduction by V. A. Eyles. Darien, Conn.: Hafner Publishing Company, 1970. Pp. xxiii + 203. $12.95. British Journal for the History of Science, 6(1):96–97, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025278.

Morrell:1972:BREb

[Mor72b]  J. B. Morrell.  Book re-

view: Eighteenth and Nineteenth Centuries: Victorian Science: A Self-Portrait from the Presidential Addresses of the British Association for the Advancement of Science. Ed. by George Basalla, William Coleman, and Robert H. Kargon. New York: Doubleday, and London: Bell, 1970. Pp. x + 510. £1.20. British Journal for the History of Science, 6(1):99–100, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025281.

Morrell:1972:ER

[Mor72c]  J. B. Morrell.  Essay re-

views. British Journal for the History of Science, 6(2):200– 202, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025291.

Morrell:1972:SSU

[Mor72d] J. B. Morrell. Science and Scottish university reform: Edinburgh in 1826. British Journal for the History of Science, 6(1):39–56, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025262.

Morris:1972:LCT

[Mor72e] Robert J. Morris. Lavoisier and the Caloric Theory. British Journal for the History of Science, 6(1):1–38, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Morrell:1973:BREa

[Mor73a]  J. B. Morrell.  Book re-

view: Eighteenth and Nineteenth Centuries: Henry Home, Lord Kames, and the Scottish Enlightenment. By William C. Lehmann. The Hague: Martinus Nijhoff, 1971. Pp. xxvi + 358. 60.15 guilders. Lord Kames and the Scotland of his Day. By Ian Simpson Ross. Oxford: Clarendon Press: Oxford University Press, 1972. Pp. ix + 420. £6. British Journal for the History of Science, 6 (3):326–327, June 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025460.

Morrell:1973:BREb

[Mor73b] J. B. Morrell. Book review:

Eighteenth Century: James Burnett. Lord Monboddo. By E. L. Cloyd. Oxford: Clarendon Press: Oxford University Press, 1972. Pp. xii + 196. £4.50. British Journal for the History of Science, 6(4):449– 450, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025520.

Morrell:1974:BRE

[Mor74a]  J. B. Morrell.  Book re-

view: Eighteenth and Nineteenth Centuries: Philosophers and Provincials: the Yorkshire Philosophical Society from 1822 to 1844. By A. D. Orange. York: Yorkshire Philosophical Society, 1973. Pp. 76. £1.75. British Journal for the History of Science, 7(2): 196–197, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025221.

Morrell:1974:BRN

[Mor74b]  J. B. Morrell.  Book re-

view: Nineteenth and Twentieth Centuries: The World of Walther Nernst: The Rise and Fall of German Science. By K. Mendelssohn. London: Macmillan, 1973. Pp. 200.

£4.95. British Journal for the History of Science, 7(3): 302, November 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025550.

Morrell:1975:BRG

[Mor75a] J. B. Morrell. Book review: General Social Stratification in Science. By Jonathan R. Cole and Stephen Cole. Chicago and London: University of Chicago Press, 1973. Pp. xiv + 283. £6.25. British Journal for the History of Science, 8(3): 250, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025561.

Morrell:1975:BRN

[Mor75b]  J. B. Morrell.  Book re-

view: Nineteenth and Twentieth Centuries: The Universities and British Industry 1850– 1970. By Michael Sanderson. London: Routledge and Kegan Paul, 1972. Pp. x + 436. £6.50. British Journal for the History of Science, 8(2): 186–187, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025655.

Morrell:1975:BRS

[Mor75c] J. B. Morrell. Book review: Sociology of Science The Sociology of Science. Theoretical and Empirical Investigations. By Robert K. Merton. Ed. by Norman W. Storer. Chicago and London: University of Chicago Press, 1973. Pp. xxxi + 605. £6.25. British Journal for the History of Science, 8(1):70–71, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025817.

Morrell:1976:BRI

[Mor76a] J. B. Morrell. Book review: Institutions. The Cavendish Laboratory 1874–1974. By J. G. Crowther. London: Macmillan, 1974. Pp. xvi + 464. £25.00. British Journal for the History of Science, 9(1): 71–72, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025710.

Morrell:1976:LIL

[Mor76b] J. B. Morrell. London institutions and Lyell’s career: 1820–1841. British Journal for the History of Science, 9 (2):132–146, July 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025801.

Morrell:1977:BRA

[Mor77a] J. B. Morrell. Book review:

American and Canadian Science The Formation of the American Scientific Community. The American Association for the Advancement of Science 1848–1860. By Sally Gregory Kohlstedt. Urbana, Chicago, and London: University of Illinois Press, 1976. Pp. xiii + 264 + Appendix. $10.95. British Journal for the History of Science, 10(1):78–80, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025593.

Morrell:1977:BRN

[Mor77b] J. B. Morrell. Book review: Nineteenth and Twentieth Centuries: Historical Studies in the Physical Sciences: Fifth annual volume. Edited by Russell Mc-Cormmach. Princeton: Princeton University Press, 1975. Pp.

187. £7.30. Physics circa 1900: Personnel, Funding, and Productivity of the Academic Establishments. By Paul Forman, John L. Heilbron, and Spencer Weart. Princeton: Princeton University Press, 1975. Pp. 187. £7.30. British Journal for the History of Science, 10(3):272– 273, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025859.

Morrell:1977:BRS

[Mor77c]  J. B. Morrell.  Book re-

view: Sociology of Science Scientific Knowledge and Sociological Theory. By Barry Barnes. London and Boston: Routledge and Kegan Paul, 1974. Pp. x + 192. £3.95; £1-95 (paperback). British Journal for the History of Science, 10 (2):161–162, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025871.

Morrell:1978:BRG

[Mor78] J. B. Morrell. Book review:

General A History of Uppsala University, 1477–1977. By Sten Lindroth. Stockholm: Almqvist and Wiksell for Uppsala University, 1976. Pp. viii + 260. Sw. Kr. 67. British Journal for the History of Science, 11 (3):280, November 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025781.

Morrell:1980:BRN

[Mor80]  J. B. Morrell.  Book re-

view: Nineteenth Century: Michael Foster and the Cambridge School of Physiology: the Scientific Enterprise in Late Victorian Society. By Gerald L. Geison. Princeton: Princeton University Press, 1978. Pp. xix + 401. £18.40. British Journal for the History of Science, 13(2):175–176, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025908.

Morrell:1981:BRE

[Mor81]  J. B. Morrell.  Book re-

view: Eighteenth to Twentieth Centuries Natural Knowledge in Social Context: the Journals of Thomas Archer Hirst FRS. By William H. Brock and Roy M. MacLeod. London: Mansell, 1980. 80 microfiches. £168.00. British Journal for the History of Science, 14(3): 293–294, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025991.

Morrell:1982:BRD

[Mor82a]  J. B. Morrell.  Book re-

view: Darwiniana Martha Mc-Mackin Garland, Cambridge before Darwin: the ideal of a liberal education, 1800–1860. Cambridge: Cambridge University Press, 1980. Pp. ix + 196. £14.50. British Journal for the History of Science, 15(1):79, March 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025929.

Morsink:1982:AGA

[Mor82b] Johannes Morsink. Aristotle: On the generation of animals: a philosophical study. University Press of America, Washington, DC, USA, 1982. ISBN 0-8191-2606-3, 0-8191-2607-1 (paperback). viii + 184 pp. LCCN QL41.A717 M67 1982. US$19.50, US$9.50.

Morrell:1983:BRJ

[Mor83a]  J. B. Morrell.  Book re-

view: Journals and Dictionaries Charles Schmitt (ed.), History of universities: Volume I: ‘Continuity and change in early modern universities’. Amersham: Avebury Publishing Company, 1981. Pp. viii + 225. £20.00 (Vol I only). Annual subscription: £16.00 (libraries and institutions); £12.00 (individuals). British Journal for the History of Science, 16(3): 315, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026182.

Morrell:1983:BRM

[Mor83b]  J.  B.

Morrell. Book review: Mathematical Sciences: Thomas L. Hankins, Sir William Rowan Hamilton. Baltimore and London: Johns Hopkins University Press, 1980. Pp. xxi + 474. $32.50/£19.50. British Journal for the History of Science, 16(3):288–289, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026159.

Morrell:1983:BRS

[Mor83c] J. B. Morrell. Book review: Science and Society James Paradis and Thomas Postlewait (eds), Victorian science and Victorian values: literary perspectives. New York: New York Academy of Sciences, 1981. Pp. xiii + 362. $70.00. British Journal for the History of Science, 16(3): 307–308, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026175.

Morse:1983:BRT

[Mor83d] JoAnn S. Morse. Book review: Technology and Instruments Stephen K. Victor (ed.) Practical geometry in the high middle ages. Artis cuiuslibet consummatio, and the Pratike de geometrie. Philadelphia: American Philosophical Society, 1979. Pp. xii + 638. (No price stated.). British Journal for the History of Science, 16(2):211–212, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026252.

Moreau:1984:CCA

[Mor84a] René Moreau. The computer comes of age: the people, the hardware, and the software. MIT Press series in the history of computing. MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-13194-3. x + 227 pp. LCCN QA76.17 .M6713 1984.

Morgan:1984:BRC

[Mor84b] N. Morgan. Book review: Chemistry and Biochemistry J. S. Fruton, A bio-bibliography for the history of the biochemical sciences since 1800.

Philadelphia: American Philosophical Society, 1982. Pp. xiii + 885. ISBN 0-87169-983-4. $20. British Journal for the History of Science, 17(2): 243, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026575.

Morrell:1984:BRN

[Mor84c] Jack Morrell. Book review:

National Traditions in Science T. W. Heyck, The transformation of intellectual life in Victorian England. London and Canberra: Croom Helm, 1982. Pp. 262. ISBN 0-7099-1206-4. £14.50. British Journal for the History of Science, 17(1): 97, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026032.

Morrell:1984:BRS

[Mor84d] Jack Morrell. Book review:

Science in Culture: Arthur J. Engel, From clergyman to don: the rise of the academic profession in nineteenthcentury Oxford. Oxford: Oxford University Press, 1983. Pp. xi + 302. ISBN 0-19-822606-3. £22.50. British Journal for the History of Science, 17 (2):247–248, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026581.

Morrell:1984:GSE

[Mor84e] Jack Morrell, editor. Gentlemen of science: early correspondence of the British Association for the Advancement of Science, volume 30 of Camden fourth series. Offices of the Royal Historical Society, University College London, London, UK, 1984. ISBN 0-86193-103-3. 382 pp. LCCN DA20 .R913 vol. 30. US$10.00.

Morris:1984:BRC

[Mor84f] Peter Morris. Book review: Chemistry and Biochemistry Frank M. McMillan, The chain straighteners. Macmillan: London, 1979. Pp. xvi +207. £17.00. British Journal for the History of Science, 17(2): 243–244, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026576.

Morris:1984:BRN

[Mor84g] Peter Morris. Book review: National Traditions in Science Helmut Tammen, Die I. G. Farbenindustrie Aktiengesellschaft (1925–1933): ein Chemiekonzern in der Weimarer Republik. Berlin: Verlag Helmuth Tammen, 1978. Pp. 468. Price 82 DM (about £20). British Journal for the History of Science, 17(1):87–88, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026022.

Morselli:1984:AAS

[Mor84h] Mario Morselli. Amedeo Avogadro, a scientific biography. Chemists and chemistry. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1624-2. xi + 375 pp. LCCN QD22.A8 M67 1984.

Morrell:1985:WWP

[Mor85] J. B. Morrell. Wissenschaft in Worstedopolis: Public science in Bradford, 1800–1850. British Journal for the History of Science, 18(1):1–23, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026263.

Moran:1987:BRE

[Mor87a] Bruce T. Moran. Book review: Edward Rosen. Three Imperial Mathematicians: Kepler Trapped Between Tycho Brahe and Ursus. New York: Abaris Books, 1986. Pp. 384. ISBN 0-89835-242-8. $20.00. British Journal for the History of Science, 20(2): 235–236, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026321.

Morrell:1987:BRM

[Mor87b]  J. B. Morrell.  Book re-

view: Martin J.S. Rudwick. The Great Devonian Controversy: The Shaping of Scientific Knowledge Among Gentlemanly Specialists. Chicago and London: University of Chicago Press, 1985. Pp. xxxiii + 494. ISBN 0-226-73101-4. £36.75. British Journal for the History of Science, 20(1): 88–89, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026445.

Morris:1989:BRJ

[Mor89a] Peter Morris. Book reviews: Jonathan Liebenau. Medical Science and Medical Industry. The Formation of the American Pharmaceutical Industry. London: Macmillan, 1987. Pp ix + 207. ISBN 0-333-41742-9, £29.50. John P. Swann. Academic Scientists and the Pharmaceutical Industry. Cooperative Research in Twentieth Century America. Baltimore: The Johns Hopkins University Press, 1988. Pp xi + 249. ISBN 0-8018-3558-5, £22.50. British Journal for the History of Science, 22(4): 442–444, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026923.

Morris:1989:ASR

[Mor89b] Peter John Turnbull Morris. The American synthetic rubber research program. The chemical sciences in society series. University of Pennsylvania Press, Philadelphia, PA, USA, 1989. ISBN 0-8122-8205-1. xii + 191 pp. LCCN TS1925 .M58 1989.

Morselli:1989:BRR

[Mor89c] Mario A. Morselli. Book review: Roberto Maiocchi, La Belle Époque dell’atomo. Milano: Franco Angeli, 1988. Pp. ii + 315. Lire 28 000 (paperback). British Journal for the History of Science, 22(2): 243–245, July 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026663.

Morgan:1990:HEI

[Mor90a] Mary S. Morgan. The history of econometric ideas. Historical perspectives on modern economics. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37398-0, 0-521-42465-8. xv + 296 pp. LCCN HB139 .M66 1990.

Morris:1990:HWS

[Mor90b] P. R. (Peter Robin) Morris. A history of the world semiconductor industry, volume 12 of IEE history of technology series. P. Peregrinus on behalf of the Institution of Electrical Engineers, London, UK, 1990. ISBN 0-86341-227-0. ix + 171 pp. LCCN HD9696.S42 M67 1990.

Moran:1991:AWG

[Mor91a] Bruce T. Moran. The alchemical world of the German court: occult philosophy and chemical medicine in the circle of Moritz of Hessen, 1572–1632, volume Heft 29 of Sudhoffs Archiv. Beihefte, 0341-0773. F. Steiner Verlag, Stuttgart, Germany, 1991. ISBN 3-515-05369-7. 193 pp. LCCN BF1434.G5 M67 1991. DM58.00.

Moran:1991:CPE

[Mor91b] Bruce T. Moran. Chemical pharmacy enters the university: Johannes Hartmann and the didactic care of chymiatria in the early Seventeenth Century, volume new ser., no. 14 of Publication / American Institute of the History of Pharmacy. American Institute of the History of Pharmacy, Madison, WI, USA, 1991. ISBN 0-931292-23-9 (paperback), 0-931292-24-7. vii + 88 pp. LCCN RS65 .M67 1991.

Moseley:1977:BRAa

[Mos77a] Russell Moseley. Book review: American and Canadian Science A Curious Field-Book. Science and Society in Canadian History. Ed. by Trevor H. Levere and Richard A. Jarrell. Toronto: Oxford University Press, 1974. Pp. x + 233. £3.50. British Journal for the History of Science, 10(1): 80, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025594.

Moseley:1977:BRAb

[Mos77b] Russell Moseley. Book review:

American and Canadian Science Let us be Honest and Modest. Technology and Society in Canadian History. Ed. by B. Sinclair, N. R. Ball, and J. O. Petersen. Toronto: Oxford University Press, 1974. Pp. xvi + 309. £3.80. British Journal for the History of Science, 10(1):80, March 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Moss:1993:NHR

[Mos93]  Jean Dietz Moss.  Novel-

ties in the heavens: rhetoric and science in the Copernican controversy. University of Chicago Press, Chicago, IL, USA, 1993. ISBN 0-226-54234-3, 0-226-54235-1 (paperback). xiv + 353 pp. LCCN QB29 .M67 1993. URL http://www.

loc.gov/catdir/
description/uchi052/
92021608.html;   http://
www.loc.gov/catdir/toc/
uchi051/92021608.html.

Moyer:1983:APT

[Moy83] Albert E. Moyer. American physics in transition: a history of conceptual change in the late Nineteenth Century, volume 3 of The History of modern physics, 1800–1950. Tomash Publishers, Los Angeles, 1983. ISBN 0-938228-06-4. xx + 218 pp. LCCN QC9.U5 M68 1983. US$30.00 (est.).

Moyal:1986:BSL

[Moy86] Ann Mozley Moyal. A bright and savage land: scientists in colonial Australia. Collins, London, UK, 1986. ISBN 0-00-217555-X. 192 pp. LCCN QH26 .M68 1986.

Maffioli:1989:ISL

[MP89] Cesare S. Maffioli and Lodewijk C. Palm, editors. Italian scientists in the low countries in the XVIIth and XVIIIth centuries, volume 34 of Nieuwe Nederlandse bijdragen tot de geschiedenis der geneeskunde en der natuurwetenschappen. Rodopi, Amsterdam and Leiden, The Netherlands, 1989. ISBN 90-5183-121-8. ISSN 0168-9827. LCCN Q127.I8 I84; Q141 .I82 1988. URL http:/

                                        /catalog.hathitrust.org/
                                        api/volumes/
                                        oclc/20849739.html. Invited

papers from the congress held in Utrecht on 25–27 may 1988 to commemorate the 350th anniversary of the publication of Galileo Galilei’s “Discorsi e dimostrazioni matematiche intorno a due nove scienze”, Leyden 1638.

Mehra:1982:DQM

[MR82a] Jagdish Mehra and Helmut Rechenberg. The discovery of quantum mechanics, 1925, volume 2 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90674-6. vi + 355 pp. LCCN QC173.98 .M44 vol. 2. See other volumes in this series [MR82c, MR82b, MR82f, MR82g, MR87a, MR00].

Mehra:1982:FMM

[MR82b] Jagdish Mehra and Helmut Rechenberg. The formulation of matrix mechanics and its modifications, 1925–1926, volume 3 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90675-4, 3-540-90675-4. viii + 334 pp. LCCN QC173.98 .M44 vol.

3. URL http://www.loc.gov/
catdir/enhancements/
fy1005/82003253-d.html;
http://www.loc.gov/
catdir/enhancements/
fy1005/82003253-t.html. See

other volumes in this series [MR82a, MR82c, MR82f, MR82g, MR87a, MR00].

Mehra:1982:FEQ

[MR82c] Jagdish Mehra and Helmut Rechenberg. The fundamental equations of quantum mechanics, 1925–1926. The reception of the new quantum mechanics, 1925–1926, volume 4 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90680-0. viii + 322 pp. LCCN QC173.98 .M44 vol. 4.

See other volumes in this series [MR82a, MR82b, MR82f, MR82g, MR87a, MR00].

Mehra:1982:HDQa

[MR82d] Jagdish Mehra and Helmut Rechenberg. The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-90675-4 (Berlin), 0-387-90675-4 (New York). vi + 334 pp. LCCN QC173.98 .M44 vol. 3.

Mehra:1982:HDQb

[MR82e] Jagdish Mehra and Helmut Rechenberg. The historical development of quantum theory. Vol. 4,1, The fundamental equations of quantum mechanics 1925– 1926. Vol. 4,2, The reception of the new quantum mechanics 1925–1926. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90680-0, 3-540-90680-0. viii + 222 pp. LCCN ????

Mehra:1982:QTPa

[MR82f] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 1) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-90642-8 (Berlin), 0-387-90642-8 (New York). 372 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82g, MR87a, MR00].

Mehra:1982:QTPb

[MR82g] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 2) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90667-3 (New York), 3-540-90667-3 (Berlin). 375– 878 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82f, MR87a, MR00].

Mehra:1987:ESR

[MR87a] Jagdish Mehra and Helmut Rechenberg. Erwin Schrödinger and the rise of wave mechanics. Part 1, Schrödinger in Vienna and Zurich 1887–1925, volume 5, part 1 of The Historical Development of Quantum Theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96284-0 (vol. 1, New York), 3-540-96284-0 (vol. 1, Berlin), 0-387-96377-4 (vol. 2, New York), 3-387-96377-4 (vol. 2, Berlin). xx + 366 pp. LCCN QC173.98 .M44 vol. 5. US$48.00, US$49.50. See other volumes in this series [MR82a, MR82c, MR82b, MR82f, MR82g, MR00].

Mehra:1987:FEQ

[MR87b] Jagdish Mehra and Helmut Rechenberg. The fundamentals equations of quantum mechanics, 1925–1926. Part 2, The reception of the new quantum mechanics, 1925–1926, volume 4 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-90680-0 (New York), 3-540-90680-0 (Berlin). viii + 322 pp. LCCN ????

MacLeod:1988:NGE

[MR88] Roy M. MacLeod and Philip F. Rehbock, editors. Nature in its greatest extent: Western science in the Pacific. University of Hawaii Press, Honolulu, HI, USA, 1988. ISBN 0-8248-1120-8. LCCN Q127.P23 I56 1985. US$34.00. URL http:

                                        //catalog.hathitrust.org/
                                        api/volumes/oclc/
                                        16901461.html. Papers in

this volume were presented at the XVIIth Internat. Congress of History of Science, University of California, Berkeley, 31 July–8 August 1985.

Mendenhall:1989:ASE

[MR89] Thomas C. (Thomas Corwin) Mendenhall and Richard Rubinger. An American scientist in early Meiji Japan: the autobiographical notes of Thomas C. Mendenhall, volume 35 of Asian studies at Hawaii. University of Hawaii Press, Honolulu, HI, USA, 1989. ISBN 0-8248-1177-1. xvii + 89 pp. LCCN QC16.M483 .A3 1989; DS21 .A831 no.35. US$11.00. URL http://

catalog.hathitrust.org/
api/volumes/oclc/
18521314.html.

Miller:1996:VEV

[MR96] David Philip Miller and Peter Hanns Reill, editors. Visions of empire: voyages, botany, and representations of nature. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-48303-4 (hardcover). xix + 370 pp. LCCN QH11 .V57

1996. URL http://www.loc.
gov/catdir/description/
cam027/95019085.html;
http://www.loc.gov/
catdir/toc/cam021/
95019085.html.

Mehra:2000:CQM

[MR00] Jagdish Mehra and Helmut Rechenberg. The completion of quantum mechanics, 1926– 1941, volume 6 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-95086-9 (part 1), 0-387-95182-2 (part 2). xxxvi + 1612 pp.

LCCN QC173.98 .M44 vol. 6. See other volumes in this series [MR82a, MR82c, MR82b, MR82f, MR82g, MR87a].

MartinezAlier:1987:EEE

[MS87] Juan Martínez Alier and Klaus Schlüpmann. Ecological economics: energy, environment, and society. Blackwell, Oxford, UK, 1987. ISBN 0-631-15739-5. ix + 286 pp. LCCN HD85.S7 M363513 1987. US$39.95.

Miller:1995:HMM

[MS95] Hugh Miller and Michael Shortland. Hugh Miller’s memoir: from stonemason to geologist. Edinburgh University Press, Edinburgh, Scotland, 1995. ISBN 0-7486-0521-5. v + 266 pp. LCCN QE22.M5 A3 1995.

McGuire:1983:CPQ

[MT83] J. E. McGuire and Martin Tamny. Certain philosophical questions: Newton’s Trinity notebook. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-23164-7. xii + 519 pp. LCCN Q175 .M417

                                        1983. URL http://www.loc.
                                        gov/catdir/description/
                                        cam031/82022200.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam031/
                                        82022200.html.

Mulkay:1979:KSS

[Mul79] Michael Mulkay. Kuhn and the sociology of science. British Journal for the History of Science, 12(3):298–301, November 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026005.

Muller:1987:BRR

[Mül87] Gerhard H. Müller. Book review: Robert P. McIntosh. The Background of Ecology. Concept and Theory. Cambridge: Cambridge University Press, 1985. Pp. xiii + 383. ISBN 0-521-24935-X. £30.00, $39.50. British Journal for the History of Science, 20(1): 111–113, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026465.

Muller:1989:BRJ

[Mül89] Gerhard Müller. Book review: J. Heniger. Hendrik Adriaan van Reede tot Drakenstein (1636–1691) and Hortus Malabaricus. A Contribution to the History of Dutch Colonial Botany. Rotterdam and Boston: A.A. Balkema, 1986. Pp. xvi + 295. ISBN 90-6191-681-X. £50.00; $75.00; Dfl 150.00. British Journal for the History of Science, 22(1): 90–91, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026690.

Murdin:1985:UNS

[Mur85] Lesley Murdin. Under Newton’s shadow: astronomical practices in the Seventeenth Century. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-456-0. viii + 152 pp. LCCN QB33.G7 M87 1985. US$18.00. URL

                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0745/85120279-d.html.

Murray:1989:GHI

[Mur89] Penelope Murray, editor. Genius: the history of an idea. Blackwell, Oxford, UK, 1989. ISBN 0-631-15785-9. ix + 235 pp. LCCN BF412 .G44 1989. US$50.00.

Musson:1974:BRN

[Mus74] A. E. Musson. Book review: Nineteenth and Twentieth Centuries: Science and the Rise of Technology since 1800. Ed. by C. A. Russell and D. C. Goodman. Bristol: John Wright and Sons in association with the Open University Press, 1972. Pp. xii + 338. £3.95. British Journal for the History of Science, 7(1):97, March 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025194.

Maffioli:1994:GSW

[MW94]  Cesare  S.

Maffioli and Richard S. Westfall. Out of Galileo: the science of waters: 1628–1718, volume 49 of Nieuwe Nerderlandse bijdragen tot de geschiedenes der geneeskunde en der natuurwetenschappen. Erasmus Publ., Rotterdam, The Netherlands, 1994. ISBN 90-5235-071-X. xx + 509 pp. LCCN GB737 .M34 1994.

Nudds:1988:SIT

[N^{+}88] John R. Nudds et al., editors.

Science in Ireland, 1800–1930: tradition and reform: proceedings of an international symposium held at Trinity College, Dublin, March 1986. Trinity College, Dublin, Ireland, 1988. ISBN 0-9513586-1-8. LCCN Q127.I6 S354 1988; Q127.I73 S35 1988. URL http://

catalog.hathitrust.org/
api/volumes/oclc/
18819538.html.

Nadeau:1983:BRA

[Nad83] Robert L. Nadeau. Book review: Americana R. E. Martin, American literature and the universe of force. Durham, North Carolina: Duke University Press, 1981. Pp. x + 284. $27.75. British Journal for the History of Science, 16 (3):321, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026190.

Nakayama:1975:BRC

[Nak75] Shigeru Nakayama. Book review: China Chinese Mathematics in the Thirteenth Century. The Shu-shu Chiuchang of Ch’in Chiu-shao. By Ulrich Libbrecht. M.I.T. East Asian Science Series, Volume 1. Cambridge, Mass., and London:

M.I.T. Press, 1973. Pp. xxxi + 555. $25.00. British Journal for the History of Science, 8(3):252–254, November 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025565.

Nakayama:1991:STS

[Nak91] Shigeru Nakayama. Science, technology, and society in postwar Japan. Japanese studies. Kegan Paul International, London, UK, 1991. ISBN 0-7103-0428-5. xv + 259 pp. LCCN Q127.J3 N27 1991.

Napier:1990:R

[Nap90] John Napier. Rabdology, volume 15 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-14046-2. xxxvi + 135 pp. LCCN QA75 .N2713 1990.

Naylor:1974:GPF

[Nay74] R. H. Naylor. Galileo and the problem of free fall. British Journal for the History of Science, 7(2):105–134, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025199.

Naylor:1975:BRS

[Nay75] Ron Naylor. Book review: Seventeenth Century: Galileo Galilei. Two New Sciences, including Centres of Gravity and Force of Percussion. Trans, by Stillman Drake. Madison: University of Wisconsin Press, 1974. Pp. xxxix + 323. No price stated. British Journal for the History of Science, 8(3): 257–258, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025569.

Naylor:1977:BRS

[Nay77] Ron Naylor. Book review: Seventeenth and Eighteenth Centuries: Galileo. By Colin A. Ronan. London: Weidenfeld and Nicholson, 1974. Pp. 264. £5.00. British Journal for the History of Science, 10(3):260– 261, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025850.

Naylor:1980:BRS

[Nay80] R. H. Naylor. Book review:

Scientific revolution: New Perspectives on Galileo. Edited by Robert E. Butts and Joseph C. Pitt. Dordrecht: D. Reidel, 1977. Pp.xvi + 262. £19.00/ £9.00 (paperback). Galileo’s Early Notebooks: the Physical Questions. By William A. Wallace. Notre Dame & London: University of Notre Dame press, 1977. Pp. xiii + 321. £10.50. British Journal for the History of Science, 13(1): 74–76, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026132.

Naylor:1985:BRS

[Nay85] R. Naylor. Book review: Stillman Drake, Cause Experiment and Science: A Galilean dialogue incorporating a new English translation of Galileo’s ‘Bodies That Stay atop Water or Move in it.’ Chicago: University of Chicago Press, 1981. Pp. xxviii + 237. ISBN 0-226-16228-1. $20.00. British Journal for the History of Science, 18(1):102, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026284.

Numbers:1995:CTC

[NBR^{+}95] Ronald L. Numbers, Arthur I. Brown, William Bell Riley, William Vance Trollinger, George McCready Price, Harold Willard Clark, and Frank Lewis Marsh, editors. Creationism in twentiethcentury America: a ten-volumes anthology of documents, 1903–1961. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1802-2 (1), 0-8153-1803-0 (2), 0-8153-1804-9 (3), 0-8153-1805-7 (4), 0-8153-1806-5 (5), 0-8153-1807-3 (6), 0-8153-1808-1 (7), 0-8153-1809-X (8), 0-8153-1810-3 (9), 0-8153-1811-1 (10). ???? pp. LCCN ????

Neuburger:1981:HDE

[NC81] Max Neuburger and Edwin Clarke. The historical development of experimental brain and spinal cord physiology before Flourens. The Johns Hopkins University Press, Baltimore, MD, USA, 1981. ISBN 0-8018-2380-3. xxv + 391 pp. LCCN QP353 .N4813.

Nebeker:1995:CWM

[Neb95] Frederik Nebeker. Calculating the weather: meteorology in the 20th Century, volume 60 of International geophysics series. Academic Press, New York, USA, 1995. ISBN 0-12-515175-6. vii + 255 pp. LCCN QC871 .N344 1995. URL http:

//www.loc.gov/catdir/
description/els032/
94037625.html;   http://
www.loc.gov/catdir/toc/
els032/94037625.html.

Needham:1969:BRC

[Nee69] Joseph Needham. Book review: China John Fryer: The Introduction of Western Science and Technology into Nineteenth Century China. By Adrian Arthur Bennett. Harvard University Press and Oxford University Press. 1967. (Harvard East Asian Monographs no. 24.) Pp. 157 + xii. 31s. British Journal for the History of Science, 4(4):416– 418, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025249.

Needham:1978:AOS

[Nee78] Joseph Needham. Address to the opening session of the XV International Congress of the History of Science, Edinburgh, 11 August 1977. British Journal for the History of Science, 11(2):103–113, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025724.

Needell:1985:BRR

[Nee85] Allan Needell. Book review: Richard F. Hirsh, Glimpsing an Invisible Universe: The Emergence of X-Ray Astronomy. Cambridge University Press, 1983. Pp. viii + 186. ISBN 0-521-25121-4. £20, $39.50. British Journal for the History of Science, 18 (2):237–238, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026336.

Nersessian:1987:PSC

[Ner87]  Nancy  J.

Nersessian, editor. The process of science: contemporary philosophical approaches to understanding scientific practice. Science and philosophy. Martinus Nijhoff, Dordrecht and The Hague, The Netherlands, 1987. ISBN 90-247-3425-8. xii + 221 pp. LCCN Q175.3 .P785 1987.

Nersessian:1989:BRJ

[Ner89] Nancy J. Nersessian. Book review: Joseph C. Pitt & Marcello Pera (eds). Rational Changes in Science: Essays on Scientific Reasoning. Dordrecht: D. Reidel, 1987. Pp. xiii + 224. ISBN 9-02772-417-2. £49.50; $64.00. British Journal for the History of Science, 22(1):101, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026701.

Neve:1978:BRG

[Nev78] Michael Neve. Book review:

Geology History of the Earth Sciences during the Scientific and Industrial Revolutions. By D. H. Hall. Amsterdam, Oxford, & New York: Elsevier, 1976. Pp. xi + 297. $19.25. British Journal for the History of Science, 11(1): 68–69, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025612.

Neve:1980:BRN

[Nev80] Michael Neve. Book review:

Nineteenth Century: Henry De la Beche: observations on an observer. By Paul J. McCartney. Cardiff: Friends of the National Museum of Wales, 1977. Pp. xiv + 76. £3.00. British Journal for the History of Science, 13(2):168–169, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025902.

Neve:1983:BRB

[Nev83] Michael Neve. Book review: Biography Charles Lyte, Sir Joseph Banks: Eighteenth Century Explorer, Botanist and Entrepreneur, Newton Abbot: David and Charles. 1980. Pp.

248. £10.50. British Journal for the History of Science, 16 (1):105–106, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026110.

Neve:1984:BRN

[Nev84] Michael Neve. Book review: National Traditions in Science J. V. Pickstone (ed.), Health, disease and medicine in Lancashire 1750–1950: four papers on sources, problems and methods. Manchester: U.M.I.S.T., 1980. Pp. 103. £2.00. British Journal for the History of Science, 17(1):98–99, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026034.

Newton:1989:PMI

[New89] Sir Isaac Newton. The preliminary manuscripts for Isaac Newton’s 1687 Principia, 1684– 1685, volume 2 of Newton manuscripts series. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-33499-3. xxi + 246 pp. LCCN QA803 .N413 1989.

Newman:1994:GFL

[New94]  William  Royall  New-

man. Gehennical fire: the lives of George Starkey, an American alchemist in the scientific revolution. Harvard University Press, Cambridge, MA, USA, 1994. ISBN 0-674-34171-6. xiv + 348 pp. LCCN QD24.S73 N49 1994.

Needham:1987:SCC

[NhTWL87] Joseph Needham, Tsuen hsuin Tsien, Ling Ed Wang, and Gwei Djen Lu. Science and civilisation in China. 05, 01, Chemistry and chemical technology, Paper and printing. Cambridge University Press, Cambridge, UK, third edition, 1987. ISBN 0-521-08690-6. xxv + 485 pp. LCCN ????

Nicholson:1985:BRE

[Nic85] Malcolm Nicholson. Book review: Edward Lee Greene, Landmarks of Botanical History, edited by Frank N. Egerton. Stanford: Stanford University Press, 1983. Part I, Part II, Pp. 1139. ISBN 0-8047-1075-9. $100.00. British Journal for the History of Science, 18(1): 117–118, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026297.

Nicholls:1988:HMP

[Nic88a] Phillip A. Nicholls. Homoeopathy and the medical profession. Croom Helm, London, UK, 1988. ISBN 0-7099-1836-4. 298 pp. LCCN RX51 .N53 1988.

Nicolson:1988:BRT

[Nic88b] Malcolm Nicolson. Book review: Thomas Söderqvist. The Ecologists, from Merry Naturalists to Saviours of the Nation: A sociologically informed narrative survey of the ecologization of Sweden 1895–1975. Stockholm: Almqvist and Wiksell, 1986. Pp. vii + 330. ISBN 91-22-00827-6. No price given. British Journal for the History of Science, 21(2): 257–259, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026988.

Newton:1991:SIN

[NJ91] Sir Isaac Newton and Peter Jones. Sir Isaac Newton: manuscripts and papers. Chadwyck-Healey, Cambridge, England, 1991. ISBN 0-85964-226-7 (guide). 43 microfilm reels pp. LCCN Q158.5; Microfilm 95/4567 (Q).

Needham:1984:SCC

[NLH^{+}84] Joseph Needham, Gwei-Djen Lu, H. T. Huang, Francesca Bray, Christian Daniels, Nicholas K. Menzies, and Nathan Sivin. Science and civilisation in China.

Vol. 6, Biology and biological technology. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-08731-7 (pt. 1), 0-521-25076-5 (pt. 2), 0-521-41999-9 (pt. 3), 0-521-65270-7 (pt. 5), 0-521-63262-5 (pt. 6). ???? pp. LCCN HD2098.NEE.

Nelson:1995:CWB

[NN95] Byron C. (Byron Christopher)

Nelson and Paul Nelson. The creationist writings of Byron C. Nelson, volume 5 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1806-5. xxvi + 505 pp. LCCN BS651 .N44 1995.

Norberg:1996:TCT

[NOF96] Arthur L. (Arthur Lawrence)

Norberg, Judy E. (Judy Elizabeth) O’Neill, and Kerry J. Freedman. Transforming computer technology: information processing for the Pentagon, 1962–1986, volume [new ser., no. 18] of Johns Hopkins studies in the history of technology. The Johns Hopkins University Press, Baltimore, MD, USA, 1996. ISBN 0-8018-5152-1. xiv + 360 pp. LCCN QA76.17 .N67

1996. URL http://www.loc.
gov/catdir/bios/jhu051/
95023820.html;
http://www.loc.gov/
catdir/description/
jhu051/95023820.html.

Nola:1988:RRS

[Nol88] Robert Nola, editor. Relativism and realism in science, volume 6 of Australasian studies in history and philosophy of science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2647-7. ix + 299 pp. LCCN Q175 .R385 1988.

Noll:1994:JCO

[Nol94] Richard Noll. The Jung cult: origins of a charismatic movement. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03724-8. xv + 387 pp. LCCN BF109.J8 N65 1994. US$24.95. URL http://

                                        www.loc.gov/catdir/
                                        description/prin031/
                                        94004831.html;   http://
                                        www.loc.gov/catdir/toc/
                                        prin031/94004831.html.

Noll:1995:FMO

[Nol95] Steven Noll. Feeble-minded in our midst: institutions for the mentally retarded in the South, 1900–1940. University of North Carolina Press, Chapel Hill, NC, USA, 1995. ISBN 0-8078-2220-5, 0-8078-4531-0 (paperback). xiii + 254 pp. LCCN HV3006.S85 N65 1995.

North:1966:WAB

[Nor66] J. D. North. Werner, Apian, Blagrave and the meteoroscope. British Journal for the History of Science, 3(1): 57–65, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025102.

North:1968:BRH

[Nor68a] J. D. North. Book review:

History of Mathematics The Mathematical Papers of Isaac Newton, Volume I: 1664–1666. Edited by D. T. Whiteside, with the assistance in publication of M. A. Hoskin. London: Cambridge University Press, 1967. £10 10s. British Journal for the History of Science, 4(1):82–84, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025000.

North:1968:BRS

[Nor68b] J. D. North. Book review: Sixteenth and Seventeenth Centuries: John Kepler. By Angus Armitage. London: Faber and Faber. 1966. Pp. 194. 21s. The Six-cornered Snowflake. By Johannes Kepler, edited and translated by Colin Hardie, with essays by L. L. Whyte and B. F. J. Mason. London: Oxford University Press, Clarendon Press. 1966. Pp. xvi + 75. 21s. British Journal for the History of Science, 4(2): 190–191, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025163.

North:1969:BRH

[Nor69a] J. D. North. Book review:

History of mathematics. The Mathematical Papers of Isaac Newton, Volume II: 1667–1670. Ed. by D. T. Whiteside, with the assistance in publication of M. A. Hoskin. London: Cambridge University Press. 1968. Pp. xxii + 520. £10 10s. British Journal for the History of Science, 4(3):289–290, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025480.

North:1969:WWH

[Nor69b] J. D. North. William Whewell and the history and philosophy of science. British Journal for the History of Science, 4(4):401–402, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025234.

North:1970:BRM

[Nor70a] J. D. North. Book review: Mathematics: The Mathematical Papers of Isaac Newton, Volume III, 1670–1673. Ed. by D. T. Whiteside, with the assistance in publication of M. A. Hoskin and A. Prag. London: Cambridge University Press. 1969. Pp. xxxvii + 576. £10 10s. British Journal for the History of Science, 5(2): 188–189, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025420.

North:1970:EHE

[Nor70b]  J. D. North.  Essays in

the history of the exact sciences. British Journal for the History of Science, 5(1): 85–86, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025355.

North:1971:BRH

[Nor71a] J. D. North. Book review:

History and Philosophy of Science Dieter v. Stephanitz, Exakte Wissenschaft und Recht, Münsterische Beiträge zur Rechts und Staatswissenschaft, Heft 15. Berlin, Walter de Gruyter. Pp. xii + 273. 1970. D.M. 48. British Journal for the History of Science, 5(3): 293, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025320.

North:1971:BRM

[Nor71b] J. D. North. Book review:

Middle Ages John Pecham and the Science of Optics. Perspectiva communis. Edited with an introduction, English translation, and critical notes, by David C. Lindberg. Madison: University of Wisconsin Press. 1970. Pp. xvii + 300. $15.00. British Journal for the History of Science, 5(4):415–416, December 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025401.

North:1972:BRM

[Nor72] J. D. North. Book review: Mathematics: Gesammelte Abhandlungen. By David Hubert. 2nd edition. Berlin: Springer-Verlag, 1970. 3 vols. Pp. xvi + 539; viii + 453; vii + 435 $27.00. Hilbert Gedenkband. Ed. by K. Reidemeister. Berlin: Springer-Verlag, 1971. Pp. v + 86. DM 22. British Journal for the History of Science, 6(2): 208–209, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025297.

North:1973:BRN

[Nor73a] J. D. North. Book review: Newtonian Studies The Mathematical Papers of Isaac Newton. Volume V, 1683–1684. Ed. by D. T. Whiteside, with the assistance in publication of M. A. Hoskin and A. Prag. London: Cambridge University Press, 1972. Pp. xxiv + 628. £20. British Journal for the History of Science, 6(4):444– 445, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025516.

North:1973:W

[Nor73b] J. D. North. L. L. Whyte. British Journal for the History of Science, 6(3):339–340, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025468.

North:1977:BRM

[Nor77] J. D. North. Book review:

Middle Ages Studies in Medieval Philosophy, Science, and Logic. Collected Papers, 1933– 1969. By Ernest A. Moody. Berkeley, Los Angeles, & London: University of California Press, 1975. Pp. xx + 454. £11.00. British Journal for the History of Science, 10(3): 258–260, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025849.

Norton:1980:BRN

[Nor80] Bernard Norton. Book review: Nineteenth and Twentieth Centuries: Thomas Hunt Morgan. The Man and his Science. By Garland E. Allen. Princeton: Princeton University Press, 1979. Pp. xvii + 447. £17.60. British Journal for the History of Science, 13(3): 277–278, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026212.

Norton:1982:BRD

[Nor82] Bernard Norton. Book review: Darwiniana William Bateson, Problems of genetics.

Historical introduction by G. Evelyn Hutchinson and Stan Rachootin. Silliman Milestones in Science. New Haven and London: Yale University Press, 1979. Pp. xxii + 258. £12.30; £3.15 (paperback). British Journal for the History of Science, 15(1):78, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025928.

Norton:1983:BRS

[Nor83] Bernard Norton. Book review: Science and Society Donald MacKenzie, Statistics in Britain 1865–1930. The social construction of scientific knowledge. Edinburgh: Edinburgh University Press, 1981. Pp. viii + 306. £12.50. British Journal for the History of Science, 16(3):304–306, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026173.

Norton:1985:BRE

[Nor85a] Bernard Norton. Book review: Edward Yoxen, The Gene Business. Who should control Biotechnology. London: Pan Books, 1983. Pp. 265. ISBN 0-330-28112-7. £3.95 (paperback). British Journal for the History of Science, 18(2): 241, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026340.

Norton:1985:BRP

[Nor85b] Bernard Norton. Book review: Peter Bowler, The eclipse of Darwinism: Anti-Darwinian evolution theories in the decades around 1900. Baltimore: The Johns Hopkins University Press, 1983. Pp. xi + 291. ISBN 0-8018-2932-1. $25.00. British Journal for the History of Science, 18(2):245, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026344.

Norst:1989:FBA

[Nor89]  Marlene J. Norst.  Ferdi-

nand Bauer: the Australian natural history drawings, volume 1 of Art in natural history. British Museum (Natural History), London, UK, 1989. ISBN 0-565-01067-0 (hardcover), 0-565-01048-4 (paperback). 120 pp. LCCN NC994.5.B34 N67 1989. US$16.95, US$8.95.

North:1994:FHA

[Nor94]  John David North.  The

Fontana history of astronomy and cosmology. Fontana history of science. Fontana Press, London, UK, 1994. ISBN 0-00-686177-6 (paperback). xxvii + 697 + 8 pp. LCCN ????

Neve:1977:ACG

[NP77] Michael Neve and Roy Porter.

Alexander Catcott: Glory and geology. British Journal for the History of Science, 10(1): 37–60, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025580.

Naumann:1992:LEE

[NPH92] Bernd Naumann, Frans Plank, and Gottfried Hofbauer, editors. Language and earth: effective affinities between the emerging sciences of linguistics and geology, volume 66 of Amsterdam studies in the theory and history of linguistic science. Series III, Studies in the history of the language sciences, 0304-0720. J. Benjamins, Amsterdam, The Netherlands, 1992. ISBN 90-272-4553-3 (Netherlands), 1-55619-361-0 (US). LCCN P69 .L36 1992. US$110.00.

Needham:1986:CCT

[NPYGDL86] Joseph Needham, Ho Ping-Yu, Lu Gwei-Djen, and Wang Ling. Chemistry and chemical technology. Part 7, Military technology: the gunpowder epic, volume 5, pt. 7 of Science and civilisation in China. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-30358-3. xxxiii + 703 pp. LCCN DS721 .N39 vol. 5 pt. 7.

Needham:1986:SSCa

[NR86] Joseph Needham and C. A. (Colin Alistair) Ronan. The shorter science and civilisation in China: an abridgement of Joseph Needham’s original text. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25272-5, 0-521-31560-3 (paperback). ix + 298 pp. LCCN ????

Newton:1984:OPI

[NS84] Sir Isaac Newton and Alan E.

(Alan Elihu) Shapiro. The optical papers of Isaac Newton. Volume I. The optical lectures 1670–1672. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25248-2. xix + 627 pp. LCCN QC353 .N52 1984.

Negrel:2002:ERE

[NS02] Eric Négrel and Jean-Paul Sermain, editors. Une expérience rhétorique: l’éloquence de la Révolution. (French) [A rhetorical experiment: the eloquence of the Revolution], volume 2002:02 of SVEC 0435-2866. Voltaire Foundation, Oxford, UK, 2002. ISBN 0-7294-0788-8. vi + 333 pp. LCCN DC155 .E9 2002.

Numbers:1987:MNW

[Num87] Ronald L. Numbers, editor.

Medicine in the New World: New Spain, New France, and New England. University of Tennessee Press, Knoxville, TN, USA, 1987. ISBN 0-87049-517-8. LCCN R465 .M43 1987. US$18.95.

Numbers:1992:C

[Num92] Ronald L. Numbers. The creationists. A. A. Knopf, New York, NY, USA, 1992. ISBN 0-679-40104-0. xvii + 458 pp. LCCN BS651 .N85 1992. US$27.50.

Numbers:1995:ABW

[Num95a] Ronald L. Numbers, editor. Antievolutionism before World War I, volume 1 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1802-2. xvii + 403 pp. LCCN BT712 .A57 1995.

Numbers:1995:CED

[Num95b]  Ronald  L.  Numbers,

editor. Creation-evolution debates, volume 2 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1803-0. xiv + 505 pp. LCCN BS651 .C6924 1995.

Nunan:1988:BRH

[Nun88] Richard Nunan. Book review: H. W. Menard. The Ocean of Truth: A Personal History of Global Tectonics. Princeton: Princeton University Press, 1986. Pp. xiv + 353 + 24 un-numbered (illustrations). ISBN 0-691-08414-9. $29.50. British Journal for the History of Science, 21(1): 120–122, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026877.

Nuttall:1982:BM

[Nut82a] R. H. Nuttall. Back matter. British Journal for the History of Science, 15(3):320, November 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026234.

Nuttall:1982:BRI

[Nut82b] R. H. Nuttall. Book review:

Instruments Anita McConnell, Geomagnetic instruments before 1900: An illustrated account of their construction and use. London: Harriet Wynter, 1980. Pp. viii + 75. £19.50. British Journal for the History of Science, 15(3): 314, November 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026231.

Nuttall:1983:BRT

[Nut83]  R. H. Nuttall.  Book re-

view: Technology and Instruments F. A. B. Ward, A catalogue of European scientific instruments in the Department of Medieval and Later Antiquities of the British Museum. London: British Museum Publications, 1981, Pp. 152 + Pl.

61. £50.00. British Journal for the History of Science, 16 (2):209–210, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026250.

Nutton:1984:BRA

[Nut84]  Vivian Nutton.  Book re-

view: Antiquity to the Renaissance Johannes Morsink, Aristotle on the generation of animals, a philosophical study. Washington: University Press of America, 1982. Pp. 192. ISBN 0-8191-2607-1. $19.50 ($9.50 paperback). Abraham Wasserstein, Galen’s Commentary on the Hippocratic Treatise, Airs, Waters, Places. Jerusalem: Proceedings of the Israel Academy of Sciences and Humanities, VI.3, 1982. Pp. 119 (paperback). (No price stated.). British Journal for the History of Science, 17 (1):102–103, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026038.

Neu:1982:ICBb

[NW82a] John Neu and Magda Whitrow, editors. ISIS cumulative bibliography. 1913–65, Formed from ISIS bibliographies 1–90. 5, Civilizations and periods, 15th to 19th centuries: a bibliography of the history of science. Mansell, London, UK, 1982. ISBN 0-7201-1643-0. xi + 593 pp. LCCN ????

Neu:1982:ICBa

[NW82b] John Neu and Magda Whitrow, editors. ISIS cumulative bibliography. 1913–65, Formed from ISIS critical bibliographies 1– 90. 4, Civilizations and periods, Prehistory to Middle Ages: a bibliography of the history of science, volume 1913–65 of ISIS cumulative bibliography. Mansell, London, UK, 1982. ISBN 0-7201-1642-2. xviii + 457 pp. LCCN ????

Nye:1976:MFM

[Nye76] Mary Jo Nye. The moral freedom of man and the determinism of nature: The Catholic synthesis of science and history in the revue des questions scientifiques. British Journal for the History of Science, 9(3): 274–292, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025683.

Nye:1986:BRW

[Nye86] Mary Jo Nye. Book review:

W. H. Brock. From Protyle to Proton. William Prout and the Nature of Matter, 1785–1985. Bristol and Boston: Adam Hilger, Ltd., 1985. Pp. xii + 252. ISBN 0-85274-801-9. £25.00, $45.00. British Journal for the History of Science, 19(2):216–217, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026602.

Nye:1990:EAS

[Nye90] David E. Nye. Electrifying America: social meanings of a new technology, 1880–1940. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-14048-9. xv + 479 pp. LCCN HD9685.U5 N94 1990.

Nye:1993:CPT

[Nye93] Mary Jo Nye. From chemical philosophy to theoretical chemistry: dynamics of matter and dynamics of disciplines, 1800– 1950. University of California Press, Berkeley, CA, USA, 1993. ISBN 0-520-08210-9. xvii + 328 pp. LCCN QD452 .N94

                                        1993. URL http://www.loc.
                                        gov/catdir/bios/ucal051/
                                        92043114.html;
                                        http://www.loc.gov/
                                        catdir/description/
                                        ucal041/92043114.html.

Nyhart:1995:BTF

[Nyh95] Lynn K. Nyhart. Biology takes form: animal morphology and the German universities, 1800– 1900. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1995. ISBN 0-226-61086-1 (hardcover), 0-226-61088-8 (paperback). xiii + 414p. pp. LCCN QL799.5 .N94

                                        1995. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi051/95003227.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0609/95003227-b.html;
                                        http://www.loc.gov/

catdir/toc/uchi051/
95003227.html.

Olby:1990:CHM

[O^{+}90] Robert C. (Robert Cecil) Olby et al., editors. Companion to the history of modern science. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1990. ISBN 0-415-01988-5. xxvi + 1081 pp. LCCN Q125 .C565

1990. URL http://www.loc.
gov/catdir/enhancements/
fy0648/89010483-d.html.

OConnor:1988:FBP

[O’C88] Walter John O’Connor, editor.

Founders of British physiology: a biographical dictionary, 1820–1885: a collection of biographical sketches. Manchester University Press, Manchester, UK, 1988. ISBN 0-7190-2537-0. ix + 278 pp. LCCN QP25 .O28 1988. US$25.00. URL http:

//catalog.hathitrust.org/
api/volumes/oclc/
17506405.html.

ODonnell:1983:WRH

[O’D83]  Seán O’Donnell.  William

Rowan Hamilton: portrait of a prodigy, volume 1 of Profiles of genius series, 0332-3218. Boole Press, Dublin, Ireland, 1983. ISBN 0-906783-06-2. xvi + 224 pp. LCCN MLCM 91/04283 (Q).

Ordonez:1990:CSP

[OE^{+}90] Javier Ordóñez, Alberto Elena, et al., editors. La ciencia y su público: perspectivas históricas. (Spanish) [Science and its public: historical perspectives], volume 12 of Estudios sobre la ciencia. Consejo Superior de Investigaciones Científicas, Madrid, Spain, 1990. ISBN 84-00-07056-9. xv + 443 pp. LCCN Q175.55 .C55 1990.

Ogilvie:1975:RCN

[Ogi75] Marilyn Bailey Ogilvie. Robert Chambers and the Nebular Hypothesis. British Journal for the History of Science, 8(3): 214–232, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025557.

Ogilvie:1986:WSA

[Ogi86]  Marilyn  Bailey

Ogilvie. Women in science: antiquity through the nineteenth century: a biographical dictionary with annotated bibliography. MIT Press, Cambridge, MA, USA, 1986. ISBN 0-262-15031-X. xi + 254 pp. LCCN Q141 .O34 1986.

Oresme:1985:NOM

[OH85]  Nicole  Oresme  and

Bert Hansen. Nicole Oresme and The marvels of nature: a study of his De causis mirabilium with critical edition, translation, and commentary, volume 68 of Studies and texts, 0082-5328. Pontifical Institute of Mediaeval Studies, Toronto, ON, Canada, 1985. ISBN 0-88844-068-5 (paperback). xi + 478 pp. LCCN Q153 .O7413 1985.

OHara:1986:BRS

[O’H86] Jim O’Hara. Book reviews:

Seá O’Donnell. William Rowan Hamilton: Portrait of a Prodigy. Dublin: Boole Press, 1984. Pp. xvi + 224. ISBN 0-906783-06-2. IR £19.95, $24.95. Desmond MacHale. George Boole: His Life and Work. Dublin: Boole Press, 1985. Pp. xiii + 304. ISBN 0-906783-05-4. IR £19.95, $24.95. British Journal for the History of Science, 19(3):360–361, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026531.

Olby:1963:CDM

[Olb63] R. C. Olby. Charles Darwin’s manuscript of pangenesis. British Journal for the History of Science, 1(3):251– 263, June 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024923.

Olby:1965:MC

[Olb65] R. C. Olby. The Mendel Centenary. British Journal for the History of Science, 2(4): 343–349, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024890.

Olby:1969:BRB

[Olb69] Robert Olby. Book review: Biology and Medicine Theodor Boveri. Life and Work of a Great Biologist. By Fritz Baltzer, translated from the German by Dorothea Rudnick. Berkeley and Los Angeles: University of California Press. 1967. Pp. xii + 165. 57s. British Journal for the History of Science, 4(4):412– 413, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025244.

Olby:1970:BRB

[Olb70] Robert Olby. Book review: Biology and Medicine Eugenics and the Progressives. By Donald K. Pickens. Pp. xii + 260. Nashville, Tennessee: Vanderbilt University Press. 1968. $7.95. British Journal for the History of Science, 5(2): 200, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025431.

Olby:1971:BRB

[Olb71a] Robert Olby. Book review: Biological Sciences and Medicine The Chemistry of Life. Lectures on the History of Biochemistry. Ed. by Joseph Needham. London: Cambridge University Press. 1970. Pp. xxx + 214. £3. British Journal for the History of Science, 5 (3):303–304, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025332.

Olby:1971:BRG

[Olb71b] Robert Olby. Book review: Geology and Biology Time’s Arrow and Evolution. By Harold F. Blum. Third Edition. Princeton, New Jersey: Princeton University Press. 1968. Pp.

232. £1.30. British Journal for the History of Science, 5(4): 409–410, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025394.

Olby:1972:BRE

[Olb72] Robert Olby. Book review: Eighteenth and Nineteenth Centuries: La Logique du Vivant. Une Histoire de l’Hérédité. By François Jacob. Paris: Gallimard, 1970. Pp. 354. £2.70. British Journal for the History of Science, 6(2):221–222, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025308.

Olby:1977:HGE

[Olb77] R. C. Olby. Human genetics, eugenics and race. British Journal for the History of Science, 10(2):156–159, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025869.

Olby:1982:BRB

[Olb82] Robert Olby. Book review: Bibliography David Bearman and John T. Edsall (ed), Archival sources for the history of biochemistry and molecular biology: a reference guide and report. Boston: American Academy of Arts and Sciences; Philadelphia: American Philosophical Society, 1980. Pp. xii + 338. (No price stated.). British Journal for the History of Science, 15(1):85–86, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025934.

Olby:1985:BRS

[Olb85]  Robert  Olby.

Book review: Sheldon Krimsky, Genetic Alchemy. The Social History of the Recombinant DNA Controversy. Cambridge, Mass., & London: MIT Press, 1982. Pp. xiii + 444. ISBN 0-262-11083-0. £17.50. British Journal for the History of Science, 18(2):241–243, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026341.

Olby:1987:WBI

[Olb87]  Robert  Olby.

William Bateson’s introduction of Mendelism to England: A reassessment. British Journal for the History of Science, 20(4): 399–420, October 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026416.

Olby:1989:DSC

[Olb89] Robert Olby. The dimensions of scientific controversy: The biometric–Mendelian debate. British Journal for the History of Science, 22(3):299– 320, September 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026898.

Oldroyd:1972:RHM

[Old72] D. R. Oldroyd. Robert Hooke’s methodology of science as exemplified in his ‘Discourse of Earthquakes’. British Journal for the History of Science, 6(2): 109–130, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025287.

Oldroyd:1977:LPD

[Old77] D. R. Oldroyd. A list of Ph.D.

theses in the history of science and related areas in Australian universities to 1976. British Journal for the History of Science, 10(1):86–87, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025598.

Oldroyd:1986:AKI

[Old86] D. R. (David Roger) Oldroyd. The arch of knowledge: an introductory study of the history of the philosophy and methodology of science. Methuen, New York, NY, USA, 1986. ISBN 0-416-01331-7, 0-416-01341-4 (paperback). 413 pp. LCCN Q174.8 .O43 1986.

Oleson:1984:GRM

[Ole84] John Peter Oleson. Greek and Roman mechanical waterlifting devices: the history of a technology, volume 16 of Phoenix: journal of the Classical association of Canada, Supplementary volume. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1693-5. ISSN 0079-1784. xiv + 458 + 149 pp. LCCN TJ840.

Olesko:1991:PCD

[Ole91] Kathryn Mary Olesko. Physics as a calling: discipline and practice in the Königsberg seminar for physics. Cornell history of science series. Cornell University Press, Ithaca, NY, USA, 1991. ISBN 0-8014-2248-5. xviii + 488 pp. LCCN QC47.G3 O43 1991.

Olmi:1992:LDM

[Olm92] Giuseppe Olmi. L’inventario del mondo: catalogazione della natura e luoghi del sapere nella prima et à moderna. (Italian) [The inventory of the world: cataloging the nature and places the knowledge from the first to modern times], volume 17 of Annali dell’Istituto storico italo-germanico. Monografia. Il Mulino, Bologna, Italy, 1992. ISBN 88-15-03647-4. 457 pp. LCCN QH83 .O46 1992. L50000.

Olson:1982:SDS

[Ols90] Richard Olson. Science deified and science defied: the historical significance of science in Western culture. University of California Press, Berkeley, CA, USA, 1982–1990. ISBN 0-520-04621-8 (vol. 1), 0-520-06846-7 (vol. 2). ???? pp. LCCN Q125 .O44 1982.

OHara-May:1975:BRM

[OM75] Jane O’Hara-May. Book review: Medicine The Medical Society of London 1773–1973. Ed. by Thomas Hunt. London: Heinemann, 1972. Pp. xviii + 141. £2.50. British Journal for the History of Science, 8 (1):74–75, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025821.

OMeara:1989:PRM

[O’M89] Dominic J. O’Meara. Pythagoras revived: mathematics and philosophy in late antiquity. Clarendon Press, New York, NY, USA, 1989. ISBN 0-19-824485-1. xii + 251 pp. LCCN B243 .O46 1989. US$49.95.

URL http://www.loc.gov/
catdir/enhancements/
fy0637/88025246-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0637/88025246-t.html.

Oakley:1989:ABS

[OO89] Brian Oakley and Kenneth Owen. Alvey, Britain’s strategic computing initiative. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-15038-7. xv + 337 pp. LCCN QA76 .O225 1989.

OHara:1987:HMS

[OP87] J. G. (James G.) O’Hara and W. (Willibald) Pricha. Hertz and the Maxwellians: a study and documentation of the discovery of electromagnetic wave radiation, 1873–1894, volume 8 of IEE History of technology series. P. Peregrinus Ltd. in association with the Science Museum, London, UK, 1987. ISBN 0-86341-101-0. xiv + 154 pp. LCCN QC670 .O47 1987.

Oppenheim:1985:OWS

[Opp85] Janet Oppenheim. The other world: spiritualism and psychical research in England, 1850– 1914. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-26505-3. xii + 503 + 16 pp. LCCN BF1242.G7

O66 1985. URL http://www.
loc.gov/catdir/
description/cam022/
84005887.html;   http://
www.loc.gov/catdir/toc/
cam029/84005887.html.

Orange:1972:OBA

[Ora72]  A. D. Orange.  The ori-

gins of the British Association for the Advancement of Science. British Journal for the History of Science, 6(2): 152–176, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025289.

Orange:1973:BRN

[Ora73] A. D. Orange. Book review:

Nineteenth Century: Sketches of the Royal Society and Royal Society Club (London, 1849). By Sir John Barrow. Reprint. London: Frank Cass, 1971. British Journal for the History of Science, 6(4):450–451, December 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025521.

Orange:1974:BRN

[Ora74] A. D. Orange. Book review:

Nineteenth and Twentieth Centuries: The Organisation of Science in England. By D. S. L. Cardwell. Revised edition. London: Heinemann, 1972. Pp. xii + 268. £1.25. British Journal for the History of Science, 7(3):298–299, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025546.

Orange:1976:BRN

[Ora76] A. D. Orange. Book review: Nineteenth and Twentieth Centuries: The Corresponding Societies of the British Association for the Advancement of Science. 1883–1929. By R. M. MacLeod, J. R. Friday, and C. Gregor. London: Mansell Information/Publishing Ltd, 1975. Pp. xxii + 147. No price stated. British Journal for the History of Science, 9(1): 84, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025718.

Orange:1977:BRE

[Ora77]  A.  D.

Orange. Book review: Eighteenth and Nineteenth Centuries: William Scoresby Arctic Scientist. By Tom and Cordelia Stamp. Whitby: Caedmon of Whitby Press, 1976. Pp. xii + 253. £2.95. British Journal for the History of Science, 10 (1):84–85, March 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025596.

Orange:1980:BRN

[Ora80] A. D. Orange. Book review: Nineteenth Century: Science in Culture: The Early Victorian Period. By Susan Faye Cannon. Folkestone: Dawson; New York: Science History Publications, 1978. Pp. xii + 296.

£12.50/$17.95. British Journal for the History of Science, 13 (2):171–172, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025905.

Ordaz:1976:FST

[Ord76] J. Ordaz. The first Spanish translation of Lyell’s elements of geology. British Journal for the History of Science, 9(2):237–240, July 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025810.

Orel:1984:M

[Ore84] Vítezslav Orel. Mendel. Past masters. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1984. ISBN 0-19-287625-2, 0-19-287624-4 (paperback). vii + 111 pp. LCCN QH31.M45 O6813 1984. US$15.90, US$3.90.

Overmier:1992:BMB

[OS92a]  Judith  A.

Overmier and John Edward Senior. Books and manuscripts of the Bakken. Scarecrow Press, Metuchen, NJ, USA, 1992. ISBN 0-8108-2570-8. xvii + 512 + 4 pp. LCCN Z6664.8.E4 O94 1992.

Owen:1992:HLC

[OS92b]  Sir  Richard

Owen and Phillip R. Sloan.

The Hunterian lectures in comparative anatomy: May–June, 1837. Natural History Museum publications. Natural History Museum, London, UK, 1992. ISBN 0-565-01106-5 (hardcover), 0-565-01144-8 (paperback). 340 pp. LCCN M93.E02368; M93.E02370; QL808 OWE 1992.

Ogee:2006:DEC

[OS06] Frédéric Ogée and Anthony Strugnell, editors. Diderot and European culture, volume 2006:09 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2006. ISBN 0-7294-0883-3. xii + 246 pp. LCCN PQ1979 .D43 2006.

Osborne:1985:BRG

[Osb85] Catherine Osborne. Book review: G. S. Kirk, J. E. Raven, M. Schofield, The Presocratic Philosophers. 2nd edition. Cambridge: Cambridge University Press, 1983. Pp. xiii and 501. ISBN 0-521-25444-2 £30.00 (Paperback £10.95). British Journal for the History of Science, 18(1): 93–94, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026276.

Osborne:1989:BRG

[Osb89a] Michael A. Osborne. Book review: Goulven Laurent, Paléontologie et évolution en France de 1800 à 1860: Une histoire des idées de Cuvier et Lamarck à Darwin. Paris: Editions du Comité des Travaux Historiques et Scientifiques, 1987. Pp. xiv + 554. ISBN 2-7355-0126-4. 350 FF. British Journal for the History of Science, 22(2):250– 252, July 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026669.

Osborne:1989:BRS

[Osb89b] Michael A. Osborne. Book review: Scott Atran et al., Histoire du Concept d’Espèce dans les Sciences de la Vie Paris: Guathier-Villars, 1987. Pp. xiii + 324. ISBN 2-900927-19-6. FF 150.00. British Journal for the History of Science, 22(1):96–97, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026695.

Osborne:1994:NES

[Osb94] Michael A. Osborne. Nature, the exotic, and the science of French colonialism. Science, technology, and society. Indiana University Press, Bloomington, IN, USA, 1994. ISBN 0-253-34266-X. xvi + 216 pp. LCCN Q127.F8 O78 1994.

Osler:1991:APT

[Osl91]  Margaret  J.

Osler, editor. Atoms, pneuma, and tranquillity: Epicurean and Stoic themes in European thought. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-40048-1. xii + 304 pp. LCCN B181 .A76

                                        1991. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/90045068.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        90045068.html.

Ospovat:1974:BRG

[Osp74] Alexander M. Ospovat. Book review: Geology Dissertations on Steno as Geologist. Ed. by Gustav Scherz. Odense: Odense University Press, 1971. Pp. 319. No price stated. British Journal for the History of Science, 7(1):84–87, March 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025182.

Ospovat:1976:DWL

[Osp76] Alexander M. Ospovat. The distortion of Werner in Lyell’s principles of geology. British Journal for the History of Science, 9(2):190–198, July 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025806.

Osterbrock:1984:JKP

[Ost84] Donald E. Osterbrock. James E. Keeler, pioneer American astrophysicist, and the early development of American astrophysics. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-26582-7. xii + 411 pp. LCCN QB36.K37 O88

1984. URL http://www.loc.
gov/catdir/description/
cam022/84007727.html;
http://www.loc.gov/
catdir/samples/cam034/
84007727.html;   http://
www.loc.gov/catdir/toc/
cam021/84007727.html.

Oster:1989:BDA

[Ost89] Malcolm R. Oster. The ‘Beame of Diuinity’: Animal suffering in the early thought of Robert Boyle. British Journal for the History of Science, 22 (2):151–180, July 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026659.

Outram:1980:BRE

[Out80a] Dorinda Outram. Book review:

Eighteenth Century: L’Institution de la science et l’expérience du vivant: méthode et expérience à l’Académie royale des sciences, 1666–1793. By Claire Salomon-Bayet. Paris: Flammarion, 1978. Pp. 464. No price stated. British Journal for the History of Science, 13 (1):78–80, March 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026134.

Outram:1980:BRH

[Out80b] Dorinda Outram. Book review: Historiography of Science Epistemologia e storia della scienza: le svolte teoriche da Duhem a Bachelard. By Pietro Redondi. Milan: Feltrinelli Editore, 1978. Pp. 256. L. 8,500. British Journal for the History of Science, 13(3): 263, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026201.

Outram:1980:DF

[Out80c] Dorinda Outram. Darwinism in Florence. British Journal for the History of Science, 13 (2):157–159, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025894.

Outram:1980:PVF

[Out80d] Dorinda Outram. Politics and vocation: French science, 1793– 1830. British Journal for the History of Science, 13(1):27– 43, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026118.

Outram:1982:BRT

[Out82a] Dorinda Outram. Book review: Technology and Development Robert Fox and George Weisz (ed), The organization of science and technology in France 1808–1914. Cambridge: Cambridge University Press. Paris: Editions de la Maison des Sciences de l’Homme, 1980. Pp. x + 355. £15.00. British Journal for the History of Science, 15(1):90–91, March 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025937.

Outram:1982:IS

[Out82b] Dorinda Outram. Italian science. British Journal for the History of Science, 15(3):293– 301, November 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026223.

Outram:1983:BRB

[Out83] Dorinda Outram. Book review: Biography Charles B. Paul, Science and immortality: the Éloges of the Paris Academy of Sciences (1699– 1791), Berkeley, Los Angeles & London: University of California Press, 1981. Pp. xii + 203. £11.75. British Journal for the History of Science, 16 (1):106–108, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026111.

Outram:1984:BRHa

[Out84a] Dorinda Outram. Book review: History of Natural History: Madeleine Barthélemy-Madaule, Lamarck the mythical precursor: a study of the relations between science and ideology. Translated by M. H. Shank. Cambridge, Mass., and London: The M.I.T. Press, 1982. Pp. xv + 174. ISBN 0-262-02179-X. £12.25. British Journal for the History of Science, 17(3):319– 320, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026633.

Outram:1984:BRHb

[Out84b] Dorinda Outram. Book review: History of Natural History: Paul Lawrence Farber, The emergence of ornithology as a scientific discipline: 1760– 1850. Dordrecht, Boston & London: D. Reidel Publishing Company, 1982. Pp. xxi+ 191. 339.50. British Journal for the History of Science, 17 (3):321, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026634.

Outram:1984:GCV

[Out84c] Dorinda Outram. Georges Cuvier: vocation, science, and authority in post-revolutionary France. Manchester University Press, Manchester, UK, 1984. ISBN 0-7190-1077-2. viii + 299 pp. LCCN QH31.C89 O95 1984.

Outram:1985:BRJ

[Out85] Dorinda Outram. Book review:

L. J. Jordanova, Lamarck. Oxford: Oxford University Press (Past Masters), 1984. Pp. x + 118. ISBN 0-19-287588-4, £7.95 (hardcover). 0-19-287587-6, £1.95 (paperback). British Journal for the History of Science, 18(3): 368, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026404.

Outram:1987:BRM

[Out87] Dorinda Outram. Book review:

Margaret Alic. Hypatia’s Heritage: A History of Women in Science from Antiquity to the Late Nineteenth Century. London: The Women’s Press, 1986. Pp. ix + 230. ISBN 0-7043-3954-4. £4.95. British Journal for the History of Science, 20 (2):224–225, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026310.

Outram:1988:BRF

[Out88a] Dorinda Outram. Book review:

A. Fourcy. Histoire de l’Ecole Polytechnique. Paris: Belin, 1987. Pp. 509 + 198. ISBN 2-7011-0640-0. FF. 140.00. British Journal for the History of Science, 21(1):135, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026890.

Outram:1988:BRJ

[Out88b] Dorinda Outram. Book review: Janis Langins. La République avait Besoin de Savants: les Débuts de l’Ecole Polytechnique: l’Ecole Centrale des Travaux Publics et les Cours Revolutionnaires de l’an III. Paris: Belin, 1987. Pp. 287. ISBN 2-7011-0641-9. FF. 90.00. British Journal for the History of Science, 21(1):135, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Outram:1995:E

[Out95] Dorinda Outram. The Enlightenment, volume 7 of New approaches to European history. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-41522-5, 0-521-42534-4 (paperback). xv + 141 pp. LCCN B802 .O98

                                        1995. URL http://www.loc.
                                        gov/catdir/description/
                                        cam026/94035014.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam022/
                                        94035014.html.

Ogborn:2010:GB

[OW10] Miles Ogborn and Charles W. J. Withers, editors. Geographies of the book. Ashgate Pub., Farnham, England, 2010. ISBN 0-7546-7850-4 (hardcover), 0-7546-9675-8 (ebook). xiv + 302 pp. LCCN Z278 .G45 2010.

Pacey:1970:BRP

[Pac70]  A. J. Pacey.  Book re-

view: Physics The Conflict between Atomism and Conservation Theory 1644 to 1860. By Wilson L. Scott. London: Mac-Donald, and New York: Elsevier. 1970. Pp. xiv + 312. Illustr. £5. British Journal for the History of Science, 5(2): 191–192, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025423.

Pacey:1971:BRT

[Pac71] A. J. Pacey. Book review:

Technology Moving the Obelisks. By Bern Dibner. (Reprint of Burndy Library Publication no. 6, 1950.) Cambridge, Mass. & London: M.I.T. Press. 1970. Pp. 61. Illus. 95p. (paperback). British Journal for the History of Science, 5(3):294, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025322.

Pacey:1972:BRC

[Pac72] A. J. Pacey. Book review:

China Science and Civilisation in China. Volume 4, Part 3: Civil Engineering and Nantics. By Joseph Needham with Wang Ling and Lu Gwei-Djen. London: Cambridge University Press, 1971. Pp. lvii + 931 £18. British Journal for the History of Science, 6(2):

210–212, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025299.

Pacey:1974:SEH

[Pac74] A. J. Pacey. Some early heat engine concepts and the conservation of heat. British Journal for the History of Science, 7(2):135–145, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025200.

Pagel:1966:BRM

[Pag66]  Walter Pagel.  Book re-

view: Michael Scot. By Lynn Thorndike. Pp. 143. 4 plates. London: Thomas Nelson, 1965. 30s. British Journal for the History of Science, 3(2):189– 190, December 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025032.

Pagel:1967:ER

[Pag67] Walter Pagel. Essay review. British Journal for the History of Science, 3(4):395–396, December 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024964.

Pagel:1969:BRC

[Pag69] Walter Pagel. Book review: Chemistry Joachim Jungius’ Experimente und Gedanken zur Begründung der Chemie als Wissenschaft. Ein Beitrag zur Geistesgeschichte des siebzehnten Jahrhunderts. By Hans Kangro. (Boethius. Texte und Abhandlungen zur Geschichte der exakten Wissenschaften, vol. VII.) Wiesbaden: Franz Steiner. Pp. xxv + 479. Plates. 1968. DM. 90. British Journal for the History of Science, 4(4):409–411, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025241.

Pagel:1971:HC

[Pag71] Walter Pagel. A history of chemistry. British Journal for the History of Science, 5(4): 397–399, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025382.

Page:1976:RBC

[Pag76] Leroy E. Page. The rivalry between Charles Lyell and Roderick Murchison. British Journal for the History of Science, 9(2):156–165, July 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025803.

Pagel:1982:JBV

[Pag82] Walter Pagel. Joan Baptista van Helmont: reformer of science and medicine. Cambridge monographs on the history of medicine. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24807-8. x + 219 pp. LCCN R524.H4 P313

                                        1982. URL http://www.loc.
                                        gov/catdir/description/
                                        cam031/81024193.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        81024193.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam031/81024193.html.

Pais:1982:SLS

[Pai82] Abraham Pais. “Subtle is the Lord”: the science and the life of Albert Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1982. ISBN 0-19-853907-X. xvi + 552 + 9 pp. LCCN QC16.E5 P26 1982.

Pais:1986:IBM

[Pai86] Abraham Pais. Inward bound: of matter and forces in the physical world. Clarendon Press, New York, NY, USA, 1986. ISBN 0-19-851971-0, 0-19-851997-4 (paperback). xiv + 666 pp. LCCN QC7 .P27 1986. US$24.95.

Pancaldi:1991:DIS

[Pan91] Giuliano Pancaldi. Darwin in Italy: science across cultural frontiers. Indiana University Press, Bloomington, IN, USA, updated and expanded edition, 1991. ISBN 0-253-34287-2. xv + 222 pp. LCCN QH31.D2 P36313 1991.

Parry:1987:PVW

[Par87] G. J. R. Parry. A Protestant vision: William Harrison and the reformation of Elizabethan England. Cambridge studies in the history and theory of politics. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-32997-3. ix + 348 pp. LCCN BX5071 .P37

1987. URL http://www.loc.
gov/catdir/description/
cam023/86017091.html;
http://www.loc.gov/
catdir/samples/cam031/
86017091.html;   http://
www.loc.gov/catdir/toc/
cam031/86017091.html.

Patterson:1969:MS

[Pat69] Elizabeth C. Patterson. Mary Somerville. British Journal for the History of Science, 4(4): 311–339, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025228.

Patterson:1983:MSC

[Pat83a] Elizabeth Chambers Patterson.

Mary Somerville and the cultivation of science, 1815–1840, volume 102 of Archives internationales d’histoire des idées = International archives of the history of ideas. Martinus Nijhoff, Dordrecht and The Hague, The Netherlands, 1983. ISBN 90-247-2823-1. xiii + 264 pp. LCCN Q143.S72 P37 1983.

Pattison:1983:BVP

[Pat83b] Iain Pattison. The British veterinary profession, 1791–1948. J. A. Allen, London, UK, 1983. ISBN 0-85131-379-5. x + 207 + 4 pp. LCCN SF657 .P37 1984.

Patzlaff:1984:BRS

[Pat84] Martin M. Patzlaff. Book review: Science in Culture: Karl Snell, Die Schöpfung des Menschen. Vorlesungen über die Abstammung des Menschen (ed. Friedrich A. Kipp). Stuttgart: Verlag Freies Geistesleben, 1981. Pp. 230. DM 28. British Journal for the History of Science, 17(2):250– 251, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026584.

Patterson:1986:BRJ

[Pat86] Elizabeth C. Patterson. Book review: Jack Morrell & A. Thackray (eds). Gentlemen of Science: Early Correspondence of the British Association for the Advancement of Science. (Camden Fourth Series, Volume 30). London: Offices of the Royal Historical Society, University College, London, Cower Street WC 1, 1984. Pp. 382. ISBN 0-86193-103-3. £7.50. British Journal for the History of Science, 19(2): 224–225, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026608.

Patai:1994:JAH

[Pat94]  Raphael Patai.  The Jew-

ish alchemists: a history and source book. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03290-4. xiv + 617 pp. LCCN QD24.A2 P38 1994. US$49.50. URL http://

www.loc.gov/catdir/
description/prin031/
93035687.html;   http://
www.loc.gov/catdir/toc/
prin031/93035687.html.

Paul:1976:BRN

[Pau76] Harry W. Paul. Book review:

Nineteenth and Twentieth Centuries: Stendhal du Cote de la Science. By Jean Théodoridès. Aran, Switzerland: Editions du Grand Chêne, 1972. Pp. xi + 303. 36 Swiss francs. British Journal for the History of Science, 9(1):81–82, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025716.

Paul:1984:GAS

[Pau84] Robert Paul. German academic science and the Mandarin ethos, 1850–1880. British Journal for the History of Science, 17(1):1–29, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026014.

Paul:1993:MWG

[Pau93] Erich Robert Paul. The Milky Way galaxy and statistical cosmology, 1890–1924. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-35363-7 (hardcover). xiv + 262 pp. LCCN QB857.7 .P38

1993. URL http://www.loc.
gov/catdir/description/
cam025/92036530.html;
http://www.loc.gov/
catdir/toc/cam029/
92036530.html.

Payen:1974:WCR

[Pay74] Jacques Payen. From Watt to Clausius. The rise of thermodynamics in the early industrial age. British Journal for the History of Science, 7 (2):171–175, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025204.

Popper:1982:OUA

[PB82] Sir Karl R. (Karl Raimund) Popper and William Warren Bartley. The open universe: an argument for indeterminism: from the Postscript to the logic of scientific discovery, volume 2 of Postscript to the logic of scientific discovery / Karl R. Popper; Postscript to “The logic of scientific discovery” / Karl R. Popper; Postscript to “The logic of scientific discovery”. Hutchinson, London, UK, 1982. ISBN 0-09-146180-4 (hardcover) 0-09-146181-2 (paperback). xxii + 185 pp. LCCN QC6.2 .P6 1982.

Popper:1983:RAS

[PB83] Sir Karl Raimund Popper and William Warren Bartley. Realism and the aim of science: from the postscript of the logic of scientific discovery, volume 1 of Postscript to the logic of scientific discovery / Karl R. Popper. Hutchinson, London, UK, 1983. ISBN 0-09-151450-9, 0-09-151451-7. xxxix + 420 pp. LCCN

Q175. URL http://catalog.
hathitrust.org/api/
volumes/oclc/9771811.html.

Parnwell:1989:NOP

[PB^{+}89]  E.  C.  Parnwell,

Raymond Burns, et al. The new Oxford picture dictionary. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, english/spanish edition, 1989. ISBN 0-19-434355-3. iii + 144 pp. LCCN PE2835 .P37

1989. URL http://www.loc.
gov/catdir/enhancements/
fy1004/88034532-d.html.

Poggi:1994:RSS

[PB94] Stefano Poggi and Maurizio Bossi, editors. Romanticism in science: science in Europe, 1790–1840, volume 152 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2336-X, 94-017-2921-2.

ISSN 0068-0346. xv + 245 pp. LCCN Q174 .B67 vol.

152. URL http://www.loc.
gov/catdir/enhancements/
fy0821/93001728-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0821/93001728-t.html;
https://link.springer.
com/book/10.1007/978-94-
017-2921-5.

Porter:1999:HCP

[PB99] Roy S. Porter and German E. Berrios, editors. A history of clinical psychiatry: the origin and history of psychiatric disorders. New Brunswick, N.J., London, UK, 1. paperback edition, 1999. ISBN 0-485-24211-7 (paperback), 0-485-24011-4 (hardcover). xx + 684 pp. LCCN ????

Paterson:1991:SWW

[PCA91]  Sir  Clifford  Paterson,

Sir Robert Clayton, and Joan Algar. A scientist’s war: the war diary of Sir Clifford Paterson, 1939–45: 1st September 1939–9th May 1945, volume 14 of IEE history of technology series. P. Peregrinus, in association with the Science Museum, London, UK, 1991. ISBN 0-86341-218-1. xxvi + 674 pp. LCCN TK140.P38 A3 1991.

Pedersen:1985:BRL

[Ped85] Lawrence Pedersen. Book review: Lorin Anderson, Charles Bonnet and the Order of the Known. (Studies in the History of Modern Science, 11.)

Dordrecht: Reidel, 1982. Pp. xiii + 155. ISBN 90-277-1389-8. DFl. 85, $37. British Journal for the History of Science, 18 (1):103–104, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026286.

Pedersen:1992:BN

[Ped92a] Olaf Pedersen. The book of nature. Vatican Observatory Publications, Vatican City State, 1992. ISBN 0-268-00690-3. vi + 92 pp. LCCN BL240.2 .P354 1992.

Pedersen:1992:LLH

[Ped92b] Olaf Pedersen. Lovers of learning: a history of the royal Danish Academy of Sciences and Letters 1742–1992. Munksgaard, Copenhagen, Denmark, 1992. ISBN 87-7304-236-6. 348 pp. LCCN ????

Pelling:1985:BRI

[Pel85a] Margaret Pelling. Book review:

Iain Pattison, The British Veterinary Profession 1791–1948. London: J. A. Allen, 1983. Pp. x + 207. ISBN 0-85131-379-5. £9.50. British Journal for the History of Science, 18 (3):369, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026405.

Pelling:1985:BRW

[Pel85b] Margaret Pelling. Book review:

W. E. van Heyningen and John R. Seal. Cholera. The American Scientific Experience, 1947– 1980. Boulder, Colorado: Westview Press, 1983. Pp. xviii + 343. £24.00. British Journal for the History of Science, 18(2):243, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026342.

Penrose:1989:AAB

[Pen89]  Harald Penrose.  An an-

cient air: a biography of John Stringfellow of Chard, the Victorian aeronautical pioneer. Smithsonian Institution Press, Washington, DC, USA, 1989. ISBN 0-87474-752-X. 183 pp. LCCN TL540.S837 P46 1989.

Pepper:1967:LNT

[Pep67] Jon V. Pepper. A letter from Nathaniel Torporley to Thomas Harriot. British Journal for the History of Science, 3(3): 285–290, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025053.

Pepper:1971:BRS

[Pep71] Jon V. Pepper. Book review: Sixteenth and Seventeenth Centuries: Simon Stevin. Science in the Netherlands around 1600. By E. J. Dijksterhuis. The Hague: Martinus Nijhoff. 1970. Pp. x + 145. 5 plates. 27 guilders. British Journal for the History of Science, 5(4):

416–417, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025403.

Pepper:1972:BRS

[Pep72] Jon V. Pepper. Book review: Sixteenth and Seventeenth Centuries: The Mathematical Papers of Isaac Newton. Volume IV, 1674–1684. Edited by D. T. Whiteside, with the assistance in publication of M. A. Hoskin and A. Prag. London: Cambridge University Press, 1971. Pp. xxxiv + 678. £18. British Journal for the History of Science, 6(2): 216–217, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025304.

Pepper:1986:BRJ

[Pep86] Jon V. Pepper. Book review: John W. Shirley. Thomas Harriot: a Biography. Oxford: Clarendon Press, 1983. Pp. xii + 508. ISBN 0-19-822901-1. £25.00. British Journal for the History of Science, 19(2): 212–216, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026601.

Peppas:1989:OHY

[Pep89] Nicholas A. Peppas, editor.

One hundred years of chemical engineering: from Lewis M. Norton (M.I.T. 1888) to present: [papers presented at the American Chemical Society symposium held in Toronto, Canada, 6/5–10], volume 9 of Chemists and chemistry. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 0-7923-0145-5. LCCN TP165 .O54 1989.

Perrin:1970:EOP

[Per70] Carleton Perrin. Early opposition to the phlogiston theory: Two anonymous attacks. British Journal for the History of Science, 5(2):128–144, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025414.

Perrin:1973:LMS

[Per73]  C. E. Perrin.  Lavoisier,

Monge, and the synthesis of water, a case of pure coincidence? British Journal for the History of Science, 6(4): 424–428, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025504.

Pereira:1989:ACA

[Per89a]  Michela Pereira.  The al-

chemical corpis attributed to Raymond Lull, volume 18 of Warburg Institute surveys and texts. Warburg Institute, University of London, London, UK, 1989. ISBN 0-85481-078-1 (paperback). ISSN 0266-1772. 118 pp. LCCN B765.L84; Z8527.7 .P47 1989. URL http://

catalog.hathitrust.org/
api/volumes/oclc/
20485312.html.

Perrin:1989:DTM

[Per89b] C. E. Perrin. Document, text and myth: Lavoisier’s crucial year revisited. British Journal for the History of Science, 22(1):3–25, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026676.

Pera:1992:AFG

[Per92] Marcello Pera. The ambiguous frog: the Galvani–Volta controversy on animal electricity. Princeton University Press, Princeton, NJ, USA, 1992. ISBN 0-691-08512-9. xxvi + 203 pp. LCCN QP341 .P4713 1992. US$29.95. URL http:

//www.loc.gov/catdir/
description/prin031/
91003077.html.

Pestre:1984:PPF

[Pes84] Dominique Pestre. Physique et physiciens en France, 1918– 1940. (French) [Physics and medical doctors in France, 1918–1940]. Histoire des sciences et des techniques, 0761-1102. Editions des archives contemporaines, Paris, France, 1984. ISBN 2-903928-08-8. xv + 356 pp. LCCN QC9.F7 P47x 1984. 225.00F.

Petit:1984:LSA

[Pet84] Annie Petit. L’esprit de la science anglaise et les français au XIXème siècle. British Journal for the History of Science, 17(3):273–293, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026623.

Petroski:1989:PHD

[Pet89] Henry Petroski. The pencil: a history of design and circumstance. Faber and Faber, London, UK, 1989. ISBN 0-571-16182-0. xi + 434 pp. LCCN ????

Peterfreund:1990:LST

[Pet90] Stuart Peterfreund, editor. Literature and science: theory and practice. Northeastern University Press, Boston, MA, USA, 1990. ISBN 1-55553-058-3. vi + 248 pp. LCCN PN55 .L5 1990. US$37.50 (est.).

Petroski:1994:DPC

[Pet94]  Henry  Pet-

roski. Design paradigms: case histories of error and judgment in engineering. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-46108-1, 0-521-46649-0 (paperback). xii + 209 pp. LCCN TA174 .P473

                                        1994. URL http://www.loc.
                                        gov/catdir/description/
                                        cam025/93032560.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam024/
                                        93032560.html.

Pacey:1967:DBV

[PF67] A. J. Pacey and S. J. Fisher.

Daniel Bernoulli and the vis viva of compressed air. British Journal for the History of Science, 3(4):388–392, December 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024962.

Phillips:1967:BRH

[Phi67] E. D. Phillips. Book review:

History and Philosophy of Science George Sarton. De mens en zijn werk uit brieven aan vrienden en kennissen. By Paul van Oye. Verhandelingen van de Koninklijke Vlaamse Academie voor Wetenschappen, Letteren en Schone Kunsten van Belgie. Klasse der Wetenschappen. Jaargang XXVII, No. 82, 1965. Pp. 166. 13 Plates. 350 Belgian francs. British Journal for the History of Science, 3 (4):397, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024966.

Phillips:1969:BRS

[Phi69]  E. D. Phillips.  Book re-

view: Science and Medicine in Classical Antiquity Ancient Medicine. Selected Papers of Ludwig Edelstein. Ed. by Owsei Temkin and C. Lilian Temkin. Johns Hopkins Press and Oxford University Press, 1967. Pp. xiv + 496. £5 19s. British Journal for the History of Science, 4(3): 295–296, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025488.

Phillips:1987:WHE

[Phi87] Vivian J. Phillips. Waveforms: a history of early oscillography. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-274-6. xii + 259 pp. LCCN TK381 .P47

                                        1987. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0745/87180307-d.html.

Pauli:1979:WPW

[PHvMW93] Wolfgang Pauli, Armin Hermann, K. von Meyenn, and Victor F. (Victor Frederick) Weisskopf. Wolfgang Pauli, wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a.. (German) [Wolfgang Pauli, scientific correspondence with Bohr, Einstein, Heisenberg, and others], volume 2, 6, 11 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979–1993. ISBN 0-387-08962-4, 3-540-08962-4, 0-387-13609-6, 3-540-13609-6, 0-387-54911-0, 3-540-54911-0. ???? pp. LCCN QC16.P37 W64.

Pickstone:1976:BRN

[Pic76] J. V. Pickstone. Book review: Nineteenth and Twentieth Centuries: American Physicians in the Nineteenth Century. From Sects to Science. By William G. Rothstein. Baltimore and London: Johns Hopkins University Press, 1972. Pp. xv + 362. £6.75. British Journal for the History of Science, 9 (1):77–80, March 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025714.

Pickstone:1978:LD

[Pic78] J. V. Pickstone. Locating Dutrochet. British Journal for the History of Science, 11 (1):49–64, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025607.

Pickstone:1979:SNB

[Pic79] J. V. Pickstone. Short notices of books: Insects, Hygiene and History. By J. R. Busvine. London: Athlone Press, 1976. Pp.

262. £6.95. British Journal for the History of Science, 12 (2):235–236, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025769.

Pickstone:1980:BRN

[Pic80] J. V. Pickstone. Book review:

Nineteenth and Twentieth Centuries: Organized Medicine in the Progressive Era. The Move toward Monopoly. By James G.

Burrow. Baltimore: Johns Hopkins University Press, 1978. Pp. ix + 218. £9.75. British Journal for the History of Science, 13 (3):276, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026211.

Pickering:1984:BRN

[Pic84a] Andrew Pickering. Book reviews: National Traditions in Science Everett Mendelsohn and Yehuda Elkana (eds), Sciences and cultures. Dordrecht, Boston and London: Reidel, 1981. Pp. xvii + 270. ISBN 90-277-1234-4 (hardcover)/1235-2 (paperback). (No price stated.) Norbert Elias, Herminio Martins and Richard Whitley (eds), Scientific Establishments and hierarchies. Reidel, 1982. Pp. xi + 368. ISBN 90-277-1322-7 (hardcover)/1323-5 (paperback). Dfl. 95.00, $42.50 (hardcover); Dfl. 50.00, $19.95 (paperback). British Journal for the History of Science, 17(1): 100–101, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026036.

Pickering:1984:CQS

[Pic84b] Andrew Pickering. Constructing quarks: a sociological history of particle physics. Edinburgh University Press, Edinburgh, Scotland, 1984. ISBN 0-85224-458-4, 0-85224-535-1 (paperback). xi + 468 pp. LCCN QC793.16 .P53 1986. US$20.00. URL http:

//catalog.hathitrust.org/
api/volumes/oclc/
11167232.html.

Pickering:1987:BRH

[Pic87a] Andy Pickering. Book review: Henk Zandvoort. Models of Scientific Development and the Case of Nuclear Magnetic Resonance. Dordrecht: D. Reidel, 1986. Pp. xiii + 305. ISBN 90-277-2351-6. £49.50, Dfl. 115.00. British Journal for the History of Science, 20(4): 481–482, October 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026422.

Pickering:1987:FLS

[Pic87b] Andy Pickering. Forms of life: Science, contingency and Harry Collins. British Journal for the History of Science, 20 (2):213–221, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026308.

Picolet:1987:JPD

[Pic87c]  Guy  Picolet,  edi-

tor. Jean Picard et les débuts de l’astronomie de précision au XVIIe siècle: actes du colloque du tricentenaire. (French) [Jean Picard and the origins of precision astronomy in the 17th Century: acts of the tricentennial conference]. Editions du Centre national de la recherche scientifique, Paris, France, 1987. ISBN 2-222-04104-X. 382 pp. LCCN QB36.P45 J43 1987. 150F.

Pickstone:1988:BRD

[Pic88a] John V. Pickstone. Book review: S. E. D. Shortt. Victorian Lunacy. Richard M. Bucke and the Practice of Late Nineteenth-Century Psychiatry. Cambridge: Cambridge University Press, 1986. Pp. xvi + 207. ISBN 0-521-30999-9. Price £25.00, $29.95. British Journal for the History of Science, 21(1):120, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026876.

Pickstonk:1988:BRJ

[Pic88b] John V. Pickstonk. Book review: John Harley Warner. The Therapeutic Perspective: Medical Practice, Knowledge, and Identity in America 1820– 1885. Cambridge, MA: Harvard University Press, 1986. Pp. 367. ISBN 0-674-88330-6. £27.50. British Journal for the History of Science, 21(1): 126–127, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026882.

Pinch:1983:BRA

[Pin83]  T.  J.

Pinch. Book review: Alchemy and Parapsychology Seymour H. Mauskopf and Michael R. McVaugh, The elusive science: origins of experimental psychical research, Baltimore & London: Johns Hopkins University Press, 1981. Pp. xvi + 368. £14.75/$30.60. British Journal for the History of Science, 16 (1):98–99, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026106.

Pinch:1987:BRJ

[Pin87a]  Trevor Pinch.  Book re-

view: J. A. Schuster & R. R. Yeo (eds). The Politics and Rhetoric of Scientific Method. Dordrecht: Reidel, 1986. Pp. xxxix + 305. ISBN 90-277-2152-1. £42.45. British Journal for the History of Science, 20(3):356–357, July 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026367.

Pinch:1987:BRL

[Pin87b] Trevor Pinch. Book review:

Lewis Pyenson. The Young Einstein: The Advent of Relativity. Bristol and Boston: Adam Hilger, 1985. Pp. xiv + 246. ISBN 0-85274-779-9. £19.95, $28.00. British Journal for the History of Science, 20 (1):89–91, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026446.

Pinch:1988:BRA

[Pin88] Trevor Pinch. Book review: Allan Franklin. The Neglect of Experiment. Cambridge: Cambridge University Press, 1986. Pp. xii + 290. ISBN 0-521-32016-X. £30.00. British Journal for the History of Science, 21(1):122–123, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026878.

Pirie:1983:LE

[Pir83] N. W. Pirie. Letter to the Editor. British Journal for the History of Science, 16(3): 273–274, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026147.

Pittenger:1993:ASE

[Pit93] Mark Pittenger. American socialists and evolutionary thought, 1870–1920. History of American thought and culture. University of Wisconsin Press, Madison, WI, USA, 1993. ISBN 0-299-13600-0, 0-299-13604-3 (paperback). x + 310 pp. LCCN HX83 .P58 1993. US$60.00, US$24.95.

Pugh:1991:IES

[PJP91] Emerson W. Pugh, Lyle R.

Johnson, and John H. Palmer. IBM’s 360 and early 370 systems. History of computing. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-16123-0. xx + 819 pp. LCCN QA76.8.I12 P84 1991.

Parvum:1972:E

[PL72] Almagestum Parvum and R. P.

Lorch. Enquiries. British Journal for the History of Science, 6(1):106, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025285.

Paret:1992:PIP

[PLP92] Peter Paret, Beth Irwin Lewis, and Paul Paret. Persuasive images: posters of war and revolution from the Hoover Institution archives. Princeton University Press, Princeton, NJ, USA, 1992. ISBN 0-691-03204-1. xv + 233 pp. LCCN D522.25 .P37 1992. US$29.95. URL http://www.

loc.gov/catdir/
description/prin031/
92004388.html;   http://
www.loc.gov/catdir/toc/
prin031/92004388.html.

Price:1995:SWG

[PN95] George McCready Price and Ronald L. Numbers. Selected works of George Mc-Cready Price, volume 7 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1808-1. xvii + 489 pp. LCCN BS651 .P77 1995.

Pnina:1982:MB

[Pni82] Abir-Am Pnina. Modern biochemistry. British Journal for the History of Science, 15(3): 301–305, November 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026224.

Pocock:1988:EBR

[Poc88]  Rowland F. Pocock.  The

early British radio industry. Manchester University Press, Manchester, UK, 1988. ISBN 0-7190-2621-0. viii + 184 pp. LCCN HD9696.R363 G747 1988. US$35.00 (est.).

Poirier:1993:ALL

[Poi93]  Jean Pierre Poirier.  An-

toine Laurent de Lavoisier, 1743–1794. Pygmalion, Paris, France, 1993. ISBN 2-85704-384-8. x + 545 pp. LCCN QD22.L4 P65 1993. 178F.

Poirson:2004:AAF

[Poi04] Martial Poirson, editor. Art et argent en France au temps des premiers modernes (XVIIe et XVIIIe siècles). (French) [Art and money in France at the time of the first modern 17th and 18th Centuries], volume 2004:10 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2004. ISBN 0-7294-0849-3. xi + 341 pp. LCCN PQ245 .A77 2004.

Poole:1998:TAC

[Poo98]  Robert Poole.  Time’s al-

teration: calendar reform in early modern England. UCL Press, London, UK, 1998. ISBN 1-85728-622-7. xix + 243 pp. LCCN CE73 .P66

1998. URL http://www.loc.
gov/catdir/enhancements/
fy0909/99159352-d.html.

Popper:1982:QTS

[Pop82] Karl R. (Karl Raimund) Popper. Quantum theory and the schism in physics. The Postscript to The logic of scientific discovery. Unwin Hyman, London, UK, 1982. ISBN 0-04-445555-0 (paperback), 0-09-146170-7 (hardcover). 230 pp. LCCN QC174.12 .P67 1982.

Porter:1976:CLP

[Por76a] Roy Porter. Charles Lyell and the principles of the history of geology. British Journal for the History of Science, 9 (2):91–103, July 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025798.

Porter:1976:SU

[Por76b] Roy Porter. Science and the universities. British Journal for the History of Science, 9(3): 320–323, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025685.

Porter:1977:BRG

[Por77] Roy Porter. Book review: General Reason, Ridicule and Religion, The Age of Enlightenment in England, 1660–1750. By John Redwood. London: Thames and Hudson, 1976. Pp.

287. £7.00. British Journal for the History of Science, 10(3): 269–270, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025856.

Porter:1978:BRN

[Por78a] Roy Porter. Book review: Nineteenth and Twentieth Centuries: Fossils and Progress. Paleontology and the Idea of Progressive Evolution in the Nineteenth Century. By Peter J. Bowler. New York: Science History Publications, 1976. Pp. viii + 191 + XIV plates. British Journal for the History of Science, 11(1):83–85, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025624.

Porter:1978:BRS

[Por78b] Roy Porter. Book review: Science and Literature The Lay of the Land. Metaphor as Experience and History in American Life and Letters. By Annette Kolodny. Chapel Hill: The University of North Carolina Press, 1975. Pp. xiv + 186. No price stated. British Journal for the History of Science, 11 (2):178–179, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025737.

Porter:1979:BRN

[Por79a] Roy Porter. Book review:

Nineteenth and Twentieth Centuries: Charles Darwin: ‘storia’ ed ‘economia’ della natura. By Giuliano Pancaldi. Firenze: La Nuova Italia, 1977. Pp. 212. L. 4,000. British Journal for the History of Science, 12(1): 99–100, March 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025673.

Porter:1979:BRS

[Por79b] Roy Porter. Book review: Seventeenth and Eighteenth Centuries: Dr. Woodward’s Shield. History, Science and Satire in Augustan England. By Joseph M. Levine. Berkeley, Los Angeles & London: University of California Press, 1977. Pp. x + 362. $19.75/£14.75. British Journal for the History of Science, 12(2):227–228, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025762.

Porter:1981:BRG

[Por81] Roy Porter. Book review: Geology Two Hundred Years of Geology in America. Proceedings of the New Hampshire Bicentennial Conference on the History of Geology. Ed. by Cecil J. Schneer. Hanover, New Hampshire: University Press of New England, 1979. Pp. xvii + 385. $20.00. British Journal for the History of Science, 14(2):209–210, July 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025952.

Porter:1982:BRB

[Por82] Roy Porter. Book review: Biography Stephen J. Pyne, Grove Karl Gilbert: a great engine of research. Austin and London: University of Texas Press, 1980. Pp. xiv + 306. £12.00. British Journal for the History of Science, 15(1):79–81, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025930.

Porter:1983:BRS

[Por83a] Roy Porter. Book review: Science and Society Margarita Bowen, Empiricism and geographical thought: from Francis Bacon to Alexander von Humboldt. Cambridge: Cambridge University Press, 1981. Pp. 351. £25.00. British Journal for the History of Science, 16(3):301–302, November 1983.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026171.

Porter:1983:ESA

[Por83b] Roy Porter. The earth sciences:

an annotated bibliography, volume 3; vol. 315 of Bibliographies of the history of science and technology; Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1983. ISBN 0-8240-9267-8. xix + 192 + 15 pp. LCCN Z6031 .P67 1983. US$30.00.

Porter:1984:BRE

[Por84a] Roy Porter. Book review: Enlightenment Marguerite Carozzi, Voltaire’s attitude toward geology. Geneva: Société de Physique et d’Histoire Naturelle, 1983. Pp. 146. SFRp 28. British Journal for the History of Science, 17(1):116– 117, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026054.

Porter:1984:BRH

[Por84b] Roy Porter. Book review: History of Natural History: John Lyon and Phillip R. Sloan, From natural history to the history of nature. Notre Dame and London: University of Notre Dame Press, 1981. Pp. xiv+ 406. £11.95. David Goodman, Buffon’s Natural History. (The Open University Arts: A Second Level Course, The Enlightenment.) Milton Keynes: The Open University Press, 1980. Pp. 74. British Journal for the History of Science, 17(3): 321–322, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026635.

Porter:1985:BRL

[Por85a] Roy Porter. Book review: Lloyd G. Stevenson (ed.), A celebration of medical history. The fiftieth anniversary of the Johns Hopkins Institute of the History of Medicine and the Welch Medical Library. Baltimore and London: The John Hopkins University Press, 1982. Pp. vi + 228. £12.75/ 522.10. British Journal for the History of Science, 18(1):83– 84, March 1985. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Porter:1985:BRS

[Por85b] Roy Porter. Book review: Sally Shuttleworth, George Eliot and Nineteenth Century Science. The Make-Believe of a Beginning. Cambridge: Cambridge University Press, 1984. Pp. xiv + 257, £20.00 Redmond O’Hanlon, Joseph Conrad and Charles Darwin. The Influence of Scientific Thought on Conrad’s Fiction. Edinburgh: The Salamander Press, 1984. Pp. 189, £17.50. British Journal for the History of Science, 18 (1):107–109, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026289.

Porter:1985:RU

[Por85c] Roy Porter. Review: [untitled]. British Journal for the History of Science, 18(1): 84, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026268.

Porter:1985:MSV

[Por85d] Theodore M. Porter. The mathematics of society: Variation and error in Quetelet’s statistics. British Journal for the History of Science, 18(1): 51–69, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026265.

Porter:1986:BRD

[Por86a] Roy Porter. Book review:

Dorothy A. Stansfield. Thomas Beddoes MD, 1760–1808. Dordrecht, Boston, Lancaster: D. Reidel, 1984. Pp. xix + 306. ISBN 90-277-1686-2. Dfl. 140. British Journal for the History of Science, 19(1):121– 122, March 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026494.

Porter:1986:RST

[Por86b] Theodore M. Porter. The rise of statistical thinking, 1820– 1900. Princeton University Press, Princeton, NJ, USA, 1986. ISBN 0-691-08416-5. xii + 333 pp. LCCN QA276.15 .P67 1986. US$35.00.

Porter:1987:BRJb

[Por87a]  Roy Porter.  Book re-

view: James Paradis & Thomas Postlewait (eds). Victorian Science and Victorian Values. Literary Perspectives. New Jersey: Rutgers University Press, 1985. Pp. xiii + 362. British Journal for the History of Science, 20(1):107–108, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026461.

Porter:1987:BRJa

[Por87b] Roy Porter. Book review: Jeffrey Moussaieff Masson, (ed. and tr.) The Complete Letters of Sigmund Freud to Wilhelm Fliess, 1887–1904. Cambridge, Mass. and London: Harvard University Press, 1985. Pp. xv + 505. ISBN 0-674-15420-7. £19.95. British Journal for the History of Science, 20(1): 100–101, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026454.

Porter:1987:BRS

[Por87c] Roy Porter. Book review:

Sander L. Gilman. Difference and Pathology. Stereotypes of Sexuality, Race and Madness. Ithaca: Cornell University Press, 1985. Pp. 292. $38.45 (hardcover), $14.25 (paperback). British Journal for the History of Science, 20(1): 104–105, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026458.

Porter:1988:BRP

[Por88] Roy Porter. Book review:

Phillip A. Nicholls. Homoeopathy and the Medical Profession, Beckenham: Croom Helm, 1988. Pp. 298. ISBN 0-7099-1836-4. £27.50. British Journal for the History of Science, 21 (3):368–369, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026841.

Porgan:1989:BRW

[Por89a] Sophie Porgan. Book review:

Wolfgang Shivelbusch. Disenchanted Night: The Industrialisation of Light in the Nineteenth Century. Oxford, New York, Hamburg: Berg Publishers Ltd., 1988. Pp. 227. ISBN 0-85496-191-7. £18.50. British Journal for the History of Science, 22(3):393–394, September 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026910.

Porter:1989:BRF

[Por89b] Dorothy Porter. Book review: F. B. Smith. The Retreat of Tuberculosis 1850– 1950. London: Croom Helm, 1988. Pp. 271. ISBN 0-7099-3383-5. £25.00. British Journal for the History of Science, 22(1):93–95, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026693.

Porter:1989:BRLa

[Por89c] Dorothy Porter. Book review: Linda Bryder. Below the Magic Mountain. A Social History of Tuberculosis in Twentieth-Century Britain. Oxford: Clarendon Press, 1988. Pp. xv + 298. ISBN 0-19-822947-X. £30.00. British Journal for the History of Science, 22(1):89–90, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026689.

Porter:1989:BRC

[Por89d] Dorothy E. Porter. Book review: Charles Webster. The Health Services Since the War. Volume 1. Problems of Health Care. The National Health Service Before 1957. London: H.M.S.O., 1988. Pp. x + 479. ISBN 0-11-630942-3. No price given. British Journal for the History of Science, 22(4):478– 479, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026948.

Porter:1989:BRD

[Por89e] Roy Porter. Book review:

Daniel Roche. Les Republicains des Lettres. Gens de Culture et Lumières au XVIII Siècle. Paris: Fayard, 1988. Pp. 393. ISBN 2-213-02122-8. 160 FF (paperback). British Journal for the History of Science, 22 (4):463–464, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026939.

Porter:1989:HSQ

[Por89f] Roy Porter. Health for sale:

quackery in England, 1660– 1850. Manchester University Press, Manchester, UK, 1989. ISBN 0-7190-1903-6. xi + 280 pp. LCCN R730 .P67 1989.

Porter:1989:BRLb

[Por89g] Theodore M. Porter. Book review: Lorraine Daston. Classical Probability in the Enlightenment. Princeton: Princeton University Press, 1988. Pp. xviii + 423. ISBN 0-691-08497-1, £27.50, $49.50. British Journal for the History of Science, 22(4):444–446, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026924.

Porter:1995:TNP

[Por95] Theodore M. Porter. Trust in numbers: the pursuit of objectivity in science and public life. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03776-0. xiv + 310 pp. LCCN Q175.5 .P67

                                        1995. URL http://www.loc.
                                        gov/catdir/description/
                                        prin031/94021440.html;
                                        http://www.loc.gov/
                                        catdir/toc/prin031/
                                        94021440.html.

Poundstone:1993:PDJ

[Pou93] William Poundstone. Prisoner’s dilemma: John von Neumann, game theory, and the puzzle of the bomb. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-286162-X (paperback), 0-385-41580-X. xi + 294 + 8 pp. LCCN QA29.V66 P68 1993. US$12.95 (US$15.95 Can.).

Powers:1982:PNP

[Pow82] Jonathan Powers. Philosophy and the new physics, volume 792 of Ideas; University paperbacks. Methuen, London, UK, 1982. ISBN 0-416-73470-7, 0-416-73480-4 (paperback). xvii + 203 pp. LCCN QC6 .P694 1982.

Power:1996:ASN

[Pow96] Michael Power, editor. Accounting and science: natural inquiry and commercial reason, volume 26 of Cambridge studies in management. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55325-3 (hardcover), 0-521-55699-6 (paperback). xvii + 288 pp. LCCN HF5657 .A2565 1996. URL http://

www.loc.gov/catdir/
description/cam027/
95038403.html;   http://
www.loc.gov/catdir/toc/
cam022/95038403.html.

Poynter:1970:BRB

[Poy70] F. N. L. Poynter. Book review: Biology and Medicine John Hunter. By Jessie Dobson. Edinburgh and London: E. & S. Livingstone. 1969. Pp. xvii + 361. Plates. 50s. British Journal for the History of Science, 5(2):199, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025430.

Pare:1982:MM

[PP82] Ambroise Paré and Janis L.

Pallister. On monsters and marvels. University of Chicago Press, Chicago, IL, USA, 1982. ISBN 0-226-64562-2. xxxii + 224 + 1 pp. LCCN QM691 .P3713 1982. US$16.50.

Pitt:1987:RCS

[PP87] Joseph C. Pitt and Marcello Pera, editors. Rational Changes in Science: Essays on Scientific Reasoning, volume 98 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1987. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346. xiii + 224 pp. LCCN Q174 .B67 vol. 98. US$64.00. URL https:

//link.springer.com/book/
10.1007/978-94-009-3779-6.

Panza:1995:SEV

[PP95] Marco Panza and Jean-Claude Pont, Dr ès sciences mathématiques, editors. Les savants et l’épistémologie vers la fin du XIXe siècle. (French) [Scientists and epistemology at the end of the 19th Century]. Bibliothèque scientifique Albert Blanchard. Libr. scientifique et technique Albert Blanchard, Paris, France, 1995. ISBN 2-85367-196-8. 282 pp. LCCN B2185 .S38 1995.

Perez-Ramos:1988:FBI

[PR88] Antonio Pérez-Ramos. Francis Bacon’s idea of science and the maker’s knowledge tradition. Clarendon Press, New York, NY, USA, 1988. ISBN 0-19-824979-9. xi + 334 pp. LCCN B1198 .P44 1988. US$32.50.

URL http://www.loc.gov/
catdir/enhancements/
fy0637/88009835-d.html.

Pietsch:1995:FCC

[PR95] Theodore W. Pietsch and Louis Renard, editors. Fishes, crayfishes, and crabs: Louis Renard’s natural history of the rarest curiosities of the seas of the Indies. Foundations of natural history. The Johns Hopkins University Press, Baltimore, MD, USA, 1995. ISBN 0-8018-4790-7 (set : hardback). ???? pp. LCCN QL334.E27

F57 1995. URL http://www.
loc.gov/catdir/bios/
jhu051/93049411.html;
http://www.loc.gov/
catdir/description/
jhu051/93049411.html.

Prag:1970:BRM

[Pra70] A. Prag. Book review: Mathematics: A History of Mathematics. By Carl B. Boyer. New York & London: John Wiley & Sons. 1968. Pp. xv + 717. 97s. British Journal for the History of Science, 5(1):89, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025357.

Price:1983:MRF

[Pri83] Roger Price. The modernization of rural France: communications networks and agricultural market structures in Nineteenth-Century France. [Hutchinson university library]. Hutchinson, London, UK, 1983. ISBN 0-09-145180-9. 503 pp. LCCN HD1945 .P737x 1983b.

Provine:1986:SWE

[Pro86] William B. Provine. Sewall Wright and evolutionary biology. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1986. ISBN 0-226-68474-1. xvi + 545 pp. LCCN QH361 .P87 1986.

Proctor:1988:RHM

[Pro88] Robert Proctor. Racial hygiene: medicine under the Nazis. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-74580-9. viii + 414 pp. LCCN RA418.3.G3 P76 1988.

Percival:1978:BM

[PS78a] Janet Percival and W. A. Smeaton. Back matter. British Journal for the History of Science, 11(2):196, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025748.

Percival:1978:LAR

[PS78b] Janet Percival and W. A. Smeaton. Library and archive resources in the history of science at University College London. British Journal for the History of Science, 11(2):191– 195, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025747.

Palm:1982:AVL

[PS82] Lodewijk C. Palm and Harry A. M. Snelders, editors. Antoni van Leeuwenhoek 1632– 1723: studies on the life and work of the Delft scientist commemorating the 350th anniversary of his birthday, volume 8 of Nieuwe Nederlandse bijdragen tot de geschiedenis der geneeskunde en der natuurwetenschappen. Rodopi, Amsterdam and Leiden, The Netherlands, 1982. ISBN 90-6203-824-7 (paperback). LCCN QH31.L55 A57

1982. URL http://catalog.
hathitrust.org/api/
volumes/oclc/9892184.html.

Prigogine:1984:OCMa

[PS84a] I. (Ilya) Prigogine and Isabelle Stengers. Order out of chaos: man’s new dialogue with nature. Heinemann, London, UK, 1984. ISBN 0-434-60395-3. xxxi + 349 pp. LCCN Q175 PRI 1984.

Prigogine:1984:OCMb

[PS84b] I. (Ilya) Prigogine and Isabelle Stengers. Order out of chaos: man’s new dialogue with nature. Bantam Books, New York, NY, USA, 1984. ISBN 0-553-34082-4. xxxi + 349 pp. LCCN Q175.P8823; Q175 .P8823 1984.

Pera:1991:PSA

[PS91] Marcello Pera and William R.

Shea, editors. Persuading science: the art of scientific rhetoric. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1991. ISBN 0-88135-071-0. xi + 212 pp. LCCN Q175 .P384 1991. US$40.00.

Ptolemy:1984:PA

[PT84] Ptolemy and G. J. Toomer. Ptolemy’s Almagest. Duckworth classical, medieval, and renaissance editions. Duckworth, London, UK, 1984. ISBN 0-7156-1588-2. ix + 693 pp. LCCN QB41 .P957x 1984b. US$55.

Porter:1986:RH

[PT86] Roy Porter and Mikuláš Teich, editors. Revolution in history. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25978-9, 0-521-27784-1 (paperback). x + 341 pp. LCCN D32 .R44

                                        1986. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/86006085.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        86006085.html.

Porter:1992:SRN

[PT92] Roy Porter and Mikuláš Teich, editors. The scientific revolution in national context. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-39510-0 (hardcover), 0-521-39699-9 (paperback). x + 305 pp. LCCN Q127.E8 S36

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/91034348.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam021/
                                        91034348.html.

Paracelsus:1996:FTT

[PTS^{+}96]  Paracelsus,  C.  Lilian

(Clarice Lilian) Temkin, Henry E. (Henry Ernest) Sigerist, et al. Four treatises of Theophrastus von Hohenheim, called Paracelsus. The Johns Hopkins University Press, Baltimore, MD, USA, johns hopkins paperbacks edition, 1996. ISBN 0-8018-5523-3 (paperback). xii + 256 pp. LCCN R128.6 .P2213

1996. URL http://www.loc.
gov/catdir/bios/jhu051/
96030930.html;
http://www.loc.gov/
catdir/description/
jhu052/96030930.html.

Pugh:1984:MSI

[Pug84] Emerson W. Pugh. Memories that shaped an industry: decisions leading to IBM System/ 360. The MIT Press series in the history of computing. MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-16094-3. x + 323 pp. LCCN QA76.8.I1015 P83 1984.

Pumfrey:1984:BRS

[Pum84] Stephen Pumfrey. Book review:

Seventeenth Century: Steven J. Dick, Plurality of worlds. The origins of the extraterrestrial life debate from Democritus to Kant. Cambridge: Cambridge University Press, 1982. ISBN 0-521-24308-4. Pp. x + 246. £19.00. British Journal for the History of Science, 17 (1):107–108, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026044.

Pumfrey:1987:BRC

[Pum87] Stephen Pumfrey. Book review: C. A. Ronan. The Shorter Science and Civilisation in China. Volume 3. Cambridge: Cambridge University Press, 1986. Pp. ix + 298. ISBN 0-521-25272-5, £30.00, $49.50 (hardcover); 0-521-31560-3, £12.50, $19.95 (paperback). British Journal for the History of Science, 20(3):354–355, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026365.

Pumfrey:1988:BRG

[Pum88] Stephen Pumfrey. Book review: G. J. R. Parry. A Protestant Vision: William Harrison and the Reformation of Elizabethan England. Cambridge: Cambridge University Press, 1987. Pp. ix + 348. ISBN 0-521-32997-3. £27.50. British Journal for the History of Science, 21(4):498–499, December 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026970.

Pumfrey:1989:TMS

[Pum89] Stephen Pumfrey. ‘O tempera, O magnes!’: A sociological analysis of the discovery of secular magnetic variation in 1634. British Journal for the History of Science, 22 (2):181–214, July 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026660.

Pauli:1985:WBB

[PvM85] Wolfgang Pauli and Karl von Meyenn. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a. 2, 1930–1939. (German) [Scientific letters with Bohr, Einstein, Heisenberg and others, 2, 1930–1939], volume 6 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-13609-6 (Berlin), 0-387-13609-6 (New York). xxxix + 783 pp. LCCN ????

Pinsky:1989:CTC

[PW89a]  Valerie  Pinsky  and

Alison Wylie, editors. Critical traditions in contemporary archaeology: essays in the philosophy, history, and socio-politics of archaeology. New directions in archaeology. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-32109-3. ix + 160 pp. LCCN CC72 .C75 1989.

Principe:1989:JBD

[PW89b] Lawrence M. Principe and Andrew Weeks. Jacob Boehme’s divine substance salitter: its nature, origin, and relationship to Seventeenth Century scientific theories. British Journal for the History of Science, 22 (1):53–61, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026678.

Pyenson:1976:BRN

[Pye76]  Lewis Pyenson.  Book re-

view: Nineteenth and Twentieth Centuries: Science at the Crossroads. By Herbert Dingle. London: Martin Brian and O’Keefe, 1972. Pp. 256. £4.00. British Journal for the History of Science, 9(3):336– 337, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025698.

Pyenson:1985:YEA

[Pye85] Lewis Pyenson. The young Einstein: the advent of relativity. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-779-9. xiv + 255 + 9 pp. LCCN QC16.E5 P93

                                        1985. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0745/85120280-d.html.

Pyenson:1987:BRG

[Pye87]  Lewis Pyenson.  Book re-

view: Gerald James Holton. The Advancement of Science, and its Burdens. Cambridge: Cambridge University Press, 1986. Pp. xi + 351. ISBN 0-521-27243-2 £9.95 (paperback). British Journal for the History of Science, 20(4):475– 477, October 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026419.

Pyenson:1993:CME

[Pye93] Lewis Pyenson. Civilizing mission: exact sciences and French overseas expansion, 1830–1940. The Johns Hopkins University Press, Baltimore, MD, USA, 1993. ISBN 0-8018-4421-5. xxi + 377 pp. LCCN Q127.F8 P94

1993. URL http://www.loc.
gov/catdir/description/
jhu051/92022577.html.

Queisser:1988:CM

[Que88] Hans J. Queisser. The conquest of the microchip. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-16296-X. x + 200 + 12 pp. LCCN TK7874 .Q3613 1988.

Russell:1985:RDH

[R^{+}85] Colin Archibald Russell et al., editors. Recent developments in the history of chemistry. Royal Society of Chemistry, London, UK, 1985. ISBN 0-85186-917-3. x + 333 pp. LCCN QD15. US$27.50.

Reardon-Anderson:1991:SCC

[RA91] James Reardon-Anderson. The study of change: chemistry in China, 1840–1949. Studies of the East Asian Institute, Columbia University. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-39150-4 (hardcover). xix + 444 pp. LCCN QD18.C5 R43

                                        1991. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/90034975.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam029/
                                        90034975.html.

Rabb:1978:PS

[Rab78] Theodore K. Rabb. Puritan science? British Journal for the History of Science, 11(2): 156–159, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025728.

Rafter:1988:WTE

[Raf88] Nicole Hahn Rafter, editor. White trash: the eugenic family studies, 1877–1919. Northeastern University Press, Boston, MA, USA, 1988. ISBN 1-55553-030-3. x + 382 pp. LCCN HQ755.5.U5 W52 1988. US$40.00.

Rainger:1991:AAH

[Rai91] Ronald Rainger. An agenda for antiquity: Henry Fairfield Osborn and vertebrate paleontology at the American Museum of Natural History, 1890– 1935. History of American science and technology series. University of Alabama Press, Tuscaloosa, AL, USA, 1991. ISBN 0-8173-0536-X. xiii + 360 pp. LCCN QE841 .R24 1991.

Ramsey:1988:PPM

[Ram88]  Matthew Ramsey.  Profes-

sional and popular medicine in France, 1770–1830: the social world of medical practice. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-30517-9. xvii + 406 pp. LCCN R504 .R36

1988. URL http://www.loc.
gov/catdir/description/
cam023/87015882.html;
http://www.loc.gov/
catdir/samples/cam031/
87015882.html;   http://
www.loc.gov/catdir/toc/
cam031/87015882.html.

Randell:1974:BRT

[Ran74]  Brian Randell.  Book re-

view: Technology The Computer from Pascal to von Neumann. By Herman H. Goldstine. Princeton: Princeton University Press, and London: Oxford University Press, 1973. Pp. xii + 378. £6.25. British Journal for the History of Science, 7(3):289–290, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025539.

Rappaport:1969:EDB

[Rap69] Rhoda Rappaport. The early disputes between Lavoisier and Monnet, 1777–1781. British Journal for the History of Science, 4(3):233–244, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025473.

Rappaport:1971:BRE

[Rap71]  R. Rappaport.  Book re-

view: Eighteenth Century: Telliamed, or Conversations between an Indian Philosopher and a French Missionary on the Diminution of the Sea. By Benoît de Maillet. Trans, and ed. by Albert V. Carezzi. Urbana and London: University of Illinois Press. 1969. Pp. xiv + 465. £4.75. British Journal for the History of Science, 5(3):312–313, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025342.

Rappaport:1973:LTE

[Rap73] Rhoda Rappaport. Lavoisier’s Theory of the Earth. British Journal for the History of Science, 6(3):247–260, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025444.

Rappaport:1975:BRG

[Rap75] Rhoda Rappaport. Book review: Geology The Meaning of Fossils. Episodes in the History of Palaeontology. By Martin J. S. Rudwick. London: Macdonald, and New York: American Elsevier, 1972. Pp. iv + 287. £6. British Journal for the History of Science, 8(1): 71–73, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025818.

Rappaport:1978:GOC

[Rap78] Rhoda Rappaport. Geology and orthodoxy: The case of Noah’s Flood in Eighteenth-Century thought. British Journal for the History of Science, 11(1):1–18, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025604.

Rappaport:1979:BRG

[Rap79] Rhoda Rappaport. Book review: Geology The Making of Geology: Earth Science in Britain, 1660–1815. By Roy Porter. Cambridge: Cambridge University Press, 1977. Pp. ii + 288. £8.95. British Journal for the History of Science, 12 (2):221–223, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025759.

Rappaport:1982:BWP

[Rap82] Rhoda Rappaport. Borrowed words: problems of vocabulary in eighteenth-century geology. British Journal for the History of Science, 15(1):27– 44, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025922.

Rappaport:1986:BRP

[Rap86a] Rhoda Rappaport. Book review: Paolo Rossi. The Dark Abyss of Time. The History of the Earth and the History of Nations from Hooke to Vico. A translation by Lydia G. Cochrane of I Segni del Tempo (1979). Chicago and London: University of Chicago Press, 1985. Pp. xvi + 338. ISBN 0-226-72835-8. £36.75. British Journal for the History of Science, 19(3): 362–363, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026534.

Rappaport:1986:HEL

[Rap86b] Rhoda Rappaport. Hooke on earthquakes: Lectures, strategy and audience. British Journal for the History of Science, 19(2):129–146, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026589.

Rattansi:1965:BRA

[Rat65a] P. M. Rattansi. Book review: Andreas Vesalius of Brussels, 1514–1564. By C. D. O’Malley. Pp. xvi + 480, with 64 plates. Berkeley and Los Angeles: University of California Press; Cambridge University Press, 1964. £4 net. British Journal for the History of Science, 2(3):267–268, June 1965.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024949.

Rattansi:1965:ER

[Rat65b] P. M. Rattansi. Essay reviews. British Journal for the History of Science, 2(4):355– 356, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024892.

Rattansi:1966:BRM

[Rat66a] P. M. Rattansi. Book review:

De Mundo nostro Sublunari Philosophia Nova. By William Gilbert. (Facsimile of the 1651 Amsterdam edition.) Pp. 320. The De Mundo of William Gilbert. By Sister Suzanne Kelly, O.S.B. Pp. 142. Amsterdam: Menno Herzberger & Co., 1965. $40 the set. British Journal for the History of Science, 3(2):191, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025033.

Rattansi:1966:BRS

[Rat66b] P. M. Rattansi. Book review: Sixteenth and Seventeenth Centuries: A Translation of John Dee’s “Monas Hieroglyphica” (Antwerp, 1564), with an Introduction and Annotations. By C. H. Josten. Ambix, Vol. xii, Nos. 2 and 3. June and October 1964. Pp.

83-222. 35s. British Journal for the History of Science, 3 (1):82, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025111.

Rattansi:1968:BRS

[Rat68] P. M. Rattansi. Book review: Sixteenth and Seventeenth Centuries: The Cambridge Economic History of Europe: Vol. IV, The Economy of Expanding Europe in the 16th and 17th Centuries. Edited by E. E. Rich and C. H. Wilson. London: Cambridge University Press. 1967. Pp. xxxii + 642. 75s. British Journal for the History of Science, 4(2): 189–190, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025162.

Rattansi:1969:BRS

[Rat69] P. M. Rattansi. Book review: Seventeenth Century: The Lives of the Professors of Gresham College. By John Ward. (Facsimile of the 1740 London edition.) New York and London: Johnson Reprint Corporation. 1967. Pp. xxiv + 338 + 156. Plates. $20.00. British Journal for the History of Science, 4(4):419–420, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025252.

Rauschenberg:1975:BRS

[Rau75] Roy A. Rauschenberg. Book review: Seventeenth and Eighteenth Centuries: Linnaeus’s Öland and Gotland Journey, 1741. By Carl Linnaeus. Trans. by Marie Asberg and William T. Stearn. London: Linnean Society of London, 1974. Pp. viii + 220. £6. British Journal for the History of Science, 8(2):185–186, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025654.

Ravetz:1965:BRG

[Rav65] J. R. Ravetz. Book review:

Geschichte der Mathematik. I. Von den Anfängen bis zum Auftreten von Fermat und Descartes. By J. E. Hofmann. Berlin: de Gruyter (Sammlung Göschen, Band 226/226a). 1963. Pp. 251. DM. 5.80. British Journal for the History of Science, 2 (4):360, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024895.

Ravetz:1966:BRM

[Rav66a] J. R. Ravetz. Book review:

Mathematics: and Logic in History and in Contemporary Thought. By Ettore Carruccio, translated by Isabel Quigly. London: Faber and Faber, 1964. 63s. British Journal for the History of Science, 3(2):

195, December 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025038.

Ravetz:1966:BRS

[Rav66b] J. R. Ravetz. Book review: Science in History. Third Edition. By J. D. Bernal. Pp. xviii + 1039. London: C. A. Watts, 1965. £4 4s. British Journal for the History of Science, 3 (2):188, December 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025030.

Ravetz:1967:BRH

[Rav67] J. R. Ravetz. Book review: History and Philosophy of Science A Guide to Case Studies of Scientific Activity. By Ben-Ami Lipetz (assisted by C. P. McLaughlin). Carlisle, Mass.: Intermedia Inc., 1965; vii + 350 pp. $12. British Journal for the History of Science, 3 (4):397, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ravetz:1968:BRS

[Rav68] J. R. Ravetz. Book review: Seventeenth Century: Galileo Galilei, his life and his works. By Raymond J. Seeger. Pp. xi + 286. Oxford: Pergamon Press. 1966. 21s. British Journal for the History of Science, 4(1):71–72, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024991.

Ravetz:1969:BRA

[Rav69] J. R. Ravetz. Book review:

Astronomy and Physics Thaddeus Hagecius ab Hayck, Dialexis de novae et prius incognitae stellae apparitione, 1574. Edited by Z. Horsky. Prague: Pragopress. Pp. 176 + 35. 1967. Editio Cimelia Bohemica, Vol. I. (Johnson Reprint Corporation. $10.00.) Johannus Marcus Marci, Thaumantias, Liber de arcu coelesti, deque colorum apparantium natura, ortu et causis …, 1648. Edited by J. Marek. Prague: Pragopress. Pp. 270 + 32. 1968. Editio Cimelia Bohemica, Vol. III. (Johnson Reprint Corporation. $12.00.). British Journal for the History of Science, 4(4): 405–406, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025237.

Ravetz:1970:BRA

[Rav70] J. R. Ravetz. Book review:

Astronomy and Navigation A History of Japanese Astronomy: Chinese Background and Western Impact. By Shigeru Nakayama. Pp. xiii + 329. Harvard University and Oxford University Press. 1969. 95s. British Journal for the History of Science, 5(1):92– 93, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025361.

Ravetz:1971:BRM

[Rav71] J. R. Ravetz. Book review: Mathematics: Lebesgue’s Theory of Integration. Its Origins and Development. By Thomas Hawkins. Madison: University of Wisconsin Press, 1970. Pp. xv + 227. £5.95. British Journal for the History of Science, 5(4):401–402, December 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025385.

Ravetz:1972:BRS

[Rav72] J. R. Ravetz. Book review: Sixteenth and Seventeenth Centuries: Galileo Studies: Personality, Tradition and Revolution. By Stillman Drake. Ann Arbor: University of Micshigan Press, 1970. Pp. 289. $8.50. Galileo, Science and the Church. By Jerome J. Langford. Ann Arbor: University of Michigan Press, 1971. Pp. 207. $2.45 (paperback). British Journal for the History of Science, 6 (1):93–94, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025274.

Ravetz:1973:BRS

[Rav73] J. R. Ravetz. Book review:

Science and Society The Scientist’s Role in Society. A Comparative Study. By Joseph Ben-David. Englewood Cliffs, N.J., and London: Prentice-Hall, 1971. Pp. xii + 207. £1.20 (paperback). British Journal for the History of Science, 6(4): 435–436, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025508.

Ravetz:1984:BRS

[Rav84] J. R. Ravetz. Book review:

Science in Culture: Marcel Chotkowski la Follette (ed.), Quality in science. Cambridge, Mass., and London: The MIT Press, 1982. Pp. xi + 169. ISBN 0-262-62040-5 (paperback). £10.00. British Journal for the History of Science, 17(2):251, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026585.

Ray:1987:ER

[Ray87] Christopher Ray. The evolution of Relativity. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-423-4. x + 211 pp. LCCN QC173.52 .R39

1987. URL http://www.loc.
gov/catdir/enhancements/
fy0745/87021137-d.html.

Rainger:1988:ADB

[RBM88] Ronald Rainger, Keith Rodney Benson, and Jane Maienschein, editors. The American development of biology. University of Pennsylvania Press, Philadelphia, PA, USA, 1988. ISBN 0-8122-8092-X. LCCN QH305.2.U6 A54

1988. URL http://catalog.
hathitrust.org/api/
volumes/oclc/17509080.
html.

Rayner-Canham:1998:WCT

[RCRC98] Marelene F. Rayner-Canham and Geoffrey Rayner-Canham. Women in chemistry: their changing roles from alchemical times to the mid-Twentieth Century. History of modern chemical sciences, 1069– 2452. American Chemical Society, Washington, DC, 1998. ISBN 0-8412-3522-8. xiv + 284 pp. LCCN QD20 .R39 1998.

Rimmer:1995:APH

[RD95] Harry Rimmer and Edward Bradford Davis, editors. The antievolution pamphlets of Harry Rimmer, volume 6 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1841-3, 0-8153-1807-3. LCCN BS659 .R56 1995.

Rechenberg:1986:BRE

[Rec86] Helmut Rechenberg. Book review: Elizabeth Crawford. The Beginning of the Nobel Institution. The Science Prizes, 1901–1915. Cambridge: Cambridge University Press, 1985.

Pp. ix + 281. ISBN 0-521-26584-3. £22.50. British Journal for the History of Science, 19(2):229, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026612.

Redondi:1983:GE

[Red83] Pietro Redondi. Galileo eretico, volume 7 of Microstorie. G. Einaudi, Torino, Italy, 1983. ISBN 88-06-05632-8. x + 463 + 16 pp. LCCN Q125.2 .R43 1983. L25000.

Redhead:1984:BRP

[Red84] Michael L. G. Redhead. Book review: Physics and its concepts: Abraham Pais, ‘Subtle is the Lord’: the science and the life of Albert Einstein. Oxford: Clarendon Press, 1982. Pp. xvi + 552. ISBN 0-19-853907-X. £15. British Journal for the History of Science, 17(2):226–227, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026560.

Redner:1987:ESE

[Red87] Harry Redner. The ends of science: an essay in scientific authority. Westview Press, Boulder, CO, USA, 1987. ISBN 0-8133-0452-0. xiv + 344 pp. LCCN Q158.5 .R43 1987.

Rees:1980:BRB

[Ree80] T. A. V. Rees. Book review: Biology The Spontaneous Generation Controversy from Descartes to Oparin. By John Farley. Baltimore and London: Johns Hopkins University Press, 1977. Pp. xiv + 225. £10.25. British Journal for the History of Science, 13 (2):161–162, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025896.

Reeves:1985:BRD

[Ree85] Nigel Reeves. Book review: Dieter Oldenburg, Romantische Naturphilosophie und Arzneimittellehre 1800– 1840, Veröffentlichungen aus dem Pharmaziegeschichtlichen Seminar der Technischen Universität Braunschweig vol. 20. Brunswick, 1979. Pp. 267. No priced stated. ISSN 0068-0729. British Journal for the History of Science, 18(1):88– 89, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026272.

Rehbock:1983:PNT

[Reh83] Philip F. Rehbock. The philosophical naturalists: themes in early Nineteenth-Century British biology, volume 3 of Wisconsin publications in the history of science and medicine. University of Wisconsin Press, Madison, WI, USA, 1983. ISBN 0-299-09430-8. xv + 281 pp.

LCCN QH305.2.G7 R43 1983. US$30.00.

Reingold:1968:BMS

[Rei68] Nathan Reingold. Babbage and Moll on the state of science in Great Britain: A note on a document. British Journal for the History of Science, 4 (1):58–64, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024987.

Reich:1985:MAI

[Rei85] Leonard S. Reich. The making of American industrial research: science and business at GE and Bell, 1876–1926. Studies in economic history and policy. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-30529-2. xvi + 309 pp. LCCN HD30.42.U5 R44

1985. URL http://www.loc.
gov/catdir/description/
cam023/85006645.html;
http://www.loc.gov/
catdir/samples/cam031/
85006645.html;   http://
www.loc.gov/catdir/toc/
cam031/85006645.html.

Reingold:1986:BRH

[Rei86] Nathan Reingold. Book review: History of Science Today, 1. Uniformity as Hidden Diversity: History of Science in the United States, 1920– 1940. British Journal for the History of Science, 19(3):243– 262, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026504.

Reingold:1987:DPG

[Rei87] Nathan Reingold. On not doing the papers of great scientists. British Journal for the History of Science, 20(1):29– 38, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026434. Conference

on documentary editing (Tarrytown, N.Y., 1985).

ReyGonzalez:1990:EMS

[Rey90] Antonio M. Rey González. Estudios médico-sociales sobre marginados en la Espaäna del siglo XIX, volume 17 of Textos clásicos espaänoles de la salud pública. Ministerio de Sanidad y Consumo, Madrid, Spain, 1990. ISBN 84-7670-204-3, 84-7670-205-1. 237 pp. LCCN RA418.3.S7 R49

1990. URL http://catalog.
hathitrust.org/api/
volumes/oclc/21796295.
html.

R:1965:AMA

[RG65] K. R. and D. L. Gardiner. André-Marie Ampère and his English acquaintances. British Journal for the History of Science, 2(3):235–245, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024937.

Roll-Hansen:1986:BRA

[RH86] Nils Roll-Hansen. Book review:

Alexander Vucinich. Empire of Knowledge. The Academy of Sciences of the USSR (1917– 1970). Berkeley: University of California Press, 1984. Pp. x + 484. ISBN 0-520-04871-7. £23.95. British Journal for the History of Science, 19(2): 222–223, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026607.

Roll-Hansen:1988:BRL

[RH88] Nils Roll-Hansen. Book review: Loren R. Graham. Science, Philosophy, and Human Behaviour in the Soviet Union. New York: Columbia University Press, 1987. Pp. XIV + 565. ISBN 0-231-06442-X. $45.00. British Journal for the History of Science, 21(3):381– 382, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026852.

Roll-Hansen:1989:GEM

[RH89] Nils Roll-Hansen. Geneticists and the eugenics movement in Scandinavia. British Journal for the History of Science, 22 (3):335–346, September 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026900.

Richards:1982:EEB

[Ric82] Robert J. Richards. The emergence of evolutionary biology of behaviour in the early Nineteenth Century. British Journal for the History of Science, 15(3):241–280, November 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026220.

Richards:1987:QPR

[Ric87a] Evelleen Richards. A question of properly rights: Richard Owen’s evolutionism reassessed. British Journal for the History of Science, 20 (2):129–171, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026305.

Richards:1987:WHP

[Ric87b] Graham Richards. Of what is history of psychology a history? British Journal for the History of Science, 20(2): 201–211, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026307.

Richardson:1987:DDD

[Ric87c]  Ruth Richardson.  Death,

dissection, and the destitute. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1987. ISBN 0-7102-0919-3. xvii + 426 pp. LCCN GT3243 .R53 1987.

Richards:1988:BRH

[Ric88a] Graham Richards. Book review: L. S. Hearnshaw. The Shaping of Modern Psychology: An Historical Introduction. London: Routledge & Kegan Paul, 1987. Pp. vii + 423. ISBN 0-7102-0576-7. £19.95. British Journal for the History of Science, 21(1): 135–136, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026891.

Richards:1988:MVP

[Ric88b] Joan L. Richards. Mathematical visions: the pursuit of geometry in Victorian England. Academic Press, New York, USA, 1988. ISBN 0-12-587445-6. xiii + 266 pp. LCCN QA443.5 .R53 1988.

Richards:1989:BRW

[Ric89] Graham Richards. Book review: William Hurst (ed.). The Making of Cognitive Science: Essays in honor of George A. Miller. Cambridge: Cambridge University Press, 1988. Pp. x + 284. ISBN 0-521-34255-4, £25, $29.95. British Journal for the History of Science, 22(3):395– 396, September 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026912.

Richards:1991:VCC

[Ric91] Evelleen Richards. Vitamin C and cancer: medicine or politics? MacMillan Publishing Company, New York, NY, USA, 1991. ISBN 0-333-44419-1. xii + 269 pp. LCCN RC271.A78 R53 1991b.

Richards:1992:MMO

[Ric92a] Graham Richards, editor. Mental machinery: the origins and consequences of psychological ideas: part 1: 1600–1850. Athlone Press, London, UK, 1992. ISBN 0-485-11394-5. xi + 490 pp. LCCN BF95 .R53 1992. US$50.00.

Richards:1992:MEM

[Ric92b]  Robert  J.

(Robert John) Richards. The meaning of evolution: the morphological construction and ideological reconstruction of Darwin’s theory. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-71202-8. xv + 205 pp. LCCN QH366.2 .R52

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi053/91019017.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0666/91019017-b.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0666/91019017-t.html.

Ridley:1982:CIN

[Rid82] Mark Ridley. Coadaptation and the inadequacy of natural selection. British Journal for the History of Science, 15 (1):45–68, March 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025923.

Ridley:1983:BRL

[Rid83] Mark Ridley. Book review:

Life Sciences: John C. Greene, Science, ideology, and world view: essays in the history of evolutionary ideas. Berkeley: University of California Press, 1981. Pp x + 202. £11.50/ $20.60; £4.00/$7.45 (paperback). British Journal for the History of Science, 16(2): 214–215, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026255.

Riley:1985:BRJ

[Ril85] Denise Riley. Book review:

John Forrester, Language and the origins of psychoanalysis. London: MacMillan, 1980. Pp xv + 285. £15. British Journal for the History of Science, 18(1):85, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026269.

Ritvo:1990:DIF

[Rit90] Lucille B. Ritvo. Darwin’s influence on Freud: a tale of two sciences. Yale University Press, New Haven, CT, USA, 1990. ISBN 0-300-04131-4. xii + 267 pp. LCCN BF173 .R515 1990.

Ramsey:1997:CBC

[RL97] J. T. (John T.) Ramsey and A. Lewis Licht. The comet of 44 B.C. and Caesar’s funeral games, volume 39 of American classical studies. Scholars Press, Atlanta, Ga., 1997. ISBN 0-7885-0273-5 (hardcover), 0-7885-0274-3 (paperback). xx + 236 pp. LCCN GT4852.R6 R36

                                        1997. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0603/96017535-d.html.

Raaashid:1996:EHA

[RM96] Rushdåi Råashid and Régis Morelon, editors. Encyclopedia of the history of Arabic science. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1996. ISBN 0-415-02063-8 (set), 0-415-12410-7 (vol. 1), 0-415-12411-5 (vol. 2), 0-415-12412-3 (vol. 3). xiv + 1105 pp. LCCN Q127.A5 E53

                                        2000. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0648/00362029-d.html.

Roberts:1975:BRN

[Rob75] Gerrylynn K. Roberts. Book review: Nineteenth and Twentieth Centuries: Science for the People. The Origins of the School Science Curriculum in England. By David Layton. London: George Allen and Unwin, 1973. Pp. 226.

£3.55. British Journal for the History of Science, 8(1): 86–87, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025831.

Roberts:1980:BRN

[Rob80a] Gerrylynn K. Roberts. Book review: Nineteenth Century: The Patronage of Science in the Nineteenth Century. Ed. by G. L’E. Turner. Science in History, i. Leyden: Noordhoff International Publishing, 1976. Pp. vi + 218. Df155.00/ $21.25. British Journal for the History of Science, 13(2): 172–173, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025906.

Roberts:1980:SCS

[Rob80b] Gerrylynn K. Roberts. The social and cultural significance of science: The Royal Institution. British Journal for the History of Science, 13(2): 154–157, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025893.

Robinson:1986:AKE

[Rob86] Joseph D. Robinson. Appreciating key experiments. British Journal for the History of Science, 19(1):51–56, March 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026484.

Roberts:1987:BRM

[Rob87] Gerrylynn K. Roberts. Book review: Marie Boas Hall. All Scientists Now. The Royal Society in the Nineteenth Century. Cambridge: Cambridge University Press, 1984. Pp. xii + 261. ISBN 0-521-26746-3. £25.00. British Journal for the History of Science, 20(1): 81–82, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Roberts:1989:BRD

[Rob89a] G. K. Roberts. Book review: David Knight. The Age of Science: The Scientific World View in the Nineteenth Century. Oxford: Basil Blackwell, 1988. Pp. 251. ISBN 0-631-16176-7, £7.95. (Paper: original hardback edition 1986.). British Journal for the History of Science, 22(4):449– 450, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026928.

Roberts:1989:JPR

[Rob89b] Shirley Roberts. Sir James Paget: the rise of clinical surgery. Eponymists in medicine. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1989. ISBN 0-905958-99-3 (hardcover), 0-905958-91-8. ISSN 0955-8640. 192 pp. LCCN RA489.R62 R62 1989. US$7.50. URL http://

catalog.hathitrust.org/
api/volumes/oclc/
22547173.html.

Robertson:1992:BSS

[Rob92]  Peter Robertson.  Beyond

southern skies: radio astronomy and the Parkes telescope. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41408-3. xi + 357 pp. LCCN QB479.2 .R63

1992. URL http://www.loc.
gov/catdir/description/
cam025/92003077.html;
http://www.loc.gov/
catdir/toc/cam025/
92003077.html.

Roche:1981:HRS

[Roc81a] John J. Roche. Harriot’s “Regiment of the Sun” and its background in sixteenth-century navigation. British Journal for the History of Science, 14(3): 245–262, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025982.

Rocke:1981:KBH

[Roc81b] A. J. Rocke. Kekulé, Butlerov, and the historiography of the theory of chemical structure. British Journal for the History of Science, 14(1):27– 57, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026071.

Rocke:1984:CAN

[Roc84] Alan J. Rocke. Chemical atomism in the Nineteenth Century: from Dalton to Cannizzaro. Ohio State University Press, Columbus, OH, USA, 1984. ISBN 0-8142-0360-4. xviii + 386 pp. LCCN QD15 .R63 1984. US$27.50.

Roche:1987:BRJ

[Roc87] John Roche. Book review: J. Z. Buchwald. From Maxwell to Microphysics. Chicago: University of Chicago Press, 1985. Pp. xv + 339. ISBN 0-226-07882-5. £59.50. British Journal for the History of Science, 20(2):229–230, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026315.

Roche:1988:RLG

[Roc88a] Daniel Roche. Les républicains des lettres: gens de culture et lumières au XVIIIe siècle. (French) [Republicans of letters: people of culture and light in the 18th Century]. Fayard, Paris, France, 1988. ISBN 2-213-02122-8. 393 pp. LCCN Z1003.5.F7 R63 1988. 160F.

Rocke:1988:BRD

[Roc88b]  A. J. Rocke.  Book re-

view: Dean Stanley Tarbell & Ann Tracy Tarbell. Essays on the History of Organic Chemistry in the United States, 1875–1955. Nashville, Tennessee: Folio Publishers, 1986. Pp. x + 434. ISBN 0-939454-03-3. $21.95; orders are to be directed to the authors, Box 1520, Station B, Vanderbilt University, Nashville, TN 37235, U.S.A. British Journal for the History of Science, 21 (1):130–131, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026886.

Roche:1990:PLB

[Roc90] John J. (John James) Roche, editor. Physicists look back: studies in the history of physics. Adam Hilger Ltd., Bristol, UK, 1990. ISBN 0-85274-001-8. xii + 393 pp. LCCN QC7 .P475

1990. URL http://www.loc.
gov/catdir/enhancements/
fy0705/91148368-d.html.

Rogers:1983:BRB

[Rog83] G. A. J. Rogers. Book review: Biography Richard S. Westfall, Never at rest: a biography of Isaac Newton, Cambridge: Cambridge University Press 1981. Pp. xviii + 908. £25.00. British Journal for the History of Science, 16(1): 101–105, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026109.

Rogers:1988:BRDa

[Rog88a] G. A. J. Rogers. Book review: David Oldroyd. The Arch of Knowledge. An Introductory Study of the History of the Philosophy and Methodology of Science. New York and London: Methuen, 1986. Pp. 413. ISBN 0-416-01341-4. £9.95. British Journal for the History of Science, 21(3): 384, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026854.

Rogers:1988:BRDb

[Rog88b] G. A. J. Rogers. Book review: Derek Gjertsen. The Newton Handbook. London and New York: Routledge and Kegan Paul, 1986. Pp. xiv + 665. ISBN 0-7102-0279-2. £25.00. British Journal for the History of Science, 21(3):385–386, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026856.

Ronan:1983:CIH

[Ron83] Colin A. Ronan. The Cambridge illustrated history of the world’s science. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-600-38423-3 (Newnes) 0-521-25844-8 (Cambridge). 543 pp. LCCN Q125 R645.

Ronan:1986:BRW

[Ron86] Colin Ronan. Book review: Wen-Yuan Qian. The Great Inertia. London: Croom Helm, 1984. Pp. xii + 155. ISBN 0-7099-2104-7. £15.95. British Journal for the History of Science, 19(3):370–371, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026542.

Rose:1968:GTB

[Ros68] Paul-Lawrence Rose. Galileo’s Theory of Ballistics. British Journal for the History of Science, 4(2):156–159, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025144.

Rosenberg:1979:BRM

[Ros79] Charles Rosenberg. Book review: Medicine The Double Face of Janus and other Essays in the History of Medicine. By Owsei Temkin. Baltimore & London: Johns Hopkins University Press, 1977. Pp. x + 543. £16.90. British Journal for the History of Science, 12 (2):223–225, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025760.

Roseveare:1982:MPV

[Ros82] N. T. Roseveare. Mercury’s perihelion, from Le Verrier to Einstein. Oxford science publications. Clarendon Press, New York, NY, USA, 1982. ISBN 0-19-858174-2. viii + 208 pp. LCCN QB611 .R67 1982. US$20.00.

Rosen:1984:CSR

[Ros84a] Edward Rosen. Copernicus and the scientific revolution. Anvil series. Krieger, Malabar, FL, USA, original edition, 1984. ISBN 0-89874-573-X (paperback). 220 pp. LCCN QB36.C8 R65 1984. US$5.95.

Rossi:1984:DAT

[Ros84b] Paolo Rossi. The dark abyss of time: the history of the earth and the history of nations from Hooke to Vico. University of Chicago Press, Chicago, IL, USA, 1984. ISBN 0-226-72835-8. xvi + 338 pp. LCCN D16.8 .R6813 1984. US$37.50.

Rosen:1986:TIM

[Ros86] Edward Rosen. Three imperial mathematicians: Kepler trapped between Tycho Brahe and Ursus. The New horizon series. Abaris Books, New York, NY, USA, 1986. ISBN 0-89835-242-8. 384 pp. LCCN QB35 .R67 1986.

Rossi:1990:MLS

[Ros90] Bruno Benedetto Rossi. Moments in the life of a scientist. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-36439-6. xiii + 181 + 32 pp. LCCN QC16.R495 R6713 1990. URL http://

                                        www.loc.gov/catdir/
                                        description/cam023/

89022283.html;   http://
www.loc.gov/catdir/toc/
cam028/89022283.html.

Ross:1991:NCA

[Ros91] Sydney Ross. Nineteenth-Century attitudes: men of science, volume 13 of Chemists and chemistry. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1308-9 (hardcover). xi + 235 pp. LCCN Q174.8 .R67 1991.

Ross:1994:MIH

[Ros94] Dorothy Ross, editor. Modernist impulses in the human sciences, 1870– 1930. The Johns Hopkins University Press, Baltimore, MD, USA, 1994. ISBN 0-8018-4744-3, 0-8018-4745-1 (paperback). xii + 379 pp. LCCN H53.U5

M63 1994. URL http://www.
loc.gov/catdir/bios/
jhu051/93038354.html;
http://www.loc.gov/
catdir/description/
jhu051/93038354.html.

Rothenberg:1982:HST

[Rot93] Marc Rothenberg. The history of science and technology in the United States: a critical and selective bibliography, volume 2, 17; vol. 308, 815 of Bibliographies of the history of science and technology; Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1982– 1993. ISBN 0-8240-9278-3 (vol.

1), 0-8240-8349-0 (vol. 2). 2 v. pp. LCCN Z7405.H6 R67 1982. US$24.00.

Rowlands:1990:NCM

[Row90a] Peter Rowlands. Newton and the concept of mass-energy, volume 4 of Liverpool historical essays. Published for the Dept. of History, University of Liverpool [by] Liverpool University Press, Liverpool, UK, 1990. ISBN 0-85323-187-7. 34 pp. LCCN QC73 .R69 1990.

Rowlands:1990:OLL

[Row90b] Peter Rowlands. Oliver Lodge and the Liverpool Physical Society, volume 4 of Liverpool historical studies. Published for the Department of History, University of Liverpool [by] Liverpool University Press, Liverpool, UK, 1990. ISBN 0-85323-027-7 (paperback). iv + 310 pp. LCCN QC16.L6 R69 1990. US$12.50.

Reingold:1981:SAD

[RR81] Nathan Reingold and Ida H. Reingold, editors. Science in America, a documentary history, 1900–1939. The Chicago history of science and medicine. University of Chicago Press, Chicago, IL, USA, 1981. ISBN 0-226-70946-9. xii + 490 pp. LCCN Q125 .S43433. US$39.50

                                        (est.). URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0608/81002584-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0608/81002584-t.html;

http://www.loc.gov/
catdir/enhancements/
fy0609/81002584-b.html.

Reingold:1987:SCC

[RR87] Nathan Reingold and Marc Rothenberg, editors. Scientific colonialism: a cross-cultural comparison. Smithsonian Institution Press, Washington, DC, USA, 1987. ISBN 0-87474-785-6. LCCN Q124.6 .S39 1987. Papers from a conference at Melbourne, Australia 25–30 May 1981.

Roberts:1992:FBE

[RT92] K. B. Roberts and J. D. W.

Tomlinson. The fabric of the body: European traditions of anatomical illustrations. Clarendon Press, New York, NY, USA, 1992. ISBN 0-19-261198-4. xx + 638 pp. LCCN R836 .R63 1990. US$180.00. URL http://www.

loc.gov/catdir/
enhancements/fy0640/
90007650-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0640/90007650-t.html.

Riley:1995:CTC

[RT95]  William  Bell  Riley

and William Vance Trollinger, editors. Creationism in Twentieth-Century America. 4: the antievolution pamphlets of William Bell Riley. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1801-4, 0-8153-1805-7. xxii + 221 pp. LCCN BS651 .R546 1995.

Rudwick:1962:HWC

[Rud62] M. J. S. Rudwick. Hutton and Werner compared: George Greenough’s geological tour of Scotland in 1805. British Journal for the History of Science, 1(2):117–135, December 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025128.

Rudwick:1963:FGS

[Rud63] M. J. S. Rudwick. The foundation of the Geological Society of London: Its scheme for co-operative research and its struggle for independence. British Journal for the History of Science, 1(4):325–355, December 1963. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024906.

Rudwick:1969:BRH

[Rud69] M. J. S. Rudwick. Book review: History of Geology Source Book in Geology, 1900– 1950. Ed. by Kirtley F. Mather. Harvard University Press and Oxford University Press. 1967. £5. British Journal for the History of Science, 4(3):294– 295, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025487.

Rudwick:1970:BRG

[Rud70] Martin Rudwick. Book review:

Geology The Earth in Decay. A History of British Geomorphology 1578–1878. By Gordon L. Davies. London: Macdonald. 1969. Pp. xvi + 390, 8 plates. £5. British Journal for the History of Science, 5(2): 195–196, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025427.

Rudwick:1971:BRG

[Rud71] Martin Rudwick. Book review: Geology and Biology Sir Charles Lyell’s Scientific Journals on the Species Question. Ed. by Leonard G. Wilson. New Haven and London: Yale University Press. 1970. Pp. lxi + 572. £7.85. British Journal for the History of Science, 5(4): 408–409, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025393.

Rudwigk:1974:PSV

[Rud74] Martin J. S. Rudwigk. Poulett Scrope on the volcanoes of Auvergne: Lyellian time and political economy. British Journal for the History of Science, 7(3):205–242, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025531.

Rudwigk:1976:CLS

[Rud76] Martin J. S. Rudwigk. Charles Lyell speaks in the lecture theatre. British Journal for the History of Science, 9(2): 147–155, July 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025802.

Rudwick:1980:BRE

[Rud80] Martin Rudwick. Book review: Eighteenth Century: James Hutton’s Theory of the Earth: the Lost Drawings. By G. Y. Craig (editor), D. B. McIntyre, and C. D. Waterson. Edinburgh: Scottich Academic Press, 1978. 67 pp + portfolio of 29 facsimiledrawings. £75.00/$175.00. British Journal for the History of Science, 13(1):82–83, March 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026137.

Rudwick:1983:BRL

[Rud83] Martin Rudwick. Book review: Life and earth sciences: William A. S. Sarjeant, Geologists and the History of Geology: An International Bibliography from the Origins to 1978. London and Basingstoke: Macmillan, 1980. Pp. iv + 4526 in 5 vols. £250. British Journal for the History of Science, 16(3):275–277, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026148.

Rudwick:1984:BRH

[Rud84] Martin Rudwick. Book reviews: History of Natural History: Mott T. Greene, Geology in the Nineteenth Century. Changing Views of a Changing World. Ithaca and London: Cornell University Press, 1983. Pp. 324. ISBN 0-8014-1467-9. £23.50, $38.35. Nicolaas A. Rupke, The Great Chain of History. William Buckland and the English School of Geology (1814–1849). Oxford: Clarendon Press, 1983. Pp. xii + 322. ISBN 0-19-822907-0. £22.50. British Journal for the History of Science, 17(3):314– 316, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026629.

Rudwick:1985:GDC

[Rud85] M. J. S. Rudwick. The great Devonian controversy: the shaping of scientific knowledge among gentlemanly specialists. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1985. ISBN 0-226-73101-4. xxxiii + 494 pp. LCCN QE665 .R83

1985. URL http://www.loc.
gov/catdir/description/
uchi052/84016199.html;
http://www.loc.gov/
catdir/enhancements/

                                        fy0608/84016199-t.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0609/84016199-b.html.

Rudwick:1986:BRF

[Rud86] Martin Rudwick. Book reviews: Frederick Burkhardt & Sydney Smith (eds). A Calendar of the Correspondence of Charles Darwin, 1821–1882. New York and London: Garland Publishing Inc., 1985. Pp. 690. ISBN 0-8240-9224-4. $100, £85. The Correspondence of Charles Darwin. Volume 1, 1821–1836. Cambridge [etc.]: Cambridge University Press, 1985. Pp. xxxii + 702. ISBN 0-521-25587-2. £30, $37.50. British Journal for the History of Science, 19(3):354–356, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026525.

Ruderman:1988:KMS

[Rud88a] David B. Ruderman. Kabbalah, magic, and science: the cultural universe of a sixteenth-century Jewish physician. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-49660-4. viii + 232 pp. LCCN BM755.J273 R83 1988.

Rudwick:1988:BRF

[Rud88b] Martin Rudwick. Book review: Frederick Burkhardt & Sydney Smith (eds). The Correspondence of Charles Darwin, Volume 2 (1837–1843). Cambridge: Cambridge University Press, 1986. ISBN 0-521-25588-0. £30.00/£37.50. British Journal for the History of Science, 21(1):129–130, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026885.

Rudwick:1989:BRW

[Rud89] Martin Rudwick. Book review: William A. S. Sarjeant. Geologists and the History of Geology. An International Bibliography from the Origins to 1978. Supplement 1979–1984 and Additions. Malabar, Florida: Robert E. Krieger Publishing Company, 1987. 2 vols. Pp. xxxii + 1691. ISBN 0-89874-939-5. No price given. British Journal for the History of Science, 22(1): 114–115, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026710.

Rudwick:1992:SDT

[Rud92] M. J. S. Rudwick. Scenes from deep time: early pictorial representations of the prehistoric world. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-73104-9 (hardcover). xiii + 280 pp. LCCN QE705.A1 R82

1992. URL http://www.loc.
gov/catdir/description/
uchi052/91047677.html;
http://www.loc.gov/

                                        catdir/enhancements/
                                        fy0608/91047677-t.html.

Rupke:1983:GCH

[Rup83] Nicolaas A. Rupke, editor. The great chain of history: William Buckland and the English school of geology (1814– 1849). Clarendon Press, New York, NY, USA, 1983. ISBN 0-19-822907-0. LCCN QE48.G7 R82 1983. URL http://

                                        catalog.hathitrust.org/
                                        api/volumes/oclc/
                                        10126243.html.

Rupke:1985:BRA

[Rup85] N. A. Rupke. Book reviews: Anthony Hallam, Great geological controversies. Oxford: Oxford University Press, 1983. Pp. vii + 182. ISBN 0-19-854431-6 (hardcover). £15. ISBN 0-19-854430-8 (paperback). £7.95. British Journal for the History of Science, 18(3):352– 353, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026393.

Rupke:1988:BRE

[Rup88] N. A. Rupke. Book review: Eric Buffetaut. A Short History of Vertebrate Palaeontology. London: Croom Helm, 1987. Pp. 223. ISBN 0-7099-3962-0. £30.00 (hardcover). British Journal for the History of Science, 21 (1):131–132, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026887.

Rupke:1989:BRW

[Rup89] N. A. Rupke. Book review:

S. Warren Carey. Theories of the Earth and Universe. A History of Dogma in the Earth Sciences. Stanford: Stanford University Press, 1988. Pp. xviii + 413. ISBN 0-8047-1364-2. $45.00. British Journal for the History of Science, 22 (2):253, July 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026671.

Russell:1964:KLP

[Rus64] J. L. Russell. Kepler’s Laws of Planetary Motion: 1609– 1666. British Journal for the History of Science, 2(1): 1–24, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025081.

Russell:1966:BRH

[Rus66] C. A. Russell. Book review:

History of Chemistry A History of the Concept of Valency to 1930. By W. G. Palmer. Pp. viii + 178. London: Cambridge University Press, 1965. 42s. British Journal for the History of Science, 3(1):89, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025118.

Russell:1967:BRH

[Rus67]  C. A. Russell.  Book re-

view: History of Chemistry Jac. Berzelius, His Life and Work. By J. Erik Jorpes, trans, from the Swedish MS. by Barbara Steele. Pp. 156, illus. 1966. Stockholm: Almqvist & Wiksell. Price not stated. British Journal for the History of Science, 3(4):403–404, December 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024973.

Russell:1968:BRN

[Rus68] C. A. Russell. Book review: Nineteenth Century: Humphry Davy. By Sir Harold Hartley. Pp. viii + 160. 9 plates. London: Nelson. 1966. 35s. British Journal for the History of Science, 4(1):77–78, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024995.

Russell:1969:BRC

[Rus69] C. A. Russell. Book review: Chemistry Source Book in Chemistry, 1900–1950. (Ed.) Henry M. Leicester. Cambridge, Mass.: Harvard University Press. 1968. Pp. xvii + 408. $11.95. British Journal for the History of Science, 4 (4):412, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025243.

Russell:1972:BRG

[Rus72a] C. A. Russell. Book review:

General The Science of Matter. A Historical Survey. Selected Readings. Ed. by M. P. Crosland. London: Penguin Books, 1971. Pp. 440. 70p. A History of Medicine. Selected Readings. Ed. by Lester S. King. London: Penguin Books, 1971. Pp. 316. 60p. British Journal for the History of Science, 6(2):205–206, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025293.

Russell:1972:BRS

[Rus72b] J. L. Russell. Book review: Sixteenth and Seventeenth Centuries: Johannes Kepler: Selbstzeugnisse. Selected by Franz Hammer. Stuttgart-Bad Cannstatt: Friedrich Frommann Verlag (Günther Holzboog), 1971. Pp. 97. No price stated. British Journal for the History of Science, 6(1): 92–93, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025273.

Russell:1974:BRS

[Rus74] J. L. Russell. Book review: Sixteenth and Seventeenth Centuries: Internationales Kepler-Symposium. Weil der Stadt 1971. Referate und Diskussionen. Ed. by Fritz Krafft, Karl Meyer, and Bernhard Sticker. Hildesheim: Verlag Dr. H. A. Gerstenberg, 1973. Pp. xii + 490. DM 160. British Journal for the History of Science, 7(3):294–295, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025544.

Ruse:1976:CLP

[Rus76a] Michael Ruse. Charles Lyell and the philosophers of science. British Journal for the History of Science, 9(2):121– 131, July 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025800.

Russell:1976:ARB

[Rus76b] John L. Russell. Action and reaction before Newton. British Journal for the History of Science, 9(1):25–38, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025704.

Russell:1977:BRA

[Rus77a]  C. A. Russell.  Book re-

view: Agriculture Crop Nutrition: Science and Practice before Liebig. By G. E. Fussell. Lawrence, Kansas: Coronado Press, 1971. Pp. 232. No price stated. British Journal for the History of Science, 10(2):

168, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025875.

Russell:1977:BRG

[Rus77b] John L. Russell. Book review:

General Science and Religious Belief. A Selection of Recent Historical Studies. Ed. by C. A. Russell. London: University of London Press in association with The Open University Press, 1973. Pp. viii + 323. £1.70 (paperback). British Journal for the History of Science, 10(1):70–71, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025585.

Russell:1977:BRS

[Rus77c] John L. Russell. Book review: Seventeenth and Eighteenth Centuries: Thomas White’s De Mundo Examined. By Thomas Hobbes. Translated by Harold Whitmore Jones. London: Bradford University Press in conjunction with Crosby Lockwood Staples, 1976. Pp. xi + 518. £18.00. British Journal for the History of Science, 10(3): 261–262, November 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025851.

Russell:1979:BRN

[Rus79a]  C. A. Russell.  Book re-

view: Nineteenth and Twentieth Centuries: Pflanzenchemie zwischen Lavoisier und Liebig. By Reinhard Löw. Straubing & München: Donau-Verlag, 1977. Pp. vii + 404 + 75. DM 29. British Journal for the History of Science, 12(2): 231–232, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025765.

Russell:1979:BRR

[Rus79b] John L. Russell. Book review: Renaissance: Copernicus: On the Revolutions of the Heavenly Spheres. A new translation from the Latin, with an introduction and notes by A. M. Duncan. Newton Abbot: David & Charles; New York: Barnes & Noble, 1976. Pp. 328. £12.50. British Journal for the History of Science, 12(1):94–95, March 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025670.

Russell:1983:SS

[Rus83a] Colin A. (Colin Archibald) Russell. Science and social, 1700–1900. MacMillan Publishing Company, New York, NY, USA, 1983. ISBN 0-333-29272-3. xvii + 307 pp. LCCN ????

Russell:1983:SSC

[Rus83b] Colin Archibald Russell. Science and social change: 1700– 1900. Themes in comparative history. MacMillan Publishing Company, New York, NY, USA, 1983. ISBN 0-333-29272-3 (hardcover), 0-333-29273-1 (paperback). xvii + 307 + 16 pp. LCCN Q175 .5. US$15.00.

Russell:1985:BRR

[Rus85] C. A. Russell. Book review:

Robert Friedel Pioneer Plastic. The making and selling of celluloid. Maddison, Wisconsin: The University of Wisconsin Press, 1983. Pp. xix + 153. ISBN 0-299-09170-8. £15.00. British Journal for the History of Science, 18(2): 240, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026339.

Russell:1986:LCE

[Rus86a] Colin Archibald Russell. Lancastrian chemist: the early years of Sir Edward Frankland. Open University Press, Milton Keynes, England, 1986. ISBN 0-335-15175-2. ix + 187 pp. LCCN QD22.F65 R87 1986.

Russell:1986:LEL

[Rus86b] Nicholas Russell. Like engend’ring like: heredity and animal breeding in early modern England. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-30657-4. ix + 271 pp. LCCN SF53 .R87

1986. URL http://www.loc.
gov/catdir/description/
cam031/85021350.html;
http://www.loc.gov/
catdir/enhancements/
fy0665/85021350-t.html.

Ruse:1987:BRR

[Rus87a] Michael Ruse. Book review: Robert M. Young. Darwin’s Metaphor: Nature’s Place in Victorian Culture. Cambridge: Cambridge University Press, 1985. Pp. xvii + 341. ISBN 0-521-31742-8. £27.50, $44.50 (hardcover), £9.95, $15.95 (paperback). British Journal for the History of Science, 20(1): 118–119, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026471.

Russell:1987:BRD

[Rus87b] Colin A. Russell. Book review: David C. Lindberg & Ronald L. Numbers (eds). God and Nature: Historical Essays on the Encounter between Christianity and Science. Berkeley, Los Angeles and London: University of California Press, 1986. Pp. xi + 516. ISBN 0-520-05538-1, £42.50 (hardcover); 0-520-04592-2, £15.25 (paperback). British Journal for the History of Science, 20(3): 355–356, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026366.

Ruse:1988:PBT

[Rus88a]  Michael Ruse.  Philosophy

of biology today. SUNY series in philosophy and biology. State University of New York Press, Albany, NY, USA, 1988. ISBN 0-88706-910-X, 0-88706-911-8 (paperback). x + 155 pp. LCCN QH331 .R8793 1988.

Russell:1988:RDB

[Rus88b]  Colin A. Russell.  ‘rude

and disgraceful beginnings’: A view of history of chemistry from the Nineteenth Century. British Journal for the History of Science, 21(3):273–294, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026830.

Russell:1988:OAO

[Rus88c] Nicholas Russell. Oswald avery and the origin of molecular biology. British Journal for the History of Science, 21(4): 393–400, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026959.

Russell:1988:THB

[Rus88d]  Nicholas Russell.  Towards

a history of biology in the twentieth century: Directed autobiographies as historical sources. British Journal for the History of Science, 21(1): 77–89, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026863.

Russ:1989:BRR

[Rus89a] Steve Russ. Book review: Rolf Herken (ed.). The Universal Turing Machine: A Half-Century Survey. Oxford: Oxford University Press, 1988. Pp. xiv + 661. ISBN 0-19-853741-7. £55.00. British Journal for the History of Science, 22(4): 451–452, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026930.

Russett:1989:SSV

[Rus89b] Cynthia Eagle Russett. Sexual science: the Victorian construction of womanhood. Harvard University Press, Cambridge, MA, USA, 1989. ISBN 0-674-80290-X. 245 pp. LCCN QP81.5 .R87 1989. US$25.00.

Robinson:1972:SSL

[RW72] F. J. G. Robinson and P. J. Wallis. Scientists and subscription lists. British Journal for the History of Science, 6(2): 227–228, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Roberts:1990:JDL

[RWD90] R. J. (Richard Julian) Roberts, Andrew G. Watson, and John Dee, editors. John Dee’s library catalogue. Bibliographical Society, London, UK, 1990. ISBN 0-19-721795-8. viii + 253 + 170 pp. LCCN Z997.D298 J64

1990. URL http://www.loc.
gov/catdir/enhancements/
fy0637/90228267-d.html.

Ryan:1970:BRS

[Rya70] W. F. Ryan. Book review:

Scientific Instruments Astronomy Gnomonics. A Catalogue of Instruments of the Fifteenth to the Nineteenth Centuries in the Collections of the National Technical Museum, Prague. By Zdenek Horsky and Otilie Skopová. Prague. Pp.

202. 43 plates. 1968. Price not stated. British Journal for the History of Science, 5(2): 187–188, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025419.

Salthe:1985:EHS

[Sal85] Stanley N. Salthe. Evolving hierarchical systems: their structure and representation. Columbia University Press, New York, NY, USA, 1985. ISBN 0-231-06016-5. x + 343 pp. LCCN QH83 .S26 1985.

Sangster:1987:BRJ

[San87] James Sangster. Book review:

J. L. Heilbron. The Dilemmas of an Upright Man: Max Planck as Spokesman for German Science. Berkeley: University of California Press, 1986. Pp. xiii + 238. ISBN 0-520-05710-4. £15.25. British Journal for the History of Science, 20(1):123–124, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026475.

Sangwan:1988:IRE

[San88] Satpal Sangwan. Indian response to European science and technology 1757–1857. British Journal for the History of Science, 21(2):211–232, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026980.

Sangwan:1991:STC

[San91] Satpal Sangwan. Science, technology and colonisation: an Indian experience, 1757–1857. Anamika Prakashan, Delhi, India, 1991. ISBN 81-85150-11-7. xviii + 196 pp. LCCN MQ.

Sapp:1990:WTL

[Sap90] Jan Sapp. Where the truth lies: Franz Moewus and the origins of molecular biology. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-36550-3, 0-521-36751-4 (paperback). ix + 340 pp. LCCN QH506 .S27 1990. URL http:

                                        //www.loc.gov/catdir/
                                        description/cam023/
                                        89032582.html;   http://

www.loc.gov/catdir/toc/
cam026/89032582.html.

Sardar:1981:BRI

[Sar81] Ziauddin Sardar. Book review: Islamic Science Al-Biruni Commemorative Volume. Proceedings of the International Congress held in Pakistan on the Occasion of Millenary of Abu Raihan Muhammad ibn Ahmad al-Biruni (c 973-1051 AD), November 26 1973 thru’ December 12 1973. Ed. By Hakim Mohammed Said, Karachi: Hamdard National Foundation, 1979. Pp. vi + 844. PRs 200/$30.00. British Journal for the History of Science, 14(3):285–286, November 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025986.

Sarjeant:1987:GHG

[Sar87] William Antony S. Sarjeant.

Geologists and the history of geology: an international bibliography from the origins to 1978. Supplement, 1979–1984 and additions. R.E. Krieger Pub. Co., Malabar, Fla., U.S.A., 1987. ISBN 0-89874-939-5. 2 v. pp. LCCN Z6031 .S28 Suppl.

Sarkar:1992:FEG

[Sar92] Sahotra Sarkar, editor. The founders of evolutionary genetics: a centenary reappraisal, volume 142 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346. 300 pp. LCCN Q174 .B67 vol. 142. URL http:

                                        //www.loc.gov/catdir/
                                        enhancements/fy0823/
                                        92012823-d.html;   http:/
                                        /www.loc.gov/catdir/
                                        enhancements/fy0823/
                                        92012823-t.html;   https:
                                        //link.springer.com/book/
                                        10.1007/978-94-011-2856-8.

Smith:1992:DRA

[SB92] Murphy D. Smith and David Borodin. Due reverence: antiques in the possession of the American Philosophical Society, volume 203 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1992. ISBN 0-87169-184-1. vi + 77 pp. LCCN Q11 .P612 vol. 203.

Shea:2009:GSN

[SB09] William R. Shea and Tiziana Bascelli. Galileo’s Sidereus Nuncius, or, a sidereal message. Science History Publications, Sagamore Beach, MA, USA, 2009. ISBN 0-88135-375-2. viii + 124 pp. LCCN QB41 .G173 2009.

Stayer:1988:ND

[SC88] Marcia Sweet Stayer and Boris Castel, editors. Newton’s dream. Published for Queen’s quarterly by McGill–Queen’s University Press, Montréal, Québec, Canada, 1988. ISBN 0-7735-0689-6. LCCN QA803 .N487

1988. URL http://www.loc.
gov/catdir/enhancements/
fy1001/89121159-d.html.

Schofield:1965:GMH

[Sch65] Christine Schofield. The geoheliocentric mathematical hypothesis in sixteenth century planetary theory. British Journal for the History of Science, 2(4):291–296, December 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024887.

Schmitt:1968:BRS

[Sch68a] C. B. Schmitt. Book review: Sixteenth and Seventeenth Centuries: A Catalogue of Sixteenth-Century Printed Books in the National Library of Medicine. Compiled by Richard J. Durling. Bethesda, Maryland. U.S. Department of Health, Education, and Welfare. Public Health Service, National Library of Medicine. 1967. Pp. xii + 698. $5.00. British Journal for the History of Science, 4(2):187–188, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025160.

Schonland:1968:A

[Sch68b] Sir Basil Schonland. The atomists (1805–1933). Clarendon Press, New York, NY, USA, 1968. 10 + 198 pp. LCCN QC173 .S2832.

Scherz:1969:BRH

[Sch69a] G. Scherz. Book review: History of Geology The Prodromus of Nicolaus Steno’s Dissertation concerning a Solid Body enclosed by Process of Nature within a Solid. An English Version with an introduction and explanatory notes by John Garret Winter. Foreword by William H. Hobbs (Facsimile of University of Michigan Humanistic Studies, vol. xi, pt. 2, 1916). Introduction by George W. White. New York and London: Hafner Publishing Company, Inc. 1968. Pp. vii + 115. $12.50. British Journal for the History of Science, 4(3):294, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025486.

Schmitt:1969:BRM

[Sch69b] C. B. Schmitt. Book review: Middle Ages Albertus Magnus, Book of Minerals. Translated by Dorothy Wyckoff. London: Oxford University Press: Clarendon Press. 1967. Pp. xlii + 309. 84s. British Journal for the History of Science, 4 (4):418, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025250.

Schmitt:1970:BRS

[Sch70] C. B. Schmitt. Book review: Sixteenth and Seventeenth Centuries: Art, Science, and History in the Renaissance. Edited by Charles S. Singleton. Baltimore: Johns Hopkins Press. 1967. Pp. vii + 446. £5 19s. British Journal for the History of Science, 5(1):98–99, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025368.

Schmitt:1971:BRB

[Sch71] C. B. Schmitt. Book review: Biological Sciences and Medicine A History of Bubonic Plague in the British Isles. By J. F. D. Shrewsbury. London: Cambridge University Press. 1970. Pp. xi + 661, 4 plates. £8. British Journal for the History of Science, 5 (3):302–303, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025331.

Schmitt:1972:BRG

[Sch72] C. B. Schmitt. Book review:

General Esquisse d’une Histoire de l’Idée de Nature. By Robert Lenoble. Ed. by Joseph Beaude. Paris: Éditions Albin Michel, 1969. Pp. 446. 9 francs. British Journal for the History of Science, 6(1):85, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025266.

Schmitt:1973:BRR

[Sch73a] C. B. Schmitt. Book review: Renaissance: La Dottrina della Scienza in Giacomo Zabarella. By Antonino Poppi. Centro per la storia della tradizione aristotelica nel Veneto, Saggi e testi, vol. 12. Padova: Editrice Antenore, 1972. Pp. 351. 5,500 lire. British Journal for the History of Science, 6(4): 440–442, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025514.

Schmitt:1973:BRS

[Sch73b] C. B. Schmitt. Book review: Sixteenth and Seventeenth Centuries: William Harvey and the Circulation of the Blood. By Gweneth Whitteridge. London: Macdonald, and New York: American Elsevier Inc., 1971. Pp. xiii + 269. £4. British Journal for the History of Science, 6(3): 320–321, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025455.

Schmitt:1974:BRSb

[Sch74a] C. B. Schmitt. Book review: Seventeenth Century: La Philosophie de Gassendi. Nominalisme, Matérialisme et Métaphysique. By Olivier René Bloch. The Hague: Martinus Nijhoff, 1971. Pp. xxx + 525. Hfl. 75.00. The Selected Works of Pierre Gassendi. Ed. and trans. by Craig B. Brush. New York and London: Johnson Reprint Corporation, 1972. Pp. xiv + 442. $25.00. British Journal for the History of Science, 7(2):188–189, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025213.

Schmitt:1974:BRSa

[Sch74b] C. B. Schmitt. Book review: Sixteenth and Seventeenth Centuries: Galileo’s Intellectual Revolution. By William R. Shea. London: Macmillan, 1972. Pp. xii + 204. £4.95. British Journal for the History of Science, 7(1):90–91, March 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025187.

Schmitt:1975:BRS

[Sch75] C. B. Schmitt. Book review:

Seventeenth Century: The New Science of Motion. A Study of Galileo’s ‘De motu locali’. By Winifred L. Wisan. Published in Archive for history of exact sciences, xiii, number 2/3. Berlin, Heidelberg, New York: Springer-Verlag, 1974. Pp. 204. No. price stated. British Journal for the History of Science, 8(3):256–257, November 1975.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025568.

Schmitt:1976:BRM

[Sch76] C. B. Schmitt. Book review: Middle Ages A Source Book in Medieval Science. Ed. by Edward Grant. Cambridge, Mass: Harvard University Press, 1974. Pp. xviii + 864. £16.25. British Journal for the History of Science, 9 (1):74–75, March 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025712.

Schmitt:1978:BRA

[Sch78a]  Charles  B.

Schmitt. Book review: Antiquity Aristotle’s Posterior Analytics. Translated with Notes by Jonathan Barnes. Oxford: Clarendon Press: Oxford University Press, 1975. Pp. xix + 277. £3.00 (paperback). British Journal for the History of Science, 11(1):73–74, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025616.

Schmitt:1978:BRR

[Sch78b] Charles B. Schmitt. Book review: Renaissance: Touches of Sweet Harmony. Pythagorean Cosmology and Renaissance Poetics. By S. K. Heninger Jr, San Marino, California: The Huntington Library, 1974. Pp. xvii + 446 + 52 plates. $19.50. Pythagorean Palaces. Magic and Architecture in the Italian Renaissance. By G. L. Hersey. Ithaca & London: Cornell University Press, 1976. Pp. 216 + plates + tables. £18. British Journal for the History of Science, 11(1):78–79, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025620.

Schneer:1978:BRS

[Sch78c] Cecil J. Schneer. Book review:

Seventeenth and Eighteenth Centuries: L’etrenne, ou la neige sexangulaire. By Johann Kepler. Traduction critique par Robert Hallenx. Paris: Vrin, 1975. Pp. 169. No price stated. British Journal for the History of Science, 11(2):181– 182, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025740.

Schmitt:1979:BRR

[Sch79a] C. B. Schmitt. Book review: Renaissance: The Cosmographical Glass: Renaissance Diagrams of the Universe. By S. K. Heninger Jr., San Marino, California: The Huntington Library, 1977. Pp. xx + 209; 117 illustrations. $17.50. British Journal for the History of Science, 12(2): 225–227, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025761.

Schmitt:1979:SNB

[Sch79b] C. B. Schmitt. Short notices of books: Ricerche sull’atomismo del seicento. By Ugo Baldini, Giancarlo Zanier, Paolo Farina, and Francesco Trevisani. Florence: La Nuova Italia, 1977. Pp. 223. L. 5,000. British Journal for the History of Science, 12(1):103, March 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025676.

Schuster:1979:KLH

[Sch79c] John A. Schuster. Kuhn and Lakatos and the history of science: Kuhn and Lakatos revisited. British Journal for the History of Science, 12(3): 301–317, November 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026006.

Schaffer:1980:HBN

[Sch80a] Simon Schaffer. Herschel in Bedlam: natural history and stellar astronomy. British Journal for the History of Science, 13(3):211–239, November 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026197.

Schmitt:1980:BRMb

[Sch80b] C. B. Schmitt. Book review:

Middle Ages and Renaissance: Archimedes in the Middle Ages. Volume iii: The Fate of the Medieval Archimedes, 1300 to 1565. By Marshall Clagett. Philadelphia: American Philosophical Society, 1978 (Memoirs of the American Philosophical Society, vol. 125). Pp. 1582 in 3 tomes. $75.00. British Journal for the History of Science, 13(2):163–164, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025898.

Schmitt:1980:BRMa

[Sch80c] Charles B. Schmitt. Book review: Middle Ages Alle origini dell’aristotelismo padovano (sec. XII-XIII). By Paolo Marangon. Padua: Editrice Antenore, 1977. Pp.

189. L. 7,000. British Journal for the History of Science, 13 (1):67–68, March 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026127.

Schmitt:1980:BRR

[Sch80d] Charles B. Schmitt. Book review: Renaissance: Man and Nature in the Renaissance. By Allen G. Debus. Cambridge, London, New York, Melbourne: Cambridge University Press, Pp. + 159. £7.95/£2.50 (paperback). British Journal for the History of Science, 13(1): 68–70, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026128.

Schaffer:1981:BRS

[Sch81a] Simon Schaffer. Book review: Seventeenth and Eighteenth Centuries: Philosophers at war: the quarrel between Newton and Leibniz. By A. Rupert Hall. Cambridge: Cambridge University Press, 1980. Pp. xiv + 338. £15.00. British Journal for the History of Science, 14(2):213–214, July 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025955.

Schmitt:1981:BRR

[Sch81b] Charles B. Schmitt. Book review: Renaissance and Seventeenth Century: Venezia e la peste, 1348–1797. By Comune di Venezia, Assessorato alla Culture e Belle Arti. Venice: Marsilio, 1979. pp.

380. L25,000. La scienza a corte. Collezionismo eclettico, natura e immagine a Mantova fra Rinascimento e Manierisme. By Dario A. Franchini et al. Rome: Bulzoni, 1979. Pp. 280. L22,000. Firenze e la Toscana dei Medici nell’ Europa del Cinquecento. Florence: Edizioni Medicee, 1980. Pp. 438. No price stated. Livorno e Pisa: due città e un territorio nella politica dei Medici.

Pisa: Nistri-Lischi & Pacini, 1980. Pp. 599. No price stated. British Journal for the History of Science, 14(3):287– 289, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025987.

Schaffer:1983:BRL

[Sch83a] Simon Schaffer. Book review:

Lives and Works H. J. M. Bos, M. J. S. Rudwick, H. A. M. Melders and R. P. W. Visser (eds), Studies on Christiaan Huygens, Lisse: Swets & Zeitlinger, 1980. Pp. vi + 321. Dfl.65. British Journal for the History of Science, 16 (2):199–201, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026243.

Schmitt:1983:BRA

[Sch83b] C. B. Schmitt. Book review: Alchemy and Parapsychology Antoine Faivre and Rolf Christian Zimmermann (eds.), Epochen der Naturmystik. Hermetische Tradition im wissenschaftlichen Fortschritt, Berlin: Erich Schmidt Verlag, 1979. Pp. 459. DM. 128. British Journal for the History of Science, 16(1):99, March 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026107.

Schmitt:1983:AR

[Sch83c] Charles B. Schmitt. Aristotle and the Renaissance, volume 27 of Martin classical lectures. Published for Oberlin College by Harvard University Press, Cambridge, MA, USA, 1983. ISBN 0-674-04525-4. viii + 187 pp. LCCN PA25 .M3 vol. 27.

Schaffer:1984:BRN

[Sch84a] S. Schaffer. Book review: National Traditions in Science Dirk J. Struik, The land of Stevin and Huygens: a sketch of science and technology in the Dutch Republic during the golden century. Dordrecht. Boston & London: D. Reidel, 1981. Pp. xx + 162. Dfl. 65/$34.00; Dfl. 38/$18.50 (paperback). British Journal for the History of Science, 17(1): 85–87, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026021.

Schaffer:1984:BRE

[Sch84b] Simon Schaffer. Book reviews: Enlightenment Robert E. Sullivan, John Toland and the deist controversy. (Harvard Historical Studies No. 101.) Cambridge, Massachusetts: Harvard University Press, 1982. Pp. viii + 355 ISBN 0-674-48050-3. £19.25. François Duchesneau. La physiologic des lumières. Empirisme, modèles et théories (International Archives of the History of Ideas; 95). The Hague: Martinus Nijhoff, 1982. Pp. xxi + 611. ISBN 90-247-2500-3. Cloth, Dfl. 185.00, US 397.00. British Journal for the History of Science, 17(1): 117–118, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026055.

Schaffer:1984:BRH

[Sch84c] Simon Schaffer. Book reviews: History of Mathematical Sciences: Huygens et la France. Foreword by René Taton. Paris: Vrin. 1982. Pp. ix + 268. ISBN 2-7116-2018-2. 210F. Henry Guerlac, Newton on the Continent. Ithaca: Cornell University Press, 1981. Pp. 169. ISBN 0-8014-1409-1. £8.75. Marie-Françoise Biarnais, Les Principia de Newton: Genèse et structure des chapitres fondamentaux avec traduction nouvelle. Foreward by A. Rupert Hall. Paris: Centre Nationale de la Recherche Scientifique, 1982. Pp. 287. ISBN 2-222-03094-3. (Cahiers d’Histoire et de Philosophie des Sciences, Nouvelle série, No. 2.) 25F. British Journal for the History of Science, 17(2): 227–230, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026562.

Schutt:1984:EIF

[Sch84d] Hans-Werner Schütt. Die Entdeckung des Isomorphismus: eine Fallstudie zur Geschichte der Mineralogie und der Chemie. (German) [The discovery of isomorphicism: a case study in the history of mineralogy and chemistry], volume 9 of Arbor scientiarum. Reihe A, Abhandlungen. Gerstenberg Verlag, Hildesheim, Germany, 1984. ISBN 3-8067-0902-5. 324 pp. LCCN QD951 .S48 1984.

Schmitt:1985:BRD

[Sch85a] Charles B. Schmitt. Book review: David C. Lindberg, Roger Bacon’s Philosophy of Nature. A Critical Edition, with English Translation, Introduction, and Notes, of De Multiplicatione speciarum and De speculis comburentibus. Oxford: Clarendon Press, 1983. Pp. lxxxi + 420. ISBN 0-19-858164-5. £40.00. British Journal for the History of Science, 18(1):95–96, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026278.

Schulze:1985:RSP

[Sch85b] Ludmilla Schulze. The Russification of the St. Petersburg Academy of Sciences and Arts in the eighteenth century. British Journal for the History of Science, 18(3):305– 335, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026383.

Schaffer:1986:BRJ

[Sch86a] Simon Schaffer. Book review: James E. Force. William Whiston: Honest Newtonian. Cambridge: Cambridge University Press, 1985. Pp. xxiii + 208. ISBN 0-521-26590-8. £25. British Journal for the History of Science, 19(2):226– 228, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026610.

Schaffer:1986:BRM

[Sch86b] Simon Schaffer. Book review: Michel Blay. La Conceptualisation newtonieene des Phénomènes de la Couleur. Paris: Vrin, 1983. Pp. 307. ISBN 2-7116-0825-5. 99F (paperback). British Journal for the History of Science, 19(1): 123–124, March 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026496.

Schaffer:1987:BRM

[Sch87] Simon Schaffer. Book review: Michael J. Crowe. The Extraterrestrial Life Debate, 1750–1900: the Idea of a Plurality of Worlds from Kant to Lowell. Cambridge: Cambridge University Press, 1986.

Pp. xix + 680. ISBN 0-521-26305-0. £40.00. British Journal for the History of Science, 20(4):485–486, October 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026425.

Schivelbusch:1988:DNI

[Sch88] Wolfgang Schivelbusch. Disenchanted night: the industrialisation of light in the Nineteenth Century. Berg, Oxford, UK, 1988. ISBN 0-85496-191-7. 227 pp. LCCN GT440 .S3613x 1988b.

Scholz:1989:SGD

[Sch89] Erhard Scholz, editor. Symmetrie, Gruppe, Dualität: zur Beziehung zwischen theoretischer Mathematik und Anwendungen in Kristallographie und Baustatik des 19. Jahrhunderts. (German) [Symmetry, group, duality: on the relationship between theoretical mathematics and applications in crystallography and structural design of the 19th Century]. Science networks historical studies. Birkhäuser Verlag, Basel, Switzerland, 1989. ISBN 3-7643-1974-7. ISSN 1421-6329. 406 pp. LCCN QD911 .S324 1989b.

Schabas:1990:WRN

[Sch90]  Margaret  Sch-

abas. A world ruled by number: William Stanley Jevons and the rise of mathematical economics. Princeton University Press, Princeton, NJ, USA, 1990. ISBN 0-691-08543-9. xii + 192 pp. LCCN HB103.J5 S33 1990. US$29.95. URL http://

www.loc.gov/catdir/
description/prin031/
89070245.html.

Schaaf:1992:SHT

[Sch92a] Larry J. (Larry John) Schaaf.

Out of the shadows: Herschel, Talbot and the invention of photography. Yale University Press, New Haven, CT, USA, 1992. ISBN 0-300-05705-9 (hardcover). xii + 188 pp. LCCN TR57 .S32 1992. US$65.00.

Schutt:1992:EMB

[Sch92b]  Hans-Werner Schütt.  Eil-

hard Mitscherlich, Baumeister am Fundament der Chemie. (German) [Eilhard Mitscherlich, master builder of the basis of chemistry], volume n.F., Bd. 8 of Abhandlungen und Berichte. In Kommission bei R. Oldenbourg, München, Germany, 1992. ISBN 3-486-26273-4. 192 pp. LCCN AM101 .M9743 Bd. 8.

Schneider:1993:JFR

[Sch93] Ivo Schneider. Johannes Faulhaber: 1580–1635; Rechenmeister in einer Welt des Umbruchs. (German) [Johannes Faulhaber: 1580–1635: Computing master in a world of radical change], volume 7 of Vita Mathematica. Birkhäuser Verlag, Basel, Switzerland, 1993.

ISBN 3-7643-2919-X. xiv + 271 pp. LCCN QA29.F38 S35

                                        1993. URL http://catalog.
                                        hathitrust.org/api/
                                        volumes/oclc/29955228.
                                        html.

Schsler:2001:DLF

[Sch01] Jörn Schösler, editor. La diffusion de Locke en France: traduction au dix-huitiéme siècle: lectures de Rousseau. (French) [The diffusion of Locke in France: 18th Century translation of the readings of Rousseau], volume 2001:04 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2001. ISBN 0-7294-0748-9. v + 364 pp. LCCN MLCM 2005/02528 (B).

Scott:1967:BRM

[Sco67] J. F. Scott. Book review: The Mathematical Practitioners of Hanoverian England 1714–1840. By E. G. R. Taylor. Cambridge University Press for the Institute of Navigation. Pp. xv + 502. Diagrams. 1966. 84s. British Journal for the History of Science, 3(3):300, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025061.

Scriba:1964:RNC

[Scr64a] C. J. Scriba. Report on the Ninth Colloquium on the History of Mathematics. British Journal for the History of Science, 2(2):162–164, December 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025015.

Scriba:1964:MKT

[Scr64b] Christoph J. Scriba. Mercator’s Kinckhuysen-translation in the Bodleian Library at Oxford. British Journal for the History of Science, 2(1): 45–58, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025083.

Scriba:1968:BRC

[Scr68a]  C.  J.

Scriba. Book review: Classical Antiquity Die Fragmente des Eudoxos von Knidos. Herausgegeben, übersetzt und kommentiert von François Lasserre. (Texte und Kommentareeine altertumswissenschaftliche Reihe, Bd. 4.) Berlin: Walter de Gruyter & Co. 1966. Pp. viii + 299. DM 52. British Journal for the History of Science, 4(2):186–187, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025159.

Scriba:1968:BRH

[Scr68b] Christoph J. Scriba. Book review: History of Mathematics Arithmetical Books from the Invention of Printing to the Present Time. By Augustus de Morgan. London, Taylor and Walton, 1847. Reprinted with an Introduction by A. Rupert Hall. Pp. (7) + xxviii + 124. London: Hugh K. Elliott Ltd. 1966. £5 5s. British Journal for the History of Science, 4(1):85–86, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025002.

Stachel:1987:CPA

[SCS^{+}87] John J. Stachel, David C. Cassidy, Robert Schulmann, Jürgen Renn, Olga Griminger, Gary Smith, and Robert Summerfield, editors. The collected papers of Albert Einstein. Vol. 1. The early years, 1879–1902. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-08407-6. lxvi + 433 pp. LCCN QC16.E5 A2 1987. US$52.50. URL ftp:/

/uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/
catdir/description/
prin021/86043132.html;
http://www.loc.gov/
catdir/toc/prin032/
86043132.html.  Published

with a translation into English by Anna Beck, ISBN 0-691-08475-0 [BH87], which is not available separately. Consultant for the translation: Peter Havas.

Siegfried:1976:HDG

[SD76] Robert Siegfried and R. H. Dott. Humphry Davy as geologist, 1805–1829. British Journal for the History of Science, 9(2):219–227, July 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025808.

Snow:1991:NIV

[SE91] John Snow and Richard H. Ellis. On narcotism by the inhalation of vapours. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1991. ISBN 1-85315-158-0. xxx + 112 pp. LCCN RD81 .S663 1991.

Secord:1983:BRL

[Sec83]  J. Secord.  Book review:

Life and earth sciences: Colin Speakman, Adam Sedgwick: geologist and dalesman. Heathfield, East Sussex: The Broad Oak Press, 1982. Pp. 145. £5.75 (sewn paperback). ISBN 0-906716-01-2. British Journal for the History of Science, 16(3):286–287, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026156.

Secord:1984:BRH

[Sec84] James A. Secord. Book review:

History of Natural History: William Glen, The Road to Jaramillo: Critical Years of the Revolution in Earth Science. Stanford, California: Stanford University Press, 1982. Pp. xx + 459. ISBN 0-8047-1119-4.

$37.50. British Journal for the History of Science, 17(3): 316–318, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026630.

Secord:1985:BRI

[Sec85]  James Secord.  Book re-

view: Ian Inkster and Jack Morrell (eds.), Metropolis and Province: Science in British Culture, 1780–1850. London: Hutchinson, 1983. Pp. 288. ISBN 0-09-145180-9. £17.50. British Journal for the History of Science, 18(1):111– 113, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026292.

Secord:1986:CVG

[Sec86] James A. Secord. Controversy in Victorian geology: the Cambrian–Silurian dispute. Princeton University Press, Princeton, NJ, USA, 1986. ISBN 0-691-08417-3. xvii + 363 pp. LCCN QE656 .S4 1986. US$49.50.

Secord:1988:BRD

[Sec88] J. A. Secord. Book review: Dorothy Stein. Ada: A Life and a Legacy. Cambridge, Mass.: The MIT Press, 1985. Pp. xix + 321. ISBN 0-262-19242-X. £17.50 (hardcover), £8.95 (paperback). British Journal for the History of Science, 21(3):

375–377, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026847.

Segre:1991:WG

[Seg91] Michael Segre. In the wake of Galileo. Rutgers University Press, New Brunswick, NJ, USA, 1991. ISBN 0-8135-1700-1, 0-8135-1701-X (paperback). xix + 192 pp. LCCN Q127.I8 S44 1991.

Segal:1994:FIM

[Seg94] Howard P. Segal. Future imperfect: the mixed blessings of technology in America. University of Massachusetts Press, Amherst, MA, USA, 1994. ISBN 0-87023-881-7 (hardcover), 0-87023-882-5 (paperback). xviii + 245 pp. LCCN T21 .S43 1994. US$40.00, US$15.95.

Seidel:1985:BRA

[Sei85] Robert W. Seidel. Book review:

Albert E. Moyer, American Physics in Transition: A History of Conceptual Change in the Late Nineteenth Century, Los Angeles: Tomash, 1983. Pp. xx + 218. ISBN 0-938228-06-4. $30.00. British Journal for the History of Science, 18(3):353–354, November 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026394.

Sellers:2010:CCV

[Sel10] Susan Sellers, editor. The Cambridge companion to Virginia Woolf. Cambridge companions to literature. Cambridge University Press, Cambridge, UK, second edition, 2010. ISBN 0-521-89694-0 (hardcover), 0-521-72167-9 (paperback). xxi + 272 pp. LCCN PR6045.O72 Z5655 2010. URL http://

assets.cambridge.org/
97805218/96948/cover/
9780521896948.jpg.

Sepper:1988:GCN

[Sep88] Dennis L. Sepper. Goethe contra Newton: polemics and the project for a new science of color. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-34254-6. xvi + 222 pp. LCCN QC495 .S43

1988. URL http://www.loc.
gov/catdir/description/
cam031/87021463.html;
http://www.loc.gov/
catdir/toc/cam031/
87021463.html.

Series:1988:BRJ

[Ser88] Caroline Series. Book review: Jeremy Gray. Linear Differential Equations and Group Theory from Riemann to Poincaré. Basel & Boston: Birkhauser, 1986. Pp. xxv + 460. ISBN 3-7643-3318-9 (Basel), 0-8176-3318-9 (Boston). $51.00. British Journal for the History of Science, 21(1): 112–114, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026869.

Seymour:1977:BRA

[Sey77] P. A. H. Seymour. Book review: Astronomy The Milky Way: An Elusive Road for Science. By Stanley L. Jaki. Newton Abbot: David and Charles, 1973. Pp. xi + 352. £5.95. British Journal for the History of Science, 10(2): 162–163, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025872.

Sadoun-Goupil:1980:RES

[SG80] Michelle Sadoun-Goupil. Revue de L’essai de statique chimique, Claude-Louis Berthollet: édition critique d’un projet de seconde édition de L’essai de statique chimique. (French) [Review of the Test of chemical statistics, Claude-Louis Berthollet: the critical edition of a project of the second edition of Test of chemical statistics]. L’École Polytechnique, Paris, France, 1980. ISBN 2-7302-0019-3. vii + 204 pp. LCCN QD28

.B533. URL http://catalog.
hathitrust.org/api/
volumes/oclc/9259834.html.

Schofield:1990:BHS

[SG90] Robert E. Schofield and Elizabeth Garber, editors. Beyond history of science: essays in honor of Robert E. Schofield. Lehigh University Press, Bethlehem, PA, USA, 1990. ISBN 0-934223-11-4. 325 pp. LCCN Q126.9 .B49 1990.

Sarasohn:1996:GEF

[SG96] Lisa T. Sarasohn and Pierre Gassendi. Gassendi’s ethics: freedom in a mechanistic universe. Cornell University Press, Ithaca, NY, USA, 1996. ISBN 0-8014-2947-1 (hardcover). xii + 236 pp. LCCN B1887 .S27 1996.

Silver:2000:FTL

[SGŚ00]  Marie-France  Alberte

Silver and Marie-Laure Girou-Świderski, editors. Femmes en toutes lettres: les épistolières du XVIIIe siècle. (French) [Women of all letters: the epistolieres of the 18th Century], volume 2000:04 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2000. ISBN 0-7294-0741-1. xiii + 277 pp. LCCN PQ711 .F45 2000.

Shapin:1973:TCH

[SH73] Steven Shapin and Susan Hill. The Turner Collection of the history of mathematics at the University of Keele. British Journal for the History of Science, 6(3):336–337, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025466.

Shapin:1973:BRS

[Sha73a]  Steven Shapin.  Book re-

view: Science and Society Science and Society 1600–1900. Ed. by Peter Mathias. London: Cambridge University Press, 1972. Pp. viii + 166. £2.80. British Journal for the History of Science, 6(4):436– 437, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025509.

Sharp:1973:BRS

[Sha73b]  Lindsay Sharp.  Book re-

view: Sixteenth and Seventeenth Centuries: A Bibliography of Sir William Petty, F.R.S. and of Observations on the Bills of Mortality by John Graunt, F.R.S. By Sir Geoffrey Keynes. Oxford: Clarendon Press: Oxford University Press, 1971. Pp. xviii + 103. £5.50. British Journal for the History of Science, 6 (3):321–322, June 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025456.

Shapin:1974:PPP

[Sha74]  Steven Shapin.  Property,

patronage, and the politics of science: The founding of the Royal Society of Edinburgh. British Journal for the History of Science, 7(1): 1–41, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025174.

Shapin:1977:BRN

[Sha77a]  Steven Shapin.  Book re-

view: Nineteenth and Twentieth Centuries: Conquest of Mind: Phrenology and Victorian Social Thought. By David de Giustino. London: Croom Helm, 1975. Pp. viii + 248. £6.00. British Journal for the History of Science, 10(2): 177–179, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025883.

Shapiro:1977:BRS

[Sha77b] Alan E. Shapiro. Book review: Seventeenth and Eighteenth Centuries: The Gresham Lectures of John Flamsteed. Ed. by Eric G. Forbes. London: Mansell, 1975. Pp. xvii + 479. £18. British Journal for the History of Science, 10(3):264–266, November 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025853.

Shaffer:1978:BRN

[Sha78]  E. S. Shaffer.  Book re-

view: Nineteenth and Twentieth Centuries: The Secularization of the European Mind in the Nineteenth Century. The Gifford Lectures in the University of Edinburgh, 1973-4. By Owen Chadwick. Cambridge: Cambridge University Press, 1975. Pp. v + 286. £6.50. (Paperback, 1978, £3.95). British Journal for the History of Science, 11(3): 292–294, November 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025791.

Shapin:1979:BRG

[Sha79] Steven Shapin. Book review: General Science, Technology and Society: A Cross-Disciplinary Perspective. Edited by Ina Spiegel-Rösing and Derek de Solla Price. London and Beverly Hills: Sage, 1977. Pp. xi + 607. £20.00. British Journal for the History of Science, 12 (1):90–91, March 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025667.

Shapin:1981:BRR

[Sha81] Steven Shapin. Book review:

Renaissance and Seventeenth Century: Witch Hunting, Magic and the New Philosophy: An Introduction to Debates of the Scientific Revolution, 1450–1750. By Brian Easlea. Brighton: Harvester, 1980. Pp. xii + 283.£25.00/£8.50 (paperback). British Journal for the History of Science, 14(3):291– 292, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025989.

Shapin:1983:BRA

[Sha83a]  Steven Shapin.  Book re-

view: Alchemy and Parapsychology Marsha P. Hannen, Margaret J. Osler, and Robert G. Weyant (eds.), Science, pseudo-science and society, Waterloo, Ontario: Wilfrid Laurier University Press, for the Calgary Institute for the Humanities, 1980. Pp. x + 303. $7.50. British Journal for the History of Science, 16 (1):99–101, March 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026108.

Shapiro:1983:PCS

[Sha83b] Barbara J. Shapiro. Probability and certainty in seventeenthcentury England: a study of the relationships between natural science, religion, history, law, and literature. Princeton University Press, Princeton, NJ, USA, 1983. ISBN 0-691-05379-0, 0-691-10146-9 (lim. paperback ed.). x + 347 pp. LCCN DA380 .S53 1983. US$32.50, US$12.50.

Shaw:1984:WPS

[Sha84] John Shaw. Water power in Scotland, 1550–1870. John Donald, Edinburgh, Scotland, 1984. ISBN 0-85976-072-3. xi + 606 pp. LCCN TJ859 .S53; TC147. US$25.00.

URL Biblioth\`;   http://
catalogue.bnf.fr/ark:/
12148/cb34808984k.

Shapin:1986:BRW

[Sha86]  Steven Shapin.  Book re-

view: Wilda C. Anderson. Between the Library and the Laboratory: the Language of Chemistry in Eighteenth-Century France. Baltimore and London: The Johns Hopkins University Press, 1984. Pp. vii + 190. ISBN 0-8018-3229-2. £20.00, $22.50. British Journal for the History of Science, 19(3):361–362, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026532.

Shapin:1987:BRT

[Sha87] Steven Shapin. Book review:

Tom Sorrell. Hobbes (The Arguments of the Philosophers, series ed. Ted Honderich). London and New York: Routledge and Kegan Paul, 1986. Pp. xii + 163. ISBN 0-7100-9845-6. £14.95. British Journal for the History of Science, 20 (2):236–237, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026322.

Shaffer:1989:BRF

[Sha89]  E. S. Shaffer.  Book re-

view: Frederick Amrine, Francis J. Zucker, and Harvey Wheeler (eds) Goethe and the Sciences: A Reappraisal. Boston Studies in the Philosophy of Science Vol. 97. Dordrecht/Boston/Lancaster/ Tokyo: D. Reidel Publishing Company, 1987. Pp. xv + 442. ISBN 90-277-2265-X. Dfl. 190, £56.75, $79.00. British Journal for the History of Science, 22 (1):120–121, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026716.

Shapiro:1993:FPP

[Sha93] Alan E. (Alan Elihu) Shapiro. Fits, passions, and paroxysms: physics, method, and chemistry and Newton’s theories of colored bodies and fits of easy reflection. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-40507-6. xvii + 400 pp. LCCN QC402 .S53

1993. URL http://www.loc.
gov/catdir/description/
cam031/92014762.html;
http://www.loc.gov/
catdir/enhancements/
fy0907/92014762-t.html.

Shamdasani:1998:CFC

[Sha98] Sonu Shamdasani. Cult fictions: C. G. Jung and the founding of analytical psychology. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1998. ISBN 0-415-18613-7 (hardcover), 0-415-18614-5 (paperback). x + 121 pp. LCCN BF173 .S486

1998. URL http://www.loc.
gov/catdir/enhancements/
fy0649/97032830-d.html.

Sheppard:1963:BRS

[She63] P. A. Sheppard. Book review: Sir George Cayley’s Aeronautics, 1796–1855. By C. H. Gibbs Smith. Pp. xxiii + 269. Her Majesty’s Stationery Office, for Science Museum. 1963. 30s. British Journal for the History of Science, 1 (3):286–287, June 1963. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024928.

Shea:1970:GCF

[She70] W. R. J. Shea. Galileo’s claim to fame: The proof that the Earth moves from the evidence of the tides. British Journal for the History of Science, 5(2): 111–127, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025413.

Shea:1985:BRP

[She85] William R. Shea. Book review: Pietro Redondi, Galileo Eretico. Turin: Einaudi, 1983. Pp. x + 463. ISBN 88-06-05632-8. 25,000 lire. British Journal for the History of Science, 18 (1):100–102, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026283.

Sheail:1986:GME

[She86] John Sheail. Grassland management and the early development of British ecology. British Journal for the History of Science, 19(3):283– 299, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026506.

Shea:1991:MNM

[She91] William R. Shea. The magic of numbers and motion: the scientific career of René Descartes. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1991. ISBN 0-88135-098-2. x + 371 pp. LCCN Q125.2 .S54 1990.

Sheehan:1995:IFW

[She95] William Sheehan. The immortal fire within: the life and work of Edward Emerson Barnard. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-44489-6 (hardcover). xiv + 429 pp. LCCN QB36.B3 S48

                                        1995. URL http://www.loc.
                                        gov/catdir/description/
                                        cam026/94026742.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam021/
                                        94026742.html.

Shirley:1983:THB

[Shi83] John William Shirley. Thomas Harriot, a biography. Clarendon Press, New York, NY, USA, 1983. ISBN 0-19-822901-1. xii + 508 pp. LCCN Q143.H36 S46 1983. US$59.00.

Shortland:1986:BRJ

[Sho86a] Michael Shortland. Book review: Janet Oppenheim. The Other World: Spiritualism and Psychical Research in England, 1850–1914. Cambridge: Cambridge University Press, 1985. Pp. xii + 503. ISBN 0-521-26505-3. £25, $44.50. British Journal for the History of Science, 19(2):219–221, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026605.

Shortt:1986:VLR

[Sho86b] S. E. D. (Samuel Edward Dole)

Shortt. Victorian lunacy: Richard M. Bucke and the practice of late Nineteenth-Century psychiatry. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-30999-9. xvi + 207 pp. LCCN RC448.O53 S56

1986. URL http://www.loc.
gov/catdir/description/
cam032/85026919.html;
http://www.loc.gov/
catdir/enhancements/
fy1009/85026919-t.html.

Short:1987:BRJ

[Sho87a] Graham Hollister Short. Book review: Joseph Needham & Francesca Bray. Science & Civilization in China, Vol. 6, Part 2. Cambridge: Cambridge University Press, 1984. xxvii + 724. ISBN 0-521-25076-5. £50.00. British Journal for the History of Science, 20(1): 101–102, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026455.

Shortland:1987:CCE

[Sho87b] Michael Shortland. Courting the cerebellum: Early organological and phrenological views of sexuality. British Journal for the History of Science, 20(2):173–199, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026306.

Shortland:1987:SMT

[Sho87c] Michael Shortland. Screen memories: Towards a history of psychiatry and psychoanalysis in the movies. British Journal for the History of Science, 20(4):421–452, October 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026417.

Shortland:1988:BRH

[Sho88] Michael Shortland. Book review: H. Tristram Engelhardt Jr., and Arthur L. Caplan (eds). Scientific Controversies: Case Studies in the Resolution and Closure of Disputes in Science and Technology. Cambridge: Cambridge University Press, 1987. Pp. x + 639. ISBN 0-521-25565-1. £35.00, $59.50 (hardcover), £15.00 (paperback). British Journal for the History of Science, 21(2): 264–267, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026995.

Shortland:1989:BRP

[Sho89] Michael Shortland. Book review: Pierre Thuillier. D’Archimède à Einstein: les Faces Cachées de l’Invention Scientifique. Paris: Fayard, 1988. Pp. xxvii + 395. FF 160.00. Pierre Thuillier. Les Passions du Savoir: Essais sur les Dimensions Culturelles de la Science. Paris: Fayard, 1988. Pp. xii + 267. FF 145.00. British Journal for the History of Science, 22 (1):102, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026702.

Shortland:1993:SN

[Sho93a] Michael Shortland, editor. Science and nature. British Society for the History of Science, Chalfont St. Giles, Bucks, England, 1993. ISBN 0-906450-08-X (paperback). viii + 291 pp. LCCN ????

Shortland:1993:SNE

[Sho93b] Michael Shortland, editor. Science and nature: essays in the history of the environmental sciences, volume 8 of BSHS monographs. British Society for the History of Science, Chalfont St. Giles, Bucks, England, 1993. ISBN 0-906450-08-X. viii + 291 pp. LCCN GF13 .S33

1993. URL http://catalog.
hathitrust.org/api/
volumes/oclc/28748717.
html.

Shortland:1996:HMC

[Sho96] Michael Shortland, editor. Hugh Miller and the controversies of Victorian science. Clarendon Press, New York, NY, USA, 1996. ISBN 0-19-854053-1. viii + 401 pp. LCCN Q143.M645 A3

1996. URL http://www.loc.
gov/catdir/enhancements/
fy0637/95041168-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0637/95041168-t.html.

Shteir:1996:CWC

[Sht96]  Ann B. Shteir.  Cultivat-

ing women, cultivating science: Flora’s daughters and botany in England, 1760–1860. The Johns Hopkins University Press, Baltimore, MD, USA, 1996. ISBN 0-8018-5141-6 (hardcover). xi + 301 pp. LCCN QK21.G7 S57

1996. URL http://www.loc.
gov/catdir/bios/jhu051/
95012736.html;
http://www.loc.gov/
catdir/description/
jhu051/95012736.html.

Shuttleworth:1996:CBV

[Shu96] Sally Shuttleworth. Charlotte Brontë and Victorian psychology, volume 7 of Cambridge studies in nineteenth-century literature and culture. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55149-8 (hardcover). xiv + 289 pp. LCCN PR4169 .S48

1996. URL http://www.loc.
gov/catdir/description/
cam026/95004681.html;
http://www.loc.gov/
catdir/toc/cam023/
95004681.html.

Shuttleworth:2004:CBV

[Shu04] Sally Shuttleworth. Charlotte Brontë and Victorian psychology, volume 7 of Cambridge studies in nineteenth-century literature and culture. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-61717-0 (paperback), 0-521-55149-8. xiv + 289 pp. LCCN PR4169 .S48

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0632/2005278193-d.html;
 http://www.loc.gov/
catdir/enhancements/
fy0632/2005278193-t.html.

Siegel:1991:IME

[Sie91]  Daniel M. Siegel.  Inno-

vation in Maxwell’s electromagnetic theory: molecular vortices, displacement current, and light. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-35365-3. x + 225 pp. LCCN QC670 .S48

                                        1991. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/90042511.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam029/
                                        90042511.html.

Simms:1984:BRA

[Sim84] D. L. Simms. Book review: Antiquity to the Renaissance Ivo Schneider, ‘Archimedes’ Erträge der Forschung, Wissenschaftliche Buchgesellschaft. Darmstadt, 1979. British Journal for the History of Science, 17(1):101–102, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026037.

Simms:1988:AWW

[Sim88a] D. L. Simms. Archimedes’ weapons of war and Leonardo. British Journal for the History of Science, 21(2):195– 210, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026979.

Simonton:1988:SGP

[Sim88b] Dean Keith Simonton. Scientific genius: a psychology of science. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-35287-8. x + 229 pp. LCCN Q172.5.C74 S57

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/87027823.html;

http://www.loc.gov/
catdir/toc/cam031/
87027823.html.

Sinclair:1974:BRN

[Sin74] S. B. Sinclair. Book review:

Nineteenth and Twentieth Centuries: Molecular Reality. A Perspective on the Scientific Work of Jean Perrin. By Mary Jo Nye. New York: Elsevier, and London: Macdonald, 1972. Pp. xii + 201. £5. British Journal for the History of Science, 7(3):300–301, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025548.

Sinclair:1976:BRN

[Sin76] S. B. Sinclair. Book review:

Nineteenth and Twentieth Centuries: H. G. J. Moseley. The Life and Letters of an English Physicist, 1887–1915. By J. L. Heilbron. Berkeley, Los Angeles, and London: University of California Press, 1974. Pp. xiv + 312. $15.00. British Journal for the History of Science, 9(3):334–335, November 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025696.

Sinclair:1977:BRN

[Sin77]  S. B. Sinclair.  Book re-

view: Nineteenth and Twentieth Centuries: Radioactivity and Atomic Theory. Presenting facsimile reproduction of the Annual Reports on Radioactivity 1904–1920 to the Chemical Society. By Frederick Soddy, F.R.S. Ed. with commentary by Thaddeus J. Trenn. London: Taylor & Francis, 1975. Pp. xv + 517. £12.00. British Journal for the History of Science, 10(2):182, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025886.

Sinclair:1978:BRN

[Sin78]  S. B. Sinclair.  Book re-

view: Nineteenth and Twentieth Centuries: Early History of Planck’s Radiation Law. By Hans Kangro. London: Taylor and Francis, 1976. Pp. xviii + 282. £25.00. British Journal for the History of Science, 11(1):88, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025627.

Sinclair:1981:BRR

[Sin81] S. B. Sinclair. Book review: Rutherford and Physics at the Turn of the Century. Edited by Mario Bunge and William R. Shea. New York: Science History Publications; Folkestone: Dawson, 1979. Pp. viii + 184.£12.50. British Journal for the History of Science, 14 (1):96–97, March 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026083.

Sinclair:1984:BRS

[Sin84] Steve Sinclair. Book review:

Science Studies Alex Keller, The Infancy of Atomic Physics. Hercules in his Cradle. Oxford: Oxford University Press, 1983. Pp. 230. £12.50. ISBN 0-19-853904-5. British Journal for the History of Science, 17 (3):335, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026651.

Sinaceur:1991:CME

[Sin91] Hourya Sinaceur. Corps et modèles: essai sur l’histoire de l’algèbre réelle. (French) [Bodies and models: essay on the history of real algebra]. Mathesis. J. Vrin, Paris, France, 1991. ISBN 2-7116-1038-1. 496 pp. LCCN QA155 .S56 1991. 198F.

Siraisi:1987:ARI

[Sir87] Nancy G. Siraisi. Avicenna in Renaissance Italy: the Canon and medical teaching in Italian universities after 1500. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-05137-2. xii + 410 + 8 pp. LCCN R791.A6 S57 1987. US$65.00.

Siraisi:1990:MER

[Sir90] Nancy G. Siraisi. Medieval and early Renaissance medicine: an introduction to knowledge and practice. University of Chicago Press, Chicago, IL, USA, 1990. ISBN 0-226-76129-0, 0-226-76130-4 (paperback). xiv + 250 pp. LCCN R141 .S546

1990. URL http://www.loc.
gov/catdir/description/
uchi052/89020368.html;
http://www.loc.gov/
catdir/toc/uchi051/
89020368.html.

Siraisi:1997:CMG

[Sir97] Nancy G. Siraisi. The clock and the mirror: Girolamo Cardano and Renaissance medicine. Princeton University Press, Princeton, NJ, USA, 1997. ISBN 0-691-01189-3 (hardcover). xiv + 361 pp. LCCN R520.C32 S57

1997. URL http://www.loc.
gov/catdir/description/
prin021/96046354.html;
http://www.loc.gov/
catdir/toc/prin031/
96046354.html.

Six:1987:DDN

[Six87] Jules Six. La découverte du neutron. (French) [The discovery of the neutron]. Editions du Centre national de la recherche scientifique, Paris, France, 1987. ISBN 2-222-04052-3. vii + 174 pp. LCCN MLCM 90/06183 (Q). 110F.

Simpson:1987:SCL

[SL87] George Gaylord Simpson and Leo F. Laporte. Simple curiosity; letters from George Gaylord Simpson to his family, 1921– 1970. University of California Press, Berkeley, CA, USA, 1987. ISBN 0-520-05792-9. x + 340 pp. LCCN QE707.S55 A4 1987.

Slaughter:1982:ULS

[Sla82] M. M. (Mary M.) Slaughter.

Universal languages and scientific taxonomy in the Seventeenth Century. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24477-3. x + 277 pp. LCCN PM8009 .S77

1982. URL http://www.loc.
gov/catdir/enhancements/
fy1002/81020147-d.html;
http://www.loc.gov/
catdir/enhancements/
fy1002/81020147-t.html.

Sloan:1979:BGB

[Slo79] Phillip R. Sloan. Buffon, German biology, and the historical interpretation of biological species. British Journal for the History of Science, 12 (2):109–153, July 1979. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025750.

Sloan:1980:OP

[Slo80] Phillip R. Sloan. Ontogeny and phylogeny. British Journal for the History of Science, 13(1):50–55, March 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026120.

Sloane:1985:DME

[Slo85] Thomas O. Sloane. Donne, Milton, and the end of humanist rhetoric. University of California Press, Berkeley, CA, USA, 1985. ISBN 0-520-05212-9, 0-520-05213-7. xv + 332 pp. LCCN PR545.H86 S58 1985.

Sanchez:1988:NKQ

[SLT88] Francisco Sánchez, Elaine Limbrick, and D. F. S. (Douglas Ferguson Scott) Thomson. That nothing is known = (Quod nihil scitur). Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-35077-8. x + 310 pp. LCCN B785.S23 Q6813

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/87033395.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        87033395.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam031/87033395.html.

Smith:1994:DTD

[SM94] Merritt Roe Smith and Leo Marx, editors. Does technology drive history?: the dilemma of technological determinism. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-19347-7, 0-262-69167-1 (paperback). xv + 280 pp. LCCN T14.5 .D64 1994.

Smeaton:1966:BRH

[Sme66] W. A. Smeaton. Book review: History of Chemistry: Chemistry and Beyond. A selection from the writings of the late Professor F. A. Paneth. Edited by Herbert Dingle and G. R. Martin, with the assistance of Eva Paneth. Pp. xxi + 285. Frontispiece and 35 figures. New York, London, Sydney: Interscience (John Wiley), 1964. 45s. British Journal for the History of Science, 3(1):88–89, June 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025117.

Smeaton:1968:BRH

[Sme68a] W. A. Smeaton. Book review: History of Chemistry Alfred Werner. Founder of Co-ordination Chemistry. By George B. Kauffman. Berlin, Heidelberg and New York: Springer-Verlag. 1966. Pp. xv + 127. DM. 24. British Journal for the History of Science, 4(2):183, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025156.

Smeur:1968:BRS

[Sme68b] A. J. E. M. Smeur. Book review: Sixteenth and Seventeenth Centuries: Petrus Ramus en de wiskunde. By J. J. Verdenk. Assen, Netherlands. Van Gorcum. 1966. Pp. x + 455. 34.90 Dutch guilders. British Journal for the History of Science, 4(2):188–189, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025161.

Smeaton:1969:BRS

[Sme69] W. A. Smeaton. Book review: Scientific Societies Mémoires de Physique et de Chimie de la Société d’ Arcueil. A Facsimile of the Paris, 1807–1817, Edition. With a new Introduction and an Analytical Table of Contents by Maurice P. Crosland. The Sources of Science, No. 36. New York and London: Johnson Reprint Corporation. 1967. Vol. I, Pp. xlvi + iv + 382; Vol. II, Pp. 498; Vol. III, Pp. 618. £23 16s. British Journal for the History of Science, 4(3): 287–288, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025478.

Smeaton:1970:BRC

[Sme70] W. A. Smeaton. Book review: Chemistry The Periodic System of Chemical Elements: A History of the First Hundred Years. By J. W. van Spronsen. Amsterdam, London and New York: Elsevier. 1969. Pp. xv + 368, with portraits, tables and 139 figures, £6. British Journal for the History of Science, 5(2):194–195, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025426.

Smeaton:1971:BRC

[Sme71] W. A. Smeaton. Book review: Chemistry A Dissertation on Elective Attractions. By Torbern Bergman. Second edition. Introduction by A. M. Duncan. London: F. Cass. 1970. Pp. xl + xv + 383 + (15). 7 folding plates and tables. £7.35. British Journal for the History of Science, 5 (4):406, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025390.

Smeaton:1978:BM

[Sme78a] W. A. Smeaton. Back matter. British Journal for the History of Science, 11(1):101– 102, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025634.

Smeaton:1978:OAA

[Sme78b] W. A. Smeaton. Obituary: Angus Armitage, 19 July 1902–8 February 1976. British Journal for the History of Science, 11 (1):99–100, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025633.

Smeaton:1979:OHD

[Sme79] W. A. Smeaton. Obituary:

Herbert Dingle, 2 August 1890– 4 September 1978. British Journal for the History of Science, 12(2):242–243, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025774.

Smeaton:1984:BRC

[Sme84a] W. A. Smeaton. Book review: Chemistry and Biochemistry Claude-Louis Berthollet, Revue de l’Essai de Statique Chimique, édition critique par Michelle Sadoun-Goupil. Paris: École Polytechnique, 1980. Pp. vii + 204. ISBN 2-7302-0019-3. (Price not stated.). British Journal for the History of Science, 17(2):242, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026573.

Smeaton:1984:OWP

[Sme84b] W. A. Smeaton. Obituary: William Persehouse Delisle Wightman: 4 June 1899–15 January 1983. British Journal for the History of Science, 17 (2):214–216, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026554.

Smeaton:1987:ONH

[Sme87] W. A. Smeaton. Obituary: Niels Hugh de Vaudrey Heathcote: 22 January 1895–23 June 1985. British Journal for the History of Science, 20(3): 350–351, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026363.

Smeaton:1988:ORB

[Sme88]  W. A. Smeaton.  Obitu-

ary: Richard Broke Freeman: 1 April 1915–1 September 1986. British Journal for the History of Science, 21(1): 101, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026866.

Smith:1966:BRH

[Smi66] Sydney Smith. Book review:

History of Biological Sciences and Medicine History of Science. An Annual Review of Literature, Research and Teaching. Volume 4. 1965. Edited by A. C. Crombie and M. A. Hoskin. Pp. 155. Cambridge: W. Heffer. Cloth, 30s. Paper bound, 21s. British Journal for the History of Science, 3 (1):91–92, June 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025120.

Smith:1971:BRI

[Smi71] J. Graham Smith. Book review: Industrial Revolution Science and Technology in the Industrial Revolution. By A. E. Musson and Eric Robinson. Manchester: Manchester University Press. 1969. Pp. viii + 534. £2.50. British Journal for the History of Science, 5 (3):296–297, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025324.

Smith:1972:ARW

[Smi72] Roger Smith. Alfred Russel Wallace: Philosophy of nature and man. British Journal for the History of Science, 6(2): 177–199, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025290.

Smith:1973:BRE

[Smi73a]  Roger Smith.  Book re-

view: Eighteenth and Nineteenth Centuries: Biology in the Nineteenth Century: Problems of Form, Function, and Transformation. By William Coleman. New York: John Wiley, 1971. Pp. x + 187. £3.50; £1.75 (paperback). British Journal for the History of Science, 6(3):330–331, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025463.

Smith:1973:BRN

[Smi73b] Roger Smith. Book review: Nineteenth Century: Wallace and Natural Selection. By H. Lewis McKinney. New Haven and London: Yale University Press, 1972. Pp. xx + 193.

£4.95. British Journal for the History of Science, 6(4):453– 455, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025524.

Smith:1974:BRS

[Smi74a] Alan G. R. Smith. Book review: Sixteenth and Seventeenth Centuries: Humanism and the Rise of Science in Tudor England. By Antonia McLean. London: Heinemann, 1972. Pp. 258. £3.75. British Journal for the History of Science, 7(1):88–89, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025185.

Smith:1974:BRG

[Smi74b] J. Graham Smith. Book review: General Perspectives in the History of Science and Technology. Ed. by Duane H. D. Roller. Norman: University of Oklahoma Press, 1972. Pp. x + 307. £5.00. British Journal for the History of Science, 7(2):182–183, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025207.

Smith:1974:DHC

[Smi74c] Roger Smith. Darwin and his critics. The reception of Darwin’s Theory of Evolution by the scientific community. British Journal for the History of Science, 7(3):278– 285, November 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025534.

Smith:1975:BRT

[Smi75a] N. A. F. Smith. Book review: Technology Roman and Islamic Water-Lifting Wheels. By Thorkild Schiøler. Odense: Odense University Press, 1973. Pp. 201. No price stated. British Journal for the History of Science, 8(2):179, July 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025647.

Smith:1975:BRM

[Smi75b] Roger Smith. Book review: Medicine An Illustrated History of Brain Function. By Edwin Clarke and Kenneth Dewhurst. Oxford, Sandford Publications, 1972. Pp. 154. £5.50. British Journal for the History of Science, 8(1): 74, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025820.

Smith:1976:BRN

[Smi76a] Crosbie Smith. Book review: Nineteenth and Twentieth Centuries: Lord Kelvin and the Age of the Earth. By Joe D. Burchfield. London: Macmillan, 1975. Pp. xii + 260. £10.00. British Journal for the History of Science, 9(1): 82–83, March 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025717.

Smith:1976:NPT

[Smi76b] Crosbie Smith. Natural philosophy and thermodynamics: William Thomson and “The dynamical theory of heat”. British Journal for the History of Science, 9(3):293–319, November 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025684.

Smith:1976:BRL

[Smi76c] Roger Smith. Book review:

Life Sciences: Studies in the Philosophy of Biology: Reduction and Related Problems. Ed. by Francisco Jose Ayala and Theodosius Dobzhansky. London: Macmillan, 1972. Pp. xix + 390. £12.00. British Journal for the History of Science, 9(3):333–334, November 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025695.

Smith:1978:BRNa

[Smi78a]  Crosbie Smith.  Book re-

view: Nineteenth and Twentieth Centuries: Catalogue of the Manuscript Collections of Sir George Gabriel Stokes and Sir William Thomson, Baron Kelvin of Largs, in Cambridge University Library. Compiled by David B. Wilson. Cambridge: Cambridge University Library, 1976. 2 vols. Pp. iii + 589; iii + 363. £14.00. British Journal for the History of Science, 11(1):85–86, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025625.

Smith:1978:BRA

[Smi78b] Norman A. F. Smith. Book review: Antiquity Gears from the Greeks: The Antikythera Mechanism — a Calendar Computer from ca. 80 B.C. by Derek de Solla Price. New York: Science History Publications, 1975. Pp.

70. $8.50. British Journal for the History of Science, 11 (1):77–78, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025619.

Smith:1978:BRNb

[Smi78c] Sydney Smith. Book review: Nineteenth and Twentieth Centuries: A Century of Zoology at the British Museum through the lives of two Keepers, 1815– 1914. By Albert E. Gunther. Folkestone: Dawson’s, 1975. Pp. 533. £17.50. British Journal for the History of Science, 11(1):86–87, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025626.

Smith:1979:DDT

[Smi79] Crosbie Smith. From design to dissolution: Thomas Chalmers’ debt to John Robison. British Journal for the History of Science, 12(1):59–70, March 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025663.

Smith:1980:BRN

[Smi80a]  Crosbie Smith.  Book re-

view: Nineteenth and Twentieth Centuries: The Temperature of History: Phases of Science and Culture in the Nineteenth Century. By Stephen G. Brush. New York: Burt Franklin, 1978. Pp. xii + 210. $18.95. British Journal for the History of Science, 13(3): 272–273, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026208.

Smith:1980:BRT

[Smi80b] Roger Smith. Book review:

Twentieth Century Freud in Germany: Revolution and Reaction in Science, 1893–1907. By Hannah S. Decker. Psychological Issues, vol. xi, No 1, Monograph 41. New York: International Universities Press, 1977. Pp. xi + 360. No price stated. British Journal for the History of Science, 13(2): 177–179, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025909;
                                        http://www.jstor.org/
                                        stable/4025910.

Smith:1982:BRN

[Smi82a] Crosbie Smith. Book review: The New Physics Enrico Bellone, A world on paper: studies on the second scientific revolution, trans. by Mirella and Riccardo Giacconi. Cambridge, Mass., & London: MIT Press, 1980. Pp. xiii + 220. £9.30. British Journal for the History of Science, 15(2): 197–199, July 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025971.

Smith:1982:EUA

[Smi82b] Robert W. (Robert William) Smith. The expanding universe: astronomy’s “Great Debate,” 1900–1931. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-23212-0 (US). xiv + 220 pp. LCCN QB981 .S69 1982.

Smith:1982:WWR

[Smi82c] Roger Smith. Wilhelm Wundt resurrected. British Journal for the History of Science, 15(3): 285–293, November 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026222.

Smith:1983:BRE

[Smi83] Crosbie Smith. Book review:

Essay Collections G. N. Cantor and M. J. S. Hodge (eds), Conceptions of ether: studies in the history of ether theories, 1740–1900. Cambridge: Cambridge University Press, 1981. Pp. x + 351. £30.00. British Journal for the History of Science, 16(2):196–198, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026241.

Smith:1985:BRG

[Smi85a] Crosbie Smith. Book review:

G. N. Cantor. Optics after Newton. Theories of light in Britain and Ireland, 1704–1840. Manchester: Manchester University Press, 1983. Pp. ix + 257. ISBN 0-7190-0938-3. £20. British Journal for the History of Science, 18(1):105– 107, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026288.

Smith:1985:BRMa

[Smi85b] Crosbie Smith. Book review:

Martin Goldman, The demon in the aether. The story of James Clerk Maxwell. Edinburgh: Paul Harris Publishing, 1983. Pp. 224. ISBN 0-86228-026-5. £18. British Journal for the History of Science, 18 (1):120–121, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026300.

Smith:1985:BRMb

[Smi85c] Denis Smith. Book review: T. M. Charlton, A History of Theory of Structures in the Nineteenth Century. Cambridge: Cambridge University Press, 1982. Pp. viii + 194. ISBN 0-521-23419-0. £22.50. British Journal for the History of Science, 18(2):235, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026334.

Smith:1985:BRMc

[Smi85d]  Roger Smith.  Book re-

view: Molly Harrower, Kurt Koffka: an unwitting selfportrait. Gainsville: University Presses of Florida, 1983. Pp. xvi + 334.ISBN 0-8130-0760-7. $30. British Journal for the History of Science, 18 (2):246, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026346.

Smith:1987:BRS

[Smi87a] Crosbie Smith. Book review: Sadi Carnot. Reflexions on the Motive Power of Fire. A Critical Edition with the Surviving Manuscripts. Translated and edited by Robert Fox. Manchester and New York: Manchester University Press and Lilian Barber Press, Inc., 1986. Pp. ix + 230. ISBN 0-7190-1741-6. £25.00. British Journal for the History of Science, 20 (2):228–229, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026314.

Smith:1987:BRC

[Smi87b] Norman A. F. Smith. Book review: Clive Hart. The Prehistory of Flight. Berkeley: University of California Press, 1985. Pp. xvii + 279. ISBN 0-520-05213-7. £29.75. British Journal for the History of Science, 20(1): 113–114, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026466.

Smith:1987:BRJ

[Smi87c] Robert W. Smith. Book review:

J. B. Hearnshaw. The Analysis of Starlight: One Hundred and Fifty Years of Astronomical Spectroscopy. Cambridge: Cambridge University Press, 1986. Pp. xv + 531. ISBN 0-521-25548-1 £50.00, $79.50. British Journal for the History of Science, 20(4): 482–483, October 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026423.

Smith:1988:RT

[Smi88] F. B. (Francis Barrymore) Smith. The retreat of tuberculosis, 1850–1950. Croom Helm, London, UK, 1988. ISBN 0-7099-3383-5. 271 pp. LCCN RC310.5 .S65 1988. US$25.00 (est.).

Smith:1989:WHS

[Smi89a] Crosbie Smith. William Hopkins and the shaping of Dynamical Geology: 1830–1860. British Journal for the History of Science, 22(1):27–52, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026677.

Smith:1989:BRG

[Smi89b] N. A. F. Smith. Book review: Gunther Garbrecht. Hydraulics and Hydraulic Research. A Historical Review. Rotterdam and Boston: A. A. Balkema, 1987. Pp. ix + 362. ISBN 90-6191-621-6. £53.50. British Journal for the History of Science, 22 (1):116–117, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026713.

Smith:1989:BRR

[Smi89c] Robert W. Smith. Book reviews: R. J. Tayler (ed.). History of the Royal Astronomical Society. Volume 2. 1920–1980. Oxford: Published for the Society by Blackwell Scientific Publications. ISBN 0-632-01792-9 (paperback), £14.50; ISBN 0-632-01791-0 (hardcover), £29.50. British Journal for the History of Science, 22(1):88–89, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026688.

Smith:1989:STS

[Smi89d] Robert W. (Robert William)

Smith. The space telescope: a study of NASA, science, technology, and politics. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-26634-3. xviii + 478 pp. LCCN QB500.267 .S55

1989. URL http://www.loc.
gov/catdir/enhancements/
fy0803/89000707-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0803/89000707-t.html.

Smith:1989:BRD

[Smi89e] Roger Smith. Book review:

David Kohn (ed.). The Darwinian Heritage. Princeton: Princeton University Press in association with Nova Pacifica, 1985. Pp. xii + 1138. ISBN 0-691-08356-8. £67.90 (hardcover), £15.50 (paperback). British Journal for the History of Science, 22 (4):452–456, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026931.

Smith:1992:IHM

[Smi92] Roger Smith. Inhibition: history and meaning in the sciences of mind and brain. Free Association Books, London, UK, 1992. ISBN 1-85343-181-8. xi + 333 pp. LCCN QP353 .S65 1992b.

Smith:1994:FFB

[Smi94a] Jonathan Smith. Fact and feeling: Baconian science and the Nineteenth-Century literary imagination. Science and literature. University of Wisconsin Press, Madison, WI, USA, 1994. ISBN 0-299-14350-3, 0-299-14354-6 (paperback). ix + 277 pp. LCCN PR468.S34 S65 1994.

Smith:1994:BAS

[Smi94b] Pamela H. Smith. The business of alchemy: science and culture in the Holy Roman Empire. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-05691-9. xii + 308 pp. LCCN Q125.2 .S58 1994. US$45.00. URL http:

                                        //www.loc.gov/catdir/
                                        description/prin031/
                                        93044856.html;   http://
                                        www.loc.gov/catdir/toc/
                                        prin031/93044856.html.

Smith:1994:SFS

[Smi94c] Tim D. (Tim Denis) Smith. Scaling fisheries: the science of measuring the effects of fishing, 1855–1955. Cambridge studies in applied ecology and resource management. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-39032-X. x + 392 pp. LCCN SH329.F56 S58

1994. URL http://www.loc.
gov/catdir/description/
cam026/93038292.html;
http://www.loc.gov/
catdir/toc/cam029/
93038292.html.

Smith:1997:FHH

[Smi97] Roger Smith. The Fontana history of the human sciences. Fontana history of science. Fontana Press, London, UK, 1997. ISBN 0-00-686178-4. xviii + 1036 + 8 pp. LCCN H51 .S64 1997.

URL http://books.google.
com/books?id=4395AAAAIAAJ.

Smoller:1994:HPS

[Smo94] Laura Ackerman Smoller. History, prophecy, and the stars: the Christian astrology of Pierre d’Ailly, 1350–1420. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-08788-1. xii + 233 pp. LCCN BF1679.8.A453 S56 1994. US$35.00. URL http:

//www.loc.gov/catdir/
description/prin021/
93044907.html;   http://
www.loc.gov/catdir/toc/
prin031/93044907.html.

Smyth:1989:AEM

[Smy89] Henry De Wolf Smyth. Atomic energy for military purposes: the official report on the development of the atomic bomb under the auspices of the United States Government, 1940–1945. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1721-4 (hardcover), 0-8047-1722-2 (paperback). xv + 324 pp. LCCN QC773.A1 S69 1989.

Sharov:1993:EHD

[SN93]  A.  S.

(Aleksandr Sergeevich) Sharov and I. D. (Igor Dmitrievich) Novikov. Edwin Hubble, the discoverer of the big bang universe. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-41617-5. xv + 187 pp. LCCN QB36.H83 S5313

1993. URL http://www.loc.
gov/catdir/description/
cam025/92024683.html;
http://www.loc.gov/
catdir/toc/cam028/
92024683.html.

Snelders:1974:BRE

[Sne74] H. A. M. Snelders. Book review: Eighteenth and Nineteenth Centuries: Goethe and the Scientific Tradition. By H. B. Nisbet. London: Institute of Germanic Studies, University of London, 1972. Pp. xii + 83. No price stated. British Journal for the History of Science, 7(2):194–195, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025219.

Snelders:1987:BRH

[Sne87] H. A. M. Snelders. Book review: History of Science Today, 2.: History of Science in the Netherlands. British Journal for the History of Science, 20(3):343–348, July 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026361.

Scharlau:1985:FML

[SO85] Winfried Scharlau and Hans Opolka. From Fermat to Minkowski: lectures on the theory of numbers and its historical development. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-90942-7. xi + 184 pp. LCCN QA241 .S2813 1985.

Soderqvist:1986:EMN

[Söd86] Thomas Söderqvist. The ecologists: from merry naturalists to saviours of the nation: a sociologically informed narrative survey of the ecologization of Sweden 1895–1975. Almqvist and Wiksell International, Stockholm, Sweden, 1986. ISBN 91-22-00827-6 (paperback), 91-22-00812-8. vii + 330 pp. LCCN QH541 .S62x

Sokal:1987:PTA

[Sok87] Michael M. (Michael Mark) Sokal, editor. Psychological testing and American society, 1890–1930. Rutgers University Press, New Brunswick, 1987. ISBN 0-8135-1193-3. LCCN BF176 .P78 1987. US$28.00. URL http:

                                        //catalog.hathitrust.org/
                                        api/volumes/oclc/
                                        13580898.html.

Soloman:1983:BRT

[Sol83]  Joan Soloman.  Book re-

view: Technology and Instruments Mary Tasker, Teaching the history of technology. London: The Historical Association. 1980. Pp. 47. £1.40 (70p to members). British Journal for the History of Science, 16 (2):207–208, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026248.

Sorabji:1983:TCC

[Sor83] Richard Sorabji. Time, creation and the continuum: theories in antiquity and the early Middle Ages. Duckworth, London, UK, 1983. ISBN 0-7156-1693-5. xviii + 473 pp. LCCN BD638 .S67 1983. US$29.50.

Sorell:1986:H

[Sor86] Tom Sorell. Hobbes. The Arguments of the philosophers. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1986. ISBN 0-7100-9845-6. xii + 163 pp. LCCN B1247 .S63 1986. US$25.00.

Sorell:1987:D

[Sor87] Tom Sorell. Descartes. Past masters; Oxford paperbacks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1987. ISBN 0-19-287635-X (paperback), 0-19-287636-8 (hardcover). 120 pp. LCCN B1875 .S67 1987. US$4.50; US$15.00. URL http://www.

loc.gov/catdir/
enhancements/fy0637/
86023662-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0724/86023662-b.html.

Southgate:1993:CTL

[Sou93] Beverley C. Southgate. “Covetous of truth”: the life and work of Thomas White, 1593– 1676, volume 134 of Archives internationales d’histoire des idées. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 0-7923-1926-5 (hardcover). xi + 189 pp. LCCN B1299.W454 S68

1993. URL http://www.loc.
gov/catdir/enhancements/
fy0821/92026439-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0821/92026439-t.html.

Sheets-Pyenson:1988:CSD

[SP88] Susan Sheets-Pyenson. Cathedrals of science: the development of colonial natural history museums during the late Nineteenth Century. McGill-Queen’s University Press, Kingston, ON, Canada, 1988. ISBN 0-7735-0655-1. xii + 144 + 20 pp. LCCN QH70.A1 S54 1988. US$24.95.

Sparks:1965:BRH

[Spa65] W. R. Sparks. Book review: The History of the Barometer. By W. E. Knowles Middleton. Pp. xx + 489. Johns Hopkins University Press; London: Oxford University Press, 1964. British Journal for the History of Science, 2(3):265– 266, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4024946.

Speakman:1982:ASG

[Spe82] Colin Speakman. Adam Sedgwick, geologist and dalesman, 1785–1873: a biography in twelve themes. Broad Oak Press, Broad Oak, Heathfield, East Sussex, UK, 1982. ISBN 0-906716-01-2. 145 pp. LCCN QE22.S5 S64 1982. US$5.75.

Springford:1997:ECV

[Spr97] Michael Springford, editor. Electron: a centenary volume. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56130-2 (hardcover). xii + 330 pp. LCCN QC793.5.E62 E39

                                        1997. URL http://www.loc.

gov/catdir/description/
cam028/96044744.html;
http://www.loc.gov/
catdir/toc/cam021/
96044744.html.

Shapin:1985:LAP

[SSH85] Steven Shapin, Simon Schaffer, and Thomas Hobbes. Leviathan and the air-pump: Hobbes, Boyle, and the experimental life: including a translation of Thomas Hobbes, Dialogus physicus de natura aeris by Simon Schaffer. Princeton University Press, Princeton, NJ, USA, 1985. ISBN 0-691-08393-2. xiv + 440 pp. LCCN QC166 .S47 1985. US$59.50.

Schmitt:1988:CHR

[SSK88] Charles B. Schmitt, Quentin Skinner, and Eckhard Kessler, editors. The Cambridge history of Renaissance philosophy. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-25104-4. xiii + 968 pp. LCCN B775 .C25

1988. URL http://www.loc.
gov/catdir/description/
cam023/87005212.html;
http://www.loc.gov/
catdir/toc/cam023/
87005212.html.

Stafford:1984:VSA

[Sta84a] Barbara Maria Stafford. Voyage into substance: art, science, nature, and the illustrated travel account, 1760– 1840. MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-19223-3. xxiii + 645 pp. LCCN G468 .S73 1984.

Stansfield:1984:TBM

[Sta84b] Dorothy A. Stansfield. Thomas Beddoes, M.D., 1760–1808: chemist, physician, democrat. Chemists and chemistry. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1686-2. xix + 306 pp. LCCN QD22.B263 S83 1984.

Staudenmaier:1985:TSR

[Sta85] John M. Staudenmaier. Technology’s storytellers: reweaving the human fabric. Society for the History of Technology and the MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-19237-3. xxvii + 282 pp. LCCN T20 .S73 1985.

Stachel:1987:MMT

[Sta87] John Stachel. “A man of my type”—editing the Einstein papers. British Journal for the History of Science, 20(1):57– 66, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026436. Conference

on documentary editing (Tarrytown, N.Y., 1985).

Stafford:1989:SES

[Sta89a] Robert A. Stafford. Scientist of empire: Sir Roderick Murchison, scientific exploration and Victorian imperialism. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-33537-X. xii + 293 pp. LCCN QE22.M8 S82

1989. URL http://www.loc.
gov/catdir/description/
cam023/88034134.html;
http://www.loc.gov/
catdir/samples/cam031/
88034134.html;   http://
www.loc.gov/catdir/toc/
cam027/88034134.html.

Star:1989:RMB

[Sta89b] Susan Leigh Star. Regions of the mind: brain research and the quest for scientific certainty. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1673-0. xiii + 278 pp. LCCN QP385 .S66 1989. US$27.50. URL http:

//www.loc.gov/catdir/
description/cam023/
89004432.html;   http://
www.loc.gov/catdir/toc/
cam025/89004432.html.

Stanton:1990:HEI

[Sta90] Charles M. Stanton. Higher education in Islam: the classical period. Rowman and Littlefield, Lanham, MD, USA, 1990. ISBN 0-8476-7637-4. ???? pp. LCCN BP.

Stafford:1991:BCI

[Sta91] Barbara Maria Stafford. Body criticism: imaging the unseen in Enlightenment art and medicine. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-19304-3. xxi + 587 pp. LCCN BF697.5.B63 S73 1992.

Stafford:1994:ASE

[Sta94] Barbara Maria Stafford. Artful science: enlightenment, entertainment, and the eclipse of visual education. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-19342-6. xxix + 350 pp. LCCN QP495 .S73 1994. US$35.00.

Stearn:1962:BRB

[Ste62a] W. T. Stearn. Book review: Beloved Son Felix: the Journal of Felix Platter, a Medical Student in Montpellier in the Sixteenth Century. Translated and introduced by Scán Jennett, with a Foreword by Jack Lindsay. 8vo. Pp. 158; 6 plates; 20 text figures. London: Frederick Muller. Price 25s. British Journal for the History of Science, 1(2):189–191, December 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025136.

Stearn:1962:BRJ

[Ste62b]  W. T. Stearn.  Book re-

view: James Lee and the Vineyard Nursery, Hammersmith. By Eleanor Joan Willson. 8 vol. Pp. x + 88; portrait, map. Hammersmith Local History Group, Central Library, Hammersmith, London, W.6. Price 15s. British Journal for the History of Science, 1 (1):89, June 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025077.

Stearn:1962:BRE

[Ste62c]  W. T. Stearn.  Book re-

view: The Encyclopedia of the Biological Sciences. Edited by Peter Gray. 8vo. Pp. xxii, 1119; numerous text-figures. New York: Reinhold Publishing Corporation; London: Chapman and Hall. Price $20 (£8). British Journal for the History of Science, 1(2):188– 189, December 1962. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025135.

Stearn:1962:ILB

[Ste62d] William T. Stearn. The influence of Leyden on botany in the Seventeenth and Eighteenth Centuries. British Journal for the History of Science, 1(2):137–158, December 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025129.

Stearn:1967:API

[Ste67] W. T. Stearn. The anatomy of plants with an idea of a philosophical history of plants and several other lectures read before the Royal Society. British Journal for the History of Science, 3(3):291–293, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025054.

Stearn:1970:BRE

[Ste70] W. T. Stearn. Book review: Eighteenth Century: Daniel Carl Solander, Naturalist on the Endeavour. By R. A. Rauschenberg. Transactions of the American Philosophical Society, New Series, Vol. lvii, part 8. Philadelphia: American Philosophical Society. 1968. Pp. 66. $2.50. British Journal for the History of Science, 5(2):205–207, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025436.

Steneck:1975:BRM

[Ste75] Nicholas H. Steneck. Book review: Middle Ages The Book of Secrets of Albertus Magnus. Ed. by Michael R. Best and Frank H. Brightman. Oxford: Clarendon Press: Oxford University Press, 1973. Pp. xlviii + 128. £3.00. British Journal for the History of Science, 8(2):181–183, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025651.

Steneck:1981:BRC

[Ste81] Nicholas H. Steneck. ‘The Ballad of Robert Crosse and Joseph Glanvill’ and the background to Plus Ultra. British Journal for the History of Science, 14(1):59–74, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026072.

Stein:1985:ALL

[Ste85] Dorothy Stein. Ada, a life and a legacy. MIT Press series in the history of computing. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-19242-X. xix + 321 + 16 pp. LCCN QA29.L72 S74 1985. US$19.95.

Stephenson:1987:KPA

[Ste87] Bruce Stephenson. Kepler’s physical astronomy, volume 13 of Studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 3-540-96541-6 (Berlin) 0-387-96541-6 (New York). v + 216 pp. LCCN ????

Sternberg:1990:LFH

[Ste90a] Charles H. (Charles Hazelius)

Sternberg. The life of a fossil hunter. Indiana University Press, Bloomington, IN, USA, 1990. ISBN 0-253-35549-4, 0-253-20571-9 (paperback). xxvii + 286 + 46 pp. LCCN QE707.S8 A3 1990.

Stewart:1990:DCC

[Ste90b]  John A. Stewart.  Drift-

ing continents and colliding paradigms: perspectives on the geoscience revolution. Science, technology, and society. Indiana University Press, Bloomington, IN, USA, 1990. ISBN 0-253-35405-6. xii + 285 pp. LCCN QE511.5 .S86 1990.

Steinberg:1991:NII

[Ste91a] Theodore Steinberg. Nature incorporated: industrialization and the waters of New England. Studies in environment and history. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-39215-2. xv + 284 pp. LCCN TS1323.15 .S74

1991. URL http://www.loc.
gov/catdir/description/
cam024/90048485.html;
http://www.loc.gov/
catdir/samples/cam031/
90048485.html;   http://
www.loc.gov/catdir/toc/
cam029/90048485.html.

Stemerding:1991:PAF

[Ste91b] Dirk Stemerding. Plants, animals and formulae: natural history in the light of Latour’s Science in action and Foucault’s The order of Things, volume 4 of WMW-publikatie. Faculteit Wijsbegeerte en Maatschappijwetenschappen, Universiteit Twente, Enschede, The Netherlands, 1991. ISBN 90-365-0379-5. ISSN 0925-0131. 202 pp. LCCN QH83 .S7 1991.

URL http://arno.unimaas.
nl/show.cgi?fid=5638.

Stewart:1992:RPS

[Ste92] Larry R. Stewart. The rise of public science: rhetoric, technology, and natural philosophy in Newtonian Britain, 1660– 1750. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41700-7. xxxiv + 453 pp. LCCN Q127.G4 STE 1992; Sci.315; Q; R/STE; Hist.

Stephenson:1994:MHK

[Ste94] Bruce Stephenson. The music of the heavens: Kepler’s harmonic astronomy. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03439-7. xi + 260 pp. LCCN QB361 .S74 1994. US$39.50. URL http://www.

loc.gov/catdir/
description/prin031/
93044916.html;   http://
www.loc.gov/catdir/toc/
prin031/93044916.html.

Stimson:1970:BRA

[Sti70] A. Stimson. Book review: Astronomy and Navigation A History of Nautical Astronomy. By Charles Cotter. London: Hollis & Carter. 1968. Pp. xii + 387. Figs. 12 Plates. 63s. British Journal for the History of Science, 5(1):93–94, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025362.

Stiefel:1984:TCM

[Sti84] Tina Stiefel. Twelfth-century matter for metaphor: the material view of Plato’s timaeus. British Journal for the History of Science, 17(2):169– 185, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026550.

Stigler:1986:HSM

[Sti86] Stephen M. Stigler. The history of statistics: the measurement of uncertainty before 1900. Belknap Press of Harvard University Press, Cambridge, MA, USA and London, UK, 1986. ISBN 0-674-40340-1. xvi + 410 pp. LCCN QA276.15 .S75 1986.

Still:1989:BRM

[Sti89] Arthur Still. Book review: Mitchell G. Ash & William W. Woodward (eds). Psychology in Twentieth-Century Thought and Society. Cambridge: Cambridge University Press, 1987. Pp. ix + 320. ISBN 0-521-32523-4. £30.00, $42.50. British Journal for the History of Science, 22(4):459–460, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026935.

Stoddart:1976:DLG

[Sto76] D. R. Stoddart. Darwin, Lyell, and the geological significance of coral reefs. British Journal for the History of Science, 9 (2):199–218, July 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025807.

Stocking:1983:OOE

[Sto83] George W. Stocking, editor.

Observers observed: essays on ethnographic fieldwork, volume 1 of History of anthropology. University of Wisconsin Press, Madison, WI, USA, 1983. ISBN 0-299-09450-2. vi + 242 pp. LCCN GN346 .O27 1983. US$19.95.

Stocking:1984:FHE

[Sto84] George W. Stocking, editor.

Functionalism historicized: essays on British social anthropology, volume 2 of History of anthropology. University of Wisconsin Press, Madison, WI, USA, 1984. ISBN 0-299-09900-8. vi + 244 pp. LCCN GN363 .F86 1984. US$19.95.

Stokes:1988:DPH

[Sto88] Raymond G. Stokes. Divide and prosper: the heirs of I. G. Farben under Allied authority, 1945–1951. University of California Press, Berkeley, CA, USA, 1988. ISBN 0-520-06248-5. xiv + 290 pp. LCCN HD9654.9.I5 S76 1988.

Stroup:1973:BRS

[Str73] Alice Stroup. Book review: Sixteenth and Seventeenth Centuries: Le Physicien Ignace Gaston Pardies S.J. (1636– 1673). By Ziggelaar S. J. August Copenhagen: Odense University Press, 1971. Pp. 242. Dan. kr. 60.00. British Journal for the History of Science, 6(3):324–325, June 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025458.

Stroup:1975:BRS

[Str75] Alice Stroup. Book review: Seventeenth Century: Critique du ‘De Mundo’ de Thomas White. By Thomas Hobbes. Ed. by Jean Jacquot and Harold Whitmore Jones. Paris: J. Vrin, 1973. Pp. 548. No price stated. British Journal for the History of Science, 8 (1):80–81, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025826.

Stroup:1987:RFP

[Str87] Alice Stroup. Royal funding of the Parisian Académie royale des sciences during the 1690s, volume 77, 4 of Transactions of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1987. ISBN 0-87169-774-2. ISSN 0065-9746. 167 pp. LCCN Q46 .S77 1987.

Sampson:1987:HRA

[STT87] R. A. (Ralph Allen) Sampson, R. J. (Roger John) Tayler, and H. H. (Herbert Hall) Turner, editors. History of the Royal Astronomical Society, Vol. 2, 1920–1980. Published for the Society by Blackwell Scientific Publications, Oxford, UK, 1987. ISBN 0-632-01791-0 (hardback:v. 2), 0-632-01792-9 (paperback:v. 2). vii + 262 + 14 pp. LCCN QB1.R5.

Stuewer:1971:BRP

[Stu71a] Roger H. Stuewer. Book review: Physics and Astronomy The Nature of Light. By Vasco Ronchi. Trans. by V. Barocas. London: Heinemann Educational Books. 1970. Pp. xii + 288. £6. British Journal for the History of Science, 5(4): 402–403, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025386.

Stuewer:1971:WHB

[Stu71b] Roger H. Stuewer. William H. Bragg’s corpuscular theory of X-rays and γ-rays. British Journal for the History of Science, 5(3):258–281, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025316.

Stuewer:1978:BRN

[Stu78] Roger H. Stuewer. Book review: Nineteenth and Twentieth Centuries: The Solvay Conferences on Physics. Aspects of the Development of Physics since 1911. By Jagdish Mehra. Dordrecht and Boston: D. Reidel, 1975. Pp. xxxii + 415. Dfl. 140.00/$56.00. British Journal for the History of Science, 11 (1):91–92, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025629.

Stuewer:1985:BRD

[Stu85] Roger H. Stuewer. Book reviews: David Wilson, Rutherford: Simple Genius. London: Hodder and Stoughton, 1983. Pp. 639. ISBN 0-340-23805-4. £14.95. Guy Hartcup and T. E. Allibone, Cockcroft and the Atom. Bristol: Adam Hilger, 1984. Pp. xii + 320. ISBN 0-85274-759-4. £18.95. John Hendry (ed), Cambridge Physics in the Thirties. Bristol: Adam Hilger, 1984. Pp. xi + 209. ISBN 0-85274-761-6. £17.50. British Journal for the History of Science, 18(3): 357–360, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026397.

Sturdy:1986:PJB

[Stu86] David Sturdy. Pierre-Jean-Baptiste Chomel (1671–1740): A case study in problems relating to the social status of scientists in the Early Modern Period. British Journal for the History of Science, 19(3): 301–322, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026507.

Stuewer:1987:BRD

[Stu87] Roger H. Stuewer. Book review: Dominique Pestre. Physique et Physiciens en France 1918–1940. Paris: Editions des Archives Contemporaines, 1984. Pp. xv + 356. ISBN 2-903928-08-8. FF 295.00. British Journal for the History of Science, 20(1): 83, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026440.

Sturdy:1988:BST

[Stu88] Steve Sturdy. Biology as social theory: John Scott Haldane and physiological regulation. British Journal for the History of Science, 21(3):315– 340, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026832.

Sturdy:1995:SSS

[Stu95] D. J. (David J.) Sturdy. Science and social status: the members of the Academie des sciences 1666–1750. Boydell Press, Woodbridge, Suffolk, UK, 1995. ISBN 0-85115-395-X. xvi + 461 pp. LCCN Q149.F7 S78 1995.

Stokes:1990:CBS

[STW90] George Gabriel Stokes, William Thomson (Baron Kelvin), and David B. Wilson. The correspondence between Sir George Gabriel Stokes and Sir William Thomson, Baron Kelvin of Largs. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-32831-4 (set). lvi + ix + 783 pp. LCCN QC7 .S79

1990. URL http://www.loc.
gov/catdir/description/
cam023/89017463.html;
http://www.loc.gov/
catdir/toc/cam028/
89017463.html.

Styles:1968:BRH

[Sty68] Brian T. Styles. Book review: History of Botany Botanical Latin. By William T. Stearn. Pp. xiv + 556. 41 figs. London: Thomas Nelson & Sons. 1966. 105s. British Journal for the History of Science, 4 (1):90, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025006.

Sullivan:1982:JTD

[Sul82a]  Robert E. Sullivan.  John

Toland and the Deist controversy: a study in adaptations, volume 101 of Harvard historical studies. Harvard University Press, Cambridge, MA, USA, 1982. ISBN 0-674-48050-3. viii + 355 pp. LCCN BL2790.T6 S94.

Sullivan:1982:CRA

[Sul82b] Woodruff Turner Sullivan, editor. Classics in radio astronomy, volume 10 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1982. ISBN 90-277-1356-1, 90-277-1363-4. xxiv + 348 pp. LCCN QB475 .C55 1982.

Sutton:1974:SJH

[Sut74] M. A. Sutton. Sir John Herschel and the development of spectroscopy in Britain. British Journal for the History of Science, 7(1):42–60, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025175.

Sutton:1995:SPS

[Sut95] Geoffrey V. Sutton. Science for a polite society: gender, culture, and the demonstration of enlightenment. Westview Press, Boulder, CO, USA, 1995. ISBN 0-8133-1575-1 (hardcover), 0-8133-1576-X (paperback). xiii + 391 pp. LCCN Q125 .S982

1995. URL http://www.loc.
gov/catdir/enhancements/
fy0832/95020157-b.html;
http://www.loc.gov/
catdir/enhancements/
fy0832/95020157-d.html.

Shortland:1989:THS

[SW89] Michael Shortland and Andrew Warwick, editors. Teaching the history of science. British Society for the History of Science, Chalfont St. Giles, Bucks, England, 1989. ISBN 0-631-16977-6, 0-631-16978-4 (paperback). 281 pp. LCCN Q126.9 .T4 1989. US$59.00 (U.S. : est.), US$19.90 (U.S. : est.).

Swann:1983:SSP

[Swa83] John Patrick Swann. The search for synthetic penicillin during World War II. British Journal for the History of Science, 16(2):154–190, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026237.

Swann:1986:BRT

[Swa86] John Patrick Swann. Book review: Trevor I. Williams. Howard Florey: Penicillin and After. Oxford: Clarendon Press, 1984. Pp. xvi + 404. ISBN 0-19-858173-4. £17.50. British Journal for the History of Science, 19(3):369–370, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026541.

Swann:1988:ASP

[Swa88] John Patrick Swann. Academic scientists and the pharmaceutical industry: cooperative research in Twentieth-Century America. The Johns Hopkins University Press, Baltimore, MD, USA, 1988. ISBN 0-8018-3558-5. xi + 249 pp. LCCN RS122 .S93 1988.

Swerdlow:1976:BRA

[Swe76] N. M. Swerdlow. Book review: Astronomy The Astronomical Tables of Levi ben Gerson. By Bernard R. Goldstein. Transactions of the Connecticut Academy of Arts and Sciences, xlv (1974). Archon Books: Hamden, Connecticut, 1974. Pp. 285. $15.00. British Journal for the History of Science, 9(3):324–325, November 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025686.

Swinburne:1967:BRH

[Swi67] R. G. Swinburne. Book review: History and Philosophy of Science: Science, Folklore, and Philosophy. By Harry Girvetz, George Geiger, Harold Hantz, and Bertram Morris. Pp. x + 547. New York and London: Harper and Row, 1966. 72s. British Journal for the History of Science, 3(4): 399, December 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024968.

Swinbank:1970:BRT

[Swi70] P. Swinbank. Book review:

Technology James Watt and the Separate Condenser. By R. J. Law. Science Museum Monograph. London: H.M.S.O. 1969. Pp. 46. 35 plates. 6s. British Journal for the History of Science, 5(2):202–203, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025433.

Swinbank:1971:BRP

[Swi71] P. Swinbank. Book review: Physics and Astronomy The Collected Works of Count Rumford. Volume III, Devices and Techniques; Volume IV, Light and Armament. Ed. by Sanborn C. Brown. Harvard University Press & Oxford University Press. 1970. Pp. vii + 504. Illus. £4.75 each. British Journal for the History of Science, 5(4):403–404, December 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025387.

Swinbank:1977:BRG

[Swi77] Peter Swinbank. Book review: General Sources for the History of Science, 1660–1914. By David Knight. London: The Sources of History Limited, distributed through Hodder and Stoughton Limited, 1975. Pp. 223. £4.25; £2.25 (paperback). British Journal for the History of Science, 10 (1):67–68, March 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025582.

Schuster:1986:PRS

[SY86] John Andrew Schuster and Richard R. Yeo, editors. The politics and rhetoric of scientific method: historical studies, volume 4 of Australasian studies in history and philosophy of science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1986. ISBN 90-277-2152-1. xxxix + 305 pp. LCCN Q175.3 .P65 1986. US$64.50.

Shortland:1996:TLS

[SY96]  Michael  Shortland

and Richard R. Yeo, editors. Telling lives in science: essays on scientific biography. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-43323-1 (hardcover). xiv + 295 pp. LCCN Q141 .T45

1996. URL http://www.loc.
gov/catdir/description/
cam027/95047399.html;
http://www.loc.gov/
catdir/toc/cam021/
95047399.html.

Thackray:1985:CAH

[T^{+}85] Arnold Thackray et al., editors.

Chemistry in America, 1876– 1976: historical indicators. Chemists and chemistry. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1985. ISBN 90-277-1720-6. xxiii + 564 pp. LCCN QD18.U6 C47 1985.

Takahashi:1992:MLT

[Tak92] Ken’ichi Takahashi. The medieval Latin traditions of Euclid’s Catoptrica: a critical edition of De speculis with an introduction, English translation and commentary. Kyushu University Press, Fukuoka-shi, Japan, 1992. ISBN 4-87378-299-6. viii + 373 pp. LCCN ????

Talbot:1962:SBC

[Tal62] C. H. Talbot. Simon Bredon (c. 1300–1372). Physician, mathematician and astronomer. British Journal for the History of Science, 1(1): 19–30, June 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025072.

Tanner:1962:REI

[Tan62] R. C. H. Tanner. On the role of equality and inequality in the history of mathematics. British Journal for the History of Science, 1(2): 159–169, December 1962. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025130.

Tann:1982:BRT

[Tan82]  Jennifer Tann.  Book re-

view: Technology and Development Anne Van Neck, Les debúts de la machine à vapeur dans l’industrie belge, 1800– 1850. Brussels: Palais des Academics, 1979. Pp. 898. No price. British Journal for the History of Science, 15(1):86– 88, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025935.

Taylor:1987:HRA

[Tay87] R. J. (Roger John) Taylor, editor. History of the Royal Astronomical Society: volume 2, 1920–1980. Reprinted for the Society by Blackwell Scientific Publications, Palo Alto, CA, USA, 1987. ISBN 0-632-01791-0, 0-632-01792-9 (paperback). vii + 262 pp. LCCN QB1.R75 H57 1987.

Taylor:1989:RG

[Tay89] Selwyn Taylor. Robert Graves.

Royal Society of Medicine Services, London, UK and New York, NY, USA, 1989. ISBN 0-905958-98-5 (hardcover), 0-905958-90-X (paperback). 160 pp. LCCN ???? US$15.00; US$7.50.

Taylor:1996:WJC

[Tay96] Eugene Taylor. William James on consciousness beyond the margin. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-01136-2. xiii + 215 pp. LCCN BF109.J28

T38 1996. URL http://www.
loc.gov/catdir/
description/prin021/
96000898.html;   http://
www.loc.gov/catdir/toc/
prin031/96000898.html.

Thomson:2010:CTF

[TBDA10] Ann Thomson, Simon Burrows, Edmond Dziembowski, and Sophie Audidière, editors. Cultural transfers: France and Britain in the long Eighteenth Century, volume 2010:04 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2010. ISBN 0-7294-0993-7 (paperback). xi + 326 pp. LCCN DA47.1 .C86 2010.

Thoren:1990:LUB

[TC90] Victor E. Thoren and J. R. (John Robert) Christianson. The Lord of Uraniborg: a biography of Tycho Brahe. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-35158-8 (hardcover). xi + 523 pp. LCCN QB36.B8 T49

1990. URL http://www.loc.
gov/catdir/description/
cam032/90001477.html;
http://www.loc.gov/
catdir/samples/cam034/
90001477.html;   http://
www.loc.gov/catdir/toc/
cam028/90001477.html.

Tauber:1991:MOI

[TC91] Alfred I. Tauber and Leon Chernyak. Metchnikoff and the origins of immunology: from metaphor to theory. Monographs on the history and philosophy of biology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-506447-X. xviii + 247 pp. LCCN QR180.72.M48 T38

1991. URL http://www.loc.
gov/catdir/enhancements/
fy0603/90026245-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0603/90026245-t.html.

Tee:1984:BRH

[Tee84]  Garry J. Tee.  Book re-

view: History of Mathematical Sciences: Stella Mills (ed.), The collected letters of Colin MacLaurin. Nantwich: Shiva Publishing Limited, 1982. Pp. xix + 496. British Journal for the History of Science, 17 (2):233–234, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026566.

Tee:1988:MPB

[Tee88] Garry J. Tee. Mathematics in the Pacific Basin. British Journal for the History of Science, 21(4):401–417, December 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026960.

Tee:1989:NBA

[Tee89] Garry J. Tee. A note on Bechmann’s approximate construction of π, suggested by a deleted sketch in Villard de Honnecourt’s manuscript. British Journal for the History of Science, 22(2):241– 242, July 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026662.

Teich:1970:PVC

[Tei70] Mikulás Teich. Purkyne and valentin on ciliary motion: An early investigation in morphological physiology. British Journal for the History of Science, 5(2):168–177, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025416.

Teich:1978:BRN

[Tei78]  Mikulás Teich.  Book re-

view: Nineteenth and Twentieth Centuries: Crystals, Fabrics, and Fields: Metaphors of Organicism in Twentieth-Century Developmental Biology. By Donna Jeanne Haraway. New Haven and London: Yale University Press, 1976. Pp. 231. £9.00. British Journal for the History of Science, 11(1):92–94, March 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025630.

Teich:1983:BRLb

[Tei83a]  Mikulás Teich.  Book re-

view: Life and earth sciences: Hans Krebs in collaboration with Roswitha Schmid, Otto Warburg Cell Physiologist Biochemist and Eccentric. Transl. by Hans Krebs and Anne Martin. Oxford: Oxford University Press, 1981. Pp. viii + 141. £10.00. British Journal for the History of Science, 16(3): 281–282, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026151.

Teich:1983:BRLa

[Tei83b] Mikulás Teich. Book review:

Life Sciences: Eduard Glas, Chemistry and physiology in their historical and philosophical relations. Delft: Delft University Press, 1979. Pp xvi + 199. (No price stated.). British Journal for the History of Science, 16(2):218–219, July 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026258.

Teich:1984:BRC

[Tei84] Milulás Teich. Book review:

Chemistry and Biochemistry Robert E. Kohler, From medical chemistry to biochemistry. The making of a biomedical discipline. Cambridge: University Press, 1982. Pp. ix + 399. £22.50. British Journal for the History of Science, 17 (2):239–241, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026572.

Teich:1988:BRK

[Tei88] M. Teich. Book review: Kenneth J. Carpenter. The History of Scurvy and Vitamin C. Cambridge: Cambridge University Press, 1986. Pp. viii + 288. ISBN 0-521-32029-1. £27.50, $39.50. British Journal for the History of Science, 21 (1):128–129, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026884.

Tropp:1993:AFM

[TFC93] E. A. (Eduard Abramovich) Tropp, Viktor Iakovlevich Frenkel, and A. D. (Artur Davidovich) Chernin. Alexander A. Friedmann: the man who made the universe expand. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-38470-2. x + 267 pp. LCCN QC16.F73 T7613

                                        1993. URL http://www.loc.
                                        gov/catdir/description/
                                        cam025/92028315.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        92028315.html.

Thackray:1966:BRJ

[Tha66a] Arnold Thackray. Book review: John Dalton. 1766–1844. A Bibliography of Works by and about him. By A. L. Smyth. Pp. xvi + 114. Manchester University Press, 1966. 42s. British Journal for the History of Science, 3(2):194, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025037.

Thackray:1966:EDC

[Tha66b] Arnold W. Thackray. The emergence of Dalton’s Chemical Atomic Theory: 1801– 1808. British Journal for the History of Science, 3(1): 1–23, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025099.

Thackray:1967:BRJ

[Tha67] Arnold Thackray. Book review: John Dalton and the Atom. By Frank Greenaway. London: Heinemann. Pp. x + 244. 1966. 42s. British Journal for the History of Science, 3(3):300–301, June 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025062.

Thackray:1974:BRG

[Tha74] Arnold Thackray. Book review: General Human Understanding. Volume 1: General Introduction and Part 1. The Collective Use and Evolution of Concepts. By Stephen Toulmin. Oxford: Clarendon Press: Oxford University Press, 1972. Pp. xii + 520. £4.75. British Journal for the History of Science, 7(1):80–81, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025179.

Thackray:1983:BRLa

[Tha83a]  John Thackray.  Book re-

view: Life and earth sciences: Roy Porter (ed.), The Earth generated and anatomized. By William Hobbs. Ithaca and London: Cornell University Press and British Museum (Natural History) 1981. Pp.

158. £24.50/$43.75; £16 (paperback). British Journal for the History of Science, 16(3): 287–288, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026157.

Thackray:1983:BRLb

[Tha83b] John Thackray. Book review: Life and earth sciences: S. M. Andrews, The discovery of fossil fishes in Scotland up to 1845. Edinburgh: Royal Scottish Museum, 1982. Pp. 87. £5.00. British Journal for the History of Science, 16(3): 288, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026158.

Thackray:1984:BRHa

[Tha84a] J. C. Thackray. Book review: History of Natural History: Adrian Desmond, Archetypes and ancestors: palaeontology in Victorian London, 1850–1870. London: Blond and Briggs, 1982. Pp. 287. ISBN 0-85634-121-5. £15.95. British Journal for the History of Science, 17(3):312–313, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026626;

http://www.jstor.org/
stable/4026627.

Thackray:1984:BRHb

[Tha84b] John Thackray. Book review:

History of Natural History: Ron J. Cleevely, World palaeontological collections. London: British Museum (Natural History) and Mansell, 1982. Pp.

365. ISBN 0-56500-850-1/0-7201-1655-4. £50. British Journal for the History of Science, 17(3):322–323, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026637.

Thackray:1985:BRR

[Tha85] John Thackray. Book review:

Roy Porter, The Earth sciences, an annotated bibliography. New York and London: Garland Publishing Inc., 1983. Pp. xviii + 192. ISBN 0-8240-9267-8, $35. British Journal for the History of Science, 18(2):247, July 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026348.

Thackray:1987:BRJ

[Tha87]  John Thackray.  Book re-

view: J. A. Secord. Controversy in Victorian Geology: the Cambrian-Silurian Dispute. Princeton: Princeton University Press, 1986. Pp. xix + 364. ISBN 0-691-08417-3.

£37.00. British Journal for the History of Science, 20(4): 483–484, October 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026424.

Theodorides:1966:HE

[Thé66] Jean Théodoridès. Humboldt and England. British Journal for the History of Science, 3(1):39–55, June 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025101.

Theisen:1978:NJB

[The78] Wilfred Theisen. A note on John of Beaumont’s version of Euclid’s de visu. British Journal for the History of Science, 11(2):151–155, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025727.

Theerman:1987:BRJ

[The87] Paul Theerman. Book review: John Hendry. James Clerk Maxwell and the Theory of the Electromagnetic Field. Bristol and Boston: Adam Hilger, 1986. Pp. xix + 305. ISBN 0-85274-563-X. £30.00. British Journal for the History of Science, 20(3):365–366, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026375.

Thomson:1974:JNU

[Tho74a] Ron B. Thomson. Jordanus de Nemore and the University of Toulouse. British Journal for the History of Science, 7 (2):163–165, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025202.

Thoren:1974:KSL

[Tho74b] Victor E. Thoren. Kepler’s Second Law in England. British Journal for the History of Science, 7(3):243–256, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://

www.jstor.org/stable/

4025532. With an appendix by Owen Gingerich and Barbara Welther: Note on Flamsteed’s lunar tables.

Thoren:1976:BRAa

[Tho76a] Victor E. Thoren. Book review: Astronomy The Reception of Copernicus’ Heliocentric Theory. Ed. by J. Dobrzycki. Dordrecht, Holland: Reidel, 1972. Pp. 367. No price stated. British Journal for the History of Science, 9(3): 326, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025687.

Thoren:1976:BRAb

[Tho76b] Victor E. Thoren. Book review: Astronomy The Scientific World of Copernicus. Ed. by B. Bienowska. Dordrecht, Holland: Reidel, 1973. Pp. xii + 142. $19.50. British Journal for the History of Science, 9 (3):326, November 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Thomas:1987:DTS

[Tho87] Donald E. Thomas. Diesel: technology and society in industrial Germany. University of Alabama Press, Tuscaloosa, AL, USA, 1987. ISBN 0-8173-0295-6. xii + 279 pp. LCCN TJ140.D5 T48 1987.

Thoma:1989:BRM

[Tho89] Deborah Thoma. Book review: Michael Chapman. Constructive evolution. Cambridge: Cambridge University Press, 1988. ISBN 0-521-33163-3, £12.95 (paperback) 0-521-36712-3, £35 (hardcover). British Journal for the History of Science, 22(4):468– 472, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026943.

Tilling:1975:BRS

[Til75a] Laura Tilling. Book review: Seventeenth Century: The Unpublished First Version of Isaac Newton’s Cambridge Lectures on Optics 1670–1672. A facsimile of the autograph. Introduction by D. T. Whiteside. Cambridge: The University Library, 1973. Pp. x + 129. £10. British Journal for the History of Science, 8(1):84, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025828.

Tilling:1975:EEG

[Til75b] Laura Tilling. Early experimental graphs. British Journal for the History of Science, 8(3):193–213, November 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025556.

Tilling:1976:BRA

[Til76]  Laura Tilling.  Book re-

view: Astronomy Isaac Newton’s Theory of the moon’s motion (1702). Ed. by I. Bernard Cohen. Folkestone: Dawson, 1975. Pp. viii + 170. £12.00. British Journal for the History of Science, 9(3):326–327, November 1976. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025688.

Todd:1983:TRA

[Tod83] Alexander R. (Alexander Robertus) Todd. A time to remember: the autobiography of a chemist. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25593-7. viii + 257 pp. LCCN QD22.T74 A35

1983. URL http://www.loc.

                                        gov/catdir/enhancements/
                                        fy1001/83005172-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy1001/83005172-t.html.

Tomkeieff:1965:LCL

[Tom65] S. I. Tomkeieff. Letters from Charles Lyell and Adam Sedgwick to H. B. Tristram. British Journal for the History of Science, 2(3):251–253, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024939.

Tomes:1984:GCT

[Tom84] Nancy Tomes. A generous confidence: Thomas Story Kirkbride and the art of asylumkeeping, 1840–1883. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-24172-3. xv + 387 pp. LCCN RC445.P4 P685 1984.

Torrens:1980:OVE

[Tor80] H. S. Torrens. Obituary: Victor Eyles 10 October 1895– 8 March 1978. British Journal for the History of Science, 13(3):282–283, November 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026214.

Torrens:1983:AAM

[Tor83] H. S. Torrens. Arthur Aikin’s mineralogical survey of Shropshire 1796–1816, and the contemporary audience for geological publications. British Journal for the History of Science, 16(2):111–153, July 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026236.

Torrens:1987:OJM

[Tor87]  H. S. Torrens.  Obituary:

Joan Mary Eyles, née Biggs: 1907–1986. British Journal for the History of Science, 20 (3):349–350, July 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026362.

Torrini:1988:OHS

[Tor88] Maurizio Torrini. Observations on the history of science in Italy. British Journal for the History of Science, 21(4): 427–446, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026962.

Tort:1992:DS

[Tor92] Patrick Tort, editor. Darwinisme et société. (French) [Darwinism and society]. Presses universitaires de France, Paris, France, 1re édition, 1992. ISBN 2-13-044774-0. 690 pp. LCCN HM106 .D26 1992. 480 FF.

Talbot:1966:SEK

[TP66] G. R. Talbot and A. J. Pacey.

Some early kinetic theories of gases: Herapath and his predecessors. British Journal for the History of Science, 3(2): 133–149, December 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025026.

sTeich:1990:FSL

[TP90] Mikuláš Teich and Roy Porter, editors. Fin de siècle and its legacy. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-34108-6 (hardcover), 0-521-34915-X (paperback). xii + 345 pp. LCCN D395 .F55 1990. URL http:

//www.loc.gov/catdir/
description/cam023/
89022347.html;   http://
www.loc.gov/catdir/toc/
cam021/89022347.html.

Timms:1992:PCC

[TR92] Edward Timms and Ritchie Robertson, editors. Psychoanalysis in its cultural context, volume 3 of Austrian studies. Edinburgh University Press, Edinburgh, Scotland, 1992. ISBN 0-7486-0359-X. xi + 209 pp. LCCN BF173 .P7756 1992.

Traynham:1987:EHO

[Tra87] James G. Traynham, editor. Essays on the history of organic chemistry. Louisiana State University Press, Baton Rouge, 1987. ISBN 0-8071-1293-3. LCCN QD248 .E87 1987.

Travis:1993:RMO

[Tra93] A. S. Travis. The rainbow makers: the origins of the synthetic dyestuffs industry in western Europe. Lehigh University Press, Bethlehem, PA, USA, 1993. ISBN 0-934223-18-1. 335 pp. LCCN HD9660.D843 E858 1993.

Trengove:1969:DE

[Tre69] Leonard Trengove. Dalton as experimenter. British Journal for the History of Science, 4(4): 394–398, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025232.

Trenn:1979:CRE

[Tre79] Thaddeus J. Trenn. The central role of energy in Soddy’s holistic and critical approach to nuclear science, economics, and social responsibility. British Journal for the History of Science, 12(3):261–276, November 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026001.

Trenn:1987:BRF

[Tre87a] T. J. Trenn. Book review:

Françoise Giroud. Marie Curie: A Life. Translated from the French by Lydia Davis. New York: Holmes and Meier, 1986. Pp. v + 291. ISBN 0-8419-0977-6. $34.50. British Journal for the History of Science, 20(2):232, April 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026317.

Trenn:1987:BRH

[Tre87b] Thaddeus J. Trenn. Book review: Hilde Levi. George de Hevesy: Life and Work. Bristol and Boston: Adam Hilger, 1985. Pp. 147. ISBN 0-85274-555-9. £15.00, $25.00. British Journal for the History of Science, 20(1):118, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026470.

Trenn:1988:BRE

[Tre88] Thaddeus J. Trenn. Book review: Elisabeth Crawford, J. L. Heilbron, Rebecca Ullrich. The Nobel Population (1901–1937): A census of the Nominators and Nominees for the Prizes in Physics and Chemistry. Berkeley: Office for History of Science and Technology, 1987. Pp. vii + 337. ISBN 0-918102-15-4. $20.00. British Journal for the History of Science, 21(4): 497, December 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026969.

Trigger:1989:HAT

[Tri89] Bruce G. Trigger. A history of archaeological thought. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-32878-0, 0-521-33818-2 (paperback). xv + 500 pp. LCCN CC100 .T75 1989. URL http:

//www.loc.gov/catdir/
description/cam023/
88016926.html;   http://
www.loc.gov/catdir/toc/
cam021/88016926.html.

Trombley:1989:SFT

[Tro89] Stephen Trombley. Sir Frederick Treves: the extra-ordinary Edwardian. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1989. ISBN 0-415-03423-X. ix + 218 pp. LCCN RD27.35.T74 T76 1989.

Turner-:1987:HRA

[TST87] H. H. (Herbert Hall) Turner, R. A. (Ralph Allen) Sampson, and R. J. (Roger John) Tayler, editors. History of the Royal Astronomical Society. Published for the Society by Blackwell Scientific Pub., Oxford, UK, 1987. ISBN 0-632-02173-X (vol. 1), 0-632-01791-0 (vol. 2), 0-632-02175-6 (paperback: vol. 1), 0-632-01792-9 (paperback: vol. 2). ???? pp. LCCN QB1.R5.

Tarbell:1986:EHO

[TT86]  D.  Stan-

ley Tarbell and Ann Tracy Tarbell. Essays on the history of organic chemistry in the United States, 1875–1955. Folio Publishers, Nashville, Tenn., 1986. ISBN 0-939454-03-3. x + 433 pp. LCCN QD248.5.U6 T37 1986. US$22.00 (est.).

Tunbridge:1992:LKH

[Tun92] Paul Tunbridge. Lord Kelvin, his influence on electrical measurements and units, volume 18 of IEE history of technology series. P. Peregrinus on behalf of the Institution of Electrical Engineers, London, UK, 1992. ISBN 0-86341-237-8. ix + 107 pp. LCCN QC16.K3 T86 1992.

Tupper:1982:BRT

[Tup82] Stephen M. Tupper. Book review: Technology and Development Robert Lewis, Science and industrialisation in the USSR: industrial research and development 1917–1940. London: Macmillan in association with the Centre for Russian and East European Studies, University of Birmingham. 1979 Pp. xiv + 211. £12.00. British Journal for the History of Science, 15(1):93–94, March 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025939.

Turner:1968:BRS

[Tur68] G. L’E. Turner. Book review: Scientific Instruments and Museums: Robert T. Gunther. A Pioneer in the History of Science, 1869–1940. By A. E. Gunther, Early Science in Oxford, Vol. xv. Pp. xiii + 520. Oxford: Printed for the Subscribers. 1967. £5 5s. British Journal for the History of Science, 4(2): 180–181, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025153.

Turner:1969:BRS

[Tur69] G. L’E. Turner. Book review: Scientific Instruments The Evolution of the Microscope. By S. Bradbury. Pp. x + 357. Oxford: Pergamon Press. 1967. £4. British Journal for the History of Science, 4(3):288–289, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025479.

Turner:1970:BM

[Tur70a] G. L’E. Turner. Back matter. British Journal for the History of Science, 5(2):216, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025441.

Turner:1970:BRS

[Tur70b] G. L’E. Turner. Book review:

Scientific Instruments The Apparatus of Science at Harvard 1765–1800. Collection of Historical Scientific Instruments, Harvard University. By David P. Wheatland, assisted by Barbara Carson. Cambridge, Mass., and London: Harvard University Press and Oxford University Press. 1968. Pp. xii + 204. £9 10s. British Journal for the History of Science, 5(1):94–95, June 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025363.

Turner:1970:BRT

[Tur70c] G. L’E. Turner. Book review: Twentieth Century Early Papers on Diffraction of X-rays by Crystals. Ed. by J. M. Bijvoet, W. G. Burgers and G. Hägg. Utrecht: published for the International Union of Crystallography by A. Oosthoek’s Uitgeversmaatschappij. 1969. Pp. xvi + 372. £5 14s. British Journal for the History of Science, 5(2):208, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025438.

Turner:1970:RU

[Tur70d] G. L’E. Turner. Review: [untitled]. British Journal for the History of Science, 5(2): 208, December 1970. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025438.

Turner:1972:BRE

[Tur72a] G. L’E. Turner. Book review: Eighteenth and Nineteenth Centuries: La Genèse de la Science des Cristaux. By Hélène Metzger. New edition. Paris: Albert Blanchard, 1969. Pp. 284. 15 francs. British Journal for the History of Science, 6(1):97, June 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025279.

Turner:1972:BRS

[Tur72b] G. L’E. Turner. Book review: Scientific Instruments II Museo di Storia della Scienza a Firenze. By Maria Luisa Righini Bonelli. Milan: Electra, 1968. Pp. 252. No price stated. British Journal for the History of Science, 6(2): 206, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025294.

Turner:1972:JJL

[Tur72c] G. L’E. Turner. J. J. Lister’s account book. British Journal for the History of Science, 6(1): 106, June 1972. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Turner:1973:BRR

[Tur73a] A. J. Turner. Book review:

Renaissance: The Rosicrucian Enlightenment. By Frances A. Yates. London: Routledge & Kegan Paul, 1971. Pp. xvi + 269. £4.50. British Journal for the History of Science, 6(4): 442–444, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025515.

Turner:1973:BRS

[Tur73b] A. J. Turner. Book review: Sixteenth and Seventeenth Centuries: Mystiques, Spirituels, Alchimistes du XVIe siècle allemand. By Alexandre Koyré. Paris: Gallimard, 1971. Pp.

184. No price stated. British Journal for the History of Science, 6(3):319, June 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025454.

Turner:1973:BRE

[Tur73c] G. L’ E. Turner. Book review: Eighteenth Century: Natural Philosophy through the 18th Century and Allied Topics. Ed. by Allan Ferguson. London: Taylor and Francis, 1972. Pp. xii + 164. £5. British Journal for the History of Science, 6(4):447–448, December 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025518.

Turner:1973:BRC

[Tur73d] G. L’E. Turner. Book review: Crystallography Early Papers on Diffraction of X-rays by Crystals. Volume II. Ed. by J. M. Bijvoet, W. G. Burgers and G. Hägg. Utrecht: A. Oosthoek’s Uitgeversmaatschappij (for the International Union of Crystallography), 1972. Pp xx + 484. No price stated. British Journal for the History of Science, 6(3):317–318, June 1973.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025452.

Turner:1974:BRS

[Tur74a] A. J. Turner. Book review: Sixteenth and Seventeenth Centuries: Polish Academy of Science. Nicolas Copernicus Complete Works. Volume I. The Manuscript of Nicolas Copernicus’ ‘On the revolutions’. London, Warsaw, and Cracow: Macmillan and Polish Scientific Publishers, 1972. Pp. xii + 24 + 23 plates + facsimile on unnumbered pages. £35. British Journal for the History of Science, 7(3):292– 293, November 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025542.

Turner:1974:HTE

[Tur74b] A. J. Turner. Hooke’s theory of the Earth’s axial displacement: Some contemporary opinion. British Journal for the History of Science, 7 (2):166–170, July 1974. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025203.

Turner:1974:BRT

[Tur74c] G. L’E. Turner. Book review: Technology A History of Glassmaking. By R. W. Douglas and Susan Frank. Henleyon-Thames: G. T. Foulis, 1972. Pp. xii + 213. £4.50. British Journal for the History of Science, 7(3):291–292, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025541.

Turner:1975:BRT

[Tur75] G. L’E. Turner. Book review: Technology Windmühlen, der Stand der Forschung über das Vorkommen und den Ursprung. By Jannis C. Notebaart. The Hague and Paris: Mouton Verlag, 1972. Pp. 405. Hfl. 69. British Journal for the History of Science, 8(1):77, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025823.

Turner:1977:BRS

[Tur77a] G. L.’E. Turner. Book review: Scientific Instruments Scientific Instruments. By Harriet Wynter and Anthony Turner. London: Studio Vista, 1975. Pp.

239. £12.50. British Journal for the History of Science, 10 (1):77–78, March 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025591.

Turner:1977:OSD

[Tur77b] R. Steven Turner. The Ohm– Seebeck dispute, Hermann von Helmholtz, and the origins of physiological acoustics. British Journal for the History of Science, 10(1):1–24, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025578.

Turner:1978:BRA

[Tur78a] G. L’E. Turner. Book review:

Architecture English Architecture: An Illustrated Glossary. By James Stevens Curl. Newton Abbot: David & Charles, 1977. Pp. 192. £9.50. British Journal for the History of Science, 11(2):174–175, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025733.

Turner:1978:BRG

[Tur78b] G. L’E. Turner. Book review: General The Nature of Science: The History of Science in Western Culture since 1600. By David Knight. London: André Deutsch, 1976. Pp. 215. £4.95. British Journal for the History of Science, 11(2):172, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025731.

Turner:1978:BRS

[Tur78c] G. L’E. Turner. Book review:

Scientific Instruments Thinkers and Tinkers: Early American Men of Science. By Silvio A.

Bedini. New York: Charles Scribner’s Sons, 1975. Pp. xix + 520. $17.50. British Journal for the History of Science, 11(3):283–284, November 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025784.

Turner:1981:BRE

[Tur81] G. L’E. Turner. Book review: Eighteenth Century: The Construction of a Thermometer. By James Six. Prefaced by an Account of his Life & Works and the Use of his Thermometer over Two Hundred Years by Jill Austin & Anita McConnell. London: Nimbus Books, 1980. Pp. 28 + xiv + 64 + xxiv. £15.00. British Journal for the History of Science, 14(1):90–91, March 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026079.

Turner:1982:BRI

[Tur82] G. L’E. Turner. Book review: Instruments D. A. Wittop Koning & G. M. M. Houben, 2000 Jaar Gewichten in de Nederlanden; stelsels, ijkwezen, vormen, makers, merken, gebruik. Lochem-Poperinge: Uitgeversmaatschappij De Tijdstroom, 1980. Pp. 222. Dfl84.50. British Journal for the History of Science, 15(3):313–314, November 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026230.

Turner:1983:BRJ

[Tur83a] G. L’E. Turner. Book review: Journals and Dictionaries Ronald Edward Zupko, Italian weights and measures from the middle ages to the ninetenth century. Philadelphia: American Philosophical Society, 1981. Pp. lxxxiv + 339. $8.00. British Journal for the History of Science, 16 (3):318, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026186.

Turner:1983:NCS

[Tur83b] Gerard L’Estrange Turner.

Nineteenth-Century scientific instruments. Sotheby Publications, London, UK, 1983. ISBN 0-520-05160-2, 0-85667-170-3. 320 + 20 pp. LCCN Q185 .T86 1983.

Turner:1983:BRM

[Tur83c] R. Steven Turner. Book review: Mathematical Sciences: Paul D. Sherman, Colour vision in the nineteenth century. The Young–Helmholtz– Maxwell theory. Bristol: Adam Hilger Ltd, 1981. Pp. xiii + 233. $77.00/£35.00. British Journal for the History of Science, 16(3):297–298, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026167.

Turner:1987:BRM

[Tur87a] G. L. E. Turner. Book review: M. J. G. Cattermole & A. F. Wolfe. Horace Darwin’s Shop: a History of the Cambridge Scientific Instrument Company 1878–1968. Bristol: Adam Hilger, 1987. Pp. xvi + 285. ISBN 0-85274-569-9. £35.00. British Journal for the History of Science, 20(4):486–487, October 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026426.

Turner:1987:ST

[Tur87b] Gerard L’E. Turner. Scientific toys. British Journal for the History of Science, 20(4): 377–398, October 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026415.

Turner:1987:TTC

[Tur87c]  Roland Turner,  editor.

Thinkers of the Twentieth Century. St. James Press, Chicago, IL, USA, second edition, 1987. ISBN 0-912289-83-X. vii + 977 pp. LCCN CT120 .T45 1987.

Turner:1988:BRJ

[Tur88] A. J. Turner. Book review: J. V. Field & J. J. Gray. The Geometrical Work of Girard Desargues. New York, Berlin, Heidelberg, London, Paris & Tokyo: Springer-Verlag, 1987. Pp. x + 237 with 69 illustrations, ISBN 0-387-96403-7 and 3-540-96403-7. DM 158.00. British Journal for the History of Science, 21 (1):111–112, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026868.

Turner:1990:SIE

[Tur90] Gerard L’Estrange Turner, editor. Scientific instruments and experimental philosophy, 1550– 1850, volume CS331 of Variorum collected studies series. Variorum, Aldershot, Hampshire, UK, 1990. ISBN 0-86078-280-8. LCCN Q185 .T87 1991.

Turner:1994:EMV

[Tur94] R. Steven (Roy Steven) Turner.

In the eye’s mind: vision and the Helmholtz–Hering controversy. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03397-8. xiv + 338 pp. LCCN QP491 .T87 1994. US$49.50. URL http:

//www.loc.gov/catdir/
description/prin031/
93037258.html;   http://
www.loc.gov/catdir/toc/
prin031/93037258.html.

Turkle:1996:LSI

[Tur96] Sherry Turkle. Life on the screen: identity in the age of the Internet. Weidenfeld and Nicolson, London, UK, 1996.

ISBN 0-297-81514-8. 347 pp. LCCN QA76.9.C66 T87 1996.

Tselos:1987:GAM

[TW87] George D. Tselos and Colleen Wickey. A guide to archives and manuscript collections in the history of chemistry and chemical technology, volume 7 of Center for History of Chemistry publication. Center for History of Chemistry, Philadelphia, PA, USA, 1987. ISBN 0-941901-05-X. viii + 198 pp. LCCN QD18.U6 T84

1987. URL http://www.loc.
gov/catdir/enhancements/
fy0630/95167507-d.html.

Taton:1989:PAR

[TW95] René Taton and Curtis Wilson, editors. Planetary astronomy from the Renaissance to the rise of astrophysics, volume 2 of The General history of astronomy. Cambridge University Press, Cambridge, UK, 1989–1995. ISBN 0-521-24254-1 (pt. A), 0-521-35168-5 (pt. B). ???? pp. LCCN QB15 .G38 1984 vol. 2. URL http://www.

loc.gov/catdir/
description/cam023/
88025817.html;   http://
www.loc.gov/catdir/toc/
cam022/88025817.html.

Tweedale:1989:BRM

[Twe89] Geoffrey Tweedale. Book review: Martin Campbell-Kelly (ed.). The Works of Charles Babbage. London: Pickering & Chatto (Publishers) Ltd, 1989.

11 vols. ISBN 1-85196-005-8. £500. British Journal for the History of Science, 22(4):481– 482, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026950.

Twyman:1988:SMD

[Twy88] Alan Twyman. In search of the mysterious Doctor Weekes: (a fragment of Sandwich history). Field’s Secretarial Services, Dover, UK, 1988. ISBN 0-9513889-0-8 (paperback). xi + 162 pp. LCCN QC515.W4 T99 1988.

Tylecote:1964:RLW

[Tyl64] R. F. Tylecote. Roman lead working in Britain. British Journal for the History of Science, 2(1):25–43, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025082.

Tylecote:1975:BRC

[Tyl75] R. F. Tylecote. Book review: China Science and Civilisation in China. Volume V, Part 2. By Joseph Needham. Cambridge: Cambridge University Press, 1974. Pp. 510. £12.00. British Journal for the History of Science, 8(3): 254–255, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025566.

Tytler:1982:PEN

[Tyt82] Graeme Tytler. Physiognomy in the European novel: faces and fortunes. Princeton University Press, Princeton, NJ, USA, 1982. ISBN 0-691-06491-1. xix + 436 pp. LCCN PN3500 .T9. US$25.00.

Ucko:1991:AR

[U^{+}91] Peter J. Ucko et al., editors. Avebury reconsidered: from the 1660s to the 1990s. Unwin Hyman, London, UK, 1991. ISBN 0-04-445919-X. xiv + 293 pp. LCCN GN776.22.G7 A94 1991. US$100.00.

Ubbelohde:1963:EST

[Ubb63] A. R. Ubbelohde. Edwardian science and technology: Their interactions. British Journal for the History of Science, 1 (3):217–226, June 1963. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024921.

Underwood:1962:F

[Und62] E. Ashworth Underwood. Foreword. British Journal for the History of Science, 1(1): 3, June 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025070.

Underwood:1963:DSH

[Und63] E. Ashworth Underwood. The Dutch Society for the History of Science. British Journal for the History of Science, 1 (4):384, December 1963. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024911.

Underwood:1964:BRP

[Und64a] E. Ashworth Underwood. Book review: The Plague of the Philistines, and other Medical-Historical Essays. By J. F. D. Shrewsbury. London: Victor Gollancz. 1964. Pp. 189. 25s. net. British Journal for the History of Science, 2(2):169– 171, December 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025017.

Underwood:1964:ERa

[Und64b] E. Ashworth Underwood. Essay review. British Journal for the History of Science, 2(2): 165–168, December 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025016.

Underwood:1964:ERb

[Und64c] E. Ashworth Underwood. Essay reviews. British Journal for the History of Science, 2 (1):68–71, June 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025086.

Underwood:1964:ERc

[Und64d] E. Ashworth Underwood. Essay reviews. British Journal for the History of Science, 2 (1):71–75, June 1964. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025087.

Underwood:1965:ODW

[Und65] E. Ashworth Underwood. Obituary: Dorothea Waley Singer (1882–1964). British Journal for the History of Science, 2 (3):260–262, June 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024942.

Underwood:1966:BRH

[Und66a] E. Ashworth Underwood. Book review: History of Biological Sciences and Medicine Circulation of the Blood, Men and Ideas. Edited by Alfred P. Fishman and Dickinson W. Richards. Pp. xiv + 859. New York, Oxford University Press, 1964. £7 7s. British Journal for the History of Science, 3 (1):93, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025122.

Underwood:1966:BRO

[Und66b] E. Ashworth Underwood. Book review: Overtures to Biology: The Speculations of Eighteenth-Century Naturalists. By Philip C. Ritterbush. Pp. x + 287. Yale University Press, New Haven and London, 1964. 56s. British Journal for the History of Science, 3(2):196–197, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025040.

Underwood:1966:BRE

[Und66c] E. Ashworth Underwood. Book review: The Evolution of Biology. By M. J. Sirks and Conway Zirkle. Pp. vi + 376. New York: Ronald Press Company, 1964. $6. British Journal for the History of Science, 3(2): 197, December 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025041.

Underwood:1971:OFH

[Und71] E. Ashworth Underwood. Obituary: Francis Herbert Culverhouse Butler (1894–1970). British Journal for the History of Science, 5(3):319–320, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025347.

Underwood:1974:DF

[Und74]  E.  Ashworth  Under-

wood. Dorothy Feyer. British Journal for the History of Science, 7(3):304, November 1974.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025552.

Underwood:1975:BRM

[Und75a] E. Ashworth Underwood. Book review: Medicine Therapeutics from the Primitives to the 20th Century. With an Appendix: History of Dietetics. By Erwin H. Ackerknecht. New York: Hafner Press, and London: Collier Macmillan, 1973. Pp. x + 194. £6.25. British Journal for the History of Science, 8(2):178, July 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025646.

Underwood:1975:BRS

[Und75b] E. Ashworth Underwood. Book review: Seventeenth Century: The Baglivi Correspondence from the Library of Sir William Osler. Ed. by Dorothy M. Schullian. Ithaca, N.Y., and London: Cornell University Press, 1974. Pp. xxii + 531. £21.40. British Journal for the History of Science, 8(3): 260–262, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025571.

Underwood:1979:BRN

[Und79a] E. Ashworth Underwood. Book review: Nineteenth and Twentieth Centuries: The Development of Microbiology. By Patrick Collard. Cambridge: Cambridge University Press, 1976. Pp. viii + 201. £6.50. British Journal for the History of Science, 12(1):102– 103, March 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025675.

Underwood:1979:BRS

[Und79b] E. Ashworth Underwood. Book review: Seventeenth Century: The Correspondence of Marcello Malpighi. Edited by Howard B. Adelmann. 5 vols. Ithaca, NY and London: Cornell University Press, 1975. Pp. xxii + 2, 228. £50.00. British Journal for the History of Science, 12(1):95–98, March 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025671.

Underwood:1980:BRN

[Und80] E. Ashworth Underwood. Book review: Nineteenth Century: The Life of William Carmichael M’Intosh, MD, FRS, of St Andrews, 1838–1931, a Pioneer in Marine Biology. By A. E. Gunther. Edinburgh: Scottish Academic Press for the University of St Andrews, 1977. Pp. 214. £5.00. British Journal for the History of Science, 13(2):176– 177, July 1980. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

VanHelden:1985:MUC

[Van85] Albert Van Helden. Measuring the universe: cosmic dimensions from Aristarchus to Halley. University of Chicago Press, Chicago, IL, USA, 1985. ISBN 0-226-84881-7, 0-226-84882-5 (paperback). viii + 203 pp. LCCN QB15 .V33

                                        1985. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi052/84016397.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0608/84016397-t.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0609/84016397-b.html.

VanHelden:1986:BRG

[Van86] Albert Van Helden. Book review: G. L. Toomer (tr). Ptolemy’s Almagest. London: Duckworth. Pp. ix + 693. ISBN 0-7156-1588-2. £55.00. British Journal for the History of Science, 19(1):115– 116, March 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026487.

VanRiper:1993:MAM

[Van93] A. Bowdoin Van Riper. Men among the mammoths: Victorian science and the discovery of human prehistory. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1993. ISBN 0-226-84991-0 (hardcover), 0-226-84992-9 (paperback). xv + 267 pp. LCCN GN805 .V36

1993. URL http://www.loc.
gov/catdir/description/
uchi052/93017006.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/93017006-t.html.

Voltaire:1980:CWV

[VB0 ] Voltaire and Theodore Besterman. The complete works of Voltaire = Les Œuvres completes de Voltaire. Voltaire Foundation at the Taylor Institution, Oxford, UK, 1980- . ISBN 0-7294-0745-4 (vol. 1A), 0-7294-0774-8 (vol. 1B), 0-7294-0833-7 (vol. 1C), 0-7294-0790-X (vol. 3A), 0-7294-0474-9 (vol. 3B), 0-7294-0507-9 (vol. 4), 0-7294-0567-2 (vol. 5), 0-7294-0359-9 (vol. 8), 0-7294-0646-6 (vol. 9), 0-7294-0325-4 (vol. 10), 0-7294-0360-2 (vol. 14), 0-7294-0374-2 (vol. 15), 0-7294-0794-2 (vol. 16), 0-7294-0421-8 (vol. 17), 0-7294-0893-0 (vol. 18A), 0-7294-0894-9 (vol. 18B), 0-7294-0508-7 (vol. 19), 0-7294-0753-5 (vol. 20A), 0-7294-0754-3 (vol. 20B), 0-7294-0874-4 (vol. 22), 0-7294-0871-X (vol. 28A), 0-7294-0899-X (vol. 28B), 0-7294-0800-0 (vol. 30A), 0-7294-0814-0 (vol. 30B), 0-7294-0828-0 (vol. 30C), 0-7294-0426-9 (vol. 31A), 0-7294-0427-7 (vol. 31B), 0-7294-0842-6 (vol. 32A), 0-7294-0843-4 (vol. 32B), 0-7294-0338-6 (vol. 33), 0-7294-0376-9 (vol. 35), 0-7294-0377-7 (vol. 36), 0-7294-0854-X (vol. 38), 0-7294-0922-8 (vol. 40), 0-7294-0942-2 (vol. 45A), 0-7294-0568-0 (vol. 46– 47), 0-7294-0717-9 (vol. 46), 0-7294-0718-7 (vol. 47), 0-7294-0118-9 (vol. 48), 0-7294-0968-6 (vol. 49B), 0-7294-0334-3 (vol. 50), 0-7294-0729-2 (vol. 56A), 0-7294-0644-X (vol. 56B), 0-7294-0744-6 (vol. 56C), 0-7294-0348-3 (vol. 62), 0-7294-0361-0 (vol. 63A), 0-7294-0896-5 (vol. 63B), 0-7294-0824-8 (vol. 63C), 0-7294-0312-2 (vol. 64), 0-7294-0476-5 (vol. 66), 0-7294-0900-7 (vol. 67), 0-7294-0823-X (vol. 68), 0-7294-0475-7 (vol. 69), 0-7294-0863-9 (vol. 71A), 0-7294-0825-6 (vol. 71B), 0-7294-0813-2 (vol. 73), 0-7294-0719-5 (vol. 74A), 0-7294-0892-2 (vol. 74B), 0-7294-0943-0 (vol. 75A), 0-7294-0982-1 (vol. 75B), 0-7294-0931-7 (vol. 80A), 0-7294-0936-8 (vol. 136), 0-7294-0937-6 (set, vol. 137A-137B), 0-7294-0963-5 (v. 137A), 0-7294-0964-3 (vol. 137B), 0-7294-0938-4 (vol. 138), 0-7294-0777-2 (vol. 141), 0-7294-0887-6 (vol. 142). ???? pp. LCCN PQ2070 1980.

vonBaer:1993:CKE

[vBDGO93] Karl Ernst von Baer, Anton Dohrn, Christiane Groeben, and Jane M. (Jane Marion) Oppenheimer. Correspondence, Karl Ernst von Baer (1792– 1876), Anton Dohrn (1840– 1909), volume 83, part 3 of Transactions of the American Philosophical Society, 0065-9746. American Philosophical Society, Philadelphia, PA, USA, 1993. ISBN 0-87169-833-1. 156 pp. LCCN QH31.B15 A4 1993.

vanderWaerden:1985:HAA

[vdBLW85] B. L. van der (Bartel Leendert) Waerden. A history of algebra: from al-Khwarizmi to Emmy Noether. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-13610-X (US). xi + 271 pp. LCCN QA151 .W343 1985.

vanderVeer:1991:UVQ

[vdVV91] René van der Veer and Jaan Valsiner. Understanding Vygotsky: a quest for synthesis. Blackwell, Oxford, UK, 1991. ISBN 0-631-16528-2. x + 450 pp. LCCN BF109.V95 V44 1991. US$35.00.

Veale:1964:BRH

[Vea64] Elspeth M. Veale. Book review: A History of the Worshipful Society of Apothecaries of London. Vol. I, 1617–1815. By C. Wall, H. C. Cameron and E. A. Underwood. Pp. xiv + 450. Publications of the Wellcome Historical Medical Museum, new series, no. 8. Oxford University Press, 1963. £2 15s. British Journal for the History of Science, 2(1): 78–79, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025090.

Verwey:1985:PAB

[Ver85] Gerlof Verwey. Psychiatry in an anthropological and biomedical context: philosophical presuppositions and implications of German psychiatry, 1820– 1870, volume 15 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1985. ISBN 90-277-1713-3. xix + 316 pp. LCCN RC450.G3 V4713 1985.

Vernon:1988:FNT

[Ver88] Keith Vernon. The founding of numerical taxonomy. British Journal for the History of Science, 21(2):143–159, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026976.

Vernon:1989:BRD

[Ver89a] Keith Vernon. Book review: David L. Hull. Science as a Process: An Evolutionary Account of the Social and Conceptual Development of Science. Chicago and London: The University of Chicago Press, 1988. Pp. xii + 586. ISBN 0-226-36050-4. £31.95. British Journal for the History of Science, 22(4):461–462, December 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026937.

Vernon:1989:BRM

[Ver89b] Keith Vernon. Book review:

Michael Ruse. Philosophy of Biology Today. Albany, New York: State University of New York Press, 1988. Pp. x + 155. ISBN 0-88706-910-X. $29.50, $9.95. British Journal for the History of Science, 22(4):460– 461, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026936.

vanGinneken:1992:CPP

[vG92] Jaap van Ginneken. Crowds, psychology, and politics, 1871– 1899. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40418-5 (hardcover). xii + 269 pp. LCCN HM281 .G46

1992. URL http://www.loc.
gov/catdir/description/
cam024/91042797.html;
http://www.loc.gov/
catdir/toc/cam029/
91042797.html.

Vallisnieri:1991:E

[VG98] Antonio Vallisnieri and Dario Generali. Epistolario, volume 9 of Filosofia e scienza nel Cinquecento e nel Seicento; Pubblicazioni del “Centro di studi del pensiero filosofico del Cinquecento e del Seicento in relazione ai problemi della scienza” del Consiglio nazionale delle ricerche. Serie III, Testi. F. Angeli, Milano, Italy, 1991– 1998. ISBN 88-204-6733-X (vol. 1). ???? pp. LCCN QH31.V35 A4 1991. L70000.

vonGuericke:1994:NCM

[vGA94] Otto von Guericke and Margaret Glover Foley Ames. The new (so-called) Magdeburg experiments of Otto von Guericke, volume 137 of Archives internationales d’histoire des idées. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2399-8. xxiv + 394 pp. LCCN Q155 .G9313 1994.

vonHumboldt:1997:CSP

[vH97] Alexander von Humboldt. Cosmos: a sketch of a physical description of the universe. Foundations of natural history. The Johns Hopkins University Press, Baltimore, MD, USA, 1997. ISBN 0-8018-5502-0 (vol. 1 : paperback), 0-8018-5503-9 (vol. 2 : paperback). ???? pp. LCCN QB981 .H8613

                                        1997. URL http://www.loc.
                                        gov/catdir/bios/jhu051/
                                        96036421.html;
                                        http://www.loc.gov/
                                        catdir/description/
                                        jhu052/96036421.html.

vonHelmholtz:1995:SCP

[vHC95] Hermann von Helmholtz and David Cahan. Science and culture: popular and philosophical essays. University of Chicago Press, Chicago, IL, USA, 1995. ISBN 0-226-32658-6, 0-226-32659-4 (paperback). xviii + 418 pp. LCCN Q174.8 .H45 1995. URL http://www.

loc.gov/catdir/
description/uchi051/
95012217.html;   http://
www.loc.gov/catdir/toc/
uchi051/95012217.html.

Vickers:1984:OSM

[Vic84] Brian Vickers, editor. Occult and scientific mentalities in the Renaissance. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25879-0, 0-521-33836-0 (paperback). xiv + 408 pp. LCCN H7.5VIC; BF1429 OCC 1984.

Vickers:1987:ESB

[Vic87] Brian Vickers, editor. English science, Bacon to Newton. Cambridge English prose texts. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-30408-3, 0-521-31683-9 (paperback). xii + 244 pp. LCCN Q127.G4 E54

1987. URL http://www.loc.
gov/catdir/enhancements/
fy0642/86019265-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0642/86019265-t.html.

Vickers:1988:BRG

[Vic88]  Brian Vickers.  Book re-

view: Graham Rees assisted by Christopher Upton. Francis Bacon’s Natural Philosophy: A New Source. A transcription of manuscript Hardwick 72A with translation and commentary. Chalfont St Giles, Bucks.: British Society for the History of Science, 1984. £7.90. British Journal for the History of Science, 21(2):256–257, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026987.

Viete:1983:AAN

[Viè83] François Viète. The analytic art: nine studies in algebra, geometry, and trigonometry from the Opus restitutae mathematicae analyseos, seu, Algebrâ novâ. Kent State University Press, Kent, OH, USA, 1983. ISBN 0-87338-282-X. 450 pp. LCCN QA33 .V4813 1983.

Vignoles:1982:CBV

[Vig82] Keith H. Vignoles. Charles Blacker Vignoles, romantic engineer. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-23930-3 (US). xii + 187 pp. LCCN TA140.V6 V59 1982.

vonLeibniz:1976:MNT

[vL76] Freiherr Gottfried Wilhelm von Leibniz. Mathematischer, naturwissenschaftlicher und technischer Briefwechsel. (German) [Mathematical, scientific, and technical correspondence]. Sämtliche Schriften und Briefe / Gottfried Wilhelm Leibniz. Akademie-Verlag, Berlin, Germany, 1976. ISBN 3-05-000235-2 (vol. 2), 3-05-000766-4 (Vol. 3), 3-05-002602-2 (vol. 4), 3-05-003471-8 (vol. 5), 3-05-004116-1 (vol. 6). ???? pp. LCCN ????

vonLeibniz:1988:DMR

[vLMB88] Freiherr Gottfried Wilhelm von Leibniz, R. N. D. (R. Niall D.) Martin, and Stuart C. Brown. Discourse on metaphysics and related writings. Classics of philosophy and science series. Manchester University Press, Manchester, UK, 1988. ISBN 0-7190-2338-6. ix + 170 pp. LCCN B2572.E5 M37 1988. US$25.00 (US$35.00 U.S.).

vanLeeuwenhoek:1982:AVL

[vLP82] Antoni van Leeuwenhoek and L. C. Palm, editors. Antoni van Leeuwenhoeck, sixteen hundred and thirty-two 1632– 1723: studies on the life and work of the Delft scientist commemorating the 350th anniversary of his birthday, volume 8 of Nieuwe Nederlandse bijdragen tot de geschiedenis der geneeskunde en der natuurwetenschappen. Rodopi, Amsterdam and Leiden, The Netherlands, 1982. ISBN 90-6203-824-7. 209 pp. LCCN ????

vonLiebig:1988:JLA

[vLvHH^{+}88] Freiherr Justus von Liebig, August Wilhelm von Hofmann, Emil Heuser, Regine Zott, and Emil Erlenmeyer. Justus von Liebig und August Wilhelm Hofmann in ihren Briefen: Nachträge 1845–1869. Justus von Liebig und Emil Erlenmeyer in ihren Briefen von 1861–1872. (German) [Justus von Liebig and August Wilhelm Hofmann in their Letters: Appendix 1845–1869. Justus von Liebig und Emil Erlenmeyer in their letters of 1861–1872]. Bionomica-Verlag, Mannheim, Germany, 1988. ISBN 3-88208-012-4. 54 + 32 pp. LCCN QD22.L7 A4 1988.

VonStaden:1989:HAM

[VoC89]  Heinrich  Von

Staden and Herophilus of Chalcedon. Herophilus: the art of medicine in early Alexandria: edition, translation and essays. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-23646-0. xliii + 666 pp. LCCN R126.H3/; 0120; N.i.224; C.Gr.H.313.

Meyenn:1988:BRG

[Von88] Karl Von Meyenn. Book review: Gereon Wolters. Mach I, Mach II, Einstein und ihre Folgen. Berlin und New York: de Gruyter, 1987. XV + 474 Seiten. ISBN 3-11-010825-9 DM 188.00. British Journal for the History of Science, 21 (3):363–365, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026835.

vonPufendorf:1996:GWG

[vPSBD^{+}04] Samuel Freiherr von Pufendorf, Wilhelm Schmidt-Biggemann, Detlef Döring, Gerald Hartung, and Frank Böhling. Gesammelte Werke. (German) [Collected works]. Akademie-Verlag, Berlin, Germany, 1996– 2004. ISBN 3-05-002887-4 (set), 3-05-001920-4 (vol. 1), 3-05-002880-7 (vol. 2), 3-05-003441-6 (vol. 3), 3-05-003181-6 (vol. 4), 3-05-003605-2 (vol. 5). ???? pp. LCCN B2603 .A3 1996.

vonProschwitz:1990:BCE

[vPvP90] Gunnar von Proschwitz and Mavis von Proschwitz, editors. Beaumarchais et le Courier de l’Europe: documents inédits ou peu connus. (French) [Beaumarchais and the Courier of Europe: unknown or little-known documents], volume 273–274 of Studies on Voltaire and the eighteenth century, 0435-2866. Voltaire Foundation, Oxford, UK, 1990. ISBN 0-7294-0397-1 (set), 0-7294-0398-X (vol. 1), 0-7294-0399-8 (vol. 2). xxvii + 1289 pp. LCCN PQ2105.A2 S8 vol. 273–274.

vanRuler:1995:CCV

[vR95] J. A. van Ruler. The crisis of causality: Voetius and Descartes on God, nature, and change, volume 66 of Brill’s studies in intellectual history, 0920-8607. E. J. Brill, Leiden, The Netherlands, 1995. ISBN 90-04-10371-6. xii + 353 pp. LCCN B4000.V64 R85 1995.

Vucinich:1984:EKA

[Vuc84] Alexander Vucinich. Empire of knowledge: the Academy of Sciences of the USSR (1917– 1970). University of California Press, Berkeley, CA, USA, 1984. ISBN 0-520-04871-7. x + 484 pp. LCCN AS262.A68 V79 1984.

Wainwright:1990:MCS

[Wai90] Milton Wainwright. Miracle cure: the story of penicillin and the golden age of antibiotics. Blackwell, Oxford, UK, 1990. ISBN 0-631-16492-8. xi + 196 pp. LCCN RM267 .W24 1990. US$30.00 (U.S. : est.).

Wallace:1969:CES

[Wal69] William A. Wallace. The “calculatores” in early Sixteenthcentury physics. British Journal for the History of Science, 4(3):221–232, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025472.

Wallis:1974:BRG

[Wal74] Peter Wallis. Book review: General Natural Science Books in English 1600–1900. By David M. Knight. London: Batsford, 1972. Pp. x + 262. £6.50. British Journal for the History of Science, 7(1): 81–82, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025180.

Wallis:1977:BRG

[Wal77] P. J. Wallis. Book review: General Sale Catalogues of Libraries of Eminent Persons.

Vol. II Scientists. General Editor: A. N. L. Munby. Edited with introduction by H. A. Feisenberger. London: Mansell with Sotheby Parke Bernet Publications, 1975. Pp. 296. £11.50. British Journal for the History of Science, 10(1): 71–73, March 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025586.

Walker:1978:K

[Wal78]  Ralph  Charles  Suther-

land Walker. Kant. Arguments of the philosophers. Routledge and K. Paul, London, UK, 1978. ISBN 0-7100-0009-X (paperback), 0-7100-8994-5. xii + 201 pp. LCCN B2798.

Wallace:1981:PGE

[Wal81a] William A. Wallace. Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought, volume 62 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1981. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346. xvi + 369 pp. LCCN Q174 .B67

vol. 62. URL https://link.
springer.com/book/10.
1007/978-94-009-8404-2.

Walton:1981:BRN

[Wal81b] John K. Walton. Book review: Nineteenth Century:

Museums of Madness: the Social Organization of Insanity in Nineteenth-Century England. By Andrew T. Scull. London: Allen Lane, 1979. Pp.

275. £8.50. British Journal for the History of Science, 14 (1):94–96, March 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026082.

Wallace:1984:GHS

[Wal84] William A. Wallace. Galileo and his sources: the heritage of the collegio romano in Galileo’s science. Princeton University Press, Princeton, NJ, USA, 1984. ISBN 0-691-08355-X. xiv + 371 pp. LCCN QB36.G2 W35 1984. US$42.50.

Walker:1988:SJE

[Wal88] Margot Walker. Sir James Edward Smith, M.D., F.R.S., P.L.S., 1759–1828: first president of the Linnean Society of London. Linnean Society of London, London, UK, 1988. ISBN 0-9506207-1-8. 60 pp. LCCN QK31.S64 W34 1988.

Wallace:1989:BRP

[Wal89] William A. Wallace. Book review: Peter Dear. Mersenne and the Learning of the Schools. Ithaca and London: Cornell University Press, 1988. Pp. xiii + 264. ISBN 0-8014-1875-5. $28.50. British Journal for the History of Science, 22(4):465–467, December 1989.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026941.

Walter:1990:SCC

[Wal90] Maila L. Walter. Science and cultural crisis: an intellectual biography of Percy Williams Bridgman (1882–1961). Stanford University Press, Stanford, CA, USA, 1990. ISBN 0-8047-1796-6. x + 362 pp. LCCN QC16.B73 W35 1990. US$39.50

(est.). URL http://www.loc.
gov/catdir/toc/cam023/
90030511.html.

Wali:1991:CBC

[Wal91] K. C. (Kameshwar C.) Wali.

Chandra: a biography of S. Chandrasekhar. University of Chicago Press, Chicago, IL, USA, 1991. ISBN 0-226-87054-5, 0-226-87055-3 (paperback). x + 341 + 32 pp. LCCN QB36.C46 W35

1991. URL http://www.loc.
gov/catdir/description/
uchi052/90010845.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/90010845-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/90010845-b.html.

Wallace:1992:GLD

[Wal92] William A. Wallace. Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics, volume 137 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346. xxiii + 323 pp. LCCN Q174 .B67

vol. 137. URL https://link.
springer.com/book/10.
1007/978-94-015-8040-3.

Warner:1986:TPM

[War86] John Harley Warner. The therapeutic perspective: medical practice, knowledge, and identity in America, 1820– 1885. Harvard University Press, Cambridge, MA, USA, 1986. ISBN 0-674-88330-6. x + 367 pp. LCCN RM47.U6 W37 1986.

Wardy:1988:BRD

[War88a] Robert Wardy. Book review: David Furley. The Greek Cosmologists: vol. 1: The Formation of the Atomic Theory and its Earliest Critics. Cambridge: Cambridge University Press, 1987. Pp. viii + 220. ISBN 0-521-33328-8. £25.00. British Journal for the History of Science, 21(1): 132–133, March 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026888.

Warner:1988:BRH

[War88b] Deborah Jean Warner. Book review: Helen Wright. James Lick’s Monument. Cambridge: Cambridge University Press, 1987. Pp. xvi + 231. ISBN 0-521-32105-0. £25.00, $32.50. British Journal for the History of Science, 21(1):117– 118, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026873.

Warner:1989:BRF

[War89] John Harley Warner. Book review: F. G. Gosling. Before Freud: Neurasthenia and the American Medical Community, 1870–1910. Urbana and Chicago: University of Illinois Press, 1987. Pp. xviii + 192. ISBN 0-252-01406-5. $25.00. British Journal for the History of Science, 22(1): 81–83, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026682.

Wardy:1990:CCS

[War90] Robert Wardy. The chain of change: a study of Aristotle’s Physics VII. Cambridge classical studies. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37327-1. x + 345 pp. LCCN Q151 .W36

1990. URL http://www.loc.
gov/catdir/description/
cam023/89009804.html;
http://www.loc.gov/
catdir/toc/cam023/
89009804.html.

Wasserstein:1969:BRS

[Was69]  A. Wasserstein.  Book re-

view: Science and Medicine in Classical Antiquity Pythagoras and Early Pythagoreanism. By J. A. Philip. University of Toronto Press and Oxford University Press. 1966. Pp. x + 222. 52s. British Journal for the History of Science, 4 (3):296–298, June 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025489.

Waters:1968:CB

[Wat68] D. W. Waters. The Cook bicentenary. British Journal for the History of Science, 4(2): 160–165, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025145.

Waters:1973:BRO

[Wat73] D. W. Waters. Book review: Oceanography Scientists and the Sea 1650–1900. By Margaret Deacon. London and New York: Academic Press, 1971. Pp. xvi + 445. £5.50. British Journal for the History of Science, 6(3):318–319, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025453.

Waters:1980:SSH

[Wat80] D. W. Waters. Seamen, scientists, historians, and strategy: Presidential address, 1978. British Journal for the History of Science, 13(3):189– 210, November 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026196.

Watanabe:1984:CTM

[Wat84] Masao Watanabe. The caloric theory of S. L. Metcalfe. British Journal for the History of Science, 17(2):210–213, July 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026553.

Wayman:1987:DOB

[Way87] Patrick A. Wayman. Dunsink Observatory, 1785–1985: a bicentennial history, volume 7 of Historical studies in Irish science and technology. Dublin Institute for Advanced Studies, Dublin, Ireland, 1987. ISBN 0-86027-020-3. xiii + 353 pp. LCCN QB82.I732 D869 1987. US$25.00.

Womersley:1997:EGB

[WBP97]  David  Womer-

sley, J. W. (John Wyon) Burrow, and J. G. A. (John Greville Agard) Pocock, editors. Edward Gibbon: bicentenary essays, volume 355 of Studies on Voltaire and the eighteenth century, 0435-2866. Voltaire Foundation, Oxford, UK, 1997. ISBN 0-7294-0552-4. xii + 429 pp. LCCN PQ2105.A2 S8 vol.

Williams:1982:SHT

[WD82]  Trevor  Illtyd  Williams

and T. K. (Thomas Kingston) Derry. A short history of Twentieth-Century technology c. 1900–c. 1950. Clarendon Press, New York, NY, USA, 1982. ISBN 0-19-858159-9. xix + 411 pp. LCCN T20 .W55 1982. US$25.00. URL http:

                                        //www.loc.gov/catdir/
                                        enhancements/fy0604/
                                        82004362-d.html.

Wightman:1971:RC

[WE71] W. D. P. Wightman and M. P. Earles. Report of Council. British Journal for the History of Science, 5(4):427–429, December 1971. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wightman:1972:RCY

[WE72] W. P. D. Wightman and M. P. Earles. Report of Council for the year 1971-2. British Journal for the History of Science, 6(2):230–232, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025312.

Wear:1978:BRM

[Wea78] Andrew Wear. Book review: Middle Ages Giles of Rome and the Medieval Theory of Conception: A Study of the De formatione corporis humani in utero. By M. Anthony Hewson. London: Athlone Press, 1975. Pp. viii + 268. £10.00. British Journal for the History of Science, 11(2):179–180, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025738.

Wear:1979:SNB

[Wea79] Andrew Wear. Short notices of books: Medicine without doctors: home health care in American history. Edited by Guenter B. Risse, Ronald L. Numbers, and Judith Waltzer Leavitt. New York: Science History Publications, 1977. Pp.

124. $7.95/$4.95 (paperback). British Journal for the History of Science, 12(2):236, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025770.

Wear:1980:BRM

[Wea80] Andrew Wear. Book review:

Middle Ages and Renaissance: Symphorien Champier and the Reception of the Occultist Tradition in Renaissance France. By Brian P. Copenhaver, The Hague, Paris & New York: Mouton, 1978. Pp. 368. DM 92. British Journal for the History of Science, 13(2):166– 168, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025900.

Wear:1983:BRJ

[Wea83] Andrew Wear. Book review:

Journals and Dictionaries J. D. Hunt (ed.), Journal of garden history. London: Taylor and Francis Ltd. £30.00 (libraries)/ £16.00 (individuals) per annum. British Journal for the History of Science, 16(3):314– 315, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026181.

Wear:1984:BRS

[Wea84]  Andrew Wear.  Book re-

view: Science Studies Jonathan Barnes, Jacques Brunschwig, Myles Burnyeat, Malcolm Schofield (eds), Science and Speculation. Cambridge: Cambridge University Press, 1982. Pp. xxiv + 351. ISBN 0-521-24689-X. £25.00. British Journal for the History of Science, 17(3):332–334, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026649.

Weart:1988:NFH

[Wea88] Spencer R. Weart. Nuclear fear: a history of images. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-62835-7. 535 pp. LCCN QC773 .W43 1988.

Weatherall:1990:SCH

[Wea90] M. (Miles) Weatherall. In search of a cure: a history of pharmaceutical discovery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-261747-8 (hardcover). xiv + 298 + 12 pp. LCCN RM45 .W43 1990. US$37.50. URL http://www.

loc.gov/catdir/
enhancements/fy0636/
90007426-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0636/90007426-t.html.

Webster:1965:BRD

[Web65a] C. Webster. Book review: Darwiniana. Essays and Reviews Pertaining to Darwinism, by Asa Gray. Edited by A. Hunter Dupree. Pp. xxiv + 327. Harvard University Press; London: Oxford University Press, 1964. £2. British Journal for the History of Science, 2(3): 270–271, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024953.

Webster:1965:BRO

[Web65b] C. Webster. Book review:

On the Origin of Species. By Charles Darwin. A facsimile of the first edition, with an introduction by Ernst Mayr. Pp. xxvii + 502. Harvard University Press; London: Oxford University Press. 1964. 48s. British Journal for the History of Science, 2(3):269– 270, June 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024952.

Webster:1966:BRH

[Web66a] C. Webster. Book review: History of Biological Sciences and Medicine William Harvey. By Kenneth D. Keele. Pp. xi + 244. Plates. London, Thomas Nelson and Sons, Ltd., 1965. 42s. British Journal for the History of Science, 3(1):94, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025123.

Webster:1966:BRO

[Web66b] C. Webster. Book review: Origins of Mendelism. By Robert C. Olby. Pp. 204; plates and illustrations. London: Constable, 1966. 30s. Foundations of Genetics. By F. A. E. Crew. Pp. xiii + 202; plates and illustrations. Pergamon Press, Oxford, London, 1966. 21s. British Journal for the History of Science, 3(2):197–198, December 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025042.

Webster:1966:BRS

[Web66c] C. Webster. Book review: Sixteenth and Seventeenth Centuries: The Correspondence of Henry Oldenburg, Volume I. By A. Rupert Hall and Marie Boas Hall. Pp. xl + 504. Madison, Wisconsin: University of Wisconsin Press, 1965. $12.50. British Journal for the History of Science, 3(1): 79–80, June 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025108.

Webster:1967:BRH

[Web67a] C. Webster. Book review:

History of Biological Sciences and Medicine The Life of William Harvey. By Sir Geoffrey Keynes. Pp. xviii + 483. Plates. London: Clarendon Press, Oxford University Press. 1966. 90s. British Journal for the History of Science, 3(4):406–407, December 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024976.

Webster:1967:HGO

[Web67b] C. Webster. Harvey’s De Generatione: Its origins and relevance to the theory of circulation. British Journal for the History of Science, 3(3): 262–274, June 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025051.

Webster:1968:BRSa

[Web68a] C. Webster. Book review: Seventeenth Century: Elias Ashmole (1617–1692). His Autobiographical and Historical Notes, his Correspondence, and Other Contemporary Sources Relating to his Life and Work.

Edited, with a biographical introduction by C. H. Josten. Oxford University Press: Clarendon Press. 5 vols. Pp. xx + 2,065. 25 plates. 1966. £18 18s. British Journal for the History of Science, 4(1):72– 73, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024992.

Webster:1968:BRSb

[Web68b] C. Webster. Book review: Sixteenth and Seventeenth Centuries: Science in Utopia. A Mighty Design. By Nell Eurich. Harvard University Press. London: Oxford University Press. 1967. Pp. x + 332. 74s. British Journal for the History of Science, 4(2):192– 193, December 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025165.

Webster:1969:BRSd

[Web69a] C. Webster. Book review: Seventeenth Century: A Bibliography of Sir Thomas Browne, Kt, M.D. By Geoffrey Keynes. Second Edition. London: Oxford University Press: Clarendon Press. 1968. Pp. xv + 293. Illustr. £7 7s. British Journal for the History of Science, 4(4):418–419, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025251.

Webster:1969:BRSc

[Web69b] C. Webster. Book review:

Seventeenth Century: Experimental Philosophy. By Henry Power. Ed. by Marie Boas Hall. Johnson Reprint Corporation, New York and London. 1966. Pp. xxvii + 193. $12.50. British Journal for the History of Science, 4(3):299–300, June 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025491.

Webster:1969:HMD

[Web69c]  C. Webster.  Henry More

and Descartes: Some new sources. British Journal for the History of Science, 4(4): 359–377, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025230.

Webster:1970:BRS

[Web70] C. Webster. Book review: Sixteenth and Seventeenth Centuries: Leonard Rauwolf: Sixteenth-century Physician, Botanist and Traveler. By Karl H. Dannenfeldt. Pp. xi + 321. 4 figs., table. Harvard University Press and Oxford University Press. 1969. 76s. British Journal for the History of Science, 5(1):99–100, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025369.

Webster:1971:BRSa

[Web71a] C. Webster. Book review: Seventeenth Century: The Life of the Honourable Robert Boyle, F.R.S. By R. E. W. Maddison. London: Taylor & Francis. 1969. Pp. xxii + 332. £9.50. British Journal for the History of Science, 5(3): 310–311, June 1971. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025340.

Webster:1971:BRSb

[Web71b] C. Webster. Book review: Sixteenth and Seventeenth Centuries: Science and Change, 1500–1700. By Hugh Kearney. London: Weidenfeld & Nicolson. 1971. Pp. 253. Illus. £1.75. British Journal for the History of Science, 5 (4):416, December 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025402.

Webster:1972:BRSc

[Web72] C. Webster. Book review: Sixteenth and Seventeenth Centuries: Science, Technology & Society in Seventeenth-Century England. By Robert K. Merton. New York: Howard Fertig, 1970. Pp. xxxii + 279. $11. British Journal for the History of Science, 6(1):94, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025275.

Weber:1975:BRN

[Web75] Gay Weber. Book review:

Nineteenth and Twentieth Centuries: Researches into the Physical History of Man. By James Gowles Prichard. Ed. by George W. Stocking Jr., Chicago and London: University of Chicago Press, 1973. Pp. cxliv + 568. £6.55. British Journal for the History of Science, 8(1):85–86, March 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025830.

Weber:1976:BRN

[Web76] Gay Weber. Book review:

Nineteenth and Twentieth Centuries: Francis Galton. The Life and Work of a Victorian Genius. By D. W. Forrest. London: Elek, 1974. Pp. x + 340. £5.50. British Journal for the History of Science, 9 (1):85–86, March 1976. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025720.

Weber:1978:BRA

[Web78] Gay Weber. Book review:

American Science No Other Gods. On Science and American Social Thought. By Charles E. Rosenberg. Baltimore & London: Johns Hopkins University Press, 1976. Pp. xiii + 273. £9.45. British Journal for the History of Science, 11 (2):175–176, July 1978. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025734.

Weber:1984:BRN

[Web84] Gay Weber. Book review: National Traditions in Science Ian Langham, The building of British social anthropology: W. H. R. Rivers and his Cambridge disciples in the development of kinship studies, 1898–1931. Dordrecht, Boston & London: D. Reidel, 1981. Pp. xxxii + 392. Dfl. 145/$72.50; Dfl. 65/ $29.50 (paperback). British Journal for the History of Science, 17(1):96, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026031.

Webster:1987:BRJ

[Web87] Charles Webster. Book review: John V. Pickstone. Medicine and Industrial Society. A History of Hospital Development in Manchester and its Region, 1752– 1946. Manchester and Dover (New Hampshire): Manchester University Press, 1985. Pp. xii + 369. ISBN 0-7190-1809-0. £35.00. British Journal for the History of Science, 20(1): 84–85, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026442.

Webster:1988:PHC

[Web88] Charles Webster. Problems of health care: the National Health Service before 1957, volume 1 of The Health services since the war; Peacetime history. Her Majesty’s Stationary Office, London, England, 1988. ISBN 0-11-630942-3. x + 479 pp. LCCN RA412.5.G7 W43 1988.

Webster:1998:NHS

[Web98] Charles Webster. The National Health Service: a political history. Oxford paperbacks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-289296-7. xiii + 241 pp. LCCN RA413.5.G7 W43 1998.

Weisz:1977:PP

[Wei77] G. Weisz. Prophets and patrons. British Journal for the History of Science, 10(2): 151–155, July 1977. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025868.

Weindling:1980:SSH

[Wei80] Paul Weindling. Science and sedition: How effective were the acts licensing lectures and meetings, 1795–1819? British Journal for the History of Science, 13(2):139–153, July 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025892. See reply

[Ink81].

Weiss:1982:MTM

[Wei82] John Hubbel Weiss. The making of technological man: the social origins of French engineering education. MIT Press, Cambridge, MA, USA, 1982. ISBN 0-262-23112-3. xviii + 377 pp. LCCN T121 .W44.

Weindling:1983:BRM

[Wei83a] Paul Weindling. Book review: Medical Sciences: Denyse Rockey, Speech disorder in nineteenth century Britain: the history of stuttering. London: Groom Helm 1980. Pp. 280. £19.95. British Journal for the History of Science, 16(1):91–92, March 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026101.

Weisz:1983:EMU

[Wei83b] George Weisz. The emergence of modern universities in France, 1863–1914. Princeton University Press, Princeton, NJ, USA, 1983. ISBN 0-691-05375-8. xii + 397 pp. LCCN LA698 .W44 1983. US$35.00.

Weindling:1984:DOP

[Wei84] Paul Weindling. Darwinism and the origin of psychoanalysis. British Journal for the History of Science, 17(1):64– 67, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026018.

Weindling:1985:BRI

[Wei85a] Paul Weindling. Book review:

Ignaz Semmelweis, The Etiology, Concept, and Prophylaxis of Childbed Fever. Translated and Edited with an Introduction by K. Codell Carter. Madison and London: The University of Wisconsin Press, 1983. Pp. xii + 262. $35 cloth, $10.95 paper. British Journal for the History of Science, 18(1):120, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026299.

Weindling:1985:BRN

[Wei85b] Paul Weindling. Book review:

Nancy Mandelker Frieden, Russian Physicians in an Era of Reform and Revolution 1856– 1905. Princeton, New Jersey: Princeton University Press, 1981. Pp. xvii + 378. ISBN 0-691-05335-0. £22.80. British Journal for the History of Science, 18(1):86, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026270.

Weindling:1986:BRE

[Wei86] Paul Weindling. Book review:

Ernst Bäumler. Paul Ehrlich. Scientist for Life. Translated by Grant Edwards. New York and London: Holmes and Meier, 1984. Pp. xvi + 288. ISBN 0-8419-0837-0. British Journal for the History of Science, 19 (2):225–226, July 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026609.

Weindling:1989:SGE

[Wei89] Paul Weindling. The ‘Sonderweg’ of German eugenics: Nationalism and scientific internationalism. British Journal for the History of Science, 22(3): 321–333, September 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026899.

Weindling:1991:DSD

[Wei91] Paul Weindling. Darwinism and social Darwinism in imperial Germany: the contribution of the cell biologist, Oscar Hertwig, 1849–1922, volume 3 of Forschungen zur neueren Medizin- und Biologiegeschichte, 0176-330X. G. Fischer, Stuttgart, Germany, 1991. ISBN 3-437-11305-4. 355 pp. LCCN QH31.H38 W38 1991. DM148.00.

Weintraub:1992:THGa

[Wei92] E. Roy (Eliot Roy) Weintraub, editor. Towards a history of game theory: annual supplement to volume 24, History of political economy, volume 24 of History of political economy. Duke University Press, Durham, NC, USA, 1992. ISBN 0-8223-1253-0. ISSN 0018-2702. vi + 306 pp. LCCN HB1 .H55 suppl. vol. 24 1992; HB144 .T68 1992.

Wells:1967:GIB

[Wel67]  George A. Wells.  Goethe

and the intermaxillary bone. British Journal for the History of Science, 3(4):348–361, December 1967. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024960.

Wells:1970:BRO

[Wel70]  G. A. Wells.  Book re-

view: Optics Goethe’s Theory of Colours. Translated by C. L. Eastlake. London: F. Cass. 1967. Pp. xlviii + 428. £6 6s. Goethe’s Theory of Colours, with Introduction by Deane B. Judd. London and Cambridge, Mass.: M.I.T. Press. 1970. Pp. lxii + 423. 93s. British Journal for the History of Science, 5(2):192–194, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025425.

Werskey:1975:SIS

[Wer75] Gary Werskey. Science and ideology in the Soviet Union. British Journal for the History of Science, 8(3):240–245, November 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025559.

Westfall:1962:FNP

[Wes62] Richard S. Westfall. The foundations of Newton’s philosophy of nature. British Journal for the History of Science, 1(2): 171–182, December 1962. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025131.

Westfall:1967:HLU

[Wes67] Richard S. Westfall. Hooke and the Law of Universal Gravitation: A reappraisal of a reappraisal. British Journal for the History of Science, 3 (3):245–261, June 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025050.

Westman:1970:BRS

[Wes70] Robert S. Westman. Book review: Sixteenth and Seventeenth Centuries: Bibliographia Kepleriana, ein Führer durch das gedruckte Schrifttum von Johannes Kepler. Im Auftrag der Bayerischen Akademie des Wissenschaften unter Mitarbeit von Ludwig Rothenfelder. Herausgegeben von Max Caspar. Zweite Auflage besorgt von Martha List. Munich: C. H. Beck’sche Verlagsbuchhandlung. 1968. Pp. xv + 181. 86 facs. 65 D.M. British Journal for the History of Science, 5(2):204–205, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025435.

Westfall:1974:BRS

[Wes74] Richard S. Westfall. Book review: Sixteenth and Seventeenth Centuries: Religion and the Rise of Modern Science. By R. Hooykaas. Edinburgh and London: Scottish Academic Press and Chatto and Windus, 1972. Pp. xiv + 162. £2.25; £1.25 (paperback). British Journal for the History of Science, 7(1):89–90, March 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025186.

Westfall:1979:NEH

[Wes79] Richard S. Westfall. Newton enthroned: His correspondence from 1718 to 1727. British Journal for the History of Science, 12(2):197–200, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025753.

West:1988:IRT

[Wes88] Graham West. Innovation and the rise of the tunnelling industry. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-33512-4. xv + 355 pp. LCCN TA803 .W47

1988. URL http://www.loc.
gov/catdir/description/
cam023/87008089.html;
http://www.loc.gov/
catdir/toc/cam028/
87008089.html.

Westfall:1989:ETG

[Wes89] Richard S. Westfall. Essays on the trial of Galileo, volume 1, no. 5 of Vatican Observatory publications. Special series. Studi Galileiani. Vatican Observatory, Vatican City State, 1989. ISBN 0-268-00923-6. xi + 103 pp. LCCN QB36.G2 W47 1989.

Weyer:1992:GWI

[Wey92] Jost Weyer. Graf Wolfgang II von Hohenlohe und die Alchemie: alchemistische Studien in Schloss Weikersheim, 1587–1610. (German) [Count Wolfgang II of Hohenlohe and Alchemy: alchemical studies in Weikersheim Castle], volume 39 of Forschungen aus Württembergisch Franken. J. Thorbecke, Sigmaringen, Germany, 1992. ISBN 3-7995-7639-8 (J. Thorbecke), 3-921429-39-0 (Historischer Verein für Württembergisch Franken). 516 pp. LCCN QD24.H64 W49 1992.

Wood:1981:CVL

[WG81] P. B. Wood and J. V. Golinski. Collections VIII: Library and archive resources in the history of science and medicine at the University of Leeds. British Journal for the History of Science, 14(3):263– 281, November 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025983.

Whiffen:1991:RSC

[WH91] D. H. (David Hardy) Whiffen and D. H. (Donald Holroyde) Hey. The Royal Society of Chemistry: the first 150 years. Royal Society of Chemistry, London, UK, 1991. ISBN 0-85186-294-2. viii + 266 pp. LCCN M91.E02912.

Wheaton:1983:TSE

[Whe83] Bruce R. Wheaton. The tiger and the shark: empirical roots of wave-particle dualism. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25098-6. xxiv + 355 pp. LCCN QC476.W38 W45

1983. URL http://www.loc.
gov/catdir/description/
cam022/82022069.html;
http://www.loc.gov/
catdir/toc/cam023/
82022069.html.

Wheaton:1989:BRG

[Whe89] Bruce R. Wheaton. Book review: T. Glick (ed.). The Comparative Reception of Relativity. Dordrecht/Boston: D. Reidel, 1987. Pp. vii + 412. ISBN 90-277-2498-9. £54.50, $79.00. British Journal for the History of Science, 22(1): 83–84, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026683.

Whiteside:1964:NET

[Whi64] Derek T. Whiteside. Newton’s early thoughts on planetary motion: A fresh look. British Journal for the History of Science, 2(2):117–137, December 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025011.

Whitrow:1965:ER

[Whi65] G. J. Whitrow. Essay reviews. British Journal for the History of Science, 2(4):356– 358, December 1965. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024893.

Whiteside:1968:BRH

[Whi68a] D. T. Whiteside. Book review: History of Mathematics Studien zur Mathematik des John Wallis (1616–1703): Winkelteilungen, Kombinationslehre und zahlentheoretische Probleme im Anhang die Bücher und Handschriften von Wallis. By C. J. Scriba (Boethius, Vol. 6). Pp. xii + 144. 4 plates. Wiesbaden: Franz Steiner Verlag, 1966. DM. 32. British Journal for the History of Science, 4(1):84–85, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025001.

Whitrow:1968:BRT

[Whi68b] G. J. Whitrow. Book review: Twentieth Century Sir Arthur Eddington. By C. W. Kilmister. Pp. vi + 279. The Commonwealth Library: Selected Readings in Physics. Oxford: Pergamon Press. 1966. 21s. British Journal for the History of Science, 4(1):81– 82, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4024999.

Whitteridge:1968:BRS

[Whi68c] Gweneth Whitteridge. Book review: Seventeenth Century: William Harvey’s Biological Ideas. Selected aspects and historical background. By Walter Pagel. Pp. 394. 44 plates. Basel/New York: S. Karger, 1967. £6 2s. 6d. British Journal for the History of Science, 4(1):68–69, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024989.

White:1970:BRG

[Whi70a] George W. White. Book review: Geology The Hydrologic Cycle and the Wisdom of God, A Theme in Geoteleology. By Yi-Fu Tuan. University of Toronto Press and Oxford University Press. Pp. xiii + 160, 9 figs. 1968. 47s. 6d. British Journal for the History of Science, 5(2):196–197, December 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025428.

Whiteside:1970:EHE

[Whi70b] D. T. Whiteside. Essays in the history of the exact sciences. British Journal for the History of Science, 5(1): 87–88, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025356.

Whitrow:1971:LM

[Whi71] G. J. Whitrow. The laws of motion. British Journal for the History of Science, 5 (3):217–234, June 1971. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025314.

Whiteside:1972:BRM

[Whi72a] D. T. Whiteside. Book review: Mathematics: The Usefulness of Mathematical Learning explained and demonstrated: Being Mathematical Lectures read in the Publick Schools at the University. … Translated by… John Kirkby (London, 1734). By Isaac Barrow. Reprint. London: Frank Cass, 1970. Pp. xxxii + 456. £7.35. British Journal for the History of Science, 6(1):

86–88, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025268.

Whiteside:1972:BRS

[Whi72b] D. T. Whiteside. Book review: Sixteenth and Seventeenth Centuries: Force in Newton’s Physics. The Science of Dynamics in the Seventeenth Century. By Richard S. Westfall. London: Macdonald, and New York: American Elsevier, 1971. Pp. xii + 579. £10. British Journal for the History of Science, 6(2):217– 218, December 1972. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025305.

Whiteside:1973:BRN

[Whi73] D. T. Whiteside. Book review: Newtonian Studies Isaac Newton’s Philosophiae Naturalis Principia Mathematica. The Third Edition (1726) with Variant Readings. Ed. by Alexandre Koyré and I. Bernard Cohen, with the assistance of Anne Whitman. 2 volumes. Cambridge: Cambridge University Press, and Cambridge, Mass: Harvard University Press, 1972. Pp. xlii + 916. £25. British Journal for the History of Science, 6(4): 445–447, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025517.

Whiteside:1975:BRS

[Whi75] D. T. Whiteside. Book review: Seventeenth Century: Leibniz in Paris, 1672–1676. His growth to mathematical maturity. By Joseph E. Hofmann. Cambridge: Cambridge University Press, 1974. Pp. xi + 372. £8.50. British Journal for the History of Science, 8(3): 258–260, November 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025570.

Whitteridge:1977:BRS

[Whi77] Gweneth Whitteridge. Book review: Seventeenth Century: Medicine in Seventeenth Century England. A Symposium held at UCLA in honor of C. D. O’Malley. Ed. by Allen G. Debus. Berkeley, Los Angeles, and London: University of California Press, 1974. Pp. xiii + 485. £8.50. British Journal for the History of Science, 10 (2):171–173, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025879.

Whitrow:1984:BRS

[Whi84] G. J. Whitrow. Book review: Science Studies P. T. Landsberg (ed.). The Enigma of Time. Bristol: Adam Hilger, 1982. pp. xii + 248. ISBN 0-85274-545-1. £13.95, $28.00. British Journal for the History of Science, 17(3):334– 335, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026650.

Whiteside:1985:BRF

[Whi85a] D. T. Whiteside. Book review:

François Viète, The Analytic Art. Nine Studies in Algebra, Geometry and Trigonometry from the ‘Opus Restitutae Mathematics Analyseos, seu Algebra Nova’. Edited by T. Richard Witmer. Kent, Ohio: State University Press [European distributor: Eurospan Ltd., 3 Henrietta Street, London WC2E] 1983. Pp. 450. ISBN 0-87338-282-X. $45. British Journal for the History of Science, 18 (1):98–100, March 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026281.

Whitrow:1985:BRR

[Whi85b] G. J. Whitrow. Book review:

Richard Sorabji, Time, Creation and the Continuum: theories in antiquity and the early middle ages. London: Duckworth, 1983. Pp. xviii + 473. ISBN 0-7156-1693-5. £29.50. British Journal for the History of Science, 18(3):349– 351, November 1985. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026391.

Whitrow:1988:BRC

[Whi88a] G. J. Whitrow. Book review: C. Ray. The Evolution of Relativity. Bristol and Philadelphia: Adam Hilger, 1987. Pp. x + 211. ISBN 0-85274-423-4. £29.50. British Journal for the History of Science, 21(3): 366–367, September 1988. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026838.

Whitrow:1988:THE

[Whi88b] G. J. Whitrow. Time in history: the evolution of our general awareness of time and temporal perspective. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. ISBN 0-19-215361-7. x + 217 pp. LCCN D21.3 .W47 1988. US$22.00. URL http://www.

                                        loc.gov/catdir/
                                        enhancements/fy0640/
                                        88004245-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0726/88004245-b.html.

Whitley:1989:BRH

[Whi89a] R. D. Whitley. Book review: Harry Redner. The Ends of Science: An Essay in Scientific Authority. Boulder, Colorado: Westview Press, 1987. Pp. xiv + 344. ISBN 978-0-8133-0452-6; $39.50. British Journal for the History of Science, 22(1): 106–108, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026705.

Whitrow:1989:BRH

[Whi89b] G. J. Whitrow. Book review: N. S. Hetherington. Science and Objectivity: Episodes in the History of Astronomy. Ames, Iowa State University Press, 1988. Pp. xii + 168. ISBN 0-8138-1159-7. No price given. British Journal for the History of Science, 22(4):448– 449, December 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026926.

Whitney:1990:PRM

[Whi90]  Elspeth  Whitney.

Paradise restored: the mechanical arts from Antiquity through the Thirteenth Century, volume 80:1 of Transactions of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1990. ISBN 0-87169-801-3. ISSN 0065-9746. 169 pp. LCCN T15 .W57 1990; Q11 .P6 n.s. vol. 80 pt.1.

Whitrow:1993:JWJ

[Whi93] Magda Whitrow. Julius von Wagner-Jauregg: (1857–1940). Smith-Gordon and Co., London, UK, 1993. ISBN 1-85463-012-1. xxiv + 221 + 8 pp. LCCN RC339.52.W34 W45

                                        1993. URL http://catalog.
                                        hathitrust.org/api/
                                        volumes/oclc/29571664.
                                        html.

Whorton:1982:CFH

[Who82] James C. Whorton. Crusaders for fitness: the history of American health reformers. Princeton University Press, Princeton, NJ, USA, 1982. ISBN 0-691-04694-8. 359 + 4 pp. LCCN RA418.3.U6 W5 1982. US$19.50.

Whiteside:1992:IDT

[WHS92] D. T. (Derek Thomas) Whiteside, P. M. (Peter Michael) Harman, and Alan E. (Alan Elihu) Shapiro, editors. The investigation of difficult things: essays on Newton and the history of the exact sciences in honour of D. T. Whiteside. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-37435-9. xvi + 531 pp. LCCN QC16.N7 I58

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        cam025/92248567.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        92248567.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam028/92248567.html.

Wiener:1993:ICF

[Wie93] Norbert Wiener. Invention: the care and feeding of ideas. MIT Press, Cambridge, MA, USA, 1993. ISBN 0-262-23167-0. xxiv + 159 pp. LCCN T15 .W65 1993.

Wightman:1969:BRD

[Wig69] William P. D. Wightman. Book review: Dictionaries of Scientists A Biographical Dictionary of Scientists. Ed. by Trevor I. Williams. London: Adam & Charles Black. 1969. Pp. xi + 592. £5. British Journal for the History of Science, 4(4): 403–404, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025235.

Wightman:1973:TA

[Wig73] William P. D. Wightman. The tyranny of abstractions. British Journal for the History of Science, 6(3):233–246, June 1973. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025443.

Wightman:1974:BRG

[Wig74] William P. D. Wightman. Book review: General Science, Medicine and Society in the Renaissance. Essays to Honor Walter Pagel. Ed. by Allen G. Debus. London: Heinemann, 1972. 2 Vols. Pp. 275; 338. £12.00. British Journal for the History of Science, 7(2):183–184, July 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025208.

Wightman:1977:BRM

[Wig77] William P. D. Wightman. Book review: Medicine Public Health and the Medical Profession in the Renaissance. By Carlo Cipolla. London: Cambridge University Press, 1976. Pp. viii + 136. £5.50. British Journal for the History of Science, 10 (1):74–75, March 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025588.

Wightman:1978:BRR

[Wig78] William P. D. Wightman. Book review: Renaissance: Essays on the Life and Works of Thomas Linacre, c. 1460–1524. Edited by Francis Maddison, Margaret Peiling, and Charles Webster. Oxford: Clarendon Press: Oxford University Press, 1977. Pp. liii + 416. £12.00. British Journal for the History of Science, 11(3):290–292, November 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025789.

Wilkie:1962:SRR

[Wil62] J. S. Wilkie. Some reasons for the rediscovery and appreciation of Mendel’s work in the first years of the present century. British Journal for the History of Science, 1(1): 5–17, June 1962. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025071.

Wilkie:1965:BRS

[Wil65a] J. S. Wilkie. Book review: A Short History of Biology. By Isaac Asimov. London and Edinburgh: Nelson, 1965. Pp. 180. 21s. net. British Journal for the History of Science, 2(4):364, December 1965. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Williams:1965:BRL

[Wil65b] Trevor I. Williams. Book review: The Lunar Society of Birmingham; A Social History of Provincial Science and Industry in Eighteenth-century England. By Robert E. Schofield. Pp. x + 491. Oxford: Clarendon Press, 1963. 70s. net. British Journal for the History of Science, 2(4):361–362, December 1965. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4024897.

Williams:1968:BRN

[Wil68] L. Pearce Williams. Book review: Nineteenth Century: The Contributions of Faraday and Maxwell to Electrical Science. By R. A. R. Tricker. Pp. ix + 289. 6 plates. Oxford: Pergamon Press. 1966. 25s. British Journal for the History of Science, 4(1):78–79, June 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4024996.

Williams:1969:BRD

[Wil69] Trevor I. Williams. Book review: Dictionaries of Scientists World Who’s Who in Science: A Biographical Dictionary of Notable Scientists from Antiquity to the Present. Ed. by Allen G. Debus. Chicago: Marquis-Who’s Who Inc. Pp. xvi + 1,855. 1968. $51.00 net. British Journal for the History of Science, 4(4):404– 405, December 1969. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025236.

Wilson:1972:GGS

[Wil72] David B. Wilson. George Gabriel Stokes on stellar aberration and the luminiferous ether. British Journal for the History of Science, 6(1): 57–72, June 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025263.

Wilkie:1975:BRM

[Wil75] J. S. Wilkie. Book review: Medicine Galenism: Rise and Decline of a Medical Philosophy. By Owsei Temkin. Ithaca and London: Cornell University Press, 1973. Pp. xvii + 240. No price stated. British Journal for the History of Science, 8(1):73, March 1975. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025819.

Williams:1977:BRA

[Wil77] A. R. Williams. Book review: Antiquity Le Problème des métaux dans la science antique. By Robert Halleux. Paris: Société d’Edition ‘Les Belles Lettres’, 1974, Pp. xvi + 236. 40 francs. British Journal for the History of Science, 10(2):169–170, July 1977. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025876.

Wilson:1978:BRS

[Wil78] Dudley Wilson. Book review:

Science and Literature Science and the Human Comedy. By Harcourt Brown. Toronto: University of Toronto Press, 1976. Pp. xx + 224. £9.75. British Journal for the History of Science, 11(2):176–177, July 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025735.

Williams:1981:HBG

[Wil81a] Trevor Illtyd Williams. A history of the British gas industry. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1981. ISBN 0-19-858157-2. xvii + 304 + 16 pp. LCCN HD9581.G72 W54.

Wilsonp:1981:DP

[Wil81b] David B. Wilsonp. Darwin and the protestants. British Journal for the History of Science, 14(2):200–202, July 1981. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025948.

Wilde:1982:MSN

[Wil82a] C. B. Wilde. Matter and spirit as natural symbols in eighteenth-century British natural philosophy. British Journal for the History of Science, 15(2):99–131, July 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4025963.

Wilson:1982:D

[Wil82b]  Margaret  Dauler

Wilson. Descartes. The Arguments of the philosophers. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1982. ISBN 0-7100-9208-3 (paperback). xvii + 255 pp. LCCN B1875 .W58 1982.

Wilson:1983:RSG

[Wil83] David Wilson. Rutherford, simple genius. Hodder and Stoughton, London, UK, 1983. ISBN 0-340-23805-4. 638 + 8 pp. LCCN QC16.R8 W54 1983. US$14.95. URL http:

                                        //catalog.hathitrust.org/
                                        api/volumes/oclc/
                                        10560773.html.

Williams:1984:BPE

[Wil84a] Mari Williams. Beyond the planets: early nineteenth-century studies of double stars. British Journal for the History of Science, 17(3):295–309, November 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026624.

Williams:1984:BRN

[Wil84b] Mari E. W. Williams. Book review: National Traditions in Science E. G. Forbes, Tobias Mayer (1723–1762). Pioneer of enlightened science in Germany. Göttingen: Vandenhoeck & Ruprecht, 1980. Pp.

248. DM 38. British Journal for the History of Science, 17 (1):89–90, March 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026024.

Williams:1984:BRP

[Wil84c] Mari E. W. Williams. Book review: Physics and its concepts: N. T. Roseveare, Mercury’s perihelion from Le Verrier to Einstein. Oxford: Oxford University Press, 1982. Pp. viii + 208. ISBN 0-19-858174-2. £20.00. British Journal for the History of Science, 17 (2):225–226, July 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026559.

Williams:1984:BRHb

[Wil84d] Roger Williams. Book review: History of Technology: Brian Wynne, Rationality and ritual: the Windscale Inquiry and nuclear decisions in Britain. Chalfont St. Giles, Bucks.: The British Society for the History of Science (BSHS Monographs 3), 1982. Pp. x + 222. ISBN 0-906450-02-0. £6.50, $13.50. British Journal for the History of Science, 17 (3):331, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026647.

Williams:1984:BRHa

[Wil84e] Trevor I. Williams. Book review: History of Technology: William H. McNeill, The Pursuit of Power: Technology Armed Force and Society since AD 1000. Oxford: Basil Blackwell, 1983. ISBN 0-631-13134-5. £15.00. British Journal for the History of Science, 17(3): 324–325, November 1984. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026639.

Williams:1984:HFP

[Wil84f] Trevor Illtyd Williams. Howard Florey, penicillin and after. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1984. ISBN 0-19-858173-4 (US). ???? pp. LCCN RB17.F57 W55 1984.

US$22.95. URL http://www.

loc.gov/catdir/
enhancements/fy0635/
84007845-t.html.

Williams:1985:BRJ

[Wil85a] Mari E. W. Williams. Book review: John Flamsteed, Preface to Historia Coelestis Britannica. Edited and introduced by Allan Chapman, based on a translation by Alison Dione Johnson. Greenwich: National Maritime Museum, Maritime Monographs and Reports, No. 52, 1982. Pp. vi + 222. ISBN 0-905555-60-0. No price given. British Journal for the History of Science, 18(1): 103, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026285.

Williams:1985:BRT

[Wil85b] Mari E. W. Williams. Book review: Trevor I. Williams, A History of the British Gas Industry. Oxford: Oxford University Press, 1981. Pp. xvii + 304. ISBN 0-19-858157-2. £18.50. British Journal for the History of Science, 18(2): 246–247, July 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026347.

Williams:1986:BRB

[Wil86a] L. Pearce Williams. Book review: I. Bernard Cohen.

Revolution in Science. Cambridge, Mass.: Harvard University Press, 1985. Pp. xx + 711. ISBN 0-674-76777-2. £22.75. British Journal for the History of Science, 19(3): 339–342, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026511.

Williams:1986:BRD

[Wil86b] Mari E. W. Williams. Book review: Dieter B. Herrmann. The History of Astronomy from Herschel to Hertzsprung. Translated by Kevin Krisciunas. Cambridge: Cambridge University Press, 1984. Pp. x + 220. ISBN 0-521-257336. £12.50. British Journal for the History of Science, 19(3): 347–348, November 1986. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026518.

Wilcox:1987:MTP

[Wil87a] Donald J. Wilcox. The measure of times past: pre-Newtonian chronologies and the rhetoric of relative time. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-89721-4, 0-226-89722-2 (paperback). ix + 292 pp. LCCN D11 .W59 1987. US$38.00, US$14.00.

URL http://www.loc.gov/
catdir/enhancements/
fy0608/87001669-d.html;
http://www.loc.gov/

catdir/enhancements/
fy0608/87001669-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/87001669-b.html.

Williams:1987:BRW

[Wil87b] Mari Williams. Book review:

Willem Hackmann. Seek and Strike: Sonar, Anti-Submarine Warfare and the Royal Navy, 1914–1954. London: Her Majesty’s Stationery Office, 1984. Pp. xvii + 487. ISBN 0-11-290423-8. £15.95. British Journal for the History of Science, 20(1):84, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026441.

Williams:1987:BRV

[Wil87c] Mari E. W. Williams. Book review: A. van Helden. Measuring the Universe. Cosmic Dimensions from Aristarchus to Halley. Chicago and London: University of Chicago Press, 1985. Pp. viii + 203. ISBN 0-226-84881-7. £31.50, $34.50. British Journal for the History of Science, 20(1): 115–116, January 1987. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026468.

Williams:1987:BRP

[Wil87d] Mari E. W. Williams. Book reviews: P. R. De Clerq (ed.). Nineteenth Century Scientific Instruments and their Makers: Papers Presented at the Fourth Scientific Instrument Symposium, Amsterdam 23-26 October 1984. Boerhaave Museum, Leiden, and Rodopi, Amsterdam, 1985. ISBN 90-6292-070-5 (Boerhaave Museum), ISBN 90-6203-917-0 (Rodopi). Pp.

275. No price given. British Journal for the History of Science, 20(3):358–359, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026369.

Williams:1987:KFA

[Wil87e]  Robert  Chad-

well Williams. Klaus Fuchs, atom spy. Harvard University Press, Cambridge, MA, USA, 1987. ISBN 0-674-50507-7. vi + 267 + 10 pp. LCCN UB271.R9 F838 1987.

Williamson:1987:MP

[Wil87f] Rajkumari Williamson, editor. The making of physicists. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-524-9. viii + 200 pp. LCCN QC47.E85 M35

                                        1987. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0745/87204779-d.html.

Wilson:1987:KSC

[Wil87g] David B. Wilson. Kelvin and Stokes: a comparative study in Victorian physics. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-526-5. xvi + 253 pp. LCCN QC16.K3 W55

1987. URL http://www.loc.
gov/catdir/enhancements/
fy0745/87181742-d.html.

Wilson:1989:BRR

[Wil89a] David B. Wilson. Book review: Robert Kargon & Peter Achinstein (eds). Kelvin’s Baltimore Lectures and Modern Theoretical Physics: Historical and Philosophical Perspectives. Cambridge, Massachusetts and London: The MIT Press, 1987. Pp. xii + 547. ISBN 0-262-11117-9. £35.95. British Journal for the History of Science, 22(1):109–110, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026706.

Wilson:1989:BRJ

[Wil89b] J. F. Wilson. Book review:

J. Liebenau (ed.). The Challenge of New Technology: Innovation in British Business since 1850. Aldershot: Gower, 1988. Pp. x + 155. ISBN 0-566-05147-8. £23.50. British Journal for the History of Science, 22(2):248–249, July 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026667.

Williams:1990:NUE

[Wil90a] Rosalind H. Williams. Notes on the underground: an essay on technology, society, and the imagination. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-23145-X. xi + 265 pp. LCCN TA712 .W55 1990.

Williams:1990:RRC

[Wil90b] Trevor Illtyd Williams. Robert Robinson, chemist extraordinary. Clarendon Press, New York, NY, USA, 1990. ISBN 0-19-858180-7. viii + 201 + 7 pp. LCCN QD22.R63 W54 1990. US$37.00. URL http:

                                        //www.loc.gov/catdir/
                                        enhancements/fy0635/
                                        89022123-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0635/89022123-t.html.

Wilkie:1991:BSP

[Wil91a]  Tom Wilkie.  British sci-

ence and politics since 1945. Making contemporary Britain. Blackwell, Oxford, UK, 1991. ISBN 0-631-16849-4, 0-631-16851-6 (paperback). vii + 142 pp. LCCN Q127.G4 W62 1991. US$39.95, US$18.95.

Wilson:1991:EDS

[Wil91b] Fred Wilson. Empiricism and Darwin’s science, volume 47 of The University of Western Ontario series in philosophy of science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1019-5 (hardcover). xiv + 358 pp. LCCN QH331 .W53 1991.

Willmoth:1993:SJM

[Wil93] Frances Willmoth. Sir Jonas Moore: practical mathematics and restoration science. Boydell Press, Woodbridge, Suffolk, UK, 1993. ISBN 0-85115-321-6. xi + 244 pp. LCCN QA29.M65 W55 1993. US$35 (US$70.00).

Williams:1994:PMH

[Wil94a] Mari E. W. Williams. The precision makers: a history of the instruments industry in Britain and France, 1870–1939. Comparative and international business series. Modern histories. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1994. ISBN 0-415-03732-8. viii + 216 pp. LCCN HD9706.6.G72 W55 1994.

Wilson:1994:MMM

[Wil94b] Adrian Wilson. The making of man-midwifery: childbirth in England, 1660–1770. UCL Press., London, UK, 1994. ISBN 1-85728-292-2. xii + 239 pp. LCCN RG518.G7. URL http://

catalog.hathitrust.org/
api/volumes/oclc/
32910640.html.

Wilson:1995:IWE

[Wil95] Catherine Wilson. The invisible world: early modern philosophy and the invention of the microscope. Studies in intellectual history and the history of philosophy. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03418-4. x + 280 pp. LCCN B67 .W55 1995.

URL http://www.h-net.org/
review/hrev-a0a0i9-aa;

http://www.loc.gov/
catdir/description/
prin031/94024556.html;
http://www.loc.gov/
catdir/toc/prin031/
94024556.html.

Winter:1975:BRA

[Win75] H. J. J. Winter. Book review: Arabic Science: A Testament of Alchemy. Being the Revelations of Morienus to Khalid ibn Yazid. Ed. and trans, by Lee Stavenhagen. Hanover, New Hampshire: University Press of New England, for the Brandeis University Press, 1974. Pp. 76. No price stated. British Journal for the History of Science, 8(3):255–256, November 1975. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025567.

Winsor:1991:RSN

[Win91] Mary P. Winsor. Reading the shape of nature: comparative zoology at the Agassiz Museum. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1991. ISBN 0-226-90214-5 (hardcover), 0-226-90215-3 (paperback). xviii + 324 pp. LCCN QL71.U62 C358

1991. URL http://www.loc.
gov/catdir/enhancements/
fy0608/91008742-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/91008742-t.html;
http://www.loc.gov/

catdir/enhancements/
fy0609/91008742-b.html.

Wise:1985:WRW

[Wis85] George Wise. Willis R. Whitney, General Electric, and the origins of U.S. industrial research. Columbia University Press, New York, NY, USA, 1985. ISBN 0-231-06044-0, 0-231-06016-5. 375 pp. LCCN T40.W45 W57 1985.

Wise:1995:VP

[Wis95] M. Norton Wise, editor. The values of precision. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03759-0 (hardcover), 0-691-01601-1 (paperback). LCCN QC39 .V26 1995. US$49.50. URL http://

www.loc.gov/catdir/
description/prin031/
94019562.html;   http://
www.loc.gov/catdir/toc/
prin031/94019562.html.

Weingart:1988:RBG

[WKB88] Peter Weingart, Jürgen Kroll, and Kurt Bayertz. Rasse, Blut und Gene: Geschichte der Eugenik und Rassenhygiene in Deutschland. (German) [Race, blood and genes: History of eugenics and racial cleanliness in Germany]. Wissenschaftsforschung. Suhrkamp, Frankfurt am Main, Germany, 1988. ISBN 3-518-57887-1, 3-518-57886-3. 746 pp. LCCN

???? URL http://www.gbv.
de/dms/bs/toc/025182471.

Wiener:1985:CSS

[WM85] Norbert Wiener and Pesi Rustom Masani. Cybernetics, science, and society: Ethics, aesthetics, and literary criticism: Book reviews and obituaries, volume 23; vol. 4 of Mathematicians of our time; Collected works with commentaries / Norbert Wiener; Collected works with commentaries. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-23123-9. xx + 1083 pp. LCCN QA3 .W53 vol. 4.

Wolters:1987:MMI

[Wol87]  Gereon Wolters.  Mach I,

Mach II, Einstein und die Relativitätstheorie: eine Fälschung und ihre Folgen. (German) [Mach I, Mach II, Einstein and their consequences]. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 1987. ISBN 3-11-010825-9. xv + 474 s pp. LCCN ???? URL 01;

                                         http://www.zentralblatt-
                                        math.org/zmath/en/search/
                                        ?an=0643.01029.

Wonders:1993:HDI

[Won93] Karen Wonders. Habitat dioramas: illusions of wilderness in museums of natural history, volume nova ser. 25 of Acta Universitatis Upsaliensis. Figura, 0071-481X. Acta Universitatis Upsaliensis, Uppsala, Sweden, 1993. ISBN 91-554-3157-7. 262 pp. LCCN N25 .F5 no. 25.

Woods:1967:BRH

[Woo67] J. A. Woods. Book review: A History of the Royal College of Physicians of London, Volume 2. By Sir George Clark. London: Clarendon Press: Oxford University Press. Pp. xix + 373. 12 plates. 1966. 63s. British Journal for the History of Science, 3(3):299, June 1967. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025060.

Woods:1973:BRS

[Woo73] John A. Woods. Book review:

Science and Society Science in American Society. A Social History. By George H. Daniels. New York: Knopf, 1971. Pp. xii + 390. $10. British Journal for the History of Science, 6(4):438, December 1973. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025511.

Wood:1980:BRA

[Woo80a] P. B. Wood. Book review: Astrology and the Popular Press: English Almanacs, 1500–1800. By Bernard Capp. London and Boston: Faber and Faber, 1979. Pp. 452. £15.00. British Journal for the History of Science, 13(3):265–266, November 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026203.

Wood:1980:BRSa

[Woo80b] P. B. Wood. Book review: Scientific revolution: Hermeticism and the Scientific Revolution. Papers read at a Clark Library Seminar, 9 March 1974. By Robert S. Westman and J. E. McGuire. Los Angeles: William Andrews Clark Memorial Library, 1977. Pp. iv + 150. No price stated. British Journal for the History of Science, 13 (1):70–72, March 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026129.

Wood:1980:BRSb

[Woo80c] P. B. Wood. Book review: Scientific revolution: The Correspondence of Henry Oldenburg. Volume x, June 1673-April 1674. Ed. and trans. by A. Rupert Hall and Marie Boas Hall. London: Mansell, 1975. Pp. xxvii + 596. No price stated. The Correspondence of Henry Oldenburg. Volume xi, May 1674-September 1675. Ed. and trans. by A. Rupert Hall and Marie Boas Hall. London: Mansell, 1977. Pp. xxiv + 543. No price stated. British Journal for the History of Science, 13(1):73–74, March 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026131.

Wood:1980:MAT

[Woo80d] P. B. Wood. Methodology and apologetics: Thomas Sprat’s history of the Royal Society. British Journal for the History of Science, 13(1):1– 26, March 1980. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026117.

Woodward:1980:SNB

[Woo80e] Calvin A. Woodward. Short notices of books: Holistic Thought in Social Science. By D. C. Phillips. London and Basingstoke: Macmillan, 1977. Pp. v + 149. £6.95. British Journal for the History of Science, 13(2):182, July 1980. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025915.

Wood:1982:ESV

[Woo82] P. B. Wood. English scientific virtuosi in the 16th and 17th centuries. British Journal for the History of Science, 15(1):69–71, March 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025924.

Wood:1983:BRM

[Woo83]  P. B. Wood.  Book re-

view: Medical Sciences: Arthur Donovan and Joseph Prentiss, James Hutton’s medical dissertation. Transactions of the American Philosophical Society. Vol. 70, pt. 6. Philadelphia: The American Philosophical Society, 1980. Pp. 57. $8.00. British Journal for the History of Science, 16(1):88– 89, March 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026099.

Wood:1984:BRN

[Woo84] P. B. Wood. Book review: National Traditions in Science R. H. Campbell and Andrew S. Skinner (eds.), The origins and nature of the Scottish enlightenment, Edinburgh: John Donald Publishers Ltd., 1982. Pp. viii + 231. £15.00. British Journal for the History of Science, 17(1):94–95, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026030.

Woolley:1985:BRT

[Woo85] A. R. Woolley. Book review: Tom Simkin and Richard S. Fiske, Krakatau 1883: the volcanic eruption and its effects. Washington, D.C.: Smithsonian Institution Press, 1983. Pp. 464. £13.95. British Journal for the History of Science, 18(1):116–117, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026296.

Wood:1989:BRA

[Woo89] P. B. Wood. Book review:

Alan Charles Kors & Paul J. Korshin (eds). Anticipations of the Enlightenment in England, France, and Germany. Philadelphia: University of Pennsylvania Press, 1987. Pp. viii + 290. ISBN 0-8122-8057-1. No price given. British Journal for the History of Science, 22(1):92–93, March 1989. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026692.

Wood:1992:HIB

[Woo92] James Wood. History of international broadcasting, volume 19< 23 > of IEE history of technology series. P. Peregrinus Ltd. in association with the Science Museum, London, UK, 1992. ISBN 0-86341-281-5, 0-85296-920-1 (vol. 2). ???? pp. LCCN HE8697.4 .W66

1992. URL http://www.loc.
gov/catdir/enhancements/
fy0631/93171088-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0631/93171088-t.html.

Worrall:1979:RDB

[Wor79] John Worrall. A reply to David Bloor: “Polyhedra and the abominations of Leviticus” [British J. Hist. Sci. 11 (1978), no. 39, part 3, 245–272; MR 80f:01027]. British Journal for the History of Science, 12(1):71–81, March 1979. CO-DEN BJHSAT. ISSN 0007- (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4025664. With a re-

sponse by David Bloor and a further reply by Worrall. See [Blo78].

Workman:1988:BRB

[Wor88] Michael E. Workman. Book review: Brooke Hindle & Steven Lubar. Engines of Change: The American Industrial Revolution, 1790–1860. Washington, D.C., and London: Smithsonian Institution Press, 1986. Pp. 285. ISBN 0-87474-540-3. (hardcover)/539-X (paperback). $25.95/514.95 (paperback). British Journal for the History of Science, 21(1): 115, March 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026870.

Wheaton:1992:ISH

[WR92]  Bruce  R.  Wheaton

and Robin E. Rider, editors. Inventory of sources for history of Twentieth-Century physics: report and microfiche index to 700,000 letters. Verlag für Geschichte der Naturwissenschaften und der Technik, Stuttgart, Germany, 1992. ISBN 3-928186-09-4. 62 microfiches (16,035 frames) pp. LCCN QC15 .I68 1992.

Wright:1987:JLM

[Wri87] Helen Wright. James Lick’s monument: the saga of Captain Richard Floyd and the building of the Lick Observatory. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-32105-0. xvi + 231 pp. LCCN QB82.U62 M689

1987. URL http://www.loc.
gov/catdir/description/
cam031/86004179.html;
http://www.loc.gov/
catdir/toc/cam031/
86004179.html.

Wright:1994:MPD

[Wri94] Susan Wright. Molecular politics: developing American and British regulatory policy for genetic engineering, 1972–1982. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-91065-2 (hardcover), 0-226-91066-0 (paperback). xxii + 591 pp. LCCN QH442 .W75

1994. URL http://www.loc.
gov/catdir/description/
uchi051/93047054.html;
http://www.loc.gov/
catdir/toc/uchi051/
93047054.html.

Wallace:1991:ARW

[WS91a] Alfred Russel Wallace and Charles H. (Charles Hyde) Smith. Alfred Russel Wallace: an anthology of his shorter writings. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-857725-7. x + 551 pp. LCCN Q158.5 .W3525 1991. US$62.50 (U.S.). URL http:

//catalog.hathitrust.org/
api/volumes/oclc/
22422509.html.

Wallace:1991:AHS

[WS91b] Alfred Russel Wallace and Charles H. (Charles Hyde) Smith. An anthology of his shorter writings. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-857725-7. x + 551 pp. LCCN ????

Wright:1982:PMK

[WT82] Peter Wright and Andrew Treacher, editors. The problem of medical knowledge: examining the social construction of medicine. Edinburgh University Press, Edinburgh, Scotland, 1982. ISBN 0-85224-417-7 (paperback). 232 pp. LCCN RA418 .P77 1982.

Wujastyk:1985:BRR

[Wuj85] Dominik Wujastyk. Book review: A. Rahman, M. A. Alvi, S. A. Khan Ghori, and K. V. Samba Murthy, Science and technology in mediaeval India — a bibliography of source materials in Sanskrit, Arabic and Persian. Delhi: Indian National Science Academy, 1982. Pp. xxxi + 719. Rs.200 (inland), $70 (foreign). British Journal for the History of Science, 18(1):96–97, March 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026279.

Wussing:1984:GAG

[Wus84] Hans Wussing. The genesis of the abstract group concept: a contribution to the history of the origin of abstract group theory. MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-23109-3. 331 pp. LCCN QA171 .W8713 1984.

Waters:1992:JHB

[WV92] C. Kenneth Waters and Albert Van Helden, editors. Julian Huxley: biologist and statesman of science: proceedings of a conference held at Rice University, 25–27 September 1987. Rice University Press, Houston, TX, USA, 1992. ISBN 0-89263-314-X. LCCN 96.E03636; HUXLj.106.

Wallis:1975:IBM

[WW75] Peter John Wallis and R. V.

Wallis. An index of British mathematicians: a check-list. Issued by the Project for Historical Biobibliography at the University of Newcastle upon Tyne, Newcastle upon Tyne, UK, 1975. ISBN 0-7017-0000-9 (paperback: vol. 2), 1-871768-01-2 (pt. 3). LCCN QA30 .W35 1976.

Wallis:1986:BBM

[WW86] Ruth V. Wallis and Peter John Wallis. Biobibliography of British mathematics and its applications. Part II, 1701–1760. Epsilon Press [for the] Project for Historical Biobibliography, Letchworth, UK, 1986. ISBN 1-85027-003-1. xxiii + 502 pp. LCCN Z6651.

Wallis:1988:ECM

[WW88a] Peter John Wallis and R. V. Wallis. Eighteenth Century medics: subscriptions, licences, apprenticeships. Project for Historical Biobibliography, Newcastle upon Tyne, UK, second edition, 1988. ISBN 1-871768-00-4. xxvi + 690 + xix pp. LCCN R713.29 .W34 1988.

Wodeham:1988:TIC

[WW88b] Adam Wodeham and Rega Wood. Tractatus de indivisibilibus: a critical edition, volume 31 of Synthese historical library. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2424-5. vii + 333 pp. LCCN B105.I53 W63 1988.

Wallis:1993:IBM

[WW93] Ruth V. Wallis and Peter John Wallis. Index of British Mathematicians Part 3, 1701–1800. PHIBB, Project for Historical Biobibliography, Newcastle upon Tyne,UK, 1993. ISBN 1-871768-01-2. xiv + 175 pp. LCCN ????

Wylie:1989:YCP

[Wyl89] Ian Wylie. Young Coleridge and the philosophers of nature. Oxford English monographs. Clarendon Press, New York, NY, USA, 1989. ISBN 0-19-812983-1. x + 176 pp. LCCN PR4487.P6 W9 1989. US$33.75. URL http://www.

loc.gov/catdir/
enhancements/fy0636/
88021084-d.html.

Wynne:1982:RRW

[Wyn82]  Brian Wynne.  Rationality

and ritual: the Windscale Inquiry and nuclear decisions in Britain, volume 3 of BSHS monographs. British Society for the History of Science, Chalfont St. Giles, Bucks, England, 1982. ISBN 0-906450-02-0 (paperback). x + 222 pp. LCCN HD9698.G72 W96x 1982.

Yates:1972:BRS

[Yat72]  Frances Yates.  Book re-

view: Sixteenth and Seventeenth Centuries: Religion and the Decline of Magic. Studies in Popular Beliefs in sixteenth and seventeenth century England. By Keith Thomas. London: Weidenfeld and Nicolson, 1971. Pp. 716. £8. British Journal for the History of Science, 6(2):213–214, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025301.

Yates:1983:RRI

[Yat83] Frances Amelia Yates. Renaissance and reform: the Italian contribution, volume 2 of Collected essays. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1983. ISBN 0-7100-9530-9. xii + 273 + 16 pp. LCCN DG445 .Y37

Young-Eisendrath:1997:CCJ

[YED97] Polly Young-Eisendrath and Terence Dawson, editors. The Cambridge companion to Jung. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-47309-8 (hardcover), 0-521-47889-8 (paperback). xxxiv + 332 pp. LCCN BF109.J8 C36

                                        1996. URL http://www.loc.
                                        gov/catdir/description/
                                        cam027/96029101.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam023/
                                        96029101.html.

Yeo:1986:PPN

[Yeo86] Richard R. Yeo. The principle of plenitude and natural theology in Nineteenth-Century Britain. British Journal for the History of Science, 19(3):263–282, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026505.

Yeo:1988:BRB

[Yeo88] Richard Yeo. Book review: Bernard Lightman. The Origins of Agnosticism. Victorian Unbelief and the Limits of Knowledge. Baltimore and London: The Johns Hopkins University Press, 1987, Pp. x and 249. ISBN 0-8018-3375-2. £20.95. British Journal for the History of Science, 21(2): 263–264, June 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026994.

Yeo:1993:DSW

[Yeo93] Richard R. Yeo. Defining science: William Whewell, natural knowledge, and public debate in early Victorian Britain. Ideas in context. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-43182-4 (hardcover). xiii + 280 pp. LCCN Q127.G4 Y46

1993. URL http://www.loc.
gov/catdir/description/
cam025/92030620.html;
http://www.loc.gov/
catdir/toc/cam028/
92030620.html.

Yoshikawa:1994:SNN

[YK94] Hideo Yoshikawa and Joanne Kauffman. Science has no national borders: Harry C. Kelly and the reconstruction of science and technology in postwar Japan. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-24037-8. xvii + 137 + 9 pp. LCCN Q127.J3 Y677 1994.

Lam:1966:COG

[yL66] Lay yong Lam. On the Chinese origin of the Galley method of arithmetical division. British Journal for the History of Science, 3(part 1, 9):66–69, 1966. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

York:1989:AOT

[Yor89] Herbert F. (Herbert Frank)

York. The advisors: Oppenheimer, Teller, and the superbomb. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1713-3, 0-8047-1714-1 (paperback). xiv + 201 pp. LCCN UG1282.A8 Y67

1989. URL http://www.loc.
gov/catdir/description/
cam023/88062671.html;
http://www.loc.gov/
catdir/toc/cam027/
88062671.html.

Young:1966:BRH

[You66] Robert M. Young. Book review: History of Biological Sciences and Medicine The Mechanistic Conception of Life. By Jacques Loeb. Edited by Donald Fleming. Pp. xlii + 216. Harvard University Press; London: Oxford University Press, 1964. £1 14s. British Journal for the History of Science, 3(1):92–93, June 1966. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025121.

Young:1983:BRS

[You83] Gregg De Young. Book review: Science and Society D. R. Stoddard (ed.), Geography, ideology and social concern. Oxford: Basil Blackwell, 1981. Pp. vi + 250. £5.50 (paperback); £12.00 (library edition). British Journal for the History of Science, 16(3):301, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026170.

Young:1985:DMN

[You85] Bob (Robert M.) Young. Darwin’s metaphor: nature’s place in Victorian culture. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-30083-5, 0-521-31742-8 (paperback). xvii + 341 pp. LCCN QH361 .Y68

1985. URL http://www.loc.
gov/catdir/description/
cam031/84021377.html.

Yoxen:1978:HMB

[Yox78] Edward Yoxen. The history of molecular biology. British Journal for the History of Science, 11(3):273–279, November 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025780.

Yoxen:1979:TRB

[Yox79] Edward Yoxen. Television review: The Body in Question. Written and presented by Dr. Jonathan Miller. BBC2 (UK), November 1978–February 1979. British Journal for the History of Science, 12(2):237– 238, July 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025772.

Yoxen:1981:SSP

[Yox81] Edward Yoxen. Socialism and scientists’ politics. British Journal for the History of Science, 14(1):78–83, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026074.

Yoxen:1982:BRS

[Yox82a] Edward Yoxen. Book review: Science in Different Countries Philip Gummett, Scientists in Whitehall. Manchester: Manchester University Press, 1980. Pp. xi + 237. £14.50. British Journal for the History of Science, 15(2): 195–197, July 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4025970.

Yoxen:1982:BRN

[Yox82b] Edward Yoxen. Book review: The New Physics: Television review of ‘Oppenheimer’, 7 part serial, shown on BBC-2 in the UK in 1980 and of Peter Goodchild, J. Robert Oppenheimer ‘Shatterer of Worlds’. London: BBC Publications, 1980. Pp 301 incl. index and illustrations. £9.95. British Journal for the History of Science, 15 (2):204–207, July 1982. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4025976.

Yoxen:1983:BRLa

[Yox83a] Edward Yoxen. Book review: Life and earth sciences: Andre Lwoff & Agnes Ullman (ed), Origins of molecular biology: a tribute to Jacques Monod. New York, San Francisco & London: Academic Press, 1979. Pp. x + 246. £15.40/$23.50. J. D. Watson. The double helix: a personal account of the discovery of the structure of DNA. A new critical edition including text, commentary, reviews and original papers, ed. by G. S. Stent. London: Weidenfeld & Nicolson, 1981. Pp.

298. £10.00. British Journal for the History of Science, 16(3): 278–281, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026150.

Yoxen:1983:BRLb

[Yox83b]  Edward Yoxen.  Book re-

view: Life and earth sciences: J. D. Watson and J. Tooze, The DNA Story: a documentary history of gene cloning. San Francisco and Oxford: W. H. Freeman and Co., 1981. Pp. xxiv + 605. £13.95. British Journal for the History of Science, 16(3): 282, November 1983. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026152.

Yoxen:1983:GBW

[Yox83c] Edward Yoxen. The gene business who should control biotechnology? Crucible. Science in society; A Channel Four book. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1983. ISBN 0-330-28112-7 (paperback). 264 pp. LCCN TP248.3.

Qian:1985:GIS

[yQ85] Wen yuan Qian. The great inertia: scientific stagnation in traditional China. Croom Helm, London, UK, 1985. ISBN 0-7099-2104-7. xii + 155 pp. LCCN Q127.C5 Q24 1985. US$27.00.

Zambelli:1986:AHSa

[Zam86] Paola Zambelli, editor. “Astrologi hallucinati”: stars and the end of the world in Luther’s time. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 1986. ISBN 3-11-010317-6. LCCN BF1711 .A88 1986; BR333.5.A8 A8.

Zambelli:1992:SAE

[Zam92] Paola Zambelli. The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries, volume 135 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346. xvi + 352 pp. LCCN Q174 .B67 vol. 135. URL http://

                                        www.loc.gov/catdir/
                                        enhancements/fy0826/
                                        91024566-d.html;   http:/

/www.loc.gov/catdir/
enhancements/fy0826/
91024566-t.html;   https:
//link.springer.com/book/
10.1007/978-94-017-3467-7.

Zandvoort:1986:MSD

[Zan86] Henk Zandvoort. Models of scientific development and the case of nuclear magnetic resonance, volume 184 of Synthese library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1986. ISBN 90-277-2351-6. xiii + 305 pp. LCCN Q175 .Z3 1986.

Zott:1992:SGJ

[ZH92] Regine Zott and Emil Heuser, editors. Die streitbaren Gelehrten: Justus Liebig und die preussischen Universitäten: kommentierte Edition eines historisches Disputes. (German) [The contentious scholars: Justus Liebig and the Prussian universities: commentated edition of a historical dispute], volume 15 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. ERS-Verlag, Berlin, Germany, 1992. ISBN 3-928577-13-1. 289 pp. LCCN QD49.G3 S77 1992.

Zinsser:2006:EDC

[ZH06] Judith P. Zinsser and Julie Candler Hayes, editors. Emilie Du Châtelet: rewriting Enlightenment philosophy and science, volume 2006:01 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2006. ISBN 0-7294-0872-8. xi + 325 pp. LCCN PQ1981.D55 Z66 2006.